A Bibliography of Publications in Applied Numerical Mathematics: Transactions of IMACS

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

17 October 2025 Version 1.159

Title word cross-reference

(−1, 1) [OR20]. (0, +∞) [DO17a]. (2 + 1) [AJ24b, TWMP20]. (λ^{∗}, λ) [Ma24]. (m, ℓ) [Fuj02]. (p, q) [WJF25]. (q, h) [GS15b]. (r + 1)^{st} [ZZW97]. 0 < α < 1 [DCJ20]. 0 < θ < 1 [TM05]. 1 [EP15, HWCF15, KM19, KD13, KKR15, LW07, Now96, SS00, Vej10, WC24b, ZL18a]. 1 + 1 [MPSS16]. 1/t [AM16b]. 1s [CPD^{+}05]. 2 [AG05a, AEMX17, ASS21, Ant13, AC96, AN22, BS21, BF15, BNS25, BDRZ19, BCV21, CNA23, Cau08, CCM02, CCM17, DMPSC16, Fou00, FS24, Gon06, HGP11, HHR12, HA21, Hey20b, KFR25, KTYY24, LPR00b, LLHC17, LJ20a, MDA24, Mur99b, OT22, PMP23, RZ00, Sch02, TCCW89, UHUL21, VA05, WBCK02, WPT19, YM24, YZ17, ZS21b, ZS18]. 2d [SL17]. 2r [MRÁSY24]. 3 [ASS21, Ant13, ASC03, BB25, BV94, BM18, BS08, CK22, CPC25, CPZ17, FMS18, FLH22, GS20, HDY21, HW15, HZ12, Jay95, Kal96, KNP16, KLSW06, LS21, Mai06, MSGM23, MKH16, MG22, MM20b, NBNTGV11, NRR06, Ost93, PSP05, PK23, PPC00, QPT23, Ran15, Sha98, Ste97, TYKK01b, WSY18, WSC09, WSC21, YH00, Zha09, ZS21b]. 4 [BPV25, DC21, EHM01, PT23, Ran15]. 6(5) [CV95]. 7 [AB15]. 7(6) [CV95]. 8 [PPS05]. 9(8) [Tsi01]. > [DLN^{+}24]. [H^{−1}, L^{2}, L^{2}] [KLY05]. ^{0} [Hop23]. ^{1} [HAC22, PZMX16, Zho18]. ^{2} [KDD23]. ^{∞} [Hua98]. ^{m} [BE99]. _{0} [THW19]. _{1} [Kan04]. _{2} [BHRY21, TK19]. _{2}F_{2} [JJ15]. _{3}F_{2} [JJ15]. _{q} [Bec02, BHRY21]. A [BC89a, Cha96, ST89, GPHAM12, SCvdH92, ZZW97]. α [KR25, lLH25, WCS21]. α ∈ (1, 2) [CDH24]. at [BB24b]. Ax = b [Cul95]. B [AFK92, FSU89, HCX03]. β [AFS00, MDD18, Sus10]. C^{0} [DGE22, LHC09, LWY20]. C^{1} [DS17, ZLY23, Ber86, RL86]. C^{2} [AR24]. C^{4} [Ver93]. C_{1} [Arc06]. H_{2} [FG13]. H_{α}^{s} [WWL21]. cos(z) [VC10]. cos π/2x^{2} [Coh90]. d [BDV17, DM97, GÖS20, LL02, MdR05, PSP04b]. δ [XZL07]. Δ^{2} [Kam25]. θ̇ [Bho11]. ℓ^{2} [WSY18]. ℓ_{1} [CXZ14]. ϵ [FR01, GGS04, Tan87]. EQ_{1}^{rot} [ZSS23]. η [BSGU94]. exp(z) [IN89]. F_{n}(z) = f_{n} ∘ ⋯ ∘ f_{1}(z) [Gil91]. G [Lev91a, BC89a, BHHS10, HJ21, NZY21, SS10]. G(ϵ) [BIJ23]. G^{2} [FPT03]. γ [JN07]. GL(E) [Ben98]. H [Wan96, HAC22, MS08b, XYZ24, ZZ24, AOW94, Mus11, SS00, TM05, Usm97, VO00a, XY19]. H(div) [HSF25]. H^{1} [KS02, SL22, WCS21]. H^{θ} [TM05]. H_{1} [ZJH18]. H_{2} [ZJH18]. Â [DLV24]. Î [DLV24]. hp [AS97, AK00, BTP96, Bür12, DF96, FM95, GHH09, Mus11, SGS00, tSqWyG16]. ILU [PRGO16]. k [DS97a, GH91, Jac88, KF97, Lem88, ZZF25]. K_{2}(P_{2}) [SHA12]. L [FRV11, KDD23, YXX24, dACR10, BHRY21]. L^{1} [GKA17, CR23b, Muo23]. L^{2} [Lee23, LCH20, Ort20, Shi25, ZWG25]. L^{2}(Γ) [YZ24b]. L^{∞} [LZ14]. L^{∞}(L^{2}) [Dek17]. L^{q} [Dol14]. l_{1} [CLL23, YP18b]. L_{2} [MB08]. L_{∞} [CC20a, LZW19]. L_{p}(R) [DMA22]. Lq [DKK94]. LU [Doi91, Phi91]. M [GP17, SHL19]. R^{2} [Kim14]. R^{3} [HOS11]. H [BB24a]. HSS [SB25]. M [JL23b]. O(ϵ^{−∞}) [KSJY25]. N [GT19a, Gil10, HOS99, Lei99, Lei02, LZW19, MV20, PRST02, DhW09, Lev91b, Li16, SA90, SWCH15]. N^{−1} ≤ ϵ [ZL21]. P [CKL03, LC20, SS09, VDVV98, AL09, AOW94, ADM22, ADFP25, AM16a, BP02, BS96b, CS09, FVGS13, Gwi09, JJ94, KJ99, KX03, LZJ21, Mai06, ML91, MS00, PPS10, PK21, SS00, VO00a, XY19, Yos00, dG91]. p − 3 [Ver06]. P_{0}^{2} [HJYL19]. P_{1} [VB07, HJYL19]. P_{∗}(κ) [YZH19b]. P_{N} [GS20]. ϕ [LZZ22]. ϕ^{4} [Sus10]. Ψ [Rah25]. Q [LTDY24, AHGM21, TJK18]. Q_{1} [WCXL09]. Qrot_{1}/Q_{0} [CH07]. Q_{n} [Pav00]. Q_{n−2} [Pav00]. Q_{p} [ST08]. qd [DS10, AC10]. QR [DS21c]. QZ [DSW96]. r [Zeg97, ZZW97]. R^{d} [LL02]. R^{n×n} [SB25]. s [DE18]. S/P [Not92]. s = 1 [Beg00]. S_{n} [YH07]. σ, μ [Ort20]. sin(z) [VC10]. sin π/2x^{2} [Coh90]. \sqrt{3} [JP17]. T [CJM88, JP19, AX20, EJRR23]. Θ [Fdi97b, CHZ14, CD20a, DFLM19, Gan09, GHHG22, LZQ22, RT20, RU25, RFMV24, Sca22, WGW15, BGT97]. V [CK98, NN10]. ε [DS97a]. φ [Gem23]. W [SW95b, Nov08, Sch95b]. x′′(t) = Λx(t) + f(t) [ST20]. y′ = f(x, y) [CP05a].

-adaptive [BTP96, Usm97]. -additive [MSGM23]. -algorithm [AC10, GGS04, Tan87]. -anisotropic [GHH09]. -averaging [FVGS13]. -based [MS08b]. -Bernstein [GS15b]. -biharmonic [ADFP25]. -blocking [AFS00]. -body [HOS99, Lei99, Lei02, PRST02]. -Brownian [HJ21]. -circulant [lLH25]. -conforming [HSF25, PZMX16, ZLY23]. -continued [Lev91a]. -continuous [LWY20]. -convergence [AFK92, HCX03]. -coupled [LZW19]. -covariance [FSU89]. -cube [BDV17]. -cubic [FPT03]. -cycle [CK98, NN10]. -cycles [KF97]. -Cyclic [PPS10, dG91]. -d [CCM17, WC24b, Kal96, Now96, YH00, AG05a, Ant13, KNP16, Sha98, TCCW89, TYKK01b, VA05]. -decreasing [DKK94]. -designs [AX20]. -dim [HA21]. -dimensional [AJ24b, BM18, DhW09, Li16, SHL19, SWCH15]. -dimensions [TWMP20]. -div [XYZ24]. -EM [LZQ22]. -energy [DE18]. -extension [TM05]. -FEM [MS00, Yos00]. -finite [AS97, Bür12, SGS00, Lee23]. -fraction [CJM88]. -fractional [Rah25]. -fractions [Lev91b]. -function [Sus10]. -functions [Gem23, LZZ22]. -Galerkin [SL22]. -Laplace [LC20]. -Laplacian [ADM22, FVGS13, LZJ21, PK21]. -lattices [AHGM21]. -mapping [XZL07]. -Maruyama [WGW15]. -matrices [Wan96]. -matrix [HB02, BB24a]. -method [CHZ14, DFLM19, RU25]. -methods [CD20a, Gan09, RFMV24, Sca22, SW95b, Nov08, Sch95b, BGT97]. -Milstein [GHHG22, RT20]. -model [JN07]. -neuron [Bho11]. -norm [CXZ14, HAC22, Dol14, LCH20, WCS21]. -order [ZZW97]. -orthogonal [DM97, MdR05, TJK18, CKL03, FRV11, dACR10]. -penalty [Zho18]. -pentadiagonal [WJF25]. -periodic [Jac88, Lem88]. -phase [TCCW89]. -point [BPV25, EHM01, LS21, ZZW97, Gil10, MV20]. -product [EJRR23]. -property [THW19]. -quadratic [LL02]. -quasilinear [Ben98]. -refinement [DF96, FM95, JP17, Zeg97]. -Robust [WCS21]. -soliton [GT19a]. -spectral [tSqWyG16]. -stability [BC89a, Cha96, DLV24, Ort20, Shi25]. -stable [ST89, GPHAM12, SCvdH92, SS09, VDVV98, YXX24, ZZW97]. -stage [NBNTGV11]. -step [GH91]. -superconvergence [BSGU94]. -Sylvester [JP19]. -symplectic [BIJ23]. -tensor [JL23b]. -tetrahedra [PPS05]. -th [AL09, AB15]. -transformation [BHHS10, NZY21]. -triangles [PSP04b]. -type [Fdi97b]. -uniform [FR01]. -version [AK00, AM16a, CS09, Gwi09, JJ94, KJ99, Mai06, ML91]. -versions [Mus11]. -virtual [DGE22]. -weighted [Muo23, LTDY24].

/Smad [MDD18].

0012 [De 02].

1 [BRS^{+}18, WL24]. 1-D [WKM04, WL24].

110 [Bic21]. 125 [YP18a]. 12th [MH89]. 16 [Kni95]. 160 [AS21]. 18 [Tsy96]. 19 [CML^{+}25]. 19th [Ano02g]. 1D [DGGM25].

2 [AAB^{+}22, BRSD91, BS94a, JLH13, SSR23, WVBM88]. 2-D [SSR23]. 2009 [Wen10a]. 2021 [Ano21p, Ano21o, Ano21s, Ano21t, Ano21n, Ano21q, Ano21m, Ano21r]. 2022 [Ano22q, Ano22t, Ano22s, Ano22w, Ano22u, Ano22p, Ano22o, Ano22v, Ano22n, Ano22x, Ano22m, Ano22r]. 2023 [Ano23t, Ano23r, Ano23m, Ano23v, Ano23s, Ano23o, Ano23n, Ano23w, Ano23q, Ano23p]. 2024 [Ano24u, Ano24o, Ano24r, Ano24n, Ano24m, Ano24v, Ano24t, Ano24q, Ano24s, Ano24p]. 2025 [Ano25s, Ano25o, Ano25u, Ano25w, Ano25y, Ano25v, Ano25n, Ano25t, Ano25r, Ano25x, Ano25q, Ano25p]. 2026 [Ano26b]. 20th [HSX18]. 2D [BP95, Chu03, FZA25, LTC03, YHT23, ZZ25]. 2D/3D [YHT23]. 2nd [FJ97, KP07].

3 [BtTBV87]. 3-dimensional [YL26]. 3d [EH05, CML05, FZA25, YHT23, ZZ25, de 92b].

4-th [KK23]. 4th [FL24b, LHX20].

4th-order [LHX20].

5 [VV95].

60th [CHM09]. 61 [ZCSH11a].

8th [vdHSW98].

90-th [EST15]. 95j [Kni95]. 96g [Tsy96].

=1 [Bra00].

A-EFIE [LCJQ12]. A-posteriori [BM13, CMP03, MWC21, SZ09, Sha21]. AB2 [LT12]. Abel [DSK12, MK21, PRS23, SVB17, SS16].

Abramowitz [Mac92]. abrupt [Sal93]. abscissae [Joh05]. absolute [CYYH21, DS20, DSS20, Guo26, JL21, MYSC17, NSCC19, YW25]. Absorbing [BG11a, EZ98, TY98, AG98, BDSG09, BP97, GD22, GD23b, HD04, PXHZ20, PG21]. abstract [RT14]. Accelerated [He24, Liu24, TWH21, BBC^{+}25, Bog12, Boy91b, HPY92, LLY24, YEZ25, Zen21]. Accelerating [PKSB10, ZB19a, ITZ17]. Acceleration [BL86, DE16, HL02b, SC08, VVR08, ACR25, Ben98, BRZ10, Fdi97b, FL01a, GM94, Hau88, Ise94, Kam25, Kru99, Kza92, Kza99, Lav94, LS99a, Lem88, Lev91b, Lev91a, LSGK15, LHT20, Sid10, Zha97]. accelerator [PCA10]. accelerators [CRHTV25, Cuy90]. acceptance [HP15]. Accuracy [BVB09, FG09, Hun02, LWD^{+}09, Tur86, Bal00, BH12a, BPTT15, DT10, DS15, EN09, GMGF02, GT18, KD25, LH21, LHC23, LLZL24, LK07, NW09, Rou20b, RGK21, Rum87, SG00, SGN08, Tsa91, Tsa92, XXYZ24, XY24, YF24]. Accurate [CH95a, Enr06, FL04, Hua19, JBLC11, JCSR03, KK17, vdES04, ACP24, AES15, ACLM22, AHB20, AS20c, BP14, BF15, Bos09, CGA93, CL85, CS08, CRU15, DRVA20, DL22b, ECHF^{+}20, Fac03, FSX25, GHS25, HHR12, HOEC86, IR22, Kop89, KLSW06, KK20b, KK22b, KDS22, LO22, Li23, LGZH24, LAH09, LBCN00, MCS16, Mat05, Nic86, PNA21, Qiu23, RZ00, RA03, Sac93, SC20, Sid23, SS10, TDPU17, WDU21, WPN^{+}24, WLCZ26, Zak20, ZWJ18]. accurately [SB18]. achievements [AB09b]. achieving [RB12]. acidic [BMP05]. Acoustic [AL87, HD88, CAAT16, GGN12, GKB^{+}22, HD04, HM86, HMP14, KJL12, KL07, LH11, Lee94, MN24, OL18, RX08, YL26, ZP24, ZL24]. acoustics [AD15, Har98, HX11, LY16]. acousto [ZZ25]. acousto-electric [ZZ25]. across [Gea93, GX93, Jéz99, Pag25]. act [Ise02]. activator [ZZJ21]. activator-inhibitor [ZZJ21]. Active [CG13, NU15, CPY20, HO16, HMW05, LWaZ24, LT05, RU07, Sch09]. Adams [ZJH18, AM95b, AGM95, IM98, JL91, Sha87]. Adams-type [AM95b]. Adaptation [BMGM12, Bai97, BS97a, But92, CLGD06, FMP04, JM94, Ram94]. adaptations [in 96]. Adapted [AEG12, Aso21, Fra04a, FM11, GHW01, HT20, KTYY24, Lin01, LE94, Mat09, RN25, TZ00, TS24b, Van00]. adapting [Kos02]. adaption [BS94b]. Adaptive [AFS11, BCS17, BGS06, BB24a, BBRBS09, BH97, CP97, CSX23, DGM22, DGD03, DM10, DJJ^{+}15, FL09, GZZ19, GHKM09, GS89, IM00, KPRU20, KESYB23, KCW16, KMS10, KS09c, Lan97, LOS03, LGH11, MS05, MKBS26, MT05, Mol95, NRWF08, PSP04a, Pet87, Sar05, SFJ^{+}05, YPD21, ZWK15, AJ24a, AS97, AF23, ABR05, AF89, AS06, Aug89, AFLG^{+}12, AFLP12, AFF^{+}15, ÁMS17, BLS94, BHJ05, BS21, BS14b, BKP09, BKR13, Beh97, BS00a, BC08a, BCU00, BF92b, BTP96, BLY16, BFA93, BDF94, BV94, Bor97, BHR05, BF95, BD11, Bür13, CHR03, Car09a, CK22, CR05, CXZ09, CSL^{+}24, CL26, CL01b, Chr96, CKM25, CY05, De 06, DF96, Dol14, DMR18, DGRS09, DH07, Ein18, FD97, FVB05, FMGN94, FM95, FLÖ^{+}97, GMZ11, GV18, GGO12, GGO16, GS20, GDS^{+}15, GM17, HKZ08, He24]. adaptive [HS22, HW97, Hop23, HWCF15, HSY18, IS23, JD09, JGK11, JOL23, Kal96, KDT17, KL23b, KS09b, LW93, Lan95, Lei99, LD10, LLY24, LLN25, LLS^{+}96, LC20, LQY25, Log04, DLM16, lLX22, Man96, MRF00, MZS10, MMT90, MR94, Mit97, MCE^{+}09, Moo95b, NK11, NMKE13, NRR06, NBP94, Now96, OT22, OZ96, PD96, PSWZ21, Pic05, Pow94, PRS20, QCW^{+}23, Qui96, Ran15, RS20, SR09, SWW17, Shy86, Shy91a, Shy91b, SG09, SB19, TLQ21, Tho85, TLV92, TDW23, Tse00, Usm97, VO00a, VO00b, VNC21, Wal00b, WKM04, WPL16, WSS97, WZ17, XY24, XZZL15, XHYM22, YXF25, YH18, YR22, ZH15, dDF^{+}94, dSFDG20]. adaptively [Du11]. Adaptivity [Söd06, Bac17a, BCJ97, BLP01, GHH09, Roz05, YT25, DGN12]. Addition [MA04, Shy86, Shy91b]. Additive [GGR97, KC03, KFOF02, Bac18, BTBR20, CGH23, CL18, DLN04, FL01a, GCHR06, HR06, HLY22, KC19b, Kha21, KZ13, LT19, Mai09, MSGM23, MH04, MT20, NN13, Pan07, ST19, TN16, WCXL09]. ADER [BBD20, ÖPT25, TH09]. ADER-DG [BBD20]. adhesion [NS12]. ADI [AD20c, APJ09, BHSW16, BHSW20, Bog16, CC20a, CQZ20, Den15, FG13, GPHA22, HEG16, itHT18, JM06, LZCF21, LGZH24, LCZ25, MG22, QXQ22, WLM21, WCS21, WCM21, WB03, WPL16, WPT19, YQCZ22, ZZjQ24, iW07, iW09, iM13]. ADI-type [GPHA22]. adiabatic [Rei99]. Adini [HLY04, LH02]. adjacent [Bre91]. Adjoint [BL05, HS97, KS09a, ARSW05, CGT13, DG10, Gul15, KOGL25, RK08, Rob10, SW09a, YLL21, ZX22]. ADMM [MBS23, SZY21, SGY22]. adsorption [SXL22]. Advanced [FDF^{+}25]. advancement [SSS21]. Advances [Pro24, SP99, BRW21, BM09, JN02, Spi99]. advected [AD15]. Advection [IHS13, AJ24b, ABR05, BN12, BM04b, BM04c, CdFN01, CL24a, CK20, CRSF19, FMP04, GPPR12, GPHA16, GS94, HM15, HAA21, Hin95, HT94, HV95, JRS20, JR18, JL17, KHB22, KOW05, KL09, KZ21, LKV01, LR18b, LWYG22, LLT20b, lLNW21, LS21, Liu09, LYZW22, MN23, MSGM23, MOU14, NLHT25, NMB10, ÖPT25, Ort20, PSWZ21, wSJP15, SA19, SvdVvD06, TS23, Tro96, Usm97, ZM19, ZG21, ZLWF21, dlHV13]. advection-diffusion [CRSF19, GS94, LS21, LYZW22, NLHT25, PSWZ21, dlHV13]. advection-diffusion-reaction [GPPR12, MN23, wSJP15]. advection-dispersion [JL17, ZM19]. advection-dominated [Hin95]. advection-reaction [HV95, KZ21]. advection-reaction-diffusion [CdFN01]. advective [FMU15, GJR03]. aeroacoustic [AF04]. aeroacoustics [HG98]. aerodynamic [Jam93]. aerodynamical [De 02]. aerosol [DS07b, SB03]. aerosols [DS07a]. aerospace [Buc04]. Affine [WD22, BMM03, BSZ99, GGO13, dRT99]. Affine-invariant [WD22]. affine-scaling [dRT99]. after [For25]. after-effect [For25]. against [Jes93]. age [AM95a, DFLM19, DZMB21, LZY24, MPV24, NK24a]. age-dependent [DFLM19, DZMB21, NK24a]. age-of-infection [MPV24]. age-structured [AM95a, LZY24]. agent [LSGK15]. agent-based [LSGK15]. aggregation [AL20, GJV08, LS16, Mar03]. aggregation/disaggregation [Mar03]. Air [MZ87, DS03]. airfoil [De 02, KTS03, OK98]. Aitken [BKP15, Fik23, Kam25]. AKNS [Ma24]. Alexisbad [vdHSW98]. algebra [BRW21, HP91, LSK12, ZJR25]. Algebraic [AHA23, DCSB25, DSV13, JS25, KS89, PS09, ST89, VV02, AB10b, Arn93, ALP^{+}96, Bai96, BL05, BCG21, BKP09, BHSW16, BHSW20, BDF89, BVV09, BG02b, BS18, BS20b, BB24b, BIS25, BC95, CC90, CG92, CCM02, CF13b, CC20b, CM04, CMS04, CMS06, DS05, DS21d, ELLE02, EGL09, FTB97, Fuh01, GMG19, GGR97, GJV08, GS92, HGP11, Has09, HY02, Hig93b, Jay95, JW01, JLZ25, KM17, Kau93, KP15, KKR15, LP05, LPZ00, LLZ^{+}22, LYLL23, LM22b, LS98, Luc05, MP96, MDT05, Mar99a, Mär95, Mär02, MT11, Mur99b, Oji88, Ost93, PPT02, Pis22, Pul09, PSL18, RGS25, RA05, SST04, San03, SA12a, SMEN04, SM93, Sch12, Sch98, Sch02, SG06, ZFC20, ZP97, ZP98, Zla85a, vB95, Ney95]. Algebraically [HH10a, LH24]. Algorithm [Jac87, RBBC85, ZW87, AG05a, AB97, AC10, AM16b, AL05, BHJ05, BC99, Baz03, Bec02, BLY16, BDFV95, Boh03, BSV21, Boy91b, BIS25, BWEP95, Cao03, CL08, CW20, Car94, Car09b, CS04, Che88, CZ90, CCY22, CLL23, CL01b, CMS06, CRSF19, DZ12b, DW15, Din19, DSAB20, DS02, DSW96, DHS05, DH94, DS10, DY03, DZMB21, DJJ^{+}15, EAS12, FLS94, FD97, Fdi97b, FBM17, Fre91, GZW22, GGN12, Gla94, GGS04, GH20, HS02, HJR22, Han19, HM15, HD23, Hua98, Hua17, HT20, Ibr25, IS22, IB24, Jia00, Jia02, JY20, JBLC11, JP93, KBK21, KCS07, KKR24, KM21, Khe91, Kni94, Kni95, KDS22, LTC03, LZ13, Leo10a, LG87, LD10, LHC23, LWLW24, LLY24, LHX20, LSWW22, Liu24, LQY25, Lo06, LWLL24, MZJ^{+}25, MK14, Man96, MD23a, MC00, NH24, NMB10, O’L87, OZHP23]. algorithm [Ou11, OMSS25, PGS10, PM91, PLI03, Pic05, PPC00, Poh93, PRS23, PMP23, PM14, QCW^{+}23, Qiu23, Qui96, RA03, Ria22, RU15, Sad97, SHL19, SAA20, SMC08, SNOK21, SWW17, She00, SXP09, SYW22, SS10, Sob24, SC22, TWMP20, Tan87, THW19, TLGC22, TLV92, TGV22, Tou97, Wag98, Wai98, WZ14, WOLY25, WB90, Woź10, WYP12, WPS18, XY25, XZL07, XZZL15, XGQ20, Yam23, YH18, YXZL24, YWH20, YL26, YZH19b, YLH20, YLW20b, ZXYW22, ZY23, ZLS20, ZP12, ZZ25, ZB19b, HM17, LFS15]. algorithmic [Sae14, BV94]. Algorithms [DGCW17, GV18, KP07, LS16, NRR06, ST86, VVV24, ZWK15, AY21, APA92, AGKK94, ABCC18, BS91, BS93, Bre96, BHHS10, BF95, CL01a, CG92, CP97, CG89, CDW13, Con89, CSM07, Cui04, DE16, DMM24a, DB08, DC09, FS19, FL93, FG01, FS88a, For02, FL01a, FN95, FCW20, GHKM09, GS89, GPMPR03, GN86, GLW^{+}25, HKZ08, HP91, HGZW21, HL19, HWY20, HH10b, IMM04, Jam93, JMS99, JW01, JMP06, Kim95, KFOF02, KW12, Lab98, Lab99, LMV25, Li11, LGH11, LLL12, LS93, MH16a, Mai06, MP98, NT20, Noo95, OP04, PSB91, PRST02, PH91, PT15, QAMX17, QM19, Riv09, RRMJ12, RN04, SP99, SY03, Sch91, SH10, SQ17, SKO19, Söd06, Spi00, SSKS21, Ste05a, SZ99, SL01b, SND19, TQY24, WH19b, WS21, Wan23, WOLY25, WG18, WWF20, XWC25, YK04a]. algorithms [YP18a, YP18b, ZC91, ZHL22, Zha07, ZS21b, dPT96]. aliasing [BSQ96]. Allen [BEH24, BCM04, GZQS23, GAX24, HPH20, HGZW21, HZZ25, JZXJ21, KSJY25, LMPS19, LCK22, TZ21, YYZ23, ZXYW22, ZYQS21, ZYQS23, ZD25a]. allows [GGS04]. alloys [WM08]. almost [But97, Vic92]. Along [Eis86, MPDB24, uIVS13]. Alpert [HAML21]. Alternating [SN22, AD20c, ACM91, BLD17, BMSZ21, BCDR25, CXZ14, DP85, DY03, GX11, JLZ25, JL94, LZH^{+}25, LLZ19, Liu21, MPHFP23, MLB97, Phi87, QWX20, SS25, SI20, WH23, ZHJ14, ZN21]. alternating-direction [ACM91, JLZ25]. alternative [BDDV12, CP10]. alternatives [ADSS17]. Althammer [LMV25]. ambiguity [CFV10]. American [ALY03, BAD13, Bi20, CXZ15, GK22, HFL13, RP17, ST11, ZD25b]. AMG [DSSC13]. AMLI [AV91]. AMNMAS [FDF^{+}25]. among [CLP15]. amorphization [ZAB15]. Ampère [JLWY25, SWR11]. amplification [SKS^{+}25a]. amplification-fitting [SKS^{+}25a]. amplifying [KSK^{+}25]. Amplitude [PM05, CJ90, Lyn99, Mur19]. Amplitude-shape [PM05]. AMR [BS94a]. AMSS [CF13a]. analyses [DL20, HHR12]. Analysis [AHT17, AS20a, AMCR17, Ano02g, Bac16, Bac19, BBD18, BK17, Bec18, BCCHM21, BBBN21, BJS12, BC23, BGHR12, BGH^{+}15, Cai15, CCQ^{+}23, CDGA17, CR23a, CH07, Cod08, CG16, CA15, Cui04, DA16, DGCW17, DW21, DG22, yDqGnJT09, ELvdHS98, EN09, Fra16, GMM09, GS18, GJV08, GGO13, HN03, HGF24, HJZ23, HW04, HS96, HS09b, IV16, JM25, Jbi25, JEG10, KTD20, KLY05, KvyS15, KKE16, LHS00, LD25, MOV24, NV23, ÖPT25, OMSS25, Par14, SZE20, SRMDRL23, SLMD21, Sid14, TMS87, VCC12, VT93, WR20, WOLY25, XLZ23, YX25, ZdBT03, vSW90, vS93, AG98, AGLRS23, ABdSG23, AA22, AKGR14, AMH24, ADNR21, ADFP25, AGP97, AL17, An20, ASV19, AGQ^{+}24, AMH03, Arn95, AC16, AAD14, ACR25, AS20c, BBV13, BMSZ21, BT97a, BS21, Ban97, BM13, BGG04]. analysis [BBS11, BCFQ19, BCFQ21, BM00, BLRGVR23, Ben17, BN12, BRVC09, BC04a, BLM17a, BDFF23, BW15, BS09, BFLR23, BM09, Buc99, BS12, Bus06, CFCH09, CCOVF22, CFKS07, CG03, CGMS21, Car23, Car94, CGG25, CST18, CSW19, CKB13, CXZ09, CHS19, CWX21, CKK10, CL18, ÇK13, CD20a, CDJT06, DMS23, DD19, DKSS24, DEPS15, Dav98, DLN^{+}24, DG10, DR09a, DA17, DA19, DK20, DYX09, DLM20, Dob05, DLN04, DLQZ23, EE20, EL01, FF20, FFY08, FV85, FTB97, FH10, FSB97, FR14, FLR08, FL24b, GAML04, GT15, GLV03, GGN12, GHH09, GD21, GAZD25, GGMP88, GRLL01, GLLY25, GKT10, GS94, GJLL20, GH20, GZHQ23, GAX24, HZ09, Har93, HH22, HZ20, HZD21, HM00, HB20, HL19, HH10b, HH18, HM22, HAC22, HO16, JMDN^{+}22, JMU^{+}25, Jac93, JKN94, Jam93, JLZ20]. analysis [JT88, KW21, Kel85, KHM^{+}19, Kok08, Kom07, KS04, KQ13a, KQ13b, KN93, KD25, KKR15, KAS17, LRS23, LHC09, LPT94, LZ13, LRV25, LWD^{+}09, LPZ00, Li01a, LW19a, LSWM19, LSP20, LW20a, LLVX20, LLY21, LA21, LWaZ24, LGC25a, LGC25b, LC25, LSG24, LZW17, LT01, Liu09, LW18b, LHX20, LS20, LCZ21, LCZ25, LD22, LARGVR23, LRT99, LRE04, MK20, MB08, Mag91, MO17, MP96, Man97, MMDH19, MT11, MN24, MM07, MCBV20, MW24b, Mir20, Mit22, MFAD23, MD25, Mok17, NN20, Odi19, OL18, Ort20, OS12, OCVW22, OXW25, PD96, PKP19, PFHL09, PA05, PH17, PSW02, PSL18, Qi24, QZH25, Quy19, Ran15, RSK24, RU25, RLHC19, RKR20, RAOC18, RGL16, RGMO19, RREP^{+}20, SHL19, SA21, SMEN04, Sch16b, SS94b, SS19, SW94, SZ12, SY18, SJ20, SL20, SW20a, SW21, SP22, SNW22, SD24b]. analysis [ST25, ST05, Sou09, SHG86, TLP18a, TLP18b, TX18, TH18, TLGC22, Tem23, Toc01, TZA13, Ton04, TLSS09, TD09, VBVA22, WDZS21, WY02, jWqW09, Wan17b, WSY18, WJW19, WG19, WW19, Wan20, WCS21, WHL19, WYY20, WL21, XZW19, XZZ19, XP23, YM24, YMD21, YZH19a, YLY19, YZC21, Yan22, YXZ18, Zak19, ZAED21, ZZ18, ZSG^{+}20, ZLW20b, ZLW20a, Zha20b, ZYQS21, ZLX22, ZYC22, ZZZ23, ZSS23, ZL23, ZBNC25, ZZO16, ZZX19b, ZH24, Zhe07, ZKO^{+}21, ZD25b, dPT96, dlC23, vdHVW01, BRW17]. Analytic [JR02, LZL14, Fat10, KMH21, Kza92, LSL11, Maj20]. Analytic-numerical [JR02]. Analytical [ADR17, BRS^{+}18, ESEKZ10, HD04, MN23, AD19a, Boy91a, BH96, CSS19, GQ89, KDH20, SPYS24]. analyzing [MMKN17]. Anderson [ACR25]. anelastic [Abr93]. anew [Mär02]. Angelesco [Lee10]. angiogenesis [DGGM25]. angiograms [QM03]. angle [ADK94, WL24]. angular [Sch23]. anharmonic [JZJ10]. Anisotropic [AL98, CDRT19, CZS04, CYM09, DMP08, Dol14, FMP04, Sim94b, AG05a, AGJM04, AKL08, DHM09, ESS15, FH08, FM11, GHH09, Li01a, LL20a, LQXK23, MCS06, MM02c, Ngu15, Obe15, OH20, QPT23, SW24, TSB10, Wei18, YW19, Yua20, dSFDG20]. annular [Phi91, YTZZ18]. anomalous [KDH20, ZJLA22]. Ansatz [GT15]. antennas [DDZK05]. anti [BMMZ06, BS20a, Has08, HJ17]. anti-diffusion [BS20a]. anti-diffusive [BMMZ06]. anti-Gauss [Has08]. anti-periodic [HJ17]. antilinear [WS21]. antitriangular [LMV17]. any [CYM09, ZZZ23]. AOR [JL23b, WWS07]. AOR-type [JL23b]. aperiodic [GPíP03]. APLA [DS02]. APNUM [LST07]. APP [Noo95]. appearance [SH91]. appearing [KM95]. Appell [AMT17]. Appl [AS21, Bic21, BtTBV87, Kni95, TLP18a, Tsy96, YP18a]. applicability [SBS24]. Application [BFH09, CKP15, CJ90, DSA20, Ewi91, FKA^{+}13, FMP04, FP24, HvdHV10, KCJP01, KRBK16, MS19, Rab94, Sid90, SDK15, AK00, AVMVMV09, ABKG21, AM16a, BKAG22, BG06, DLS22, Cao10, CAD03, CZS04, DF11, Die15, DG22, DSSC13, DC18b, FMS24, FMS18, FHM^{+}02, Gan96, GS89, Gug05, HS22, HDS20, HWY20, HT20, HK09, Jor11, Kam25, KM25, Kim12, LZZ22, LAZ20, LIPT18, LWCT07, MB08, MG97, MII13, MD21, Mur98, NLZB23, Odi19, PKP19, PYD21, PG02, QM10, RSK24, RDH^{+}12, RKR20, RW87, SST12, Sam94, SD25, SS16, Sod91, SC22, Sza94, VG04, Vas17, WS21, WSP04, XWC25, YY24, YLW20b, ZWK15, BS14a, BY22, BGG12, Guo00, HBB25, LP01, Mur99b]. Applications [AB12a, BJS12, BW06, CHM09, DGCW17, DGRS09, FF06, FFMZ13, GD09, Gro94, Jbi25, KP07, LP00, MA09, VB99, ZBNC25, AKM^{+}22, ADNR21, AEA23, ASS21, AES13, ABCC18, ABRW18, ACR25, BAA22, BRW21, BDF^{+}25, BV96, By01, BR01, BKP14, BV94, BDFF23, BDRZ04, Buc04, Cao98a, CW20, CDRT19, CSL^{+}24, Che12b, Coh90, DK11, DMP08, DKL24, DKK94, EH07a, FRRZ25, FR14, FBM17, GM93, GS21, EEJB22, yGqWsWC05, GSW09, GLW^{+}25, HGM^{+}21, HHYD20, HS98, HL23, HZCZ23, JM16, JM94, KNN03, KF97, KKLD21, Lai09, LWLW24, LGH11, LW20b, LSWW22, LO96, DLM16, MZJ^{+}25, MAD23, MMP20, Ost02, OTK04, PCA10, ROB17, Riv09, RV05b, Ror06, SSS^{+}23, Sae14, SYW22, TLQ21, TLGC22, WOLY25, Wen05, WWLS08, YXX24, ZYQS21, vdHS01, LD02]. Applied [FDF^{+}25, Rei85, ZCSH11a, AMT13, AMH03, AC96, Bac14, BMGM12, BTMT08, BJ00, BR20, Cas06, CBHM19, DZ12a, DS05, Den15, Fer14, Gol86, HHR12, HO10, HM00, HMW05, IHS13, JM25, KDAK16, KKR25, LKV01, Lei99, LZ17, LZ22, LCM24, LXLY25, MOS02, MI03, Mol95, MM20b, RU21, SR09, SP22, SZE^{+}92, SW85, VA05, VSeYD02, Wan96, WM07, ZZ19b, iW07, iW09]. applying [LWW22]. Approach [TMS87, AKM^{+}22, AMT17, AS20b, AAL^{+}26, AD99, AMV03, ARSW05, BMR^{+}17a, Beh93, BMM03, BMGGG12, BDDV12, BBO03, CGT13, CCBGV08, Cao98b, Car94, CLR11, Cha25, CGCMTR02, CXZ15, CD18, CCJ99, CDP^{+}25, DPPR16, DD21, DGGM25, DN21, DS21d, DTGN23, Dou91, DS15, EB12, FJ17, FV85, For11, FCW21, GQ89, GHH20, GRGJ02, GKL07, GL17, GD23b, Har98, HP18, HD88, HP14, IHS13, JMU^{+}25, Jes93, JCJP21, KHM^{+}14, KK11, KHA12, KDH20, KS22, KL09, KKR25, KZ21, KR15, Kür23, KDKW20, LO23, LG21, LBLT13, LS25, LMSW17, LKJ20, LN08, LP00, Lte24, Maj20, MKDC25, MP97, Mit97, MMM19, Mou03, Nes16, NV23, NT16, OT22, OAHN22, PB21, RV09, RMS17, RGA19, RGK21, SKS^{+}25a, SC11, SK97, SH09, SW95d, SWR11, SSS21, TS23, VBH96, WCW14]. approach [WJF25, WPL16, Yan22, Yüz22, Zan01, ZPT92, ZW19b]. Approaches [CHR03, Min87, AAD^{+}08, Buc99, CC23a, CW25, CKS05, Dat99b, DMH18, JUAZ22, LWD^{+}09, LW21b, LL20b, SMTHE22b, SMTHE22a, Sv95, WL10]. Approaching [DL13]. approximant [MV20]. approximants [Ari87, BW15, Cat10, CPD^{+}05, CV88, Dar90, DMGVO05, Gil10, GVP93, IN89, KS91, Kid90a, Mat91, MdR05, OGV92b, OGV92a, Pre90, Pré95, Sad97, Zul25, van86a, vI87]. Approximate [AM09, BLS^{+}17, Huc99, MK21, SdSC99, AKM^{+}22, AD18a, BF92a, BT99, BSV21, BSvdV99, BWY17, BVB10, BVRB14, BG02b, CG05, CKM10, CD18, DS97b, Dun18, ELvdHS98, Enr06, EGH01, FV99, GV02, HG98, JMPY10, KCJP01, Mag91, MKN23, MSS21, Not92, SDG20, SY05, SHG86, TZ00, Tar98, WZ02, WMC09, Zha00, Zha14, ZPT92, GPHA16]. approximated [WC02]. Approximating [DJL04, DS01, GS17, MMRV20, Wan17a, BRTB19, DSM22, GKS20, Ito17, JJL^{+}24, Kos02, LR01, QNA23, Rab94, RMH20, Sim94a]. Approximation [ARGA00, AK00, AR24, AB14, BM89, BS24, DO17a, EL94, Fik23, HP97, IKR^{+}22, KL98, KCS07, KS01, KP01, Lyo12, OT24, Wen10a, ACKV24, AQJ18, AQ20, ASGGRGW25, ABJ12, ABK12, AY22, AEG12, An16, ASZ18, AC23, BC12, BLD17, BDGP96, BDP99, BD07, BBJ25, BM05, BFQ22, BKR13, Bel91, By01, BGGG13, BM04c, Bho11, Bho12, Bla00, BG14, BDV17, BO11, BP85, BMV06, Bre02b, Bre10, BE99, CY23, CFS13, CF13a, CCdlH20, Cha25, CGW20, CLX21, CWP21, CYWH22, Che12b, Cho13, Cod08, CGK25, Con20, Cop03, CF05, CD05, CCM17, CLS04, Coy12, CJ22, CMP23, DZ12a, DDP12, DO17b, DLN^{+}24, DR09b, DESZ25, DJR25, DDNZ18, DMGVPO09, DB95, DCJ20, DL06, DAMA23, DR93, EAS12, EGH01, EHV24, FF20, FT06, FJ95, FJS99, Fun94, GS24]. approximation [GS15a, GP00, GKKM21, GL17, GN86, yGqWsWC05, Gwi09, HGM^{+}21, HB02, HPS12, HHT97, HSS04, HR14, HQAZ24, Hin97, HW93, HJX^{+}19, Jad94, JWZ21, JM05, Jun06, JT06b, KP18, Kał22, KBS11, KK11, KO92, KESYB23, Kha21, Kid90b, Kie17, Kim14, KPR12a, KR25, KX03, Lee23, LW04, LHWF08, Li16, LSY17, LWW20, LS12, LC99, LCH20, LSW23, LMS08, Lor10, Lu98a, MRF00, MN20, MS02, MN24, MS13, Mit22, MD21, MG22, MHL18, Mot17, Mur19, NY13, NWL^{+}22, Nke07, NS16, ORT24, OT25, Odi19, PL25, PR09, Pas91, PH17, PM14, PWX24, Ril92, Rog19, RMS17, RX08, Ror06, RA09, Sad96, SRK22, SSZ16, SYL^{+}20, SYW22, SA20, Sol15, SD24b, SA19, TKN11, TY03, Vab21, Vas17, VGL24, WTB24, WQ17, WSY18, WW19, WZZ21, WWM22, WWC^{+}25, WYYL19, XY24]. approximation [XL09b, YM24, Yam23, YY13, YR09, ZRA23, ZZ19a, ZD25a, ZW24, ZEW20, Zho17, ZB19b, Zou10, ZCC11]. Approximations [Ghe97, Sub04, AL95, AOW94, AMCM09, AD23, AD01b, BKM13, BM00, Bok03, BBD24, Bre06, Bre96, Buc06, BGS02, BC00a, But09, CCDJ20, CC90, CHSS01, CJX11, Che12a, CPY20, CL18, Coh90, CSM07, DRVA20, DDZK05, Dek17, FJP17, Fun90, GT15, GG22, GT00, GS20, GJ00, Gol00, GO21, GOLJ25, GGG16, GND19, HCY18, HN22, JM25, JV09, KTD20, KPY15, KZ13, KW93, KK22a, LS25, Len00, Li11, Lia22, LX09, Lyn99, MKBS26, MM22, MS99a, MT06, MV17, MP15, MD23b, NB01, Nov08, PGS10, PS19, QAE^{+}09, QZM25, RF16, Ren14, SA12b, SS19, Ske99, SN04, Tol04, TC19, VR01, WAV12, WWLS08, Xu13, YLL09, ZM19, ZP24, ZCZ15, ZL18b, ZLWF21, ZXW17, Zup03, dSFDG20]. April [TLP18a, Ano21p, Ano22q, Ano23t, Ano24u, Ano25s]. aqueous [BMP05]. aquifer [MNR14]. aquifers [AMRR18]. Arbitrarily [HGZW21, JWG20, CL24b, XXYZ24]. Arbitrary [ABI22, JQSC22, ÖT20, PGM86, Bal00, BS00a, CLS04, CG14, DM09a, FBM17, GNUV24, GRGJ02, LZZ25, RVD00, Tol03, WJF25]. arbitrary-order [Tol03]. arbitrary-stride [WJF25]. arc [YZH19b]. arc-search [YZH19b]. architectures [BC99, JP93, WWS^{+}93]. arcs [DSV13, HLL09]. area [AMV03, LCZ23, Sal93]. areas [CCM17]. argument [Bic16, DGS24, ZLJ20]. arguments [BF20, Hu99, LZ17, XZ19]. arise [DL01, ZNK02]. arising [Aff94, AA22, ABFV09, AC08, BK06, BVT14, BR94, DSM22, COPPS25, CG92, CKM15, Elm02, FBS09, Guo96, HAN23, Hua98, Iga85, KO96, KAS17, Le 12, LMA18, LGZH24, NA21, Pea16, Pło22, Por17, SS99, Sch87, WPT19, YLW20a, ZR21]. arithmetic [AA04, HFL13, LMO24, LSK12, MA04, PH91, Sch87, Sch89, Ske99]. arity [ZZZ23]. ARK [BR05]. ARKN [FW07, Fra06, Li19, SWL20, YW08]. ARMS [SST04]. Arnoldi [AR23, AB14, BNKR20, GNNR19, HJR22, Hey10, Jia02, RS08b, SW95b]. Arnoldi-type [AB14]. arrows [AB12a]. ART [San03]. arteries [TDC13]. Artificial [AEK23, MP85, AK21, DK20, EK96, LW04, LWYG22, Nor97, QL15, RGÖS18, Rya00, SGN06, WL18, WS04, XL23, XL09b]. artificially [LT23]. artificially-damped [LT23]. ascent [SYW22]. Asian [BNV06, ZO14]. aspect [ML91, NR97, Pic05]. Aspects [Gar87, AR23, BV94, BLL24, Dal00, Fun90, GEGG^{+}20, GKT10, KPR06, LA12, MdR05, RLMG24, vdSvdH95]. asphalt [ZBNC25]. assembling [NRR06]. assembly [BMR^{+}17a]. assessing [LMPS19]. assessment [BDSG09, FP24, LS07b]. asset [Bi20, CF08, ZJH^{+}23, ZD25b]. asset-price-dependent [ZJH^{+}23]. assigned [CMP15]. assignment [LB21, OP04]. assimilation [DC21, RW87, ZMF25]. assisted [GIS23]. Associated [RdAP96, BHSW20, BCC16, Bic21, CDG19, DB95, DR01, ELR^{+}15, Gas92, Hua19, JK21, Jou05, Mar94, MYSC17, NSCC19, Wu03, dACR10]. Association [MH89]. assumption [PRS20]. assumptions [DK21]. astronomical [CG03]. astronomy [BCF^{+}13]. asymmetric [GNZ21]. asymptomatic [ABdSG23]. Asymptotic [AV91, BGT97, Bor10, CFCH09, Han93, IKM23, Leo10b, LC99, LX09, LTT19, LLZ^{+}22, PT09, Wan01, ZP97, ZP98, AL24, AGK24, By01, BGG^{+}20, Bre88, BS12, DC09, Fel06, HH98, Has13, KM17, KS07, LZL14, LS24a, Sch04, SS21, Wal90, Wan17a, WW24, Yi12]. Asymptotic-numerical [IKM23, LTT19]. Asymptotically [AB10a, AW14, Bac21b, BSV09, Kim21, VA21, XYHM20, ZWJ18]. asymptotics [CMR12, Sus10]. Asynchronous [Per99, AAI^{+}93, TDPU17]. Atangana [Hey20b]. Atmosphere [DRC85]. Atmospheric [Beh97, AB97, Bou02, BN03, BBO03, KW98, SKAW12, VS95, VBH96]. atomic [CCP04, Lei99, LW04, PCA10, XL09b]. atoms [MT05]. attraction [BS24, CN16]. attractors [CGPT19, QR03]. Augmented [CSM07, ABKG21, AL05, BBG14, EAV16, Hua21, KKR25, Liu24, SW85]. August [Ano21o, Ano22t, Ano23r, Ano24o, Ano25o]. Author [Ano01a, Ano01b, Ano01c, Ano01d, Ano02a, Ano02b, Ano02c, Ano02d, Ano03a, Ano03b, Ano03c, Ano03d, Ano04c, Ano04a, Ano04b, Ano05a, Ano05b]. auto [LM22a]. auto-convolution [LM22a]. autocorrelation [AGKS25b]. Automata [CFCH09]. automated [SPS20, SW94]. Automatic [BSW93, RH92, Bür13, CL24b, DS02, MCE^{+}09, PRST02]. autonomous [CL01a, DM11b, LMTW20, MT20, SB14]. autoresonance [CSSZ25]. auxiliary [HMD21]. avalanches [DEGDdL25]. avascular [LBLT13]. average [CMP06, Has08, YX25]. averaged [DRS19, DDRS24, RS21]. averages [AB12b, DJL04]. Averaging [DD97, BP12b, CCMSS11, CSSZ20, CSSZ25, CH90, Dal13, EK95, FVGS13]. AVF [HL21]. Avoiding [CGPT19, Hin95, PJB04]. awareness [SAMSB20a]. AWENO [FGGL22, WLG22, WDL23]. axially [DII15, LDIW16]. axis [AX19, GHKM09, GS09, IMM04, LMO24, MI03]. axis-symmetric [AX19]. Axisymmetric [Mac86, AC08, FJS25, Nke07, Sel14]. axons [LFL14]. AZTEC [SW86].

B [Bas21, CGMS21, DK11, GV18, GS25a, Joh05, Li05, MST09, RK08, RTA19, Rou20a, SRK22, SYG^{+}05]. B-series [DK11]. B-spline [Bas21, Joh05, MST09, RK08, RTA19, Rou20a, SRK22]. B-splines [CGMS21, GV18, GS25a, SYG^{+}05]. B-theory [Li05]. back [BF99]. Bäcklund [Car19]. Backward [AGP97, BS21, Ber86, CGH23, CWC24, CHK99, DMA22, Fre98, HM01, HJ17, HJ21, JL23a, LW18b, LMW23, LQY25, MFAD23, MD20c, OXW25, RO16, Saz22, SWW16, SM20, Ske89a, Ske89b, SD24b, SND21, TLG20, TOD11, WL09a, YXN21, Zhe19]. backward-facing [CHK99]. bacteria [CST18]. bacterial [SL22]. bad [CEL25, FVB05, Mul99]. Baer [TKN11]. Bakhvalov [NV23, XLHL25, ZL21, ZL22]. Bakhvalov-type [NV23, XLHL25, ZL21, ZL22]. balance [AMT13, AL20, GBBC^{+}23, KR12, LS16, Mon21, QAE^{+}09, SMW21]. balanced [AHT17, AG25, BDMG12, CCCC^{+}25, DEGDdL25, FGGL22, GBBC^{+}23, KTK20, KD13, RBdM25, TK15, Udd20, WL09a, WLG22, WDL23, WC24a, WC25, YTC24, Zha20a, ZGDL17]. balancing [Chr96, DBH^{+}05, FLÖ^{+}97, dDF^{+}94]. ball [GS21, SG04]. Banach [AHAS21, ABM17, CCOVF22, CP03b, DE16, DMA22, GIS23, GH21, PGA93, VA21, Xu23]. band [BG06, LLM19]. banded [Con89]. bank [Wal00a, Wal00b]. bar [Meh08]. Barbamarco [RI02]. baroclinic [Bou02]. barrier [BP14, LRT99, NV23]. barrier-function [NV23]. barriers [FV01, Hin95]. Bartels [SMC08]. barycentric [ABH22, BFK11, Elg17, HA16]. Barzilai [AX20]. Base [dG91]. Base- [dG91]. based [AD19a, AD20c, AHJM19, AJ19, ABH22, AQJ18, AQ20, AGLRS23, AJ24a, AK21, AK09, AEMX17, AAH21, ABKG21, AM95b, AGM95, AD99, AGJM04, Aro96, AAD14, AD18b, AD04, AFF^{+}15, AEN22, BKAG22, BMGM12, BY09, BCS17, BES18, BKP09, BBW19, BNKR20, BCGS24, BM04b, Ber86, BLY16, Boh03, BDFF23, BZ96, Bru93, BT97c, Bür12, CGS19, Cai15, CCOVF22, CHZZ06, CGMS21, CDGA17, CDD^{+}17, CDW19, CW22, CKS05, ÇK13, CGK25, CNS00, CDP12, CST97, CSM07, DdSF07, DS21c, De 02, DA18a, DA19, DSAB20, Dol14, DY03, DZMB21, DT89, EFLFP09, FK23, FZA25, Fra06, FV87, FM07, FCW21, GMG02, GIS23, GHKM09, GM18, GHH20, GK09, GG19, GAH25, GD21, Gje07, GLLY25, GPHA16, GGR97, GH21, GD22, Han19, Har98, HSS04, HL08, HDY21, HV22, HNP17, HFL12, HCW16, HT20]. based [HS21b, IB24, JMU^{+}25, Jam95, JM25, JY20, JLH13, KKT16, Kam25, Kam16, KDT17, Kan04, KM19, KB21, KHYY21, Kok08, KM11, KW12, Kru99, KK20b, KGZ24, KAS17, Kür23, LG21, LHH96, LCHR03, LS16, LY10, LWD^{+}09, LSK12, LR20a, LFQH21, LWLW24, LZZ25, lLH25, LHÖ13, LSGK15, LFS21, LSWW22, Lo06, LYOI99, LP97, LS98, LZY09, Lyo12, MH16a, MZJ^{+}25, MB10, MSZ^{+}24, MWC21, MZXX24, MO01, Mat91, MS08b, MKH16, MSS^{+}15, Min04, MD10, MV20, Mit24a, MD20c, MKJ23, MHL18, MM25, Moo04, MJS23, MVBS25, Nes16, NWL^{+}22, Nov03, OR18, PSP05, Par04, PGS10, PB21, PP24, Pat98, Pat00, PK21, PC00, PT15, PB10, PRS23, Pow94, Pro24, QDA^{+}26, QM03, ROL19, RSY12, Ric08, Ril92, RMK09, SMTHE22b, SST12, Sae14, SAA20, Sch08a, SW09a, SG96, SH10, SWW17]. based [SK22, SXP09, SD25, SP22, SNW22, SR24, SL15, SSKS21, SLZ10, SL17, SH21b, SW86, TWMP20, TYKK01a, TQA^{+}25, TLG20, TH23, TC19, ÜSHT03, UHUL21, VV02, VBD93, VRC21, Wal00b, WL10, WJM22, WWC^{+}25, WJF25, WG18, XY25, XG22, XGQ20, XZT21, YCY12, Yi12, YÇ16, YR22, Yu99, YP18a, YP18b, YH07, YRV21b, ZDM18, ZBD24, Zen21, ZHJ14, ZO14, ZR21, ZLL22, ZZJ21, ZS21a, ZH15, ZZLL21, Zla85b, dAF17, dFN00, dSFDG20, BMSZ21]. bases [AA94, AFS11, DDHS97, GPP04, PR12, PRS23, Win04]. Bashforth [ZJH18]. basic [AR93, HSW99]. basins [BS24, CN16]. basis [Alb96, AD18a, BAD13, BG11c, CPZ17, Gar96, GS92, HA21, Jun07, JD09, KKT16, Kwe00, LH11, MH14, MKJ23, RÁM23, Roz05, SRK22, Sar05, SB18, SJ11, uIVS13, SBS24, Ste97, TLSS09, War92, XB14, ZM19, ZT06, ZHL03]. basket [DLM16]. batch [Die15, QAE^{+}09]. Bates [itHT18]. BBKS [ÁKM20]. BCH [WC02]. BCH-formula [WC02]. BDDC [PSWZ21]. BDF [AS04, CC90, Hin95, KWY24, LMY18, LS20, PH17, QXQ22, Sha87, SND21, Wan20, WWM22, YJ21, YQCZ22, ZZF25, ZH24]. BDF-type [PH17]. BDF2 [DK20, HS96, LT12, LWW23, LGC25a, LSG24, MSZ^{+}24, Shi25, WWWW25, XL23, ZCQ25, ZZjQ24]. BDF2-ADI [ZZjQ24]. BDF2-FEM [LSG24]. BDF3 [GZHQ23]. BDM [LN21, XYZ24, ZZ24]. BDM-like [LN21]. be [Mau08, Mul99]. Beam [CRR03, DRC85, TMS87, AC18, BH20, CZY18, CF13c, EKT19, HB20, KWY24, LT07, Ma03]. beams [AMH03, CCS02, GLQ26, LR01]. Beavers [BSZ22]. Becker [DS01]. beds [VCN20, VNC21]. Behavior [CY98, Sch02, BGG^{+}20, Cao97, KDS22, LPR00a, MPV24, Sim94a, Wal95, Wan01, WXY24]. behavioral [COPPS25]. behaviors [By01]. behaviour [Bre88, BIMV19, Bur91, CCP04, DLM02, LO96, MRS03, Sch04, SK01]. Behavioural [SMC08]. behaviours [Leo10b]. Bell [Rah25, Yüz22]. Bellman [BGS06, FLS94, For11, RF16]. Bel’tyukov [SV00]. bem [MOS12, AD08, ADSS17, AFLG^{+}12, AFLP12, AFF^{+}15, BF15, CML05, CS09, DS07c, DG22, EH05, GMM09, GÖS20, GMS12, Gon06, HY01, JS09, KMS10, MS08b, NS12, PRS20, RV09, Sel14, ST14b]. Benchmark [GGO12, GGO16, AS00, NER95]. Benchmarking [BT19]. bending [HLC01, LR01]. Bendixson [BW21]. Benjamin [ZJ25, AJK20, AEN22, SJ20, WLY24, ZRA23]. Berenger [BP97]. Bernoulli [AC10, FZA25, GLQ26, MDASAO21, Nap16, ROB17, SSC23]. Bernstein [GS15b, BRIP08, BWY17, BWS21, DSM11, HS19a, JBLC11, Win04]. Bertalanffy [RA17]. Bessel [Che12a, Har00, KM17, KW21, SS10, XGM08, XFG19]. best [JM05, PA91]. Beta [YC00]. Beta-spline [YC00]. between [BY22, Fdi97a, FLL11, Jia12, Kid90b, LLD18, MS08a, Mar99b, PSP05, QR03, Win04]. beyond [BWM21, SW95d]. Bézier [BRIP08, GS15b, JBLC11, ZSJ04]. BFGS [ABKG21, BKAG22, CD18, CSM07, WOLY25]. BGK [Ale11, JN07]. bi [DZZZ24, DDK19, MMDS21, JP17, Ria22, SW21, SSvG10]. Bi-CGSTAB [SSvG10]. Bi-Conjugate [Ria22]. bi-directional [DDK19]. bi-fractional [MMDS21]. Bi-frames [JP17]. bi-Laplace [DZZZ24]. Bi-wave [SW21]. biased [HL97]. BiCG [Sv95]. BICGSTAB [MC00, Fuj02]. Bicompact [AR15, BR20, Bra22, CR19, Rog19]. bicontinuous [RS22]. bidiagonalization [JY20, KBG04]. bidomain [BK09, PS09]. BIE [ZS18]. Biennial [Ano02g]. Bifurcation [KF97, ZG92a, ZG92b, HH10b, KBK21, Mar93, RSL89, YT03]. Bifurcations [RAS99, Eir99, RV05b]. big [ECHF^{+}20]. Biharmonic [AS21, ADFP25, Chr01, CH25, HAC22, KM19, LHH08, LZZ25, MB10, NMB10, YZ22, ZBY19, ZM25, HS21a]. Bilateral [CC23b, MKBS26]. bilinear [GGR97, Yan22]. bilinear-constant [Yan22]. bin [AM16b]. binary [CPY20, DLQZ23, HPW21, KBK21, SMW21]. Bingham [HS22, Mur15, TT03]. binning [DS07b]. binomial [DL01, MZN21]. bio [SK96]. bio-chemical [SK96]. biochemical [BRBM08, BBKS07, MCD20]. biodegradation [BM05]. biofilm [CL02b, HJZ23]. biofilms [CCK03]. biological [PGC01]. biology [DSM22, Dal00]. bioluminescence [CdCV03]. biometric [Sae14]. Biorthogonal [Bre02b, GTS20, MB08]. biorthogonality [BZ96]. Biot [BLY17, GLV03, GLV06, PQ25, WSHC20, Wen25]. biotechnology [AMV03]. biperiodic [Rat13]. biplane [QM03]. bipolar [Mol95]. biquadratic [HLZ06]. Birkhoff [NBNTGV11, CN17, FMS24, MS24, Mil17]. Birth [BMR17b, BLM17b]. Birthday [CHM09, EST15]. bisection [PSP05, Riv09, RRMJ12]. bistable [BS24, EV96]. Bits [WWS^{+}93]. bivariate [AR24, BES18, BWS21, CDRT19, JBLC11, KP03a, MST07, OGV92b]. Björck [MP98]. Black [AAM03, ALZ^{+}21, Bis11, Rou20b, RG21, iV09]. blades [ARSW05]. Blair [LVW21]. Blended [BM02, CL85]. blending [DSV13, LY01, MMP02a]. Bloch [DA18a, HLR18]. Block [Cao10, CV95, DB95, JM16, LZ17, LZ22, LN92, MM02c, Sim04, ZC10, AB14, BN99, BL21, BHB23, BLY17, BRRS15, BSV21, BDFF23, BT98, BM02, BM06b, BT00, Cao07, CL06, CZ12, DGD03, DS10, EJS04, FW22, GY94, GAZD25, Guo96, Guo01, GS08, HJR22, JCY25, LVfP14, LR18a, LR18b, LHY24, LAZ20, lLH25, Mag91, Mor05, Nak12, Noo95, Not99, RS08b, RGL16, RCR24, Sch89, SGY22, SXL22, SS13a, SCvdH92, TYJ11, WPS18, XXF22, YK04a, YG99, Yua93, Zen21, ZC91, Zha00, ZLJ20, Zha21a, ZYC22, ZZW97, Zha01, ZZ19b, de 95a, vC93, vdHS01]. block-by-block [ZYC22]. block-centered [LR18a, LR18b, LHY24, SXL22, XXF22]. Block-Cholesky [LN92]. block-diagonal [BL21, BLY17]. block-GMRES [Mor05]. block-iterative [ZC91]. block-structured [DGD03]. block-triangularly [vdHS01]. blocking [AFS00]. blocks [BMR17b, MCBV20]. Blood [PVH25, AI19, PCRR17, WPAZ24]. Blow [ALMM96, ALMM98, ALMM01, BDKM92, Cho13, LYA^{+}19, PVM22, SG92]. Blow-up [ALMM96, ALMM98, ALMM01, BDKM92, Cho13, PVM22, SG92]. blur [MRS10]. blur- [MRS10]. Board [Ano18g, Ano18h, Ano87, Ano91, Ano93, Ano03e, Ano03f, Ano04d, Ano04e, Ano04f, Ano04g, Ano04h, Ano04i, Ano04j, Ano04k, Ano05d, Ano05e, Ano11a, Ano11b, Ano11c, Ano11d, Ano12a, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano12g, Ano12h, Ano12i, Ano12j, Ano12k, Ano12l, Ano13a, Ano13b, Ano13c, Ano13d, Ano13e, Ano13f, Ano13g, Ano13h, Ano13i, Ano13j, Ano13k, Ano14a, Ano14b, Ano14c, Ano14d, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano15k, Ano15l, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano16k, Ano16l, Ano17k, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17j, Ano18a, Ano18b, Ano18c, Ano18d, Ano18e]. Board [Ano18f, Ano18i, Ano18j, Ano18k, Ano19a, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l, Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano22a, Ano22b, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano23a, Ano23b, Ano23c, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano24l, Ano24a, Ano24b, Ano24c, Ano24d, Ano24e, Ano24f, Ano24g, Ano24h, Ano24i, Ano24j, Ano24k, Ano25a, Ano25b, Ano25c, Ano25d, Ano25e, Ano25f, Ano25g, Ano25h, Ano25i, Ano25j, Ano25k, Ano25l]. Board [Ano25m, Ano26a]. boards [BLW02]. Body [PGM86, Arn95, CFKS07, HOS99, Lei99, Lei02, Mur98, PRST02, SK01]. Boltzmann [LLL08, MDA24, VVR08, ZCGS21]. Bona [AJK20, AEN22, SJ20, WLY24, ZRA23, ZJ25, ADM10]. bonded [Kok08]. bonds [ALY03]. BoomerAMG [HY02]. Booster [PSB91]. Borwein [AX20]. Bose [ZB07]. boson [MPMD21, ZH21]. both [BB98, CFLW22, CL07, Kop89, ZZHS18]. Boubaker [RR21]. Bound [TD09, BMM03, GHS25, GJlL23, HZZ25, LLVX20, PL25, Yam18, YX25, Yao24, Yu08, ZYH23, dC18a]. bound-constrained [BMM03]. bound-constraints [ZYH23]. bound-preserving [GHS25]. Boundaries [Gus88, ADFR18, BHJ05, BDSG09, LR03, MMKN17, Nor97, Nor99, Rya00]. Boundary [AQS94, AM95b, BG11a, CHM09, GPHA25, Hsi06, ILS19, LT93, PT09, Pec09, SM85, TMS87, TLSS09, AS11, AL09, AyLqW18, ADG^{+}24, AMH24, AMP03, AMCR17, AZ23, AMR14, AGM95, AD18b, AFLG^{+}12, ABRW18, Bac18, Bac19, BTBR19, BY22, BC02, BPN24, BP14, BBD18, BDSG09, BM00, BCL15, BK21b, BCC16, Bic21, BGH08, BBRS97, BBD08, Bog00, BP97, Bor02, BP85, BMV06, BT93a, BT93b, BT95, BT98, Cai09, CA21, CY23, CDV00, CL08, CGA93, Cas96, CW98, CJ18, CR23a, CH01, CZ12, CHH15, CQZ20, CS09, Chr01, CRR03, CN15, CN17, DY17, DRVA20, DB97, DS24, Dea11, DA16, DMPSC16, DYZ20, DP85, DY03, Ehr08, EGV25, EH88, EZ98, EH06, EFLFP09, EH91, FS15, FMS18, FJS25, Fat12, FE93, FL15, FL20, FMPP24, Fer93, FJ09]. boundary [Fun94, GZZ20, GP23, GAW09, GGM95, GM95, Ghe97, GP98, GO19, GO23, GH02, Gwi09, HGM^{+}21, HG98, HH98, HN03, HJ09, HD04, HM15, HDY21, HO24b, HK93, Heu00, HMY19, HL03, HDS20, HLW25, HX11, HT00, HLL09, HS19b, HBJ09, HLIS16, IO18, Ito22, JP08b, Jeo09, JL23a, JCJP21, Jun06, KDAK13, KDT17, KHLV22, Kat89, KK20a, Kel85, KOR18, KG90, KS09a, KvyS15, KS01, KW93, KK02, KK20c, Kur98, LHH96, LRS09, Lam13, LM00, LP05, LY01, LW04, Li11, LA11, LZ17, LZ20, LXZ21, LZJ21, LH21, LWYG22, LWW22, LHC23, Lie01, LC99, LO03, LMSW17, LKJ20, LM22a, LZIZ23, LL02, LOM98, LXLY25, Ma03, MOS02, Mai06, MS03, MKBS26, MS90, MG18, Nes16, NY13, O’L87, OS08, Odi24, OGS20, PXHZ20, PZMX16, PG21, PPS10, Pap95, PNA21, PTV20]. boundary [PLB22, PT95, Pul12, QMLC15, QAMX17, Quy19, RNG22, RZS21, RK08, RVdCVR02, Roo20, RTU15, RGA19, RTA19, Rou20a, Rya00, SN22, Sch96, Sch93, Sch16b, SNOK21, Set24, SSC23, SWFK13, SC08, SSR23, SG00, Sod91, SDK15, Sof17, SS02, SHG86, SW85, ST08, Sub04, Tau09, TY98, UWY22, VO00b, VV02, Wag85, Wan07a, Wan07b, jWqW09, Wan11, jW15, WL16, Wan17b, WL18, jWS20, WDU21, War92, WW14, WT17, WS04, XF22, XL23, XL09b, XZH19, YM24, YGY15, YL13, YZ24b, Ye04, YL26, Yos00, hYqW12, ZTZ15, Zak19, ZLY23, ZG92b, Zha96, ZC10, ZLJ20, ZR21, ZW19a, ZPT92, ZLS20, Zho17, Zho18, ZZ19b, ZWL11, ZS18, ZR15, vR04, AGM09]. boundary-corrected [HLW25]. Boundary-Fitted [TMS87]. boundary-initial [GO19]. boundary-value [Bac18, Bac19, BTBR19, BPN24, Fun94, LC99]. boundary/initial [GO23]. bounded [Ber05, BL91, CLTA18, Che12b, CD20b, Dun18, Hin97, MDRR11, MS03, Wal00a]. boundedness [CHM22, FS05, GGT24, GLMY17, MDP10, MV17, Spi13, XCHW22]. boundedness-preserving [MDP10].

Bounding [BRRS15, Coo89, Gil10]. Bounds [MSP10, BY22, Bou16, BH20, Che16, CF13b, CFM^{+}24, Dra97, Fac03, FW22, FXY22, GY94, GH02, JR02, KS10, LSK12, LVfP14, LC24, LW92a, Nak12, NN10, Roz05, The17, WWX13, YB10, ZM25, ZL24]. Boussinesq [ADM10, CEL25, DW21, EZ03, HC22, JY23, KD13, LC19, LZW17, LZW19, Lte24, XZL19, ZWH^{+}17, ZWG25]. box [AM04, CLL23, GLW^{+}25, MKBS26]. box-constrained [GLW^{+}25]. brain [DEPS15, LWCT07, MPSS16]. branch [Oji88]. branched [BRTB19, CV88, TBRBM20]. Bratu [AY21, BRTB19, KOR18, SSS21, TBRBM20, WPN^{+}24]. Bratu-type [KOR18, SSS21]. breakage [DKSS24, SMW21]. Breakdown [HKS86, Mac86, Ta’86]. breakdowns [RS08b]. breaking [LY24b]. breast [AFIS24]. breather [CSS19]. breathing [DG96]. breathing-mode [DG96]. Bregman [AHAS21, ABR23, LWLL24]. brick [YB10]. Bricks [DV20]. brightness [CP06]. Brinkman [CD23, CGG25, ÇK13, HJ09, JJL^{+}24, LYZJ23, ZFZ19]. broader [SL08]. Broer [NAF24]. Brouwer [RB12]. Brownian [AACP20, HJ21, JK21, SHL19]. Broyden [SW09a]. Brunner [Bru97]. bubble [Dar00]. buffer [KKE16]. Buildings [BO87]. bulk [CRTU15]. Buoyant [RBBC85]. Burg [CDI91]. Burgers’ [AZHD23, GZHQ23, Hus20, Kha21, Mit24b, SAMSB20b, VRC21, XP23, Yan21a, ZRA23, ZY25, AJK20, AEN22, BDKM92, BHJ13, CLX21, CBHM19, KKR24, KSHB21, KK86, LZY09, PWY21, PR90, Rob01, SSA24, WL10, WLY24, ZJH18, ZZL17]. Burgers-type [BHJ13]. buried [QL16]. Butcher [JKN94, VS94]. BV [LT93]. BV-stability [LT93]. BVM [BMT93]. BVMs [MS02]. BVPs [AAB^{+}22, BP12b, BRW17, BRS^{+}18, GHKM09, HM09, MS02, MST09, PPS10, SAA20, SM93, ZLG15].

C [AR93, Hop23, PZMX16]. C-fraction [AR93]. cable [AMK18, DA16, LR20a, YJZ18, ZZ17, ZXW17]. cables [IJ14]. Cache [HKZ08, LW92b]. CADNA [AV96, Tou97]. Cahn [WaZW21, ASGGRGW25, AZ23, BCM04, BEH24, BMWH20, CGH23, CCZZ18, CW22, CCL04, DLQZ23, GLML20, GZQS23, GHS25, GGT24, GJLL20, GAX24, HLT07, HPH20, HGZW21, HZZ25, JZXJ21, JJL^{+}24, KK09a, KSJY25, LMPS19, LCK22, LXPY25, SC25b, TZ21, WWL21, WWZJ22, YYZ23, ZXYW22, ZCY20, ZYQS21, ZYQS23, ZD25a]. Cahn-type [ZYQS23]. calcium [NRWF08]. Calculating [HC01, Tar98, AY15, DSW96]. Calculation [JZJ10, Bec02, KCW16, MAG13, MV20, SW06, WSC09]. Calculations [Mac86, RBBC85, BNH01, BC93, CSCM96, KL87, LTC03, Mun00, Var92]. calculus [Lub92, MC17, MAD23, MVBS25, Yam23]. calibration [RW87]. Camassa [AS06, CLP15, JWG20, NYPW21, QR24, RLSS06, ZZ18]. can [Var92]. Canada [CFTW08]. cancer [AFIS24]. Canonical [BLM17b, LYLL23, CMT25, WL21]. cantilever [LT07]. Capacitance [DP85, GOGF03, O’L87]. Capillary [VK17, NAF24]. capital [DZMB21, OHA25]. capturing [Aca12, KD13]. Caputo [CA21, CDW13, KDH20, KAD24, LW19a, LW20a, LIPT18, MDASAO21, OB20, OCVW22, RMH20, SC20, SA19, ZHS22, ZEW20, ZZ19b]. Caputo-type [LW19a, LW20a, OB20, SA19]. Carathéodory [Sae14]. cardiac [BK09, CS19, CKP15, CK22, FPS15, NK11, YV17]. cardinal [Hey19, HBB25]. cardinality [KKR25]. Carleman [Kli15]. Carlo [FSX25, JL24, LHÖ13, MD06, DE18, Sob24]. Carreau [KHM^{+}19]. carriers [ABdSG23]. Cartesian [ASCM02, ASC03, Bac17c, Bac21b, BP12a, FD16, FdSB02, Fre04, Kim14, LQY25, THW19, TY00, WE99]. cascade [AB15]. Cascadic [MRS10, YXX19]. case [Beg00, BB24b, BRW17, CMR12, CY05, GANT02, Jes93, MN03, Med96, SvV22, ZL21]. cases [PCA10, YEZ25]. catenary [AS00]. Cauchy [AK95, BBBN21, CJ18, Cum95, FFQ09, GLLY25, JP08a, JL24, Kli15, MM18, MP98, Mat91, ORT24, Rab94, Shi20, Vab22]. Cauchy-type [Mat91]. Caustics [SR88b]. cavitating [XLK07]. cavitation [GNUV24]. cavities [DdCVR03, QC12, QL15, dB03]. cavity [Du11, Guo00, JP08a, RSK24, RBC02, WWLS08]. Cayley [LP00, MO01]. CBEM [JS09]. CDG [YZ22]. CE [QM10]. CE/SE [QM10]. cell [Aya09, Bar09, Ber15, BW95, BG03, Car94, Dal00, EK95, FS88b, ID19, JP08b, Jeo09, MSS^{+}15, MM20b, OH20, Ush18, YR22, ZL18a, PVH25]. cell-adaptive [YR22]. cell-averaging [EK95]. cell-centered [Bar09, BW95, BG03, FS88b, OH20]. cells [Ber15, LARGVR23]. cellular [CKP15]. CEM [AG98, DK14]. centennial [FJ97]. center [CPD^{+}05, ZB07]. Centered [Tol04, Bar09, BW95, BG03, EN09, FS88b, LR18a, LR18b, LHY24, OH20, SXL22, XXF22, ZY19]. centers [GH20]. Central [AGKS25a, LPV24, ASCM02, ASC03, BL06, CXNF14, CR19, CKM25, DL16, FCX06, JTB15, KTK20, LPR00b, LGSW25, Li25, NRR06, Pap95, PGDB08, QW04, Tou10, TJ12, TK15, YTC24, YR22, iV09, BM12b]. central-upwind [CXNF14, CKM25, FCX06]. centrality [EHNR24]. centred [SGN08]. centrifugation [BBCS05]. centrifuges [BCS06]. centrosymmetric [SYW18]. Certain [DL01, BKP15, DMGVPO09, Dra97, Fuh01, GVSL96, GS05, LS86, LFP04]. certification [JS25]. certified [Che16, RÁM23]. CFD [BTMT08, BS97a, QM03, SST04]. CFIEs [CM14]. CFL [GK19, MZ04]. CG [Sch99, dv95a]. CGLS [Baz03]. CGPM [LSWW22]. CGPM-based [LSWW22]. CGSTAB [SSvG10]. chain [Bec02]. chained [FRRZ25]. chains [BT02, Buc99, DNW18, Mar03]. challenges [DBH^{+}05]. change [CH15, Sal93, Yao24]. changed [AACP20, LMTW20]. changes [Wag98]. changing [DCC14, SY08, VVV24]. channel [BJ00, GKKM21, Gla93, Jun06, Li25, PK91, RKVZ15, jWyG08]. chaos [CJL13, Pul09, Pul12, SSL93, SZ12, SZ17, Shy91a, Shy91b, SH91, ZWK15]. chaotic [SK01]. character [Mar99b]. characterisation [AL24]. Characteristic [BF20, ZSG^{+}20, Bre06, CHLA21, EAS12, FLR08, HO24b, JZJ10, LY08, LHY24, SAG86, SWW11, ZMC13]. characteristics [DMM24b, JQYM23, LCVG01, Mar03, OEAS21, PG02, YY13, Zió99]. characteristics-mixed [YY13]. characterization [ST89]. Characterizations [LX08]. Charge [GG95, AZA22, NER95, NR97, SG05]. Charge-oriented [GG95]. charge-preserving [AZA22]. charged [LW22]. charged-particle [LW22]. charges [HNP17]. Cheap [SZ97, GHW01, WGB99]. Chebychev [TGB08]. Chebyshev [ARGA00, AAEMY21, BC08a, Boy06, Boy07, BO11, Coh90, DDS89, DRS19, DDRS24, DSW96, EK95, FFY08, GW20, HM00, Hey19, HBB25, Ibr25, JOL23, KK09b, KP19, KYI17, LM21, MQO17, Mac92, Maj17b, MM07, MRFF17, OT21, OGS20, PL20, PH17, SWW17, SWX00, She00, TC19, Xu16, Yan18, ZB19a, ZK00, vS97]. Chebyshev-collocation [MQO17, SWX00]. Cheeger [ILS19]. Chelyshkov [ROL19]. Chemical [ND85, BBV05, CKB13, DS02, DS03, KCJP01, PM05, SK96, ZdBT03, dB03]. chemicals [HH22]. chemistry [ACR25, BGG12, VS95, VBH96, WK02]. chemo [GGRBRG22]. chemo-repulsion [GGRBRG22]. chemoattraction [LARGVR23]. chemotaxis [BLRGVR23, BGG^{+}20, CST18, HH22, WDL23]. chemotaxis-Navier [BLRGVR23]. chemotaxis-swimming [CST18]. Cherenkov [YD22]. chess [WWS^{+}93]. Chile [BGHR12, BGH^{+}15]. Chilean [BGHR12, BGH^{+}15]. China [LST07]. chip [SG05]. chiral [WG19]. Choice [GC15, BHRY21, BS00b, CDD00, EG88, Lay08, Neu88, PA91]. Choices [San89, GAW09, SWFK13]. Cholesky [HPS12, KK20b, LN92, Noo95]. Cholesky-factorized [KK20b]. Choose [Rob10]. choreographies [RS14]. Christoffel [Bre91]. chromatography [DGM18, DMM24b, vHA98]. Ciarlet [ZBY19]. CIR [BDOG19, TWL23]. circle [BRS16, CCBGV08, DGV00, DBCBPP10, DIR13, SL01a]. circuit [AKGR14, BLW02, GG95, MT11, PGS10, ST05]. circuits [BBS11, GR93]. circulant [lLH25, NR14, Zen21]. circulant-matrix-based [Zen21]. circular [AA20, Duf90, HT00, IR22, LMS08, LBCN00, RGK21, Son91]. Circulation [ZW87, LG02]. CIRM [Wen10a]. Cittert [Dun18]. clamped [BBD18, Lam13, SWY^{+}23]. class [AL95, ADR17, ADFR18, ADM22, Ant25, ABF09, AD18b, Bai96, BHB23, BDSG09, Beg00, BMPR15, BJ20, BIJ23, BLL24, BC23, BJM01, BKP15, BIMV19, Bur91, Bus06, Cam99, CZ19, CLLM21, CN17, CG14, Cum95, Cve02, De 06, DGN12, DYX09, DM09b, DZ12b, DIR13, DZW24, EM05a, EVO04, FM21, FL05, FH22, Fer09, Fer96, FR18, FJ95, FJS99, FS24, GT15, GMG19, GJlL23, Gon06, GFPG18, HS20, Hey20a, HJYL19, JCL18, JQYM23, Kat89, KM21, KLR25, LP24, Li12, LL20a, LM21, Lot19, MH16a, MAH18, Maj14, MKN23, Man96, MMM19, NDM20, NTHC21, Ost93, PV24, PRSS24, PD01, RY13, RGA19, RT95, Sad97, SEGV02, Sid23, SG05, SL08, SSPZ20, Tsy96, Udd20, VDVV98, VRC21, Wan11, WZ14, WZ22, Wan23, WSW96, WdG92, XWW19, XZ19, YMD21, YLY19, YFLX20, YLLZ21, Zha19a]. class [ZZL01, Zha20a, Zha20b, ZYC22, ZZW97, in 92]. classes [KM17, MK21, UHUL21]. Classical [Lee10, AS11, Beg00, BDRZ04, CCJ99, DJM09, FR18, GMG02, GP01, LS12, LMTW20, MZZ17, MOS02, Mar94, Mar99b, MM02a, MMP02b, PA18, Sal03, Sin23, TJK18, WWL21, Wel10a, Wel10b, ZS18]. classification [BPV25, KKR15, Ush18]. cleaning [GDS^{+}15]. Clenshaw [CL14, Maj17a, LYZ25]. climatology [Pło22]. Cloned [FZM20]. closed [TMM15]. closure [JR18]. cloud [RGS25, WK02]. Clough [Bar12]. cluster [DS01]. clustering [GLW^{+}25, LTDY24]. clusters [Nür09]. CM [BS94a, VV95]. CM-2 [BS94a]. CM-5 [VV95]. CM2 [Pet92]. CN [ZYLL20]. CNLF [LT12]. CNTs [EKT19]. co [CZY08, LLY11, TK19]. co-coercive [LLY11]. co-existing [CZY08]. coagulation [DS07b, Sin23]. coalescence [SNW22]. Coarse [SS99, BBD24, FdSB02, NN13]. coarse-mesh [FdSB02]. coarsening [GGLR09, GGR97, LWT07, de 96]. coastal [KDAK16]. coaxial [IJ14]. cochlear [BF09]. Code [Ara99, ZW87, AHJM19, AJ19, AA87, AB98, CST97, GPPR12, Sal89, Set24, SW86, WSP97, WB92a]. codes [BMR^{+}17a, Ber86, BT97c, GMG02, GMG04, Sha85b, Sha87]. coefficient [BJ01, BJ03, BRW17, CDW19, CCL04, DP12, DL01, DMA22, FL23, HP14, HAY20, HLIS16, HN22, IT16, JR02, KS07, Lu98a, LJ20b, PWY21, PJB04, Quy19, RL21, SHB^{+}25, Su94, SW17, TK05, WZZ21, ZEW20, ZLWF21]. Coefficients [Hig93b, AS21, AL09, BF17, BB25, BC05, CJ90, CC04b, Chn17, CK20, CL88, DYX09, DCC14, DKK94, DC09, GHW20, HS21a, HLMKZ06, JV09, KP92, LPS25, LDC24, LS21, LYZ25, LYZW22, LMW23, LSW23, LP01, Maj17b, ML16, Mul19, PM14, Spi13, SW20b, Top21, VL19, VSeYD02, WCXL09, WR20, WYYL19, WG22, YLFT20, YT00, ZM17, Zha20a, Zha20b, ZW24]. coercive [DHS05, LLY11, SW24]. coexisting [PK21]. coherent [CCP04]. Cole [XK25]. Collaborating [MR92]. collapse [TYKK01b]. collapsible [LT00, TYKK01a]. collective [BIMV19]. collisional [DKSS24]. collisions [CLP15]. Collocation [AB88, All24, BR97, BS20b, CMPP24, CN15, HM17, LM22a, SDG20, SWB20, ACDP22, AN25, AV96, AB24, AFS02, AJK20, AAB^{+}22, Bal00, BAD13, BZ17a, BIM15, BIW24, Bru92, BMM97b, BMM97a, Bru07, COPPS25, CA21, CHLX07, CCQ^{+}23, CP05a, CP09, CDP12, CN17, Cum95, DP12, DS07a, DN21, DDF24, DHWL22, DHL00, DLPV17, DBBH14, DLV24, ERS00, EHM01, EEAS25, FK23, FID18a, FWL18, FH20, FH22, GS24, GPHAM12, GT93a, Gu19, Gu20, Gug05, yGpY09, GW20, Han93, HM15, HM00, HSWW24, Hou23, Hu99, HR96, HY24b, HKP89, IO18, JZK06, Joh05, JCJP21, KO08, KSHB21, KLR25, KGZ24, KK23, LHHR94, LHH96, LFB00, LWL18, LZW17, LM21, LZCF21, LYZW22, MQO17, MK20, MAHZ21, MS86, MD19b, MSS21, Mir20, MKJ23, MDP23, MRFF17, MPDB24, NDM20, NLS20, PL20, Pat00]. collocation [PM91, PT11, PTV16, PA91, PK91, QWX20, QM20, ROL19, RE19, RK08, RTW21, RN22, RAOC18, RTA19, Rou20a, Sch93, SSC23, SLW17, SWX00, tSqWyG16, SJ11, Shi20, uIVS13, SK10, TC19, WKM04, WMF17, WW24, WC14, XZZ19, YMD21, Yan18, YT21, Zak20, ZAED21, ZHS22, ZWJ18, ZL18b, ZD20, ZCSH11a, ZCSH11b, ZP24]. collocation-based [Pat00]. collocation-type [Han93]. color [Pro24, ZN21]. colored [Kom07]. column [GY94]. column-block [GY94]. combination [HEG16, LKV01, MM02b, SYY20]. combinations [DJM09, LH02, WM08]. combinatorial [Ren99]. combine [AN15]. Combined [KOGL25, SD22a, APJ09, HJYL19, Lot19, MK14, MD20c, NLS20, PA18, SL22, SPYS24, SA18, Yua20]. Combining [DM12, CYWH22, LYZF25]. Combustion [BH85, BBV05, GS89, KCJP01, NER95, NR97]. coming [QPT23]. Comments [Guo26]. commodity [ZGO12]. common [Alb96, BWY17, BWS21, SC22]. Communication [DJNR22, dv95a]. commutativity [Yam18]. Commutator [MM25, BM06a]. Commutator-based [MM25]. commutator-free [BM06a]. commuting [BB98]. Compact [AA24, GLQ26, SM20, AZA22, BIO24, CFLW22, CGA93, CC19, CC20a, CQZ20, CRU15, CS18, GX11, GLLW14, GZHQ23, HHL23, HC22, HCC24, HJKW17, HL21, HL24, HLJ20, JHGZ20, JZZH22, Jor11, KOGL25, LXZ21, LW21a, LXZS22, LYZF25, LZW19, LYC24, MZXX24, MG22, Pir09, QXQ22, RZ18, RG21, SG16, SXL22, SBS24, SA18, Ven15, Wan17b, WR20, WLM21, WCM21, WZ22, WPT19, XWC25, YZG23, YT00, ZG21, ZQZ23, ZCQ25, ZJ25, ZFX25, ZL24, ZGDL17]. compact-WENO [SG16, ZGDL17]. compactification [DB95, MW93]. compactified [KS22]. compactly [Ehr08]. Companion [BR01, BW06, Win04]. Comparative [BdFPSdSC08, BT99, CG03, DDK19, EGV25, GRLL01, PSP05, SA21]. Comparison [ACM09, BDE22, CKPS15, EVO06, GGRBRG22, LLV18, Moo95b, NN10, PCRR17, Aya09, BD07, Boy91a, CL01a, CS19, EK97, GP01, HT94, KPR06, KW12, LWD^{+}09, LPZ00, LWL18, MD96, NN13, Par21, PZ20, VSeYD02, ZP24, LCM24]. Comparisons [YTZZ18, WSS97]. compartmental [Par21]. Compatible [GP00, ZBD24, ZWJ18]. Compensated [RT20, WG10]. Compensation [KRBK16, LCW20]. competition [EE20]. competitive [BLRGVR23]. complement [Cao09, Cao10, CG89, KMS10, LN24]. complementarity [CZ19, HV22, HZZ24, Ius97, KP03b, LWZ22, LLL12, LMP99, MH16a, THW19, WZ14, WG18, YZH19b, ZR21]. complements [RZ04, GM87]. complete [CGGGS11, Meu14]. Completion [Sei02, Shi20, Liu21]. Complex [CFCH09, AQJ18, AC08, AES13, ÁMS17, DJNR22, CC24, CDW19, CLLM21, Dar90, DS21d, DJR25, DKK94, DGS24, FZA25, Fun94, Gab02, Kar89, KDK17, LAZ20, LZIZ23, Lu98a, MZN21, PH91, QDA^{+}26, SD22b, TC03, TY00, YWW23, ZWFX22]. complex-symmetric [AC08]. Complexity [LRT99, BMSZ21, HZCZ23, MZS10, MMT90]. compliance [CFRA08, FH10]. compliant [DII15, TDC13]. complicated [AEA23, DKK94]. component [GGG16, MM14, OZ96]. component-wise [MM14]. components [BLW02, CG14]. Composite [DP90, Maj14, jWyG08, AGKS25a, AMH03, Dra91, FSB97, He24, HY01, Jor11, MMT90, MR94, NBP94, NU15, PS19, ROL19, SSA24, WDZS21, ZMY21]. composites [RS22]. Composition [DK11, Bla01, BCT19, BCET24, CCC08]. compound [PWY21, RS08a]. comprehensive [ÁKM20]. compressed [ABKG21]. Compressible [BW86, CFX08, CKK10, CGS20, DS97a, DGRS09, DII15, Fai00, FL01a, JKW12, JN07, KC94, KCL00, Kie17, Kwe00, Kwe01, Kwe03, LDIW16, LHY24, Mou03, QR03, RCGM98, RZ15, SZE^{+}92, Sod91, Tan24, Tur86, ZYSZ14, ZSG^{+}20, ZQLK11, vdHVW01]. compression [AEK23, LYC24, PGS10]. compressive [BKAG22, KKLD21]. computable [BKK25]. Computation [ABD^{+}25, BJS12, BPTT15, BGVHN10, CMP15, DV95b, DGS24, ERS00, EV96, HGR01, Jbi25, MH89, MR06, SL01a, Sau00, Sch23, Waa88, YT03, BCU00, BG06, Boh21, BHJ13, BWY17, BWS21, CDV00, CCBGV08, CJ90, CC23b, CD00, CJ22, DNW18, DS02, EJRR23, EH88, FvdMS17, FLÖ^{+}97, GP01, GSR00, EJR25, Har98, Has20, HR14, Hin97, Jéz04, Kop89, LM00, Mar93, MST07, Nes16, Nov03, Ram12, Rob10, RTT01, SW06, Shy86, Sv95, SS10, Vas17, VGL24, Wee01, XFG19, YH00, ZCGS21, vHA98, MH89]. Computational [AFIS24, BLW02, BBPR05, BF09, Fun90, Jam93, LFL14, LY24b, MMT90, MOZ87, Spi99, ADR17, AEA23, CFKS07, Car09a, CHNN20, DCN^{+}19, Elm02, FL05, FKA^{+}13, FMP04, GEGG^{+}20, HG98, HAN23, HD88, Hey20a, JRT90, KD25, Kür23, LMSW17, LY16, MNSS22, MdR05, MZ04, Pul86, Qui96, RVM23, SYG^{+}05, Shy91a, Shy91b, SD93, SPYS24, VS91, VT91, ZW09]. Computationally [YK07, AA87, IR22]. Computations [BDKM92, Li01b, AS97, Asc12, BC02, BB24a, Bou25, Chr96, Dat99a, FL01a, GT15, Gat91, KH91, Kor95, LSGK15, MS08a, MM16, MPG^{+}16, NLLG20, WPS18, ZP12, van95, vdHMdS99]. compute [FvdMS20, MDRR11]. computed [EH08, WT08]. computer [AGKS25b, BRSD91, Fre91, Her91, Ney95, de 95a, IMMS20]. computerized [San03]. computers [SK91, Som93, VS94, dH95, de 92b, dv95a, vvdV97]. Computing [Alt85, ALP^{+}96, Boy06, CKL03, CFTW08, CF18, HM09, HP91, KBG04, LMO24, LST07, LZZ22, LW07, Lu98b, SBBC21, SBS24, Tan87, Yos00, ARY23, AB24, AT13, Bai25b, BC99, CKM10, CB99, CCS02, Din93, FHM^{+}02, Gil91, GAOB20, HCS20, Ibr25, Jia02, JY23, KBK21, KW98, LMV17, LMV25, Li00b, LfX15, MB20, MNSS22, Mar03, Mat86, Moo95a, PK21, ST09b, Tru00, WB90, Yam23, ZC91]. concave [SC11]. Concepción [BGHR12, BGH^{+}15]. concept [CGK25]. concepts [HSW99]. concerning [Han06, JL24]. condensates [ZB07]. Condition [GK93, Kor95, LL20a, LLW22, Win01, Ang06, BM12a, BP97, BG11c, CQZ20, CMS06, DA16, Dob05, GZZ20, GO21, HS19b, KBS11, Lam13, LHH08, LH09, LW18a, NC16, NCYC22, OMP98, RVdCVR02, RBC02, Vul95, jWqW09, jW15, WL16, Yam18, YL13, hYqW12, Zho18, dC18a, vR04]. conditional [BBBK22, GO18, MJS23]. conditioned [IKS25]. Conditioning [Mag91, Not92, BF09, Li12, WM22]. Conditions [Gus88, JL86, JL87, PT09, AS11, AQS94, ADG^{+}24, AMH24, AZ23, AAL^{+}26, AK95, BY22, BKK25, BBD18, BDSG09, BCL15, BMS89, BSZ22, BK21b, Bor02, BMV06, BB98, BS96b, BT97d, CL02a, CH95b, CPP02, CAAT16, Chn17, CN15, CH90, Cve02, DY17, Dea11, DYZ20, DP85, Ehr08, EGV25, EH88, EZ98, FJ09, Gar10, GP98, GMS12, Gu01, HG98, HH98, HN03, HHAA22, HJ09, HD04, HJ17, HMY19, HL03, HBJ09, JW01, JS25, KK20a, KvyS15, Kom07, KM18, LA11, LZ20, LX21, LWYG22, Lie01, Lin01, Ma03, MR20, MVVA09a, MVVA09b, MW93, MZ04, Meu14, MG18, Muo23, Nes16, NMB10, O’L87, OGS20, PXHZ20, PPS10, PTV20, Pel20, PLB22, QMLC15, QAMX17, RZS21, Rya00, SN22, Sch96, Sch16b, SC08, SSR23, SG00, SDK15, Sof17, SS02]. conditions [SHG86, Spi97, Str98a, TM24, UWY22, VV02, Wag85, WL18, WT17, WS04, XF22, XL23, YM24, YBW20, ZBNC25, ZW19a, ZR15, BG11a]. conduction [AX19, BLS94, BCFQ19, DCN^{+}19, JL23a, LW18b, LGZH24, MVVA09b, RMK09, SM20, YY13]. conductive [BBV13]. conductivities [YV17]. conductivity [CML05, DM09a, TT20]. conductor [LSV22]. conduit [IR22, RGK21]. cone [BSZ99, LWZ22, MYSC17, NSCC19, THW19]. cones [MMP02a]. Conference [Ano02g, LST07, FJ97, Wen10a, vdHSW98, Ano00a]. configuration [RGK21]. configurations [BH20, DCN^{+}19]. confined [MT05]. conflict [Fuj99]. conflict-free [Fuj99]. confluent [KW21]. conformable [SA20]. Conformal [LZ24, Ari03, BBO03, CLR11, DDGN23, LWCT07, Mur19]. Conforming [CCZZ18, MM22, AW03, CYM09, DDP12, DN08, GGO16, HSF25, PZMX16, PK91, WCSQ18, XL09b, ZLY23]. Congress [MH89, HSX18]. conic [LDH^{+}24]. conics [DS17]. conjecture [TQA^{+}25]. Conjugate [BGH08, IMMS20, AF23, AKA19, CW20, CSL^{+}24, Cui24, DW00, DW15, EAS12, GR02, HVY91, HWY20, HZC22, HZCZ23, KKLD21, LWLW24, LZW20, MK19, Pfl08, SSW20, SSS^{+}23, SKS^{+}25b, San03, SW95d, SD24a, Wan23, WQS^{+}24, WOLY25, YFLX20, YS24b, YLH20, YLW20b, vdES04, Ria22]. Conjugate-symplecticity [IMMS20]. connected [LZIZ23]. Connection [Pet92, VV95]. conquer [DH94, Jes93]. conservation [AM09, AKG14, Aff94, AL20, Ang06, Bac14, BMGM12, BJ02, BTP96, BFA93, BDF94, BM12b, Bor16, BGP11, BGS02, CFLW22, CCL22, CSW19, CH15, Dav92, DF96, DS23, DGRS09, EK96, FL01b, GZZ19, GT00, GJ17, HSS07, Hor02, HWCF15, KS09b, KXR^{+}04, KLSW06, LPR00a, LPR00b, LCW20, Luo18, NK24b, PGDB08, Pel20, RGB20, SG17, Tan01, TDW23, TJ12, XWZ21, XF06, XCHW22, Ye04]. Conservative [BR20, WDH20, XWW19, BDM03, CHP19, CW25, CCOVF22, CHSS01, CS08, CCL04, CD20b, Fou00, FHX22, FGGL22, FLL11, HCS20, HC22, HLMKZ06, HJL18, JPP19, Kim07, LO22, LDIW16, LR18b, LSWM19, LW21a, LZW19, LD22, MZXX24, MMD20, MQ03, ÖT20, Wan21, WCJ23, XZL19, XWX21, YZWQ25, ZJ25, ZQLK11, MCD20]. conserved [ZXYW22]. conserves [MMDS21]. conserving [BRBM08, BBKS07, HTSZ23, HMD21, Hor93, Ito17, MD20a, MMDH19, SMB23, UWY22, ZFX25]. consideration [SPYS24]. considerations [HFL13]. Consistency [Hor93, Nak24, CMMR23, DCJ20, HZ20, Kel85, LL06, MW24b, SKR^{+}16]. Consistent [HBJ09, Xu13, GHK16, KW20, LS23, Sid90, SMW21, Sin23, SWJ^{+}24, VRK25, WKP12]. consistently [Not92]. consolidation [GLV03, GLV06, PQ25]. constant [Fan11, FP02, GC15, KCJP01, LPS25, LYZ25, MP20, RF16, SH97, TL07, TSF24, WL24, Yan22, ZLJ20, ZWL11, VK17]. Constant-Convection [VK17]. Constants [ND85, Mat86]. constituted [HS22]. constitutive [Die15]. constrained [AKM^{+}22, Arn98, BL21, BMM03, BH93, CW25, CXZ14, CZHX19, CD18, DZW24, EM05b, FLMR14, GO18, GLW^{+}25, Han87, KKLD21, KKR25, KLSW10, LWLW24, LZH^{+}25, LSWW22, LS93, LWCT07, lLX22, MZJ^{+}25, Meh08, NH24, RP01, SP99, SGY22, SD24a, Tou10, ÜSHT03, Yu08, YP18a, YP18b, ZP12, dOF20, ZYH23]. Constraint [Cao09, RP01, ZW87, Che16, HGZW21, LCH20, LJYS20, LCZ21, WWH24, YD07]. constraints [Aca12, AMP20, AKA19, BSZ15, BGIW18, CDI91, CLL23, He24, HdSRI17, KP03a, LDP^{+}14, Lot19, MDT05, MKBS26, PS21, Sch09, SZY21, TWD23, YLS^{+}09, ZYH23]. constructed [AGKS25b]. Constructing [CCDJ20, KDKW20, LS93, Ran15, Ren14, Sad97, SKO19, Tol03, WWS^{+}93]. Construction [BZ91, BJ96, BJ98, CCP17, GH02, HZ02, IJ21, JM17, LWL18, Mat91, MW24b, Nic86, PR09, PT95, SM93, YR09, AG98, ABH14, AX20, CV88, FJ17, JL86, JL87, Lem02, LX08, Man96, MP97, SAH24, SYY20, TY00, WLZ25]. Constructive [AAD^{+}08, JNPC03]. contact [AC18, BBD18, CFRA08, CH13, Cop03, CF05, CF13c, CF14, CA15, CA16, DPPR16, GS15a, Gwi09, HS07, HL02a, HMW05, KR25, KM11, KW12, Por17, WWH24, ZBNC25]. containing [Meh22]. contaminant [SXL22]. contaminants [BM04a, BM04b]. Contents [Ano01e, Ano05c]. continua [Hin97]. continuation [BZ17b, GS89, HH10b, Jac88, JMP06, KMH21, KLSW10, lLXhLZ21, lLX22]. continuations [Lyo12]. Continued [JT88, Bre88, CJV88, CV88, GGMP88, Gil91, Hau88, Jac88, Lem88, Lev91a, Lor10, Njå88, Waa88]. continuity [AW03, LSWW22]. Continuous [BDGP96, BDP99, CW98, CM07, CLGD06, EH97, TYJ11, Arc06, BT97a, BV96, Bel97, BN12, BGS02, CCP17, DS21d, DV95b, DMR18, Ere19, FS19, Fra16, GFB99, JLZ25, JR02, JNPC03, KK17, KK20b, KK22b, LHC09, LOS03, LZ17, LW19b, LWY20, LW22, LDC24, MST09, Pic05, SA12b, VZ93, XZ19, XY19, XZ22, ZL18a, Zha20a, Zha20b, ZX14]. continuous-discontinuous [Fra16]. continuous-discrete [KK20b, KK22b, LOS03]. continuous-stage [LW22]. continuous-time [DS21d, JLZ25, KK17, KK20b]. continuous/discontinuous [ZX14]. continuously [CST97, Zha09]. continuum [Dal00, JEG10]. contours [FPT03, Hin97]. contract [GQ08, MQ00]. contraction [Cer25, PK91, SLMD21]. Contractivity [Bel97, GCHR06, JL94, Zen93, BZ92, Hor99, Kra92, ONL89, San89, Wan17a, in 02]. Contractivity/monotonicity [GCHR06]. contrast [ZCC11]. contribution [Jou05, WM22]. contributions [AFS96, GR02, vS96]. control [AH09, AA25, AS20b, AB09a, AVMVMV09, AW09, BL21, BKM13, BCT16, BGH08, BS96a, BSvdV99, BSZ15, BS24, BC00a, CKP15, CLY19, CZHX19, CL26, CHK99, CL18, CY05, Con01, CF05, Dah02, Dat99a, Dat99b, DMP08, Deh01, De 06, ECB07, EGL09, FS23b, GP23, Han06, HA21, Hey20b, Hig93a, HL19, Hua98, HFL13, HO16, JL25, Jéz04, KP18, KN08, KR25, KR15, LAZ20, LD21, LCH20, LY03, LZIZ23, LT05, Lot19, lLX22, MDHK06, MKBS26, MC00, MP94, NK11, NH24, NH15, NU15, OAHN22, Pea16, PS21, RSD^{+}06, SS19, Sha05, SG06, SW94, Sim94b, Söd06, TLQ21, WZZ21, WCL22, WJM22, WKP12, XL11, YGY15, YXF25, YZ24b, YÇ16, YBW20, ZZ19a, ZLW20b, ZZH25, ZSJ04]. controllable [Leo10a]. controlled [EK95]. controller [Ein18, LT07]. controlling [For02, MJS23]. Convection [RBBC85, AJ24a, BNS25, BBJ25, BO04, BC08a, BBCS05, BC04a, BW96a, BW97, Bla00, BS20a, BTDV10, BFLR23, CW21, CP05b, CR23a, CT93, CPC25, ÇK13, ÇY22, CJLS98, CJ23, CJ24, CJ25, CGK25, DDS89, DYX09, DM11a, Die15, DdCVR03, DL06, EGV25, FZA25, Fra14, Fra16, GAML04, GZZ20, GD23a, GD21, GD09, GO21, GO23, HO24a, HM00, Hua17, JT18, Jun06, KT05, KC03, KS01, KTYY24, KD25, Len00, LY09, LGC25b, LGSW25, LS99b, LS20, LWC24, LD02, Luc05, MOS02, Mar05, MPHFP23, Mat09, MG18, Nag22, NS03, NWL^{+}22, OQ15, OEAS21, PMP23, RG22, Sac93, SRK21, SRK22, Sch91, SZ09, Sha21, SSR23, SKS23, SG07, Str98a, ST08, SWW11, SFZ21, TS24a, Top21, TS24b, Tse00, VCN20, VNC21, VL08, VN21, WL09b, WK00, WPT19, WWF20, XZW19]. convection [YX25, YXF25, YZ17, YZ19, ZMC13, ZZPJ23, ZMF25, ZML^{+}12, ZX14, iW07, VK17]. convection-diffusion [AJ24a, BBJ25, BW97, Bla00, CW21, CT93, Die15, Fra14, Fra16, HO24a, KS01, KTYY24, KD25, Len00, LGSW25, LS99b, OQ15, Sac93, SRK21, SRK22, Sha21, SSR23, SKS23, TS24b, VN21, WPT19, YZ17, YZ19, ZMC13, ZML^{+}12, ZX14]. convection-diffusion-reaction [GD21, VNC21]. convection-dominated [GZZ20, OEAS21, YXF25, ZMF25]. convection-reaction [BW96a]. convective [CGG25, Yao24]. conventional [Rab25]. Convergence [ACMR06, APJ10, AAB^{+}22, ABCC18, AFLG^{+}12, AS20c, BASC17, BM00, BK09, Ben17, BGG^{+}21, BFK11, BW15, Boy15, BD11, Bür13, CGT13, CC90, Car09a, CC04a, CXZ09, CZ12, CC19, CWX21, CWC24, CHS17, CCS17b, DG10, DS05, DYX09, DFLM19, DLM20, DLN04, DH07, Fel06, FM11, GMZ11, GGN12, GD21, GPHA22, GS05, GGRN17, Hau88, Ise94, Ito22, IS22, JK21, JTB15, Jay95, KO08, KB21, Kok08, KFR25, KK86, Kru99, KYI17, Kza92, Kza99, Lem88, Lev91b, Li01a, LSP20, LGC25a, Liu02, LD22, MZS10, MT11, NN20, OT02, PdV99, Ran16, RT14, Sac93, SHL19, San20, SA12b, SHLY19, SNW22, TLGC22, Wan96, WH19a, WS04, YM24, YMD21, YZH19a, ZT06, Zha20a, ZZZ23, ZZ19b, vdSvdH95, ALMM98, AHT17, AM99, AM00, AMP03, AHR12, AF23, ABM17, AGQ^{+}24, AFLP12, AFK92]. convergence [AL05, BWY03, BW23a, BS21, BGO13, BGM19, Bho12, BDFF23, BRZ10, BM09, CGGGS11, CGH23, Cao97, Cao98a, Cao01, CL07, CJX11, CC20a, CH13, CCZ22, Con04, CRHTV25, Cum95, Cuy90, CH90, Cve02, Dar90, DD19, DKSS24, DA17, DMGVO05, Doi91, DLQZ23, EE20, EGH01, EHV19, FFY08, Fdi97b, FW22, FS88b, FR01, FSWZ19, GH91, Gan96, GAZD25, GM94, GWLN22, GH20, HPY92, HO10, HZC22, HLY22, HCX03, ITZ17, JMDN^{+}22, JMU^{+}25, JW01, KS91, KME20, Kie15, KAS17, LR93, Len00, Lev91a, LR20a, LDC24, LZW19, LS21, Lin01, LW20b, LMP99, MK20, MOS02, MO17, Mar03, MT94, MS90, MR94, MT20, Nke07, Oos95, OGV92b, Poh93, PCA10, PRS20, Pré95, QW25, Rou20b, RM25, SKS^{+}25b, SSV89, SW09a, SI20, SXP09, SL09, SW21, SP22, Sid10, SMW21, Sin23, SS13b, TN16]. convergence [TWL23, TX18, TH18, TC22, Vul95, Wal90, WW19, WCS21, WMC09, WZ17, Xu13, YXT17, Yua93, ZAED21, ZZL01, Zha19b, ZLW20a, ZMY21, ZL21, ZZW97, ZYX20, ZZjQ24, in 95, ACR25, GPHAM12, RU25]. Convergent [Fid17, AÁ21, AB07, BM01, BO11, CDW13, CJ23, CJ25, HB20, HW22, KKN^{+}13, KKN^{+}17, KD25, MPDB24, MPMD21, RG22, RTU15, SG04, SKS23, SC22, TQA^{+}25, VV02, Wen10a, Woź10, YFLX20, ZJ19a, ZLG15, ZX14]. converging [BH20, Jéz04]. convex [AM10b, AKA19, BDD^{+}20, CP97, CLTA18, CKK10, CF18, EB12, GFPG18, GH21, He24, HT20, HS21b, KOS20, KKLD21, LZH^{+}25, LS24c, Mik97, PCR17, PS19, RL86, SC11, SZY21, SGY22, SYY20, YJJ^{+}24, YLS^{+}09, ZGF^{+}25]. convex/concave [SC11]. Convolution [SWW16, Bho11, HR14, LM22a, LFS15, Lub92, SND19]. convolutional [RMS17]. Convolutions [ADG11]. convolved [Con04]. Cook [CCZZ18]. cooperate [AC10]. cooperative [CZY08, PK21, SCLL21]. Coordinate [TMS87, Bai25a, Bai25b, CFX08, FJH^{+}01, LT01, ZB19a, ZG20]. coordinate-free [FJH^{+}01]. coordinates [DS15, KDAK13, LWT07, LLT20a, Pet00, WE99, YH00]. cope [KSK^{+}25]. copolymers [ZCY20]. cordial [AN25]. core [MMBB07, MM16]. cores [HK22]. corner [AQ00, CGG02, CKK10]. corners [BP12a, Chu03, FL15, FL20, Lau17a, MPTT17]. correct [Pru00]. Corrected [Con99, AFS00, CKB12, HLW25]. correction [AK21, AB12b, BDFV95, BM06b, CP94, CW98, DSZ15b, DSZ15a, FR01, GDS^{+}15, HS11, Hu99, KM21, Kre07, Lay08, Lay09, LFQH21, LXPY25, NN13, ÖT20, QMLC15, SLW24, VV02, XWW19, XHYM22, YXX24, ZHJ14, Zha19b, ZS21b]. corrections [Agu15, BG03, CH90, GPHA25, HFL12, Hun02, Min04, WY22]. corrective [KN08]. corrector [BK21b, Bur91, Car09b, CXZ15, IM02, JL91, LLL12, MS24, MPPR22, THW19, Wai98, Zla85b, de 95b, vC93, Bur93b]. correctors [Lay08]. correlation [WL21]. Correspondence [Jia12]. corresponding [LfX15]. Corrigendum [AS21, Bic21, BtTBV87, JL87, TLP18a, YP18a]. corrosion [BMP05]. cortical [DEPS15, FOMC05]. coseismic [AA25]. cosine [AMCM09, CM13, Har00, KAD24, SBBC21, ZO14]. Cost [TSFB01, BBG14, Dra97, ECHF^{+}20, FV01, OAHN22]. Cost-effective [TSFB01]. costs [AD99]. Cosymmetry [KT05]. Cotes [Agu15, SKS^{+}25a]. Couette [MD00, SH91]. Coulomb [KPR06, Sch23]. countably [Sob24]. Coupled [ABD16, DS97a, AZA22, AGJ12, AA22, AB24, ACP23, BCG21, BSV09, Bog16, BSP04, Cao98b, CD23, CH19, CPY20, CC20a, CKK25, Den07, DW21, DLM20, DTQ^{+}20, HCS20, HC22, HCC24, HZ20, HAA21, HAR21, HL24, Hus20, IB24, JCSR03, JNPC03, KKR24, KD25, KLSW10, LN26, DBH25, LMA18, LLHC18, LL19, LWW23, LZW17, LZW19, LD22, MD20b, MP05, NS21b, PK23, QAE^{+}09, QH19, QCW^{+}23, SA12a, Sea09, SD25, ST05, SW05, TS23, TT20, Tro93, WPT19, XLZ20, XZL07, YF24, ZZHS18, ZSS23, ZFS24, ZD25a, ZZZX26, ZYH23, Zhe07]. Coupling [EJS11, GMS12, UNGD08, Ben17, BMS89, BEH24, DZ12a, Den07, DMQ02, Dor91, FPS15, FLL11, GQ89, GMM09, GÖS20, HY01, HCW16, JS09, KMS10, MOS12, MT11, Rob10, RVM23, Sus10, UHLZ24, XZH19]. coupon [ALY03]. covariance [FSU89, KK22b]. COVID [CML^{+}25]. COVID-19 [CML^{+}25]. Covolume [Kan04]. Covolume-based [Kan04]. Cowell [LZ22, vdHMdS99]. Cox [AHT17, KL23b, WMC09]. CPR [DSZ15b, DSZ15a]. crack [JK20, MAG13, Par14]. crack-like [Par14]. Crank [ZR15, AD20c, ATAW25, DLM20, FP02, FSWZ19, HJYL19, KTYY24, LN26, LFB00, LL19, LLZL24, LZCF21, LJ20a, QH22, QXG21, TH18, WH19a, WZ22, YCWH23, ZL17, ZJH18]. Cray [CB99, HVY91, Noo95, WVBM88, van95]. Cray-2 [WVBM88]. creeping [GP00]. criss [AR24]. criss-cross [AR24]. Criteria [Ush18, BCV21, Gol00, Gul15]. criterion [CDD00, LPT16, SL09]. critical [Ari04, BH97, CJL13, CH21, MP11]. Cross [BWEP95, AR24, BKR13, BEH24, GS15a, GLS09, IJ14, Meh08, RLMG24, van86a]. cross-diffusion [GS15a]. cross-kinetic [BEH24]. cross-section [Meh08]. cross-sections [RLMG24]. crossed [NT16]. Crouzeix [BC08b, Nud24]. crystal [LMY18, LQXK23, LL20b, LC21, NS13, PHY19, QZH25, WaZW23, ZY19, ZLSZ22, ZH24, ZAB15]. crystallization [ABP95]. crystallizers [QAE^{+}09]. crystals [CPY20, FMW18]. CSCMO [QDA^{+}26]. cubature [CBHY11, DSV13, KK17, KK22b]. cube [BDV17, Nak24]. cubed [CNT07]. cubed-sphere [CNT07]. cubic [ABY22, BY22, CGP15, DMPSC16, DTQ^{+}20, DF92, FPT03, LHN24, LLL08, LZW17, MN03, PPT02, PT19, PRS23, QW25, TLG20, VR01, Yan23, ZY19]. cubic-quintic [DTQ^{+}20, LHN24]. curl [BVV09, HS97]. curl-curl [BVV09]. curl-systems [Lee23]. Current [AL87, AGLRS23, BKP14, BM18, DDZK05, MS08b, RV09, Sel14]. Curtis [CL14, LYZ25, Maj17a]. Curvature [WT17, ALP^{+}96, LP25, Mik97, MŠ99b, YCY12]. curvature-based [YCY12]. Curvature-induced [WT17]. curve [CH22, CGG02, Mai09, YC00]. curve/surface [YC00]. Curved [hYK86, FMW05, GWZ25, HLW25, KX03, LL02, RA09, SFJ^{+}05, XXYZ24, YZZ25, ZKO^{+}21, dSFDG20]. curvelet [SGS20]. Curves [Eis86, Att97, BDV17, BCDR25, CGGM17, GS15b, Han06, LS10, Mik97, MŠ99b, ZS18]. curvilinear [BB25, DS15, FID18b, YH00]. cut [Ber15, BH12b, HLZ14, HD22]. cuts [RU15]. cutting [IMM04, MI03]. CVBEM [HW93]. CVTs [SUP^{+}12]. CWENO [FCX06, LPR00a, SW12]. CWENO-type [FCX06]. CWI [vS96]. cycle [CK98, NN10, NER95, NR97]. cycles [KF97]. Cyclic [PPS10, Alb96, BMR17b, Sch89, dG91]. cyclides [MMP02a]. cylinder [Duf90, DGS24, LMS08, PP00]. cylinders [Son91]. cylindrical [BCS06, IR22, Kwa09, LWT07, Pet00, RGK21, TT03].

D [ZS21b, CCM17, HW15, WC24b, AG05a, AEMX17, ASS21, Ant13, ASC03, AN22, BS21, BF15, BNS25, BB25, BV94, BS08, BCV21, CNA23, Cau08, CK22, CPC25, CPZ17, DC21, DMPSC16, EP15, FMS18, FLH22, Fou00, FS24, GS20, Gon06, HDY21, Hey20b, HWCF15, HZ12, Kal96, KM19, KD13, KNP16, KFR25, KLSW06, KTYY24, LHN24, LPR00b, LLHC17, LW07, LJ20a, Mai06, MKH16, MDA24, MG22, MM20b, NRR06, Now96, OT22, PSP05, PK23, PT23, PPC00, PMP23, QPT23, RZ00, SS00, Sha98, SSR23, Ste97, TCCW89, TYKK01b, UHUL21, VA05, Vej10, WKM04, WSY18, WL24, WPT19, WSC21, YM24, YH00, YZ17, ZZL01, Zha09, ZL18a, ZS21b, ZS18]. D-convergence [ZZL01]. D-flows [BS08]. D/ [ZS21b]. DAE [CY98, NR97]. DAE-aspect [NR97]. DAEs [AFS00, AC96, BDF^{+}25, BHSW20, CCMSS11, Cam99, CZ97, GPHAM12, GS09, HMT03a, HMT03b, San02, ST05, Wen05, YT03]. Dai [YFLX20]. Dai-Liao [YFLX20]. dam [CGN03]. damage [CFKS07, CFRA08]. Damped [KH91, AD19a, AD21, CC23a, CD20a, LFS21, LT23, MDP10, Moe98, MDA24, SD24b, WL22, WTF25]. damping [CRR03, GLQ26, Jes85, JZS20, LHN24, LN26, LZ24, LR01, QM19, QZM25, XLZ20, YLY19, ZS21b]. Dantzig [YP18a, YP18b]. Darboux [Bre91]. Darcy [BC12, CD23, Chi21, ÇK13, DZ12a, EJS11, EL01, GG22, HS17, JL25, Kie17, LLHC18, LH20, LFQH21, LHY24, QH19, QCW^{+}23, UNGD08, ZKO^{+}21]. DASPK [MD96]. Data [DR93, LP05, LRRÁY25, Vas17, WCGW95, AAB^{+}22, BG02a, BCDR25, BRW17, BRS^{+}18, CY23, CPZ17, DC21, DMM24a, Ehr08, ECHF^{+}20, FS19, FvdMS17, GO19, GO23, Har10, HKZ08, HM15, HM17, Jun97, KKN^{+}13, KOS20, KKN^{+}17, KDS22, LO96, PS02, PS03, PS19, RS09, RL86, RCGM98, RW87, RV05b, Sch08a, SS19, SHB^{+}25, SND21, TT20, TWD23, TD09, WL21, ZJH18, ZMF25, ZHL08, ZXW17]. Data-dependent [DR93, LRRÁY25]. data-driven [TWD23]. Data-parallel [WCGW95]. Data-sparse [LP05, Vas17]. databases [WWS^{+}93]. datum [NTHC21]. Davidson [Gen10, MS99a]. DC [HD23]. DDE [TS06]. DDEs [RFMV24, ST01, Sha05]. dead [HK22]. Death [BMR17b, BLM17b]. deblurring [BS10, DE16]. Debye [HJX^{+}19, YWH20]. DeC [ÖPT25]. decay [BDKM92, Maj17b]. December [Ano21s, Ano22s, Ano23m, Ano24r, Ano25u]. decomposed [GCP91]. Decomposition [CFC03, TMS87, AAD^{+}08, AF04, AM10b, Att97, ABCC18, BN99, BRRS15, BM89, BRVC09, Bog00, BP90, BS91, CG89, CES91, CJ90, Chn17, CDG19, CG16, Cve02, DC21, DS21c, De 93b, DA18b, DDS89, Dor91, Dou91, EZ98, Ewi91, FL05, GDEdLD23, Gas92, GNNR19, GK09, Gen10, GG19, GHF00, GS08, Haa97, HPS12, HSS07, IVA93, JM06, Jun97, KCS07, KG90, Kok08, Kop89, KJ99, KR20, Kuz90, LY08, LY09, LX21, LZH^{+}25, LR00, LOM98, LRT99, LZY09, MH14, ML91, Meu91, MLB97, MVBS25, NX22, NRWF08, NTT22, NMB10, Par04, PKSB10, Pas91, Pav00, PR90, RV04, RV05a, RBdM25, SWW17, SG09, Ste05b, Su94, SLZ10, TS08, Tse00, VBVA22, VN21, WM07, WJM22, Yu99, ZS21a, ZHL03, ZZLL21]. decomposition-type [BN99, Cve02]. decompositions [Phi91, PT15]. deconvolution [Dun18, FLMR14]. Decoupled [CHLA21, HCS20, LMA18, BSZ22, CLL25, CH25, GLML20, JY23, LH20, LC21, QCW^{+}23, SFZ21, WWZ25, ZZHS18, ZFS24, ZFX25]. decreasing [DKK94]. Dedicated [CHM09, FFMZ13, LVW21]. Dedication [Ger94, GT93b]. deduced [AD19b]. deep [OXW25, TK19, WTB24]. default [CC23b]. defeating [BO11]. Defect [AK21, FR01, QMLC15, ZS21b]. defect-correction [QMLC15, ZS21b]. Defect-deferred [AK21]. deferred [AK21, BM06b, CW98, HS11, Kre07, Lay08, Lay09, LXPY25, Min04, ÖT20, VV02, YXX24]. defibrillation [CKP15]. deficient [Baz03, SY07]. defined [Att97, BWS21, CFS13]. definite [AA05, Ara99, Cao01, DL01, LWZ22, Liu02, Lu98b, MAD23, Sha98, SH21a, TM15, Zha21a]. definiteness [Mar08]. definition [KW93, ZS18]. Deflated [EAV16, Du11, WPS18]. deflating [AA94]. deflation [BWEP95, Mor05, VSeYD02]. deflections [YTZZ18]. deformation [GKT10, KM11, TYKK01b]. deformations [PFHL09]. deformed [GKB^{+}22]. Degasperis [GZZ19]. degeneracy [Li08, PPS05, SZQH23, SHB^{+}25]. Degenerate [DFZ16, KXK92, ASGGRGW25, AJ24a, AD23, BBRBS09, BBCS05, BCS06, DSSC13, GGT24, yGyZ07, KL98, KK09a, LC99, PK23, Pło22, Spi90, WYYL19]. degradation [DSSC13]. degree [AA04, AL09, ATW20b, Beg00, BWY17, DZZZ24, DhW09, Has08, Kin94, LZZ18, Mil17, PM91, PLB22, Sal03, Xiu08]. degrees [AHGM21, Mau08]. DEIM [YV17].

Delaunay [LS93]. Delay [SG06, ZFX17, AD19a, AD21, AHT17, AB24, BF92a, BP06a, BZ92, Bel97, BR94, BFdS10, Bre06, But92, Cah92, CSSZ20, CHZ14, CZ12, CWC24, CFM^{+}24, Den15, DFLM19, EH97, Enr06, Fan11, FJ97, Fel06, FNT06, For25, Gan09, GHHG22, GeO24, Gug05, Guo01, HAML21, HZD21, HJ17, HS95, Hu99, JVZ97, JZK06, JP19, JCY25, KMS19, KV95, KO96, KF97, KK20c, LZ14, LZJ21, LZ22, LW92a, MZK05, MPDB24, NP21, NLS20, OB24, OZHP23, Pau92, PH17, PWX24, RA17, SMJ12, SDK24, SH97, SOB20, SZ17, SA20, Spi97, SAMSB20a, SAMSB20b, TWL23, TYJ11, Van00, WZL08, WC11, WGW15, WCM23, WB92a, WB92b, WMC09, XY19, YXT17, YRV21a, ZZL01, ZH09, ZC10, ZHL22, ZZH25, ZZO16, ZYX20, ZP97, ZP98, ZX09, in 92, in 96]. Delay-dependent [ZFX17]. delay-differential [BR94, WB92a, WB92b]. Delay-differential-algebraic [SG06]. delay-integro-differential [BF92a, ZH09]. delay-type [NLS20]. delayed [LLD18, PMP23, RREP^{+}20, SS25]. delays [AKS21, BT97a, BCT16, HHT97, Hig93a, Hua00, Kür23, LRE04, NT92, OHA25, Pis22, RE19, tSqWyG16]. delivery [FGP23]. DELSOL [WB92a]. delta [BW15, GGS04]. Deming [Mon09]. demographic [OHA25]. dendritic [LQXK23]. denoising [BS10, Han19, YCY12]. denominators [PA91]. dense [AZHD23, FMS24, LW92b, ZNK02]. density [ASGGRGW25, An20, AKT97, CCDJ20, CKS05, LA21, SW06, WWC^{+}25, WWF20, ZFS24, BNH01]. Dependence [BN03]. Dependent [FG96, AB09a, AL17, AL22, AF89, ARS97, ABRW18, ÁMS17, BL21, BHJ05, BP14, BDF89, BK21b, BtTBV87, Bre02a, BK12, BJ01, BJ03, BJ06, Cah92, CP05b, CL24a, dCCSR03, CCS17b, CST97, CT21, DFLM19, Dia95, DY03, DZMB21, DMA22, DR93, EBEHAS24, EEAS25, FNT06, GS19, GLPW09, GÖS20, GGR97, HHT97, HP14, HM15, HM17, HILK13, HR96, HLIS16, JCSR03, KK20a, KH91, LKV01, Lay09, LRRÁY25, LGC25b, LLT07, LHX20, Lua17, MG97, MS08b, MD20c, MD25, MC21, NT92, NK24a, NC16, Nor97, Pea16, PJB04, PAJ12, RZ00, RY13, RGA19, Rya00, Sar05, SM13, Sim10, SS17, ST25, SvdVvD06, SZW19, TER03, Toc01, Wan09, WZW13, WCM23, Won08, WdG92, XZW19, XL23, YS09, Yan22, ZLW22, ZJH^{+}23, ZFX17, ZYS17, ZJ10, tV87, vdVS08]. depending [BY22]. deposition [CKB13]. derefinement [PPC00]. Derivation [Chr01, CMRV11, CAAT16, EGM25]. derivative [AY21, AH11, ABH14, AH15, ACDP22, AHA23, AHO16, BP02, BB94, BMV06, CSLY19, DCJ20, EH07b, FH20, GJ00, Hey20b, IAH25, KGR08, KAD24, KKLD21, LR20b, Moo04, MAF20, MSA20, MAH22, PB21, Ram12, RGA19, Rus95, SC20, SSA^{+}22, SL15, SND19, WL16, WB92b, Wu09, ZY14, iW07, iW09, iM13]. derivative-free [IAH25, KKLD21]. derivatives [AGY08, APJ09, BFK11, CHZZ06, Che12b, CHS17, CD00, Dal13, DO17b, DR01, FP24, GND19, HEG16, JLH13, KCS07, KSHB21, KS01, LHWF08, LHÖ13, LfX15, MD23b, OB20, RMH20, Rat13, SSA^{+}22, Tan87, XC20, ZLG15, ZZ19b]. deriving [LIPT18]. descent [Bai25b, HZC22, SD25, Wan23, ZG20]. descent-based [SD25]. describing [CML^{+}25, GRGJ02, Rou20a]. descriptor [EK97, GKS20, Sim93, Udd20]. DESI [Dia95]. Design [Ale03, BZ17b, JMDN^{+}22, PSW02, RGMO19, RREP^{+}20, SW95a, AC08, ARSW05, BdFPSdSC08, CH04, IPL02, JPP19, Jam93, PLI03, SA08, SG05]. designs [AGKS25a, AGKS25b, AX20, KKL^{+}25, KS10]. destruction [BDM03]. detect [GGS04, MMKN17]. Detection [Eir99, HK22, AS21, AC10, BCV25, HS21a, JD09, JGK11, KM95, KXR^{+}04, Lo06, SSW04]. Determinantal [Pre90]. determinants [WB90]. Determination [Deh01, ZH20, AED12, CJL13, HK09, KYI17, LV12]. determined [SK10]. Determining [AGY08, LN08, OAHN22, SOA^{+}25, Su94, WL09c]. Deterministic [GS15a]. detonation [ZG92a]. Developing [BDF89, Pau92]. Development [AL87, Bor97, RA03, de 96, AFS96, DdSF07, FPPS00, vS96, van96]. Developments [Bak89, AQS94, ABFV09, BC89c, Dat99b]. deviating [BF20]. deviations [CWC24, HD23, Yam23]. device [LS07b, YY13]. devices [AC08, BBS11]. DFPM [MZJ^{+}25]. DFPM-based [MZJ^{+}25]. DG [BBD20, Ort20, YSZG25]. DG-VEM [YSZG25]. diabetes [MLK06]. diagonal [BL21, BLY17, BKS07, Fuh01, TW00, ZH15]. diagonalization [HC01, WL24, YW24]. Diagonalized [AyLqW18, LLJY20]. Diagonally [But93, DR09a, JR00, KC19a, ABH14, BKAG22, BJ96, BJ98, BW03, CDP12, FS08, HH10a, Hor98, JVZ95, KW20, MAF20, MSA20, RBdM25, SAH24, vSC92, vS93, WBCK02]. Diagonally-implicit [But93, BW03, CDP12]. diagrams [Mar93]. diameter [CL07]. Diamond [LC19]. diblock [ZCY20]. dichotomically [ELLE02]. dielectric [AA20, HJX^{+}19, YWH20]. diesel [SM08]. Difference [BKM95, BCE04, BGG^{+}20, FLH22, KKP17, Pot85, RBBC85, RS00, TS24a, AD20c, AES15, AEMX17, AAH21, AA24, AAI^{+}93, AFS00, APJ10, BM00, BRTB19, BM12b, BIO24, BS18, BS20b, BCS06, CFLW22, CCDJ20, CCG13, CGA93, CHSS01, CP05b, CC19, Chi12, Cho13, CCL04, CS18, Con99, CBHM19, CMCGTR02, DS21c, DA18a, DL20, DW21, DL22a, Dor01, DS15, DCJ20, EP15, Est95, FE93, FWHM20, FG98, FKA^{+}13, Fou00, FGGL22, GX11, GNUV24, GLV03, GHKM09, Gla94, Gol00, GZHQ23, GLQ26, GHF00, Hag15, HHAA22, HZ09, HHL23, HC22, HCC24, HZD21, HSWW24, HL21, HL24, HLJ20, JL86, JL87, JT18, JZS20, Jor11, JV09, Kam00, KT05, KBS11, Kat89, KD13, Kni94, Kni95, Kop89, KDK17, KW93, KLSW06, LO22, LHN24, LN26, LCK22, LY08, LHWF08, LY09, LR18a, LR18b, LR20a]. difference [LW21a, LWYG22, LXZS22, LHY24, LYZF25, LZW19, LS99b, LCL18, Lte24, LR20b, Lyn99, LLW20, MDP10, MP11, MW24a, MM14, MT06, MM20a, Mic03, MFAD23, MdD04, Nag22, PXHZ20, PWY21, Pap95, PCL^{+}25, Pir09, QXQ22, RZ00, RZ18, RL21, RTU15, RGMO19, RG21, RT95, RU07, RU15, SS25, Sam94, SXL22, SDK15, SBS24, SA18, SL15, SMA01, SSKS21, Sto96, Str98b, SWW11, zSW06, SLZ10, TZ21, Tan93, TYKK01b, TDW23, Tol03, Tsy96, TY00, UWY22, Vab22, VVD95, Ven15, VRC21, WL09b, Wan17b, WL18, WR20, WDU21, WZ22, WLG22, WDL23, WS04, WSC21, XWW19, XZL19, XLZ20, XWZ21, XWX21, XGQ20, XXF22, YZG23, YDWW17, YH07, ZG21, ZQZ23, ZJ25, ZW19a, ZLL22, ZZL17, ZZLL21, ZFX25, iV09, vSW90]. difference-quadrature [AAH21]. difference/compact [QXQ22]. difference/finite [DA18a]. difference/operational [SSKS21]. differences [Gul15, Zeg97]. differencing [AEF^{+}14, Dea11, DLQZ23, RK91, SC25a, van86b]. different [BW23a, CHLA21, DCN^{+}19, DJM09, GGG16, HS02, HHL23, JUAZ22, JT09, KW21, LWD^{+}09, MM16, OAHN22, Set24, SZ22a, SWFK13, SC08, SFZ21, Yu99]. differentiable [Zha09]. Differential [BGHR12, BGH^{+}15, FR18, Jac87, Mär02, Pet87, Rei85, RU07, AKM^{+}21, AJ19, ABZ21, AGZD22, Abu04, AAL21, AB88, ABdSG23, ASA20, AAH21, ADM22, ACLM22, AM99, AM00, AMCM08, AAL^{+}26, AT93, ABI22, AL24, ABF09, AF89, Arn93, ARS97, ALP^{+}96, AKS21, AD19b, AAEMY21, ADH00, AV00, Bac16, Bac17b, BTBR20, Bac21a, BS14a, BHJ05, BHJJ06, BKM95, BF92a, BT97a, BBPR05, BP06a, BF20, BL05, BCG21, BKP09, BJ05, Bas21, BZ92, Bel97, By01, BMGGG12, BGGG13, BDF89, BF92b, BBC^{+}25, Bho12, BCC16, Bic21, Bla01, BtTBV87, BR94, Bok03, BJ11, BMPR15, BJ20, Bre06, BP92, BMM97a, BO21, Buc06, BS12, BS18, BS20b, BC89b, BB96, BDP96, BB98, But92, BC95, BJ96, BJ98, CGEV19, CC90, CGA96, CL01a, CHZ14, CNA23, CDP19, CD95]. differential [CSS87, CCM02, Cha98, CDH24, CZ12, CSLY19, CGW20, CDW23, CWC24, CL01b, Chi12, CC20b, Con20, CST97, CN11, CN15, CP03b, DD21, DV20, DS05, DR09b, Den93, DL13, DLPV17, DKL24, DB08, DCY20, DAMA23, ELCWS98, ESEKZ10, EGV25, ER18, ELLE02, EGL09, EMMK01, EH97, Enr06, FM21, FK23, FID18a, FMMK01, Fan11, FL93, FH20, FH22, FTB97, Fel06, FT06, FNT06, FGP23, Gan09, Gan96, GS99a, GS19, GHHG22, GMG19, GAZD25, GPHA06, GLM09, GM17, GS25b, Gu19, Gu20, GNAS^{+}20, GAOB20, GN86, yGqWsWC05, yGyZ07, yGpY09, GGO13, GW20, Guo01, HM87, HZBM05, HP18, HAML21, HJ05, HK25, HZD21, HJ17, HS19a, Hig93b, HXW15, HCY18, HJ21, HR96, HN22, HL89, IM98, Ise02, IJ17a, JL91, Jac93, JVZ95, JVZ96, JVZ97, Jac02, JRW06]. differential [JZK06, JAH21, JK21, Jay95, JW01, JCY25, JCN94, JMPY10, KMS19, KPY15, KV95, KK11, Kau95, KC19b, KO96, KLR25, KS08, Koz94, KP15, KW20, KGZ24, KKR15, Kür23, KDKW20, LHHR94, LZQ22, LP24, LPS25, Lay09, LLKJ21, LPZ00, Li05, LZ14, LZ17, LYF17, LW19a, LWY20, LW20a, LZJ21, LZ22, LSL11, LPV24, LMSW17, LZL14, LYK17, LLD18, LTT19, LMTW20, LLZ^{+}22, LMW23, LYLL23, LCZ23, LSW23, LP00, LM22b, LS98, Luc05, MDD14, MAH18, MHA19, MD19b, MP96, MKN23, MCS16, MWC21, Mär95, MMRV20, MT11, MZK05, MT06, Mit22, Mit24b, MMM19, MFAD23, MD20c, Mok17, MP94, MJS23, MN08, Mur99b, NP21, NMKE13, NLS18, NLS20, NT92, Odi19, OB24, ÖPT25, Olv92, Ost93, PM05, Pau92, PT11, PTV16, PTV20, PVM22, PH17, PAJ12, Pot97, PM14, PWX24, PG02, Pul05, Pul09]. differential [PSL18, QWX20, QXG21, RR21, ROB17, ROL19, Rah25, RE19, RA05, RSR23, RTW21, RT20, RKVZ15, RN22, RTV00, RTV02, RU25, SS08a, SS25, SA90, San89, Sar05, SA12a, SMEN04, Sch02, Sch87, SOB20, SD13b, SG06, SGS20, SDG20, tSqWyG16, SZ12, SZ17, SWB20, SWB21, SSA^{+}22, SP22, SZQH23, SA20, uIVS13, Sob24, Sol15, SvdHN86, SCvdH92, Spi97, SSKS21, SSPZ20, SL01b, SG17, TN16, Tan93, TX18, Tem15, TLG20, Tho85, Tia15, TB01, TYJ11, TZA13, TS24b, TDPU17, TV91, Tro93, VV09, Van00, Ver96b, WGKS12, WZL08, WG10, WC11, WWX13, WGW15, WMF17, WH19b, WLM21, WCM21, WZ22, WCL22, WCM23, WAV12, WC14, WYL11, WHL19, WB92a, WB92b, WSS97, WdG92, XFLC00, XY19, Xu13, XZZ19, YXT17, YMD21, YBL13, Yan18, YS22, YC13, YZH24, YDWW17, YY24, YRV21a, YRV21b]. differential [Yüz22, Zak20, ZHS22, Zha19a, ZG92b, ZH09, ZC10, ZZ19a, Zha20a, ZD20, Zha20b, ZZO16, ZFX17, ZW19a, ZLX19, ZYX20, ZSJ04, ZZ19b, ZJ19c, ZSZZ20, ZP97, ZP98, ZX09, ZC99, ZAB15, ZK00, Zuu95, dDF^{+}94, in 92, in 95, in 96, tV87, vSW90, van96, vS97, vvdV97, vdHSW98, vB95]. differential-algebraic [Arn93, BL05, BCG21, BKP09, BDF89, DS05, EGL09, Hig93b, Jay95, JW01, KKR15, LPZ00, LLZ^{+}22, Luc05, MP96, Mär95, Mur99b, Ost93, RA05, SA12a, ZP97, vB95]. differential-delay [KO96]. differential-functional [ZK00]. differentiated [KP92]. differentiation [Bal00, Ber86, DGD03, EES05, Jam95, LLT20a, RH92, Ske89a, Ske89b, ZL11b, Zha21b, ZYJD25]. DiffMan [EMMK01]. Diffraction [CJ22]. diffractive [DHL00]. diffuse [LS24b, Sch16b, YLW20a]. Diffusion [GM85, GNZ21, AS11, AD19a, AD20c, AD21, ALMM96, ALMM98, ALMM01, AJ24a, ABJ12, ASA20, AKBF19, AJ24b, AC18, AL98, ABR05, AGJCP25, AN22, BPN24, Bar05, BNS25, BL15, BBJ25, BAD13, Bec18, BM01, BB25, BO04, BBRBS09, BC08a, BBCS05, BC04a, BM04b, BW97, BK21a, Bla00, Bi20, BS20a, BIO24, BTDV10, BBD24, Bra22, BFdS10, BFLR23, BK12, BJTZ20, BC97a, BCDP17, CGGGS11, CHZ21, CHM22, CFCH09, CC18, CGEV19, CdFN01, CW21, CCQ^{+}23, CM02, ÇD17, CR23a, CT93, CDW13, CLTA18, CW22, CL24a, CQZ20, CPC25, Chn17, CK20, ÇY22, CJLS98, CJ23, CJ24, CJ25, CF08, CGK25, CA16, CRSF19, DRVA20, Deh01, DYX09, Den15, DM11a, Die15, Din19, DHM09, DL06, EGV25, EV96, FZA25, FMP04, Fou00, FV87, Fra14, Fra16, FJ95, FS24]. diffusion [FSWZ19, GAML04, GS15a, GLS09, GHW20, GZZ20, Gas92, GIS23, GD23a, GV02, GG19, GD21, GT19b, GPPR12, GPHA16, GO21, GO23, GOLJ25, GS94, HM15, HO24a, HW06, HEG16, HQAZ24, HAA21, HS95, HK22, HFL13, HS19b, HAY20, HZAT21, HS21b, HM22, IKM23, JRS20, JR18, JT18, JWZ21, JHGZ20, Jia12, JL24, JT06a, Jun06, JOL23, KM95, KKT16, KP18, Kał22, KV07, KK20a, KAS22, KC03, KDH20, KOW05, KL09, KS01, KS04, KZ21, KTYY24, KD25, LWT07, LW93, Lan95, Lan97, Len00, Li01a, LY09, LR18a, LSP20, LCHW20, LWYG22, LGC25b, LGSW25, LLT20b, LQS21, lLNW21, LS21, LS99b, Liu09, LYY15, LCLW17, LCL18, LLZ19, LZCF21, LL21, LYZW22, LSW23, LWC24, LZY24, LXCM21, LO96, MOS02, MD19a, MMKN17, MN23, MSZ^{+}24, MPSS16, MSGM23, Mar05, Mat09, MM20a, MG18, MOV24]. diffusion [MOSW00, MHL18, MRFF17, Mul19, Nag22, NS03, NLHT25, NWL^{+}22, NMB10, OQ15, ÖPT25, OZHP23, OH20, Ort20, OCVW22, OAHN22, Par04, PGYF20, PSWZ21, PS09, PMP23, Que21, RSL89, RZ18, RL21, RG22, RY13, RREP^{+}20, Sac93, SRK21, SRK22, SDK24, ST11, ST14a, SZE20, SZ09, SRMDRL23, SM13, SK22, Sha21, SY18, SC20, wSJP15, SSR23, SJ18, SG09, SKS23, SG07, SW13, SA18, SA19, SL15, SOA^{+}25, Str98a, ST08, Su94, SWW11, SW24, SvdVvD06, zSW06, SW17, SZW19, TS23, TS24a, TSB10, Top21, TS24b, TH09, Tro96, TMM15, VA05, VVR08, VCN20, VNC21, Vej10, VBVA22, VL19, VN21, Wan01, WL09b, Wan09, WZW13, WR20, Wan20, WZZ21, WCS21, WY22, WC24b, WW14, Wei09, WB03, Won08, WG11, WPT19, WYYL19, WG22, XZW19, XF22, YX25, YXF25, YL13, YZ17, YLFT20]. diffusion [YXN21, YZ19, YLLZ21, YWSL20, ZMC13, ZCZ15, ZWJ18, ZJ19a, ZYLL20, ZG21, ZZPJ23, ZWN23, ZZX20, ZLL22, ZML^{+}12, Zhe19, ZZ20, ZEW20, ZLWF21, ZJLA22, ZSY20, ZD25b, ZX14, dlHV13, iW07, iW09, iM13, VK17]. diffusion-advection [LWYG22]. diffusion-reaction [MD19a, VK17]. diffusion-wave [AD19a, AD21, BK21a, CC18, Din19, HZAT21, LR18a, LCL18, LLZ19, OCVW22, SA18, zSW06, YLLZ21, ZJ19a, ZZ20]. diffusions [AY22, Dah02, DS23, Spi90]. diffusive [BL21, BMMZ06, ÇK13, CCM17, KAD24, MMP20, QPT23, Zha19b, ZYJZ23]. diffusive-viscous [ZYJZ23]. diffusivity [MG18, WZZ21, YW19]. digital [GG95]. dilation [GP17]. dim [HA21]. dimension [CBHM19, HL89, LL98, MP11, SSvG10, TM04, WSS97]. Dimensional [Per88, Pet87, SR88b, TMS87, TD09, AS11, AD20c, AJT19, AA24, AW03, AJ24b, AMR12, ASCM02, Bac14, Bak89, Bar09, BLN25, BBRBS09, BG11b, BZ17a, BFA93, BS94b, BBLT15, BM18, Bri85, BS12, BF95, CR23a, CH01, CDW13, CLTA18, CQZ20, CAAT16, CJ23, CJ24, CJ25, CY05, dCCSR03, CPOGO17, CMCGTR02, DhW09, DE06, Deh01, DA18a, DA18b, DHWL22, DCC14, Din93, Din19, DN08, ELCWS98, EEAS25, FP02, FD97, FdSB02, FCW20, FXCW21, GX11, GM16, GMM09, GNX19, HBA^{+}24, HZ09, HPH20, Hor99, HK85, HR97, HZAT21, HS21b, HY24b, HN22, JTB15, JZXJ21, JHGZ20, JZZH22, JCJP21, JT06a, JGK11, KTK20, KV07, KME20, KM16, KF97, KCC04, KS09b, LM00, Lan95, LCHR03, Li16, LC19, LH21, LLZL24, LGSW25, LS07b, LLZ19, LZCF21, LGZH24, LCZ25, LAH09, LS93, DLM16]. dimensional [LMWZ07, MQO17, MS19, MM22, Med96, Mit24b, MH16b, Moo95b, Moo95c, Mus11, NAD25, NY13, OFY^{+}23, OGS20, Pan21, PR90, QM10, QWX20, QXQ22, QNA23, RV22, SHL19, Sch91, SLJ86, SLW17, SR09, SC25a, SXP09, SC20, SJ11, Shi25, SZE^{+}92, SBS^{+}20, Sob24, SvdHK94, SK96, SSA24, SW12, Su94, SWCH15, TBRBM20, Tem23, Ter22, TJ12, TV91, WZL13, WQ17, WMLB19, WLM21, WCS21, WCM21, WHW21, WPN^{+}24, WKWC26, XC85, XL09a, XL23, XGQ20, XZT21, XLZ23, XWM26, YQCZ22, YWW23, YL26, ZAED21, ZJ19a, ZQZ23, ZZX20, ZLG24, ZCSH11a, ZQLK11, ZCSH11b, dlHV13, iW09, ZJR25, vSK97]. dimensionality [WL21]. Dimensions [MOZ87, BHJJ06, BD11, CMP03, CH25, DL20, EFLFP09, FL09, Gar03, HL21, Hus20, JWZ21, KZ13, MZN21, MPSS16, Mar05, MD96, Ran20, Ros93, Str98a, TWMP20, TM05, WCJ23, ZLCH20, ZW19a, ZLWF21]. diminishing [FL24a, JL86, JL87]. DIMSIMs [FJ17, Jac02, JM17]. dipole [Nes16]. dipoles [HNP17]. Dirac [BB25, CL20, FXY22, Gia12, GGS04, LW21a, Li22, WCJ23]. Direct [ECB07, Leo10a, VV95, YC13, AT93, ABI22, BCT16, BJM01, Car23, CM02, CGCMTR02, CCJ99, CMCGTR02, CRSF19, DB08, HFL13, HS19b, HT94, LCVG01, LZ20, Map05, RCR24, ZS18, Zla85a, MS13]. direction [AD20c, ACM91, BMSZ21, BJ00, CSXL14, FMU15, GX11, JLZ25, JL94, KH91, LLZ19, LZW20, Phi87, QWX20, SN22, XZZL15, ZN21]. direction-dependent [KH91]. directional [BKR13, Dal13, DDK19]. Directions [Bie87, CXZ14, Sch99, dRT99]. Dirichlet [Ant13, Bie12, CY23, DA16, DYZ20, DP85, DY03, GP23, HJ09, HBJ09, LZIZ23, MH14, Mau08, MG18, OMP98, PA18, Sch16b, VBVA22, Zho18, vR04]. Dirichlet-to-Neumann [Bie12, OMP98]. DIRK [Ale03, Cam99, HO10]. disadvantages [Car09b]. disaggregation [Mar03]. discontinuities [AG05a, AGY08, BBW19, CFXZ06, Hig93a, JD09, KS07, NTT22, SB18, WB92b]. discontinuity [DL13, MQO17]. Discontinuous [BP06a, CGJ16, Cas06, Fre98, GHH09, Hop23, Kam16, Lia22, LYK17, Por17, SW05, VCN20, WSHC20, YZ17, ZYJZ23, AD20a, AKG14, AL09, AW14, AS20a, ASA20, ABI22, AFS11, Aug89, Bac14, Bac16, Bac17a, Bac17b, Bac17c, Bac18, Bac19, BTBR19, BTBR20, Bac21a, Bac21b, BVT14, BB25, BTP96, BS20a, Bus06, CY23, CP06, CGN03, CL20, CCK08, ÇY22, DGN12, DN13, DL13, DMR18, DL16, DL21b, ER07, EKT19, FD16, Fra16, FJ95, GP23, GM18, GS20, GÖ20, GO21, HLMP09, HGF24, HSS07, HMP14, HJL18, HH18, HS19b, HAY20, HJ03, ID19, KPY15, KB21, KwS19, KWLK00, KXR^{+}04, KQ13a, KQ13b, LH11, LHS00, LZ14, LW19a, LW20a, LB23, LGSW25, LS07b, LWC24, LYMD24, LM22b, ML16, MFAD23, MD25, Mus11, NS21a, NWL^{+}22, NSD23, PCRR17]. discontinuous [RC18, RSR23, SRK21, SC19, SA12b, SZ09, wSJP15, SvdVvD06, Tem15, Tem23, VSeYD02, Wal19, WCXL09, WTY21, WJM22, WWLL23, WaZ24, Wen25, XF06, XWM26, YW19, YT00, Yua20, ZY14, ZYZJ24, ZZ17, ZZ20, ZX14, ZSQ20, ZSQ21, vR04, vdVS08, vB95, CHZ21]. discontinuously [VT91]. Discrepancy [Sch04, DE18]. Discrete [BKK25, DHWL22, DLM05, Ehr08, FHK05, Har93, KR18, LMS08, PP24, PZ20, WY22, XG22, AZHD23, AM09, Ale11, AD18b, BS14a, BM12a, BHL^{+}21, BKM13, BRIP08, BCE04, BF09, BH96, BL06, BGS02, CXNF14, CRS05, CGW20, CLL25, CCZ22, CMS04, Dal00, DV95b, DLM20, Dun18, EV96, Enr06, EEE22, FGPR12, GFB99, GNNR19, GKS20, Gon06, HAA21, Hu99, KPR12a, KR20, KK17, KK20b, KK22b, LRS23, LOS03, LSVT25, LSK12, LLY21, LYZJ23, LCL18, LXCM21, Lyn92, MD22, MW24a, MC17, MW24b, MKJ23, MRH14, MWYZ18, Nak24, OHA25, PR09, PT11, PD01, PLB22, RKR20, RT14, Rog19, SZ22b, SMW21, Sin24, SW24, zSW06, SND21, TZ21, VVV24, Vej10, jWjJ17, WS21, jWC22, WQS^{+}24, WWZ25, XP23, YT25, YK04b, ZQY18, Zha19b, ZFC20, ZD21, ZLSZ22, ZWK15, ZR15]. discrete-ordinate [Sin24]. discrete-time [GKS20, LSVT25, WS21, ZFC20]. discrete/continuous [GFB99]. discretely [FV01]. discretisation [BDM03, HvdHV10, JR18, LT19, QPT23, TGB08]. discretisations [GÖ20, JPP19, Mat09]. discretised [HS95]. discretising [PSL18]. Discretization [AB09a, BLN25, FI03, LS98, Roo20, AMT13, AZ23, ASZ15, AKT97, AB12b, BM12a, BCS17, BW95, BK12, BEH24, CGT13, CLT97, CWZ23, DY17, DM11a, DT10, FMS18, GHK16, GAW09, GTS20, HL99, HT94, KDAK13, KM19, KPR06, KvyS15, Kra92, LPT94, LSV22, Li25, LCLW17, LO96, MS86, MSZ^{+}24, MR20, MMDS21, MMP20, MT20, NBNTGV11, PSWZ21, Que21, Rei99, Rob01, SW12, TN16, Tau09, VB07, VK17, VL08, WJM22, YB10, ZBY19, ZFS24, vR04]. discretizations [AMC02, APJ10, BVV09, BJ01, BJ06, CGMS21, CHOR19, DLZ21, Fun04, GLV06, Gje07, GT93a, HS98, HW97, HST14, HS97, LO95, Mul99, Osw97, PJB04, Sch02, Sch87, SC25b, Str98a, XG22, YX25, vSK97]. Discretize [SC11, CD18]. Discretize-then-relax [SC11]. Discretized [Imo00, ADH00, BL21, BN12, DJ10, Guo15, Ise97, Poh93]. discretizing [HAML21]. Disease [HL97, ABdSG23, BS24]. disorder [Den07]. disordered [ABD16]. dispersion [AB97, JL17, LR18b, ÖPT25, ZM19, van86b]. Dispersive [Rog19, BBD20, DM09b, DGM18, DMM24b, DL16, Est95, IJ21, JUAZ22, LSG24, LT23, PD01, WXY24, XK25]. displacement [FBS09, LY08, LHY24, RVD00, ZYSZ14, ZYZJ24, ZGR23]. dissimilar [MAG13]. dissipated [GM08]. dissipates [MD22]. Dissipation [AZ23, BFQ22, CC23a, CSW19, CKM25, Leo10a, LFS21, RGÖS18, SRMDRL23, SGN06, VRK25, ZH21]. Dissipation-preserving [AZ23, CC23a, LFS21, ZH21]. dissipative [CW25, Gar96, HO10, JZS20, Rog19, WXY24, XLZ20]. dissipatively [HL99]. Dissipativity [Gan09, Hua00, WYL11, MDD14]. distances [CP07]. distorted [WG11]. Distributed [SW20b, AD20c, AD21, ABG^{+}15, BGH08, BSZ15, BJTZ20, Eij95, KK20a, LR18a, LSGK15, LSY21, LQY25, LRE04, MRV93, MMM19, MRFF17, Per99, ROL19, RZ18, XL11, XZZ19, YRV21b, ZZX20, ZLL22, dv95a]. distributed-order [AD20c, AD21, BJTZ20, KK20a, LR18a, ROL19, XZZ19, YRV21b, ZZX20, ZLL22]. distributing [Mit97]. distribution [DNW18, Shy91b]. distributions [AEA23, BD07, DMGVPO09, Maj20, Rus95, SdSC99]. div [BVV09, Lee23, LFQH21, MV18, XYZ24, YZ21, ZZ24, MD23b]. Div-Free [MD23b]. div-grad [BVV09]. div/curl [Lee23]. div/curl-systems [Lee23]. divergence [CHSS01, GDS^{+}15, HZ12, HSF25, HS97, KFR25, LD25, Mat08, MP15, MD21, MD23b, MWYZ18, Pas91, Sin23, WaZW23, YZ21, YZ22, ZBD24, ZFZ19, ZMF25, ZZZX26]. divergence-free [HZ12, HSF25, LD25, MP15, MD21, MD23b, Pas91, WaZW23, ZBD24, ZFZ19, ZMF25, ZZZX26]. divergent [BW15, CFV10, Wen10a, Wen10b, ZJ10]. diverse [DCN^{+}19]. diversification [AD99]. diversity [DESZ25]. divide [DH94, Jes93]. divide-and-conquer [DH94]. division [FT96]. divisor [BWY17, BWS21]. DKLAG6 [CST97]. DLN [QCW^{+}23]. DMLPG [MS13]. document [MMBB07]. Domain [AF04, Bog00, BS91, CES91, IVA93, JM06, KG90, Kop89, KJ99, LY08, ML91, Pas91, Pav00, TMS87, ADG^{+}24, AMR14, ABCC18, ABG^{+}15, BM89, BRVC09, BWS21, BP90, BH12b, CG89, CM14, Chn17, CDG19, CRR03, CG16, DF11, DDZK05, De 93b, DDS89, DHL00, Dor91, Dou91, DY03, EZ98, Ewi91, FWL18, FL05, Fat12, FPPS00, Gas92, GK09, Gen10, GCP91, GD22, GHF00, Haa97, Hab08, HZ09, Har98, HSS07, HLJ20, Jun97, Kok08, KRBK16, Kuz90, LLL08, LY09, LX21, LZZ25, LR00, LOM98, LS05, MAHZ21, MS86, MH14, Meu91, Mit97, MG22, MLB97, MVBS25, Mul19, NX22, NRWF08, NTT22, NMB10, Par04, PR09, PR90, RAS99, RV04, RV05a, SH10, SWW17, Ste05b, SM89, Su94, SW20b, Tse00, VBVA22, WL18, WJW19, WJM22, Yu99, YDWW17, ZPT92, ZS21a, ZHL03]. domain [Zho17, Zho18]. domain-based [Har98]. Domains [EHV19, AEN22, BN12, BBD24, CL08, CM02, CSX23, DS17, Dun18, FID18b, FMS18, Fer09, FL15, FL20, FBM17, FS24, yGyZ07, GWZ25, HHYD20, HZBM05, HP18, HY01, HS21b, JT06a, KDK17, Lau17a, LHWF08, LZH19, LC25, LZIZ23, MPTT17, Nke07, Pec09, PS19, Roz05, RA09, RS00, RU15, SFJ^{+}05, SvdVvD06, Tsy98, TY00, Wal00a, WS04, Yos00, YLW21, Zha14, ZLW22, ZZL17]. Dominant [LXCM21, DB95, SMJ12]. dominated [BC08a, CP05b, GAML04, GZZ20, Hin95, OEAS21, YXF25, ZMF25]. Döring [DS01]. Dormand [EL97]. Double [MZXX24, SD22b, BHB23, ÇK13, DS97b, EEAS25, GL17, IAH25, MK14, MCS06, MMD20, PG21, Wel10b]. double-diffusive [ÇK13]. double-fractional [MMD20]. double-sided [GL17]. doubly [BW06, KW20, PR22, SKO19, WKP12]. down [CH04]. downdating [YK04a]. Downwind [BW97]. DP [DHS05]. DPMHSS [CDW19]. DPMHSS-based [CDW19]. DPS [SD22b]. drag [RGK21]. drift [ACLM22, BK12, BSTT22, DR09a, Fou00, GHW20, GT19b, LSW23]. drift-diffusion [BK12, Fou00]. drift-implicit [DR09a]. drift-oscillatory [ACLM22]. driven [AKT97, Bac18, BTBR20, DS21a, DHM09, Guo00, HHW18, HJ21, HLY22, Mik97, MŠ99b, MT20, RSK24, RBC02, RTT01, ST19, Sob24, TWD23, WYY20]. driven-cavity [Guo00]. Drug [BCFR25, FGP23]. Drummond [BRZ10]. dry [BDMG12, DEGDdL25, WC24a, WC25]. DtN [CAAT16, LLN25]. Dual [CLS04, GHT05, BCFQ19, BCFQ21, BSZ99, BSZ22, CHLA21, CW21, Car09b, CCK03, He24, HT20, HMW05, MDT05, NYPW21, NRR06, SZ22a, SZ99, TDW23, WL25, ZZX19a, dRT99]. dual-phase-lag [BCFQ19, BCFQ21]. dual-porosity-Navier [CHLA21]. dual-porosity-Stokes [BSZ22, SZ22a]. dual-power [NYPW21]. dual-species [CCK03]. Duality [BBW19, DE16, DL21a, LRS23]. duals [MB08]. duct [Gla94]. Due [PGM86, BMP05, Fai00]. Duffing [EK95]. Dundee [Ano02g]. dunes [FF20]. Dupin [MMP02a]. DY [LWLW24]. Dynamic [HL89, MC00, Ahn07, AZ23, AC18, AGQ^{+}24, BBD18, BS94b, CS01, Chr96, CMMR23, CF13c, CF14, DS07a, De 06, DBH^{+}05, EW08, FLÖ^{+}97, GL17, HS07, HLR18, Ito17, Ito22, KR25, KW12, LA12, LZ13, MW24b, OHA25, OXW25, RP01, RN04, SMC08, SYL^{+}20, Shy91a, Shy91b, SM89, Tan23, UWY22, Wee01]. Dynamical [CLT97, Con20, Jéz04, Arn95, Bai02, BM06a, BCF^{+}13, BS24, CH95a, Car94, CCP04, CYYH21, DV95b, GEGG^{+}20, Hua00, Ise94, LBCN00, MR06, Pul12, Ram96, RW87, Sch95a, SD25]. dynamically [CPY20, Tho85, TLV92]. dynamically-adaptive [Tho85]. Dynamics [MOZ87, SH97, ZB07, AZHD23, ABdSG23, AB12b, AS00, ADH00, Aya09, BASC17, Buc04, CKP15, DG96, DL22b, DM11b, DC18b, Elm02, EK97, FL05, FXY22, FJL21, FMP04, FJ95, Gus87, HCS20, IKMM23, JRT90, Leo10a, LW22, LSGK15, LY24b, Lub04, LS05, MCBV20, NRWF08, Pow94, Pul86, Rob10, RA17, SB03, SD93, Tad86, TK19, Tur93, VS91, VT91, WPAZ24, ZML^{+}12, dPT96, FZM20].

each [GGG16]. early [AFIS24, ZGO12, ZO14]. early-exercise [ZGO12, ZO14]. early-stage [AFIS24]. Earth [Nak05]. earthquake [FKA^{+}13]. Economical [BM06b, CN11, WH19b]. economization [Som86]. eddy [AGLRS23, BKP14, LCVG01, MS08b, RV09, Sel14]. eddy-current [RV09]. Edge [CS04, BG06, CXZ09, KM95, OK98, PPS05, Sou09]. edges [JGK11]. Editorial [Ano87, Ano11a, Ano18g, Ano18h, BRW21, GW02, Vic05, Ano91, Ano93, Ano03e, Ano03f, Ano04d, Ano04e, Ano04f, Ano04g, Ano04h, Ano04i, Ano04j, Ano04k, Ano05d, Ano05e, Ano11b, Ano11c, Ano11d, Ano12a, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano12g, Ano12h, Ano12i, Ano12j, Ano12k, Ano12l, Ano13a, Ano13b, Ano13c, Ano13d, Ano13e, Ano13f, Ano13g, Ano13h, Ano13i, Ano13j, Ano13k, Ano14a, Ano14b, Ano14c, Ano14d, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano15k, Ano15l, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano16k, Ano16l, Ano17k, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i, Ano17j, Ano18a, Ano18b]. Editorial [Ano18c, Ano18d, Ano18e, Ano18f, Ano18i, Ano18j, Ano18k, Ano19a, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l, Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano21a, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano22a, Ano22b, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano23a, Ano23b, Ano23c, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano24l, Ano24a, Ano24b, Ano24c, Ano24d, Ano24e, Ano24f, Ano24g, Ano24h, Ano24i, Ano24j, Ano24k, Ano25a, Ano25b, Ano25c, Ano25d, Ano25e, Ano25f, Ano25g, Ano25h, Ano25i]. Editorial [Ano25j, Ano25k, Ano25l, Ano25m, Ano26a]. EEG [LWD^{+}09]. eeKdV [NK24b]. ef [Car23]. Effect [CRTU15, LL23, LR03, AD15, AM16b, ADH00, BKM13, BSQ96, EK96, For25, HD04, Spi95, WDZS21, dv95a]. Effective [CG92, LHH08, LH09, NCYC22, SPS20, AA87, BS96b, But98, IB24, LS12, NLS18, Ror06, RT95, TSFB01, Tsy96, ZN21]. Effects [BO87, GT19a, FI03, IAH25, KX03, SWJ^{+}24, TS23, vB95]. efficacy [IKS25]. efficiency [CZ90, Fuj02, Lay09, MNSS22]. Efficient [BF15, BHSW20, BMR17b, BMT25, BIM15, CC23a, CC24, CHOR19, CH01, CJ24, Deh05, DMH18, DB08, DK14, EH06, FTB97, FJL21, Gem23, GKB^{+}22, GAH25, GLW^{+}25, HLMP09, Has20, HC22, HLJ20, HCGW22, HST14, Kwa09, LV12, LAZ20, LD21, LS86, MAH18, MCS16, MR01, NS20, SS08a, SWL20, SKO19, SSKS21, WMF17, WT20, Woź10, XF22, YK04a, YV17, YLFT20, YS24a, YLW21, YW24, ZXYW22, ZGO12, ZCY20, ZLW22, Zou11, de 95b, AG05a, AGJ12, AA20, AGQ^{+}24, AEN22, BD85, BMR^{+}17a, BvG19, BCET24, BS97b, BJ20, BJM01, BZ17b, CFX08, CJL13, CKB13, CSL^{+}24, CJ25, CKK25, CGPT19, CPZ17, DR09b, DA18a, DZW24, DII15, Elg17, FXCW21, Gla93, Gla94, GPHAM12, Guo26, HBA^{+}24, HJR22, HAN23, HKZ08, HMN20, HS86, HOS11, HJ21, Hua98, Hua17, IR22, JZXJ21, JJL^{+}24, KK20a]. efficient [KCY19, KZ21, KGZ24, LWT07, LTC03, Lei02, LRC19, LWW22, LL23, LTDY24, LL20b, LBCN00, MKDC25, MVVA09a, MMD20, MK19, Nes16, NMB10, PB21, PRS23, QNA23, Ric08, ST14b, SWR11, TS24b, Wal19, WCM23, XY25, XGHM22, YK07, YWW23, YLLZ21, YW25, ZH21, ZY19, ZZF25, Zha01, ZLL22]. Efficiently [BC89b, WW19]. EFG [DA16, DA18b]. EFIE [LCJQ12]. eigen [LYC24]. eigen-problems [LYC24]. eigenbasis [GT15]. eigenfunction [BMV19]. eigenfunctions [Bie12, CF18, HM09, Pru00]. eigenmodes [LLL08]. eigenpairs [ARY23, Bai25b, Jia02, YSBL14]. eigenproblem [JP93]. eigenproblems [Jia00, Pet92, RA09]. eigensolvers [AG05b]. eigenspace [LfX15]. eigensystems [Tan87]. Eigenvalue [Nak12, ZGL98, AGM09, AMN24, ARY23, ASZ18, AG05b, AC08, AP08, Bai25a, BGO13, BNH01, BLY16, Bou16, BMV19, CL08, CJX11, CWP21, DDP12, FdSB02, Gen10, GGO12, GGO16, GCP91, EJR25, Jes93, JLZ20, Khe91, Kim19, LRV25, LVfP14, LB23, LX09, LfX15, LB21, Mat05, MZM20, Men23, PL25, PV93, SLJ11, Sol15, SL01b, TM15, TC19, XC20, XYHM20, YLL09, YSBL14, YXX19, ZBY19]. eigenvalues [An16, AT13, BC99, Bie12, BG14, BMV06, CF18, Con99, Ghe97, HC01, LfX15, Mor05, Ore93, Sch23, YBL13, ZLHW19]. eigenvector [BGO13, GGO16, XC20]. eigenvectors [ADNR21, HC01, LfX15]. eight [Wen98]. eighth [EL97]. eighth-order [EL97]. eikonal [ACMR06]. Einstein [EHNR24, EEJB22, ZB07]. Elastic [CLP15, Abr93, AMV17, BFQ22, CFKS07, Che88, DG96, DT89, FJH^{+}01, GWZ25, HS07, IT07, KK06, LH11, LDIW16, LB23, LR01, LR03, RL06, RLHC19, SWY^{+}23, Sim98, WQ17, XC85, YR22]. elastic-visco-plastic [HS07]. elastica [Ito17, Ito22]. elasticity [BF01, BBG14, BB24a, BRS05, BLJ21, BCV21, CGS19, CHX13, De 93b, DN13, HH18, LHC09, LH21, MOS12, MG00, Por17, SS99, Wan05, XZL07]. elastodynamic [DPPR16, DS15]. elastohydrodynamic [AGJ12]. elastoplasticity [BBS25, BS02, SW11]. elastostatics [AOW94, FMGN94]. electric [AGLRS23, Dav98, ESS15, GR93, LW07, ZZ25]. electrical [BBS11]. electricity [CGEV19]. electro [FPS15, NS16, RV09, RGK21]. electro-hydrodynamic [RGK21]. electro-magnetostatics [RV09]. electro-mechanical [FPS15]. electro-stress [NS16]. electroelastic [MZ04]. electroencephalographic [KRBK16]. electroencephalography [FOMC05]. electrohydrodynamic [IR22, Rou20a]. Electromagnetic [DRC85, CM14, Du11, EBEHAS24, EH05, GAW09, Hab08, HM86, IJ14, LSV22, LGH11, NT16, Par14, PB21, Ric08, WN12, WWLS08]. electromagneto [BGM^{+}09]. Electron [DRC85, TMS87, CPD^{+}05, PCA10]. electronic [BLW02, SW06]. Electrophoresis [VK17]. electrophysiology [CK22, NK11]. electrostatic [HNP17]. Element [BS14a, BW86, BTDV10, CF86, Rei85, ZW87, AD19a, AD20b, ABZ21, AMRR18, AP20, ANN19, AMN24, AL09, AMT13, ADG^{+}24, AS97, AK00, ADK94, AMH24, ATW20a, ATW20b, ATAW25, AA87, ADFR18, ADM22, AJ24b, AKL08, ABR05, ACP23, Arc06, AM16a, AMH03, APJ09, BSGU94, BC12, BKK25, BHJ05, BP14, BPS19, BY00, BF01, BK17, BR25, BLY17, BM01, Beh97, BB15, BLJ21, BC08a, BCGS24, BMS89, BC08b, BGH08, BKP14, BDF94, Bla00, BC01, BBD08, Bof06, BG14, Boh03, Bra00, BM18, BJM01, BJTZ20, BD11, Bür12, Bür13, BL08, BH12b, BD22, CGS19, CHP19, CY23, CLT97, CHLA21, CW21, CDGA17, Car09a, CGG25, Cau08, CS94, ÇD17, CGRT18, CCZZ18, CH01, CX01, CC04b, CZ04, CH07, CXZ09, CCLT10, CL10, CJX11, CHX13, CSXL14, CZHX19, CWHF19, CYWH22, CWZ23]. element [CL24b, CLL25, CPC25, CL18, CK20, ÇK13, Cod08, CGK25, dCCSR03, Cop03, Coy12, CH25, DPPR16, DZZZ24, DV20, DGM22, DB97, DA16, DA17, DA18a, DA18b, DA19, DSA20, DGE22, Dek17, DK21, DTGN23, DDS89, DM11a, DLP06, DN08, DL16, DCL23, yDqGnJT09, DYF23, DL06, DR93, EH07a, EBEHAS24, ER07, EFLFP09, FID18b, FS15, Fat12, FLH22, FMGN94, FLÖ^{+}97, Fre98, FJS99, FWW^{+}21, GAML04, GM10, GP23, GIS23, GHH20, GH25, GAW09, GLPW09, GD21, GGO12, GGO16, GP01, Gol86, GHT05, GCZZ23, GGG16, GGRN17, GGRBRG22, GS94, Gwi09, HLZ14, Har10, HKZ08, HLMP09, HSS04, HH22, HL08, HS20, HD22, HTSZ23, HGF24, HCJW24, HO24a, HS22, HMdV03, HL02a, HL23, HL03, HvdHV10, HW97, HCY18, HJYL19, Hsi06, HLC01, HZ02, HZ12, Hua17, HJX^{+}19, HAC22, Hus20, HSY18]. element [JPP19, JJ94, JP08b, Jeo09, JK14, JCL18, JLZ20, JWZ21, JEG10, Joh01, Kam16, KDAK13, KTD20, Kar89, Kie17, KK09b, Kim12, Kim21, KJ99, KFOF02, KZ13, KK22a, KN93, KDD23, Kur98, KX03, Kwe01, Kwe03, LHC09, Lam13, LCHR03, LP05, Lee23, LRV25, LWD^{+}09, Li98, Li00a, Li01a, LHH08, LH09, LA11, Li12, Li16, LHW17, LLHC18, LW19a, LL20a, LSP20, LW20a, LWW20, LCHW20, LLY21, LH21, LL23, LL24, LWaZ24, LLZL24, LD25, LGC25a, LGC25b, LP25, LR24, LX09, LLS^{+}96, Liu09, LCLW17, LC20, LN21, LZIZ23, LY24a, LS24b, LD22, LT19, LAH09, LL02, LW95, Luc95, LZY09, LJ20a, LJ20b, MD00, MK14, Mai06, MM22, MCS06, MZS10, MP20, MZM20, Men23, MF23, MH16b, MD96, Mou03, MWYZ18, MC21, MT20, Mul99, MM20b, Nak05, NLHT25, NY13, Nke07, Nud24, NS16, OL18, OH20, OEAS21, PL25, PZMX16, Pec09]. element [PSWZ21, PQ25, PP00, PT95, PS21, PS19, Qi24, QW25, QMLC15, Quy19, RN25, RZS21, Ran20, RSK24, RS20, Ric91, Ric94, Ris05, Rob10, RV15, RX08, RBT15, RA09, SSZ16, SS00, SGS00, SD11, SD13b, SS19, SH10, SQ17, SC25a, SWX00, SWY^{+}23, SA00, SJ20, SL20, SW20a, SW21, SG07, SD09, SCLL21, Sol15, SW20b, TLP18a, TLP18b, TWMP20, TN16, TC22, TW00, TM21, TTP24, Top21, Tou25, TM04, TM05, VO00a, VNC21, VMS07, WZL13, WCSQ18, WSY18, WZ19, WaZW21, WY22, WL22, WaZW23, WLY24, WK00, Won08, WYYL19, WWH24, WWWW25, XZW19, XL23, XL09b, XZH19, XHYM22, XP23, XYZ24, YGY15, YLL09, YS09, YY13, YSBL14, YQCZ22, YYZ23, YZ24b, YT25, YZZ25, YZ21, YCWH23, YZ19, YHT23, YW24, Yua20, ZBD24, ZL17, ZLY23, ZOZ09, ZCZ15, ZSG^{+}20, ZLW20b, ZL21, ZL22, ZZPJ23, ZSS23]. element [ZZ24, ZYZJ24, ZZH25, ZWG25, ZZZX26, ZZZ19, ZPT92, Zhe07, ZWK15, ZZ17, ZZ20, ZLWF21, ZS21a, Zho17, ZZC^{+}18, dFN00, dSFDG20, van95, BTDV10, CLR11]. element-free [CPC25, Li16, LL23, LD25, ZOZ09]. Element/Spectral [BTDV10]. Elements [CHM09, AW03, Arc06, Aso21, BHJJ06, BBS25, Ber04, Ber05, BG06, Bot97, Bra00, BL08, BH12b, CHZ21, CZS04, CYM09, DDGN23, EJS11, FM11, Fun94, HLMP09, HDS20, HLY04, HLZ06, HMW05, HJ03, Jun06, KV20, KOS^{+}12, KX03, LPT94, LM00, LH02, LMWZ10, ML91, MM02b, Pav00, PT23, Pic05, SV24, ST08, TW00, Tro96, TMM15, VRK25, WB90, YBL13, Yi12, Zeg97, Zha09, dSFDG20]. eliminated [Mau08]. Eliminating [ACM91]. elimination [AGP97, GP93, GY94, YH18]. elliptic [ANN19, AyLqW18, AK09, AEMX17, AB09a, AD23, AMR14, Ars20, AC16, Bac21b, BKK25, BS14b, Bar09, BGG04, BGH08, Bog20, BHR05, BMV19, BSZ15, BH97, CY23, CL08, Cas06, CSS87, CES91, CC04b, CZ04, CYM09, CZHX19, CDW23, CWZ23, CL24b, CL26, Che12b, CL20, DDHS97, DT15, DS17, DS21c, DDF24, DP90, EH24, EP15, EH08, EGH01, FG98, FLR08, GSR00, GRLL01, Gon06, GGR97, GFPG18, HZ21, Har10, HP18, HM01, HD22, HCJW24, HO24a, HO24b, HMdV03, HOS11, HLW25, HSY18, JP08b, JS09, Jeo09, JK14, JCL18, KL98, KG90, Kim07, KS09a, KwS19, KS07, LRS23, LHHR94, LHH96, LH02, LMWZ10, LX09, LY03, LAH09, LL02, LP01, ML16, MM02c, NY13, OS08, OT22, PL25, PAP17, PZ20, PS21, RTV00, RTV02, Roz05, Sch16b, Sch87, SNOK21, SD13b, SSR23, Shi20]. elliptic [SD09, SZL18, TL07, VO00b, VSeYD02, Wan07b, WCXL09, WCW14, Wan17b, WCSQ18, WZ19, War92, WKN20, XZT21, XHJM21, Xu21, XHYM22, YGY15, YSBL14, YZ24b, YSZG25, Yos00, YXB95, ZTZ15, ZWL11, ZX22, Zou10, Zou11, dSFDG20]. elliptical [BDSG09, DSV13]. elliptical- [BDSG09]. Elsässer [AEK23]. EMAC [CHOR19]. Embedded [KNT13, BDGP96, BDP99, BCT19, vR04, Kam16]. embryogenesis [MSS^{+}15]. Emden [SSC23, AY21, Cha25, DSAB20, ER18, GS24, Güm20, IB24, LO03, RTA19]. emphasis [GR02, HDS20]. empirical [LZZ18]. enclosed [DS17]. enclosing [DMH18]. Enclosure [RBBC85]. enclosures [RNG22]. Encoding [DMM24a]. end [Agu15, Wel10b]. end-point [Wel10b]. endgame [WWS^{+}93]. endpoint [ZMY21]. ends [Ito17, RL06, RLHC19]. Energetic [LY24a]. energies [FJP17]. Energy [AFF^{+}15, BUL23, BMWH20, CWY20, EFLFP09, GGT24, HS20, LW19b, Li22, UWY22, WJW19, WWZS25, ZHL08, AGK24, AG25, Bar12, BFGP08, BCCHM21, CHP19, CFLW22, CCL04, DGE22, DL20, DW21, DL22b, FXCW21, GM08, GLML20, HTSZ23, HMD21, HL21, HWZ22, HL24, HCGW22, Ito17, JWG20, JJL^{+}24, KWY24, LMY18, LCW20, LW22, LL20b, LC21, LYA^{+}19, MD20a, MD22, DE18, MMDH19, MMD20, MMDS21, NK24b, PHY19, QH22, Qi24, QZH25, SRMDRL23, SL21, TZ21, Wan20, WLY24, WWZ25, WKWC26, XXYZ24, XLZ20, XWZ21, XCHW22, YZH19a, Yan23, YF24, ZCY20, ZFS24, ZZF25, ZH24, ZFX25]. energy- [DW21]. energy-conservative [CHP19]. Energy-conserving [UWY22, HTSZ23, HMD21, MD20a, MMDH19, ZFX25]. energy-optimized [ZZF25]. energy-preservation [DL22b]. Energy-preserving [BUL23, HS20, LW19b, Li22, DL20, FXCW21, HL21, HL24, JWG20, LW22, MMD20, WKWC26, YF24]. Energy-stable [GGT24, QH22, Qi24, WLY24, XXYZ24, Yan23]. energy-transport [BCCHM21]. engine [NER95, NR97, SM08, van98]. engineering [SW95d]. engines [Fer14]. Enhanced [BB94, CCC08, GT00, Aff94, CHX13, SBS24]. enhancement [BGO13, WZ02]. Enlarging [Zha07]. ENO [AMR12, BL06, HS98, KS09b, SZE^{+}92].

ENO-discretizations [HS98]. enough [Enr06]. enriched [CGRT18, CL20, DDF24, OEAS21, PT23, WCSQ18]. enrichment [DTGN23]. enrichments [Nud24, WLZ25]. ensemble [BGM19, JY23]. ensembles [AB17]. ensures [Rob01]. Enthalpy [Jes85, Ros93]. entire [Dés08]. entropy [AJ24a, CDI91, CSW19, HJP10, SWJ^{+}24, Tad86]. environment [GGNP02, KLS13, WG19]. environmental [DMP08]. environments [LS99a, RN04]. epidemic [CHM22, CT21, MPV24, TSF24]. epidemics [HJ06]. epitaxial [KWY24, SL21]. epitaxy [CZY18]. Epsilon [Lot19]. equal [GHH20, KOS21]. equal-order [GHH20]. equalities [MMP09]. equality [LDP^{+}14, LJYS20, LCZ21, ZP12]. Equation [Ant13, PB10, PGM86, AS11, AD19a, AD20b, AD20a, AD20c, AD21, AZHD23, ABJ12, ADK94, AS04, AMK18, AMH24, ATAW25, AÁ21, AHO16, Ale11, AT15, ADFP25, AA24, AZ23, ACMR06, ABKG21, AB10b, ALZ^{+}21, Ang06, AM10a, AD01a, ABR05, AR15, AJK20, APJ09, AM04, AED12, ADH00, ÁMS17, AEN22, Bac17c, BF92a, BS06, BOEP00, BCS17, Bas21, BM01, BGT97, BCM04, BMGGG12, BN12, BM04c, BBC^{+}25, BLM17a, BK21a, Bho12, BDKM92, DSM22, Bor02, BDN^{+}97, BIO24, Boy91a, Boy07, Bra22, BFdO07, BP92, BPTT15, BR97, BMWH20, BC97a, BB98, COPPS25, CGT13, Cai09, CHP19, CC23a, CGH23, CFS13, CGP15, CNA23, CCG13, CM09, ÇD17, CHX26, CCZZ18, CjW18, CJL13, CJ18, CEL25, Che88, CH01, CDW13, CHH15, CXZ17, CWX21, CLX21, CCST22, CL24a, CQZ20, CWY20]. equation [CG21, CLP15, CF13b, Cho13, CCL04, CCK08, CK20, CRR03, CMMR23, CH89, CFL^{+}20, CGK25, CBHM19, CRSF19, CH25, DD21, DKSS24, Dav98, DS07a, Deh01, Deh05, DA16, DA17, DA19, DS20, DS21d, DGE22, Dek17, DL22a, DL22b, DS97b, Din19, DSK12, DCY20, Duj09, DAMA23, DC18b, EZ03, EEAS25, ESE20, EVO04, EVO06, Est95, FWL18, FP02, FRRJT10, Fdi96, FWHM20, FT06, FXY22, FL09, FL15, FvdMS20, FI03, Fil25, FG13, Fre98, FFQ09, FCW20, FXCW21, FCW21, GZZ19, GZQS23, Geb24, GGN12, Gia12, GLLY25, GWLN22, GGT24, Guo96, GGO13, GD23b, GAX24, HBA^{+}24, HT24, HZ09, HAN23, HP14, HM15, HM17, HAML21, HMN20, HILK13, HLT07, HPH20, HDY21, HS95, HW04, HMY19, HJKW17, HJL18, HvdHV10, Hor99, Hor02, HS23, HZZ25, HK85, HCGW22, HLY04, HLZ06, HS19b, HAY20]. equation [HAC22, HO16, Ise97, Isk89, IKM23, JPP19, JR00, JRW06, JP19, JT18, JL25, Jéz99, JZS20, JZXJ21, JHGZ20, JWG20, JL21, JQSC22, JL23b, JLZ25, JL24, JLWY25, JR02, KMS19, KDAK13, KOS21, KBS11, KAS22, KO96, Kha21, Kim95, KCY19, KRBK16, Kra92, KK86, KZ21, KTYY24, KK23, KR15, LO22, LO23, LMPS19, Lau17a, LCK22, Len00, LHWF08, Li08, LHW17, LLHC17, LR18a, LZ18, LR18b, LWCH19, LZH19, LRC19, LSWM19, LSP20, LCW20, LR20a, LW21a, Li22, LWYG22, LXZS22, LHC23, Li23, LH24, LZ24, LC24, LLZL24, LYZF25, LGSW25, LSG24, LD97, LO03, Liu97, LCLW17, LCL18, LL20b, LSY21, LYMD24, LCZ25, LXPY25, LP01, LMWZ07, LZY09, LYA^{+}19, LR20b, LJ20b, Lyn99, LLW20, Ma03, MZZ17, MDP10, MDRR11, MD22, MB10, MPTT17, Mai09, MN23, Maj20, MP11, MZXX24, MR20, Mar93, MVVA09a, Mar05]. equation [MG97, MM20a, MYSC17, MW24b, Mit22, Moe98, MG22, MAD23, MPDB24, Mul99, MPMD21, Nak24, NP21, NYPW21, NAD25, NTHC21, Ngu15, NSCC19, NSD23, NK24b, NTT22, NMB10, OMP98, OL18, OAHN22, PXHZ20, PWY21, PG21, PR89, Pau92, PR90, Phi91, Pło22, Pło23, PRS23, QH22, Qi24, QWX20, QNA23, QXG21, QZM25, RZ18, RSR23, RV22, Ric94, Rob01, RT14, RGMO19, RREP^{+}20, Rou20b, RMK09, RS00, SS08a, SRK21, SRK22, SN22, SC19, Saz24, Sch23, SS94b, SRMDRL23, SLW17, SY18, SJ20, SL20, SW21, SD22a, SJ18, Shi20, SP22, SZQH23, Sim91, SG09, SMW21, SNW22, Sin23, Sin24, SPYS24, SS17, SDK15, SR24, ST25, SA18, SL15, SSA24, SG92, Su94, SWR11, SW17, SZW19, SND21, TWMP20, TN16, TS24a, TMD92, Tar98, Tem23, Ter22, TOD11]. equation [TDMT21, Tia15, Usm97, Van92, Vas17, VRC21, jWyG08, WL10, WN12, WZW13, jWjJ17, Wan17a, WQ17, WH18, WH19a, WMLB19, WT20, Wan20, WZZ21, WDU21, Wan21, WWL21, WCS21, WCM21, WL22, WCL22, WLY24, WW24, WWZS25, WTF25, WLCZ26, WC14, WW14, WL24, WS04, WSC21, WG23b, WWWW25, XXYZ24, XW19, XC85, XZL19, XFL22, XL23, XZZ19, XLKY19, XCHW22, XXF22, XP23, YZQ^{+}22, YX25, YJZ18, YLY19, YLFT20, YLX21, YXN21, Yan21a, YJ21, YS22, YQCZ22, YZG23, YYZ23, Yan23, YZ22, YF24, YW24, YXZ18, Yua20, ZH21, ZRA23, ZL17, ZXYW22, ZZ18, ZWH^{+}17, ZJH18, ZLHW19, ZB19a, ZZ19a, ZJ19a, ZYLL20, ZLCH20, ZYQS21, ZG21, ZLSZ22, ZYJZ23, ZZPJ23, ZD25a, ZY25, ZM25, ZJ25, ZZX19b, ZZX20, ZYX20, ZLL22, ZZjQ24, ZLG24, ZML^{+}12, ZZ17, ZZL17, ZZ20, ZEW20, ZJLA22, ZZC^{+}18, ZJ19c, ZZLL21, ZFX25, ZXW17, ZR15, ZCGS21, dVA02, iV09]. Equation-based [PB10]. Equations [BGHR12, BGH^{+}15, CP09, Jac87, Pet87, Rei85, ST86, AKM^{+}21, AS21, AY21, AHJM19, ABH22, AZA22, AN25, AS13, ABZ21, ALMM96, ALMM98, AGZD22, AKM^{+}22, Abu04, AJ24a, AAL21, AQS94, AB88, AS20a, ASA20, AMT13, AK09, ADR17, All24, AAH21, ADM22, ACLM22, AM99, AM00, AMCM08, ASS21, AAL^{+}26, AKBF19, AHR12, AX19, AHB20, AT93, AAM03, AB07, AL17, AB24, AW03, AL22, AL24, ABF09, AR18, AMP20, AF89, Arn93, ARS97, ALP^{+}96, AKS21, AAD14, AD18a, AD18b, AD19b, AM10b, AAEMY21, AFF^{+}15, AO91, AV00, AS20c, AKA19, Aze22, BSGU94, Bac16, Bac17b, BTBR20, Bac21a, BKM19, BL21, BUL23, BHJ05, BHJJ06, BKM95, BBPR05, BP06a, BF20, BL05, BKM13, BLW07, BES18, BBCR22, BKP09, BF17, BJ05, BBJ25, BGS06, BLN25, BLY17]. equations [BRSD91, Beh97, BNKR20, BZ92, Bel97, BO04, By01, BB15, BHSW16, BHSW20, BG11b, BGGG13, BG24, BCGS24, BF92b, BW95, BG02a, BZ17a, BCC16, Bic21, BG03, BT19, BLM17b, Bla01, BC01, BtTBV87, BDFV95, Bog16, Bok03, DLS22, BVV09, BTC23, BFH09, BBLT15, BJ11, BHJ13, BS20a, BCGI13, BTDV10, BVB10, BVRB14, BRBB18, BDE22, BR20, BFdS10, BMPR15, BJ20, BM18, Bre06, BLL24, BC23, BS93, BS08, BK12, Bru92, Bru97, BMM97b, BMM97a, Bru07, BL06, BJTZ20, BO21, Buc06, BH97, BS18, BS20b, BB24b, BIS25, BDM03, Bür12, Bür13, BCS06, BIMV19, BC89b, BB96, BD22, But92, BC95, BJ96, BJ98, Cah92, CFXZ06, CXNF14, CLT97, CC24, CC18, CGEV19, CdFN01, CC90, CGA96, CL01a, CP04, CHLX07, Cao10, CHZ14, CCP17, CDP17, CDP19, Car23]. equations [Car19, CKPS15, CD95, CSS87, CCM02, Cha25, CHOR19, Cha98, CDH24, CJ90, CT93, CH01, CL07, CXZ09, CZ12, CL14, CSXL14, CHH15, CLTA18, CC19, CLY19, CSLY19, CGW20, CYYH21, CW22, CDW23, CWC24, CL26, CS09, CL01b, CRU15, CR19, Chi12, CC20b, Chi93, CL18, CBD16, CCZ22, ÇK13, ÇY22, CGS20, CCJ99, CY05, Cod08, CNS00, CP17, Con20, CMPP24, CFM^{+}24, CGTTN24, CST97, CN11, CN15, CCM17, CG14, CP03b, Cui24, Cum95, DT15, DS21b, DD97, DS17, DG10, DS05, DR09b, DN21, DA18a, DSA20, DMH18, DK21, DJ10, DYX09, DM09b, DZ12b, Den15, DFLM19, DL20, DW21, DHWL22, Den93, Der92, DM11a, DCC14, DS97a, Die15, DLPV17, DSAB20, DKL24, Dob05, DB08, DBBH14, DLN04, DGRS09, DSSC13, DC09, DYZ20, DSS20, Dor01, DCL23]. equations [DMA22, DJL04, DS01, EAV16, Ehr08, ELCWS98, ESEKZ10, EGV25, ER18, EV96, ELLE02, EGL09, EMMK01, EH24, EH97, Enr06, EEE22, EK06, EH09, EHV19, EHV24, FM21, FK23, FID18a, FLS94, FS15, FMS18, FMMK01, FD16, Fan11, FJS25, FL93, FH20, FH22, FJ97, Fel06, FNT06, FCX06, FLH22, Fer09, FPR12, FJP17, FR18, FV87, FJ95, FJS99, Fre04, FSWZ19, FGGL22, Fuh01, Fun04, GS24, Gan09, Gan96, GS99a, GX11, GLLW14, GM16, GS19, GZZ20, GHHG22, GHS25, Gar96, GDEdLD23, GG19, GAH25, GMG19, GAZD25, Gol86, GJlL23, GLM09, GM17, GKA17, GNX19, GS25b, GGS16, Gu19, Gu20, Gug05, GGRN17, GNAS^{+}20, GAOB20, Güm20, GN86, GS94, Guo00, yGqWsWC05, GH07, yGyZ07, yGpY09, Guo15, GW20, GZHQ23, GWZ25, Guo26, Guo01, HGP11, HS21a, HP85, HN03, HM87]. equations [HHAA22, Han93, HZBM05, HO10, HZ21, HP18, HHT97, HJ05, HLR18, HHC08, HL08, HS20, HC22, HTSZ23, HCC24, HK25, HS86, HEG16, HZ20, HZD21, HMD21, HQAZ24, HJ17, HS19a, HS17, Heu00, Hey19, Hey20a, HAA21, HAR21, HBB25, Hey10, HK22, HL03, HLMKZ06, HXW15, HW15, Hop23, HA16, HDS20, HCY18, HL19, HJYL19, HL21, Hou23, HL24, HX11, Hu99, HLJ20, HJ21, HR96, Hua98, HLL09, HZ12, HFL12, Hua17, HZAT21, HS21b, HY24b, HSF25, HV95, Hus20, HN22, HL89, IM98, IKR^{+}22, IAH25, Iga85, ID19, Ise02, ITZ17, IB24, IRC12, IJ17a, JMDN^{+}22, JL91, Jac93, JVZ95, JVZ96, JVZ97, JZK06, JAH21, JK21, JRS20, JMS99, JKW12, Jes85, JWZ21, JW01, JEG10, JZZH22, JY23, JCY25, JCN94, Joh01, JMPY10, JT06a, Jun06, JOL23, JT02, Kal96, Kam25]. equations [KTD20, KL21, KV07, Kan89, KX91, KXK92, KNN03, KO08, KPY15, KS00, KT05, KME20, KV95, KK11, KKR24, Kau93, Kau97, KM25, KK20a, KM16, Kel85, KCL00, KC03, KC19b, KOR18, KSHB21, KO92, KLR25, KHB22, KDH20, KF97, KS02, KLY05, KS08, KHYY21, KS07, KS22, KOW05, KNT13, KSSS16, KOGL25, KKLD21, KCW16, Koz94, KCB02, KP15, KW20, KR18, KN93, KDD23, KGZ24, KKK25, KLSW10, KKR15, Kür23, KDKW20, LH11, LHHR94, LZQ22, Lan95, Lan97, LP05, LP24, LPS25, Lau17b, Lay09, LT12, LHN24, LN26, LS16, LLKJ21, DBH25, LHS00, LPZ00, LH02, Li05, LHH08, LY09, Li11, LA11, LZ14, LZ17, LYF17, LW19a, LCS19, LWY20, LW20a, LCHW20, LZJ21, LLY21, LW21b, LA21, LZ22, LS23, LYZJ23, LH23, LWW23, LL24, LWLW24, LD25, LGC25a, LWWZ25, LSL11, Lia22, LS12, LZW17, LZW19, Lie01]. equations [LPV24, LKJ07, LMSW17, LS13, LY16, LQS21, lLNW21, LS21, lLH25, LW92a, LW92b, LY03, LW07, Liu09, LZL14, LYK17, LLD18, LLZ19, LTT19, LHX20, LW20b, LMTW20, LZCF21, LL21, LN21, LLZ^{+}22, LM22a, LYZW22, LSWW22, LT23, LMW23, LYLL23, LCZ23, LSW23, LS24a, LNZ12, LM22b, LR00, LCM22, LCM24, LXCM21, LO95, LO96, LS98, LJ20a, LRE04, MDD14, MQO17, MM18, MK20, MAHZ21, MZN21, Ma24, MZJ^{+}25, MD19a, MD20b, MAH18, MHA19, Maj14, MKDC25, MD19b, MO17, MS19, MK21, ML16, MS08a, MKN23, Man96, Man97, MCS16, MWC21, MW24a, Mar99a, MVVA09b, Mär95, MS25, MZK05, Med96, MT06, MN24, MV17, MG18, Mie89, MOU14, MM02c, MSS21, MD10, Mit24b, MMM19, MFAD23, MG22, MH16b, MD25, Mok17, MKJ23, MDASAO21, MP94, MJS23, MN08, MRFF17, MWYZ18, MC21, Muo23, NX22, NK11]. equations [Nag22, Naj20, NMKE13, NDM20, NLS18, NLS20, NT92, Ngu15, NT16, NWL^{+}22, Nke07, Nor97, NLZB23, OT22, ORT24, Odi19, Odi24, OB24, ÖPT25, Oji88, OFY^{+}23, Olv92, Ort20, OT02, OS12, OCVW22, Pan21, PK23, PM05, Pas91, PP24, PM91, PT11, PTV16, PTV20, PV24, PD01, PGYF20, PA18, PVM22, Pet00, PP92, PH17, Pis22, PAJ12, Pot97, Pow94, PS19, PM14, PWX24, PYD21, PG02, Pul05, Pul09, PSL18, QLL^{+}08, QM10, QR24, QR03, QM20, QDA^{+}26, QMLC15, QAMX17, QM19, Qiu23, RR21, Rab94, RZ00, ROB17, ROL19, Rah25, RG20, RE19, RL21, RZS21, RA05, RR14, RSK24, RTW21, RG05, RV05a, RF16, RA03, RT20, RN22, RTV00, RTV02, Ric91, RU25, Ril92, RS08b, RLSS06, RAOC18, RK91, Rog19, RSK14, Roz05, RT95, SMTHE22b]. equations [SMTHE22a, SSW20, SSS^{+}23, SST12, SS25, SHL19, SSZ16, SA90, San89, Sar05, SZE20, Saz22, SH09, SKBAS08, SA12a, SS99, Sch98, SZ09, Sch87, Sch95b, Sea09, SB14, SOB20, SD11, SD13b, SG06, SH10, SK22, SC25a, SGS20, SV00, SDG20, She96, She00, SY05, SY07, tSqWyG16, SZ12, SWW16, SZ17, SC20, SW20a, SXL22, SL22, SWB20, SWB21, SD24a, SD25, Shi25, SSA^{+}22, SVB17, SP22, SLW24, SHB^{+}25, Sid02, SA20, SBS^{+}20, SG07, uIVS13, SS16, SS13a, Sob24, SvdHN86, SCvdH92, SW13, SD24b, SA19, Spi97, SMA01, STS00, SOA^{+}25, SSKS21, SCT05, SW12, Sti03, Str98a, SG05, SBG09, SS12, SWW11, SFZ21, SW24, SSPZ20, SLZ10, SG17, SSS21, SAMSB20b, Tad86, Tah96, TZ21, Tan93, gTpM07, TX18, TH18, Tan23, Tau09, TJK18, Tem15, TLG20, Tho85]. equations [TLV92, TB01, TYJ11, TZA13, Ton04, TS24b, TH23, TH09, TDPU17, TLSS09, TV91, Tro93, Tro96, Tsy96, TY98, UWY22, UNGD08, Vab22, VG04, VA05, VB07, VBD93, VV09, Van00, VCN20, VNC21, VGL24, VK17, VCC12, Ver96b, Vic87b, VL19, WTB24, Wan25, WGKS12, Wan01, WY02, WZL08, WL09b, WG10, WC11, WWX13, WZL13, WCW14, WGW15, WZ16, WMF17, WL18, WSY18, WW19, WH19b, WR20, WDH20, WTY21, WS21, WLM21, WHW21, WZ22, jWC22, WLG22, WWM22, WCM23, WXY24, WC24a, WWC^{+}25, WKWC26, WAV12, WaZ24, Wei09, WYL11, WHL19, Wen25, WB92a, WB92b, WB03, WPL16, WSS97, WG11, WPT19, WdG92, XFLC00, XK25, XL09a, XZW19, XY19, XLZ20, XF22, XZ22, XWX21, XL11, XWC25, XGQ20, XHJM21, XHYM22, XWM26, YXT17, YPD21, YMD21, YTC24, YZWQ25, YXF25, YH00, YS09, Yan18, YZH19a, YT21]. equations [Yan22, YJ23, YS24a, YT25, YZ21, YT00, YÇ16, YLLZ21, YC13, YXB95, YZH24, YLS^{+}09, YDWW17, YWSL20, YW25, YLH20, YRV21a, YRV21b, Yüz22, ZM19, Zak20, ZAED21, ZHS22, ZBD24, ZdBT03, Zha19a, ZG92b, ZH09, ZC10, Zha14, ZL18b, ZFZ19, Zha20a, ZLJ20, ZD20, ZFC20, Zha20b, ZYC22, ZHL22, ZQZ23, ZYQS23, ZSS23, ZL23, ZZ24, ZFS24, ZCQ25, ZWG25, ZMF25, ZZW97, Zha01, ZZO16, ZFX17, ZW19a, ZLX19, ZSJ04, ZLWF21, ZS21a, ZS21b, ZZ19b, ZSY20, ZSZZ20, ZP97, ZP98, ZX09, ZCSH11a, ZCSH11b, ZGDL17, ZC99, ZL24, ZAB15, Zou10, ZK00, Zuu95, dAF17, dH95, dDF^{+}94, de 92b, dG91, de 93a, dPT96, dOF20, dlHV13, ebKMZ24, in 92, in 95, in 96, iW07, iW09, iM13, tV87, vSW90, van96, vS97, vvdV97, vdHSW98, vdVS08, CDP12, RMM12]. equatorial [Fan19]. equi [BKR13]. equi-directional [BKR13]. equidistant [AC23, Bec18, DS97b, GÖ20, Kał22]. equidistributed [BM00, DRVA20]. equidistribution [BM01]. equilibrated [DGM22, Hop23, Yi12]. equilibria [Eir99, GPHA06, Sch02]. equilibrium [ACR25, BAA22, CZ19, DGM18, DMM24b, DJL04, GKKM21, ITZ17, TQY24]. equispaced [BS09]. equivalence [BC89a, LLD18]. equivalent [PSP04b, RMCG04]. ERES [CKM10]. ERKN [LW20b, SWL20, WWX13]. Ernst [CHM09]. Errata [Tsy96]. Erratum [Kni95, ZCSH11a]. Error [ADK94, AL17, An20, AD01a, AD01b, BS06, BJ02, Bel91, BC04b, BCJP18, CLY19, CC08, CL18, CBHM19, Coo89, DA19, DK21, DM11a, DHM09, GGM95, GAX24, Hig93a, HB20, Kre07, KSJY25, LL98, Li16, LYF17, LZ18, LH20, LS23, LL24, LGC25b, LM21, LW92a, LT01, LZIZ23, LCZ25, Mit22, MD23b, NS21a, OL18, OS12, PD01, PS19, QZH25, Rei85, RL06, Sch16b, SW11, SS94b, SC25b, SZ12, SG07, Spi96, SWCH15, Tem23, Ton04, jWqW09, WWX13, Wan17b, WZZ21, WWWW25, XZW19, XZZ19, YXZ18, ZLX22, ZCQ25, Zhe07, Zup04, AQJ18, AW14, AC15, AA25, AOW94, AGP97, AD23, AKT97, AJ24b, AGJM04, AR18, ASV19, AB12b, AP08, AM16a, Ars20, AAD14, AFF^{+}15, BLS94, Bac14, Bac16, Bac17a, BTBR19, Bac21b, Ban97, BGO13, BKM13, BS14b, BCS17, BM13, BBG14, BBW19, BRVC09]. error [BC04a, BLM17a, BCT19, BC01, Bor97, BHR05, BGP11, Boy15, BM06b, Bür12, Bur91, BP06b, CHZ21, CY23, CCOVF22, CP04, CL06, CDGA17, CGG25, Cha17, CMP03, CCS17a, CZHX19, CCST22, CDW23, CKK10, CL20, CKK25, CH89, CY05, CGK25, CF05, DDHS97, DDP12, DMP08, DA18b, DSA20, Dek17, DM09b, Dol14, DYZ20, Dra97, DLZ21, Dun18, EFLFP09, Fac03, FHM^{+}02, FV85, Fdi96, Fel06, FXY22, FSX25, FH10, FM11, GHH09, GMS12, Goo90, GHT05, GGRN17, GZHQ23, HBA^{+}24, HZ96, Han93, HOS11, HLMKZ06, HJKW17, HJL18, Hop23, HLY04, HLZ06, HH18, HAC22, HMW05, Ito17, JR02, Joh01, JM94, KDT17, KW21, KV20, Kim07, Kim12, Kim21, KS89, KMS10, KK22a, Kwe00, Kwe03, LCHR03, LSK12, LZ14, LLVX20, LW21a, LA21, LYZJ23, LC24, LX09, LY03, Liu09, LS20, LO96, Lyo12, MWC21, MW24a]. error [MN24, MSP10, MD25, Moo04, Neu88, NV23, NY13, OZ96, OXW25, PRSS24, PH17, QWW24, Roz05, SZ09, SS19, Sha21, Sim94b, ST25, The17, TTP24, TLSS09, TM04, TM05, VNC21, Wal19, WH13, WG19, WCM21, WCJ23, WC24b, WWZS25, WKWC26, WKP12, Wei18, XL11, XYHM20, Xu21, Yam18, YSBL14, Yan21a, YJ21, YS22, YJ23, Yan23, YT25, Yi12, ZCZ15, ZZHS18, ZFW20, ZQZ23, ZL23, ZZH25, ZM25, ZWG25, ZH24, ZYS17, ZEW20, Zho17, ZXW17, ZL24, Zou10, Zou11, Zup03, de 92a]. error-based [Moo04]. error-minimizing [Bor97]. errors [AC96, ACM91, BMGM12, BSQ96, Bor10, HV95, KRBK16, MKH16, RB12, Sid14]. ESDIRK [Ran15]. ESIRK [BC98]. ESPRIT [PT15]. essential [GS15b, LZ20, jWqW09, hYqW12]. essentially [DSZ15b, DSZ15a, HOEC86, PCL^{+}25]. estimate [AA04, AC15, Bel91, BGP11, CY23, CP04, Cha17, CCST22, CL20, DA18b, FP02, HBA^{+}24, HJKW17, HJL18, Ito17, Kwe03, LW21a, LS23, LM21, WZZ21, WCM21, WC24b, WWZS25, Yan21a, Yan23, ZEW20, Zho17]. Estimates [BG02a, Dun18, AGM09, AW14, AA25, ADK94, AD23, AJ24b, AD01b, AP08, Ars20, ABCC18, AV91, Bac14, Bac16, Bac17a, BTBR19, BS06, Ban97, BS14b, BJ02, BCS17, BMT25, BC01, BC04b, BG06, BM06b, CL06, CMP03, CLY19, CZHX19, CDW23, CC08, CKK25, CGK25, CBHM19, DDP12, Dek17, DK21, DM09b, DM11a, Dol14, DYZ20, DHM09, FHM^{+}02, FSX25, FM11, GGM95, GGO16, GGRN17, HLY04, HLZ06, HMW05, Joh01, JM94, KPY15, Kli15, Kre07, KK22a, KDD23, KSJY25, Kwe00, LCHR03, LL98, LMWZ10, LZ14, Li16, LYF17, LZ18, LH20, LYZJ23, LL24, LY03, LZIZ23, LO95, LO96, MW24a, MD23b, NS21a, NY13, PD01, PS19, RL06, SW11, Sha21, SC25b, SG07, ST25, Spi93, SWCH15, TTP24, TM04, TM05, VNC21, WH13, WCJ23, WKWC26, WWWW25, XL11, Xu21, YSBL14, YJ21, YS22]. estimates [YJ23, YT25, ZMC13, ZCZ15, ZZHS18, ZFW20, ZQZ23, ZCQ25, ZWG25, ZYS17, ZXW17, Zou10, Zou11, Zup03, Zup04]. Estimating [SS25, KK17, PRSS24, YK07]. Estimation [BCMV03, Fuj99, KV20, Rei85, RVdCVR02, AOW94, An16, ASZ15, AMV03, AB12b, BLS94, BGO13, BD07, BBW19, BHR05, BP92, BP06b, CDGA17, CCS17a, DDHS97, DSA20, DT89, EFLFP09, FZM20, GHT05, HLMKZ06, KDT17, KPRU20, KB21, Kim12, KK20b, LSK12, Moo04, OZ96, Tia15, Wei18, YV17, ZZH25]. estimations [AM16a, Fdi96, MWC21]. Estimator [AFLP12, AKT97, AR18, Bac21b, BKS07, Bür12, HOS11, Hop23, Kim21, LS12, SZ09, Wal19, XYHM20]. estimators [AFF^{+}15, BBG14, BCT19, CHZ21, DLZ21, Kim07, SS00, Yi12]. Eta [MS25]. ETRs [BT97c]. Euclidean [BS00a, CP07]. Euler [ALMM01, AGK24, BRSD91, Bor10, BW15, BCGI13, Boy91b, CGH23, CHPV09, CGW20, CWC24, CMP23, DLN04, DGRS09, Fel06, Fre04, FGGL22, GDEdLD23, GLMY17, GLM18, GLQ26, HN03, HS86, IMMS20, JKW12, Jes85, KTK20, KKW00, Kor95, LA21, LDC24, LW07, LLD18, LMTW20, LMW23, LYLL23, LSW23, MT20, PP92, PM14, PWX24, RZ00, RA03, San02, SMA01, TWL23, Tan23, Wan25, WC11, WHW21, WMC09, YZC21, Zha20a, ZL23, vS97]. Euler-accelerated [Boy91b]. Euler-flow [Kor95]. Euler-Riemann [YZC21]. Eulerian [Aso21, KCC04, RVD00, Sea09]. European [CPOGO17, itHT18, Rou20b]. evaluate [CPD^{+}05, LL15]. Evaluating [EC07, ACM09, BVB09, LG21]. Evaluation [BO87, BP02, BK06, BZ17b, DFC09, EW97, JBLC11, KW21, LS16, NRZR12, Pel15, ST14b, SGN08]. evaluations [Dah02]. Even [AS05, BCC16, Bic21, BtTBV87, tV87]. event [Aca12, ABI22, LM22b]. evolution [AHO16, ADM22, ADFP25, BG02a, BLL24, CLT97, CFKS07, CNS00, DA17, EH91, Hag15, HO10, KNT13, LT12, LWWZ25, LT23, LGZH24, LMWZ07, MQO17, MPHFP23, Mik97, MŠ99b, NA21, OS12, PPC00, QXQ22, QM20, RT14, SG92, Tah96, XL11, XGQ20]. evolutionary [BSZ22, BJ01, Chi21, CJLS98, CG14, Gar96, JY23, LH20, lLNW21, PM91, YQCZ22, Zeg97, ZBNC25]. evolving [FM07]. Ewald [HNP17]. Exact [CH15, FJ09, Ise97, LSVT25, AB10a, AW14, Bac21b, FV87, GHKM09, HH98, Hag15, Kim21, LA12, LWYG22, SL17, XYHM20]. exactly [CCCC^{+}25]. exactness [Mil17, Yi12]. examination [Sch98]. example [QM03, Ran15, Ran16, VVR08]. Examples [CAD03, HS98, Waa88]. exchange [BDOG19]. exchanger [Wag98]. excitation [BKP14, MCBV20]. exclusive [Kür23]. execution [For11]. exercise [ZGO12, ZO14]. exhibiting [Bec18]. exhibits [SG92]. Existence [Fuh01, HS95, KS25, KP15, Ma03, MD20a, Zak19, DD21, FL01b, Gu01, HHT97, LYA^{+}19, SNOK21, Spi13]. existing [CZY08]. exotic [DLM16]. EXP [BBT25]. EXP-RBF [BBT25]. expanded [Ars20, CWHF19, HCY18, LCHW20]. Expanding [HW93]. expansion [Boy07, Cer25, CKK10, Fel06, GGS16, Han93, Hig93b, HH10b, LX09, MK21, NS13, Wal90, WW24, WE99, Zan01]. expansions [AR93, Bor10, CJM88, DKL24, DC09, Har00, KM17, KP92, LC99, Mac92, Mie89, ZO14]. Experience [Ney95, SD93, BR97]. experiences [BSFDM02, HvdHV10]. Experiment [TYKK01a]. Experiment-based [TYKK01a]. experimental [Aya09, FHM^{+}02, KKN^{+}13, KKN^{+}17, SA08]. Experiments [BH85, HG98, Sch99, Ste97, AGKS25b, Bas21, BG11c, CFKS07, JMP06, PWS98, SY03, SH02, SG92, TOD11, VP91]. Explicit [ABH22, AM99, Ant23, CMRdlT24, Cha25, Cha98, CS03, Con01, DM11b, DGM18, HLR01, lLXhLZ21, Noo95, Ric91, ST20, Som93, VS95, Ver96b, WVBM88, YF24, AHJ^{+}23, ACP24, AS04, ARS97, BC08a, Bos09, BCJP18, BH93, BIMV19, Bur85, BB96, BW03, CH95b, CMRV11, CZ97, CCG13, CV95, CGS20, DZMB21, FW07, FW08, FWL18, FD97, FHV97, FHX22, GM16, Gje07, GND19, HMR25, HS09a, Hoa15, HZ02, HJ03, IJ17b, KKT16, KKP07, KHLV22, KCL00, KSMMM16, KW98, KDKW20, LCK22, LGC25b, Lua17, MK20, MD20a, MD20b, MMDH19, MOU14, MD25, MSA20, MAH22, MVBS25, NS21a, NMKE13, NK24a, ÖPT25, Olv92, Ost93, Pag25, Ric94, San20, SA21, SWJ09, SG96, SV00, uIVS13, Som86, ST25, SS09, TGV22, Tsi01, Ver96a, VBH96, WWX13, WKP12, WG11, WdG92, YZWQ25]. explicit [ZYQS21, ZCQ25, ZZO16, ZSZZ20, ZL24, in 02, vSW90]. explicit-extended [KSMMM16]. explicit-implicit [DZMB21, WdG92]. exploding [NMKE13]. Exploiting [GV04, MNSS22, BCV25]. exploration [RKR20]. Exploring [Her91]. Exponential [CGP15, Duj09, HHYD20, HO05, LZQ22, QR03, RK08, TGB08, WEA12, WPT19, XCHW22, ZMY21, ABK12, AB14, ACM09, AEF^{+}14, BP12b, Bor10, BDE22, CO09, CC24, CL07, CGW20, CS17, DL22b, DLQZ23, EEAS25, FXY22, FHX22, GS25b, Har00, Has13, HK25, KHYY21, KNT13, LMO24, LY10, LZZ22, Li22, LW22, LC24, LZL14, LT19, Lua17, MS99a, NS20, NS21b, SC25a, SW09b, TN16, WT20, WPS18, WL21, ZZPJ23, ZZF25, ZZO16, dC18a]. exponential-type [CC24]. Exponentially [BO11, FSX25, VV05, VV09, Zar17, AW03, CDP17, GS94, HMR25, VV07]. Exponentially-convergent [BO11]. Exponentially-fitted [VV09, VV07]. exponents [BH97, DV95b, FZM20]. expressed [HS24]. expression [OZHP23, Pre90]. expressions [Maj20]. Extended [Hey10, KCI03, SV00, Sim10, TM21, WX06, ZLJ20, AMH24, BW23b, Bar05, BL86, BT97c, CL85, CZ12, CD20b, FW22, GC25, HJR22, JMU^{+}25, KD13, KSMMM16, KK17, KK22b, LZ18, LW19b, LXZS22, Lte24, QNA23, Sim93, WWZS25, WX22, YT03, ZH09, van86a, OR18, CLR11]. extended-cubature [KK22b]. Extending [AMV17]. Extension [MP20, AGM09, AFK92, EL97, Haa97, HZ02, LZZ25, Lot19, MST09, Roo20, SB25, TM05, Ye04, Zha21b, ZYJD25]. Extensions [BRZ10, AFS02, BS05, CW98, VZ93]. exterior [BM13, EB12, FL15, Gon06, Har98, HK85, Lau17a, MM02b, RS00, YWW23, ZLHW19, Zhe07]. external [Cos25, Tom24]. extra [WL24]. extra-wide [WL24]. extracted [Dah02]. Extragradient [TQY24, BAA22, DP21]. Extrapolated [Bou25, Cao98a, DK20, DL22a, DLM20, LJ20a, LR20b, LJ20b, SW03, ZYLL20, ZZLL21]. Extrapolation [BZ94a, Bre96, CM00, WC02, Bre85, BZ91, DSS15, Fik23, FS88a, GG95, GC15, HN03, JM16, LD97, LX09, LL02, NS03, Nor97, Nor99, Now96, OMSS25, Rha97, SMEN04, Sid90, VVR08, Wan07a, WH19a, Wen10a, YBL13, YXZL24, ZHL22]. extrapolations [BLN25, Cha96, CHH15, HLL09]. Extremal [DIR13]. extreme [Bai25b, WLZ25]. extremum [WY22]. extrinsic [ebKMZ24]. Eyring [PT09].

FAATNA20 [DLN^{+}24]. Fabrizio [LIPT18, MDASAO21, SC20]. FAC [MMT90]. face [CW22, QPT23, ZY19]. face-based [CW22]. face-centered-cubic [ZY19]. facilities [TER03]. facing [CHK99]. factor [BDOG19, GZQS23, JHGZ20, LM25, LCM22, LCM24, Mar09, RP17, SCT05, Tan23, YJJ^{+}24, ZYQS21, ZYQS23, ZD25a]. factored [WZ02]. factorial [Bor10, Wen10b]. factorially [BW15]. Factorizable [Sid02]. Factorization [vdHS01, BD85, BSV21, BCSH16, CG05, ELvdHS98, GV02, GNX19, Guo96, LMV17, LTDY24, LDP^{+}14, Not99, ST09b, SHG86, SYW18, GPHA16, GLW^{+}25]. factorizations [BLM17b, Doi91, For02, Gu01, Mag91, Not92]. factorized [KK20b]. factors [CGEV19, KK22b]. fADE [GKKM21]. Failure [Gar92, KN08]. failure-prone [KN08]. fair [RN04]. falsi [CL07]. Families [BCET24, DR09b, BCF^{+}13, Bre91, BDRZ04, CCdlH20, CGPT19, FR18, Hin97, Nud24, ZZJ21, vdES04]. Family [CR19, PSR04, AK09, AÁ21, ADM10, BFK11, CN16, Fra16, GPHAM12, KSSS16, Kom07, MPPR22, NWL^{+}22, OGV92a, Pan07, Ver93, YG95, YG99, Zha09]. fan [GWZ25]. fan-shaped [GWZ25]. Far [EH88, MRS03, Chu03, QL16]. FAS [Spi00]. Fast [BCDR25, Cao03, DNW18, EH05, HZ21, HV22, HR14, JHGZ20, JL23a, LMV25, LWYG22, LYC24, MP98, Pea16, PT15, Ram12, RN04, STS00, SL17, SND19, XWC25, XFG19, ZJH^{+}23, ZW19a, ZLL22, AKS21, BP14, BDN^{+}97, Cai09, CGJ16, CLTA18, CL24a, DM12, Eng11, FJS25, GZW22, GL93, Ibr25, IT07, JWZ21, LS16, LLKJ21, LCJQ12, LCLW17, LL20b, LNZ12, LLW20, Mat05, MMT90, MR94, MRS10, RL21, She00, SKW17, TGB08, VK17, XL09a, XL23, XWX21, ZXYW22, ZY23, Mar03, ST14b]. fast-wave-slow-wave [SKW17]. faster [AJW23]. fastest [Dar00]. FCC [ZY19]. FD [OT22, ST25]. FE [HHR12]. FE-analyses [HHR12]. FEA [SA05]. feasibility [AHAS21, vR04]. feasible [dOF20]. Feb [AS21]. February [Ano22w, Ano23v, Ano24n, Ano25w]. feedback [HJR22, MDHK06, SA05]. Feedforward [MDHK06]. Fejèr [Nov03]. Feller [Zhe19]. fem [MOS12, AGLRS23, AMK18, AN22, BS14b, BGG04, BBG14, CCOVF22, CD23, CR23a, DS07c, DG22, Fra16, FS24, GT15, GMZ11, Gas92, GMM09, GKMS09, GÖS20, GMS12, Gon06, GWLN22, HY01, HSF25, KR25, KMS10, KTYY24, Lan97, LSG24, LLN25, LX24, MS00, MS08b, NS12, RDH^{+}12, RV09, Sel14, SL22, TS24b, Usm97, WCXL09, Wan20, YXF25, YJ23, Yos00, ZXW17]. FEM-BEM [GÖS20]. FEMs [Bec18, EKT19, LY01, LH02, LSWM19, LWWZ25, YJ21]. Fer [Zan01]. Fermi [ZB19a]. ferrohydrodynamics [CLL25]. ferromagnetic [Fra04b, SC03]. FETI [DHS05, Ste05a]. few [ARY23, DL21a, DV95b]. Feynman [SND21]. FGMRES [MRH14]. Fibonacci [BL86, GS24, KSSS16]. fibres [NS13]. Fick [KHM^{+}19]. Fickian [BFdS10]. Fictitious [BH12b, AMR14, ABG^{+}15, NCYC22, SW20b, WJW19, Zho17, Zho18]. fidelity [CKS05, Rob01, SYG^{+}05]. Field [DRC85, AGLRS23, AAD^{+}08, BC02, BF17, BLY17, BP85, CL02a, CMP20, CjW18, CPY20, CKS05, Cos25, Dav98, EH88, ESS15, GD09, HSF25, JK17, LM00, LMY18, LW22, LQXK23, LL20b, LC21, LY24a, Maj20, MRS03, MD20c, PHY19, PA18, QZH25, QL16, Ram94, Ric08, RS22, SC25b, SWJ^{+}24, TOCV02, WJW19, YLL21, Yao24, YXX24, YK04b, ZY19, ZCY20, ZLSZ22, ZH24]. fields [DFC09, FGGL22, JR18, Lei99, LW07, MAG13, WSC09]. Fifth [QR24, WDL23, AAL^{+}26, AAEMY21, Chr01, CN16, Tan23, Tan24]. fifth-kind [AAEMY21]. Fifth-order [QR24, WDL23, AAL^{+}26, Tan23, Tan24]. filling [FGPR12, LS10]. film [AJW23, SL21]. films [BCMV03, SSW04]. Filon [DS21a, Has09, Has13, Maj17a]. Filon-type [Has09, Has13]. Filter [Wal00a, GT18, HC01, Hua20, KK22b, QH19, SXP09, TZ00, Wal00b, ZP12]. filter-type [Hua20]. filtered [OT21, ORT24, OT25]. filtering [Bar05, BGS02, KK17, KK22b, LOS03, MMRV20, RGÖS18, SM20]. filters [KK20b, YR09]. filtration [CGN03, CD13]. final [HP14, Muo23, YL13]. finance [Lai09, Le 12, RO16, SS08a]. financial [CKM15, MHL18, RKR20]. find [CGTTN24, NSD23]. Finding [CZ19, AKM^{+}22, BASC17, Car94, CGG02, CZY08, CH21, HdSRI17, Iga85, Meh08, OFY^{+}23]. Fine [RTV02, Gus87]. Finite [AMH24, ADFR18, BW86, BGG^{+}20, BKP14, BTDV10, ÇD17, CLR11, CF86, ÇK13, Cop03, DL06, EGH01, FJS99, GAML04, GAZD25, GS94, HH22, JV09, KZ13, KR12, LRV25, LMWZ07, MT06, NB01, NFAE03, Pot85, Quy19, RBBC85, Rei85, RV15, RA09, Sid23, SD09, Sol15, WYYL19, YGY15, ZW87, ZZH25, AD20c, ABZ21, AB17, AMRR18, AP20, AL09, AMT13, AS97, AK00, ADK94, ATW20a, ATW20b, ATAW25, AA87, ADM22, AAI^{+}93, AKT97, AW03, AJ24b, AKL08, ABR05, ACP23, AM16a, ASCM02, ASC03, APJ09, APJ10, Aso21, BSGU94, BC12, BKK25, BHJ05, BHJJ06, BBS25, BPS19, BY00, BF01, BK17, BLY17, BM00, BM01, Beh97, BRTB19, BK09, BB15, BLJ21, BC08a, BCGS24, BMS89, BM04a, BM04b, BM04c, BW95, Ber04, Ber05, BC08b, BG03]. finite [BDF94, BC01, Bof06, BG06, BG14, Boh03, BP97, Bot97, Boy15, Bra00, BM18, BJM01, BJTZ20, BD11, Bür12, Bür13, BL08, BH12b, BD22, CHP19, CFLW22, CY23, CCDJ20, CLT97, CHLA21, CW21, CCG13, CGA93, CDGA17, Car09a, CHSS01, CGG25, Cau08, CS94, CP05b, CCZZ18, CMP03, CS03, CC04b, CZ04, CCLT10, CL10, CJX11, CHX13, CXZ15, CXZ17, CC19, CH19, CWHF19, CYWH22, CL24b, CLL25, CML05, CC08, CNT07, Cho13, CL18, CK20, CCZ22, CS18, CDJT06, Cod08, CGK25, dCCSR03, Con99, CBHM19, CMCGTR02, Coy12, CH25, DPPR16, DAH24, DZZZ24, DKSS24, DS21c, DA18a, Dek17, DK21, DJ10, DDS89, DL20, DW21, DL22a, DM11a, DLP06, DT10, DN08, DL16, DS15, DHM09, DCL23, yDqGnJT09, DCJ20, DYF23, DII15, DR93, EH07a, EBEHAS24, ER07, EN09, EJS11, Eva94, FLH22, FL04]. finite [FG98, FMGN94, FLÖ^{+}97, FKA^{+}13, FM11, Fre98, FM07, FMU15, FWW^{+}21, FSWZ19, FGGL22, FL01b, GM10, GX11, GLLW14, GNUV24, GP23, GLV03, GIS23, GHH20, GH25, GLPW09, GD21, GGO12, GGO16, GP01, Gol00, GBBC^{+}23, GHT05, GCZZ23, GGG16, GGRN17, GGRBRG22, Gul15, GJ17, GZHQ23, HHR12, HHAA22, HZ09, HLZ14, Har10, HKZ08, HLMP09, HHL23, HSS04, HHC08, HL08, HS20, HD22, HTSZ23, HGF24, HO24a, HS22, HJZ23, HZD21, HMdV03, HL02a, HSWW24, HL23, HL03, HvdHV10, HW97, HCY18, HJYL19, HL21, HL24, HLW25, HLJ20, HZ02, HZ12, Hua17, HJX^{+}19, HAC22, HMW05, HJ03, Hus20, HSY18, Ioa89, ID19, JPP19, JJ94, JK14, JCL18, JZS20, JWZ21, JEG10, Joh01, Jor11, Kal96, Kam16, KDAK13, KPRU20, KT05, Kat89, KD13, Kie17, KF97, Kim12, Kim21, Kop89, KJ99, KFOF02]. finite [KOS^{+}12, KDK17, KK22a, KN93, KDD23, KX03, Kwe01, Kwe03, LRS23, LHC09, Lam13, LPT94, LM00, LCHR03, LHN24, LN26, LCK22, Lee23, LWD^{+}09, Li98, Li00a, Li01a, LY08, LHH08, LH09, LMWZ10, LA11, Li12, LHW17, LR18a, LR18b, LLHC18, LW19a, LL20a, LSP20, LW20a, LZ20, LR20a, LWW20, LCHW20, LLY21, LW21a, LXZS22, LL24, LWaZ24, LLZL24, LHY24, LGC25a, LGC25b, LYZF25, LP25, LZW19, LX09, Lin10, LLS^{+}96, Liu09, LCLW17, LCL18, LC20, LZIZ23, LS24b, LD22, LT19, LAH09, Lte24, LL02, LW95, LP01, Luc95, LZY09, LJ20a, LR20b, LJ20b, Lyn99, LLW20, MDP10, MD00, MK14, MA09, MP11, MS08a, ML91, MZS10, MW24a, Mar99b, MM14, MR94, MM02b, MVG14, Mic03, MOU14, MF23, MD96, MdD04, Mou03, MWYZ18, MC21, MT20, Mul99, MM20b, NAF24, Nag22, Nak05, NLHT25, NY13, Nke07, Nud24, NS16, OH20, OEAS21, PZMX16]. finite [PCL^{+}25, Pel20, PGYF20, PSWZ21, PQ25, Pic05, Pir09, PS21, PS19, Qi24, QW25, QMLC15, Que21, RN25, RZS21, Ran20, RSK24, RS20, Ric91, Ric94, RVD00, Ris05, RGMO19, RG21, RBT15, RU07, SSZ16, SS00, SGS00, SD11, SD13b, SS19, SH10, SQ17, SWX00, SYY20, SG16, SJ20, SL20, SW20a, Sin23, SG07, SCLL21, SDK15, SBS24, SL15, SMA01, SSA24, SSKS21, Str98b, SH91, SW24, SLZ10, SW20b, SN04, SGN06, SGN08, TLP18a, TLP18b, TWMP20, TN16, Tan93, TYKK01b, TC22, TDW23, TM21, Tol04, TTP24, Top21, TH09, Tou25, TK15, Tro96, TM04, TM05, TMM15, UWY22, VNC21, VRK25, Ven15, VRC21, VMS07, WL09b, WZL13, Wan17b, WCSQ18, WL18, WZ19, WR20, WDU21, WaZW21, WY22, WL22, WLG22, WDL23, WLY24, Won08, WSC21, XZW19, XWZ21, XL23, XL09b, XZH19, XGQ20]. finite [XHYM22, XXF22, XP23, XYZ24, YLL09, YS09, YY13, YBL13, YSBL14, YW19, YQCZ22, YZG23, YYZ23, YT25, YZZ25, YZ21, Yi12, YCWH23, YZ19, YHT23, ZBD24, Zeg97, ZL17, ZLY23, Zha09, ZCZ15, ZSG^{+}20, ZLW20b, ZL21, ZL22, ZZPJ23, ZSS23, ZZ24, ZYZJ24, ZWG25, ZZZX26, ZW19a, ZLL22, Zhe07, ZZ17, ZZL17, ZZ20, ZLWF21, ZS21a, Zho17, ZZC^{+}18, ZSY20, ZZLL21, ZFX25, Zou10, dSFDG20, iV09, van95, BS14a, BGP11]. finite-conductivity [CML05]. finite-difference [AAI^{+}93, CCG13, CGA93, CMCGTR02, DS15, HZ09, KT05, MDP10, Mic03, MdD04, RU07, SDK15, SMA01, WLG22]. finite-dimensional [KF97]. finite-element [CWHF19, GGRN17, Mul99, Nak05]. finite-length [SH91]. finite-part [Ioa89]. finite-time [KPRU20]. finite-type [Mar99b]. finite-volume [BP97, DT10, GBBC^{+}23, HJZ23, Kal96, RVD00, Tol04]. fins [HHL23]. First [HS17, HvdHV10, AACP20, ACMR06, AL17, AAEMY21, BKM19, BC01, Boy07, BMV19, Bru97, BMM97a, CHH15, CS09, DN21, HP85, HDS20, HLL09, IT16, IKMM23, JJ94, KM19, KLY05, KW10, Moo04, MPDB24, NDM20, PP24, PCR17, gTpM07, VGL24, Wu09, Xu16, ZL18b, Zha19b]. first-kind [MPDB24, ZL18b]. first-order [AL17, AAEMY21, HP85, JJ94, KLY05, PCR17, gTpM07]. Fisher [AMH24, BFdO07, CMMR23, LZ18, LXZS22, MSZ^{+}24, QNA23, jWjJ17, WWZS25]. fit [Mon09]. Fitted [HO24a, TMS87, AW03, AHS03, CDP17, CSLY19, FW08, GS94, HMR25, HS09a, Hoa15, Li19, Nag22, NLHT25, NS16, Pat00, RSR23, VV05, VV07, VV09, YW08]. fitting [AN15, BCDR25, KDAK13, PS02, PS03, SKS^{+}25a, ZZPJ23]. FitzHugh [BSTT22, LYZF25, ZLW20a]. five [IMM04, MI03]. five-axis [IMM04, MI03]. fixed [Gil91, GKS20, HGM^{+}21, HS95, HT20, RV05b, SL08, Tow16, VA21, Wai98, WYP12, XY24, YP18a, YP18b]. fixed-accuracy [XY24]. fixed-point [WYP12, YP18a, YP18b]. flame [RAS99, RSL89, Son91]. flat [Nak05]. flexible [HP15, Lam24, LDIW16]. Floater [CHS17]. floating [LMO24, Ske99]. floating-point [Ske99]. flocculated [BBCS05]. flood [CCM17]. Flory [GJLL20]. Flow [BW86, PP00, PGM86, AH09, AA22, ASZ15, AKT97, BDMG12, BFS17, BJ00, BBT25, BIMV19, BL08, Cai15, CML05, CRTU15, DZ12a, DM09a, DYF23, Duf90, ER07, EJS11, Ewi91, EWW99, FMW18, FL01a, FLL11, GM08, GMZ08, GM10, Gar92, GHK16, GG22, Gat91, GOGF03, GJR03, HJR22, HJ09, HS22, HCW16, HST14, IK24, IMC22, IR22, KBK21, KHM^{+}19, KHA12, KSMMM16, KH91, Kor95, Lam24, LTC03, LC02, LP25, LMS08, MD00, MM07, Min04, Mou03, NS21a, Obe15, PNA21, PT09, PK91, PCRR17, RSK24, RR00, RVD00, RZ15, Rou20a, RGK21, RBC02, Sch16a, SQ17, SG16, SZE^{+}92, Shy86, Sid02, SED21, SY08, SH91, SW05, TYKK01a, TYKK01b, TOCV02, TC03, VO00a, Zan91, ZOZ09, ZZZX26, ZYS17, ZQLK11, ZS18, dB03, vdHVW01]. Flows [CF86, Mac86, hYK86, AI19, AJW23, ASS21, An20, BL91, BC04b, BBL02, BS08, CIZ96, CCD^{+}20, CHPV09, CCLT10, CHK99, DCSB25, DK20, Din93, DMQ02, DdCVR03, DL21b, DII15, EH07a, EAS12, EL01, Fai00, FS23a, FPRA09, GP00, GKKM21, Gla93, Gla94, HS24, HH10b, JN07, KMR09, Kar89, Kie17, KW98, Kop89, KDK17, KK22a, Kwe00, LDIW16, Li25, LT00, LS24b, LR87, LP97, MLJ19, Mur15, NC16, PBC08, QW04, RCGM98, SC25b, SWJ^{+}24, TLP18a, TLP18b, TER03, Tan24, TKN11, Tur86, WaZW23, WPT19, XLK07, YZ24a, ZZF25, ZB19b]. Flowsheet [Bie87]. Fluid [CFXZ06, AK21, AA22, Aso21, CFX08, CL09, DL21a, DLQZ23, DYF23, DJL04, Elm02, Ewi91, FJL21, FMP04, FLL11, GMM09, Gat91, GÖS20, GS18, GJR03, HS22, HT00, HCW16, JRT90, KHM^{+}19, KR25, LTC03, LGS21, LX21, LC25, LK07, LS24b, MP05, Mat08, MR01, Mur19, NC16, Nür09, OGS20, PT09, Pul86, RDH^{+}12, Rou20a, RGK21, Sid02, SD93, SED21, SM89, Tow16, Tur86, Tur93, VS91, VT91, WJW19, Wu09, ZOZ09, ZYS17, dPT96, vBvdZdB08]. fluid-dynamics [dPT96]. fluid-fluid [AK21, LGS21, LX21]. fluid-particle [Tow16]. fluid-structure [DL21a, GÖS20, GS18, LC25, MR01, RDH^{+}12, WJW19]. fluid-surfactant [DLQZ23]. fluidized [VCN20, VNC21]. fluids [Ari04, Bir87, CMP06, CPY20, FPRA09, HS22, Kie17, LWaZ24, LR19, MZN21, QR03, ZGR23]. Flux [FM07, Kim12, AKG14, GD21, Gla94, HSS07, JK14, KL21, Kni94, Kni95, LGSW25, Li25, Nor99, RGÖS18, RS20, RA03, Tor06, ZEW20]. Flux-based [FM07, GD21]. flux-extrapolation [Nor99]. flux-splitting [RA03]. fluxes [DGM22, FS23a, XWM26]. fly [BMR^{+}17a]. Foerster [HT24]. Fok [NT16]. Fokas [AC16, MPSS16]. Fokker [CjW18, Van92, jWyG08, jWC22]. following [SZ99]. FOM [JMS99, WPS18]. force [Kie15]. force-gradient [Kie15]. forced [CFC03, LBCN00]. forces [DS21a, Sch16a]. Forchheimer [CD23, CGG25, CD13, Kie17, LHY24]. Forchheimer/Darcy [CD23]. forcing [CSSZ25, HW04, IKM23, WDU21, Zen93]. forecast [WCGW95]. forecasting [RW87]. foreign [BDOG19]. Foreword [FJ92, IN01, Tah19, Tah24, Vic87a]. form [AN25, Boy06, CMT25, EK97, GND19, JY20, JBLC11, Kam00, MN03, MZM20, OZHP23, PWS06, Ree03, Sim93]. Formal [CAAT16, Mur99a, Mur99b, WB90]. formalism [EHNR24]. formally [QXQ22]. format [RS22, SMTHE22b, SMTHE22a]. formation [HPW21]. forming [AGJCP25]. forms [BS14a, Beg00, BMT25, CJ90, Eir95, Sal03, SST09, Zan91]. formula [DhW09, Kza99, MV20, MD20c, SHA12, Ske89a, WC02, Zla85b, Zup04]. formulae [BZ93, Ber86, CV95, Dia95, RMH20, WX06]. formulas [BRS16, BDMGVO05, BGVHN10, CCBGV08, Chr01, DGV00, DBCBPP10, DIR13, GPHA16, Has08, HJ17, JKN94, KS09c, Kza92, Maj17a, Meh22, SG96, Sid14, Sid23, Ske89b, SBS24, Xiu08, ZZJ21]. formulation [AA87, AFS02, BC02, BK17, BRS05, BSQ96, BM18, BP95, CGS19, CHOR19, CSW19, FS24, Gat91, GP00, GS18, HP85, HH98, HGF24, HL03, KKP07, KS02, Li98, LZY09, MG00, MS08b, Med96, NK24b, SLJ86, SMW21, Wan05, Wen98, ZdBT03, DM12]. formulations [CDG19, LRS09, LLL08, SM13, WK00]. Fortran [GBDB97, TS06]. forward [BS21, Fre98, HJ21, JM94, LQY25, MS25, MII13, SD24b, YLW20a]. forward-backward [BS21, HJ21]. forward-backward-forward [LQY25]. forward/backward [SD24b]. found [Duf90]. Foundations [FWW^{+}21]. four [BLY17, CPD^{+}05, LZJ21, WBCK02]. four-center [CPD^{+}05]. four-field [BLY17]. four-point [LZJ21]. four-stage [WBCK02]. Fourier [ST14b, AQ20, BH96, CjW18, CLX21, CPOGO17, DM09b, DL22b, DN24, FJS25, FFQ09, GANT02, HM15, HM17, Has20, JZS20, KHM^{+}19, Li22, LZZ25, LT23, LZIZ23, Lyo12, MM07, Moo95c, NK24a, Nke07, Pan21, PLB22, RGL16, RO16, She00, WZL13, WMF17, WDH20, WWM22, YLX21, YWW23, YR09, ZO14, ZLW20a, ZYJD25]. Fourier-finite-element [Nke07]. Fourier-type [Has20]. Fourierization [SW07]. Fourth [AKBF19, BM06a, CHSS01, EH09, GX11, Lua17, XLKY19, YT00, AP20, BTBR19, BIO24, CFLW22, CS19, CES91, DhW09, DZ12b, Den15, DB08, DCL23, FK23, yGqWsWC05, HPH20, HTSZ23, Kat89, KZ13, LO22, LLJY20, LW21a, LXZS22, LWC24, MFAD23, NP21, PZMX16, PGDB08, QXG21, RN25, RZ18, Rog19, RTA19, Rou20a, SC25a, SXL22, SP22, SKS23, TQA^{+}25, Ven15, Wan07b, Wan17b, WXY24, Wen98, WYP12, XWX21, XZT21, ZLY23, Zha14, ZJ25, ZZLL21, BGH^{+}15]. Fourth- [BM06a, CHSS01]. Fourth-order [AKBF19, EH09, GX11, Lua17, AD20c, BTBR19, BIO24, CFLW22, CS19, CES91, DZ12b, Den15, DB08, DCL23, FK23, yGqWsWC05, HPH20, HTSZ23, Kat89, KZ13, LO22, LLJY20, LW21a, LXZS22, LWC24, MFAD23, NP21, PGDB08, QXG21, RN25, RZ18, RTA19, Rou20a, SXL22, SP22, SKS23, TQA^{+}25, Ven15, Wan07b, Wan17b, WYP12, XWX21, XZT21, ZLY23, Zha14, ZJ25, ZZLL21]. Fowler [GS24, LO03, SSC23]. Fox [Meh22]. FR [MK19]. Fractal [BPN24, BPV25, ABdSG23, DdCVR03, Hey20b, LCZ25, SSA^{+}22]. fractal-fractional [SSA^{+}22]. fraction [AR93, CJM88, CV88]. Fractional [AGZD22, ABD24, BJ03, CD18, FZA25, HAML21, MD19b, RR21, ROB17, AD19a, AD20c, AD21, AKM^{+}21, ACKV24, ABZ21, AACP20, ABdSG23, AS20a, ASA20, AMK18, AHO16, AKBF19, ALZ^{+}21, AL22, APJ10, AKS21, AD19b, BL21, BFS17, BK21a, BZ17a, BK21b, Boh21, BBBN21, BS20a, BCGI13, BMWH20, BJTZ20, BO21, BJ01, BJ06, BCDP17, CML^{+}25, CA21, CC18, CDP19, CCdlH20, ÇD17, CDH24, CDW13, CLTA18, CLX21, CCST22, CQZ20, CWY20, CG21, CJLS98, CP03b, DA16, DA17, DA18a, Din19, DCY20, DCJ20, ESEKZ10, FM21, FK23, FWHM20, FZM20, FP24, FSWZ19, FCW20, FXCW21, FCW21, FHX22, GG19, GMG19, GAZD25, Gu20, GWLN22, GND19, GW20, HHAA22, Han19, HZ21, HMN20, HA21, HPH20, HZ20, HZD21, HMD21, HQAZ24, HJ17, Hey19, Hey20a, Hey20b, HAA21, HAR21, HCGW22, HS19b, HAY20]. fractional [HZAT21, HS21b, HAC22, Hus20, IB24, JK21, JRS20, JZXJ21, JWZ21, JHGZ20, JL17, JL24, JCJP21, KBS11, KS25, KK20a, KAS22, KSHB21, KLR25, KHB22, KDH20, KAD24, KP19, KZ21, KGZ24, LP24, LPS25, LLKJ21, LYF17, LHW17, LR18a, LR18b, LRC19, LSP20, LR20a, LCHW20, LZJ21, LPV24, LQS21, LS21, LYY15, LYK17, LCLW17, LCL18, LLZ19, LZCF21, LL21, LSY21, LYZW22, LWLL24, Lot19, LXCM21, LR20b, LLW20, MD20a, MD20b, MD22, MN23, MCS16, MWC21, MMDH19, MMD20, MMDS21, MC17, MT06, MM20a, Mit22, MMM19, MAD23, Mok17, MDASAO21, MMP20, NP21, NLS18, NLS20, NTHC21, NA21, NWL^{+}22, Odi19, OB20, Odi24, OB24, OGS20, OCVW22, OAHN22, PB21, PTV16, PTV20, PH17, PJB04, PAJ12, PMP23, QWX20, QWW24, ROL19, Rah25, RE19, RMH20, RZ18, RL21, RV22, Roo20, RGMO19]. fractional [RREP^{+}20, Rou20b, RG21, SRK21, SRK22, SKBAS08, SOB20, SK22, SY18, SC20, SWB20, SWB21, SJ18, SSA^{+}22, SA20, SS17, SD24b, SA18, SA19, SL15, SOA^{+}25, SSKS21, SWW11, SSPZ20, SW17, SND19, SZW19, SND21, TDMT21, TTP24, Top21, UHUL21, Vab21, VA05, VL19, WZW13, WH18, WMLB19, WR20, WZZ21, Wan21, WLM21, WCS21, WCM21, WHW21, WCL22, WW14, Wu09, WWF20, XWW19, XW19, XZL19, XLZ20, XWZ21, XF22, XWX21, XZZ19, XGQ20, YMD21, YL13, YJZ18, YLFT20, YLX21, YXN21, YLLZ21, YDWW17, YWSL20, YRV21a, YRV21b, Yüz22, ZM19, ZDM18, Zak19, Zak20, ZHS22, ZL17, ZQY18, ZZ19a, ZJ19a, ZYLL20, ZYC22, ZJH^{+}23, ZW19a, ZLX19, ZZ17, Zhe19, ZZ20, ZEW20, ZZJ21, ZLWF21, ZJLA22, ZZ19b, ZSZZ20, ZXW17]. fractional-in-space [BMWH20, XZL19]. Fractional-order [RR21, ROB17, ACKV24, Han19, IB24, KGZ24, LLKJ21, YRV21b]. fractional-step [ZQY18]. fractionally [HLY22]. fractions [Bre88, CJV88, GGMP88, Gil91, Hau88, Jac88, JT88, Lem88, Lev91b, Lev91a, Lor10, Njå88, Waa88]. fracture [CH19, KR25]. Frame [Liu21, HS24, IPL02, Jia12, PLI03, Shi20]. frames [GS25a, JP17]. framework [Aso21, BGO13, DSZ15b, DSZ15a, GR02, HZD21, KN19, Kim19, KS07, SB03, ZDM18]. frameworks [MM16]. Fredholm} [All24, AKM^{+}21, AHB20, AS20c, Aze22, BKM19, BES18, BBCR22, BMGGG12, CHLX07, CC20b, DAMA23, EGV25, EHV19, Fer09, HS19a, HDS20, JRW06, Kan89, KX91, KM25, KPR12b, KKK25, LW21b, LNZ12, MKDC25, MKN23, Mie89, MSS21, Pan21, PP24, PT11, ROB17, RG20, RSR23, RN22, SP22, SBS^{+}20, TH23, VBD93, Vas17, XL09a, ZD20]. free [AD19a, AD20b, ATW20a, ATW20b, AMH03, BBD20, BSV09, BM06a, CDV00, CYYH21, CPC25, CS24, DA16, DA18b, DA19, DN13, EH06, FPRA09, FJH^{+}01, Fuj99, HD22, HZ12, HSF25, HLIS16, IAH25, Ito17, KTD20, KKLD21, Li16, LL23, LB23, LD25, LMSW17, LR87, LYOI99, MD22, MNR14, Mat08, MP15, MD21, MD23b, MWYZ18, NNJ23, NW09, Pas91, SW95b, SD25, SV24, Ver93, WaZW23, XLK07, YZ21, ZBD24, ZG92b, ZOZ09, ZFZ19, ZR21, ZMF25, ZZZX26, MD23b]. free-surface [BBD20]. freedom [Mau08]. Frenet [HS24]. Frenkel [YEZ25]. frequencies [CSSZ25]. Frequency [MK20, BvG19, BCV25, HM86, KSK^{+}25, KKN^{+}17, KCW16, LT07, LCJQ12, LW20b, Par14, PG02, Pul05, SKS^{+}25a, SWL20, WWX13, ebKMZ24]. Frequency-explicit [MK20]. friction [BBD18, Gwi09, KPR06, LX21, Por17, QMLC15, QAMX17]. friction-type [LX21]. frictional [ZBNC25]. frictionless [Ahn07]. Friedrichs [LGSW25]. friendly [TS06]. Frobenius [Gab02, MdR05]. frog [TH18]. front [Gro94, SW86, WC24a, WC25]. Fronts [SM13, GOLJ25, LMPS19]. Froude [BTC23, BD22]. frozen [ZGF^{+}25]. Full [Spi00, Ben02, CGW20, Mit97, SH10, XZZL15]. full-discrete [CGW20]. Full-FAS [Spi00]. full-Newton [XZZL15]. Fully [CPR93, ER07, FID18a, Gon06, Wan17a, ZFS24, AGJ12, ALM04, Aso21, BKM13, BGG^{+}20, CCCC^{+}25, CLL25, CL01b, DS17, DLM20, DL16, GIS23, Guo15, HHL23, IS23, JLWY25, Kni94, Kni95, Lan95, LR18b, LLY21, MD20b, MW24a, Now96, PD01, Rog19, SZ22b, zSW06, SND21, jWjJ17, jWC22, WWZ25, XP23, YT25, YJJ^{+}24, ZQY18, ZLSZ22, ZYH23, ZS21a, vSK97]. fully-coupled [AGJ12]. fully-discrete [SZ22b, ZLSZ22]. Fully-geometric [Wan17a]. fully-mixed [GIS23]. Function [MRF00, AQJ18, AQ20, ABI22, AC23, Boy15, BRZ17, BP95, CCDJ20, FK23, GS17, HSS07, JD09, KKT16, KW21, KESYB23, KAS17, LN08, LSY21, Man96, Mon21, NV23, NRZR12, Pré10, RS09, Sar05, SB18, SJ11, SZQH23, uIVS13, SS09, SZW19, Sus10, Sza94, Ver93, WL09c, YWSL20, Yüz22, ZYJD25]. Functional [BNH01, COPPS25, Ere19, Jac87, AL24, BKM95, BBPR05, BCC16, Bic21, Bru07, Buc06, Cah92, CFM^{+}24, CST97, CH90, DD21, DLN^{+}24, FT06, GM08, JL91, KO92, Li05, LZ17, LLZ^{+}22, LYLL23, NT92, NH24, Par14, PH17, RTW21, RAOC18, tSqWyG16, Wan17a, WCM23, WYL11, ZK00]. functional-differential [CST97, JL91]. Functionally [HS09a, Hoa15]. Functionally-fitted [Hoa15]. functionals [AY22, BRRS15, LS86, OAHN22]. Functions [ADNR21, AG05a, AMCM09, ABY22, AR24, AGY08, ACM09, AD18a, BF20, Beh93, BO11, BG11c, BE99, CMP15, CFV10, CPD^{+}05, CWC24, CJM88, Coo89, CPZ17, Dar90, DW00, Dés08, DMGVO05, Die15, DKL24, Dra97, DGS24, EGM25, GFB99, Gem23, GS25a, Gil10, GK93, GP04, GPP04, GP17, GGS04, yGyZ07, GSW09, HZ96, Has20, HA21, HBB25, HN22, Iva07, IJ21, Jac88, Jad94, JZJ10, JCY25, JM05, Jun07, JGK11, KCS07, Kza92, LW21b, LZZ22, LWW22, DLM16, Luo18, LJ20a, Mac92, MD06, MS19, MCS16, Mar94, MS25, Meh22, MKJ23, NLS18, NSCC19, Nov03, PGS10, Pas91, ROB17, Sad96, Sad97, SRK22, Sch08a, SBS^{+}20, SW09b, SS10, SBS24, TLSS09, Uty08, VVV24, WZ14, War92, XB14, ZM19, ZDM18, ZT06, Zha07, ZHL03, Zul25]. Fundamental [ASS21, Lyn92, AMV17, GJR03, KJL12, LLHC17, LWCH19, ZLHW19, ZLCH20]. furnace [BGM^{+}09]. Further [CH95b, NN13, ABFV09]. Future [Bie87, Son00]. Fuzzy [PGC01, DHWL22, GNAS^{+}20, GAOB20]. FV [ZWN23]. FVEG [LMWZ07].

G [BtTBV87]. Gabor [GS25a]. gain [GT19a, KPRU20]. gain/loss [GT19a]. Galerkin [ADSS17, BBRS97, Cas06, DM09a, LX24, MS13, AD19a, AD20b, AD20a, AD21, ABZ21, ABJ12, AW14, AS20a, ASA20, AD08, AA22, ATW20a, ATW20b, ATAW25, An16, ASZ18, Ant13, AFS11, AD18b, AD19b, AAEMY21, ÁMS17, AV00, Bac14, Bac16, Bac17a, Bac17b, Bac17c, Bac18, Bac19, BTBR19, BTBR20, Bac21a, Bac21b, BBV13, Bec18, BVT14, BNV06, BTP96, BS20a, Bus06, Cai24, CGJ16, CW21, CGRT18, CK98, CDW23, CPC25, CL20, CCK08, ÇY22, Cul95, DA16, DA18b, DA19, DK21, DGN12, DN13, DB08, DMR18, DL16, DLZ21, DYF23, EAS12, EKT19, FM21, FD16, FJS25, Fat10, Fra16, FJ95, Fre98, GZZ20, GG22, GP23, GM18, GHH09, GD21, GS20, GÖ20, GT93a, GCZZ23, GWLN22, GJV08, HHYD20, HZ21, HLMP09, HGF24, HZ20, HQAZ24, HMP14, HJL18]. Galerkin [Hop23, HLW25, HCGW22, HH18, HS19b, HAY20, HST14, Hus20, ID19, Ito22, JRT90, JLWY25, Kam16, KDT17, KTD20, KNN03, KPY15, KB21, KSMMM16, KwS19, KWLK00, KXR^{+}04, KQ13a, KQ13b, KDD23, Kwa09, LH11, LPT94, LHS00, LH02, Li11, LZ14, Li16, LHW17, LW19a, LSWM19, LWY20, LW20a, LWW20, LL23, LB23, LD25, LGC25a, LGC25b, LGSW25, Lia22, LM21, LCH20, LT01, LS07b, LYK17, LWC24, LYMD24, LNZ12, LMWZ07, MQO17, Mai06, MSZ^{+}24, MFAD23, MD25, MWYZ18, MC21, Mus11, NN20, NS21a, NWL^{+}22, OHA25, OEAS21, PL25, Pan21, PQ25, Pło22, Pło23, Por17, PCRR17, QW25, QM20, QXG21, RP17, RN25, Sac93, SRK21, SC19, SZ09, Sha21, SW07, SL22, wSJP15, Sin24, SZL18, SW85, SvdVvD06, SW05, TWMP20, Tem15, Tem23, Top21, TS24b, VCN20, Wal19, mWyG00, WMLB19, WTY21]. Galerkin [WL22, WCL22, WJM22, WWLL23, WaZ24, WSHC20, Wen25, Won08, XZW19, XY19, XL23, XF06, Xu21, XWM26, YZ17, Yan21a, YJ21, YS22, YQCZ22, YJ23, Yan23, YZZ25, YWW23, Yua20, ZH21, ZP24, ZOZ09, ZY14, ZFZ19, ZZ19a, ZLX22, ZYJZ23, ZZ24, ZYZJ24, ZMF25, ZZX20, ZGR23, ZLG24, ZZ17, ZZC^{+}18, ZX14, ZSQ20, ZSQ21, ZX22, vR04, vdVS08]. Galerkin-characteristic [EAS12]. Galerkin-characteristics [OEAS21]. Galerkin-finite [DL16]. Galerkin-Legendre [HCGW22]. Galerkin-mixed [YJ23]. Galerkin-multigrid [CK98]. Galerkin-spectral [LT01]. game [FVGS13]. games [CZ19, SCLL21]. gap [BG06]. Gardner [ZZX19b]. Gargantini [CP94]. GARK [SGR21]. gas [EWW99, Gla94, LK07, MRS03, Pow94, Tad86, Tan23]. gases [CCP04, CS01]. gauge [Cau08, SED21, WWC^{+}25]. Gauss [KK09b, SLW17, ABD^{+}25, BLS^{+}17, CG13, DBBH14, ELR^{+}15, FMS24, GP93, GY94, yGpY09, Has08, Has09, KSK^{+}25, KS09c, Kza99, Lab99, LHT20, MSP10, NLS20, OL18, PRSS24, PR22, RE19, RS21, The17, Tom24, VV05, Wel10a, Wel10b, WX22, XZZ19, ZMY21, CKB12]. Gauss-type [ABD^{+}25, Tom24]. Gaussian [AT13, BBCR22, Boy15, Car23, CBHY11, DRS19, DDRS24, GY94, HHW18, Ioa89, Kza92, LS24c, MMP20, ST19, SdSC99, SMJ24]. Gaussian-impulse [LS24c]. Gaussian-localized [Boy15]. Gautschi [KR15]. GB [RMS17]. GB-splines [RMS17]. gBBKS [IKMM23]. GCD [CKM10]. GCRO [SMTHE22b]. GCRO-based [SMTHE22b]. GCV [Baz03]. GDE [DS07a]. GDN [ZZL01]. GDN-stability [ZZL01]. GeCo [IKMM23, MCD20]. Gegenbauer [By01, DSM11, Elg17, ER18, FK23, Has08, LLJY20, RV22]. gene [OZHP23]. General [BBO03, BB98, But85, CJM88, HM21, LHHR94, WLZ25, Zla85a, AH11, AJ19, AR93, AA87, Bac21a, BCE04, BCJW17, BCJP18, BJ20, BIJ23, BCSH16, Buc17, Bur91, BC95, BJ96, BS00b, BW03, BP06b, BD17, CA21, CIJ17, CY98, CMR12, CST18, CMCGTR02, Dav98, DS07a, EG88, FW07, Fdi97b, FS05, Gan96, GM16, GDEdLD23, GAH25, GP17, GS21, HZD21, HH10a, HS24, HZZ25, HZ02, HCX03, JMU^{+}25, JY20, JLZ25, Leo10b, Li05, LYA^{+}19, MAH18, MHA19, MAH22, Mur98, Que21, SA90, San20, Sch12, TWMP20, Wal00b, XFLC00, YH00, Zha00, ZT06, ZLX22]. general-form [JY20]. Generalisations [RS08a]. generalised [EC07]. Generalization [OXW25, ACP24, BKR13, BHHS10, BC97b, CGA96]. Generalizations [Bre91, BKP15, BNKR20]. Generalized [AHAS21, De 88, GM87, GPHAPPR23, GSW09, IJ17a, LYZ25, LFS15, Mil17, SZ17, SS16, TCCW89, TYKK01b, AD20b, AD20a, AN25, ARY23, ÁKM20, AEN22, BW21, BNH01, Bis11, BDKM92, BBBN21, BFdO07, BRBM08, But09, Cao03, CGMS21, CjW18, CJL13, Che96, CCLT10, CEW00, DS20, DGE22, DMH18, DSK12, DRS19, DBBH14, DMM24b, DSM11, FH04, yGyZ07, GR02, Har93, HW22, HD23, IM98, ILS19, JP93, KDAK13, KOS21, KAS22, KHM^{+}19, KTS03, KR25, LHH96, Li12, LZ22, LO03, Liu02, Luo18, MNSS22, MDRR11, MD22, MCS16, MS13, MD20c, MD21, MD23b, NYPW21, NR14, OB20, OGS20, OZHP23, PV93, PCA10, Pou00, RE19, RSY12, RS21, SW95d, Sch99, SLW17, SZ12, SD22a, SKO19, TW00, VRC21, jWyG08, WL16, Wan17a, WH23, Wu09, XWM26, YHT23, YR22, YWSL20]. generalized [Yua20, YRV21b, ZGL98, ZD20, ZZF25, ZJ25, ZYH23, BGG^{+}20, BWEP95, TD09]. generalized-Jacobi-function [YWSL20]. generally [WX22]. generate [CEW00, RS22]. Generated [DRC85, CGCMTR02, FKA^{+}13, GS25a, GRGJ02, GPP04, Hey20b, War92, YXB95]. generating [AEA23, CHR03, LS25, Ver93]. Generation [Bak86, Eis86, QM03, Ari03, AD04, BBV13, Bak89, BSFDM02, BS00a, Boh03, CM02, CL24b, CM04, DH94, EM05a, EM05b, MI03, MCE^{+}09, PR12, SFJ^{+}05, SYG^{+}05, Son00, Spi99, Spi00, YC00]. generator [WE99]. generators [GPíP03]. generic [Eir99]. genetics [MS25]. genus [LWCT07, NAF24]. geodesic [ZML^{+}12].

Geological [SR88b]. Geometric [CR04, Fra04b, GQ08, HL23, LW22, MQ03, Moo95a, VV07, AH09, BCR01, BG02b, CLMSS98, DF92, GGLR09, GGR97, Ost02, Wan17a, Xu13, Yan21b, MCD20]. Geometrical [BCM04]. geometrically [LA12, SS99]. geometries [FZA25, Fun94, Kar89, Kwa09, She00]. Geometry [SYG^{+}05, FdSB02, GRGJ02, MD00, MMP02b]. Geophysical [Per88, NC16]. geostrophic [MAD23, WZL13]. Germany [vdHSW98]. Gersgorin [CP10]. GEW [KOS21]. GFEM [WLZ25]. GHDMR [TD09]. ghost [CFX08, LK07, CFXZ06]. Gibbs’ [SRMDRL23, Jun07]. Gilbert [BS06]. Ginzburg [CC24, HAR21, DBH25, LHW17, LWW23, SL20]. given [CP07, Gwi09, Jac88]. glioma [LARGVR23]. GLMs [AAH21]. GLMs-based [AAH21]. Global [AMP20, FJ95, HZ20, HZC22, JMS99, LY01, LH02, LYA^{+}19, PGP03, SW09a, SXP09, AF23, BRRS15, Car94, CN11, EAV16, EHV19, FSB97, HdSRI17, JEG10, KP18, Kał22, Kim19, LS10, LMS08, RM25, SKS^{+}25b, WKM04, WKP12, ZD21, dv95a]. global-local [FSB97]. globalization [CH22, Li25]. Globally [HLMKZ06, AÁ21, AB07, BF95, HW22, Ise02, KKN^{+}13, KKN^{+}17, LDC24, NMKE13, SG04, WG22, YFLX20, ZFZ19, Zha20a, Zha20b, ZMF25, ZZZX26]. glow [FMSV07]. glow-plug [FMSV07]. glucose [MLK06]. glucose-insulin [MLK06]. GMLS [MD23b]. GMRES [AB07, Bad20, Cao97, DB97, Du11, GNNR19, GL93, JMS99, KYC03, Meu14, Mor05, PKSB10, SMTHE22a, Tou97, WZ16, ZD21, dv95a, van95]. GMRES-like [van95]. goal [AS05, XHYM22]. goal-oriented [AS05, XHYM22]. Godunov [KMR09, LE94, Mat05, SW86]. Godunov-inverse [Mat05]. Goldstein [LLY11]. Gompertz [NK24a]. Good [Zup03, EZ03, ZWH^{+}17]. Gordon [CC19, FCW20, KCY19, LSWM19, MD20a, MMDH19, MMD20, MMDS21, WDH20, WCJ23, XWZ21, AS20a, ATAW25, Bac17c, CH89, DL22b, DC18b, FCW21, HZ09, HMN20, LYA^{+}19, Lyn99, Nak24, Odi24, RV22, SLW17, WW19, WG23b, XWX21, YZQ^{+}22]. governed [AZHD23, CLY19, FS23b, HL19, LYZJ23, LY03, WZZ21, WCL22, XL11, YÇ16, ZZ19a]. governing [Rou20b]. GP [GPP04]. GPBiCG [AS13, Fuj02]. GPST [Che88, CZ90]. GPU [BBC^{+}25]. GPU-accelerated [BBC^{+}25]. GPUs [CP17]. grad [BVV09, LFQH21, MV18]. grade [Cim25, Wu09]. graded [BLM17a, DL06, Mai09, YZ17, YZ19, Zar17]. Gradient [SD25, ATW20b, AF23, ABY22, AKA19, BSGU94, BBBK22, BGH08, BIMV19, CW20, CHSS01, CR05, CSL^{+}24, CCL04, Cui24, DW00, DW15, EAS12, EKT19, EGH01, GM08, GO18, GH21, GR02, HVY91, HWY20, HZC22, HZCZ23, HS97, Kie15, KKLD21, KLS13, LWW20, LWLW24, LZW20, LY24a, MK19, Pfl08, Ria22, SSW20, SSS^{+}23, SKS^{+}25b, San03, SW95d, SYW22, Wan23, WQS^{+}24, WOLY25, WL25, YFLX20, YS24b, Yu08, YLS^{+}09, YLH20, YLW20b, ZY23, ZGF^{+}25, ZZF25, ZZ25, vdES04]. gradient-boundary [BGH08]. gradients [AGJM04, LYZJ23, Ren13, SB18]. Gradual [FT06]. Grail [Bur93b]. granular [BBL02, TSFB01]. Graph [BT94, ABD24, Gar03, GM17, LTDY24, MM20a]. graphene [LY24b, NT20]. Graphics [CB99, Zar99]. Graphics/Cray [CB99]. graphs [BDFK95, DE06, FRRZ25, Obe15, ZJR25]. grating [KN19]. gratings [NT16, Rat13]. gravitational [FGGL22, Lei02, Maj20]. gravity [AD15, GD09, HD88, KTK20, NC16]. gray [GS20]. GRBF [Hou23]. greatest [BWY17, BWS21]. greedy [Bai25a, LG19, Liu24, DE18, ZG20]. green [KKK25, Arc06, CA15, DL16].

Green-Naghdi [DL16]. Gregory [CL06]. Greville [Joh05]. Grid [Bai97, Eis86, MI03, Pet87, Son00, AZHD23, AGM95, Ari03, AD04, ACM91, BSFDM02, BV94, Boh03, Bor97, BB10, CM02, CL10, CWHF19, CYWH22, CDW23, CWZ23, CM04, CMS06, DH94, EM05a, EM05b, FD97, FL04, Fer96, FSB97, Fou00, GLV06, GKT10, HS86, HCY18, HL19, HJYL19, HY01, HCW16, IM00, Jun97, Kni94, Kni95, LKV01, LCHW20, LLZL24, LHY24, LSG24, LGH11, LYOI99, LE94, MMT90, MR94, MT05, MM20b, NN13, OZ96, QLL^{+}08, Qui96, Ram94, SJ20, SW20a, Shy86, Shy91a, Shy91b, SR24, Spi99, Spi00, Tow16, Tro93, TY00, VT93, Wag98, WH19b, WTF25, WSS97, WWWW25, XY25, XGQ20, XXF22, YC00, YWH20, YCWH23, ZTZ15, ZZHS18, ZBY19, ZC92]. grid-based [LYOI99]. grid-orientation [ACM91]. Grids [SM85, AG05a, ASCM02, ASC03, APJ09, Bac17c, Bac21b, BKP09, BM00, BCE04, BS94b, BS96a, BTDV10, CGCMTR02, CH19, Chu03, CS18, DF11, DT10, DGD03, DSSC13, EL94, FD16, FS88b, Fre04, GGLR09, HP97, HZ02, HS97, Imo00, JTB15, Kal96, KWLK00, KTS03, Koz94, KKE16, KQ13b, LRC19, LR20a, Lin01, MRF00, MS90, Mit97, MKS12, PRGO16, PC00, RGL16, Rus95, Ste05b, SN04, SGN06, Tho85, Tol04, TJ12, VL08, Vic87b, VT91, WC24a, Zha09, ZLL22, ZML^{+}12, dAF17, dVA02, iV09]. gridsize [Hol01]. Gross [LCW20, WLCZ26]. ground [Bec02]. groundwater [DMQ02]. group [BK17, TM21, YF24]. groups [LP00, MA09, WGKS12, WC02]. growing [GHW20, ZW24]. growth [DLS22, CL02a, CL02b, CDP^{+}25, GeO24, HZ96, JK17, KSJY25, LM25, LBLT13, LQXK23, LARGVR23, NK24a, RA17, SS25, SL21]. Grünwald [CHS19]. GSAV [LS23]. GSOR [Zen21]. Guaranteed [Kim21, ZCZ15]. guarantees [HZCZ23]. guess [Boy07]. guide [BLJ21]. guided [GKB^{+}22]. Gurtin [YZC21].

H [BtTBV87, BFP11, Hua98, Ree03, YZ21, Zho18]. H-div [YZ21]. H-form [Ree03]. Haar [AAL^{+}26, JCJP21, LS16, PK23, SSS21, VRC21]. Habetler [SB25]. Hadamard [AC10, BAA22, DO17a, GS21, KS10, LM25, OCVW22, Sid23, TQY24, WWLS08, ZHS22]. Hadamard-type [GS21]. Hagen [MM07]. Hager [SSW20]. Hahn [BMPR15]. Hakopian [SL17]. half [Abr93, AHO16, BH93, BT00, GHKM09, HMY19, HY24a, IV16, KM19, MG22, OT24, OT25, Ost93]. half-axis [GHKM09]. half-block [BT00]. half-explicit [BH93, Ost93]. half-line [AHO16, HMY19, HY24a, OT24, OT25]. half-space [IV16]. half-spaces [Abr93]. half-step [KM19, MG22]. Halley [EH07b, KYI17]. Halton [Sch08b]. Hamburger [Njå88]. Hamel [AI19]. Hamilton [BGS06, BL06, CGT13, CXNF14, CFS13, CL01a, CY05, FLS94, For11, JZZH22, KHYY21, LHS00, RF16, SM13, Sti03, YXZ18]. Hamiltonian [ABK12, Ant23, Ant25, ABD16, BL05, BDF^{+}25, BF99, BGIW18, BD17, CM00, CS01, CBD16, Eir99, Fil25, FHX22, GBDB97, HL99, HHW18, LW19b, LFS21, LXLY25, MM25, PA05, Ror06, SG17, ZJ19b]. Hammerstein [BBCR22, DS21b, EH09, EHV24, Han93, HY24b, KXK92, KNN03, KM25]. Hammerstein-type [EHV24]. hand [AED12, EJS04, Shi20, ZD21]. Hansen [PSR04]. harmonic [FJ09, GH07, Har98, Jia02, MM18, Pla08, QDA^{+}26, RV09, ST19, WSY18]. Harvard [RW87]. Hasimoto [HS24]. having [CL88, DGS24]. HDG [ASV19, CL26, MLJ19, SZE20, ZZZX26]. Heat [MCM12, AX19, AL22, BLS94, BS21, BCFQ19, BN12, Bor02, CCDJ20, CM09, CHNN20, DCN^{+}19, DMA22, FP02, Fre98, HILK13, Hor99, Hor02, Jéz99, JL23a, KV07, KBS11, Kra92, LW18b, LGZH24, MVVA09b, MW24b, RMK09, SKBAS08, SM20, TOD11, TY03, Wag98, jW15, WS04, YY13, ZZC^{+}18]. heat-diffusion [KV07]. heated [Wu09]. heating [BGM^{+}09, WDZS21, YJ23]. hedging [Bis11]. Heisenberg [Bec02]. Hele [GLML20, GJLL20, WWZJ22]. Helmholtz [AAD^{+}08, BOEP00, BPTT15, CH01, DS07c, EVO04, EVO06, FFQ09, Gol86, GLLY25, Guo96, HDY21, HO16, Kim94, Kim95, KR20, LRS09, Li08, LWCH19, LHC23, MPTT17, MP20, NTT22, OMP98, OKS10, PG21, PM03, PB10, RS00, WSC21, Yua20, ZLHW19, dVA02]. Helmholtz/Schrödinger [BOEP00]. hemodynamics [DZ12a, RDH^{+}12, RVM23]. Hermann [Bru97]. Hermite [AJK20, Bar12, Boy15, CXNF14, CjW18, CSX23, CCS17b, CN15, CS17, FMS24, GS25a, GS21, Iva07, KHLV22, KK23, LHHR94, LHH96, LHH08, Luo18, MS24, MST07, NN20, NBNTGV11, PRS23, ZL11a, ZZZ23, ZL11b, Zha21b, ZHL08]. Hermitian [DS21d, GGMP88, LVfP14, LXCM21, Nak12, Pet92, SH21a, Zha21a]. heroin [LZY24]. hesitancy [CML^{+}25]. Hessenberg [GS08, ZP98]. Hessian [MKH16]. Hessian-based [MKH16]. Hestenes [HWY20, Wan23]. Heston [BDOG19, GGO13]. heterogeneity [Chn17]. heterogeneous [AK09, Ben17, CM09, DCSB25, DN13, DG22, IJ14, JL25, LP01, MPSS16, MPV24, MLJ19, OL18, OH20, RLMG24, RN04, SS94b, VBVA22, WCW14, XZL07]. heterovalent [vHA98]. hexagonal [DD20]. hexahedral [BS97a, Tob14, Ush18]. HHO [KSJY25]. HHT [KR25]. HHT- [KR25]. hidden [RR14]. Hierarchical [AB15, AD04, CM09, CZ90, FSB97, Haa97, Kwe00, NBP94, Zou10, BLS94, CLT97, Che96, EH08, GV18, Gar96, GKB^{+}22, GS92, KJ99, Ste97, Zou11, GMS12]. hierarchy [Che88]. Higdon [BG11a, Dea11]. Higgs’ [MPMD21, ZH21]. High [AJ24a, AES15, AEMX17, BBD20, Bla01, BCR01, Bra22, BT95, Bru07, BB96, CXNF14, CW25, CSSZ20, CM13, CDP17, CP05b, CKS05, CRU15, CS18, CB99, CDR20, CRHTV25, DEGDdL25, Dit21, DRC85, DS15, DCL23, EBEHAS24, Elg17, ER18, EP15, FS23a, Fou00, FHX22, FGGL22, GS19, GZQS23, GMG04, GLPW09, Hag15, HEG16, JS24, JZZH22, KL07, LXZ21, MPTT17, MG22, MPDB24, MK99, NW09, NAD25, NT20, PLB22, Pir09, RR00, SST12, SMB23, SB14, SZE^{+}92, SND21, TSF24, TDC13, TDW23, TD09, Ver96a, Wan21, WLG22, YZWQ25, YZG23, YYZ23, ZSQ20, ZSQ21, Ale11, AAL^{+}26, ABI22, BMGM12, BG11a, BOEP00, BMR^{+}17a, BM01, BSV21, BVV09, BTC23, BJ11, BR20, BPTT15, BO21, BDM03, BJ98, CC24, CGA93, CC19, CL24b, CS08, CD00, CN15, CN17, DV20, Din19, EW97, FJ17]. high [FHM^{+}02, GLLW14, GHS25, GÖ20, GML00, GHF00, HOEC86, HZD21, HGZW21, HL21, HR97, HN22, INR01, JM17, JUAZ22, JCL18, Jia12, JWG20, JQSC22, KWY24, KN19, KS00, KM19, KK09b, KwS19, KSK^{+}25, KOGL25, KDKW20, Kwe01, LCVG01, Lee94, LLKJ21, LX08, LYF17, LR18b, LH21, LHC23, LH24, LL21, LS24c, MS91, ML91, MD96, NT16, ÖT20, PSB91, PBC08, Pic05, QW04, RMH20, RGA19, Rou20b, RGK21, SKS^{+}25a, San02, SD13b, SQ17, SYG^{+}05, SL21, SDK15, Som93, STS00, Str98b, Sub04, TX18, TY00, VV02, Var92, WKM04, WR20, WDU21, WLM21, WTF25, WC14, WS22, XWW19, XG22, XFL22, YF24, YR22, YZH24, YXZ18, Zha96, ZYQS21, ZG21, ZYQS23, ZFX25, ebKMZ24, van98, vdHMdS99, NNJ23]. High-accuracy [DS15, LH21]. High-degree [PLB22]. high-dimensional [HN22]. high-frequency [KSK^{+}25, SKS^{+}25a, ebKMZ24]. high-index [San02]. High-level [CRHTV25, PSB91]. High-order [AJ24a, AES15, AEMX17, BCR01, Bra22, BT95, Bru07, CW25, CSSZ20, CM13, CRU15, DEGDdL25, Dit21, EBEHAS24, Elg17, ER18, EP15, FHX22, GS19, GZQS23, GLPW09, HEG16, JZZH22, KL07, LXZ21, MPTT17, NAD25, NT20, SST12, SMB23, SB14, SZE^{+}92, SND21, TSF24, Ver96a, Wan21, WLG22, YZWQ25, YZG23, YYZ23, ZSQ20, ZSQ21, AAL^{+}26, ABI22, BG11a, BR20, BO21, BDM03, CC24, CGA93, CC19, DV20, Din19, GLLW14, GHS25, GÖ20, GHF00, HOEC86, HZD21, HL21, INR01, JCL18, Jia12, JWG20, JQSC22, KWY24, KN19, KK09b, Lee94, LLKJ21, LR18b, LH24, LL21, LS24c, NT16, ÖT20, San02, SL21, SDK15, Som93, STS00, Str98b, TX18, VV02, WKM04, WR20, WDU21, WTF25, WC14, WS22, XWW19, XFL22, YZH24, ZG21, ZYQS23, ZFX25]. High-performance [CB99]. high-precision [Var92, vdHMdS99]. High-rate [MPDB24]. High-resolution [CXNF14, Hag15, Pir09]. high-speed [MS91]. Higher [Aca12, Agu15, BM05, DRVA20, EES05, Gar96, Joh05, KC19b, Kha21, NK11, Vej10, WL09b, ZX14, Arn93, BBS25, BKR13, BNV06, BC05, BD11, CZ97, CJX11, DKSS24, ELCWS98, ECB07, FM11, GJ00, GND19, HT24, HAN23, HO24b, HvdHV10, HST14, KOS^{+}12, KK22a, LMWZ10, LYZF25, MNSS22, MS08a, Mat09, Nic86, RA03, SDK24, SM93, SS13b, Tem15, Tob14, WDH20, WB03, XY24, YM24, ZL24, ZJR25, Coy12]. higher-dimensional [ZJR25]. higher-index [ECB07]. Higher-order [Agu15, Gar96, KC19b, Vej10, WL09b, BBS25, BKR13, DKSS24, LYZF25, MS08a, RA03, SM93, Tem15, WB03, XY24, YM24, ZL24]. Highly [Ben96, Fac03, IRC12, IJ17b, JP17, JJL^{+}24, LAH09, WPN^{+}24, ACP24, AK09, AS20c, BBV13, Cai24, CSSZ20, CC04b, Che12a, DIJ12, DYX09, Den93, ECHF^{+}20, GHHG22, Has13, Has20, Ise02, IKM23, IR22, KDS22, Li01a, LWWX10, Li23, LVW21, LTT19, LW20b, MK20, Maj17a, RCGM98, WLCZ26, XFG19, YH18, ZYX20, ZJ19c]. highly-oscillatory [Ise02]. highly-stable [ACP24]. Hilbert [AMCM09, BLM17a, DO17a, DO17b, DP21, GNAS^{+}20, GAOB20, JL94, OT24, SAA20]. Hilfer [SSA^{+}22]. Hille [Lub92]. Hilliard [ASGGRGW25, BEH24, GLML20, GHS25, GJLL20, JJL^{+}24, LXPY25, WaZW21, WWZJ22, AZ23, BMWH20, CGH23, CCZZ18, CW22, CCL04, DLQZ23, GGT24, HLT07, KK09a, SC25b, WWL21, ZCY20]. Hilliard-Cook [CCZZ18]. Hilliard-Hele [GLML20]. Hilliard/Allen [BEH24]. hillslopes [IMC22]. hinged [Ito22]. Hirota [Nak24]. histopolation [KOS20]. historical [Bre96, BW96b]. history [But96, TCCW89, ZC92]. HJB [AS20b]. HJB-POD [AS20b]. HJM [Mar09]. HLL [Sch16a]. HLL-type [Sch16a]. Hodge [AM10b, BS14a]. Hodgkin [Par21]. HOFiD bvp [Set24]. Hohenberg [DGE22, QH22, Qi24, XXYZ24]. Hölder [COPPS25, LS10, LSW23]. Hole [FGPR12]. Holistic [Rob01]. hollow [LR19]. Holm [AS06, CLP15, JWG20, NYPW21, QR24, RLSS06, ZZ18]. holomorphic [Dar90, LFP04]. holonomic [BGIW18]. Holy [Bur93b]. homogeneous [AMV17, BP12b, MH14, PFHL09, RLMG24, WG19, YM24, ZCGS21]. Homogenization [MS00, LP01, RCGM98, Shi20, YXZ18]. homotopies [BZ17b]. Homotopy [AI19, YCY12, CH21, LW18b, Odi19, SL08, TWH21]. HOOI [XY24]. Hopf [JMDN^{+}22]. HOPS [NT16]. hopscotch [BtTBV87, Duf90, HV89, tV87]. hoRA [GT18]. horizons [MVG14]. Horizontal [CMP06, CML05, GRGJ02, QR03]. Hormann [CHS17]. HOTV [San18]. hp [ADGM08, BF15, BS14b, BPS19, BH10, BD11, Bür13, CZHX19, DN21, Dol14, DMR18, DJ08, DH07, DM10, GS20, GW20, HH18, KPY15, KWLK00, KFOF02, KR10, LWY20, MS05, Moo04, NDM20, PZMX16, SA00, WTY21, dSFDG20, LX24, YMD21, ZMY21, BCV25]. hp-adaptive [BS14b, Bür13, Dol14, dSFDG20]. hp-BEM [BF15]. hp-estimates [KPY15]. hp-finite [BPS19]. hp-mesh [DMR18]. hp-refinement [Moo04]. HP-splines [BCV25]. hp-version [DN21, GW20, HH18, KFOF02, LWY20, NDM20, WTY21, YMD21, ZMY21]. HSS [BY09]. HSS-based [BY09]. Huggins [GJLL20]. Hull [GGO13]. human [OHA25]. hundred [Bru97]. Huxley [Par21]. Hybrid [AC98, CHLX07, KPR06, PH91, Rha99, VA21, ZYZJ24, AKM^{+}22, AP16, AP20, AA87, AD04, BDMG12, Bai96, BNS25, Baz03, BCSH16, BD22, CH22, CXZ17, CSL^{+}24, CPC25, Dal00, DD19, De 06, FHK05, Fuj02, GFB99, GNAS^{+}20, Hua09, JT18, Kal96, KD13, KWLK00, LY01, LH02, LK14, LCS19, LWLW24, LS99b, Lte24, MZJ^{+}25, MS19, MK21, MK19, Nak05, NV23, Ou11, PH15, PMP23, RY13, SA00, wSJP15, Sv95, TC22, TYJ11, Tou97, VB07, Wan25, WM08, WTF25, WL25, XFL22, ZBD24, ZZPJ23, ZLG15, ZGDL17]. hybrid-line-and-curve [CH22]. Hybridization [BPS19]. hydraulics [HS98]. hydro [CPY20, LY16]. hydro-dynamically [CPY20]. hydrodynamic [BGM^{+}09, DSW96, GNUV24, JT02, JT06b, RGK21, WaZ24, YZH19a]. hydrodynamics [CDD00, LL06, Qiu23, Qui96, RI02, SKR^{+}16]. hydrogen [CPD^{+}05]. hydrologic [RVdCVR02]. hydrostatic [DL21b, FGGL22, WLG22, WDL23]. Hyers [KS25]. Hyperbolic [De 88, AM09, AW14, AA22, AAEMY21, BJ02, BF17, BW95, Ber04, Ber05, BTP96, BM12b, Bor16, BGP11, BR20, CCL22, CJ18, CSW19, CRU15, CR19, Chi93, CL09, Dav92, DGRS09, FE93, FJS99, GQ89, Gas92, GK19, GDS^{+}15, GJ17, GHF00, HMD21, HT19, HWCF15, IT16, JJ94, JNPC03, KPY15, Kop86, KXR^{+}04, KR12, KW10, LJ20b, MD20b, MS08a, MG22, Mon21, MC21, NB01, NFAE03, PGDB08, QZM25, RZ00, RGB20, Ric91, Rog19, RMK09, Sea09, SYL^{+}20, SZQH23, uIVS13, Sof17, Str98b, gTpM07, TDW23, TJ12, TM15, Vic87b, Wal00b, WDL23, XF06]. hyperbolic-heat [RMK09]. hyperbolic-parabolic [FJS99]. hyperbolic/parabolic [AA22]. Hyperbolicity [LY16]. hypercomplex [AMT17]. hypercube [CSS87, De 93b, IKS25, LS86]. hypergeometric [AR93, CJM88, DIR13, KW21, Zul25]. hyperparameter [DESZ25, LV12]. hyperplane [KLSW10]. hyperplane-constrained [KLSW10]. hyperplanes [Mon09]. hypersingular [AFF^{+}15, BVB09, BVB10, BVRB14, BRBB18, Cai09, Mai09, Sid14]. hypersonic [KH91]. hyperspectral [Lo06]. hypoacusia [BF09]. hypoelliptic [AY22]. hysteresis [LR01]. hysteretic [FS23a].

IA [AC96]. IBVP [Str98b]. ice [CDV00]. Iced [KTS03]. ICOSAHOM [Ano00a]. ideal [CFXZ06, GDS^{+}15, QM10, TDW23, Tou10]. idealization [SW94]. identical [BASC17]. Identification [ABP95, BBPR05, CAD03, ND85, SZQH23, SW17, VSG17, AX19, AGKK94, BBV05, BBCS05, BR94, Die15, HM17, HLIS16, HK09, KN19, KJL12, KK06, LDP^{+}14, MCM12, NYPW21, PKP19, SM08, Saz22, SHB^{+}25]. identify [WZW13]. Identifying [TDMT21, HP14, HM15, Quy19]. identities [AR93]. identity [Bre91, MV20]. IEFG [AD19a, DA19]. IEQ [FCW21]. IFE [ZFW20]. II [BDDV12, BDMGVO05, BH12b, CL24b, DO98, GHH09, GPHAM12, KQ13b, LH02, LFS15, Mar03, MSP10, Mur99b, SMTHE22b, Shy86, Shy91b, Tsa92]. IIA [GPHA16]. III [DGN12, VC10]. ill [BHL^{+}21, BBBK22, CRS05, EG88, GNNR19, HDY21, KMH21, Kli15, LHT20, Luc95, MRH14, Sam94, TWH21, TWD23, WQS^{+}24, XXQ17, ZD21, ZGF^{+}25]. ill-posed [BHL^{+}21, BBBK22, CRS05, EG88, GNNR19, HDY21, KMH21, Kli15, LHT20, Luc95, MRH14, Sam94, TWH21, TWD23, WQS^{+}24, XXQ17, ZD21, ZGF^{+}25]. Illustration [BSTT22]. ILQR [XG22]. ILU [Zha00]. IMACS [CFTW08, HSX18, MH89]. Image [MMKN17, ABD24, BCM04, BS10, BHRY21, CP06, CW20, CF13a, CH87, CSL^{+}24, DE16, FP24, EEJB22, GLW^{+}25, Han19, HWY20, KM95, LK14, QM03, RU21, Sae14, SC08, SMA01, SL17, SH21b, SC22, Wan23, WQS^{+}24, WOLY25, WYP12, YLH20, YLW20b, Zen21, ZC91, ZZX19a, ZN21]. image-based [QM03]. images [GOGF03, TSB10]. imaginary [SD22a]. imaging [CG03, HLR18, KNP16, KRBK16, Lo06, Par14, TT20]. imbedded [CST97]. IMEX [BTC23, CCS17a, CSW19, Dit21, JM17, LT12, LWWZ25, MPPR22, Ort20, RBdM25, ST14a, SA21, SKW17, SW18, TTP24, WCM23, ZL23, PYD21]. IMEX- [Ort20]. IMEX-L1 [TTP24]. IMEX-SAV [ZL23]. Immersed [CYWH22, ZLW20b, AL09, JK14, KV07, Li98, SWFK13, WZ19, YCWH23]. immigration [GeO24]. immiscible [Nür09]. immobile [GWLN22, JL17, LCZ25, NWL^{+}22]. immune [BR94]. impact [Ahn07]. impacts [Aca12]. impedance [YL26]. imperfect [ZSY20]. implementable [BC89b]. Implementation [AH15, CO09, GT93a, Jac02, LPT94, Pot85, AHJ^{+}23, AQJ18, Ale03, BD85, BF15, BS92, BS08, BMT93, BT98, BM02, BIM15, BWEP95, CLMSS98, CCP17, CK06, Coy12, DVV93, FV99, GV18, GPíP03, HS96, IM98, IMMS20, KN19, LT07, MP97, NS20, OMP98, PGS10, PRST02, Pel20, PWS06, SW94, SWL20, Uty08, WMF17, XWX21, Zha01]. implementational [GKT10]. implementations [DLP06, PV93]. Implementing [HP91, AEA23, Ney95]. Implicit [AS04, ARS97, BBT25, BIMV19, CR05, DGM18, FV01, GM16, Gje07, GBBC^{+}23, GHF00, HJ03, KBS11, Kau93, Kau95, KW98, MAH22, OK98, Pul86, SWE05, TB01, WBCK02, ZD25a, ZZX20, ZSZZ20, AD20c, ABH14, AHT17, AMCM08, ACM91, BT97a, BC08a, Bos09, BSvdV99, BJ11, Bou02, BN03, BCJP18, Bru93, BM02, BM06b, BJ06, BD22, But93, BJ96, BC97b, BJ98, BC00b, BW03, BSW93, CCCC^{+}25, CdFN01, CG05, CS04, Cha96, CWX21, CS24, CL09, CDP12, CRSF19, CGGM17, DR09a, DK11, DZMB21, DII15, ELCWS98, Ein18, ER07, FD16, FE93, FWHM20, FS08, FHV97, FMU15, FCW21, FL01b, GX11, GLPW09, GPMR95, GPMPR03, GPHA06, Gug05, Guo01, HV22, HH10a, HLJ20, IM02, ID19, IJ17b, JMDN^{+}22, JVZ95, JR00, JM25, JKW12, JHGZ20]. implicit [JLZ25, KKT16, KM19, KC19a, KMG09, Kni94, Kni95, KKW00, KS09c, KW20, Lab98, Lay08, LL15, LX08, LMA18, LL19, LA21, LGC25a, LGC25b, LLZ19, LL20b, LYZW22, LXPY25, MMDS21, Min04, MD20c, MG22, MAF20, MSA20, MT20, MPMD21, NS21a, ÖPT25, Pan07, Phi87, PWS06, QWX20, RZ00, RBdM25, RGMO19, San02, San20, SGR21, SA21, SMB23, SH09, SMEN04, Sch98, SS21, SAH24, SZ97, ST25, Spi95, Sza94, TWL23, TYJ11, TZA13, VA21, VDVV98, VBH96, WC11, WWM22, WSW96, WdG92, XWW19, XWZ21, YX25, YZWQ25, YZC21, YT00, ZLX22, ZCQ25, ZZLL21, in 02, vSC92, vS93, vSK97, vvdV97, vdHS01]. Implicit-Explicit [DGM18, BIMV19, GM16, ZSZZ20, Bos09, IJ17b, KKT16, NS21a, San20, ST25, VBH96, YZWQ25, ZCQ25]. implicit/outflow [MOU14]. Implicitly [Bru92, HY24b, Att97, HS22, HC01, Jia02, KBG04]. importance [SA05]. imposed [GZZ20, HJ09]. Imposing [ZW87]. imposition [ADG^{+}24, LZ20, jWqW09, hYqW12]. Improved [CHZ21, CX08, DKSS24, DP21, FPRA09, LC24, Pot85, Ver06, WG23a, AC15, BZ17a, BRBM08, CL07, GL93, LWWX10, MZZ17, ROL19, VN21, WC24b]. improvement [CP94, CZ90, DDGN23, Kor95, MPPR22, dFN00]. Improvements [MS86, Tan23, Gil10, WKN20]. Improving [Bal00, Dea11, Gen10, KS10, KP03b, Sal89, MM14]. impulse [CCS02, LS24c]. impulses [LFL14]. impulsive [ABRW18, BKM95, HW04, LSL11, LZL14]. impulsively [TOCV02]. incidence [WN12]. including [BBS11, BFH09, Cao98b, TS23]. inclusion [CP94, PH91, RTH23, SC22, TLGC22]. inclusions [Par14]. incoherent [CCP04]. incompatible [GO19, GO23]. Incomplete [Guo96, ZNK02, de 95a, BD85, BCSH16, Doi91, EVO06, Gu01, Not99, Phi91, RV05b, ST09b, SS10, XC20]. incomplete-data [RV05b]. incompressible [AD20b, An20, BRS05, BC01, BC04b, BBL02, BL08, Cai15, CHOR19, CH07, CHX13, CSXL14, CCZ22, DK20, DN13, DLM20, FD16, GM08, GHK16, Gat91, GNX19, Guo15, HJR22, HSS04, HS22, HH10b, HST14, KSMMM16, KLY05, KDK17, LR24, LD22, LR00, Min04, Mou03, QAMX17, QM19, RK91, SQ17, She96, SED21, TH18, TLV92, VG04, WWM22, YT25, YHT23, ZBD24, Zan91, ZOZ09, Zha14, ZYZJ24, ZZZX26, ZGR23, ZS21a]. inconsistent [BW23b, Liu24]. Incorporating [Tom24]. increases [GT18, RAS99]. Increasing [DDNZ18, GMGF02, Wal90]. Incremental [CT93, Gar03, GGM07, FS19, Gar96, MM02c, Pou00, PB10, SW13, Che96]. indefinite [CK98, FL05, LMV17, LW18a, MBS23, SS99, Sha98, WM22, WL24, ZX22]. independent [CH95b, KP18, ST19]. Index [Ano00b, Ano00c, Ano00d, Ano00e, Ano01a, Ano01b, Ano01c, Ano01d, Ano02a, Ano02b, Ano02c, Ano02d, Ano02f, Ano03a, Ano03b, Ano03c, Ano03d, Ano04c, Ano04a, Ano04b, Ano04l, Ano05a, Ano05b, PSL18, An16, Ano92, AFS00, Arn93, AC96, BHSW20, CZ97, CCM02, ECB07, HGP11, HMT03a, HMT03b, Jay95, KKR15, MNSS22, Mur99b, Ost93, RA05, San02, Sch02, SG05, SBG09, Wen98, Wen05, WBCK02]. index- [AC96, CCM02, HGP11, Mur99b]. index-1 [HMT03a, SG05, SBG09]. index-2 [AFS00, BHSW20, HMT03b]. index-3 [Wen98]. Index-analysis [PSL18]. indexing [MMBB07, TS08]. indicator [KMS10, RGB20, WG23a]. indicators [GMS12, Tan24]. indices [FP24, LS98, LD02]. indifference [GK22]. indirect [WZZ21]. Induced [BC02, BBD20, DDZK05, DRS19, KCJP01, SSvG10, WT17, XLK07, YT03]. induction [BGM^{+}09, SC19]. inductionless [LD22, XY25]. industrial [BKP14, LDP^{+}14]. iNEOS [GGNP02]. Inequalities [DR01, AK95, BNV06, BHR05, DSM11, DP21, GS21, GH20, IS23, JMDN^{+}22, JMU^{+}25, LLY11, LQY25, MMP09, RTH23, SI20, TLQ21, TGV22, VA21, WH13, YXZL24, ZHJ14]. inequality [CXZ14, DHS05, Gwi09, Lot19, MG00, OMSS25]. inequivalent [KS10]. Inertia [For02, AC18, Ran20]. Inertia-controlling [For02]. inertial [IKR^{+}22, IAH25, IS22, JMDN^{+}22, JMU^{+}25, MZJ^{+}25, OMSS25, SI20, SLMD21, TLQ21, TLGC22, WH23]. Inexact [GH21, HFL13, LZH^{+}25, Zha21a, dOF20, AMP20, BHSW16, BHSW20, CH22, CLL23, CMP23, GO18, HMW05, MBS23, PKSB10, SL09, THW19, WZ16, YH18, Zha97, ZP12]. inexact-Newton [WZ16]. inf [Che16]. inf-sup [Che16]. infection [MPV24]. infiltration [BDNV19]. infinite [AD08, DLS22, DT15, Dés08, EC07, Has13, KKP17, LRT99, RG20, jWyG08, jW15, Zhe07, ZR15]. infinity [By01, IMMS20]. Inflow [MOU14]. Inflow-implicit [MOU14]. Inflow-implicit/outflow-explicit [MOU14]. Influence [Win92, AQ00, BP95, SS02]. information [CMP23, KNP16, Pro24, SA05]. information-based [Pro24]. informed [CDP^{+}25, SOA^{+}25]. infrared [SS16]. Ingersoll [AHT17, KL23b, WMC09]. inherent [AJ19, KFR25]. inherits [CFLW22]. inhibition [GS15a]. inhibitor [ZZJ21]. Inhomogeneous [Gus88, CGN03, DZZZ24, DE06, EVO06, Kur98, OS12, QL16, ST20, jW15]. Initial [Bic16, De 88, TOCV02, AB15, AMP03, AMCR17, AT93, Bac17a, Bac17b, BY00, Bur93a, BT00, BC89c, Cah89, CL85, CHS19, CKK25, Ehr08, EHM01, FW08, FD97, Fer93, GGM95, GO19, GO21, GO23, GH02, yGpY09, HHAA22, Hig93a, JR02, JCSR03, KHM^{+}14, KOR18, KKW00, KW93, KK20c, KDS22, KDKW20, LW17, LDC24, LS07a, LL21, LT93, LOM98, LYA^{+}19, MS03, MS90, Muo23, Nap16, NNJ23, Odi19, OB20, Odi24, RCGM98, SAG86, SWE05, XFLC00, ZJH18]. initial-boundary [GH02, KK20c, LOM98, MS90, Odi24]. Initial-Value [De 88, Bac17a, Bac17b, CHS19, CKK25]. initial/boundary [Fer93]. initial/periodic [HHAA22]. initialization [HdSRI17]. initio [DC21]. Inner [GCHR06, VVV24]. Innovative [UHLZ24, GS24]. input [Har10]. inspired [Yam23]. instabilities [Cul95]. Instability [CJV88, SD24b, CSS19, DG96, Lyn99, WT17]. instance [KSP10]. Instantaneous [CHK99]. insulin [MLK06]. integer [ACKV24, KS91, Mit22]. integer-order [ACKV24]. integrable [Ma24, NAF24]. Integral [AS21, BBLT15, BB24b, CP09, HS21a, PGM86, AHJM19, ABH22, AN25, AL95, ADR17, All24, AAL^{+}26, AHB20, AM10a, AAD14, AD18a, AD18b, AD19b, AFF^{+}15, AS20c, Aze22, BKM19, BC02, BES18, BBCR22, BLN25, BLM17a, DSM22, DLS22, BVB10, BVRB14, BRBB18, BGIW18, Bru97, BMM97b, BIS25, Cah92, Cai09, CHLX07, CCP17, CDP17, Car23, CJ18, CL14, CHH15, CZHX19, CS09, Chi12, Chr01, CP17, CMPP24, CFM^{+}24, CJ22, Cum95, DS21b, DD97, Dav98, DN21, DHWL22, Der92, DSK12, EGV25, ER18, EH09, EHV19, EHV24, FS15, FMS18, FJS25, Fer09, FL15, FV87, Fun04, GAH25, GKA17, GGS16, HB02, HGP11, Han93, HP14, Heu00, Hey19, Hey20a, HBB25, HA16, HDS20, HS23, HK85, HX11, Hu99, HLL09, HY24b, IV16, IRC12, Kan89, KX91, KO08]. integral [KME20, KM25, Kel85, KPR12b, KR18, KKK25, LRS09, Lau17b, Li11, LW21b, Lia22, LD97, LKJ20, LM22a, LNZ12, LYC24, LRE04, MK20, MAHZ21, Mai09, Maj20, Maj14, MKDC25, MO17, MS19, MK21, Man96, Man97, MR20, MV17, Mie89, MSS21, MD10, MKJ23, MPDB24, Naj20, NDM20, Ngu15, NT16, NSD23, NLZB23, ORT24, Pan21, PR89, PP24, PTV20, PV24, PA18, Pis22, PS21, PRS23, Rab94, RG20, Ril92, RAOC18, SST12, SHL19, SN22, SV00, SVB17, SP22, SPYS24, SS16, SW85, TMD92, Tau09, TH23, TLSS09, VGL24, WN12, WQ17, WTY21, WHW21, WW24, WWLS08, XL09a, XZH19, YT21, YS24a, Zak20, ZAED21, ZL18b, ZHL22, dAF17, de 93a, CLR11, CDP12]. Integral-algebraic [BB24b, BIS25, Pis22]. integral-differential [Chi12]. integrals [AA05, BCU00, BVB09, CPD^{+}05, Che12a, CEW00, DN24, Eva94, EW97, EC07, Fat10, Fat12, Har00, Has09, Has13, Has20, Ioa89, Jad94, KM17, LG21, LMO24, LS16, LWWX10, Maj17a, Mot17, PCA10, SL01a, Sau00, Sid14, Sid23, Yam23, ZMY21, ZPT92]. integrand [CDR20, Tom24]. integrands [KCI03, VV05, XFG19]. Integrated [FM95, CM02, HLY22, SK91, WYY20]. Integrating [LCM24, AMCR17, FMMK01, GZQS23, LCM22, Par21, ZYQS21, ZYQS23]. Integration [TM04, ABH14, ALMM01, AD08, ALM04, AD01a, AAD14, AM16b, BCF^{+}13, BtTBV87, BDE22, BRBM08, BBKS07, BB24b, BF95, But93, BJ98, CMP15, CH95a, CGP15, CL85, CBD16, CR04, DhW09, Den93, Die20, FS15, FW07, Fra04a, Fra04b, FWW^{+}21, GMG04, GPHA16, GPHA25, GVSL96, GML00, GMGF02, GKA17, GQ08, GS21, HS02, Hai97, HOS99, HMdV03, Hig96, HMT03a, HMT03b, HLR01, JVZ95, JR00, JHGZ20, KGR08, KNT13, Leo10a, LD10, LZ20, LL23, LLT20b, Log04, LMG02, DLM16, DE18, MF99, MC17, MQ03, MAF20, MSA20, Naj20, OR20, PT23, RCR24, SSV97, SA12a, ST19, Sim93, Ske99, Sol15, SvdHK94, SCT05, TMD92, TGB08, VV07, WK02, Xiu08, ZD25a, ZLX19, ZSZZ20, tV87]. integrations [KSK^{+}25, LLT20a]. integrator [BCGI13, DL22b, ELLE02, FXY22, FHX22, HK25, Lei02, LZ18, Li22, LC24, Lub04, MD20b, SS94a, TN16, WE99, XCHW22]. integrators [Ale03, BF99, BDF89, Bla01, BM06a, BP12b, BDE22, CO09, CLMSS98, CZ97, CY98, CDR20, EK06, LWL18, LZZ22, LW22, LW20b, LT19, Lua17, MQ00, MPG^{+}16, Mur99a, Mur99b, PWS98, SWL20, SZ97, WWX13, YC13, ZZO16]. integrity [JP08a]. integro [AKM^{+}21, AAL21, AB88, AAH21, ABF09, AKS21, ADH00, AV00, BF92a, BT97a, BMGGG12, BGGG13, Bho12, Bog20, BP92, BMM97a, CGEV19, CNA23, CZ12, CP03b, DLPV17, DCY20, DAMA23, EGV25, EH97, FK23, FID18a, FH22, Gan96, GS99a, GMG19, GAZD25, Gu19, GAOB20, HM87, HZBM05, HS19a, HCY18, JRW06, JCY25, KPY15, Kau95, Kür23, LZQ22, LP24, LPS25, LWY20, MAH18, MHA19, MKN23, MMRV20, Mit22, MFAD23, Mok17, NLS20, PT11, PTV16, PTV20, Pot97, QWX20, QXG21, ROB17, Rah25, RSR23, RN22, SS08a, SOB20, SD13b, SDG20, tSqWyG16, SCvdH92, SSPZ20, Tan93, TX18, WH19b, WLM21, WCM21, WZ22, WC14, WHL19, YS22, ZH09, Zha20a, ZD20, Zha20b, ZYX20, vS97]. integro-differential [AKM^{+}21, AAL21, AB88, AAH21, ABF09, AKS21, ADH00, AV00, BT97a, BMGGG12, BGGG13, Bho12, BP92, BMM97a, CNA23, CZ12, CP03b, DLPV17, DCY20, DAMA23, EGV25, EH97, FK23, FID18a, FH22, Gan96, GMG19, GAZD25, Gu19, GAOB20, HM87, HZBM05, HS19a, HCY18, JRW06, JCY25, KPY15, Kau95, Kür23, LZQ22, LP24, LPS25, LWY20, MAH18, MHA19, MKN23, MMRV20, Mit22, MFAD23, Mok17, NLS20, PT11, PTV16, PTV20, Pot97, QWX20, QXG21, ROB17, Rah25, RSR23, RN22, SS08a, SOB20, SD13b, SDG20, tSqWyG16, SCvdH92, SSPZ20, Tan93, TX18, WH19b, WLM21, WCM21, WZ22, WC14, WHL19, YS22, Zha20a, ZD20, Zha20b, vS97]. integro-elliptic [Bog20]. integrodifferential [Kau93, QZM25, Vab22, VSG17, Wan01]. integrodifferential-algebraic [Kau93]. intensity [CC23b]. inter [Rob10]. inter-element [Rob10]. Interacting [SR88b, MT05]. interaction [AK21, Aso21, BSP04, EZ03, GMM09, GÖS20, GS18, LGS21, RDH^{+}12, Sal93, Sod91, Tow16, XLK07, vBvdZdB08]. interactions [BGG^{+}21, CBD16, DL21a, GT19a, Nes16, WJW19]. interactive [GGNP02]. interconnecting [Pec09]. interdisciplinary [BRW21]. interdomain [Dor91]. interest [KGR08, Mar09]. Interface [Aso21, PGM86, RTV00, ABG^{+}15, BSZ22, CES91, CYWH22, CL24b, Chn17, Dav92, DS21c, Dek17, GMS12, HLZ14, HD22, HCJW24, HLW25, HSY18, JCL18, JV09, KL21, KV07, Li98, LMWZ10, LX21, LS24b, LAH09, MAG13, NMB10, PAP17, PH15, RS20, RTV02, SSV89, Sch16b, SJ11, SD09, SZL18, SW20b, WCSQ18, WZ19, WLZ25, XZT21, XGHM22, YZZ25, YCWH23, ZSY20, ZKO^{+}21, ZCC11]. interfaces [AEMX17, CL24b, FL04, FMW05, FM07, FMU15, HLW25, KS07, NLHT25, TS24a, ZLW20b]. interfacial [EAS12, ZQLK11]. interference [BCV25]. Intergrid [Osw97, Kan04]. Interior [AD20a, EKT19, LO95, AD99, BSGU94, Bec18, BH97, Car09b, CR23a, EB12, HC01, Ius97, Jia02, LHHR94, LSY21, MOSW00, OQ15, QL15, SZQH23, SHB^{+}25, XZZL15, YZ17, YZH19b, ZL18a, Zha07, Hop23]. interior-exterior [EB12]. interior-point [AD99, XZZL15, YZH19b, Zha07]. Interlacing [DJM09, JT09, TJK18]. interlayer [ZBNC25]. Intermediate [Lai09, SHG86, VA21]. internal [CRR03, Mur19, NC16, Rus95, TT20, VS91]. Internality [DRS19]. International [CFTW08, LST07, MH89, FJ97]. interpolant [Ber86, KHLV22]. interpolants [ABH22, AR24, BY22, BES18, BFK11, BS92, CHS17, LMS09, MST07, NLZB23, SST12, SST09, SST15]. interpolated [QPT23]. interpolating [AD19a, AD20b, CPC25, CJV88, DA19, DF92, KP03a, KPR12a, SWCH15, WT93]. Interpolation [AM16a, BG06, Dés08, Moo04, Spe12, AA04, AG05a, AMR12, AN15, ABY22, BMGM12, Bar12, BRIP08, BGS06, BH12a, BS00a, BCJ97, BS09, BB10, BO11, BG11c, Bre02a, BMM97a, CDD00, CDD^{+}17, CGGM17, DMPP99, Dol14, DR93, ECHF^{+}20, FM11, Fra14, FWW^{+}21, GS19, GGR97, GJ17, HB02, Iva07, KV20, Kid90b, KP01, LY10, LPV24, OT21, PD96, Pla08, PB10, QWW24, RL86, Rus95, SE93, SJ18, SL17, VBD93, VZ93, WW05, ZL11a, ZY14, ZL18a, ZY25, ZHL08, Zup04, de 92a, Boy15, FMS24, OR18]. interpolations [CZS04]. interpolatory [BPV25, DBCBPP10, Fid17, GS05, BDMGVO05]. interpretation [Bot97]. Interval [Sch87, Ben02, Boy15, DBCBPP10, DMH18, GPP04, KKP17, LZ13, Mar99a, MT94, NN20, Sch89, TD09, BO11]. intervals [DLS22, BDMGVO05, RG20, Tar98]. interwoven [SSA24]. intracellular [NRWF08]. intrinsic [PWY21]. Introducing [RA17]. Introduction [BKM05, CFS94, Had13, LP93, Sal91, But97, CDG19, Wen10a]. intrusion [AMRR18, MNR14]. invariance [BLP01]. Invariant [HL99, AB10b, BHJJ06, BC97a, CGW20, DB95, ERS00, GSR00, LMW23, Moo95a, Tru00, Wan25, WD22]. invariants [HCC24, HXW15, Rei99, SMB23]. invariants-preserving [HCC24]. invasion [MPSS16]. Inverse [FOMC05, IT16, LLT20a, Per88, Tah96, YD07, AS21, BT99, BG24, BLY16, BMT25, BG02b, CR05, CYYH21, CR23b, CD18, CPOGO17, FBS09, Hab08, HS21a, HILK13, Huc99, JLZ20, JM94, KJL12, KNP16, KKN^{+}17, KCW16, KRBK16, KK02, LM21, LN24, LYY15, LDH^{+}24, MG00, MN03, Mat05, Men23, MCM12, QC12, QL15, ST09a, ST09b, SPS20, Saz24, SLJ11, pSLqJcY16, SHLY19, SJ18, Shi20, SHB^{+}25, SOA^{+}25, TL07, TK05, WZ02, WL09c, WL16, WW14, XC85, YJZ18, YD22, YLW20a, YY24, Zha00, ZW19b]. inverse-free [CYYH21]. inverses [Kos02, MNSS22]. inversion [Che88, CZ90, DSK12, Gem23, LFP04, Moo95c, Per99, VA05, vI87]. inverted [KKN^{+}17]. invertibility [WKN20]. Invertible [CRS05, RU21]. Investigation [AD20b, AFIS24, Chr01, EZ03, Liu97, MDASAO21, Pot97]. investigations [AD19a, CSS19, KOS21]. inviscid [Mou03, Tow16]. Invited [Ano05f, Ano02e]. involution [Sei02]. involving [CFRA08, CF13c, MV18, NYPW21, TLQ21, ZJLA22]. ion [PS00, RGK21]. ion-drag [RGK21]. IPDG [BS14b, HSF25]. IPDG-FEM [BS14b]. iPSC [BRSD91, de 93a]. iPSC/2 [BRSD91]. IRK [KSK^{+}25, LQS21]. IRK-WSGD [LQS21]. iron [BMP05]. irrational [MG22]. irreducibility [AHA23]. irregular [AY22, APJ09, AEN22, BW95, DT10, DP90, Eva94, EW97, DLM16, PM14, Vic87b]. irregularly [Sau00]. irrotational [Fan19]. ISLIM [Lo06]. Isogeometric [CGMS21, HM22, GV18, ADSS17, ZP24]. isolines [BS00a]. isometric [Chu03, Gar05]. Isoparametric [JT06a, PT23, LL02]. isospectral [CIZ96, LP97]. Isothermal [DdCVR03]. isotherms [DMM24b]. Issue [BGHR12, BGH^{+}15, CHM09, FFMZ13, ADG^{+}16, CD95, DLN^{+}24, DGCW17, HKNV16, KP07, LST07, SSV97]. Issues [TÖR22, BF09, DS23, LMV17, Sid10]. Itô [DR09a, DR09b, MEGW23, SHL19]. Itô-Volterra [SHL19]. Italy [RI02]. iterate [ST09a]. Iterated [LNZ12, YXN21, Bic16, BCC16, Bic21, BDFV95, Buc17, CM07, CX08, DVV93, Fil25, MZK05]. iterates [Bog12]. Iteration [BMSZ21, Bad20, BY09, Bai25a, BLY16, CKB13, CDW19, CLLM21, DS21d, ELvdHS98, Gil88, GRLL01, GJlL23, Gug05, Guo26, HV22, HL02b, Jia00, KM21, LYC24, LXCM21, Mat05, RM25, ST09a, ST09b, SW95b, SD22b, SB25, Spi95, TWH21, WPL16, XY25, XY24, YL26, YW25, Zha01, vS93, vdSvdH95, Ney95]. iterations [BSvdV99, BBLT15, DB95, DSS15, EH09, JKN94, Lab99, LLV18, Phi87, TT20]. Iterative [Bea04, BNKR20, BG24, CFTW08, GPMR95, HGM^{+}21, HHL23, JP19, JGK11, MP05, Mar03, SD24a, SS13a, SL01b, TSB10, WS21, dH95, AC98, AGJ12, AHR12, AG05b, BRTB19, Bog16, Bog20, BRBB18, Cao98a, Cao98b, Cao01, CH01, CGPT19, CRHTV25, Cui04, DS20, DE16, DL22a, DP90, DZW24, Eij95, Ein18, ESE20, FL05, FLMR14, FL24a, GH91, GM18, GK09, GT02, HHAA22, HS22, HvdHV10, HW97, HZZ24, IT07, JY20, JD09, Kim94, Kim95, KF97, Kin94, KYC03, KSSS16, Kru99, KCB02, KR20, KYI17, LPT16, LZ13, LWD^{+}09, LGS21, Li23, LTDY24, LAZ20, LR00, LR20b, MA09, MRV93, Mar99a, Mit24a, MDASAO21, Mon21, Nak05, NH24, NTT22, NH15, OMP98, Pea16, PM03, PT95, RV05b, RSK14, SW95a, SMTHE22b, SMTHE22a, SHL19, ST11, Saz24, SH02, SH10, SVB17]. iterative [SW03, SFZ21, TQA^{+}25, TK05, TWD23, TH23, VMP03, Wan96, Wan07b, WL09b, WWS07, WYP12, WSC21, XXQ17, YG95, YHT23, YR92, YXB95, YH07, ZC91, vSK97]. iteratively [GZW22, HSY18]. IVPs [BP12b, CM07, CX08, GM95, Jac96, JN02, RCR24, SKS^{+}25a, SS09].

J [BtTBV87]. Jacobi [BGS06, FLS94, For11, RF16, AGZD22, AM16a, BP02, BL06, CGT13, CXNF14, CFS13, CY05, DMR10, DB08, DBBH14, Gen10, yGqWsWC05, GH07, GSW09, HGP11, HHYD20, Has20, HY24b, JZZH22, KHYY21, LHS00, MB20, MCS16, Mit22, Mok17, NLS20, RN22, SM13, Sti03, The17, YS24a, YZH24, hYqW12, YWSL20, YXZ18, ZDM18, ZHS22, ZLW22]. Jacobi-spherical [GH07]. Jacobi-type [Has20]. Jacobi-weighted [AM16a]. Jacobian [CZ97, CS24, GHW01, KCW16, NNJ23, SY05, WGB99, ZWFX22]. Jacobian-free [CS24, NNJ23]. Jacobians [KCJP01]. January [BGHR12, Ano22u, Ano23s, Ano24m, Ano25y, Ano26b]. Jeffery [AI19]. jet [AHS03, KHA12]. joint [DNW18, JY20]. jointly [PTW19]. joints [MMP02b]. Jordan [GY94]. Joseph [BSZ22]. Joule [YJ23]. JS [Tan23]. July [Ano21t, Ano22o, Ano23n, Ano24v, Ano25n]. Jump [CGEV19, AG05a, BAD13, Bi20, CF08, DS23, GT19b, HFL13, JD09, KKT16, KP18, Kał22, KS07, MHL18, RS20, ST11, ST14a, SB19, SW20b, ZD25b].

Jump-diffusion [CGEV19, BAD13, Bi20, GT19b, KKT16, KP18, Kał22, ST11, ST14a, ZD25b]. jumping [MII13]. jumps [AO05, DMS23, KK11, LL15, MDD14, Moo04, RP17, RT20, WG10]. June [Ano22p, Ano23o, Ano25v].

Kaas [AB17, MEGW23]. Kac [SND21]. Kacanov [GMZ11, HJS97]. Kaczmarz [BW23a, BW23b, GC25, LLY24, LG19, Liu24, LHT20, TWD23, Yan21b, YL26, ZGF^{+}25, ZZ25]. Kaczmarz-type [BW23a, ZZ25]. Kahan [SS94a]. Kalandiya [JK20]. Kalman [BGM19, KK17, KK20b, KK22b, TZ00]. Kantorovich [SL09]. Kantorovich-type [SL09]. Kármán [Hop23]. Kaup [NAF24]. Kawahara [Geb24]. KCE [VK17]. KdV [Car19, CCST22, EEE22, FvdMS20, GD22, HAN23, HCC24, JPP19, KSHB21, LYMD24, LXLY25, Nak24, PWY21, WDU21, XWM26]. KdV-type [Car19, EEE22]. Keller [DvHM19]. Kellogg [VN21]. Kelvin [DYF23]. Kepler [Boy07]. Kernel [FRV11, ZRA23, AD21, ABdSG23, AEN22, BLM17a, Boh21, BZ94b, CDD00, CDD^{+}17, Che12a, CMPP24, DCY20, FK23, FZA25, Has13, HBB25, KXK92, KAS22, KSHB21, KZ21, LLKJ21, LW21b, LWW22, Lia22, LD97, MAHZ21, Mir20, MP15, MHL18, MDASAO21, NSD23, QXG21, QZM25, SAA20, Sch08a, SKR^{+}16, SPYS24, TMD92, Tan93, TY03, VSG17, XZH19, XZT21, ZD20]. kernel-based [AEN22, CDD^{+}17, MHL18, Sch08a]. kernels [Cai24, CHLX07, CNA23, CL14, GMG19, GNZ21, KR18, LNZ12, Maj14, MKDC25, MO17, MAD23, OB24, PT11, Vab22, WTY21, WCM21, YT21]. Kerr [DTQ^{+}20]. kind [All24, AHB20, AD18a, AAEMY21, Aze22, BKM19, BG24, BHR05, Bru97, Cai24, CHH15, CS09, DS21b, DN21, Fer09, Gwi09, HDS20, HLL09, Kan89, KX91, KM19, KCY19, LW21b, Lia22, LNZ12, Man97, Mie89, MKJ23, MPDB24, NDM20, NSD23, Pan21, PP24, RG20, Set24, SV00, SDG20, TH23, VGL24, WH13, WW24, XL09a, Xu16, ZL18b, ZLX19]. Kinetic [LMPS19, Ale11, BEH24, FJP17, KW95, RA03, RGMO19, WTB24, VK17]. kinetics [BLRGVR23, DS02, PM05]. Kirchhoff [Wei18, YLY19]. Klein [CC19, CH89, DL22b, FCW20, HZ09, HMN20, KCY19, LSWM19, LYA^{+}19, Lyn99, MD20a, MMDH19, MMD20, MMDS21, Odi24, RV22, WW19, WDH20, WCJ23, WG23b, XWZ21, YZQ^{+}22]. Klibanov [NK24a]. knots [GZZ19]. known [ADNR21]. Koiter [SYL^{+}20]. Kolgan [VCC12]. Kolgan-type [VCC12]. Kolmogorov [BFdO07, AMH24, LZ18, LXZS22, MSZ^{+}24, MS25, QNA23, WWZS25]. Kontorova [YEZ25]. Koren [Hol01]. Korteweg [HY24a, AM04, Bas21, BDKM92, BR97, FWL18, Isk89, KS22, WH19a, ZJ25]. Korteweg-de [HY24a, Bas21, ZJ25]. KPP [CMMR23]. Krein [KKN^{+}13]. Kreiss [AK95]. Kronecker [CMT25]. Krylov [HH10b, ABR23, Bai02, BBLT15, CM97, EAV16, Fac03, GTS20, EEJB22, EJR25, LS99a, LWZ22, MS99a, MN08, PR12, PWS98, RSY12, RU21, SW98, Sim10, SSX14, SS12, Wal95, WSP97, Wei95, Wen05, ZGL98, ZFC20]. Krylov-ROW [Wen05]. Krylov-subspace [BBLT15]. Kung [TQA^{+}25]. Kuntzmann [VS94]. Kuramoto [AS04, KK23, LO23, Mar93, MDA24]. Kurchatov [CGTTN24]. Kutt [Ioa89]. Kutta [AMCM08, BDP99, BC08a, BCR01, BCET22, But97, CH95b, CG13, CSLY19, CPR93, DM11b, GMG04, GML00, Hoa15, Lab98, LW17, LW19b, LS05, MEGW23, Mur98, Pat98, Pat00, Som93, VDVV98, VV02, vSC92, AH11, AHJM19, AJ19, AHJ^{+}23, AM95a, AB17, Alb96, AMP03, AH17, Ant23, ARS97, AFK92, BMQW16, BDGP96, BT97a, BJ05, BZ93, BV96, Bel97, Ben96, BG02a, BS92, Bos09, Bot97, BJ11, Bru93, BJ03, BB96, BB98, But96, BS96b, BW96b, BC97b, BT97d, BC00b, BSW93, BS02, CMR94, CH95b, CIZ96, CdFN01, CG05, CCG13, CV95, Cas96, Cha96, CCM02, CCS17a, CS03, CGS20, Coo89, CST97, CN11, DIJ12, De 88, DVV93, DS05, DR09a, DR09b, Dia95, Duj09, DLV24, Eir95, EH97, Ere19, FMS24, FS05, FS08, FJL21, Fra04a, GHW20, GZQS23]. Kutta [GCHR06, Gar10, GJ00, GPMR95, GPMPR03, GPHAM12, HZ96, Hig96, Hig97, HR06, HMR25, HS09a, HO05, HXW15, HHW18, Hor98, Hor05, HEJ96, IRC12, IJ17b, Jac93, JVZ96, JVZ97, JKN94, Jay95, JM25, JS25, Kau93, Kau95, KCL00, KC03, KC19a, KC19b, KMG09, KS08, KW98, Kom07, KM18, KS89, KS09c, KKR15, KNO96, LX08, LLZ^{+}22, LP97, LO96, MK99, NS21a, NS20, Ore93, Ost93, OT02, Pan07, PCR17, PJB04, PAJ12, PCRR17, Ran16, SA21, ST89, Sch02, SC25a, Som86, ST25, Spi96, Tsi01, Van00, VZ93, Ver96a, Ver06, Ver96b, WGKS12, WYL11, WHL19, WSW96, WBCK02, WX06, XG22, YZWQ25, YBW20, YC13, Zen93, ZZL01, ZJ19b, ZYQS21, ZFX17, ZZX20, ZSQ20, ZSQ21, ZC99, in 92, in 96, vS93, vC93, vS96, van96, vdSvdH95]. Kutta [vB95]. Kutta-like [PCR17, WX06].

l [YP18a, CGG02, CD18]. L-BFGS [CD18]. L-curve [CGG02]. L1 [CCST22, HZ20, HLY22, TTP24, WCM21, WYY20]. L1-ADI [WCM21]. lacunary [Mat91]. lag [BCFQ19, BCFQ21, KV95, SS09]. lagged [MG18]. Lagoon [RI02, BDES12]. Lagrange [OR18, AAL21, ABG^{+}15, BGG04, BNV06, BCGI13, CN17, HLZ06, KwS19, KGZ24, LH02, PAP17, PP92, RM25, SW20b, ZYH23]. Lagrange-collocation [CN17]. Lagrange-spectral [AAL21]. Lagrangian [AL05, BB15, BC08a, BBC^{+}25, Bou02, CFX08, CF13a, CBHM19, FVGS13, KCC04, KKR25, MZZ17, PP00, RVD00, Sea09, WL10]. Laguerre [AT15, BMPR15, Fun90, yGyZ07, JCN94, Kza99, Ter22, jWyG08, YWW23]. lake [TK19]. Lambda [Leo10a]. Lamé [LRV25]. Laminar [hYK86, AHS03, Kar89]. laminate [GP01]. laminated [FSB97]. Lanczos [BRRS15, BS93, EJS04, FW22, FN95, JP93, Khe91, KBG04, RG02, YR09]. Lanczos-type [BS93, RG02]. Landau [BS06, CC24, CWX21, HAR21, DBH25, LHW17, LWW23, SL20, YJ21]. Landmark [LWCT07]. Lane [ER18, AY21, Cha25, DSAB20, Güm20, IB24, RTA19, SSC23]. Langmuir [DMM24b]. language [EE20, PSB91]. Laplace [BSGU94, BF15, Bie12, BDN^{+}97, DZZZ24, EVO06, HvdHV10, KFR25, KvyS15, Lau17a, LLHC17, LC20, LFP04, Moo95c, VA05, ZLCH20, vI87]. Laplacian [AK00, ABD24, ADM22, CCdlH20, CF18, FVGS13, JL24, LZJ21, MZM20, MAD23, PK21, SN04, XW19, ZJLA22]. Laplacians [Lyn92]. Large [BDD^{+}20, Dat99a, Lyn99, ND85, PS03, WVBM88, van95, ARY23, AB10b, AB07, Bai02, BHL^{+}21, BW23b, BC99, BLW07, BNH01, BHSW16, BHSW20, BT02, BJ11, BCDR25, Buc99, CWC24, CLL23, CR23b, DW15, Ewi91, FTB97, FBM17, FS24, GS17, GC25, Gen10, GT02, Hab08, EJR25, HLT07, Hey10, IKR^{+}22, Jac96, Jia00, Jia02, JY20, KM21, KS07, KW95, KM11, KK20c, LCVG01, LP05, LS13, LW92b, LCJQ12, LS24a, Mat86, Mur19, NNJ23, NMSF94, Pea16, Per99, RGS25, RS08b, SDK24, SEGV02, SM89, TYKK01b, Toc01, Wan23, WQS^{+}24, WSP97, Yam23, Yan21b, YR92, YLH20, ZNK02, ZG20, ZD21]. large-amplitude [Mur19]. Large-Scale [WVBM88, BDD^{+}20, Dat99a, AB10b, Bai02, BHL^{+}21, BHSW16, BHSW20, BT02, BJ11, CLL23, CR23b, DW15, Ewi91, FTB97, FBM17, GT02, IKR^{+}22, LCJQ12, SM89, Toc01, Wan23, WQS^{+}24, YLH20, ZD21]. larger [MBS23]. Laslett [RMCG04]. Lasso [AC23, LW25]. latency [GR93]. latent [TS08]. Latin [IKS25]. lattice [BT19, EK06, Lai09, MDA24, MKS12, Nak24, VVR08, ZAB15, LLL08]. lattices [AHGM21, ABD16]. Laurent [BDMGVO05, Njå88]. law [CCL22, KPRU20, NYPW21, RMCG04, RB12, SPYS24]. laws [AM09, AKG14, Aff94, Bac14, BMGM12, BJ02, BTP96, BFA93, BDF94, BM12b, Bor16, BGP11, BGS02, CSW19, CH15, Dav92, DF96, DGRS09, EK96, FL01b, GZZ19, GT00, GBBC^{+}23, GJ17, HSS07, HWCF15, KHM^{+}19, KS09b, KXR^{+}04, KR12, KLSW06, LPR00a, LPR00b, Luo18, Mon21, PGDB08, Pel20, RGB20, Tan01, TDW23, TJ12, XF06, Ye04]. Lawson [AMCR17]. Lax [De 88, LGSW25, LVW21]. Layer [ÜSHT03, Bec18, CCCC^{+}25, DRVA20, DS97b, DK14, FM11, GGN12, GD22, GD23b, Jun06, KDT17, KL21, KC94, KTYY24, Lin01, Mat09, MOSW00, OQ15, PNA21, RN25, Sod91, ST08, TS24b, ZFW20]. layer-adapted [FM11, Lin01, Mat09, TS24b]. layered [IT07]. layers [AG98, BBV13, Bog00, CR23a, EH91, FLR08, HO24b, KOW05, LBCN00, MOS02, MS91, OS08, Pet00, TY98, Ye04, ZTZ15, ZMC13]. LBB [Aso21, BM12a, Dob05]. LBB-stable [Aso21]. LDG [Cas06, AC15, CR23a, MFAD23, SRK22, WC24b]. LDG-FEM [CR23a]. leading [AB15, GS17, OK98]. Lean [SA08]. leap [TH18]. leap-frog [TH18]. Leapfrog [NH15]. learned [CDI^{+}24]. learning [HGM^{+}21, IKS25, KSP10, WTB24, WOLY25, WLZ25, Xu23]. Least [BM04b, CCL22, CF86, CCLT10, KR10, MH16b, PS02, ZSJ04, AN15, AD01b, AD19b, BLS^{+}17, BB94, Ben02, BIW24, CP07, CZ04, CC20b, DCY20, yDqGnJT09, DZW24, HP85, Han87, HS17, Hua19, KN19, KS02, KLY05, KS09a, LG21, LV12, LS25, LRRÁY25, Li11, Li16, LSY17, LW18a, LJYS20, LCZ21, LLW22, MOS12, MS08a, MSS21, MD23a, MD20c, MD21, MD23b, Mon09, Mou03, NZY21, PS03, PR22, Ren13, SD13a, SLJ86, ST25, SWCH15, WM22, WQS^{+}24, XZT21, YK04a, Yua93, ZG20, Zha07, Zup03]. least-square [Li16]. Least-squares [CCL22, CCLT10, KR10, MH16b, CZ04, yDqGnJT09, Han87, KS02, KS09a, MOS12, MS08a, Mon09, Mou03, PR22, SWCH15, XZT21, Yua93, ZG20]. Lectures [Ano05f, Ano02e]. Leffler [ABdSG23, Boh21, KZ21, OB24]. left [SY05]. leg [Hua00, WZL08, Wan17a, ZH09]. Legendre [AKM^{+}21, AY21, AN25, AyLqW18, AA22, An16, DSAB20, DKL24, FMS24, GeO24, yGpY09, HAA21, HCGW22, IB24, KME20, KSK^{+}25, LM21, LIPT18, MQO17, NLS20, OL18, Pan21, PP24, Pis22, Pło22, QM20, RE19, SLW17, SW07, SBS^{+}20, gTpM07, jWyG08, jWqW09, WW24, WC14, XZZ19, YLW21, YRV21a, ZH21, ZMY21, ZZJ21]. Legendre-collocation [AN25]. Legendre-tau-Galerkin [QM20]. lemma [DLM05]. lemniscates [FH04]. length [SH91]. Lepp [Riv09, RRMJ12]. Lepp-bisection [Riv09, RRMJ12]. Leray [LRS23]. less [KD25]. Letnikov [CHS19]. Level [KM11, BvG19, BSV21, BIO24, CCS17b, CRHTV25, DdSF07, DLP06, FM07, FMU15, GKL07, HL02b, HFL12, Hua17, Joh01, KM19, KL09, KK22a, Kuz90, LA11, LCS19, LXZS22, LYC24, Nak05, PSB91, PXHZ20, PH15, PRS20, QM19, TL07, Tob14, ZOZ09, ZS21b, ZWL11]. level-block [BSV21]. level-dependent [CCS17b]. level-set [PH15]. level-set/moving-mesh [PH15]. levels [LYA^{+}19]. Levenberg [ZGF^{+}25]. Levin [LWWX10, PCA10]. Levy [LCZ23, ZO14]. Liao [YFLX20]. library [AV96, Eij95, SW95a, Tou97]. LIDAR [BH20, HB20]. Lidstone [CD05]. Lie [WGKS12, WC02, YF24]. Lie-group [YF24]. Lifshitz [BS06, CWX21, YJ21]. lifting [MB08]. light [Boh03, Den07, FGP23, GV04]. light-triggered [FGP23]. like [CJL13, DS21d, DSS20, GO18, IKM23, Li01b, LCM22, LCM24, Man97, Mur15, Par14, PCR17, SOB20, SLJ11, SW13, SH21a, TY98, WT17, WX06, Zha21a, ZZW97, JKN94, LN21, van95]. Limit [Gil88, AGK24, BB24b, Jac88, Lem88, MW24b]. limitations [CCM17]. limited [BKAG22, SM13]. limited-memory [BKAG22]. limiter [DSZ15b, DSZ15a]. limiters [MS08a, YR22, ZQLK11, ZSQ20, ZSQ21]. Limiting [BM04c, BR20, ID19, KXR^{+}04, Sim94a, ZBD24]. limits [BB24b]. Lindsay [CA15]. Line [BGIW18, HP15, Rei85, AHO16, CH22, DS24, DL22a, HMY19, HY24a, KS22, Kür23, LLJY20, MOV24, MD23a, OT24, OT25, SXP09, WZ16, jWjJ17, jWC22, YLW20b, ZH15, ZP12]. Linear [CSS19, HV89, KK02, MS91, Pło22, TZA13, XWZ21, AC98, AH11, AJ19, AS13, AAL21, AW14, AK00, Alb96, AMC02, AMP03, AMCM08, AMCR17, AT13, AL24, Ara99, AR15, Arn98, AES13, AGKK94, AC16, AAEMY21, Bac21b, BHL^{+}21, BW23b, BHB23, BF01, BBG14, BF17, BK06, BBD20, BB24a, BG24, BGM19, BM04c, BM06a, BP12b, DLS22, BRBB18, BCJW17, BCJP18, BJ20, BIJ23, BS93, BM09, Bru92, BCSH16, Buc04, BS12, BH97, BJ06, BS18, BRW17, BWEP95, But85, BC95, BJ96, BS00b, BW03, BP06b, BD17, CGS19, CIJ17, CRS05, Cao98b, Car09b, CL85, CDI91, CHX13, CZY18, CLLM21, CYWH22, CLL25, CKK25, CG16, Cve02, DMS23, DB97, DS05, DHWL22, DN13, DJM09, Eij95, EJS04, EG88, Fac03, FM21, FHK05, FSX25, DSV13, FH08, FMGN94]. linear [FG09, FHV97, GC25, GMZ11, GHK16, Gas92, GNNR19, GAH25, GMG19, GD09, GGM07, GS09, GHT05, GFPG18, GCZZ23, Güm20, Guo96, HM21, Han87, He24, HGF24, HK25, HVY91, HS19a, HH10a, HP91, HA16, HCY18, HL24, HZZ25, HZZ24, Hua98, HCX03, HH18, HS21b, HY24b, HMW05, HT94, IO18, IJ17a, JQSC22, KPY15, KM21, KZ13, KP03b, KCB02, KLSW06, Lam24, Lem02, Lev91a, Li05, LSK12, LYF17, LLV18, LCS19, LWW22, LB23, LZ24, LAZ20, LC99, LKJ20, lLNW21, LWZ22, LYZ25, LS99b, LW92b, LLL12, LYK17, LW20b, LJYS20, LL20b, LC21, LCZ21, LMW23, LSW23, Liu24, LMG02, LD02, LRT99, MAH18, MHA19, MOS12, ML16, MW24a, MSGM23, Mar08, MF99, MP98, MMDS21, Mär95, MS90, MST09, MH04, MVG14, Mit24b, MG22, MM25, Mon21, MAH22, MRH14, Mot17, MT20, Nap16, NRZR12]. linear [ONL89, Odi19, Ort20, OKS10, OTK04, PGA93, PK23, PGS10, PTV20, Pic05, Pir09, Pło23, PT95, PJB04, Por17, QH22, QL16, RA05, RC18, RU21, RM25, Rum87, SS25, SRK21, San20, SS99, Sch12, SH02, SEGV02, SI20, SZY21, SWY^{+}23, SWW16, SL21, SD22b, Sid90, ST20, SS13a, SvdHN86, SW03, SSX14, Spi95, Spi13, SS09, Su94, TDPU17, VMP03, WL09c, WW19, WT20, WWZJ22, WWZ25, WJF25, WKN20, WWS07, XFLC00, YG95, YLY19, Yan21b, YR92, YZH19b, Yüz22, Zen21, ZXYW22, ZNK02, ZG20, Zha21a, ZR21, ZD21, ZHL22, ZFS24, Zha07, ZFX25, ZP97, ZC99, Zla85a, dH95, in 02]. linear-decoupled [WWZ25]. linearised [HS22, OQ15]. linearization [AÁ21, Aro96, BCS17, HCW16, LZH^{+}25, NNJ23]. linearizations [Mär95]. Linearized [CXZ14, UHUL21, BG11b, BIO24, BP85, BCV21, CHP19, CCST22, DG10, DJ10, FWL18, GWLN22, HN03, KM16, LSWM19, LLW20, PXHZ20, SL20, WH23, WLY24, XWM26, YJ21, Zha19b, ZQZ23]. Linearly [Bru93, CdFN01, CG05, ELCWS98, JKW12, PWS06, SGR21, WWM22, BJ06, BSW93, DK20, DW21, EM05b, FCW21, GHW20, GLPW09, lLX22, PHY19, SMB23, SMEN04, SGY22, WSW96, XWW19, YZC21, ZW24]. Linearly-implicit [Bru93, PWS06, BSW93, SMEN04, WSW96, XWW19]. Lines [FG96, BL15, BDNV19, BDF89, BW96a, BCJ97, HT94, PSL18, Tah96, TT03]. link [MS08a]. linked [Car19]. LINSOL [HSW99, HS02, SH02]. Lions [LRS23]. Liouville [AGM09, BBC^{+}25, Con99, DCJ20, Ghe97, Gu20, Hey20b, LZCF21, LR20b, Pru00, SL15, SND19, VVD95]. Lipschitz [BSTT22, LS10, LDC24, LSWW22, Mat86, RA09, TLQ21, WG22, ZM17, Zha20a, Zha20b]. Lipschitzian [WZ14]. liquid [BSP04, CPY20, FMW18, TK19, WaZW23]. liquid-structure [BSP04]. liquidity [GK22]. Lissajous [BDV17]. List [Ano02e]. Load [dDF^{+}94, Chr96, DBH^{+}05, FLÖ^{+}97, RN04]. load-balancing [Chr96]. loading [CS09, Che12b, Fai00]. Lobatto [OL18, BGVHN10, FMS24, KK09b, SLW17, Wel10a]. Local [ASA20, AGQ^{+}24, BE99, CCST22, DM09a, Eis86, HAC22, KL09, LS07b, LWC24, Mat09, MS13, NNJ23, NWL^{+}22, PC00, RS09, RM25, Sim94a, uIVS13, Tro93, WWLL23, XHJM21, AJ24a, AS20a, ACMR06, AMR14, Ang06, ABM17, AF89, AD18b, AD19b, Bac17c, Bac18, Bac19, Bac21b, BDSG09, BBRBS09, BS20a, BMV06, Bus06, DDHS97, DF11, DJL04, FSB97, Fre91, GG22, GP98, HH98, Han06, HD04, HL08, HJL18, HS23, HAY20, JD09, KDT17, LCHR03, LL15, LH09, LW19a, LW20a, LGSW25, LYMD24, LMP99, MB08, MVVA09a, MFAD23, MMP20, Mus11, NS21a, OHA25, PK21, RP17, RGL16, Roo20, SRK21, SG00, Sti03, Tan24, Tob14, VT93, XWM26, YH18, Yua20, ZZPJ23, Zup04, Cas06]. Local/global [RM25]. Localization [Rab25, Sza94]. Localized [MHL18, Boy15, CNA23, Cha17, NT20, Wal19]. Locally [GKMS09, BSGU94, BSTT22, DN13, EL94, FJ09, HMP14, HN22, Ise02, Kim07, RZ00, Zha14, ZM17]. locally-refined [Zha14]. Location [EHV24, ABI22, CHR03, LM22b, Rya00]. locations [AGY08]. locking [DN13, LB23]. locking-free [DN13, LB23]. locomotion [Ran20]. Loewner [ACKV24]. log [CXZ15]. log-stable [CXZ15]. logarithm [CKL03, CL06, Lu98b]. Logarithmic [ZHS22, CG21, GJLL20, HBB25, KM17, LZH19, LRT99, MD19a, SPYS24, TMD92, TM24, XLKY19, YZQ^{+}22, YLY19, ZCQ25]. logarithmic-power-law [SPYS24]. logarithms [DMPP99]. logic [PGC01]. logically [HS97]. logistic [KF97]. Lommel [Zul25]. Long [CD20a, Mur19, RTT01, TK19, WWL21, WXY24, ZFW20, AD01a, CH95a, CFRA08, CBD16, DMS23, DL22b, EH88, FXY22, IK24, JZS20, JQSC22, LH24, LO96, MPV24, Shi25, XXF22, YZG23, ZLG24]. Long-term [CD20a, CH95a]. Long-time [RTT01, WXY24, DL22b, FXY22, LO96, MPV24, Shi25]. long-wave [IK24, JQSC22, LH24, YZG23, ZLG24]. longest [PPS05]. longest-edge [PPS05]. look [Bot97]. Loose [SMTHE22a]. Lorentz [AHGM21]. loss [BDD^{+}20, GT19a]. lossy [IJ14]. Lotka [BLRGVR23, GLS09, Mic03, MdD04]. Low [BBG14, HM86, HMR25, KCL00, KS08, VGL24, BLD17, Baz03, BHSW16, BHSW20, Bla01, BCET24, Cam99, CMP06, CKM25, ÇY22, Con20, DK21, DESZ25, ECHF^{+}20, FL09, GZW22, Gem23, HPS12, HLMP09, HR14, Hey10, LZZ18, LLT20a, LCJQ12, DE18, Ran20, Ris05, SY18, SZL18, VRK25, Ver96a, XY24, YZWQ25, ZCGS21]. low-degree [LZZ18]. low-discrepancy [DE18]. low-dissipation [CKM25, VRK25]. low-frequency [LCJQ12]. low-order [HLMP09]. low-rank [BHSW16, BHSW20, ÇY22, Con20, DESZ25, GZW22, HPS12, HR14, Hey10, LLT20a]. low-rank-deficient [Baz03]. Low-storage [HMR25, KCL00, KS08]. Lower [PL25, YB10, BB24b, Che16, KS10, dC18a]. Lowest [DZZZ24, ATW20a, GHH20, Kim21, YJ23]. Lowest-degree [DZZZ24]. lowest-order [ATW20a, Kim21, YJ23]. LPS [AR18, RÁM23]. LPV [SM08]. LQP [BMSZ21]. LQP-based [BMSZ21]. LQR [XG22]. LSNN [CCL22]. LTI [KPRU20]. LTL [CW22]. LU [BSV21, EVO06, Gu01, HS02, SH02, de 95a]. lubrication [AGJ12]. Lucas [HAR21]. Luminy [Wen10a]. lumped [AMK18, GS94]. lumping [DS02]. Lyapunov [DV95b, FZM20, JP19, LZZ22, LS13, SD22a, SS12, Wan17a, WPL16].

M [BtTBV87, dv95a]. MAC [DYZ20, LRC19, XLZ23]. MacCamy [YZC21]. Machine [Pet92, VV95, Duf90, HGM^{+}21, IKS25, IMM04, MI03, WOLY25, WLZ25, Xu23]. machines [Alt85, MRV93, SSW04]. Maclaurin [IMMS20]. macro [XYZ24]. magnetic [BC02, HLR18, LM00, LSV22, LW22, PVH25, SC19, WSC09]. magnetization [BC93]. magneto [Qiu23, WaZ24, YZH19a]. magneto-hydrodynamic [WaZ24, YZH19a]. magneto-hydrodynamics [Qiu23]. magnetohydrodynamic [DM09a, DLM20, SD11, SC25b, WPT19, XY25, ZFS24]. magnetohydrodynamics [AL17, BM18, CFXZ06, CCZ22, DA19, GDS^{+}15, LMA18, LL19, LA21, Pow94, Tou10, YHT23]. magnetoquasistatic [BBS11]. magnetostatic [BP85, Kur98]. magnetostatics [RV09]. magnetostriction [BS06]. Magnus [BM06a, INR01]. Mahony [AJK20, AEN22, WLY24, ZRA23, SJ20, ZJ25]. maintenance [KN08]. Malliavin [Yam23]. Manakov [GT19a]. manifold [PTW19]. manifolds [BAA22, EMMK01, FMMK01, DE18, Moo95a, MEGW23, MK99, TQY24]. manipulator [BdFPSdSC08]. manufacturing [KN08]. many [AA20, KW21]. MAOR [HPY92, MT94]. map [KF97, WE99]. Mapped [BL06, YT21, ARGA00, DE18, YS24a]. mapping [Bri85, Car94, GWZ25, LWCT07, Mul19, Mur19, SL08, VA21, XZL07, ZJ10]. mappings [AHAS21, BBO03, LFP04]. maps [AM10a, ABD16, DE16]. March [Ano22v, Ano23w, Ano24t, Ano25t]. marching [DD97, Dav98, LY24a]. margin [KSP10]. market [GK22]. markets [CGEV19]. Markov [BT02, Buc99, DNW18, DSM11, Mar03]. Markovian [DFLM19]. Markowitz [HW22]. Marquardt [ZGF^{+}25]. Marsden [FL24a]. martensitic [WM07]. Maruyama [GLMY17, GLM18, LDC24, LLD18, LMTW20, LMW23, LSW23, WGW15, WMC09]. MAS [KFR25]. Mass [DLP06, AMK18, BRBM08, BBKS07, CGGGS11, CFLW22, CGJ16, DW21, FLL11, KL21, LCW20, Maj20, SNW22, TW00, WKWC26, XCHW22, YZWQ25, Yan23, ZB07, ZFX25]. mass- [WKWC26, Yan23, ZFX25]. mass-conservative [YZWQ25]. mass-conserving [BBKS07]. mass-preserving [DW21, KL21]. Massive [Gea93]. massively [BMR^{+}17a, de 95a]. Master [Ano92, Ano02f, Ano04l, FL09]. matched [DK14]. matching [BMS89, BTDV10, CLR11, CH19, LN24, Ste05b, TCCW89, ZC92]. material [AEMX17, MAG13, NTT22]. materials [CFRA08, DCN^{+}19, Fra04b, HS07, HJX^{+}19, SC03, Wee01, YWH20]. Math [AS21, Bic21, BtTBV87, Kni95, TLP18a, Tsy96, YP18a]. Mathematical [BGG12, CL02a, DMQ02, For25, FMSV07, FDF^{+}25, IJ14, MLK06, MDD18, OCVW22, AKGR14, DSM22, CDJT06, ESS15, EE20, EL01, FF20, LW20a, LFL14, PVH25, Riv09, Sae14, Sam94, SAMSB20a, Var92]. Mathematics [MH89, ZCSH11a]. MATLAB [EMMK01, RH92, ST01]. matrices [ABK12, ADNR21, BN99, BWY03, BLD17, Bai25b, BC99, BG11c, BW06, CL06, CDW19, DP12, DL01, DMPP99, EES05, EH08, GK93, GS08, HA21, Hua19, Jia02, KS10, LM25, LMV17, LVfP14, LLT20a, LLT20b, Liu02, LB21, MS19, Nak12, Not92, Nov03, PV93, SKO19, SH21a, SYW18, UHUL21, Wan96, YD07, Zha00, ZWFX22, dC18a, dlHV13]. matrices-based [UHUL21]. Matrix [AMT17, DS20, FS24, GPHA16, GLW^{+}25, SW95b, ZJLA22, AKBF19, AB14, ACM09, AQ00, Bal00, BB24a, BKS07, BRRS15, BLM17b, BMT25, BZ96, BHRY21, BRW17, BP95, CG05, CCBGV08, CP04, CKL03, CCJ99, Dat99a, DS21d, DMH18, DMPP99, DP85, EAV16, ESE20, Fuj99, GS17, Gab02, GZW22, Gem23, GGMP88, GGR97, HB02, HV22, Hey10, Huc99, IB24, Jam95, JMS99, JCN94, KCS07, KM21, Kru99, KGZ24, LZZ22, LTDY24, LDP^{+}14, LIPT18, LP00, Lu98b, MH16a, MB20, MS99a, NRZR12, O’L87, PB21, Pea16, Per03, Ram96, SW95a, ST09b, SK97, SW06, SBBC21, SD22a, SD24a, SD25, SSA^{+}22, SW09b, SSKS21, TW00, TM15, WS21, Win04, WG18, WPS18, YP18a, YP18b, Zen21, ZY23]. matrix-dependent [GGR97]. Matrix-free [SW95b]. matrix-inverse [BMT25]. Matrix-oriented [FS24]. matrix-valued [LZZ22]. Maximal [AH11, AA94, Khe91, Mil17]. maximization [KSP10, SS25]. maximized [LS05]. Maximum [DYZ20, Hor99, Kra92, ZQZ23, BKK25, FP02, FHK05, GZQS23, GJlL23, GPHA22, HPH20, HZZ25, SYY20, SG16, SW24, TZ21, Vej10, YYZ23, Yao24, ZYQS21, ZYQS23, ZSY20]. maximum-norm [FP02]. maximum-principle [YYZ23]. maximum-principle-preserving [GZQS23, ZYQS23]. maximum-principle-satisfying [SG16]. Maxwell [AA22, AG05b, AM10b, BC12, BG03, BFH09, Bür12, Bür13, CXZ09, DCC14, FJP17, Hou23, HZ12, KCW16, MM18, MF23, MD25, Ngu15, Nke07, OGS20, Pet00, QDA^{+}26, RR14, RA09, SC25a, WSY18, XK25, YH00, YT00]. May [CFTW08, Mul99, Ano21n, Ano22n, Ano23q, Ano25r, FJ97]. McCracken [Pot85]. McKean [LSW23]. McKendrick [HT24]. MDRK [CS24]. Mean [Abu04, AL24, Fan19, MDD14, WC11, BS12, DMS23, RU25, TZA13, YCY12, dVA02]. Mean-square [MDD14, WC11, BS12, RU25, TZA13]. means [EH08, MS02, VV07, Wen10b]. measure [CGW20, DESZ25, Has08, LW04, PS19, SS19, Sob24, XL09b, Zha07, ZWK15]. measured [HM17, NTHC21]. measurement [KKN^{+}17, LSY21]. measurements [Cos25, HM15, KK17, KK20b, Muo23, QL16]. measures [DRS19, DDRS24, EHNR24, LMW23, dACR10]. Measuring [Sha85a]. Mechanical [CHH15, HLL09, Arn95, Arn98, BH93, FPS15]. mechanics [Car09a, NA21]. mechanism [JZS20]. mechanisms [BFQ22]. mechanobiochemical [MM20b]. media [AK09, AKT97, BM04a, BM04b, BBL02, CCD^{+}20, CHNN20, CCK03, CML05, CD13, DCSB25, DD19, DE06, DN13, DG22, EH07a, ESS15, EWW99, FBS09, FJ09, GS20, GJR03, HJ09, HCW16, IT07, JL25, KMR09, LLN25, MK14, MLJ19, MCM12, NS21a, OL18, SS94b, SY08, TSFB01, VBVA22, WCW14, XZL07, YR22, ZYSZ14, ZGR23]. medical [TSB10]. medium [AD15, Cho13, DTQ^{+}20, KT05, KKN^{+}17, Mur15, QL16, SR88a, XK25, ZQLK11]. Mehrotra [Car09b, LLL12].

Mehrotra-type [Car09b, LLL12]. Meiko [HP91]. Mellin [Lau17b]. membrane [Boz11, YSZG25]. membranes [CGJ16]. Memory [LVW21, ADM22, BKAG22, CFRA08, FVB05, Fre91, JP93, KAS22, LW92b, LCJQ12, LGZH24, MRV93, MFAD23, VSG17, WM08, WH19b, Win92, dv95a]. memoryless [ABKG21]. MEMS [HWZ22]. Menten [Wan09]. Menten-type [Wan09]. merging [Ber15]. Meromorphic [Jac88, DMGVO05]. Merton [Bi20]. Mesh [BBV13, Bak86, BS96a, FL93, AL98, AGJM04, AKL08, AF89, BMGM12, BKK25, BHJ05, Bak89, BKM13, BM01, BLM17a, BFA93, BS97a, Bre02a, BT97b, BC97a, CHX26, CST18, CL24b, CJLS98, CLGD06, DF11, DEPS15, DMR18, DF92, FdSB02, FMP04, FR01, HW97, HWCF15, HR97, JP08b, JM94, JOL23, Kał22, KTYY24, LBLT13, Len00, LH09, LLT07, LS99b, Man96, MCE^{+}09, Nag22, NS03, NRR06, PD96, PH15, QM03, RN25, Ren14, Rou20a, SFJ^{+}05, SYG^{+}05, Sim94b, SS08b, Wan17a, WTF25, XLHL25, YSZG25, Zar17, ZL22, ZGR23]. mesh-dependent [LLT07]. mesh-moving [BFA93]. meshes [ANN19, BSGU94, Bec18, BW95, BS05, BBRS97, Bof06, BP97, BTC23, CHR03, CS08, DRVA20, DSZ15b, DL06, FHK05, FVB05, FMGN94, FL01a, FM11, GNUV24, GHH09, HMP14, JHGZ20, KFOF02, KDD23, KSJY25, Li01a, LL20a, LVW21, Mai09, Mat09, MST09, MKH16, NV23, NFAE03, PSP04a, PT19, Pic05, QM03, Que21, QPT23, RG22, SE93, SY08, SA19, SW24, TC22, Tob14, TS24b, Tur86, WG11, WZ17, Xu21, XYZ24, YZ17, YZ22, YZ19, ZL21, ZWN23, ZSQ20]. meshfree [CCD^{+}20, GM17, LZ20, ZLX22]. meshing [ÜSHT03]. Meshless [Hou23, Li11, MS13, AD21, AAD14, AD18b, AD19b, BGG^{+}21, BK21a, BHJ13, CNA23, DS21c, GZZ19, HBA^{+}24, HS23, KDT17, MD10, MD20c, RP17, SD13a, Sch08a, SR09, SJ18, Shi20, SG17, ZT06, DM09a]. meshless-type [SR09]. meso [MW24b]. meso-scale [MW24b]. message [BC99]. meta [KS10]. meta-programming [KS10]. metaheuristics [BdFPSdSC08]. metastatic [DLS22]. Method [BL15, BO87, DT15, De 88, DC18b, FG96, HT94, Pet87, Ta’86, TMS87, TT03, AS11, AD19a, AD21, AZHD23, AS21, AH09, AZA22, AS13, ABZ21, ALMM01, AGZD22, AY15, ARGA00, AMRR18, AGM09, AP16, AP20, ANN19, AC15, AK21, AEK23, AHT17, ABK12, AS20a, ASA20, AMT13, AA94, ADG^{+}24, AK00, AY22, ATW20a, ATW20b, AT15, ARY23, ADM22, AFIS24, AMV17, AHR12, AMR14, AF23, ABKG21, AB10b, AT93, AB07, AL17, An20, AX20, ALZ^{+}21, Ang06, AJ24b, AL22, AR18, ASV19, ACP23, Arc06, AB14, ABM17, AMP20, AM16a, AMH03, AF89, AJK20, Ars20, AC16, AKS21, AAD14, AD18b, AD19b, Aug89, ABG^{+}15, AEF^{+}14, ÁMS17, AV00, AKA19, AEN22, Aze22, BAA22, Bac14, Bac16, Bac17b, Bac17c, Bac18, Bac19, BTBR19, BTBR20, Bac21a]. method [Bac21b, Bad20, BBV13, BLD17, BMSZ21, BHL^{+}21, BWM21, BW23b, Bai25b, BP14, Bal00, BS21, BLW07, BF01, BFS17, BNS25, BBCR22, BR25, BBJ25, Bas21, BAD13, BLY17, BVT14, BO04, BRTB19, BBBK22, BGG^{+}21, BHSW16, BHSW20, BDNV19, BGGG13, BC08a, BCGS24, BDOG19, BM04a, BDF89, BF92b, BW96a, BCJ97, BBC^{+}25, BTP96, BLM17a, BK21a, BZ17a, BC08b, BSZ22, BGH08, BC01, BC04b, BBRS97, BBD08, Bof06, Bog16, DLS22, BVV09, Bor97, Bos09, BJ00, BHJ13, BS20a, BRBB18, BR20, BJTZ20, BSP04, BHRY21, BSTT22, BCV21, Bür12, BH12b, Bus06, BS96b, BC00b, CGS19, CGT13, CFX08, Cai15, CHP19, CA21, CCL22, CGH23, Cai24, Cao98a, CHZZ06, CXZ14, CHZ14, CHLA21, CDP19, CP94, CGG02, CM02, Cas06, CGG25, Cau08, CS94, ÇD17, CM97, CHX26, Cer25, CjW18, CHNN20]. method [CJ18, CMP03, CKB13, CS01, CX01, CC04b, CZY08, CXZ09, CHX13, CSXL14, Che16, CXZ17, CLTA18, CH19, CWX21, CW22, CL24b, CWC24, CSL^{+}24, CL26, CC08, CS08, CWY20, CG21, CPC25, CC20b, CCK08, CH13, CK20, CKM10, CCZ22, CEW00, CL09, CR23b, ÇK13, CJLS98, CJ23, CJ25, CY05, dCCSR03, CG16, CNS00, CMS04, CGTTN24, Coy12, CSCM96, Cui24, CH25, Cum95, CH90, DAH24, DD19, DM12, DV20, DS21b, DB97, DW00, Dav92, DS21c, DS24, De 93b, DG10, DM09a, DA16, DA18a, DA18b, DA19, DSA20, DS20, DJ10, DDNZ18, DD20, DDGN23, DDF24, DK20, DYX09, DM09b, Den15, DFLM19, DHWL22, DL22b, Den93, DLP06, DN13, DS97a, DS97b, Din93, DHL00, DN08, DBBH14, Dol14, DL16, DCN^{+}19, DSS20, DP85, DP90, DCY20, DCL23, yDqGnJT09, DYF23, DZW24, DII15]. method [EBEHAS24, EJS04, EEAS25, Elg17, ER18, EK95, EL97, ESE20, EE20, EW97, EH07b, FM21, FK23, FID18a, FID18b, Fan11, FW08, FWL18, FJS25, FZA25, FL05, FFY08, FV85, Fdi96, Fel06, FLH22, FW22, FS23b, FSX25, Fer09, FvdMS20, FL20, FdSB02, FSB97, FG13, FPPS00, FV99, FR01, FM07, FMU15, FSWZ19, FXCW21, FLL11, GS24, GS15a, GLS09, GM10, GZZ19, GZZ20, GHHG22, GC25, GNUV24, Gas92, GM18, GHH20, GH25, GK09, GD23a, Gen10, GG19, GD21, GAZD25, GÖS20, Gla93, Gol86, GRLL01, GDS^{+}15, GO18, Gon06, GS09, GLM09, GT93a, GT02, GM17, GOP06, GKT10, GKA17, GGS16, Gu19, Gu20, GH21, GCZZ23, GJR03, GJV08, GH07, yGyZ07, yGpY09, Guo15, GLMY17, GJ17, GLM18, GW20, GAX24, GWZ25, GLQ26, GL93, Gwi09, GK22, Hab08, HBA^{+}24, HS21a]. method [HPY92, HHR12, HHYD20, Han93, HJS97, HZ09, HLZ14, HZ21, Har10, HLMP09, HM01, Has13, HHL23, HHC08, HL08, HDY21, HD22, HTSZ23, HO24b, HS22, HSS07, HS19a, HS17, Heu00, Hey20a, HJL18, Hop23, HA16, HDS20, HS23, HLW25, HWCF15, HJ21, HZC22, HW22, HZCZ23, HZZ24, HR96, HC01, HLC01, HZ12, HFL12, HP15, HCW16, HH18, HS19b, Hua20, HAY20, HT20, HZAT21, Hua21, HAC22, HS09b, Hus20, HSY18, IAH25, ILS19, Ius97, IS23, Jad94, JJ94, JRT90, JM25, Jeo09, JK14, JS24, JCL18, JZS20, Jia02, JHGZ20, JL17, JL23b, JLZ25, JL24, JLH13, JCJP21, JM06, JT06a, JD09, JOL23, JK20, KKT16, KMS19, Kal96, Kam25, Kam16, KDT17, KV07, Kan89, KXK92, KNN03, Kan04, KJL12, KKN^{+}13, KV95, Kat89, KM19, KL23a, KM25, KK20a, KL23b, KOR18, KSHB21]. method [KO92, KLR25, KSMMM16, KS02, KK09a, Kim12, KwS19, Kim21, KYC03, KWLK00, KKW00, KKN^{+}17, KKLD21, Kop86, KJ99, KPR12b, KP03b, KS09b, KLS13, KQ13a, KQ13b, KR20, KR18, KSJY25, KK23, KLSW10, KAS17, KX03, Kwe01, Kwe03, LH11, LHH96, LWT07, LHC09, LZQ22, LW93, LCHR03, Lau17b, Lau17a, Lee94, LL15, LLKJ21, LCK22, Leo10a, LHS00, LWD^{+}09, Li98, LFB00, Li00b, LY08, Li08, LH09, LY09, LWWX10, LK14, Li16, LYF17, LHW17, LLHC17, LR18a, LZ18, LR18b, LLHC18, LMQZ18, LWCH19, LCS19, LWY20, LZ20, LR20a, LCHW20, LGS21, LFQH21, LH21, LL23, LB23, LQXK23, LL24, LTDY24, LC24, LWaZ24, LLZL24, LDC24, LHY24, LD25, LGC25b, LC25, LP25, LR24, LM21, LD97, LR01, LX09, LWZ22, lLH25, LYZ25, Lin10, LT01, LK07, LN08, Liu09, LHÖ13, LYK17, LW18b, LLZ19, LC20, LMTW20, LZCF21, Liu21, LLZ^{+}22, LYZW22, LT23, LMW23]. method [LYLL23, LZIZ23, LWC24, LS24b, LYMD24, LHT20, LMG02, Lot19, LAH09, LfX15, LXCM21, LB21, LDH^{+}24, LOM98, Luc95, Luo18, LJ20a, LR20b, LJ20b, lLX22, LBCN00, LLW20, MZZ17, MQO17, MM18, MB20, MAHZ21, MZN21, MDRR11, MD00, MMKN17, MH14, Mai06, Mai09, MOS12, MP11, MD19b, MKBS26, MPSS16, MZS10, MCS16, MZXX24, MW24a, Map05, Mar99a, Mar05, MPHFP23, MT94, MP20, MMRV20, MMT90, MC17, MZK05, MS99a, MZM20, Men23, MM07, Meu91, MG18, MOV24, MYSC17, MAG13, MOSW00, MF23, MSS21, MD10, MS13, Mit22, Mit24b, Moe98, MDA24, MG22, MH16b, Mok17, MKJ23, MDASAO21, MLJ19, MD96, MN08, MRFF17, MLB97, MVBS25, Mou03, MK19, MWYZ18, MC21, MT20, Mul19, MPMD21, MM20b, NX22, Nag22, Nak05, NS21a, NMKE13, NDM20, NLS18, NLS20, Ney95, NK24a, Ngu15, NLHT25]. method [NWL^{+}22, NCYC22, NH15, Nov03, NS16, OT22, Odi19, Oji88, OFY^{+}23, OL18, OGS20, OH20, OHA25, OEAS21, PL20, PWY21, PZMX16, PG21, Pap95, PK23, Par04, PM05, PNA21, PBC08, PAP17, Pat00, PK21, PT11, PQ25, PR90, PH15, PR22, Pet92, Pfl08, Phi87, PP00, Pło23, PS21, Pot97, PRS23, PG02, PSL18, QM10, Qi24, QWX20, QL16, QM20, QDA^{+}26, QMLC15, QXG21, QPT23, RR21, RP17, RMCG04, ROL19, Rah25, RN25, RE19, RZS21, RG22, Ran20, RK08, RSR23, RTW21, RS20, RV04, RV05a, RMM12, Ree03, Ren14, RR00, RN22, Ric94, RVD00, RU25, RS08b, RAOC18, RG02, RREP^{+}20, Rou20a, Rou20b, RMK09, RM25, RO16, RU15, SSS^{+}23, SRK21, SD13a, SKS^{+}25b, ST11, SN22, SB18, SZE20, Saz24, Sch16b, SS00, Sch98, Sch99, SW11, Sea09]. method [Sel14, SG04, SOB20, SSC23, SD11, SD13b, SLW17, SZ22a, SWFK13, SWW17, SC25a, SGS20, Sha21, SDG20, SLMD21, She96, SY05, SW07, SXP09, SLJ11, SWY^{+}23, tSqWyG16, SWW16, SJ20, SL20, SW20a, wSJP15, SD22a, SD22b, SD24a, SJ18, Shi20, SVB17, SP22, Shy86, Shy91a, Shy91b, SLW24, Sim91, Sin24, uIVS13, SW09b, SSvG10, SB25, Sod91, SS08b, SY08, SZL18, SR24, ST25, SL15, SMA01, SSA24, SS09, Ste05b, SW85, Sti03, SG05, Su94, SWW11, SvdVvD06, SWCH15, SW20b, SH21b, SWJ^{+}24, SW86, SSS21, Sza94, TLP18a, TLP18b, Tah96, TL07, TWMP20, TT20, TS24a, TYKK01b, gTpM07, TM24, TC22, TW00, Tem23, TM21, TTP24, TWH21, TWD23, Top21, TH23, TV91, TM04, TM05, Tse00, TC19, VO00a, VRC21, VBVA22, Wal19, Wan96, WKM04, jWyG08, jWqW09, WC11, Wan11]. method [WN12, WZL13, jW15, WZ16, Wan17b, jWjJ17, WQ17, WCSQ18, WJW19, WZ19, WMLB19, WR20, jWS20, WTY21, WaZW21, WZ22, jWC22, WL22, WCL22, WCM23, WH23, WWLL23, WC24b, WPAZ24, WW24, WQS^{+}24, WWC^{+}25, WTF25, WC14, WW14, WaZ24, WSHC20, Wen25, Wen98, WBCK02, Won08, WWS07, WPT19, WL21, WX22, WG23b, WWH24, WdG92, XXYZ24, XL09a, XY19, XC20, XFL22, XZ22, XL23, XXQ17, XZH19, XGQ20, XZT21, XHJM21, XGHM22, XHYM22, XXF22, XLHL25, XWM26, YXT17, YMD21, YS09, YCY12, YY13, Yan18, YJZ18, YW19, YLX21, Yan21b, YXN21, YZC21, Yan21a, YS22, YQCZ22, YYZ23, Yan23, YS24a, YZ24a, YT25, YZZ25, YLL21, YWW23, Yao24, YZ22, YT00, Yi12, YCWH23, YÇ16, YBW20, YZ19, YHT23, YJJ^{+}24, YS24b, Yu08, YLS^{+}09, hYqW12, YWSL20, YLW21, Yua93, YH07, Yua20, YXX19, YRV21a, YRV21b, ZTZ15, Zak20, ZH21, ZAED21, ZHS22, ZRA23, ZL17, Zen21, ZLY23]. method [ZGL98, ZT06, ZOZ09, ZY14, Zha14, ZHJ14, ZYSZ14, ZQY18, ZL18a, ZFZ19, ZLHW19, ZZX19a, Zha19b, ZBY19, ZJ19a, Zha20a, ZYLL20, ZCY20, ZD20, ZLW20b, ZLCH20, ZLW20a, ZN21, ZR21, ZG21, ZWFX22, ZYC22, ZL22, ZYJZ23, ZZPJ23, ZSS23, ZYZJ24, ZGF^{+}25, ZD25a, ZZF25, ZM25, ZMF25, ZZZX26, ZZZ19, ZYX20, Zha21b, ZLL22, ZYH23, ZGR23, ZLG24, ZYJD25, Zhe07, ZZ17, Zhe19, ZZ20, ZS21a, ZHL03, ZHL08, ZH15, Zho17, Zho18, ZZC^{+}18, ZSY20, ZC92, ZWL11, Zió99, ZK00, ZCC11, dFN00, dOF20, dlC23, vdVS08, vB95, AI19, BGG^{+}20, CFXZ06, PT09, RB12]. methodology [AGJM04, CG14, FV87, KSK^{+}25, PRST02]. Methods [BH85, CFTW08, DGCW17, FF06, FFMZ13, MP85, Per88, Rei85, dH95, AH11, ABH14, AH15, AHJM19, AJ19, ABH22, ACDP22, AHJ^{+}23, AN25, AM95a, AB17, Abu04, AMN24, AAL21, AB88, AyLqW18, AHA23, ADK94, AS04, AA22, ATAW25, Alb96, AES15, ADR17, All24, AM99, AM00, AMP03, AMCM08, AMCR17, ABR23, AV96, AKBF19, AH17, AAI^{+}93, AM95b, AGM95, ABI22, AHS03, Ant23, Ant25, Ant13, AKL08, ABF09, AFS11, AFS00, AFS02, ASCM02, ASC03, Arn93, Arn98, ARS97, ACM09, AC96, AAB^{+}22, AFK92, AFS96, ABRW18, ÁMS14, AV91, BKAG22, BMQW16, BMGM12, Bai96, BN99, BY09, BW23a, BUL23, Bai25a, BKM95, BDGP96, BT97a, BDP99, BD07, BPS19, BJ02, BLS^{+}17, BOEP00, BY00, BES18, BJ05, BK06, BASC17, BLN25, Bea04, BZ92, BZ93, BV96, Bel97]. methods [Ben96, BB15, BLJ21, BNV06, BM89, BMS89, BW95, BG02a, BG03, Bie12, BT19, BDF94, BSQ96, Bla01, BCR01, BCF^{+}13, BCT19, BCET22, BCET24, Bog20, BDFF23, DSM22, Bot97, Bou25, BGP11, BJ11, BBD24, Boy91a, Boy91b, BVB09, BFdO07, Bra00, BCJW17, BCJP18, BJ20, BIJ23, BH93, Bre85, BZ94a, Bru93, BMT93, BT93a, BT93b, BT95, BT98, BM02, BM06b, BIM15, BIW24, Bru92, BMM97b, BMM97a, Bru07, BCDR25, BH97, BJ03, BIMV19, BH12b, Bur85, BC89b, Bur91, Bur93a, Bur93b, BB96, BB98, BT00, But85, But93, BC95, BJ96, But96, BW96b, BC97b, But97, BT97d, BJ98, BC98, BS00b, BW03, BR05, BP06b, BD17, BSW93, BS02, Cah89, CH22, CC24, CIJ17, CH95b, CIZ96, CdFN01, CG05, CMRV11, CSSZ20, CMRdlT24, CSSZ25, Cam99]. methods [CPP02, CGJ16, CGA96, CM13, Cao98b, Cao01, CHLX07, CW21, CCP17, CDP17, Car23, CG92, CP05a, CDGA17, Car09a, CL85, CD95, Cas96, CP05b, CS19, CCZZ18, CJL13, CG89, CM00, CCM02, CG13, CCC08, Cha98, Che96, CR05, CK98, CS03, CH07, CL07, CCLT10, CL10, CZ12, CHH15, CDW19, CZHX19, CSLY19, CWHF19, CLLM21, CWP21, CYWH22, CDW23, CLL25, CK06, Chn17, CL20, CSX23, CS24, CN16, ÇY22, CD20a, CD20b, CGS20, CJ24, CFL^{+}20, DCSB25, Con01, CM07, CX08, CP09, CDP12, CP17, CMPP24, CBHM19, Coo89, CGPT19, CRHTV25, CST97, CN11, CN17, CPR93, CKM15, Cul95, Cve02, DIJ12, DP12, DMS23, DGM22, DVV93, DS05, DR09a, DR09b, DK11, DS07a, DS07b, DK21, DGN12, DL20, DW21, DL21a, DM11b, DGRS09, DMPSC16, DSW96, Dou91, Duj09, DLV24]. methods [DP21, EAV16, EH07a, EM05a, ELvdHS98, ELCWS98, Ein18, Eir95, EHM01, EGV25, Eng11, EG88, EK97, EH24, EH97, ER07, EP15, EFLFP09, Ere19, Eva94, EC07, FMMK01, FD16, FS23a, FW07, FPRA09, Fat12, FLMR14, FH20, FH22, FXY22, FL04, FMPP24, FS08, Fer93, Fer96, FJL21, FM95, FMW05, FS88a, Fou00, Fra04a, Fra06, FG09, FHV97, Fre98, FWW^{+}21, Fuj02, GH91, Gan09, GHW20, GS99a, GM16, GV18, GANT02, GCHR06, GMG02, GP23, Gar10, GGLR09, GM93, GIS23, GHH09, GAW09, GY94, GLPW09, GG95, GAH25, GGM95, GM95, Ghe97, GD09, GKKM21, GS20, GK19, Gje07, GTS20, GÖ20, GP01, GBBC^{+}23, GPMR95, GPMPR03, GPHA06, GPHAM12, GPHA16, GPHA22, GPHAPPR23, GPHA25, GML00, GMGF02, GO19, GM94]. methods [GKL07, GQ08, Gug05, EEJB22, GR93, GHW01, Guo26, Guo01, GHF00, GR02, HM21, EJR25, Hag15, HM87, HZ96, Hai97, Has09, HJ05, HLT07, HS20, HV22, He24, HGF24, HO24a, HW06, HMdV03, HK93, HM00, Hes00, HH10a, Hey10, Hig96, HR06, HMR25, HL02a, HMP14, HS09a, Hoa15, HO05, HSWW24, HL23, HXW15, HHW18, Hor93, Hor98, Hor05, HT19, HCY18, HJYL19, Hou23, Hsi06, HCGW22, Hua00, HCX03, HLL09, Hua09, HFL13, HJX^{+}19, HY24b, HKP89, HL89, IM98, IM02, IMMS20, IKR^{+}22, INR01, IVA93, IRC12, IJ17a, IJ17b, IJ21, JMDN^{+}22, JL91, Jac93, JVZ95, JVZ96, JVZ97, JR00, JZK06, JN02, JP19, Jay95, JM25, JM16, JKW12, JP08b, JT18, JLZ20, JQYM23, JEG10, JL23a, JLWY25, Joh01, JS25, Jun97, JGK11, Kam16, KT05, Kau93, Kau95]. methods [KC94, KC19a, KMG09, KB21, KESYB23, Kie95, Kie15, Kim07, KS08, KS09a, KK09b, Kin94, KW95, KW98, KNT13, KCC04, Kok08, Kop89, Kos02, KKL^{+}25, KR25, KS89, KM11, KK02, Kre07, KXR^{+}04, Kru99, KK17, KDD23, KKR15, KDKW20, Kuz90, Kwa09, KYI17, Lab98, Lab99, LHHR94, Lam13, LPT94, LPT16, LP05, Lay08, Lay09, LT12, Lei99, LPR00a, Li00a, LPZ00, Li01a, Li05, LX08, LHH08, LLY11, LA11, Li12, LZ14, LZ17, LW17, LWL18, LW19b, LW19a, Li19, LL20a, LSP20, LLJY20, LW20a, LCW20, LWW20, LCHW20, LLY21, LX21, LZ22, LH24, LZ24, LGSW25, LZW17, LKJ07, LFL14, LS13, LQS21, LVW21, LLS^{+}96, Liu02, LS07b, LZW20, LC21, LM22a, LY24a, LNZ12, LMP99, LT93, LP97, LCM24, LW95, LMWZ07, LS05, LRT99, LXLY25, MDD14, MK20, MOS02, MMP09, MAH18, MHA19, MKDC25, MP96, MKN23, MDT05]. methods [MSGM23, Mar93, Mar03, MRV93, MO01, Mat09, MPPR22, MST09, MR94, MG97, MII13, MP97, MOU14, MSS^{+}15, Min04, Mir20, Mit24a, MFAD23, Moo95b, Moo95a, MAF20, MSA20, MAH22, MDP23, MRS10, MEGW23, MK99, MKS12, Mur15, Mur98, Mus11, NRWF08, NN20, Naj20, NNJ23, NS20, NS21b, NBNTGV11, Nic86, NFAE03, NS13, Nov08, OS08, ONL89, ÖPT25, OZ96, Oos95, OG08, Ore93, Ost93, OT02, Ost02, OTK04, OS12, OCVW22, Pag25, PGA93, PL25, Pan21, Par21, PP24, Pat98, Pav00, Pec09, PLB22, PA91, PSR04, PP92, PS00, PA05, Phi91, PGP03, Pir09, PSW02, PWS05, PWS06, PT95, PCR17, PAJ12, Por17, PCRR17, Pul86, QLL^{+}08, QW25, QWW24, RG20, Ran15, Ran16, RSY12, RU21, RT20, RKVZ15, Rha97, RTV00]. methods [RTV02, Ric91, RFMV24, RX08, RTA19, RVM23, Roz05, RBT15, SMTHE22b, SSW20, Sae14, ST09a, SPS20, San20, SH97, Sar05, Sca22, Sch08a, ST89, SW09a, SW95b, SW98, SWE05, SWJ09, Sch12, SW95d, SGS00, SZ09, Sch95a, Sch87, Sch95b, SEGV02, SB14, SR09, Sha98, SAH24, SI20, SWX00, SL09, pSLqJcY16, SHLY19, SA00, SC08, SJ11, SWB20, SD25, Sid90, ST20, SG07, SD09, Sv95, SKW17, SW18, Som86, SvdHN86, SCvdH92, Som93, SK96, SW03, SdSC99, Spi95, Spi96, Spi13, SS13b, SW12, Sto96, SS12, SL08, SFZ21, SW05, SGN08, TSF24, TLQ21, TMD92, TSB10, TDC13, TOD11, TB01, TYJ11, TK15, Tru00, Tur93, UHLZ24, VB07, VDVV98, VV02, VV07, VV09, Van00, VS94, VZ93, Ver06, VS95, Ver96b, VS91, VMS07]. methods [VMP03, VN21, Wal95, WGKS12, mWyG00, WZL08, WM08, WL09b, WL09c, WL09a, WG10, WL10, WZW13, WGW15, WMF17, Wan17a, WH18, WT20, WCJ23, WSP04, WEA12, WKP12, Wei95, WYL11, WHL19, WSW96, Wen05, WCGW95, WSS97, WZ17, WG22, XFLC00, XZ19, XZW19, XWZ21, XG22, XF22, XZZ19, Xu21, YGY15, YZWQ25, YK07, YG99, YW08, YZ17, YLFT20, YT21, YZ24b, YFLX20, YXX24, YLLZ21, YF24, YR92, YXB95, Yu99, YW24, YW25, YTZZ18, Zen93, ZZ18, ZZL01, ZH09, ZC10, ZZHS18, ZJ19b, ZLJ20, ZFC20, ZFW20, ZG20, ZL21, ZLW22, ZYQS23, ZZ24, ZWG25, ZZW97, ZL11b, ZFX17, ZW19a, ZZX20, ZSJ04, ZZJ21, ZZ19b, ZKO^{+}21, ZP97, ZX09, ZCSH11a, ZQLK11, ZCSH11b, ZX14, ZSQ20, ZSQ21, ZX22, ZC99, Zla85a, Zla85b, dAF17, de 92b, dRT99, de 95b, in 92, in 95, in 96, in 02, ZJR25, van95, vdES04, vdHVW01, vSC92]. methods [vS93, vC93, van93, vS96, van96, vS97, vdHMdS99, vdHS01, vdSvdH95, AGM09, BGT97]. metric [HT20, MM20a]. metrics [CLGD06]. Metzner [ZB19b]. MFE [ZYSZ14]. MFEM [CLY19, HL19, SY18]. MFS [SK10]. MG [JMP06]. MGT [BFQ22]. MGT-type [BFQ22]. MHD [AEK23, AI19, AA22, DM09a, DA19, HSS04, LD22, QM10, RSK24, SD11, WWC^{+}25, WPT19, ZBD24, ZZZX26]. MIC [GKMS09]. Michaelis [Wan09]. Mickens [MdD04]. microbiota [CDP^{+}25]. microenvironments [MDD18]. microgravity [CMP06]. micromagnetic [TSFB01]. micromagnetics [LW04, MPPR22, XL09b]. micropolar [HSF25, SED21, ZL23, SLW24]. microscale [CL02b, LC02]. microstructure [GP01, MS00]. microstructures [Li00b, Li01b, RS22]. microtemperatures [BCFQ21]. midpoint [SR97]. migration [MM20b, NS16]. milling [IMM04, MI03]. Milstein [GHHG22, RT20, RKVZ15, WL09a, WG22, Yam18]. MIMD [SD93]. mimetic [APJ10]. MIMO [ACKV24]. min [PK21]. min-orthogonal [PK21]. Mindlin [KX03, Lam13]. Minimal [Bar12, BFGP08, CP10, LS13, CKB12, Cul95, Hin97, Kie95, DE18, PL20, PR22, Ren14, Zha97]. minimal-norm [PR22]. minimally [BRIP08]. minimax [WL25]. Minimization [Dra97, Gab02, WW99, BHRY21, CXZ14, DW00, FBM17, GZW22, HT20, KKR25, Sch95a, WH23, YP18a, YP18b, ZHL08]. minimizing [Bor97, HJP10, MKH16]. minimum [BSvdV99, GL17, Sch96, Tad86]. MINLP [EW08]. MINRES [lLH25]. mirabilis [Aya09]. miscible [LHY24, NS21a, SY08, ZYSZ14, ZYZJ24, ZGR23]. missing [AGKS25a]. mites [CL02a]. Mittag [ABdSG23, Boh21, KZ21, OB24]. Mittag-Leffler [ABdSG23, Boh21, KZ21, OB24]. Mixed [BBS25, BMM97a, CjW18, GH07, HSS04, LS24c, MN24, MZ04, Pan21, TJK18, VB07, VMS07, jW15, YZ24b, dSFDG20, ADM22, AD23, APJ09, Ars20, AFLG^{+}12, BGG04, BBG14, BLY17, BRS05, BCGS24, BK21a, Bra00, BJM01, BMM97b, CGS19, CCOVF22, CL08, CP05a, CD23, CGG25, CS03, CZ04, CJX11, CWHF19, Che12b, CMS04, DY17, DRVA20, DV20, FL20, GMM09, GIS23, GHH20, GD21, HDY21, HEG16, HL03, HW97, HCY18, HJYL19, HLW25, HAC22, JNPC03, Kie17, Kim21, Kwe01, Kwe03, LM00, Li00a, LCHW20, LLZL24, LS12, LX09, LLZ19, LLW22, LS24b, Lot19, MMP09, MKDC25, MSZ^{+}24, MS90, MZM20, MF23, O’L87, OR20, Par04, RTW21, SW11, SY08, SW85, SG05, VBD93, Wan05, Wan20, Wen25, XYZ24, YY13, YJ23, YXZL24, ZZHS18, ZBY19, ZSG^{+}20, ZZPJ23, ZYZJ24, ZBNC25, ZZC^{+}18, iW07, iW09]. mixed [iM13, BTDV10]. mixed-FEM [BGG04, CCOVF22, GMM09]. Mixed-hybrid [VB07]. Mixing [dB03, Gro94, MS91, MPV24]. mixture [CPY20, CG14]. mixtures [HPW21]. MLDG [AD18b]. MLPG [DM09a, AS11]. MLS [MS13]. MMOCAA [ZYSZ14]. mobile [GWLN22, JL17, LCZ25, NWL^{+}22]. mobile-immobile [JL17]. mobile/immobile [GWLN22, LCZ25, NWL^{+}22]. mobility [ASGGRGW25, GGT24, GJLL20, HZZ25, KK09a]. mock [Ibr25]. mock-Chebyshev [Ibr25]. Modal [FID18b, Dat99b, RGÖS18, XC20]. mode [DG96, EH91, ZY19, dPT96]. Model [BO87, DC21, HK09, Mac86, MVG14, MPtM16, TD09, WM07, AD20b, ASGGRGW25, AK21, AHT17, AB15, AJ24b, AL98, AD04, BKAG22, BC12, BFS17, BM05, BCFQ21, BVT14, BK09, BF09, BDOG19, Bho11, BSZ22, BTMT08, Bi20, Bou02, BSTT22, CGGGS11, CML^{+}25, CHM22, CHLA21, CCQ^{+}23, CF13a, CKP15, CST18, Cha17, CXZ15, CZY18, CH19, CPY20, CLL25, Chi21, Chr96, CL09, CH89, CF13c, CA15, CR04, CT21, DvHM19, DGGM25, DMR18, DGM18, DMM24b, DL16, DLQZ23, DYF23, EE20, FF20, FKA^{+}13, Fou00, GNUV24, GHK16, GLV03, GM18, GHH20, GH25, GKB^{+}22, GT19a, GeO24, GLLY25, GD22, Han19, HJ09, HGF24, HJZ23, HMD21, HGZW21, HWZ22, itHT18, IK24, IMC22, ITZ17, JK17, JN07, JL17, JT06b, KWY24, KCW16, KSJY25, KAS17, LSVT25, LMY18, LFQH21, LQXK23, LWaZ24, LC25]. model [LR24, LFL14, LN08, LC21, LS24b, LZY24, LG02, MD22, MPSS16, MT11, MF23, MT05, MAD23, MHL18, MLJ19, MM20b, NK24a, NA21, OZHP23, OHA25, Par21, PB21, PHY19, PQ25, PVH25, QAE^{+}09, QZH25, QH19, QCW^{+}23, RMCG04, RV22, RÁM23, RLHC19, RA17, RKR20, RW87, RG21, SMJ12, SH97, SZ22a, SYL^{+}20, SWY^{+}23, SC25b, SG16, SZ22b, SL21, SB19, SWJ^{+}24, SAMSB20a, TWL23, TKN11, TLV92, Udd20, WL09c, WWLL23, WPAZ24, WWZ25, WTF25, WS22, WSHC20, WCGW95, WMC09, WYP12, YZQ^{+}22, YCY12, YV17, YZC21, YD22, YSZG25, YEZ25, ZW09, ZQY18, ZZHS18, ZY19, Zha19b, ZCY20, ZFW20, ZLW20a, ZJH^{+}23, ZH24, ZZLL21, ZD25b, ZAB15, van98]. Modeling [BH85, LC02, Per88, AL20, AEA23, AFIS24, BLS94, Bai02, BLW02, Beh97, BDNV19, BGG12, BMP05, BBT25, BBO03, CM09, CKM15, ESS15, GKB^{+}22, GJR03, HM00, IJ14, IHS13, IM00, KGR08, KKL^{+}25, KDAK16, LCJQ12, LARGVR23, MDD18, MCM12, Mur15, Neu88, NBP94, Ram94, SKAW12, SWFK13, TDPU17, YC00, ZBNC25]. Modelling [BW86, RCGM98, BBPR05, BDES12, BBCS05, BR94, BCS06, BIMV19, CC23b, DS03, For25, FMSV07, Gwi09, JRW06, KRBK16, LBLT13, Rob10, SMC08]. Models [FDF^{+}25, AM95a, ALM04, ABD16, BL15, BAD13, BBV05, BCCHM21, DLS22, BN03, BBD24, Cai15, CFCH09, CMP20, CGEV19, CDV00, CD13, CF08, dCCSR03, CDI^{+}24, Dal00, DEGDdL25, DMQ02, DS03, Dun18, DT89, EP15, EL01, FV01, GDEdLD23, GV02, Goo90, HL97, KKT16, KD13, KK20b, LS07b, Lte24, MLK06, MSGM23, MM14, MPV24, MOV24, MD20c, Mus11, PKP19, PCRR17, RP17, RR14, RS22, ST11, ST14a, SLJ86, SvdHK94, TSF24, VVR08, WDL23, WC25, WK02, YK07, YXX24, ZC92, vSK97]. modes [CCS02, Mul99]. Modification [Bas21, BKM13, CJV88, NLS18, TWMP20, VO00b, Waa88]. Modifications [Rei85]. Modified [CLLM21, CWP21, EH24, GZZ20, GC25, Guo15, Has08, LS99a, MH16a, Mar94, SSS^{+}23, SY03, SL08, Yan18, ZWFX22, All24, AF23, ABKG21, AKA19, BHL^{+}21, Bog12, CHLA21, CL14, CL09, Cui24, DP12, DDRS24, DL21b, Est95, Geb24, HPY92, HP85, HDY21, HWY20, JPP19, KDH20, KM18, KR18, LY09, LHC23, LGSW25, LK07, LN21, LC21, LB21, Men23, MK19, QZH25, SW09b, SH21a, VVD95, WL10, WW14, WYP12, XWW19, XC85, XXQ17, XZZL15, XP23, YL26, YS24b, YLW20b, ZHJ14, ZLHW19, ZR21, ZS18, Zup04]. Modifying [CGTTN24, VVV24]. modular [LFQH21, SA12a]. modulated [Pul05]. module [SW95a]. Modulus [WG18, HV22, MH16a, SH21b, Wal90, ZR21]. Modulus-based [WG18, HV22, MH16a, SH21b, ZR21]. MOESP [SMC08]. MOL [HS98, Hin95, Now96, WSP04, Zha96]. MOL-applications [HS98]. MOL-systems [WSP04]. MOL-treatment [Now96]. Molecular [CS01, AB12b, CZY18, KWY24, Lub04]. mollification [AM09, HDY21]. moment [AL20, Bis11, CXZ15, GDEdLD23, LR01, LS07b, Njå88]. moments [CS17, DFC09, MJS23, SMW21, Sin23]. momentum [CCOVF22]. Monge [JLWY25, SWR11]. mono [VDVV98]. mono-implicit [VDVV98]. Monod [BM05]. monodomain [NK11]. monomial [Has20]. monostable [YEZ25]. Monotone [ABY22, Bog16, HHAA22, Wan07b, AKM^{+}22, AKA19, BAA22, Bog12, Bog20, CRU15, FBM17, GRLL01, GJ17, HN22, IKR^{+}22, Kam25, KKLD21, LSWW22, Pel20, SSW20, SSS^{+}23, TGV22, WL09b, YLS^{+}09]. monotonic [AEF^{+}14, DKK94]. Monotonicity [CMR12, DMR10, Eir95, BY22, GCHR06, Hor02, IS23, mWyG00]. Monte [FSX25, JL24, LHÖ13, MD06, DE18, Sob24]. monument [DSSC13]. Morley [HLC01, SW21]. Morocco [BDES12]. morphodynamics [KDAK16]. morphology [QPT23]. Morrison [Mit24a]. mortality [HT24]. Mortar [BC08b, FMW05, Ars20, CX01, CH07, EJS11, HLC01, JLZ20, WCXL09]. mortar-type [CH07, HLC01, WCXL09]. mortars [Ste05a]. Motion [PGM86, AACP20, FMU15, HJ21, HT00, JK21, KOW05, SHL19]. motivated [FGP23, VCC12]. mound [SR88a]. mountain [DH12b]. movement [Bai97, DF11]. Moving [BW86, MSS21, Pet87, ADFR18, AD01b, AD19b, BHJ05, BHJJ06, BP14, BFA93, BBD08, BC97a, CHR03, CRR03, dCCSR03, FE93, Fer96, GZZ19, GH20, HR96, HR97, HBJ09, LCHR03, LSV22, LBLT13, LRRÁY25, Li11, Li16, LSY17, LXZ21, LR03, MD20c, MD21, MD23b, MM20b, NLHT25, RL06, RLHC19, RTT01, SD13a, SS08b, SWCH15, Wag98, Zha96, Zho17, Zup03]. moving-boundary [BBD08, Zho17]. moving-mesh [PH15]. MP [HVY91, van95]. MPE [AC98]. MPE-iterative [AC98].

MPI [AAD14, Nak05]. MR [Kni95, Tsy96]. MRMIL [SKS^{+}25b]. MSABC [PT09]. MSOR [Wan96]. Multi [BC00a, GS20, Kom07, Kuz90, LW04, LW17, Log04, MS24, Pul05, QM19, AHAS21, Arn95, BW23b, BvG19, BDOG19, BZ17a, BS12, BF95, But93, CNA23, Cer25, CCS17a, CZ19, CaAL96, CDG19, CJ25, DDS89, ESEKZ10, FPPS00, Har93, HPH20, HZD21, HLMKZ06, HZAT21, Jun97, KL21, KXK92, KCC04, KKN^{+}17, KCW16, KM11, LX08, LLV18, LH24, LTDY24, LW07, LLZ19, LW20b, LYC24, DLM16, LLW20, MAHZ21, MZN21, MKDC25, MSGM23, Mar09, MM16, Mit24b, MPG^{+}16, Mur98, NN20, NAD25, OFY^{+}23, OGS20, Par14, Poh93, PCA10, QWX20, RKVZ15, SGY22, SWL20, SBS^{+}20, SA18, SM89, SZW19, TLG20, Tor06, WWX13, XL09b, YLLZ21, ZAED21, ZH20, ZJ19b, ZFW20, ZYQS23, ZWK15, ZCSH11a, ZQLK11, ZCSH11b, ZSQ20, ZSQ21, iW09, BO11, ZY25]. Multi-adaptive [Log04]. Multi-atomic [LW04, XL09b]. multi-block [SGY22]. multi-body [Arn95, Mur98]. multi-channel [RKVZ15]. Multi-colored [Kom07]. multi-core [MM16]. multi-dimensional [BZ17a, BS12, BF95, HPH20, KCC04, DLM16, Mit24b, NAD25, OFY^{+}23, OGS20, SBS^{+}20, ZAED21, iW09]. multi-domain [DDS89, FPPS00, MAHZ21, SM89]. multi-electron [PCA10]. multi-element [ZWK15]. multi-factor [BDOG19, Mar09]. multi-frequency [BvG19, KKN^{+}17, KCW16, LW20b, Par14, SWL20, WWX13]. multi-Galerkin [NN20]. multi-grid [Jun97]. Multi-hp [GS20]. Multi-interval [BO11]. multi-layer [KL21, ZFW20]. Multi-level [Kuz90, QM19, LYC24]. multi-linear [LLV18]. multi-medium [ZQLK11]. multi-objective [ZH20]. multi-parameter [SGY22]. multi-person [CZ19]. multi-physical [LH24]. Multi-Quadrics [ZY25]. multi-resolution [Har93, MZN21, ZSQ20, ZSQ21]. multi-revolution [LX08]. multi-scale [KM11, MPG^{+}16]. multi-scaled [CJ25]. multi-splitting [Poh93]. multi-stage [But93, Cer25, CCS17a, Tor06]. Multi-step [BC00a, LW17, MS24, BW23b, MSGM23, TLG20, ZYQS23]. multi-symplectic [HLMKZ06, ZJ19b, ZCSH11a, ZCSH11b]. multi-term [CNA23, ESEKZ10, HZD21, HZAT21, LLZ19, LLW20, QWX20, SA18, SZW19, YLLZ21]. multi-time [DA18a]. multi-trace [CDG19]. multi-valued [AHAS21, LW07]. multi-variable [KXK92]. multi-view [LTDY24]. multi-wavelets [CaAL96]. Multiblock [Li00a, Ars20, DS15, FV99, Oos95, dPT96]. multibody [Aca12, EK97, HMW05, Sim93, Sim98, WSW96, Wen98]. Multicomponent [BCV25, DD19, EWW99, QW04, YH18]. multiderivative [Dit21, SS21]. Multidimensional [BG03, KW10, MKS12, AR15, AM16b, BOEP00, BBBK22, Bra22, CXNF14, CRU15, CJLS98, Cuy90, FCX06, GPHA25, GS21, EJR25, Hag15, HHYD20, HKZ08, LWL18, LLW22, MD20b, OP04, RBT15, Saz22, Saz24, SZQH23, YTC24, ZC92, ZL24, iM13]. Multidomain [LLT20b, MQO17, Tro96, YH00, BO11, JOL23, Kim95, Kop86]. Multifluid [DRC85]. multifluids [CFX08]. Multigrid [De 02, DMPSC16, FH08, OG08, Sha98, Ta’86, WZ02, Zha97, ABCC18, BMR^{+}17a, BRSD91, BW97, BDFF23, BVV09, Bre02a, BG02b, CDRT19, CZ90, CK98, CH07, DCSB25, DS97a, EVO06, FL01a, GGLR09, GGN12, GGR97, GJV08, HY02, HST14, JT06a, Kan04, KH91, Kor95, LWT07, LP05, Mic95, Mol95, NN10, Oos95, PS09, Pfl08, Phi87, RDH^{+}12, SWX00, Spi99, Spi00, SMA01, SH21b, TLV92, VP91, WZ17, XHJM21, XGHM22, YXX19, ZR21, vBvdZdB08]. Multilayer [GDEdLD23, EHNR24].

Multilayer-moment [GDEdLD23]. multilayered [NBP94]. Multilevel [AG05b, BM12a, BC04a, GT19b, HW97, Hu99, HN22, Ney95, XHYM22, ZC92, Bai96, BJ00, BCSH16, CLT97, DM12, FMGN94, FPS15, GKT10, GGR97, HKZ08, lLNW21, MB10, Mai09, Moe98, MRS10, Not99, Osw97, SST04, Sob24, Zha00, ZW19b]. Multilinear [LLVX20, ZJR25, Bou25, XY24]. multilinear-rank [XY24]. multimaterial [JN07]. Multimesh [PD96]. multimode [GKB^{+}22]. multinode [DDF24]. multioperators [Tol03]. multiphase [KMR09, WK02, ZC92]. multiphysics [GM18, GHH20, GH25, NS21b]. Multiple [ELLE02, HS86, HLIS16, KSP10, LZJ21, SAG86, WGB99, hYK86, BN12, BGH08, BG14, Boy91a, CZY08, CGTTN24, DRVA20, EJS04, JWZ21, KS04, KD25, KYI17, Lee10, Lei02, MS86, MDA24, Mon09, PK21, PH91, RMK09, SDK24, SW95b, pSLqJcY16, SHLY19, SWL20, SAMSB20a]. multiplication [FT96, MA04]. Multiplicative [FL01a, Mai06, DE16, Hou23, KKP07, KR12, LLY21, MT20, NN13, TN16]. Multiplier [ONL89, ABG^{+}15, PAP17, RM25, SW20b, ZZX19a]. multiplier/fictitious [SW20b]. multipliers [BMSZ21, BGG04, KwS19, ZN21]. multiply [BS96a, RSK24]. multiply-refined [BS96a]. multipoint [HS19a, JL86, JL87, MW93]. multipole [DM12, Eng11]. multiprocessor [CSS87]. multiprocessors [AAI^{+}93, IVA93]. multiquadric [Jun07, JD09]. Multirate [GM18, GR93, BCG21, BKP09, Chi21, EL97, HS09b, KP15, PG02, Pul09, PSL18, Ske89b, SG05, SBG09]. multiresolution [BBRBS09, DGRS09, FR14, GDS^{+}15, JP17, PD96, PGS10]. multiresolution-based [PGS10]. multirevolution [MP97]. Multiscale [CL08, CHNN20, DC09, DCN^{+}19, EH07a, MLJ19, WCW14, BBD24, CKB13, CC04b, CK20, CGK25, DDHS97, DYX09, HSY18, Jeo09, JEG10, LNZ12, RZS21, RSK24, Rob10, SKAW12, SQ17, WPAZ24, YÇ16, ZLS20, ZS21a]. Multispectral [CdCV03]. multisplitting [DS20, Liu02]. multistage [ABH14, BJ98, JVZ95, JR00, MAF20, MSA20]. Multistep [CP09, FMMK01, VZ93, ZL18b, AFS00, AFS02, Arn98, BCS17, BT95, CM13, CL85, Cha98, Den15, FHV97, GM16, Gje07, IJ17a, LMG02, MD19b, MST09, MPDB24, ONL89, OTK04, PGA93, SvdHN86, Spi95, Spi13, in 02, vB95]. Multisymplectic [AM04]. Multithread [RRMJ12]. Multithreaded [Chr96]. Multivalue [ACDP22, Bur85, BC89b, CD20b, MDP23]. multivalued [RGS25]. multivariables [Ari87]. multivariate [AL20, AEG12, CC04a, CV88, EGM25, MD06, MRÁSY24, OGV92a, PGS10, WT93, de 92a]. multiwavelet [DJJ^{+}15]. multiwavelets [AFS11, CS17, HAML21, MB08]. Munthe [MEGW23, AB17]. Müntz [BKM19]. MUR8 [EL97]. Murakami [CN16]. MUSTA [Tor06]. mutualism [RKR20]. myelinated [LFL14].

NACA [De 02]. Nador [BDES12]. Naghdi [DL16]. Nagumo [BSTT22, LYZF25, ZLW20a]. Nanjing [LST07]. Nanostructures [KNP16]. nanotubes [NT20]. narrow [Con89]. NASCA [BJS12]. Nash [CZ19]. natural [AHJM19, CdCV03, Che16, DdCVR03, DY03, HY01, Hua17, HS97, LS20, Sch91, Ste05b, SFZ21, VZ93, WWF20]. natural-norm [Che16]. nature [DRVA20, VS91]. Navier [ASGGRGW25, GHS25, KM16, Kni95, LH23, LXPY25, AD20b, AR18, BLRGVR23, BB15, BG11b, BCGS24, BC01, CCOVF22, CHLA21, Cau08, CHOR19, CSXL14, CD13, CGS20, DY17, DJ10, DS97a, DN08, DJL04, FD16, GM10, GNX19, Guo00, GH07, Guo15, HJR22, HHC08, HL08, HW15, HFL12, HSF25, ITZ17, Joh01, Kal96, KS00, KCL00, Kni94, KDK17, KN93, Kwe03, LRV25, LA11, LLY21, LS23, LH23, LL24, LD25, LHX20, LN21, LR00, LO96, LJ20a, Med96, Nor97, OK98, QR03, QMLC15, QAMX17, QM19, RZS21, RV05a, RK91, SH10, She96, SA00, Shi25, Shy86, SLW24, ST86, TH18, TLV92, Ton04, VG04, Wan25, WY02, WaZW21, WWM22, YT25, YÇ16, Zha14, ZL23, ZS21a, ZS21b]. Navier-Stokes [Shy86]. NCP [PGP03]. Near [PGM86, Ari04, BRIP08, KLSW10, PL20, RS08b, Sch02]. near-breakdowns [RS08b]. near-critical [Ari04]. near-minimal [PL20]. near-minimally [BRIP08]. nearby [DMPP99, SL01a]. nearest [GKS20]. Nearly [CD20b, BRS05, DN13, ELR^{+}15, Fat10, Mou03, Win01]. nearly-singular [Fat10]. Necessary [Meu14]. necessity [BC05, BDDV12, DBDV10]. necking [NMSF94]. need [Mau08]. needle [Li01b]. needle-like [Li01b]. negative [BC05, CLS04, LDP^{+}14]. negativity [AHT17]. neighborhood [Bar05, YZH19b]. neighborhoods [Zha07]. nematic [CPY20, WaZW23]. Nemytskii [EHV19]. Nernst [LL24, PYD21, YPD21]. nerve [LFL14]. Nessyahu [ASC03]. nested [Imo00, KS09c, Li23, PSP04a]. NETNA2015 [DGCW17]. network [ADNR21, Boh03, CCL22, CFS13, CG03, CDP^{+}25, DVV93, EM05a, EHNR24, PCL^{+}25, Pel20, SG05, SBG09]. networks [Bho11, Bor16, BCK22, DJNR22, CG16, DJR25, DC18b, JRW06, SMC08, SG16, SOA^{+}25]. Neumann [AS11, AMH24, Bie12, CQZ20, DP85, DY03, EH24, FL15, HM17, HMY19, KFR25, KvyS15, Lau17a, LLHC17, MH14, MKBS26, O’L87, OMP98, PA18, RNG22, RBC02, VBVA22, jWqW09, jW15, jWS20, XF22, hYqW12]. Neumann-to-Dirichlet [Bie12]. Neumann-type [HM17]. neural [Boh03, CCL22, CG03, CDP^{+}25, JRW06, PCL^{+}25, SMC08, SOA^{+}25]. neuron [Bho11]. Neutral [Jac87, AD19a, AD21, BP06a, BLL24, Bru07, EH97, Gan09, HHT97, HS19a, JL91, LLD18, RN22, WZL08, WC11, WC14, YXT17, ZH09, ZLJ20, ZYX20, ZJ19c, ZP98]. neutral-type [ZJ19c]. neutron [FdSB02, RGMO19, RREP^{+}20]. neutrons [RREP^{+}20]. Neville [AGP97, GP93]. Newell [MDRR11]. Newmark [KW12]. Newton [LHT20, Agu15, AB07, ABM17, AMP20, Aro96, BLS^{+}17, BHSW16, BHSW20, CH22, CKB12, Car94, CCY22, CLL23, Fre91, GM93, GO18, GH21, Gug05, HH10b, HCW16, HST14, JKN94, JLZ25, JLL90, KR25, KYI17, LMP99, MYSC17, PKSB10, PR22, PGP03, SKS^{+}25a, SW09a, SY05, SL09, pSLqJcY16, SHLY19, Spi95, ST14b, WZ16, XY25, XZZL15, YH18, ZWFX22, ZZW97, Zha01, ZYH23, dOF20]. Newton-like [JKN94, GO18, ZZW97]. Newton-multigrid [HST14]. Newton-PSBTS [ZWFX22]. Newton-type [PGP03, pSLqJcY16, SHLY19]. Newton/GMRES [PKSB10]. Newtonian [CCLT10]. NFFT [HNP17, Nes16]. Nicolson [ZJH18, ZR15, AD20c, ATAW25, DLM20, FP02, FSWZ19, HJYL19, KTYY24, LN26, LFB00, LL19, LLZL24, LZCF21, LJ20a, QH22, QXG21, TH18, WH19a, WZ22, YCWH23, ZL17]. Nicolson-type [QXG21, WZ22]. Nicolson/Adams [ZJH18]. Nicolson/fourth [AD20c]. Nicolson/fourth-order [AD20c]. Nine [Pot85, Lyn92, WG11]. nine-point [Lyn92, WG11]. ninth [ST20]. NIPG [RG22, XLHL25]. NIRK [KK20b]. NIRK-based [KK20b]. Nitsche [BDDV12, BH12b, DZ12a, DBDV10, MAG13, ZKO^{+}21]. no [BtTBV87, Die20, Kni95, Tsy96]. nodal [HMdV03]. node [Bai97, DTGN23, EN09, SGN08]. node-centered [EN09]. node-centred [SGN08]. nodes [CMP15, CHS17, Ibr25, KK09b, OT21, PL20, SR09, SMJ24]. Noether [Dor01]. Noether-type [Dor01]. noise [AHO16, CGH23, CWC24, CL18, CDI^{+}24, Hou23, HLY22, Kha21, KZ13, KR12, KKR15, LLY21, LS24c, LT19, MH04, MRS10, MT20, SD24b, TN16, WYY20]. noise-removal [MRS10]. noises [Bac18, BTBR20, ST19]. noisy [KLS13]. Non [ADSS17, AMR12, BS00a, BFdS10, FR14, FBM17, LY10, DBH25, LDP^{+}14, NTT22, NLZB23, PSP04b, PPS05, RTA19, SCLL21, TJ12, XL09b, AD15, AS11, ABH22, AHT17, AK00, ACMR06, AMV17, Ang06, ASCM02, AGKK94, AM10b, BAA22, BUL23, BBD18, BGM19, BRVC09, BM06a, BP12b, BCFR25, BTDV10, BO11, BMV06, Buc04, Buc17, BDD^{+}20, BJ06, BCS06, BB98, CL01a, CP97, CM02, Cat10, CLR11, CC04a, CKK10, CCJ99, CLS04, DDP12, Dea11, DDGN23, DS97b, DM11b, DN08, DSZ15b, DSZ15a, DTQ^{+}20, Ehr08, EL01, FJS25, FvdMS17, Fou00, FM11, GKKM21, GS20, GP98, Gul15, GS94, HM01, HK22, HL02b, JHGZ20, Jun97, KM25, KSHB21, KOGL25, KLSW06, LH11, LPS25, LSV22, LLKJ21, LRC19, LR20a, LGS21, LDC24, LC99, LVW21, LMTW20, LS24c, LOM98, LJ20a, MDRR11, MVVA09a, MST09, Mon21, MT20, MRÁSY24]. non [PK23, PCL^{+}25, PGDB08, PFHL09, Pis22, QH19, QZM25, Rab25, RS09, RV04, RV05a, RGMO19, Rou20a, SRK21, Sch93, Sch04, SNOK21, SB14, SYG^{+}05, SG00, Sin23, SH21a, SA18, SA19, Ste05b, TSF24, TLQ21, TW00, TTP24, TWH21, TWD23, UHLZ24, Vab21, VL08, WTB24, WZ14, WT17, WG22, XYHM20, YM24, Zha20a, Zha20b, Zha21a, ZX22, iV09, Pet92]. non-asymptotic [Sch04]. non-autonomous [CL01a, DM11b, LMTW20, MT20, SB14]. non-axisymmetric [FJS25]. non-clamped [BBD18]. non-classical [AS11, CCJ99, Sin23]. non-commuting [BB98]. non-compactly [Ehr08]. non-conformal [CLR11]. Non-conforming [XL09b, DDP12, DN08]. non-constant [LPS25, TSF24]. non-conventional [Rab25]. non-convex [AM10b, BDD^{+}20, CP97, CKK10, LS24c]. Non-cooperative [SCLL21]. non-cylindrical [BCS06]. non-Darcy [EL01]. Non-degeneracy [PPS05]. non-equidistant [DS97b]. Non-equivalent [PSP04b]. non-existence [SNOK21]. Non-Fickian [BFdS10]. non-globally [LDC24, WG22, Zha20a, Zha20b]. non-gray [GS20]. non-Hermitian [SH21a, Zha21a, Pet92]. non-homogeneous [AMV17, BP12b, PFHL09, YM24]. Non-iterative [NTT22]. non-Kerr [DTQ^{+}20]. non-linear [AK00, BGM19, BM06a, Buc04, BJ06, KLSW06, LC99, Mon21, PK23, SRK21]. non-Lipschitz [TLQ21]. non-Lipschitzian [WZ14]. non-local [ACMR06, Ang06, BMV06, MVVA09a]. non-lumped [GS94]. non-magnetic [LSV22]. non-matching [BTDV10, Ste05b]. Non-monotone [FBM17, BAA22]. Non-negative [LDP^{+}14, CLS04]. non-negativity [AHT17]. Non-optimal [RTA19]. Non-oscillatory [TJ12, ASCM02, DSZ15b, DSZ15a, PCL^{+}25, PGDB08]. non-overlapping [BRVC09, Jun97, LOM98, RV04, RV05a]. non-parametric [AGKK94]. non-periodic [BO11]. non-perturbative [Cat10]. Non-polynomial [ADSS17]. non-positive [QZM25]. non-quadrilateral [TW00]. non-rationally [Sch93]. non-reflecting [Dea11, GP98, SG00]. non-self [ZX22]. non-self-adjoint [Gul15, KOGL25]. non-selfadjoint [HM01, XYHM20]. Non-separable [AMR12, MRÁSY24]. Non-Sibsonian [BS00a]. non-singular [KSHB21, LLKJ21]. non-smooth [BUL23, FvdMS17, HK22, KM25, TWH21, TWD23, UHLZ24]. non-spatial [LGS21]. non-standard [FM11, MDRR11, RGMO19, WT17]. Non-stationary [LY10, Buc17, CC04a, LJ20a, QH19, Vab21, WTB24]. non-stiff [ABH22]. non-symmetrized [HL02b]. non-trivial [CM02]. Non-uniform [FR14, NLZB23, AD15, Fou00, JHGZ20, LH11, LRC19, LR20a, LVW21, MST09, RS09, Rou20a, SYG^{+}05, SA18, SA19, TTP24, VL08, iV09]. non-vanishing [Pis22]. nonautonomous [GHHG22]. noncompact [Maj14, YT21]. Nonconforming [AMN24, BMS89, JP08b, LS24b, YLL09, CK98, CZS04, CH25, GHT05, HD22, HCJW24, JK14, Kan04, Kim12, KOS^{+}12, LWW23, LR24, LN21, MCS06, MZS10, MP20, Nud24, Osw97, Par04, Ris05, SY18, VMS07, WSY18, ZCZ15, ZSS23, ZZZ19]. nonconstant [CCL04]. nonconvex [CCY22, GZW22, GT02, HZC22, HZCZ23, LWLL24, PS19, SL08, WH23]. noncooperative [CZ19]. nondegenerate [KL98]. nondifferentiable [MH16a]. nonempty [Lee23]. nonexistence [BSV09]. nonexpansive [AHAS21, VA21]. nonhomogeneous [BJ00, jWS20]. Nonlinear [Bar05, BL91, De 93b, DSS15, EH91, LW20b, MSS^{+}15, Que21, RBBC85, Ren13, SM85, SW13, SCT05, Tan01, WZL08, AC98, AZA22, AN25, AKM^{+}22, AGLRS23, AAL21, AI19, AS20b, ADR17, ABI22, AAM03, AB07, ASZ18, ALM04, AJ24b, AR24, AGQ^{+}24, AMP20, AD18a, AD18b, AD19b, AAB^{+}22, AO91, ADH00, ÁMS14, AKA19, AEN22, Aze22, Bac14, Bac16, Bac17a, Bac17b, Bac19, Bac21a, BKK25, BY09, BUL23, BHJJ06, BLS^{+}17, BGG04, Bas21, BBD20, BLN25, BMM03, BNH01, BS10, BK21a, BK21b, BKP14, BSV09, BSQ96, BBRS97, Bog16, Bog20, BGP11, BVRB14, BRBB18, BMV19, BC23, BT98, Bru92, BC97a, BKW06, BIMV19, BRS^{+}18, Bus06, CHP19, CCL22, CC23a, CNA23, CCG13, CW98, CS94, CHNN20, CJL13, Cha17, CR05, CL07, CZY08, CDW19, CWHF19, CC20a, CWP21, CWY20, CKK25]. nonlinear [CNS00, CMPP24, CGPT19, CGTTN24, CRHTV25, Cui24, CMP23, DS21b, DW00, DS17, DN21, DA17, DM09b, DZ12b, Den15, DL20, DHWL22, DL22b, Die15, DKL24, DSSC13, DL16, DHM09, DCY20, DTQ^{+}20, DAMA23, ESEKZ10, ESE20, EJS11, EHV19, EHV24, FID18a, Fdi96, FWHM20, Fer96, FI03, FZM20, FV87, FJS99, FL01b, GLS09, Gan09, Gan96, GS99a, GLLW14, GM16, GHHG22, Gar96, GHKM09, GAZD25, GMS12, GO18, Gon06, GGNP02, GT93a, GKA17, Gu20, GWLN22, Güm20, GW20, GD23b, GLQ26, HM87, HT24, HJS97, HMN20, HA21, HCS20, HS20, HTSZ23, HCC24, HdSRI17, HZ20, HZD21, HMD21, HK93, HM09, Hey19, Hey20a, Hey20b, HAA21, HAR21, HBB25, HS95, HMY19, HJKW17, HJL18, HJYL19, HL21, HL24, HLJ20, HCGW22, Hua98, HJX^{+}19, Hua20, HZAT21, IKR^{+}22, IAH25, Iga85, IK24, Ius97, JZK06]. nonlinear [JUAZ22, JRS20, JRT90, JHGZ20, JW01, Jia12, JZZH22, JLWY25, KM95, KN19, KK06, Kim07, KNT13, KSSS16, KOGL25, KKLD21, KS04, KKP17, KK17, KK20b, KTYY24, KLSW10, KDKW20, LO22, LRS23, LPT94, LZQ22, LW93, LHN24, LN26, LOS03, LLKJ21, LA11, LHW17, LSWM19, LR20a, LW21a, LGS21, LZ22, Li22, Li23, LZ24, LC24, LLZL24, LWLW24, LGC25b, LGSW25, LWWZ25, LSG24, LKJ20, LY03, Liu09, LZL14, LL21, LLZ^{+}22, LSWW22, LY24a, LS24a, LZY24, LHT20, LMP99, Luo18, LYA^{+}19, LLW20, Ma03, MH16a, MZJ^{+}25, MMP09, MDP10, MD20b, MD22, MK14, MP11, MS19, MK21, MKN23, MS02, MN24, Mik97, Moe98, MD23a, MEGW23, Naj20, NYPW21, NDM20, NC16, Odi19, OB24, Oji88, OFY^{+}23, OT02, OTK04, OXW25, PXHZ20, PK23, PM91, PV24, PD01, Per99, PA91, PVM22, PR22, Pło22, QNA23, QM20, QZM25, RE19]. nonlinear [RZS21, RV22, RN22, Rha99, RLHC19, RT14, RV05b, RGA19, Rou20a, SSW20, SRK22, SHL19, SH97, Saz24, SH09, Sch09, SS99, Sch98, Sch87, SRMDRL23, SC25a, SY05, SY07, tSqWyG16, SYY20, SZ12, SZ17, SJ20, SW20a, SXL22, SZ22b, SWB21, SVB17, SC03, SS09, SG92, SL01b, Tah96, TS24a, TQA^{+}25, TWH21, TH23, TH09, UWY22, UHUL21, UHLZ24, VSG17, VP91, VMP03, mWyG00, Wan01, Wan07a, Wan07b, WL09b, WC11, Wan11, WZ14, WGW15, WZ16, Wan17b, WL18, WW19, WMLB19, WH19b, Wan20, Wan21, WY22, jWC22, WL22, WCM23, WXY24, WLY24, WTF25, Wei09, WP99, WSHC20, WB03, WYYL19, WG23b, WL25, XZ19, XY19, XLZ20, XZL07, XZZ19, XLKY19, XGQ20, XCHW22, XXF22, XP23, YMD21, YH18, YLY19, YLFT20, YZC21, YS22, YWH20, YF24, YLS^{+}09, YLH20, Zak20, ZAED21, Zha96, ZH09, ZJ19a, ZYLL20, ZG21]. nonlinear [ZWFX22, ZQZ23, ZSS23, ZM25, ZZW97, Zha01, ZZX20, ZYX20, ZPT92, ZW19b, ZZ19b, ZJ19c, ZSZZ20, ZZLL21, ZFX25, dlC23, in 95]. nonlinear-nonlocal [QZM25]. nonlinearities [GS25b, LYA^{+}19, NYPW21]. nonlinearity [Fuh01, HK22, LN26, LC24, QW25, WWL21, XLKY19, Yan21a, Yan23, ZCQ25]. nonlinearly [MŠ99b]. nonlocal [ADFR18, Bor02, BIMV19, CCQ^{+}23, CC25, CFL^{+}20, DAH24, Deh05, DM09b, FL23, GZQS23, GLQ26, HGZW21, KBS11, LP25, LL20b, Ma24, MSZ^{+}24, MVVA09b, MOV24, NY13, PD01, Pło22, QXQ22, QZM25, VS91, WWLL23, YQCZ22, ZXYW22, ZWJ18, ZD25a, ZW19a, Zhe19]. nonlocally [EEE22]. nonmatching [CH19]. nonmodal [Dat99b]. Nonmonotone [BKAG22, ZH15, BBD18, MD23a, YLL21, Yu08].

Nonnegative [HZZ24, BWY03, DESZ25, LTDY24, GLW^{+}25]. nonnegatively [FLMR14]. nonnegativity [Hor93]. nonnegativity-conserving [Hor93]. nonoscillatory [HOEC86]. nonparaxial [MP11]. nonquadratic [NH24]. Nonreflecting [Rya00]. nonselfadjoint [CX01]. Nonsingular [XZH19]. Nonsmooth [LO96, ZXW17, AGZD22, Aca12, AAB^{+}22, BG02a, BRW17, BRS^{+}18, CLX21, LZH^{+}25, LWLL24, Nke07, Ril92, SND21, WH23, Zak20, dOF20]. Nonsmooth-data [LO96]. nonsolvable [CC90]. Nonstaggered [Li25]. Nonstandard [KCC04, BF01, CMMR23, LYZF25, MCD20, Mic03, VRC21]. nonstationary [Cao01, CHLA21, Qiu23]. nonstiff [AJ19, AFS00, BF92b, Cha98, CM07, CX08, Pan07, VS94]. Nonsymmetric [GH20, AL05, Cao98b, Cao09, CK98, Jes93, KG90, KCB02, MRS10, SEGV02, YG95, YSBL14, YXX19]. Nonuniform [JM94, CS18, DSSC13, Li01a, LWW20, MS90, Tur86, KQ13b]. nonunique [CY98]. nonviscous [LKJ07]. nonzero [RS20]. Nordsieck [AH15, BJ05, PWS06]. norm [AFF^{+}15, CKB12, CXZ14, Che16, Dol14, EFLFP09, FP02, Gab02, GZW22, GPHA22, Hor99, HAC22, JM05, Kra92, LCH20, PR22, WCS21]. normal [CFRA08, CJ90, FH10, FMU15, GS09, LW22, dPT96]. normalized [Auz03, Tan23]. normed [BRIP08]. norms [Bre02a, GCHR06, HJP10, KK22a, Liu21, TM05]. Note [JMU^{+}25, ATW20b, Ano04m, AHS03, BY22, Ber15, Cao97, CS17, EK96, FS19, GKS20, GLM18, HV95, Iva07, JW01, Jun07, JK20, KM16, Li12, Lon88, Mar93, MVVA09a, MdD04, Mul19, NR14, O’L87, RSK14, TY03, Wel10a, YSBL14]. Notes [DT10, BW96b, DN24]. Novel [BLD17, HA21, LC21, NMSF94, TWMP20, WWF20, AN22, BHL^{+}21, Cha25, DSAB20, ESE20, FM21, Kam25, KSK^{+}25, KCY19, KR20, LS25, LMSW17, LCL18, MC17, NSD23, SAMSB20a, Tem23, UHUL21, WLM21, XLZ20, YZQ^{+}22, Yan22, ZZF25, ZZL17, ZZJ21]. November [Ano21q, Ano22x, Ano23p, Ano24q, Ano25x]. NtD [CAAT16]. nuclear [GZW22]. NUFFT [MZZ17]. null [Lee23, LS25]. nullspaces [SY05]. Num [AS21]. NumAn [ADG^{+}16, HKNV16]. Number [FG01, PGM86, BG11c, BS96b, CH95b, CG14, FHM^{+}02, GK19, GS09, GPíP03, HD04, KO96, Kwe01, LHH08, LH09, MP20, NCYC22, PBC08, SBBC21, dC18a]. numbering [BW97]. numbers [BBLT15, BD22, CMS06, GK93, LW18a, LL20a, LLW22, SQ17, SFZ21, Win01]. Numer [Bic21, Kni95, TLP18a, Tsy96, YP18a]. Numer. [BtTBV87]. numeric [CBHY11]. Numerical [AD19a, AN25, AGLRS23, Aff94, AD08, AJT19, Ale11, ALY03, ADFP25, AO05, ASS21, AL87, AX20, ALM04, AM10a, Ano02g, ADM10, Ari04, AS06, AC16, BP02, BCFQ19, BCFQ21, BFQ22, Bec02, BDES12, BNV06, BGM^{+}09, BBCS05, Bho11, Bho12, BR94, Bog20, Bor16, BMP05, BBL02, BG11c, Boz11, BFdO07, BFLR23, BGHR12, BGH^{+}15, BCDP17, BH85, Cah89, Cah92, CSS19, CDV00, CCMSS11, CCG13, CCdlH20, CHPV09, CST18, CEL25, CCK03, Che12a, CPY20, CBD16, CF08, CFL^{+}20, CFM^{+}24, CF05, DZ12a, Dal00, DS24, DS07a, DN21, DGCW17, DG96, Den07, DS23, DL13, Die15, DRC85, DLM02, EGM25, Eis86, EZ03, FF06, FFMZ13, FMW18, FDF^{+}25, FL24b, GM85, GNUV24, GKT10, GNAS^{+}20, GJLL20, HKS86, HM87, HLR18, HM00, Hey20b, HSWW24, HPW21]. Numerical [HJ06, IMC22, Jac87, JRW06, JUAZ22, JRS20, Jbi25, JLZ20, JCY25, JMP06, Jun06, JT06b, KX91, KW21, KOS21, KME20, KO92, KHM^{+}19, KW95, KP07, KS04, KPR12a, KKL^{+}25, KDAK16, KP19, LRS23, LW25, LA12, LP24, LMV17, Le 12, LZH19, LLY21, LHC23, LWaZ24, LC25, LLT07, LZW17, LKJ07, LO03, Liu97, LHX20, LZY24, LP97, LM22b, LRE04, MO17, MP96, ML16, MSZ^{+}24, MNR14, MF99, MS25, Mat08, MS02, MQ00, MH04, MCBV20, MMM19, Mok17, MDASAO21, Moo95c, MRFF17, Mot17, NP21, NS12, NER95, NA21, ND85, Nür09, Obe15, OB20, Odi24, OB24, OZHP23, OHA25, PT09, PTV20, Pel15, Per88, PS00, PA05, PWS98, Pot97, QAE^{+}09, Qi24, QZM25, RP17, RC18, RMM12, RLHC19, RAOC18, RMS17, RX08, RO16, RBT15, SA90, SKBAS08, SKAW12, SH02]. Numerical [SR88a, SWW17, SYL^{+}20, SSR23, SC03, Son91, SR88b, ST05, Spi90, Spi93, Spi97, SG92, Str98b, SH91, SSPZ20, SAMSB20a, SAMSB20b, TER03, TKN11, Tia15, Tsy98, Vab22, VA05, Van92, VBD93, VP91, Wan09, WWC^{+}25, Wee01, WL24, Wu09, XGM08, Xiu08, YZC21, YLW20a, Ye04, ZM19, ZYQS21, ZG21, ZYC22, ZBNC25, ZYJD25, ZJ19c, ZD25b, ZCSH11a, ZAB15, de 92b, de 93a, vHA98, AD21, ALMM01, Abu04, AACP20, ASGGRGW25, AQS94, AL95, AA22, AEMX17, AA20, AR23, AEA23, AFIS24, APA92, Ang06, AD01a, Arn93, AS00, AKS21, AAD14, AM16b, Auz03, ÁMS14, BDMG12, BBV13, BRW21, BBCR22, BGG04, Bas21, BZ92, BO04, BLRGVR23, BB15, BGG^{+}21, BDNV19, BCCHM21, BT19, Bla01, BCF^{+}13, BtTBV87, BDFV95, DLS22, Bor02, BBBN21, BHJ13, Boy91a, BFdS10, BLL24, Bre10]. numerical [BS08, BMWH20, Buc04, BIS25, BGS02, BC89b, BC89c, CHM22, CFKS07, CL08, CG92, CL85, CD95, CW98, CP05b, CHX26, CGN03, CK22, CJ18, CJ90, CZY08, CLTA18, CAAT16, CJ24, CKK25, CMMR23, CDR20, CNS00, CM04, CRSF19, Cul95, CDP^{+}25, DD21, DD19, DS21b, DDZK05, DS17, DLN^{+}24, DGGM25, DG10, DA19, DZ12b, Den93, Din19, DMQ02, DSK12, Dob05, DBBH14, DC09, DZMB21, DAMA23, ESEKZ10, EGV25, EEE22, EE20, EL01, Fai00, FS23a, FW07, FHM^{+}02, FJ97, FL15, FvdMS20, FGP23, FdSB02, FMSV07, Fra04a, FLL11, GS24, GS19, GLML20, GHS25, GQ89, GD23a, GAH25, GMG19, GAZD25, Gla93, GeO24, GML00, GMGF02, GLM09, GOP06, GO19, GKL07, GL17, GOLJ25, GS21, GGT24, GN86, Guo96, HBA^{+}24, HT24, HZBM05, Han19, Har09, HHT97, HM01, HHL23]. numerical [HMN20, HJ05, HO24b, HZD21, Hin97, Hor99, Hor02, HA16, HDS20, HWZ22, HJ21, HZAT21, HBJ09, IV16, Isk89, IKM23, IMM04, JL91, Jac96, Jad94, JK21, JK17, JS24, JQYM23, JL17, JR02, JCSR03, JT88, KMS19, KHLV22, KV95, KS25, KKR24, KD13, KAS22, Kel85, KC94, KHB22, KPR06, KK09a, KNP16, KKW00, KKN^{+}17, KK20c, KAS17, LO23, LPT94, Lau17b, Lau17a, Lee94, LLKJ21, Leo10a, LD10, LSK12, LYF17, LMY18, LZ20, LL23, LSL11, LC99, LT00, LZL14, LLD18, LTT19, LL21, LC21, LR87, LARGVR23, LFP04, LFS15, LP00, LD02, DLM16, Ma03, MDD14, MD19a, MPTT17, MHA19, MN23, MS03, MD19b, MS19, Man97, MRF00, Map05, MVVA09b, MV17, MPV24, MSS^{+}15, MP15, MDD18, Mur15, Mur99a, Mur99b, NRWF08, NK11, Naj20, NMKE13, NMSF94, NLS18]. numerical [NAD25, NT92, NSD23, NLZB23, NS16, OS08, Oji88, OCVW22, PR89, Par21, PB21, PK21, PVM22, Pet00, PT23, PP92, PFHL09, Pło23, PCA10, PMP23, PYD21, QNA23, Rab25, Rah25, Ram94, RI02, RSR23, RV22, RG05, RN22, Ril92, RLMG24, RGA19, RREP^{+}20, Rou20b, RGK21, RMK09, RT95, RTT01, SHL19, SDK24, San89, SB03, SY03, Saz22, Saz24, Sch91, SS94b, SNOK21, ST19, SOB20, SGS20, Sid10, Sid14, Sid23, Sim93, Sim91, SKO19, Sol15, SvdHK94, SS02, SSKS21, SWR11, SWJ^{+}24, Tah96, TSF24, TYKK01a, TBRBM20, TOD11, Tho85, TK05, Tor06, Tou97, Tsy96, Uty08, VO00b, VV07, VRC21, VV95, VBVA22, VN21, Wag85, Wan01, Wan07b, WL10, WQ17, WH18, WB92a, WCGW95, XL09a, XF22, XWM26, YBL13, YLFT20, YZ21, YRV21b, Zak19, ZP24, ZRA23]. numerical [ZXYW22, ZWH^{+}17, ZY19, ZJ19a, Zha20b, ZJH^{+}23, ZY25, ZL11b, Zha21b, ZML^{+}12, ZZJ21, ZP97, ZB19b, ZCC11, tV87, van86b, van98, vdHSW98, BJS12, LW19a]. Numerically [War92, CH04, MD20c, MAD23, WT08]. numerics [FJH^{+}01, HW04]. Numerov [AGM09, FW08, Sim91, Wan07a, Wan11, ZR15]. Numerov-type [Sim91, FW08]. NUMOL [Dia95]. Nunziato [TKN11]. NURBS [ZS18]. Nyström [Pat98, AMCM08, AS20c, BDP99, BBCR22, BCR01, BCET22, DSM22, CL14, CSLY19, CPR93, DM11b, Fer09, FL20, GMG04, GML00, Hoa15, Lab98, LW17, LW19b, LS05, Mur98, Pat00, PA05, RG20, Som93, Tau09, TH23, VDVV98, VV02, vSC92, CH95b]. Nyström-based [TH23].

Object [LLS^{+}96, EMMK01]. Object-oriented [LLS^{+}96, EMMK01]. objective [ZH20]. objects [DH12a, ESS15, IT07, QL16]. oblique [Sod91, WN12, WL16]. Obrechkoff [VV09]. observation [FL23, SHB^{+}25]. Observations [Pet87, AGKS25a, Quy19, YGY15, YK07, YZ24b]. observed [FV01]. observers [PGC01]. obstacle [BS14b, Imo00, KJL12, Map05, QL15, YL26, Zou11]. obstacles [AA20, WG19, YWW23]. obtain [PLB22]. Obviating [AM16b]. Occasion [CHM09, EST15]. Ocean [AL87, Beh97, Lee94, LG02, RG05, RMM12, RW87, TK19]. October [Ano23u, Wen10a, Ano21m, Ano22m, Ano24s, Ano25q]. ODAE [PP92]. odd [AS05, BtTBV87, tV87]. odd-even [BtTBV87, tV87]. ODE [Aro96, BM09, MPG^{+}16, SSV89, SKS23]. ODEIVPs [Bur93b]. ODEs [ABH14, AH15, AB97, AV96, AB98, AB09b, BK06, BT97b, BM02, BM06b, BKW06, BP06b, CCMSS11, CM97, CS24, CPR93, CMP23, Gea93, GX93, HSWW24, HS09b, Jac96, JN02, JM25, KC19a, KKP17, LS07a, LYZ25, LW20b, MW93, MQ03, Mur98, Mur99a, PK23, Pat98, SC11, Sal89, SA12b, SK16, Sha85b, Sha05, VS95, WSP97]. off [FV85, Fuh01]. off-diagonal [Fuh01]. oil [Aff94, FBS09]. Old [Cuy90]. Oldroyd [AD20b, ZQY18]. olive [BS24]. omitted [BB10]. on-the-fly [BMR^{+}17a]. One [Buc06, De 88, Gil10, KL09, Pet87, ABI22, Bac14, Bho12, BFA93, Bok03, CP05a, CWC24, Chi93, CJ23, CJ24, CY05, CBHM19, CN11, DR09a, DL13, Eir99, EK97, FP02, FD97, FdSB02, FR18, Guo01, HZ09, Hin97, Hor99, Hua00, HL89, JZZH22, JCJP21, Kam00, KTK20, Kie95, KS09b, KDKW20, LL98, LGSW25, LAH09, MQO17, MB20, MII13, Moo95b, Mus11, Nud24, Saz22, SMEN04, SLJ86, SM93, Sha85b, SJ11, SW12, Ter22, TYJ11, TM04, UHLZ24, Ver93, WZL08, Wan17a, WJW19, WSS97, XZ19, XL23, XZT21, XWM26, YDWW17, ZH09, ZZW97, ZQLK11, Bru97]. one- [JZZH22, Saz22]. one-block [Guo01, ZZW97]. One-Dimensional [Pet87, Bac14, BFA93, CJ23, CY05, FP02, FD97, HZ09, Hor99, JCJP21, KS09b, LAH09, MQO17, Moo95b, Mus11, SLJ86, SJ11, SW12, XL23, XZT21, XWM26, ZQLK11]. one-field [WJW19]. one-leg [Hua00, WZL08, Wan17a, ZH09]. One-level [KL09]. one-parameter [Hin97, Nud24, XZ19]. One-point [Gil10]. one-sided [ABI22, MB20]. one-speed [FdSB02]. one-stage [Chi93]. One-Step [De 88, Buc06, Bok03, EK97, Kie95, KDKW20, SMEN04, SM93, Sha85b, TYJ11, UHLZ24]. one-way [Ter22]. Ontario [CFTW08]. Oono [WaZW21]. Open [Gus88, Gla93, HGR01, HLL09, Kim14, Li25, Mai09, Man97, Mar94, RW87]. Operational [MZ87, HA21, KGZ24, LIPT18, Lub92, MS19, Mok17, SSA^{+}22, SSKS21, UHUL21, dlHV13]. operations [HR14, IMM04, MI03]. Operator [Bi20, CT21, GV02, LCW20, SvV22, AKM^{+}22, ABdSG23, AMT13, ADFP25, AHS03, AB14, BF15, BMR^{+}17a, Bie12, BC05, BC04b, Boh21, BFH09, Bre06, CHP19, CM02, CS19, CL24a, CC20b, FL24a, GK09, HPH20, HDY21, HCGW22, HD23, IHS13, Kan04, Kie15, KNT13, KvyS15, LO22, LS25, LZJ21, LZZ22, LC24, LIPT18, LMS08, Lte24, MS99a, MV18, MG22, MM25, Odi19, Ost02, QDA^{+}26, Roo20, RM25, SCT05, SN04, Vab21, WD22, WS22, WK00, WL24, WWLS08, YXZL24, ZZ18, ZWH^{+}17, ZZX19b]. Operator-compensation [LCW20]. operator-splitting [AMT13, BC04b, CL24a, Lte24]. operators [ACP24, AL95, Auz03, BRIP08, BK06, BCE04, BMV06, BLL24, Chr01, DV20, EH07b, GGMP88, GSR00, GP04, GS05, Haa97, HB02, HH98, HM01, HZ02, HS97, LLT20a, LLT20b, LYC24, MMP20, Osw97, PGA93, Pel15, RGÖS18, Str98b, SGN06, TLQ21, WKN20, Xu13, YBL13, YY24, vSW90]. Optical [BFS17, ASZ15, AES13, BCMV03, Den07, YLW20a]. optics [DHL00, RVdCVR02]. Optimal [AMP03, AMCM08, APA92, AW09, BTBR19, Bac21b, BL21, BS05, CC18, Chi93, CL20, CGS20, DM09b, GS99a, GP98, GS21, HQAZ24, HJKW17, HJL18, IPL02, JL25, KP18, Kał22, KMG09, KD25, Kwe03, LMWZ10, LW21a, LYZJ23, MG97, Nke07, ST09a, SHA12, TTP24, YS22, YJ23, ZZHS18, ZM25, ZWG25, ZLWF21, AA04, AA25, ATW20b, AJW23, AM99, AM00, AVMVMV09, ARSW05, Bac19, BCT16, BGH08, BGP11, BVB09, BSZ15, BH20, CKP15, CLY19, CZHX19, CL26, CL18, ECB07, EGL09, FS23b, FG13, For11, GK19, GWLN22, Has08, HA21, Hey20b, HL19, KMH21, KP15, KR15, LZ20, LAZ20, LCH20, LZIZ23, Lot19, LDH^{+}24, MZS10, MKH16, NX22, NK11, NH24, NSD23, Odi19, OAHN22, PLI03, PS21, PRS20, RTA19, Rou20a, RM25, RSD^{+}06, SSW20, SSS^{+}23, SS19, SG00, Sim94b]. optimal [TLQ21, WZZ21, WCL22, WCJ23, XL11, YGY15, YXF25, YJ21, Yan23, YZ24b, YK04b, YÇ16, YBW20, ZZ19a, ZLW20b, ZZH25]. optimal- [FG13]. Optimal-order [ZLWF21]. Optimality [BCV21, NSD23]. Optimally [KLSW06, Qiu23]. Optimally-stable [KLSW06]. Optimization [Bie87, CDI91, Jor11, LHÖ13, SS09, AH09, AD99, CH87, CLL23, CD18, CMS06, Cui24, DESZ25, DW15, EM05a, EM05b, Fer14, For02, GM08, GMZ08, GM10, GGNP02, He24, HZC22, JQYM23, KK11, Kim19, Kok08, KLS13, KL87, LS10, LS86, LZH^{+}25, LT05, LWLL24, LB21, lLXhLZ21, lLX22, ML16, Meh08, Min87, MK19, NK11, Ou11, PTW19, PCR17, Rab25, RP01, Ren99, Rha97, Rha99, SP99, Sch09, SWJ09, SZY21, SGY22, SA08, SdSC99, TLP18a, TLP18b, WM08, Wan23, YZ24a, Yu08, ZH20, ZYH23, ZWL11, ZP12]. optimization-based [Kok08]. optimization/simulation [Min87]. optimizations [LYZJ23]. optimize [CD18, IMM04, MI03, SH02]. optimize-then-discretize [CD18]. Optimized [Chn17, LCM22, OH20, QLL^{+}08, Tsi01, AKGR14, CM02, GKMS09, KKR24, Mar05, SEGV02, Ven15, ZZF25]. optimum [Meh08]. option [AD99, CF08, FV01, GGO13, GK22, KKT16, LL15, RG21, ZW09, ZJH^{+}23, ZD25b]. options [ALY03, AO05, BP14, BAD13, BNV06, BDOG19, Bis11, Bi20, CGEV19, Cer25, CXZ15, CPOGO17, itHT18, HFL13, DLM16, MHL18, RP17, Rou20b, ST11, ST14a, SB19, TGB08, ZGO12, ZO14]. orbit [DG10, vdHMdS99]. orbital [GML00]. orbits [Zha19a]. Order [BCJW17, Bur85, BT97d, But98, But09, Gar10, IN89, Jac87, NNJ23, Sim98, ZM17, ZJ10, AD20c, AD21, ACKV24, AH11, AB12a, Abu04, Aca12, AP20, AJ24a, AKG14, AB88, Agu15, ASA20, AB15, ATW20a, AES15, AEMX17, Ale11, AM99, AM00, AMC02, AMCM08, AMCR17, ACMR06, AAL^{+}26, AKBF19, AHR12, ABI22, AL17, AT13, AAEMY21, Bac17a, Bac17b, Bac19, BTBR19, Bac21a, BMGM12, BG11a, Bai02, BBS25, BCG21, BOEP00, BFS17, BBD20, BM05, BKR13, BM01, BNV06, Ber04, Ber05, BZ17a, BCC16, Bic21, Bla01, BCR01, BC05, BM06a, BCET22, BCET24, BC01, BVV09, Bos09, BTC23, BJ11, BIO24, BR20, Bra22, Bra00, BMPR15, BDRZ19, BPTT15, BRBM08, BBKS07, BT95, BEH24, BMM97a, Bru07, BJTZ20, BO21, BDM03, BB96, BB98, BS96b]. order [BJ98, CML^{+}25, CFLW22, CW25, CC24, CH95b, CSSZ20, Cam99, CPP02, CGA96, CM13, CDP17, CGA93, CHSS01, CP05b, CS19, CES91, CCC08, CDH24, CS03, CYM09, CJX11, CXZ17, CC19, CSLY19, CWX21, CL24b, CWC24, CS08, Che12b, Chi21, CRU15, Chr01, CN16, CS18, CCJ99, CDR20, CRHTV25, CST97, CD00, CN11, CN15, CN17, CLS04, Coy12, CP03b, DRVA20, DKSS24, DV20, DR09a, DR09b, Deh05, DDNZ18, DZ12b, Den15, DL22a, Den93, DEGDdL25, Din19, Dit21, DB08, DC09, DL21b, DCL23, DJJ^{+}15, DAMA23, EBEHAS24, EHM01, EJRR23, EES05, Elg17, ER18, EL97, EP15, EN09, EW97, EH09, FK23, FJ17, FS23a, FW08, FWHM20, Fer93, FJP17, FZM20, Fou00, FG09, FM11, FL24b, FHX22, FGGL22, GS99a, GX11, GLLW14, GS19, GLML20, GZQS23, GHS25, Gar96, GMG04, GP23, GHH20]. order [Gem23, GLPW09, GÖ20, GJ00, GPMPR03, GML00, GOP06, GGR97, GND19, Guo00, yGqWsWC05, yGpY09, GAX24, GHF00, HGM^{+}21, HP85, HT24, HHR12, Han19, HS11, HKO12, HAN23, HP18, HOEC86, HLMP09, HAML21, HMN20, HA21, HPH20, HTSZ23, HO24b, HEG16, HZD21, HMdV03, HS17, Hey19, Hey20a, HAA21, HAR21, HGZW21, HvdHV10, HL21, HZZ25, HST14, INR01, ITZ17, IT16, IB24, IKMM23, JM17, Jad94, JUAZ22, JJ94, JS24, JCL18, JZXJ21, JWZ21, JN07, Jia12, JL17, JWG20, JQSC22, JZZH22, JY23, JJL^{+}24, JCN94, Joh05, KWY24, KN19, KS00, KMH21, KS25, Kat89, KM19, KK20a, KC19b, Kha21, KHB22, KLY05, KL07, KK09b, KwS19, Kim21, KHYY21, Kom07, KOGL25, KM18, KCW16, KZ13, KOS^{+}12, KKP17, KLSW06, KW10, KDD23, KGZ24, KK23, KDKW20, LO22, LCVG01, LH11, LHHR94, LM25]. order [Lee94, LLKJ21, Lev91a, LPR00b, LX08, LMWZ10, LYF17, LW17, LR18a, LR18b, LMY18, LL19, LLJY20, LH20, LWY20, LXZ21, LW21a, LZ22, LXZS22, LQXK23, LH23, LH24, LYZF25, LX09, LWZ22, LTT19, LHX20, LS20, LW20b, LL21, LSY21, LCZ23, LGZH24, LWC24, LS24c, LXPY25, LMG02, LE94, LL02, LDH^{+}24, Lua17, Luc05, LJ20a, LR20b, LJ20b, MPTT17, MS08a, MZXX24, MVVA09a, Mat09, MYSC17, Mit22, MMM19, MPtM16, MFAD23, MD20c, MD96, MRFF17, MK99, Mur98, NK11, NP21, Nap16, NAD25, NSCC19, NV23, NT16, NT20, Nic86, NWL^{+}22, ÖT20, PZMX16, PCL^{+}25, PNA21, PGDB08, PHY19, PCR17, PJB04, PMP23, QW04, Qi24, QXQ22, QR24, QXG21, Qiu23, RR21, RZ00, ROB17, ROL19, RN25, RMH20, RZ18, Ran15, RA03, RR00, Ris05, RT14, Rog19, RTU15, RGA19, RTA19, Rou20a, RCR24, RO16, SMJ12, SST12]. order [Sac93, SAA20, SDK24, SA90, San02, SMB23, SM93, Set24, SB14, SSC23, SD13b, Sha87, SC25a, SY18, SC20, SXL22, SL21, SP22, SZE^{+}92, ST20, SKS23, Ske89a, SDK15, Sof17, SvdHN86, Som93, SR24, SA18, STS00, SS09, SS13b, Ste08, Str98b, Sub04, SW17, SND21, SAMSB20a, SAMSB20b, TSF24, TZ21, gTpM07, TX18, Tan23, Tan24, TQA^{+}25, Tem15, TDW23, Tob14, TZA13, Tol03, Tol04, Ton04, TS24b, TDPU17, TY00, Udd20, VV02, VV09, Vej10, Ven15, Ver96a, Ver06, VL19, Wai98, WDZS21, WKM04, Wan07b, WL09b, WWX13, Wan17b, WMF17, WR20, WDH20, WDU21, Wan21, WLM21, WHW21, WLG22, WWZJ22, WDL23, WG23a, WXY24, WLY24, WTF25, WKN20, WAV12, WC14, WS22, Wen98, WB03, WYP12, XXYZ24, XWW19, XY24, XK25, XY19, XG22, XFL22, XWX21, XZZ19, XLKY19, XZT21, Xu21, YM24, YMD21]. order [YZWQ25, YV17, YJ21, YQCZ22, YZG23, YYZ23, YJ23, YT00, YBW20, YC13, YZH24, YDWW17, YXZ18, YRV21b, ZLY23, Zha96, Zha14, ZWH^{+}17, Zha19b, ZYQS21, ZG21, ZLX22, ZLW22, ZJH^{+}23, ZYQS23, ZJ25, ZZW97, ZFX17, ZZX20, ZLL22, ZH24, ZLS20, ZZJ21, ZLWF21, ZZLL21, ZFX25, ZX14, ZSQ20, ZSQ21, ZL24, ebKMZ24, iW09]. Order-dependent [ZJ10]. Order-preserving [ZM17]. ordered [HP97, Not92]. ordering [ZY19]. orderings [Ara99]. orders [AMP03, ESEKZ10, SV00, Tsi01]. ordinary [AT93, Bac16, Bac17b, Bac21a, BJ05, BZ92, Bok03, BJ11, BJ20, BC89b, BB96, BB98, BC95, BJ96, BJ98, CL01a, CD95, Den93, Enr06, FMMK01, FH20, GGMP88, yGpY09, HJ05, IM98, Ise02, IJ17a, JVZ95, JVZ96, JAH21, KC19b, KW20, KDKW20, LMSW17, MP94, NMKE13, ÖPT25, PVM22, RT95, San89, SZ12, SCvdH92, SL01b, Tem15, TYJ11, Tsy96, WCL22, XFLC00, YC13, ZG92b, ZLX19, ZC99, in 95]. ordinate [Sin24]. orientation [ACM91, LQXK23, ZML^{+}12]. orientation-field [LQXK23]. oriented [AS05, EMMK01, FS24, GG95, LLS^{+}96, MPG^{+}16, SG05, WSS97, XHYM22]. originated [DRVA20]. Orlicz [LRS23]. Ornstein [ZGO12]. orthoexponential [Kür23]. Orthogonal [AGKS25b, Ari03, BDMGVO05, Cer25, AyLqW18, AB24, ADG11, BSFDM02, BR01, BRS16, Bre91, BDRZ04, BDRZ19, CIZ96, CKL03, CMR12, CCdlH20, Chu03, Cod08, DA18b, DR01, DMR10, DIR13, DM97, DJM09, ERS00, FRV11, FMPP24, FR18, FBM17, GG19, Hig96, JT09, Jou05, KS10, Lee10, LS25, LFB00, LZW17, LZY09, Mar94, MM02a, MdR05, Mie89, PK21, QWX20, SYW22, SG09, SKO19, SLZ10, TJK18, VVV24, WM07, XY24, ZZLL21, dACR10, RMM12]. orthogonality [BDRZ04, JJ15]. orthogonalization [SS99, ZNK02]. Orthonormal [HAA21, CaAL96]. orthotropic [LLN25, RLMG24]. OSC [ZYLL20]. oscillating [CPR93, DYX09, DC09, FW08, Mar94, RCGM98, Sau00]. oscillation [ZYC22]. oscillations [KFR25, ebKMZ24]. oscillator [EK95, MH04, ST19]. oscillators [CD20a, DS21a, FW07, GMGF02, JZJ10, YW08]. oscillatory [ACLM22, ASCM02, Cai24, CSSZ20, CC04b, Che12a, CSLY19, CEW00, CMPP24, Den93, DSZ15b, DSZ15a, Eva94, EW97, EC07, Fra04a, Fra06, GMG02, GMG04, Has09, Has13, Has20, Ise02, IKM23, KM17, KW21, KCI03, Lam24, LWWX10, LW19b, Li23, LTT19, LW20b, LFS21, MK20, Maj17a, Mot17, PCL^{+}25, PGDB08, SWL20, TJ12, VV05, WWX13, XFG19]. Oseen [AKL08, ASV19, Cao10, Cod08, KLY05, ZMF25, vdVS08]. Osher [HS86, LDIW16]. Osher-type [LDIW16]. other [CSSZ25, DLM16, Ske99]. otherwise [BB10]. Out-of-core [MMBB07]. outer [MNSS22]. outflow [Nor97, Nor99]. outflow-explicit [MOU14]. output [CDGA17, FMS24]. output-based [CDGA17]. overdeterminations [HP14]. Overflow [FT96, FT06]. Overhauser [Arc06]. overlap [Pav00]. Overlapped [DS07c, DG22]. Overlapping [Coy12, TMS87, Wan05, ZHL03, ABCC18, BRVC09, GG22, Jun97, LOM98, MVBS25, RV04, RV05a, RGL16, SA12a, ZS21a, de 95a]. overrelaxation [HPY92, KP03b]. oversampling [LZZ25]. overset [QLL^{+}08]. Overview [WP99, FNT06, Hsi06, MLK06, Sv95, Wei95]. own [KV07].

P [CHM09, Kan04, THW19]. P2 [Kim12]. P SPARSLIB [SW95a]. package [HSW99, HS02, HK93, HVY91, SH02, dH95]. Padé [Ari87, Kid90a, Sad96, AEG12, Bel91, BW15, Bre96, Bre02b, Bre10, But09, CL06, Cat10, Cha25, Dar90, DMGVO05, DMGVPO09, GLLW14, Gil10, GVP93, GM94, KS91, Kid90a, Kid90b, Lor10, MdR05, MV20, OGV92b, Pre90, Pré95, Sad97, ZD25a, Zul25, van86a, vI87]. Padé-type [Ari87, Kid90a, Dar90, Kid90b, OGV92b, vI87]. PageRank [Bou25, LLVX20]. Pages [Ano21p, Ano21o, Ano21s, Ano21t, Ano21n, Ano21q, Ano21m, Ano21r, Ano22q, Ano22t, Ano22s, Ano22w, Ano22u, Ano22p, Ano22o, Ano22v, Ano22n, Ano22x, Ano22m, Ano22r, Ano23t, Ano23r, Ano23m, Ano23v, Ano23s, Ano23o, Ano23n, Ano23w, Ano23q, Ano23p, Ano23u, Ano23x, Ano24u, Ano24o, Ano24r, Ano24n, Ano24m, Ano24v, Ano24t, Ano24q, Ano24s, Ano24p, Ano25s, Ano25o, Ano25u, Ano25w, Ano25y, Ano25v, Ano25n, Ano25t, Ano25r, Ano25x, Ano25q, Ano25p, Ano26b]. Painlevé [AY15, Fil25]. pair [Cul95, FW07, RCR24, Tsi01]. pairs [ABD^{+}25, BL05, BS92, Hig97, KW20, SV00, TM15, Ver96a]. panels [NBP94]. Pantograph [AS00, BGT97, DBBH14, HAML21, Ise97, Liu97, NLS18, RU25, Wan17a, Yan18]. pantograph-type [Yan18]. papers [FJ97, vdHSW98]. Para [BRS16]. Para-orthogonal [BRS16]. parabolic [AP16, AP20, ADK94, AA22, AES15, ADFR18, AMC02, AAI^{+}93, APJ10, AED12, AN22, BKM13, BNS25, BBJ25, Ben17, BGG^{+}20, BC08b, BK21b, Bog12, Bog16, BJ03, BJ06, BCS06, CC25, CS94, CJ18, CX01, CL10, CWHF19, CGW20, CYWH22, CL01b, CJ23, CJ24, CJ25, Con20, Deh05, DK21, DSSC13, DCL23, DGS24, ELCWS98, EL94, FID18a, FHK05, FS23b, FM95, FJS99, GQ89, GD23a, Gol00, GDS^{+}15, GJlL23, GPHA22, GPHA25, Gon06, GO19, GO23, HHAA22, HZBM05, HKO12, HM01, HP14, HW04, HO05, HCY18, HJYL19, IVA93, JCSR03, KL98, KDAK13, KL21, KS07, KZ13, KD25, Le 12, LCS19, LXZ21, LGC25a, LAZ20, LM21, LYLL23, LZIZ23, LARGVR23, LW95, LOM98, LO95, MD22, MS03, MKBS26, MW24a, MR20, MVVA09a, MS90, Meu91, Mit24b, Moo95b, MD96, Muo23]. parabolic [NTHC21, Now96, OS08, OT02, OTK04, PK23, PAP17, Pło22, PJB04, PS19, RN25, RC18, RS20, SDK24, Saz22, Saz24, Sch87, Sch95b, Sea09, SRMDRL23, SB14, SS19, SYL^{+}20, SWW16, SHB^{+}25, SD09, SR24, ST25, Sti03, SLZ10, SW20b, Tau09, TDMT21, Tou25, TV91, VSG17, Ver96b, WKM04, WH19b, WZ22, WJM22, WL24, WWWW25, XWC25, YM24, YS22, YCWH23, Yu99, YLW21, ZLW20b, ZLX22, ZYQS23, ZCQ25, ZZH25, ZZZ19, ZZjQ24, Zho17, ZK00, Zuu95, Ant13]. parabolic-hyperbolic [GDS^{+}15, Sea09]. parabolised [BS08]. Parallel [AB09b, BOEP00, BDF94, BS92, BTDV10, BMT93, BT98, BF95, Bur93a, BT00, BC95, CK22, CSXL14, Con89, CP17, Der92, DF96, FLÖ^{+}97, FV99, FPS15, GM93, HH10b, IM02, Kie95, Kim95, LCS19, MW93, Mic95, Mie89, Nak05, Ore93, Pet92, PV93, RGS25, SK97, Sch91, SH10, SQ17, Ste05a, WSP04, YH07, ZC91, ZYH23, ZX09, vSC92, vC93, van93, vdHMdS99, AT93, AB98, AGJCP25, BN99, BC99, BMR^{+}17a, BRSD91, Ben96, BNH01, BT02, BTP96, BDFV95, BS91, BM18, BT93a, BT97c, BDP96, BS00b, CKB13, Chr96, CM07, CX08, CPR93, Cve02, DdSF07, DVV93, FL93, FL01a, Fre91, Gen10, GY94, GGM95, Hab08, HK25, HY02, HVY91, Jéz99, JL23a, KG90, LS99a, LGS21, LHX20, LYOI99, LN92, Noo95, OP04, Oos95, PSB91, PSW02]. parallel [QR03, Qui96, RZS21, SW95a, SWE05, SWJ09, SGY22, SK91, SCvdH92, Som93, Tsa91, Tsa92, VG04, VS94, VP91, WKP12, WWS^{+}93, WCGW95, XHJM21, YJJ^{+}24, Zar99, ZS21a, dH95, dDF^{+}94, de 92b, de 95a, dv95a, de 95b, vS93, vS96, vvdV97, vdSvdH95, Ria22, dH95]. parallel-in-time [JL23a]. parallel-iterated [CM07, CX08]. Parallelism [Che88, GX93, KP07, Doi91, Gea93, Pag25, PWY21, SD93]. parallelizable [Kim94]. parallelization [Jun97, RRMJ12]. ParalleloGAM [AB98]. Parameter [BBV05, BP92, GO19, GO21, KK06, KK20c, OS08, PKP19, SDK24, San18, SWJ09, SSW04, AMV03, AD04, BNS25, BO04, BR94, BCV25, CH21, Deh01, DT89, Eir99, EG88, FHM^{+}02, Fik23, GD23a, GOP06, HD22, Hin97, KMS19, KPRU20, LR24, LN08, MCS06, MOSW00, Nag22, Neu88, NCYC22, Nud24, SPS20, SGY22, SD22b, SD25, Sim10, XZ19, ZH20]. parameter-free [SD25]. Parameter-uniform [GO19, GO21, KK20c, BNS25, FHM^{+}02, GD23a, MOSW00]. parameterisation [CM04]. parameterization [Gar05, LG02]. parameters [AB09a, ABP95, BAP^{+}06, BD07, BBCS05, BCMV03, BS00b, Cat10, CJL13, He24, HK09, LLY24, MW93, Pul09, SOA^{+}25, WL09c, XLHL25, Zar17, ZL22, ZFX17]. Parametric [GFB99, LP25, Lo06, AGKK94, BP85, Che16, CGPT19, GZQS23, GS99b, LS07a, Ren14, SC11, YHT23]. parametrization [GLM09]. parametrized [HM09]. parasitism [CD20b, NH15]. Part [BV94, Mur99a, BO11, GGO16, Ioa89, Kru99, LPV24, MSA20, MAH22, XB14, AAB^{+}22, BDDV12, BRS^{+}18, CL24b, CY05, DGN12, GPHAM12, JLH13, LFS15, Mur99b, NRR06, SMTHE22b, SMTHE22a, Sid23]. Partial [BGHR12, BGH^{+}15, Luc05, Pet87, Rei85, AGZD22, AM99, AM00, AMCM08, AF89, ARS97, AKS21, AAEMY21, BMSZ21, BHJ05, BHJJ06, BKP09, BRRS15, BtTBV87, BO21, CGEV19, CSS87, Cha98, CGW20, CL01b, Dat99b, DS05, DMGVO05, DKL24, FID18a, FL93, FGP23, GAZD25, GVP93, GM17, GN86, GGO13, HP18, HZD21, HR96, HN22, HL89, JZK06, JMPY10, Koz94, KP15, KGZ24, LHHR94, Lay09, LW18a, LW19a, LW20a, LYK17, LYLL23, LB21, LS98, MT11, MZK05, MT06, Meu14, Mit24b, MMM19, MFAD23, Olv92, PM05, PB21, PAJ12, PT15, Pre90, PG02, Pul09, PSL18, QXG21, RA05, SS08a, Sar05, Sch87, SD13b, SGS20, SP22, SZQH23, SA20, uIVS13, SSKS21, SG17, TN16, Tan93, Tho85, Tia15, TS08, TDPU17, TV91, Tro93, VC10, Ver96b, WSS97, WdG92]. partial [YDWW17, YRV21b, ZW19a, dDF^{+}94, tV87, vSW90, van96]. partial-integro [CGEV19]. partially [ADNR21, GC25, GLMY17, GLM18, HP97, SZY21]. Particle [CKM15, IK24, MP85, AD21, CDD00, Cat10, CK06, GS15a, LW22, LL06, Nic86, RA03, SKR^{+}16, Tow16, ZML^{+}12]. Particular [PCA10, Pré95, ZLCH20, ZPT92]. Partition [CDD^{+}17, BSP04, Mit97, PPS05, SH10]. Partitioned [BD17, CH95b, NS21b, SZ22a, WPAZ24, DM11b, HR06, NS20, Ree03, RVM23, WGKS12, Wen98]. Partitioning [DS02, HW06, BSW93, FL93, GY94, RP01, TD09]. partitions [DS97b, LMS09, LWW20, PSP05, PSP04b, SST09, SST15]. partnership [HL97]. parts [RGÖS18, SD22a, Str98b, Zha21a]. passage [AACP20]. passing [BC99]. passive [CG13, CPY20]. passive/active [CPY20]. past [CRTU15, LMS08, PP00, TOCV02, Son00]. Patankar [BDM03, KM18, ÖT20, TÖR22]. Patankar-type [BDM03, ÖT20, TÖR22]. patch [LSGK15, VR01, ZY14]. patches [WM08]. path [GT19b, KP18, LDIW16, RP01, SZ99, Yam23]. path-conservative [LDIW16]. path-constrained [RP01]. path-following [SZ99]. path-independent [KP18]. pathology [AQ00]. paths [HHW18]. pathway [MDD18]. pathwise [Kha21]. Patrick [PSR04]. patterns [AGJCP25, Huc99, LC19, LR19, NMSF94]. pavements [ZBNC25]. PC [CM07]. PDAEs [LD02]. PDDEs [RFMV24]. PDE [AACP20, CCDJ20, CCD^{+}20, CC23b, Chr96, CD18, DKK94, MR92, MSS^{+}15, PR09, PZ20, Quy19, SSV89, Sch96, VG04, WYP12, YH18, Zha96, ZYH23]. PDE-constrained [ZYH23]. PDE/stiff [SSV89]. PDEs [AS20b, AAI^{+}93, AC16, AN22, Ber04, Ber05, BK21b, BV94, BCK22, BKW06, CY23, CMCGTR02, CKM15, Dia95, FS23b, FSX25, FS24, FHX22, GLPW09, GPPR12, GPHA16, GGR97, GFPG18, HT19, IVA93, JUAZ22, JR18, KP19, Lua17, Mir20, OXW25, Toc01, TTP24, Wal00b, WKM04, YM24, Zeg97, ZJ19b]. peak [Wal95]. peak/plateau [Wal95]. peakon [CLP15]. peakons [AS06, QR24]. Peaks [Cul95]. Pearson [DS23]. Peer [PWS05, SW12, WSP04, AHJ^{+}23, CMRV11, CMRdlT24, GLPW09, JKW12, KW20, Pag25, SWE05, SWJ09, Sch12, SAH24, SKW17, SW18, WEA12, WKP12]. Penalization [BCL15, PBC08]. penalized [KvyS15]. penalties [Imo00, YZ17]. Penalty [DESZ25, GK22, Hop23, SFZ21, AD20a, BKAG22, BB94, Buc17, BDD^{+}20, CH13, CSM07, DL21a, EKT19, GH21, HHC08, HD22, Hes00, HFL13, KSMMM16, LY01, LZH^{+}25, LS24c, Lot19, LDH^{+}24, TC19, WaZW21, YZ24a, ZL18a, ZGF^{+}25, Zho17, Zho18, ZKO^{+}21, ZCC11]. penalty-duality [DL21a]. penalty-projection [WaZW21]. penalty-type [LDH^{+}24]. pencils [AA94, AT13, CMT25]. penetrable [BBLT15]. penetrative [Tse00]. pentadiagonal [WJF25]. perception [Pro24]. Peregrine [CSS19]. Pereyra [MP98]. perfectly [CFLW22, DK14]. perforated [BBD24, CL08]. Performance [BDSG09, GY94, LS07b, MRV93, WT08, BMR^{+}17a, BSV21, BDP96, CB99, Dea11, Du11, Fre91, Gen10, HD04, JJ94, KHB22, MMBB07, Sal89, Win92, van98]. Performances [CCM17, CG03]. performed [BCE04]. peridynamics [CFL^{+}20]. Periodic [GM85, AC23, AB24, BGG^{+}20, BKP14, BDKM92, BO11, BMM97b, BMM97a, CDP17, CCP04, CjW18, CHNN20, Con99, DG10, DB95, DCN^{+}19, FJ09, GT19a, Gil88, HHAA22, HJ17, Jac88, KN19, Lem88, Li00b, LLM19, MN08, NAF24, Nes16, Ngu15, NT16, PPS10, Pat98, Sid14, Sid23, SvdHN86, SS09, VVD95, Vic92, WDU21, WS21, Zha14]. periodical [Ito22, WDZS21]. Periodically [YXB95]. Periodized [XB14]. permeability [CGN03]. permeable [CGJ16]. permittivity [QDA^{+}26]. permittivity-based [QDA^{+}26]. permutation [FV85]. permutation-perturbation [FV85]. Perron [Fer14]. Perry [SSS^{+}23]. Persistence [CMMR23]. person [CZ19]. personal [Han87]. perspective [MMP02b]. perspectives [AB09b, Car19]. Perturbation [LCZ21, RA05, Arn95, BF99, FHM^{+}02, FV85, FS23b, HP18, KHM^{+}14, KN19, LVfP14, LLVX20, Nak12, NT16, NT20, SdSC99, Sto96, TWH21, TM15, Vul92, Vul95, ZGL98, AI19]. perturbations [BT94]. perturbative [Cat10]. Perturbed [RSL89, Rei85, AY21, AL98, AL24, AN22, ÁMS14, BPN24, BNS25, BBJ25, BM00, BM01, BO04, BCGS24, BBRS97, Bog00, BFLR23, CHX26, CR23a, CJ23, CJ24, CJ25, CKK25, Cum95, DLZ21, DAMA23, FW07, FJ09, FL24b, GD23a, GD21, GMGF02, GOP06, GO19, GO21, GO23, HL99, HO24a, HO24b, KMS19, Kau95, Kau97, KS04, KK20c, KD25, Li00a, Li01a, LX24, LT01, LWC24, LMG02, LW95, MF99, MOSW00, MPtM16, Mus11, OS08, OQ15, PZMX16, Pap95, PMP23, RN25, RG22, RK08, RC18, RSR23, RTU15, SDK24, SK16, SWW17, SW21, SSR23, SKS23, TS24b, VN21, WC24b, XLHL25, YM24, YW08, YZ17, Zar17, ZCZ15, ZL18a, ZL22, ZLG15, ZX14]. Peterlin [WWZ25, Zha19b]. Petrov [DM09a, MS13, ABZ21, AD08, BBRS97, DGN12, DLZ21, FM21, GJV08, HZ21, KDT17, KNN03, LWY20, RP17, Sac93, mWyG00, WCL22, XY19, ZLG24]. Petrovskii [BFdO07]. PGEM [ABFV09]. Phase [JK17, BCFQ19, BCFQ21, Boz11, CMP20, CH15, CPY20, CLL25, ER07, GHK16, GOGF03, Has20, KSMMM16, KNP16, LL98, LY08, LMY18, LL20b, LC21, LY24a, LS24b, MZN21, MR06, MD20c, NS21a, Pat00, PHY19, QZH25, RS22, RZ15, Sch16a, SC25b, SY08, SS09, SWJ^{+}24, TCCW89, TKN11, Wag98, WM08, Yao24, YXX24, ZY19, ZCY20, ZLSZ22, ZH24]. Phase-field [JK17, CMP20, CPY20, LY24a, RS22, SC25b, SWJ^{+}24, ZCY20, ZH24]. phase-fitted [Pat00]. phase-lag [SS09]. phenomena [Pir09, Ven15, ZG92a, ZG92b]. Phenomenology [Bri85]. phenomenon [BCJW17, FRRJT10, Jun07, RSL89, BO11]. Phillips [Lub92]. photonic [LY24b, NS13]. Physical [HCC24, LH24, YK04b]. physics [AK09, Cat10, CDP^{+}25, MA09, PCA10, Sam94, SOA^{+}25, TT20]. physics-based [AK09]. physics-informed [CDP^{+}25, SOA^{+}25]. physiological [CKP15]. physiologically [BL15]. Picard [DS20, HN22, Lay08, Li23]. Picard-iterative [DS20]. PIDEs [MFAD23, TTP24]. Piece [LSY17]. Piece-wise [LSY17]. pieces [WWS^{+}93]. Piecewise [AB24, HBB25, PGS10, RF16, AR24, Bec18, CCQ^{+}23, CL14, DS17, GVSL96, JQYM23, KESYB23, KDS22, LZ17, MP20, Pic05, Pla08, Sti03, TL07, WL24, XZ19, ZLJ20, ZYJD25, ZWL11, CL24b]. piecewise-smooth [KDS22, CL24b]. piezoelectricity [MZ04]. PIM [dH95]. pine [ABdSG23]. pipe [DM09a, RR00]. pipelines [GKB^{+}22]. Pipes [hYK86, TT03]. Piskunov [BFdO07]. Pitaevskii [LCW20, WLCZ26]. pitching [OK98]. PiTSBiCG [Ria22]. pivoted [DS21c, HPS12]. pivoting [GP93]. placement [YK04b]. planar [BF01, EM05a, EM05b, Gab02, RAS99, SK01]. Planck [CjW18, LL24, PYD21, Van92, jWyG08, jWC22, YPD21]. Plane [PGM86, LHC09, MM02b, Mik97, MŠ99b, PA18, Wan05, Yua20]. plane-wave [PA18]. Planewave [NS13]. Planning [MZ87]. plant [BBT25]. Plasma [DRC85]. plasmon [NT20]. plastic [HS07]. plate [BLS94, BH12a, CM09, yDqGnJT09, Fai00, HLC01, KX03, SWY^{+}23, YTZZ18]. plateau [Wal95, Har09]. plateaus [Cul95]. platelets [SWFK13]. plates [CCS02, GLQ26, KK06, Lam13, NMSF94, QR03, Wei18]. platforms [BCFR25]. plotting [FH04]. plug [FMSV07]. plus [Haa97, LY01]. PML [CWM09, Kim14, TY98]. PMLs [BP12a]. POD [AD20c, AS20b, DA18b, FS19, SR24, YV17]. POD-based [AD20c]. Poincaré [DLM05, Hau88, Lev91a, WE99]. Point [Pot85, SM85, WKWC26, AAL^{+}26, AD99, AL05, Bac18, Bac19, BM12a, BHB23, Bec18, BCGS24, BCC16, Bic21, BB10, BPV25, CA21, Cai24, Cao03, Cao07, Cao09, Cao10, Car09b, Cas96, CW98, CJL13, CL07, CWC24, CC08, Cui04, DJ10, DMGVO05, DMGVPO09, DMPSC16, EHM01, Elm02, GD23a, Gil10, GVP93, GH21, GH20, HGM^{+}21, HK93, HM09, HT20, Ius97, IS22, Jac93, KHLV22, Kat89, KS01, Kor95, LMO24, LZJ21, LW21a, LS21, Lyn92, MV20, MRÁSY24, NCYC22, OG08, OGV92b, OGV92a, PPT02, RV05b, RGMO19, SJ18, Sim04, Ske99, SSA24, Ste97, VA21, Wan07a, Wan11, Wel10b, WG11, WYP12, XZZL15, YW19, YP18a, YP18b, YZH19b, ZGF^{+}25, ZZW97, Zha07, ZZ25, Zup03, dSFDG20]. point-based [BCGS24]. point-like [CJL13]. Point-wise [WKWC26, LW21a]. Points [Eis86, AC23, Bal00, BS00a, BE99, BS20b, CN15, FG13, Gil91, GÖ20, HGM^{+}21, HM09, HS95, LZZ18, Maj17a, DE18, Nov03, Oji88, OL18, PSP04b, SK10, SL08, Wai98, WT08, Xu16, YT03, YDWW17, ZSJ04]. Pointwise [GZHQ23, Len00, Wei09, ZMC13, BSZ15, CL26, Fra14, HSF25, TM04, WC24b]. poisedness [WW05]. Poiseuille [MM07]. Poisson [LL24, AGK24, AO91, BSGU94, BS14a, CHPV09, EK06, HLY04, HLZ06, KRBK16, LX08, LHWF08, LW07, Maj20, NAD25, Phi91, PYD21, RV04, RT20, She00, Sob24, Tem23, WKWC26, YPD21, Zhe07]. Poisson-type [AO91, She00]. Poisson/symplectic [LX08]. Polar [SST09, Kwa09, LLT20a, She00]. polarization [HJX^{+}19, YWH20]. Polarizations [SR88b]. pole [Jad94]. poles [AQJ18, SMJ12, SL01a, Tom24]. policy [RF16]. Pollaczek [MN20]. Pollaczek-type [MN20]. pollution [AVMVMV09, Bou16, DS03, TM04]. poly [KB21, KESYB23]. Poly-Sinc [KESYB23]. poly-Sinc-based [KB21]. Pólya [AHGM21]. polycrystalline [HPW21]. polydisperse [MM14]. polyfractonomials [ABZ21]. polygamma [Meh22]. polygon [AM10b, CKK10]. polygonal [ANN19, Hin97, KDD23, KSJY25, LHWF08, PR09, PS19, SW24, WTF25, Xu21, XYZ24, ZGR23]. polygons [CF18, GFPG18, LD10, VMS07]. polyhedral [RA09, SV24, XYZ24, Yos00]. polymer [ABP95, Lam24]. polymeric [BCFR25]. Polynomial [BSZ99, BDV17, MN20, Pul09, Pul12, Rus95, SW98, AA04, ADSS17, AC23, BM04c, Boy06, Boy07, BS09, BO11, CCQ^{+}23, CP94, CJL13, Che12b, CS17, DDGN23, DJJ^{+}15, FZA25, GVSL96, GJ17, Han06, HL08, Iva07, JBLC11, KLR25, Kid90b, KSJY25, KYI17, LZZ18, Mar99b, MN03, MQ03, Mie89, NLLG20, Nov03, OT21, OT25, PH91, Pla08, QW25, SD13a, SZ12, SZ17, VVV24, WWL21, Win01, Win04, Wu03, XC20, ZWK15, de 92a]. polynomial-based [FZA25]. Polynomialization [Boy07]. polynomials [AMT17, AyLqW18, AHGM21, Alt85, ADG11, BR01, Ber05, Bor10, BWY17, BWS21, Boy15, BRS16, Bre91, BDRZ04, BDRZ19, BDMGVO05, CMR12, CKM10, DGV00, DR01, DMR10, DIR13, DRS19, DB08, DM97, DJM09, FRV11, FMPP24, FR18, GSW09, HS19a, HAA21, HAR21, JCN94, JJ15, JT09, Jou05, KP92, KHYY21, KP19, KGZ24, Kür23, LMV25, Lee10, LY10, Mar94, MM02a, MG97, MPTV25, NN20, Nap16, Pat98, Per03, Pis22, Rah25, Sim94a, SMJ24, TJK18, WHW21, dACR10]. polynomials/functions [GSW09]. Polytopes [FWW^{+}21]. polyvalent [CMR94]. population [AM95a, AMT13, AL20, ALM04, Aya09, DFLM19, GeO24, LS16, MS25, QAE^{+}09, SH97, SMW21]. poro [BFQ22, YR22]. poro-elastic [BFQ22]. poroelasticity [GM18, GHH20, GH25, GCZZ23, HGF24, LR24]. porosity [BSZ22, CHLA21, SZ22a]. porous [AKT97, BM04a, BM04b, BBL02, CCD^{+}20, CHNN20, CCK03, CML05, Cho13, CD13, DCSB25, DD19, DCN^{+}19, EH07a, EWW99, FBS09, GJR03, HJ09, HHL23, HCW16, IMC22, JL25, KT05, KMR09, MK14, MLJ19, MCM12, NS21a, SR88a, SY08, ZYSZ14, ZFW20, ZGR23]. port [BDF^{+}25, MM25]. port-Hamiltonian [BDF^{+}25, MM25]. portable [PSB91]. portfolio [HW22]. posed [BHL^{+}21, BBBK22, CRS05, EG88, GNNR19, HDY21, KMH21, Kli15, Lie01, LHT20, Luc95, MRH14, Sam94, TWH21, TWD23, WQS^{+}24, XXQ17, ZD21, ZGF^{+}25]. posedness [Geb24, MPPR22, MN24, SSZ16]. position [CM06]. Positive [BW95, Cao07, DBCBPP10, AHGM21, Ara99, BRBM08, BBKS07, Cao01, Che12b, CL88, DL01, DZMB21, GP93, GS25a, GP04, GP17, KCB02, LZJ21, LWZ22, Liu02, Lu98b, MAD23, QZM25, SH21a, Zha21a]. Positivity [BK12, CHM22, Hor98, HWZ22, Sca22, Ber04, Ber05, GM87, HKO12, Hor05, KP03a, ÖT20, PGYF20, QZH25, SRMDRL23, TÖR22, WLG22, WDL23, WC24a, WC25, YTC24, ZWN23]. Positivity-preserving [HWZ22, PGYF20, QZH25, TÖR22, WLG22, WDL23, WC24a, WC25, YTC24, ZWN23]. Post [BSZ15, CCM02, YWH20]. Post-processing [BSZ15, YWH20]. Post-projected [CCM02]. Posteriori [Rei85, AS05, AB10a, AOW94, AD23, AR18, ASV19, AP08, AM16a, BLS94, Bac14, Bac16, Bac17a, Bac21b, Ban97, BS14b, BCS17, BM13, BBG14, BRVC09, BC04a, BHR05, Bür12, CHZ21, CCOVF22, CHX26, CMP03, CCS17a, CSW19, CZHX19, CDW23, CKK25, CGK25, DDP12, EG88, EFLFP09, FH10, GHH09, GHT05, Hop23, ID19, Joh01, JM94, KDT17, Kim07, Kim12, Kim21, Kwe00, LCHR03, LS12, LY03, MWC21, MW24a, Moo04, Neu88, Roz05, SS00, SZ09, SS19, Sha21, VNC21, WH13, WG19, XL11, XYHM20, Xu21, YSBL14, YT25, ZBD24, ZCZ15, CGG25, DLZ21]. postprocess [dFN00]. Postprocessing [CJX11, GANT02, MWC21, YS09, ZLY23]. Potential [Man96, BP85, CHPV09, CJL13, DD97, DS97b, FS15, FJP17, FL23, GJLL20, JLZ20, KK02, LWD^{+}09, Li23, LN08, ST14b, SZW19, VZ93, WL24, dRT99]. Potentiality [Set24, BT97c]. potentials [AES15, AEMX17, CCP04, EP15, FXY22, FvdMS17, GT19a, Gia12, Kam00, LLM19, MZZ17, RV15, RS00, RU15, TS24a]. Poussin [ORT24, OT25]. Powell [BFGP08, LMS09, MB10, PT09, SST09, SST15, Spe12]. Power [DRC85, BWM21, BL06, CFV10, CL88, Lon88, LYA^{+}19, Mat91, NYPW21, ST09a, SPYS24, Wen10b, Win04, YLY19]. power-type [LYA^{+}19, YLY19]. powers [Vab21]. Practical [Iga85, CK06]. Prandtl [CMP06, FHM^{+}02, PR89, RLSS06]. preassigned [SMJ24]. precision [Var92, vdHMdS99]. Preconditioned [WWS07, AGJ12, AG05b, Bad20, CR05, DW00, HVY91, JWZ21, JL23b, lLH25, WSC21, YG95, YH18, Zen21]. Preconditioner [BTMT08, Aro96, BvG19, BLY17, BG11b, BSV21, BCSH16, DO95, DO98, Dor91, Fuj99, GNX19, GS92, HY02, HvdHV10, HS21b, KKP07, KM16, PSWZ21, PM03, SK22, SWX00, SH21a, Ste97, Wan05, WCXL09]. Preconditioners [Elm02, lLNW21, AK09, BHB23, BT99, BS10, Buc17, Cai15, CRS05, Cao07, Cao09, Cao10, CES91, CH01, DdSF07, DJ10, EVO04, EVO06, GGM07, GCP91, GGR97, Haa97, Kos02, MB10, MD96, NLLG20, NR14, Osw97, PS09, Sim04, ZNK02, Zha21a, de 95a]. Preconditioning [AES13, KK09b, NRZR12, OKS10, San03, ASZ15, Bai96, BL21, BT93a, CD18, DB97, DDS89, DSSC13, GNNR19, Gen10, GKMS09, Guo96, HLMP09, Huc99, JP19, KJ99, LD21, LLT20b, LN24, ML91, Mun00, Not99, QDA^{+}26, SW98, SH02, SvV22, Tur93, WZ02, Zha00]. preconditionings [BD85]. predator [BGG^{+}21, CL02a, KS25, SZ22b, Wan09, ZSS23]. predator-prey [ZSS23]. predicted [KDS22]. Prediction [Bre85, KM21, Cat10, CDP^{+}25, YH07, ZHJ14]. Prediction-correction [KM21, ZHJ14]. Predictor [BK21b, THW19, ASC03, Bur91, CXZ15, IM02, JL91, LLL12, MS24, MPPR22, Wai98, Zla85b, de 95b, vC93, Bur93b]. predictor-corrector [Bur91, JL91, LLL12, MS24, MPPR22, Zla85b, vC93, Bur93b]. predictors [HR06]. Preface [AMT05, AGH^{+}10, ADG^{+}16, Ano00f, Ano00g, Ano10, AM05, AM08, BGS14, BKV08, BC94, BRZW10, BR09, BGHR12, But94, CPS02, CP03a, DFJ^{+}17, DFL^{+}25, GHR06, GCJ^{+}12, GKS04, HHV03, HKNV16, Jac06, JRS14, LST07, LDF^{+}20, PS04, PS05, SW95c, SG02, SSW02, SVSW05, SVAW09, Spi03, Spi04, VS04, VHAW12, tRDvdV91a, tRDvdV91b, BRW21]. preferential [QAE^{+}09]. prefractal [BBV13]. preliminary [QLL^{+}08]. prescribed [Liu21, Tol04]. prescribed-order [Tol04]. presence [BBW19, Maj17a, Neu88]. present [Son00]. Presentation [KW12]. presented [Wen10a]. Preservation [HXW15, Rei99, BMQW16, Ber04, DL22b, Mat08, WS22]. preserve [Rob10]. preserves [SRMDRL23]. Preserving [Ber05, AZA22, ASGGRGW25, ABK12, AZ23, AB14, AGK24, BUL23, BK12, BEH24, CHM22, CC23a, CW25, CWY20, DL20, DW21, DLP06, DF92, FCW20, FXCW21, FCW21, GZQS23, GHS25, Gje07, GJlL23, GGT24, HKO12, HS20, HPH20, HCC24, HMT03a, HMT03b, HGZW21, HL21, HWZ22, Hou23, HL24, HZZ25, HCGW22, Hua09, JK14, JWG20, JQSC22, KGR08, KL21, KT05, KMG09, KP01, LW19b, Li22, LW22, LH24, LZ24, LLZL24, LFS21, LYMD24, MB20, MDP10, MDRR11, MMD20, MPV24, MAF20, MSA20, MAH22, NBNTGV11, ÖT20, PGYF20, PA05, QZH25, SA21, Sca22, SS21, SNW22, SW24, TÖR22, Udd20, VL08, WH18, WLG22, WDL23, WC24a, WC25, WLCZ26, WKWC26, XCHW22, YTC24, YX25, YYZ23, Yao24, YF24, ZH21, ZM17, ZYQS21, ZWN23, ZYQS23, ZWG25, ZSY20]. pressure [AMN24, BC01, CKS05, GP00, GRGJ02, GS18, HL08, HGF24, KS02, LA11, LGS21, LFQH21, LY16, PQ25, RÁM23, RBC02, XZ22, Zha19b]. pressure-correction [LFQH21]. pressure-velocity [LY16]. preventive [KN08]. prey [BGG^{+}21, CL02a, KS25, SZ22b, Wan09, ZSS23]. prey-predator [BGG^{+}21, CL02a, SZ22b]. price [ZJH^{+}23]. prices [LL15]. Pricing [SB19, ZO14, ALY03, AD99, BAD13, BNV06, BDOG19, CGEV19, CXZ15, CC23b, CPOGO17, FV01, GGO13, GK22, KKT16, MHL18, RP17, RG21, ST11, ST14a, TGB08, ZW09, ZGO12, ZJH^{+}23, ZD25b]. Primal [AP16, AP20, HT20, dRT99, BSZ99, CW21, Car09b, He24, HMW05, MDT05, Par04, SZ99, TC22, WL25, ZZX19a]. Primal-dual [HT20, CW21, Car09b, He24, HMW05, WL25, ZZX19a]. primitive [AGLRS23, GGRN17, LLL08, RG05, RMM12]. Prince [EL97]. principle [CL01a, FHK05, GZQS23, GJlL23, HPH20, HZZ25, SYY20, SG16, SW24, Tad86, TZ21, Tol03, Vej10, YYZ23, Yao24, ZYQS21, ZYQS23, ZSY20]. principle-preserving [HZZ25]. principles [BKK25, PZ20, WY22]. printed [BLW02]. priori [AA25, Ars20, Cha17, CKK25, Dek17, HMW05, JR02, KPY15, KK22a, LMWZ10, MS03, TM05, MN24]. prism [LMQZ18]. prismatic [KV20]. probabilistic [dlC23]. probability [BD07, CCDJ20, DNW18, Shy91b, Yan21b, ZCGS21]. Problem [BH85, YK04b, AS21, AH09, AMRR18, AGLRS23, AMN24, AA25, AI19, AB15, AHAS21, AHS03, Ant13, AC18, AL98, AKL08, ASV19, ACP23, AMV03, AP08, AS00, AFLG^{+}12, BS21, BF01, BBD18, BM13, BGG04, BCFQ19, BM00, BRS05, BRTB19, Ben02, BGM^{+}09, BBCS05, BLY16, Bis11, BR94, DSM22, BBBN21, BJ00, Boz11, BS97b, BP85, BFLR23, BSZ15, BSP04, BP95, CCOVF22, CFRA08, CL08, CC25, CD23, CGG25, Cau08, CGRT18, CGN03, CHNN20, CKP15, CR23a, CMP03, CX01, CH07, CJX11, CZ19, CWZ23, CL26, CL18, CH04, Cop03, CF05, CF13c, CF14, CA15, CA16, CG14, DAH24, DY17, DT15, Du11, yDqGnJT09, DZW24, EM05b, EGV25, EW08, FHM^{+}02, FE93, FS23b, FL15, FBS09, FH10, Fer93, Fer96, FBM17, FL24b, FFQ09, GS15a, GP23, GLV06]. problem [GMM09, GD21, GMS12, GLLY25, GS18, GOP06, GO21, GO23, GL17, GGG16, Guo00, HS21a, HS07, HLZ14, Han87, Har09, HILK13, HDY21, HJ06, HK85, HX11, HW22, HLY22, HZCZ23, HZZ24, HLC01, HT20, IV16, ILS19, Ius97, JP08a, Jes93, JS24, JLZ20, JL24, JK20, KS91, KMH21, KL23a, KPR06, Kim94, KS09a, KNP16, KKN^{+}17, KS01, KS04, KCW16, KRBK16, KR25, KW93, Kwe01, LSV22, Le 12, LL98, Lei02, LRV25, LPZ00, LMA18, LW18a, LLHC18, LZJ21, LB23, LHC23, LTDY24, LHY24, LGC25b, LD21, LM21, LO03, LWZ22, LLN25, LK07, LLL12, LYY15, LW18b, LC20, LS20, LWC24, LD22, LDH^{+}24, LP01, Luc95, MM18, MSZ^{+}24, MG00, Man97, MM22, Map05, MNR14, MN03, MPHFP23, MP20, Mat05, MM02b, MS08b, MZM20, Men23, MOSW00, Mol95, MCM12, Muo23, NS03, NS12, NLHT25, NY13]. problem [Njå88, OQ15, Odi24, OP04, OGS20, OMSS25, PS00, Pic05, Por17, PS21, QXQ22, QC12, QL15, Ram96, RSL89, RG22, RLMG24, RTU15, Rou20a, RBT15, SSZ16, Saz24, SGS00, SR88a, Sel14, SSC23, Sha21, SC08, SSR23, SJ18, Shi20, SM20, SHB^{+}25, SOA^{+}25, SK01, Sto96, ST08, TBRBM20, Tia15, TWH21, TT03, Tou25, TM15, TC19, Vab22, VSG17, VR01, Wal19, WW05, WN12, WL16, WQ17, WZZ21, WCL22, WPN^{+}24, WC24b, WQS^{+}24, WW14, WT17, Wu09, WYY20, WWH24, XYHM20, XLZ23, XLHL25, YLL09, YL13, YJZ18, YD22, Yao24, YZ19, YZH19b, YXX19, ZTZ15, ZH20, Zar17, ZG92a, ZCZ15, ZZ19a, ZL22, ZY23, ZYZJ24, Zhe07, Zhe19, Zho17, Zho18, Zou11, dSFDG20, vBvdZdB08]. Problems [De 88, Gus88, Per88, SM85, AY21, ALMM01, ACP24, AP16, AP20, ANN19, AL09, AB97, AGJ12, AyLqW18, AES15, AEMX17, AA20, ARY23, AMC02, AMP03, AMCR17, AB09a, AD23, AMV17, AMR14, AX19, AT93, ABI22, ASZ18, AD99, AG05b, AC08, APJ10, AGJCP25, Ars20, Aso21, Aug89, ABG^{+}15, ABRW18, ÁMS14, AL05, BAA22, Bac17a, Bac17b, Bac18, Bac19, BTBR19, Bac21b, BC12, BKK25, BHL^{+}21, BL21, Bai25a, BPN24, BS14b, BLS^{+}17, BY00, Bar09, BNS25, BvG19, Baz03, BFQ22, Bec18, BBV05, BB25, BBBK22, BNH01, Ben17, BG24, BC08a, BGG12, BC04a, BDF89, BW96a, BCT16, BW97, BC08b, Bic16, BCC16, Bic21, BGH08, BMR17b, BTMT08, BKP14, BS97a, Bla00, BBRS97, BBD08, Bog00, Bog12, BMMZ06, BDFF23, Bos09, BBLT15, BP90, BJM01, BT95, BT98]. problems [BM09, BJ01, BJ03, BJ06, BCV21, Bur93a, BT00, Bus06, BC89c, BS00b, BR05, BD17, Cah89, CA21, CHZ21, CRS05, CSSZ20, CSSZ25, Cao03, Cao07, Cao09, Cao10, CW21, CP97, CL85, Cas96, CW98, CM02, Cas06, CS94, CP05b, CES91, CM00, CP07, CJ18, Che96, CR05, CK98, CC04b, CZ04, CYM09, CL10, CM14, CHS19, CLY19, CZHX19, CWHF19, CWP21, CYWH22, CL24b, CSL^{+}24, CML05, CPC25, Che12b, Chn17, CL20, CH13, CR23b, CJLS98, CKK25, CG16, Con99, CGPT19, CRHTV25, CN17, Cui04, DC21, DZ12a, DPPR16, DDHS97, DZZZ24, DDP12, DS21c, DS24, DA18b, DDF24, DW15, DS07c, DMPSC16, DCN^{+}19, DSW96, DS15, DP90, DY03, DLZ21, DL06, EHM01, Elm02, ECB07, EG88, EGL09, EH06, Ewi91, EL94, EGH01, FMS18, FW08, FHK05, FL05, FD97, FHM^{+}02, FL20]. problems [FMPP24, FdSB02, For25, FMP04, Fra04a, FG09, FLR08, Fra14, Fra16, Fun94, GAML04, GMZ11, GMG02, GMG04, GG22, GIS23, GNNR19, GZW22, GD23a, Gen10, GAW09, GS89, GK19, GRLL01, GPMPR03, GPHA22, GPHAPPR23, GPHA25, GML00, Gon06, GO19, GH02, GCZZ23, yGpY09, GHF00, Hab08, HGM^{+}21, HJR22, HHYD20, HJS97, HKO12, Har98, Har10, HOS99, HM01, HSS04, HA21, HV22, HD22, HCJW24, HO24a, HO24b, HS22, HW06, HMdV03, HK93, HM00, HS19a, Hey20b, Hig93a, HL02a, Hin95, HO05, HOS11, HL19, HLW25, HY01, HCX03, Hua19, Hua21, HM22, HBJ09, HMW05, HLIS16, HK09, HSY18, IO18, Iga85, Imo00, IT16, JJ94, JP08b, JS09, Jeo09, JK14, JCL18, JL23a, JR02, JCSR03, JNPC03, JM94, JCJP21, JV09, KDT17, KHLV22, Kat89, KM19, KOR18, KG90]. problems [Khe91, Kim07, KwS19, Kim19, KSK^{+}25, Kli15, KL09, KKW00, KCC04, KFR25, KP03b, KM11, KKR25, KK02, KK20c, KDS22, Kur98, KAS17, KDKW20, LRS23, LHH96, LWT07, LHC09, LRS09, LV12, LKV01, Lau17a, Lee94, Lei99, LFB00, Li00a, Li01a, LY08, LMWZ10, LW17, LLHC17, LLJY20, LXZ21, LH21, LWW22, LDC24, LLT07, LAZ20, LC99, LMSW17, LS07a, LX09, LCH20, LN24, LZH^{+}25, LS99b, LX24, LT01, LY03, LCJQ12, LJYS20, LCZ21, LLW22, LZIZ23, LGZH24, LT05, LHT20, LWLL24, LYC24, LMP99, LT93, LMG02, Lot19, LAH09, LfX15, LW95, LOM98, lLX22, MH16a, MOS02, MK14, MH14, MA09, MS03, ML16, MKBS26, Mar94, MS00, MMRV20, Mat86, MS90, Mat09, Meh08, MR01, Meu91, MV18, Min87, MRH14, MK19, Nap16, NNJ23, NTHC21, Ngu15, NH24, NZY21, NB01, NFAE03, Now96, OS08, OB20]. problems [OH20, OTK04, OEAS21, Pag25, PL25, PZMX16, Pap95, PNA21, PAP17, Pea16, PSWZ21, Per99, PA91, PR22, PV93, PPC00, PJB04, Pou00, PMP23, Pul12, QWW24, Que21, RP01, RN25, Ram94, RNG22, RK08, RS20, RV04, RTH23, Rha99, Ric08, RV09, RV05b, Roo20, Ros93, RGA19, RTA19, RVM23, RU07, SST04, Sac93, SDK24, SAG86, Sam94, SPS20, Saz22, Sch93, Sch09, SKAW12, Sch16b, SWE05, Sch95a, SNOK21, Set24, SS19, SWW17, Sha98, SLJ11, pSLqJcY16, SHLY19, SYW22, SWW16, SJ11, wSJP15, SWB20, SWB21, Sim04, ST20, SD09, SKW17, Sol15, SZL18, SD24b, SW85, Sub04, SvdVvD06, SL01b, SW05, SW20b, SC22, TL07, TLQ21, TQY24, THW19, TLGC22, TDC13, TLV92, TK05, TWD23, Top21, Tsy98, UHUL21, UHLZ24, Vab21, VO00b, VV02, VVD95, VS94]. problems [Vej10, VBH96, VBVA22, VSeYD02, Vul92, Vul95, VN21, Wan07a, Wan07b, WCXL09, WL09c, Wan11, WZ14, Wan17b, WCSQ18, WZ19, WR20, jWS20, WM22, WY22, WJM22, Wan23, WH23, WLZ25, War92, WWLS08, WG18, WYYL19, WWF20, WL25, XFLC00, XC85, XC20, XXQ17, XZT21, Xu21, XGHM22, YGY15, YXF25, YSBL14, YZ17, YZ24b, YZZ25, YLW20a, YWW23, YCWH23, YÇ16, YBW20, Yos00, Yu99, YLW21, YY24, Yua93, YD07, YLH20, YLW20b, ZDM18, Zak19, ZP24, Zen21, ZLY23, ZG92b, ZMC13, ZWJ18, ZL18a, ZLHW19, ZBY19, ZLW20b, ZG20, ZR21, ZD21, ZLX22, ZLW22, ZWN23, ZZH25, ZGF^{+}25, ZBNC25, ZZZ19, ZYH23, ZPT92, ZLS20, ZW19b, ZWL11, ZX14, ZSQ21, ZX22, ZS18, Zla85a, ZCC11, dlC23, FG96]. Procedure [Pot85, BFA93, BS94b, CKB12, DSZ15b, DSZ15a, HP15, Kim94, LR87, Mon21, Par04, SW85, TQA^{+}25, TK05, ZQLK11]. procedures [BSZ15, CIJ17, FS05, GGO12, GGO16, LP97, Nor97, Rha99]. Proceedings [Ano02g, FG96]. Procesi [GZZ19]. Process [Bie87, BF17, BNKR20, BRZ10, BKP15, DMS23, GT19b, KL23b, KL87, LYY15, OR18, PGC01, RW87, RSD^{+}06, Sob24, Zha96, ZGO12]. processed [BCET24]. processes [ABD16, BLM17b, BZ91, CKB13, KSSS16, QPT23, RY13, TS23, ZO14]. processing [Alt85, Bla01, BCR01, BSZ15, CF13a, Duf90, EEJB22, JP17, LN92, RU21, Sae14, Söd06, YWH20]. processor [VP91, Win92]. Procrustes [FBM17]. Product [TMD92, AAD14, BS05, EHNR24, EJRR23, Fuj99, GCHR06, GVSL96, GKA17, EEJB22, MN03, MC17, Naj20, OR20, RU21, RG02, VVV24, ZLX19]. production [BDM03, GGRBRG22, KN08]. products [DKL24, Gil91, Goo90, Mar09, SBBC21, WGB99]. Professor [CHM09, EST15]. profile [BN03, HHL23]. profiles [CDV00]. program [HSW99, HS02]. programming [BSZ99, Car09b, EB12, GS99b, GT02, GL17, HD23, Hua20, Hua21, JS25, KS91, KS10, LLS^{+}96, LDH^{+}24, LRT99, MM16, OXW25, Ren99, SZ99, VB99, YJJ^{+}24, Zha07, dRT99, ZNK02]. programs [SAMSB20a]. progression [AFIS24]. Progressive [FL24a, EW97, IKS25, MRÁSY24]. Projected [TMM15, CCM02, KP03b, SS12, Yu08]. Projection [MKN23, SI20, dAF17, AMR14, AL17, BJ02, BLW07, BS91, Cai15, CM97, CWX21, CR23b, ÇK13, Die20, GG22, GV04, Guo15, HM21, HL08, IAH25, JMU^{+}25, Kam25, Kan89, KL09, KKLD21, KR18, LA11, LGS21, LWLW24, Liu21, MM18, MKDC25, Mat09, Min04, PP24, SEGV02, SLMD21, SLW24, TLQ21, Tob14, WaZW21, WaZ24, YZH19a, YÇ16, YLS^{+}09, Zha14, ZHJ14, Zha19b]. projection-based [ÇK13, JMU^{+}25, Kam25, LWLW24, PP24, YÇ16]. projections [CH87, dOF20]. Projectors [DV20]. prolate [BDSG09, FRRJT10]. prolongations [Ste08]. prominent [MAD23]. prone [KN08]. proof [SNOK21]. propagating [GOLJ25, RAS99, SM13]. Propagation [AL87, DE06, Gus88, AD15, AD01a, BR25, BvG19, Boh03, BCJP18, CH89, Den07, DG22, FMS18, FJ09, GKB^{+}22, GD22, GL93, HB20, HJX^{+}19, IJ14, KS89, KKE16, LFL14, Spi96, Ven15, Vic87b, Vic92, Wag85, Wee01, YWH20, YR22, ZSG^{+}20]. propagations [Son91]. propagator [NW09]. proper [DA18b, GG19, LZY09, SG09, SLZ10, WM07, ZZLL21]. Properties [GVSL96, ARGA00, Ale11, BN99, BF92a, BZ92, BGT97, Bre85, BT93b, BPV25, BS18, Bur85, Chi12, DIR13, EGL09, Gia12, Gil10, GS15b, GS05, GPíP03, HLMKZ06, IM98, IMMS20, JVZ96, JVZ97, LW25, LIPT18, LBCN00, MD19a, MCM12, Phi87, RZ04, Riv09, Rog19, SKS^{+}25b, Sch93, Ske89b, SG17, TJK18, WW99, XFLC00, ZD21]. Property [ASGGRGW25, BY22, CR19, CMS04, LH24, Nak24, THW19, WS22, XWZ21]. Property-preserving [ASGGRGW25, LH24]. proportional [AD99, BBBN21, MPDB24, SA20, SAMSB20b]. Proposed [Mac86]. Proteus [Aya09]. Prothero [Ran15, Ran16]. proximal [GZW22, GH20, Hua21, IS22, LZH^{+}25, LHT20, LWLL24, MBS23, SZY21, WH23, YJJ^{+}24, YP18a, YP18b, ZY23]. proximal-indefinite [MBS23]. proximity [Zha07]. PRP [Cui24, HZC22, LWLW24, MK19, YLW20b]. PRP-DY [LWLW24]. PRP-FR [MK19]. PSBLAS [DdSF07]. PSBLAS-based [DdSF07]. PSBTS [ZWFX22]. Pseudo [RZ04, ST86, Ant23, Boy15, Con01, CX08, DL22b, ER18, HS17, HS09a, Hoa15, HH10b, JZS20, LZ18, Li22, lLXhLZ21, MR20, MKS12, NTHC21, OGS20, PBC08, QR24, YF24, ZLW20a, ZCQ25]. pseudo-parabolic [MR20, NTHC21]. pseudo-peakons [QR24]. pseudo-penalization [PBC08]. pseudo-RBFs [Boy15]. Pseudo-Schur [RZ04]. pseudo-spectral [DL22b, ER18, HS17, JZS20, LZ18, Li22, MKS12, OGS20, YF24, ZLW20a]. pseudo-symplectic [Ant23]. Pseudo-Time [ST86]. pseudo-transient [HH10b, lLXhLZ21]. pseudocompressibility [She96]. pseudomonotone [DP21]. pseudoparabolic [FL23]. pseudorandom [GPíP03]. Pseudospectral [BMV06, HT00, AT15, Aug89, Boy91b, CD00, DDS89, DHL00, EES05, Elg17, Fun90, GT93a, JR00, KS09a, LLT20a, LLT20b, Mar93, MZK05, Mit22, Mit24b, TX18, jWyG08, jWjJ17, jWC22, WWM22, YH00, YLX21, ZM25, ZL11b, ZK00, PRS23]. pseudostress [CGS19, CLY19, GIS23, HL19]. pseudostress-assisted [GIS23]. pseudostress-based [CGS19]. pseudostress-velocity [CLY19, HL19]. Publisher [Ano04m]. pulse [JCY25, TER03, TCCW89, XC85]. pulse-spectrum [TCCW89, XC85]. pump [RVD00]. Pure [BB15, Usm97]. purpose [AA87]. pushbelt [SUP^{+}12]. put [ALY03]. PVM [Mic95].

q [MM02a]. QED [Sus10]. QHSS [CLLM21]. QLM [IB24]. QLM-based [IB24]. QMR [Cao98b, FN95]. QN [Ree03]. QR [GS08, KP19]. Quadratic [FS88b, Han06, HL02a, IO18, LMS09, LT05, RS14, SE93, AHR12, Bic16, BMT25, Bra00, BDRZ19, CCQ^{+}23, CL07, Cos25, DW00, Eir95, EB12, GT02, GGRBRG22, Hey19, Hey20a, HXW15, HMW05, JQYM23, Kür23, LZCF21, LP00, LL02, LfX15, LB21, MMP09, RL86, SMB23, SST09, SST15]. Quadratic/linear [IO18]. quadratically [YP18a, YP18b]. Quadrature [KM17, MO01, SMJ24, AAH21, ABD^{+}25, ADH00, BBCR22, Bas21, Bre10, BDMGVO05, BGVHN10, CMP15, CCBGV08, Car23, CL14, CHH15, Chr01, CEW00, DS21a, DGV00, DBCBPP10, Der92, DIR13, DRS19, DDRS24, EC07, Fid17, HA16, HLL09, Ioa89, KCI03, Kza92, Kza99, LWWX10, LYZ25, LFS15, Lub92, Maj14, MMP20, NLS20, OR18, PL20, PRSS24, Pel15, RdAP96, RS21, SST12, SHA12, Sid14, Sid23, SS16, SV24, SND19, The17, Tom24, VV05, WT08, XGM08, ZMY21]. quadratures [DMGVPO09, Elg17, FL24a, MD06, MSP10, Mil17, Wel10a, Wel10b]. Quadrics [ZY25]. quadrilateral [Bof06, MCS06, SN04, TC22, TW00, Yi12]. quadrilaterals [BSGU94, GHT05, jWS20]. Qualitative [WS22, YLY19]. Qualitatively [Pru00]. quality [BGG12, BS96a, CGCMTR02, FP24, Zup03]. quantification [BF17, DCSB25, PKP19].

Quantum [AHGM21, CFLW22, CCP04, JT06b, Lub04, Pro24, XWW19]. quarter [dSFDG20]. quarter-point [dSFDG20]. quartic [AR24, HW22]. Quasi [BDRZ04, CPP02, Gar05, JJ15, Lem02, MD06, DE18, RVM23, SW21, AHAS21, AR24, AMP20, BWY03, BRIP08, BES18, BDRZ19, Car94, CLL23, Chu03, Cop03, CA16, Fre91, GS19, GMZ11, GLV06, GH25, GS05, IS23, JMDN^{+}22, JMU^{+}25, JL23a, Jou05, KW20, LMS09, LMP99, LD02, Mit24b, MG22, MAD23, NLZB23, RR14, SST12, SST09, SST15, SW09a, WZL13, WW14, WKP12, WG18, ZY25, BMR17b, BLM17b, FMS24, Fil25, GM93]. Quasi-Birth [BMR17b, BLM17b]. quasi-boundary [JL23a, WW14]. quasi-cell [Car94]. quasi-complementarity [WG18]. quasi-consistent [KW20, WKP12]. quasi-geostrophic [MAD23, WZL13]. quasi-interpolants [AR24, BES18, LMS09, NLZB23, SST12, SST09, SST15]. quasi-interpolation [BRIP08, GS19, ZY25, FMS24]. quasi-interpolatory [GS05]. Quasi-isometric [Gar05, Chu03]. Quasi-linear [Lem02, GMZ11, LD02, Mit24b, MG22]. quasi-monotonicity [IS23]. Quasi-Monte [MD06, DE18]. quasi-Newton [AMP20, CLL23, Fre91, LMP99, SW09a, GM93]. quasi-nonexpansive [AHAS21]. quasi-nonnegative [BWY03]. quasi-orthogonal [BDRZ19, Jou05]. Quasi-orthogonality [BDRZ04, JJ15]. Quasi-simultaneous [RVM23]. quasi-static [Cop03, CA16, GLV06, GH25, RR14]. Quasi-uniform [SW21]. quasi-variational [JMDN^{+}22, JMU^{+}25]. quasilinear [Ben98, Bra22, BIS25, EH24, LL98, Pło23, PZ20, Sch87, Tou25, Vul92, ZLG15]. quasilinearization [AN25, Naj20]. quasineutral [AGK24]. Quasistatic [CFKS07, WAV12, CFRA08, MF23]. quaternion [MB20]. quenching [LLT07]. quintic [AJK20, BPN24, Bas21, DS17, DTQ^{+}20, LHN24, Spe12]. QX [SYW18].

R [BE99, CCdlH20, WSC09]. RA [XY24]. RA-HOOI [XY24]. rabies [AJT19]. Radau [AC96, BIM15, GPHA16, Wel10a, Wel10b]. Radau-IIA [GPHA16]. radial [AT15, AD18a, BAD13, BG11c, CPZ17, Jun07, JD09, KKT16, MKJ23, Sar05, SB18, SJ11, SJ18, Sim91, uIVS13, SBS24, TLSS09, XB14, ZM19, ZT06, ZHL03]. radially [AX19, MPMD21]. radiation [CAAT16, HG98, NC16, OMP98, PGYF20, YD22]. radiative [GS20, SK97, Sin24, Vas17]. radiator [IV16]. radio [PG02]. radius [BK06]. Ramanujan [AR93]. Ramanujan-type [AR93]. ramp [MPTV25]. Random [BO87, AEA23, BJ02, BF17, BBC^{+}25, CCDJ20, CS09, Che12b, DFC09, HP18, HM22, Lav94, Pul09, SZ12, SZ17, Sob24, SD24b, SdSC99, TD09, Zha19a, dlC23]. random-data [TD09]. Randomized [JLL90, OP04, BW23b, Bai25a, GC25, LG19, Liu24, PM14, PWX24, WL21, WX22, Yan21b, ZG20]. randomly [DCSB25, DH12a]. range [Auz03, CBD16, Eva94, EC07, Has13, JP08a]. ranges [Spi93]. Rank [XY24, BLD17, Baz03, BHSW16, BHSW20, Ben02, BMR17b, CMT25, ÇY22, Con20, DESZ25, Dor91, DSS15, GZW22, GKS20, HPS12, HR14, Hey10, Hua19, Lai09, LLT20a, SY07, VGL24, ZY23]. Rank-adaptive [XY24]. rank-deficient [SY07]. rank-structured [Hua19]. rank-updating [CMT25]. Rannacher [GKMS09, HCJW24]. Raphson [JLL90]. Rapid [LR93]. rapidly [DC09]. Rate [ND85, Pré95, Sin23, SS13b, AL05, CWC24, KGR08, KB21, Maj17b, Mar09, MPDB24, TSF24, TWL23]. rates [BW23a, BFK11, CHS17, KN08, QW25]. ratio [Pic05, Wan09]. ratio-dependent [Wan09]. Rational [AM00, CPD^{+}05, KOS20, ABH22, AQ20, AMCM09, BFK11, BGVHN10, CAD03, CLS04, CV88, DBCBPP10, DBBH14, DTQ^{+}20, HA16, IO18, IN89, Iva07, JM05, LLJY20, LWZ22, Lu98a, MN08, OGV92a, PA91, RL86, SB18, SYG^{+}05, Vab21, VVV24, WT93, YR92, YLW21, ZB19a, ZLW22]. rational-Gauss [DBBH14]. rationally [Sch93]. ratios [AR93, CJM88, ML91]. Raviart [BC08b, Bra00, Kim21, Nud24, ZBY19]. ray [SSW04]. Rayleigh [OGS20, SFZ21]. RBF [BS21, BBT25, BB10, CGGM17, GJ17, JUAZ22, JGK11, KCS07, KP19, NCYC22, OT22, ST25]. RBF-adaptive [BS21]. RBF-FD [OT22, ST25]. RBF-QR [KP19]. RBFs [Boy15]. RC [AKGR14]. RD [ZWN23]. RD-FV [ZWN23]. Re [CM04]. Re-parameterisation [CM04]. reacting [TER03, dB03]. Reaction [GM85, ND85, ALMM96, ALMM98, ALMM01, AKBF19, AL98, ABR05, AGJCP25, BPN24, BJ02, BNS25, BM01, BBRBS09, BC08a, BC04a, BW96a, BIO24, BFdS10, BJTZ20, CFCH09, CdFN01, CR23a, Chn17, CK20, CGK25, DRVA20, Den15, EV96, FMP04, FJ95, FS24, GAML04, GV02, GD21, GPPR12, GPHA16, GOLJ25, HW06, HQAZ24, HAA21, HS95, HK22, HV95, JR18, JHGZ20, JT06a, JOL23, KK20a, KC03, KHB22, KCJP01, KOW05, KL09, KS04, KZ21, KTYY24, LW93, Lan95, Lan97, Li01a, LCHW20, LWYG22, Liu09, LL21, LZY24, LO96, MD19a, MMKN17, MN23, MSGM23, MG18, MOV24, MOSW00, Nag22, OZHP23, PS09, Quy19, SDK24, SZE20, wSJP15, SG09, SG07, SW13, Str98a, Top21, TH09, TMM15, VVR08, VCN20, VNC21, Vej10, Wan01, WL09b, Wan20, WC24b, Wei09, Won08, YX25, YLLZ21]. reaction [YWSL20, ZdBT03, ZCZ15, ZYLL20, ZG21, ZZPJ23, ZLWF21, VK17]. reaction-diffusion [ALMM96, ALMM01, AKBF19, BM01, Chn17, Den15, EV96, FS24, HQAZ24, HS95, HK22, JHGZ20, LW93, Lan95, Li01a, LO96, MOV24, MOSW00, OZHP23, SW13, TMM15, Wan01, ZCZ15]. reactive [CGJ16, KW98, KCC04, SW05]. reactor [MA09]. Real [LT07, MZ87, BC99, Boy06, CKL03, DO17b, DS24, DMPP99, Gla94, JJ15, Kim19, KS22, KCB02, Kür23, LMO24, MPTV25, Ore93, SD22a, dC18a]. Real-Time [MZ87, LT07]. realization [HSW99]. rebate [BP14]. receptances [LB21]. reciprocity [BGH08]. Recognition [SS17]. reconstructed [DL16]. Reconstruction [CPZ17, CGGM17, GOGF03, PTW19, ABJ12, ABD24, CH87, CdCV03, DL21b, DH12b, DSZ15b, DSZ15a, FGGL22, KHM^{+}14, Kim12, LR24, LYY15, LN21, RS09, RGÖS18, San03, SL17, WLG22, WDL23, WC24a, WC25, XZL07, ZC91]. Reconstructions [AN15, BM04c, BL06, MS08a, ZZ25]. Recovering [LSY21, SZW19, Muo23]. Recovery [FL23, Sch08a, AC23, BB94, BGS02, CCY22, Cos25, DLZ21, LN08, Nke07, SC22, YL26, ZY14]. recovery-type [DLZ21]. recruitment [TSF24]. rectangular [BWS21, CYM09, HS97, LWW20, PH91, SW21, ST08, WM07, WC24a]. Recurrence [Pré10, BRS16, Cha25, Leo10b, Lev91a, TJK18, VVV24]. recurrences [Cao98b, WW99]. recurrent [Boh21]. Recursive [FS88a, MST07, BCSH16, Koz94, SST04, SS10]. Recycle [WSC09]. recycling [SSX14]. Red [PVH25]. Redistribution [Eis86, AGM95, DPPR16, RN04]. Reduce [WSC09]. Reduced [BCG21, Dal13, GG19, Roz05, AD20c, AW09, Bai02, Boy91a, DF11, DA18b, DL22a, DS03, Dor91, DSS15, GAX24, HMdV03, LLZL24, LZY09, LJ20a, LR20b, LJ20b, Ma24, MH14, PCRR17, RÁM23, SBBC21, SR24, SLZ10, WTF25, ZZLL21]. Reduced-basis [Roz05]. reduced-dimensional [LLZL24]. reduced-order [AD20c, Bai02, DL22a, LJ20a, LR20b, LJ20b, SR24, WTF25, ZZLL21]. reducible [Der92]. Reducing [RB12, dv95a, DS02]. reduction [AMC02, AMCR17, AFLP12, BMR17b, BCJW17, CH95b, CHZZ06, Cha17, DC21, Dah02, DY03, GKB^{+}22, HK09, JLH13, KCW16, LHÖ13, MZXX24, MVG14, MPtM16, MLJ19, PJB04, SMJ12, Sch96, Sch89, Sim98, SSvG10, Udd20, WM07, WJF25, WL21, YV17, YT03, dG91, dRT99]. reentrant [MPTT17]. reference [BN03, WG23a]. Refinable [GP04, CMP15, GPP04, GS05, GP17, Pel15]. Refined [FW22, RM25, SM85, BSGU94, BS96a, EL94, HMP14, Jia00, Jia02, Zha14]. Refinement [PT19, AA94, AS97, AL98, AGJM04, AKL08, AF89, BMGM12, BW21, BFA93, DF96, FM95, HW97, JP17, LCHR03, MT05, Moo04, OT22, PC00, PPC00, RSK14, SS13a, Tro93, VT93, Zeg97]. refinement/derefinement [PPC00]. refinements [LH09]. reflecting [Dea11, GP98, SG00]. reflection [Vic87b, VS91]. Reflectionless [Pet00]. reflections [Waa88]. reflectivity [SSW04]. reformulations [PGP03]. refraction [An16]. refractions [Gro94]. regime [BAD13, MHL18, RY13, Sch04, WT20]. regime-switching [BAD13, RY13]. regimes [GOGF03, LW22]. region [AMV03, BMM03, CLTA18, FW22, lLXhLZ21, MMP09, Ou11, Ren14, SP99, ZH15]. regional [Bou02]. regions [CP05a, CMCGTR02, CSM07, DP90, NU15, Phi91, Yu99]. regression [BDD^{+}20, DEPS15, HZCZ23, LW25, LG19]. regridding [Fer93, TV91]. regula [CL07]. Regular [AFS00, Hig97, AR93, AA94, Chu03, HP97, MOS02]. Regularisation [Fil25]. Regularity [JVZ96, JVZ97, XFLC00, AW09, BM05, DK21, PV24, SZL18, YZWQ25]. Regularization [YY24, AR23, BFS17, BNKR20, BHRY21, BS18, CR23b, DH12b, EG88, Fik23, FFQ09, GH21, HSS04, HJP10, JY20, KMH21, Kli15, LPT16, Lei99, LTDY24, lLX22, MM18, Muo23, Neu88, PP24, RSY12, Sam94, San18, SWW16, SM20, Sid14, Sid23, SD24b, TWD23, WL09c, WZW13, WQS^{+}24, XXQ17, YL13, YXN21, ZH20, ZY23, Zhe19, GLLY25]. Regularized [Bad20, CZY18, CG21, AD01a, EEE22, JZS20, JQSC22, KK22a, LV12, LZH19, LH24, LHT20, Lyo12, SPS20, TT20, XXF22, YZQ^{+}22, YZWQ25, YZG23, ZLG24]. Regularizing [Buc17, Chi12, LN08, ZD21, ZW19b]. regulatory [MLK06]. Reissner [KX03, Lam13]. related [ABD^{+}25, BG24, BHHS10, DMGVPO09, FL24a, KS91, MMP20, PT19, SM13, SW09b, Zak20, dACR10]. relating [BW03]. Relation [Kid90b, IV16, Kam16, MC17]. relations [Leo10b, Lev91a, Mar99b, VVV24]. Relationship [CL02a]. Relative [TM15, Goo90, QDA^{+}26]. Relativistic [DRC85]. relax [SC11]. relaxation [AKGR14, AL22, AEF^{+}14, BN99, BZ93, CC18, CWY20, Cve02, Fan11, GGMP88, GLLY25, ITZ17, Jac93, JZK06, JW01, JT02, Kor95, LMPS19, LR93, Li00b, Mar05, MDA24, OG08, Poh93, RTV00, RTV02, RZ15, SS08b, VP91, YZWQ25, Ye04, YJJ^{+}24, ZYC22, ZY23, ZFX25, ZK00, in 95]. relaxation-oscillation [ZYC22]. relaxation-type [ZFX25]. relaxations [BDP96, SC11]. relaxed [Gu01, Liu02, PR22, SLMD21, SZL18, ZG20]. relaxing [MRS03, Tan01]. release [BCFR25]. reliability [BLW02]. Reliable [AC10, CCDJ20, CH04, HOS11, LMV25, TSF24, Wal19, Zou11]. ReLU [CCL22]. remapping [CS08]. Remarks [BK06, BP12a, Car19, BG14, DL21a]. remediation [AVMVMV09]. removal [CDI^{+}24, DFZ16, JCL18, LS24c, MRS10]. removals [ZLHW19]. renewal [BLL24]. representation [AD04, BRIP08, BJ05, DGD03, DJJ^{+}15, GNZ21, KHLV22, Moo95c, NW09, RS21, TD09]. representations [KAD24]. represented [CFV10]. representing [EGL09, Kam00]. reproducing [AD21, BLM17a, DCY20, MDASAO21, SAA20, SD13a, XZT21, ZD20]. repulsion [CPD^{+}05, GGRBRG22]. requiring [BBLT15]. research [HEJ96, KNO96, LWCT07, Var92, CB99]. reservoir [ZC92]. reservoirs [Aff94]. Residual [CF13b, Joh01, Wal95, Wei18, BB94, BSvdV99, Bür12, Cul95, GC25, GMS12, Kam16, KAS17, LS13, RBdM25, Sha05, YSBL14, Yi12, Zha97]. residual-based [Bür12, Kam16]. residue [DCY20]. resistive [BM18]. resolution [DLS22, CXNF14, CR19, CJLS98, CSCM96, Hag15, Har93, Jor11, MZN21, MM14, Moe98, MG22, Pir09, TDC13, YR22, ZSQ20, ZSQ21]. resolved [WK02]. Resolvent [Str98a, AK95, Spi97]. resolvents [EH08]. resolving [SB18]. resonance [CDJT06, HLR18]. resonances [Kim14, NT20]. resonators [Den07]. respect [CMP15, Jac93, TM05, Zen93]. response [BR94, CCS02, GRGJ02, dlC23]. responses [CFC03]. Restarted [Mor05, HC01, Jia02, KBG04]. restarting [Du11, WPS18]. restoration [BS10, BHRY21, CW20, CSL^{+}24, HWY20, Kam25, SC08, Wan23, WQS^{+}24, WOLY25, WYP12, YLH20, YLW20b, Zen21, ZZX19a, ZN21]. Restoring [LL06]. restricted [SK01]. restrictions [FS05, Ste08]. result [KS25, Vul95]. resultant [Win04]. resulting [Tol04]. Results [Bie87, AV96, BM89, BT98, BJ01, BJ06, BRS^{+}18, BC89b, CKPS15, DMA22, Fdi97b, FNT06, GH91, GGO12, GGO16, GPHAM12, GPP04, HOEC86, IS22, JL94, LLV18, LS07b, LFS15, Ma03, MRF00, NSD23, QLL^{+}08, San89, San20, Tou97, Van00]. retarded [Bic16, DD97, FS15, MD19b, ZP98]. retinex [SH21b]. retrieving [KDS22]. Reuse [WSC09, CZ97, Ere19]. reused [CMRdlT24]. reverse [AGM95, JM25]. Reversible [Lei99, HLR01, Lei02]. reverting [DMS23]. Review [Gat91, Tur93, ABFV09, Cas06, CV88, GM94, Pla08, Tsy98]. revised [SKS^{+}25b]. revisited [LJYS20]. Revisiting [ZB19a]. revolution [LX08]. reweighted [GZW22, HZZ24]. Reynolds [FHM^{+}02, Kwe01, PBC08, PGM86, SQ17]. rezone [HL89]. Riccati [AB10b, BHSW16, BHSW20, CHX26, CF13b, DS21d, JLZ25, MDASAO21, RR21, ZFC20]. Richards’ [AÁ21, APJ09, BCS17]. Richardson [LX09, NS03]. ridge [LG19]. Riemann [Hey20b, CH15, CG14, DCJ20, Gu20, LK07, LZCF21, LR20b, RZ15, Sch16a, SL15, SND19, YZC21]. Riemannian [SKS^{+}25b, SYW22, YLL21]. Riesz [AD20c, AHO16, CDW13, CLTA18, HMD21, HS21b, JHGZ20, MCS16, DE18, SSKS21, XF22, XWX21, ZM19, ZZ19a, ZZX20, Zhe19]. Riesz-fractional [AHO16]. Riesz-space [SSKS21]. right [AED12, EJS04, SY05, Shi20, ZD21]. right-hand [AED12, EJS04, Shi20, ZD21]. rigid [MCBV20]. Rigorous [RNG22, CY05, JR18, MPG^{+}16]. ring [WB90]. rings [DG96, MPTV25]. Ripa [AG25, TK15, WC25]. Ritz [BGO13, LH02, Lot19]. river [AVMVMV09, HS98, IHS13]. RK [AFS96, CX08, GPHAPPR23, SG96, SAH24]. RK-methods [AFS96]. RKDG [ZQLK11]. RKMK2e [WWZS25]. RKN [Con01, CM07, FG09, Lab99, WMF17]. RKN-type [CM07, WMF17]. RLW [TWMP20]. RMIL [KKLD21, YS24b]. Robin [AS21, Uty08, BCL15, GD23a, HS21a, HS19b, LHC23, NX22, OGS20, SDK24, SSR23, SHB^{+}25, WR20]. Robin-Robin [NX22]. Robin-type [Uty08, SSR23]. Robinson [Ran15, Ran16]. robotic [BdFPSdSC08]. Robust [BBD24, DLZ21, FH04, DLM16, MV20, PFHL09, Sha21, WCS21, ZMF25, ZZZX26, BGO13, BO04, BW97, BZ17b, CPZ17, DPPR16, DZZZ24, Eva94, GOP06, GWZ25, HOS11, JJL^{+}24, KOS^{+}12, LTC03, LR24, LB21, Nag22, SZ09, PLB22]. robustness [EH08, Fuj02, PRGO16, WZ02]. rock [RMCG04]. rods [Cop03, LA12]. rogue [DTQ^{+}20]. Romanovski [AGZD22, YZH24]. root [AC10, BASC17, CHM22, CKL03, Car94, KK20b, KK22b, OFY^{+}23, Waa88, Win01, Lu98a]. root-finding [Car94, OFY^{+}23]. rooted [Kom07]. Rootfinding [Boy07]. roots [Alt85, BBT25, Boy06, Bre06, CP94, CGTTN24, Iga85]. Rosenau [AZHD23, CCK08, MZXX24, TWMP20, ZZLL21]. Rosenbrock [AMC02, AH17, CO09, GTS20, HHR12, Nov08, PWS05, Ran16, Wen98]. Rosenbrock-type [CO09, HHR12, PWS05]. Ross [AHT17, KL23b, WMC09]. rotary [RVD00]. rotated [BL21, WCXL09]. Rotating [LR19, TC03, ZB07]. rotation [Zan91]. Rotational [LFQH21, AC18, LGS21, SLW24]. Rothe [LW93]. rough [By01, CWM09, HHW18, Mul19]. round [FV85]. round-off [FV85]. rounding [CH89, RB12]. Roundoff [QWW24]. row [BS91, SEGV02, GR93, GHW01, SW98, SBG09, WSP97, Wen05, ZX09]. ROW-code [WSP97]. ROW-methods [SW98, ZX09]. ROW-scheme [SBG09]. ROWMAP [WSP97]. rows [BW23a, KS91]. Rüdiger [GR02]. rule [CP03b, DN24, Ioa89, RS08a, Str98b, van86a]. rules [Agu15, ABD^{+}25, BW23a, BBCR22, BT97c, BDMGVO05, CMP15, CEW00, CBHY11, Der92, DRS19, DDRS24, ELR^{+}15, FLMR14, GVSL96, KM17, KCI03, Lai09, Maj14, OR20, PRSS24, Pel15, RdAP96, RS21, SKS^{+}25a, SST12, SMJ24, The17, Tom24, VV05]. run [LWD^{+}09]. run-time [LWD^{+}09]. Runge [But97, Gar10, KM18, WBCK02, AH11, AHJM19, AJ19, AHJ^{+}23, AM95a, AB17, Alb96, AMP03, AMCM08, AH17, Ant23, ARS97, AFK92, BMQW16, BDGP96, BT97a, BDP99, BJ05, BZ93, BV96, Bel97, Ben96, BC08a, BG02a, BCR01, BCET22, BS92, Bos09, Bot97, BJ11, BO11, Bru93, BJ03, BB96, BB98, But96, BS96b, BW96b, BC97b, BT97d, BC00b, BSW93, BS02, CMR94, CH95b, CIZ96, CdFN01, CG05, CCG13, CV95, Cas96, Cha96, CCM02, CG13, CCS17a, CS03, CSLY19, CGS20, Coo89, CST97, CN11, CPR93, DIJ12, De 88, DVV93, DS05, DR09a, DR09b, Dia95, DM11b, Duj09, DLV24, Eir95, EH97, Ere19, FMS24, FS05, FS08, FJL21, Fra04a, GHW20, GZQS23, GCHR06, GMG04, GJ00, GPMR95, GPMPR03, GPHAM12, GML00, HZ96, Hig96, Hig97, HR06]. Runge [HMR25, HS09a, Hoa15, HO05, HXW15, HHW18, Hor98, Hor05, HEJ96, IRC12, IJ17b, Jac93, JVZ96, JVZ97, JKN94, Jay95, JM25, JS25, Kau93, Kau95, KCL00, KC03, KC19a, KC19b, KMG09, KS08, KW98, Kom07, KS89, KS09c, KKR15, KNO96, Lab98, LX08, LW17, LW19b, LLZ^{+}22, LP97, LO96, LS05, MEGW23, MK99, Mur98, NS21a, NS20, Ore93, Ost93, OT02, Pan07, Pat98, Pat00, PCR17, PJB04, PAJ12, PCRR17, Ran16, SA21, ST89, Sch02, SC25a, Som86, Som93, ST25, Spi96, Tsi01, VDVV98, VV02, Van00, VZ93, Ver96a, Ver06, Ver96b, WGKS12, WYL11, WHL19, WSW96, WX06, XG22, YZWQ25, YBW20, YC13, Zen93, ZZL01, ZJ19b, ZYQS21, ZFX17, ZZX20, ZSQ20, ZSQ21, ZC99, in 92, in 96, vSC92, vS93, vC93, vS96, van96, vdSvdH95]. Runge [vB95]. running [CRU15]. rural [CCM17]. Rutishauser [AC10]. RWPM [HK93]. Ryaben′kiı [Tsy96]. Ryaben’kii [Ano00h, EST15].

S [EST15, MZ87, Tsy96, UHLZ24, SZY21]. S-ADMM [SZY21]. s-stage [UHLZ24]. Sabin [BFGP08, LMS09, MB10, SST09, SST15, Spe12]. Saddle [DJ10, AL05, BM12a, BHB23, BCGS24, Cao03, Cao07, Cao09, Cao10, Cui04, Elm02, OG08, Sim04]. saddle-point [Sim04]. saddles [SSL93]. Saint [FCX06, IMC22]. Saint-Venant [FCX06, IMC22]. sample [YXZL24]. sample-size [YXZL24]. sampled [RS09]. sampling [AQJ18, AQ20, IKS25, KK02, QL16]. sampling-based [AQJ18]. sand [FF20]. sandwiches [NBP94]. SAOR [Yua93]. Satisfaction [PT09]. satisfying [BRS16, SYY20, SG16, Str98b]. saturable [LN26]. saturated [MCM12]. saturation [PRS20]. SAV [HMD21, ZL23, ZZF25]. saving [LCJQ12]. scalable [DHS05]. scalar [AKG14, Bac14, BP85, CCL22, CP97, Die15, HMD21, KKR15, Mon21, Pel20, Ric91, Tan01, TDC13, TZA13]. scalars [MA04]. Scale [BHJJ06, WVBM88, AB10b, Bai02, BHL^{+}21, BMR^{+}17a, BHSW16, BHSW20, BT02, BJ11, BDD^{+}20, CLL23, CR23b, DRVA20, Dat99a, DW15, Ewi91, FTB97, FBM17, Gen10, GT02, Hab08, IKR^{+}22, JY20, KM11, KD25, LP05, Lei02, LS13, LCJQ12, MW24b, MPG^{+}16, Per99, PS03, PS19, Pul05, RMK09, SDK24, SM89, Toc01, WDZS21, Wan25, Wan23, WQS^{+}24, Yan21b, YLH20, ZD21, van95]. Scale-invariant [BHJJ06, Wan25]. Scaled [GP93, BKAG22, CJ25]. Scales [RMM12, Boy91a, JEG10, SZ22a, SWL20]. Scaling [BLP01, BMM03, BSZ99, FK23, Mon21, NZY21, PCR17, PRS23, SBS^{+}20, SW09b, Wei09, dRT99]. scattered [BE99, CPZ17, DR93, GÖ20, LN08, PS02, PS03, ZHL08]. Scattering [FvdMS17, AA20, BBLT15, BPV25, CWM09, CM14, CAAT16, DH12a, FvdMS20, GAW09, HD04, HM86, HMP14, IT07, KJL12, KNP16, LKJ20, LLN25, LGH11, MP05, Map05, Men23, Ngu15, NT16, QC12, QL15, Rat13, Tah96, VT91, WN12, WL16, WQ17, WG19, YL26]. schema [MR20]. schemata [MPTV25]. scheme [AA05, AD20c, AKG14, AAH21, AA24, ACMR06, ABR05, AR15, APJ09, AGK24, AG25, AS20c, BF92a, Bar09, BK09, BtTBV87, BMMZ06, BP97, BTC23, BIO24, BC23, BRBM08, BBKS07, BC97a, BS18, BCS06, BD22, COPPS25, CGT13, CXNF14, CFLW22, CHM22, CCDJ20, CFS13, CEL25, CHS19, CC19, CC20a, CCST22, CL24a, CQZ20, CNT07, Chi21, CRU15, CCL04, CMCGTR02, CMP23, DKSS24, Dav98, DGE22, DL22a, DLM20, DGM18, DMM24b, DYZ20, DLQZ23, DHM09, Duf90, DJJ^{+}15, EEE22, FVGS13, FP02, FE93, FWHM20, FGP23, FCW21, FGGL22, GLLW14, GLML20, GHS25, GM18, GS94, Guo00, GZHQ23, HT24, HKO12, HPH20, HC22, HK25, HJZ23, HS86, Hol01, HLMKZ06, HJKW17, HJYL19, HL21, HZZ25, HLJ20, HLY22, HV89, IMC22, Ito17, Ito22, JK21, JK17, JZXJ21, JN07, JZZH22, JT02, KL21, KS00, KMH21]. scheme [KD13, KHB22, KHYY21, KCY19, KW12, KKK25, LO22, LRS23, LHN24, LN26, LDIW16, LPR00b, LMA18, LMY18, LL19, LH20, LA21, LWYG22, LXZS22, LS23, LH23, LGC25a, LYZF25, Li25, LZW19, LZH^{+}25, LS99b, LSGK15, LCL18, LLD18, LS20, LFS21, LL21, LCZ23, LGZH24, LCZ25, LXPY25, LD22, LE94, Lte24, LP01, MDP10, MD19a, MD20a, MSZ^{+}24, MMDH19, MMD20, MM20a, MPV24, Mic03, MD25, MdD04, NV23, OR18, OR20, OFY^{+}23, PXHZ20, PCL^{+}25, PKSB10, PGDB08, PHY19, Pel20, PD01, PGYF20, PVM22, PH17, Pło22, Pow94, PWX24, PYD21, QH22, QXQ22, RZ00, RZ18, RL21, RGB20, RT14, RK91, Ror06, Ros93, RGMO19, RG21, RO16, Sac93, SC19, SRMDRL23, SYY20, SG16, SC20, SXL22, SL22, SL21, Shi25, SNW22, Sin23, SKS23, SED21, SB19, SA18, SBG09, SR97, SW24, zSW06, SG17, SND21, TZ21]. scheme [TWL23, Tan93, TX18, TH18, Tan23, TLG20, TDW23, Ton04, Tou25, TDPU17, Tow16, TY00, UHUL21, VA21, VVD95, VCC12, VRC21, VL19, Wan25, Wan07a, WL18, WH19a, WDH20, Wan20, WDU21, WWL21, WLM21, WCS21, WCM21, WLG22, WWZJ22, WDL23, WG23a, WaZW23, WLY24, WWZS25, WWZ25, WC25, WLCZ26, WS04, WG11, WYY20, WSC21, WWWW25, WL25, XWW19, XW19, XK25, XZL19, XWX21, XWC25, XLZ23, Yam18, YTC24, YZH19a, Yan22, YZ21, YCWH23, YXZ18, Yua20, ZBD24, ZW09, ZWH^{+}17, ZWJ18, ZJH18, ZLX22, ZLSZ22, ZJH^{+}23, ZWN23, ZL23, ZFS24, ZY25, ZZF25, Zha01, ZLX19, ZH24, ZZL17, ZZLL21, ZFX25, ZGDL17, Zla85a, ZR15, ZL24, tV87]. schemes [AC98, Aca12, AJ24a, AD08, AEMX17, AÁ21, ADFR18, ACLM22, AGM95, AM04, ACM91, ÁKM20, BDMG12, BCG21, BBD20, BBRBS09, BCE04, Ben17, BG24, BCCHM21, BK21b, BC05, Bi20, BM12b, Bor16, BSvdV99, BO11, BDE22, BR20, Bra22, BPV25, BL06, BMWH20, BO21, BS20b, BL08, CCCC^{+}25, CLT97, CMP20, CCG13, CGA93, CW98, CC04a, CCS17a, CZY18, CR19, Chi12, Chi93, CKM25, CS18, CKK25, CMMR23, Con04, CM04, CCS17b, DZZZ24, DD97, DEGDdL25, DT10, Dit21, DKK94, DMR18, DS15, EN09, Est95, FMS24, FCX06, FG98, Fid17, FHX22, FL01b, GX11, GZQS23, GHKM09, GPMR95, GGRN17, GGRBRG22, GGT24, Gul15, HO10, HOEC86, HCS20, HCC24, HEG16, HZ20, HW15, HWZ22, HL24, itHT18, ID19, IKMM23, JL86, JL87, JTB15, JWG20, JQSC22, JY23]. schemes [JJL^{+}24, JL94, KGR08, KTK20, KWY24, KMR09, KAS22, KCL00, KC03, KC19b, KL07, KOGL25, KM18, Kra92, KR12, KKP17, KLSW06, LCVG01, LP24, LMPS19, LY10, LRC19, LXZ21, LW21a, Li22, Li23, LPV24, LS21, LYC24, MS03, MS24, MZXX24, MM14, MCD20, MV18, MJS23, NWL^{+}22, NRR06, ÖT20, Ort20, Pan07, QW04, QZH25, QNA23, Ram94, RGÖS18, RBdM25, Rog19, RV05b, RTU15, Rus95, ST14a, Sam94, SGR21, SA21, SMB23, Sch93, SUP^{+}12, SM93, SS94b, SZ22b, SZE^{+}92, Sid02, SDK15, SHG86, SSPZ20, SLZ10, SGN06, Tah96, Tan01, Tan24, TGB08, TZA13, Tol03, Tol04, TÖR22, TH09, Tou10, TJ12, Tur86, UWY22, Vab21, VV02, VRK25, Ven15, Vul92, Wan21, WWM22, WPAZ24, WKWC26, WB03, XLZ20, XP23, YPD21, YZG23, YR22, YZH24, YDWW17, YH07, ZM17, ZY19, ZYQS21, ZZZ23]. schemes [ZQZ23, ZCQ25, ZJ25, ZZjQ24, ZW24, ZLG15, Zla85b, iW07, iV09, iW09, iM13, vSW90]. Schnakenberg [OZHP23]. Schoenberg [FL24a]. Schole [Rou20b]. Scholes [AAM03, ALZ^{+}21, Bis11, RG21, iV09]. Schröder [CP94]. Schrödinger [ZCSH11a, AZA22, AT15, AL22, ÁMS14, ÁMS17, BUL23, BOEP00, BZ17a, BSV09, BC23, CHP19, CC23a, CGP15, CCG13, CJL13, CC19, CC20a, CWY20, CG21, DW21, DTQ^{+}20, Duj09, Ehr08, EEAS25, FWHM20, FI03, FCW20, FXCW21, GT15, GX11, GLLW14, GM16, GD23b, HBA^{+}24, HCS20, HC22, HCC24, HZ20, HMY19, HY24a, HLMKZ06, HJKW17, HJL18, HLJ20, HCGW22, JLZ20, KCY19, KOGL25, KLSW10, LO22, LHN24, LN26, LZH19, LSWM19, Li23, LZ24, LC24, LZW17, LZW19, LS24a, LCM22, LCM24, MZZ17, MP11, MG97, Moe98, NAD25, PXHZ20, Sim91, WL18, WH18, WMLB19, WDH20, WT20, Wan21, WL22, WKWC26, WL24, XW19, XL23, XCHW22, YZWQ25, YLX21, Yan23, YF24, ZM25, ZWG25, ZFX25, ZCSH11b, ZR15]. Schrödinger-like [LCM22, LCM24]. Schrödinger-type [Ehr08]. Schrödinger/heat [AL22]. Schur [Cao09, Cao10, CG89, GM87, KMS10, LN24, RZ04]. Schur-complements [GM87]. Schwarz [AL22, AK95, CC18, DdSF07, DLN04, DSS15, Heu00, HH10b, KKP07, KFOF02, LHH96, LCS19, LX21, MOS02, Mai06, Mai09, Mar05, MPHFP23, MOSW00, MLB97, NMB10, OH20, QLL^{+}08, Wan05, WCXL09, ZT06]. science [KAS17]. Sciences [FDF^{+}25, COPPS25]. Scientific [BJS12, CFTW08, Jbi25, LST07, MH89, MM16]. Scott [BL08]. scrambled [Sch08b]. SDE [AH17]. SDEs [BSTT22, CDR20, DR09a, GHW20, KP18, Kał22, MEGW23, RO16, Sca22, WG22, Yam23, ZM17, ZW24]. SDFEM [FLR08, ST08, ZMC13]. SDIMSIMs [JAH21]. SDIRK [CPP02, IJ21]. SE [QM10]. Search [DIJ12, Bur93b, CH22, Fer14, HP15, MD23a, Sch99, SXP09, WZ16, YZH19b, YLW20b, ZH15, ZP12]. seawater [AMRR18, MNR14]. secant [ABKG21, CSM07, Nov08]. Second [AKG14, Beg00, BMPR15, FH20, FJP17, yGqWsWC05, ITZ17, LL19, TDPU17, WDZS21, YQCZ22, AH11, ABH14, AH15, ACDP22, Abu04, AB88, AHA23, All24, AHB20, AT13, AD18a, Aze22, Bac17a, Bac17b, Bac19, BG24, BHR05, BRBM08, BBKS07, BEH24, BMM97a, Cai24, CGA96, CXZ17, CSLY19, CWX21, Chi21, CCJ99, DR09b, Deh05, Den93, DC09, DL21b, DAMA23, EHM01, EH91, EH07b, FW08, FWHM20, Fer93, Fil25, FG09, GLML20, GP23, GOP06, GGR97, Guo00, yGpY09, Gwi09, HGM^{+}21, HKO12, HP18, HZZ25, IKMM23, JZXJ21, JN07, JY23, JJL^{+}24, JCN94, Kan89, KX91, KwS19, Kin94, KKP17, KLSW06, KW10, KDD23, LHHR94, LW17, LMY18, LLJY20, LH20, LWY20, LW21b, LZ22, LQXK23, Lia22, LX09, LWZ22, LTT19, LS20, LW20b, LCZ23, LGZH24, LXPY25, LNZ12, LE94, LL02, LDH^{+}24, Luc05, Man97, MVVA09a, MYSC17]. second [Mie89, MD20c, MKJ23, Moo04, MAF20, MSA20, MAH22, Mur98, Nap16, NSCC19, NV23, NWL^{+}22, NSD23, Pan21, PHY19, PMP23, Qi24, QXQ22, Qiu23, RZ00, RG20, RT14, RO16, SMJ12, Sac93, SAA20, Set24, SV00, SC20, Ske89a, Sof17, SvdHN86, TZ21, Ton04, TH23, Udd20, VV02, VV09, VL19, WH13, WWX13, WMF17, WWZJ22, WLY24, WW24, WKN20, WAV12, Wu09, XXYZ24, XK25, XL09a, XY19, Xu21, YJ21, ZWH^{+}17, ZLX22, ZLW22, ZFX17, ZH24, ZLS20, ZZJ21, ebKMZ24, iW09]. second- [LLJY20]. second-degree [Kin94]. second-kind [Lia22]. second-mode [EH91]. Second-order [BMPR15, LL19, WDZS21, YQCZ22, Abu04, AB88, AT13, Bac17a, Bac17b, Bac19, BBKS07, BEH24, CGA96, CXZ17, CSLY19, CWX21, CCJ99, Deh05, Den93, DL21b, DAMA23, EHM01, FW08, FWHM20, Fer93, FG09, GLML20, GGR97, HGM^{+}21, HKO12, JN07, JCN94, KwS19, KLSW06, KDD23, LHHR94, LW17, LMY18, LH20, LWY20, LQXK23, LWZ22, LTT19, LW20b, LCZ23, LGZH24, LXPY25, LE94, LDH^{+}24, MVVA09a, MYSC17, MD20c, NSCC19, NV23, NWL^{+}22, PHY19, Qi24, QXQ22, Qiu23, RO16, Sac93, SAA20, SC20, Ske89a, SvdHN86, TZ21, Udd20, VV09, WWX13, WMF17, WWZJ22, WLY24, WKN20, XXYZ24, XK25, XY19, YJ21, ZLX22, ZH24, ZZJ21, ebKMZ24, iW09]. Secondary [DRC85]. secret [MPTV25]. section [IJ14, Lin10, Meh08]. sections [RLMG24]. Sectorial [Auz03, BG02a, Chn17, PGA93]. securing [GL93]. sediment [BDES12, KDAK16]. sedimentation [Die15, MM14, RBT15]. Segel [DvHM19, MDRR11]. segmentation [BCM04, LK14, SMA01]. segregated [vdHVW01]. Seidel [WX22]. Seismic [SR88b, BR25, BBD20]. seismograms [Abr93]. Selected [FJ97, ND85, vdHSW98, KW12, Sca22]. selection [BW23a, BBPR05, BT97b, CZY18, Fdi97a, HL97, KWY24, KKR25, NCYC22, Odi19, San18, SPS20, SL21, Söd06, WPL16]. selector [YP18a, YP18b]. Self [TLQ21, BS18, GLS09, Gul15, KOGL25, LQY25, NZY21, RK08, Rob10, SL08, YLL21, ZX22]. Self-adaptive [TLQ21, LQY25]. self-adjoint [RK08, Rob10, YLL21]. self-diffusion [GLS09]. self-mapping [SL08]. self-regularization [BS18]. self-scaling [NZY21]. selfadjoint [HM01, XYHM20]. semantic [TS08]. Semi [Bou02, CCCC^{+}25, Fat10, FMU15, GPHA06, KMR09, Mar99b, Min04, AGQ^{+}24, Beg00, BB15, BC08a, BN03, BL06, BH97, BD22, CGT13, CXNF14, CGJ16, Cao01, CWX21, CYWH22, CL09, CBHM19, DTQ^{+}20, DII15, EH24, EEE22, FVGS13, FD16, FR18, GBBC^{+}23, GS09, HCY18, HLJ20, HS21b, ID19, KDH20, KR25, KKP17, LWT07, Lay08, LMA18, LL19, LA21, LW20b, LL20b, LXPY25, LfX15, LRT99, MZZ17, ML16, MMDS21, MD20c, Pan07, PP00, RT14, SH09, SS21, TZA13, WL10, WC11, WW19, WJM22, ZR15, de 96, CF13a]. semi- [EH24]. Semi-analytic [Fat10]. semi-analytical [KDH20]. semi-axis [GS09]. Semi-classical [Mar99b, Beg00, FR18, MZZ17]. semi-coarsening [LWT07, de 96]. semi-discrete [BL06, CXNF14, EEE22, RT14]. semi-discretization [CGT13, WJM22]. Semi-Godunov [KMR09]. Semi-implicit [CCCC^{+}25, FMU15, GPHA06, Min04, Bou02, BN03, BD22, CWX21, CL09, DII15, FD16, GBBC^{+}23, HLJ20, ID19, Lay08, LMA18, LL19, LA21, LL20b, LXPY25, MD20c, Pan07, SH09, SS21, TZA13, WC11]. semi-infinite [KKP17, LRT99, ZR15]. Semi-Lagrangian [Bou02, BB15, BC08a, CBHM19, FVGS13, MZZ17, PP00, WL10, CF13a]. semi-linear [BH97, CYWH22, HCY18, HS21b, LW20b, ML16, MMDS21, WW19]. semi-local [AGQ^{+}24]. semi-permeable [CGJ16]. semi-rational [DTQ^{+}20]. semi-simple [LfX15]. semi-smooth [KR25]. semi-stable [GPHA06]. semiaxis [DO17b]. semiclassical [WT20]. semiconductor [AW03, BCCHM21, BK12, MT11, ST05, YY13]. semiconductors [JT02, JT06b, Sac93]. semiconvergence [SW03]. Semidefinite [Ren99, BSZ99, GS99b, Hua20, Hua21, JS25, SZ99, VB99, dRT99, ZNK02]. semidiscretisations [Sei02]. semidiscretizations [ALMM96, ALMM98]. semilinear [ANN19, AA24, APJ10, Bog12, CL10, CDW23, CWZ23, CJ24, GJlL23, HKO12, HN22, LCHW20, LXZ21, LS98, MR20, MT20, RS20, Sch95b, SD09, SR24, Tia15, Vul95, WWWW25, XHJM21, XGHM22, XHYM22, YM24, YCWH23]. Semilocal [AHR12, ABM17]. seminormal [RSK14]. semiorthogonal [MB08]. semiperiodic [VVD95]. semiregular [KV20]. semismooth [PGP03]. semistrip [DT15]. Sense [BAP^{+}06, SSA^{+}22, VA21, ZAB15]. sensing [ABKG21, BKAG22, KKLD21]. Sensitivity [BB10, HO16, LPZ00, SMEN04, Toc01, ABdSG23, BBPR05, BAP^{+}06, CHZZ06, FTB97, GFB99, JLH13, LHÖ13, MP96, TLP18a, TLP18b]. sensor [RSD^{+}06]. sensorineural [BF09]. separable [AMR12, BD17, GBDB97, MRÁSY24, SB14, SY07, SZY21, SGY22, YJJ^{+}24]. separated [Eng11]. Separation [KL87, PM03, CDH24, EW08, EHV24, Gus87, OK98]. Separation-of-variables [PM03]. September [Ano23x, Wen10a, Ano21r, Ano22r, Ano24p, Ano25p, vdHSW98]. septic [KK23]. Sequence [Bak86, Ber85, BZ94b, BZ96, Bre02b, Dra91, Fdi97a, Gil91, GM94, Hua98, Lem02]. sequences [AGKS25b, BL86, BKP15, Dés08, DJM09, GGM07, Jéz04, JT09, Lav94, Mar99b, Sch08b, SSX14, TJK18, Wen10a, Zuu95]. sequential [BMGGG12, BT97c]. Series [GGS16, RB12, AR93, BP02, Bor10, BW15, Boy06, Bre02b, CFV10, CM06, CL88, DK11, Gil91, HW93, JKN94, LG87, Lon86, Lon88, Mat91, MdR05, Meh22, Moo95c, Mur99a, Mur99b, PCA10, Pré95, Sch98, Wen10a, Wen10b, WB90, Woź10, YR09, ZB19a, ZJ10]. set [CP07, DMH18, DLP06, FM07, FMU15, GT02, GKL07, HA21, HMW05, KM11, KK22a, MCS06, MPG^{+}16, Sch09, TL07, VA21, ZWL11]. set-oriented [MPG^{+}16]. set/moving [PH15]. sets [BCDR25, CP10, DMM24a, Eng11, HP97, RTH23, SL08, Zup03]. setting [AD04, GC15]. settling [BCS06]. Seventh [CFTW08]. Several [KC94, BS20b, CKM10, Dar90, Kid90a, Kid90b, MDD14, Ros93, Str98a, Xu13, ZD21]. SFDEs [XZ19]. SGI [Zar99]. Shabat [FvdMS17]. Shading [CP06, GV04]. Shadowing [Zha19a, HJ05]. Shadows [SSL93]. shallow [AQS94, BDMG12, Beh97, BTC23, BDE22, BD22, CCCC^{+}25, CFXZ06, CCM17, DEGDdL25, DL21b, FCX06, HAN23, HL03, ID19, IM00, KOS21, Lie01, PCL^{+}25, QLL^{+}08, STS00, SCT05, SW12, Tou10, VCC12, WLG22, WC24a, YTC24, ZGDL17, de 92b, vdHS01]. shallow-water [Beh97, QLL^{+}08, SW12]. Shanks [BRZ17]. Shannon [BDOG19, CPOGO17]. Shape [CP06, DF92, GM08, GMZ08, PGM86, Rat13, TLP18a, TLP18b, YZ24a, ZWL11, GM10, KP01, LYZJ23, Meh08, NCYC22, PM05, Rab25, Rya00, WM08, GV04]. Shape-from-Shading [CP06]. shaped [BDSG09, DH12a, GWZ25, LC19, Pel20, SW24]. shapes [HHL23, ZS18]. shaping [EH05, LT07]. shared [JP93]. shared-memory [JP93]. sharing [MPTV25]. Sharp [WCM21, GOLJ25]. Sharpening [FP02]. Sharpness [FNT06]. Shaw [GLML20, GJLL20, WWZJ22]. shear [IK24, KC94, LBCN00]. shear-layer [KC94]. sheet [CDV00]. shell [SYL^{+}20, YSZG25]. shells [FSB97]. Shepard [DDNZ18, DD20, DDGN23, DDF24, Zup04]. Sherman [Mit24a]. shielding [RU07]. Shift [BLM17b, BFLR23, FL24b, WPL16]. Shifted [AAEMY21, BLY16, BB10, CCQ^{+}23, ER18, EVO06, HAA21, HvdHV10, LIPT18, PH17, QDA^{+}26, SSX14, Tow16, WPS18, vdES04]. shifted-Laplace [EVO06, HvdHV10]. ship [Mot17]. Shishkin [FR01, Len00, LS99b, NS03, RN25, RG22]. Shock [KXR^{+}04, Sal93, CSCM96, Gro94, KD13, KCJP01, KOW05, Qui96, Sod91, XLK07]. shock-capturing [KD13]. shock-induced [KCJP01]. shocks [Kop89]. Shooting [ABRW18, ELLE02, GM93, GGM95, HK93]. short [Mun00, WW99]. shortening [FG01]. Shortley [LHWF08]. shrinkage [Jia12]. Sibsonian [BS00a]. side [AED12, DL13, Shi20]. sided [ABI22, GH02, GL17, LS21, MB20, MWC21, MT06, YLFT20, YS24a, YDWW17, ZEW20, ZLWF21]. sides [EJS04, ZD21]. sigmoid [Aze22]. sign [DCC14, QZM25]. sign-changing [DCC14]. Signal [ABJ12, BPV25, Kam25, SM13, Söd06]. signaling [MDD18]. signals [BCV25, Pul05]. Signorini [BRS05, KPR06, Wal19]. Silicon [CB99]. SIMD [Ney95, SD93]. SIMD-type [Ney95]. similarities [Gab02]. Simple [RGB20, AAL^{+}26, Ban97, DSZ15b, DSZ15a, Lam13, LS12, LS93, LfX15, Tan24, Van92, Win01]. Simpler [SMTHE22a, SMTHE22b]. simplest [SvV22]. simplex [DhW09]. simplicial [MKH16, PT23, PC00]. simplification [DEPS15]. simplified [CLL25, FN95, GS20, LH02, SKS^{+}25b, Zha01, dlC23]. SimRank [NLLG20]. simulate [DA18b]. simulated [Dah02, GOGF03]. simulating [FPRA09, Pul05, QPT23, SWFK13, SWJ^{+}24, Tan24].

Simulation [CF86, CL02b, DRC85, MH89, MD00, QW04, Ta’86, Tse00, WM08, AQS94, AA22, AEA23, Ari04, Arn95, AS06, BBD20, BGM^{+}09, Boh03, BBL02, CC24, CK22, CCK03, CCM17, DS07a, DA19, Den07, EWW99, FMW18, FKA^{+}13, FMSV07, FJH^{+}01, GG95, GT19b, Gro94, GD22, HLR18, HPW21, KL07, KTS03, LCVG01, LA12, LT00, MH04, Min87, MDD18, NRWF08, NT20, NA21, OB20, Odi24, OB24, OK98, OHA25, QM03, RI02, RR00, RVD00, Sch91, SUP^{+}12, Son91, ST05, Spi90, SH91, TER03, TYKK01a, VV95, Wag85, XLK07, ZdBT03, Zha96, ZS21b, dlC23]. Simulations [WVBM88, AJT19, AF04, BGG^{+}21, CS19, CS01, CKS05, De 02, DGGM25, FJL21, Kar89, KC94, KKE16, LS07b, MNR14, NC16, Nür09, SYG^{+}05, SM89, Tia15, TSFB01, WK00, Zan91, ZQLK11, dB03]. Simulator [Nak05]. Simultaneous [ASZ15, KN08, CP94, DO17b, KYI17, PSR04, RVM23]. simultaneously [SBBC21]. Sinc [BBJ25, AQ20, AT13, Boy91b, CJ22, MLB97, Ril92, EEAS25, KB21, KESYB23, QXG21, RG20]. sinc-Gaussian [AT13]. sine [AS20a, ATAW25, Bac17c, Boy07, DC18b, FCW21, Har00, HS21b, KAD24, LZ18, Nak24, SBBC21, SLW17, SK22, YF24, ZM25, XWX21]. sine-Gordon [AS20a, ATAW25, Bac17c, DC18b, Nak24, SLW17]. sine-pseudospectral [ZM25]. Single [KKN^{+}17, AM99, BC12, BN12, CDR20, GGN12, HJZ23, LL98, TT20, TLQ21]. single-phase [LL98]. single-species [HJZ23]. singly [BC97b, BC00b, FS08, KW20, RBdM25]. singly-diagonally-implicit [FS08, RBdM25]. singly-implicit [BC97b, BC00b]. Singular [LPS25, WW24, AY21, Aff94, AL95, ADR17, AFS00, AAD14, AM10b, Att97, AAB^{+}22, BWY03, BLM17a, BPTT15, Bru07, BKW06, BRW17, BRS^{+}18, CHLX07, CJ90, CHS19, CCST22, Chi12, CMT25, CH21, CP17, DN21, DLPV17, DCY20, ELR^{+}15, FK23, FHM^{+}02, Fat10, GS17, GMG19, GLM09, GGS16, Gu19, HM87, HT24, HM09, HBB25, HSWW24, HY24b, Jad94, KV07, KX91, KHLV22, Kat89, KM25, KSHB21, KLR25, KKW00, KBG04, KDKW20, LG21, LP24, LLKJ21, LDC24, Lia22, LD97, LO03, LMSW17, LCZ25, LNZ12, MAHZ21, Maj14, MKDC25, MO17, Man96, Mok17, Naj20, ORT24, PPT02, PT11, PTV16, PTV20, PT15, QXG21, Rab94, RN22, RGA19, Rou20a, RCR24, SWW17, SY05, pSLqJcY16, SHLY19, SVB17, Sid90, Sid14, SPYS24, SS16, SW03, Sto96, TMD92, Tan93, TS08, Vas17, Vul92, Vul95, WTY21]. singular [WCM21, WZ22, Win01, Yos00, ZG92b, ZD20, dAF17, dSFDG20]. singularities [BH97, DDF24, Has20, KM17, LH02, LHWF08, MK20, PV24, YS24a, ZMY21]. Singularity [RLSS06, CKK10, Has09, HY01, KLSW10, LL21, PA18, YT03]. Singularly [Rei85, AL98, AN22, ÁMS14, BPN24, BNS25, BBJ25, BM00, BM01, BO04, BBRS97, Bog00, BFLR23, CHX26, CR23a, CJ23, CJ24, CJ25, CKK25, DLZ21, DAMA23, FL24b, GD23a, GD21, GOP06, GO19, GO21, GO23, HO24a, HO24b, KMS19, Kau95, Kau97, KS04, KK20c, KD25, Li00a, Li01a, LX24, LT01, LWC24, LW95, MOSW00, MPtM16, Mus11, OS08, OQ15, PZMX16, Pap95, PMP23, RN25, RG22, RK08, RC18, RSR23, RTU15, SDK24, SK16, SW21, SSR23, SKS23, TS24b, VN21, WC24b, XLHL25, YM24, YZ17, Zar17, ZCZ15, ZL18a, ZL22, ZLG15, ZX14]. sintering [Zha96]. SIPG [MKBS26]. SIR [CML^{+}25, LSVT25, WS22]. SISO [De 06]. Sitter [MPMD21, ZH21]. Sivashinsky [AS04, KK23, Mar93, MDA24]. Six [MS03, Bri85, Par21]. six-dimensional [Bri85]. Sixth [SA18, AHR12, BM06a, CHSS01, CST97, RTA19]. Sixth-order [SA18, BM06a, CHSS01, CST97, RTA19]. size [ALM04, AFS02, DS07b, Den07, Ein18, GPPR12, HOS99, HLR01, Hor05, KP18, MJS23, PAJ12, RAS99, Wai98, WWM22, WK02, YXZL24]. size-binning [DS07b]. size-resolved [WK02]. size-structured [ALM04]. sizes [Sha85b, ZZjQ24]. skeletal [MD25]. skeleton [PC00]. skew [DS21d, Kru99, KCB02, Zan91]. skew-Hermitian [DS21d]. skew-symmetric [Kru99, KCB02, Zan91]. Slater [CPD^{+}05]. Sliding [Die20]. slightly [BB10, Kie17, LHY24, de 95a]. slip [AA25, LA11, RZS21]. slope [CZY18, KWY24, SL21]. slots [CDJT06]. slow [KOW05, SKW17]. slowly [CSSZ25, Woź10]. Smad [MDD18]. Smagorinsky [RÁM23]. Small [PGM86, AKGR14, AA20, BBLT15, CWC24, FXY22, FS23b, Pav00, XLHL25, Zar17]. smallest [KBG04]. smectic [FMW18]. Smith [ADM10]. Smooth [JR18, AR24, BUL23, BCFR25, BBLT15, CL14, CL24b, Dar00, FvdMS17, GANT02, HK22, KM25, KR25, KR18, KDS22, MD06, Mat86, Ram12, TWH21, TWD23, UHLZ24, ZYJD25, dOF20]. smoothed [CDD00, GJV08, LL06, SKR^{+}16]. smoother [BGM19, BS97b]. Smoothers [Bre02a, ABCC18, BG02b, PRGO16, RGL16]. Smoothing [DLPV17, GSR00, Naj20, ZY23, CRS05, Haa97, KM95, MYSC17, NSCC19, Phi87, RN22, THW19, Tan23, Wal95, ZRA23, Zha97, vSW90]. smoothness [RGB20, Tan24, WG23a]. Sobolev [AD20a, AyLqW18, CMR12, Che12b, DSA20, DMR10, FMPP24, MD23b, Ren13, XFL22, Yan21a, YW24, ZQZ23]. Soft [GAOB20, KJL12]. Software [Enr06, FN95, NT92, SG96, BDF89, HK93, MLK06, MP96]. soil [CDP^{+}25, MCM12]. soils [BBT25]. solenoidal [MM07]. solid [GMM09, MP05]. solidification [CMP06]. solids [Cim25, MR06]. solitary [AD01a, CGP15, DLM02, EZ03]. soliton [GT19a, Ma24]. solute [DA18b, DFC09, FLL11, WPAZ24]. Solution [AJK20, AO91, BBD08, Bre06, Che96, CC20b, Fre04, GM85, Jac87, KM95, KS91, LH21, MK14, Mik97, MŠ99b, Njå88, PGM86, Rum87, SH09, SBS^{+}20, SM85, AD21, AN25, AACP20, AM09, AHT17, AGJ12, AA20, AHB20, AB09b, ADM10, Ara99, AS00, AAD14, AD18a, BHJ05, BF92a, BOEP00, BFS17, BBCR22, BR25, Bea04, BRTB19, BB15, BHSW20, BNV06, BF92b, BLM17a, BGH08, BKP14, BDFV95, Bor02, BHR05, BMP05, BBBN21, BHJ13, BTDV10, BVB10, BVRB14, BGIW18, BMM97b, BMM97a, BSP04, BIS25, BCDP17, BP95, CLT97, CDP17, CG92, Cas96, CW98, CJ18, Cha25, CEL25, CH01, CH04, CF08, dCCSR03, CNS00, Con89, Con20, CFM^{+}24, CST97, Coy12, CG14, CRSF19, CMP23, DD21, DS21b, DB97, DS17, DN21, DA18a, DMH18, Der92, Dia95]. solution [DS97a, DL13, DB08, DBBH14, DMPSC16, DZMB21, ECB07, ESE20, FL93, FL05, FJ97, FL09, FL15, FvdMS20, FH08, FV01, FV99, GS19, GNUV24, GMG19, GeO24, GLM09, Güm20, GN86, Guo96, HGP11, HZBM05, Har09, HS86, Hey20b, Hig93b, Hin95, HvdHV10, HW97, HJ06, Hor99, HDS20, HBJ09, HST14, Isk89, JL91, JRW06, Jac96, Jéz99, JCSR03, JNPC03, JMPY10, KL21, KHLV22, KS25, KKR24, Kel85, KG90, KNP16, KW95, KKW00, Kop86, KP19, Kur98, LO23, Lau17b, Lau17a, Le 12, Lee23, LSVT25, Lev91a, LHWF08, LZH19, LLT07, LO03, LW92a, LCLW17, LT93, LP00, LD02, LYA^{+}19, MPTT17, MS03, MD19b, MO17, MS19, MK21, ML16, MKN23, Mar99a, MVVA09b, MS25, MMRV20, Mat05, MR01, Mie89, MSS21, MD10, MD96, MRFF17, NK11, Naj20, NAD25]. solution [NS12, NT92, NLZB23, OMP98, Obe15, PR89, PM91, PTV20, Per99, Pet00, PP92, Pow94, Pul86, Rab94, RGS25, RC18, RV22, RMM12, RTH23, RN22, Ric08, Ril92, RBT15, SW95a, SST04, SS08a, SA90, SKAW12, SR88a, SGS20, SDG20, Sim91, SK91, SMA01, Sub04, SWR11, TBRBM20, Tho85, TK05, Tsy98, Usm97, Vab22, Van92, VBD93, VCN20, VNC21, VN21, WPN^{+}24, WB92a, Won08, WL24, Wu09, XL09a, YPD21, ZG21, ZY25, ZYH23, ZML^{+}12, ZJ19c, ZAB15, Zou10, dH95, dDF^{+}94, dPT96, dlHV13, vdHVW01, vSW90, vSK97, vvdV97]. solution-flux [KL21]. Solutions [FF06, FFMZ13, Nap16, AGZD22, AKM^{+}22, Ale11, AMV17, ASS21, AAM03, AB24, Arc06, BSGU94, Bac17a, BP06a, BPN24, BM05, BM01, BSV09, Bog20, Bok03, BDKM92, DSM22, BCFR25, Boz11, BCK22, BPTT15, Cah92, CHPV09, CDH24, CZY08, CHS19, CWHF19, CLX21, CCST22, CLP15, Chi93, Cho13, CPR93, DvHM19, DDS89, EK96, ESEKZ10, Enr06, EHV24, FW08, FHM^{+}02, FJS99, Fuh01, FL01b, GANT02, GJR03, GNAS^{+}20, HJ05, HdSRI17, Hor02, Hu99, HT00, Hua19, IV16, JR00, JRS20, JCY25, JR02, KX91, KJL12, KME20, KM25, KO92, KFR25, KS04, KK86, KP15, LW25, Lan95, DBH25, LLHC17, LWCH19, LZJ21, LSL11, LS07a, LS24a, LCZ25, LJ20b, LRE04, Lyn92, Ma03, Ma24, MDRR11, MD20a, MN23, ML16, Mat86, MPMD21, NAF24, Olv92, PB21, Pat98, PK21, PV24]. solutions [PVM22, PM14, PWX24, RNG22, RY13, Ril92, RTT01, SS25, SC11, Saz22, SNOK21, SvdHN86, STS00, Tem15, ÜSHT03, UHLZ24, VO00b, WTB24, Wan01, Wan07b, Wan09, WHW21, WY22, Wu03, WMC09, XF06, YK04a, YLY19, YLW20a, Yos00, YEZ25, Zak19, Zak20, ZL18b, ZLHW19, ZLCH20, ZPT92, ZLG15, dVA02]. Solvability [HJX^{+}19, COPPS25, Sch98]. solvable [CL24a, LMY18]. solve [ABZ21, AMCM08, AAL^{+}26, AM10b, Chn17, DG10, Hua98, IB24, Sch95a, SG09, SL08, TH23, Yüz22, Zha96]. solved [Bic16, BCC16, Bic21, BDN^{+}97, De 93b, FHK05]. solvents [Per03]. solver [AA20, AJW23, BRSD91, Beh97, BT02, BV94, BM18, BJM01, Cai09, CCD^{+}20, GBDB97, HY02, IT07, KDK17, LZZ25, LYZF25, LCJQ12, LYOI99, Pau92, PM03, RG05, RZ15, Sch16a, SH02, SS21, Ter22, TS06, VG04, Wal00b, WSC21]. solvers [AGJCP25, CH15, ÇY22, CDG19, Eij95, Fac03, FPS15, HKZ08, HZD21, IKM23, KPR06, KCB02, LWD^{+}09, LAZ20, LTT19, MZN21, MP05, MR92, Nak05, Pea16, SST04, SA00, Sim98, WP99, YG95]. Solving [AKM^{+}21, AS13, AD18b, BKM19, BNH01, BGG^{+}20, Ben02, BW96a, BCT16, CSS87, DS07b, DS03, EEAS25, EW08, HS19a, HY01, IM98, JL21, Koz94, LWZ22, LW92b, LW18b, MDA24, Pis22, RS22, SRK21, SAA20, ST01, Sha05, SY07, SSA^{+}22, YL13, Yu99, Yu08, Zhe19, ebKMZ24, AA05, AD20a, AD20c, AS21, ABH22, AZA22, AGLRS23, AAL21, AK09, AÁ21, All24, AT93, AC08, AMV03, APJ09, AD19b, AEN22, BW23b, Bai25a, BLW07, BLS^{+}17, BES18, BBJ25, BBBK22, BMGGG12, BGGG13, BM04b, BBC^{+}25, Bog12, Bog16, BRBB18, BS93, BM09, BCSH16, BC89b, Bur93b, Cao98b, CNA23, CL07, CHH15, CXZ17, CDW19, CZ19, CYYH21, CW22, CGPT19, CMCGTR02, Cul95, DT15, DA16, DSA20, DS20, DS21d, DGE22, DDF24, DHWL22, DSAB20, DSS20, DS15, DCY20, DZW24, DP21, EAS12]. solving [EBEHAS24, EHM01, ER18, EG88, ELLE02, EMMK01, EVO04, EE20, EH09, FK23, FJS25, Fdi96, Fer96, FG09, FV87, FFQ09, GC25, GD21, GKA17, GNX19, GFPG18, GD23b, Guo26, HS21a, HHAA22, HA21, HVY91, HAA21, HW22, HR96, Hua20, HT20, Hus20, IAH25, IMC22, IVA93, IR22, JMU^{+}25, JL17, JLZ25, JCJP21, Kan89, KOR18, KSHB21, KSMMM16, KS07, KSSS16, KFOF02, KPR12b, KS09b, KW20, KGZ24, KAS17, Kür23, Lee94, LYF17, LLV18, LCS19, LCW20, LTDY24, Li25, LLY24, LT01, LYK17, LLZ19, LW20b, LL21, LLZ^{+}22, LP97, Lot19, LL02, LDH^{+}24, MCS16, Mar08, MP98, Meh08, MOU14, Mit24b, MFAD23, MD20c, MAD23, MKJ23, Mon21, MN08, MK19, NLS18, NLS20, NAD25, NH24, ORT24, OXW25, PM05, PR90, PV93, Phi91, PPC00, PT95, Pou00, PB10, PRS23]. solving [PG02, QM10, QDA^{+}26, RR21, RZ00, ROB17, RE19, RZ18, Rha99, RS08b, Rou20a, RT95, SSW20, SSS^{+}23, SHL19, Sal89, Sch09, SEGV02, Set24, SOB20, SD11, SD13b, SWW17, SD22a, SD22b, SD25, SJ18, SVB17, ST20, SS16, SS13a, SC22, TQY24, TLGC22, TLG20, TOD11, TGV22, Tsa91, Tsa92, Tsy96, WZL13, WZ16, WW19, jWS20, WG23b, WL25, XC85, XWC25, YXN21, YZG23, YR92, YXB95, YW25, YTZZ18, YRV21b, ZH20, ZNK02, ZD20, ZG20, ZD21, ZWFX22, ZHL22, ZFS24, ZZW97, ZLX19, ZSJ04, ZLS20, ZSQ21, Zla85a, ZB19b, dOF20, Ise02]. Some [AR23, BN99, BSFDM02, BV96, BG14, BC89c, BW96b, Car09b, CH90, Cve02, DN24, Fdi97b, Goo90, Gu01, HOEC86, HT19, HWY20, JN02, LLY11, LZW20, Mac86, MM14, NY13, Nov08, QNA23, RMH20, SMTHE22b, SMTHE22a, SLZ10, The17, TM05, Vul92, WKN20, WSS97, AB09b, AGKK94, AFLG^{+}12, Aze22, BY22, Ban97, BFQ22, BNKR20, BR01, Bre85, BDRZ04, BT93b, Cai15, CC25, CH87, DS07c, EHM01, FRRZ25, GH91, GM95, HJS97, HS98, Hor93, JMDN^{+}22, JP08a, JJ15, JM05, JT09, JV09, Kau93, KPR06, KAS17, Kza92, MK21, MRF00, Mar03, MAD23, ORT24, PSP05, PSW02, PT95, Rha99, Saz22, Sid14, WB90, Won08, Woź10, dSFDG20]. Sommerfeld [IV16]. SOR [DO95, DO98, DSS20, PPS10]. SOR-like [DSS20]. Sound [Bir87, KJL12, LT05, NU15]. sound-soft [KJL12]. Source [AX19, ASZ15, BM12b, BFH09, CCD^{+}20, Cho13, DCSB25, HP14, HM15, HM17, HILK13, JRS20, JCL18, KHM^{+}14, KRBK16, LWD^{+}09, LDP^{+}14, LN24, OAHN22, PK23, RC18, RSR23, Saz24, SJ18, SZQH23, SS17, TS23, TDMT21, WZW13, WW14, YLY19, YD22]. source-type [CCD^{+}20, DCSB25]. sources [BB25, CL20, FOMC05, KV07, RTT01, SK10, YK04b]. Space [AGJCP25, BC01, BJTZ20, DGRS09, KL23a, SLW17, SvdVvD06, vBvdZdB08, AD20c, AD21, AS20a, AJT19, AKBF19, ALZ^{+}21, AM10a, AL22, ABM17, BS00a, BLM17a, BZ17a, BK21b, Bou02, BFLR23, BRZ17, BR97, BMWH20, BD11, CC25, CS94, CK22, CDW13, CLTA18, CWY20, CG21, CT21, DC21, Dah02, DA18a, Din19, DCY20, DMA22, ELCWS98, EK97, EL94, FMS18, FD16, FWHM20, FM95, Fra04b, FWW^{+}21, FSWZ19, Gar03, Gea93, GNAS^{+}20, GAOB20, HZ20, HZD21, HMD21, HQAZ24, HZAT21, HS21b, HST14, HL89, IV16, JRS20, Jéz99, JWZ21, JHGZ20, KK20a, KZ13, KP19, KGZ24, LSV22, LL98, Lee23, LS25, LS10, LCS19, LQS21, LS21, LYK17, LCLW17, LZCF21, LYZW22, LXCM21, MZN21, MD22, Man97, MT06, Mit24b, MD96, MPMD21, NTHC21, NLHT25, PKP19, QXQ22, QNA23, SAA20, SS99, SHA12, SK22, SW07]. space [SC20, SSKS21, SW12, Str98a, SWCH15, TDMT21, Tou25, ÜSHT03, UHUL21, VL19, WZW13, WMLB19, Wan21, WSS97, WYY20, WG23b, XWW19, XZL19, XLZ20, XWZ21, XF22, XWX21, YJZ18, YLFT20, YQCZ22, Yu99, YDWW17, YWSL20, YLW21, ZM19, ZD20, ZLW22, ZW19a, ZZX20, ZLG24, ZZ20, ZLWF21, vdVS08, ASZ15]. space-dependent [CT21, WZW13]. space-dimensional [ELCWS98]. space-dimensions [BD11]. space-filling [LS10]. space-fractional [AD20c, AKBF19, BK21b, CWY20, CG21, HMD21, JRS20, JHGZ20, KK20a, LZCF21, LXCM21, MT06, WMLB19, XF22, XWX21, YLFT20, ZLWF21]. space-like [Man97]. Space-time [KL23a, SLW17, ALZ^{+}21, CK22, FWW^{+}21, KGZ24, LSV22, Mit24b, MPMD21, NLHT25, Tou25, WG23b, YLW21, ZLG24, ASZ15]. space/multi [DA18a]. space/multi-time [DA18a]. Space/time [vBvdZdB08]. spacecraft [CRTU15]. spaced [CHS17]. spaces [Abr93, AHAS21, All24, AMCM09, AM16a, COPPS25, CCOVF22, Che12b, CP03b, DE16, DP21, GIS23, GP00, GH21, GGG16, JL94, LRS23, Li16, MD23b, PGA93, VA21, Xu23, ZBD24]. spaces-based [GIS23]. spacetime [ALP^{+}96, ZH21]. spacewise [HP14, HM15, HM17]. spacewise-dependent [HM15]. Sparse [BG02b, CS09, Che12b, Cos25, DJR25, JMPY10, Xu23, Zha00, AB07, Ara99, BW23b, BT99, CCY22, DNW18, GC25, HVY91, HZZ24, Jia00, Kam25, KM21, LP05, LKV01, LGH11, Mag91, Mar08, Muo23, PTW19, PV93, SW95a, ST09b, Vas17, WZ02, YR92]. sparse-grid [LKV01]. sparse-sparse [ST09b]. sparsity [Huc99]. Spatial [FPPS00, SUP^{+}12, AKS21, BCJ97, Boy91a, CCZ22, DEPS15, ECHF^{+}20, HZBM05, HPH20, HJ03, JR18, JEG10, JM94, LGS21, LE94, Rog19, SZQH23, ZHL22]. spatial-temporal [ECHF^{+}20]. spatial-time [AKS21]. spatially [EV96, LSGK15, WKM04]. spatio [Won08]. spatio-temporally [Won08]. SPD [ST09b]. SPDE [KZ13]. SPDEs [LT19, MT20, ZWK15]. Special [ADG^{+}16, BGHR12, BGH^{+}15, CHM09, CD95, DLN^{+}24, DGCW17, FFMZ13, FDF^{+}25, HKNV16, Jbi25, KP07, SSV97, BBRS97, CJLS98, GR02, JQYM23, LPZ00, LST07, LW17, RT95, SG05, Tsy96, WB90, YC13]. species [CCK03, HH22, HJZ23]. specifications [Deh05]. Speckle [CDI^{+}24]. speckled [TSB10]. Spectra [BLL24, DSW96, LCVG01]. Spectral [AMC02, AQ00, BR25, BN12, Bla00, Bou16, BTDV10, BCK22, Cai24, DY17, DvHM19, DA17, EJRR23, FD16, Fun94, GD09, GKKM21, Gia12, Gu19, Gu20, yGyZ07, Hes00, HKP89, JZK06, JLWY25, Kar89, KS22, Luo18, SJ11, SM89, Tru00, YLS^{+}09, ZZ19a, ZCGS21, AZHD23, AY21, AAL21, AAD^{+}08, AyLqW18, AA22, An16, ASZ18, ALZ^{+}21, BK06, BM89, BMS89, BSQ96, BBD08, BP95, CA21, CGMS21, CDP19, CG92, CjW18, CG89, CSXL14, CZHX19, CLX21, CG21, CNT07, CR19, CSX23, CSCM96, DSA20, DM09b, DL22b, DSAB20, DB08, ER18, EK95, FID18a, FID18b, FWL18, FSX25, FMPP24, FPPS00, GANT02, GT00, GGM07, GM17, GH07, GWZ25, HGP11, HZ21, HZ20, HQAZ24, HS17, HCGW22, HY24b, Ito17, Ito22, JZS20, JLZ20, KL23a, KSHB21, KHA12, KWLK00, Kop86, Kop89, Kür23, Kwa09]. spectral [Lay09, LWL18, LZ18, LMQZ18, LLJY20, Li22, LCH20, LT01, LLZ19, LY24a, LXPY25, LR00, LJ20a, LBCN00, MZZ17, MAHZ21, MS86, MD00, MCS16, MMRV20, MM02b, MM07, Min04, MSS21, MD23a, MH16b, Mun00, MKS12, NH24, OL18, OGS20, PL25, Pan21, Pas91, PP24, PM91, Pav00, PD01, PK91, PP00, Pło22, QM20, RTW21, RX08, SD11, SC25a, SW07, tSqWyG16, SA00, SJ18, SVB17, gTpM07, TW00, Tse00, TC19, UHLZ24, VO00a, VS91, jWqW09, jW15, WMLB19, WDH20, WT20, jWS20, WZZ21, WCL22, Wan23, WC14, WK00, WG23b, YMD21, YB10, YJZ18, YT21, YS24a, YLL21, YWW23, YXX24, YF24, YZH24, Yu08, hYqW12, YWSL20, YLW21, YW24, Yua20, ZDM18, Zak20, ZH21, ZAED21, ZLW20a, ZLW22, ZL11b, ZZX20, ZEW20, ZZJ21, dFN00]. spectral-collocation [HY24b]. spectral-element [CSXL14, LY24a]. Spectral-Galerkin [Cai24, JLWY25, ASZ18, DB08, Kwa09, YWW23]. spectral/hp [KWLK00, SA00]. Spectral/Rosenbrock [AMC02]. Spectrally [Sid23]. spectrum [FZM20, GCP91, KQ13a, KQ13b, LMS08, TCCW89, XC85]. speed [FdSB02, MS91, SM13, TSFB01]. Speeding [NLLG20]. SPH [FPRA09]. sphere [CNT07, CRSF19, TOCV02, TLSS09, WZL13, ZL11a]. spheres [MD20c]. Spherical [Har00, AX20, Bar12, GH07, LMS09, Mir20, Pet00, RA17, STS00]. spherically [YXN21]. spheroidal [BDSG09, FRRJT10, Sch23]. spheroidal-shaped [BDSG09]. spherulites [HPW21]. spin [AB17, Bec02]. spin-Heisenberg [Bec02]. spinning [CS04]. Spline [PTV16, ABJ12, ADSS17, ABY22, AR24, BS14a, BY22, BPN24, BES18, Bas21, BGS06, Beh93, BH12a, DLPV17, DMPSC16, DCL23, DT89, EBEHAS24, EHM01, FMS24, GS99a, GG19, IO18, Jam95, Joh05, KO08, KOS20, LMS09, LFB00, LZW17, LZCF21, MB10, MST07, MST09, NLZB23, PRS23, QWX20, RL86, RK08, RTA19, Rou20a, SST12, SRK22, SE93, SST09, SST15, TLG20, VR01, WE99, YC00]. spline-based [DT89, Jam95]. Splines [Sto96, AJK20, Bar12, BBCR22, Bic16, BCC16, Bic21, BCV25, CGMS21, DS17, DF92, FGPR12, GV18, GS25a, GS09, KP01, KP03a, KPR12a, PR09, PLB22, Rab94, Ram12, SA90, SYG^{+}05, Spe12, Ver93, WT93, WW05, RMS17]. Split [CHZ14, RKVZ15, RU25, WL09a, WMLB19, WG22, AHAS21, ABR23, BG11b, CS01, DFLM19, Hor98, IHS13, KK20a, KM16, MP11, RTH23, Roo20, YXT17, ZYX20, LCM24]. split-Hamiltonian [CS01]. split-operator [IHS13]. Split-step [CHZ14, RKVZ15, RU25, WL09a, WMLB19, WG22, DFLM19, KK20a, MP11, YXT17, ZYX20, LCM24]. Splitting [BDF^{+}25, BCT19, CMP20, GPHA16, LL02, MV18, SK96, Vab21, ZHL22, AZA22, AMT13, An20, ABF09, BC05, BCF^{+}13, BCET22, BC04b, Bi20, BFH09, Bou25, Bou02, BC23, BSTT22, CS19, CSXL14, CL24a, CG21, CJ23, CT21, DS21d, ELCWS98, FLS94, GK09, GV02, Gla94, GPHA25, GGRN17, HPH20, HV22, HT19, HD23, HLL09, HV95, Hun02, JK17, KS00, KM21, KL23b, Kie15, Kni94, Kni95, KNT13, KW10, LHH96, LLV18, LS23, LFS21, LYLL23, Lte24, LXCM21, Lub04, MH16a, MZZ17, MG97, Mit24a, MM25, Ost02, OS12, Poh93, RA03, SK16, SB14, SD22b, SCT05, WT20, WS22, WaZ24, WG18, XZL19, YJJ^{+}24, YP18a, YP18b, ZZ18, ZWH^{+}17, ZSG^{+}20, ZM25, ZZX19b, ZZLL21]. splitting-like [DS21d]. Splitting-methods [GPHA16]. splittings [BWY03, BM09]. sponge [Pet00]. spread [AJT19]. Spurious [Mul99, VT91, Wai98, CM06, ZLHW19]. SQP [BdFPSdSC08, HP15, SXP09]. Square [KK22b, Abu04, AL24, AD01b, BB94, BM04b, BS12, CHM22, CKL03, JLZ20, KK20b, Li11, Li16, MDD14, PL20, RU25, SD13a, SHLY19, TZA13, Waa88, WM07, WC11, ZH24, Zup03, dB03, dVA02]. square-based [BM04b]. Square-root [KK22b, CHM22, KK20b, Waa88]. Squares [CF86, AN15, AD19b, BLS^{+}17, Ben02, BIW24, CCL22, CP07, CZ04, CCLT10, CC20b, yDqGnJT09, DZW24, HP85, Han87, HS17, Hua19, KN19, KS02, KLY05, KS09a, KR10, LG21, LV12, LS25, LRRÁY25, LSY17, LW18a, LJYS20, LCZ21, LLW22, MOS12, MS08a, MSS21, MD23a, MD20c, MD21, MD23b, MH16b, Mon09, Mou03, NZY21, PS02, PS03, PR22, Ren13, SLJ86, ST25, SWCH15, WM22, WQS^{+}24, XZT21, YK04a, Yua93, ZG20, ZSJ04]. squares-based [LG21]. squares-total [LLW22]. squaring [CH04, SW09b]. squaring-down [CH04]. SSC [ABCC18]. SSOR [BD85, SH21a, Zha21a]. SSOR-like [SH21a, Zha21a]. SSP [GK19]. ST [SY03]. stabilisation [RÁM23]. Stability [Arn93, BF92a, BT97a, BLM17a, Bor02, BSvdV99, BT93b, BJ01, BJ06, CGGGS11, Car23, CP05a, CCZ22, DLQZ23, DMA22, FSWZ19, FL01b, GM85, GM95, GAZD25, Gol00, Gul15, Guo01, HMT03a, HMT03b, Jac93, KKR15, LT12, LL15, LRC19, LR20a, LSL11, LS21, LLD18, LLM19, Mac86, MB08, MS90, MV17, NTHC21, OZ96, Ost02, OTK04, PGA93, RGÖS18, SSV89, Sch93, SG00, Ske89b, ST25, SHG86, SN04, TH18, WHL19, Zha19b, ZLW20a, ZZO16, ZZjQ24, Zió99, iW07, iV09, iM13, vdHVW01, AH11, AJ19, AHJ^{+}23, Abu04, AHA23, AMCM09, AL24, ABF09, BG11a, BC89a, BS21, BP12a, BGT97, Ben17, Ber15, BM12b, BN03, BS08, BS12, Bur85, But09, CCG13, Cha96, CZ12, Dav98, DDZK05, DvHM19, DR09a, DA17, DSW96, DLV24, EE20, FP02, FS08, FG09, FHV97]. stability [GD09, Gje07, Gu01, GGG16, GLMY17, GT18, HM00, HS95, Hin95, Hor93, HA16, HDS20, Hua09, HV89, Hun02, Ise97, IJ21, JJ94, JS25, KS25, KMG09, KS89, KW93, LZQ22, LMPS19, LZ13, LZL14, LW20b, LL20b, LLZ^{+}22, LZY24, LT93, LS05, LRE04, MS91, MHA19, MG97, MZ04, MM07, MAF20, MSA20, MAH22, NMSF94, NBNTGV11, NB01, NFAE03, Ort20, Pot97, PYD21, QR03, RU25, RFMV24, RGMO19, Sam94, SA21, Sch12, Sch98, SAH24, Shi25, Sid10, SW13, Sou09, Spi93, Spi97, TZ21, Tan01, TZA13, Van00, Vul95, WZL08, WC11, WW19, WWL21, WWZS25, ZZL01, ZL23, ZFX17, ZYX20, ZZ19b, ZP97, ZP98, ZL24, dPT96, in 92, in 96, iW09, MFAD23]. Stabilization [ABFV09, De 88, YXF25, AMR14, Arn98, BPS19, ÇK13, GG22, HJR22, KL09, LFQH21, RG05, Tia15, Tob14, XF06]. Stabilized [AMT13, AKL08, BL08, GNX19, LW95, MF23, Pag25, PAP17, SC19, SA12a, SED21, AD19a, AD20b, AS13, ABR05, ACP23, BC01, BH12b, CK20, CCZ22, Cod08, GM10, GHH20, HL08, HFL12, JZXJ21, Kam16, KTD20, LLHC18, PSWZ21, Pic05, QMLC15, QM19, RSK24, Sch93, SGS00, WZ19, WaZW21, WaZW23, ZLSZ22, ZS21a, ZS21b]. stabilizer [ATW20a, ATW20b]. stabilizing [Hun02]. Stable [CGA93, DDHS97, GÖ20, KAS22, Ria22, SGS00, SGN06, WK00, ACP24, AJ24a, AGK24, AG25, Aso21, Ben96, Bok03, BMWH20, CCD^{+}20, Cao07, CDD^{+}17, CGRT18, CXZ15, CRSF19, DIJ12, DGE22, DSK12, DCY20, DLM02, DK14, Elg17, ELLE02, FL04, GLML20, GHS25, GHK16, GAW09, GKS20, GPHA06, GPHAM12, GGT24, HH10a, HWZ22, IRC12, IJ17b, JY23, JJL^{+}24, KWY24, KLSW06, LCK22, LK14, LMY18, LWW22, LQXK23, LH24, Lin10, LC21, MW93, PHY19, QH22, Qi24, QZH25, Ran20, ST89, SK16, SL21, Ske89a, Sv95, SB19, SCvdH92, SS09, SWJ^{+}24, TQA^{+}25, TOD11, VDVV98, VK17, VZ93, Wan20, WWZJ22, WLY24, WLZ25, WWZ25, XXYZ24, XK25, YZH19a, Yan23, YXX24, ZCY20, ZFS24, ZZW97, ZH24]. Stage [HR06, AFIS24, BJ11, But93, CPP02, Cer25, CCS17a, Chi93, FJ17, GPMR95, JM17, LW19b, LW22, Lua17, NBNTGV11, Tor06, UHLZ24, Ver96a, Ver06, Wen98, WBCK02]. stage-order [Ver06]. stages [CMRdlT24, Kie95, Pag25]. Staggered [GLV06, Pou00, ASCM02, BD22, CS08, FD16, ID19, JTB15, LR20a, ZLL22]. stagnation [Meu14]. standard [Bra00, DTGN23, FM11, KDAK13, MDRR11, MZS10, RGMO19, Vul95, WT17]. standing [LC19]. star [MM20a, Pel20, SW24]. star-shaped [Pel20, SW24]. stars [IN89]. started [TOCV02]. starter [vB95]. Starting [CIJ17, Lab98, Lab99, Sha87, GPMPR03, JM25, NH15, Ver06]. state [ABRW18, BSZ15, Cah92, CM02, CZHX19, CML05, CPC25, Chi93, CY05, CST97, Dah02, EK97, FNT06, HHT97, HK09, ITZ17, KK86, KK20b, Li25, lLNW21, LRE04, MZN21, NT92, PKP19, PS21, RY13, VVR08, WCM23, ZSQ21]. state-dependent [ABRW18, Cah92, CST97, HHT97, NT92, RY13, WCM23]. states [Bec02, LLM19, SAG86]. static [AM10b, Cop03, CA16, DT89, GLV06, GH25, HB20, RR14, TV91, YTZZ18]. static-regridding [TV91]. Stationary [AAM03, Kin94, AR15, BM18, Buc17, Cai24, CGG25, CC04a, CW22, DNW18, DN08, EH06, HL08, HFL12, KBK21, LY10, LJ20a, Maj17a, QH19, QAMX17, QM19, SH10, SFZ21, TMM15, Vab21, WTB24, XY25, YXF25, YLX21, YÇ16, YHT23]. Statistical [GPíP03, Ber85]. Steady [CM02, RV05a, BHJ13, CPC25, Chi93, CY05, DS97a, FLH22, GM10, HS86, HT00, ITZ17, Jes85, KK86, Li25, lLNW21, LRE04, SAG86, Wan25, ZZZX26, ZS21b, ZSQ21]. steady-state [CPC25, ITZ17, Li25, lLNW21, ZSQ21]. steel [IPL02, PLI03]. steep [SB18]. steepest [Bai25b]. Stefan [LL98, Ros93]. Stein [BJ11]. Steiner [PSP04b]. Steklov [AP08, BLY16, LB23, XYHM20, YLL09]. stellar [Vas17]. stencil [Tow16]. stenoses [LT00]. stenotic [TYKK01a, TYKK01b]. Step [De 88, Jac87, AHJ^{+}23, AM99, AGM95, AFS02, APJ10, AN22, BW23b, BJ05, Ben17, Bho12, Bok03, Buc06, BJ01, BJ03, BJ06, BT97d, BC00a, CLMSS98, CMRV11, CMRdlT24, CHZ14, CCP17, CP05a, CHK99, CJLS98, Con01, CX08, CDP12, DIJ12, DP12, DFLM19, Ein18, EK97, FH20, FH22, FJL21, GH91, GLPW09, GPPR12, GT18, Hai97, HOS99, HS09a, Hoa15, HLR01, Hor05, HP15, IHS13, IS22, KP18, KM19, KK20a, Kie95, KW20, KDKW20, Leo10a, LW17, LWW23, LGC25a, LWWZ25, lLX22, MP11, MS24, MSGM23, MG22, MJS23, OMSS25, PSW02, PWS05, PWS06, PJB04, PAJ12, PMP23, QAMX17, RKVZ15, RU25, SMEN04, SM93, SWJ09, Sha85b, SAH24, ST20, Söd06, SS09, TLG20, TYJ11, UHLZ24, Ver06, Wai98, WL09a, WMLB19, WZ22, WWM22, WSP04, WG22, XZZL15, YXT17]. step [ZQY18, ZYQS23, ZZF25, ZYX20, ZZjQ24, ZX09, ZP12, Zuu95, van93, vdSvdH95, AD04, LCM24]. step-by-step [van93]. step-parallel [vdSvdH95]. step-size [AFS02, HLR01, KP18, MJS23, PAJ12, WWM22]. Stephan [CHM09]. stepping [Aca12, AKBF19, BBRBS09, BSvdV99, HLT07, HJ03, KR15, LHC09, LYK17, LFS15, QCW^{+}23, Qiu23, SUP^{+}12, SZ22a, Shi25, uIVS13, WTY21, WWF20, XY19, YPD21, YLLZ21]. steps [CHLA21, LLT07, SI20, Yu99]. Stepsize [FS05, BG02a, BC98, BC00a, Con01, GC15, JVZ95, JAH21, KW20, LMG02, MBS23, MP94, MDP23, SH97, Spi13, WKP12, Zla85b]. stepsize-coefficients [Spi13]. stepsizes [CGA96]. stepwise [LLT20a]. Stewart [SMC08]. Stiefel [HWY20, Wan23]. Stieltjes [DMGVO05, Gil10, Not92]. Stiff [Kie15, AH15, ABH22, ACP24, Ale03, Aro96, BJ02, BF92b, Bos09, BDM03, BC89b, BC95, BS00b, BR05, BP06b, CL85, Cas96, CS24, FG09, GPMPR03, GPHAM12, GPHAPPR23, HCX03, HS09b, Jac02, Jac96, KC19a, KSK^{+}25, KW20, MP94, NS20, Pag25, PSW02, RKVZ15, Sal89, SSV89, Sch93, SM93, SWE05, Sim98, TB01, VS95, WL09a, WAV12, WSP97, ZX09, in 95]. stiffly [HHR12]. stiffness [Sha85a]. stimulate [Var92]. Stirling [SG04]. Stochastic [AHO16, AH17, Bis11, ÇY22, FV85, FS23b, LT19, MA04, ZJ19b, AA04, AA05, AZA22, ABZ21, AB17, Abu04, ACLM22, Ant23, Ant25, AL24, Bac18, BTBR20, BF17, Bok03, BC23, Buc06, BS12, BB96, BB98, CHM22, CGH23, CGEV19, CHZ14, CDP19, Cer25, CCZZ18, CGW20, CC23b, CWC24, CF13b, CL18, CD20a, CN11, DS21a, DMS23, DR09b, DK11, DFLM19, DZMB21, Fan11, GHHG22, Har10, HXW15, Hou23, HJ21, HLY22, JK21, JCY25, KGR08, KK11, Kha21, KS08, Kom07, KK17, KK20b, KKR15, LZQ22, Lav94, LSK12, LLY21, LWaZ24, LDC24, LLD18, LMTW20, LMW23, LCZ23, LSW23, LYMD24, LQY25, MDD14, MH04, MSS21, Moe98, MJS23, MVBS25, MEGW23, OMSS25, PKP19, Pot97, PM14, PWX24, RP17, RT20, RKVZ15, RU25, SHL19, Sca22, SKO19, Sob24, SB19, TN16, TLG20]. stochastic [TB01, TZA13, WL09a, WG10, WC11, WCW14, WGW15, WYY20, XZ19, YXT17, YXZL24, YBW20, Zha20a, Zha20b, ZYX20, ZJ19c, ZAB15]. Stochastically [Bok03]. Stoke [KTD20, MH16b]. Stokes [KM16, Kni95, LH23, LL24, TLP18a, WaZW21, DZ12a, AD20b, AH09, ASGGRGW25, AMN24, ASS21, AR18, ACP23, BC12, BM13, BLRGVR23, BB15, BG11b, BCGS24, BSZ22, BC01, BDFF23, BS97b, BP90, BP95, CCOVF22, CHLA21, CKPS15, Cau08, CGRT18, CHOR19, CMP03, Che96, CH07, CSXL14, CHH15, CXZ17, CLY19, Chi21, CKK10, CL18, CD13, CGS20, DY17, DJ10, DS97a, DN08, Dob05, DYZ20, DJL04, EJS11, FD16, FRRJT10, FLH22, FMP04, GMZ08, GM10, GHS25, GG22, GNX19, GGG16, Guo00, GH07, Guo15, HJR22, HLZ14, HHC08, HL08, HS17, HW15, HL19, HK85, HFL12, HSF25, ITZ17, Joh01, Kal96, KS00, KL23a, KCL00, KS02, Kni94, KDK17, KN93, Kwe00, Kwe01, Kwe03, LLL08, LA11, LLHC18, LRC19, LH20, LLY21, LFQH21, LS23, LYZJ23, LH23, LD25, LD21, LS12, LHX20, LN21, LXPY25]. Stokes [LR00, LO96, LJ20a, MM22, Med96, MWYZ18, NX22, Nor97, OK98, Pas91, Pea16, PK91, Pic05, Pou00, QR03, QH19, QCW^{+}23, QMLC15, QAMX17, QM19, RZS21, RSK24, RV04, RV05a, RK91, SSZ16, SGS00, SZ22a, SH10, She96, SA00, Shi25, Shy86, SLW24, SW20b, ST86, TLP18b, TH18, TLV92, Ton04, TC19, UNGD08, VG04, VO00a, Wan25, WY02, WWM22, WT17, Wu09, XZ22, XLZ23, YS09, Yan22, YZ24a, YT25, YZZ25, YZ21, YÇ16, Zha14, ZZHS18, ZL23, ZZ24, ZS21a, ZS21b, Zho18, ZKO^{+}21, ZS18].

Stokes-like [WT17]. Stokes-Temperature [ACP23]. Stokes’/Darcy [DZ12a, LFQH21, QH19]. Stokes/Forchheimer [CD13]. Stokes/Navier [LH23]. Stokes/parabolic [SW20b]. Stopping [FLMR14, LPT16, Spi95]. storage [HMR25, KCL00, KS08]. Störmer [CGA96, LZ22, vdHMdS99]. straight [TT03]. strain [CHX13, HHR12, XLKY19]. Strang [KW10, Zan01, ZR15]. Strang-type [Zan01]. strange [CGPT19]. Strategies [Bie87, BM04c, BO11, GC15, PT19, WSC09]. strategy [AS97, BT97b, BD11, Bür13, CZ90, CSL^{+}24, DF11, DEPS15, DB97, FS15, GGLR09, HR97, HMW05, lLXhLZ21, MCE^{+}09, MP94, PK91, Sch09, SK91, VO00b, ZGF^{+}25, ZD25a]. Stratified [TC03, LTC03, TOCV02]. stratigraphy [BVT14]. Stratonovich [CDR20, ZAB15]. Streakline [Din93]. stream [BP95, LJ20a]. streamfunction [KN93]. streamfunction-vorticity [KN93]. streamline [AN22, Bec18, DLZ21, Gas92, KN93, Par04, YZ19]. streamline-diffusion [AN22, Bec18, YZ19]. strengthened [AK95]. stress [GP00, GS18, KS02, MAG13, NS16]. stretched [KKE16]. stretching [VT91]. strict [NB01, NFAE03]. stride [WJF25, But92]. string [Ahn07, LR03, RL06]. strings [RLHC19]. strip [DDZK05, jW15, ZR15]. strips [DP85]. Stroboscopic [CSSZ25, CCMSS11, CSSZ20]. stroke [van98]. Strong [BZ92, FS08, Hua09, LDC24, LCZ23, MAF20, MT20, MEGW23, NBNTGV11, PM14, QW25, RY13, Sus10, TWL23, YXT17, YBW20, ZYX20, ABY22, BB96, Chn17, CN11, DMGVPO09, Gje07, KMG09, KS07, LW22, MSA20, MAH22, Njå88, NTT22, SA21, WY22, WMC09, YLY19, ZM17, Zha21a]. Strong-order [YBW20]. Strong-stability-preserving [NBNTGV11]. Strongly [VRK25, BBCS05, BBRS97, BCS06, GPHAM12, HCS20, JNPC03, KCB02, LG21, Rou20a, SC22, Wan11, WTF25, XLZ20, YW19]. Structural [BBS11, GEGG^{+}20, BS12, FL05, Leo10a, LS05, MD19a, NBP94]. structurally [BASC17]. Structure [FCW20, KGR08, Shy91a, Shy91b, Udd20, WH18, ABK12, AS20b, AB14, ABM17, Aso21, AC96, BEH24, BSP04, CW25, DL21a, FCW21, GÖS20, GS18, Gus87, HGZW21, Hou23, JQSC22, Jun97, Kok08, LZ24, LC25, LLM19, LZW20, LYMD24, MB20, MR01, PA05, RP01, RDH^{+}12, SW06, WDZS21, WJW19, ZY19, ZWG25, vBvdZdB08]. structure-interaction [vBvdZdB08]. Structure-preserving [FCW20, WH18, AB14, BEH24, CW25, FCW21, HGZW21, Hou23, JQSC22, LZ24, LYMD24, MB20, ZWG25]. Structured [Buc99, MPTV25, MD23a, Ram94, AM95a, ALM04, Aya09, BLD17, BL15, BMR17b, DP12, DMPP99, DGD03, FH08, Hua19, LZY24, Ram96, Rus95, Tol04, YC00, ZHJ14]. Structures [SR88b, BL91, CAD03, DT89, Fai00, FLÖ^{+}97, HKZ08, IPL02, Mar09, PLI03, Vic92]. studies [LBCN00, Ran16, RW87, Str98b]. Study [HKS86, KK23, MFAD23, PCA10, RV04, RV05a, SR88b, AZHD23, Aff94, BH12a, BT99, BdFPSdSC08, BFdS10, Buc04, CL02a, CSSZ25, CGN03, CaAL96, DD21, Den15, DDK19, EGV25, Fai00, GLLY25, GKL07, GOLJ25, HD04, HHL23, HDS20, JUAZ22, LDP^{+}14, MP15, MPG^{+}16, NMSF94, PSP05, PPS05, PYD21, RMCG04, SWFK13, SWW17, SC03, SAMSB20a, SAMSB20b, vR04]. studying [PVH25, SD25]. Sturm [AGM09, Con99, Ghe97, Pru00, VVD95]. Sub [RMM12, ASA20, CQZ20, GG19, ID19, KDH20, RZ18, Roz05, WR20]. sub-cell [ID19]. sub-diffusion [ASA20, CQZ20, GG19, KDH20, RZ18, WR20]. sub-domains [Roz05]. Sub-Scales [RMM12]. subcell [CSCM96]. subcritical [Gla93]. subdiffusion [HLY22, JS24, KHB22, Pło23, WYY20, ZL17]. subdiffusive [ABD16]. subdivision [Boy06, BPV25, CC04a, CDRT19, Con04, CM04, CCS17b, LY10, ZZZ23]. subdivisions [Wal00a, Wal00b]. subdomain [CHLA21, FV99]. subdomains [DCN^{+}19, de 95a]. subgradient [DP21, LWLL24]. subgrid [AK21, CK20, DK20, JR18]. subject [Ang06, Deh05, MVVA09a, MVVA09b]. subjected [MCBV20]. submatrices [DL01]. submatrix [YD07]. subproblem [FW22]. subscales [Cod08]. subsequences [Lin10]. subsidence [SR88a]. subsonic [Kor95]. subspace [ABR23, AB10b, Bai02, BBLT15, EAV16, EEJB22, Gul15, EJR25, Jia00, LWZ22, PKSB10, Pet92, PR12, RSY12, RU21, Sim10, SSX14, SS12, Wal95, WZ16, Wei95, ZGL98, ZFC20]. subspaces [GSR00]. substructuring [GN86, MDT05, MVBS25, O’L87]. subtraction [PA18]. successive [Che16, CWP21, DW00, KP03b, TT20]. Successively [SM85]. sudden [HH10b]. Sufficient [Lin01, CH90, LLL12, Meu14]. Suitable [Has09, PGS10]. Sum [Boy91b, HJP10]. Sum-accelerated [Boy91b]. Summation [Wen10b, ZJ10, BW15, CL88, HNP17, LG87, Lon86, Lon88, Meh22, RGÖS18, Str98b, Woź10]. summation-by-parts [RGÖS18]. sums [VC10]. Sumudu [AAL21]. sup [Che16]. super [ASA20, GHW20, LMW23, LSW23, ZW24]. super-diffusion [ASA20]. super-linear [LMW23, LSW23]. super-linearly [GHW20, ZW24]. Supercloseness [XLHL25, ZL18a, ZL22, FPR12]. supercomputers [Fuj99]. superconducting [GT15]. Superconsistent [Fun04, FFY08].

Superconvergence [Bac17b, Bac17c, Bra00, CZ04, FLR08, Fra14, HCJW24, HL19, LHWF08, LWW20, LSG24, Ris05, SY18, SJ20, Tem15, WY02, Wan20, XZ22, XWM26, Yan21a, AC15, BSGU94, Bac19, BTBR19, Bac21b, Li01a, LY01, LH02, LRC19, LSWM19, LSP20, MKDC25, PL25, SW20a, SZ22b, WSY18, XP23, Yan22, ZY14, ZSS23]. Superconvergent [KM25, KKK25, LW21b, SST15, SW18, ZLY23, ZX22, LR01, MCS06, SL20]. Supercritical [Gar87, Gla93]. superlinear [GT19b]. Superlinearly [CDW13]. superoptimal [NR14]. superposition [FMGN94]. supersaturated [KKL^{+}25, SA08]. supersonic [CRTU15, EH91, Nor99]. supervised [IKS25]. SUPG [Usm97, ZB19b]. SUPG-FEM [Usm97]. support [MCBV20, SSW04, CM02]. supported [Ehr08]. Supraconvergence [Fer96, FPR12, Fer93, FG98]. Supraconvergent [Bar09]. surface [Ari03, BBD20, CWM09, CAAT16, DEPS15, Dav98, DL13, DMQ02, FPRA09, Fat10, Gar05, Hin97, JP17, LY10, LR87, LYOI99, LWCT07, LG02, Maj20, MAD23, NT20, RZ15, Sch16a, XLK07, YC00, YSZG25, HP91]. surfaces [BFGP08, CS04, CW22, CGGM17, FJS25, FGPR12, FR14, GT15, HGR01, HM22, KN19, LY01, NT16, NT20, PR09, Ren14, TMM15, XXYZ24, YXF25]. surfactant [DLQZ23]. surgery [MDHK06]. Surprising [Asc12]. Survey [Sid10, AAI^{+}93, Bre96, But85, BC89c, FSU89, HJ05, Kau97, MN20, Mun00, Tho85]. susceptibility [PVH25]. suspensions [BBCS05, BCS06]. SVD [BS09, KK22b, Lyo12, MB20, MMBB07]. SVD-based [Lyo12]. SVDs [PS02]. SVM [LZ24]. Swartz [LVW21]. SWEs [QW25]. Swift [DGE22, QH22, Qi24, XXYZ24]. swimming [CST18]. swing [CGEV19]. switching [BAD13, DFLM19, GK22, MHL18, RY13, Vul92]. Sylvester [BLW07, DMH18, FG13, Hey10, JP19, RS08b, SMTHE22b, SMTHE22a, WS21]. Symbol [BDFF23, CGMS21]. Symbol-based [CGMS21]. Symbolic [CBHY11, CCS02]. symbols [CLS04]. Symmetric [ADSS17, LRS09, SvdHN86, SZ99, WGKS12, AW14, AX19, AC08, AES13, BM12a, Bai25a, Bai25b, BC99, Bec02, Beg00, CM13, Cao01, Cao07, ÇD17, CZ04, CLLM21, DMGVPO09, DII15, JZS20, JP17, JP93, Kim19, Kru99, KCB02, KW10, LMV17, LDIW16, LX08, Li19, LWZ22, Liu02, Lu98b, Luc05, Mar08, Mat05, MPMD21, PV93, RdAP96, San03, Sel14, SYW22, SD22b, Sim04, SW24, THW19, WZ16, YXN21, YD07, Zan91, ZWFX22, ZFX17]. symmetrization [CG13, GC15]. Symmetrized [BH93, HL02b]. Symmetry [Olv92, VL08, INR01, LY24b, MDRR11, Mat08, Zan01]. Symmetry-preserving [VL08, MDRR11]. Symplectic [HOS99, HHW18, Li19, Ske99, XW19, ZCSH11a, ZCSH11b, Ant23, Ant25, ABD16, BF99, BCF^{+}13, BCET22, BIJ23, CH95b, CL01a, CM00, DM11b, Eir95, Hai97, HL99, HLMKZ06, KCY19, LX08, Mur99a, Rei99, SS94a, SZ97, WW19, WLCZ26, WE99, ZJ19b]. symplectic-preserving [WLCZ26]. symplecticity [IMMS20]. Symposium [CFTW08]. Synthetic [Abr93]. System [TMS87, AAL21, Aff94, AEK23, AA22, AA87, AMP20, AGK24, AG25, BS14a, BBJ25, BS00a, BO04, BLRGVR23, BGG^{+}20, BSV09, BMV19, BEH24, BS24, CCCC^{+}25, CFLW22, CC20a, CYYH21, CKK10, CH04, DL20, DLM20, DN08, DTQ^{+}20, DZMB21, Eij95, EK97, Fac03, FCX06, FvdMS17, FGP23, FL23, GLS09, GLML20, GBDB97, GGRBRG22, GAOB20, GJLL20, HP85, HH22, HCS20, HS17, HAA21, HY24a, HX11, Ise94, IT16, IB24, JJL^{+}24, KME20, KS25, KN08, KLY05, KK09a, KCB02, KKK25, Kwe03, Lee10, DBH25, LMA18, LC19, LL19, LH20, LX21, LAZ20, Liu09, LARGVR23, LB21, LMWZ07, MD20a, MMKN17, MLK06, MO17, MMDH19, MMD20, MMDS21, MT11, Mic03, MKJ23, MdD04, Mur19, NAF24, Nag22, Oji88, PK23, PS09, PVM22, RC18, SSS^{+}23, SAA20, zSW06, TZ00, TK15, VNC21, VMS07]. system [Wan01, Wan09, WS21, WaZW21, WWZJ22, WCJ23, Wu03, WPT19, XY25, XWZ21, YT03, ZdBT03, ZZJ21, ZKO^{+}21, MZ87]. Systems [GM85, ND85, Pet87, AC98, AZHD23, ACKV24, AJ19, Aca12, AW14, Ale03, ADFR18, AH17, AB07, Ant23, Ant25, ADM10, Ara99, Arn95, Arn98, AES13, AGKK94, AEF^{+}14, BM12a, Bai02, BY09, BW23b, BT97a, BHB23, BL05, BCG21, BK06, BMM03, BBRBS09, BGGG13, BBC^{+}25, BM06a, BR94, Bog16, Bog20, BTDV10, BR20, BH93, BS93, BRBM08, BBKS07, BGIW18, BCSH16, BS12, BDM03, BWEP95, BDP96, BB98, BCDP17, CW25, CFCH09, CH95a, CL01a, Cao98b, Cao01, CG92, CM97, CHPV09, CZY08, CDW19, CLLM21, CS24, CBD16, CFC03, CJ23, CJ24, CJ25, CKK25, Con89, DT15, DRVA20, DB97, Dav92, De 06, DB95, DV95b, DL13, DDK19, DKK94, EJS04, FM21, FZA25, FTB97, Fil25, FH08, FZM20, FM95, Fra06, Fuh01, GFB99, GC25, GEGG^{+}20, GP93]. systems [GQ89, Gas92, GKMS09, GKS20, GGM07, Gol00, GBBC^{+}23, GO18, GPHA06, GPHAM12, GOLJ25, Guo96, HL99, HHAA22, HdSRI17, HVY91, Heu00, Hey20a, Hig93b, HNP17, HHW18, Hua00, Jac93, Jac02, Jac96, Jay95, JCN94, KL98, Kał22, KTK20, KPRU20, Kau95, KM21, KG90, Kim14, KW95, Kop86, KFOF02, KPR12b, KKP17, KW10, KK17, KD25, LW93, LT12, Lee23, LZ13, LLV18, LW19b, LLY24, lLNW21, LYZ25, LW92b, LFS21, Liu24, LY24a, LP00, LL02, Luc05, LXLY25, MMP09, MD20b, MKDC25, MP96, Man97, Mar08, MF99, MP98, MCD20, Mär02, MVG14, MM16, MPtM16, MM25, MP94, Moo95b, MD96, Mur98, Mur99b, NNJ23, NS21b, NMKE13, NT92, NRZR12, OFY^{+}23, OG08, OKS10, Ost93, Pan07, PK21, Pea16, PA05, PSW02, PT95, Pul12, RZ00, RGS25, Ram96, RKVZ15]. systems [Rum87, SMJ12, Sal89, SH09, SMEN04, SS99, Sch95a, Sch02, Sch87, SEGV02, SWL20, SWB21, SD22b, SD24a, SD25, Sid90, Sim93, Sim98, Sim10, SS13a, Sof17, SW03, SSX14, Tar98, TQA^{+}25, TDC13, Tro93, Tsa91, Tsa92, TMM15, Udd20, VG04, mWyG00, WL09b, WL09a, WWX13, WJF25, WSP04, WP99, WSW96, Wen98, WB92a, WB92b, Win92, Won08, WWS07, XWW19, YG95, YH18, Yan21b, YR92, Zak20, ZAED21, Zen21, ZZL01, ZNK02, Zha21a, ZWFX22, Zha01, dH95, dG91, in 96, vdES04, vB95, Mur99a]. Szego [BGVHN10, DGV00].

T [AR93]. T-fraction [AR93]. table [BH96]. tackling [SKS^{+}25a]. Tadmor [ASC03]. Tadmor-type [ASC03]. tailored [SSA24, YW19]. tails [ZCGS21]. Taking [SK10]. Tamed [GHW20, LSW23, LDC24]. tangent [LMS08, YLL21]. tangential [WPL16]. tank [Buc04]. tapered [FT06]. target [Lo06]. TASE [GPHAPPR23]. TASE-RK [GPHAPPR23]. tau [AGZD22, DSW96, KO92, QM20, SWW17, YJZ18, MMRV20, Mok17, SOB20, SP22]. tau- [DSW96]. Tau-like [SOB20]. TauToolbox [VGL24]. taxis [GV02]. Taylor [CM06, Dar00, DK11, Güm20, Hig93b, KOR18, MD00, RB12, RO16, Sch98, SH91, TB01, TZA13, YRV21b]. Tchebycheff [HJP10]. Tchebychev [DM97]. tearing [Pec09]. technique [AD20b, AD20a, AKBF19, AT13, AGQ^{+}24, AF04, AD19b, BP90, BT93a, CH22, CNA23, CMT25, CPOGO17, CM04, DA17, DA18a, DSAB20, DT89, ECHF^{+}20, EH08, FHM^{+}02, Fdi97a, Gen10, GG19, GJlL23, GAOB20, HHAA22, HAN23, HMN20, HJP10, HEG16, IR22, KGZ24, LKV01, LWW22, LL23, LZZ25, LN24, LCLW17, LYC24, LR00, LCM22, LJ20b, MS86, MWC21, MD20c, NP21, NSD23, PA18, Rab25, RMK09, RT95, San03, SR09, SM20, SR24, SLZ10, TCCW89, Tsy96, Wan07b, WZ16, XC85, YWH20, YW24, ZY14, ZZPJ23, ZJLA22]. techniques [Bai02, BLW02, BDF^{+}25, BRVC09, BLM17b, BBT25, BWEP95, CDD^{+}17, CH87, CPZ17, DD97, Deh05, DF96, DS02, EW08, Ewi91, FH08, Gar03, GM93, GSR00, GFPG18, GN86, HST14, IM98, KG90, KS10, LD21, LM22b, LCM24, MRF00, MM14, MVVA09a, MPDB24, Mun00, Now96, PSP04a, PCR17, RDH^{+}12, RCR24, SS13a, TC19, WSP97, YH07, Zha00, de 96]. technology [Zar99]. telegraph [ebKMZ24]. telegraphic [MG22]. Tempe [FJ97]. temperature [BN03, PGYF20, ACP23]. temperatures [CF14]. tempered [BO21, ÇD17, Din19, LGZH24, LR20b, RV22, SWB20, SSPZ20, YDWW17, Zak19]. templates [Jor11]. Temporal [Cha17, GAW09, LA21, ZYQS23, ZH24, ECHF^{+}20, LH23, LLT07]. temporally [Won08]. tension [NMSF94, RZ15]. Tensor [AL20, EEJB22, RU21, BNKR20, BS05, CZ19, DHM09, DZW24, EHNR24, Han19, JL21, JL23b, LLY24, RS22, SMTHE22b, SMTHE22a]. tensor-driven [DHM09]. tensor-product [BS05]. Tensorial [BBBK22, ABJ12]. tensors [BKR13, EJRR23, SYW22, XY24]. term [AF23, BFH09, BDDV12, BRS16, Buc17, CHM22, CH95a, Cao98b, CNA23, CdCV03, CD20a, DMS23, DW15, DBDV10, ESEKZ10, FL24b, HZD21, HZCZ23, HZAT21, IKR^{+}22, JRS20, KTYY24, LZ24, LLZ19, LZW20, LSWW22, LLW20, MBS23, Mar09, OAHN22, QWX20, RC18, RSR23, RG05, SA18, SZW19, TDMT21, Wan23, WW24, WL25, YLY19, YQCZ22, YFLX20, YLLZ21]. term-by-term [RG05]. term-structures [Mar09]. Terminal [SWB21, NTHC21, SWB20, SHB^{+}25]. terms [BJ02, BSQ96, BM12b, CY98, GT19a, GH21, IKM23, MFAD23, NYPW21, WXY24, YWSL20, Zen93, ZGF^{+}25, iW07, iW09, iM13]. ternary [KK09a]. terrain [TC03]. test [GM87, JP08a]. Testing [CGCMTR02, JP08a, GGO12, GGO16, PSW02]. Tethered [Lam24]. TetraFreeQ [SV24]. tetrahedra [DFZ16, PPS05, SV24]. Tetrahedra-free [SV24]. Tetrahedral [BS97a, AW03, ASC03, BS96a, MCE^{+}09, PT19, YZ22, Zha09]. tetranichidae [CL02a]. texture [PKP19]. TGF [MDD18]. TGF- [MDD18]. TGM [SL22]. th [AL09, AB15, BP02, EST15, KK23, SA90]. th-order [SA90]. their [AGM95, BLN25, CHH15, CM06, DL20, EH07a, FSX25, GS21, GSW09, HLL09, IMMS20, JS25, KCS07, Lem02, Ma24, MD19b, MP97, NLZB23, PWS06, ROB17, RZ04, Sae14, SSW04, TLQ21, Uty08, WOLY25, WSP04, WWLS08, XWC25, Xu13, ZLHW19]. theorem [AHGM21, BW21, Cao98a, Sza94, Wu03]. theorems [Boy15, Dor01, JT09]. Theoretical [BFdS10, JQYM23, LARGVR23, MMDH19, Zha20b, BT98, BC89b, GGMP88, San89, VBVA22, Wei95, BLRGVR23, BLL24, KPR06, RLMG24]. theories [Alb96]. Theory [CHM09, DGCW17, Gar87, Sam94, WLZ25, XB14, AK00, AVMVMV09, ÁKM20, BWY03, BF99, CA16, CV88, DLN^{+}24, Gan96, HJ06, KO96, KK02, Li05, MDT05, MR94, Men23, Mot17, Ost02, PdV99, Pro24, San89, SG06, Sus10, TM15, WM22, BNH01, HL23, MdR05]. Thermal [DDS89, GD09, HLIS16, HK09]. Thermally [RBC02, DLM20, LD22, ZFS24, ZZZX26]. thermistor [DAH24]. thermo [BGM^{+}09, YR22]. thermo-electromagneto-hydrodynamic [BGM^{+}09]. thermo-poro-elastic [YR22]. thermocapillary [SWJ^{+}24]. thermochronologycal [RMCG04]. thermodynamically [GHK16, SWJ^{+}24]. Thermodynamics [Cim25]. thermoelastic [AC18, BCFQ19, Cop03, CF05]. thermophysical [MCM12]. thermoviscoelastic [Cim25, CA16, YJ23]. thermoviscoelasticity [CF14, CA15]. theta [BF92b, DMS23, HS09b, WG10, WG22, YXT17, ZYX20]. theta-method [HS09b]. thick [CL02b]. Thiele [CJV88]. thin [AJW23, BH12a, BCMV03, CDJT06, DDZK05, yDqGnJT09, KHA12, Par14, SL21, SSW04, YTZZ18]. Think [Ise02]. Third [HMdV03, Sal03, YDWW17, Bac21a, Bos09, DS21b, EJRR23, Fer09, KHYY21, Lev91a, LPR00b, PCL^{+}25, PNA21, RTU15, RCR24, SSC23, SDG20, TS24b, WG23a, YC13, BGHR12]. third-kind [DS21b, SDG20]. third-order [Bac21a, EJRR23, KHYY21, Lev91a, PCL^{+}25, PNA21, RCR24, SSC23, TS24b, WG23a, YC13]. Thomas [Bra00, Kim21, ZB19a]. Three [AW03, LM00, MOZ87, MMP02a, Nak05, Per88, SR88b, TMS87, AGKS25a, AF23, ASCM02, BC12, Bak89, BS94b, BIO24, BWS21, BRS16, Cao98b, CH01, CH25, DTGN23, DW15, Gar03, GGLR09, GPMR95, HD23, HZCZ23, HFL12, KHLV22, KZ13, LXZS22, LZW20, LSWW22, LGZH24, LMWZ07, LYA^{+}19, MZN21, MH16b, MD96, PXHZ20, PGYF20, QXQ22, Ran20, Sch91, SR09, SZE^{+}92, SvdHK94, SK96, Ste97, SK01, Ter22, Wan23, XLZ23, YQCZ22, YFLX20, ZLCH20, ZFX17, ZW19a, ZLWF21, vSK97]. three-coarsening [GGLR09]. Three-Dimensional [Per88, SR88b, TMS87, AW03, LM00, ASCM02, Bak89, BS94b, CH01, LGZH24, LMWZ07, QXQ22, Sch91, SZE^{+}92, SvdHK94, SK96, Ter22, YQCZ22, vSK97]. Three-level [Nak05, LXZS22, PXHZ20]. three-node [DTGN23]. three-operator [HD23]. three-point [KHLV22, Ste97]. three-stage [GPMR95]. three-temperature [PGYF20]. three-term [AF23, Cao98b, DW15, HZCZ23, LZW20, LSWW22, Wan23, YFLX20]. threshold [Hor05, LZY24]. Tikhonov [AR23, BNKR20, Buc17, Fik23, JY20, Neu88, PP24, RSY12, WQS^{+}24, YXN21]. Time [BY00, BP12b, CM14, FG96, INR01, MR20, MMP20, MZ87, Söd06, SvdHK94, ST86, WJM22, WK02, XGQ20, Zuu95, AD21, ALMM98, AGZD22, AACP20, Aca12, ACP24, AGLRS23, AS20a, ASA20, ADG^{+}24, AMK18, AMCM08, AB09a, AKBF19, AB09b, AL17, An20, ALZ^{+}21, ASZ15, AL22, AGQ^{+}24, AF89, AGJCP25, ARS97, AKS21, AEF^{+}14, ÁMS17, BL21, BHJ05, BP14, BKM13, BCS17, BBRBS09, Ben17, BDF89, BK21a, BZ17a, BK21b, BC01, BtTBV87, BP97, BVV09, BSvdV99, Bot97, BDE22, BK12, BR97, BJTZ20, BJ01, BJ03, BJ06, BCDP17, CGP15, CHLA21, CC25, CS94, CP05b, CK22, CDW13, CLX21, CCST22, CL24a, dCCSR03, DS21a, DC21, DD97, Dav98, DDZK05, DA18a, DS21d, DL22b, Dia95, DHL00, Din19, DGRS09, DLQZ23, DY03, DMA22, EBEHAS24, EEAS25, EH88, EL94, FK23, FMS18]. time [FD16, FS19, FWHM20, FXY22, FM95, FJ09, Fra04b, FWW^{+}21, FL01b, GS19, GHK16, GX93, GLPW09, GAZD25, GKS20, GÖS20, Gje07, GTS20, GeO24, GPPR12, GPHA16, GPHA25, GL17, GWLN22, GT18, Hab08, Hai97, HHAA22, Har98, HMN20, HILK13, HLT07, HZ20, HZD21, HQAZ24, HAA21, HLR01, HL21, HR96, HS19b, HAY20, HZAT21, HAC22, HJ03, HST14, HLIS16, Hus20, ITZ17, IKM23, Jac93, Jéz99, JZXJ21, JWZ21, JL17, JL23a, JLZ25, JCSR03, KPRU20, KBS11, KL23a, KK20a, KAS22, KHB22, KDH20, KNT13, KZ13, Kre07, KK17, KK20b, KP19, KZ21, KGZ24, KD25, KR15, LHC09, LKV01, Lay09, LSV22, Lei02, LSVT25, Leo10a, LWD^{+}09, LT07, LR18a, LR18b, LCS19, LSP20, LCHW20, LWW23, LGC25a, LGC25b, LWWZ25, LYK17, LCLW17, LCL18, LLZ19, LHX20, LMTW20, LL21, LY24a, LS24a, Log04, LFS15]. time [Lua17, LO95, LO96, lLX22, LLW20, MZZ17, MM18, MD22, MN23, MG97, MS08b, MM20a, MVG14, MPV24, Mit24b, MMM19, MD20c, MDA24, MD25, MN08, MRFF17, MC21, MPMD21, NTHC21, NLHT25, NC16, NWL^{+}22, Nor97, Odi24, OZHP23, PNA21, Pea16, PVM22, PJB04, PAJ12, PS19, PMP23, PM14, Pul05, QWX20, QH19, QCW^{+}23, QDA^{+}26, Qiu23, QWW24, RMCG04, RZ00, RE19, RZ18, RL21, Ria22, RV09, Rou20b, Rya00, RTT01, SMJ12, SDK24, SSV97, Sar05, Saz22, SA12a, SUP^{+}12, SLW17, SZ22a, SC25a, SW07, SC25b, SY18, SWL20, SC20, Shi25, SJ18, SA20, uIVS13, SS17, SS08b, ST25, SA18, Su94, SvdVvD06, SW17, SZW19, SAMSB20a, TER03, gTpM07, TGB08, TDMT21, Toc01, TTP24, Top21, Tou25, ÜSHT03, UHUL21, WZW13, WSY18, WR20, WDU21, WTY21, WS21, WWL21]. time [WCS21, WCM21, WZ22, WXY24, WW14, WYY20, WWF20, WG23b, WdG92, XZW19, XY19, XL23, YPD21, YX25, YS09, YJZ18, YXN21, Yan22, YT25, YCWH23, YLLZ21, Yu99, YWSL20, YLW21, Zan01, ZL17, Zha19b, ZJ19a, ZYLL20, ZFC20, ZFW20, ZLW22, ZJH^{+}23, ZFS24, ZZH25, ZM25, ZW19a, ZLG24, ZYS17, ZZ20, ZSZZ20, ZXW17, ZL24, tV87, vBvdZdB08, vdVS08, NBNTGV11]. time-accurate [ACP24]. Time-averaging [BP12b]. time-changed [AACP20, LMTW20]. time-delay [GeO24, SMJ12]. Time-Dependent [FG96, AL17, AL22, AF89, ARS97, ÁMS17, BL21, BHJ05, BP14, BDF89, BK21b, BtTBV87, CP05b, CL24a, dCCSR03, Dia95, DY03, DMA22, EBEHAS24, EEAS25, GS19, GLPW09, GÖS20, HILK13, HLIS16, KK20a, LKV01, Lay09, LGC25b, LHX20, Lua17, MG97, MD20c, MD25, NC16, Nor97, Pea16, PAJ12, RZ00, Rya00, SS17, ST25, SvdVvD06, SZW19, TER03, WdG92, XZW19, XL23, YS09, Yan22, ZLW22, tV87, vdVS08]. time-differencing [AEF^{+}14]. Time-discretization [MR20, NBNTGV11]. time-discretizations [BVV09]. time-domain [DHL00]. time-efficient [AGQ^{+}24]. time-fractional [AGZD22, AMK18, ALZ^{+}21, FK23, GAZD25, GWLN22, HS19b, HAY20, HAC22, Hus20, JWZ21, KAS22, KHB22, KDH20, KZ21, LR18a, LR18b, LLW20, MN23, MM20a, Odi24, QWW24, RL21, Rou20b, SY18, SJ18, SS17, SZW19, WZW13, WR20, WCS21, WCM21, WW14, YXN21, YLLZ21, ZL17, ZJH^{+}23]. time-harmonic [FJ09, Har98, MM18, QDA^{+}26, RV09, WSY18]. time-implicit [FL01b, YX25]. Time-integration [WK02]. time-irregular [PM14]. time-like [IKM23]. Time-local [MMP20]. time-marching [DD97]. time-periodic [MN08, WDU21]. time-point [Jac93]. time-primitive [AGLRS23]. time-reversible [HLR01]. time-scale [Lei02]. time-space [AD21, AS20a, Din19, FWHM20, HZD21, HZAT21, MD22, NTHC21, TDMT21, UHUL21, YJZ18]. time-splitting [An20, ZM25]. Time-step [Söd06, Zuu95]. time-step-size [GPPR12]. time-stepping [Aca12, HLT07, KR15, QCW^{+}23, Qiu23, SUP^{+}12, WTY21, WWF20, YPD21, YLLZ21]. time-symmetry [Zan01]. time-varying [DS21a, KPRU20, MVG14]. Timestepping [HW15]. Timoshenko [CF13c]. tissue [BK09, JK17]. Tocher [Bar12]. Toda [BT19]. Toeplitz [lLNW21, Sae14, WWLS08]. Toeplitz-type [WWLS08]. tomography [AES13, GOGF03, RSD^{+}06, San03, SS16, YLW20a, ZZ25]. toolbox [EMMK01]. tools [Ber85, MLK06]. topics [Wen10a]. Topographical [HdSRI17]. topography [Ant13, DL21b, DH12b]. Topological [AH09, Tan87]. topology [ZWL11]. tori [DB95, ERS00, HL99, Tru00]. toroidal [FJS25, Maj20]. Toronto [CFTW08, HEJ96]. Torrey [DA18a]. torsion [RLMG24]. Total [DH12b, FS05, GM87, HGF24, JL86, JL87, LV12, LPR00a, LJYS20, LCZ21, LLW22, MMDS21, PQ25, ZZX19a]. total-variation-boundedness [FS05]. Totally [GP17, GP93, GS25a]. TPFA [AFIS24]. TR [HS96]. TR-BDF2 [HS96]. trace [CDG19]. Tracing [Att97]. Tracking [CFXZ06, CL26, Dav92, Gro94, LD21, MSS^{+}15, PH15, RLSS06, SW86, WB92b]. trade [For11]. traditional [BDE22]. Traffic [MZ87, SG16]. train [RS22]. trains [AL20]. trajectories [IK24]. transaction [AD99, FV01]. transcendental [Boy07]. transcendents [AY15]. transcription [BCT16, ECB07]. transfer [AKBF19, CGJ16, GS20, Jéz99, Kan04, MCM12, Osw97, Sin24, Vas17, jW15, ZJLA22]. transform [BPV25, DO17b, GFPG18, HS21b, LFP04, MPSS16, MO01, OT24, RU21, SK22, Tah96, TZ00, XFG19, vI87, ST14b]. transformation [BW15, BHHS10, BW03, DLPV17, HS24, HL23, LT01, Naj20, NZY21, RN22, SS10, Win04]. Transformations [Bak86, AEA23, Ben98, Ber85, BZ94b, BZ96, Bre02b, BRZ17, Car19, Dra91, EEAS25, Fdi97a, Lem02, Mat91, Sim94b, WM07, XGM08, ZB19a]. Transformed [MSA20]. Transforming [ZPT92]. transforms [AQ20, BH96, DO17a, FV87, KW21, Moo95c, PLB22, VA05]. transient [AGLRS23, DZ12a, DCN^{+}19, Fai00, GRGJ02, HHC08, HH10b, LKJ20, lLXhLZ21, Que21, RSK24, SWW11, WSC09]. transistor [Mol95]. transition [BL91, Her91, RR00, RZ15]. transitions [MR06, Sch16a]. Transmission [ABdSG23, An16, CHNN20, DL22a, DS07c, GMS12, HX11, HL97, LRS09, LZY24, Men23, Pel20]. Transonic [CF86, Gar92]. Transparent [HL03, Ehr08, Lie01, SDK15, Sof17, ZR15]. transport [ABJ12, AR15, BDES12, BM04a, BCCHM21, CNT07, CG16, DD19, DA18b, DFC09, FdSB02, FLL11, GKKM21, GWLN22, KCC04, KDAK16, LC02, LCZ25, OEAS21, PdV99, PS00, SK97, SXL22, SvdHK94, SK96, SW05, TDC13, Tou10, VBH96, YH07, vSK97]. transport-chemistry [VBH96]. transpose [SD24a]. transputer [DVV93]. transversal [BDNV19]. transverse [HMdV03]. trapezoidal [BT97c, CP03b, DN24, RS08a]. Traub [TQA^{+}25]. Traveling [YEZ25, EV96, FF20, LMPS19]. travelling [DvHM19]. Treatment [AL87, BB25, Fat12, AM10a, AAEMY21, BDMG12, BR94, EH06, Fat10, HSWW24, KK20c, Now96, PT09, PGC01, SDK24, SB03, SKBAS08, SSR23, Tou10, WZ02, de 93a, vdHSW98]. treatments [CGA93]. tree [AS20b, KBK21, Kom07, Pow94]. tree-based [Pow94]. Treffetz [Li08]. Trefftz [HMP14, MP20]. Trends [DGCW17, Bak89]. Tresca [Por17]. tri [SXP09]. tri-dimensional [SXP09]. Triangle [BDFK95, CGG02, CD05, Ren14]. triangle-mesh [Ren14]. triangles [PSP04b]. Triangular [KCB02, BP97, Bra00, Cao07, Cao10, CL06, CST18, DDNZ18, DDGN23, DTGN23, GGLR09, JTB15, JP08b, KFOF02, Kru99, LMQZ18, LE94, MRF00, PRGO16, RGL16, SC25a, Sim04, SN04, Tsa91, Tsa92, YZ21, YZ22, ZWN23, ZSQ20]. triangularly [vdHS01]. triangulation [LS93]. triangulations [AR24, Bar12, BFGP08, Dal13, DR93]. tridiagonal [GS08, Jac93, LVfP14, Mat05, Nak12, dG91]. tridiagonalization [WJF25]. tridiagonalization-based [WJF25]. triggered [FGP23]. trigonometric [AC23, BCDR25, Han06, LZ18, Pat98, SMJ24]. Trigonometrically [CSLY19, YW08, FW08, Li19]. Trigonometrically-fitted [YW08, Li19]. triple [BH20, CMR94, HB20]. triplets [GS17, KBG04]. trivial [CM02, DBH25]. Trondheim [KNO96]. TRSVD [BHL^{+}21]. Truncated [LMTW20, Sch89, Sof17, CCY22, CM06, DRS19, GHHG22, GLMY17, GLM18, HZCZ23, PS02, SMTHE22a, TM24, TD09]. truncation [Bor10, NV23, Udd20]. truncation-error [NV23]. trusses [AMH03]. Trust [MMP09, AMV03, BMM03, CSM07, FW22, lLXhLZ21, Ou11, Ren14, SP99, ZH15]. Trust-region [MMP09, BMM03, FW22, lLXhLZ21]. trust-region-approach [AMV03]. Tsan [VN21]. Tseng [OMSS25]. tsunami [KHM^{+}14]. tsunamis [FKA^{+}13]. Tsuzuki [LO23]. Tsynkov [Tsy96]. Tubelike [MMP02b]. tubes [DII15, LDIW16, LT00, TYKK01a, TYKK01b]. tumor [DLS22, DGGM25, MPSS16, MDD18, NK24a, Rab25]. tumor-angiogenesis [DGGM25]. tumour [LBLT13, RA17]. tuning [RTV02]. tupleware [Dou91]. Turán [MSP10]. Turán-type [MSP10]. turbine [ARSW05]. turbocharged [SM08]. Turbulence [BC93, LCVG01, DS97a, Dun18, SLJ86, VV95]. turbulent [GKKM21, Kar89]. Turek [GKMS09, HCJW24]. Turing [AGJCP25]. turning [Bec18, GD23a]. TV [LS24c]. Two [AY21, AL95, BHB23, BES18, BCC16, Bic21, CL10, CWHF19, CWZ23, CPOGO17, CDP12, DP12, DCC14, Eva94, GS15b, GVP93, Guo26, HCY18, HJYL19, HL24, HFL12, KM19, Lam13, Lan95, Leo10a, LA11, LCHW20, LL20b, LS93, MOZ87, Nud24, QAMX17, RV22, SSW20, SMW21, SM85, TOD11, TS08, WZW13, WH19b, WHW21, XLZ20, YW19, YW25, ZJ25, ZS21b, AS11, AD20c, AZHD23, AHJ^{+}23, AJT19, AA24, AAL^{+}26, AAM03, AMR12, ASCM02, AN22, Bac18, Bac19, BHJJ06, Bar09, BJ05, BMR^{+}17a, BvG19, BLN25, BO04, BRTB19, BBRBS09, BRVC09, BC05, Bi20, BBLT15, BWY17, BO11, Boz11, BT97d, CCCC^{+}25, CA21, CGEV19, CMRV11, CMRdlT24, CCP17, Cas96, CW98, CR23a, CMP03, CDW13, CLTA18, CSLY19, CYWH22, CDW23, CLL25, CQZ20, Chi93, CAAT16, CL09, CJ25, CF08, dCCSR03]. two [Con01, CX08, Cop03, CF14, CMCGTR02, CN15, DdSF07, DIJ12, DR09a, DS07a, DE06, Deh01, DA18a, DA18b, DL20, DHWL22, DMGVO05, DMGVPO09, Din93, Din19, DN08, DMPSC16, EEAS25, ER07, EFLFP09, FMS18, FH20, FH22, FJL21, FdSB02, FCW20, FXCW21, GX11, GM16, GHK16, GMM09, GLPW09, Gil10, GOGF03, GOP06, GH02, GH21, GGRBRG22, HBA^{+}24, HGM^{+}21, HH22, HGF24, HK93, HM09, HS09a, Hoa15, HL19, HL21, HK85, HL02b, HCW16, Hua17, HZAT21, HS21b, HY24b, Hus20, IS22, Jad94, JTB15, JZXJ21, JHGZ20, JZZH22, Joh01, JT06a, JGK11, KMS19, KTK20, KV07, KM17, KME20, Kat89, KSMMM16, KL09, KS01, KZ13, KW20, LCHR03, LT12, LY08, LC19, LCS19, LW21a, LH21, LH23, LLZL24, LHY24, LAZ20, LSG24, LS21, LS07b, LLZ19, LZCF21, LS24b, LCZ25, Lua17, MZN21, MS19]. two [Man97, MM22, MWC21, Mar05, Med96, MT06, MT05, Moo95c, Mur19, Nag22, NS21a, NY13, OGV92b, OGV92a, Pan21, PR90, PSW02, PWS05, PWS06, PRS20, PS19, QM10, QWX20, QNA23, QR03, RP17, RZ15, Sch16a, SWJ09, SLW17, SC25a, SAH24, SC25b, SJ20, SC20, SW20a, Shi25, SZE^{+}92, ST20, Sin23, SY08, SR24, SSA24, SS09, Su94, SWJ^{+}24, SAMSB20b, TBRBM20, Tem23, TKN11, Tob14, TJ12, TV91, TM05, Ver06, WDZS21, Wan07a, Wan11, WZL13, WQ17, WMLB19, WLM21, WCS21, WCM21, WCJ23, WPN^{+}24, WW24, WTF25, WKWC26, WSP04, WWWW25, XY25, XC85, XL09a, Xiu08, XGQ20, XXF22, XP23, XLHL25, YLFT20, YS24a, YWH20, YWW23, YCWH23, YY24, Yua93, ZTZ15, Zar17, Zen21, ZOZ09, ZZHS18, ZY19, ZBY19, ZJ19a, ZL22, ZQZ23, ZGF^{+}25, ZWG25, ZZW97, ZZX20, ZLG24, ZEW20, ZZJ21, ZLWF21, ZD25b, ZX09]. two [ZCSH11a, ZCSH11b, ZZ25, dlHV13, van98]. two- [ASCM02]. two-and [SZE^{+}92]. two-asset [Bi20, ZD25b]. two-block [Yua93]. two-branched [BRTB19, TBRBM20]. two-by-two [LAZ20, Zen21]. two-derivative [CSLY19]. Two-dimensional [CPOGO17, DCC14, Lan95, RV22, WHW21, AS11, AD20c, AJT19, AA24, AMR12, BLN25, BBRBS09, BBLT15, CR23a, CDW13, CLTA18, CQZ20, CAAT16, dCCSR03, CMCGTR02, DE06, Deh01, DA18a, DA18b, DHWL22, Din93, Din19, DN08, EEAS25, FdSB02, FCW20, FXCW21, GX11, GM16, GMM09, HBA^{+}24, HK85, HZAT21, HS21b, HY24b, JTB15, JHGZ20, JZZH22, JT06a, JGK11, KTK20, KV07, KME20, LCHR03, LC19, LH21, LS07b, LZCF21, LCZ25, LS93, MS19, MM22, Med96, Moo95c, NY13, Pan21, PR90, QM10, SLW17, SC20, Shi25, Su94, TBRBM20, Tem23, TJ12, TV91, WZL13, WQ17, WMLB19, WCS21, WCM21, WPN^{+}24, WKWC26, XC85, XGQ20, YWW23, ZJ19a, ZQZ23, ZZX20, ZLG24, ZCSH11a, ZCSH11b, dlHV13]. two-fluid [CL09, Mur19]. Two-grid [CL10, CWHF19, CWZ23, HCY18, HJYL19, LCHW20, WH19b, AZHD23, CYWH22, CDW23, HL19, LLZL24, LHY24, LSG24, SJ20, SW20a, SR24, WTF25, WWWW25, XY25, XGQ20, XXF22, YWH20, YCWH23, ZTZ15, ZZHS18, ZBY19]. two-layer [CCCC^{+}25]. Two-level [HFL12, KM19, LA11, ZS21b, BvG19, DdSF07, HL02b, Hua17, Joh01, KL09, LCS19, PRS20, Tob14, ZOZ09]. two-mode [ZY19]. two-order [LH23]. two-parameter [GOP06, KMS19]. two-phase [Boz11, CLL25, ER07, GHK16, GOGF03, KSMMM16, LY08, LS24b, MZN21, NS21a, RZ15, Sch16a, SC25b, SY08, SWJ^{+}24, TKN11]. Two-Point [SM85, BCC16, Bic21, GVP93, AAL^{+}26, Bac18, Bac19, CA21, Cas96, CW98, DMGVO05, DMGVPO09, Gil10, GH21, HGM^{+}21, HK93, HM09, Kat89, KS01, OGV92b, OGV92a, Wan07a, Wan11, ZGF^{+}25, ZZ25]. two-points [CN15]. two-scale [BMR^{+}17a, PS19, WDZS21]. two-sided [GH02, LS21, MWC21, MT06, YLFT20, YS24a, ZEW20, ZLWF21]. two-space [QNA23]. two-species [HH22]. two-stage [Chi93, Lua17]. Two-step [CDP12, DP12, Leo10a, QAMX17, AHJ^{+}23, AN22, BJ05, BT97d, CMRV11, CMRdlT24, CCP17, Con01, CX08, DIJ12, FH20, FH22, FJL21, GLPW09, HS09a, Hoa15, IS22, KW20, PSW02, PWS05, PWS06, SWJ09, SAH24, SS09, Ver06, WSP04, ZX09]. two-stroke [van98]. two-term [WW24]. twofold [BCGS24]. Type [Ma24, ABD^{+}25, AH17, AM95b, AAM03, AB14, Ari87, ASC03, AO91, BN99, BW23a, BFGP08, BFQ22, Bla00, BHJ13, BS93, Bru07, BDM03, BJ96, BC00b, CO09, CC24, CCD^{+}20, Car19, Car09b, CMR12, CPD^{+}05, Cha25, CH07, CJ25, DCSB25, CM07, Cui04, Cve02, Dar90, DRVA20, DN24, DMGVO05, DMR10, DN08, Dor01, DLZ21, Ehr08, ER18, EP15, EEE22, EHV19, EHV24, Fac03, FW08, Fdi97b, FCX06, Fun04, Gan09, GD23a, Gol86, GPHA22, GKA17, GS21, yGyZ07, HHR12, Han93, HGR01, HM17, Has09, Has13, Has20, HAML21, Hau88, HO24a, HS19a, Hig96, HZC22, HLC01, HZ12, HCW16, Hua20, IMC22, JL23b, KM17, KOR18, KDH20, Kid90a, Kid90b, KLY05, KS09c, LRS23, Lau17b, LDIW16, Lev91a, LLY11, LW19a, LW20a, LX21, LWLW24, LLL12, LWC24, LS24a, LDH^{+}24, LYA^{+}19, LJ20b, Mar99b]. type [MP98, MN20, Meh22, MOV24, MAG13, MSP10, NLS20, Ney95, NK24a, NV23, OB20, ÖT20, OG08, OGV92b, Par21, PSR04, PGP03, PH17, PWS05, QR24, QXG21, RG22, RG02, RTT01, SDK24, Sch16a, SR09, SC25a, She00, SL09, pSLqJcY16, SHLY19, SSR23, Shi20, SVB17, SS16, SA19, SMJ24, SSS21, Tan23, Tom24, TÖR22, Uty08, VV05, VCC12, WZL08, WCXL09, Wan09, WMF17, WZ22, WWLS08, XLHL25, YSBL14, Yan18, YLY19, Yan21a, YC13, YR09, Zan01, ZL21, ZL22, ZYQS23, ZJ19c, ZFX25, ZP98, ZQLK11, ZX14, ZSQ20, ZZ25, vI87, vC93, AR93, Mat91, Sim91]. types [SAMSB20b].

U [Bec02]. U. [MZ87]. UE [NLZB23].

UE-spline [NLZB23]. Uhlenbeck [ZGO12]. Ulam [KS25]. Ulm [SLJ11]. Ulm-like [SLJ11]. ultra [BTBR19, DM12]. Ultra-Weak [DM12]. ultrasound [MDHK06]. ultraspherical [KP92]. unbounded [BT97a, BDMGVO05, CSX23, DG22, DY03, EH07b, FMS18, FWL18, Fer09, Gia12, yGyZ07, GD22, HHYD20, HZBM05, HZ09, HY01, HLJ20, LZH19, LC25, NN20, Pec09, Tsy98, WL18, WS04, YLW21, ZLW22, ZZL17]. uncertain [FV01, ZW09]. uncertainties [FS23b]. Uncertainty [BF17, BH20, DCSB25, LGH11, PKP19, Tar98]. Unconditional [CC20a, Geb24, HO10, LSWM19, LZW19, MPPR22, SL20, SW20a, SZ22b, WCJ23, XXYZ24, XP23, YJ21, ZSS23, ZL23, ZYS17, iW09, LL20b, Shi25]. Unconditionally [DGE22, GWLN22, JY23, KWY24, Yan23, ZJ19a, BRBM08, BBKS07, CRSF19, GZQS23, GHS25, GQ08, HWZ22, HZZ25, JJL^{+}24, LK14, LQXK23, LC21, QH22, SL21, Ske89a, WWZJ22, WWZ25, XK25, YZH19a, YYZ23, ZYQS23, ZFS24, ZH24]. unconfined [AMRR18]. unconstrained [Cui24, DW15, Fre91, Kim19, KLS13, lLXhLZ21, MK19, Ou11, Rha99, Wan23]. unconventional [SS94a]. uncoupling [LT12]. undamped [ST19]. under-determined [SK10]. under-field [CL02a]. underdetermined [BLS^{+}17, CKB12, PR22]. underflow [FT96, FT06]. underlying [AO05]. underpin [JR18]. underwater [XLK07]. undulatory [Ran20]. unfitted [CL24b, HCJW24, ZCC11]. Unified [GP23, NSCC19, AC16, Bus06, DD21, DTGN23, GFPG18, Har98, HP14, Mou03, SK97, Tou25, ZAED21]. Uniform [FXY22, WZ17, ZL21, AD15, BNS25, BH12a, BCE04, BB10, Cum95, DL22a, FHM^{+}02, FR14, Fou00, FR01, GD23a, GO19, GO21, JHGZ20, JM05, KMS19, KK20c, LH11, LRC19, LR20a, LC24, Lin01, LVW21, MST09, MOSW00, NLZB23, OS08, RS09, Rou20a, SDK24, SE93, SW21, SYG^{+}05, SA18, SA19, TTP24, Tro93, VL08, VT93, Vul95, WZ22, YW19, iV09]. uniformity [CMS04]. Uniformly [ACLM22, BM01, Li23, RG22, RTU15, SKS23, ZLG15, CJ23, CJ25, DL22b, GH21, HOEC86, KD25, ZX14]. Unifying [ABM17]. Unilateral [SSZ16, Aca12, CH13, Gwi09, HL02a, WWH24]. uniquely [LMY18]. Uniqueness [LLVX20, Fuh01, Gar92, HHT97, KS25, KP15, LYY15, Zak19]. unit [BRS16, CCBGV08, DGV00, DBCBPP10, DIR13, KPRU20, SL01a]. unit-vector [KPRU20]. united [DMH18]. unity [CDD^{+}17]. univariate [BH12a]. univariately [TD09]. universal [QR03]. universe [ZH21]. Universidad [BGHR12, BGH^{+}15]. University [Ano02g, CFTW08]. unknown [HP14, LJ20b, Mar08, SZQH23, TDMT21]. unknowns [CT93, Gar96, Gar03, MM02c, Pou00, PB10, SW13, Che96]. unsaturated [BBT25]. unscented [KK20b]. unstable [DLM02, Gro94, WK00]. Unstaggered [Tou10]. Unsteady [Duf90, SA00, AA22, CML05, CCZ22, DM09a, Gla94, KM19, KTYY24, LS20, Mar05, MV18, Mur15, PBC08, SSA24, TYKK01a, TH18, UHUL21, VNC21, WPT19, ZFW20, ZFS24]. unsteady-state [CML05]. Unstructured [FVB05, ASC03, BS94b, BTC23, Dal13, DF11, DSZ15b, FL01a, JTB15, Kni94, Kni95, LYOI99, NFAE03, OK98, PSP04a, Ram94, SGN06, SGN08, TJ12]. unsymmetric [Khe91, Nov03, Sch08a]. update [BLD17]. updated [YK07]. updated-observations [YK07]. updates [BTMT08, LLT20a, SW09a]. Updating [BS10, GS08, Bai25a, CMT25, lLXhLZ21, TS08, XZZL15, ZH15]. upon [AGJM04]. upstream [DN08]. uptake [BBT25]. upwind [BGP11, BCS06, CXNF14, CKM25, DLZ21, FCX06, Hol01, JT18, LY09, Pir09, RK91, SY08, SR97, VL08, van86b]. upwind-mixed [SY08]. upwind/Petrov [DLZ21]. upwinding [AS06, KN93]. urban [CCM17]. Urysohn [BLN25, KR18]. USA [FJ97]. Use [FJH^{+}01, RS08b, AA04, Att97, BLS94, BZ93, Gil91, JKN94, Man96, SH02, Tou97, YR92]. used [BBO03, BP95, MS08a, Sha85b]. user [WSS97]. user-oriented [WSS97]. uses [SA05, TS08]. Using [CMS04, MS99a, RSD^{+}06, Sch95a, SG09, ST08, TMS87, Zha96, AKM^{+}21, AN25, ABZ21, ABJ12, AyLqW18, AFIS24, Arc06, AJK20, AD18a, AD19b, BKM19, BC02, BOEP00, BM01, BGG^{+}20, BDF89, BW96a, BCJ97, Ber05, BCT16, BGH08, Bla01, BL08, BH12b, BWEP95, CC24, CAD03, CDD^{+}17, CCdlH20, CJL13, CKP15, CGCMTR02, Che96, CS01, CKS05, Cod08, CMS06, CJ22, Dah02, DB97, Dea11, Der92, Dia95, DB08, DP85, DL06, EHNR24, EJRR23, EEAS25, Elg17, EW08, ELLE02, EJS11, EC07, FZM20, FPPS00, Fra14, Gab02, GV18, GNUV24, GGLR09, Gat91, GGM95, Ghe97, GKKM21, GeO24, GDS^{+}15, GGG16, yGyZ07, HH98, HS19a, HSWW24, JUAZ22, JM94, Jun97, Jun06, KO92, KK09b, Kni94, Kni95, KTS03, KR25, Koz94, KP19, KX03, LS10, Li98, LH09, LH21, LLT07]. using [LS24c, LL02, LB21, MZZ17, MMKN17, MK21, MRF00, MCS16, MS25, MM16, Mic95, MW24b, MPtM16, MDASAO21, MMP20, MD96, Moo04, MRFF17, Mur19, NRWF08, NN20, Naj20, Nap16, NS03, NLS18, PL20, PT09, Pel15, PP00, Pic05, PRS23, PCRR17, QW04, RP17, Ram12, RGÖS18, RG05, Roo20, RBT15, SRK21, SRK22, SAA20, SK97, Sch08a, SMC08, SW95b, SW11, SY05, SA00, SYG^{+}05, SSA^{+}22, SD93, SBS^{+}20, SA08, SW13, SOA^{+}25, TLP18a, TLP18b, WM07, WS04, WYP12, YC00, YL26, Yos00, YR22, YDWW17, ZT06, ZG21, ZYJD25, ZPT92, ZSJ04, ZWL11, dRT99]. Utility [AD99]. Uzawa [Cao03, Cui04, Mur15, SED21, WWC^{+}25]. Uzawa-like [Mur15].

V [EST15, Tsy96]. V.S [Ano00h]. vaccination [CML^{+}25]. vadose [BDNV19]. vague [MMKN17]. Validated [LS07a, JN02, Jéz99]. validating [Tou97]. Validation [AV96, CAAT16, TLV92, BWEP95]. Vallée [ORT24, OT25]. valuation [AO05, CF08, Mar09]. Value [AGM09, De 88, SM85, AA04, AL09, AyLqW18, AB15, AMP03, AMCR17, AMR14, AT93, AM95b, AGM95, Att97, AFLG^{+}12, ABRW18, Bac17a, Bac17b, Bac18, Bac19, BTBR19, BPN24, BGO13, BY00, BM00, Bic16, BCC16, Bic21, BBRS97, BT93a, BT93b, BT95, BT98, Bur93a, BT00, BC89c, Cah89, CA21, CL85, Cas96, CW98, CJ90, CZ12, CHS19, CYYH21, CKK25, CH21, CN17, DS24, DS20, DMPSC16, DSS20, EHM01, FW08, FD97, FL20, FMPP24, Fer93, Fun94, GGM95, GM95, Ghe97, GH02, yGpY09, Guo26, HGM^{+}21, HK93, Hig93a, HR06, IO18, JL21, JL23a, JR02, JCSR03, JCJP21, KHLV22, Kat89, KOR18, KG90, KS09a, KKW00, KS01, KW93, KK02, KK20c, KDKW20, LHH96, LZ17, LW17, LZJ21, LWW22, LHC23, LDC24, LC99, LO03, LS07a, LM22a, LT93, LDH^{+}24, LOM98, LXLY25]. value [MS03, MS90, MYSC17, MRÁSY24, Nap16, NNJ23, NSCC19, OB20, Odi24, PZMX16, Pap95, PT15, Pul12, RNG22, RK08, Roo20, RTU15, RGA19, RTA19, Rou20a, SAG86, Sch93, SWE05, SNOK21, Set24, SSC23, pSLqJcY16, SHLY19, SWB20, SWB21, SW85, Sub04, TS08, VO00b, VV02, Wan07a, Wan07b, Wan11, Wan17b, jWS20, War92, WW14, XFLC00, YL13, Yos00, YW25, Zak19, ZLY23, ZG92b, ZC10, ZLJ20, ZLS20, ZZ19b, ZWL11, dVA02]. valued [AHAS21, LZZ22, LAZ20, LW07]. values [CH21, Pré10, pSLqJcY16, SHLY19, Zha07]. valuing [itHT18]. Vandermonde [MP98, dC18a]. vanishing [CS17, KV95, LZZ18, MPDB24, Pis22]. vanishing-lag [KV95]. vapor [CKB13]. Variable [AGM95, Ber04, CLMSS98, GPMPR03, Hai97, Jac87, JVZ95, JAH21, KW20, LWW23, LWWZ25, MDP23, PAJ12, WKP12, YXZL24, Zla85b, ASGGRGW25, ASA20, AEA23, An20, Ant13, AFS02, BCGS24, Ber05, BG02a, BZ17a, BB24b, BRW17, BC98, CGA96, CK20, CR23b, CFM^{+}24, DFLM19, Fel06, GD09, GPPR12, GJLL20, HOS99, HP14, HMN20, HA21, HMD21, Hey19, Hey20a, HAA21, HAR21, HLR01, Hol01, HLMKZ06, IJ14, JWZ21, JL17, JV09, KXK92, KS00, KS25, KHB22, KKR25, LWT07, LLL08, LA21, LGC25a, LLT07, LS21, LYZW22, LFS15, LMG02, MJS23, NP21, PWY21, PM91, PQ25, QZM25, RL21, SH97, SRMDRL23, Sha87, SS08b, SAMSB20a, SAMSB20b, Top21, VL19, Wai98, WR20, WZZ21, WWM22, WWC^{+}25, WWF20, YMD21, YLFT20, ZHL22, ZJH^{+}23, ZFS24, ZZF25, ZZjQ24, ZEW20, ZLWF21]. variable-coefficient [PWY21, ZLWF21]. Variable-Order [Jac87, Ber04, GPMPR03, Ber05, BZ17a, Hey19, Hey20a, HAA21, HAR21, JWZ21, KHB22, LMG02, NP21, Sha87, SAMSB20a, SAMSB20b, YMD21]. variable-sign [QZM25]. Variable-Step [Jac87, AGM95, ZZF25]. Variable-stepsize [KW20, WKP12, LMG02]. Variable-time-step [LWW23, LWWZ25, LGC25a]. variables [AEK23, BC93, Dar90, Kid90a, Kid90b, Lav94, Man97, PM03]. Variance [Dah02, JLH13, CHZZ06, LHÖ13]. Variant [LG19, LLY11, MC17, QM10, Ree03, SMTHE22b]. variants [DRS19, HS02]. variates [Dah02]. variation [BAP^{+}06, BS20b, DH12b, FS05, FL24a, ILS19, JL86, JL87, LPR00a, ZZX19a]. Variational [BFS17, BCGI13, DM12, KP03a, RSK24, RTH23, BNV06, BHR05, Bot97, CP97, CBD16, CGK25, CDI^{+}24, DC21, DHS05, DP21, FGPR12, GH20, Gwi09, Har98, IS23, JMDN^{+}22, JMU^{+}25, KHM^{+}14, KPR12a, KPR12b, LLY11, LWL18, LQY25, Lub04, MD20b, MG00, OMSS25, RZS21, RM25, SQ17, SI20, SMA01, SC22, TLQ21, TGV22, VA21, WH13, WPAZ24, YXZL24, YÇ16, ZDM18, ZHJ14, ZBNC25, ZMF25, ZS21a]. Variations [KYC03, RG02, SST04]. various [Nes16, VSeYD02]. varying [CSSZ25, DS21a, KPRU20, MVG14, RE19, Rya00]. VCSEL [AC08]. Vector [BZ96, DTQ^{+}20, NT16, AAD^{+}08, Alt85, Bel91, Bre02b, BRZ17, CC04a, Con04, Duf90, Fer14, FS88a, Fuj99, Gia12, HvdHV10, KPRU20, MZZ17, Sad96, Sad97, Sid90, SSW04, WGB99, WWS^{+}93, Win92, Wu03, YLL21, ZL24, de 92b, van86a]. vectorial [JPP19]. vectorizable [BV94]. vectors [CP07, GM94, LJ20b, PTW19, VSeYD02]. vegetated [IMC22]. velocities [DJL04, Fan19]. Velocity [BM18, AMN24, Ale11, CHR03, CLY19, DFC09, GLS09, Gat91, GP00, GS18, GGG16, HL19, HSF25, KS02, LD21, LY16, LW07, Med96, RR00, SLW24]. velocity-correction [SLW24]. Velocity-current [BM18]. velocity-pressure [AMN24]. velocity-pressure-stress [GP00, KS02]. VEM [LWW23, YSZG25]. Venant [FCX06, IMC22]. Verification [LZZ18, RBBC85, SK01]. verifications [WKN20]. verified [Fac03, Rum87]. version [AK00, AM16a, CS09, DN21, GGO13, GW20, Gwi09, HH18, JJ94, KJ99, KFOF02, KX03, LWY20, Mai06, ML91, NDM20, SS00, SK10, WTY21, XY19, YMD21, YS24b, ZMY21]. versions [Mus11, SLMD21]. versus [CHR03]. Vertex [ZL18a]. Vertex-cell [ZL18a]. vertical [BDNV19, GDEdLD23]. vertices [Dal13]. Verwer [BtTBV87]. very [BC99, BP14, BCMV03]. vessels [BCS06]. via [AD20b, ABdSG23, AA87, ALP^{+}96, Bai97, BBCR22, BGG04, BRRS15, BGG^{+}21, BHSW20, Bor10, BM06b, CFXZ06, CCDJ20, Car19, Cha25, CZ19, CM04, CPZ17, CGGM17, CDI^{+}24, DSZ15b, DSZ15a, FVGS13, EEJB22, HA21, HR14, HXW15, JMDN^{+}22, JCJP21, JS25, KP18, KW21, KOS21, KME20, KNP16, KK22b, KSP10, Li25, Liu21, MD19b, MS00, MMRV20, NRZR12, OHA25, Pas91, PTW19, PM14, PWX24, SA90, SSA^{+}22, Sid14, Udd20, WZ16, WQ17, YV17, YT03, YR22, Yu08, ZM19, ZH20, Zha07, ZZ25, dlHV13]. viability [BMMZ06]. vibrating [Ahn07, CCJ99, CFC03, LR03]. Vibration [LHC09, AMH03, Dat99b, LFB00, MR01, dlC23]. vibrationally [MRS03]. vibrations [RL06]. video [EEJB22]. Vienna [AFS96]. VIEs [Cai24]. Vieta [HAR21, Wu03]. view [Han87, LTDY24, WH13]. viewpoint [Shy91a, Shy91b]. views [MMP02a]. violating [Vul95]. Virtual [ANN19, Men23, SWY^{+}23, AMN24, Aya09, BLJ21, CGS19, CWZ23, DV20, DGM22, DGE22, ECB07, LR24, LW92b, LN21, MM22, MP20, MZM20, WaZW23, WWH24, WWWW25, YZ24b, ZZZ19]. visco [HS07, LDIW16]. visco-elastic [LDIW16]. viscoelastic [BR25, CFRA08, CCLT10, DYF23, KK06, MF23, NA21, WWZ25, Zha19b, ZYS17, ZB19b]. viscoelasticity [HHR12, Wen05]. viscoplastic [Mur15]. Viscosity [FF06, FFMZ13, MP85, AJ24a, AK21, BCGS24, CSW19, CRTU15, DK20, EK96, GT00, GGRN17, KS00, Luo18, VA21, YWSL20]. viscosity-splitting [GGRN17]. Viscous [XF06, BM18, CPY20, CCL04, Duf90, FPRA09, FL01b, GM08, GZHQ23, HT00, JT06b, LKJ07, LT00, LR87, PR90, TYKK01a, TYKK01b, WWLL23, ZYJZ23]. visual [Spi99]. visualization [Zar99]. Vlasov [BS14a, LSW23]. VLUGR3 [BV94]. VMS [RMM12, RÁM23]. Vogelius [BL08]. void [NS16]. Voigt [DYF23]. volatility [Cer25, FV01, LL15, SB19, ZW09]. volatility/transaction [FV01]. Volterra [AHJM19, ABH22, AN25, AAL21, AB88, AAH21, AM10a, AD18a, AD19b, BT97a, BLN25, BLRGVR23, BGGG13, BLM17a, BDFV95, DLS22, BP92, Bru92, Bru97, BMM97b, BMM97a, Bru07, CCP17, CDP17, Car23, CZ12, CP09, CDP12, CP17, CMPP24, CFM^{+}24, DS21b, DN21, DHWL22, Der92, DLPV17, EH97, EK06, FH22, FJ97, FPR12, GLS09, Gar10, GAH25, GGS16, Gu19, GAOB20, Han93, HZBM05, HS19a, HA16, Hu99, HY24b, IRC12, KO08, KME20, Kau97, LZQ22, Li05, LWY20, Lia22, LD97, LW92a, LM22a, MK20, MAHZ21, MAH18, MHA19, Maj14, MO17, Man97, MV17, Mic03, MSS21, MdD04, MPDB24, Naj20, NDM20, NSD23, PV24, ROB17, Ril92, SHL19, SV00, SDG20, tSqWyG16, SVB17, SP22, SPYS24, Vab22, WTY21, WHW21, WW24, WC14, WYL11, WHL19, YT21, ZL18b, Zha20a, Zha20b]. Volterra [ZHL22]. Volume [Ano01c, Ano01d, Ano02a, Ano02b, Ano02c, Ano02d, Ano02f, Ano03a, Ano03b, Ano03c, Ano03d, Ano04c, Ano04a, Ano04b, Ano05a, Ano05b, BMQW16, Jbi25, AKT97, ASCM02, ASC03, BK09, BM04a, BM04c, BW95, BC08b, BG03, BP97, BD22, CMP03, CL10, CXZ17, CH19, CNT07, CCZ22, DAH24, DKSS24, DJ10, DM11a, DT10, DN08, DHM09, DII15, EN09, EGH01, FL04, FDF^{+}25, FM07, FMU15, FSWZ19, FL01b, GLLW14, GBBC^{+}23, GQ08, GJ17, HHC08, HJZ23, ID19, Kal96, KD13, KCJP01, KvyS15, KR25, KR12, LRS23, LLHC18, LZ20, LLS^{+}96, LZIZ23, Lte24, LP01, LMWZ07, MK14, MS08a, Mat08, MR94, MQ00, MOU14, Ngu15, NB01, NFAE03, PGYF20, Que21, RVD00, RBT15, SD11, SD13b, SYY20, SG16, SNW22, Sin23, SG07, SW24, SN04, SGN06]. volume [SGN08, Tol04, TH09, TK15, WZL13, WZ19, WYYL19, YS09, ZBD24, ZXYW22, ZSY20, Zou10, Ano01a, Ano01b, BGP11]. volume-element [RBT15]. volume-finite [KD13, MK14]. volume/finite [BD22, Lte24, ZBD24]. Volumes [Ano04l, BM04b, Pel20, Ano92]. Volumetric [BH20]. Voronoi [BTC23]. Vortex [BO87, Gus87, HKS86, Mac86, MOZ87, Ta’86, CR04, Nic86, Sod91, VR01, ZdBT03]. vortical [Din93]. vorticity [AQ00, BP95, Gat91, KN93, LJ20a, Med96, RR00]. vorticity-stream [BP95, LJ20a]. vorticity-velocity [Gat91, Med96, RR00]. voting [Han19]. Vries [BDKM92, ZJ25, AM04, Bas21, BR97, FWL18, HY24a, Isk89, KS22, WH19a]. vs [KL09, Nak05].

W [Ost02, PSW02, WSP04]. W-methods [Ost02, PSW02, WSP04]. W. [SS94a]. W4 [OFY^{+}23]. Wachspress [SB25]. wakes [FPPS00]. wall [BL91, DM09a, TYKK01b, Uty08, ZFW20]. wall-bounded [BL91]. Wanner [Ran16]. was [BDN^{+}97]. wastewater [PGC01]. water [AQS94, BDMG12, BBD20, Beh97, BGG12, BTC23, BBT25, BDE22, BD22, CCCC^{+}25, CFXZ06, CCM17, DL21b, FCX06, FBS09, HAN23, HL03, IMC22, ID19, IM00, KOS21, Lie01, QLL^{+}08, STS00, SCT05, SW12, Tou10, VCC12, WLG22, WWLL23, WC24a, YTC24, ZGDL17, de 92b, vdHS01]. waters [CdCV03]. Wave [AD15, DH12a, DC18b, Gus88, Vic87b, WVBM88, YR22, AD19a, AD21, AA24, AMV17, Ang06, AD01a, BR25, BvG19, BK21a, Boy91a, CHP19, CC18, Che88, CFL^{+}20, CR04, CSCM96, DvHM19, Dek17, DM09b, DZ12b, DL20, DL22b, Din19, DG22, DC09, EH91, FMS18, Fan19, FRRJT10, FXY22, FJP17, FJ09, FJH^{+}01, GKB^{+}22, GD22, GL93, HD04, HM17, HS20, HTSZ23, HvdHV10, HL21, HL24, HCGW22, HJX^{+}19, HZAT21, IK24, IJ14, JR00, JZS20, JEG10, JQSC22, JR02, Kim95, KKE16, KDD23, KR15, LO22, LH11, LR18a, LC19, LH24, LC24, LLZL24, LSG24, LKJ07, LKJ20, lLH25, LN08, LCL18, LLZ19, LW20b, LMWZ07, LG02, LLW20, MDP10, MP20, Mot17, Mul99, Mur19, OL18, OCVW22, PB21, PD01, PA18, Pir09, RL21, RA03, Ric94, RTT01, Sal93, SN22, Sch23, SS94b]. wave [SW20a, SW21, SS17, Sod91, SDK15, SKW17, SD24b, SA18, zSW06, Ter22, TK19, TY98, UWY22, Ven15, Vic92, Wag85, WTB24, WQ17, WWLL23, WXY24, WTF25, Wee01, XC85, XLZ20, XLKY19, XXF22, YZG23, YWH20, YL26, YLLZ21, YEZ25, Yua20, ZP24, ZJ19a, ZYJZ23, ZLG24, ZZ20, ZL24]. wave-like [TY98]. wave-type [RTT01]. wave-vortex [CR04]. wave/particle [RA03]. Waveform [Fan11, AKGR14, AL22, BZ93, BDP96, CC18, JZK06, JW01, KL07, LR93, Mar05, Poh93, VP91, ZK00, in 95]. wavefunction [LY24b]. waveguides [SDK15]. Wavelet [AV00, BKP09, KNN03, TZ00, AAL^{+}26, BDOG19, BPV25, Cer25, CL01b, CPOGO17, CNS00, DGD03, EH05, FK23, GS24, GPP04, GM17, Güm20, Jam95, JCJP21, LS16, LYZW22, NLS20, PK23, Ric08, SW13, SSS21, VRC21, WPN^{+}24, YR22, YRV21a, ZCSH11a, ZCSH11b]. Wavelet-based [BKP09, CNS00, Ric08, YR22]. Wavelet-Galerkin [AV00]. wavelet-like [SW13]. Wavelets [SW06, AKM^{+}21, BKM19, BCU00, CAD03, CaAL96, DSAB20, GeO24, Har93, Hey19, KME20, KOR18, RR21, ROL19, RV22, YRV21b]. Waves [SR88b, AD01a, BBD20, Bir87, CGP15, DTQ^{+}20, DLM02, EBEHAS24, EZ03, EV96, Hag15, HAN23, HD88, KOS21, MRS03, MZ04, Mur19, NAF24, NC16, RX08]. way [Ter22, TSFB01]. Weak [AB17, ADG^{+}24, BWY03, CGH23, GCZZ23, HLY22, MJS23, PL25, PQ25, TN16, Yam18, ZW24, ZZC^{+}18, Abu04, AY22, ATW20a, ATW20b, ATAW25, BTBR19, CW21, CDW23, CRR03, DR09a, DR09b, DK21, DYF23, GZZ20, GD21, Gol86, HLW25, Hus20, KK11, Kom07, KDD23, LWW20, LC24, LGC25a, LGC25b, MK20, MWYZ18, MC21, RN25, Sch08a, Sha21, Sin24, SZL18, TZA13, Top21, TS24b, XZW19, Xu21, Yam23, YLY19, YZZ25, YZ22, ZFZ19, ZZ24, ZMF25, ZGR23, ZX22, dlC23, DM12, LX24, SI20]. Weak-convergence [BWY03]. weak-internal [CRR03]. weaker [TM24]. Weakly [HJ09, SPYS24, AL95, AAD14, BY09, BLM17a, CHLX07, CHS19, CDW19, CCST22, CP17, DN21, DLPV17, DL16, DCY20, FK23, GZZ20, GMG19, GGS16, Gu19, HY24b, IK24, KX91, KM25, LZ24, Lia22, LD97, LCZ25, LNZ12, MH16a, MAHZ21, Maj14, MKDC25, MO17, Moe98, Mok17, Naj20, PT11, PTV16, PTV20, QXG21, RN22, Tan93, Vas17, WTY21, WZ22, WW24, ZD20, dAF17]. wear [WWH24]. weather [WCGW95]. weight [DSM11, LMO24, LSY21]. Weighted [BIW24, DDRS24, FRRZ25, All24, AMR12, AM16a, BL06, Dea11, DN08, DSZ15b, DSZ15a, GHT05, HSS04, Hua19, JZZH22, LTDY24, LWWZ25, MS08a, Mil17, Muo23, NZY21, PCL^{+}25, SL15]. weights [ELR^{+}15, MN20, TDW23, WD22, Wel10b, WG11]. Weiss [GR02]. Well [KTK20, Lie01, MN24, TK15, WC25, ZGDL17, AG25, BDMG12, CCCC^{+}25, CHS17, DEGDdL25, Eng11, EL01, FGGL22, Geb24, GBBC^{+}23, KD13, MPPR22, SSZ16, WLG22, WDL23, WC24a, YTC24]. Well-balanced [KTK20, TK15, WC25, ZGDL17, AG25, BDMG12, CCCC^{+}25, DEGDdL25, FGGL22, GBBC^{+}23, KD13, WLG22, WDL23, WC24a, YTC24]. Well-posed [Lie01]. Well-posedness [MN24, Geb24, MPPR22, SSZ16]. well-separated [Eng11]. well-spaced [CHS17]. wellbore [CML05]. Weller [LHWF08]. Wellposedness [ZEW20]. wells [GRGJ02]. Wendroff [De 88, LVW21]. Weniger [PCA10]. WENO [AJ24a, BL06, DGM18, DMM24b, GOLJ25, GJ17, KHYY21, LPR00b, MM14, MC21, MRÁSY24, RGB20, SG16, Tan23, Tan24, TDW23, VRK25, Wan25, WD22, WG23a, XZL07, ZQLK11, ZGDL17, ZSQ20, ZSQ21]. WENO- [MRÁSY24]. WENO-JS-type [Tan23]. WENO-Z [RGB20]. wet [BDMG12, HHL23, WC24a, WC25]. wet-dry [WC24a, WC25]. wet/dry [BDMG12]. WG [KTYY24, XZ22]. WG-FEM [KTYY24]. which [MMDS21, SRMDRL23, SG92, ZNK02]. white [AHO16, CL18, KZ13, ST19, WYY20, GGO13, ZB19b]. Whitehead [MDRR11]. Whitham [DDK19]. whole [LLJY20, jWjJ17, jWC22]. whose [MB08, Mar08]. wide [ADK94, WL24, YZH19b]. wide-angle [ADK94]. width [AM16b, KOS21]. Wiener [JMDN^{+}22, Sob24]. Williams [GT18]. Willmore [HGR01, Obe15]. Willmore-type [HGR01]. Wilson [YB10]. wilt [ABdSG23]. Wind [BO87, Fan19]. Wing [Gar87]. wire [DDZK05]. wise [LSY17, LW21a, MM14, WKWC26]. within [AD15, DS15, FR18, SWFK13]. without [AQJ18, AMC02, ASC03, AQ00, BM12a, Boy07, KWY24, KNP16, LSWW22, LCZ23, SL21, mWyG00, WZ16, Yam18, YDWW17]. WLS [ZB19b]. WLS/SUPG [ZB19b]. WONAPDE [BGHR12, BGH^{+}15]. wood [QPT23]. Woodbury [Mit24a]. work [AV91]. working [BW23a, GT02]. working-set [GT02]. Workshop [BGHR12, BGH^{+}15, FG96]. World [HSX18, MH89]. wormhole [ZSG^{+}20]. WR [BK06]. Wright [Meh22]. Wright-type [Meh22]. WSGD [ZYLL20, LQS21, LS21]. WWP [YLW20b]. Wynn [MV20, Pré95].

x [Lu98a, Ram12, SSW04]. X-ray [SSW04]. X-splines [Ram12]. Xfem [CLR11]. XVA [CC23b].

Y-MP [HVY91, van95]. Yanosik [Pot85]. years [Bru97, Gil10]. yields [PRS20]. Young [LW04, Mil17, XL09b].

Z [RGB20]. Zakharov [CFLW22, FvdMS17, MD20a, MMDH19, MMD20, MMDS21, XWW19, XWZ21]. zebrafish [MSS^{+}15]. zero [AGKS25b, ALY03, BC00a, LWCT07, SHLY19, Ske89a]. zerocoupon [ALY03]. zero-stable [Ske89a]. Zeros [BDRZ19, VC10, ADG11, CMR12, CM06, DR01, DMR10, DIR13, DJM09, JJ15, JT09, KYI17, LMV25, PH91, Sim94a, TJK18, Tom24, Wu03]. zeroth [SW17]. zerothorder [SW17]. zeta [Pré10]. Zhang [RG02, SSW20]. zone [BDNV19]. zones [BDMG12, KKE16]. ZZ [CHZ21].

References

Alaylioglu:1987:GPF

[AA87] Ayse Alaylioglu and H. Alaylioglu. A general purpose finite element system code via computationally effective hybrid formulation. Applied Numerical Mathematics: Transactions of IMACS, 3(4):291–303, August 1987. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ahues:1994:RMM

[AA94] Mario Ahués and Kwami Attoh. A refinement method for maximal deflating bases of regular pencils. Applied Numerical Mathematics: Transactions of IMACS, 14(4): 367–382, June 10, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=14&
issue=4&aid=478.

Abbasbandy:2004:USA

[AA04] S. Abbasbandy and M. A. Fariborzi Araghi. The use of the stochastic arithmetic to estimate the value of interpolation polynomial with optimal degree. Applied Numerical Mathematics: Transactions of IMACS, 50(3–4): 279–290, September 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927404000066.

Abbasbandy:2005:SSS

[AA05] Saeid Abbasbandy and Mohammad Ali Fariborzi Araghi.

A stochastic scheme for solving definite integrals. Applied Numerical Mathematics: Transactions of IMACS, 55(2):125–136, October 2005. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Alharbi:2020:ESN

[AA20] Sayer O. Alharbi and Saleh M. Alzahrani. An efficient solver for the numerical solution of EM scattering problems by many small dielectric circular obstacles. Applied Numerical Mathematics: Transactions of IMACS, 147(??):88– 100, January 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419302120.

Albuja:2021:FNG

[AÁ21] Guillermo Albuja and Andrés I. Ávila. A family of new globally convergent linearization schemes for solving Richards’ equation. Applied Numerical Mathematics: Transactions of IMACS, 159(??): 281–296, January 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302889.

Akyildiz:2022:NAG

[AA22] F. Talay Akyildiz and Fehaid Alshammari. A new analysis of Galerkin Legendre spectral methods for coupled hyperbolic/parabolic system arising in unsteady MHD flow of Maxwell fluid and numerical simulation. Applied Numerical Mathematics: Transactions of IMACS, 176(??):83–103, June 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422000423.

Aloraini:2024:CDS

[AA24] Najla M. Aloraini and Talha Achouri. Compact difference scheme for the twodimensional semilinear wave equation. Applied Numerical Mathematics: Transactions of IMACS, 202(??): 173–188, August 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001119.

Aguayo:2025:PEE

[AA25] Jorge Aguayo and Rodolfo Araya. A priori error estimates for a coseismic slip optimal control problem. Applied Numerical Mathematics: Transactions of IMACS, 209(??):84–99, March 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424003192.

Auer:2022:NSB

[AAB^{+}22] Felix Karl Auer, Winfried Auzinger, Jana Burkotová, Irena Rachunková, and Ewa B. Weinmüller. On nonlinear singular BVPs with nonsmooth data. Part 2: Convergence of collocation methods. Applied Numerical Mathematics: Transactions of IMACS, 171(??): 149–175, January 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421002440.

Abundo:2020:FPF

[AACP20] M. Abundo, G. Ascione, M. F. Carfora, and E. Pirozzi. A fractional PDE for first passage time of time-changed Brownian motion and its numerical solution. Applied Numerical Mathematics: Transactions of IMACS, 155(??):103–118, ???? 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419301928.

Ahusborde:2008:CSA

[AAD^{+}08] E. Ahusborde, M. Azaïez, M. O. Deville, R. Gruber, and E. H. Mund. Constructive spectral approaches for the Helmholtz decomposition of a vector field. Applied Numerical Mathematics: Transactions of IMACS, 58(7):955–967, July 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Assari:2014:NSW

[AAD14] Pouria Assari, Hojatollah Adibi, and Mehdi Dehghan. The numerical solution of weakly singular integral equations based on the meshless product integration (MPI) method with error analysis. Applied Numerical Mathematics: Transactions of IMACS, 81(??): 76–93, July 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414000427.

Atta:2021:SFK

[AAEMY21] A. G. Atta, W. M. Abd-Elhameed, G. M. Moatimid, and Y. H. Youssri. Shifted fifth-kind Chebyshev Galerkin treatment for linear hyperbolic first-order partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 167(??): 237–256, September 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421001392.

Almasoodi:2021:GBD

[AAH21] A. Y. J. Almasoodi, A. Abdi, and G. Hojjati. A GLMsbased difference-quadrature scheme for Volterra integrodifferential equations. Applied Numerical Mathematics: Transactions of IMACS, 163(??):292–302, May 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742100043X.

Amitai:1993:SAF

[AAI^{+}93] D. Amitai, A. Averbuch, M. Israeli, S. Itzikowitz, and E. Turkel. A survey of asynchronous finite-difference methods for parabolic PDEs on multiprocessors. Applied Numerical Mathematics: Transactions of IMACS, 12(1–3):27–45, May 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/016892749390110D.

Adewumi:2021:SLS

[AAL21] Adebayo Olusegun Adewumi, Saheed Ojo Akindeinde, and Ramoshweu Solomon Lebelo. Sumudu Lagrange-spectral methods for solving system of linear and nonlinear Volterra integro-differential equations. Applied Numerical Mathematics: Transactions of IMACS, 169(??): 146–163, November 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742100180X.

Alwuthaynani:2026:HOH

[AAL^{+}26] Maher Alwuthaynani, Muhammad Ahsan, Weidong Lei, Muhammad Abuzar, Masood Ahmad, and Aditya Sharma. A high-order Haar wavelet approach to solve differential equations of fifthorder with simple, two-point and two-point integral conditions. Applied Numerical Mathematics: Transactions of IMACS, 219(??): 122–144, January 2026. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425001783.

Amster:2003:SST

[AAM03] P. Amster, C. G. Averbuj, and M. C. Mariani. Stationary solutions for two nonlinear Black–Scholes type equations. Applied Numerical Mathematics: Transactions of IMACS, 47(3– 4):275–280, December 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Aguilar:1988:CMS

[AB88] M. Aguilar and H. Brunner. Collocation methods for second-order Volterra integro-differential equations. Applied Numerical Mathematics: Transactions of IMACS, 4(6):455– 470, November 1988. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ahmad:1997:AOA

[AB97] I. Ahmad and M. Berzins. An algorithm for ODEs from atmospheric dispersion problems. Applied Numerical Mathematics: Transactions of IMACS, 25(2–3): 137–149, October 17, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=25&
                                        issue=2-3&aid=810.

Amodio:1998:PPC

[AB98] Pierluigi Amodio and Luigi Brugnano. ParalleloGAM: a parallel code for ODEs. Applied Numerical Mathematics: Transactions of IMACS, 28(2–4):95–106, October 1, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/

store/apnum/sub/1998/28/
2-4/904.pdf.

An:2007:GCN

[AB07] Heng-Bin An and Zhong-Zhi Bai. A globally convergent Newton–GMRES method for large sparse systems of nonlinear equations. Applied Numerical Mathematics: Transactions of IMACS, 57(3):235–252, March 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Alt:2009:DEC

[AB09a] Walter Alt and Nils Bräutigam.

Discretization of elliptic control problems with time dependent parameters. Applied Numerical Mathematics: Transactions of IMACS, 59(3–4):410–423, March/ April 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Amodio:2009:PST

[AB09b] Pierluigi Amodio and Luigi Brugnano. Parallel solution in time of ODEs: some achievements and perspectives. Applied Numerical Mathematics: Transactions of IMACS, 59(3–4): 424–435, March/April 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Adjerid:2010:AEP

[AB10a] Slimane Adjerid and Mahboub Baccouch. Asymptotically exact a posteriori. Applied Numerical Mathematics: Transactions of IMACS, 60(9):903–914, September 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Amodei:2010:ISM

[AB10b] L. Amodei and J.-M. Buchot. An invariant subspace method for large-scale algebraic Riccati equation. Applied Numerical Mathematics: Transactions of IMACS, 60(11):1067–1082, November 2010. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Abdi:2012:AOA

[AB12a] A. Abdi and J. C. Butcher. Applications of order arrows. Applied Numerical Mathematics: Transactions of IMACS, 62(5): 556–566, May 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927412000098.

Arizumi:2012:ECD

[AB12b] Nana Arizumi and Stephen D. Bond. On the estimation and correction of discretization error in molecular dynamics averages. Applied Numerical Mathematics: Transactions of IMACS, 62(12): 1938–1953, December 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412001493.

Archid:2014:AME

[AB14] Atika Archid and Abdeselem Hafid Bentbib. Approximation of the matrix exponential operator by a structure-preserving block Arnoldi-type method. Applied Numerical Mathematics: Transactions of IMACS, 75(??):37–47, January 2014. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412002048.

Akram:2015:HCM

[AB15] Ghazala Akram and Christian Beck. Hierarchical cascade model leading to 7th order initial value problem. Applied Numerical Mathematics: Transactions of IMACS, 91(??): 89–97, May 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414002128.

Ableidinger:2017:WSR

[AB17] M. Ableidinger and E. Buckwar. Weak stochastic Runge–Kutta Munthe–Kaas methods for finite spin ensembles. Applied Numerical Mathematics: Transactions of IMACS, 118(??): 50–63, August 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417300326.

Ando:2024:POC

[AB24] Alessia Andò and Dimitri Breda. Piecewise orthogonal collocation for computing periodic solutions of coupled delay equations. Applied Numerical Mathematics: Transactions of IMACS, 200(??):58–79, June 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423001307.

Aulisa:2018:CEM

[ABCC18] E. Aulisa, G. Bornia, S. Calandrini, and G. Capodaglio. Convergence estimates for multigrid algorithms with SSC smoothers and applications to overlapping domain decomposition. Applied Numerical Mathematics: Transactions of IMACS, 131(??):16–38, September 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0168927418301004.

Antonopoulos:2016:CSM

[ABD16] Chris G. Antonopoulos, Tassos Bountis, and Lambros Drossos. Coupled symplectic maps as models for subdiffusive processes in disordered Hamiltonian lattices. Applied Numerical Mathematics: Transactions of IMACS, 104(??):110–119, June 2016. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415001105.

Aleotti:2024:FGL

[ABD24] Stefano Aleotti, Alessandro Buccini, and Marco Donatelli. Fractional graph Laplacian for image reconstruction. Applied Numerical Mathematics: Transactions of IMACS, 200(??): 43–57, June 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001277.

Alqahtani:2025:CPR

[ABD^{+}25] H. Alqahtani, C. F. Borges, D. Lj. Djukić, R. M. Mutavdžić Djukić, L. Reichel, and M. M. Spalević. Computation of pairs of related Gauss-type quadrature rules. Applied Numerical Mathematics: Transactions of IMACS, 208 (part A) (??):32–42, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927424000515.

Ahmad:2023:TDS

[ABdSG23] Zubair Ahmad, Giuliano Bonanomi, Daniela di Serafino, and Francesco Giannino. Transmission dynamics and sensitivity analysis of pine wilt disease with asymptomatic carriers via fractal–fractional differential operator of Mittag-Leffler kernel. Applied Numerical Mathematics: Transactions of IMACS, 185(??): 446–465, March 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422003269.

Araujo:2009:SCS

[ABF09] A. Araújo, J. R. Branco, and J. A. Ferreira. On the stability of a class of splitting methods for integrodifferential equations. Applied Numerical Mathematics: Transactions of IMACS, 59(3–4):436–453, March/ April 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Araya:2009:SAP

[ABFV09] Rodolfo Araya, Gabriel R.

Barrenechea, Leopoldo P. Franca, and Frédéric Valentin. Stabilization arising from PGEM: a review and further developments. Applied Numerical Mathematics: Transactions of IMACS, 59(9):2065–2081, September 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Auricchio:2015:FDM

[ABG^{+}15] Ferdinando Auricchio, Daniele Boffi, Lucia Gastaldi, Adrien Lefieux, and Alessandro Reali. On a fictitious domain method with distributed Lagrange multiplier for interface problems. Applied Numerical Mathematics: Transactions of IMACS, 95(??):36– 50, September 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414001238.

Abdi:2014:CSD

[ABH14] A. Abdi, M. Braś, and G. Hojjati. On the construction of second derivative diagonally implicit multistage integration methods for ODEs. Applied Numerical Mathematics: Transactions of IMACS, 76(??):1– 18, February 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927413001360.

Abdi:2022:EMB

[ABH22] A. Abdi, J.-P. Berrut, and S. A. Hosseini. Explicit methods based on barycentric rational interpolants for solving non-stiff Volterra integral equations. Applied Numerical Mathematics: Transactions of IMACS, 174(??):127–141, April 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422000046.

Amodio:2022:AHO

[ABI22] Pierluigi Amodio, Luigi Brugnano, and Felice Iavernaro. Arbitrary high-order methods for one-sided direct event location in discontinuous differential problems with nonlinear event function. Applied Numerical Mathematics: Transactions of IMACS, 179(??):39– 49, September 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422001064.

Addam:2012:SRD

[ABJ12] M. Addam, A. Bouhamidi, and K. Jbilou. Signal reconstruction for the diffusion transport equation using tensorial spline Galerkin approximation. Applied Numerical Mathematics: Transactions of IMACS, 62(9):1089–1108, September 2012. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411000729.

Agoujil:2012:SPA

[ABK12] S. Agoujil, A. H. Bentbib, and A. Kanber. A structure preserving approximation method for Hamiltonian exponential matrices. Applied Numerical Mathematics: Transactions of IMACS, 62(9):1126–1138, September 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411000535.

Aminifard:2021:AMB

[ABKG21] Zohre Aminifard, Saman Babaie-Kafaki, and Saeide Ghafoori. An augmented memoryless BFGS method based on a modified secant equation with application to compressed sensing. Applied Numerical Mathematics: Transactions of IMACS, 167(??):187–201, September 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421001239.

Argyros:2017:USL

[ABM17] Ioannis K. Argyros, Ramandeep Behl, and S. S. Motsa. Unifying semilocal and local convergence of Newton’s method on Banach space with a convergence structure. Applied Numerical Mathematics: Transactions of IMACS, 115(??): 225–234, May 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417300107.

Andreucci:1995:IPP

[ABP95] Daniele Andreucci, Monica Bianchini, and Aldo Pasquali. Identification of parameters in polymer crystallization. Applied Numerical Mathematics: Transactions of IMACS, 17(3): 191–211, September 4, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=17&
                                        issue=3&aid=576. Numeri-

cal methods for ordinary differential equations (Atlanta, GA, 1994).

Abramovici:1993:SSE

[Abr93] Flavian Abramovici. Synthetic seismograms for elastic and anelastic halfspaces. Applied Numerical Mathematics: Transactions of IMACS, 12(1– 3):3–25, May 1993. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0168927493901095.

Araya:2005:ASF

[ABR05] Rodolfo Araya, Edwin Behrens, and Rodolfo Rodríguez. An adaptive stabilized finite element scheme for the advection–reaction–diffusion equation. Applied Numerical Mathematics: Transactions of IMACS, 54(3– 4):491–503, August 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Alotaibi:2023:KSS

[ABR23] Majed Alotaibi, Alessandro Buccini, and Lothar Reichel. Krylov subspace split Bregman methods. Applied Numerical Mathematics: Transactions of IMACS, 184(??):371– 390, February 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422002793.

Auzinger:2018:SMS

[ABRW18] Winfried Auzinger, Jana Burkotová, Irena Rachunková, and Victor Wenin. Shooting methods for state-dependent impulsive boundary value problems, with applications. Applied Numerical Mathematics: Transactions of IMACS, 128(??): 217–229, June 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418300515.

Abukhaled:2004:MSS

[Abu04] Marwan I. Abukhaled. Mean square stability of secondorder weak numerical methods for stochastic differential equations. Applied Numerical Mathematics: Transactions of IMACS, 48(2):127– 134, February 2004. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Arandiga:2022:MCS

[ABY22] Francesc Aràndiga, Antonio Baeza, and Dionisio F. Yáñez. Monotone cubic spline interpolation for functions with a strong gradient. Applied Numerical Mathematics: Transactions of IMACS, 172(??): 591–607, February 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421003159.

Abedini:2021:PGF

[ABZ21] Nazanin Abedini, Ali Foroush Bastani, and Bijan Zohouri Zangeneh. A Petrov– Galerkin finite element method using polyfractonomials to solve stochastic fractional differential equations. Applied Numerical Mathematics: Transactions of IMACS, 169(??):64–86, November 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0168927421001689.

Aubry:1996:SER

[AC96] A. Aubry and P. Chartier.

On the structure of errors for Radau IA methods applied to index-2 DAEs. Applied Numerical Mathematics: Transactions of IMACS, 22(1–3):23–34, December 4, 1996. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=22&
issue=1-3&aid=705. Special

issue celebrating the centenary of Runge–Kutta methods.

Abbasian:1998:HMI

[AC98] Reza O. Abbasian and Graham F. Carey. Hybrid MPE-iterative schemes for linear and nonlinear systems. Applied Numerical Mathematics: Transactions of IMACS, 26(3):277–291, March 2, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1998/26/
                                        3/855.pdf.

Arbenz:2008:SCS

[AC08] Peter Arbenz and Oscar Chinellato. On solving complex-symmetric eigenvalue problems arising in the design of axisymmetric VCSEL devices. Applied Numerical Mathematics: Transactions of IMACS, 58(4):381–394, April 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Allouche:2010:RRD

[AC10] Hassane Allouche and Annie Cuyt. Reliable root detection with the qd-algorithm: When Bernoulli, Hadamard and Rutishauser cooperate. Applied Numerical Mathematics: Transactions of IMACS, 60(12):1188–1208, December 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Adjerid:2015:ISE

[AC15] Slimane Adjerid and Nabil Chaabane. An improved superconvergence error estimate for the LDG method. Applied Numerical Mathematics: Transactions of IMACS, 98(??):122–136, December 2015. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415001130.

Ashton:2016:NAF

[AC16] A. C. L. Ashton and K. M.

Crooks. Numerical analysis of Fokas’ unified method for linear elliptic PDEs. Applied Numerical Mathematics: Transactions of IMACS, 104(??):120–132, June 2016. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415000975.

Aouadi:2018:DCP

[AC18] M. Aouadi and M. I. M.

Copetti. A dynamic contact problem for a thermoelastic diffusion beam with the rotational inertia. Applied Numerical Mathematics: Transactions of IMACS, 126(??):113–137, April 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741730257X.

An:2023:LTP

[AC23] Congpei An and Mou Cai. Lasso trigonometric polynomial approximation for periodic function recovery in equidistant points. Applied Numerical Mathematics: Transactions of IMACS, 194(??):115–130, December 2023. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423002398.

Acary:2012:HOE

[Aca12] Vincent Acary. Higher order event capturing timestepping schemes for nonsmooth multibody systems with unilateral constraints and impacts. Applied Numerical Mathematics: Transactions of IMACS, 62(10): 1259–1275, October 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927412001122.

Abdi:2022:MSD

[ACDP22] A. Abdi, D. Conte, R. D’Ambrosio, and B. Paternoster. Multivalue second derivative collocation methods. Applied Numerical Mathematics: Transactions of IMACS, 182(??):344–355, December 2022. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422002161.

Abdalla:2024:LIO

[ACKV24] Hassan Mohamed Abdelalim Abdalla, Daniele Casagrande, Wiesław Krajewski, and Umberto Viaro. Loewner integer-order approximation of MIMO fractional-order systems. Applied Numerical Mathematics: Transactions of IMACS, 198(??): 112–121, April 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423003197.

Almuslimani:2022:UAS

[ACLM22] Ibrahim Almuslimani, Philippe Chartier, Mohammed Lemou, and Florian Méhats. Uniformly accurate schemes for drift-oscillatory stochastic differential equations. Applied Numerical Mathematics: Transactions of IMACS, 181(??):468–482, November 2022. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001751.

Augenbaum:1991:EGO

[ACM91] J. Augenbaum, S. E. Cohn, and D. Marchesin. Eliminating grid-orientation errors in alternating-direction implicit schemes. Applied Numerical Mathematics: Transactions of IMACS, 8(1):1– 10, August 1991. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ashi:2009:CME

[ACM09] H. A. Ashi, L. J. Cummings, and P. C. Matthews. Comparison of methods for evaluating functions of a matrix exponential. Applied Numerical Mathematics: Transactions of IMACS, 59(3–4): 468–486, March/April 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Alvarez:2006:CFO

[ACMR06] O. Alvarez, E. Carlini, R. Monneau, and E. Rouy. Convergence of a first order scheme for a non-local eikonal equation. Applied Numerical Mathematics: Transactions of IMACS, 56(9):1136–1146, September 2006. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Araya:2023:SFE

[ACP23] Rodolfo Araya, Cristian Cárcamo, and Abner H. Poza. A stabilized finite element method for the Stokes-Temperature coupled problem. Applied Numerical Mathematics: Transactions of IMACS, 187(??): 24–49, May 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423000338.

Aceto:2024:GTA

[ACP24] Lidia Aceto, Dajana Conte, and Giovanni Pagano. On a generalization of timeaccurate and highly-stable explicit operators for stiff problems. Applied Numerical Mathematics: Transactions of IMACS, 200(??): 2–17, June 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423000867.

Awada:2025:AAC

[ACR25] Rawaa Awada, Jérôme Carrayrou, and Carole Rosier. Anderson acceleration. Convergence analysis and applications to equilibrium chemistry. Applied Numerical Mathematics: Transactions of IMACS, 208 (part A) (??):60–75, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927424000229.

Andersen:1999:UBO

[AD99] Erling D. Andersen and Anders Damgaard. Utility based option pricing with proportional transaction costs and diversification problems: an interior-point optimization approach. Applied Numerical Mathematics: Transactions of IMACS, 29(3):395– 422, March 1, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1999/29/
3/969.pdf.

Araujo:2001:EPN

[AD01a] A. Araújo and A. Durán. Error propagation in the numerical integration of solitary waves. The regularized long wave equation. Applied Numerical Mathematics: Transactions of IMACS, 36(2– 3):197–217, February 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.nl../30/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/65/31/30/article.
pdf.

Armentano:2001:EEM

[AD01b] María G. Armentano and Ricardo G. Durán. Error estimates for moving least square approximations. Applied Numerical Mathematics: Transactions of IMACS, 37(3):397–416, May 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/73/31/
35/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/73/31/35/article.
pdf.

Athanasiadis:2004:HPS

[AD04] Aristotelis N. Athanasiadis and Herman Deconinck. Hierarchical parameter setting for hybrid grid generation based on the STEP model representation. Applied Numerical Mathematics: Transactions of IMACS, 51(4):427–436, December 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Aimi:2008:NIS

[AD08] A. Aimi and M. Diligenti. Numerical integration schemes for Petrov– Galerkin infinite BEM. Applied Numerical Mathematics: Transactions of IMACS, 58(8):1084–1102, August 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Abarbanel:2015:WPA

[AD15] S. Abarbanel and A. Ditkowski.

Wave propagation in advected acoustics within a non-uniform medium under the effect of gravity. Applied Numerical Mathematics: Transactions of IMACS, 93(??):61–68, July 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414000543.

Assari:2018:ASN

[AD18a] Pouria Assari and Mehdi Dehghan. The approximate solution of nonlinear Volterra integral equations of the second kind using radial basis functions. Applied Numerical Mathematics: Transactions of IMACS, 131(??): 140–157, September 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418301119.

Assari:2018:SCN

[AD18b] Pouria Assari and Mehdi Dehghan. Solving a class of nonlinear boundary integral equations based on the meshless local discrete Galerkin (MLDG) method. Applied Numerical Mathematics: Transactions of IMACS, 123(??):137–158, January 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417301836.

Abbaszadeh:2019:NAI

[AD19a] Mostafa Abbaszadeh and Mehdi Dehghan. Numerical and analytical investigations for neutral delay fractional damped diffusionwave equation based on the stabilized interpolating element free Galerkin (IEFG) method. Applied Numerical Mathematics: Transactions of IMACS, 145(??): 488–506, November 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927419301126.

Assari:2019:MLG

[AD19b] Pouria Assari and Mehdi Dehghan. A meshless local Galerkin method for solving Volterra integral equations deduced from nonlinear fractional differential equations using the moving least squares technique. Applied Numerical Mathematics: Transactions of IMACS, 143(??): 276–299, September 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419301011.

Abbaszadeh:2020:IPD

[AD20a] Mostafa Abbaszadeh and Mehdi Dehghan. Interior penalty discontinuous Galerkin technique for solving generalized Sobolev equation. Applied Numerical Mathematics: Transactions of IMACS, 154(??): 172–186, August 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300866.

Abbaszadeh:2020:IOM

[AD20b] Mostafa Abbaszadeh and Mehdi Dehghan. Investigation of the Oldroyd model as a generalized incompressible Navier–Stokes equation via the interpolating stabilized element free Galerkin technique. Applied Numerical Mathematics: Transactions of IMACS, 150(??): 274–294, April 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741930234X.

Abbaszadeh:2020:PBR

[AD20c] Mostafa Abbaszadeh and Mehdi Dehghan. A PODbased reduced-order Crank– Nicolson/fourth-order alternating direction implicit (ADI) finite difference scheme for solving the two-dimensional distributed-order Riesz spacefractional diffusion equation. Applied Numerical Mathematics: Transactions of IMACS, 158(??): 271–291, December 2020.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302208.

Abbaszadeh:2021:GMR

[AD21] Mostafa Abbaszadeh and Mehdi Dehghan. A Galerkin meshless reproducing kernel particle method for numerical solution of neutral delay time-space distributedorder fractional damped diffusion-wave equation. Applied Numerical Mathematics: Transactions of IMACS, 169(??):44–63, November 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0168927421001410.

Alvarez:2023:PEE

[AD23] María Luz Alvarez and Ricardo G. Durán. A posteriori error estimates for mixed approximations of degenerate elliptic problems. Applied Numerical Mathematics: Transactions of IMACS, 188(??):146–159, June 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423000673.

Almeida:2025:NAE

[ADFP25] Rui M. P. Almeida, José C. M. Duque, Jorge Ferreira, and Willian S. Panni.

Numerical analysis for an evolution equation with the p-biharmonic operator. Applied Numerical Mathematics: Transactions of IMACS, 216(??):164– 186, October 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425001047.

Almeida:2018:FES

[ADFR18] Rui M. P. Almeida, José C. M. Duque, Jorge Ferreira, and Rui J. Robalo. Finite element schemes for a class of nonlocal parabolic systems with moving boundaries. Applied Numerical Mathematics: Transactions of IMACS, 127(??):226–248, May 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418300163.

Area:2011:CZO

[ADG11] Iván Area, Dimitar K. Dimitrov, and Eduardo Godoy. Convolutions and zeros of orthogonal polynomials. Applied Numerical Mathematics: Transactions of IMACS, 61(7):868–878, July 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Akrivis:2016:PSI

[ADG^{+}16] Georgios D. Akrivis, Vassilios A. Dougalis, Efstratios Gallopoulos, Apostolos Hadjidimos, Ilias S. Kotsireas, Dimitrios Noutsos, Yiannis G. Saridakis, and Michael N. Vrahatis. Preface [Special issue: NumAn 2014]. Applied Numerical Mathematics: Transactions of IMACS, 104(??): 99–102, June 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416000301.

Aimi:2024:WIB

[ADG^{+}24] A. Aimi, G. Di Credico, H. Gimperlein, C. Guardasoni, and G. Speroni. Weak imposition of boundary conditions for the boundary element method in the time domain. Applied Numerical Mathematics: Transactions of IMACS, 200(??): 18–42, June 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001319.

Azaiez:2008:NH

[ADGM08] M. Azaïez, M. O. Deville, R. Gruber, and E. H. Mund. A new hp. Applied Numerical Mathematics: Transactions of IMACS, 58(7):985–998, July 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Aves:2000:EQD

[ADH00] Mark A. Aves, Penny J. Davies, and Desmond J. Higham. The effect of quadrature on the dynamics of a discretized nonlinear integro-differential equation. Applied Numerical Mathematics: Transactions of IMACS, 32(1):1–20, January 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/59/28/
27/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/59/28/27/article.
pdf.

Akrivis:1994:EEF

[ADK94] Georgios D. Akrivis, Vassilios A. Dougalis, and Nikolaos A. Kampanis. Error estimates for finite element methods for a wideangle parabolic equation. Applied Numerical Mathematics: Transactions of IMACS, 16(1–2):81–100, December 1, 1994. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=16&
issue=1-2&aid=522.  A

Festschrift to honor Professor Robert Vichnevetsky on his 65th birthday.

Antonopoulos:2010:NSB

[ADM10] D. C. Antonopoulos, V. A.

Dougalis, and D. E. Mitsotakis. Numerical solution of Boussinesq systems of the Bona–Smith family. Applied Numerical Mathematics: Transactions of IMACS, 60(4):314–336, April 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Almeida:2022:MFE

[ADM22] Rui M. P. Almeida, José C. M. Duque, and Belchior C. X. Mário. A mixed finite element method for a class of evolution differential equations with p-Laplacian and memory. Applied Numerical Mathematics: Transactions of IMACS, 181(??): 534–551, November 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001787.

AlMugahwi:2021:FEP

[ADNR21] Mohammed Al Mugahwi, Omar De la Cruz Cabrera, Silvia Noschese, and Lothar Reichel. Functions and eigenvectors of partially known matrices with applications to network analysis. Applied Numerical Mathematics: Transactions of IMACS, 159(??):93– 105, January 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420302555.

Allaei:2017:ACM

[ADR17] Sonia Seyed Allaei, Teresa Diogo, and Magda Rebelo. Analytical and computational methods for a class of nonlinear singular integral equations. Applied Numerical Mathematics: Transactions of IMACS, 114(??): 2–17, April 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416300940.

Aimi:2017:NPS

[ADSS17] A. Aimi, M. Diligenti, M. L. Sampoli, and A. Sestini. Non-polynomial spline alternatives in Isogeometric Symmetric Galerkin BEM. Applied Numerical Mathematics: Transactions of IMACS, 116(??):10–23, June 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416301118.

Almalki:2023:NMC

[AEA23] Nawal Hassan Almalki, Abdelhafeez Ahmed Elshekhipy, and Ibtesam Shafi Almutairi.

A numerical modeling and its computational implementing simulation for generating distributions of the complicated random variable transformations with applications. Applied Numerical Mathematics: Transactions of IMACS, 186(??):274–288, April 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423000193.

Ashyralyev:2012:DRH

[AED12] A. Ashyralyev, A. S. Erdogan, and O. Demirdag. On the determination of the right-hand side in a parabolic equation. Applied Numerical Mathematics: Transactions of IMACS, 62(11): 1672–1683, November 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412001134.

Aursand:2014:ETD

[AEF^{+}14] Peder Aursand, Steinar Evje, Tore Flåtten, Knut Erik Teigen Giljarhus, and Svend Tollak Munkejord. An exponential time-differencing method for monotonic relaxation systems. Applied Numerical Mathematics: Transactions of IMACS, 80(??):1–21, June 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414000129.

Allouche:2012:AMP

[AEG12] Hassane Allouche, Ebby Mint El Agheb, and Noura Ghanou. Adapted multivariate Padé approximation. Applied Numerical Mathematics: Transactions of IMACS, 62(9): 1061–1076, September 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927411001255.

Aggul:2023:ACM

[AEK23] Mustafa Aggul, Fatma G. Eroglu, and Songül Kaya. Artificial compression method for MHD system in Elsässer variables. Applied Numerical Mathematics: Transactions of IMACS, 185(??): 72–87, March 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422002951.

Albright:2017:HON

[AEMX17] Jason Albright, Yekaterina Epshteyn, Michael Medvinsky, and Qing Xia. Highorder numerical schemes based on difference potentials for 2D elliptic problems with material interfaces. Applied Numerical Mathematics: Transactions of IMACS, 111(??):64–91, January 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301672.

Azarnavid:2022:EKB

[AEN22] Babak Azarnavid, Mahdi Emamjomeh, and Mohammad Nabati. An efficient kernel-based method for solving nonlinear generalized Benjamin–Bona–Mahony– Burgers equation in irregular domains. Applied Numerical Mathematics: Transactions of IMACS, 181(??): 518–533, November 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001775.

Arridge:2013:PCS

[AES13] S. R. Arridge, H. Egger, and M. Schlottbom. Preconditioning of complex symmetric linear systems with applications in optical tomography. Applied Numerical Mathematics: Transactions of IMACS, 74(??):35– 48, December 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413000949.

Albright:2015:HOA

[AES15] Jason Albright, Yekaterina Epshteyn, and Kyle R. Steffen. High-order accurate difference potentials methods for parabolic problems. Applied Numerical Mathematics: Transactions of IMACS, 93(??):87–106, July 2015. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892741400141X.

Arney:1989:ALM

[AF89] David C. Arney and Joseph E. Flaherty. An adaptive local mesh refinement method for time-dependent partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 5(4):257–274, July 1989. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Arina:2004:DDT

[AF04] Renzo Arina and Marco Falossi. Domain decomposition technique for aeroacoustic simulations. Applied Numerical Mathematics: Transactions of IMACS, 49(3–4):263–275, June 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Amini:2023:AMT

[AF23] Keyvan Amini and Parvaneh Faramarzi. An adaptive modified threeterm conjugate gradient method with global convergence. Applied Numerical Mathematics: Transactions of IMACS, 190(??): 187–199, August 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001101.

Affouf:1994:NSS

[Aff94] M. Affouf. Numerical study of a singular system of conservation laws arising in enhanced oil reservoirs. Applied Numerical Mathematics: Transactions of IMACS, 15(1):1–11, August 1, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=15&
issue=1&aid=486.

Aurada:2015:ENB

[AFF^{+}15] Markus Aurada, Michael Feischl, Thomas Führer, Michael Karkulik, and Dirk Praetorius. Energy norm based error estimators for adaptive BEM for hypersingular integral equations. Applied Numerical Mathematics: Transactions of IMACS, 95(??):15– 35, September 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414000178.

Alotaibi:2024:CME

[AFIS24] Manal Alotaibi, Françoise Foucher, Moustafa Ibrahim, and Mazen Saad. Computational modeling of earlystage breast cancer progression using TPFA method: a numerical investigation. Applied Numerical Mathematics: Transactions of IMACS, 198(??):236–257, April 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424000096.

Auzinger:1992:ECR

[AFK92] W. Auzinger, R. Frank, and G. Kirlinger. An extension of B-convergence for Runge–Kutta methods. Applied Numerical Mathematics: Transactions of IMACS, 9(2):91–109, February 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Aurada:2012:CAB

[AFLG^{+}12] M. Aurada, S. Ferraz-Leite, P. Goldenits, M. Karkulik, M. Mayr, and D. Praetorius. Convergence of adaptive BEM for some mixed boundary value problem. Applied Numerical Mathematics: Transactions of IMACS, 62(4):226–245, April 2012.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411000602.

Aurada:2012:ERC

[AFLP12] Markus Aurada, Samuel Ferraz-Leite, and Dirk Praetorius. Estimator reduction and convergence of adaptive BEM. Applied Numerical Mathematics: Transactions of IMACS, 62(6): 787–801, June 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411001085.

Auzinger:1996:VCD

[AFS96] Winfried Auzinger, Reinhard Frank, and Hans J. Stetter. Vienna contributions to the development of RKmethods. Applied Numerical Mathematics: Transactions of IMACS, 22(1–3): 35–49, December 4, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=22&
issue=1-3&aid=706. Special

issue celebrating the centenary of Runge–Kutta methods.

Arevalo:2000:RSB

[AFS00] Carmen Arévalo, Claus Führer, and Gustaf Söderlind. Regular and singular βblocking of difference corrected multistep methods for nonstiff index-2 DAEs. Applied Numerical Mathematics: Transactions of IMACS, 35(4):293–305, December 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.nl../27/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/64/32/27/article.
pdf.

Arevalo:2002:CFM

[AFS02] Carmen Arévalo, Claus Führer, and Mónica Selva. A collocation formulation of multistep methods for variable step-size extensions. Applied Numerical Mathematics: Transactions of IMACS, 42(1–3):5–16, August 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Archibald:2011:ADG

[AFS11] Rick Archibald, George Fann, and William Shelton. Adaptive discontinuous Galerkin methods in multiwavelets bases. Applied Numerical Mathematics: Transactions of IMACS, 61(7):879–890, July 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Abarbanel:1998:CAA

[AG98] Saul Abarbanel and David Gottlieb. On the construction and analysis of absorbing layers in CEM. Applied Numerical Mathematics: Transactions of IMACS, 27(4):331–340, August 15, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1998/27/
4/884.pdf.

Aguilar:2005:EIA

[AG05a] Juan C. Aguilar and Jonathan B.

Goodman. An efficient interpolation algorithm on anisotropic grids for functions with jump discontinuities in 2-D. Applied Numerical Mathematics: Transactions of IMACS, 55(2):137– 153, October 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Arbenz:2005:MPI

[AG05b] Peter Arbenz and Roman Geus. Multilevel preconditioned iterative eigensolvers for Maxwell eigenvalue problems. Applied Numerical Mathematics: Transactions of IMACS, 54(2):107–121, July 2005. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Arun:2025:ESW

[AG25] K. R. Arun and R. Ghorai. An energy stable and wellbalanced scheme for the Ripa system. Applied Numerical Mathematics: Transactions of IMACS, 216(??): 187–209, October 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742500114X.

Akrivis:2010:P

[AGH^{+}10] Georgios Akrivis, Efstratios Gallopoulos, Apostolos Hadjidimos, Ilias S. Kotsireas, Dimitrios Noutsos, and Michael N. Vrahatis. Preface. Applied Numerical Mathematics: Transactions of IMACS, 60(4): 293–297, April 2010. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ahmed:2012:EPI

[AGJ12] Sarfraz Ahmed, Christopher E. Goodyer, and Peter K. Jimack. An efficient preconditioned iterative solution of fully-coupled elastohydrodynamic lubrication problems. Applied Numerical Mathematics: Transactions of IMACS, 62(5): 649–663, May 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412000281.

Arraras:2025:STP

[AGJCP25] Andrés Arrarás, Francisco J.

Gaspar, Iñigo Jimenez-Ciga, and Laura Portero. Space–time parallel solvers for reaction–diffusion problems forming Turing patterns. Applied Numerical Mathematics: Transactions of IMACS, 218(??):91– 108, December 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425001539.

Apel:2004:NMA

[AGJM04] Thomas Apel, Sergei Grosman, Peter K. Jimack, and Arnd Meyer. A new methodology for anisotropic mesh refinement based upon error gradients. Applied Numerical Mathematics: Transactions of IMACS, 50(3–4): 329–341, September 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Arun:2024:APE

[AGK24] K. R. Arun, Rahuldev Ghorai, and Mainak Kar. An asymptotic preserving and energy stable scheme for the Euler–Poisson system in the quasineutral limit. Applied Numerical Mathematics: Transactions of IMACS, 198(??):375–400, April 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000187.

Asarin:1994:SAN

[AGKK94] E. Asarin, E. Gorin, M. Krasnosel′skiĭ, and N. Kuznetsov. On some algorithms for nonparametric identification of linear systems. Applied Numerical Mathematics: Transactions of IMACS, 16(1–2):265–269, December 1, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=16&
issue=1-2&aid=514.  A

Festschrift to honor Professor Robert Vichnevetsky on his 65th birthday.

Alanazi:2025:CCD

[AGKS25a] K. Alanazi, S. D. Georgiou, C. Koukouvinos, and S. Stylianou. Central composite designs with three missing observations. Applied Numerical Mathematics: Transactions of IMACS, 208 (part A)(??): 2–21, February 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927423003215.

Alhelali:2025:ODC

[AGKS25b] Omar A. Alhelali, S. D.

Georgiou, C. Koukouvinos, and S. Stylianou. Orthogonal designs for computer experiments constructed from sequences with zero autocorrelation. Applied Numerical Mathematics: Transactions of IMACS, 208 (part A) (??):22–31, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927423002544.

Acevedo:2023:NAF

[AGLRS23] Ramiro Acevedo, Christian Gómez, Bibiana López-Rodríguez, and Pilar Salgado. Numerical analysis of a FEM based on a time-primitive of the electric field for solving a nonlinear transient eddy current problem. Applied Numerical Mathematics: Transactions of IMACS, 192(??): 261–279, October 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001745.

Amodio:1995:VSB

[AGM95] P. Amodio, W. L. Golik, and F. Mazzia. Variablestep boundary value methods based on reverse Adams schemes and their grid redistribution. Applied Numerical Mathematics: Transactions of IMACS, 18(1–3): 5–21, September 27, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=18&
issue=1-3&aid=592. Sev-

enth Conference on the Numerical Treatment of Differential Equations (Halle, 1994).

Aceto:2009:BVM

[AGM09] L. Aceto, P. Ghelardoni, and C. Magherini. Boundary Value Methods as an extension of Numerov’s method for Sturm–Liouville eigenvalue estimates. Applied Numerical Mathematics: Transactions of IMACS, 59(7):1644–1656, July 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Alonso:1997:BEA

[AGP97] P. Alonso, M. Gasca, and J. M. Peña. Backward error analysis of Neville elimination. Applied Numerical Mathematics: Transactions of IMACS, 23(2):193–204, March 21, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1997/23/
2/733.pdf.

Argyros:2024:LSL

[AGQ^{+}24] Ioannis K. Argyros, Krzysztof Gdawiec, Sania Qureshi, Amanullah Soomro, Evren Hincal, and Samundra Regmi. Local and semi-local convergence and dynamic analysis of a time-efficient nonlinear technique. Applied Numerical Mathematics: Transactions of IMACS, 201(??): 446–464, July 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000837.

Aguilar:2015:HON

[Agu15] Juan C. Aguilar. Higherorder Newton–Cotes rules with end corrections. Applied Numerical Mathematics: Transactions of IMACS, 88(??):66–77, February 2015. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414001664.

Archibald:2008:DLD

[AGY08] Rick Archibald, Anne Gelb, and Jungho Yoon. Determining the locations and discontinuities in the derivatives of functions. Applied Numerical Mathematics: Transactions of IMACS, 58(5):577–592, May 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Abo-Gabal:2022:FRJ

[AGZD22] Howayda Abo-Gabal, Mahmoud A. Zaky, and Eid H. Doha. Fractional Romanovski– Jacobi tau method for timefractional partial differential equations with nonsmooth solutions. Applied Numerical Mathematics: Transactions of IMACS, 182(??): 214–234, December 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422001908.

Abdelwahed:2009:TOM

[AH09] Mohamed Abdelwahed and Maatoug Hassine. Topological optimization method for a geometric control problem in Stokes flow. Applied Numerical Mathematics: Transactions of IMACS, 59(8):1823–1838, August 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Abdi:2011:MOS

[AH11] A. Abdi and G. Hojjati. Maximal order for second derivative general linear methods with Runge–Kutta stability. Applied Numerical Mathematics: Transactions of IMACS, 61(10): 1046–1058, October 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Abdi:2015:INS

[AH15] A. Abdi and G. Hojjati. Implementation of Nordsieck second derivative methods for stiff ODEs. Applied Numerical Mathematics: Transactions of IMACS, 94(??):241–253, August 2015. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415000628.

Amiri:2017:SRK

[AH17] Sadegh Amiri and S. Mohammad Hosseini. Stochastic Runge–Kutta Rosenbrock type methods for SDE systems. Applied Numerical Mathematics: Transactions of IMACS, 115(??): 1–15, May 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416302355.

Akbari:2023:ASI

[AHA23] R. Akbari, G. Hojjati, and A. Abdi. Algebraic stability and irreducibility of second derivative methods. Applied Numerical Mathematics: Transactions of IMACS, 191(??):45– 54, September 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423001332.

Al-Homidan:2021:GSF

[AHAS21] Suliman Al-Homidan, Bashir Ali, and Yusuf I. Suleiman. Generalized split feasibility problem for multi-valued Bregman quasi-nonexpansive mappings in Banach spaces. Applied Numerical Mathematics: Transactions of IMACS, 161(??):437–451, March 2021. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420303743.

Amiri:2020:ASF

[AHB20] Sadegh Amiri, Mojtaba Hajipour, and Dumitru Baleanu. On accurate solution of the Fredholm integral equations of the second kind. Applied Numerical Mathematics: Transactions of IMACS, 150(??): 478–490, April 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419303046.

Ait-Haddou:2021:QLD

[AHGM21] Rachid Ait-Haddou, Ron Goldman, and Marie-Laurence Mazure. Quantum Lorentz degrees of polynomials and a Pólya theorem for polynomials positive on q-lattices. Applied Numerical Mathematics: Transactions of IMACS, 165(??):553–577, July 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000799.

Abdi:2023:IET

[AHJ^{+}23] A. Abdi, G. Hojjati, Z. Jackiewicz, H. Podhaisky, and M. Sharifi. On the implementation of explicit twostep peer methods with Runge–Kutta stability. Applied Numerical Mathematics: Transactions of IMACS, 186(??):213–227, April 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423000156.

Abdi:2019:NCV

[AHJM19] A. Abdi, G. Hojjati, Z. Jackiewicz, and H. Mahdi. A new code for Volterra integral equations based on natural Runge–Kutta methods. Applied Numerical Mathematics: Transactions of IMACS, 143(??):35– 50, September 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419300698.

Ahn:2007:VSD

[Ahn07] Jeongho Ahn. A vibrating string with dynamic frictionless impact. Applied Numerical Mathematics: Transactions of IMACS, 57(8):861– 884, August 2007. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Alejandre:2016:SEE

[AHO16] Martin P. Arciga Alejandre, Francisco J. Ariza Hernandez, and Jorge Sanchez Ortiz. Stochastic evolution equation with Rieszfractional derivative and white noise on the halfline. Applied Numerical Mathematics: Transactions of IMACS, 104(??): 103–109, June 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927415000884.

Amat:2012:SCS

[AHR12] S. Amat, M. A. Hernández, and N. Romero. Semilocal convergence of a sixth order iterative method for quadratic equations. Applied Numerical Mathematics: Transactions of IMACS, 62(7):833–841, July 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412000463.

Ansari:2003:NFO

[AHS03] A. R. Ansari, A. F. Hegarty, and G. I. Shishkin. A note on fitted operator methods for a laminar jet problem. Applied Numerical Mathematics: Transactions of IMACS, 45(4):353–365, June 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Aghda:2017:ANN

[AHT17] A. S. Fatemion Aghda, Seyed Mohammad Hosseini, and Mahdieh Tahmasebi. Analysis of non-negativity and convergence of solution of the balanced implicit method for the delay Cox–Ingersoll–Ross model. Applied Numerical Mathematics: Transactions of IMACS, 118(??):249–265, August 2017. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741730079X.

Ahmad:2019:HPM

[AI19] Iftikhar Ahmad and Hira Ilyas. Homotopy Perturbation Method for the nonlinear MHD Jeffery–Hamel blood flows problem. Applied Numerical Mathematics: Transactions of IMACS, 141(??):124–132, ???? 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418301533.

Abdi:2019:TCN

[AJ19] A. Abdi and Z. Jackiewicz. Towards a code for nonstiff differential systems based on general linear methods with inherent Runge–Kutta stability. Applied Numerical Mathematics: Transactions of IMACS, 136(??): 103–121, February 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418302265.

Acosta:2024:HOE

[AJ24a] Claudia Acosta and Silvia Jerez. High-order entropy stable schemes based on a local adaptive WENO viscosity for degenerate convection-diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 197(??): 322–343, March 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423002945.

Ankur:2024:NEE

[AJ24b] Ankur and Ram Jiwari. A new error estimates of finite element method for (2 + 1)-dimensional nonlinear advection–diffusion model. Applied Numerical Mathematics: Transactions of IMACS, 198(??): 22–42, April 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423003136.

Arora:2020:SBB

[AJK20] Shelly Arora, Rajiv Jain, and V. K. Kukreja. Solution of Benjamin–Bona–Mahony– Burgers equation using collocation method with quintic Hermite splines. Applied Numerical Mathematics: Transactions of IMACS, 154(??):1–16, August 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300829.

Alanazi:2019:NSS

[AJT19] Khalaf M. Alanazi, Zdzislaw Jackiewicz, and Horst R. Thieme. Numerical simulations of the spread of rabies in two-dimensional space. Applied Numerical Mathematics: Transactions of IMACS, 135(??): 87–98, January 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418301776.

Aljohani:2023:FOS

[AJW23] M. A. Aljohani, P. K. Jimack, and M. A. Walkley. A faster optimal solver for thin film flows. Applied Numerical Mathematics: Transactions of IMACS, 184(??): 357–370, February 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422002823.

Auzinger:1995:KRC

[AK95] W. Auzinger and G. Kirlinger. Kreiss resolvent conditions and strengthened Cauchy–Schwarz inequalities. Applied Numerical Mathematics: Transactions of IMACS, 18(1–3): 57–67, September 27, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=18&
                                        issue=1-3&aid=591. Sev-

enth Conference on the Numerical Treatment of Differential Equations (Halle, 1994).

Ainsworth:2000:ATV

[AK00] Mark Ainsworth and David Kay. Approximation theory for the hp-version finite element method and application to the non-linear Laplacian. Applied Numerical Mathematics: Transactions of IMACS, 34(4):329– 344, August 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/62/32/
27/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/62/32/27/article.
pdf.

Aksoylu:2009:FPB

[AK09] Burak Aksoylu and Hector Klie. A family of physicsbased preconditioners for solving elliptic equations on highly heterogeneous media. Applied Numerical Mathematics: Transactions of IMACS, 59(6):1159–1186, June 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Aggul:2021:DDC

[AK21] Mustafa Aggul and Songül Kaya. Defect-deferred correction method based on a subgrid artificial viscosity model for fluid-fluid interaction. Applied Numerical Mathematics: Transactions of IMACS, 160(??): 178–191, February 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420303111.

Awwal:2019:MCG

[AKA19] Aliyu Muhammed Awwal, Poom Kumam, and Auwal Bala Abubakar. A modified conjugate gradient method for monotone nonlinear equations with convex constraints. Applied Numerical Mathematics: Transactions of IMACS, 145(??): 507–520, November 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419301199.

Alzahrani:2019:FOT

[AKBF19] S. S. Alzahrani, A. Q. M. Khaliq, T. A. Biala, and K. M. Furati. Fourthorder time stepping methods with matrix transfer technique for spacefractional reaction-diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 146(??): 123–144, December 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419301783.

Adimurthi:2014:SOS

[AKG14] Adimurthi, K. Sudarshan Kumar, and G. D. Veerappa Gowda. Second order scheme for scalar conservation laws with discontinuous flux. Applied Numerical Mathematics: Transactions of IMACS, 80(??):46–64, June 2014. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741400018X.

Al-Khaleel:2014:MAO

[AKGR14] Mohammad Al-Khaleel, Martin J. Gander, and Albert E. Ruehli. A mathematical analysis of optimized waveform relaxation for a small RC circuit. Applied Numerical Mathematics: Transactions of IMACS, 75(??):61–76, January 2014. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413000068.

Apel:2008:SFE

[AKL08] Thomas Apel, Tobias Knopp, and Gert Lube. Stabilized finite element methods with anisotropic mesh refinement for the Oseen problem. Applied Numerical Mathematics: Transactions of IMACS, 58(12): 1830–1843, December 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Avila:2020:CTG

[ÁKM20] Andrés I. Ávila, Stefan Kopecz, and Andreas Meister. A comprehensive theory on generalized BBKS schemes. Applied Numerical Mathematics: Transactions of IMACS, 157(??):19– 37, November 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420301719.

Abbaszadeh:2021:SFF

[AKM^{+}21] D. Abbaszadeh, M. Tavassoli Kajani, M. Momeni, M. Zahraei, and M. Maleki. Solving fractional Fredholm integro-differential equations using Legendre wavelets. Applied Numerical Mathematics: Transactions of IMACS, 166(??):168–185, August 2021. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001070.

Abubakar:2022:HAF

[AKM^{+}22] Auwal Bala Abubakar, Poom Kumam, Hassan Mohammad, Abdulkarim Hassan Ibrahim, and Aliyu Ibrahim Kiri. A hybrid approach for finding approximate solutions to constrained nonlinear monotone operator equations with applications. Applied Numerical Mathematics: Transactions of IMACS, 177(??):79–92, July 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422000630.

Aslan:2021:FNM

[AKS21] Ersin Aslan, Ömür Kivanç Kürkçü, and Mehmet Sezer. A fast numerical method for fractional partial integro-differential equations with spatial-time delays. Applied Numerical Mathematics: Transactions of IMACS, 161(??): 525–539, March 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303809.

Angermann:1997:EEF

[AKT97] Lutz Angermann, Peter Knabner, and Kathrin Thiele. An error estimator for a finite volume discretization of density driven flow in porous media. Applied Numerical Mathematics: Transactions of IMACS, 26(1–2):179–191, December 22, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1997/26/
1-2/845.pdf.

Ames:1987:CDN

[AL87] William F. Ames and Ding Lee. Current development in the numerical treatment of ocean acoustic propagation. Applied Numerical Mathematics: Transactions of IMACS, 3(1–2): 25–47, May 1987. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0168927487900043.

Ahues:1995:TNA

[AL95] Mario Ahués and Alain Largillier. Two numerical approximations for a class of weakly singular integral operators. Applied Numerical Mathematics: Transactions of IMACS, 17(4):347– 362, September 11, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=17&
issue=4&aid=620.

Apel:1998:AMR

[AL98] Thomas Apel and Gert Lube. Anisotropic mesh refinement for a singularly perturbed reaction diffusion model problem. Applied Numerical Mathematics: Transactions of IMACS, 26(4):415–433, April 1, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cas/
tree/store/apnum/sub/1998/
26/4/862.pdf.

Awanou:2005:CRA

[AL05] G. M. Awanou and M. J.

Lai. On convergence rate of the augmented Lagrangian algorithm for nonsymmetric saddle point problems. Applied Numerical Mathematics: Transactions of IMACS, 54(2):122–134, July 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Adjerid:2009:DIF

[AL09] Slimane Adjerid and Tao Lin. A p-th degree immersed finite element for boundary value problems with discontinuous coefficients. Applied Numerical Mathematics: Transactions of IMACS, 59(6):1303–1321, June 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

An:2017:EAF

[AL17] Rong An and Yuan Li.

Error analysis of firstorder projection method for time-dependent magnetohydrodynamics equations. Applied Numerical Mathematics: Transactions of IMACS, 112(??): 167–181, February 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416302045.

Ahrens:2020:TTM

[AL20] Robin Ahrens and Sabine Le Borne. Tensor trains and moment conservation for multivariate aggregation in population balance modeling. Applied Numerical Mathematics: Transactions of IMACS, 153(??):473–491, July 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300696.

Antoine:2022:SWR

[AL22] Xavier Antoine and Emmanuel Lorin. A Schwarz waveform relaxation method for time-dependent space fractional Schrödinger/heat equations. Applied Numerical Mathematics: Transactions of IMACS, 182(??): 248–264, December 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001854.

Appleby:2024:MSA

[AL24] John A. D. Appleby and Emmet Lawless. Mean square asymptotic stability characterisation of perturbed linear stochastic functional differential equations. Applied Numerical Mathematics: Transactions of IMACS, 200(??):80–109, June 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001885.

Albrecht:1996:CBT

[Alb96] Peter Albrecht. The common basis of the theories of linear cyclic methods and Runge–Kutta methods. Applied Numerical Mathematics: Transactions of IMACS, 22(1–3):3–21, December 4, 1996. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=22&
issue=1-3&aid=728. Special

issue celebrating the centenary of Runge–Kutta methods.

Alexander:2003:DID

[Ale03]  Roger Alexander.  De-

sign and implementation of DIRK integrators for stiff systems. Applied Numerical Mathematics: Transactions of IMACS, 46(1):1–17, July 2003. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Alekseenko:2011:NPH

[Ale11] A. M. Alekseenko. Numerical properties of high order discrete velocity solutions to the BGK kinetic equation. Applied Numerical Mathematics: Transactions of IMACS, 61(4):410–427, April 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Allouch:2024:CMC

[All24] Chafik Allouch. Collocation and modified collocation methods for solving second kind Fredholm integral equations in weighted spaces. Applied Numerical Mathematics: Transactions of IMACS, 198(??):202–216, April 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424000114.

Angulo:2004:NIF

[ALM04] O. Angulo and J. C. López-Marcos. Numerical integration of fully nonlinear size-structured population models. Applied Numerical Mathematics: Transactions of IMACS, 50(3–4): 291–327, September 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Abia:1996:BSR

[ALMM96] Luis M. Abia, J. C. López-Marcos, and Julia Martínez. Blow-up for semidiscretizations of reaction-diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 20(1–2): 145–156, February 29, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=1-2&aid=670. Work-

shop on the method of lines for time-dependent problems (Lexington, KY, 1995).

Abia:1998:BTC

[ALMM98] Luis M. Abia, J. C. López-Marcos, and Julia Martínez. On the blow-up time convergence of semidiscretizations of reaction–diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 26(4):399– 414, April 1, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1998/26/
4/861.pdf.

Abia:2001:EMN

[ALMM01] L. M. Abia, J. C. López-Marcos, and J. Martínez. The Euler method in the numerical integration of reaction-diffusion problems with blow-up. Applied Numerical Mathematics: Transactions of IMACS, 38(3): 287–313, August 2001. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/gej-

                                        ng/10/10/28/77/33/30/abstract.
                                        html.

Ashby:1996:CSC

[ALP^{+}96] S. F. Ashby, S. L. Lee, L. R. Petzold, P. E. Saylor, and E. Seidel. Computing spacetime curvature via differential algebraic equations. Applied Numerical Mathematics: Transactions of IMACS, 20(1–2): 221–234, February 29, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=20&
                                        issue=1-2&aid=676. Work-

shop on the method of lines for time-dependent problems (Lexington, KY, 1995).

Alt:1985:CRP

[Alt85] René Alt. Computing roots of polynomials on vector processing machines. Applied Numerical Mathematics: Transactions of IMACS, 1(4):299–308, July 1985. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Allegretto:2003:NPA

[ALY03] Walter Allegretto, Yanping Lin, and Hongtao Yang. Numerical pricing of American put options on zero-coupon bonds. Applied Numerical Mathematics: Transactions of IMACS, 46(2):113– 134, August 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

An:2021:STS

[ALZ^{+}21] Xingyu An, Fawang Liu, Minling Zheng, Vo V. Anh, and Ian W. Turner. A space-time spectral method for time-fractional Black– Scholes equation. Applied Numerical Mathematics: Transactions of IMACS, 165(??):152–166, July 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000593.

Abia:1995:RKM

[AM95a] L. M. Abia and J. C. López Marcos. Runge–Kutta methods for age-structured population models. Applied Numerical Mathematics: Transactions of IMACS, 17(1):1–17, May 1, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=17&
issue=1&aid=558.

Amodio:1995:BVM

[AM95b] P. Amodio and F. Mazzia.

Boundary value methods based on Adams-type methods. Applied Numerical Mathematics: Transactions of IMACS, 18(1–3):23–35, September 27, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=18&
issue=1-3&aid=589. Sev-

enth Conference on the Numerical Treatment of Differential Equations (Halle, 1994).

Alonso-Mallo:1999:ESS

[AM99] Isaías Alonso-Mallo. Explicit single step methods with optimal order of convergence for partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 31(2): 117–131, October 23, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1999&volume=31&
issue=2&aid=997;   http:
//www.sciencedirect.com/
science/article/pii/S0168927498001329.

Alonso-Mallo:2000:RMO

[AM00] Isaías Alonso-Mallo. Rational methods with optimal order of convergence for partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 35(4): 265–292, December 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.nl../26/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/64/32/26/article.
pdf.

Ascher:2004:MBS

[AM04] Uri M. Ascher and Robert I.

McLachlan. Multisymplectic box schemes and the Korteweg–de Vries equation. Applied Numerical Mathematics: Transactions of IMACS, 48(3–4):255–269, March 2004. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Apel:2005:P

[AM05] Thomas Apel and Arnd Meyer. Preface. Applied Numerical Mathematics: Transactions of IMACS, 54(3–4):311–313, August 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Azaiez:2008:P

[AM08] Mejdi Azaiez and Ernest H.

Mund. Preface. Applied Numerical Mathematics: Transactions of IMACS, 58(7):933–934, July 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Acosta:2009:ASH

[AM09] Carlos D. Acosta and Carlos E. Mejía. Approximate solution of hyperbolic conservation laws by discrete mollification. Applied Numerical Mathematics: Transactions of IMACS, 59(9): 2256–2265, September 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Annunziato:2010:NTV

[AM10a] Mario Annunziato and Eleonora Messina. Numerical treatment of a Volterra integral equation with space maps. Applied Numerical Mathematics: Transactions of IMACS, 60(8):809– 815, August 2010. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Assous:2010:HDS

[AM10b] Franck Assous and Michael Michaeli. Hodge decomposition to solve singular static Maxwell’s equations in a non-convex polygon. Applied Numerical Mathematics: Transactions of IMACS, 60(4):432–441, April 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Armentano:2016:IJW

[AM16a] María G. Armentano and Verónica Moreno. Interpolation in Jacobi-weighted spaces and its application to a posteriori error estimations of the pversion of the finite element method. Applied Numerical Mathematics: Transactions of IMACS, 109(??): 184–207, November 2016. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301039.

Atisattapong:2016:OWE

[AM16b] Wanyok Atisattapong and Pasin Maruphanton. Obviating the bin width effect of the 1/t algorithm for multidimensional numerical integration. Applied Numerical Mathematics: Transactions of IMACS, 104(??): 133–140, June 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415001373.

Alonso-Mallo:2002:SRD

[AMC02] I. Alonso-Mallo and B. Cano.

Spectral/Rosenbrock discretizations without order reduction for linear parabolic problems. Applied Numerical Mathematics: Transactions of IMACS, 41(2): 247–268, May 2002. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Alonso-Mallo:2008:OTO

[AMCM08] I. Alonso-Mallo, B. Cano, and M. J. Moreta. Optimal time order when implicit Runge–Kutta–Nyström methods solve linear partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 58(5):539–562, May 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Alonso-Mallo:2009:SRA

[AMCM09] I. Alonso-Mallo, B. Cano, and M. J. Moreta. The stability of rational approximations of cosine functions on Hilbert spaces. Applied Numerical Mathematics: Transactions of IMACS, 59(1):21–38, January 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Alonso-Mallo:2017:AOR

[AMCR17] I. Alonso-Mallo, B. Cano, and N. Reguera. Analysis of order reduction when integrating linear initial boundary value problems with Lawson methods. Applied Numerical Mathematics: Transactions of IMACS, 118(??):64–74, August 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417300557.

Arndt:2003:CEM

[AMH03] Marcos Arndt, Roberto Dalledone Machado, and Mildred Ballin Hecke. The composite element method applied to free vibration analysis of trusses and beams. Applied Numerical Mathematics: Transactions of IMACS, 47(2):59–73, November 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Al-Musawi:2024:FEA

[AMH24] Ghufran A. Al-Musawi and Akil J. Harfash. Finite element analysis of extended Fisher–Kolmogorov equation with Neumann boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 201(??): 41–71, July 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000321.

Al-Maskari:2018:LMF

[AMK18] Mariam Al-Maskari and Samir Karaa. The lumped mass FEM for a timefractional cable equation. Applied Numerical Mathematics: Transactions of IMACS, 132(??):73–90, October 2018. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418301223.

Adak:2024:NVE

[AMN24] Dibyendu Adak, Gianmarco Manzini, and Sundararajan Natarajan. Nonconforming virtual element methods for velocity-pressure Stokes eigenvalue problem. Applied Numerical Mathematics: Transactions of IMACS, 202(??):42–66, August 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742400093X.

Alonso-Mallo:2003:OOC

[AMP03] I. Alonso-Mallo and C. Palencia. Optimal orders of convergence for Runge– Kutta methods and linear, initial boundary value problems. Applied Numerical Mathematics: Transactions of IMACS, 44(1–2): 1–19, January 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Arias:2020:GIQ

[AMP20] C. A. Arias, H. J. Martínez, and R. Pérez. Global inexact quasi-Newton method for nonlinear system of equations with constraints. Applied Numerical Mathematics: Transactions of IMACS, 150(??):559–575, April 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419303101.

Arandiga:2012:NST

[AMR12] Francesc Aràndiga, Pep Mulet, and Vicent Renau. Non-separable twodimensional weighted ENO interpolation. Applied Numerical Mathematics: Transactions of IMACS, 62(8): 975–987, August 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412000530.

Amdouni:2014:LPS

[AMR14] S. Amdouni, M. Moakher, and Y. Renard. A local projection stabilization of fictitious domain method for elliptic boundary value problems. Applied Numerical Mathematics: Transactions of IMACS, 76(??):60– 75, February 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413001256.

Abudawia:2018:FEM

[AMRR18] A. Abudawia, A. Mourad, J. H. Rodrigues, and C. Rosier. A finite element method for a seawater intrusion problem in unconfined aquifers. Applied Numerical Mathematics: Transactions of IMACS, 127(??):349–369, May 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418300229.

Avila:2014:NMN

[ÁMS14] A. I. Ávila, A. Meister, and M. Steigemann. On numerical methods for nonlinear singularly perturbed Schrödinger problems. Applied Numerical Mathematics: Transactions of IMACS, 86(??):22– 42, December 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414001202.

Avila:2017:AGM

[ÁMS17] A. I. Ávila, A. Meister, and M. Steigemann. An adaptive Galerkin method for the time-dependent complex Schrödinger equation. Applied Numerical Mathematics: Transactions of IMACS, 121(??):149–169, November 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417301496.

Adjerid:2005:P

[AMT05] Slimane Adjerid, Peter Moore, and James D.

Teresco. Preface. Applied Numerical Mathematics: Transactions of IMACS, 52(2–3):129–132, February 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ahmed:2013:SFE

[AMT13] Naveed Ahmed, Gunar Matthies, and Lutz Tobiska. Stabilized finite element discretization applied to an operator-splitting method of population balance equations. Applied Numerical Mathematics: Transactions of IMACS, 70(??): 58–79, August 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413000524.

Aceto:2017:MAH

[AMT17] Lidia Aceto, Helmut Robert Malonek, and Graça Tomaz. Matrix approach to hypercomplex Appell polynomials. Applied Numerical Mathematics: Transactions of IMACS, 116(??): 2–9, June 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301325.

Ardenghi:2003:TRA

[AMV03] Juan Ignacio Ardenghi, María Cristina Maciel, and Adriana Beatriz Verdiell. A trust-region-approach for solving a parameter estimation problem from the biotechnology area. Applied Numerical Mathematics: Transactions of IMACS, 47(3–4):281–294, December 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Alves:2017:EMF

[AMV17] Carlos J. S. Alves, Nuno F. M. Martins, and Svilen S. Valtchev. Extending the method of fundamental solutions to non-homogeneous elastic wave problems. Applied Numerical Mathematics: Transactions of IMACS, 115(??):299–313, May 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416300952.

Arandiga:2015:RCI

[AN15] Francesc Aràndiga and José Jaime Noguera. Reconstructions that combine interpolation with least squares fitting. Applied Numerical Mathematics: Transactions of IMACS, 97(??):30– 41, November 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415000999.

An:2016:LGS

[An16]  Jing An.  A Legendre–

Galerkin spectral approximation and estimation of the index of refraction for transmission eigenvalues. Applied Numerical Mathematics: Transactions of IMACS, 108(??):171– 184, October 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416300873.

An:2020:EAT

[An20] Rong An. Error analysis of a time-splitting method for incompressible flows with variable density. Applied Numerical Mathematics: Transactions of IMACS, 150(??): 384–395, April 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419302909.

Avijit:2022:NTS

[AN22] D. Avijit and S. Natesan.

A novel two-step streamlinediffusion FEM for singularly perturbed 2D parabolic PDEs. Applied Numerical Mathematics: Transactions of IMACS, 172(??): 259–278, February 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421002956.

Abdulghafoor:2025:NSG

[AN25] Salwan Tareq Abdulghafoor and Esmaeil Najafi. Numerical solution for a generalized form of nonlinear cordial Volterra integral equations using quasilinearization and Legendre-collocation methods. Applied Numerical Mathematics: Transactions of IMACS, 207(??): 384–399, January 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424002496.

Ang:2006:NMW

[Ang06] Whye-Teong Ang. A numerical method for the wave equation subject to a nonlocal conservation condition. Applied Numerical Mathematics: Transactions of IMACS, 56(8):1054–1060, August 2006. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Adak:2019:VEM

[ANN19] D. Adak, S. Natarajan, and E. Natarajan. Virtual element method for semilinear elliptic problems on polygonal meshes. Applied Numerical Mathematics: Transactions of IMACS, 145(??): 175–187, November 2019.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927419301369.

Anonymous:1987:EB

[Ano87]  Anonymous.  Editorial

board. Applied Numerical Mathematics: Transactions of IMACS, 3(1– 2):ii, May 1987. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0168927487900018.

Anonymous:1991:EB

[Ano91]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 7(1):ii– iv, January 1991. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/016892749190099L.

Anonymous:1992:MIV

[Ano92]  Anonymous.  Master in-

dex to Volumes 1–10. Applied Numerical Mathematics: Transactions of IMACS, 10(6):511–532, November 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:1993:EB

[Ano93]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 12(1– 3):ii, May 1993. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0168927493901073.

Anonymous:2000:IC

[Ano00a] Anonymous. ICOSAHOM 1998 Conference. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4):1, May 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/61/27/
27/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/61/27/27/article.
pdf.

Anonymous:2000:Ia

[Ano00b] Anonymous. Index. Applied Numerical Mathematics: Transactions of IMACS, 32(4):495–496, April 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/59/32/
36/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/59/32/36/article.
pdf.

Anonymous:2000:Ib

[Ano00c] Anonymous. Index. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4):541–543, May 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/61/27/
84/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/61/27/84/article.
pdf.

Anonymous:2000:Id

[Ano00d] Anonymous. Index. Applied Numerical Mathematics: Transactions of IMACS, 34(4):421–422, August 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/62/32/
33/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/62/32/33/article.
pdf.

Anonymous:2000:Ie

[Ano00e] Anonymous. Index. Applied Numerical Mathematics: Transactions of IMACS, 35(4):357–358, December 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.nl../31/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/64/32/31/article.
pdf.

Anonymous:2000:Pa

[Ano00f] Anonymous. Preface. Applied Numerical Mathematics: Transactions of IMACS, 32(4):359–360, April 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/21/59/32/
                                        27/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/21/59/32/27/article.
                                        pdf.

Anonymous:2000:Pb

[Ano00g] Anonymous. Preface. Applied Numerical Mathematics: Transactions of IMACS, 34(2–3):143, July 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/21/62/29/
                                        27/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/21/62/29/27/article.
                                        pdf.

Anonymous:2000:VR

[Ano00h] Anonymous. V.S. Ryaben’kii. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4):479, May 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/21/61/27/
                                        76/abstract.html;   http:
                                        //www.elsevier.nl/gej-

ng/29/17/21/61/27/76/article.
pdf.

Anonymous:2001:AIVa

[Ano01a] Anonymous. Author index — Volume 36 (2001). Applied Numerical Mathematics: Transactions of IMACS, 36(4):491–492, March 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/65/32/
32/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/65/32/32/article.
pdf.

Anonymous:2001:AIVb

[Ano01b] Anonymous. Author index — Volume 37 (2001). Applied Numerical Mathematics: Transactions of IMACS, 37(4):551–552, June 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/73/33/
34/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/73/33/34/article.
pdf.

Anonymous:2001:AIVc

[Ano01c] Anonymous. Author index — volume 38 (2001). Applied Numerical Mathematics: Transactions of IMACS, 38(4):499–500, September 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.com/
gej-ng/10/10/28/77/34/
32/abstract.html.

Anonymous:2001:AIVd

[Ano01d] Anonymous. Author index — volume 39 (2001). Applied Numerical Mathematics: Transactions of IMACS, 39(3–4):461–462, December 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/79/34/39/abstract.
html.

Anonymous:2001:C

[Ano01e] Anonymous. Contents. Applied Numerical Mathematics: Transactions of IMACS, 37(1–2):v–vi, April 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/73/26/
26/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/73/26/26/article.
pdf.

Anonymous:2002:AIVa

[Ano02a] Anonymous. Author index — volume 40 (2002). Applied Numerical Mathematics: Transactions of IMACS, 40(4):483–485, March 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/gej-

ng/10/10/28/85/34/30/abstract.
html.

Anonymous:2002:AIVb

[Ano02b] Anonymous. Author index — volume 41 (2002). Applied Numerical Mathematics: Transactions of IMACS, 41(4):529–530, July 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2002:AIVc

[Ano02c] Anonymous. Author index — volume 42 (2002). Applied Numerical Mathematics: Transactions of IMACS, 42(4):545–546, September 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2002:AIVd

[Ano02d] Anonymous. Author index — volume 43 (2002). Applied Numerical Mathematics: Transactions of IMACS, 43(4):501–502, December 2002. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2002:LIL

[Ano02e] Anonymous. List of invited lectures. Applied Numerical Mathematics: Transactions of IMACS, 42(1–3): 3–4, August 2002. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2002:MIV

[Ano02f]  Anonymous.  Master in-

dex — volume 31–40. Applied Numerical Mathematics: Transactions of IMACS, 40(4):487–508, March 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/gej-
                                        ng/10/10/28/85/34/31/abstract.
                                        html.

Anonymous:2002:PBC

[Ano02g] Anonymous. Proceedings of the 19th Biennial Conference on Numerical Analysis, University of Dundee. Applied Numerical Mathematics: Transactions of IMACS, 43(1–2):3–7, October 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2003:AIVa

[Ano03a] Anonymous. Author index — volume 44 (2003). Applied Numerical Mathematics: Transactions of IMACS, 44(4):527–528, March 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2003:AIVb

[Ano03b] Anonymous. Author index — volume 45 (2003). Applied Numerical Mathematics: Transactions of IMACS, 45(4):513–514, June 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2003:AIVc

[Ano03c] Anonymous. Author index — volume 46 (2003). Applied Numerical Mathematics: Transactions of IMACS, 46(3–4):445–446, September 2003. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2003:AIVd

[Ano03d] Anonymous. Author index — volume 47 (2003). Applied Numerical Mathematics: Transactions of IMACS, 47(3–4):587–589, December 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2003:EBa

[Ano03e]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 47(2): ??, November 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2003:EBb

[Ano03f]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 47(3–4): CO2, December 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2004:AIVb

[Ano04a] Anonymous. Author index — volume 50 (2004). Applied Numerical Mathematics: Transactions of IMACS, 50(3–4):521–522, September 2004. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2004:AIVc

[Ano04b] Anonymous. Author index — volume 51 (2004). Applied Numerical Mathematics: Transactions of IMACS, 51(4):581–583, December 2004. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2004:AIVa

[Ano04c] Anonymous. Author index— volume 49 (2004). Applied Numerical Mathematics: Transactions of IMACS, 49(3–4):441–442, June 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2004:EBa

[Ano04d]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 48(2): CO2, February 2004. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2004:EBb

[Ano04e]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 48(3–4): CO2, March 2004. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2004:EBc

[Ano04f]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 49(2):CO2, May 2004. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2004:EBd

[Ano04g]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 49(3– 4):CO2, June 2004. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2004:EBe

[Ano04h]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 50(2): CO2, August 2004. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2004:EBf

[Ano04i]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 50(3– 4):CO2, September 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2004:EBg

[Ano04j]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 51(2–3): CO2, November 2004. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2004:EBh

[Ano04k]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 51(4): CO2, December 2004. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2004:MIV

[Ano04l]  Anonymous.  Master in-

dex to volumes 41–50. Applied Numerical Mathematics: Transactions of IMACS, 51(1):127–149, October 2004. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2004:PN

[Ano04m]  Anonymous.  Publisher’s

note. Applied Numerical Mathematics: Transactions of IMACS, 48(3–4): v, March 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2005:AIVa

[Ano05a] Anonymous. Author index — volume 52 (2005). Applied Numerical Mathematics: Transactions of IMACS, 52(4):459–460, March 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2005:AIVb

[Ano05b] Anonymous. Author index — volume 53 (2005). Applied Numerical Mathematics: Transactions of IMACS, 53(2–4):543–545, May 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2005:C

[Ano05c] Anonymous. Contents. Applied Numerical Mathematics: Transactions of IMACS, 52(2–3):iii–iv, February 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2005:EBa

[Ano05d]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 52(4): CO2, March 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2005:EBb

[Ano05e]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 53(2– 4):CO2, May 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2005:IL

[Ano05f] Anonymous. Invited lectures. Applied Numerical Mathematics: Transactions of IMACS, 53(2–4): 91–92, May 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2010:P

[Ano10] Anonymous. Preface. Applied Numerical Mathematics: Transactions of IMACS, 60(11):1053, November 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2011:EBa

[Ano11a]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 61(8):??, August 2011. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2011:EBb

[Ano11b]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 61(9): ??, September 2011. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2011:EBc

[Ano11c]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 61(11): i, November 2011. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2011:EBd

[Ano11d]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 61(12): ??, December 2011. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411001747.

Anonymous:2012:EBa

[Ano12a]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 62(1): ??, January 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411001887.

Anonymous:2012:EBb

[Ano12b]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 62(2): ??, February 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927411001978.

Anonymous:2012:EBc

[Ano12c]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 62(3): ??, March 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927412000037.

Anonymous:2012:EBd

[Ano12d]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 62(4): ??, April 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927412000141.

Anonymous:2012:EBe

[Ano12e]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 62(5): ??, May 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412000323.

Anonymous:2012:EBf

[Ano12f]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 62(6): ??, June 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412000384.

Anonymous:2012:EBg

[Ano12g]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 62(7): ??, July 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412000566.

Anonymous:2012:EBh

[Ano12h]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 62(8): ??, August 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412000669.

Anonymous:2012:EBi

[Ano12i]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 62(9): ??, September 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927412000797.

Anonymous:2012:EBj

[Ano12j]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 62(10): ??, October 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927412001262.

Anonymous:2012:EBk

[Ano12k]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 62(11): ??, November 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927412001353.

Anonymous:2012:EBl

[Ano12l]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 62(12): ??, December 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412001535.

Anonymous:2013:EBa

[Ano13a]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 63(1): ??, January 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741200181X.

Anonymous:2013:EBb

[Ano13b]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 64(1): ??, February 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412001870.

Anonymous:2013:EBc

[Ano13c]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 65(??): ??, March 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412002176.

Anonymous:2013:EBd

[Ano13d]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 66(??): ??, April 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927413000081.

Anonymous:2013:EBe

[Ano13e]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 67(??): ??, May 2013. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927413000202.

Anonymous:2013:EBf

[Ano13f]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 68(??): ??, June 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927413000317.

Anonymous:2013:EBg

[Ano13g]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 69(??): ??, July 2013. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413000469.

Anonymous:2013:EBh

[Ano13h]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 70(??): ??, August 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741300055X.

Anonymous:2013:EBi

[Ano13i]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 71(??): ??, September 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741300069X.

Anonymous:2013:EBj

[Ano13j]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 72(??): ??, October 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413000998.

Anonymous:2013:EBk

[Ano13k]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 73(??): ??, November 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927413001086.

Anonymous:2014:EBa

[Ano14a]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 83(??): ifc, September 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414000890.

Anonymous:2014:EBb

[Ano14b]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 84(??): ifc, October 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414001111.

Anonymous:2014:EBc

[Ano14c]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 85(??): IFC, November 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414001342.

Anonymous:2014:EBd

[Ano14d]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 86(??): IFC, December 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414001445.

Anonymous:2015:EBa

[Ano15a]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 87(??): IFC, January 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414001688.

Anonymous:2015:EBb

[Ano15b]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 88(??): IFC, February 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414001809.

Anonymous:2015:EBc

[Ano15c]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 89(??): IFC, March 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414002025.

Anonymous:2015:EBd

[Ano15d]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 90(??): IFC, April 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415000069.

Anonymous:2015:EBe

[Ano15e]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 91(??): IFC, May 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415000240.

Anonymous:2015:EBf

[Ano15f]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 92(??): IFC, June 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415000318.

Anonymous:2015:EBg

[Ano15g]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 93(??): IFC, July 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415000410.

Anonymous:2015:EBh

[Ano15h]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 94(??): ifc, August 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415000653.

Anonymous:2015:EBi

[Ano15i]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 95(??): ifc, September 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741500080X.

Anonymous:2015:EBj

[Ano15j]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 96(??): ifc, October 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415001026.

Anonymous:2015:EBk

[Ano15k]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 97(??): ifc, November 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415001191.

Anonymous:2015:EBl

[Ano15l]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 98(??): ifc, December 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415001397.

Anonymous:2016:EBm

[Ano16a]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 99(??): ifc, January 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741500149X.

Anonymous:2016:EBn

[Ano16b]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 100(??): ifc, February 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415001634.

Anonymous:2016:EBo

[Ano16c]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 101(??): ifc, March 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415001749.

Anonymous:2016:EBp

[Ano16d]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 102(??): ifc, April 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416000076.

Anonymous:2016:EBq

[Ano16e]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 103 (??):ifc, May 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416000167.

Anonymous:2016:EBr

[Ano16f]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 104 (??):ifc, June 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416300125.

Anonymous:2016:EBs

[Ano16g]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 105 (??):ifc, July 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416300459.

Anonymous:2016:EBt

[Ano16h]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 106(??): ifc, August 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416300630.

Anonymous:2016:EBu

[Ano16i]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 107(??): ifc, September 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416300745.

Anonymous:2016:EBv

[Ano16j]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 108(??): ifc, October 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301222.

Anonymous:2016:EBw

[Ano16k]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 109(??): ifc, November 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301489.

Anonymous:2016:EBx

[Ano16l]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 110(??): ifc, December 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301805.

Anonymous:2017:EBa

[Ano17a]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 111(??): ifc, January 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416302069.

Anonymous:2017:EBb

[Ano17b]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 112(??): ifc, February 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416302252.

Anonymous:2017:EBc

[Ano17c]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 113(??): ifc, March 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416302410.

Anonymous:2017:EBd

[Ano17d]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 114(??): ifc, April 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417300120.

Anonymous:2017:EBe

[Ano17e]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 115 (??):ifc, May 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417300442.

Anonymous:2017:EBf

[Ano17f]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 116 (??):ifc, June 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417300697.

Anonymous:2017:EBg

[Ano17g]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 117 (??):ifc, July 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417300934.

Anonymous:2017:EBh

[Ano17h]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 118(??): ifc, August 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417301125.

Anonymous:2017:EBi

[Ano17i]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 119(??): ifc, September 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417301265.

Anonymous:2017:EBj

[Ano17j]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 120(??): ifc, October 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417301538.

Anonymous:2017:EB

[Ano17k]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 122(??): ifc, December 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417301903.

Anonymous:2018:EBa

[Ano18a]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 123(??): ifc, January 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741730212X.

Anonymous:2018:EBb

[Ano18b]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 125(??): ifc, March 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417302441.

Anonymous:2018:EBc

[Ano18c]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 126(??): ifc, April 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418300023.

Anonymous:2018:EBd

[Ano18d]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 127 (??):ifc, May 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418300357.

Anonymous:2018:EBe

[Ano18e]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 128 (??):ifc, June 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418300618.

Anonymous:2018:EBf

[Ano18f]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 129 (??):ifc, July 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418300795.

Anonymous:2018:EBg

[Ano18g]  Anonymous.  Editorial

board. Applied Numerical Mathematics: Transactions of IMACS, 130(??): ifc, August 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741830103X.

Anonymous:2018:EBh

[Ano18h]  Anonymous.  Editorial

board. Applied Numerical Mathematics: Transactions of IMACS, 131(??): ifc, September 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418301259.

Anonymous:2018:EBi

[Ano18i]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 132(??): ifc, October 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418301417.

Anonymous:2018:EBj

[Ano18j]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 133(??): ifc, ???? 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418301636.

Anonymous:2018:EBk

[Ano18k]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 134(??): ifc, December 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418301806.

Anonymous:2019:EBa

[Ano19a]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 135(??): ifc, January 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418302162.

Anonymous:2019:EBb

[Ano19b]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 136(??): ifc, February 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418302447.

Anonymous:2019:EBc

[Ano19c]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 137(??): ifc, March 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418302733.

Anonymous:2019:EBd

[Ano19d]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 138(??): ifc, April 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419300121.

Anonymous:2019:EBe

[Ano19e]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 139 (??):ifc, May 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419300303.

Anonymous:2019:EBf

[Ano19f]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 140 (??):ifc, June 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419300492.

Anonymous:2019:EBg

[Ano19g]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 141(??): ifc, ???? 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419300728.

Anonymous:2019:EBh

[Ano19h]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 142(??): ifc, August 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741930090X.

Anonymous:2019:EBi

[Ano19i]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 143(??): ifc, September 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419301242.

Anonymous:2019:EBj

[Ano19j]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 144(??): ifc, October 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927419301448.

Anonymous:2019:EBk

[Ano19k]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 145(??): ifc, November 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927419301965.

Anonymous:2019:EBl

[Ano19l]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 146(??): ifc, December 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S016892741930220X.

Anonymous:2020:EBa

[Ano20a]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 147(??): ifc, January 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419302594.

Anonymous:2020:EBb

[Ano20b]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 148(??): ifc, February 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419302715.

Anonymous:2020:EBc

[Ano20c]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 149(??): ifc, ???? 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419303484.

Anonymous:2020:EBd

[Ano20d]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 150(??): ifc, April 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419303575.

Anonymous:2020:EBe

[Ano20e]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 151 (??):ifc, May 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300283.

Anonymous:2020:EBf

[Ano20f]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 152 (??):ifc, June 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300477.

Anonymous:2020:EBg

[Ano20g]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 153 (??):ifc, July 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300969.

Anonymous:2020:EBh

[Ano20h]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 154(??): ifc, August 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420301343.

Anonymous:2020:EBi

[Ano20i]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 155(??): ifc, ???? 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420301021.

Anonymous:2020:EBj

[Ano20j]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 156(??): ifc, October 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420301756.

Anonymous:2020:EBk

[Ano20k]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 157(??): ifc, November 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302324.

Anonymous:2020:EBl

[Ano20l]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 158(??): ifc, December 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302701.

Anonymous:2021:EBa

[Ano21a]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 159(??): ifc, January 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303020.

Anonymous:2021:EBb

[Ano21b]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 160(??): ifc, February 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742030338X.

Anonymous:2021:EBc

[Ano21c]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 161(??): ifc, March 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303974.

Anonymous:2021:EBd

[Ano21d]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 162(??): ifc, April 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000143.

Anonymous:2021:EBe

[Ano21e]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 163 (??):ifc, May 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000520.

Anonymous:2021:EBf

[Ano21f]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 164(??): ifc, ???? 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000258.

Anonymous:2021:EBg

[Ano21g]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 165 (??):ifc, July 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000891.

Anonymous:2021:EBh

[Ano21h]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 166(??): ifc, August 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001252.

Anonymous:2021:EBi

[Ano21i]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 167(??): ifc, September 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001598.

Anonymous:2021:EBj

[Ano21j]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 168(??): ifc, October 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421001859.

Anonymous:2021:EBk

[Ano21k]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 169(??): ifc, November 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421002178.

Anonymous:2021:EBl

[Ano21l]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 170(??): ifc, December 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742100249X.

Anonymous:2021:PO

[Ano21m]  Anonymous.  Pages 1–

292 (October 2021). Applied Numerical Mathematics: Transactions of IMACS, 168(??):1–292, October 2021. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2021:PM

[Ano21n] Anonymous. Pages 1–316 (May 2021). Applied Numerical Mathematics: Transactions of IMACS, 163(??): 1–316, May 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2021:PAb

[Ano21o]  Anonymous.  Pages 1–

368 (August 2021). Applied Numerical Mathematics: Transactions of IMACS, 166(??):1–368, August 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2021:PAa

[Ano21p]  Anonymous.  Pages 1–

404 (April 2021). Applied Numerical Mathematics: Transactions of IMACS, 162(??):1–404, April 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2021:PN

[Ano21q] Anonymous. Pages 1–414 (November 2021). Applied Numerical Mathematics: Transactions of IMACS, 169(??):1–414, November 2021. CODEN ANMAEL.

ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2021:PS

[Ano21r] Anonymous. Pages 1–420 (September 2021). Applied Numerical Mathematics: Transactions of IMACS, 167(??):1–420, September 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2021:PD

[Ano21s] Anonymous. Pages 1–434 (December 2021). Applied Numerical Mathematics: Transactions of IMACS, 170(??):1–434, December 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2021:PJ

[Ano21t] Anonymous. Pages 1–622 (July 2021). Applied Numerical Mathematics: Transactions of IMACS, 165(??): 1–622, July 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2022:EBa

[Ano22a]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 171(??): ifc, January 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421002816.

Anonymous:2022:EBb

[Ano22b]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 172(??): ifc, February 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421003226.

Anonymous:2022:EBc

[Ano22c]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 173(??): ifc, March 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422000162.

Anonymous:2022:EBd

[Ano22d]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 174(??): ifc, April 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422000332.

Anonymous:2022:EBe

[Ano22e]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 175 (??):ifc, May 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422000551.

Anonymous:2022:EBf

[Ano22f]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 176 (??):ifc, June 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422000800.

Anonymous:2022:EBg

[Ano22g]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 177 (??):ifc, July 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422000897.

Anonymous:2022:EBh

[Ano22h]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 178(??): ifc, August 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001180.

Anonymous:2022:EBi

[Ano22i]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 179(??): ifc, September 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001362.

Anonymous:2022:EBj

[Ano22j]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 180(??): ifc, October 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001544.

Anonymous:2022:EBk

[Ano22k]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 181(??): ifc, November 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422002057.

Anonymous:2022:EBl

[Ano22l]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 182(??): ifc, December 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422002343.

Anonymous:2022:PO

[Ano22m]  Anonymous.  Pages 1–

120 (October 2022). Applied Numerical Mathematics: Transactions of IMACS, 180(??):1–120, October 2022. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2022:PMb

[Ano22n] Anonymous. Pages 1–132 (May 2022). Applied Numerical Mathematics: Transactions of IMACS, 175(??): 1–132, May 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2022:PJc

[Ano22o] Anonymous. Pages 1–170 (July 2022). Applied Numerical Mathematics: Transactions of IMACS, 177(??): 1–170, July 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2022:PJb

[Ano22p] Anonymous. Pages 1–182 (June 2022). Applied Numerical Mathematics: Transactions of IMACS, 176(??): 1–182, June 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2022:PAa

[Ano22q]  Anonymous.  Pages 1–

236 (April 2022). Applied Numerical Mathematics: Transactions of IMACS, 174(??):1–236, April 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2022:PS

[Ano22r] Anonymous. Pages 1–272 (September 2022). Applied Numerical Mathematics: Transactions of IMACS, 179(??):1–272, September 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2022:PD

[Ano22s] Anonymous. Pages 1–370 (December 2022). Applied Numerical Mathematics: Transactions of IMACS, 182(??):1–370, December 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2022:PAb

[Ano22t]  Anonymous.  Pages 1–

404 (August 2022). Applied Numerical Mathematics: Transactions of IMACS, 178(??):1–404, August 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2022:PJa

[Ano22u]  Anonymous.  Pages 1–

452 (January 2022). Applied Numerical Mathematics: Transactions of IMACS, 171(??):1–452, January 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2022:PMa

[Ano22v]  Anonymous.  Pages 1–

452 (March 2022). Applied Numerical Mathematics: Transactions of IMACS, 173(??):1–452, March 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2022:PF

[Ano22w]  Anonymous.  Pages 1–

628 (February 2022). Applied Numerical Mathematics: Transactions of IMACS, 172(??):1–628, February 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2022:PN

[Ano22x] Anonymous. Pages 1–664 (November 2022). Applied Numerical Mathematics: Transactions of IMACS, 181(??):1–664, November 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2023:EBa

[Ano23a]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 183(??): ifc, January 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742200263X.

Anonymous:2023:EBb

[Ano23b]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 184(??): ifc, February 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422003142.

Anonymous:2023:EBc

[Ano23c]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 185(??): ifc, March 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423000223.

Anonymous:2023:EBd

[Ano23d]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 186(??): ifc, April 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423000478.

Anonymous:2023:EBe

[Ano23e]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 187 (??):ifc, May 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423000727.

Anonymous:2023:EBf

[Ano23f]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 188 (??):ifc, June 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423000880.

Anonymous:2023:EBg

[Ano23g]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 189 (??):ifc, July 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423001137.

Anonymous:2023:EBh

[Ano23h]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 190(??): ifc, August 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001435.

Anonymous:2023:EBi

[Ano23i]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 191(??): ifc, September 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001629.

Anonymous:2023:EBj

[Ano23j]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 192(??): ifc, October 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423002131.

Anonymous:2023:EBk

[Ano23k]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 193(??): ifc, November 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423002295.

Anonymous:2023:EBl

[Ano23l]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 194(??): ifc, December 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423002489.

Anonymous:2023:PD

[Ano23m] Anonymous. Pages 1–130 (December 2023). Applied Numerical Mathematics: Transactions of IMACS, 194(??):1–130, December 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2023:PJc

[Ano23n] Anonymous. Pages 1–150 (July 2023). Applied Numerical Mathematics: Transactions of IMACS, 189(??): 1–150, July 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2023:PJb

[Ano23o] Anonymous. Pages 1–160 (June 2023). Applied Numerical Mathematics: Transactions of IMACS, 188(??): 1–160, June 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2023:PN

[Ano23p] Anonymous. Pages 1–262 (November 2023). Applied Numerical Mathematics: Transactions of IMACS, 193(??):1–262, November 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2023:PMb

[Ano23q] Anonymous. Pages 1–294 (May 2023). Applied Numerical Mathematics: Transactions of IMACS, 187(??): 1–294, May 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2023:PAb

[Ano23r]  Anonymous.  Pages 1–

344 (August 2023). Applied Numerical Mathematics: Transactions of IMACS, 190(??):1–344, August 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2023:PJa

[Ano23s]  Anonymous.  Pages 1–

368 (January 2023). Applied Numerical Mathematics: Transactions of IMACS, 183(??):1–368, January 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2023:PAa

[Ano23t]  Anonymous.  Pages 1–

410 (April 2023). Applied Numerical Mathematics: Transactions of IMACS, 186(??):1–410, April 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2023:PO

[Ano23u]  Anonymous.  Pages 1–

472 (October 2023). Applied Numerical Mathematics: Transactions of IMACS, 192(??):1–472, October 2023. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2023:PF

[Ano23v]  Anonymous.  Pages 1–

580 (February 2023). Applied Numerical Mathematics: Transactions of IMACS, 184(??):1–580, February 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2023:PMa

[Ano23w]  Anonymous.  Pages 1–

592 (March 2023). Applied Numerical Mathematics: Transactions of IMACS, 185(??):1–592, March 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2023:PSa

[Ano23x] Anonymous. Pages 1–74 (September 2023). Applied Numerical Mathematics: Transactions of IMACS, 191(??):1–74, September 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2024:EBa

[Ano24a]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 195(??): ii, January 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423002647.

Anonymous:2024:EBb

[Ano24b]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 196(??): ii, February 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423002878.

Anonymous:2024:EBc

[Ano24c]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 197(??): ii, March 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423003082.

Anonymous:2024:EBd

[Ano24d]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 198 (??):ii, April 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000400.

Anonymous:2024:EBe

[Ano24e]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 199 (??):ii, May 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000680.

Anonymous:2024:EBf

[Ano24f]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 200 (??):ii, June 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000746.

Anonymous:2024:EBg

[Ano24g]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 201 (??):ii, July 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001028.

Anonymous:2024:EBh

[Ano24h]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 202(??): ii, August 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001314.

Anonymous:2024:EBi

[Ano24i]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 203(??): ii, September 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001557.

Anonymous:2024:EBj

[Ano24j]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 204(??): ii, October 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742400182X.

Anonymous:2024:EBk

[Ano24k]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 205(??): ii, November 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424002150.

Anonymous:2024:EB

[Ano24l]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 206(??): ii, December 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424002393.

Anonymous:2024:PJa

[Ano24m]  Anonymous.  Pages 1–

156 (January 2024). Applied Numerical Mathematics: Transactions of IMACS, 195(??):1–156, January 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2024:PF

[Ano24n]  Anonymous.  Pages 1–

218 (February 2024). Applied Numerical Mathematics: Transactions of IMACS, 196(??):1–218, February 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2024:PAb

[Ano24o]  Anonymous.  Pages 1–

264 (August 2024). Applied Numerical Mathematics: Transactions of IMACS, 202(??):1–264, August 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2024:PS

[Ano24p] Anonymous. Pages 1–276 (September 2024). Applied Numerical Mathematics: Transactions of IMACS, 203(??):1–276, September 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2024:PN

[Ano24q] Anonymous. Pages 1–308 (November 2024). Applied Numerical Mathematics: Transactions of IMACS, 205(??):1–308, November 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2024:PD

[Ano24r] Anonymous. Pages 1–358 (December 2024). Applied Numerical Mathematics: Transactions of IMACS, 206(??):1–358, December 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2024:PO

[Ano24s]  Anonymous.  Pages 1–

380 (October 2024). Applied Numerical Mathematics: Transactions of IMACS, 204(??):1–380, October 2024. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2024:PM

[Ano24t]  Anonymous.  Pages 1–

388 (March 2024). Applied Numerical Mathematics: Transactions of IMACS, 197(??):1–388, March 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2024:PAa

[Ano24u]  Anonymous.  Pages 1–

508 (April 2024). Applied Numerical Mathematics: Transactions of IMACS, 198(??):1–508, April 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2024:PJb

[Ano24v] Anonymous. Pages 1–642 (July 2024). Applied Numerical Mathematics: Transactions of IMACS, 201(??): 1–642, July 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2025:EBa

[Ano25a]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 207(??): ii, January 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424002836.

Anonymous:2025:EBb

[Ano25b]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 208 (part A)(??):ii, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927424003064.

Anonymous:2025:EBc

[Ano25c]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 208 (part B)(??):ii, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424003295.

Anonymous:2025:EBd

[Ano25d]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 209(??): ii, March 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424003398.

Anonymous:2025:EBe

[Ano25e]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 210 (??):ii, April 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425000194.

Anonymous:2025:EBf

[Ano25f]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 211 (??):ii, May 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742500042X.

Anonymous:2025:EBg

[Ano25g]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 212 (??):ii, June 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425000595.

Anonymous:2025:EBh

[Ano25h]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 213 (??):ii, July 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425000765.

Anonymous:2025:EBi

[Ano25i]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 214(??): ii, August 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425000911.

Anonymous:2025:EBj

[Ano25j]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 215(??): ii, September 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425001096.

Anonymous:2025:EBk

[Ano25k]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 216(??): ii, October 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425001369.

Anonymous:2025:EBl

[Ano25l]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 217(??): ii, November 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425001631.

Anonymous:2025:EBm

[Ano25m]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 218(??): ii, December 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425001874.

Anonymous:2025:PJc

[Ano25n] Anonymous. Pages 1–108 (July 2025). Applied Numerical Mathematics: Transactions of IMACS, 213(??): 1–108, July 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2025:PAb

[Ano25o]  Anonymous.  Pages 1–

160 (August 2025). Applied Numerical Mathematics: Transactions of IMACS, 214(??):1–160, August 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2025:PS

[Ano25p] Anonymous. Pages 1–176 (September 2025). Applied Numerical Mathematics: Transactions of IMACS, 215(??):1–176, September 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2025:PO

[Ano25q]  Anonymous.  Pages 1–

222 (October 2025). Applied Numerical Mathematics: Transactions of IMACS, 216(??):1–222, October 2025. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2025:PMb

[Ano25r] Anonymous. Pages 1–228 (May 2025). Applied Numerical Mathematics: Transactions of IMACS, 211(??): 1–228, May 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2025:PAa

[Ano25s]  Anonymous.  Pages 1–

244 (April 2025). Applied Numerical Mathematics: Transactions of IMACS, 210(??):1–244, April 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2025:PMa

[Ano25t]  Anonymous.  Pages 1–

284 (March 2025). Applied Numerical Mathematics: Transactions of IMACS, 209(??):1–284, March 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2025:PD

[Ano25u] Anonymous. Pages 1–298 (December 2025). Applied Numerical Mathematics: Transactions of IMACS, 218(??):1–298, December 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2025:PJb

[Ano25v] Anonymous. Pages 1–344 (June 2025). Applied Numerical Mathematics: Transactions of IMACS, 212(??): 1–344, June 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2025:PF

[Ano25w]  Anonymous.  Pages 1–

364 (February 2025). Applied Numerical Mathematics: Transactions of IMACS, 208 (part B)(??):1–364, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2025:PN

[Ano25x] Anonymous. Pages 1–476 (November 2025). Applied Numerical Mathematics: Transactions of IMACS, 217(??):1–476, November 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2025:PJa

[Ano25y]  Anonymous.  Pages 1–

656 (January 2025). Applied Numerical Mathematics: Transactions of IMACS, 207(??):1–656, January 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Anonymous:2026:EBa

[Ano26a]  Anonymous.  Editorial

Board. Applied Numerical Mathematics: Transactions of IMACS, 219(??): ii, January 2026. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425002028.

Anonymous:2026:PJa

[Ano26b]  Anonymous.  Pages 1–

238 (January 2026). Applied Numerical Mathematics: Transactions of IMACS, 219(??):1–238, January 2026. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Antonopoulou:2013:GMP

[Ant13] D. C. Antonopoulou. Galerkin methods for the ‘Parabolic Equation’ Dirichlet problem in a variable 2-D and 3-D topography. Applied Numerical Mathematics: Transactions of IMACS, 67(??): 17–34, May 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411000699.

Anton:2023:EPS

[Ant23] Cristina Anton. Explicit pseudo-symplectic Runge– Kutta methods for stochastic Hamiltonian systems. Applied Numerical Mathematics: Transactions of IMACS, 185(??):18–37, March 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422003002.

Anton:2025:NCS

[Ant25] Cristina Anton. A new class of symplectic methods for stochastic Hamiltonian systems. Applied Numerical Mathematics: Transactions of IMACS, 208 (part A) (??):43–59, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927424000217.

Averick:1991:SNP

[AO91] B. M. Averick and J. M.

Ortega. Solution of nonlinear Poisson-type equations. Applied Numerical Mathematics: Transactions of IMACS, 8(6):443– 455, December 1991. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Almendral:2005:NVO

[AO05] Ariel Almendral and Cornelis W. Oosterlee. Numerical valuation of options with jumps in the underlying. Applied Numerical Mathematics: Transactions of IMACS, 53(1):1–18, April 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ainsworth:1994:PEE

[AOW94] M. Ainsworth, J. T. Oden, and W. Wu. A posteriori error estimation for hp approximations in elastostatics. Applied Numerical Mathematics: Transactions of IMACS, 14(1– 3):23–54, April 14, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=14&
issue=1-3&aid=458. Pro-

ceedings of the Third ARO Workshop on Adaptive Methods for Partial Differential Equations (Troy, NY, 1992).

Armentano:2008:PEE

[AP08] María G. Armentano and Claudio Padra. A posteriori error estimates for the Steklov eigenvalue problem. Applied Numerical Mathematics: Transactions of IMACS, 58(5):593–601, May 2008. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Acharya:2016:PHM

[AP16] Sanjib Kumar Acharya and Ajit Patel. Primal hybrid method for parabolic problems. Applied Numerical Mathematics: Transactions of IMACS, 108(??): 102–115, October 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416300812.

Acharya:2020:PHF

[AP20] Sanjib Kumar Acharya and Kamana Porwal. Primal hybrid finite element method for fourth order parabolic problems. Applied Numerical Mathematics: Transactions of IMACS, 152(??): 12–28, June 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300222.

Ames:1992:ONA

[APA92] W. F. Ames, F. V. Postell, and E. Adams. Optimal numerical algorithms. Applied Numerical Mathematics: Transactions of IMACS, 10(3–4):235–259, September 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). A Festschrift to honor Professor Garrett Birkhoff on his eightieth birthday.

Arraras:2009:CMF

[APJ09] A. Arrarás, L. Portero, and J. C. Jorge. A combined mixed finite element ADI scheme for solving Richards’ equation with mixed derivatives on irregular grids. Applied Numerical Mathematics: Transactions of IMACS, 59(3–4):454–467, March/ April 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Arraras:2010:CFS

[APJ10] A. Arrarás, L. Portero, and J. C. Jorge. Convergence of fractional step mimetic finite difference discretizations for semilinear parabolic problems. Applied Numerical Mathematics: Transactions of IMACS, 60(4): 473–485, April 2010. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Auteri:2000:SIM

[AQ00] F. Auteri and L. Quartapelle. Spectral influence matrix for vorticity without corner pathology. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4):135–142, May 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/21/61/27/
                                        37/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/21/61/27/37/article.
                                        pdf.

Abrarov:2020:RAS

[AQ20] Sanjar M. Abrarov and Brendan M. Quine. A rational approximation of the sinc function based on sampling and the Fourier transforms. Applied Numerical Mathematics: Transactions of IMACS, 150(??):65–75, April 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419302399.

Abrarov:2018:SBA

[AQJ18] Sanjar M. Abrarov, Brendan M. Quine, and Rajinder K. Jagpal. A samplingbased approximation of the complex error function and its implementation without poles. Applied Numerical Mathematics: Transactions of IMACS, 129(??): 181–191, July 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418300710.

Agoshkov:1994:RDN

[AQS94] V. I. Agoshkov, A. Quarteroni, and F. Saleri. Recent developments in the numerical simulation of shallow water equations. I. boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 15(2):175– 200, September 13, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=15&
issue=2&aid=490. Innova-

tive methods in numerical analysis (Bressanone, 1992).

Achuthan:1993:RCF

[AR93] P. Achuthan and R. Rangarajan. On regular Cfraction and general Tfraction expansions for ratios of basic hypergeometric series and Ramanujan-type identities. Applied Numerical Mathematics: Transactions of IMACS, 12(4): 297–313, June 1993. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Aristova:2015:BSM

[AR15] E. N. Aristova and B. V. Rogov. Bicompact scheme for the multidimensional stationary linear transport equation. Applied Numerical Mathematics: Transactions of IMACS, 93(??): 3–14, July 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414000695.

Araya:2018:PEE

[AR18] Rodolfo Araya and Ramiro Rebolledo. An a posteriori error estimator for a LPS method for Navier–Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 127(??): 179–195, May 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418300096.

Alkilayh:2023:SNA

[AR23] Maged Alkilayh and Lothar Reichel. Some numerical aspects of Arnoldi– Tikhonov regularization. Applied Numerical Mathematics: Transactions of IMACS, 185(??):503–515, March 2023. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422003312.

Arandiga:2024:APS

[AR24] Francesc Aràndiga and Sara Remogna. Approximation of piecewise smooth functions by nonlinear bivariate C^{2} quartic spline quasi-interpolants on crisscross triangulations. Applied Numerical Mathematics: Transactions of IMACS, 203(??):69–83, September 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0168927424001247.

Arantes:1999:COS

[Ara99] Ricardo D. Arantes. Code orderings for the solution of sparse positive definite linear systems. Applied Numerical Mathematics: Transactions of IMACS, 30(1):137– 140, May 10, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1999/30/
1/955.pdf.

Archer:2006:CSG

[Arc06] Rosalind Archer. C_{1} continuous solutions from the Green element method using Overhauser elements. Applied Numerical Mathematics: Transactions of IMACS, 56(2):222–229, February 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Abril-Raymundo:2000:APM

[ARGA00] M. R. Abril-Raymundo and B. García-Archilla. Approximation properties of a mapped Chebyshev method. Applied Numerical Mathematics: Transactions of IMACS, 32(2):119– 136, February 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/10/28/59/27/
                                        27/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/10/28/59/27/27/article.
                                        pdf.

Arioka:1987:PTA

[Ari87] Shigekatsu Arioka. Padétype approximants in multivariables. Applied Numerical Mathematics: Transactions of IMACS, 3(6):497– 511, November 1987. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Arina:2003:OCS

[Ari03] Renzo Arina. Orthogonal and conformal surface grid generation. Applied Numerical Mathematics: Transactions of IMACS, 46(3–4):

249–262, September 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Arina:2004:NSN

[Ari04] Renzo Arina. Numerical simulation of near-critical fluids. Applied Numerical Mathematics: Transactions of IMACS, 51(4): 409–426, December 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Arnold:1993:SNM

[Arn93] Martin Arnold. Stability of numerical methods for differential-algebraic equations of higher index. Applied Numerical Mathematics: Transactions of IMACS, 13(1–3):5–14, September 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Sixth Conference on the Numerical Treatment of Differential Equations (Halle, 1992).

Arnold:1995:PAD

[Arn95] M. Arnold. A perturbation analysis for the dynamical simulation of mechanical multi-body systems. Applied Numerical Mathematics: Transactions of IMACS, 18(1–3): 37–56, September 27, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=18&
issue=1-3&aid=590. Sev-

enth Conference on the Numerical Treatment of Differential Equations (Halle, 1994).

Arnold:1998:SLM

[Arn98] M. Arnold. The stabilization of linear multistep methods for constrained mechanical systems. Applied Numerical Mathematics: Transactions of IMACS, 28(2–4):143–159, October 1, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1998/28/
2-4/907.pdf.

Aro:1996:SOP

[Aro96]  Colin J. Aro.  A stiff

ODE preconditioner based on Newton linearization. Applied Numerical Mathematics: Transactions of IMACS, 21(4):335–352, November 22, 1996. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=21&
issue=4&aid=699.

Ascher:1997:IER

[ARS97] Uri M. Ascher, Steven J. Ruuth, and Raymond J. Spiteri. Implicit–explicit Runge– Kutta methods for timedependent partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 25(2–3): 151–167, October 17, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=25&
issue=2-3&aid=811.

Arshad:2020:PEE

[Ars20]  Muhammad Arshad.  A

priori error estimates of multiblock mortar expanded mixed method for elliptic problems. Applied Numerical Mathematics: Transactions of IMACS, 157(??): 670–686, November 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300787.

Arens:2005:AAO

[ARSW05] K. Arens, P. Rentrop, S. O.

Stoll, and U. Wever. An adjoint approach to optimal design of turbine blades. Applied Numerical Mathematics: Transactions of IMACS, 53(2–4):93–105, May 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Alkilayh:2023:MCF

[ARY23] Maged Alkilayh, Lothar Reichel, and Qiang Ye. A method for computing a few eigenpairs of large generalized eigenvalue problems. Applied Numerical Mathematics: Transactions of IMACS, 183(??): 108–117, January 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422002239.

Ainsworth:1997:ARS

[AS97] Mark Ainsworth and Bill Senior. An adaptive refinement strategy for hp-finite element computations. Applied Numerical Mathematics: Transactions of IMACS, 26(1–2):165–178, December 22, 1997. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1997/26/
                                        1-2/844.pdf.

Arnold:2000:PCD

[AS00] M. Arnold and B. Simeon. Pantograph and catenary dynamics: a benchmark problem and its numerical solution. Applied Numerical Mathematics: Transactions of IMACS, 34(4):345– 362, August 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/62/32/
28/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/62/32/28/article.
pdf.

Akrivis:2004:IEB

[AS04] Georgios Akrivis and Yiorgos-Sokratis Smyrlis. Implicit– explicit BDF methods for the Kuramoto–Sivashinsky equation. Applied Numerical Mathematics: Transactions of IMACS, 51(2–3): 151–169, November 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Adjerid:2005:EOG

[AS05] Slimane Adjerid and Mohamed Salim. Even–odd goal-oriented a posteriori. Applied Numerical Mathematics: Transactions of IMACS, 55(4):384–402, December 2005. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Artebrant:2006:NSC

[AS06] Robert Artebrant and Hans Joachim Schroll. Numerical simulation of Camassa–Holm peakons by adaptive upwinding. Applied Numerical Mathematics: Transactions of IMACS, 56(5): 695–711, May 2006. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Abbasbandy:2011:MMT

[AS11] S. Abbasbandy and A. Shirzadi. MLPG method for twodimensional diffusion equation with Neumann’s and non-classical boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 61(2):170– 180, February 2011. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Abe:2013:SLE

[AS13] Kuniyoshi Abe and Gerard L. G. Sleijpen. Solving linear equations with a stabilized GPBiCG method. Applied Numerical Mathematics: Transactions of IMACS, 67(??):4–16, May 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927411001048.

Ahmadinia:2020:ALD

[AS20a] M. Ahmadinia and Z. Safari. Analysis of local discontinuous Galerkin method for time-space fractional sine-Gordon equations. Applied Numerical Mathematics: Transactions of IMACS, 148(??):1–17, February 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419302053.

Alla:2020:HPA

[AS20b] Alessandro Alla and Luca Saluzzi. A HJB-POD approach for the control of nonlinear PDEs on a tree structure. Applied Numerical Mathematics: Transactions of IMACS, 155(??): 192–207, ???? 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419303332.

Awawdeh:2020:CAH

[AS20c] Fadi Awawdeh and Linda Smail. Convergence analysis of a highly accurate Nyström scheme for Fredholm integral equations. Applied Numerical Mathematics: Transactions of IMACS, 152(??):231–242, June 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741930337X.

Abdelhak:2021:CIE

[AS21] Hadj Abdelhak and Hacene Saker. Corrigendum to “Integral equations method for solving a Biharmonic inverse problem in detection of Robin coefficients” [Appl. Num. Math. 160 (Feb. 2021), 436–450]. Applied Numerical Mathematics: Transactions of IMACS, 161(??):147, March 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420303536.

See [HS21a].

Ahmadinia:2020:LDG

[ASA20] M. Ahmadinia, Z. Safari, and M. Abbasi. Local discontinuous Galerkin method for time variable order fractional differential equations with sub-diffusion and superdiffusion. Applied Numerical Mathematics: Transactions of IMACS, 157(??): 602–618, November 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420302154.

Arminjon:2003:NTT

[ASC03] P. Arminjon and A. St-Cyr. Nessyahu–Tadmor-type central finite volume methods without predictor for 3D Cartesian and unstructured tetrahedral grids. Applied Numerical Mathematics: Transactions of IMACS, 46(2):135–155, August 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ascher:2012:SC

[Asc12] Uri M. Ascher. Surprising computations. Applied Numerical Mathematics: Transactions of IMACS, 62(10):1276–1288, October 2012. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412001110.

Arminjon:2002:NTT

[ASCM02] P. Arminjon, A. St-Cyr, and A. Madrane. New two- and three-dimensional non-oscillatory central finite volume methods on staggered Cartesian grids. Applied Numerical Mathematics: Transactions of IMACS, 40(3):367–390, February 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.com/
gej-ng/10/10/28/85/33/
27/abstract.html.

Acosta-Soba:2025:PPN

[ASGGRGW25] Daniel Acosta-Soba, Francisco Guillén-González, J. Rafael Rodríguez-Galván, and Jin Wang. Property-preserving numerical approximation of a Cahn–Hilliard–Navier– Stokes model with variable density and degenerate mobility. Applied Numerical Mathematics: Transactions of IMACS, 209(??): 68–83, March 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424003027.

Asok:2021:IAL

[Aso21] Harshin Kamal Asok. Interface adapted LBB-stable finite elements on fluid structure interaction problems in fully Eulerian framework. Applied Numerical Mathematics: Transactions of IMACS, 162(??): 283–300, April 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421000039.

Alves:2021:FSS

[ASS21] Carlos J. S. Alves, Rodrigo G. Serrão, and Ana L. Silvestre. Fundamental solutions for the Stokes equations: Numerical applications for 2D and 3D flows. Applied Numerical Mathematics: Transactions of IMACS, 170(??):55– 73, December 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421002026.

Araya:2019:PEA

[ASV19] Rodolfo Araya, Manuel Solano, and Patrick Vega. A posteriori error analysis of an HDG method for the Oseen problem. Applied Numerical Mathematics: Transactions of IMACS, 146(??): 291–308, December 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927419301898.

Andreev:2015:SOF

[ASZ15] R. Andreev, O. Scherzer, and W. Zulehner. Simultaneous optical flow and source estimation: Space-time discretization and preconditioning. Applied Numerical Mathematics: Transactions of IMACS, 96(??):72– 81, October 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415000781.

An:2018:SGA

[ASZ18] Jing An, Jie Shen, and Zhimin Zhang. The spectral-Galerkin approximation of nonlinear eigenvalue problems. Applied Numerical Mathematics: Transactions of IMACS, 131(??):1– 15, September 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418301016.

Amodio:1993:PDM

[AT93] P. Amodio and D. Trigiante.

A parallel direct method for solving initial value problems for ordinary differential equations. Applied Numerical Mathematics: Transactions of IMACS, 11(1–3): 85–93, January 1993. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Parallel methods for ordinary differential equations (Grado, 1991).

Annaby:2013:SGT

[AT13] M. H. Annaby and M. M. Tharwat. A sinc-Gaussian technique for computing eigenvalues of second-order linear pencils. Applied Numerical Mathematics: Transactions of IMACS, 63(1): 129–137, January 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927412001730.

Alici:2015:LPM

[AT15] H. Alici and H. Taseli. The Laguerre pseudospectral method for the radial Schrödinger equation. Applied Numerical Mathematics: Transactions of IMACS, 87(??):87–99, January 2015. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414001536.

Al-Taweel:2025:CNW

[ATAW25] Ahmed Al-Taweel, Jumana Alkhalissi, and Xiaoshen Wang. The Crank–Nicolson weak Galerkin finite element methods for the sine-Gordon equation. Applied Numerical Mathematics: Transactions of IMACS, 212(??): 77–91, June 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425000248.

Attili:1997:TID

[Att97]  Basem S. Attili.  Trac-

ing implicitly defined curves and the use of singular value decomposition. Applied Numerical Mathematics: Transactions of IMACS, 25(1):1–11, September 16, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=25&
issue=1&aid=803.

Al-Taweel:2020:LOS

[ATW20a] Ahmed Al-Taweel and Xiaoshen Wang. The lowestorder stabilizer free weak Galerkin finite element method. Applied Numerical Mathematics: Transactions of IMACS, 157(??):434–445, November 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742030194X.

Al-Taweel:2020:NOD

[ATW20b] Ahmed Al-Taweel and Xiaoshen Wang. A note on the optimal degree of the weak gradient of the stabilizer free weak Galerkin finite element method. Applied Numerical Mathematics: Transactions of IMACS, 150(??):444–451, April 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419302843.

Augenbaum:1989:APM

[Aug89] J. M. Augenbaum. An adaptive pseudospectral method for discontinuous problems. Applied Numerical Mathematics: Transactions of IMACS, 5(6):459–480, 1989. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Auzinger:2003:SON

[Auz03] W. Auzinger. Sectorial operators and normalized numerical range. Applied Numerical Mathematics: Transactions of IMACS, 45(4): 367–388, June 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Axelsson:1991:AWE

[AV91] O. Axelsson and P. S.

Vassilevski. Asymptotic work estimates for AMLI methods. Applied Numerical Mathematics: Transactions of IMACS, 7(5): 437–451, June 1991. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/016892749190012O.

Alt:1996:VRC

[AV96] R. Alt and J. Vignes. Validation of results of collocation methods for ODEs with the CADNA library. Applied Numerical Mathematics: Transactions of IMACS, 21(2):119–139, July 8, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=21&
issue=2&aid=686.

Avudainayagam:2000:WGM

[AV00] A. Avudainayagam and C. Vani. Wavelet-Galerkin method for integro-differential equations. Applied Numerical Mathematics: Transactions of IMACS, 32(3): 247–254, March 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/

                                        gej-ng/29/17/21/59/30/
                                        27/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/21/59/30/27/article.
                                        pdf.

Alvarez-Vazquez:2009:AOC

[AVMVMV09] L. J. Alvarez-Vázquez, A. Martínez, M. E. Vázquez-Méndez, and M. A. Vilar. An application of optimal control theory to river pollution remediation. Applied Numerical Mathematics: Transactions of IMACS, 59(5): 845–858, May 2009. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Angermann:2003:TDE

[AW03] Lutz Angermann and Song Wang. Three-dimensional exponentially fitted conforming tetrahedral finite elements for the semiconductor continuity equations. Applied Numerical Mathematics: Transactions of IMACS, 46(1):19–43, July 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Apel:2009:OCU

[AW09] Thomas Apel and Gunter Winkler. Optimal control under reduced regularity. Applied Numerical Mathematics: Transactions of IMACS, 59(9): 2050–2064, September 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Adjerid:2014:AED

[AW14] Slimane Adjerid and Thomas Weinhart. Asymptotically exact discontinuous Galerkin error estimates for linear symmetric hyperbolic systems. Applied Numerical Mathematics: Transactions of IMACS, 76(??):101– 131, February 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413000925.

Amin:2019:SIP

[AX19] Mohammed Elmustafa Amin and Xiangtuan Xiong. Source identification problems for radially symmetric and axissymmetric heat conduction equations. Applied Numerical Mathematics: Transactions of IMACS, 138(??): 1–18, April 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418302873.

An:2020:NCS

[AX20] Congpei An and Yuchen Xiao. Numerical construction of spherical tdesigns by Barzilai–Borwein method. Applied Numerical Mathematics: Transactions of IMACS, 150(??):

295–302, April 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419302831.

Abramov:2015:MCP

[AY15] A. A. Abramov and L. F. Yukhno. A method for calculating the Painlevé transcendents. Applied Numerical Mathematics: Transactions of IMACS, 93(??): 262–269, July 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414000737.

Abdelhakem:2021:TSL

[AY21] M. Abdelhakem and Y. H. Youssri. Two spectral Legendre’s derivative algorithms for Lane–Emden, Bratu equations, and singular perturbed problems. Applied Numerical Mathematics: Transactions of IMACS, 169(??):243–255, November 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421001975.

Akiyama:2022:WAM

[AY22] Naho Akiyama and Toshihiro Yamada. A weak approximation method for irregular functionals of hypoelliptic diffusions. Applied Numerical Mathematics: Transactions of IMACS, 172(??):27–49, February 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421002749.

Ayati:2009:CDS

[Aya09] Bruce P. Ayati. A comparison of the dynamics of the structured cell population in virtual and experimental Proteus mirabilis. Applied Numerical Mathematics: Transactions of IMACS, 59(3–4):487–494, March/April 2009. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ai:2018:DLS

[AyLqW18] Qing Ai, Hui yuan Li, and Zhong qing Wang. Diagonalized Legendre spectral methods using Sobolev orthogonal polynomials for elliptic boundary value problems. Applied Numerical Mathematics: Transactions of IMACS, 127(??): 196–210, May 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418300126.

Altmann:2023:DPD

[AZ23] R. Altmann and C. Zimmer. Dissipation-preserving discretization of the Cahn– Hilliard equation with dynamic boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 190(??):254–269, August 2023. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001058.

Abdolabadi:2022:CPC

[AZA22] F. Abdolabadi, A. Zakeri, and A. Amiraslani. A charge-preserving compact splitting method for solving the coupled stochastic nonlinear Schrödinger equations. Applied Numerical Mathematics: Transactions of IMACS, 181(??): 293–319, November 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001672.

Azevedo:2022:SMS

[Aze22]  Juarez S. Azevedo.  A

sigmoid method for some nonlinear Fredholm integral equations of the second kind. Applied Numerical Mathematics: Transactions of IMACS, 181(??): 125–134, November 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001453.

Abbaszadeh:2023:TGS

[AZHD23] Mostafa Abbaszadeh, Mahmoud A. Zaky, Ahmed S. Hendy, and Mehdi Dehghan. A two-grid spectral method to study of dynamics of dense discrete systems governed by Rosenau–Burgers’ equation. Applied Numerical Mathematics: Transactions of IMACS, 187(??): 262–276, May 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423000454.

Babu:2022:EMN

[BAA22] Feeroz Babu, Akram Ali, and Ali H. Alkhaldi. An extragradient method for nonmonotone equilibrium problems on Hadamard manifolds with applications. Applied Numerical Mathematics: Transactions of IMACS, 180(??):85–103, October 2022. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001301.

Baccouch:2014:PEE

[Bac14] Mahboub Baccouch. A posteriori error estimates for a discontinuous Galerkin method applied to onedimensional nonlinear scalar conservation laws. Applied Numerical Mathematics: Transactions of IMACS, 84(??):1–21, October 2014. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741400052X.

Baccouch:2016:APE

[Bac16] Mahboub Baccouch. Analysis of a posteriori error estimates of the discontinuous Galerkin method for nonlinear ordinary differential equations. Applied Numerical Mathematics: Transactions of IMACS, 106(??): 129–153, August 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416300381.

Baccouch:2017:PEE

[Bac17a]  Mahboub Baccouch.  A

posteriori error estimates and adaptivity for the discontinuous Galerkin solutions of nonlinear secondorder initial-value problems. Applied Numerical Mathematics: Transactions of IMACS, 121(??):18– 37, November 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741730137X.

Baccouch:2017:SDG

[Bac17b] Mahboub Baccouch. Superconvergence of the discontinuous Galerkin method for nonlinear second-order initial-value problems for ordinary differential equations. Applied Numerical Mathematics: Transactions of IMACS, 115(??): 160–179, May 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417300090.

Baccouch:2017:SLD

[Bac17c] Mahboub Baccouch. Superconvergence of the local discontinuous Galerkin method for the sine-Gordon equation on Cartesian grids. Applied Numerical Mathematics: Transactions of IMACS, 113(??):124–155, March 2017. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741630232X.

Baccouch:2018:SLD

[Bac18]  Mahboub Baccouch.  A

stochastic local discontinuous Galerkin method for stochastic two-point boundaryvalue problems driven by additive noises. Applied Numerical Mathematics: Transactions of IMACS, 128(??):

43–64, June 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418300321.

Baccouch:2019:AOS

[Bac19] Mahboub Baccouch. Analysis of optimal superconvergence of a local discontinuous Galerkin method for nonlinear second-order twopoint boundary-value problems. Applied Numerical Mathematics: Transactions of IMACS, 145(??): 361–383, November 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419301102.

Baccouch:2021:DGM

[Bac21a] Mahboub Baccouch. The discontinuous Galerkin method for general nonlinear thirdorder ordinary differential equations. Applied Numerical Mathematics: Transactions of IMACS, 162(??): 331–350, April 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421000052.

Baccouch:2021:OSA

[Bac21b] Mahboub Baccouch. Optimal superconvergence and asymptotically exact a posteriori error estimator for the local discontinuous Galerkin method for linear elliptic problems on Cartesian grids. Applied Numerical Mathematics: Transactions of IMACS, 162(??): 201–224, April 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303925.

Bastani:2013:RBC

[BAD13] Ali Foroush Bastani, Zaniar Ahmadi, and Davood Damircheli. A radial basis collocation method for pricing American options under regime-switching jumpdiffusion models. Applied Numerical Mathematics: Transactions of IMACS, 65(??):79–90, March 2013. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412001961.

Badahmane:2020:RPG

[Bad20] A. Badahmane. Regularized preconditioned GMRES and the regularized iteration method. Applied Numerical Mathematics: Transactions of IMACS, 152(??): 159–168, June 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420300027.

Bai:1996:CHA

[Bai96]  Zhongzhi Bai.  A class

of hybrid algebraic multilevel preconditioning methods. Applied Numerical Mathematics: Transactions of IMACS, 19(4):389–399, February 12, 1996. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=19&
                                        issue=4&aid=654.

Baines:1997:GAN

[Bai97] M. J. Baines. Grid adaptation via node movement. Applied Numerical Mathematics: Transactions of IMACS, 26(1–2):77–96, December 22, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1997/26/
                                        1-2/838.pdf.

Bai:2002:KST

[Bai02] Zhaojun Bai. Krylov subspace techniques for reducedorder modeling of large-scale dynamical systems. Applied Numerical Mathematics: Transactions of IMACS, 43(1–2):9–44, October 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bai:2025:GRC

[Bai25a] Zhong-Zhi Bai. On greedy randomized coordinate updating iteration methods for solving symmetric eigenvalue problems. Applied Numerical Mathematics: Transactions of IMACS, 216(??): 76–97, October 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425000972.

Bai:2025:SCD

[Bai25b] Zhong-Zhi Bai. On steepest coordinate descent method for computing extreme eigenpairs of symmetric matrices. Applied Numerical Mathematics: Transactions of IMACS, 217(??): 126–134, November 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425001187.

Baker:1986:MGS

[Bak86] Timothy J. Baker. Mesh generation by a sequence of transformations. Applied Numerical Mathematics: Transactions of IMACS, 2(6):515–528, December 1986. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Baker:1989:DTT

[Bak89] Timothy J. Baker. Developments and trends in three-dimensional mesh generation. Applied Numerical Mathematics: Transactions of IMACS, 5(4):275– 304, July 1989. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Baltensperger:2000:IAM

[Bal00] Richard Baltensperger. Improving the accuracy of the matrix differentiation method for arbitrary collocation points. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4): 143–149, May 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/61/27/
38/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/61/27/38/article.
pdf.

Bank:1997:SAS

[Ban97] Randolph E. Bank. A simple analysis of some a posteriori error estimates. Applied Numerical Mathematics: Transactions of IMACS, 26(1–2):153–164, December 22, 1997. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1997/26/
1-2/843.pdf.

Baker:2006:SSV

[BAP^{+}06] C. T. H. Baker, E. O.

Agyingi, E. I. Parmuzin, F. A. Rihan, and Yihong Song. Sense from sensitivity and variation of parameters. Applied Numerical Mathematics: Transactions of IMACS, 56(3–4): 397–412, March/April 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Barash:2005:NDF

[Bar05]  Danny Barash.  Nonlin-

ear diffusion filtering on extended neighborhood. Applied Numerical Mathematics: Transactions of IMACS, 52(1):1–11, January 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Barbeiro:2009:SCC

[Bar09] S. Barbeiro. Supraconvergent cell-centered scheme for two dimensional elliptic problems. Applied Numerical Mathematics: Transactions of IMACS, 59(1):56– 72, January 2009. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Baramidze:2012:MES

[Bar12]  V. Baramidze.  Minimal

energy spherical splines on Clough–Tocher triangulations for Hermite interpolation. Applied Numerical Mathematics: Transactions of IMACS, 62(9): 1077–1088, September 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411001024.

Bashan:2021:MQB

[Bas21] Ali Bashan. Modification of quintic B-spline differential quadrature method to nonlinear Korteweg-de Vries equation and numerical experiments. Applied Numerical Mathematics: Transactions of IMACS, 167(??): 356–374, September 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001446.

Basto:2017:CDS

[BASC17] Mário Basto, Teresa Abreu, Viriato Semiao, and Francisco L. Calheiros. Convergence and dynamics of structurally identical root finding methods. Applied Numerical Mathematics: Transactions of IMACS, 120(??): 257–269, October 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417301423.

Bazan:2003:CGH

[Baz03] Fermín S. V. Bazán. CGLS– GCV: a hybrid algorithm for low-rank-deficient problems. Applied Numerical Mathematics: Transactions of IMACS, 47(2):91– 108, November 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Belytschko:1994:EDR

[BB94] Ted Belytschko and Ted D.

Blacker. Enhanced derivative recovery through least square residual penalty. Applied Numerical Mathematics: Transactions of IMACS, 14(1–3):55–68, April 14, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
apnum/cas_sub/browse/browse.
cgi?year=1994&volume=14&
issue=1-3&aid=460. Pro-

ceedings of the Third ARO Workshop on Adaptive Methods for Partial Differential Equations (Troy, NY, 1992).

Burrage:1996:HSO

[BB96] K. Burrage and P. M. Burrage. High strong order explicit Runge–Kutta methods for stochastic ordinary differential equations.

Applied Numerical Mathematics: Transactions of IMACS, 22(1–3):81–101, December 4, 1996. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=22&
issue=1-3&aid=709. Special

issue celebrating the centenary of Runge–Kutta methods.

Burrage:1998:GOC

[BB98] K. Burrage and P. M. Burrage. General order conditions for stochastic Runge– Kutta methods for both commuting and non-commuting stochastic ordinary differential equation systems. Applied Numerical Mathematics: Transactions of IMACS, 28(2–4):161–177, October 1, 1998. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1998/28/
2-4/908.pdf.

Boyd:2010:SRI

[BB10] John P. Boyd and Lauren R. Bridge. Sensitivity of RBF interpolation on an otherwise uniform grid with a point omitted or slightly shifted. Applied Numerical Mathematics: Transactions of IMACS, 60(7):659–672, July 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Benitez:2015:PLS

[BB15] M. Benítez and A. Bermúdez.

Pure Lagrangian and semi-Lagrangian finite element methods for the numerical solution of Navier–Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 95(??):62– 81, September 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414000221.

Bauer:2024:AMC

[BB24a] Maximilian Bauer and Mario Bebendorf. Adaptive Hmatrix computations in linear elasticity. Applied Numerical Mathematics: Transactions of IMACS, 201(??):1–19, July 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000291.

Bulatov:2024:IAE

[BB24b] M. V. Bulatov and M. N.

Botoroeva. Integral-algebraic equations with variable limits of integration. The case when the lower limit is at. Applied Numerical Mathematics: Transactions of IMACS, 203(??):225–234, September 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424001107.

Bejaoui:2025:TDP

[BB25] E. Bejaoui and F. Ben Belgacem. Treatment of 3D diffusion problems with discontinuous coefficients and Dirac curvilinear sources. Applied Numerical Mathematics: Transactions of IMACS, 207(??):400–413, January 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424002502.

Benchettou:2022:TCG

[BBBK22] O. Benchettou, A. H. Bentbib, A. Bouhamidi, and K. Kreit. Tensorial conditional gradient method for solving multidimensional ill-posed problems. Applied Numerical Mathematics: Transactions of IMACS, 173(??): 222–238, March 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421003354.

Boucenna:2021:ANS

[BBBN21] Djalal Boucenna, D. Baleanu, Abdellatif Ben Makhlouf, and A. M. Nagy. Analysis and numerical solution of the generalized proportional fractional Cauchy problem. Applied Numerical Mathematics: Transactions of IMACS, 167(??): 173–186, September 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001148.

Bevia:2025:GAL

[BBC^{+}25] V. J. Bevia, S. Blanes, J. C. Cortés, N. Kopylov, and R. J. Villanueva. A GPU-accelerated Lagrangian method for solving the Liouville equation in random differential equation systems. Applied Numerical Mathematics: Transactions of IMACS, 208 (part B) (??):231–255, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424002587.

Barrera:2022:NSF

[BBCR22] D. Barrera, M. Barton, I. Chiarella, and S. Remogna. On numerical solution of Fredholm and Hammerstein integral equations via Nyström method and Gaussian quadrature rules for splines. Applied Numerical Mathematics: Transactions of IMACS, 174(??): 71–88, April 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422000095.

Berres:2005:NIP

[BBCS05] S. Berres, R. Bürger, A. Coronel, and M. Sepúlveda. Numerical identification of parameters for a strongly degenerate convection–diffusion problem modelling centrifugation of flocculated suspensions. Applied Numerical Mathematics: Transactions of IMACS, 52(4): 311–337, March 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bodard:2008:SMB

[BBD08] Nicolas Bodard, Roland Bouffanais, and Michel O. Deville. Solution of movingboundary problems by the spectral element method. Applied Numerical Mathematics: Transactions of IMACS, 58(7):968–984, July 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Barboteu:2018:ADC

[BBD18] Mikael Barboteu, Krzysztof Bartosz, and David Danan. Analysis of a dynamic contact problem with nonmonotone friction and nonclamped boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 126(??): 53–77, April 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417302556.

Bassi:2020:HOA

[BBD20] C. Bassi, S. Busto, and M. Dumbser. High order ADER-DG schemes for the simulation of linear seismic waves induced by nonlinear dispersive free-surface water waves. Applied Numerical Mathematics: Transactions of IMACS, 158(??): 236–263, December 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302300.

Boutilier:2024:RMM

[BBD24] Miranda Boutilier, Konstantin Brenner, and Victorita Dolean. Robust methods for multiscale coarse approximations of diffusion models in perforated domains. Applied Numerical Mathematics: Transactions of IMACS, 201(??): 561–578, July 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424000916.

Barrios:2014:LCP

[BBG14] Tomás P. Barrios, Edwin M. Behrens, and María González. Low cost a posteriori error estimators for an augmented mixed FEM in linear elasticity. Applied Numerical Mathematics: Transactions of IMACS, 84(??):46–65, October 2014. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414000993.

Barzehkar:2025:SAM

[BBJ25] N. Barzehkar, A. Barati, and R. Jalilian. Sinc approximation method for solving system of singularly perturbed parabolic convection-diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 216(??): 127–139, October 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425001035.

Bruggeman:2007:SOU

[BBKS07] Jorn Bruggeman, Hans Burchard, Bob W. Kooi, and Ben Sommeijer. A secondorder, unconditionally positive, mass-conserving integration scheme for biochemical systems. Applied Numerical Mathematics: Transactions of IMACS, 57(1):36– 58, January 2007. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bourchtein:2002:NSI

[BBL02] Andrei Bourchtein, Lioudmila Bourchtein, and Joao Paulo Lukaszczyk. Numerical simulation of incompressible flows through granular porous media. Applied Numerical Mathematics: Transactions of IMACS, 40(1–2):291–306, January 2002. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/85/27/50/abstract.
html.

Boubendir:2015:IER

[BBLT15] Yassine Boubendir, Oscar Bruno, David Levadoux, and Catalin Turc. Integral equations requiring small numbers of Krylov-subspace iterations for two-dimensional smooth penetrable scattering problems. Applied Numerical Mathematics: Transactions of IMACS, 95(??):82– 98, September 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741500015X.

Bourchtein:2003:GAC

[BBO03] A. Bourchtein, L. Bourchtein, and E. R. Oliveira. General approach to conformal mappings used in atmospheric modeling. Applied Numerical Mathematics: Transactions of IMACS, 47(3–4): 305–324, December 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Baker:2005:CMF

[BBPR05] C. T. H. Baker, G. A. Bocharov, C. A. H. Paul, and F. A. Rihan. Computational modelling with functional differential equations: Identification, selection, and sensitivity. Applied Numerical Mathematics: Transactions of IMACS, 53(2–4):107–129, May 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bendahmane:2009:AMS

[BBRBS09] Mostafa Bendahmane, Raimund Bürger, Ricardo Ruiz-Baier, and Kai Schneider. Adaptive multiresolution schemes with local time stepping for two-dimensional degenerate reaction–diffusion systems. Applied Numerical Mathematics: Transactions of IMACS, 59(7):1668–1692, July 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Blatov:1997:GPM

[BBRS97] Igor A. Blatov, Victoria V.

Blatova, Yurii B. Rozhec, and Vadim V. Strygin. Galerkin–Petrov method for strongly nonlinear singularly perturbed boundary value problems on special meshes. Applied Numerical Mathematics: Transactions of IMACS, 25(4):321– 332, November 10, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=25&
issue=4&aid=808.

Bartel:2011:SAE

[BBS11] Andreas Bartel, Sascha Baumanns, and Sebastian Schöps. Structural analysis of electrical circuits including magnetoquasistatic devices. Applied Numerical Mathematics: Transactions of IMACS, 61(12): 1257–1270, December 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411001358.

Bammer:2025:MFE

[BBS25] Patrick Bammer, Lothar Banz, and Andreas Schröder. Mixed finite elements of higher-order in elastoplasticity. Applied Numerical Mathematics: Transactions of IMACS, 209(??): 38–54, March 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424003155.

Boujoudar:2025:IER

[BBT25] Mohamed Boujoudar, Abdelaziz Beljadid, and Ahmed Taik. Implicit EXP-RBF techniques for modeling unsaturated flow through soils with water uptake by plant roots. Applied Numerical Mathematics: Transactions of IMACS, 208 (part B) (??):79–97, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424002691.

Becker:2005:PIC

[BBV05] R. Becker, M. Braack, and B. Vexler. Parameter identification for chemical models in combustion problems. Applied Numerical Mathematics: Transactions of IMACS, 54(3–4):519–536, August 2005. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bagnerini:2013:MGN

[BBV13] Patrizia Bagnerini, Annalisa Buffa, and Elisa Vacca. Mesh generation and numerical analysis of a Galerkin method for highly conductive prefractal layers. Applied Numerical Mathematics: Transactions of IMACS, 65(??):63–78, March 2013. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412001936.

Beckers:2019:DBE

[BBW19] Susanne Beckers, Jörn Behrens, and Winnifried Wollner. Duality based error estimation in the presence of discontinuities. Applied Numerical Mathematics: Transactions of IMACS, 144(??): 83–99, October 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S016892741930131X.

Baiocchi:1989:ESS

[BC89a] Claudio Baiocchi and Michel Crouzeix. On the equivalence of A-stability and Gstability. Applied Numerical Mathematics: Transactions of IMACS, 5(1–2):19– 22, February 1989. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Recent theoretical results in numerical ordinary differential equations.

Burrage:1989:EIM

[BC89b] K. Burrage and F. H. Chipman. Efficiently implementable multivalue methods for solving stiff ordinary differential equations. recent theoretical results in numerical ordinary differential equations. Applied Numerical Mathematics: Transactions of IMACS, 5(1–2):23– 40, February 1989. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Butcher:1989:SRD

[BC89c] J. C. Butcher and J. R. Cash. Some recent developments on numerical initial value problems: a survey. Applied Numerical Mathematics: Transactions of IMACS, 5(1–2):3–18, February 1989. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Recent theoretical results in numerical ordinary differential equations.

Buttke:1993:TCM

[BC93] Thomas F. Buttke and Alexandre J. Chorin. Turbulence calculations in magnetization variables. Applied Numerical Mathematics: Transactions of IMACS, 12(1–3):47–54, May 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0168927493901114.

Special issue to honor Professor Saul Abarbanel on his sixtieth birthday (Neveh, 1992).

Brezinski:1994:P

[BC94] C. Brezinski and M. Morandi Cecchi. Preface. Applied Numerical Mathematics: Transactions of IMACS, 15(2):99–100, September 13, 1994. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=15&
issue=2&aid=489.

Butcher:1995:PGL

[BC95] J. C. Butcher and P. Chartier.

Parallel general linear methods for stiff ordinary differential and differential algebraic equations. Applied Numerical Mathematics: Transactions of IMACS, 17(3): 213–222, September 4, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=17&
issue=3&aid=577. Numeri-

cal methods for ordinary differential equations (Atlanta, GA, 1994).

Budd:1997:IMM

[BC97a] C. J. Budd and G. J.

Collins. An invariant moving mesh scheme for the nonlinear diffusion equation. Applied Numerical Mathematics: Transactions of IMACS, 26(1–2):23–39, December 22, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1997/26/
1-2/834.pdf.

Butcher:1997:GSI

[BC97b] J. C. Butcher and P. Chartier. A generalization of singlyimplicit Runge–Kutta methods. Applied Numerical Mathematics: Transactions of IMACS, 24(2–3): 343–350, June 19, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=24&
issue=2-3&aid=787. Volterra

centennial (Tempe, AZ, 1996).

Butcher:1998:EMV

[BC98] J. C. Butcher and D. J. L. Chen. ESIRK methods and variable stepsize. Applied Numerical Mathematics: Transactions of IMACS, 28(2–4):193–207, October 1, 1998. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/

store/apnum/sub/1998/28/
2-4/941.pdf.

Balle:1999:PAC

[BC99] Susanne Balle and Jane Cullum. A parallel algorithm for computing eigenvalues of very large real symmetric matrices on message passing architectures. Applied Numerical Mathematics: Transactions of IMACS, 30(2–3):341–365, June 10, 1999. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1999&volume=30&
issue=2-3&aid=984.

Butcher:2000:MSZ

[BC00a] J. C. Butcher and T. M. H.

Chan. Multi-step zero approximations for stepsize control. Applied Numerical Mathematics: Transactions of IMACS, 34(2–3): 167–177, July 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/62/29/
29/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/62/29/29/article.
pdf.

Butcher:2000:NTS

[BC00b] J. C. Butcher and D. J. L.

Chen. A new type of singly-implicit Runge–Kutta method. Applied Numerical Mathematics: Transactions of IMACS, 34(2–3): 179–188, July 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/62/29/
30/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/62/29/30/article.
pdf.

Blasco:2001:STE

[BC01] J. Blasco and R. Codina. Space and time error estimates for a first order, pressure stabilized finite element method for the incompressible Navier–Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 38(4):475–497, September 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/77/34/31/abstract.
html.

Balac:2002:IMF

[BC02] S. Balac and G. Caloz. Induced magnetic field computations using a boundary integral formulation. Applied Numerical Mathematics: Transactions of IMACS, 41(3):345–367, June 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Berrone:2004:MPE

[BC04a] Stefano Berrone and Claudio Canuto. Multilevel a posteriori error analysis for reaction–convection– diffusion problems. Applied Numerical Mathematics: Transactions of IMACS, 50(3–4):371–394, September 2004. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Blasco:2004:EEO

[BC04b] J. Blasco and R. Codina. Error estimates for an operatorsplitting method for incompressible flows. Applied Numerical Mathematics: Transactions of IMACS, 51(1):1–17, October 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Blanes:2005:NNC

[BC05] Sergio Blanes and Fernando Casas. On the necessity of negative coefficients for operator splitting schemes of order higher than two. Applied Numerical Mathematics: Transactions of IMACS, 54(1):23–37, June 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bermejo:2008:AFE

[BC08a] R. Bermejo and J. Carpio.

An adaptive finite element semi-Lagrangian implicit– explicit Runge–Kutta–Chebyshev method for convection dominated reaction–diffusion problems. Applied Numerical Mathematics: Transactions of IMACS, 58(1):16– 39, January 2008. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bi:2008:MFV

[BC08b] Chunjia Bi and Wenbin Chen. Mortar finite volume element method with Crouzeix–Raviart element for parabolic problems. Applied Numerical Mathematics: Transactions of IMACS, 58(11):1642–1657, November 2008. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Badia:2012:SMD

[BC12] Santiago Badia and Ramon Codina. Stokes, Maxwell and Darcy: a single finite element approximation for three model problems. Applied Numerical Mathematics: Transactions of IMACS, 62(4):246–263, April 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411001097.

Brehier:2023:ASS

[BC23] Charles-Edouard Bréhier and David Cohen. Analysis of a splitting scheme for a class of nonlinear stochastic Schrödinger equations. Applied Numerical Mathematics: Transactions of IMACS, 186(??):57–83, April 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423000028.

Bica:2016:TPB

[BCC16] Alexandru Mihai Bica, Mircea Curila, and Sorin Curila. Two-point boundary value problems associated to functional differential equations of even order solved by iterated splines. Applied Numerical Mathematics: Transactions of IMACS, 110(??): 128–147, December 2016. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301441.

See corrigendum [Bic21].

Bessemoulin-Chatard:2021:ANS

[BCCHM21] Marianne Bessemoulin-Chatard, Claire Chainais-Hillairet, and Hélène Mathis. Analysis of numerical schemes for semiconductor energytransport models. Applied Numerical Mathematics: Transactions of IMACS, 168(??):143–169, October 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421001677.

Burrage:2017:NST

[BCDP17] Kevin Burrage, Angelamaria Cardone, Raffaele D’Ambrosio, and Beatrice Paternoster. Numerical solution of time fractional diffusion systems. Applied Numerical Mathematics: Transactions of IMACS, 116(??):82–94, June 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417300405.

Buccini:2025:FAF

[BCDR25] Alessandro Buccini, Fei Chen, Omar De la Cruz Cabrera, and Lothar Reichel. Fast alternating fitting methods for trigonometric curves for large data sets. Applied Numerical Mathematics: Transactions of IMACS, 208 (part A) (??):104–134, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424000011.

Bendito:2004:DSU

[BCE04] Enrique Bendito, Ángeles Carmona, and Andrés M. Encinas. Difference schemes on uniform grids performed by general discrete operators. Applied Numerical Mathematics: Transactions of IMACS, 50(3–4): 343–370, September 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Blanes:2022:RKN

[BCET22] S. Blanes, F. Casas, and A. Escorihuela-Tomàs. Runge– Kutta–Nyström symplectic splitting methods of order 8. Applied Numerical Mathematics: Transactions of IMACS, 182(??):14– 27, December 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001842.

Blanes:2024:FEL

[BCET24] S. Blanes, F. Casas, and A. Escorihuela-Tomàs. Families of efficient low order processed composition methods. Applied Numerical Mathematics: Transactions of IMACS, 204(??):86– 100, October 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001429.

Blanes:2013:NFS

[BCF^{+}13] S. Blanes, F. Casas, A. Farrés, J. Laskar, J. Makazaga, and A. Murua. New families of symplectic splitting methods for numerical integration in dynamical astronomy. Applied Numerical Mathematics: Transactions of IMACS, 68(??):58–72, June 2013. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413000135.

Bazarra:2019:NAT

[BCFQ19] N. Bazarra, M. Campo, J. R. Fernández, and R. Quintanilla. Numerical analysis of a thermoelastic problem with dual-phase-lag heat conduction. Applied Numerical Mathematics: Transactions of IMACS, 140(??): 76–90, June 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419300212.

Bazarra:2021:NAD

[BCFQ21] N. Bazarra, M. I. M. Copetti, J. R. Fernández, and R. Quintanilla. Numerical analysis of a dual-phase-lag model with microtemperatures. Applied Numerical Mathematics: Transactions of IMACS, 166(??): 1–25, August 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000866.

Borges:2025:DRP

[BCFR25] J. S. Borges, G. C. M. Campos, J. A. Ferreira, and G. Romanazzi. Drug release from polymeric platforms for non smooth solutions. Applied Numerical Mathematics: Transactions of IMACS, 213(??):12–37, July 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425000467.

Bannenberg:2021:ROM

[BCG21] M. W. F. M. Bannenberg, A. Ciccazzo, and M. Günther. Reduced order multirate schemes for coupled differential-algebraic systems. Applied Numerical Mathematics: Transactions of IMACS, 168(??): 104–114, October 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001525.

Bourdin:2013:VIF

[BCGI13] Loïc Bourdin, Jacky Cresson, Isabelle Greff, and Pierre Inizan. Variational integrator for fractional Euler–Lagrange equations. Applied Numerical Mathematics: Transactions of IMACS, 71(??):14– 23, September 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892741300041X.

Bermudez:2024:PTS

[BCGS24] Isaac Bermúdez, Claudio I. Correa, Gabriel N. Gatica, and Juan P. Silva. A perturbed twofold saddle point-based mixed finite element method for the Navier–Stokes equations with variable viscosity. Applied Numerical Mathematics: Transactions of IMACS, 201(??):465–487, July 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424000849.

Berzins:1997:SAI

[BCJ97] Martin Berzins, Philip J. Capon, and Peter K. Jimack. On spatial adaptivity and interpolation when using the method of lines. Applied Numerical Mathematics: Transactions of IMACS, 26(1–2):117–133, December 22, 1997. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1997/26/
                                        1-2/841.pdf.

Bras:2018:EPI

[BCJP18] M. Braś, A. Cardone, Z. Jackiewicz, and P. Pierzchala. Error propagation for implicit– explicit general linear methods. Applied Numerical Mathematics: Transactions of IMACS, 131(??): 207–231, September 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418301144.

Bras:2017:ORP

[BCJW17] Michal Braś, Angelamaria Cardone, Zdzislaw Jackiewicz, and Bruno Welfert. Order reduction phenomenon for general linear methods. Applied Numerical Mathematics: Transactions of IMACS, 119(??): 94–114, September 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417300910.

Brio:2022:SSP

[BCK22] M. Brio, J.-G. Caputo, and H. Kravitz. Spectral solutions of PDEs on networks. Applied Numerical Mathematics: Transactions of IMACS, 172(??):99– 117, February 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421002750.

Bensiali:2015:PRB

[BCL15] Bouchra Bensiali, Guillaume Chiavassa, and Jacques Liandrat. Penalization of Robin boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 96(??):134– 152, October 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927415000902.

Benes:2004:GIS

[BCM04] Michal Beneš, Vladimír Chalupecký, and Karol Mikula. Geometrical image segmentation by the Allen–Cahn equation. Applied Numerical Mathematics: Transactions of IMACS, 51(2–3):187–205, November 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Birgin:2003:EOP

[BCMV03] Ernesto G. Birgin, Ivan E. Chambouleyron, José Mario Martínez, and Sergio D. Ventura. Estimation of optical parameters of very thin films. Applied Numerical Mathematics: Transactions of IMACS, 47(2): 109–119, November 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Blanes:2001:HOR

[BCR01] S. Blanes, F. Casas, and J. Ros. High-order Runge– Kutta–Nyström geometric methods with processing. Applied Numerical Mathematics: Transactions of IMACS, 39(3–4):245–259, December 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/79/34/28/abstract.
html.

Burger:2006:UDS

[BCS06] Raimund Bürger, Aníbal Coronel, and Mauricio Sepúlveda. On an upwind difference scheme for strongly degenerate parabolic equations modelling the settling of suspensions in centrifuges and non-cylindrical vessels. Applied Numerical Mathematics: Transactions of IMACS, 56(10–11):1397–1417, October/November 2006. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Baron:2017:AMT

[BCS17] V. Baron, Y. Coudière, and P. Sochala. Adaptive multistep time discretization and linearization based on a posteriori error estimates for the Richards equation. Applied Numerical Mathematics: Transactions of IMACS, 112(??): 104–125, February 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416301994.

Bu:2016:HRM

[BCSH16] Yiming Bu, Bruno Carpentieri, Zhaoli Shen, and Ting-Zhu Huang. A hybrid recursive multilevel incomplete factorization preconditioner for solving general linear systems. Applied Numerical Mathematics: Transactions of IMACS, 104(??):141–157, June 2016. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416000271.

Betts:2016:SOC

[BCT16] John T. Betts, Stephen L. Campbell, and Karmethia C. Thompson. Solving optimal control problems with control delays using direct transcription. Applied Numerical Mathematics: Transactions of IMACS, 108(??): 185–203, October 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416300848.

Blanes:2019:SCM

[BCT19] Sergio Blanes, Fernando Casas, and Mechthild Thalhammer. Splitting and composition methods with embedded error estimators. Applied Numerical Mathematics: Transactions of IMACS, 146(??):400–415, December 2019. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927419301941.

Bertoluzza:2000:ACI

[BCU00] Silvia Bertoluzza, Claudio Canuto, and Karsten Urban. On the adaptive computation of integrals of wavelets. Applied Numerical Mathematics: Transactions of IMACS, 34(1):13–38, June 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/62/27/
28/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/62/27/28/article.
pdf.

Burazin:2021:OCM

[BCV21] Kresimir Burazin, Ivana Crnjac, and Marko Vrdoljak. Optimality criteria method in 2D linearized elasticity problems. Applied Numerical Mathematics: Transactions of IMACS, 160(??): 192–204, February 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302993.

Bruni:2025:MSI

[BCV25] Vittoria Bruni, Rosanna Campagna, and Domenico Vitulano. Multicomponent signals interference detection exploiting HP-splines frequency parameter. Applied Numerical Mathematics: Transactions of IMACS, 209(??):20–37, March 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424003003.

Bank:1985:EIS

[BD85] Randolph E. Bank and Craig C. Douglas. An efficient implementation for SSOR and incomplete factorization preconditionings. Applied Numerical Mathematics: Transactions of IMACS, 1(6):489–492, November 1985. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Banks:2007:CAM

[BD07] H. T. Banks and Jimena L. Davis. A comparison of approximation methods for the estimation of probability distributions on parameters. Applied Numerical Mathematics: Transactions of IMACS, 57(5–7): 753–777, May 2007. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Burg:2011:CAH

[BD11] M. Bürg and W. Dörfler.

Convergence of an adaptive hp finite element strategy in higher space-dimensions. Applied Numerical Mathematics: Transactions of IMACS, 61(11):1132–1146, November 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Butcher:2017:PGL

[BD17] John C. Butcher and Raffaele D’Ambrosio. Partitioned general linear methods for separable Hamiltonian problems. Applied Numerical Mathematics: Transactions of IMACS, 117(??):69–86, July 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417300375.

Busto:2022:SSI

[BD22] S. Busto and M. Dumbser.

A staggered semi-implicit hybrid finite volume/finite element scheme for the shallow water equations at all Froude numbers. Applied Numerical Mathematics: Transactions of IMACS, 175(??): 108–132, May 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422000319.

Buccini:2020:LSR

[BDD^{+}20] Alessandro Buccini, Omar De la Cruz Cabrera, Marco Donatelli, Andrea Martinelli, and Lothar Reichel. Large-scale regression with non-convex loss and penalty. Applied Numerical Mathematics: Transactions of IMACS, 157(??): 590–601, November 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420302063.

Boufflet:2012:NNT

[BDDV12] J. P. Boufflet, M. Dambrine, G. Dupire, and P. Villon. On the necessity of Nitsche term. Part II: an alternative approach. Applied Numerical Mathematics: Transactions of IMACS, 62(5): 521–535, May 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927411002157.

Brachet:2022:CEI

[BDE22] Matthieu Brachet, Laurent Debreu, and Christopher Eldred. Comparison of exponential integrators and traditional time integration schemes for the shallow water equations. Applied Numerical Mathematics: Transactions of IMACS, 180(??):55–84, October 2022. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001295.

Benkhaldoun:2012:NMS

[BDES12] Fayssal Benkhaldoun, Salah Daoudi, Imad Elmahi, and Mohammed Seaid. Numerical modelling of sediment transport in the Nador lagoon (Morocco). Applied Numerical Mathematics: Transactions of IMACS, 62(12):1749–1766, December 2012. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412001183.

Berzins:1989:DST

[BDF89] M. Berzins, P. M. Dew, and R. M. Furzeland. Developing software for timedependent problems using the method of lines and differential-algebraic integrators. Applied Numerical Mathematics: Transactions of IMACS, 5(5):375– 397, September 1989. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Biswas:1994:PAF

[BDF94] Rupak Biswas, Karen D.

Devine, and Joseph E. Flaherty. Parallel, adaptive finite element methods for conservation laws. Applied Numerical Mathematics: Transactions of IMACS, 14(1–3):255–283, April 14, 1994. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=14&
issue=1-3&aid=459. Pro-

ceedings of the Third ARO Workshop on Adaptive Methods for Partial Differential Equations (Troy, NY, 1992).

Bartel:2025:STD

[BDF^{+}25] Andreas Bartel, Malak Diab, Andreas Frommer, Michael Günther, and Nicole Marheineke. Splitting techniques for DAEs with port-Hamiltonian applications. Applied Numerical Mathematics: Transactions of IMACS, 214(??):28–53, August 2025. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425000558.

Bolten:2023:SBC

[BDFF23] Matthias Bolten, Marco Donatelli, Paola Ferrari, and Isabella Furci. Symbol based convergence analysis in block multigrid methods with applications for Stokes problems. Applied Numerical Mathematics: Transactions of IMACS, 193(??): 109–130, November 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423002003.

Benantar:1995:TG

[BDFK95] Messaoud Benantar, Uḡur Doḡrusöz, Joseph E. Flaherty, and Mukkai S. Krishnamoorthy. Triangle graphs. Applied Numerical Mathematics: Transactions of IMACS, 17(2):85– 96, June 13, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=17&
issue=2&aid=559.

Bergamaschi:2008:CSS

[BdFPSdSC08] Paulo Roberto Bergamaschi, Sezimária de Fátima Pereira Saramago, and Leandro dos Santos Coelho. Comparative study of SQP and metaheuristics for robotic manipulator design. Applied Numerical Mathematics: Transactions of IMACS, 58(9):1396–1412, September 2008. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bocher:1995:PIC

[BDFV95] P. Bocher, H. De Meyer, V. Fack, and G. Vanden Berghe. A parallel iterated correction algorithm for the numerical solution of Volterra equations. Applied Numerical Mathematics: Transactions of IMACS, 17(1):19–32, May 1, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=17&
                                        issue=1&aid=557.

Baker:1996:CAE

[BDGP96] T. S. Baker, J. R. Dormand, J. P. Gilmore, and P. J. Prince. Continuous approximation with embedded Runge–Kutta methods. Applied Numerical Mathematics: Transactions of IMACS, 22(1–3):51–62, December 4, 1996. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=22&
                                        issue=1-3&aid=707. Special

issue celebrating the centenary of Runge–Kutta methods.

Bona:1992:CBD

[BDKM92] Jerry L. Bona, Vassilios A. Dougalis, Ohannes A. Karakashian, and William R. McKinney. Computations of blow-up and decay for periodic solutions of the generalized Korteweg–de Vries– Burgers equation. Applied Numerical Mathematics: Transactions of IMACS, 10(3–4):335–355, September 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). A Festschrift to honor Professor Garrett Birkhoff on his eightieth birthday.

Burchard:2003:HOC

[BDM03] Hans Burchard, Eric Deleersnijder, and Andreas Meister. A high-order conservative Patankar-type discretisation for stiff systems of production–destruction equations. Applied Numerical Mathematics: Transactions of IMACS, 47(1): 1–30, October 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Baeza:2012:NTW

[BDMG12] A. Baeza, R. Donat, and A. Martinez-Gavara. A numerical treatment of wet/dry zones in well-balanced hybrid schemes for shallow water flow. Applied Numerical Mathematics: Transactions of IMACS, 62(4): 264–277, April 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927411001243.

Bultheel:2005:OLP

[BDMGVO05] A. Bultheel, C. Díaz-Mendoza, P. González-Vera, and R. Orive. Orthogonal Laurent polynomials and quadrature formulas for unbounded intervals: II. Interpolatory rules. Applied Numerical Mathematics: Transactions of IMACS, 54(1):39–63, June 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Botta:1997:HFL

[BDN^{+}97] E. F. F. Botta, K. Dekker, Y. Notay, A. van der Ploeg, C. Vuik, F. W. Wubs, and P. M. de Zeeuw. How fast the Laplace equation was solved in 1995. Applied Numerical Mathematics: Transactions of IMACS, 24(4): 439–455, August 12, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=24&
                                        issue=4&aid=797.

Berardi:2019:TML

[BDNV19] Marco Berardi, Fabio Difonzo, Filippo Notarnicola, and Michele Vurro. A transversal method of lines for the numerical modeling of vertical infiltration into the vadose zone. Applied Numerical Mathematics: Transactions of IMACS, 135(??): 264–275, January 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418301892.

Berthe:2019:SWM

[BDOG19] Edouard Berthe, Duy-Minh Dang, and Luis Ortiz-Gracia. A Shannon wavelet method for pricing foreign exchange options under the Heston multi-factor CIR model. Applied Numerical Mathematics: Transactions of IMACS, 136(??):1–22, February 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418302101.

Burrage:1996:PPW

[BDP96] Kevin Burrage, Carolyn Dyke, and Bert Pohl. On the performance of parallel waveform relaxations for differential systems. Applied Numerical Mathematics: Transactions of IMACS, 20(1–2): 39–55, February 29, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/

                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=20&
                                        issue=1-2&aid=664. Work-

shop on the method of lines for time-dependent problems (Lexington, KY, 1995).

Baker:1999:CAE

[BDP99] T. S. Baker, J. R. Dormand, and P. J. Prince. Continuous approximation with embedded Runge–Kutta–Nyström methods. Applied Numerical Mathematics: Transactions of IMACS, 29(2):171–188, February 1, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1999/29/
                                        2/931.pdf.

Brezinski:2004:QOA

[BDRZ04] C. Brezinski, K. A. Driver, and M. Redivo-Zaglia. Quasiorthogonality with applications to some families of classical orthogonal polynomials. Applied Numerical Mathematics: Transactions of IMACS, 48(2):157– 168, February 2004. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Brezinski:2019:ZQQ

[BDRZ19] Claude Brezinski, Kathy A. Driver, and Michela Redivo-Zaglia. Zeros of quadratic quasi-orthogonal order 2 polynomials. Applied Numerical Mathematics: Transactions of IMACS, 135(??): 143–145, January 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418301880.

Barucq:2009:PAN

[BDSG09] H. Barucq, R. Djellouli, and A. Saint-Guirons. Performance assessment of a new class of local absorbing boundary conditions for elliptical- and prolate spheroidal-shaped boundaries. Applied Numerical Mathematics: Transactions of IMACS, 59(7):1467–1498, July 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bos:2017:PAL

[BDV17] L. Bos, S. De Marchi, and M. Vianello. Polynomial approximation on Lissajous curves in the d-cube. Applied Numerical Mathematics: Transactions of IMACS, 116(??):47–56, June 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417300284.

Brown:1999:LAF

[BE99] J. L. Brown and W. L.

Etheridge. Local approximation of functions over points scattered in R^{m}. Applied Numerical Mathematics: Transactions of IMACS, 29(2):189–199, February 1, 1999. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1999/29/
                                        2/932.pdf.

Beauwens:2004:ISM

[Bea04] Robert Beauwens. Iterative solution methods. Applied Numerical Mathematics: Transactions of IMACS, 51(4):437–450, December 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Beccaria:2002:NAC

[Bec02] Matteo Beccaria. Numerical algorithm for the calculation of the ground states in the U_{q}SU(2) symmetric spin-Heisenberg chain. Applied Numerical Mathematics: Transactions of IMACS, 41(2):269–281, May 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Becher:2018:AGS

[Bec18]  Simon Becher.  Analysis

of Galerkin and streamlinediffusion FEMs on piecewise equidistant meshes for turning point problems exhibiting an interior layer. Applied Numerical Mathematics: Transactions of IMACS, 123(??): 121–136, January 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417301873.

Beghdadi:2000:SDS

[Beg00] Driss Beghdadi. Second degree semi-classical forms of class s = 1. The symmetric case. Applied Numerical Mathematics: Transactions of IMACS, 34(1):1–11, June 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/62/27/
27/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/62/27/27/article.
pdf.

Behforooz:1993:NAS

[Beh93] G. Hossein Behforooz. A new approach to spline functions. Applied Numerical Mathematics: Transactions of IMACS, 13(4): 271–276, November 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Behrens:1997:AOM

[Beh97] Jörn Behrens. Atmospheric and ocean modeling with an adaptive finite element solver for the shallow-water equations. Applied Numerical Mathematics: Transactions of IMACS, 26(1–2):217–226, December 22, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1997/26/
                                        1-2/848.pdf.

Brunk:2024:SOS

[BEH24] Aaron Brunk, Herbert Egger, and Oliver Habrich. A second-order structurepreserving discretization for the Cahn–Hilliard/Allen– Cahn system with crosskinetic coupling. Applied Numerical Mathematics: Transactions of IMACS, 206(??):12–28, December 2024. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742400196X.

Belantari:1991:EEV

[Bel91] A. Belantari. Error estimate in vector Padé approximation. Applied Numerical Mathematics: Transactions of IMACS, 8(6):457– 468, December 1991. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bellen:1997:CCR

[Bel97] A. Bellen. Contractivity of continuous Runge–Kutta methods for delay differential equations. Applied Numerical Mathematics: Transactions of IMACS, 24(2– 3):219–232, June 19, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=24&
issue=2-3&aid=778. Volterra

centennial (Tempe, AZ, 1996).

Bendtsen:1996:HSP

[Ben96]  Claus Bendtsen.  Highly

stable parallel Runge–Kutta methods. Applied Numerical Mathematics: Transactions of IMACS, 21(1): 1–8, June 6, 1996. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=21&
issue=1&aid=685.

Benazzouz:1998:QTA

[Ben98]  A. Benazzouz.  GL(E)-

quasilinear transformations and acceleration. Applied Numerical Mathematics: Transactions of IMACS, 27(2):109–122, June 10, 1998. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/

                                        store/apnum/sub/1998/27/
                                        2/870.pdf.

Bentbib:2002:SFR

[Ben02]  A. H. Bentbib.  Solv-

ing the full rank interval least squares problem. Applied Numerical Mathematics: Transactions of IMACS, 41(2):283–294, May 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Benes:2017:CSA

[Ben17] Michal Benes. Convergence and stability analysis of heterogeneous time step coupling schemes for parabolic problems. Applied Numerical Mathematics: Transactions of IMACS, 121(??): 198–222, November 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417301617.

Berlinet:1985:STS

[Ber85]  A. Berlinet.  Sequence

transformations as statistical tools. Applied Numerical Mathematics: Transactions of IMACS, 1(6):531– 544, November 1985. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Berzins:1986:ICB

[Ber86] M. Berzins. A C^{1} interpolant for codes based on backward differentiation formulae. Applied Numerical Mathematics: Transactions of IMACS, 2(2):109–118, April 1986. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Berzins:2004:VOF

[Ber04] M. Berzins. Variable-order finite elements and positivity preservation for hyperbolic PDEs. Applied Numerical Mathematics: Transactions of IMACS, 48(3–4): 271–292, March 2004. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Berzins:2005:PPH

[Ber05] M. Berzins. Preserving positivity for hyperbolic PDEs using variable-order finite elements with bounded polynomials. Applied Numerical Mathematics: Transactions of IMACS, 52(2– 3):197–217, February 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Berger:2015:NSC

[Ber15] Marsha Berger. A note on the stability of cut cells and cell merging. Applied Numerical Mathematics: Transactions of IMACS, 96(??):180–186, October 2015. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927415000896.

Barrera:2018:TMB

[BES18] D. Barrera, F. Elmokhtari, and D. Sbibih. Two methods based on bivariate spline quasi-interpolants for solving Fredholm integral equations. Applied Numerical Mathematics: Transactions of IMACS, 127(??): 78–94, May 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417302660.

Baker:1992:SPS

[BF92a] Christopher T. H. Baker and Neville J. Ford. Stability properties of a scheme for the approximate solution of a delay-integro-differential equation. Applied Numerical Mathematics: Transactions of IMACS, 9(3–5): 357–370, April 1992. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/016892749290027B.

International Conference on the Numerical Solution of Volterra and Delay Equations (Tempe, AZ, 1990).

Berzins:1992:ATM

[BF92b] M. Berzins and R. M. Furzeland. An adaptive theta method for the solution of stiff and nonstiff differential equations. Applied Numerical Mathematics: Transactions of IMACS, 9(1):1– 19, January 1992. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bull:1995:PGA

[BF95] J. M. Bull and T. L.

Freeman. Parallel globally adaptive algorithms for multi-dimensional integration. Applied Numerical Mathematics: Transactions of IMACS, 19(1–2): 3–16, December 4, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=1-2&aid=624.

Benettin:1999:HPT

[BF99] Giancarlo Benettin and Francesco Fassò. From Hamiltonian perturbation theory to symplectic integrators and back. Applied Numerical Mathematics: Transactions of IMACS, 29(1):73–87, January 1, 1999. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1999/29/
1/939.pdf.

Barbeiro:2001:NLF

[BF01] S. Barbeiro and J. A. Ferreira. A nonstandard linear finite element method for a planar elasticity problem. Applied Numerical Mathematics: Transactions of IMACS, 37(3):331–340, May 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/10/28/73/31/
                                        31/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/10/28/73/31/31/article.
                                        pdf.

Bertaccini:2009:CCI

[BF09] Daniele Bertaccini and Stefano Fanelli. Computational and conditioning issues of a discrete model for cochlear sensorineural hypoacusia. Applied Numerical Mathematics: Transactions of IMACS, 59(8): 1989–2001, August 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bantle:2015:EAI

[BF15] Markus Bantle and Stefan Funken. Efficient and accurate implementation of hp-BEM for the Laplace operator in 2D. Applied Numerical Mathematics: Transactions of IMACS, 95(??):51– 61, September 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414000282.

Barth:2017:UQL

[BF17] Andrea Barth and Franz G.

Fuchs. Uncertainty quantification for linear hyperbolic equations with stochastic process or random field coefficients. Applied Numerical Mathematics: Transactions of IMACS, 121(??):38– 51, November 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417301459.

Baker:2020:CFD

[BF20] Christopher T. H. Baker and Neville J. Ford. Characteristic functions of differential equations with deviating arguments. Applied Numerical Mathematics: Transactions of IMACS, 149(??):17–29, ???? 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419300972.

Biswas:1993:AMM

[BFA93] Rupak Biswas, Joseph E.

Flaherty, and David C. Arney. An adaptive meshmoving and refinement procedure for one-dimensional conservation laws. Applied Numerical Mathematics: Transactions of IMACS, 11(4):259–282, February 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Branco:2007:NMG

[BFdO07] J. R. Branco, J. A. Ferreira, and P. de Oliveira. Numerical methods for the generalized Fisher–Kolmogorov– Petrovskii–Piskunov equation. Applied Numerical Mathematics: Transactions of IMACS, 57(1):89– 102, January 2007. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Branco:2010:NFD

[BFdS10] J. R. Branco, J. A. Ferreira, and P. da Silva. Non-Fickian delay reaction– diffusion equations: Theoretical and numerical study. Applied Numerical Mathematics: Transactions of IMACS, 60(5):531–549, May 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Barrera:2008:MES

[BFGP08] D. Barrera, M. A. Fortes, P. González, and M. Pasadas. Minimal energy surfaces on Powell–Sabin type triangulations. Applied Numerical Mathematics: Transactions of IMACS, 58(5): 635–645, May 2008. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Botchev:2009:AOS

[BFH09] M. A. Botchev, I. Faragó, and R. Horváth. Application of operator splitting to the Maxwell equations including a source term. Applied Numerical Mathematics: Transactions of IMACS, 59(3–4): 522–541, March/April 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Berrut:2011:CRD

[BFK11] Jean-Paul Berrut, Michael S.

Floater, and Georges Klein. Convergence rates of derivatives of a family of barycentric rational interpolants. Applied Numerical Mathematics: Transactions of IMACS, 61(9):989–1000, September 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Brdar:2023:NAS

[BFLR23] Mirjana Brdar, Sebastian Franz, Lars Ludwig, and Hans-Görg Roos. Numerical analysis of a singularly perturbed convection diffusion problem with shift in space. Applied Numerical Mathematics: Transactions of IMACS, 186(??): 129–142, April 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742300003X.

Barbeiro:2011:H

[BFP11] S. Barbeiro, J. A. Ferreira, and L. Pinto. H. Applied Numerical Mathematics: Transactions of IMACS, 61(2):201–215, February 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bazarra:2022:NAS

[BFQ22] N. Bazarra, J. R. Fernández, and R. Quintanilla. Numerical approximation of some poro-elastic problems with MGT-type dissipation mechanisms. Applied Numerical Mathematics: Transactions of IMACS, 177(??): 123–136, July 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422000708.

Bardeji:2017:OFF

[BFS17] Somayeh Gh. Bardeji, Isabel N. Figueiredo, and Ercília Sousa. Optical flow with fractional order regularization: Variational model and solution method. Applied Numerical Mathematics: Transactions of IMACS, 114(??):188–200, April 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416300599.

Beyn:2002:EVS

[BG02a] Wolf-Jürgen Beyn and Barnabas M. Garay. Estimates of variable stepsize Runge–Kutta methods for sectorial evolution equations with nonsmooth data. Applied Numerical Mathematics: Transactions of IMACS, 41(3):369–400, June 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Broker:2002:SAI

[BG02b] Oliver Bröker and Marcus J. Grote. Sparse approximate inverse smoothers for geometric and algebraic multigrid. Applied Numerical Mathematics: Transactions of IMACS, 41(1):61– 80, April 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/86/27/31/abstract.
html.

Bidegaray:2003:MCC

[BG03] B. Bidégaray and J.-M.

Ghidaglia. Multidimensional corrections to cellcentered finite volume methods for Maxwell equations. Applied Numerical Mathematics: Transactions of IMACS, 44(3):281– 298, February 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Boffi:2006:IEE

[BG06] Daniele Boffi and Lucia Gastaldi. Interpolation estimates for edge finite elements and application to band gap computation. Applied Numerical Mathematics: Transactions of IMACS, 56(10–11):1283–1292, October/November 2006. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Baffet:2011:SHO

[BG11a] Daniel Baffet and Dan Givoli. On the stability of the high-order Higdon Absorbing Boundary Conditions. Applied Numerical Mathematics: Transactions of IMACS, 61(6): 768–784, June 2011. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Benzi:2011:DSP

[BG11b] Michele Benzi and Xue-Ping Guo. A dimensional split preconditioner for Stokes and linearized Navier–Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 61(1):66– 76, January 2011. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Boyd:2011:NEC

[BG11c] John P. Boyd and Kenneth W. Gildersleeve. Numerical experiments on the condition number of the interpolation matrices for radial basis functions. Applied Numerical Mathematics: Transactions of IMACS, 61(4):443–459, April 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Boffi:2014:SRF

[BG14] Daniele Boffi and Lucia Gastaldi. Some remarks on finite element approximation of multiple eigenvalues. Applied Numerical Mathematics: Transactions of IMACS, 79(??):18–28, May 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741200150X.

Berenguer:2024:ISL

[BG24] M. I. Berenguer and M. Ruiz Galán. Iterative schemes for linear equations of the second kind and related inverse problems. Applied Numerical Mathematics: Transactions of IMACS, 200(??): 110–123, June 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423002015.

Barrios:2004:NAN

[BGG04] Tomás P. Barrios, Gabriel N. Gatica, and Luis F. Gatica. On the numerical analysis of a nonlinear elliptic problem via mixed-FEM and Lagrange multipliers. Applied Numerical Mathematics: Transactions of IMACS, 48(2):135– 155, February 2004. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bermudez:2012:MMC

[BGG12] Alfredo Bermúdez and Luz M. García-García. Mathematical modeling in chemistry. Application to water quality problems. Applied Numerical Mathematics: Transactions of IMACS, 62(4): 305–327, April 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927411000845.

Benito:2020:SFP

[BGG^{+}20] J. J. Benito, A. García, L. Gavete, M. Negreanu, F. Ureña, and A. M. Vargas. Solving a fully parabolic chemotaxis system with periodic asymptotic behavior using Generalized Finite Difference Method. Applied Numerical Mathematics: Transactions of IMACS, 157(??): 356–371, November 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420301938.

Benito:2021:CNS

[BGG^{+}21] J. J. Benito, A. García, L. Gavete, M. Negreanu, F. Ureña, and A. M. Vargas. Convergence and numerical simulations of prey-predator interactions via a meshless method. Applied Numerical Mathematics: Transactions of IMACS, 161(??): 333–347, March 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303676.

Berenguer:2013:AMS

[BGGG13] M. I. Berenguer, A. I.

Garralda-Guillem, and M. Ruiz Galán. An approximation method for solving systems of Volterra integrodifferential equations. Applied Numerical Mathematics: Transactions of IMACS, 67(??):126–135, May 2013. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411000547.

Bingjie:2008:CGB

[BGH08] Li Bingjie, Kou Guangxing, and Zhao Huiwen. Conjugate gradient-boundary element solution using multiple reciprocity method for distributed elliptic optimal control problems. Applied Numerical Mathematics: Transactions of IMACS, 58(10):1534–1552, October 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Burger:2015:SIF

[BGH^{+}15] Raimund Bürger, Gabriel N. Gatica, Norbert Heuer, Rodolfo Rodríguez, and Mauricio Sepúlveda. Special issue: Fourth Chilean Workshop on Numerical Analysis of Partial Differential Equations (WONAPDE 2013), Universidad de Concepción, Chile. Applied Numerical Mathematics: Transactions of IMACS, 95(??): 1, September 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927415000756.

Burger:2012:PSI

[BGHR12] Raimund Bürger, Gabriel N. Gatica, Norbert Heuer, and Rodolfo Rodríguez. Preface: Special issue: Third Chilean Workshop on Numerical Analysis of Partial Differential Equations (WONAPDE), 11–15 January 2010, Universidad de Concepción, Chile. Applied Numerical Mathematics: Transactions of IMACS, 62(4):225, April 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412000128.

Brugnano:2018:LIS

[BGIW18] Luigi Brugnano, Gianmarco Gurioli, Felice Iavernaro, and Ewa B. Weinmüller. Line integral solution of Hamiltonian systems with holonomic constraints. Applied Numerical Mathematics: Transactions of IMACS, 127(??):56–77, May 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417302647.

Bermudez:2009:NST

[BGM^{+}09] A. Bermúdez, D. Gómez, M. C. Muñiz, P. Salgado, and R. Vázquez. Numerical simulation of a thermo-electromagneto-hydrodynamic problem in an induction heating furnace. Applied Numerical Mathematics: Transactions of IMACS, 59(9): 2082–2104, September 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bergou:2019:CNL

[BGM19] El Houcine Bergou, Serge Gratton, and Jan Mandel. On the convergence of a non-linear ensemble Kalman smoother. Applied Numerical Mathematics: Transactions of IMACS, 137(??): 151–168, March 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418302575.

Bank:2013:FRE

[BGO13] Randolph E. Bank, Luka Grubisić, and Jeffrey S. Ovall. A framework for robust eigenvalue and eigenvector error estimation and Ritz value convergence enhancement. Applied Numerical Mathematics: Transactions of IMACS, 66(??): 1–29, April 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927412001997.

Bouche:2011:OEE

[BGP11] Daniel Bouche, Jean-Michel Ghidaglia, and Frédéric P. Pascal. An optimal error estimate for upwind Finite Volume methods for nonlinear hyperbolic conservation laws. Applied Numerical Mathematics: Transactions of IMACS, 61(11): 1114–1131, November 2011.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Burgel:2002:CRD

[BGS02] A. Bürgel, T. Grahs, and Th. Sonar. From continuous recovery to discrete filtering in numerical approximations of conservation laws. Applied Numerical Mathematics: Transactions of IMACS, 42(1–3):47–60, August 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bauer:2006:ASI

[BGS06] Florian Bauer, Lars Grüne, and Willi Semmler. Adaptive spline interpolation for Hamilton–Jacobi–Bellman equations. Applied Numerical Mathematics: Transactions of IMACS, 56(9): 1196–1210, September 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bermudez:2014:P

[BGS14] Alfredo Bermúdez, Dolores Gómez, and Pilar Salgado. Preface. Applied Numerical Mathematics: Transactions of IMACS, 79(??): 1–2, May 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414000300.

Bellen:1997:ASP

[BGT97] A. Bellen, N. Guglielmi, and L. Torelli. Asymptotic stability properties of Θmethods for the pantograph equation. Applied Numerical Mathematics: Transactions of IMACS, 24(2– 3):279–293, June 19, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=24&
issue=2-3&aid=782. Volterra

centennial (Tempe, AZ, 1996).

Bultheel:2010:CRS

[BGVHN10] Adhemar Bultheel, Pablo González-Vera, Erik Hendriksen, and Olav Njåstad. Computation of rational Szegő–Lobatto quadrature formulas. Applied Numerical Mathematics: Transactions of IMACS, 60(12): 1251–1263, December 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Byrne:1985:ENM

[BH85] George D. Byrne and Alan C. Hindmarsh. Experiments in numerical methods for a problem in combustion modeling. Applied Numerical Mathematics: Transactions of IMACS, 1(1):29– 57, January 1985. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Brasey:1993:SHE

[BH93] V. Brasey and E. Hairer.

Symmetrized half-explicit methods for constrained mechanical systems. Applied Numerical Mathematics: Transactions of IMACS, 13(1–3):23–31, September 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Sixth Conference on the Numerical Treatment of Differential Equations (Halle, 1992).

Briggs:1996:TAD

[BH96] William L. Briggs and Van Emden Henson. A table of analytical discrete Fourier transforms. Applied Numerical Mathematics: Transactions of IMACS, 21(4):375– 384, November 22, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=21&
issue=4&aid=700.

Budd:1997:AMS

[BH97] C. J. Budd and A. R.

Humphries. Adaptive methods for semi-linear elliptic equations with critical exponents and interior singularities. Applied Numerical Mathematics: Transactions of IMACS, 26(1–2):227–240, December 22, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1997/26/
                                        1-2/849.pdf.

Bespalov:2010:H

[BH10] Alexei Bespalov and Norbert Heuer. The hp. Applied Numerical Mathematics: Transactions of IMACS, 60(7):705–718, July 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bejancu:2012:SUA

[BH12a] Aurelian Bejancu and Simon Hubbert. A study of the uniform accuracy of univariate thin plate spline interpolation. Applied Numerical Mathematics: Transactions of IMACS, 62(12): 1781–1789, December 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927412001213.

Burman:2012:FDF

[BH12b] Erik Burman and Peter Hansbo. Fictitious domain finite element methods using cut elements: II. A stabilized Nitsche method. Applied Numerical Mathematics: Transactions of IMACS, 62(4):328–341, April 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411000298.

Brooms:2020:VUB

[BH20] Anthony C. Brooms and Theodore C. Holtom. Volumetric uncertainty bounds and optimal configurations for converging beam triple LIDAR. Applied Numerical Mathematics: Transactions of IMACS, 153(??): 276–299, July 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300593.

Balani:2023:TBP

[BHB23] Fariba Bakrani Balani, Masoud Hajarian, and Luca Bergamaschi. Two block preconditioners for a class of double saddle point linear systems. Applied Numerical Mathematics: Transactions of IMACS, 190(??): 155–167, August 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742300106X.

Brezinski:2010:GTR

[BHHS10] Claude Brezinski, Yi He, Xing-Biao Hu, and Jian-Qing Sun. A generalization of the G-transformation and the related algorithms. Applied Numerical Mathematics: Transactions of IMACS, 60(12):1221–1230, December 2010. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Baines:2005:MMF

[BHJ05] M. J. Baines, M. E. Hubbard, and P. K. Jimack. A moving mesh finite element algorithm for the adaptive solution of timedependent partial differential equations with moving boundaries. Applied Numerical Mathematics: Transactions of IMACS, 54(3– 4):450–469, August 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bouhamidi:2013:MMN

[BHJ13] A. Bouhamidi, M. Hached, and K. Jbilou. A meshless method for the numerical computation of the solution of steady Burgers-type equations. Applied Numerical Mathematics: Transactions of IMACS, 74(??):95– 110, December 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413001165.

Baines:2006:SIM

[BHJJ06] M. J. Baines, M. E. Hubbard, P. K. Jimack, and A. C. Jones. Scale-invariant moving finite elements for nonlinear partial differential equations in two dimensions. Applied Numerical Mathematics: Transactions of IMACS, 56(2):230– 252, February 2006. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bai:2021:NMT

[BHL^{+}21] Xianglan Bai, Guang-Xin Huang, Xiao-Jun Lei, Lothar Reichel, and Feng Yin. A novel modified TRSVD method for large-scale linear discrete ill-posed problems. Applied Numerical Mathematics: Transactions of IMACS, 164(??): 72–88, ???? 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302543.

Bhowmik:2011:NAC

[Bho11] Samir Kumar Bhowmik. Numerical approximation of a convolution model of [θ̇]- neuron networks. Applied Numerical Mathematics: Transactions of IMACS, 61(4):581–592, April 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bhowmik:2012:NCO

[Bho12] Samir Kumar Bhowmik. Numerical convergence of a one step approximation of an integro-differential equation. Applied Numerical Mathematics: Transactions of IMACS, 62(12): 1880–1892, December 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927412001444.

Bostan:2005:PEE

[BHR05] Viorel Bostan, Weimin Han, and B. D. Reddy. A posteriori error estimation and adaptive solution of elliptic variational inequalities of the second kind. Applied Numerical Mathematics: Transactions of IMACS, 52(1):13–38, January 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Buccini:2021:CRM

[BHRY21] Alessandro Buccini, Guangxin Huang, Lothar Reichel, and Feng Yin. On the choice of regularization matrix for an l _{2}- l _{q} minimization method for image restoration. Applied Numerical Mathematics: Transactions of IMACS, 164(??):211–221, ???? 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303548.

Benner:2016:ILR

[BHSW16] Peter Benner, Matthias Heinkenschloss, Jens Saak, and Heiko K. Weichelt. An inexact low-rank Newton– ADI method for largescale algebraic Riccati equations. Applied Numerical Mathematics: Transactions of IMACS, 108(??): 125–142, October 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416300824.

Benner:2020:ESL

[BHSW20] Peter Benner, Matthias Heinkenschloss, Jens Saak, and Heiko K. Weichelt. Efficient solution of largescale algebraic Riccati equations associated with index-2 DAEs via the inexact low-rank Newton–ADI method. Applied Numerical Mathematics: Transactions of IMACS, 152(??): 338–354, June 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419303241.

Boen:2020:OSS

[Bi20] Lynn Boen and Karel J. in ’t Hout. Operator splitting schemes for American options under the twoasset Merton jump-diffusion model. Applied Numerical Mathematics: Transactions of IMACS, 153(??): 114–131, July 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300398.

Bica:2016:IVP

[Bic16] Alexandru Mihai Bica. Initial value problems with retarded argument solved by iterated quadratic splines. Applied Numerical Mathematics: Transactions of IMACS, 101(??):18–35, March 2016. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0168927415001592.

Bica:2021:CTP

[Bic21] Alexandru Mihai Bica. Corrigendum to “Two-point boundary value problems associated to functional differential equations of even order solved by iterated splines” [Appl. Numer. Math. 110 (2016) 128–147]. Applied Numerical Mathematics: Transactions of IMACS, 165(??):620–621, July 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/

science/article/pii/S0168927421000738.

See [BCC16].

Biegler:1987:PFO

[Bie87] L. T. Biegler. Process flowsheet optimization strategies: Recent results and future directions. Applied Numerical Mathematics: Transactions of IMACS, 3(5):393–408, October 1987. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bielski:2012:DNN

[Bie12] Sebastian Bielski. Dirichletto-Neumann and Neumannto-Dirichlet methods for eigenvalues and eigenfunctions of the Laplace operator. Applied Numerical Mathematics: Transactions of IMACS, 62(11): 1605–1619, November 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412000712.

Bras:2023:NCS

[BIJ23] Michal Braś, Giuseppe Izzo, and Zdzislaw Jackiewicz. A new class of G(ϵ)-symplectic general linear methods. Applied Numerical Mathematics: Transactions of IMACS, 183(??):1–14, January 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742200215X.

Brugnano:2015:EIR

[BIM15] Luigi Brugnano, Felice Iavernaro, and Cecilia Magherini. Efficient implementation of Radau collocation methods. Applied Numerical Mathematics: Transactions of IMACS, 87(??):100– 113, January 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741400155X.

Burger:2019:IEM

[BIMV19] Raimund Bürger, Daniel Inzunza, Pep Mulet, and Luis M. Villada. Implicitexplicit methods for a class of nonlinear nonlocal gradient flow equations modelling collective behaviour. Applied Numerical Mathematics: Transactions of IMACS, 144(??):234–252, October 2019. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927419301059.

Boujlida:2024:TLL

[BIO24] Hanen Boujlida, Kaouther Ismail, and Khaled Omrani. A three level linearized compact difference scheme for a fourth-order reaction–diffusion equation.

Applied Numerical Mathematics: Transactions of IMACS, 195(??):126–141, January 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423002428.

Birkhoff:1987:SWF

[Bir87] Garrett Birkhoff. Sound waves in fluids. Applied Numerical Mathematics: Transactions of IMACS, 3(1–2):3–24, May 1987. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0168927487900031.

Bishwal:2011:SMP

[Bis11] Jaya P. N. Bishwal. Stochastic moment problem and hedging of generalized Black– Scholes options. Applied Numerical Mathematics: Transactions of IMACS, 61(12):1271–1280, December 2011. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411001504.

Bulatov:2025:ANS

[BIS25] Mikhail Bulatov, Tatiana Indutskaya, and Liubov Solovarova. On an algorithm for the numerical solution of quasilinear integral-algebraic equations. Applied Numerical Mathematics: Transactions of IMACS, 208 (part B) (??):348–355, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424002745.

Brugnano:2024:WLS

[BIW24] Luigi Brugnano, Felice Iavernaro, and Ewa B. Weinmüller. Weighted least squares collocation methods. Applied Numerical Mathematics: Transactions of IMACS, 203(??): 113–128, September 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001235.

Butcher:1996:CDI

[BJ96] J. C. Butcher and Z. Jackiewicz. Construction of diagonally implicit general linear methods of type 1 and 2 for ordinary differential equations. Applied Numerical Mathematics: Transactions of IMACS, 21(4):385– 415, November 22, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=21&
issue=4&aid=725.

Butcher:1998:CHO

[BJ98] J. C. Butcher and Z. Jackiewicz. Construction of high order diagonally implicit multistage integration methods for ordinary differential equations. Applied Numerical Mathematics: Transactions of IMACS, 27(1):1– 12, May 1, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1998/27/
1/865.pdf.

Bouchon:2000:MMA

[BJ00] F. Bouchon and F. Jauberteau.

A multilevel method applied in the nonhomogeneous direction of the channel flow problem. Applied Numerical Mathematics: Transactions of IMACS, 36(1):1– 34, November 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.nl../26/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/65/26/26/article.
pdf.

Bujanda:2001:SRF

[BJ01] B. Bujanda and J. C. Jorge.

Stability results for fractional step discretizations of time dependent coefficient evolutionary problems. Applied Numerical Mathematics: Transactions of IMACS, 38(1–2):69–86, July 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/10/28/77/27/
                                        30/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/10/28/77/27/30/article.
                                        pdf.

Bao:2002:EER

[BJ02] Weizhu Bao and Shi Jin. Error estimates on the random projection methods for hyperbolic conservation laws with stiff reaction terms. Applied Numerical Mathematics: Transactions of IMACS, 43(4):315–333, December 2002. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bujanda:2003:FSR

[BJ03] B. Bujanda and J. C. Jorge. Fractional step Runge–Kutta methods for time dependent coefficient parabolic problems. Applied Numerical Mathematics: Transactions of IMACS, 45(2–3): 99–122, May 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bartoszewski:2005:NRT

[BJ05] Z. Bartoszewski and Z. Jackiewicz. Nordsieck representation of two-step Runge– Kutta methods for ordinary differential equations. Applied Numerical Mathematics: Transactions of IMACS, 53(2–4):149–163, May 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bujanda:2006:SRL

[BJ06] B. Bujanda and J. C. Jorge.

Stability results for linearly implicit fractional step discretizations of non-linear time dependent parabolic problems. Applied Numerical Mathematics: Transactions of IMACS, 56(8): 1061–1076, August 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bouhamidi:2011:SIR

[BJ11] A. Bouhamidi and K. Jbilou.

Stein implicit Runge–Kutta methods with high stage order for large-scale ordinary differential equations. Applied Numerical Mathematics: Transactions of IMACS, 61(2):149– 159, February 2011. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bras:2020:NCE

[BJ20] M. Braś and Z. Jackiewicz. A new class of efficient general linear methods for ordinary differential equations. Applied Numerical Mathematics: Transactions of IMACS, 151(??):282–300, May 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419303733.

Brown:2001:EDS

[BJM01] B. M. Brown, P. K. Jimack, and M. D. Mihajlović. An efficient direct solver for a class of mixed finite element problems. Applied Numerical Mathematics: Transactions of IMACS, 38(1– 2):1–20, July 2001. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/77/27/
27/00001188.pdf;   http:
//www.elsevier.nl/gej-
ng/10/10/28/77/27/27/abstract.
html.

Bouhamidi:2012:NAS

[BJS12] Abderrahman Bouhamidi, Khalide Jbilou, and Hassane Sadok. Numerical Analysis and Scientific Computation with Applications (NASCA). Applied Numerical Mathematics: Transactions of IMACS, 62(9): 1025, September 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412000864.

Bu:2020:STF

[BJTZ20] Weiping Bu, Lun Ji, Yifa Tang, and Jie Zhou. Space– time finite element method for the distributed-order time fractional reaction diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 152(??):446–465, June 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419303186.

Bartoszewski:2006:RES

[BK06] Z. Bartoszewski and M. Kwapisz.

Remarks on evaluation of spectral radius of operators arising in WR methods for linear systems of ODEs. Applied Numerical Mathematics: Transactions of IMACS, 56(3–4): 332–344, March/April 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bendahmane:2009:CFV

[BK09] Mostafa Bendahmane and Kenneth H. Karlsen. Convergence of a finite volume scheme for the bidomain model of cardiac tissue. Applied Numerical Mathematics: Transactions of IMACS, 59(9):2266–2284, September 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Brunk:2012:PPD

[BK12] Markus Brunk and Anne Kværnø. Positivity preserving discretization of time dependent semiconductor drift-diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 62(10): 1289–1301, October 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892741200102X.

Barrenechea:2017:AGF

[BK17] Gabriel R. Barrenechea and Petr Knobloch. Analysis of a group finite element formulation. Applied Numerical Mathematics: Transactions of IMACS, 118(??): 238–248, August 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417300806.

Bhardwaj:2021:MMT

[BK21a] Akanksha Bhardwaj and Alpesh Kumar. A meshless method for time fractional nonlinear mixed diffusion and diffusion-wave equation. Applied Numerical Mathematics: Transactions of IMACS, 160(??): 146–165, February 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302956.

Biala:2021:PCS

[BK21b] T. A. Biala and A. Q. M.

Khaliq. Predictor–corrector schemes for nonlinear spacefractional parabolic PDEs with time-dependent boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 160(??): 1–22, February 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302919.

Babaie-Kafaki:2022:NDS

[BKAG22] Saman Babaie-Kafaki, Zohre Aminifard, and Saeide Ghafoori. Nonmonotone diagonally scaled limited-memory BFGS methods with application to compressive sensing based on a penalty model. Applied Numerical Mathematics: Transactions of IMACS, 181(??):618–629, November 2022. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001829.

Bahlibi:2025:DMP

[BKK25] M. T. Bahlibi, J. Karátson, and S. Korotov. Discrete maximum principles with computable mesh conditions for nonlinear elliptic finite element problems. Applied Numerical Mathematics: Transactions of IMACS, 210(??):222–244, April 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424003507.

Bainov:1995:DMI

[BKM95] Drumi D. Baĭnov, Zdzisław Kamont, and Emil Minchev. Difference methods for impulsive differential–functional equations. Applied Numerical Mathematics: Transactions of IMACS, 16(4):401– 416, April 6, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=16&
issue=4&aid=554.

Billups:2005:I

[BKM05] Stephen Billups, Andrew Knyazev, and Jan Mandel. Introduction. Applied Numerical Mathematics: Transactions of IMACS, 54(2):105–106, July 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bansch:2013:EMM

[BKM13] E. Bänsch, F. Karakatsani, and Ch. Makridakis. The effect of mesh modification in time on the error control of fully discrete approximations for parabolic equations. Applied Numerical Mathematics: Transactions of IMACS, 67(??):35–63, May 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411001619.

Bahmanpour:2019:SFI

[BKM19] Maryam Bahmanpour, Majid Tavassoli Kajani, and Mohammad Maleki. Solving Fredholm integral equations of the first kind using Müntz wavelets. Applied Numerical Mathematics: Transactions of IMACS, 143(??): 159–171, September 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419300881.

Bartel:2009:WBA

[BKP09] A. Bartel, S. Knorr, and R. Pulch. Wavelet-based adaptive grids for multirate partial differential-algebraic equations. Applied Numerical Mathematics: Transactions of IMACS, 59(3–4): 495–506, March/April 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Biro:2014:FES

[BKP14] Oszkár Bíró, Gergely Koczka, and Kurt Preis. Finite element solution of nonlinear eddy current problems with periodic excitation and its industrial applications. Applied Numerical Mathematics: Transactions of IMACS, 79(??):3–17, May 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413000779.

Buoso:2015:GAP

[BKP15] Davide Buoso, Anna Karapiperi, and Stefano Pozza. Generalizations of Aitken’s process for a certain class of sequences. Applied Numerical Mathematics: Transactions of IMACS, 90(??): 38–54, April 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414001949.

Bebendorf:2013:EDG

[BKR13] M. Bebendorf, A. Kühnemund, and S. Rjasanow. An equi-directional generalization of adaptive cross approximation for higher-order tensors. Applied Numerical Mathematics: Transactions of IMACS, 74(??):1– 16, December 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413000950.

Bekas:2007:EDM

[BKS07] C. Bekas, E. Kokiopoulou, and Y. Saad. An estimator for the diagonal of a matrix. Applied Numerical Mathematics: Transactions of IMACS, 57(11– 12):1214–1229, November/ December 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Brandts:2008:P

[BKV08] Jan Brandts, Barry Koren, and Jan Verwer. Preface. Applied Numerical Mathematics: Transactions of IMACS, 58(12):1759–1760, December 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Budd:2006:NPS

[BKW06] Chris Budd, Othmar Koch, and Ewa Weinmüller. From nonlinear PDEs to singular ODEs. Applied Numerical Mathematics: Transactions of IMACS, 56(3–4): 413–422, March/April 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Brezinski:1986:AEF

[BL86] C. Brezinski and A. Lembarki. Acceleration of extended Fibonacci sequences.

Applied Numerical Mathematics: Transactions of IMACS, 2(1):1–8, February 1986. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Biringen:1991:NST

[BL91] S. Biringen and E. Laurien. Nonlinear structures of transition in wall-bounded flows. Applied Numerical Mathematics: Transactions of IMACS, 7(1):129– 150, January 1991. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/016892749190106A.

Balla:2005:APD

[BL05] Katalin Balla and Vu Hoang Linh. Adjoint pairs of differential-algebraic equations and Hamiltonian systems. Applied Numerical Mathematics: Transactions of IMACS, 53(2–4): 131–148, May 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bryson:2006:MWW

[BL06] Steve Bryson and Doron Levy. Mapped WENO and weighted power ENO reconstructions in semidiscrete central schemes for Hamilton–Jacobi equations. Applied Numerical Mathematics: Transactions of IMACS, 56(9):1211–1224, September 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Burman:2008:SFE

[BL08] E. Burman and A. Linke.

Stabilized finite element schemes for incompressible flow using Scott–Vogelius elements. Applied Numerical Mathematics: Transactions of IMACS, 58(11): 1704–1719, November 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bartlomiejczyk:2015:MLP

[BL15] Agnieszka Bartlomiejczyk and Henryk Leszczyński. Method of lines for physiologically structured models with diffusion. Applied Numerical Mathematics: Transactions of IMACS, 94(??):140–148, August 2015. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415000537.

Bai:2021:ORB

[BL21] Zhong-Zhi Bai and Kang-Ya Lu. Optimal rotated blockdiagonal preconditioning for discretized optimal control problems constrained with fractional time-dependent diffusive equations. Applied Numerical Mathematics: Transactions of IMACS, 163(??):126–146, May 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000192.

Black:2000:SEA

[Bla00] Kelly Black. Spectral element approximation of convection-diffusion type problems. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4): 373–379, May 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/61/27/
65/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/61/27/65/article.
pdf.

Blanes:2001:HON

[Bla01] Sergio Blanes. High order numerical integrators for differential equations using composition and processing of low order methods. Applied Numerical Mathematics: Transactions of IMACS, 37(3):289–306, May 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/73/31/
28/abstract.html;   http:

//www.elsevier.nl/gej-
ng/10/10/28/73/31/28/article.
pdf.

Bai:2017:NAU

[BLD17] Jianchao Bai, Jicheng Li, and Pingfan Dai. Novel alternating update method for low rank approximation of structured matrices. Applied Numerical Mathematics: Transactions of IMACS, 121(??):223– 233, November 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417301599.

Berbatov:2021:GFV

[BLJ21] K. Berbatov, B. S. Lazarov, and A. P. Jivkov. A guide to the finite and virtual element methods for elasticity. Applied Numerical Mathematics: Transactions of IMACS, 169(??): 351–395, November 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421002014.

Breda:2024:SEO

[BLL24] Dimitri Breda, Davide Liessi, and Sjoerd M. Verduyn Lunel. Spectra of evolution operators of a class of neutral renewal equations: Theoretical and numerical aspects. Applied Numerical Mathematics: Transactions of IMACS, 200(??): 124–137, June 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001836.

Beyrami:2017:SEA

[BLM17a] Hossein Beyrami, Taher Lotfi, and Katayoun Mahdiani. Stability and error analysis of the reproducing kernel Hilbert space method for the solution of weakly singular Volterra integral equation on graded mesh. Applied Numerical Mathematics: Transactions of IMACS, 120(??): 197–214, October 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417301344.

Bini:2017:STQ

[BLM17b] D. A. Bini, G. Latouche, and B. Meini. Shift techniques for Quasi-Birth and Death processes: Canonical factorizations and matrix equations. Applied Numerical Mathematics: Transactions of IMACS, 116(??): 24–36, June 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301684.

Bazm:2025:DMT

[BLN25] Sohrab Bazm, Pedro Lima, and Somayeh Nemati. Discretization methods and their extrapolations for two-dimensional nonlinear Volterra–Urysohn integral equations. Applied Numerical Mathematics: Transactions of IMACS, 208 (part B) (??):323–337, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424002733.

Budd:2001:SIA

[BLP01] C. J. Budd, B. Leimkuhler, and M. D. Piggott. Scaling invariance and adaptivity. Applied Numerical Mathematics: Transactions of IMACS, 39(3–4):261–288, December 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/79/34/29/abstract.
html.

Beltran-Larrotta:2023:CNS

[BLRGVR23] Carlos M. Beltrán-Larrotta, Diego A. Rueda-Gómez, and Élder J. Villamizar-Roa. On a chemotaxis-Navier– Stokes system with Lotka– Volterra competitive kinetics: Theoretical and numerical analysis. Applied Numerical Mathematics: Transactions of IMACS, 184(??):77– 100, February 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422002495.

Babuska:1994:PEM

[BLS94] I. Babuška, I. Lee, and C. Schwab. On the a posteriori estimation of the modeling error for the heat conduction in a plate and its use for adaptive hierarchical modeling. Applied Numerical Mathematics: Transactions of IMACS, 14(1–3):5– 21, April 14, 1994. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=14&
issue=1-3&aid=453. Pro-

ceedings of the Third ARO Workshop on Adaptive Methods for Partial Differential Equations (Troy, NY, 1992).

Bao:2017:AGN

[BLS^{+}17] Ji-Feng Bao, Chong Li, Wei-Ping Shen, Jen-Chih Yao, and Sy-Ming Guu. Approximate Gauss–Newton methods for solving underdetermined nonlinear least squares problems. Applied Numerical Mathematics: Transactions of IMACS, 111(??):92–110, January 2017. CODEN ANMAEL.

ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S016892741630157X.

Bailey:2002:CMT

[BLW02] C. Bailey, H. Lu, and D. Wheeler. Computational modeling techniques for reliability of electronic components on printed circuit boards. Applied Numerical Mathematics: Transactions of IMACS, 40(1–2):101–117, January 2002. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/85/27/36/abstract.
html.

Bao:2007:NPM

[BLW07] Liang Bao, Yiqin Lin, and Yimin Wei. A new projection method for solving large Sylvester equations. Applied Numerical Mathematics: Transactions of IMACS, 57(5–7):521–532, May 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bi:2016:AAB

[BLY16] Hai Bi, Hao Li, and Yidu Yang. An adaptive algorithm based on the shifted inverse iteration for the Steklov eigenvalue problem. Applied Numerical Mathematics: Transactions of IMACS, 105(??):64–81, July 2016.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416000258.

Bean:2017:BDP

[BLY17] Maranda Bean, Konstantin Lipnikov, and Son-Young Yi. A block-diagonal preconditioner for a fourfield mixed finite element method for Biot’s equations. Applied Numerical Mathematics: Transactions of IMACS, 122(??):1– 13, December 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417301654.

Bernardi:1989:ARS

[BM89] Christine Bernardi and Yvon Maday. Approximation results for spectral methods with domain decomposition. Applied Numerical Mathematics: Transactions of IMACS, 6(1–2):33– 52, December 1989. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Spectral multi-domain methods (Paris, 1988).

Beckett:2000:CAF

[BM00] G. Beckett and J. A. Mackenzie. Convergence analysis of finite difference approximations on equidistributed grids to a singularly perturbed boundary value problem. Applied Numerical Mathematics: Transactions of IMACS, 35(2):87– 109, October 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/64/30/
27/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/64/30/27/article.
pdf.

Beckett:2001:UCH

[BM01] G. Beckett and J. A.

Mackenzie. Uniformly convergent high order finite element solutions of a singularly perturbed reaction-diffusion equation using mesh equidistribution. Applied Numerical Mathematics: Transactions of IMACS, 39(1):31– 45, October 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/79/27/29/abstract.
html.

Brugnano:2002:BIB

[BM02] Luigi Brugnano and Cecilia Magherini. Blended implementation of block implicit methods for ODEs. Applied Numerical Mathematics: Transactions of IMACS, 42(1–3):29–45, August 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bertolazzi:2004:FVM

[BM04a] Enrico Bertolazzi and Gianmarco Manzini. A finite volume method for transport of contaminants in porous media. Applied Numerical Mathematics: Transactions of IMACS, 49(3–4): 291–305, June 2004. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bertolazzi:2004:LSB

[BM04b] Enrico Bertolazzi and Gianmarco Manzini. Least square-based finite volumes for solving the advection– diffusion of contaminants in porous media. Applied Numerical Mathematics: Transactions of IMACS, 51(4):451–461, December 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bertolazzi:2004:LSP

[BM04c] Enrico Bertolazzi and Gianmarco Manzini. Limiting strategies for polynomial reconstructions in the finite volume approximation of the linear advection equation. Applied Numerical Mathematics: Transactions of IMACS, 49(3–4): 277–289, June 2004. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bause:2005:HOR

[BM05] Markus Bause and Willi Merz. Higher order regularity and approximation of solutions to the Monod biodegradation model. Applied Numerical Mathematics: Transactions of IMACS, 55(2):154–172, October 2005. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Blanes:2006:FSO

[BM06a] S. Blanes and P. C. Moan.

Fourth- and sixth-order commutator-free Magnus integrators for linear and nonlinear dynamical systems. Applied Numerical Mathematics: Transactions of IMACS, 56(12):1519–1537, December 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Brugnano:2006:EEE

[BM06b] Luigi Brugnano and Cecilia Magherini. Economical error estimates for block implicit methods for ODEs via deferred correction. Applied Numerical Mathematics: Transactions of IMACS, 56(5):608–617, May 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Brugnano:2009:RAL

[BM09] Luigi Brugnano and Cecilia Magherini. Recent advances in linear analysis of convergence for splittings for solving ODE problems. Applied Numerical Mathematics: Transactions of IMACS, 59(3–4): 542–557, March/April 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bacuta:2012:MDS

[BM12a] Constantin Bacuta and Peter Monk. Multilevel discretization of symmetric saddle point systems without the discrete LBB condition. Applied Numerical Mathematics: Transactions of IMACS, 62(6):667–681, June 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411001280.

Borisov:2012:SDS

[BM12b] V. S. Borisov and M. Mond. On stability of difference schemes. Central schemes for hyperbolic conservation laws with source terms. Applied Numerical Mathematics: Transactions of IMACS, 62(8):895–921, August 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412000487.

Barrientos:2013:PEA

[BM13] Mauricio A. Barrientos and Matthias Maischak. Aposteriori error analysis to the exterior Stokes problem. Applied Numerical Mathematics: Transactions of IMACS, 63(1):25– 44, January 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412001638.

Brauss:2018:PDF

[BM18] K. D. Brauss and A. J.

Meir. On a parallel, 3dimensional, finite element solver for viscous, resistive, stationary magnetohydrodynamics equations: Velocitycurrent formulation. Applied Numerical Mathematics: Transactions of IMACS, 133(??):130–143, ???? 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418300230.

Berenguer:2012:SAS

[BMGGG12] M. I. Berenguer, M. V. Fernández Muñoz, A. I. Garralda-Guillem, and M. Ruiz Galán. A sequential approach for solving the Fredholm integro-differential equation. Applied Numerical Mathematics: Transactions of IMACS, 62(4):

297–304, April 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411000614.

Baeza:2012:ABI

[BMGM12] Antonio Baeza, Anna Martínez-Gavara, and Pep Mulet. Adaptation based on interpolation errors for high order mesh refinement methods applied to conservation laws. Applied Numerical Mathematics: Transactions of IMACS, 62(4): 278–296, April 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411001711.

Brunner:1997:MIC

[BMM97a] H. Brunner, A. Makroglou, and R. K. Miller. Mixed interpolation collocation methods for first and second order Volterra integro-differential equations with periodic solution. Applied Numerical Mathematics: Transactions of IMACS, 23(4):381– 402, May 26, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1997/23/
4/757.pdf.

Brunner:1997:MCM

[BMM97b] H. Brunner, A. Makroglou, and R. K. Miller. On mixed collocation methods for Volterra integral equations with periodic solution. Applied Numerical Mathematics: Transactions of IMACS, 24(2–3): 115–130, June 19, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=24&
issue=2-3&aid=771. Volterra

centennial (Tempe, AZ, 1996).

Bellavia:2003:AST

[BMM03] Stefania Bellavia, Maria Macconi, and Benedetta Morini. An affine scaling trust-region approach to bound-constrained nonlinear systems. Applied Numerical Mathematics: Transactions of IMACS, 44(3):257– 280, February 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bokanowski:2006:ADS

[BMMZ06] O. Bokanowski, S. Martin, R. Munos, and H. Zidani. An anti-diffusive scheme for viability problems. Applied Numerical Mathematics: Transactions of IMACS, 56(9):1147–1162, September 2006. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Botte:2005:NMI

[BMP05] V. Botte, D. Mansutti, and A. Pascarelli. Numerical modeling of iron corrosion due to an acidic aqueous solution. Applied Numerical Mathematics: Transactions of IMACS, 55(3): 253–263, November 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Branquinho:2015:SOD

[BMPR15] A. Branquinho, A. Foulquié Moreno, A. Paiva, and M. N. Rebocho. Second-order differential equations in the Laguerre–Hahn class. Applied Numerical Mathematics: Transactions of IMACS, 94(??):16–32, August 2015. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927415000392.

Bader:2016:VPR

[BMQW16] Philipp Bader, David I. McLaren, G. R. W. Quispel, and Marcus Webb. Volume preservation by Runge–Kutta methods. Applied Numerical Mathematics: Transactions of IMACS, 109(??):123–137, November 2016. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301040.

Bauer:2017:TSA

[BMR^{+}17a] S. Bauer, M. Mohr, U. Rüde, J. Weismüller, M. Wittmann, and B. Wohlmuth. A twoscale approach for efficient on-the-fly operator assembly in massively parallel high performance multigrid codes. Applied Numerical Mathematics: Transactions of IMACS, 122(??):14– 38, December 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417301642.

Bini:2017:ECR

[BMR17b] Dario A. Bini, Stefano Massei, and Leonardo Robol. Efficient cyclic reduction for Quasi-Birth–Death problems with rank structured blocks. Applied Numerical Mathematics: Transactions of IMACS, 116(??): 37–46, June 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301362.

Bernardi:1989:NMC

[BMS89] Christine Bernardi, Yvon Maday, and Giovanni Sacchi Landriani. Nonconforming matching conditions for coupling spectral and finite element methods. Applied Numerical Mathematics: Transactions of IMACS, 6(1–2):65–84, December 1989. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Spectral multidomain methods (Paris, 1988).

Bai:2021:ICA

[BMSZ21] Jianchao Bai, Yuxue Ma, Hao Sun, and Miao Zhang. Iteration complexity analysis of a partial LQP-based alternating direction method of multipliers. Applied Numerical Mathematics: Transactions of IMACS, 165(??): 500–518, July 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000842.

Brugnano:1993:PIB

[BMT93] L. Brugnano, F. Mazzia, and D. Trigiante. Parallel implementation of BVM methods. Applied Numerical Mathematics: Transactions of IMACS, 11(1– 3):115–124, January 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Parallel methods for ordinary differential equations (Grado, 1991).

Bizas:2025:EEM

[BMT25] Emmanouil Bizas, Marilena Mitrouli, and Ondrej Turek. Efficient estimates for matrix-inverse quadratic forms. Applied Numerical Mathematics: Transactions of IMACS, 208 (part A) (??):76–91, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927424000138.

Breda:2006:PAE

[BMV06] Dimitri Breda, Stefano Maset, and Rossana Vermiglio. Pseudospectral approximation of eigenvalues of derivative operators with non-local boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 56(3–4): 318–331, March/April 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bozorgnia:2019:FEE

[BMV19] Farid Bozorgnia, Seyyed Abbas Mohammadi, and Tomás Vejchodský. The first eigenvalue and eigenfunction of a nonlinear elliptic system. Applied Numerical Mathematics: Transactions of IMACS, 145(??): 159–174, November 2019.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419301564.

Bu:2020:ESN

[BMWH20] Linlin Bu, Liquan Mei, Ying Wang, and Yan Hou. Energy stable numerical schemes for the fractionalin-space Cahn–Hilliard equation. Applied Numerical Mathematics: Transactions of IMACS, 158(??): 392–414, December 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420302385.

Bai:1999:SPB

[BN99] Zhong-Zhi Bai and Reinhard Nabben. Some properties of the block matrices in the parallel decomposition-type relaxation methods. Applied Numerical Mathematics: Transactions of IMACS, 29(2):167–170, February 1, 1999. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1999/29/
                                        2/928.pdf.

Bourchtein:2003:DSS

[BN03] A. Bourchtein and M. Naumov. Dependence of the stability of semi-implicit atmospheric models on reference temperature profile. Applied Numerical Mathematics: Transactions of IMACS, 47(3–4):325–343, December 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Berg:2012:SAC

[BN12] Jens Berg and Jan Nordström. Spectral analysis of the continuous and discretized heat and advection equation on single and multiple domains. Applied Numerical Mathematics: Transactions of IMACS, 62(11): 1620–1638, November 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412000724.

Bendtsen:2001:SLN

[BNH01] Claus Bendtsen, Ole H.

Nielsen, and Lars B. Hansen. Solving large nonlinear generalized eigenvalue problems from Density Functional Theory calculations in parallel. Applied Numerical Mathematics: Transactions of IMACS, 37(1–2): 189–199, April 2001. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/73/26/
37/abstract.html;   http:
//www.elsevier.nl/gej-

                                        ng/10/10/28/73/26/37/article.
                                        pdf.

Beik:2020:ITR

[BNKR20] Fatemeh Panjeh Ali Beik, Mehdi Najafi-Kalyani, and Lothar Reichel. Iterative Tikhonov regularization of tensor equations based on the Arnoldi process and some of its generalizations. Applied Numerical Mathematics: Transactions of IMACS, 151(??):425–447, May 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742030012X.

Barman:2025:PUH

[BNS25] Mrityunjoy Barman, Srinivasan Natesan, and Ali Sendur. A parameteruniform hybrid method for singularly perturbed parabolic 2D convection– diffusion–reaction problems. Applied Numerical Mathematics: Transactions of IMACS, 207(??):111–135, January 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424002307.

Bermudez:2006:NSV

[BNV06] Alfredo Bermúdez, Maria R. Nogueiras, and Carlos Vázquez. Numerical solution of variational inequalities for pricing Asian options by higher order Lagrange–Galerkin methods. Applied Numerical Mathematics: Transactions of IMACS, 56(10–11):1256– 1270, October/November 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bui:1987:EWE

[BO87] T. D. Bui and A. K. Oppenheim. Evaluation of wind effects on model buildings by the random vortex method. Applied Numerical Mathematics: Transactions of IMACS, 3(1–2): 195–207, May 1987. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0168927487900122.

Bellew:2004:PRN

[BO04] S. Bellew and E. O’Riordan.

A parameter robust numerical method for a system of two singularly perturbed convection–diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 51(2–3): 171–186, November 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Boyd:2011:ECS

[BO11] John P. Boyd and Jun Rong Ong. Exponentially-convergent strategies for defeating the Runge Phenomenon for the approximation of nonperiodic functions, part two: Multi-interval polynomial schemes and multidomain Chebyshev interpolation. Applied Numerical Mathematics: Transactions of IMACS, 61(4): 460–472, April 2011. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bu:2021:HOS

[BO21] Linlin Bu and Cornelis W. Oosterlee. On high-order schemes for tempered fractional partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 165(??): 459–481, July 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000787.

Bar-On:2000:PSM

[BOEP00] Ilan Bar-On, Åke Edlund, and Uri Peskin. Parallel solution of the multidimensional Helmholtz/ Schrödinger equation using high order methods. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4):95–104, May 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/

gej-ng/29/17/21/61/27/
33/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/61/27/33/article.
pdf.

Boffi:2006:FEM

[Bof06] Daniele Boffi. On the finite element method on quadrilateral meshes. Applied Numerical Mathematics: Transactions of IMACS, 56(10–11):1271–1282, October/November 2006. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Boglaev:2000:DDB

[Bog00] Igor Boglaev. Domain decomposition in boundary layers for singularly perturbed problems. Applied Numerical Mathematics: Transactions of IMACS, 34(2–3): 145–166, July 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/62/29/
28/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/62/29/28/article.
pdf.

Boglaev:2012:MAM

[Bog12] Igor Boglaev. On modified accelerated monotone iterates for solving semilinear parabolic problems. Applied Numerical Mathematics: Transactions of IMACS, 62(12):1849–1863, December 2012. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927412001420.

Boglaev:2016:MIA

[Bog16] Igor Boglaev. Monotone iterative ADI method for solving coupled systems of nonlinear parabolic equations. Applied Numerical Mathematics: Transactions of IMACS, 108(??): 204–222, October 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416300927.

Boglaev:2020:NSN

[Bog20] Igor Boglaev. Numerical solutions of nonlinear integroelliptic systems by monotone iterative methods. Applied Numerical Mathematics: Transactions of IMACS, 147(??):31–42, January 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419302168.

Bohn:2003:NNB

[Boh03] Christian-Arved Bohn. A neural network based finite element grid generation algorithm for the simulation of light propagation. Applied Numerical Mathematics: Transactions of IMACS, 46(3–4):263–277, September 2003. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bohaienko:2021:RCF

[Boh21] Vsevolod Bohaienko. On the recurrent computation of fractional operator with Mittag-Leffler kernel. Applied Numerical Mathematics: Transactions of IMACS, 162(??):137–149, April 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303895.

Bokor:2003:SSO

[Bok03] Rózsa Horváth Bokor. Stochastically stable one-step approximations of solutions of stochastic ordinary differential equations. Applied Numerical Mathematics: Transactions of IMACS, 44(3):299–312, February 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Borsboom:1997:DEM

[Bor97] Mart Borsboom. Development of an error-minimizing adaptive grid method. Applied Numerical Mathematics: Transactions of IMACS, 26(1–2):13–21, December 22, 1997. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1997/26/
                                        1-2/833.pdf.

Borovykh:2002:SNS

[Bor02] Natalia Borovykh. Stability in the numerical solution of the heat equation with nonlocal boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 42(1–3): 17–27, August 2002. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Borghi:2010:AFE

[Bor10] Riccardo Borghi. Asymptotic and factorial expansions of Euler series truncation errors via exponential polynomials. Applied Numerical Mathematics: Transactions of IMACS, 60(12):1242–1250, December 2010. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Borsche:2016:NSN

[Bor16]  Raul Borsche.  Numeri-

cal schemes for networks of hyperbolic conservation laws. Applied Numerical Mathematics: Transactions of IMACS, 108(??): 157–170, October 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741630085X.

Boscarino:2009:ATO

[Bos09] Sebastiano Boscarino. On an accurate third order implicitexplicit Runge–Kutta method for stiff problems. Applied Numerical Mathematics: Transactions of IMACS, 59(7):1515–1528, July 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bottasso:1997:NLF

[Bot97] Carlo L. Bottasso. A new look at finite elements in time: a variational interpretation of Runge–Kutta methods. Applied Numerical Mathematics: Transactions of IMACS, 25(4):355– 368, November 10, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=25&
issue=4&aid=828.

Bourchtein:2002:SLS

[Bou02]  A. Bourchtein.  Semi-

Lagrangian semi-implicit space splitting regional baroclinic atmospheric model. Applied Numerical Mathematics: Transactions of IMACS, 40(1–2):307–326, January 2002. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/85/27/51/abstract.
html.

Boulton:2016:SPE

[Bou16]  Lyonell Boulton.  Spec-

tral pollution and eigenvalue bounds. Applied Numerical Mathematics: Transactions of IMACS, 99(??): 1–23, January 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415001270.

Boubekraoui:2025:ESM

[Bou25] Maryam Boubekraoui. Extrapolated splitting methods for multilinear PageRank computations. Applied Numerical Mathematics: Transactions of IMACS, 208 (part A)(??):92–103, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927423002982.

Boyd:1991:CNA

[Boy91a] John P. Boyd. A comparison of numerical and analytical methods for the reduced wave equation with multiple spatial scales. Applied Numerical Mathematics: Transactions of IMACS, 7(6):453–479, June 1991. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Boyd:1991:SAP

[Boy91b]  John P. Boyd.  Sum-

accelerated pseudospectral methods: the Euler-accelerated sinc algorithm. Applied Numerical Mathematics: Transactions of IMACS, 7(4): 287–296, April 1991. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0168927491900658.

Boyd:2006:CRR

[Boy06] John P. Boyd. Computing real roots of a polynomial in Chebyshev series form through subdivision. Applied Numerical Mathematics: Transactions of IMACS, 56(8):1077– 1091, August 2006. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Boyd:2007:RTE

[Boy07] John P. Boyd. Rootfinding for a transcendental equation without a first guess: Polynomialization of Kepler’s equation through Chebyshev polynomial expansion of the sine. Applied Numerical Mathematics: Transactions of IMACS, 57(1):12–18, January 2007.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Boyd:2015:CET

[Boy15]  John P. Boyd.  Con-

vergence and error theorems for Hermite function pseudo-RBFs: Interpolation on a finite interval by Gaussian-localized polynomials. Applied Numerical Mathematics: Transactions of IMACS, 87(??):125– 144, January 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414001573.

Bozorgnia:2011:NST

[Boz11] F. Bozorgnia. Numerical solutions of a two-phase membrane problem. Applied Numerical Mathematics: Transactions of IMACS, 61(1):92–107, January 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bramble:1985:BPA

[BP85] James H. Bramble and Joseph E. Pasciak. A boundary parametric approximation to the linearized scalar potential magnetostatic field problem. Applied Numerical Mathematics: Transactions of IMACS, 1(6):493– 514, November 1985. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bramble:1990:DDT

[BP90] James H. Bramble and Joseph E. Pasciak. A domain decomposition technique for Stokes problems. Applied Numerical Mathematics: Transactions of IMACS, 6(4):251–261, May 1990. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Brewer:1992:PEV

[BP92] Dennis W. Brewer and Robert K. Powers. Parameter estimation for a Volterra integro-differential equation. Applied Numerical Mathematics: Transactions of IMACS, 9(3–5): 307–320, April 1992. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0168927492900248.

International Conference on the Numerical Solution of Volterra and Delay Equations (Tempe, AZ, 1990).

Bwemba:1995:IMU

[BP95] R. Bwemba and R. Pasquetti. On the influence matrix used in the spectral solution of the 2D Stokes problem (vorticitystream function formulation). Applied Numerical Mathematics: Transactions of IMACS, 16(3):299–315, February 16, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=16&
issue=3&aid=532.

Bonnet:1997:BAB

[BP97] Frédéric Bonnet and Frédéric Poupaud. Berenger absorbing boundary condition with time finite-volume scheme for triangular meshes. Applied Numerical Mathematics: Transactions of IMACS, 25(4):333–354, November 10, 1997. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=25&
issue=4&aid=830.

Barrio:2002:NED

[BP02] R. Barrio and J. M. Peña. Numerical evaluation of the p th derivative of Jacobi series. Applied Numerical Mathematics: Transactions of IMACS, 43(4): 335–357, December 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Baker:2006:DSN

[BP06a] Christopher T. H. Baker and Christopher A. H. Paul. Discontinuous solutions of neutral delay differential equations. Applied Numerical Mathematics: Transactions of IMACS, 56(3–4): 284–304, March/April 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Butcher:2006:EEG

[BP06b] J. C. Butcher and H. Podhaisky. On error estimation in general linear methods for stiff ODEs. Applied Numerical Mathematics: Transactions of IMACS, 56(3–4): 345–357, March/April 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Becache:2012:RSC

[BP12a] Eliane Bécache and Andrés Prieto. Remarks on the stability of Cartesian PMLs in corners. Applied Numerical Mathematics: Transactions of IMACS, 62(11): 1639–1653, November 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412000748.

Blanes:2012:TAE

[BP12b] S. Blanes and E. Ponsoda.

Time-averaging and exponential integrators for nonhomogeneous linear IVPs and BVPs. Applied Numerical Mathematics: Transactions of IMACS, 62(8):875– 894, August 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741200027X.

Ballestra:2014:VFA

[BP14] Luca Vincenzo Ballestra and Graziella Pacelli. A very fast and accurate boundary element method for options with moving barrier and time-dependent rebate. Applied Numerical Mathematics: Transactions of IMACS, 77(??):1–15, March 2014. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413001384.

Balasubramani:2024:FQS

[BPN24] N. Balasubramani, M. Guru Prem Prasad, and S. Natesan. Fractal quintic spline solutions for singularly perturbed reaction–diffusion boundary-value problems. Applied Numerical Mathematics: Transactions of IMACS, 202(??):89–99, August 2024. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001065.

Banz:2019:HSH

[BPS19] Lothar Banz, Jan Petsche, and Andreas Schröder. Hybridization and stabilization for hp-finite element methods. Applied Numerical Mathematics: Transactions of IMACS, 136(??):66– 102, February 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418302149.

Britt:2015:CSS

[BPTT15] S. Britt, S. Petropavlovsky, S. Tsynkov, and E. Turkel. Computation of singular solutions to the Helmholtz equation with high order accuracy. Applied Numerical Mathematics: Transactions of IMACS, 93(??): 215–241, July 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414001779.

Bruni:2025:FPP

[BPV25] V. Bruni, F. Pelosi, and D. Vitulano. Fractal properties of 4-point interpolatory subdivision schemes and wavelet scattering transform for signal classification. Applied Numerical Mathematics: Transactions of IMACS, 208 (part B) (??):256–270, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424002599.

Bocharov:1994:NTP

[BR94] G. A. Bocharov and A. A. Romanyukha. Numerical treatment of the parameter identification problem for delay-differential systems arising in immune response modelling. Applied Numerical Mathematics: Transactions of IMACS, 15(3): 307–326, October 5, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=15&
issue=3&aid=483. Interna-

tional Conference on Scientific Computation and Differential Equations (Auckland, 1993).

Brunner:1997:CST

[BR97] Hermann Brunner and Helmut Roth. Collocation in space and time: experience with the Korteweg–de Vries equation. Applied Numerical Mathematics: Transactions of IMACS, 25(4):369– 390, November 10, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/

cas_sub/browse/browse.
cgi?year=1997&volume=25&
issue=4&aid=854.

Berti:2001:COP

[BR01] Andrea C. Berti and A. Sri Ranga. Companion orthogonal polynomials: some applications. Applied Numerical Mathematics: Transactions of IMACS, 39(2):127–149, November 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/79/33/27/abstract.
html.

Butcher:2005:AMS

[BR05] John C. Butcher and Nicolette Rattenbury. ARK methods for stiff problems. Applied Numerical Mathematics: Transactions of IMACS, 53(2–4):165–181, May 2005. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Burger:2009:P

[BR09] Raimund Bürger and Rodolfo Rodríguez. Preface. Applied Numerical Mathematics: Transactions of IMACS, 59(9):2035, September 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bragin:2020:CLM

[BR20] Michael D. Bragin and Boris V. Rogov. Conservative limiting method for high-order bicompact schemes as applied to systems of hyperbolic equations. Applied Numerical Mathematics: Transactions of IMACS, 151(??): 229–245, May 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300064.

Barzegar:2025:SEM

[BR25] Feze Barzegar and Jalil Rashidinia. Spectral element method for the solution of viscoelastic seismic wave propagation. Applied Numerical Mathematics: Transactions of IMACS, 212(??):92–109, June 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425000157.

Brandts:2000:STO

[Bra00] Jan H. Brandts. Superconvergence for triangular order k =1 Raviart–Thomas mixed finite elements and for triangular standard quadratic finite element methods. Applied Numerical Mathematics: Transactions of IMACS, 34(1):39–58, June 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/

gej-ng/29/17/21/62/27/
29/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/62/27/29/article.
pdf.

Bragin:2022:HOB

[Bra22] Michael D. Bragin. Highorder bicompact schemes for the quasilinear multidimensional diffusion equation. Applied Numerical Mathematics: Transactions of IMACS, 174(??): 112–126, April 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422000125.

Boykov:2018:NIM

[BRBB18] I. V. Boykov, V. A. Roudnev, A. I. Boykova, and O. A. Baulina. New iterative method for solving linear and nonlinear hypersingular integral equations. Applied Numerical Mathematics: Transactions of IMACS, 127(??):280–305, May 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418300199.

Broekhuizen:2008:IGS

[BRBM08] N. Broekhuizen, Graham J.

Rickard, J. Bruggeman, and A. Meister. An improved and generalized second order, unconditionally positive, mass conserving integration scheme for biochemical systems. Applied Numerical Mathematics: Transactions of IMACS, 58(3): 319–340, March 2008. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Brezinski:1985:PPS

[Bre85] Claude Brezinski. Prediction properties of some extrapolation methods. Applied Numerical Mathematics: Transactions of IMACS, 1(6):457–462, November 1985. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Brezinski:1988:ABC

[Bre88] Claude Brezinski. On the asymptotic behaviour of continued fractions. Applied Numerical Mathematics: Transactions of IMACS, 4(2–4):231–239, June 1988. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Brezinski:1991:GCD

[Bre91] Claude Brezinski. Generalizations of the Christoffel– Darboux identity for adjacent families of orthogonal polynomials. Applied Numerical Mathematics: Transactions of IMACS, 8(3):193–199, October 1991. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Brezinski:1996:EAP

[Bre96] C. Brezinski. Extrapolation algorithms and Padé approximations: a historical survey. Applied Numerical Mathematics: Transactions of IMACS, 20(3):299– 318, May 22, 1996. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=3&aid=658. Selected

keynote papers presented at 14th IMACS World Congress (Atlanta, GA, 1994).

Brenner:2002:SMD

[Bre02a] Susanne C. Brenner. Smoothers, mesh dependent norms, interpolation and multigrid. Applied Numerical Mathematics: Transactions of IMACS, 43(1–2):45–56, October 2002. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Brezinski:2002:BVS

[Bre02b] C. Brezinski. Biorthogonal vector sequence transformations and Padé approximation of vector series. Applied Numerical Mathematics: Transactions of IMACS, 41(4):437–442, July 2002.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Breda:2006:SOA

[Bre06] Dimitri Breda. Solution operator approximations for characteristic roots of delay differential equations. Applied Numerical Mathematics: Transactions of IMACS, 56(3–4):305–317, March/April 2006. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Brezinski:2010:NQP

[Bre10] C. Brezinski. From numerical quadrature to Padé approximation. Applied Numerical Mathematics: Transactions of IMACS, 60(12): 1209–1220, December 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Briggs:1985:PSD

[Bri85] William L. Briggs. Phenomenology of a six-dimensional mapping. Applied Numerical Mathematics: Transactions of IMACS, 1(3):239– 259, May 1985. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Barrera-Rosillo:2008:BBR

[BRIP08] Domingo Barrera-Rosillo and María José Ibáñez-Pérez. Bernstein–Bézier representation and nearminimally normed discrete quasi-interpolation operators. Applied Numerical Mathematics: Transactions of IMACS, 58(1):59–68, January 2008. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bellalij:2015:BMF

[BRRS15] M. Bellalij, L. Reichel, G. Rodriguez, and H. Sadok. Bounding matrix functionals via partial global block Lanczos decomposition. Applied Numerical Mathematics: Transactions of IMACS, 94(??):127–139, August 2015. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415000549.

Belgacem:2005:MFS

[BRS05] F. Ben Belgacem, Y. Renard, and L. Slimane. A mixed formulation for the Signorini problem in nearly incompressible elasticity. Applied Numerical Mathematics: Transactions of IMACS, 54(1):1–22, June 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bracciali:2016:POP

[BRS16] C. F. Bracciali, A. Sri Ranga, and A. Swaminathan. Paraorthogonal polynomials on the unit circle satisfying three term recurrence formulas. Applied Numerical Mathematics: Transactions of IMACS, 109(??):19– 40, November 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416300964.

Burkotova:2018:NSB

[BRS^{+}18] Jana Burkotová, Irena Rachunková, Svatoslav Stanek, Ewa B. Weinmüller, and Stefan Wurm. On nonlinear singular BVPs with nonsmooth data. Part 1: Analytical results. Applied Numerical Mathematics: Transactions of IMACS, 130(??): 23–50, August 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418300734.

Beernaert:1991:MSE

[BRSD91] L. Beernaert, D. Roose, R. Struys, and H. Deconinck. A multigrid solver for the Euler equations on the iPSC/2 parallel computer. Applied Numerical Mathematics: Transactions of IMACS, 7(5): 379–398, June 1991. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/016892749190009O.

Ben-Romdhane:2019:IFD

[BRTB19] Mohamed Ben-Romdhane, Helmi Temimi, and Mahboub Baccouch. An iterative finite difference method for approximating the twobranched solution of Bratu’s problem. Applied Numerical Mathematics: Transactions of IMACS, 139(??): 62–76, May 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419300054.

Brunner:1992:ILC

[Bru92] Hermann Brunner. Implicitly linear collocation methods for nonlinear Volterra equations. Applied Numerical Mathematics: Transactions of IMACS, 9(3–5): 235–247, April 1992. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0168927492900189.

International Conference on the Numerical Solution of Volterra and Delay Equations (Tempe, AZ, 1990).

Bruder:1993:LIR

[Bru93] Jürgen Bruder. Linearlyimplicit Runge–Kutta methods based on implicit Runge– Kutta methods. Applied Numerical Mathematics: Transactions of IMACS, 13(1–3):33–40, September 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Sixth Conference on the Numerical Treatment of Differential Equations (Halle, 1992).

Brunner:1997:HBO

[Bru97] Hermann Brunner. Hermann Brunner, 1896–1996: One hundred years of Volterra integral equations of the first kind. Applied Numerical Mathematics: Transactions of IMACS, 24(2– 3):83–93, June 19, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=24&
issue=2-3&aid=769. Volterra

centennial (Tempe, AZ, 1996).

Brunner:2007:HOC

[Bru07] Hermann Brunner. Highorder collocation methods for singular Volterra functional equations of neutral type. Applied Numerical Mathematics: Transactions of IMACS, 57(5–7):533–548, May 2007. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bernardi:2009:PEA

[BRVC09] C. Bernardi, T. Chacón Rebollo, E. Chacón Vera, and D. Franco Coronil. A posteriori error analysis for two non-overlapping domain decomposition techniques. Applied Numerical Mathematics: Transactions of IMACS, 59(6):1214–1236, June 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Burkotova:2017:SBN

[BRW17] Jana Burkotová, Irena Rachunková, and Ewa B. Weinmüller. On singular BVPs with nonsmooth data: Analysis of the linear case with variable coefficient matrix. Applied Numerical Mathematics: Transactions of IMACS, 114(??):77–96, April 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301015.

Bai:2021:EPR

[BRW21] Zhong-Zhi Bai, Lothar Reichel, and Zeng-Qi Wang. Editorial preface: Recent advances in numerical algebra with interdisciplinary applications. Applied Numerical Mathematics: Transactions of IMACS, 164(??): 1–2, ???? 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420303214.

Brezinski:2010:EDP

[BRZ10] C. Brezinski and M. Redivo-Zaglia. Extensions of Drummond’s process for convergence acceleration. Applied Numerical Mathematics: Transactions of IMACS, 60(12):1231–1241, December 2010. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Brezinski:2017:SFT

[BRZ17] Claude Brezinski and Michela Redivo-Zaglia. Shanks function transformations in a vector space. Applied Numerical Mathematics: Transactions of IMACS, 116(??): 57–63, June 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892741630109X.

Brezinski:2010:P

[BRZW10] Claude Brezinski, Michela Redivo-Zaglia, and Ernst Joachim Weniger. Preface. Applied Numerical Mathematics: Transactions of IMACS, 60(12):1183, December 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bramley:1991:DDP

[BS91] R. Bramley and A. Sameh.

Domain decomposition for parallel row projection algorithms. Applied Numerical Mathematics: Transactions of IMACS, 8(4–5): 303–315, November 1991. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bogacki:1992:PII

[BS92] P. Bogacki and L. F.

Shampine. Parallel implementation of interpolants for Runge–Kutta pairs. Applied Numerical Mathematics: Transactions of IMACS, 9(3–5):371–383, April 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/016892749290028C.

International Conference on the Numerical Solution of Volterra and Delay Equations (Tempe, AZ, 1990).

Brezinski:1993:LTA

[BS93] Claude Brezinski and Hassane Sadok. Lanczos-type algorithms for solving systems of linear equations. Applied Numerical Mathematics: Transactions of IMACS, 11(6):443–473, April 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Berger:1994:AC

[BS94a] Marsha J. Berger and Jeff S. Saltzman. AMR on the CM-2. Applied Numerical Mathematics: Transactions of IMACS, 14(1–3): 239–253, April 14, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=14&
issue=1-3&aid=452. Pro-

ceedings of the Third ARO Workshop on Adaptive Methods for Partial Differential Equations (Troy, NY, 1992). Also, RIACS Report No. 92-16.

Biswas:1994:NPD

[BS94b] R. Biswas and R. C. Strawn. A new procedure for dynamic adaption of threedimensional unstructured grids. Applied Numerical Mathematics: Transactions of IMACS, 13(6):437– 452, February 1994. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Biswas:1996:MQC

[BS96a] Rupak Biswas and Roger C. Strawn. Mesh quality control for multiply-refined tetrahedral grids. Applied Numerical Mathematics: Transactions of IMACS, 20(4):337– 348, June 3, 1996. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=4&aid=647.

Butcher:1996:NCR

[BS96b] J. C. Butcher and J. M.

Sanz-Serna. The number of conditions for a Runge– Kutta method to have effective order p. Applied Numerical Mathematics: Transactions of IMACS, 22(1–3): 103–111, December 4, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=22&
issue=1-3&aid=710. Special

issue celebrating the centenary of Runge–Kutta methods.

Biswas:1997:THM

[BS97a] Rupak Biswas and Roger C.

Strawn. Tetrahedral and hexahedral mesh adaptation for CFD problems. Applied Numerical Mathematics: Transactions of IMACS, 26(1–2):135–151, December 22, 1997. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/

                                        store/apnum/sub/1997/26/
                                        1-2/842.pdf.

Braess:1997:ESS

[BS97b] D. Braess and R. Sarazin. An efficient smoother for the Stokes problem. Applied Numerical Mathematics: Transactions of IMACS, 23(1):3–19, March 7, 1997. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1997/23/

1/741.pdf. Multilevel methods (Oberwolfach, 1995).

Belikov:2000:NSI

[BS00a] Vitali V. Belikov and Andrei Yu. Semenov. Non-Sibsonian interpolation on arbitrary system of points in Euclidean space and adaptive isolines generation. Applied Numerical Mathematics: Transactions of IMACS, 32(4):371–387, April 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/21/59/32/
                                        29/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/21/59/32/29/article.
                                        pdf.

Butcher:2000:CPP

[BS00b] J. C. Butcher and A. D. Singh. The choice of parameters in parallel general linear methods for stiff problems. Applied Numerical Mathematics: Transactions of IMACS, 34(1):59–84, June 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/62/27/
30/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/62/27/30/article.
pdf.

Buttner:2002:RKM

[BS02] Jörg Büttner and Bernd Simeon. Runge–Kutta methods in elastoplasticity. Applied Numerical Mathematics: Transactions of IMACS, 41(4):443–458, July 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Beuchler:2005:OET

[BS05] Sven Beuchler and Joachim Schöberl. Optimal extensions on tensor-product meshes. Applied Numerical Mathematics: Transactions of IMACS, 54(3– 4):391–405, August 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Banas:2006:EEL

[BS06] L’ubomír Baňas and Marián Slodička. Error estimates for Landau–Lifshitz–Gilbert equation with magnetostriction. Applied Numerical Mathematics: Transactions of IMACS, 56(8):1019–1039, August 2006. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Broadhurst:2008:PSE

[BS08] Michael S. Broadhurst and Spencer J. Sherwin. The parabolised stability equations for 3D-flows: implementation and numerical stability. Applied Numerical Mathematics: Transactions of IMACS, 58(7): 1017–1029, July 2008. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Boyd:2009:SAE

[BS09] John P. Boyd and Alan M. Sousa. An SVD analysis of equispaced polynomial interpolation. Applied Numerical Mathematics: Transactions of IMACS, 59(10): 2534–2547, October 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bertaccini:2010:UPN

[BS10] Daniele Bertaccini and Fiorella Sgallari. Updating preconditioners for nonlinear deblurring and denoising image restoration. Applied Numerical Mathematics: Transactions of IMACS, 60(10):994– 1006, October 2010. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Buckwar:2012:SAA

[BS12] Evelyn Buckwar and Thorsten Sickenberger. A structural analysis of asymptotic meansquare stability for multidimensional linear stochastic differential systems. Applied Numerical Mathematics: Transactions of IMACS, 62(7):842–859, July 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412000475.

Back:2014:FEH

[BS14a] Aurore Back and Eric Sonnendrücker. Finite Element Hodge for spline discrete differential forms. Application to the Vlasov–Poisson system. Applied Numerical Mathematics: Transactions of IMACS, 79(??): 124–136, May 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414000117.

Banz:2014:PEE

[BS14b] Lothar Banz and Ernst P.

Stephan. A posteriori error estimates of hp-adaptive IPDG-FEM for elliptic obstacle problems. Applied Numerical Mathematics: Transactions of IMACS, 76(??):76–92, February 2014. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927413001372.

Bulatov:2018:SRP

[BS18] Mikhail Bulatov and Liubov Solovarova. On selfregularization properties of a difference scheme for linear differential–algebraic equations. Applied Numerical Mathematics: Transactions of IMACS, 130(??): 86–94, August 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418300771.

Bouharguane:2020:LDG

[BS20a] Afaf Bouharguane and Nour Seloula. The local discontinuous Galerkin method for convection–diffusion– fractional anti-diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 148(??):61– 78, February 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419302405.

Bulatov:2020:CVD

[BS20b] Mikhail Bulatov and Liubov Solovarova. Collocation— variation difference schemes with several collocation points for differential–algebraic equations. Applied Numerical Mathematics: Transactions of IMACS, 149(??): 153–163, ???? 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419301667.

Banei:2021:CAS

[BS21] Siamak Banei and Kamal Shanazari. On the convergence analysis and stability of the RBF-adaptive method for the forwardbackward heat problem in 2D. Applied Numerical Mathematics: Transactions of IMACS, 159(??):297– 310, January 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302506.

Bulai:2024:ABA

[BS24] Iulia Martina Bulai and Martina Salvia. Approximation of basins of attraction for bistable dynamical system for olive disease control. Applied Numerical Mathematics: Transactions of IMACS, 200(??):138–147, June 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001253.

Barrera-Sanchez:2002:SEO

[BSFDM02] Pablo Barrera-Sánchez, Guilmer F. González Flores, and Francisco Javier Domínguez-Mota. Some experiences on orthogonal grid generation. Applied Numerical Mathematics: Transactions of IMACS, 40(1–2):179–190, January 2002. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        10/28/85/27/42/abstract.
                                        html.

Babuska:1994:SIL

[BSGU94] I. Babuška, T. Strouboulis, S. K. Gangaraj, and C. S. Upadhyay. η-superconvergence in the interior of locally refined meshes of quadrilaterals: superconvergence of the gradient in finite element solutions of Laplace’s and Poisson’s equations. Applied Numerical Mathematics: Transactions of IMACS, 16(1–2):3–49, December 1, 1994. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1994&volume=16&
                                        issue=1-2&aid=505.  A

Festschrift to honor Professor Robert Vichnevetsky on his 65th birthday.

Bucchignani:2004:PMS

[BSP04] Edoardo Bucchignani, Fulvio Stella, and Fabio Paglia. A partition method for the solution of a coupled liquidstructure interaction problem. Applied Numerical Mathematics: Transactions of IMACS, 51(4):463–475, December 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Blaisdell:1996:EFN

[BSQ96] G. A. Blaisdell, E. T. Spyropoulos, and J. H. Qin. The effect of the formulation of nonlinear terms on aliasing errors in spectral methods. Applied Numerical Mathematics: Transactions of IMACS, 21(3): 207–219, August 20, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=21&
issue=3&aid=687.

Buckwar:2022:SMS

[BSTT22] Evelyn Buckwar, Adeline Samson, Massimiliano Tamborrino, and Irene Tubikanec. A splitting method for SDEs with locally Lipschitz drift: Illustration on the FitzHugh–Nagumo model. Applied Numerical Mathematics: Transactions of IMACS, 179(??): 191–220, September 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422001118.

Bisognin:2009:NAF

[BSV09] Vanilde Bisognin, Mauricio Sepúlveda, and Octavio Vera. On the nonexistence of asymptotically free solutions for a coupled nonlinear Schrödinger system. Applied Numerical Mathematics: Transactions of IMACS, 59(9):2285–2302, September 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bollhofer:2021:HPL

[BSV21] Matthias Bollhöfer, Olaf Schenk, and Fabio Verbosio. A high performance levelblock approximate LU factorization preconditioner algorithm. Applied Numerical Mathematics: Transactions of IMACS, 162(??): 265–282, April 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420304049.

Botchev:1999:SCA

[BSvdV99] Mikhail A. Botchev, Gerard L. G. Sleijpen, and Henk A. van der Vorst. Stability control for approximate implicit time stepping schemes with minimum residual iterations. Applied Numerical Mathematics: Transactions of IMACS, 31(3): 239–253, November 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/46/19/
17/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/46/19/17/article.
pdf.

Buttner:1993:APL

[BSW93] Mario Büttner, Bernhard A.

Schmitt, and Rüdiger Weiner. Automatic partitioning in linearly-implicit Runge–Kutta methods. Applied Numerical Mathematics: Transactions of IMACS, 13(1– 3):41–55, September 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Sixth Conference on the Numerical Treatment of Differential Equations (Halle, 1992).

Berkelaar:1999:PPD

[BSZ99] Arjan B. Berkelaar, Jos F.

Sturm, and Shuzhong Zhang. Polynomial primal–dual cone affine scaling for semidefinite programming. Applied Numerical Mathematics: Transactions of IMACS, 29(3):317–333, March 1, 1999. CODEN ANMAEL.

ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.com/
                                        cas/tree/store/apnum/sub/
                                        1999/29/3/965.pdf.

Brenner:2015:PPP

[BSZ15] Susanne C. Brenner, Li-Yeng Sung, and Yi Zhang. Post-processing procedures for an elliptic distributed optimal control problem with pointwise state constraints. Applied Numerical Mathematics: Transactions of IMACS, 95(??): 99–117, September 2015. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927415000379.

Bi:2022:NDM

[BSZ22] Yuxin Bi, Li Shan, and Haicheng Zhang. New decoupled method for the evolutionary dual-porosity-Stokes model with Beavers–Joseph interface conditions. Applied Numerical Mathematics: Transactions of IMACS, 175(??):73–97, May 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422000265.

Brugnano:1993:PPT

[BT93a] L. Brugnano and D. Trigiante. A parallel preconditioning technique for boundary value methods. Applied Numerical Mathematics: Transactions of IMACS, 13(4):277–290, November 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Brugnano:1993:SPS

[BT93b] L. Brugnano and D. Trigiante. Stability properties of some boundary value methods. Applied Numerical Mathematics: Transactions of IMACS, 13(4): 291–304, November 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Beauwens:1994:GP

[BT94] R. Beauwens and B. Tombuyses. Graph perturbations. Applied Numerical Mathematics: Transactions of IMACS, 16(1–2):251–264, December 1, 1994. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=16&
issue=1-2&aid=521.  A

Festschrift to honor Professor Robert Vichnevetsky on his 65th birthday.

Brugnano:1995:HOM

[BT95] L. Brugnano and D. Trigiante. High-order multistep methods for boundary value problems. Applied Numerical Mathematics: Transactions of IMACS, 18(1–3): 79–94, September 27, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=18&
issue=1-3&aid=593. Sev-

enth Conference on the Numerical Treatment of Differential Equations (Halle, 1994).

Baker:1997:SAC

[BT97a] Christopher T. H. Baker and Arsalang Tang. Stability analysis of continuous implicit Runge–Kutta methods for Volterra integrodifferential systems with unbounded delays. Applied Numerical Mathematics: Transactions of IMACS, 24(2–3):153–173, June 19, 1997. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=24&
issue=2-3&aid=774. Volterra

centennial (Tempe, AZ, 1996).

Brugnano:1997:NMS

[BT97b] Luigi Brugnano and Donato Trigiante. A new mesh selection strategy for ODEs. Applied Numerical Mathematics: Transactions of IMACS, 24(1):1–21, June 9, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=24&
issue=1&aid=763.

Brugnano:1997:PSP

[BT97c] Luigi Brugnano and Donato Trigiante. On the potentiality of sequential and parallel codes based on extended trapezoidal rules (ETRs). Applied Numerical Mathematics: Transactions of IMACS, 25(2–3): 169–184, October 17, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=25&
issue=2-3&aid=812.

Butcher:1997:OCT

[BT97d] J. C. Butcher and S. Tracogna.

Order conditions for twostep Runge–Kutta methods. Applied Numerical Mathematics: Transactions of IMACS, 24(2–3): 351–364, June 19, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-

                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=24&
                                        issue=2-3&aid=788. Volterra

centennial (Tempe, AZ, 1996).

Brugnano:1998:PIB

[BT98] Luigi Brugnano and Donato Trigiante. Parallel implementation of block boundary value methods on nonlinear problems: theoretical results. Applied Numerical Mathematics: Transactions of IMACS, 28(2–4):127–141, October 1, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1998/28/
                                        2-4/906.pdf.

Benzi:1999:CSS

[BT99] Michele Benzi and Miroslav Tuma. A comparative study of sparse approximate inverse preconditioners. Applied Numerical Mathematics: Transactions of IMACS, 30(2–3):305– 340, June 10, 1999. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1999&volume=30&
                                        issue=2-3&aid=983.

Burrage:2000:PHB

[BT00] Kevin Burrage and Tianhai Tian. Parallel half-block methods for initial value problems. Applied Numerical Mathematics: Transactions of IMACS, 32(3): 255–271, March 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/59/30/
28/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/59/30/28/article.
pdf.

Benzi:2002:PSL

[BT02] Michele Benzi and Miroslav Tůma. A parallel solver for large-scale Markov chains. Applied Numerical Mathematics: Transactions of IMACS, 41(1):135–153, April 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.com/
gej-ng/10/10/28/86/27/
36/abstract.html.

Bilman:2019:BNM

[BT19] Deniz Bilman and Thomas Trogdon. Benchmarking numerical methods for lattice equations with the Toda lattice. Applied Numerical Mathematics: Transactions of IMACS, 141(??): 19–35, ???? 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418302356.

Baccouch:2019:OEE

[BTBR19] Mahboub Baccouch, Helmi Temimi, and Mohamed Ben-Romdhane. Optimal error estimates and superconvergence of an ultra weak discontinuous Galerkin method for fourth-order boundaryvalue problems. Applied Numerical Mathematics: Transactions of IMACS, 137(??):91–115, March 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418302605.

Baccouch:2020:DGM

[BTBR20] Mahboub Baccouch, Helmi Temimi, and Mohamed Ben-Romdhane. The discontinuous Galerkin method for stochastic differential equations driven by additive noises. Applied Numerical Mathematics: Transactions of IMACS, 152(??): 285–309, June 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419303290.

Boscheri:2023:AFH

[BTC23] Walter Boscheri, Maurizio Tavelli, and Cristóbal E. Castro. An all Froude high order IMEX scheme for the shallow water equations on unstructured Voronoi meshes. Applied Numerical Mathematics: Transactions of IMACS, 185(??): 311–335, March 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422003099.

Boursier:2010:PSM

[BTDV10] Isabelle Boursier, Damien Tromeur-Dervout, and Yuri Vassilevski. Parallel solution of Mixed Finite Element/Spectral Element systems for convection–diffusion equations on non-matching grids. Applied Numerical Mathematics: Transactions of IMACS, 60(11): 1131–1147, November 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Birken:2008:PUA

[BTMT08] Philipp Birken, Jurjen Duintjer Tebbens, Andreas Meister, and Miroslav Tůma. Preconditioner updates applied to CFD model problems. Applied Numerical Mathematics: Transactions of IMACS, 58(11): 1628–1641, November 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bey:1996:PAD

[BTP96] Kim S. Bey, J. Tinsley Oden, and Abani Patra. A parallel hp-adaptive discontinuous Galerkin method for hyperbolic conservation laws. Applied Numerical Mathematics: Transactions of IMACS, 20(4):321–336, June 3, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=4&aid=649. Adaptive

mesh refinement methods for CFD applications (Atlanta, GA, 1994).

Blom:1987:COE

[BtTBV87] J. G. Blom, J. H. M. ten Thije Boonkkamp, and J. G. Verwer. Corrigendum: “On the odd-even hopscotch scheme for the numerical integration of timedependent partial differential equations” [Appl. Numer. Math. 3 (1987), no. 1, 183–193] by J. H. M. ten Thije Boonkkamp and J. G. Verwer. Applied Numerical Mathematics: Transactions of IMACS, 3(4):361– 362, August 1987. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). See [tV87].

Buchholz:1999:SAA

[Buc99]  Peter Buchholz.  Struc-

tured analysis approaches for large Markov chains. Applied Numerical Mathematics: Transactions of IMACS, 31(4):375–404, December 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/46/20/
17/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/46/20/17/article.
pdf.

Bucchignani:2004:NSN

[Buc04] Edoardo Bucchignani. A numerical study of nonlinear dynamics in a tank for aerospace applications. Applied Numerical Mathematics: Transactions of IMACS, 49(3–4):307–318, June 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Buckwar:2006:OSA

[Buc06] Evelyn Buckwar. One-step approximations for stochastic functional differential equations. Applied Numerical Mathematics: Transactions of IMACS, 56(5): 667–681, May 2006. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Buccini:2017:RPN

[Buc17] Alessandro Buccini. Regularizing preconditioners by nonstationary iterated Tikhonov with general penalty term. Applied Numerical Mathematics: Transactions of IMACS, 116(??):64–81, June 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0168927416301350.

Bai:2023:EPM

[BUL23] Jiejing Bai, Hassan Ullah, and Chun Li. Energypreserving methods for nonsmooth nonlinear Schrödinger equations. Applied Numerical Mathematics: Transactions of IMACS, 185(??): 188–202, March 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742200304X.

Burrage:1985:OSP

[Bur85]  Kevin Burrage.  Order

and stability properties of explicit multivalue methods. Applied Numerical Mathematics: Transactions of IMACS, 1(5):363–379, September 1985. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Burrage:1991:EBG

[Bur91] Kevin Burrage. The error behaviour of a general class of predictor-corrector methods. Applied Numerical Mathematics: Transactions of IMACS, 8(3):201– 216, October 1991. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Burrage:1993:PMI

[Bur93a]  Kevin Burrage.  Parallel

methods for initial value problems. Applied Numerical Mathematics: Transactions of IMACS, 11(1–3): 5–25, January 1993. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Parallel methods for ordinary differential equations (Grado, 1991).

Burrage:1993:SHG

[Bur93b] Kevin Burrage. The search for the Holy Grail, or: Predictor-corrector methods for solving ODEIVPs. Applied Numerical Mathematics: Transactions of IMACS, 11(1–3):125–141, January 1993. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Parallel methods for ordinary differential equations (Grado, 1991).

Burg:2012:RBP

[Bür12] M. Bürg. A residual-based a posteriori error estimator for the hp-finite element method for Maxwell’s equations. Applied Numerical Mathematics: Transactions of IMACS, 62(8):922–940, August 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927412000505.

Burg:2013:CAH

[Bür13] M. Bürg. Convergence of an automatic hp-adaptive finite element strategy for Maxwell’s equations. Applied Numerical Mathematics: Transactions of IMACS, 72(??):188–204, October 2013. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927413000895.

Bustinza:2006:UAL

[Bus06] Rommel Bustinza. A unified analysis of the local discontinuous Galerkin method for a class of nonlinear problems. Applied Numerical Mathematics: Transactions of IMACS, 56(10–11):1293– 1306, October/November 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Butcher:1985:GLM

[But85] J. C. Butcher. General linear methods: a survey. Applied Numerical Mathematics: Transactions of IMACS, 1(4):273–284, July 1985. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Butcher:1992:ASD

[But92] J. C. Butcher. The adaptation of STRIDE to delay differential equations. Applied Numerical Mathematics: Transactions of IMACS, 9(3–5):415–425, April 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0168927492900318.

International Conference on the Numerical Solution of Volterra and Delay Equations (Tempe, AZ, 1990).

Butcher:1993:DIM

[But93] J. C. Butcher. Diagonallyimplicit multi-stage integration methods. Applied Numerical Mathematics: Transactions of IMACS, 11(5):347–363, March 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Butcher:1994:P

[But94] John Butcher. Preface. Applied Numerical Mathematics: Transactions of IMACS, 15(3):305, October 5, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=15&
issue=3&aid=528.

Butcher:1996:HRK

[But96] J. C. Butcher. A history of Runge–Kutta methods. Applied Numerical Mathematics: Transactions of IMACS, 20(3):247–260, May 22, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=3&aid=656. Selected

keynote papers presented at 14th IMACS World Congress (Atlanta, GA, 1994).

Butcher:1997:IAR

[But97] J. C. Butcher. An introduction to “almost Runge– Kutta” methods. Applied Numerical Mathematics: Transactions of IMACS, 24(2–3):331–342, June 19, 1997. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=24&
issue=2-3&aid=786. Volterra

centennial (Tempe, AZ, 1996).

Butcher:1998:OEO

[But98]  J. C. Butcher.  Order

and effective order. Applied Numerical Mathematics: Transactions of IMACS, 28(2–4):179–191, October 1, 1998. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1998/28/
2-4/909.pdf.

Butcher:2009:OSG

[But09]  J. C. Butcher.  Order

and stability of generalized Padé approximations. Applied Numerical Mathematics: Transactions of IMACS, 59(3–4):558–567, March/ April 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Blom:1994:VVA

[BV94] J. G. Blom and J. G. Verwer. VLUGR3: a vectorizable adaptive grid solver for PDEs in 3D. part I: Algorithmic aspects and applications. Applied Numerical Mathematics: Transactions of IMACS, 16(1–2): 129–156, December 1, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=16&
issue=1-2&aid=523.  A

Festschrift to honor Professor Robert Vichnevetsky on his 65th birthday.

Bellen:1996:SAC

[BV96] A. Bellen and R. Vermiglio. Some applications of continuous Runge–Kutta methods. Applied Numerical Mathematics: Transactions of IMACS, 22(1–3):63–80, December 4, 1996. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=22&
                                        issue=1-3&aid=708. Special

issue celebrating the centenary of Runge–Kutta methods.

Boykov:2009:AOM

[BVB09] I. V. Boykov, E. S. Ventsel, and A. I. Boykova. Accuracy optimal methods for evaluating hypersingular integrals. Applied Numerical Mathematics: Transactions of IMACS, 59(6): 1366–1385, June 2009. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Boykov:2010:ASH

[BVB10] I. V. Boykov, E. S. Ventsel, and A. I. Boykova. An approximate solution of hypersingular integral equations. Applied Numerical Mathematics: Transactions of IMACS, 60(6):607–628, June 2010. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Baumann:2019:ETL

[BvG19] Manuel Baumann and Martin B. van Gijzen. An efficient two-level preconditioner for multi-frequency wave propagation problems. Applied Numerical Mathematics: Transactions of IMACS, 135(??): 316–332, January 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741830196X.

Boykov:2014:ASN

[BVRB14] I. V. Boykov, E. S. Ventsel, V. A. Roudnev, and A. I. Boykova. An approximate solution of nonlinear hypersingular integral equations. Applied Numerical Mathematics: Transactions of IMACS, 86(??):1– 21, December 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414001287.

Becker:2014:DGM

[BVT14] Roland Becker, Guy Vallet, and Abdelaziz Taakili. A discontinuous Galerkin method for a model arising from stratigraphy. Applied Numerical Mathematics: Transactions of IMACS, 78(??):68–79, April 2014. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892741300161X.

Boonen:2009:AMM

[BVV09] Tim Boonen, Jan Van lent, and Stefan Vandewalle. An algebraic multigrid method for high order time-discretizations of the div-grad and the curl-curl equations. Applied Numerical Mathematics: Transactions of IMACS, 59(3–4): 507–521, March/April 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Baines:1986:MFE

[BW86] M. J. Baines and A. J. Wathen. Moving finite element modelling of compressible flow. Applied Numerical Mathematics: Transactions of IMACS, 2(6):495– 514, December 1986. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Berzins:1995:PCC

[BW95] M. Berzins and J. M. Ware. Positive cell-centered finite volume discretization methods for hyperbolic equations on irregular meshes. Applied Numerical Mathematics: Transactions of IMACS, 16(4):417–438, April 6, 1995.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=16&
issue=4&aid=555.

Berzins:1996:SCC

[BW96a] M. Berzins and J. M.

Ware. Solving convection and convection-reaction problems using the method of lines. Applied Numerical Mathematics: Transactions of IMACS, 20(1–2): 83–99, February 29, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=1-2&aid=666. Work-

shop on the method of lines for time-dependent problems (Lexington, KY, 1995).

Butcher:1996:RKM

[BW96b] J. C. Butcher and G. Wanner. Runge–Kutta methods: Some historical notes. Applied Numerical Mathematics: Transactions of IMACS, 22(1–3):113–151, December 4, 1996. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.

                                        cgi?year=1996&volume=22&
                                        issue=1-3&aid=730. Special

issue celebrating the centenary of Runge–Kutta methods.

Bey:1997:DNR

[BW97] Jürgen Bey and Gabriel Wittum. Downwind numbering: robust multigrid for convection-diffusion problems. Applied Numerical Mathematics: Transactions of IMACS, 23(1):177–192, March 7, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1997/23/

1/749.pdf. Multilevel methods (Oberwolfach, 1995).

Butcher:2003:TRE

[BW03] J. C. Butcher and W. M. Wright. A transformation relating explicit and diagonally-implicit general linear methods. Applied Numerical Mathematics: Transactions of IMACS, 44(3):313–327, February 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Butcher:2006:ADC

[BW06] J. C. Butcher and W. M. Wright. Applications of doubly companion matrices. Applied Numerical Mathematics: Transactions of IMACS, 56(3–4):358–373, March/April 2006. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Borghi:2015:CAS

[BW15] Riccardo Borghi and Ernst Joachim Weniger. Convergence analysis of the summation of the factorially divergent Euler series by Padé approximants and the delta transformation. Applied Numerical Mathematics: Transactions of IMACS, 94(??):149– 178, August 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415000562.

Bai:2021:RGB

[BW21] Zhong-Zhi Bai and Wen-Ting Wu. On refinement of the generalized Bendixson theorem. Applied Numerical Mathematics: Transactions of IMACS, 164(??): 125–138, ???? 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742030297X.

Bai:2023:CRK

[BW23a] Zhong-Zhi Bai and Lu Wang.

On convergence rates of Kaczmarz-type methods with different selection rules of working rows. Applied Numerical Mathematics: Transactions of IMACS, 186(??):289–319, April 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423000132.

Bai:2023:MSR

[BW23b] Zhong-Zhi Bai and Lu Wang. On multi-step randomized extended Kaczmarz method for solving large sparse inconsistent linear systems. Applied Numerical Mathematics: Transactions of IMACS, 192(??):197–213, October 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001733.

Burrage:1995:IGC

[BWEP95] K. Burrage, A. Williams, J. Erhel, and B. Pohl. The implementation of a Generalized Cross Validation algorithm using deflation techniques for linear systems. Applied Numerical Mathematics: Transactions of IMACS, 19(1–2): 17–31, December 4, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=1-2&aid=564. Mas-

sively parallel computing and applications (Amsterdam, 1993–1994).

Bai:2021:PMB

[BWM21] Zhong-Zhi Bai, Wen-Ting Wu, and Galina V. Muratova. The power method and beyond. Applied Numerical Mathematics: Transactions of IMACS, 164(??): 29–42, ???? 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742030088X.

Bourne:2021:CGC

[BWS21] Martin Bourne, Joab R.

Winkler, and Yi Su. The computation of the greatest common divisor of three bivariate Bernstein polynomials defined in a rectangular domain. Applied Numerical Mathematics: Transactions of IMACS, 166(??): 348–368, August 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001100.

Bai:2003:WCT

[BWY03] Zhong-Zhi Bai, Li Wang, and Jin-Yun Yuan. Weakconvergence theory of quasinonnegative splittings for singular matrices. Applied Numerical Mathematics: Transactions of IMACS, 47(2):75–89, November 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bourne:2017:CDA

[BWY17] Martin Bourne, Joab R. Winkler, and Su Yi. The computation of the degree of an approximate greatest common divisor of two Bernstein polynomials. Applied Numerical Mathematics: Transactions of IMACS, 111(??):17–35, January 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416301465.

Bar-Yoseph:2000:TFE

[BY00] Pinhas Z. Bar-Yoseph. Time finite element methods for initial value problems. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4):435–445, May 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/21/61/27/
                                        71/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/21/61/27/71/article.
                                        pdf.

Ben-yu:2001:GAA

[By01] Guo Ben-yu. Gegenbauer approximation and its applications to differential equations with rough asymptotic behaviors at infinity. Applied Numerical Mathematics: Transactions of IMACS, 38(4):403–425, September 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/77/34/28/abstract.
html.

Bai:2009:HBI

[BY09] Zhong-Zhi Bai and Xi Yang.

On HSS-based iteration methods for weakly nonlinear systems. Applied Numerical Mathematics: Transactions of IMACS, 59(12): 2923–2936, December 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Baeza:2022:NSB

[BY22] Antonio Baeza and Dionisio F. Yáñez. A note on some bounds between cubic spline interpolants depending on the boundary conditions: Application to a monotonicity property. Applied Numerical Mathematics: Transactions of IMACS, 181(??): 320–325, November 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001702.

Brezinski:1991:CEP

[BZ91] Claude Brezinski and Michela Redivo Zaglia. Construction of extrapolation processes. Applied Numerical Mathematics: Transactions of IMACS, 8(1):11–23, August 1991. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Bellen:1992:SCP

[BZ92] A. Bellen and M. Zennaro. Strong contractivity properties of numerical methods for ordinary and delay differential equations. Applied Numerical Mathematics: Transactions of IMACS, 9(3–5):321–346, April 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0168927492900259.

International Conference on the Numerical Solution of Volterra and Delay Equations (Tempe, AZ, 1990).

Bellen:1993:URK

[BZ93] Alfredo Bellen and Marino Zennaro. The use of Runge–Kutta formulae in waveform relaxation methods. Applied Numerical Mathematics: Transactions of IMACS, 11(1–3):95–114, January 1993. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Parallel methods for ordinary differential equations (Grado, 1991).

Brezinski:1994:EM

[BZ94a] C. Brezinski and M. Redivo Zaglia. Extrapolation methods. Applied Numerical Mathematics: Transactions of IMACS, 15(2):123– 131, September 13, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=15&
issue=2&aid=491. Innova-

tive methods in numerical analysis (Bressanone, 1992).

Brezinski:1994:KST

[BZ94b] C. Brezinski and M. Redivo Zaglia. On the kernel of sequence transformations. Applied Numerical Mathematics: Transactions of IMACS, 16(1–2): 239–244, December 1, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=16&
issue=1-2&aid=530.  A

Festschrift to honor Professor Robert Vichnevetsky on his 65th birthday.

Brezinski:1996:VMS

[BZ96] C. Brezinski and M. Redivo Zaglia. Vector and matrix sequence transformations based on biorthogonality. Applied Numerical Mathematics: Transactions of IMACS, 21(4):353– 373, November 22, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=21&
issue=4&aid=701.

Bhrawy:2017:ICM

[BZ17a] A. H. Bhrawy and M. A. Zaky. An improved collocation method for multidimensional space–time variableorder fractional Schrödinger equations. Applied Numerical Mathematics: Transactions of IMACS, 111(??): 197–218, January 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301763.

Brown:2017:DEH

[BZ17b] David A. Brown and David W. Zingg. Design and evaluation of homotopies for efficient and robust continuation. Applied Numerical Mathematics: Transactions of IMACS, 118(??): 150–181, August 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417300594.

Copetti:2015:ACP

[CA15] M. I. M. Copetti and M. Aouadi. Analysis of a contact problem in thermoviscoelasticity under the Green–Lindsay model. Applied Numerical Mathematics: Transactions of IMACS, 91(??):60–74, May 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415000033.

Copetti:2016:QSC

[CA16] M. I. M. Copetti and M. Aouadi. A quasistatic contact problem in thermoviscoelastic diffusion theory. Applied Numerical Mathematics: Transactions of IMACS, 109(??): 157–183, November 2016. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301052.

Cai:2021:FSC

[CA21] Haotao Cai and Qiguang An.

A fractional spectral collocation method for general Caputo two-point boundary value problems. Applied Numerical Mathematics: Transactions of IMACS, 163(??):43–56, May 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421000088.

Chui:1996:SOM

[CaAL96] Charles K. Chui and Jian ao A. Lian. A study of orthonormal multi-wavelets. Applied Numerical Mathematics: Transactions of IMACS, 20(3):273–298, May 22, 1996. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=20&
                                        issue=3&aid=659. Selected

keynote papers presented at 14th IMACS World Congress (Atlanta, GA, 1994).

Chniti:2016:DNS

[CAAT16] Chokri Chniti, Sharefa Eisa Ali Alhazmi, Sami H. Altoum, and Moncef Toujani. DtN and NtD surface radiation conditions for twodimensional acoustic scattering: Formal derivation and numerical validation. Applied Numerical Mathematics: Transactions of IMACS, 101(??):53–70, March 2016. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927415001701.

Castro:2003:ISU

[CAD03] Liliana R. Castro, Osvaldo E. Agamennoni, and Carlos E. D’Attellis. Identification structures using rational wavelets: Examples of application. Applied Numerical Mathematics: Transactions of IMACS, 47(3–4): 345–363, December 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cahlon:1989:NMI

[Cah89] Baruch Cahlon. Numerical methods of initial value problems. Applied Numerical Mathematics: Transactions of IMACS, 5(5): 399–407, September 1989. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cahlon:1992:NSF

[Cah92] Baruch Cahlon. Numerical solutions for functional integral equations with statedependent delay. Applied Numerical Mathematics: Transactions of IMACS, 9(3–5):291–305, April 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0168927492900237.

International Conference on the Numerical Solution of Volterra and Delay Equations (Tempe, AZ, 1990).

Cai:2009:FSH

[Cai09] Haotao Cai. A fast solver for a hypersingular boundary integral equation. Applied Numerical Mathematics: Transactions of IMACS, 59(8):1960–1969, August 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cai:2015:ASP

[Cai15]  Mingchao Cai.  Analysis

of some projection method based preconditioners for models of incompressible flow. Applied Numerical Mathematics: Transactions of IMACS, 90(??): 77–90, April 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414001962.

Cai:2024:SGM

[Cai24]  Haotao Cai.  Spectral-

Galerkin method for second kind VIEs with highly oscillatory kernels of the stationary point. Applied Numerical Mathematics: Transactions of IMACS, 201(??): 102–117, July 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424000461.

Cameron:1999:CLO

[Cam99] Frank Cameron. A class of low order DIRK methods for a class of DAEs.

Applied Numerical Mathematics: Transactions of IMACS, 31(1):1–16, September 1, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1999&volume=31&
issue=1&aid=988.

Cao:1997:NCB

[Cao97] Zhi-Hao Cao. A note on the convergence behavior of GMRES. Applied Numerical Mathematics: Transactions of IMACS, 25(1): 13–20, September 16, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=25&
issue=1&aid=795.

Cao:1998:CTE

[Cao98a] Zhi-Hao Cao. A convergence theorem on an extrapolated iterative method and its applications. Applied Numerical Mathematics: Transactions of IMACS, 27(3):203–209, 1998. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cao:1998:QAI

[Cao98b] Zhi-Hao Cao. On the QMR approach for iterative methods including coupled threeterm recurrences for solving nonsymmetric linear systems. Applied Numerical Mathematics: Transactions of IMACS, 27(2):123–140, June 10, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1998/27/
2/871.pdf.

Cao:2001:CNI

[Cao01] Zhi-Hao Cao. On the convergence of nonstationary iterative methods for symmetric positive (semi)definite systems. Applied Numerical Mathematics: Transactions of IMACS, 37(3): 319–330, May 2001. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/73/31/
30/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/73/31/30/article.
pdf.

Cao:2003:FUA

[Cao03] Zhi-Hao Cao. Fast Uzawa algorithm for generalized saddle point problems. Applied Numerical Mathematics: Transactions of IMACS, 46(2):157–171, August 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cao:2007:PSB

[Cao07] Zhi-Hao Cao. Positive stable block triangular preconditioners for symmetric saddle point problems. Applied Numerical Mathematics: Transactions of IMACS, 57(8):899–910, August 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cao:2009:CSC

[Cao09] Zhi-Hao Cao. Constraint Schur complement preconditioners for nonsymmetric saddle point problems. Applied Numerical Mathematics: Transactions of IMACS, 59(1):151–169, January 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cao:2010:BTS

[Cao10] Zhi-Hao Cao. Block triangular Schur complement preconditioners for saddle point problems and application to the Oseen equations. Applied Numerical Mathematics: Transactions of IMACS, 60(3):193–207, March 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Carniel:1994:QCM

[Car94] Roberto Carniel. A quasicell mapping approach to the global dynamical analysis of Newton’s root-finding algorithm. Applied Numerical Mathematics: Transactions of IMACS, 15(2):133– 152, September 13, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=15&
issue=2&aid=492. Innova-

tive methods in numerical analysis (Bressanone, 1992).

Carstensen:2009:CAF

[Car09a] Carsten Carstensen. Convergence of adaptive finite element methods in computational mechanics. Applied Numerical Mathematics: Transactions of IMACS, 59(9):2119–2130, September 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cartis:2009:SDM

[Car09b] Coralia Cartis. Some disadvantages of a Mehrotratype primal-dual corrector interior point algorithm for linear programming. Applied Numerical Mathematics: Transactions of IMACS, 59(5):1110–1119, May 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Carillo:2019:KTE

[Car19]  Sandra Carillo.  KdV-

type equations linked via Bäcklund transformations: Remarks and perspectives. Applied Numerical Mathematics: Transactions of IMACS, 141(??):81–90, ???? 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0168927418302277.

Cardone:2023:SAE

[Car23] Angelamaria Cardone. Stability analysis of ef Gaussian direct quadrature methods for Volterra integral equations. Applied Numerical Mathematics: Transactions of IMACS, 186(??): 241–251, April 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423000181.

Cash:1996:RKM

[Cas96]  J. R. Cash.  Runge–

Kutta methods for the solution of stiff two-point boundary value problems. Applied Numerical Mathematics: Transactions of IMACS, 22(1–3):165–177, December 4, 1996. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=22&

                                        issue=1-3&aid=712. Special

issue celebrating the centenary of Runge–Kutta methods.

Castillo:2006:RLD

[Cas06] Paul Castillo. A review of the Local Discontinuous Galerkin (LDG) method applied to elliptic problems. Applied Numerical Mathematics: Transactions of IMACS, 56(10–11):1307– 1313, October/November 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cata:2010:PAP

[Cat10] Oscar Catà. Padé approximants and the prediction of non-perturbative parameters in particle physics. Applied Numerical Mathematics: Transactions of IMACS, 60(12):1273–1285, December 2010. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Caussignac:2008:GFE

[Cau08] Ph. Caussignac. A gauge finite element method for the 2D Navier–Stokes problem. Applied Numerical Mathematics: Transactions of IMACS, 58(10):1413–1433, October 2008. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cisneros:1999:HPC

[CB99] Gerardo Cisneros and Jeff P.

Brooks. High-performance computing at Silicon Graphics/Cray Research. Applied Numerical Mathematics: Transactions of IMACS, 30(1):125–135, May 10, 1999. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1999/30/
1/954.pdf.

Christodoulidi:2016:NIV

[CBD16] Helen Christodoulidi, Tassos Bountis, and Lambros Drossos. Numerical integration of variational equations for Hamiltonian systems with long range interactions. Applied Numerical Mathematics: Transactions of IMACS, 104(??): 158–165, June 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415001312.

Cook:2019:EES

[CBHM19] Stephen Cook, Chris Budd, Adrian Hill, and Thomas Melvin. Error estimates for semi-Lagrangian finite difference methods applied to Burgers’ equation in one dimension. Applied Numerical Mathematics: Transactions of IMACS, 145(??):

261–282, November 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419301643.

Cuyt:2011:SNG

[CBHY11] Annie Cuyt, Brahim Benouahmane, Hamsapriye, and Irem Yaman. Symbolic– numeric Gaussian cubature rules. Applied Numerical Mathematics: Transactions of IMACS, 61(8):929– 945, August 2011. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Campbell:1990:CBA

[CC90] Stephen L. Campbell and Kenneth D. Clark. Convergence of BDF approximations for nonsolvable differential algebraic equations. Applied Numerical Mathematics: Transactions of IMACS, 6(3):153– 158, March 1990. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Charina:2004:CMN

[CC04a] Maria Charina and Costanza Conti. Convergence of multivariate non-stationary vector subdivision schemes. Applied Numerical Mathematics: Transactions of IMACS, 49(3–4):343–354, June 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Chen:2004:MFE

[CC04b] Jinru Chen and Junzhi Cui.

A multiscale finite element method for elliptic problems with highly oscillatory coefficients. Applied Numerical Mathematics: Transactions of IMACS, 50(1):1–13, July 2004. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cheng:2008:EEF

[CC08] Rongjun Cheng and Yumin Cheng. Error estimates for the finite point method. Applied Numerical Mathematics: Transactions of IMACS, 58(6):884–898, June 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Califano:2018:OSW

[CC18] Giovanna Califano and Dajana Conte. Optimal Schwarz waveform relaxation for fractional diffusion-wave equations. Applied Numerical Mathematics: Transactions of IMACS, 127(??): 125–141, May 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418300114.

Chen:2019:CHO

[CC19] Juan Chen and Fangqi Chen. Convergence of a highorder compact finite difference scheme for the Klein– Gordon–Schrödinger equations. Applied Numerical Mathematics: Transactions of IMACS, 143(??): 133–145, September 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419300571.

Chen:2020:UCC

[CC20a] Juan Chen and Fangqi Chen. Unconditional L_{∞} convergence of a compact ADI scheme for coupled nonlinear Schrödinger system. Applied Numerical Mathematics: Transactions of IMACS, 153(??):430–442, July 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300738.

Chistyakova:2020:SDA

[CC20b] E. V. Chistyakova and V. F. Chistyakov. Solution of differential algebraic equations with the Fredholm operator by the least squares method. Applied Numerical Mathematics: Transactions of IMACS, 149(??): 43–51, ???? 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741930100X.

Cai:2023:EDP

[CC23a] Jiaxiang Cai and Juan Chen. Efficient dissipationpreserving approaches for the damped nonlinear Schrödinger equation. Applied Numerical Mathematics: Transactions of IMACS, 183(??): 173–185, January 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422002264.

Chen:2023:BXP

[CC23b] Yuwei Chen and Christina C.

Christara. Bilateral XVA pricing under stochastic default intensity: PDE modelling and computation. Applied Numerical Mathematics: Transactions of IMACS, 185(??): 236–259, March 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422003014.

Caliari:2024:ESC

[CC24] Marco Caliari and Fabio Cassini. Efficient simulation of complex Ginzburg– Landau equations using high-order exponential-type methods. Applied Numerical Mathematics: Transactions of IMACS, 206(??): 340–357, December 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424002058.

Carillo:2025:SNT

[CC25] Sandra Carillo and Michel Chipot. On some nonlocal in time and space parabolic problem. Applied Numerical Mathematics: Transactions of IMACS, 208 (part B) (??):314–322, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424002976.

Cantero:2008:MAC

[CCBGV08] María José Cantero, Ruymán Cruz-Barroso, and Pablo González-Vera. A matrix approach to the computation of quadrature formulas on the unit circle. Applied Numerical Mathematics: Transactions of IMACS, 58(3):296–318, March 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Chang:2008:EOC

[CCC08] J. C. Chang, T. M. H. Chan, and D. J. L. Chen. Enhanced order composition methods. Applied Numerical Mathematics: Transactions of IMACS, 58(3):

223–235, March 2008. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Caballero-Cardenas:2025:SIF

[CCCC^{+}25] C. Caballero-Cárdenas, M. J.

Castro, C. Chalons, T. Morales de Luna, and M. L. Muñoz-Ruiz. Semi-implicit fully exactly well-balanced schemes for the two-layer shallow water system. Applied Numerical Mathematics: Transactions of IMACS, 218(??): 128–147, December 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425001552.

Campagna:2020:SMP

[CCD^{+}20] R. Campagna, S. Cuomo, S. De Marchi, E. Perracchione, and G. Severino. A stable meshfree PDE solver for source-type flows in porous media. Applied Numerical Mathematics: Transactions of IMACS, 149(??):30–42, ???? 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741930217X.

Calatayud:2020:CRA

[CCDJ20] J. Calatayud, J.-C. Cortés, J. A. Díaz, and M. Jornet. Constructing reliable approximations of the probability density function to the random heat PDE via a finite difference scheme. Applied Numerical Mathematics: Transactions of IMACS, 151(??):413–424, May 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300131.

Cayama:2020:NAF

[CCdlH20] Jorge Cayama, Carlota M. Cuesta, and Francisco de la Hoz. Numerical approximation of the fractional Laplacian on R using orthogonal families. Applied Numerical Mathematics: Transactions of IMACS, 158(??): 164–193, December 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302245.

Caplan:2013:NSE

[CCG13] R. M. Caplan and R. Carretero-González. Numerical stability of explicit Runge–Kutta finite-difference schemes for the nonlinear Schrödinger equation. Applied Numerical Mathematics: Transactions of IMACS, 71(??):24– 40, September 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413000627.

Claeyssen:1999:DAS

[CCJ99] Julio Ruiz Claeyssen, Germán Canahualpa, and Claudio Jung. A direct approach to second-order matrix nonclassical vibrating equations. Applied Numerical Mathematics: Transactions of IMACS, 30(1):65–78, May 10, 1999. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1999/30/
1/950.pdf.

Chen-Charpentier:2003:NSD

[CCK03] Benito M. Chen-Charpentier and Hristo V. Kojouharov. Numerical simulation of dual-species biofilms in porous media. Applied Numerical Mathematics: Transactions of IMACS, 47(3– 4):377–389, December 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Choo:2008:DGM

[CCK08] S. M. Choo, S. K. Chung, and K. I. Kim. A discontinuous Galerkin method for the Rosenau equation. Applied Numerical Mathematics: Transactions of IMACS, 58(6):783–799, June 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Choo:2004:CDS

[CCL04] S. M. Choo, S. K. Chung, and Y. J. Lee. A conservative difference scheme for the viscous Cahn–Hilliard equation with a nonconstant gradient energy coefficient. Applied Numerical Mathematics: Transactions of IMACS, 51(2– 3):207–219, November 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cai:2022:LSR

[CCL22] Zhiqiang Cai, Jingshuang Chen, and Min Liu. Leastsquares ReLU neural network (LSNN) method for scalar nonlinear hyperbolic conservation law. Applied Numerical Mathematics: Transactions of IMACS, 174(??):163–176, April 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422000022.

Chen:2010:LSF

[CCLT10] T. F. Chen, C. L. Cox, H. C. Lee, and K. L. Tung. Leastsquares finite element methods for generalized Newtonian and viscoelastic flows. Applied Numerical Mathematics: Transactions of IMACS, 60(10):1024–1040, October 2010. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Chan:2002:PPR

[CCM02] R. P. K. Chan, P. Chartier, and A. Murua. Postprojected Runge–Kutta methods for index-2 differential– algebraic equations. Applied Numerical Mathematics: Transactions of IMACS, 42(1–3):77–94, August 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Costabile:2017:PLD

[CCM17] Pierfranco Costabile, Carmelina Costanzo, and Francesco Macchione. Performances and limitations of the diffusive approximation of the 2-d shallow water equations for flood simulation in urban and rural areas. Applied Numerical Mathematics: Transactions of IMACS, 116(??):141–156, June 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301106.

Calvo:2011:NSA

[CCMSS11] M. P. Calvo, Ph. Chartier, A. Murua, and J. M. Sanz-Serna. Numerical stroboscopic averaging for ODEs and DAEs. Applied Numerical Mathematics: Transactions of IMACS, 61(10): 1077–1095, October 2011.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Camano:2022:PEA

[CCOVF22] Jessika Camaño, Sergio Caucao, Ricardo Oyarzúa, and Segundo Villa-Fuentes. A posteriori error analysis of a momentum conservative Banach spaces based mixed-FEM for the Navier–Stokes problem. Applied Numerical Mathematics: Transactions of IMACS, 176(??): 134–158, June 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422000502.

Cerimele:2004:CID

[CCP04] M. M. Cerimele, M. L. Chiofalo, and F. Pistella. From coherent to incoherent dynamical behaviour of quantum atomic gases in periodic potentials. Applied Numerical Mathematics: Transactions of IMACS, 49(3–4):319–330, June 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Capobianco:2017:CIT

[CCP17] Giovanni Capobianco, Dajana Conte, and Beatrice Paternoster. Construction and implementation of twostep continuous methods for Volterra integral equations. Applied Numerical Mathematics: Transactions of IMACS, 119(??): 239–247, September 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417300533.

Cao:2023:ASP

[CCQ^{+}23] Rongjun Cao, Minghua Chen, Yingfan Qi, Jiankang Shi, and Xiaobo Yin. Analysis of (shifted) piecewise quadratic polynomial collocation for nonlocal diffusion model. Applied Numerical Mathematics: Transactions of IMACS, 185(??): 120–140, March 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422003038.

Claeyssen:2002:IRS

[CCS02] Julio R. Claeyssen, Leonardo D.

Chiwiacowsky, and German C. Suazo. The impulse response in the symbolic computing of modes for beams and plates. Applied Numerical Mathematics: Transactions of IMACS, 40(1–2):119–135, January 2002. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/

                                        10/28/85/27/37/abstract.
                                        html.

Chaudhry:2017:PEE

[CCS17a] Jehanzeb H. Chaudhry, J. B. Collins, and John N. Shadid. A posteriori error estimation for multi-stage Runge– Kutta IMEX schemes. Applied Numerical Mathematics: Transactions of IMACS, 117(??):36–49, July 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417300363.

Conti:2017:CLD

[CCS17b] Costanza Conti, Mariantonia Cotronei, and Tomas Sauer. Convergence of leveldependent Hermite subdivision schemes. Applied Numerical Mathematics: Transactions of IMACS, 116(??):119–128, June 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417300569.

Chen:2022:LEE

[CCST22] Hu Chen, Mengyi Chen, Tao Sun, and Yifa Tang. Local error estimate of L1 scheme for linearized time fractional KdV equation with weakly singular solutions. Applied Numerical Mathematics: Transactions of IMACS, 179(??):

183–190, September 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001143.

Cheng:2022:TNA

[CCY22] Wanyou Cheng, Hongsheng Chen, and Jinyun Yuan. A truncated Newton algorithm for nonconvex sparse recovery. Applied Numerical Mathematics: Transactions of IMACS, 178(??): 356–371, August 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742200099X.

Chu:2022:SCS

[CCZ22] Xiaochen Chu, Chuanjun Chen, and Tong Zhang. Stability and convergence of spatial discrete stabilized finite volume method for the unsteady incompressible magnetohydrodynamics equations. Applied Numerical Mathematics: Transactions of IMACS, 181(??): 436–467, November 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001611.

Chai:2018:CFE

[CCZZ18] Shimin Chai, Yanzhao Cao, Yongkui Zou, and Wenju Zhao. Conforming finite element methods for the stochastic Cahn–Hilliard-Cook equation. Applied Numerical Mathematics: Transactions of IMACS, 124(??): 44–56, February 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417302040.

Cash:1995:SIN

[CD95] J. Cash and L. Dieci. Special issue on numerical methods for ordinary differential equations. Applied Numerical Mathematics: Transactions of IMACS, 17(3):189– 346, 1995. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Costa:2000:CHO

[CD00] Bruno Costa and Wai Sun Don. On the computation of high order pseudospectral derivatives. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4):151–159, May 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/61/27/
39/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/61/27/39/article.
pdf.

Costabile:2005:LAT

[CD05] F. A. Costabile and F. Dell’Accio.

Lidstone approximation on the triangle. Applied Numerical Mathematics: Transactions of IMACS, 52(4): 339–361, March 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cimolin:2013:NSF

[CD13] F. Cimolin and M. Discacciati. Navier–Stokes/ Forchheimer models for filtration through porous media. Applied Numerical Mathematics: Transactions of IMACS, 72(??):205– 224, October 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413000913.

Celik:2017:FEM

[ÇD17] Cem Çelik and Melda Duman. Finite element method for a symmetric tempered fractional diffusion equation. Applied Numerical Mathematics: Transactions of IMACS, 120(??): 270–286, October 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417301368.

Cipolla:2018:FPC

[CD18] Stefano Cipolla and Fabio Durastante. Fractional PDE constrained optimization: an optimize-then-discretize approach with L-BFGS and approximate inverse preconditioning. Applied Numerical Mathematics: Transactions of IMACS, 123(??): 43–57, January 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417301824.

Citro:2020:LTA

[CD20a] Vincenzo Citro and Raffaele D’Ambrosio. Longterm analysis of stochastic θ-methods for damped stochastic oscillators. Applied Numerical Mathematics: Transactions of IMACS, 150(??):18–26, April 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419302132.

Citro:2020:NCM

[CD20b] Vincenzo Citro and Raffaele D’Ambrosio. Nearly conservative multivalue methods with extended bounded parasitism. Applied Numerical Mathematics: Transactions of IMACS, 152(??): 221–230, June 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419303393.

Caucao:2023:MFC

[CD23] Sergio Caucao and Marco Discacciati. A mixed FEM for the coupled Brinkman– Forchheimer/Darcy problem. Applied Numerical Mathematics: Transactions of IMACS, 190(??): 138–154, August 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001071.

Chalhoub:2003:MRB

[CdCV03] Ezzat S. Chalhoub and Haroldo F. de Campos Velho. Multispectral reconstruction of bioluminescence term in natural waters. Applied Numerical Mathematics: Transactions of IMACS, 47(3–4):365–376, December 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Capuzzo-Dolcetta:2000:CCI

[CDD00] Roberto Capuzzo-Dolcetta and Roberto Di Lisio. A criterion for the choice of the interpolation kernel in smoothed particle hydrodynamics. Applied Numerical Mathematics: Transactions of IMACS, 34(4):363– 371, August 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/62/32/
29/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/62/32/29/article.
pdf.

Cavoretto:2017:PUI

[CDD^{+}17] R. Cavoretto, S. De Marchi, A. De Rossi, E. Perracchione, and G. Santin. Partition of unity interpolation using stable kernel-based techniques. Applied Numerical Mathematics: Transactions of IMACS, 116(??): 95–107, June 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301313.

Calvo:2001:LIR

[CdFN01] M. P. Calvo, J. de Frutos, and J. Novo. Linearly implicit Runge–Kutta methods for advection-reaction-diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 37(4):535–549, June 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/73/33/
33/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/73/33/33/article.
pdf.

Claeys:2019:IMT

[CDG19] X. Claeys, V. Dolean, and M. J. Gander. An introduction to multi-trace formulations and associated domain decomposition solvers. Applied Numerical Mathematics: Transactions of IMACS, 135(??):69–86, January 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418301545.

Carson:2017:AOB

[CDGA17] Hugh A. Carson, David L.

Darmofal, Marshall C. Galbraith, and Steven R. Allmaras. Analysis of output-based error estimation for finite element methods. Applied Numerical Mathematics: Transactions of IMACS, 118(??): 182–202, August 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417300648.

Chaudhary:2024:SSF

[CDH24] Renu Chaudhary, Kai Diethelm, and Safoura Hashemishahraki. On the separation of solutions to fractional differential equations of order α ∈ (1, 2). Applied Numerical Mathematics: Transactions of IMACS, 203(??):84– 96, September 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424001260.

Censor:1991:OBE

[CDI91] Yair Censor, Alvaro R. De Pierro, and Alfredo N. Iusem. Optimization of Burg’s entropy over linear constraints. Applied Numerical Mathematics: Transactions of IMACS, 7(2):151– 165, February 1991. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cuomo:2024:SNR

[CDI^{+}24] Salvatore Cuomo, Mariapia De Rosa, Stefano Izzo, Francesco Piccialli, and Monica Pragliola. Speckle noise removal via learned variational models. Applied Numerical Mathematics: Transactions of IMACS, 200(??):162–178, June 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742300168X.

Clausel:2006:MAR

[CDJT06] M. Clausel, M. Duruflé, P. Joly, and S. Tordeux. A mathematical analysis of the resonance of the finite thin slots. Applied Numerical Mathematics: Transactions of IMACS, 56(10–11):1432– 1449, October/November 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Conte:2012:TSD

[CDP12] Dajana Conte, Raffaele D’Ambrosio, and Beatrice Paternoster. Twostep diagonally-implicit collocation based methods for Volterra Integral Equations. Applied Numerical Mathematics: Transactions of IMACS, 62(10): 1312–1324, October 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412000931.

Cardone:2017:HOE

[CDP17] A. Cardone, R. D’Ambrosio, and B. Paternoster. High order exponentially fitted methods for Volterra integral equations with periodic solution. Applied Numerical Mathematics: Transactions of IMACS, 114(??): 18–29, April 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741630071X.

Cardone:2019:SMS

[CDP19] Angelamaria Cardone, Raffaele D’Ambrosio, and Beatrice Paternoster. A spectral method for stochastic fractional differential equations. Applied Numerical Mathematics: Transactions of IMACS, 139(??): 115–119, May 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419300200.

Cuomo:2025:NAS

[CDP^{+}25] Salvatore Cuomo, Mariapia De Rosa, Francesco Piccialli, Laura Pompameo, and Vincenzo Vocca. A numerical approach for soil microbiota growth prediction through physics-informed neural network. Applied Numerical Mathematics: Transactions of IMACS, 207(??):97– 110, January 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424002290.

Cohen:2020:HON

[CDR20] David Cohen, Kristian Debrabant, and Andreas Rößler. High order numerical integrators for single integrand Stratonovich SDEs. Applied Numerical Mathematics: Transactions of IMACS, 158(??):264–270, December 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302270.

Charina:2019:ABS

[CDRT19] M. Charina, M. Donatelli, L. Romani, and V. Turati. Anisotropic bivariate subdivision with applications to multigrid. Applied Numerical Mathematics: Transactions of IMACS, 135(??): 333–366, January 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418302046.

Calvo:2000:NCI

[CDV00] N. Calvo, J. Durany, and C. Vázquez. Numerical computation of ice sheet profiles with free boundary models. Applied Numerical Mathematics: Transactions of IMACS, 35(2):111– 128, October 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/64/30/
28/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/64/30/28/article.
pdf.

Chen:2013:SCA

[CDW13] Minghua Chen, Weihua Deng, and Yujiang Wu. Superlinearly convergent algorithms for the twodimensional space–time Caputo– Riesz fractional diffusion equation. Applied Numerical Mathematics: Transactions of IMACS, 70(??): 22–41, August 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927413000512.

Chen:2019:DBI

[CDW19] Min-Hong Chen, Wei Dou, and Qing-Biao Wu. DPMHSSbased iteration methods for solving weakly nonlinear systems with complex coefficient matrices. Applied Numerical Mathematics: Transactions of IMACS, 146(??):328–341, December 2019. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419301904.

Chen:2023:PEE

[CDW23] Luoping Chen, Jiajia Dai, and Yiming Wen. A posteriori error estimates of twogrid weak Galerkin methods for semilinear elliptic differential equations. Applied Numerical Mathematics: Transactions of IMACS, 187(??):277–294, May 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423000612.

Charlier:2025:NSS

[CEL25] Christophe Charlier, Daniel Eriksson, and Jonatan Lenells. Numerical scheme for the solution of the “bad” Boussinesq equation. Applied Numerical Mathematics: Transactions of IMACS, 217(??): 216–233, November 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742500128X.

Cerna:2025:OWM

[Cer25] Dana Cerná. Orthogonal wavelet method for multistage expansion and contraction options under stochastic volatility. Applied Numerical Mathematics: Transactions of IMACS, 212(??): 155–175, June 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742500025X.

Chan:1991:DDI

[CES91] Tony F. Chan, Wei Nan E, and Jia Chang Sun. Domain decomposition interface preconditioners for fourth-order elliptic problems. Applied Numerical Mathematics: Transactions of IMACS, 8(4–5): 317–331, November 1991. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Chung:2000:MGG

[CEW00] K. C. Chung, G. A. Evans, and J. R. Webster. A method to generate generalized quadrature rules for oscillatory integrals. Applied Numerical Mathematics: Transactions of IMACS, 34(1):85–93, June 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/21/62/27/
                                        31/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/21/62/27/31/article.
                                        pdf.

Chen:1986:LSF

[CF86] T. F. Chen and G. J. Fix. Least squares finite element simulation of transonic flows. Applied Numerical Mathematics: Transactions of IMACS, 2(3–5):399– 408, October 1986. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Copetti:2005:NAE

[CF05] M. I. M. Copetti and D. A. French. Numerical approximation and error control for a thermoelastic contact problem. Applied Numerical Mathematics: Transactions of IMACS, 55(4): 439–457, December 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Clift:2008:NST

[CF08] Simon S. Clift and Peter A.

Forsyth. Numerical solution of two asset jump diffusion models for option valuation. Applied Numerical Mathematics: Transactions of IMACS, 58(6): 743–782, June 2008. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Carlini:2013:SLA

[CF13a] Elisabetta Carlini and Roberto Ferretti. A Semi-Lagrangian approximation for the AMSS model of image processing. Applied Numerical Mathematics: Transactions of IMACS, 73(??):16– 32, November 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412001328.

Chiang:2013:RBS

[CF13b] Chun-Yueh Chiang and Hung-Yuan Fan. Residual bounds of the stochastic algebraic Riccati equation. Applied Numerical Mathematics: Transactions of IMACS, 63(1):78–87, January 2013. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412001675.

Copetti:2013:DCP

[CF13c] M. I. M. Copetti and J. R. Fernández. A dynamic contact problem involving a Timoshenko beam model. Applied Numerical Mathematics: Transactions of IMACS, 63(1):117– 128, January 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412001729.

Copetti:2014:DCP

[CF14] M. I. M. Copetti and J. R. Fernández. A dynamic contact problem in thermoviscoelasticity with two temperatures. Applied Numerical Mathematics: Transactions of IMACS, 77(??): 55–71, March 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413001529.

Colbrook:2018:CEE

[CF18] Matthew J. Colbrook and Athanasisos S. Fokas. Computing eigenvalues and eigenfunctions of the Laplacian for convex polygons. Applied Numerical Mathematics: Transactions of IMACS, 126(??):1–17, April 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417302490.

Claeyssen:2003:DFR

[CFC03] Julio R. Claeyssen, Ines Moraes Ferreira, and Rosemaira Dalcin Copetti. Decomposition of forced responses in vibrating systems. Applied Numerical Mathematics: Transactions of IMACS, 47(3–4): 391–405, December 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Caiazzo:2009:AAC

[CFCH09] Alfonso Caiazzo, Jean-Luc Falcone, Bastien Chopard, and Alfons G. Hoekstra. Asymptotic analysis of complex automata models for reaction–diffusion systems. Applied Numerical Mathematics: Transactions of IMACS, 59(8):2023–2034, August 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Campo:2007:QED

[CFKS07] M. Campo, J. R. Fernández, K. L. Kuttler, and M. Shillor. Quasistatic evolution of damage in an elastic body: numerical analysis and computational experiments. Applied Numerical Mathematics: Transactions of IMACS, 57(9):975–988, September 2007. CODEN ANMAEL.

ISSN 0168-9274 (print), 1873-5460 (electronic).

Coclite:2020:NMN

[CFL^{+}20] G. M. Coclite, A. Fanizzi, L. Lopez, F. Maddalena, and S. F. Pellegrino. Numerical methods for the nonlocal wave equation of the peridynamics. Applied Numerical Mathematics: Transactions of IMACS, 155(??): 119–139, ???? 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418302563.

Cai:2022:FOC

[CFLW22] Yongyong Cai, Jinxue Fu, Jianfeng Liu, and Tingchun Wang. A fourth-order compact finite difference scheme for the quantum Zakharov system that perfectly inherits both mass and energy conservation. Applied Numerical Mathematics: Transactions of IMACS, 178(??):1–24, August 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742200071X.

Conte:2024:NSD

[CFM^{+}24] Dajana Conte, Eslam Farsimadan, Leila Moradi, Francesco Palmieri, and Beatrice Paternoster. Numerical solution of delay Volterra functional integral equations with variable bounds. Applied Numerical Mathematics: Transactions of IMACS, 200(??): 148–161, June 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001216.

Campo:2008:QCP

[CFRA08] M. Campo, J. R. Fernández, and Á. Rodríguez-Arós. A quasistatic contact problem with normal compliance and damage involving viscoelastic materials with long memory. Applied Numerical Mathematics: Transactions of IMACS, 58(9): 1274–1290, September 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Clark:1994:I

[CFS94] K. Clark, J. E. Flaherty, and M. S. Shephard. Introduction. Applied Numerical Mathematics: Transactions of IMACS, 14(1–3): 1–3, April 14, 1994. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=14&
issue=1-3&aid=450.

Camilli:2013:ASH

[CFS13] Fabio Camilli, Adriano Festa, and Dirk Schieborn. An approximation scheme for a Hamilton–Jacobi equation defined on a network. Applied Numerical Mathematics: Transactions of IMACS, 73(??):33– 47, November 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927413000792.

Christara:2008:SII

[CFTW08] Christina Christara, Peter Forsyth, Tamas Terlaky, and Justin W. L. Wan. Seventh IMACS International Symposium on Iterative Methods in Scientific Computing May 5–8, 2005, The University of Toronto, Toronto, Ontario, Canada. Applied Numerical Mathematics: Transactions of IMACS, 58(4): 377–380, April 2008. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Caprini:2010:AFR

[CFV10] Irinel Caprini, Jan Fischer, and Ivo Vrkoč. On the ambiguity of functions represented by divergent power series. Applied Numerical Mathematics: Transactions of IMACS, 60(12): 1264–1272, December 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cai:2008:EGF

[CFX08] Li Cai, Jian-Hu Feng, and Wen-Xian Xie. An efficient ghost fluid method for compressible multifluids in Lagrangian coordinate. Applied Numerical Mathematics: Transactions of IMACS, 58(6):859–870, June 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cai:2006:TDS

[CFXZ06] Li Cai, Jian-Hu Feng, Wen-Xian Xie, and Jun Zhou. Tracking discontinuities in shallow water equations and ideal magnetohydrodynamics equations via Ghost Fluid Method. Applied Numerical Mathematics: Transactions of IMACS, 56(12): 1555–1569, December 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Chan:1989:SCD

[CG89] Tony F. Chan and Danny Goovaerts. Schur complement domain decomposition algorithms for spectral methods. Applied Numerical Mathematics: Transactions of IMACS, 6(1–2):53– 64, December 1989. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Spectral multi-domain methods (Paris, 1988).

Carlenzoli:1992:ENA

[CG92] C. Carlenzoli and P. Gervasio. Effective numerical algorithms for the solution of algebraic systems arising in spectral methods. Applied Numerical Mathematics: Transactions of IMACS, 10(2):87–113, July 1992. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cancelliere:2003:CAN

[CG03] Rossella Cancelliere and Mario Gai. A comparative analysis of neural network performances in astronomical imaging. Applied Numerical Mathematics: Transactions of IMACS, 45(1):87–98, April 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Calvo:2005:LIR

[CG05] M. P. Calvo and A. Gerisch. Linearly implicit Runge– Kutta methods and approximate matrix factorization. Applied Numerical Mathematics: Transactions of IMACS, 53(2–4): 183–200, May 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Chan:2013:APS

[CG13] R. P. K. Chan and A. Gorgey.

Active and passive symmetrization of Runge–Kutta Gauss methods. Applied Numerical Mathematics: Transactions of IMACS, 67(??):64–77, May 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411001073.

Crochet:2014:RPS

[CG14] M. W. Crochet and K. A.

Gonthier. A Riemann problem solution methodology for a class of evolutionary mixture equations with an arbitrary number of components. Applied Numerical Mathematics: Transactions of IMACS, 76(??):145– 165, February 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413001220.

Collins:2016:ADD

[CG16] J. B. Collins and P. A. Gremaud. Analysis of a domain decomposition method for linear transport problems on networks. Applied Numerical Mathematics: Transactions of IMACS, 109(??):61– 72, November 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416300988.

Cheng:2021:RSS

[CG21] Bianru Cheng and Zhenhua Guo. Regularized splitting spectral method for space-fractional logarithmic Schrödinger equation. Applied Numerical Mathematics: Transactions of IMACS, 167(??): 330–355, September 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421001306.

Carpenter:1993:SAB

[CGA93] Mark H. Carpenter, David Gottlieb, and Saul Abarbanel. Stable and accurate boundary treatments for compact, high-order finitedifference schemes. Applied Numerical Mathematics: Transactions of IMACS, 12(1–3):55–87, May 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0168927493901125.

Special issue to honor Professor Saul Abarbanel on his sixtieth birthday (Neveh, 1992).

Cano:1996:GVS

[CGA96] Begoña Cano and Bosco García-Archilla. A generalization to variable stepsizes of Störmer methods for second-order differential equations. Applied Numerical Mathematics: Transactions of IMACS, 19(4): 401–417, February 12, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=19&
issue=4&aid=655.

Chavez-Gonzalez:2002:TQG

[CGCMTR02] A. Chávez-González, A. Cortés-Medina, and J. G. Tinoco-Ruiz. Testing the quality of grids generated using the direct approach. Applied Numerical Mathematics: Transactions of IMACS, 40(1–2):191–206, January 2002. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/85/27/43/abstract.
html.

Calvo-Garrido:2019:JDM

[CGEV19] M. Carmen Calvo-Garrido, Matthias Ehrhardt, and Carlos Vázquez. Jump-diffusion models with two stochastic factors for pricing swing options in electricity markets with partial-integro differential equations. Applied Numerical Mathematics: Transactions of IMACS, 139(??):77–92, May 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419300030.

Castellanos:2002:TMF

[CGG02] J. Longina Castellanos, Susana Gómez, and Valia Guerra. The triangle method for finding the corner of the L-curve. Applied Numerical Mathematics: Transactions of IMACS, 43(4): 359–373, December 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Caucao:2025:PEA

[CGG25] Sergio Caucao, Gabriel N. Gatica, and Luis F. Gatica. A posteriori error analysis of a mixed finite element method for the stationary convective Brinkman– Forchheimer problem. Applied Numerical Mathematics: Transactions of IMACS, 211(??):158–178, May 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742500008X.

Cabrales:2011:SCC

[CGGGS11] R. C. Cabrales, F. Guillén-González, and J. V. Gutiérrez-Santacreu. Stability and convergence for a complete model of mass diffusion. Applied Numerical Mathematics: Transactions of IMACS, 61(11):1161–1185, November 2011. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cuomo:2017:RIC

[CGGM17] Salvatore Cuomo, Ardelio Galletti, Giulio Giunta, and Livia Marcellino. Reconstruction of implicit curves and surfaces via RBF interpolation. Applied Numerical Mathematics: Transactions of IMACS, 116(??): 157–171, June 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416302173.

Cai:2023:WCB

[CGH23] Meng Cai, Siqing Gan, and Yaozhong Hu. Weak convergence of the backward Euler method for stochastic Cahn–Hilliard equation with additive noise. Applied Numerical Mathematics: Transactions of IMACS, 188(??):1–20, June 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423000570.

Cangiani:2016:DGM

[CGJ16] Andrea Cangiani, Emmanuil H.

Georgoulis, and Max Jensen.

Discontinuous Galerkin methods for fast reactive mass transfer through semi-permeable membranes. Applied Numerical Mathematics: Transactions of IMACS, 104(??): 3–14, June 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414001081.

Codina:2025:PEE

[CGK25] Ramon Codina, Hauke Gravenkamp, and Sheraz Ahmed Khan. A posteriori error estimates for the finite element approximation of the convection–diffusion– reaction equation based on the variational multiscale concept. Applied Numerical Mathematics: Transactions of IMACS, 218(??): 238–260, December 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742500159X.

Cardinali:2021:IDG

[CGMS21] Maria Lucia Cardinali, Carlo Garoni, Carla Manni, and Hendrik Speleers. Isogeometric discretizations with generalized B-splines: Symbol-based spectral analysis. Applied Numerical Mathematics: Transactions of IMACS, 166(??): 288–312, August 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001082.

Chakib:2003:NSF

[CGN03] Abdelkrim Chakib, Touria Ghemires, and Abdeljalil Nachaoui. A numerical study of filtration problem in inhomogeneous dam with discontinuous permeability. Applied Numerical Mathematics: Transactions of IMACS, 45(2–3):123–138, May 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cano:2015:ETI

[CGP15] B. Cano and A. González-Pachón. Exponential time integration of solitary waves of cubic Schrödinger equation. Applied Numerical Mathematics: Transactions of IMACS, 91(??): 26–45, May 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415000045.

Cordero:2019:ASA

[CGPT19] A. Cordero, I. Giménez-Palacios, and J. R. Torregrosa. Avoiding strange attractors in efficient parametric families of iterative methods for solving nonlinear problems. Applied Numerical Mathematics: Transactions of IMACS, 137(??): 1–18, March 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418302708.

Chaabane:2018:SEG

[CGRT18] Nabil Chaabane, Vivette Girault, Beatrice Riviere, and Travis Thompson. A stable enriched Galerkin element for the Stokes problem. Applied Numerical Mathematics: Transactions of IMACS, 132(??):1–21, October 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418300977.

Caceres:2019:MVE

[CGS19] Ernesto Cáceres, Gabriel N. Gatica, and Filánder A. Sequeira. A mixed virtual element method for a pseudostress-based formulation of linear elasticity. Applied Numerical Mathematics: Transactions of IMACS, 135(??): 423–442, January 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418301995.

Citro:2020:OER

[CGS20] V. Citro, F. Giannetti, and J. Sierra. Optimal explicit Runge–Kutta methods for compressible Navier–Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 152(??): 511–526, June 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419303137.

Cagnetti:2013:CSD

[CGT13] F. Cagnetti, D. Gomes, and H. V. Tran. Convergence of a semi-discretization scheme for the Hamilton–Jacobi equation: a new approach with the adjoint method. Applied Numerical Mathematics: Transactions of IMACS, 73(??):2–15, November 2013. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413000809.

Cordero:2024:MKM

[CGTTN24] Alicia Cordero, Neus Garrido, Juan R. Torregrosa, and Paula Triguero-Navarro. Modifying Kurchatov’s method to find multiple roots of nonlinear equations. Applied Numerical Mathematics: Transactions of IMACS, 198(??):11–21, April 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423003173.

Chen:2020:FDE

[CGW20] Ziheng Chen, Siqing Gan, and Xiaojie Wang. A fulldiscrete exponential Euler approximation of the invariant measure for parabolic stochastic partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 157(??): 135–158, November 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420301525.

Censor:1987:SOT

[CH87] Yair Censor and Gabor T. Herman. On some optimization techniques in image reconstruction from projections. Applied Numerical Mathematics: Transactions of IMACS, 3(5):365– 391, October 1987. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cloot:1989:MPR

[CH89] A. Cloot and B. M. Herbst. A model for the propagation of rounding error in a Klein–Gordon equation. Applied Numerical Mathematics: Transactions of IMACS, 5(6):481–494, October 1989. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cvetkovic:1990:SSC

[CH90] Ljiljana Cvetković and Dragoslav Herceg. Some sufficient conditions for the convergence of the method of averaging functional corrections. Applied Numerical Mathematics: Transactions of IMACS, 6(3):187–194, March 1990. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Calvo:1995:ALT

[CH95a] M. P. Calvo and E. Hairer.

Accurate long-term integration of dynamical systems. Applied Numerical Mathematics: Transactions of IMACS, 18(1–3):95– 105, September 27, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=18&
issue=1-3&aid=594. Sev-

enth Conference on the Numerical Treatment of Differential Equations (Halle, 1994).

Calvo:1995:FRN

[CH95b] M. P. Calvo and E. Hairer.

Further reduction in the number of independent order conditions for symplectic, explicit partitioned Runge–Kutta and Runge– Kutta–Nyström methods. Applied Numerical Mathematics: Transactions of IMACS, 18(1–3):107–114, September 27, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=18&
issue=1-3&aid=595. Sev-

enth Conference on the Numerical Treatment of Differential Equations (Halle, 1994).

Chen:2001:EPI

[CH01] Ke Chen and Paul J. Harris. Efficient preconditioners for iterative solution of the boundary element equations for the three-dimensional Helmholtz equation. Applied Numerical Mathematics: Transactions of IMACS, 36(4):475–489, March 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/65/32/
31/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/65/32/31/article.
pdf.

Chu:2004:NRS

[CH04] Delin Chu and Y. S. Hung. A numerically reliable solution for the squaringdown problem in system design. Applied Numerical Mathematics: Transactions of IMACS, 51(2–3): 221–241, November 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Chen:2007:AMT

[CH07] Jinru Chen and Peiqi Huang.

Analysis of mortar-type Qrot_{1}/Q_{0} element and multigrid methods for the incompressible Stokes problem. Applied Numerical Mathematics: Transactions of IMACS, 57(5–7):562–576, May 2007. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Chouly:2013:CPM

[CH13] Franz Chouly and Patrick Hild. On convergence of the penalty method for unilateral contact problems. Applied Numerical Mathematics: Transactions of IMACS, 65(??):27–40, March 2013. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412001924.

Chen:2015:ERS

[CH15] Chunguang Chen and Harumi Hattori. Exact Riemann solvers for conservation laws with phase change. Applied Numerical Mathematics: Transactions of IMACS, 94(??):222–240, August 2015. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927415000574.

Chen:2019:FVM

[CH19] Shuangshuang Chen and Qiumei Huang. A finite volume method for a coupled fracture model with matching and nonmatching grids. Applied Numerical Mathematics: Transactions of IMACS, 145(??):28– 47, November 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419301321.

Collins:2021:SVH

[CH21] J. B. Collins and Jonathan D. Hauenstein. A singular value homotopy for finding critical parameter values. Applied Numerical Mathematics: Transactions of IMACS, 161(??): 233–243, March 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420303597.

Cai:2022:HLC

[CH22] Shang-Rong Cai and Feng-Nan Hwang. A hybridline-and-curve search globalization technique for inexact Newton methods. Applied Numerical Mathematics: Transactions of IMACS, 173(??):79–93, March 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421003202.

Cui:2025:DNF

[CH25] Xuewei Cui and Xuehai Huang. A decoupled nonconforming finite element method for biharmonic equation in three dimensions. Applied Numerical Mathematics: Transactions of IMACS, 212(??):300–311, June 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425000364.

Chan:1996:SIR

[Cha96]  R. P. K. Chan.  A-

stability of implicit Runge– Kutta extrapolations. Applied Numerical Mathematics: Transactions of IMACS, 22(1–3):179–203, December 4, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=22&
issue=1-3&aid=713. Special

issue celebrating the centenary of Runge–Kutta methods.

Chatzipantelidis:1998:EMM

[Cha98] Panagiotis Chatzipantelidis. Explicit multistep methods for nonstiff partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 27(1):13–31, May 1, 1998. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1998/27/
                                        1/868.pdf.

Chaturantabut:2017:TLN

[Cha17] Saifon Chaturantabut. Temporal localized nonlinear model reduction with a priori error estimate. Applied Numerical Mathematics: Transactions of IMACS, 119(??):225–238, September 2017. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417300600.

Charkrit:2025:ESL

[Cha25] Sita Charkrit. Explicit solution of Lane–Emden type equations via a novel recurrence and Padé approximation approach. Applied Numerical Mathematics: Transactions of IMACS, 218(??):159–181, December 2025. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425001497.

Chen:1988:PHG

[Che88] Y. M. Chen. Parallelism by hierarchy of GPST inversion algorithm for elastic wave equation. Applied Numerical Mathematics: Transactions of IMACS, 4(1):83–95, March 1988. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Chehab:1996:SGS

[Che96] Jean-Paul Chehab. Solution of generalized Stokes problems using hierarchical methods and Incremental Unknowns. Applied Numerical Mathematics: Transactions of IMACS, 21(1):9– 42, June 6, 1996. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=21&
issue=1&aid=678.

Chen:2012:NAI

[Che12a]  Ruyun Chen.  Numeri-

cal approximations to integrals with a highly oscillatory Bessel kernel. Applied Numerical Mathematics: Transactions of IMACS, 62(5):636–648, May 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927412000268.

Chernov:2012:SPA

[Che12b] Alexey Chernov. Sparse polynomial approximation in positive order Sobolev spaces with bounded mixed derivatives and applications to elliptic problems with random loading. Applied Numerical Mathematics: Transactions of IMACS, 62(4): 360–377, April 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927411001103.

Chen:2016:CNN

[Che16] Yanlai Chen. A certified natural-norm successive constraint method for parametric inf-sup lower bounds. Applied Numerical Mathematics: Transactions of IMACS, 99(??):98–108, January 2016. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927415001361.

Chen:2015:MQM

[CHH15] Chong Chen, Xiaoming He, and Jin Huang. Mechanical quadrature methods and their extrapolations for solving the first kind boundary integral equations of Stokes equation. Applied Numerical Mathematics: Transactions of IMACS, 96(??):165– 179, October 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415000914.

Chiu:1993:OOS

[Chi93] Chichia Chiu. Optimal onestage and two-stage schemes for steady state solutions of hyperbolic equations. Applied Numerical Mathematics: Transactions of IMACS, 11(6):475–496, April 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Chistyakova:2012:RPD

[Chi12] E. V. Chistyakova. Regularizing properties of difference schemes for singular integral-differential equations. Applied Numerical Mathematics: Transactions of IMACS, 62(10): 1302–1311, October 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741200092X.

Chidyagwai:2021:SOM

[Chi21] Prince Chidyagwai. A second order multirate scheme for the evolutionary Stokes– Darcy model. Applied Numerical Mathematics: Transactions of IMACS, 170(??): 364–383, December 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421002324.

Choi:1999:ICB

[CHK99] Haecheon Choi, Michael Hinze, and Karl Kunisch. Instantaneous control of backward-facing step flows. Applied Numerical Mathematics: Transactions of IMACS, 31(2):133–158, October 23, 1999. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1999&volume=31&
issue=2&aid=996;   http:
//www.sciencedirect.com/
science/article/pii/S0168927498001317.

Cao:2021:DMC

[CHLA21] Luling Cao, Yinnian He, Jian Li, and Md. Abdullah Al Mahbub. Decoupled modified characteristic finite element method with different subdomain time steps for nonstationary dual-porosity-Navier–Stokes model. Applied Numerical Mathematics: Transactions of IMACS, 166(??): 238–271, August 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001112.

Cao:2007:HCM

[CHLX07] Yanzhao Cao, Min Huang, Liping Liu, and Yuesheng Xu. Hybrid collocation methods for Fredholm integral equations with weakly singular kernels. Applied Numerical Mathematics: Transactions of IMACS, 57(5–7):549–561, May 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Carstensen:2009:SIB

[CHM09] Carsten Carstensen, Norbert Heuer, and Matthias Maischak. Special issue: Boundary elements – theory and applications dedicated to Professor Ernst P. Stephan on the occasion of his 60th birthday. Applied Numerical Mathematics: Transactions of IMACS, 59(11): 2695–2697, November 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cai:2022:PBP

[CHM22] Yongmei Cai, Junhao Hu, and Xuerong Mao. Positivity and boundedness preserving numerical scheme for the stochastic epidemic model with square-root diffusion term. Applied Numerical Mathematics: Transactions of IMACS, 182(??): 100–116, December 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001921.

Chniti:2017:OIC

[Chn17]  Chokri Chniti.  Opti-

mized interface conditions in domain decomposition methods to solve reactiondiffusion problems with strong heterogeneity in the coefficients in a sectorial domain. Applied Numerical Mathematics: Transactions of IMACS, 118(??): 117–130, August 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417300612.

Chakib:2020:MCM

[CHNN20] A. Chakib, A. Hadri, A. Nachaoui, and M. Nachaoui. Multiscale computational method for nonlinear heat transmission problem in periodic porous media. Applied Numerical Mathematics: Transactions of IMACS, 150(??):164–181, April 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419302545.

Cho:2013:FDA

[Cho13] Chien-Hong Cho. On the finite difference approximation for blow-up solutions of the porous medium equation with a source. Applied Numerical Mathematics: Transactions of IMACS, 65(??):1–26, March 2013. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412001948.

Charnyi:2019:EDE

[CHOR19] Sergey Charnyi, Timo Heister, Maxim A. Olshanskii, and Leo G. Rebholz. Efficient discretizations for the EMAC formulation of the incompressible Navier–Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 141(??): 220–233, ???? 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418302629.

Cai:2019:LEC

[CHP19] Wentao Cai, Dongdong He, and Kejia Pan. A linearized energy-conservative finite element method for the nonlinear Schrödinger equation with wave operator. Applied Numerical Mathematics: Transactions of IMACS, 140(??):183–198, June 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892741930042X.

Chainais-Hillairet:2009:NSE

[CHPV09] Claire Chainais-Hillairet, Yue-Jue Peng, and Ingrid Violet. Numerical solutions of Euler–Poisson systems for potential flows. Applied Numerical Mathematics: Transactions of IMACS, 59(2):301–315, February 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Chrisochoides:1996:MMD

[Chr96] Nikos Chrisochoides. Multithreaded model for the dynamic load-balancing of parallel adaptive PDE computations. Applied Numerical Mathematics: Transactions of IMACS, 20(4):349– 365, June 3, 1996. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=20&
                                        issue=4&aid=652.

Christiansen:2001:DIF

[Chr01] Søren Christiansen. Derivation and investigation of fifth order quadrature formulas for biharmonic boundary integral operators. Applied Numerical Mathematics: Transactions of IMACS, 37(1–2):145–159, April 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/73/26/
34/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/73/26/34/article.
pdf.

Cao:2003:AGM

[CHR03] Weiming Cao, Weizhang Huang, and Robert D. Russell. Approaches for generating moving adaptive meshes: location versus velocity. Applied Numerical Mathematics: Transactions of IMACS, 47(2): 121–138, November 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cirillo:2017:CRD

[CHS17] Emiliano Cirillo, Kai Hormann, and Jean Sidon. Convergence rates of derivatives of Floater–Hormann interpolants for well-spaced nodes. Applied Numerical Mathematics: Transactions of IMACS, 116(??): 108–118, June 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416301349.

Chen:2019:AGL

[CHS19] Hu Chen, Finbarr Holland, and Martin Stynes. An analysis of the Grünwald– Letnikov scheme for initialvalue problems with weakly singular solutions. Applied Numerical Mathematics: Transactions of IMACS, 139(??):52–61, May 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419300066.

Castillo:2001:FSO

[CHSS01] J. E. Castillo, J. M. Hyman, M. Shashkov, and S. Steinberg. Fourth- and sixth-order conservative finite difference approximations of the divergence and gradient. Applied Numerical Mathematics: Transactions of IMACS, 37(1–2): 171–187, April 2001. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/10/28/73/26/
                                        36/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/10/28/73/26/36/article.
                                        pdf.

Chumakov:2003:QIR

[Chu03] Gennadii A. Chumakov. On 2D quasi-isometric regular grids that are orthogonal far from corners. Applied Numerical Mathematics: Transactions of IMACS, 46(3–4):279–294, September 2003. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Chen:2013:ESF

[CHX13] Xingding Chen, Qiya Hu, and Junmin Xiao. On the enhanced strain finite element method for incompressible linear elasticity. Applied Numerical Mathematics: Transactions of IMACS, 72(??):131– 142, October 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413000858.

Cen:2026:NMP

[CHX26] Zhongdi Cen, Jian Huang, and Aimin Xu. A numerical method on a posteriori mesh for a singularly perturbed Riccati equation. Applied Numerical Mathematics: Transactions of IMACS, 219(??):86–95, January 2026. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425001795.

Cao:2014:SSM

[CHZ14] Wanrong Cao, Peng Hao, and Zhongqiang Zhang. Split-step θ-method for stochastic delay differential equations. Applied Numerical Mathematics: Transactions of IMACS, 76(??):19– 33, February 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413001359.

Cai:2021:IZP

[CHZ21] Zhiqiang Cai, Cuiyu He, and Shun Zhang. Improved ZZ a posteriori error estimators for diffusion problems: Discontinuous elements. Applied Numerical Mathematics: Transactions of IMACS, 159(??): 174–189, January 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302816.

Cao:2006:VRM

[CHZZ06] Yanzhao Cao, M. Y. Hussaini, T. Zang, and A. Zatezalo. A variance reduction method based on sensitivity derivatives. Applied Numerical Mathematics: Transactions of IMACS, 56(6):800–813, June 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Califano:2017:SPG

[CIJ17] G. Califano, G. Izzo, and Z. Jackiewicz. Starting procedures for general linear methods. Applied Numerical Mathematics: Transactions of IMACS, 120(??): 165–175, October 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417301241.

Cimmelli:2025:TTS

[Cim25] Vito Antonio Cimmelli.

Thermodynamics of thermoviscoelastic solids of grade

3. Applied Numerical Mathematics: Transactions of IMACS, 208 (part B)(??): 301–313, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424002897.

Calvo:1996:RKM

[CIZ96] M. P. Calvo, A. Iserles, and A. Zanna. Runge– Kutta methods for orthogonal and isospectral flows. Applied Numerical Mathematics: Transactions of IMACS, 22(1–3):153–163, December 4, 1996. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.

                                        cgi?year=1996&volume=22&
                                        issue=1-3&aid=711. Special

issue celebrating the centenary of Runge–Kutta methods.

Chen:1990:ASV

[CJ90] Kang Ping Chen and Daniel D. Joseph. Application of the singular value decomposition to the numerical computation of the coefficients of amplitude equations and normal forms. Applied Numerical Mathematics: Transactions of IMACS, 6(6):425– 430, October 1990. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Chapko:2018:BIE

[CJ18] Roman Chapko and B. Tomas Johansson. A boundary integral equation method for numerical solution of parabolic and hyperbolic Cauchy problems. Applied Numerical Mathematics: Transactions of IMACS, 129(??): 104–119, July 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418300667.

Cubillos:2022:DIC

[CJ22] Max Cubillos and Edwin Jimenez. Diffraction integral computation using sinc approximation. Applied Numerical Mathematics: Transactions of IMACS, 178(??):69–83, August 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422000472.

Clavero:2023:SUC

[CJ23] C. Clavero and J. C. Jorge.

A splitting uniformly convergent method for onedimensional parabolic singularly perturbed convection– diffusion systems. Applied Numerical Mathematics: Transactions of IMACS, 183(??):317–332, January 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422002525.

Clavero:2024:ENM

[CJ24] C. Clavero and J. C. Jorge.

Efficient numerical methods for semilinear one dimensional parabolic singularly perturbed convection– diffusion systems. Applied Numerical Mathematics: Transactions of IMACS, 198(??):461–473, April 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742400028X.

Clavero:2025:EUC

[CJ25] C. Clavero and J. C.

Jorge. An efficient uniformly convergent method for multi-scaled two dimensional parabolic singularly perturbed systems of convection–diffusion type. Applied Numerical Mathematics: Transactions of IMACS, 207(??):174–192, January 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424002344.

Chakraborty:2013:EDC

[CJL13] Debananda Chakraborty, Jae-Hun Jung, and Emmanuel Lorin. An efficient determination of critical parameters of nonlinear Schrödinger equation with a point-like potential using generalized polynomial chaos methods. Applied Numerical Mathematics: Transactions of IMACS, 72(??):115– 130, October 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413000810.

Clavero:1998:FSM

[CJLS98] C. Clavero, J. C. Jorge, F. Lisbona, and G. I. Shishkin. A fractional step method on a special mesh for the resolution of multidimensional evolutionary convection–diffusion problems. Applied Numerical Mathematics: Transactions of IMACS, 27(3):211–231, July 1, 1998. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1998/27/
3/880.pdf.

Cooper:1988:GFE

[CJM88] S. Clement Cooper, William B.

Jones, and Arne Magnus. General T -fraction expansions for ratios of hypergeometric functions. Applied Numerical Mathematics: Transactions of IMACS, 4(2–4):241–251, June 1988. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cuyt:1988:IMT

[CJV88] Annie A. M. Cuyt, Lisa Jacobsen, and Brigitte M. Verdonk. Instability and modification of Thiele interpolating continued fractions. Applied Numerical Mathematics: Transactions of IMACS, 4(2–4):253–262, June 1988. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Chai:2018:MGH

[CjW18] Guo Chai and Tian jun Wang. Mixed generalized Hermite–Fourier spectral method for Fokker– Planck equation of periodic field. Applied Numerical Mathematics: Transactions of IMACS, 133(??): 25–40, ???? 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417302234.

Chen:2011:PHO

[CJX11] Hongtao Chen, Shanghui Jia, and Hehu Xie. Postprocessing and higher order convergence for the mixed finite element approximations of the eigenvalue problem. Applied Numerical Mathematics: Transactions of IMACS, 61(4):615–629, April 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Chen:1998:CGM

[CK98] Zhangxin Chen and Do Y. Kwak. V -cycle Galerkinmultigrid methods for nonconforming methods for nonsymmetric and indefinite problems. Applied Numerical Mathematics: Transactions of IMACS, 28(1):17–35, September 1, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1998/28/
                                        1/902.pdf.

Chertock:2006:PIP

[CK06] Alina Chertock and Alexander Kurganov. On a practical implementation of particle methods. Applied Numerical Mathematics: Transactions of IMACS, 56(10–11):1418– 1431, October/November 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cibik:2013:FEA

[ÇK13] Aytekin Çibik and Songül Kaya. Finite element analysis of a projection-based stabilization method for the Darcy–Brinkman equations in double-diffusive convection. Applied Numerical Mathematics: Transactions of IMACS, 64(1):35– 49, February 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412001766.

Chowdhury:2020:SMS

[CK20] Manisha Chowdhury and B. V. Rathish Kumar. On subgrid multiscale stabilized finite element method for advection–diffusion–reaction equation with variable coefficients. Applied Numerical Mathematics: Transactions of IMACS, 150(??): 576–586, April 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419303083.

Chamakuri:2022:PST

[CK22] Nagaiah Chamakuri and Philipp Kügler. Parallel space-time adaptive numerical simulation of 3D cardiac electrophysiology. Applied Numerical Mathematics: Transactions of IMACS, 173(??):295–307, March 2022. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421003184.

Campbell:2012:MNC

[CKB12] Stephen L. Campbell, Peter Kunkel, and Karen Bobinyec. A minimal norm corrected underdetermined Gauß–Newton procedure. Applied Numerical Mathematics: Transactions of IMACS, 62(5): 592–605, May 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927412000232.

Cheimarios:2013:EPI

[CKB13] N. Cheimarios, G. Kokkoris, and A. G. Boudouvis. An efficient parallel iteration method for multiscale analysis of chemical vapor deposition processes. Applied Numerical Mathematics: Transactions of IMACS, 67(??):78–88, May 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411001115.

Choi:2010:EAC

[CKK10] Sungkyu Choi, Young Pyo Kim, and Jae Ryong Kweon. An error analysis for the corner singularity expansion of a compressible Stokes system on a non-convex polygon. Applied Numerical Mathematics: Transactions of IMACS, 60(8):843–861, August 2010. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Clavero:2025:PPE

[CKK25] Carmelo Clavero, Shashikant Kumar, and Sunil Kumar. A priori and a posteriori error estimates for efficient numerical schemes for coupled systems of linear and nonlinear singularly perturbed initial-value problems. Applied Numerical Mathematics: Transactions of IMACS, 208 (part B) (??):123–147, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424002721.

Cardoso:2003:CSR

[CKL03] João R. Cardoso, Charles S.

Kenney, and F. Silva Leite. Computing the square root and logarithm of a real P - orthogonal matrix. Applied Numerical Mathematics: Transactions of IMACS, 46(2):173–196, August 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Christou:2010:EMC

[CKM10] D. Christou, N. Karcanias, and M. Mitrouli. The ERES method for computing the approximate GCD of several polynomials. Applied Numerical Mathematics: Transactions of IMACS, 60(1–2):94–114, January/ February 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cui:2015:PMP

[CKM15] Shumo Cui, Alexander Kurganov, and Alexei Medovikov. Particle methods for PDEs arising in financial modeling. Applied Numerical Mathematics: Transactions of IMACS, 93(??): 123–139, July 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414000658.

Chu:2025:NAL

[CKM25] Shaoshuai Chu, Alexander Kurganov, and Igor Menshov. New adaptive lowdissipation central-upwind schemes. Applied Numerical Mathematics: Transactions of IMACS, 209(??): 155–170, March 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424003180.

Chamakuri:2015:AOC

[CKP15] Nagaiah Chamakuri, Karl Kunisch, and Gernot Plank. Application of optimal control to the cardiac defibrillation problem using a physiological model of cellular dynamics. Applied Numerical Mathematics: Transactions of IMACS, 95(??): 130–139, September 2015. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415000161.

Carstensen:2015:CRS

[CKPS15] C. Carstensen, K. Köhler, D. Peterseim, and M. Schedensack. Comparison results for the Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 95(??): 118–129, September 2015. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741400021X.

Cheng:2005:HFF

[CKS05] Gary C. Cheng, Roy P. Koomullil, and Bharat K. Soni. High fidelity field simulations using density and pressure based approaches. Applied Numerical Mathematics: Transactions of IMACS, 55(3): 264–282, November 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cash:1985:BEL

[CL85] J. R. Cash and N. Lambrou. Blended extended linear multistep methods for the accurate numerical integration of stiff initial value problems. Applied Numerical Mathematics: Transactions of IMACS, 1(3):195– 216, May 1985. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cohen:1988:SPS

[CL88] A. M. Cohen and I. M. Longman. The summation of power series having positive coefficients. Applied Numerical Mathematics: Transactions of IMACS, 4(6):471– 476, November 1988. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cano:2001:CSH

[CL01a] Begoña Cano and H. Ralph Lewis. A comparison of symplectic and Hamilton’s principle algorithms for autonomous and nonautonomous systems of ordinary differential equations. Applied Numerical Mathematics: Transactions of IMACS, 39(3–4):289–306, December 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/79/34/30/abstract.
html.

Chiavassa:2001:FAW

[CL01b] G. Chiavassa and J. Liandrat. A fully adaptive wavelet algorithm for parabolic partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 36(2– 3):333–358, February 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.nl../38/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/65/31/38/article.
pdf.

Cabrera:2002:MSG

[CL02a] Ileana Miranda Cabrera and Mayra Ramos Lima. Mathematical study of the growth of tetranichidae mites under-field conditions: Relationship prey-predator. Applied Numerical Mathematics: Transactions of IMACS, 40(1–2):137–142, January 2002. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        10/28/85/27/38/abstract.
                                        html.

Chen:2002:STB

[CL02b] Benito Chen and Yaqing Li. Simulation of thick biofilm growth at the microscale. Applied Numerical Mathematics: Transactions of IMACS, 40(1–2):261–271, January 2002. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        10/28/85/27/48/abstract.
                                        html.

Cardoso:2006:PGE

[CL06] João R. Cardoso and F. Silva Leite. Padé and Gregory error estimates for the logarithm of block triangular matrices. Applied Numerical Mathematics: Transactions of IMACS, 56(2):253– 267, February 2006. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Chen:2007:IER

[CL07] Jinhai Chen and Weiguo Li. An improved exponential regula falsi methods with quadratic convergence of both diameter and point for solving nonlinear equations. Applied Numerical Mathematics: Transactions of IMACS, 57(1):80– 88, January 2007. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cao:2008:MNA

[CL08] Li-Qun Cao and Jian-Lan Luo. Multiscale numerical algorithm for the elliptic eigenvalue problem with the mixed boundary in perforated domains. Applied Numerical Mathematics: Transactions of IMACS, 58(9):1349–1374, September 2008. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Chung:2009:MSI

[CL09] Moon-Sun Chung and Sung-Jae Lee. A modified semiimplicit method for a hyperbolic two-fluid model. Applied Numerical Mathematics: Transactions of IMACS, 59(10):2475–2488, October 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Chen:2010:TGF

[CL10] Chuanjun Chen and Wei Liu. Two-grid finite volume element methods for semilinear parabolic problems. Applied Numerical Mathematics: Transactions of IMACS, 60(1–2):10– 18, January/February 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Chen:2014:MNC

[CL14] Qiong-Sheng Chen and Fu-Rong Lin. A modified Nyström–Clenshaw–Curtis quadrature for integral equations with piecewise smooth kernels. Applied Numerical Mathematics: Transactions of IMACS, 85(??):77– 89, November 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414001263.

Choi:2018:EAF

[CL18] Youngmi Choi and Hyung-Chun Lee. Error analysis of finite element approximations of the optimal control problem for stochastic Stokes equations with additive white noise. Applied Numerical Mathematics: Transactions of IMACS, 133(??):144–160, ???? 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418300576.

Choi:2020:OEE

[CL20] Woocheol Choi and Sanghyun Lee. Optimal error estimate of elliptic problems with Dirac sources for discontinuous and enriched Galerkin methods. Applied Numerical Mathematics: Transactions of IMACS, 150(??): 76–104, April 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419302491.

Chen:2024:FSO

[CL24a] Chengyu Chen and Xue-Lei Lin. A fast solvable operator-splitting scheme for time-dependent advection diffusion equation. Applied Numerical Mathematics: Transactions of IMACS, 204(??):48–59, October 2024. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001387.

Chen:2024:AHO

[CL24b] Zhiming Chen and Yong Liu. An arbitrarily high order unfitted finite element method for elliptic interface problems with automatic mesh generation, Part II. Piecewise-smooth interfaces. Applied Numerical Mathematics: Transactions of IMACS, 206(??): 247–268, December 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424002083.

Chen:2026:AHM

[CL26] Yanping Chen and Haitao Leng. An adaptive HDG method for the pointwise tracking optimal control problem of elliptic equations. Applied Numerical Mathematics: Transactions of IMACS, 219(??): 73–85, January 2026. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742500176X.

Courty:2006:CMM

[CLGD06] François Courty, David Leservoisier, Paul-Louis George, and Alain Dervieux. Continuous metrics and mesh adaptation. Applied Numerical Mathematics: Transactions of IMACS, 56(2):117– 145, February 2006. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cheng:2023:IQN

[CLL23] Wanyou Cheng, Zhuanghan LinPeng, and Donghui Li. An inexact quasi-Newton algorithm for large-scale l_{1} optimization with box constraints. Applied Numerical Mathematics: Transactions of IMACS, 193(??): 179–195, November 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001873.

Chen:2025:LDF

[CLL25] Xiaoyong Chen, Rui Li, and Jian Li. The linear, decoupled and fully discrete finite element methods for simplified two-phase ferrohydrodynamics model. Applied Numerical Mathematics: Transactions of IMACS, 210(??):123–146, April 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424003453.

Chen:2021:MQI

[CLLM21] Fang Chen, Tian-Yi Li, Kang-Ya Lu, and Galina V. Muratova. Modified QHSS iteration methods for a class of complex symmetric linear systems. Applied Numerical Mathematics: Transactions of IMACS, 164(??): 3–14, ???? 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300192.

Calvo:1998:VSI

[CLMSS98] M. P. Calvo, M. A. López-Marcos, and J. M. Sanz-Serna. Variable step implementation of geometric integrators. Applied Numerical Mathematics: Transactions of IMACS, 28(1):1–16, September 1, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1998/28/
                                        1/901.pdf.

Chertock:2015:ECA

[CLP15] Alina Chertock, Jian-Guo Liu, and Terrance Pendleton. Elastic collisions among peakon solutions for the Camassa–Holm equation. Applied Numerical Mathematics: Transactions of IMACS, 93(??): 30–46, July 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414000105.

Chahine:2011:NCE

[CLR11] Elie Chahine, Patrick Laborde, and Yves Renard. A nonconformal eXtended Finite Element approach: Integral matching Xfem. Applied Numerical Mathematics: Transactions of IMACS, 61(3):322–343, March 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cotronei:2004:DNN

[CLS04] Mariantonia Cotronei, Maria Laura Lo Cascio, and Thomas Sauer. Dual non-negative rational symbols with arbitrary approximation order. Applied Numerical Mathematics: Transactions of IMACS, 51(4):497–510, December 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Calgaro:1997:DMS

[CLT97] C. Calgaro, J. Laminie, and R. Temam. Dynamical multilevel schemes for the solution of evolution equations by hierarchical finite element discretization. Applied Numerical Mathematics: Transactions of IMACS, 23(4):403–442, May 26, 1997. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1997/23/
4/756.pdf.

Chen:2018:FNM

[CLTA18] S. Chen, F. Liu, I. Turner, and V. Anh. A fast numerical method for twodimensional Riesz space fractional diffusion equations on a convex bounded region. Applied Numerical Mathematics: Transactions of IMACS, 134(??):66– 80, December 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418301557.

Chen:2021:FSA

[CLX21] Li Chen, Shujuan Lü, and Tao Xu. Fourier spectral approximation for time fractional Burgers equation with nonsmooth solutions. Applied Numerical Mathematics: Transactions of IMACS, 169(??): 164–178, November 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421001513.

Chen:2019:EEP

[CLY19] Yanping Chen, Haitao Leng, and Wendi Yang. Error estimates of pseudostressvelocity MFEM for optimal control problems governed by Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 135(??): 407–422, January 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892741830206X.

Celledoni:1997:KPM

[CM97] E. Celledoni and I. Moret. A Krylov projection method for systems of ODEs. Applied Numerical Mathematics: Transactions of IMACS, 24(2–3):365–378, June 19, 1997. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=24&
issue=2-3&aid=789. Volterra

centennial (Tempe, AZ, 1996).

Chan:2000:ESM

[CM00] R. P. K. Chan and A. Murua.

Extrapolation of symplectic methods for Hamiltonian problems. Applied Numerical Mathematics: Transactions of IMACS, 34(2–3): 189–205, July 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/62/29/
31/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/62/29/31/article.
pdf.

Castillo:2002:SSD

[CM02] José Castillo and Terrence McGuinness. Steady state diffusion problems on non-trivial domains: Support operator method integrated with direct optimized grid generation. Applied Numerical Mathematics: Transactions of IMACS, 40(1–2):207–218, January 2002. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        10/28/85/27/44/abstract.
                                        html.

Conti:2004:RPT

[CM04] C. Conti and R. Morandi. Re-parameterisation technique in algebraic numerical grid generation via subdivision schemes. Applied Numerical Mathematics: Transactions of IMACS, 51(4):487–496, December 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Christiansen:2006:TTS

[CM06] Søren Christiansen and Per A. Madsen. On truncated Taylor series and the position of their spurious zeros. Applied Numerical Mathematics: Transactions of IMACS, 56(1):91– 104, January 2006. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cong:2007:CPI

[CM07] Nguyen Huu Cong and Nguyen Van Minh. Continuous parallel-iterated RKNtype PC methods for nonstiff IVPs. Applied Numerical Mathematics: Transactions of IMACS, 57(10): 1097–1107, October 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Carius:2009:HMH

[CM09] Ana Carolina Carius and Alexandre L. Madureira. Hierarchical modeling of the heat equation in a heterogeneous plate. Applied Numerical Mathematics: Transactions of IMACS, 59(9): 2105–2118, September 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cano:2013:HOS

[CM13] B. Cano and M. J. Moreta.

High-order symmetric multistep cosine methods. Applied Numerical Mathematics: Transactions of IMACS, 66(??):30–44, April 2013. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412002000.

Chen:2014:TDC

[CM14] Qiang Chen and Peter Monk.

Time domain CFIEs for electromagnetic scattering problems. Applied Numerical Mathematics: Transactions of IMACS, 79(??): 62–78, May 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413000433.

Cortes-Medina:2002:DFD

[CMCGTR02] A. Cortés-Medina, A. Chávez-González, and J. G. Tinoco-Ruiz. A direct finitedifference scheme for solving PDEs over general twodimensional regions. Applied Numerical Mathematics: Transactions of IMACS, 40(1–2):219–233, January 2002. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/85/27/45/abstract.
html.

Cheng:2005:BUS

[CML05] Yueming Cheng, Duane A. McVay, and W. John Lee. BEM for 3D unsteady-state flow problems in porous media with a finite-conductivity horizontal wellbore. Applied Numerical Mathematics: Transactions of IMACS, 53(1):19–37, April 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Caetano:2025:FOS

[CML^{+}25] Constantino Caetano, Luísa Morgado, Pedro Lima, Niel Hens, and Baltazar Nunes. A fractional order SIR model describing hesitancy to the COVID-19 vaccination. Applied Numerical Mathematics: Transactions of IMACS, 207(??): 608–620, January 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742400268X.

Clemence-Mkhope:2023:PDC

[CMMR23] D. P. Clemence-Mkhope, S. Mabuza, and M. A. Rivas. Persistence of dynamic consistency of nonstandard numerical schemes for the Fisher–KPP equation. Applied Numerical Mathematics: Transactions of IMACS, 185(??):38–55, March 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422002975.

Chatzipantelidis:2003:PEE

[CMP03] P. Chatzipantelidis, Ch.

Makridakis, and M. Plexousakis. A-posteriori error estimates for a finite volume method for the Stokes problem in two dimensions. Applied Numerical Mathematics: Transactions of IMACS, 46(1):45–58, July 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cerimele:2006:HSA

[CMP06] M. M. Cerimele, D. Mansutti, and F. Pistella. Horizontal solidification of average and low Prandtl fluids in microgravity. Applied Numerical Mathematics: Transactions of IMACS, 56(5):682–694, May 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Calabro:2015:CQR

[CMP15] Francesco Calabrò, Carla Manni, and Francesca Pitolli. Computation of quadrature rules for integration with respect to refinable functions on assigned nodes. Applied Numerical Mathematics: Transactions of IMACS, 90(??):168–189, April 2015. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414002001.

Calo:2020:SSP

[CMP20] Victor Calo, Peter Minev, and Vladimir Puzyrev. Splitting schemes for phase-field models. Applied Numerical Mathematics: Transactions of IMACS, 156(??): 192–209, October 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742030132X.

Czyzewska:2023:ESA

[CMP23] Natalia Czyzewska, Paweł M. Morkisz, and Paweł Przybyłowicz. Euler scheme for approximation of solution of nonlinear ODEs under inexact information. Applied Numerical Mathematics: Transactions of IMACS, 193(??): 226–241, November 2023.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423002209.

Conte:2024:CMN

[CMPP24] Dajana Conte, Leila Moradi, Beatrice Paternoster, and Helmut Podhaisky. Collocation methods for nonlinear Volterra integral equations with oscillatory kernel. Applied Numerical Mathematics: Transactions of IMACS, 203(??):1–15, September 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001089.

Calvo:1994:PRK

[CMR94] M. Calvo, J. I. Montijano, and L. Rández. A polyvalent Runge–Kutta triple. Applied Numerical Mathematics: Transactions of IMACS, 15(1):13–26, August 1, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=15&
issue=1&aid=484.

Castillo:2012:MAZ

[CMR12] Kenier Castillo, Mirela V.

Mello, and Fernando R. Rafaeli. Monotonicity and asymptotics of zeros of Sobolev type orthogonal polynomials: a general case. Applied Numerical Mathematics: Transactions of IMACS, 62(11): 1663–1671, November 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927412000773.

Calvo:2024:ETS

[CMRdlT24] M. Calvo, J. I. Montijano, L. Rández, and A. Saenz dela Torre. Explicit two-step peer methods with reused stages. Applied Numerical Mathematics: Transactions of IMACS, 195(??): 75–88, January 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423002465.

Calvo:2011:DET

[CMRV11] M. Calvo, J. I. Montijano, L. Rández, and M. Van Daele. On the derivation of explicit two-step peer methods. Applied Numerical Mathematics: Transactions of IMACS, 61(4): 395–409, April 2011. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Conti:2004:UDU

[CMS04] C. Conti, R. Morandi, and D. Scaramelli. Using discrete uniformity property in a mixed algebraic method. Applied Numerical Mathematics: Transactions of IMACS, 49(3–4):355–366, June 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Conti:2006:AGO

[CMS06] Costanza Conti, Rossana Morandi, and Rosa Maria Spitaleri. An algebraic grid optimization algorithm using condition numbers. Applied Numerical Mathematics: Transactions of IMACS, 56(8):1123–1133, August 2006. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Christou:2025:RUT

[CMT25] Dimitrios Christou, Marilena Mitrouli, and Dimitrios Triantafyllou. A rank-updating technique for the Kronecker canonical form of singular pencils. Applied Numerical Mathematics: Transactions of IMACS, 208 (part A) (??):135–145, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927424000151.

Costabile:2011:ERK

[CN11] F. Costabile and A. Napoli.

Economical Runge–Kutta methods with strong global order one for stochastic differential equations. Applied Numerical Mathematics: Transactions of IMACS, 61(2):160–169, February 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Costabile:2015:CHO

[CN15] Francesco Aldo Costabile and Anna Napoli. Collocation for high order differential equations with twopoints Hermite boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 87(??):157– 167, January 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414001615.

Chun:2016:BAM

[CN16] Changbum Chun and Beny Neta. The basins of attraction of Murakami’s fifth order family of methods. Applied Numerical Mathematics: Transactions of IMACS, 110(??):14–25, December 2016. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416301398.

Costabile:2017:CBL

[CN17] Francesco A. Costabile and Anna Napoli. A class of Birkhoff–Lagrange-collocation methods for high order boundary value problems. Applied Numerical Mathematics: Transactions of IMACS, 116(??):129–140, June 2017. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416302392.

Cao:2023:LMT

[CNA23] Y. Cao, O. Nikan, and Z. Avazzadeh. A localized meshless technique for solving 2D nonlinear integro-differential equation with multi-term kernels. Applied Numerical Mathematics: Transactions of IMACS, 183(??): 140–156, January 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001933.

Comincioli:2000:WBM

[CNS00] Valeriano Comincioli, Giovanni Naldi, and Terenzio Scapolla. A wavelet-based method for numerical solution of nonlinear evolution equations. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4): 291–297, May 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/21/61/27/
                                        56/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/21/61/27/56/article.
                                        pdf.

Cheruvu:2007:SFV

[CNT07] Vani Cheruvu, Ramachandran D. Nair, and Henry M. Tufo. A spectral finite volume transport scheme on the cubed-sphere. Applied Numerical Mathematics: Transactions of IMACS, 57(9):1021–1032, September 2007. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Caliari:2009:IER

[CO09] Marco Caliari and Alexander Ostermann. Implementation of exponential Rosenbrock-type integrators. Applied Numerical Mathematics: Transactions of IMACS, 59(3–4):568–581, March/April 2009. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Codina:2008:ASF

[Cod08] Ramon Codina. Analysis of a stabilized finite element approximation of the Oseen equations using orthogonal subscales. Applied Numerical Mathematics: Transactions of IMACS, 58(3):264–283, March 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cohen:1990:CAA

[Coh90] A. M. Cohen. Chebyshev approximations for cos π/2x^{2} and sin π/2x^{2} with applications. Applied Numerical Mathematics: Transactions of IMACS, 6(3):159– 167, March 1990. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Conroy:1989:PAS

[Con89] John M. Conroy. Parallel algorithms for the solution of narrow banded systems. Applied Numerical Mathematics: Transactions of IMACS, 5(5): 409–421, September 1989. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Condon:1999:CFD

[Con99] D. J. Condon. Corrected finite difference eigenvalues of periodic Sturm–Liouville problems. Applied Numerical Mathematics: Transactions of IMACS, 30(4):393– 401, July 1, 1999. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.

                                        cgi?year=1999&volume=30&
                                        issue=4&aid=958.

Cong:2001:EPT

[Con01] Nguyen Huu Cong. Explicit pseudo two-step RKN methods with stepsize control. Applied Numerical Mathematics: Transactions of IMACS, 38(1–2): 135–144, July 2001. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/10/28/77/27/
                                        34/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/10/28/77/27/34/article.
                                        pdf.

Conti:2004:CCV

[Con04] Costanza Conti. On the convergence of convolved vector subdivision schemes. Applied Numerical Mathematics: Transactions of IMACS, 51(4):477–486, December 2004. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Conte:2020:DLR

[Con20]  Dajana Conte.  Dynam-

ical low-rank approximation to the solution of parabolic differential equations. Applied Numerical Mathematics: Transactions of IMACS, 156(??): 377–384, October 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420301550.

Cooper:1989:EBF

[Coo89] G. J. Cooper. Error bounding functions for Runge– Kutta methods. Applied Numerical Mathematics: Transactions of IMACS, 5(1–2):41–50, February 1989. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Recent theoretical results in numerical ordinary differential equations.

Copetti:2003:FEA

[Cop03] M. I. M. Copetti. Finite element approximation to a quasi-static thermoelastic problem to the contact of two rods. Applied Numerical Mathematics: Transactions of IMACS, 44(1–2): 31–47, January 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Caballero:2025:FEA

[COPPS25] Josefa Caballero, Hanna Okrasińska-Płociniczak, Lukasz Płociniczak, and Kishin Sadarangani. Functional equation arising in behavioral sciences: solvability and collocation scheme in Hölder spaces. Applied Numerical Mathematics: Transactions of IMACS, 212(??): 268–282, June 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425000340.

Cosse:2025:SRQ

[Cos25] Augustin Cosse. Sparse recovery from quadratic measurements with external field. Applied Numerical Mathematics: Transactions of IMACS, 208 (part A) (??):146–169, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424000965.

Coyle:2012:OSF

[Coy12] Joe Coyle. Overlapping solution finite element method

— Higher order approximation and implementation. Applied Numerical Mathematics: Transactions of IMACS, 62(12): 1910–1924, December 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927412001468.

Carstensen:1994:IGS

[CP94] Carsten Carstensen and Miodrag S. Petković. An improvement of Gargantini’s simultaneous inclusion method for polynomial roots by Schröder’s correction. Applied Numerical Mathematics: Transactions of IMACS, 13(6):453– 468, February 1994. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Carstensen:1997:AAS

[CP97] Carsten Carstensen and Petr Plecháč. Adaptive algorithms for scalar non-convex variational problems. Applied Numerical Mathematics: Transactions of IMACS, 26(1–2):203–216, December 22, 1997. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1997/26/
1-2/847.pdf.

Castillo:2003:P

[CP03a] José Castillo and Victor Pereyra. Preface. Applied Numerical Mathematics: Transactions of IMACS, 47(3–4):273–274, December 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cuesta:2003:FTR

[CP03b] E. Cuesta and C. Palencia. A fractional trapezoidal rule for integro-differential equations of fractional order in Banach spaces. Applied Numerical Mathematics: Transactions of IMACS, 45(2–3):139–159, May 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cao:2004:EEM

[CP04] Yang Cao and Linda Petzold. An error estimate for matrix equations. Applied Numerical Mathematics: Transactions of IMACS, 50(3–4): 395–407, September 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Carpentieri:2005:SRO

[CP05a] M. Carpentieri and B. Paternoster. Stability regions of one step mixed collocation methods for y′ = f(x, y). Applied Numerical Mathematics: Transactions of IMACS, 53(2–4):201–212, May 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cecchi:2005:HOF

[CP05b] Maria Morandi Cecchi and Maria Antonietta Pirozzi. High order finite difference numerical methods for time-dependent convection– dominated problems. Applied Numerical Mathematics: Transactions of IMACS, 55(3):334–356, November 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Camilli:2006:SSD

[CP06] Fabio Camilli and Emmanuel Prados. Shape-from-shading with discontinuous image brightness. Applied Numerical Mathematics: Transactions of IMACS, 56(9): 1225–1237, September 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Chang:2007:EDL

[CP07] Xiao-Wen Chang and Christopher C. Paige. Euclidean distances and least squares problems for a given set of vectors. Applied Numerical Mathematics: Transactions of IMACS, 57(11– 12):1240–1244, November/ December 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Conte:2009:MCM

[CP09] D. Conte and B. Paternoster. Multistep collocation methods for Volterra integral equations. Applied Numerical Mathematics: Transactions of IMACS, 59(8): 1721–1736, August 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cvetkovic:2010:MSA

[CP10] Lj. Cvetković and J. M.

Peña. Minimal sets alternative to minimal Geršgorin sets. Applied Numerical Mathematics: Transactions of IMACS, 60(4):442–451, April 2010. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Conte:2017:PMW

[CP17] Dajana Conte and Beatrice Paternoster. Parallel methods for weakly singular Volterra integral equations on GPUs. Applied Numerical Mathematics: Transactions of IMACS, 114(??): 30–37, April 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892741630054X.

Cheng:2025:HIE

[CPC25] Yujun Cheng, Miaojuan Peng, and Yumin Cheng. A hybrid interpolating elementfree Galerkin method for 3D steady-state convection diffusion problems. Applied Numerical Mathematics: Transactions of IMACS, 208 (part B)(??): 21–37, February 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424002605.

Cesco:2005:RAE

[CPD^{+}05] Juan C. Cesco, Jorge E. Pérez, Claudia C. Denner, Graciela O. Giubergiaand, and Ana E. Rosso. Rational approximants to evaluate four-center electron repulsion integrals for 1s hydrogen Slater type functions. Applied Numerical Mathematics: Transactions of IMACS, 55(2):173– 190, October 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Colldeforns-Papiol:2017:TDS

[CPOGO17] G. Colldeforns-Papiol, L. Ortiz-Gracia, and C. W. Oosterlee. Two-dimensional Shannon wavelet inverse Fourier technique for pricing European options. Applied Numerical Mathematics: Transactions of IMACS, 117(??): 115–138, July 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417300624.

Cameron:2002:QSO

[CPP02] Frank Cameron, Mikko Palmroth, and Robert Piché. Quasi stage order conditions for SDIRK methods. Applied Numerical Mathematics: Transactions of IMACS, 42(1–3):61–75, August 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Crisci:1993:FPR

[CPR93] M. R. Crisci, B. Paternoster, and E. Russo. Fully parallel Runge–Kutta–Nyström methods for ODEs with oscillating solutions. Applied Numerical Mathematics: Transactions of IMACS, 11(1–3):143–158, January 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Parallel methods for ordinary differential equations (Grado, 1991).

Castillo:2002:P

[CPS02] José Castillo, Victor Pereyra, and Stanly Steinberg. Preface. Applied Numerical Mathematics: Transactions of IMACS, 40(1–2):1, January 2002. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/85/27/27/abstract.
html.

Chen:2020:NAH

[CPY20] Chuanjun Chen, Kejia Pan, and Xiaofeng Yang. Numerical approximations of a hydro-dynamically coupled phase-field model for binary mixture of passive/active nematic liquid crystals and viscous fluids. Applied Numerical Mathematics: Transactions of IMACS, 158(??):1– 21, December 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302142.

Crivellaro:2017:RSD

[CPZ17] Alberto Crivellaro, Simona Perotto, and Stefano Zonca. Reconstruction of 3D scattered data via radial basis functions by efficient and robust techniques. Applied Numerical Mathematics: Transactions of IMACS, 113(??):93–108, March 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416302227.

Cheng:2020:CAS

[CQZ20] Xiujun Cheng, Hongyu Qin, and Jiwei Zhang. A compact ADI scheme for two-dimensional fractional sub-diffusion equation with Neumann boundary condition. Applied Numerical Mathematics: Transactions of IMACS, 156(??): 50–62, October 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420301240.

Cotter:2004:GIW

[CR04] C. J. Cotter and S. Reich. Geometric integration of a wave-vortex model. Applied Numerical Mathematics: Transactions of IMACS, 48(3–4):293–305, March 2004. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Chehab:2005:IAI

[CR05] Jean-Paul Chehab and Marcos Raydan. Implicit and adaptive inverse preconditioned gradient methods for nonlinear problems. Applied Numerical Mathematics: Transactions of IMACS, 55(1):32– 47, September 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Chikitkin:2019:FCB

[CR19] A. V. Chikitkin and B. V. Rogov. Family of central bicompact schemes with spectral resolution property for hyperbolic equations. Applied Numerical Mathematics: Transactions of IMACS, 142(??): 151–170, August 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419300601.

Chaturvedi:2023:ALF

[CR23a] Abhay Kumar Chaturvedi and S. Chandra Sekhara Rao. Analysis of an LDG-FEM for a two-dimensional singularly perturbed convection– reaction–diffusion problem with interior and boundary layers. Applied Numerical Mathematics: Transactions of IMACS, 190(??): 84–109, August 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001095.

Chung:2023:VPM

[CR23b] Matthias Chung and Rosemary A. Renaut. A variable projection method for large-scale inverse problems with l^{1} regularization. Applied Numerical Mathematics: Transactions of IMACS, 192(??): 297–318, October 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001800.

Cordero:2025:HLC

[CRHTV25] Alicia Cordero, Renso V.

Rojas-Hiciano, Juan R. Torregrosa, and Maria P. Vassileva. High-level convergence order accelerators of iterative methods for nonlinear problems. Applied Numerical Mathematics: Transactions of IMACS, 217(??):390–411, November 2025. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425001448.

Clark:2003:BEW

[CRR03] H. R. Clark, M. A. Rincon, and R. D. Rodrigues. Beam equation with weakinternal damping in domain with moving boundary. Applied Numerical Mathematics: Transactions of IMACS, 47(2):139–157, November 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Calvetti:2005:ISP

[CRS05] D. Calvetti, L. Reichel, and A. Shuibi. Invertible smoothing preconditioners for linear discrete ill-posed problems. Applied Numerical Mathematics: Transactions of IMACS, 54(2):135–149, July 2005. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cruz-Rodriguez:2019:IDU

[CRSF19] Roberto C. Cruz-Rodríguez, Yuri N. Skiba, and Denis M. Filatov. An implicit direct unconditionally stable numerical algorithm for the solution of advection-diffusion equation on a sphere. Applied Numerical Mathematics: Transactions of IMACS, 142(??):1–15, August 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419300431.

Chikitkin:2015:EBV

[CRTU15] A. V. Chikitkin, B. V. Rogov, G. A. Tirsky, and S. V. Utyuzhnikov. Effect of bulk viscosity in supersonic flow past spacecraft. Applied Numerical Mathematics: Transactions of IMACS, 93(??):47–60, July 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414000130.

Chikitkin:2015:HOA

[CRU15] A. V. Chikitkin, B. V. Rogov, and S. V. Utyuzhnikov. High-order accurate monotone compact running scheme for multidimensional hyperbolic equations. Applied Numerical Mathematics: Transactions of IMACS, 93(??):150–163, July 2015. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414000245.

Cecchi:1994:STF

[CS94] M. Morandi Cecchi and A. Scarpa. A space–time finite element method for nonlinear parabolic problems. Applied Numerical Mathematics: Transactions of IMACS, 15(2):247–258, September 13, 1994. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1994&volume=15&
                                        issue=2&aid=499. Innova-

tive methods in numerical analysis (Bressanone, 1992).

Chen:2001:MDS

[CS01] Benito Chen and Jeffrey H. Schenker. Molecular dynamic simulations of gases using a split-Hamiltonian method. Applied Numerical Mathematics: Transactions of IMACS, 38(1– 2):21–48, July 2001. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/10/28/77/27/
                                        28/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/10/28/77/27/28/article.
                                        pdf.

Chen:2003:EMF

[CS03] Benito Chen and Francisco Solis. Explicit mixed finite order Runge–Kutta methods. Applied Numerical Mathematics: Transactions of IMACS, 44(1–2):21–30, January 2003. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cermak:2004:ESA

[CS04] M. Cermak and V. Skala. Edge spinning algorithm for implicit surfaces. Applied Numerical Mathematics: Transactions of IMACS, 49(3–4):331–342, June 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cheng:2008:HOA

[CS08] Juan Cheng and Chi-Wang Shu. A high order accurate conservative remapping method on staggered meshes. Applied Numerical Mathematics: Transactions of IMACS, 58(7): 1042–1060, July 2008. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Chernov:2009:SVB

[CS09] Alexey Chernov and Christoph Schwab. Sparse p-version BEM for first kind boundary integral equations with random loading. Applied Numerical Mathematics: Transactions of IMACS, 59(11):2698–2712, November 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cotronei:2017:NHM

[CS17] Mariantonia Cotronei and Nada Sissouno. A note on Hermite multiwavelets with polynomial and exponential vanishing moments. Applied Numerical Mathematics: Transactions of IMACS, 120(??): 21–34, October 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417301058.

Ciegis:2018:HOC

[CS18] R. Ciegis and O. Suboc. High order compact finite difference schemes on nonuniform grids. Applied Numerical Mathematics: Transactions of IMACS, 132(??): 205–218, October 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418301399.

Cervi:2019:CFO

[CS19] Jessica Cervi and Raymond J. Spiteri. A comparison of fourth-order operator splitting methods for cardiac simulations. Applied Numerical Mathematics: Transactions of IMACS, 145(??): 227–235, November 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419301540.

Chouchoulis:2024:JFI

[CS24] Jeremy Chouchoulis and Jochen Schütz. Jacobianfree implicit MDRK methods for stiff systems of ODEs. Applied Numerical Mathematics: Transactions of IMACS, 196(??):45– 61, February 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423002672.

Crossley:1996:SMS

[CSCM96] P. S. Crossley, R. Saunders, D. M. Causon, and C. G. Mingham. A spectral method with subcell resolution for shock wave calculations. Applied Numerical Mathematics: Transactions of IMACS, 21(2):141– 153, July 8, 1996. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=21&
issue=2&aid=682.

Chen:2024:EHC

[CSL^{+}24] Zibo Chen, Hu Shao, Pengjie Liu, Guoxin Li, and Xianglin Rong. An efficient hybrid conjugate gradient method with an adaptive strategy and applications in image restoration problems. Applied Numerical Mathematics: Transactions of IMACS, 204(??): 362–379, October 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424001685.

Chen:2019:TFT

[CSLY19] Zhaoxia Chen, Lei Shi, Siyao Liu, and Xiong You. Trigonometrically fitted two-derivative Runge– Kutta–Nyström methods for second-order oscillatory differential equations. Applied Numerical Mathematics: Transactions of IMACS, 142(??):171–189, August 2019. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419300595.

Croceri:2007:APA

[CSM07] Graciela M. Croceri, Graciela N. Sottosanto, and María Cristina Maciel. Augmented penalty algorithms based on BFGS secant approximations and trust regions. Applied Numerical Mathematics: Transactions of IMACS, 57(3):320– 334, March 2007. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Chan:1987:SEP

[CSS87] Tony F. Chan, Youcef Saad, and Martin H. Schultz. Solving elliptic partial differential equations on the hypercube multiprocessor. Applied Numerical Mathematics: Transactions of IMACS, 3(1–2):81–88, May 1987. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0168927487900079.

Calini:2019:LIP

[CSS19] A. Calini, C. M. Schober, and M. Strawn. Linear instability of the Peregrine breather: Numerical and analytical investigations. Applied Numerical Mathematics: Transactions of IMACS, 141(??):36–43, ???? 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741830254X.

Calvo:2020:HOS

[CSSZ20] M. P. Calvo, J. M. Sanz-Serna, and Beibei Zhu. Highorder stroboscopic averaging methods for highly oscillatory delay problems. Applied Numerical Mathematics: Transactions of IMACS, 152(??):466–479, June 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419303174.

Calvo:2025:SAM

[CSSZ25] M. P. Calvo, J. M. Sanz-Serna, and Beibei Zhu. Stroboscopic averaging methods to study autoresonance and other problems with slowly varying forcing frequencies. Applied Numerical Mathematics: Transactions of IMACS, 215(??):15– 24, September 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425000844.

Corwin:1997:DCB

[CST97] S. P. Corwin, D. Sarafyan, and S. Thompson. DKLAG6: a code based on continuously imbedded sixthorder Runge–Kutta methods for the solution of state-dependent functionaldifferential equations. Applied Numerical Mathematics: Transactions of IMACS, 24(2–3):319–330, June 19, 1997. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=24&
issue=2-3&aid=785. Volterra

centennial (Tempe, AZ, 1996).

Chamoun:2018:NAC

[CST18] Georges Chamoun, Mazen Saad, and Raafat Talhouk. Numerical analysis of a chemotaxis-swimming bacteria model on a general triangular mesh. Applied Numerical Mathematics: Transactions of IMACS, 127(??): 324–348, May 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418300266.

Chaudhry:2019:PAI

[CSW19] Jehanzeb H. Chaudhry, John N. Shadid, and Timothy Wildey. A posteriori analysis of an IMEX entropy–viscosity formulation for hyperbolic conservation laws with dissipation. Applied Numerical Mathematics: Transactions of IMACS, 135(??): 129–142, January 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418301788.

Chou:2023:AHS

[CSX23] Tom Chou, Sihong Shao, and Mingtao Xia. Adaptive Hermite spectral methods in unbounded domains. Applied Numerical Mathematics: Transactions of IMACS, 183(??):201–220, January 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422002276.

Chen:2014:PSE

[CSXL14] Lizhen Chen, Jie Shen, Chuanju Xu, and Li-Shi Luo. Parallel spectral-element direction splitting method for incompressible Navier– Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 84(??):66–79, October 2014. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414001019.

Chen:1993:IUC

[CT93] Min Chen and Roger Temam. Incremental unknowns for convection-diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 11(5):365–383, March 1993. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Csomos:2021:OSS

[CT21] Petra Csomós and Bálint Takács. Operator splitting for space-dependent epidemic model. Applied Numerical Mathematics: Transactions of IMACS, 159(??):259–280, January 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420302865.

Cui:2004:AIA

[Cui04] M.-R. Cui. Analysis of iterative algorithms of Uzawa type for saddle point problems. Applied Numerical Mathematics: Transactions of IMACS, 50(2):133–146, August 2004. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cui:2024:MPC

[Cui24]  Haijuan Cui.  A modi-

fied PRP conjugate gradient method for unconstrained optimization and nonlinear equations. Applied Numerical Mathematics: Transactions of IMACS, 205(??): 296–307, November 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001946.

Cullum:1995:PPN

[Cul95] Jane K. Cullum. Peaks, plateaus, numerical instabilities in a Galerkin minimal residual pair of methods for solving Ax = b. Applied Numerical Mathematics: Transactions of IMACS, 19(3):255–278, December 15, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.

                                        cgi?year=1996&volume=19&
                                        issue=3&aid=634;   http:
                                        //www.sciencedirect.com/
                                        science/article/pii/0168927495000860.

Special issue on iterative methods for linear equations (Atlanta, GA, 1994).

Cuminato:1995:UCP

[Cum95] J. A. Cuminato. On the uniform convergence of a perturbed collocation method for a class of Cauchy integral equations. Applied Numerical Mathematics: Transactions of IMACS, 16(4):439– 455, April 6, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=16&
                                        issue=4&aid=552.

Cuyt:1990:ONM

[Cuy90]  Annie Cuyt.  Old and

new multidimensional convergence accelerators. Applied Numerical Mathematics: Transactions of IMACS, 6(3):169–185, March 1990. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cuyt:1988:RBC

[CV88] Annie A. M. Cuyt and Brigitte M. Verdonk. A review of branched continued fraction theory for the construction of multivariate rational approximants. Applied Numerical Mathematics: Transactions of IMACS, 4(2–4):263–271, June 1988. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cash:1995:BER

[CV95] J. R. Cash and I. Vieira.

Block 6(5) and 7(6) explicit Runge–Kutta formulae. Applied Numerical Mathematics: Transactions of IMACS, 17(3):223–234, September 4, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=17&
issue=3&aid=578. Numeri-

cal methods for ordinary differential equations (Atlanta, GA, 1994).

Cvetkovic:2002:SCC

[Cve02] Ljiljana Cvetković. Some convergence conditions for a class of parallel decompositiontype linear relaxation methods. Applied Numerical Mathematics: Transactions of IMACS, 41(1):81– 87, April 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/86/27/32/abstract.
html.

Cash:1998:CED

[CW98] J. R. Cash and R. W. Wright. Continuous extensions of deferred correction schemes for the numerical solution of nonlinear twopoint boundary value problems. Applied Numerical Mathematics: Transactions of IMACS, 28(2–4):227–244, October 1, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1998/28/
                                        2-4/911.pdf.

Cao:2020:CGA

[CW20] Junyue Cao and Jinzhao Wu. A conjugate gradient algorithm and its applications in image restoration. Applied Numerical Mathematics: Transactions of IMACS, 152(??):243–252, June 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419303344.

Cao:2021:NPD

[CW21] Waixiang Cao and Chunmei Wang. New primal-dual weak Galerkin finite element methods for convectiondiffusion problems. Applied Numerical Mathematics: Transactions of IMACS, 162(??):171–191, April 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303858.

Chen:2022:FBL

[CW22] Sheng-Gwo Chen and Jyh-Yang Wu. A face-based LTL method for solving diffusion equations and Cahn– Hilliard equations on stationary surfaces. Applied Numerical Mathematics: Transactions of IMACS, 173(??): 13–37, March 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421003160.

Cai:2025:HOS

[CW25] Jiaxiang Cai and Yushun Wang. High-order structurepreserving approaches for constrained conservative or dissipative systems. Applied Numerical Mathematics: Transactions of IMACS, 217(??):340–354, November 2025. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425001412.

Chen:2024:COO

[CWC24] Ziheng Chen, Daoyan Wang, and Lin Chen. Convergence order of one point large deviations rate functions for backward Euler method of stochastic delay differential equations with small noise. Applied Numerical Mathematics: Transactions of IMACS, 202(??): 67–88, August 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424000977.

Chen:2019:TGM

[CWHF19] Yanping Chen, Yang Wang, Yunqing Huang, and Longxia Fu. Two-grid methods of expanded mixed finiteelement solutions for nonlinear parabolic problems. Applied Numerical Mathematics: Transactions of IMACS, 144(??):204–222, October 2019. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419301023.

Chandler-Wilde:2009:PRS

[CWM09] Simon N. Chandler-Wilde and Peter Monk. The PML for rough surface scattering. Applied Numerical Mathematics: Transactions of IMACS, 59(9): 2131–2154, September 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Chen:2021:MSA

[CWP21] Xiao-Ping Chen, Wei Wei, and Xiao-Ming Pan. Modified successive approximation methods for the nonlinear eigenvalue problems. Applied Numerical Mathematics: Transactions of IMACS, 164(??):190–198, ???? 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303573.

Chen:2021:CAS

[CWX21] Jingrun Chen, Cheng Wang, and Changjian Xie. Convergence analysis of a secondorder semi-implicit projection method for Landau– Lifshitz equation. Applied Numerical Mathematics: Transactions of IMACS, 168(??):55–74, October 2021. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001562.

Cheng:2020:EPR

[CWY20] Bianru Cheng, Dongling Wang, and Wei Yang. Energy preserving relaxation method for space-fractional nonlinear Schrödinger equation. Applied Numerical Mathematics: Transactions of IMACS, 152(??): 480–498, June 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419303162.

Chen:2023:TGV

[CWZ23] Fengxin Chen, Qiming Wang, and Zhaojie Zhou. Two-grid virtual element discretization of semilinear elliptic problem. Applied Numerical Mathematics: Transactions of IMACS, 186(??): 228–240, April 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423000090.

Chen:2001:MEM

[CX01] Jinru Chen and Xuejun Xu. The mortar element method for nonselfadjoint parabolic problem. Applied Numerical Mathematics: Transactions of IMACS, 37(3): 341–358, May 2001. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/10/28/73/31/
                                        32/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/10/28/73/31/32/article.
                                        pdf.

Cong:2008:IPI

[CX08] Nguyen Huu Cong and Le Ngoc Xuan. Improved parallel-iterated pseudo twostep RK methods for nonstiff IVPs. Applied Numerical Mathematics: Transactions of IMACS, 58(2):160– 170, February 2008. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cai:2014:HRS

[CXNF14] Li Cai, Wenxian Xie, Yufeng Nie, and Jianhu Feng. Highresolution semi-discrete Hermite central-upwind scheme for multidimensional Hamilton– Jacobi equations. Applied Numerical Mathematics: Transactions of IMACS, 80(??):22–45, June 2014. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414000154.

Chen:2009:CAA

[CXZ09] Junqing Chen, Yifeng Xu, and Jun Zou. Convergence analysis of an adaptive edge element method for Maxwell’s equations. Applied Numerical Mathematics: Transactions of IMACS, 59(12):2950–2969, December 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cao:2014:LAD

[CXZ14] Shuhan Cao, Yunhai Xiao, and Hong Zhu. Linearized alternating directions method for ℓ_{1}-norm inequality constrained ℓ_{1}-norm minimization. Applied Numerical Mathematics: Transactions of IMACS, 85(??): 142–153, November 2014. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414001214.

Chen:2015:PCA

[CXZ15] Wenting Chen, Xiang Xu, and Song-Ping Zhu. A predictor–corrector approach for pricing American options under the finite moment log-stable model. Applied Numerical Mathematics: Transactions of IMACS, 97(??):15–29, November 2015. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927415000987.

Chen:2017:SOH

[CXZ17] Zhongying Chen, Yuesheng Xu, and Jiehua Zhang. A second-order hybrid finite volume method for solving the Stokes equation. Applied Numerical Mathematics: Transactions of IMACS, 119(??): 213–224, September 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417300983.

Campbell:1998:BNT

[CY98] Stephen L. Campbell and Kevin D. Yeomans. Behavior of the nonunique terms in general DAE integrators. Applied Numerical Mathematics: Transactions of IMACS, 28(2–4):209–226, October 1, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1998/28/
2-4/910.pdf.

Cockburn:2005:AMR

[CY05] Bernardo Cockburn and Bayram Yenikaya. An adaptive method with rigorous error control for the Hamilton– Jacobi equations. Part I: The one-dimensional steady state case. Applied Numerical Mathematics: Transactions of IMACS, 52(2– 3):175–195, February 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ciloglu:2022:SDG

[ÇY22] Pelin Çiloğlu and Hamdullah Yücel. Stochastic discontinuous Galerkin methods with low-rank solvers for convection diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 172(??): 157–185, February 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421002920.

Cai:2023:EEF

[CY23] Zhiqiang Cai and Jing Yang. An error estimate for finite element approximation to elliptic PDEs with discontinuous Dirichlet boundary data. Applied Numerical Mathematics: Transactions of IMACS, 193(??):83– 92, November 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423002039.

Chen:2009:ACR

[CYM09] Shaochun Chen, Yongqin Yang, and Shipeng Mao. Anisotropic conforming rectangular elements for elliptic problems of any order. Applied Numerical Mathematics: Transactions of IMACS, 59(5):1137–1148, May 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Chen:2022:IFE

[CYWH22] Yanping Chen, Huaming Yi, Yang Wang, and Yunqing Huang. Immersed finite element approximation for semi-linear parabolic interface problems combining with two-grid methods. Applied Numerical Mathematics: Transactions of IMACS, 175(??):56–72, May 2022.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422000307.

Chen:2021:IFD

[CYYH21] Cairong Chen, Yinong Yang, Dongmei Yu, and Deren Han. An inverse-free dynamical system for solving the absolute value equations. Applied Numerical Mathematics: Transactions of IMACS, 168(??): 170–181, October 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001707.

Chen:1990:HMS

[CZ90] Y. M. Chen and F. G.

Zhang. Hierarchical multigrid strategy for efficiency improvement of the GPST inversion algorithm. Applied Numerical Mathematics: Transactions of IMACS, 6(6):431–446, October 1990. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Campbell:1997:JRE

[CZ97] Stephen L. Campbell and Yangchun Zhong. Jacobian reuse in explicit integrators for higher index DAEs. Applied Numerical Mathematics: Transactions of IMACS, 25(4):391– 412, November 10, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=25&
issue=4&aid=807.

Chen:2004:SLS

[CZ04] Yanping Chen and Manping Zhang. Superconvergence of least-squares mixed finite element for symmetric elliptic problems. Applied Numerical Mathematics: Transactions of IMACS, 48(2):195– 204, February 2004. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Chen:2012:CSE

[CZ12] Hao Chen and Chengjian Zhang. Convergence and stability of extended block boundary value methods for Volterra delay integrodifferential equations. Applied Numerical Mathematics: Transactions of IMACS, 62(2):141–154, February 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0168927411001954.

Chen:2019:FNE

[CZ19] Chiyu Chen and Liping Zhang. Finding Nash equilibrium for a class of multi-person noncooperative games via solving tensor complementarity problem. Applied Numerical Mathematics: Transactions of IMACS, 145(??): 458–468, November 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927419301205.

Chen:2019:PEE

[CZHX19] Yanping Chen, Jinling Zhang, Yunqing Huang, and Yeqing Xu. A posteriori error estimates of hp spectral element methods for integral state constrained elliptic optimal control problems. Applied Numerical Mathematics: Transactions of IMACS, 144(??): 42–58, October 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927419301229.

Chen:2004:AIA

[CZS04] Shaochun Chen, Yongcheng Zhao, and Dongyang Shi. Anisotropic interpolations with application to nonconforming elements. Applied Numerical Mathematics: Transactions of IMACS, 49(2):135–152, May 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Chen:2008:NMF

[CZY08] Xianjin Chen, Jianxin Zhou, and Xudong Yao. A numerical method for finding multiple co-existing solutions to nonlinear cooperative systems. Applied Numerical Mathematics: Transactions of IMACS, 58(11): 1614–1627, November 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Chen:2018:RLS

[CZY18] Lizhen Chen, Jia Zhao, and Xiaofeng Yang. Regularized linear schemes for the molecular beam epitaxy model with slope selection. Applied Numerical Mathematics: Transactions of IMACS, 128(??):139–156, June 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418300436.

Dehghan:2016:AEF

[DA16] Mehdi Dehghan and Mostafa Abbaszadeh. Analysis of the element free Galerkin (EFG) method for solving fractional cable equation with Dirichlet boundary condition. Applied Numerical Mathematics: Transactions of IMACS, 109(??): 208–234, November 2016. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301088.

Dehghan:2017:SET

[DA17] Mehdi Dehghan and Mostafa Abbaszadeh. Spectral element technique for nonlinear fractional evolution equation, stability and convergence analysis. Applied Numerical Mathematics: Transactions of IMACS, 119(??):51–66, September 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0168927417300818.

Dehghan:2018:ETB

[DA18a] Mehdi Dehghan and Mostafa Abbaszadeh. An efficient technique based on finite difference/finite element method for solution of twodimensional space/multitime fractional Bloch–Torrey equations. Applied Numerical Mathematics: Transactions of IMACS, 131(??): 190–206, September 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418300989.

Dehghan:2018:RPO

[DA18b] Mehdi Dehghan and Mostafa Abbaszadeh. A reduced proper orthogonal decomposition (POD) element free Galerkin (POD–EFG) method to simulate twodimensional solute transport problems and error estimate. Applied Numerical Mathematics: Transactions of IMACS, 126(??): 92–112, April 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417302520.

Dehghan:2019:EAN

[DA19] Mehdi Dehghan and Mostafa Abbaszadeh. Error analysis and numerical simulation of magnetohydrodynamics (MHD) equation based on the interpolating element free Galerkin (IEFG) method. Applied Numerical Mathematics: Transactions of IMACS, 137(??): 252–273, March 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418302290.

deAndrade:2010:OPA

[dACR10] E. X. L. de Andrade, M. S. Costa, and A. Sri Ranga. Lorthogonal polynomials associated with related measures. Applied Numerical Mathematics: Transactions of IMACS, 60(10): 1041–1052, October 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

dAlmeida:2017:PMB

[dAF17] Filomena D. d’ Almeida and Rosário Fernandes. Projection methods based on grids for weakly singular integral equations. Applied Numerical Mathematics: Transactions of IMACS, 114(??):47–54, April 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416302008.

Dahl:2002:VRS

[Dah02] Fredrik A. Dahl. Variance reduction for simulated diffusions using control variates extracted from state space evaluations. Applied Numerical Mathematics: Transactions of IMACS, 43(4): 375–381, December 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Dahi:2024:FVM

[DAH24] Ibrahim Dahi, Moulay Rchid Sidi Ammi, and Montasser Hichmani. A finite volume method for a nonlocal thermistor problem. Applied Numerical Mathematics: Transactions of IMACS, 206(??):298–321, December 2024. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424002204.

Dallon:2000:NAD

[Dal00] John C. Dallon. Numerical aspects of discrete and continuum hybrid models in cell biology. Applied Numerical Mathematics: Transactions of IMACS, 32(2):137– 159, February 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/10/28/59/27/
                                        28/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/10/28/59/27/28/article.
                                        pdf.

Dalik:2013:RAD

[Dal13] Josef Dalík. Reduced averaging of directional derivatives in the vertices of unstructured triangulations. Applied Numerical Mathematics: Transactions of IMACS, 67(??):89–97, May 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927411001000.

Durmaz:2023:SON

[DAMA23] Muhammet Enes Durmaz, Ilhame Amirali, Jugal Mohapatra, and Gabil M. Amiraliyev. A secondorder numerical approximation of a singularly perturbed nonlinear Fredholm integro-differential equation. Applied Numerical Mathematics: Transactions of IMACS, 191(??):17– 28, September 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001289.

Daras:1990:CPT

[Dar90] Nicholas J. Daras. The convergence of Padé-type approximants to holomorphic functions of several complex variables. Applied Numerical Mathematics: Transactions of IMACS, 6(5):341– 360, July 1990. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Daripa:2000:FST

[Dar00] Prabir Daripa. The fastest smooth Taylor bubble. Applied Numerical Mathematics: Transactions of IMACS, 34(4):373–379, August 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/62/32/
30/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/62/32/30/article.
pdf.

Datta:1999:LSM

[Dat99a] Biswa Nath Datta. Largescale matrix computations in control. Applied Numerical Mathematics: Transactions of IMACS, 30(1):53– 63, May 10, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1999/30/
                                        1/949.pdf.

Datta:1999:RDN

[Dat99b] Biswa Nath Datta. Recent developments on nonmodal and partial modal approaches for control of vibration. Applied Numerical Mathematics: Transactions of IMACS, 30(1):41– 52, May 10, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1999/30/
                                        1/948.pdf.

Davis:1992:ITM

[Dav92] Stephen F. Davis. An interface tracking method for hyperbolic systems of conservation laws. Applied Numerical Mathematics: Transactions of IMACS, 10(6): 447–472, November 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Davies:1998:SAT

[Dav98] Penny J. Davies. A stability analysis of a time marching scheme for the general surface electric field integral equation. Applied Numerical Mathematics: Transactions of IMACS, 27(1):33– 57, May 1, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1998/27/
1/863.pdf.

Dieci:1995:BIC

[DB95] Luca Dieci and Georg Bader. Block iterations and compactification for periodic block dominant systems associated to invariant tori approximation. Applied Numerical Mathematics: Transactions of IMACS, 17(3):251–274, September 4, 1995. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=17&
issue=3&aid=580. Numeri-

cal methods for ordinary differential equations (Atlanta, GA, 1994).

Davey:1997:PSS

[DB97] K. Davey and S. Bounds.

A preconditioning strategy for the solution of linear boundary element systems using the GMRES method. Applied Numerical Mathematics: Transactions of IMACS, 23(4):443–456, May 26, 1997. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1997/23/
                                        4/761.pdf.

deBortoli:2003:MCR

[dB03] A. L. de Bortoli. Mixing and chemical reacting flow simulations inside square cavities. Applied Numerical Mathematics: Transactions of IMACS, 47(3–4): 295–303, December 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Doha:2008:ESG

[DB08] E. H. Doha and A. H. Bhrawy. Efficient spectral-Galerkin algorithms for direct solution of fourthorder differential equations using Jacobi polynomials. Applied Numerical Mathematics: Transactions of IMACS, 58(8):1224–1244, August 2008. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Doha:2014:NJR

[DBBH14] E. H. Doha, A. H. Bhrawy, D. Baleanu, and R. M. Hafez. A new Jacobi rational-Gauss collocation method for numerical solution of generalized pantograph equations. Applied Numerical Mathematics: Transactions of IMACS, 77(??): 43–54, March 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413001530.

Deckers:2010:PRI

[DBCBPP10] Karl Deckers, Adhemar Bultheel, Ruymán Cruz-Barroso, and Francisco Perdomo-Pío. Positive rational interpolatory quadrature formulas on the unit circle and the interval. Applied Numerical Mathematics: Transactions of IMACS, 60(12):1286–1299, December 2010. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Dupire:2010:NNT

[DBDV10] G. Dupire, J. P. Boufflet, M. Dambrineand, and P. Villon. On the necessity of Nitsche term. Applied Numerical Mathematics: Transactions of IMACS, 60(9):888–902, September 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Devine:2005:NCD

[DBH^{+}05] Karen D. Devine, Erik G.

Boman, Robert T. Heaphy, Bruce A. Hendrickson, James D. Teresco, Jamal Faik, Joseph E. Flaherty, and Luis G. Gervasio. New challenges in dynamic load balancing. Applied Numerical Mathematics: Transactions of IMACS, 52(2– 3):133–152, February 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Leo:2025:NTS

[DBH25] Mariano De Leo, Juan Pablo Borgna, and Cristian Huenchul. Non trivial solutions for a system of coupled Ginzburg– Landau equations. Applied Numerical Mathematics: Transactions of IMACS, 208 (part B)(??):271–289, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424002770.

Dong:2009:MAE

[DC09] Qiao-Li Dong and Li-Qun Cao. Multiscale asymptotic expansions and numerical algorithms for the wave equations of second order with rapidly oscillating coefficients. Applied Numerical Mathematics: Transactions of IMACS, 59(12): 3008–3032, December 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

deCamargo:2018:ELB

[dC18a] André Pierro de Camargo.

An exponential lower bound for the condition number of real Vandermonde matrices. Applied Numerical Mathematics: Transactions of IMACS, 128(??): 81–83, June 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418300291.

Dutykh:2018:WDN

[DC18b] Denys Dutykh and Jean-Guy Caputo. Wave dynamics on networks: Method and application to the sine-Gordon equation. Applied Numerical Mathematics: Transactions of IMACS, 131(??):54– 71, September 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418300722.

DAmore:2021:MRS

[DC21] L. D’Amore and R. Cacciapuoti. Model reduction in space and time for ab initio decomposition of 4D variational data assimilation problems. Applied Numerical Mathematics: Transactions of IMACS, 160(??): 242–264, February 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303007.

Dhia:2014:TDM

[DCC14] A.-S. Bonnet-Ben Dhia, L. Chesnel, and P. Ciarlet, Jr. Two-dimensional Maxwell’s equations with sign-changing coefficients. Applied Numerical Mathematics: Transactions of IMACS, 79(??):29–41, May 2014. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0168927413000767.

Coimbra:2003:MFE

[dCCSR03] Maria do Carmo Coimbra, Carlos Sereno, and Alírio Rodrigues. A moving finite element method for the solution of two-dimensional time-dependent models. Applied Numerical Mathematics: Transactions of IMACS, 44(4):449–469, March 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Duan:2020:CFD

[DCJ20] Hui Duan, Xinjuan Chen, and Jae-Hun Jung. On the consistency of the finite difference approximation with the Riemann– Liouville fractional derivative for 0 < α < 1. Applied Numerical Mathematics: Transactions of IMACS, 153(??):35–51, July 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300210.

Du:2023:HOS

[DCL23] Shaohong Du, Yongping Cheng, and Mingjun Li. High order spline finite element method for the fourth-order parabolic equations. Applied Numerical Mathematics: Transactions of IMACS, 184(??): 496–511, February 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422002902.

Dong:2019:MCM

[DCN^{+}19] Hao Dong, Junzhi Cui, Yufeng Nie, Zihao Yang, Qiang Ma, and Xiaohan Cheng. Multiscale computational method for transient heat conduction problems of periodic porous materials with diverse periodic configurations in different subdomains. Applied Numerical Mathematics: Transactions of IMACS, 136(??): 215–234, February 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418302381.

Cola:2025:AMM

[DCSB25] Vincenzo Schiano Di Cola, Salvatore Cuomo, Gerardo Severino, and Marco Berardi. Algebraic multigrid methods for uncertainty quantification of sourcetype flows through randomly heterogeneous porous media. Applied Numerical Mathematics: Transactions of IMACS, 218(??):58– 72, December 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425001321.

Du:2020:SLR

[DCY20] Hong Du, Zhong Chen, and Tiejun Yang. A stable least residue method in reproducing kernel space for solving a nonlinear fractional integro-differential equation with a weakly singular kernel. Applied Numerical Mathematics: Transactions of IMACS, 157(??): 210–222, November 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420301860.

Davies:1997:ATT

[DD97] Penny J. Davies and Dugald B. Duncan. Averaging techniques for time-marching schemes for retarded potential integral equations. Applied Numerical Mathematics: Transactions of IMACS, 23(3):291–310, April 25, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.com/
cas/tree/store/apnum/sub/
1997/23/3/751.pdf.

Daripa:2019:CAH

[DD19] Prabir Daripa and Sourav Dutta. On the convergence analysis of a hybrid numerical method for multicomponent transport in porous media. Applied Numerical Mathematics: Transactions of IMACS, 146(??): 199–220, December 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927419301813.

DellAccio:2020:HSM

[DD20] Francesco Dell’Accio and Filomena Di Tommaso. On the hexagonal Shepard method. Applied Numerical Mathematics: Transactions of IMACS, 150(??): 51–64, April 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419302442.

Dang:2021:UAS

[DD21] Quang A. Dang and Quang Long Dang. A unified approach to study the existence and numerical solution of functional differential equation. Applied Numerical Mathematics: Transactions of IMACS, 170(??): 208–218, December 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421002154.

Doescher:2003:INC

[DdCVR03] Erwin Doescher, Haroldo F. de Campos Velho, and Fernando M. Ramos. Isothermal and natural convection flows in fractal cavities. Applied Numerical Mathematics: Transactions of IMACS, 47(3–4):407–419, December 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

deCougny:1994:LBP

[dDF^{+}94] H. L. deCougny, K. D. Devine, J. E. Flaherty, R. M. Loy, C. Özturan, and M. S. Shephard. Load balancing for the parallel adaptive solution of partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 16(1–2): 157–182, December 1, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/

cas_sub/browse/browse.
cgi?year=1994&volume=16&
issue=1-2&aid=515.  A

Festschrift to honor Professor Robert Vichnevetsky on his 65th birthday.

DellAccio:2024:EMS

[DDF24] Francesco Dell’Accio, Filomena Di Tommaso, and Elisa Francomano. The enriched multinode Shepard collocation method for solving elliptic problems with singularities. Applied Numerical Mathematics: Transactions of IMACS, 205(??):87– 100, November 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001776.

DellAccio:2023:ITS

[DDGN23] Francesco Dell’Accio, Filomena Di Tommaso, Allal Guessab, and Federico Nudo. On the improvement of the triangular Shepard method by non conformal polynomial elements. Applied Numerical Mathematics: Transactions of IMACS, 184(??): 446–460, February 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422002872.

Dahlke:1997:SMB

[DDHS97] Stephan Dahlke, Wolfgang Dahmen, Reinhard Hochmuth, and Reinhold Schneider. Stable multiscale bases and local error estimation for elliptic problems. Applied Numerical Mathematics: Transactions of IMACS, 23(1):21– 47, March 7, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1997/23/

1/742.pdf. Multilevel methods (Oberwolfach, 1995).

Dinvay:2019:CSB

[DDK19] Evgueni Dinvay, Denys Dutykh, and Henrik Kalisch. A comparative study of bidirectional Whitham systems. Applied Numerical Mathematics: Transactions of IMACS, 141(??): 248–262, ???? 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418302137.

DellAccio:2018:IAO

[DDNZ18] F. Dell’Accio, F. Di Tommaso, O. Nouisser, and B. Zerroudi. Increasing the approximation order of the triangular Shepard method. Applied Numerical Mathematics: Transactions of IMACS, 126(??): 78–91, April 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417302568.

Dari:2012:PEE

[DDP12] E. A. Dari, R. G. Durán, and C. Padra. A posteriori error estimates for nonconforming approximation of eigenvalue problems. Applied Numerical Mathematics: Transactions of IMACS, 62(5):580–591, May 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412000220.

Djukic:2024:WAG

[DDRS24] Dusan Lj. Djukić, Rada M. Mutavdzić Djukić, Lothar Reichel, and Miodrag M. Spalević. Weighted averaged Gaussian quadrature rules for modified Chebyshev measures. Applied Numerical Mathematics: Transactions of IMACS, 200(??): 195–208, June 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001344.

Demaret:1989:TCS

[DDS89] P. Demaret, M. O. Deville, and C. Schneidesch. Thermal convection solutions by Chebyshev pseudospectral multi-domain decomposition and finite element preconditioning. Applied Numerical Mathematics: Transactions of IMACS, 6(1–2): 107–121, December 1989. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Spectral multi-domain methods (Paris, 1988).

DAmbra:2007:DPB

[DdSF07] Pasqua D’Ambra, Daniela di Serafino, and Salvatore Filippone. On the development of PSBLASbased parallel two-level Schwarz preconditioners. Applied Numerical Mathematics: Transactions of IMACS, 57(11–12):1181– 1196, November/December 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Davies:2005:SNA

[DDZK05] Penny J. Davies, Dugald B. Duncan, and Barbara Zubik-Kowal. The stability of numerical approximations of the time domain current induced on thin wire and strip antennas. Applied Numerical Mathematics: Transactions of IMACS, 55(1):48– 68, September 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

DeGoede:1988:SLW

[De 88] E. D. De Goede. Stabilization of the Lax–Wendroff method and a generalized one-step Runge–Kutta method for hyperbolic initialvalue problems. Applied Numerical Mathematics: Transactions of IMACS, 4(5):439–453, July 1988. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

deBoor:1992:EMP

[de 92a] C. de Boor. On the error in multivariate polynomial interpolation. Applied Numerical Mathematics: Transactions of IMACS, 10(3–4): 297–305, September 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). A Festschrift to honor Professor Garrett Birkhoff on his eightieth birthday.

deGoede:1992:NMS

[de 92b] Erik D. de Goede. Numerical methods for the 3D shallow water equations on vector and parallel computers. Applied Numerical Mathematics: Transactions of IMACS, 10(1):3–18, June 1992. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

deLaBourdonnaye:1993:NTI

[de 93a] Armel de La Bourdonnaye.

Numerical treatment of integral equations on iPSC. Applied Numerical Mathematics: Transactions of IMACS, 12(5):473–483, July 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0168927493901062.

DeRoeck:1993:NES

[De 93b] Yann-Hervé De Roeck. Nonlinear elasticity solved by a domain decomposition method on a hypercube. Applied Numerical Mathematics: Transactions of IMACS, 12(5):459–471, July 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/016892749390105Z.

deSturler:1995:IBL

[de 95a] E. de Sturler. Incomplete block LU preconditioners on slightly overlapping subdomains for a massively parallel computer. Applied Numerical Mathematics: Transactions of IMACS, 19(1–2): 129–146, December 4, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=1-2&aid=625. Mas-

sively parallel computing and applications (Amsterdam, 1993–1994).

deSwart:1995:EPP

[de 95b] J. J. B. de Swart. Efficient parallel predictor–corrector methods. Applied Numerical Mathematics: Transactions of IMACS, 18(1–3):387–396, September 27, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=18&
issue=1-3&aid=609. Sev-

enth Conference on the Numerical Treatment of Differential Equations (Halle, 1994).

deZeeuw:1996:DSC

[de 96] Paul M. de Zeeuw. Development of semi-coarsening techniques. Applied Numerical Mathematics: Transactions of IMACS, 19(4): 433–465, February 12, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=19&
issue=4&aid=643.

DeBortoli:2002:MBA

[De 02] A. L. De Bortoli. Multigrid based aerodynamical simulations for the NACA 0012 airfoil. Applied Numerical Mathematics: Transactions of IMACS, 40(1–2):337–349, January 2002. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        10/28/85/27/53/abstract.
                                        html.

DelaSen:2006:ACC

[De 06] M. De la Sen. On the adaptive control of a class of SISO dynamic hybrid systems. Applied Numerical Mathematics: Transactions of IMACS, 56(5):618–647, May 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Deckelnick:2006:PGT

[DE06] Klaus Deckelnick and Charles M. Elliott. Propagation of graphs in two-dimensional inhomogeneous media. Applied Numerical Mathematics: Transactions of IMACS, 56(9):1163–1178, September 2006. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

DellAcqua:2016:AMI

[DE16] Pietro Dell’Acqua and Claudio Estatico. Acceleration of multiplicative iterative algorithms for image deblurring by duality maps in Banach spaces. Applied Numerical Mathematics: Transactions of IMACS, 99(??):121– 136, January 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741500135X.

Marchi:2018:QMC

[DE18] Stefano De Marchi and Giacomo Elefante. Quasi-Monte Carlo integration on manifolds with mapped low-discrepancy points and greedy minimal Riesz senergy points. Applied Numerical Mathematics: Transactions of IMACS, 127(??): 110–124, May 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417302672.

Dea:2011:IPH

[Dea11] John R. Dea. Improving the performance of Higdon non-reflecting boundary conditions by using weighted differencing. Applied Numerical Mathematics: Transactions of IMACS, 61(11): 1186–1197, November 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Diaz:2025:HOW

[DEGDdL25] M. J. Castro Díaz, C. Escalante, J. Garres-Díaz, and T. Morales de Luna. Highorder well-balanced schemes for shallow models for dry avalanches. Applied Numerical Mathematics: Transactions of IMACS, 215(??): 138–156, September 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742500087X.

Dehghan:2001:DCP

[Deh01] Mehdi Dehghan. Determination of a control parameter in the two-dimensional diffusion equation. Applied Numerical Mathematics: Transactions of IMACS, 37(4):489–502, June 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/10/28/73/33/
                                        30/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/10/28/73/33/30/article.
                                        pdf.

Dehghan:2005:ETS

[Deh05] Mehdi Dehghan. Efficient techniques for the secondorder parabolic equation subject to nonlocal specifications. Applied Numerical Mathematics: Transactions of IMACS, 52(1):39– 62, January 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Deka:2017:PEE

[Dek17] Bhupen Deka. A priori L^{∞}(L^{2}) error estimates for finite element approximations to the wave equation with interface. Applied Numerical Mathematics: Transactions of IMACS, 115(??):142–159, May 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417300065.

Denk:1993:NNM

[Den93] G. Denk. A new numerical method for the integration of highly oscillatory secondorder ordinary differential equations. Applied Numerical Mathematics: Transactions of IMACS, 13(1– 3):57–67, September 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Sixth Conference on the Numerical Treatment of Differential Equations (Halle, 1992).

Deng:2007:NSO

[Den07] Shaozhong Deng. Numerical simulation of optical coupling and light propagation in coupled optical resonators with size disorder. Applied Numerical Mathematics: Transactions of IMACS, 57(5–7):475–485, May 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Deng:2015:SFO

[Den15] Dingwen Deng. The study of a fourth-order multistep ADI method applied to nonlinear delay reaction-diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 96(??):118– 133, October 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415000951.

Dassi:2015:MSS

[DEPS15] Franco Dassi, Bree Ettinger, Simona Perotto, and Laura M. Sangalli. A mesh simplification strategy for a spatial regression analysis over the cortical surface of the brain. Applied Numerical Mathematics: Transactions of IMACS, 90(??):111–131, April 2015. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414001950.

Derakhshan:1992:PSV

[Der92] Mir S. Derakhshan. Parallel solution of Volterra integral equations using reducible quadrature rules. Applied Numerical Mathematics: Transactions of IMACS, 9(3–5):283–290, April 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0168927492900226.

International Conference on the Numerical Solution of Volterra and Delay Equations (Tempe, AZ, 1990).

Desideri:2008:IIS

[Dés08] Jean-Antoine Désidéri. Interpolation of infinite sequences by entire functions. Applied Numerical Mathematics: Transactions of IMACS, 58(12):1918–1932, December 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

DelBuono:2025:PHO

[DESZ25] Nicoletta Del Buono, Flavia Esposito, Laura Selicato, and Rafał Zdunek. Penalty hyperparameter optimization with diversity measure for nonnegative lowrank approximation. Applied Numerical Mathematics: Transactions of IMACS, 208 (part A)(??):189–204, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927424002708.

Dupin:1992:SPI

[DF92] Jean-Claude Dupin and Arnaud Fréville. Shape preserving interpolating cubic splines with geometric mesh. Applied Numerical Mathematics: Transactions of IMACS, 9(6):447– 459, May 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Devine:1996:PAR

[DF96] Karen D. Devine and Joseph E. Flaherty. Parallel adaptive hp-refinement techniques for conservation laws. Applied Numerical Mathematics: Transactions of IMACS, 20(4):367–386, June 3, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=20&
                                        issue=4&aid=651. Adaptive

mesh refinement methods for CFD applications (Atlanta, GA, 1994).

Darbandi:2011:RDS

[DF11] Masoud Darbandi and Nematollah Fouladi. A reduced domain strategy for local mesh movement application in unstructured grids. Applied Numerical Mathematics: Transactions of IMACS, 61(9): 1001–1016, September 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Dorini:2009:EMS

[DFC09] F. A. Dorini, F. Furtado, and M. C. C. Cunha. On the evaluation of moments for solute transport by random velocity fields. Applied Numerical Mathematics: Transactions of IMACS, 59(12): 2994–2998, December 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Diogo:2017:P

[DFJ^{+}17] Teresa Diogo, Neville Ford, Zdzislaw Jackiewicz, Pedro Lima, Luisa Morgado, and Magda Rebelo. Preface. Applied Numerical Mathematics: Transactions of IMACS, 114(??):1, April 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416302331.

Diogo:2025:P

[DFL^{+}25] Teresa Diogo, Neville J.

Ford, Pedro Lima, Luisa Morgado, and Magda Rebelo. Preface. Applied Numerical Mathematics: Transactions of IMACS, 214(??):80–81, August 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425000522.

Deng:2019:CSS

[DFLM19] Shounian Deng, Weiyin Fei, Yong Liang, and Xuerong Mao. Convergence of the split-step θmethod for stochastic agedependent population equations with Markovian switching and variable delay. Applied Numerical Mathematics: Transactions of IMACS, 139(??):15–37, May 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419300017.

deFrutos:2000:PBI

[dFN00] Javier de Frutos and Julia Novo. A postprocess based improvement of the spectral element method. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4):217–223, May 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/61/27/
47/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/61/27/47/article.
pdf.

Dassi:2016:DTR

[DFZ16] Franco Dassi, Luca Formaggia, and Stefano Zonca. Degenerate tetrahedra removal. Applied Numerical Mathematics: Transactions of IMACS, 110(??):1– 13, December 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301404.

deGroen:1991:BCR

[dG91] Pieter P. N. de Groen. Basep-cyclic reduction for tridiagonal systems of equations. Applied Numerical Mathematics: Transactions of IMACS, 8(2):117– 125, September 1991. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0168927491900463.

Dempsey:1996:NIB

[DG96] K. M. Dempsey and I. Gladwell. Numerical instability in the breathing-mode dynamics of elastic rings. Applied Numerical Mathematics: Transactions of IMACS, 20(1–2):211–220, February 29, 1996. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=1-2&aid=675. Work-

shop on the method of lines for time-dependent problems (Lexington, KY, 1995).

DeWitte:2010:CAN

[DG10] V. De Witte and W. Govaerts. Convergence analysis of a numerical method to solve the adjoint linearized periodic orbit equations. Applied Numerical Mathematics: Transactions of IMACS, 60(10): 1007–1023, October 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Dominguez:2022:AAO

[DG22] V. Domínguez and M. Ganesh. Analysis and application of an overlapped FEM–BEM for wave propagation in unbounded and heterogeneous media. Applied Numerical Mathematics: Transactions of IMACS, 171(??):76– 105, January 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421002439.

DellAccio:2017:SIN

[DGCW17] Francesco Dell’Accio, Maria Italia Gualtieri, Stefano Serra Capizzano, and Gerhard Wanner. Special issue on New Trends in Numerical Analysis: Theory, Methods, Algorithms and Applications (NETNA2015). Applied Numerical Mathematics: Transactions of IMACS, 116(??):1, June 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417300521.

Domingues:2003:AWR

[DGD03] M. O. Domingues, S. M. Gomes, and L. M. A. Díaz.

Adaptive wavelet representation and differentiation on block-structured grids. Applied Numerical Mathematics: Transactions of IMACS, 47(3–4):421–437, December 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Dehghan:2022:UES

[DGE22] Mehdi Dehghan, Zeinab Gharibi, and Mohammad Reza Eslahchi. Unconditionally energy stable C^{0}-virtual element scheme for solving generalized Swift–Hohenberg equation. Applied Numerical Mathematics: Transactions of IMACS, 178(??): 304–328, August 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422000757.

DeLuca:2025:NAT

[DGGM25] P. De Luca, A. Galletti, G. Giunta, and L. Marcellino. A numerical approach for a 1D tumorangiogenesis simulations model. Applied Numerical Mathematics: Transactions of IMACS, 210(??):83–94, April 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0168927424003258.

Donat:2018:IEW

[DGM18] R. Donat, F. Guerrero, and P. Mulet. Implicit-explicit WENO scheme for the equilibrium dispersive model of chromatography. Applied Numerical Mathematics: Transactions of IMACS, 123(??):22–42, January 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417301812.

Dassi:2022:AVE

[DGM22] F. Dassi, J. Gedicke, and L. Mascotto. Adaptive virtual element methods with equilibrated fluxes. Applied Numerical Mathematics: Transactions of IMACS, 173(??):249–278, March 2022. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421003342.

Demkowicz:2012:CDP

[DGN12] Leszek Demkowicz, Jay Gopalakrishnan, and Antti H. Niemi. A class of discontinuous Petrov–Galerkin methods. Part III: Adaptivity. Applied Numerical Mathematics: Transactions of IMACS, 62(4): 396–427, April 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411001656.

Domingues:2009:STA

[DGRS09] Margarete O. Domingues, Sônia M. Gomes, Olivier Roussel, and Kai Schneider. Space–time adaptive multiresolution methods for hyperbolic conservation laws: Applications to compressible Euler equations. Applied Numerical Mathematics: Transactions of IMACS, 59(9): 2303–2321, September 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Dunster:2024:CPC

[DGS24] T. M. Dunster, A. Gil, and J. Segura. Computation of parabolic cylinder functions having complex argument. Applied Numerical Mathematics: Transactions of IMACS, 197(??): 230–242, March 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423002969.

Daruis:2000:SPQ

[DGV00] Leyla Daruis and Pablo González-Vera. Szegő polynomials and quadrature formulas on the unit circle. Applied Numerical Mathematics: Transactions of IMACS, 36(1):79– 112, November 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.nl../29/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/65/26/29/article.
pdf.

Dougherty:1994:DCA

[DH94] Randall L. Dougherty and James M. Hyman. A divide-and-conquer algorithm for grid generation. Applied Numerical Mathematics: Transactions of IMACS, 14(1–3):125–134, April 14, 1994. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=14&
issue=1-3&aid=456. Pro-

ceedings of the Third ARO Workshop on Adaptive Methods for Partial Differential Equations (Troy, NY, 1992).

daCunha:1995:PIM

[dH95] Rudnei Dias da Cunha and Tim Hopkins. The Parallel Iterative Methods (PIM) package for the solution of systems of linear equations on parallel computers. Applied Numerical Mathematics: Transactions of IMACS, 19(1–2):33–50, December 4, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=1-2&aid=565.

Dorfler:2007:CAH

[DH07] W. Dörfler and V. Heuveline. Convergence of an adaptive hp. Applied Numerical Mathematics: Transactions of IMACS, 57(10): 1108–1124, October 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ditkowski:2012:WSR

[DH12a] A. Ditkowski and Y. Harness. Wave scattering by randomly shaped objects. Applied Numerical Mathematics: Transactions of IMACS, 62(12): 1819–1836, December 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412001249.

Dou:2012:TVR

[DH12b] Yi-Xin Dou and Bo Han. Total variation regularization for the reconstruction of a mountain topography. Applied Numerical Mathematics: Transactions of IMACS, 62(1):1–20, January 2012.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892741100167X.

Dinesen:2000:PCT

[DHL00] P. G. Dinesen, J. S. Hesthaven, and J. P. Lynov. A pseudospectral collocation time-domain method for diffractive optics. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4):199–206, May 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/21/61/27/
                                        45/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/21/61/27/45/article.
                                        pdf.

Drblikova:2009:EEF

[DHM09] Olga Drblíková, Angela Handlovičová, and Karol Mikula. Error estimates of the finite volume scheme for the nonlinear tensordriven anisotropic diffusion. Applied Numerical Mathematics: Transactions of IMACS, 59(10):2548–2570, October 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Dostal:2005:SFD

[DHS05] Zdeněk Dostál, David Horák, and Dan Stefanica. A scalable FETI–DP algorithm for a coercive variational inequality. Applied Numerical Mathematics: Transactions of IMACS, 54(3– 4):378–390, August 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Dan:2009:FDI

[DhW09] Wei Dan and Ren hong Wang. A fourth degree integration formula for the ndimensional simplex. Applied Numerical Mathematics: Transactions of IMACS, 59(12):2990–2993, December 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Deng:2022:DCM

[DHWL22] Ting Deng, Jin Huang, Xiaoxia Wen, and Hongyan Liu. Discrete collocation method for solving twodimensional linear and nonlinear fuzzy Volterra integral equations. Applied Numerical Mathematics: Transactions of IMACS, 171(??): 389–407, January 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421002683.

Diamantakis:1995:NST

[Dia95] M. T. Diamantakis. The NUMOL solution of timedependent PDEs using DESI Runge–Kutta formulae. Applied Numerical Mathematics: Transactions of IMACS, 17(3):235–249, September 4, 1995. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=17&
issue=3&aid=579. Numeri-

cal methods for ordinary differential equations (Atlanta, GA, 1994).

Diehl:2015:NIC

[Die15] S. Diehl. Numerical identification of constitutive functions in scalar nonlinear convection-diffusion equations with application to batch sedimentation. Applied Numerical Mathematics: Transactions of IMACS, 95(??): 154–172, September 2015. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414000531.

Dieci:2020:SIN

[Die20] Luca Dieci. Sliding integration with no projection. Applied Numerical Mathematics: Transactions of IMACS, 155(??):3–15, ???? 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419301850.

Dumbser:2015:ESI

[DII15] Michael Dumbser, Uwe Iben, and Matteo Ioriatti. An efficient semi-implicit finite volume method for axially symmetric compressible flows in compliant tubes. Applied Numerical Mathematics: Transactions of IMACS, 89(??):24–44, March 2015. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414001780.

DAmbrosio:2012:SHS

[DIJ12] R. D’Ambrosio, G. Izzo, and Z. Jackiewicz. Search for highly stable two-step Runge–Kutta methods. Applied Numerical Mathematics: Transactions of IMACS, 62(10):1361–1379, October 2012. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412000980.

Dinavahi:1993:SMC

[Din93] Surya Prasad G. Dinavahi.

Streakline method for computing two-dimensional vortical flows. Applied Numerical Mathematics: Transactions of IMACS, 11(4):283– 308, February 1993. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ding:2019:HON

[Din19]  Hengfei Ding.  A high-

order numerical algorithm for two-dimensional timespace tempered fractional diffusion-wave equation. Applied Numerical Mathematics: Transactions of IMACS, 135(??):30–46, January 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418301739.

Dimitrov:2013:CHP

[DIR13] D. K. Dimitrov, M. E. H. Ismail, and A. Sri Ranga. A class of hypergeometric polynomials with zeros on the unit circle: Extremal and orthogonal properties and quadrature formulas. Applied Numerical Mathematics: Transactions of IMACS, 65(??):41–52, March 2013. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892741200195X.

Dittmann:2021:HOM

[Dit21] Alexander J. Dittmann. High-order multiderivative IMEX schemes. Applied Numerical Mathematics: Transactions of IMACS, 160(??):205–216, February 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303147.

Dorao:2008:H

[DJ08] C. A. Dorao and H. A.

Jakobsen. hp. Applied Numerical Mathematics: Transactions of IMACS, 58(5):563–576, May 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Delcourte:2010:SPP

[DJ10] Sarah Delcourte and Delphine Jennequin. Saddle point preconditioners for linearized Navier–Stokes equations discretized by a finite volume method. Applied Numerical Mathematics: Transactions of IMACS, 60(11):1054–1066, November 2010. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Durdek:2015:AOP

[DJJ^{+}15] Antoine Durdek, Stig Rune Jensen, Jonas Juselius, Peter Wind, Tor Flå, and Luca Frediani. Adaptive order polynomial algorithm in a multiwavelet representation scheme. Applied Numerical Mathematics: Transactions of IMACS, 92(??): 40–53, June 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414001998.

Dunca:2004:ALA

[DJL04] Adrian Dunca, Volker John, and William Layton. Approximating local averages of fluid velocities: The equilibrium Navier–Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 49(2):187–205, May 2004. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Driver:2009:IZL

[DJM09] Kathy Driver, Kerstin Jordaan, and Norbert Mbuyi. Interlacing of zeros of linear combinations of classical orthogonal polynomials from different sequences. Applied Numerical Mathematics: Transactions of IMACS, 59(10):2424–2429, October 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Cabrera:2022:CCN

[DJNR22] Omar De la Cruz Cabrera, Jiafeng Jin, Silvia Noschese, and Lothar Reichel. Communication in complex networks. Applied Numerical Mathematics: Transactions of IMACS, 172(??):

186–205, February 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421002907.

DelaCruzCabrera:2025:SAC

[DJR25] Omar De la Cruz Cabrera, Jiafeng Jin, and Lothar Reichel. Sparse approximation of complex networks. Applied Numerical Mathematics: Transactions of IMACS, 208 (part A) (??):170–188, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927424000023.

Debrabant:2011:CSB

[DK11] Kristian Debrabant and Anne Kværnø. Composition of stochastic B-series with applications to implicit Taylor methods. Applied Numerical Mathematics: Transactions of IMACS, 61(4):501–511, April 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Duru:2014:ESP

[DK14] Kenneth Duru and Gunilla Kreiss. Efficient and stable perfectly matched layer for CEM. Applied Numerical Mathematics: Transactions of IMACS, 76(??):34– 47, February 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927413001347.

Demir:2020:ALE

[DK20] Medine Demir and Songül Kaya. An analysis of a linearly extrapolated BDF2 subgrid artificial viscosity method for incompressible flows. Applied Numerical Mathematics: Transactions of IMACS, 156(??): 140–157, October 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420301252.

Deka:2021:EEW

[DK21] Bhupen Deka and Naresh Kumar. Error estimates in weak Galerkin finite element methods for parabolic equations under low regularity assumptions. Applied Numerical Mathematics: Transactions of IMACS, 162(??):81–105, April 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420303767.

Dnestrovskaya:1994:DMS

[DKK94] E. Yu. Dnestrovskaya, N. N. Kalitkin, and L. V. Kusmina. Lq-decreasing monotonic schemes with complex coefficients and applications to complicated PDE systems. Applied Numerical Mathematics: Transactions of IMACS, 15(3): 327–340, October 5, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=15&
issue=3&aid=510. Interna-

tional Conference on Scientific Computation and Differential Equations (Auckland, 1993).

Djellouli:2024:LEP

[DKL24] Rabia Djellouli, David Klein, and Matthew Levy. Legendre expansions of products of functions with applications to nonlinear partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 201(??):301–321, July 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000631.

Das:2024:IHO

[DKSS24] Arijit Das, Prakrati Kushwah, Jitraj Saha, and Mehakpreet Singh. Improved higher-order finite volume scheme and its convergence analysis for collisional breakage equation.

Applied Numerical Mathematics: Transactions of IMACS, 196(??):118–132, February 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423002714.

Demetriou:2001:CPD

[DL01] I. C. Demetriou and E. A. Lipitakis. Certain positive definite submatrices that arise from binomial coefficient matrices. Applied Numerical Mathematics: Transactions of IMACS, 36(2–3):219–229, February 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.nl../31/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/10/28/65/31/31/article.
                                        pdf.

Duran:2006:FEA

[DL06] Ricardo G. Durán and Ariel L. Lombardi. Finite element approximation of convection diffusion problems using graded meshes. Applied Numerical Mathematics: Transactions of IMACS, 56(10–11):1314–1325, October/November 2006. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Dieci:2013:NSD

[DL13] Luca Dieci and Luciano Lopez. Numerical solution of discontinuous differential systems: Approaching the discontinuity surface from one side. Applied Numerical Mathematics: Transactions of IMACS, 67(??):98–110, May 2013. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411001644.

Dong:2016:RCD

[DL16] Haiyun Dong and Maojun Li. A reconstructed central discontinuous Galerkinfinite element method for the fully nonlinear weakly dispersive Green-Naghdi model. Applied Numerical Mathematics: Transactions of IMACS, 110(??):110–127, December 2016. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301581.

Deng:2020:EPF

[DL20] Dingwen Deng and Dong Liang. The energy-preserving finite difference methods and their analyses for system of nonlinear wave equations in two dimensions. Applied Numerical Mathematics: Transactions of IMACS, 151(??):172–198, May 2020.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419303757.

Destuynder:2021:FRP

[DL21a] Philippe Destuynder and Erwan Liberge. A few remarks on penalty and penaltyduality methods in fluidstructure interactions. Applied Numerical Mathematics: Transactions of IMACS, 167(??):1–30, September 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0168927421001161.

Dong:2021:NSO

[DL21b] Jian Dong and Ding Fang Li. A new second-order modified hydrostatic reconstruction for the shallow water flows with a discontinuous topography. Applied Numerical Mathematics: Transactions of IMACS, 161(??): 408–424, March 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742030369X.

Deng:2022:ROE

[DL22a] Qiuxiang Deng and Zhendong Luo. A reducedorder extrapolated finite difference iterative scheme for uniform transmission line equation. Applied Numerical Mathematics: Transactions of IMACS, 172(??): 514–524, February 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742100310X.

Deng:2022:UAE

[DL22b] Shuo Deng and Jiyong Li. A uniformly accurate exponential wave integrator Fourier pseudo-spectral method with energy-preservation for longtime dynamics of the nonlinear Klein–Gordon equation. Applied Numerical Mathematics: Transactions of IMACS, 178(??): 166–191, August 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422000873.

delaCruz:2023:SWS

[dlC23] H. de la Cruz. A simplified weak simulation method for the probabilistic response analysis of nonlinear random vibration problems. Applied Numerical Mathematics: Transactions of IMACS, 183(??): 186–200, January 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422002252.

delaHoz:2013:STD

[dlHV13] Francisco de la Hoz and Fernando Vadillo. The solution of two-dimensional advection-diffusion equations via operational matrices. Applied Numerical Mathematics: Transactions of IMACS, 72(??):172– 187, October 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927413000883.

Duran:2002:NBS

[DLM02] A. Durán and M. A. López-Marcos. Numerical behaviour of stable and unstable solitary waves. Applied Numerical Mathematics: Transactions of IMACS, 42(1–3):95–116, August 2002. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Desbrun:2005:DPL

[DLM05] Mathieu Desbrun, Melvin Leok, and Jerrold E. Marsden. Discrete Poincaré lemma. Applied Numerical Mathematics: Transactions of IMACS, 53(2–4): 231–248, May 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Luigi:2016:RAN

[DLM16] Christophe De Luigi, Jérôme Lelong, and Sylvain Maire. Robust adaptive numerical integration of irregular functions with applications to basket and other multi-dimensional exotic options. Applied Numerical Mathematics: Transactions of IMACS, 100(??):14– 30, February 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415001580.

Ding:2020:CAC

[DLM20] Qianqian Ding, Xiaonian Long, and Shipeng Mao. Convergence analysis of Crank–Nicolson extrapolated fully discrete scheme for thermally coupled incompressible magnetohydrodynamic system. Applied Numerical Mathematics: Transactions of IMACS, 157(??):522–543, November 2020. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302002.

Dolean:2004:CAA

[DLN04] Victorita Dolean, Stéphane Lanteri, and Frédéric Nataf. Convergence analysis of additive Schwarz for the Euler equations. Applied Numerical Mathematics: Transactions of IMACS, 49(2): 153–186, May 2004. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

DeBonis:2024:SIF

[DLN^{+}24] Maria Carmela De Bonis, Concetta Laurita, Incoronata Notarangelo, Donatella Occorsio, Maria Grazia Russo, and Yuesheng Xu. Special issue on functional analysis, approximation theory and numerical analysis (FAATNA20>22). Applied Numerical Mathematics: Transactions of IMACS, 200(??):1, June 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423002787.

DiPietro:2006:MPF

[DLP06] Daniele A. Di Pietro, Stefania Lo Forte, and Nicola Parolini. Mass preserving finite element implementations of the level set method. Applied Numerical Mathematics: Transactions of IMACS, 56(9): 1179–1195, September 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Diogo:2017:STS

[DLPV17] T. Diogo, P. M. Lima, A. Pedas, and G. Vainikko. Smoothing transformation and spline collocation for weakly singular Volterra integro-differential equations. Applied Numerical Mathematics: Transactions of IMACS, 114(??): 63–76, April 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301593.

Dong:2023:SCA

[DLQZ23] Yuzhuo Dong, Xiao Li, Zhonghua Qiao, and Zhengru Zhang. Stability and convergence analysis of the exponential time differencing scheme for a Cahn–Hilliard binary fluidsurfactant model. Applied Numerical Mathematics: Transactions of IMACS, 190(??):321–343, August 2023. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001241.

Bonis:2022:NML

[DLS22] M. C. De Bonis, C. Laurita, and V. Sagaria. A numerical method for linear Volterra integral equations on infinite intervals and its application to the resolution of metastatic tumor growth models. Applied Numerical Mathematics: Transactions of IMACS, 172(??): 475–496, February 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421003007.

Dujardin:2024:SRK

[DLV24] Guillaume Dujardin and Ingrid Lacroix-Violet. Âand Î-stability of Runge– Kutta collocation methods. Applied Numerical Mathematics: Transactions of IMACS, 202(??): 158–172, August 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424001132.

Du:2021:RRT

[DLZ21] Shaohong Du, Runchang Lin, and Zhimin Zhang. Robust recovery-type a posteriori error estimators for streamline upwind/Petrov Galerkin discretizations for singularly perturbed problems. Applied Numerical Mathematics: Transactions of IMACS, 168(??): 23–40, October 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421001495.

Douak:1997:OTP

[DM97] K. Douak and P. Maroni.

On d-orthogonal Tchebychev polynomials. I. Applied Numerical Mathematics: Transactions of IMACS, 24(1):23–53, June 9, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=24&
issue=1&aid=762.

Dehghan:2009:MLP

[DM09a] Mehdi Dehghan and Davoud Mirzaei. Meshless Local Petrov–Galerkin (MLPG) method for the unsteady magnetohydrodynamic (MHD) flow through pipe with arbitrary wall conductivity. Applied Numerical Mathematics: Transactions of IMACS, 59(5):1043–1058, May 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Deng:2009:OEE

[DM09b] Zhenguo Deng and Heping Ma. Optimal error estimates of the Fourier spectral method for a class of nonlocal, nonlinear dispersive wave equations. Applied Numerical Mathematics: Transactions of IMACS, 59(5):988–1010, May 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Dorsek:2010:AH

[DM10] P. Dörsek and J. M. Melenk. Adaptive hp. Applied Numerical Mathematics: Transactions of IMACS, 60(7):689–704, July 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Deuring:2011:EEF

[DM11a] Paul Deuring and Marcus Mildner. Error estimates for a finite element– finite volume discretization of convection–diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 61(6):785–801, June 2011. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Diele:2011:ESP

[DM11b] Fasma Diele and Carmela Marangi. Explicit symplectic partitioned Runge–Kutta– Nyström methods for nonautonomous dynamics. Applied Numerical Mathematics: Transactions of IMACS, 61(7):832–843, July 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Darrigrand:2012:CUW

[DM12] E. Darrigrand and P. Monk. Combining the Ultra-Weak Variational Formulation and the multilevel fast multipole method. Applied Numerical Mathematics: Transactions of IMACS, 62(6): 709–719, June 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741100119X.

Duc:2022:SRB

[DMA22] Nguyen Van Duc, Pham Quy Muoi, and Nguyen Thi Van Anh. Stability results for backward heat equations with time-dependent coefficient in the Banach space L_{p}(R). Applied Numerical Mathematics: Transactions of IMACS, 175(??): 40–55, May 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422000290.

Diaz-Mendoza:2005:CTP

[DMGVO05] C. Díaz-Mendoza, P. González-Vera, and R. Orive. On the convergence of twopoint partial Padé approximants for meromorphic functions of Stieltjes type. Applied Numerical Mathematics: Transactions of IMACS, 53(1):39–56, April 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Diaz-Mendoza:2009:CSS

[DMGVPO09] C. Díaz-Mendoza, P. González-Vera, M. Jiménez Paiz, and R. Orive. On certain symmetric strong distributions, two-point Padé approximation and related quadratures. Applied Numerical Mathematics: Transactions of IMACS, 59(8):2002– 2014, August 2009. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Dehghani-Madiseh:2018:EAE

[DMH18] Marzieh Dehghani-Madiseh and Milan Hladík. Efficient approaches for enclosing the united solution set of the interval generalized Sylvester matrix equations. Applied Numerical Mathematics: Transactions of IMACS, 126(??):18–33, April 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417302519.

Doctor:2024:EDS

[DMM24a] Katarina Doctor, Tong Mao, and Hrushikesh Mhaskar. Encoding of data sets and algorithms. Applied Numerical Mathematics: Transactions of IMACS, 200(??): 209–235, June 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001952.

Donat:2024:WSC

[DMM24b] R. Donat, M. C. Martí, and P. Mulet. WENO scheme on characteristics for the equilibrium dispersive model of chromatography with generalized Langmuir isotherms. Applied Numerical Mathematics: Transactions of IMACS, 201(??): 247–264, July 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000552.

Dede:2008:AEC

[DMP08] Luca Dedè, Stefano Micheletti, and Simona Perotto. Anisotropic error control for environmental applications. Applied Numerical Mathematics: Transactions of IMACS, 58(9):1320–1339, September 2008. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Dieci:1999:RLN

[DMPP99] Luca Dieci, Benedetta Morini, Alessandra Papini, and Aldo Pasquali. On real logarithms of nearby matrices and structured matrix interpolation. Applied Numerical Mathematics: Transactions of IMACS, 29(1):145–165, January 1, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1999/29/
                                        1/893.pdf.

Donatelli:2016:MMC

[DMPSC16] Marco Donatelli, Matteo Molteni, Vincenzo Pennati, and Stefano Serra-Capizzano. Multigrid methods for cubic spline solution of two point (and 2D) boundary value problems. Applied Numerical Mathematics: Transactions of IMACS, 104(??):15–29, June 2016. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414000555.

Discacciati:2002:MNM

[DMQ02] Marco Discacciati, Edie Miglio, and Alfio Quarteroni. Mathematical and numerical models for coupling surface and groundwater flows. Applied Numerical Mathematics: Transactions of IMACS, 43(1–2): 57–74, October 2002. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Dimitrov:2010:MZJ

[DMR10] Dimitar K. Dimitrov, Mirela V. Mello, and Fernando R. Rafaeli. Monotonicity of zeros of Jacobi–Sobolev type orthogonal polynomials. Applied Numerical Mathematics: Transactions of IMACS, 60(3):263–276, March 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Dolejsi:2018:CHM

[DMR18] Vít Dolejsí, Georg May, and Ajay Rangarajan. A continuous hp-mesh model for adaptive discontinuous Galerkin schemes. Applied Numerical Mathematics: Transactions of IMACS, 124(??):1– 21, February 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741730209X.

DAmbrosio:2023:LTA

[DMS23] Raffaele D’Ambrosio, Afsaneh Moradi, and Carmela Scalone. A long term analysis of stochastic theta methods for mean reverting linear process with jumps. Applied Numerical Mathematics: Transactions of IMACS, 185(??): 516–529, March 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422003336.

Djadel:2008:NCF

[DN08] K. Djadel and S. Nicaise.

A non-conforming finite volume element method of weighted upstream type for the two-dimensional stationary Navier–Stokes system. Applied Numerical Mathematics: Transactions of IMACS, 58(5):615–634, May 2008. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

DiPietro:2013:LFD

[DN13] Daniele A. Di Pietro and Serge Nicaise. A lockingfree discontinuous Galerkin method for linear elasticity in locally nearly incompressible heterogeneous media. Applied Numerical Mathematics: Transactions of IMACS, 63(1):105– 116, January 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412001717.

Dehbozorgi:2021:NSN

[DN21] Raziyeh Dehbozorgi and Khadijeh Nedaiasl. Numerical solution of nonlinear weakly singular Volterra integral equations of the first kind: an hp-version collocation approach. Applied Numerical Mathematics: Transactions of IMACS, 161(??):111–136, March 2021. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303470.

Denich:2024:SNT

[DN24] Eleonora Denich and Paolo Novati. Some notes on the trapezoidal rule for Fourier type integrals. Applied Numerical Mathematics: Transactions of IMACS, 198(??):160–175, April 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000047.

Ding:2018:FCS

[DNW18] Weiyang Ding, Michael Ng, and Yimin Wei. Fast computation of stationary joint probability distribution of sparse Markov chains. Applied Numerical Mathematics: Transactions of IMACS, 125(??):68–85, March 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417302258.

DeLong:1995:SP

[DO95] M. A. DeLong and J. M.

Ortega. SOR as a preconditioner. Applied Numerical Mathematics: Transactions of IMACS, 18(4): 431–440, November 9, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=18&
issue=4&aid=628.

DeLong:1998:SPI

[DO98] M. A. DeLong and J. M. Ortega. SOR as a preconditioner II. Applied Numerical Mathematics: Transactions of IMACS, 26(4):465– 481, April 1, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1998/26/
4/860.pdf.

DeBonis:2017:AHH

[DO17a] Maria Carmela De Bonis and Donatella Occorsio. Approximation of Hilbert and Hadamard transforms on (0, +∞). Applied Numerical Mathematics: Transactions of IMACS, 116(??): 184–194, June 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416302379.

DeBonis:2017:SAH

[DO17b] Maria Carmela De Bonis and Donatella Occorsio. On the simultaneous approximation of a Hilbert transform and its derivatives on the real semiaxis. Applied Numerical Mathematics: Transactions of IMACS, 114(??):132–153, April 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416302380.

Dobrowolski:2005:LCN

[Dob05] Manfred Dobrowolski. On the LBB condition in the numerical analysis of the Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 54(3– 4):314–323, August 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

deOliveira:2020:INM

[dOF20] F. R. de Oliveira and O. P. Ferreira. Inexact Newton method with feasible inexact projections for solving constrained smooth and nonsmooth equations. Applied Numerical Mathematics: Transactions of IMACS, 156(??):63–76, October 2020. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420301276.

Doi:1991:PCI

[Doi91] Shun Doi. On parallelism and convergence of incomplete LU factorizations. Applied Numerical Mathematics: Transactions of IMACS, 7(5):417–436, June 1991.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/016892749190011N.

Dolejsi:2014:AHA

[Dol14] Vít Dolejsí. Anisotropic hp-adaptive method based on interpolation error estimates in the L^{q}-norm. Applied Numerical Mathematics: Transactions of IMACS, 82(??):80–114, August 2014. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414000567.

Dorr:1991:DDP

[Dor91]  Milo R. Dorr.  A do-

main decomposition preconditioner with reduced rank interdomain coupling. Applied Numerical Mathematics: Transactions of IMACS, 8(4–5):333–352, November 1991. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Dorodnitsyn:2001:NTT

[Dor01] Vladimir Dorodnitsyn. Noethertype theorems for difference equations. Applied Numerical Mathematics: Transactions of IMACS, 39(3–4):307–321, December 2001. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.

elsevier.com/gej-ng/10/
10/28/79/34/31/abstract.
html.

Douglas:1991:TAD

[Dou91] Craig C. Douglas. A tupleware approach to domain decomposition methods. Applied Numerical Mathematics: Transactions of IMACS, 8(4–5):353–373, November 1991. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Dryja:1985:CMM

[DP85] M. Dryja and W. Proskurowski.

Capacitance matrix method using strips with alternating Neumann and Dirichlet boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 1(4):285–298, July 1985. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Dryja:1990:CIM

[DP90] M. Dryja and W. Proskurowski.

Composite iterative method for elliptic problems in irregular regions. Applied Numerical Mathematics: Transactions of IMACS, 6(4):263– 279, May 1990. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

DAmbrosio:2012:TSM

[DP12] Raffaele D’Ambrosio and Beatrice Paternoster. Twostep modified collocation methods with structured coefficient matrices. Applied Numerical Mathematics: Transactions of IMACS, 62(10):1325–1334, October 2012. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412000943.

Duong:2021:ISE

[DP21] Viet Thong Duong and Tu Vuong Phan. Improved subgradient extragradient methods for solving pseudomonotone variational inequalities in Hilbert spaces. Applied Numerical Mathematics: Transactions of IMACS, 163(??): 221–238, May 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000398.

Dabaghi:2016:RFE

[DPPR16] Farshid Dabaghi, Adrien Petrov, Jérôme Pousin, and Yves Renard. A robust finite element redistribution approach for elastodynamic contact problems. Applied Numerical Mathematics: Transactions of IMACS, 103(??):48–71, May 2016. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416000052.

deNicola:1996:NMS

[dPT96] C. de Nicola, G. Pinto, and R. Tognaccini. A normal mode stability analysis of multiblock algorithms for the solution of fluid-dynamics equations. Applied Numerical Mathematics: Transactions of IMACS, 19(4): 419–431, February 12, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=19&
issue=4&aid=644.

Dyn:1993:DDT

[DR93] Nira Dyn and Shmuel Rippa.

Data-dependent triangulations for scattered data interpolation and finite element approximation. Applied Numerical Mathematics: Transactions of IMACS, 12(1–3):89–105, May 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0168927493901136.

Special issue to honor Professor Saul Abarbanel on his sixtieth birthday (Neveh, 1992).

Dimitrov:2001:IZA

[DR01] Dimitar K. Dimitrov and André Ronveaux. Inequalities for zeros of associated polynomials and derivatives of orthogonal polynomials. Applied Numerical Mathematics: Transactions of IMACS, 36(2– 3):321–331, February 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.nl../37/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/65/31/37/article.
pdf.

Debrabant:2009:DDI

[DR09a] Kristian Debrabant and Andreas Rößler. Diagonally drift-implicit Runge– Kutta methods of weak order one and two for Itô SDEs and stability analysis. Applied Numerical Mathematics: Transactions of IMACS, 59(3–4):595–607, March/April 2009. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Debrabant:2009:FES

[DR09b] Kristian Debrabant and Andreas Rößler. Families of efficient second order Runge–Kutta methods for the weak approximation of Itô stochastic differential equations. Applied Numerical Mathematics: Transactions of IMACS, 59(3–4): 582–594, March/April 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Draux:1991:CST

[Dra91] André Draux. On composite sequence transformations. Applied Numerical Mathematics: Transactions of IMACS, 7(3):217– 226, March 1991. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Drachman:1997:MCC

[Dra97] Byron Drachman. Minimization of certain cost functions with error bounds. Applied Numerical Mathematics: Transactions of IMACS, 23(3):375–379, May 26, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Dolique:1985:MNS

[DRC85] J.-M. Dolique, J.-R. Roche, and F. Chatelin. Multifluid numerical simulation of secondary plasma and electromagnetic field generated in atmosphere by a high power relativistic electron beam. Applied Numerical Mathematics: Transactions of IMACS, 1(4):325– 337, July 1985. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Djukic:2019:IGA

[DRS19] Dusan Lj. Djukić, Lothar Reichel, and Miodrag M. Spalević. Internality of generalized averaged Gaussian quadrature rules and truncated variants for measures induced by Chebyshev polynomials. Applied Numerical Mathematics: Transactions of IMACS, 142(??): 190–205, August 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419300613.

deKlerk:1999:PDP

[dRT99] E. de Klerk, C. Roos, and T. Terlaky. Primal–dual potential reduction methods for semidefinite programming using affine-scaling directions. Applied Numerical Mathematics: Transactions of IMACS, 29(3):335– 360, March 1, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1999/29/
3/966.pdf.

Das:2020:HOA

[DRVA20] Pratibhamoy Das, Subrata Rana, and Jesus Vigo-Aguiar. Higher order accurate approximations on equidistributed meshes for boundary layer originated mixed type reaction diffusion systems with multiple scale nature. Applied Numerical Mathematics: Transactions of IMACS, 148(??):79– 97, February 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419302375.

Dick:1997:CSS

[DS97a] E. Dick and J. Steelant.

Coupled solution of the steady compressible Navier– Stokes equations and the k-ε turbulence equations with a multigrid method. Applied Numerical Mathematics: Transactions of IMACS, 23(1):49–61, March 7, 1997. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1997/23/

1/743.pdf. Multilevel methods (Oberwolfach, 1995).

Didenko:1997:AMN

[DS97b] V. D. Didenko and B. Silbermann. An approximate method on non-equidistant partitions for double layer potential equation. Applied Numerical Mathematics: Transactions of IMACS, 26(1–2):41–48, December 22, 1997. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1997/26/
                                        1-2/835.pdf.

Duncan:2001:ABD

[DS01] Dugald B. Duncan and Ali R. Soheili. Approximating the Becker–Döring cluster equations. Applied Numerical Mathematics: Transactions of IMACS, 37(1–2):1–29, April 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/10/28/73/26/
                                        27/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/10/28/73/26/27/article.
                                        pdf.

Djouad:2002:PTL

[DS02] Rafik Djouad and Bruno Sportisse. Partitioning techniques and lumping computation for reducing chemical kinetics. APLA: an automatic partitioning and lumping algorithm. Applied Numerical Mathematics: Transactions of IMACS, 43(4):383–398, December 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Djouad:2003:SRC

[DS03] Rafik Djouad and Bruno Sportisse. Solving reduced chemical models in air pollution modelling.

Applied Numerical Mathematics: Transactions of IMACS, 44(1–2):49–61, January 2003. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Debrabant:2005:CRK

[DS05] K. Debrabant and K. Strehmel.

Convergence of Runge– Kutta methods applied to linear partial differentialalgebraic equations. Applied Numerical Mathematics: Transactions of IMACS, 53(2–4):213–229, May 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Debry:2007:NSG

[DS07a] Edouard Debry and Bruno Sportisse. Numerical simulation of the general dynamic equation (GDE) for aerosols with two collocation methods. Applied Numerical Mathematics: Transactions of IMACS, 57(8):885– 898, August 2007. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Debry:2007:SAC

[DS07b] Edouard Debry and Bruno Sportisse. Solving aerosol coagulation with size-binning methods. Applied Numerical Mathematics: Transactions of IMACS, 57(9): 1008–1020, September 2007.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Dominguez:2007:OBF

[DS07c] Víctor Domínguez and Francisco-Javier Sayas. Overlapped BEM–FEM for some Helmholtz transmission problems. Applied Numerical Mathematics: Transactions of IMACS, 57(2):131–146, February 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Draux:2010:BA

[DS10] André Draux and Mohamed Sadik. qd block algorithm. Applied Numerical Mathematics: Transactions of IMACS, 60(12): 1300–1308, December 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Dovgilovich:2015:HAF

[DS15] Leonid Dovgilovich and Ivan Sofronov. High-accuracy finite-difference schemes for solving elastodynamic problems in curvilinear coordinates within multiblock approach. Applied Numerical Mathematics: Transactions of IMACS, 93(??): 176–194, July 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414001068.

Davydov:2017:QSD

[DS17] Oleg Davydov and Abid Saeed. C^{1} quintic splines on domains enclosed by piecewise conics and numerical solution of fully nonlinear elliptic equations. Applied Numerical Mathematics: Transactions of IMACS, 116(??):172–183, June 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301969.

Dehghan:2020:MMP

[DS20] Mehdi Dehghan and Akbar Shirilord. Matrix multisplitting Picard-iterative method for solving generalized absolute value matrix equation. Applied Numerical Mathematics: Transactions of IMACS, 158(??): 425–438, December 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302269.

DAmbrosio:2021:FQS

[DS21a] Raffaele D’Ambrosio and Carmela Scalone. Filon quadrature for stochastic oscillators driven by time-varying forces. Applied Numerical Mathematics: Transactions of IMACS, 169(??):21–31, November 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0168927421001732.

Dastjerdi:2021:NMS

[DS21b] H. Laeli Dastjerdi and F. Shayanfard. A numerical method for the solution of nonlinear Volterra Hammerstein integral equations of the third-kind. Applied Numerical Mathematics: Transactions of IMACS, 170(??):353–363, December 2021. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421002348.

Davydov:2021:MFD

[DS21c] Oleg Davydov and Mansour Safarpoor. A meshless finite difference method for elliptic interface problems based on pivoted QR decomposition. Applied Numerical Mathematics: Transactions of IMACS, 161(??): 489–509, March 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420303688.

Dehghan:2021:HSH

[DS21d] Mehdi Dehghan and Akbar Shirilord. On the Hermitian and skew-Hermitian splitting-like iteration approach for solving complex continuous-time algebraic Riccati matrix equation. Applied Numerical Mathematics: Transactions of IMACS, 170(??): 109–127, December 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001926.

DiGiovacchino:2023:NCI

[DS23] Stefano Di Giovacchino and Carmela Scalone. Numerical conservation issues for jump Pearson diffusions. Applied Numerical Mathematics: Transactions of IMACS, 191(??):55– 61, September 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001320.

DeBonis:2024:NMB

[DS24] Maria Carmela De Bonis and Valeria Sagaria. Numerical method for boundary value problems on the real line. Applied Numerical Mathematics: Transactions of IMACS, 200(??): 179–194, June 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001368.

Dehghan:2020:ASE

[DSA20] Mehdi Dehghan, Nasim Shafieeabyaneh, and Mostafa Abbaszadeh. Application of spectral element method for solving Sobolev equations with error estimation. Applied Numerical Mathematics: Transactions of IMACS, 158(??): 439–462, December 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420302452.

Dizicheh:2020:NAB

[DSAB20] A. Karimi Dizicheh, S. Salahshour, A. Ahmadian, and D. Baleanu. A novel algorithm based on the Legendre wavelets spectral technique for solving the Lane–Emden equations. Applied Numerical Mathematics: Transactions of IMACS, 153(??):443–456, July 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420300623.

deSiqueira:2020:MFE

[dSFDG20] Denise de Siqueira, Agnaldo M. Farias, Phillipe R. B. Devloo, and Sônia M. Gomes. Mixed finite element approximations of a singular elliptic problem based on some anisotropic and hp-adaptive curved quarter-point elements. Applied Numerical Mathematics: Transactions of IMACS, 158(??):85–102, December 2020. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742030221X.

Dixit:2012:SNI

[DSK12] Sandeep Dixit, Om P. Singh, and Sunil Kumar. A stable numerical inversion of generalized Abel’s integral equation. Applied Numerical Mathematics: Transactions of IMACS, 62(5): 567–579, May 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412000104.

Draux:2011:MBI

[DSM11] André Draux, Mohamed Sadik, and Borhane Moalla. Markov–Bernstein inequalities for generalized Gegenbauer weight. Applied Numerical Mathematics: Transactions of IMACS, 61(12): 1301–1321, December 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411001668.

Bonis:2022:NMA

[DSM22] Maria Carmela De Bonis, Marija P. Stanić, and Tatjana V. Tomović Mladenović. Nyström methods for approximating the solutions of an integral equation arising from a problem in mathematical biology. Applied Numerical Mathematics: Transactions of IMACS, 171(??): 193–211, January 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421002543.

Duminil:2015:NSI

[DSS15] Sébastien Duminil, Hassane Sadok, and Daniel B. Szyld. Nonlinear Schwarz iterations with reduced rank extrapolation. Applied Numerical Mathematics: Transactions of IMACS, 94(??):209– 221, August 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415000616.

Dong:2020:NSL

[DSS20] Xu Dong, Xin-Hui Shao, and Hai-Long Shen. A new SOR-like method for solving absolute value equations. Applied Numerical Mathematics: Transactions of IMACS, 156(??): 410–421, October 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420301574.

Donatelli:2013:APN

[DSSC13] M. Donatelli, M. Semplice, and S. Serra-Capizzano. AMG preconditioning for nonlinear degenerate parabolic equations on nonuniform grids with application to monument degradation. Applied Numerical Mathematics: Transactions of IMACS, 68(??):1–18, June 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413000160.

Fies:2013:ACL

[DSV13] Gaspare Da Fies, Alvise Sommariva, and Marco Vianello. Algebraic cubature by linear blending of elliptical arcs. Applied Numerical Mathematics: Transactions of IMACS, 74(??):49– 61, December 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413000974.

Dongarra:1996:CTA

[DSW96] J. J. Dongarra, B. Straughan, and D. W. Walker. Chebyshev tau-QZ algorithm methods for calculating spectra of hydrodynamic stability problems. Applied Numerical Mathematics: Transactions of IMACS, 22(4):399– 434, December 16, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=22&
issue=4&aid=731.

Du:2015:SWEb

[DSZ15a] Jie Du, Chi-Wang Shu, and Mengping Zhang. A simple weighted essentially nonoscillatory limiter for the correction procedure via reconstruction (CPR) framework. Applied Numerical Mathematics: Transactions of IMACS, 95(??): 173–198, September 2015. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414000269.

Du:2015:SWEa

[DSZ15b] Jie Du, Chi-Wang Shu, and Mengping Zhang. A simple weighted essentially nonoscillatory limiter for the correction procedure via reconstruction (CPR) framework on unstructured meshes. Applied Numerical Mathematics: Transactions of IMACS, 90(??):146–167, April 2015. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414001974.

Dutt:1989:SBP

[DT89] P. Dutt and S. Ta’asan. A spline-based parameter estimation technique for static models of elastic structures. Applied Numerical Mathematics: Transactions of IMACS, 5(3):161–175, May 1989. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Diskin:2010:NAF

[DT10] Boris Diskin and James L.

Thomas. Notes on accuracy of finite-volume discretization schemes on irregular grids. Applied Numerical Mathematics: Transactions of IMACS, 60(3): 224–226, March 2010. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Dang:2015:MIS

[DT15] Quang A. Dang and Dinh Hung Tran. Method of infinite systems of equations for solving an elliptic problem in a semistrip. Applied Numerical Mathematics: Transactions of IMACS, 87(??):114– 124, January 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414001561.

DellAccio:2023:UEA

[DTGN23] Francesco Dell’Accio, Filomena Di Tommaso, Allal Guessab, and Federico Nudo. A unified enrichment approach of the standard threenode triangular element. Applied Numerical Mathematics: Transactions of IMACS, 187(??):1–23, May 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423000326.

Du:2020:VRS

[DTQ^{+}20] Zhong Du, Bo Tian, Qi-Xing Qu, Xiao-Yu Wu, and Xue-Hui Zhao. Vector rational and semi-rational rogue waves for the coupled cubic-quintic nonlinear Schrödinger system in a non-Kerr medium. Applied Numerical Mathematics: Transactions of IMACS, 153(??):179–187, July 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300374.

Du:2011:GAD

[Du11]  Kui Du.  GMRES with

adaptively deflated restarting and its performance on an electromagnetic cavity problem. Applied Numerical Mathematics: Transactions of IMACS, 61(9): 977–988, September 2011.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Duffy:1990:UVF

[Duf90] Dean G. Duffy. Unsteady viscous flow around a circular cylinder found by the hopscotch scheme on a vector processing machine. Applied Numerical Mathematics: Transactions of IMACS, 6(3):195–208, March 1990. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Dujardin:2009:ERK

[Duj09] Guillaume Dujardin. Exponential Runge–Kutta methods for the Schrödinger equation. Applied Numerical Mathematics: Transactions of IMACS, 59(8):1839–1857, August 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Dunca:2018:EDV

[Dun18] Argus A. Dunca. Estimates of the discrete van Cittert deconvolution error in approximate deconvolution models of turbulence in bounded domains. Applied Numerical Mathematics: Transactions of IMACS, 134(??):1– 10, December 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418301478.

deSturler:1995:REG

[dv95a] E. de Sturler and H. A. van der Vorst. Reducing the effect of global communication in GMRES(m) and CG on parallel distributed memory computers. Applied Numerical Mathematics: Transactions of IMACS, 18(4): 441–459, November 9, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=18&
issue=4&aid=627.

Dieci:1995:CFL

[DV95b] Luca Dieci and Erik S. Van Vleck. Computation of a few Lyapunov exponents for continuous and discrete dynamical systems. Applied Numerical Mathematics: Transactions of IMACS, 17(3): 275–291, September 4, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=17&
issue=3&aid=581. Numeri-

cal methods for ordinary differential equations (Atlanta, GA, 1994).

Dassi:2020:BMH

[DV20] F. Dassi and G. Vacca.

Bricks for the mixed highorder virtual element method: Projectors and differential operators. Applied Numerical Mathematics: Transactions of IMACS, 155(??): 140–159, ???? 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419300674.

doVal:2002:MVS

[dVA02] João B. R. do Val and Marinho G. Andrade. On mean value solutions for the Helmholtz equation on square grids. Applied Numerical Mathematics: Transactions of IMACS, 41(4): 459–479, July 2002. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Davis:2019:SST

[DvHM19] P. N. Davis, P. van Heijster, and R. Marangell. Spectral stability of travelling wave solutions in a Keller–Segel model. Applied Numerical Mathematics: Transactions of IMACS, 141(??):54–61, ???? 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418301181.

DeMeyer:1993:IPI

[DVV93] H. De Meyer, M. Van Daele, and G. Vanden Berghe. On the implementation of parallel iterated Runge–Kutta methods on a transputer network. Applied Numerical Mathematics: Transactions of IMACS, 13(1–3): 155–163, September 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Sixth Conference on the Numerical Treatment of Differential Equations (Halle, 1992).

Davey:2000:SPC

[DW00] K. Davey and M. J. Ward. A successive preconditioned conjugate gradient method for the minimization of quadratic and nonlinear functions. Applied Numerical Mathematics: Transactions of IMACS, 35(2):129– 156, October 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/64/30/
29/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/64/30/29/article.
pdf.

Deng:2015:TTC

[DW15] Songhai Deng and Zhong Wan. A three-term conjugate gradient algorithm for large-scale unconstrained optimization problems. Applied Numerical Mathematics: Transactions of IMACS, 92(??):70–81, June 2015. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415000185.

Deng:2021:ALE

[DW21] Dingwen Deng and Qiang Wu. Analysis of the linearly energy- and mass-preserving finite difference methods for the coupled Schrödinger– Boussinesq equations. Applied Numerical Mathematics: Transactions of IMACS, 170(??):14–38, December 2021. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742100204X.

Du:2003:DNA

[DY03] Qikui Du and Dehao Yu.

Dirichlet–Neumann alternating algorithm based on the natural boundary reduction for time-dependent problems over an unbounded domain. Applied Numerical Mathematics: Transactions of IMACS, 44(4): 471–486, March 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Daikh:2017:SDN

[DY17] Yasmina Daikh and Driss Yakoubi. Spectral discretization of the Navier– Stokes problem with mixed boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 118(??):33–49, August 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417300387.

Duan:2023:WGF

[DYF23] Mengmeng Duan, Yan Yang, and Minfu Feng. A weak Galerkin finite element method for the Kelvin– Voigt viscoelastic fluid flow model. Applied Numerical Mathematics: Transactions of IMACS, 184(??): 406–430, February 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422002835.

Deng:2009:CAM

[DYX09] Weibing Deng, Xulai Yun, and Chunhong Xie. Convergence analysis of the multiscale method for a class of convection–diffusion equations with highly oscillating coefficients. Applied Numerical Mathematics: Transactions of IMACS, 59(7): 1549–1567, July 2009. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Dong:2020:MEE

[DYZ20] Haixia Dong, Wenjun Ying, and Jiwei Zhang. Maximum error estimates of a MAC scheme for Stokes equations with Dirichlet boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 150(??): 149–163, April 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419302569.

DAngelo:2012:NAN

[DZ12a] Carlo D’Angelo and Paolo Zunino. Numerical approximation with Nitsche’s coupling of transient Stokes’/ Darcy’s flow problems applied to hemodynamics. Applied Numerical Mathematics: Transactions of IMACS, 62(4):378–395, April 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411000481.

Deng:2012:NFO

[DZ12b] Dingwen Deng and Chengjian Zhang. A new fourthorder numerical algorithm for a class of nonlinear wave equations. Applied Numerical Mathematics: Transactions of IMACS, 62(12): 1864–1879, December 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927412001432.

Du:2021:PNS

[DZMB21] Yanyan Du, Qimin Zhang, and Anke Meyer-Baese. The positive numerical solution for stochastic agedependent capital system based on explicit-implicit algorithm. Applied Numerical Mathematics: Transactions of IMACS, 165(??): 198–215, July 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421000659.

Duan:2024:EIM

[DZW24] Xue-Feng Duan, Yong-Shen Zhang, and Qing-Wen Wang. An efficient iterative method for solving a class of constrained tensor least squares problem. Applied Numerical Mathematics: Transactions of IMACS, 196(??): 104–117, February 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423002702.

Dai:2024:LDR

[DZZZ24] Bin Dai, Huilan Zeng, Chen-Song Zhang, and Shuo Zhang. Lowest-degree robust finite element schemes for inhomogeneous bi-Laplace problems. Applied Numerical Mathematics: Transactions of IMACS, 203(??): 235–254, September 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001168.

El-Amrani:2012:CGA

[EAS12] Mofdi El-Amrani and Mohammed Seaïd. A conjugate gradient algorithm for solving the Galerkincharacteristic approximation of interfacial flows. Applied Numerical Mathematics: Transactions of IMACS, 62(9):1197–1214, September 2012. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927410001820.

Ebadi:2016:DAG

[EAV16] G. Ebadi, N. Alipour, and C. Vuik. Deflated and augmented global Krylov subspace methods for the matrix equations. Applied Numerical Mathematics: Transactions of IMACS, 99(??):137– 150, January 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415001336.

ElYassini:2012:IEA

[EB12] Khalid El Yassini and Safae El Haj Ben Ali. An interior-exterior approach for convex quadratic programming. Applied Numerical Mathematics: Transactions of IMACS, 62(9): 1139–1155, September 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927411000511.

El-Barkani:2024:HOS

[EBEHAS24] Imad El-Barkani, Imane El-Hadouti, Mohamed Addam, and Mohammed Seaid. Highorder spline finite element method for solving timedependent electromagnetic waves. Applied Numerical Mathematics: Transactions of IMACS, 206(??):48– 74, December 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424001983.

elbhira:2024:SSO

[ebKMZ24] H. Ait el bhira, M. Kzaz, F. Maach, and J. Zerouaoui. Solving second-order telegraph equations with highfrequency extrinsic oscillations. Applied Numerical Mathematics: Transactions of IMACS, 195(??):89– 104, January 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423002453.

Evans:2007:EIR

[EC07] G. A. Evans and K. C.

Chung. Evaluating infinite range oscillatory integrals using generalised quadrature methods. Applied Numerical Mathematics: Transactions of IMACS, 57(1):73– 79, January 2007. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Engelsone:2007:DTS

[ECB07] A. Engelsone, S. L. Campbell, and J. T. Betts. Direct transcription solution of higher-index optimal control problems and the virtual index. Applied Numerical Mathematics: Transactions of IMACS, 57(3): 281–296, March 2007. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Esmaeilbeigi:2020:LCH

[ECHF^{+}20] M. Esmaeilbeigi, O. Chatrabgoun, A. Hosseinian-Far, R. Montasari, and A. Daneshkhah. A low cost and highly accurate technique for big data spatialtemporal interpolation. Applied Numerical Mathematics: Transactions of IMACS, 153(??):492–502, July 2020.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420300763.

Eslahchi:2020:CSA

[EE20] M. R. Eslahchi and Sakine Esmaili. The convergence and stability analysis of a numerical method for solving a mathematical model of language competition. Applied Numerical Mathematics: Transactions of IMACS, 151(??):119–140, May 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419303629.

Elgharbi:2025:STD

[EEAS25] S. Elgharbi, M. Essaouini, B. Abouzaid, and H. Safouhi. Solving the two-dimensional time-dependent Schrödinger equation using the Sinc collocation method and double exponential transformations. Applied Numerical Mathematics: Transactions of IMACS, 208 (part A) (??):222–231, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424000862.

Erbay:2022:SDN

[EEE22] H. A. Erbay, S. Erbay, and A. Erkip. A semi-discrete numerical scheme for nonlocally regularized KdV-type equations. Applied Numerical Mathematics: Transactions of IMACS, 175(??): 29–39, May 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422000289.

Guide:2022:TKS

[EEJB22] M. El Guide, A. El Ichi, K. Jbilou, and F. P. A. Beik. Tensor Krylov subspace methods via the Einstein product with applications to image and video processing. Applied Numerical Mathematics: Transactions of IMACS, 181(??): 347–363, November 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001684.

Elbarbary:2005:HOP

[EES05] Elsayed M. E. Elbarbary and Salah M. El-Sayed. Higher order pseudospectral differentiation matrices. Applied Numerical Mathematics: Transactions of IMACS, 55(4):425–438, December 2005. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Erath:2009:ENB

[EFLFP09] C. Erath, S. Ferraz-Leite, S. Funken, and D. Praetorius. Energy norm based a posteriori error estimation for boundary element methods in two dimensions. Applied Numerical Mathematics: Transactions of IMACS, 59(11):2713–2734, November 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Engl:1988:PPC

[EG88] Heinz W. Engl and Helmut Gfrerer. A posteriori parameter choice for general regularization methods for solving linear ill-posed problems. Applied Numerical Mathematics: Transactions of IMACS, 4(5):395– 417, July 1988. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Eymard:2001:FVA

[EGH01] R. Eymard, T. Gallouët, and R. Herbin. Finite volume approximation of elliptic problems and convergence of an approximate gradient. Applied Numerical Mathematics: Transactions of IMACS, 37(1–2):31–53, April 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/10/28/73/26/

28/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/73/26/28/article.
pdf.

England:2009:PDA

[EGL09] Roland England, Susana Gómez, and René Lamour. The properties of differential-algebraic equations representing optimal control problems. Applied Numerical Mathematics: Transactions of IMACS, 59(10):2357–2373, October 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Egidi:2025:NDM

[EGM25] Nadaniela Egidi, Josephin Giacomini, and Pierluigi Maponi. Numerical derivation of multivariate functions. Applied Numerical Mathematics: Transactions of IMACS, 210(??): 165–176, April 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424003520.

Elango:2025:CSN

[EGV25] Sekar Elango, L. Govindarao, and R. Vadivel. A comparative study on numerical methods for Fredholm integro-differential equations of convection–diffusion problem with integral boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 207(??): 323–338, January 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424002320.

Engquist:1988:FFB

[EH88] Björn Engquist and Laurence Halpern. Far field boundary conditions for computation over long time. Applied Numerical Mathematics: Transactions of IMACS, 4(1):21–45, March 1988. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Erlebacher:1991:NES

[EH91] Gordon Erlebacher and M. Y. Hussaini. Nonlinear evolution of a second-mode wave in supersonic boundary layers. Applied Numerical Mathematics: Transactions of IMACS, 7(1):73– 91, January 1991. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0168927491901048.

Enright:1997:CRK

[EH97] W. H. Enright and Min Hu. Continuous Runge–Kutta methods for neutral Volterra integro-differential equations with delay. Applied Numerical Mathematics: Transactions of IMACS, 24(2– 3):175–190, June 19, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=24&
issue=2-3&aid=775. Volterra

centennial (Tempe, AZ, 1996).

Eppler:2005:FWB

[EH05] Karsten Eppler and Helmut Harbrecht. Fast wavelet BEM for 3d electromagnetic shaping. Applied Numerical Mathematics: Transactions of IMACS, 54(3– 4):537–554, August 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Eppler:2006:ETS

[EH06] Karsten Eppler and Helmut Harbrecht. Efficient treatment of stationary free boundary problems. Applied Numerical Mathematics: Transactions of IMACS, 56(10–11):1326–1339, October/November 2006. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Efendiev:2007:MFE

[EH07a] Y. Efendiev and T. Hou.

Multiscale finite element methods for porous media flows and their applications. Applied Numerical Mathematics: Transactions of IMACS, 57(5–7): 577–596, May 2007. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ezquerro:2007:HMO

[EH07b] J. A. Ezquerro and M. A. Hernández. Halley’s method for operators with unbounded second derivative. Applied Numerical Mathematics: Transactions of IMACS, 57(3):354–360, March 2007. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Espig:2008:RER

[EH08] M. Espig and W. Hackbusch. On the robustness of elliptic resolvents computed by means of the technique of hierarchical matrices. Applied Numerical Mathematics: Transactions of IMACS, 58(12): 1844–1851, December 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ezquerro:2009:FOI

[EH09] J. A. Ezquerro and M. A. Hernández. Fourth-order iterations for solving Hammerstein integral equations. Applied Numerical Mathematics: Transactions of IMACS, 59(6):1149–1158, June 2009.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Engstrom:2024:MNN

[EH24] Emil Engström and Eskil Hansen. Modified Neumann– Neumann methods for semiand quasilinear elliptic equations. Applied Numerical Mathematics: Transactions of IMACS, 206(??): 322–339, December 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424002071.

El-Hawary:2001:SPS

[EHM01] H. M. El-Hawary and S. M.

Mahmoud. On some 4-point spline collocation methods for solving second-order initial value problems. Applied Numerical Mathematics: Transactions of IMACS, 38(1–2):223–236, July 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/77/27/
38/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/77/27/38/article.
pdf.

El-Halouy:2024:TFM

[EHNR24] Smahane El-Halouy, Silvia Noschese, and Lothar Reichel. A tensor formalism for multilayer network centrality measures using the Einstein product. Applied Numerical Mathematics: Transactions of IMACS, 200(??):236–253, June 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423001782.

Ehrhardt:2008:DTB

[Ehr08] Matthias Ehrhardt. Discrete transparent boundary conditions for Schrödinger-type equations for non-compactly supported initial data. Applied Numerical Mathematics: Transactions of IMACS, 58(5):660–673, May 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ezquerro:2019:DGC

[EHV19] J. A. Ezquerro and M. A. Hernández-Verón. Domains of global convergence for a type of nonlinear Fredholm– Nemytskii integral equations. Applied Numerical Mathematics: Transactions of IMACS, 146(??): 452–468, December 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419302065.

Ezquerro:2024:LSA

[EHV24] J. A. Ezquerro and M. A. Hernández-Verón. Location, separation and approximation of solutions of nonlinear Hammerstein-type integral equations. Applied Numerical Mathematics: Transactions of IMACS, 198(??): 1–10, April 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423003185.

Eijkhout:1995:LDI

[Eij95]  Victor Eijkhout.  A li-

brary of distributed iterative linear system solvers. Applied Numerical Mathematics: Transactions of IMACS, 19(3):359–373, December 15, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=19&
issue=3&aid=639;   http:
//www.sciencedirect.com/
science/article/pii/0168927495000917.

Special issue on iterative methods for linear equations (Atlanta, GA, 1994).

Einkemmer:2018:ASS

[Ein18]  Lukas Einkemmer.  An

adaptive step size controller for iterative implicit methods. Applied Numerical Mathematics: Transactions of IMACS, 132(??): 182–204, October 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418301387.

Eirola:1995:MQF

[Eir95] Timo Eirola. Monotonicity of quadratic forms with symplectic Runge–Kutta methods. Applied Numerical Mathematics: Transactions of IMACS, 17(3):293–298, September 4, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=17&
                                        issue=3&aid=582. Numeri-

cal methods for ordinary differential equations (Atlanta, GA, 1994).

Eirola:1999:DGO

[Eir99] Timo Eirola. Detection of generic one parameter bifurcations of Hamiltonian equilibria. Applied Numerical Mathematics: Transactions of IMACS, 29(1):129–143, January 1, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1999/29/
                                        1/894.pdf.

Eiseman:1986:LRP

[Eis86] Peter R. Eiseman. Local redistribution of points along curves for numerical grid generation. Applied Numerical Mathematics: Transactions of IMACS, 2(3–5):427– 437, October 1986. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hachimi:2025:KSM

[EJR25] Anas El Hachimi, Khalide Jbilou, and Ahmed Ratnani. Krylov subspace methods for large multidimensional eigenvalue computation. Applied Numerical Mathematics: Transactions of IMACS, 208 (part A) (??):205–221, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927424000175.

ElHachimi:2023:SCT

[EJRR23] Anas El Hachimi, Khalide Jbilou, Ahmed Ratnani, and Lothar Reichel. Spectral computation with thirdorder tensors using the tproduct. Applied Numerical Mathematics: Transactions of IMACS, 193(??):1– 21, November 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001964.

ElGuennouni:2004:BLM

[EJS04] A. El Guennouni, K. Jbilou, and H. Sadok. The block Lanczos method for linear systems with multiple righthand sides. Applied Numerical Mathematics: Transactions of IMACS, 51(2– 3):243–256, November 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ervin:2011:CNS

[EJS11] V. J. Ervin, E. W. Jenkins, and S. Sun. Coupling nonlinear Stokes and Darcy flow using mortar finite elements. Applied Numerical Mathematics: Transactions of IMACS, 61(11): 1198–1222, November 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Elnagar:1995:CAC

[EK95] Gamal N. Elnagar and M. A. Kazemi. A cell-averaging Chebyshev spectral method for the controlled Duffing oscillator. Applied Numerical Mathematics: Transactions of IMACS, 18(4): 461–471, November 9, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=18&
                                        issue=4&aid=623.

Efraimsson:1996:NEA

[EK96] Gunilla Efraimsson and Gunilla Kreiss. A note on the effect of artificial viscosity on solutions of conservation laws. Applied Numerical Mathematics: Transactions of IMACS, 21(2):155– 173, July 8, 1996. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=21&
issue=2&aid=688.

Engstler:1997:COS

[EK97] Christian Engstler and Peter Kaps. A comparison of one-step methods for multibody system dynamics in descriptor and state space form. Applied Numerical Mathematics: Transactions of IMACS, 24(4): 457–468, August 12, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=24&
issue=4&aid=796.

Ergenc:2006:PIV

[EK06] T. Ergenç and B. Karasözen.

Poisson integrators for Volterra lattice equations. Applied Numerical Mathematics: Transactions of IMACS, 56(6):879–887, June 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Eptaimeros:2019:IPD

[EKT19] K. G. Eptaimeros, C. Chr. Koutsoumaris, and G. J. Tsamasphyros. Interior penalty discontinuous Galerkin FEMs for a gradient beam and CNTs. Applied Numerical Mathematics: Transactions of IMACS, 144(??): 118–139, October 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927419301357.

Ewing:1994:APP

[EL94] R. E. Ewing and R. D. Lazarov. Approximation of parabolic problems on grids locally refined in time and space. Applied Numerical Mathematics: Transactions of IMACS, 14(1–3): 199–211, April 14, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=14&
issue=1-3&aid=454. Pro-

ceedings of the Third ARO Workshop on Adaptive Methods for Partial Differential Equations (Troy, NY, 1992).

Engstler:1997:MME

[EL97] C. Engstler and C. Lubich.

MUR8: a multirate extension of the eighth-order Dormand–Prince method. Applied Numerical Mathematics: Transactions of IMACS, 25(2–3):185–192, October 17, 1997. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=25&
issue=2-3&aid=813.

Ewing:2001:MAN

[EL01] Richard E. Ewing and Yanping Lin. A mathematical analysis for numerical well models for non-Darcy flows. Applied Numerical Mathematics: Transactions of IMACS, 39(1):17–30, October 2001. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/79/27/28/abstract.
html.

Eichler-Liebenow:1998:LIS

[ELCWS98] C. Eichler-Liebenow, N. H.

Cong, R. Weiner, and K. Strehmel. Linearly implicit splitting methods for higher space-dimensional parabolic differential equations. Applied Numerical Mathematics: Transactions of IMACS, 28(2–4):259–274, October 1, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1998/28/
                                        2-4/913.pdf.

Elgindy:2017:HOS

[Elg17] Kareem T. Elgindy. Highorder, stable, and efficient pseudospectral method using barycentric Gegenbauer quadratures. Applied Numerical Mathematics: Transactions of IMACS, 113(??):1–25, March 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892741630215X.

England:2002:MSU

[ELLE02] Roland England, René Lamour, and Jesús López-Estrada. Multiple shooting using a dichotomically stable integrator for solving differential–algebraic equations. Applied Numerical Mathematics: Transactions of IMACS, 42(1– 3):117–131, August 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Elman:2002:PSP

[Elm02] Howard C. Elman. Preconditioners for saddle point problems arising in computational fluid dynamics. Applied Numerical Mathematics: Transactions of IMACS, 43(1–2):75–89, October 2002. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Esnaola:2015:GRA

[ELR^{+}15] E. Berriochoa Esnaola, A. Cachafeiro López, F. Cala Rodríguez, J. Illán González, and J. M. Rebollido Lorenzo. Gauss rules associated with nearly singular weights. Applied Numerical Mathematics: Transactions of IMACS, 91(??):1–10, May 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414002104.

Eichler-Liebenow:1998:AAF

[ELvdHS98] C. Eichler-Liebenow, P. J.

van der Houwen, and B. P. Sommeijer. Analysis of approximate factorization in iteration methods. Applied Numerical Mathematics: Transactions of IMACS, 28(2–4):245–258, October 1, 1998. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1998/28/
2-4/912.pdf.

Egidi:2005:CNO

[EM05a] N. Egidi and P. Maponi. A class of network optimization methods for planar grid generation. Applied Numerical Mathematics: Transactions of IMACS, 52(4):363–379, March 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Egidi:2005:LCO

[EM05b] N. Egidi and P. Maponi. A linearly constrained optimization problem for planar grid generation. Applied Numerical Mathematics: Transactions of IMACS, 55(3):283–294, November 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Engo:2001:DOO

[EMMK01] Kenth Engø, Arne Marthinsen, and Hans Z. Munthe-Kaas. DiffMan: an objectoriented MATLAB toolbox for solving differential equations on manifolds. Applied Numerical Mathematics: Transactions of IMACS, 39(3–4):323–347, December 2001. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        10/28/79/34/32/abstract.
                                        html.

Eriksson:2009:AOA

[EN09] Sofia Eriksson and Jan Nordström. Analysis of the order of accuracy for node-centered finite volume schemes. Applied Numerical Mathematics: Transactions of IMACS, 59(10): 2659–2676, October 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Engblom:2011:WSS

[Eng11] Stefan Engblom. On wellseparated sets and fast multipole methods. Applied Numerical Mathematics: Transactions of IMACS, 61(10):1096–1102, October 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Enright:2006:SOD

[Enr06] W. H. Enright. Software for ordinary and delay differential equations: Accurate discrete approximate solutions are not enough. Applied Numerical Mathematics: Transactions of IMACS, 56(3–4):459–471, March/April 2006. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Epshteyn:2015:HOD

[EP15] Yekaterina Epshteyn and Spencer Phippen. Highorder difference potentials methods for 1D elliptic type models. Applied Numerical Mathematics: Transactions of IMACS, 93(??): 69–86, July 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414000191.

Epshteyn:2007:FID

[ER07] Y. Epshteyn and B. Rivière. Fully implicit discontinuous finite element methods for two-phase flow. Applied Numerical Mathematics: Transactions of IMACS, 57(4):383–401, April 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Elgindy:2018:HOS

[ER18] Kareem T. Elgindy and Hareth M. Refat. High-order shifted Gegenbauer integral pseudo-spectral method for solving differential equations of lane–Emden type. Applied Numerical Mathematics: Transactions of IMACS, 128(??):98–124, June 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418300278.

Eremin:2019:FCR

[Ere19] Alexey S. Eremin. Functional continuous Runge– Kutta methods with reuse.

Applied Numerical Mathematics: Transactions of IMACS, 146(??):165–181, December 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927419301849.

Edoh:2000:CIT

[ERS00] K. D. Edoh, R. D. Russell, and W. Sun. Computation of invariant tori by orthogonal collocation. Applied Numerical Mathematics: Transactions of IMACS, 32(3):273–289, March 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/59/30/
29/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/59/30/29/article.
pdf.

Erfanifar:2020:NIM

[ESE20] Raziyeh Erfanifar, Khosro Sayevand, and Hamid Esmaeili. A novel iterative method for the solution of a nonlinear matrix equation. Applied Numerical Mathematics: Transactions of IMACS, 153(??): 503–518, July 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300799.

El-Sayed:2010:ANS

[ESEKZ10] A. M. A. El-Sayed, I. L. El-Kalla, and E. A. A. Ziada. Analytical and numerical solutions of multi-term nonlinear fractional orders differential equations. Applied Numerical Mathematics: Transactions of IMACS, 60(8):788–797, August 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Epov:2015:MME

[ESS15] M. I. Epov, E. P. Shurina, and N. V. Shtabel. The mathematical modeling of the electric field in the media with anisotropic objects. Applied Numerical Mathematics: Transactions of IMACS, 93(??):164–175, July 2015. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414001226.

Estep:1995:MED

[Est95] Donald Estep. A modified equation for dispersive difference schemes. Applied Numerical Mathematics: Transactions of IMACS, 17(3):299–309, September 4, 1995. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.

cgi?year=1995&volume=17&
issue=3&aid=583. Numeri-

cal methods for ordinary differential equations (Atlanta, GA, 1994).

Epshteyn:2015:PVR

[EST15] Yekaterina Epshteyn, Ivan Sofronov, and Semyon Tsynkov. Professor V. S. Ryaben’kii. On the occasion of the 90th birthday. Applied Numerical Mathematics: Transactions of IMACS, 93(??): 1–2, July 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415000227.

Elmer:1996:CTW

[EV96] Christopher E. Elmer and Erik S. Van Vleck. Computation of traveling waves for spatially discrete bistable reaction-diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 20(1–2):157–169, February 29, 1996. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=1-2&aid=671. Work-

shop on the method of lines for time-dependent problems (Lexington, KY, 1995).

Evans:1994:TRM

[Eva94] G. A. Evans. Two robust methods for irregular oscillatory integrals over a finite range. Applied Numerical Mathematics: Transactions of IMACS, 14(4): 383–395, June 10, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1994&volume=14&
                                        issue=4&aid=479.

Erlangga:2004:CPS

[EVO04] Y. A. Erlangga, C. Vuik, and C. W. Oosterlee. On a class of preconditioners for solving the Helmholtz equation. Applied Numerical Mathematics: Transactions of IMACS, 50(3–4): 409–425, September 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Erlangga:2006:CMI

[EVO06] Y. A. Erlangga, C. Vuik, and C. W. Oosterlee. Comparison of multigrid and incomplete LU shifted-Laplace preconditioners for the inhomogeneous Helmholtz equation. Applied Numerical Mathematics: Transactions of IMACS, 56(5):648–666, May 2006. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Evans:1997:HOP

[EW97] G. A. Evans and J. R. Webster. A high order, progressive method for the evaluation of irregular oscillatory integrals. Applied Numerical Mathematics: Transactions of IMACS, 23(2):205–218, March 21, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1997/23/
2/740.pdf.

Emet:2008:SDS

[EW08] Stefan Emet and Tapio Westerlund. Solving a dynamic separation problem using MINLP techniques. Applied Numerical Mathematics: Transactions of IMACS, 58(4): 395–406, April 2008. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ewing:1991:ADD

[Ewi91] Richard E. Ewing. Application of domain decomposition techniques in large-scale fluid flow problems. Applied Numerical Mathematics: Transactions of IMACS, 8(4–5):375–388, November 1991. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ewing:1999:SMG

[EWW99] Richard E. Ewing, Junping Wang, and Suzanne L. Weekes. On the simulation of multicomponent gas flow in porous media. Applied Numerical Mathematics: Transactions of IMACS, 31(4):405–427, December 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/46/20/
18/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/46/20/18/article.
pdf.

Engquist:1998:ABC

[EZ98] Bjorn Engquist and Hong-Kai Zhao. Absorbing boundary conditions for domain decomposition. Applied Numerical Mathematics: Transactions of IMACS, 27(4):341–365, August 15, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1998/27/
4/885.pdf.

El-Zoheiry:2003:NIS

[EZ03] H. El-Zoheiry. Numerical investigation for the solitary waves interaction of the “good” Boussinesq equation. Applied Numerical Mathematics: Transactions of IMACS, 45(2–3):161–173, May 2003. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Facius:2003:HAV

[Fac03] Axel Facius. Highly accurate verified error bounds for Krylov type linear system solvers. Applied Numerical Mathematics: Transactions of IMACS, 45(1): 41–58, April 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Fairlie:2000:NSL

[Fai00] R. Fairlie. A numerical study of loading on plate structures due to transient compressible flows. Applied Numerical Mathematics: Transactions of IMACS, 34(4):381–403, August 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/62/32/
31/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/62/32/31/article.
pdf.

Fan:2011:WRM

[Fan11]  Zhencheng Fan.  Wave-

form relaxation method for stochastic differential equations with constant delay. Applied Numerical Mathematics: Transactions of IMACS, 61(2):229–240, February 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Fan:2019:MVI

[Fan19] Lili Fan. Mean velocities in an irrotational equatorial wind wave. Applied Numerical Mathematics: Transactions of IMACS, 141(??): 158–166, ???? 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419300546.

Fata:2010:SAT

[Fat10] S. Nintcheu Fata. Semianalytic treatment of nearlysingular Galerkin surface integrals. Applied Numerical Mathematics: Transactions of IMACS, 60(10):974– 993, October 2010. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Fata:2012:TDI

[Fat12] S. Nintcheu Fata. Treatment of domain integrals in boundary element methods. Applied Numerical Mathematics: Transactions of IMACS, 62(6):720–735, June 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927410001200.

Francisco:2017:NMA

[FBM17] J. B. Francisco, F. S. Viloche Bazán, and M. Weber Mendonça. Non-monotone algorithm for minimization on arbitrary domains with applications to large-scale orthogonal Procrustes problem. Applied Numerical Mathematics: Transactions of IMACS, 112(??):51– 64, February 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301945.

Fernandez-Berdaguer:2009:IPA

[FBS09] Elena M. Fernández-Berdaguer and Gabriela B. Savioli. An inverse problem arising from the displacement of oil by water in porous media. Applied Numerical Mathematics: Transactions of IMACS, 59(10): 2452–2466, October 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Fu:2020:SPA

[FCW20] Yayun Fu, Wenjun Cai, and Yushun Wang. Structurepreserving algorithms for the two-dimensional fractional Klein–Gordon–Schrödinger equation. Applied Numerical Mathematics: Transactions of IMACS, 156(??): 77–93, October 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420301264.

Fu:2021:LIS

[FCW21] Yayun Fu, Wenjun Cai, and Yushun Wang. A linearly implicit structurepreserving scheme for the fractional sine–Gordon equation based on the IEQ approach. Applied Numerical Mathematics: Transactions of IMACS, 160(??): 368–385, February 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420303160.

Feng:2006:CTC

[FCX06] Jian-Hu Feng, Li Cai, and Wen-Xian Xie. CWENOtype central-upwind schemes for multidimensional Saint-Venant system of shallow water equations. Applied Numerical Mathematics: Transactions of IMACS, 56(7):1001–1017, July 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Farrell:1997:EAG

[FD97] Kevin Farrell and Luke O’C. Drury. An explicit, adaptive grid algorithm for onedimensional initial value problems. Applied Numerical Mathematics: Transactions of IMACS, 26(1–2):3–12, December 22, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1997/26/
1-2/832.pdf.

Fambri:2016:SSI

[FD16] Francesco Fambri and Michael Dumbser. Spectral semiimplicit and space–time discontinuous Galerkin methods for the incompressible Navier–Stokes equations on staggered Cartesian grids. Applied Numerical Mathematics: Transactions of IMACS, 110(??):41– 74, December 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301416.

Francomano:2025:SVA

[FDF^{+}25] Elisa Francomano, Stefano De Marchi, Galina Filipuk, Giuliana Ramella, and Federico Zullo. Special volume on Advanced Mathematical and Numerical Models in Applied Sciences (AM-NMAS). Applied Numerical Mathematics: Transactions of IMACS, 210(??): 82, April 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424003441.

Fdil:1996:NMS

[Fdi96] A. Fdil. A new method for solving a nonlinear equation with error estimations. Applied Numerical Mathematics: Transactions of IMACS, 21(4):417– 429, November 22, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=21&
                                        issue=4&aid=703.

Fdil:1997:NTS

[Fdi97a]  A. Fdil.  A new tech-

nique of selection between sequence transformations. Applied Numerical Mathematics: Transactions of IMACS, 25(1):21–40, September 16, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.com/
                                        cgi-bin/cas/tree/store/
                                        apnum/cas_sub/browse/browse.
                                        cgi?year=1997&volume=25&
                                        issue=1&aid=798.

Fdil:1997:SRC

[Fdi97b]  A. Fdil.  Some results

of convergence acceleration for a general Θ-type algorithm. Applied Numerical Mathematics: Transactions of IMACS, 23(2):219–240, March 21, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1997/23/
2/738.pdf.

Filho:2002:CMN

[FdSB02] Hermes Alves Filho, Fernando Carvalho da Silva, and Ricardo C. Barros. A coarsemesh numerical method for one-speed neutron transport eigenvalue problems in twodimensional Cartesian geometry. Applied Numerical Mathematics: Transactions of IMACS, 40(1–2):167–177, January 2002. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/85/27/41/abstract.
html.

Fazio:1993:IDS

[FE93] Riccardo Fazio and David J.

Evans. An implicit difference scheme for a moving boundary hyperbolic problem. Applied Numerical Mathematics: Transactions of IMACS, 12(6):485–496, August 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Feldstein:2006:CAE

[Fel06]  Alan Feldstein.  Conver-

gence and asymptotic error expansion for Euler’s method for variable delay differential equations. Applied Numerical Mathematics: Transactions of IMACS, 56(3–4): 271–283, March/April 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ferreira:1993:SOI

[Fer93] J. A. Ferreira. The secondorder initial/boundary value problem: supraconvergence of regridding methods. Applied Numerical Mathematics: Transactions of IMACS, 13(1–3):69–81, September 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Sixth Conference on the Numerical Treatment of Differential Equations (Halle, 1992).

Ferreira:1996:SCM

[Fer96] J. A. Ferreira. Supraconvergence of a class of moving grid methods for solving a nonlinear problem. Applied Numerical Mathematics: Transactions of IMACS, 21(1):43–56, June 6, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=21&
issue=1&aid=683.

Fermo:2009:NMC

[Fer09] Luisa Fermo. A Nyström method for a class of Fredholm integral equations of the third kind on unbounded domains. Applied Numerical Mathematics: Transactions of IMACS, 59(12): 2970–2989, December 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Fercoq:2014:PVO

[Fer14]  Olivier Fercoq.  Perron

vector optimization applied to search engines. Applied Numerical Mathematics: Transactions of IMACS, 75(??):77–99, January 2014. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413000147.

Falcone:2006:NMV

[FF06] Maurizio Falcone and Roberto Ferretti. Numerical methods for viscosity solutions and applications. Applied Numerical Mathematics: Transactions of IMACS, 56(9):1135, September 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Falcone:2020:NMM

[FF20] M. Falcone and S. Finzi Vita. A new mathematical model for traveling sand dunes: analysis and approximation. Applied Numerical Mathematics: Transactions of IMACS, 155(??): 208–225, ???? 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892741930368X.

Falcone:2013:SID

[FFMZ13] Maurizio Falcone, Roberto Ferretti, Ian M. Mitchell, and Hongkai Zhao. Special issue dedicated to numerical methods for viscosity solutions and applications. Applied Numerical Mathematics: Transactions of IMACS, 73(??):1, November 2013. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927413001062.

Fu:2009:FRS

[FFQ09] Chu-Li Fu, Xiao-Li Feng, and Zhi Qian. The Fourier regularization for solving the Cauchy problem for the Helmholtz equation. Applied Numerical Mathematics: Transactions of IMACS, 59(10):2625–2640, October 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Fatone:2008:CAS

[FFY08] Lorella Fatone, Daniele Funaro, and Gang Joon Yoon. A convergence analysis for the superconsistent Chebyshev method. Applied Numerical Mathematics: Transactions of IMACS, 58(1):88–100, January 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Fairweather:1996:PWM

[FG96] G. Fairweather and I. Gladwell, editors. Proceedings of the 1995 Workshop on the Method of Lines for Time-Dependent Problems, volume 20(1–2) of Applied Numerical Mathematics: Transactions of IMACS. North-Holland Publishing Co., Amsterdam, The Netherlands, February 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ferreira:1998:SEF

[FG98] J. A. Ferreira and R. D.

Grigorieff. On the supraconvergence of elliptic finite difference schemes. Applied Numerical Mathematics: Transactions of IMACS, 28(2–4):275–292, October 1, 1998. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1998/28/
2-4/914.pdf.

Feldstein:2001:NSA

[FG01] Alan Feldstein and Richard H.

Goodman. Number shortening algorithms. Applied Numerical Mathematics: Transactions of IMACS, 36(2–3):249–274, February 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.nl../33/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/10/28/65/31/33/article.
                                        pdf.

Franco:2009:ALS

[FG09] J. M. Franco and I. Gómez. Accuracy and linear stability of RKN methods for solving second-order stiff problems. Applied Numerical Mathematics: Transactions of IMACS, 59(5):959–975, May 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Flagg:2013:AMS

[FG13] Garret M. Flagg and Serkan Gugercin. On the ADI method for the Sylvester equation and the optimal-H_{2} points. Applied Numerical Mathematics: Transactions of IMACS, 64(1):50– 58, February 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892741200178X.

Fu:2022:HOW

[FGGL22] Qingcheng Fu, Zhen Gao, Yaguang Gu, and Peng Li. High order well-balanced conservative finite difference AWENO scheme with hydrostatic reconstruction for the Euler equations under gravitational fields. Applied Numerical Mathematics: Transactions of IMACS, 180(??):1–15, October 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001271.

Ferreira:2023:NSP

[FGP23] J. A. Ferreira, H. P. Gómez, and L. Pinto. A numerical scheme for a partial differential system motivated by light-triggered drug delivery. Applied Numerical Mathematics: Transactions of IMACS, 184(??): 101–120, February 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742200229X.

Fortes:2012:HFS

[FGPR12] M. A. Fortes, P. González, M. Pasadas, and M. L. Rodríguez. Hole filling on surfaces by discrete variational splines. Applied Numerical Mathematics: Transactions of IMACS, 62(9):1050–1060, September 2012. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927411002054.

Farouki:2004:RPG

[FH04] Rida T. Farouki and Chang Yong Han. Robust plotting of generalized lemniscates. Applied Numerical Mathematics: Transactions of IMACS, 51(2–3):257–272, November 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Fischer:2008:MST

[FH08] Rainer Fischer and Thomas Huckle. Multigrid solution techniques for anisotropic structured linear systems. Applied Numerical Mathematics: Transactions of IMACS, 58(4):407–421, April 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Fernandez:2010:PEA

[FH10] J. R. Fernández and P. Hild. A posteriori error analysis for the normal compliance problem. Applied Numerical Mathematics: Transactions of IMACS, 60(1–2):64– 73, January/February 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Fazeli:2020:SDT

[FH20] S. Fazeli and G. Hojjati. Second derivative two-step collocation methods for ordinary differential equations. Applied Numerical Mathematics: Transactions of IMACS, 156(??): 514–527, October 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742030163X.

Fazeli:2022:CTS

[FH22] S. Fazeli and G. Hojjati. A class of twostep collocation methods for Volterra integro-differential equations. Applied Numerical Mathematics: Transactions of IMACS, 181(??):59– 75, November 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001489.

Farago:2005:DMP

[FHK05] I. Faragó, R. Horváth, and S. Korotov. Discrete maximum principle for linear parabolic problems solved on hybrid meshes. Applied Numerical Mathematics: Transactions of IMACS, 53(2–4):249–264, May 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Farrell:2002:ETC

[FHM^{+}02] P. A. Farrell, A. F. Hegarty, J. J. H. Miller, E. O’Riordan, and G. I. Shishkin. An experimental technique for computing parameter-uniform error estimates for numerical solutions of singular perturbation problems, with an application to Prandtl’s problem at high Reynolds number. Applied Numerical Mathematics: Transactions of IMACS, 40(1–2):143–149, January 2002. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/85/27/39/abstract.
html.

Frank:1997:SIE

[FHV97] J. Frank, W. Hundsdorfer, and J. G. Verwer. On the stability of implicit–explicit linear multistep methods. Applied Numerical Mathematics: Transactions of IMACS, 25(2–3):193–205, October 17, 1997. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=25&
issue=2-3&aid=814.

Fu:2022:HOE

[FHX22] Yayun Fu, Dongdong Hu, and Zhuangzhi Xu. Highorder explicit conservative exponential integrator schemes for fractional Hamiltonian PDEs. Applied Numerical Mathematics: Transactions of IMACS, 172(??):315–331, February 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421002968.

Fibich:2003:DEN

[FI03] Gadi Fibich and Boaz Ilan.

Discretization effects in the nonlinear Schrödinger equation. Applied Numerical Mathematics: Transactions of IMACS, 44(1–2):63–75, January 2003. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Fidalgo:2017:CIQ

[Fid17]  U. Fidalgo.  Convergent

interpolatory quadrature schemes. Applied Numerical Mathematics: Transactions of IMACS, 111(??): 111–143, January 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301702.

Fakhar-Izadi:2018:FSC

[FID18a] Farhad Fakhar-Izadi and Mehdi Dehghan. Fully spectral collocation method for nonlinear parabolic partial integro-differential equations. Applied Numerical Mathematics: Transactions of IMACS, 123(??):99– 120, January 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417301800.

Fakhar-Izadi:2018:MSE

[FID18b] Farhad Fakhar-Izadi and Mehdi Dehghan. Modal spectral element method in curvilinear domains. Applied Numerical Mathematics: Transactions of IMACS, 128(??):157–182, June 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418300205.

Fika:2023:ATR

[Fik23] Paraskevi Fika. Approximation of the Tikhonov regularization parameter through Aitken’s extrapolation. Applied Numerical Mathematics: Transactions of IMACS, 190(??): 270–282, August 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423001010.

Filipuk:2025:RIR

[Fil25] Galina Filipuk. Regularisation and iterated regularisation of Hamiltonian systems of the second quasi–Painlevé equation. Applied Numerical Mathematics: Transactions of IMACS, 208 (part B)

(??):290–300, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424002794.

Feldstein:1992:F

[FJ92] Alan Feldstein and Zdzislaw Jackiewicz. Foreword. Applied Numerical Mathematics: Transactions of IMACS, 9(3–5):183–185, April 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0168927492900134.

French:1995:GDD

[FJ95] Donald A. French and Søren Jensen. Global dynamics of a discontinuous Galerkin approximation to a class of reaction–diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 18(4):473–487, November 9, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=18&
issue=4&aid=641.

Feldstein:1997:VCS

[FJ97] A. Feldstein and Z. Jackiewicz. Volterra centennial. selected papers. 2nd international conference on the numerical solution of Volterra and delay equations, May 1996, Tempe, AZ, USA. Applied Numerical Mathematics: Transactions of IMACS, 24(2–3):81–438, 1997. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Fliss:2009:EBC

[FJ09] Sonia Fliss and Patrick Joly. Exact boundary conditions for time-harmonic wave propagation in locally perturbed periodic media. Applied Numerical Mathematics: Transactions of IMACS, 59(9):2155–2178, September 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Famelis:2017:NAC

[FJ17] I. Th. Famelis and Z. Jackiewicz. A new approach to the construction of DIM-SIMs of high order and stage order. Applied Numerical Mathematics: Transactions of IMACS, 119(??):79– 93, September 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417300892.

Friis:2001:UCF

[FJH^{+}01] Helmer André Friis, Tor Arne Johansen, Magne Haveraaen, Hans Munthe-Kaas, and Åsmund Drottning. Use of coordinate-free numerics in elastic wave simulation. Applied Numerical Mathematics: Transactions of IMACS, 39(2):151–171, November 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/79/33/28/abstract.
html.

Figueroa:2021:ETS

[FJL21] Alejandro Figueroa, Zdzisław Jackiewicz, and Rainald Löhner. Efficient twostep Runge–Kutta methods for fluid dynamics simulations. Applied Numerical Mathematics: Transactions of IMACS, 159(??): 1–20, January 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302488.

Ferreira:2017:SOA

[FJP17] J. A. Ferreira, D. Jordão, and L. Pinto. Second order approximations for kinetic and potential energies in Maxwell’s wave equations. Applied Numerical Mathematics: Transactions of IMACS, 120(??): 125–140, October 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417301204.

French:1999:FEA

[FJS99] Donald A. French, Søren Jensen, and Thomas I. Seidman. Finite element approximation of solutions to a class of nonlinear hyperbolicparabolic equations. Applied Numerical Mathematics: Transactions of IMACS, 31(4):429–450, December 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/10/28/46/20/
                                        19/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/10/28/46/20/19/article.
                                        pdf.

Fang:2025:FFG

[FJS25] Yiying Fang, Ying Jiang, and Jiafeng Su. A fast Fourier–Galerkin method for solving boundary integral equations on nonaxisymmetric toroidal surfaces. Applied Numerical Mathematics: Transactions of IMACS, 218(??):73– 90, December 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425001540.

Faheem:2023:WCM

[FK23] Mo Faheem and Arshad Khan. A wavelet collocation method based on Gegenbauer scaling function for solving fourth-order timefractional integro-differential equations with a weakly singular kernel. Applied Numerical Mathematics: Transactions of IMACS, 184(??): 197–218, February 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422002732.

Flouri:2013:AFD

[FKA^{+}13] Evangelia T. Flouri, Nikos Kalligeris, George Alexandrakis, Nikolaos A. Kampanis, and Costas E. Synolakis. Application of a finite difference computational model to the simulation of earthquake generated tsunamis. Applied Numerical Mathematics: Transactions of IMACS, 67(??):111–125, May 2013. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411000973.

Farhat:1993:MPA

[FL93] C. Farhat and M. Lesoinne.

Mesh partitioning algorithms for the parallel solution of partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 12(5): 443–457, July 1993. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/016892749390104Y.

Fournier:2001:MAP

[FL01a] L. Fournier and S. Lanteri. Multiplicative and additive parallel multigrid algorithms for the acceleration of compressible flow computations on unstructured meshes. Applied Numerical Mathematics: Transactions of IMACS, 36(4):401–426, March 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/10/28/65/32/
                                        28/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/10/28/65/32/28/article.
                                        pdf.

Fuhrmann:2001:SES

[FL01b] Jürgen Fuhrmann and Hartmut Langmach. Stability and existence of solutions of time-implicit finite volume schemes for viscous nonlinear conservation laws. Applied Numerical Mathematics: Transactions of IMACS, 37(1–2):201–230, April 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/10/28/73/26/
                                        38/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/10/28/73/26/38/article.
                                        pdf.

Ferm:2004:ASG

[FL04] Lars Ferm and Per Lötstedt.

Accurate and stable grid interfaces for finite volume methods. Applied Numerical Mathematics: Transactions of IMACS, 49(2): 207–224, May 2004. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Farhat:2005:IDD

[FL05] Charbel Farhat and Jing Li.

An iterative domain decomposition method for the solution of a class of indefinite problems in computational structural dynamics. Applied Numerical Mathematics: Transactions of IMACS, 54(2):150–166, July 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ferm:2009:ASM

[FL09] Lars Ferm and Per Lötstedt.

Adaptive solution of the master equation in low dimensions. Applied Numerical Mathematics: Transactions of IMACS, 59(1):187– 204, January 2009. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Fermo:2015:NSB

[FL15] L. Fermo and C. Laurita. On the numerical solution of a boundary integral equation for the exterior Neumann problem on domains with corners. Applied Numerical Mathematics: Transactions of IMACS, 94(??):179– 200, August 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415000598.

Fermo:2020:NMM

[FL20] L. Fermo and C. Laurita. A Nyström method for mixed boundary value problems in domains with corners. Applied Numerical Mathematics: Transactions of IMACS, 149(??):65–82, ???? 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419303058.

Fu:2023:RPC

[FL23] Jun-Liang Fu and Jijun Liu. Recovery of a potential coefficient in a pseudoparabolic system from nonlocal observation. Applied Numerical Mathematics: Transactions of IMACS, 184(??): 121–136, February 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422002549.

Fornaca:2024:PIS

[FL24a] Elena Fornaca and Paola Lamberti. Progressive iterative Schoenberg–Marsden variation diminishing operator and related quadratures. Applied Numerical Mathematics: Transactions of IMACS, 206(??): 269–282, December 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424002101.

Franz:2024:NAS

[FL24b] Sebastian Franz and Kleio Liotati. Numerical analysis of a singularly perturbed 4th order problem with a shift term. Applied Numerical Mathematics: Transactions of IMACS, 201(??):514–530, July 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000886.

Feng:2022:DFE

[FLH22] Xinlong Feng, Xiaoli Lu, and Yinnian He. Difference finite element method for the 3D steady Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 173(??): 418–433, March 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421003445.

Fuhrmann:2011:NMM

[FLL11] Jürgen Fuhrmann, Alexander Linke, and Hartmut Langmach. A numerical method for mass conservative coupling between fluid flow and solute transport. Applied Numerical Mathematics: Transactions of IMACS, 61(4):530–553, April 2011. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Favati:2014:SRI

[FLMR14] P. Favati, G. Lotti, O. Menchi, and F. Romani. Stopping rules for iterative methods in nonnegatively constrained deconvolution. Applied Numerical Mathematics: Transactions of IMACS, 75(??):154–166, January 2014. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0168927413001189.

Flaherty:1997:PSD

[FLÖ^{+}97] J. E. Flaherty, R. M. Loy, C. Özturan, M. S. Shephard, B. K. Szymanski, J. D. Teresco, and L. H. Ziantz. Parallel structures and dynamic load balancing for adaptive finite element computation. Applied Numerical Mathematics: Transactions of IMACS, 26(1–2):241–263, December 22, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1997/26/
1-2/850.pdf.

Franz:2008:SAS

[FLR08] S. Franz, T. Linß, and H.-G. Roos. Superconvergence analysis of the SDFEM for elliptic problems with characteristic layers. Applied Numerical Mathematics: Transactions of IMACS, 58(12):1818–1829, December 2008. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Falcone:1994:SAH

[FLS94] Maurizio Falcone, Piero Lanucara, and Alessandra Seghini. A splitting algorithm for Hamilton–Jacobi– Bellman equations. Applied Numerical Mathematics: Transactions of IMACS, 15(2):207–218, September 13, 1994. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=15&
issue=2&aid=493. Innova-

tive methods in numerical analysis (Bressanone, 1992).

Flaherty:1995:IST

[FM95] Joseph E. Flaherty and Peter K. Moore. Integrated space–time adaptive hp-refinement methods for parabolic systems. Applied Numerical Mathematics: Transactions of IMACS, 16(3):317–341, February 16, 1995. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=16&
                                        issue=3&aid=535.

Frolkovic:2007:FBL

[FM07] Peter Frolkovič and Karol Mikula. Flux-based level set method: a finite volume method for evolving interfaces. Applied Numerical Mathematics: Transactions of IMACS, 57(4): 436–454, April 2007. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Franz:2011:CLA

[FM11] Sebastian Franz and Gunar Matthies. Convergence on layer-adapted meshes and anisotropic interpolation error estimates of nonstandard higher order finite elements. Applied Numerical Mathematics: Transactions of IMACS, 61(6): 723–737, June 2011. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Faghih:2021:NPG

[FM21] A. Faghih and P. Mokhtary.

A novel Petrov–Galerkin method for a class of linear systems of fractional differential equations. Applied Numerical Mathematics: Transactions of IMACS, 169(??):396–414, November 2021. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421002038.

Fish:1994:AMS

[FMGN94] J. Fish, S. Markolefas, R. Guttal, and P. Nayak. On adaptive multilevel superposition of finite element meshes for linear elastostatics. Applied Numerical Mathematics: Transactions of IMACS, 14(1–3): 135–164, April 14, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=14&
issue=1-3&aid=462. Pro-

ceedings of the Third ARO Workshop on Adaptive Methods for Partial Differential Equations (Troy, NY, 1992).

Faltinsen:2001:MMI

[FMMK01] Stig Faltinsen, Arne Marthinsen, and Hans Z. Munthe-Kaas. Multistep methods integrating ordinary differential equations on manifolds. Applied Numerical Mathematics: Transactions of IMACS, 39(3–4):349–365, December 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        10/28/79/34/33/abstract.
                                        html.

Formaggia:2004:AMA

[FMP04] Luca Formaggia, Stefano Micheletti, and Simona Perotto. Anisotropic mesh adaptation in computational fluid dynamics: Application to the advection–diffusion– reaction and the Stokes problems. Applied Numerical Mathematics: Transactions of IMACS, 51(4): 511–533, December 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Fernandez:2024:SOP

[FMPP24] Lidia Fernández, Francisco Marcellán, Teresa E. Pérez, and Miguel A. Piñar. Sobolev orthogonal polynomials and spectral methods in boundary value problems. Applied Numerical Mathematics: Transactions of IMACS, 200(??):

254–272, June 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423002106.

Falletta:2018:DAT

[FMS18] S. Falletta, G. Monegato, and L. Scuderi. On the discretization and application of two space–time boundary integral equations for 3D wave propagation problems in unbounded domains. Applied Numerical Mathematics: Transactions of IMACS, 124(??):22– 43, February 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417302106.

Falini:2024:HBS

[FMS24] Antonella Falini, Francesca Mazzia, and Alessandra Sestini. Hermite–Birkhoff spline Quasi-Interpolation with application as dense output for Gauss–Legendre and Gauss–Lobatto Runge– Kutta schemes. Applied Numerical Mathematics: Transactions of IMACS, 200(??): 273–285, June 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423002064.

Formaggia:2007:MMN

[FMSV07] L. Formaggia, S. Micheletti, R. Sacco, and A. Veneziani. Mathematical modelling and numerical simulation of a glow-plug. Applied Numerical Mathematics: Transactions of IMACS, 57(10): 1125–1144, October 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Frolkovic:2015:SIF

[FMU15] Peter Frolkovic, Karol Mikula, and Jozef Urbán. Semiimplicit finite volume level set method for advective motion of interfaces in normal direction. Applied Numerical Mathematics: Transactions of IMACS, 95(??): 214–228, September 2015. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414001196.

Flemisch:2005:MMC

[FMW05] B. Flemisch, J. M. Melenk, and B. I. Wohlmuth. Mortar methods with curved interfaces. Applied Numerical Mathematics: Transactions of IMACS, 54(3– 4):339–361, August 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Fallahpour:2018:NSF

[FMW18] Merlin Fallahpour, Sean Mc-Kee, and Ewa B. Weinmüller. Numerical simulation of flow in smectic liquid crystals. Applied Numerical Mathematics: Transactions of IMACS, 132(??): 154–162, October 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418301326.

Freund:1995:SSL

[FN95] Roland W. Freund and Noël M. Nachtigal. Software for simplified Lanczos and QMR algorithms. Applied Numerical Mathematics: Transactions of IMACS, 19(3):319–341, December 15, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=19&
issue=3&aid=637;   http:
//www.sciencedirect.com/
science/article/pii/0168927495000895.

Special issue on iterative methods for linear equations (Atlanta, GA, 1994).

Feldstein:2006:SRS

[FNT06] Alan Feldstein, Kenneth W.

Neves, and Skip Thompson. Sharpness results for state dependent delay differential equations: An overview. Applied Numerical Mathematics: Transactions of IMACS, 56(3–4): 472–487, March/April 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Fraguela:2005:IEC

[FOMC05] A. Fraguela, J. Oliveros, M. Morín, and L. Cervantes. Inverse electroencephalography for cortical sources. Applied Numerical Mathematics: Transactions of IMACS, 55(2):191– 203, October 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Forsgren:2002:ICF

[For02] Anders Forsgren. Inertiacontrolling factorizations for optimization algorithms. Applied Numerical Mathematics: Transactions of IMACS, 43(1–2):91–107, October 2002. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Forsyth:2011:HJB

[For11]  Peter A. Forsyth.  A

Hamilton–Jacobi–Bellman approach to optimal trade execution. Applied Numerical Mathematics: Transactions of IMACS, 61(2):241– 265, February 2011. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ford:2025:MMP

[For25] Neville J. Ford. Mathematical modelling of problems with delay and aftereffect. Applied Numerical Mathematics: Transactions of IMACS, 208 (part B) (??):338–347, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424002757.

Fournie:2000:HOC

[Fou00] Michel Fournié. High order conservative difference methods for 2D drift-diffusion model on non-uniform grid. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4):381–392, May 2000. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/29/
17/21/61/27/66/abstract.
html;   http://www.elsevier.
nl/gej-ng/29/17/21/61/
27/66/article.pdf.

Farago:2002:SES

[FP02] I. Faragó and C. Palencia. Sharpening the estimate of the stability constant in the maximum-norm of the Crank–Nicolson scheme for the one-dimensional heat equation. Applied Numerical Mathematics: Transactions of IMACS, 42(1– 3):133–140, August 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Frackiewicz:2024:AFD

[FP24] Mariusz Frackiewicz and Henryk Palus. Application of fractional derivatives in image quality assessment indices. Applied Numerical Mathematics: Transactions of IMACS, 204(??): 101–110, October 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001442.

Forestier:2000:SDW

[FPPS00] M. Y. Forestier, R. Pasquetti, R. Peyret, and C. Sabbah. Spatial development of wakes using a spectral multi-domain method. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4): 207–216, May 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/61/27/
46/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/61/27/46/article.
pdf.

Ferreira:2012:SSV

[FPR12] J. A. Ferreira, L. Pinto, and G. Romanazzi. Supraconvergence and supercloseness in Volterra equations. Applied Numerical Mathematics: Transactions of IMACS, 62(12):1718–1739, December 2012. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741200116X.

Fang:2009:ISM

[FPRA09] Jiannong Fang, Aurèle Parriaux, Martin Rentschler, and Christophe Ancey. Improved SPH methods for simulating free surface flows of viscous fluids. Applied Numerical Mathematics: Transactions of IMACS, 59(2):251– 271, February 2009. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Franzone:2015:PMS

[FPS15] P. Colli Franzone, L. F.

Pavarino, and S. Scacchi. Parallel multilevel solvers for the cardiac electromechanical coupling. Applied Numerical Mathematics: Transactions of IMACS, 95(??):140–153, September 2015. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414001858.

Figueroa:2003:CC

[FPT03] Giovanni Figueroa, Marco Paluszny, and Francisco Tovar. G^{2}-cubic contours. Applied Numerical Mathematics: Transactions of IMACS, 47(3–4):439–448, December 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Frohner:2001:UCD

[FR01] Anja Fröhner and Hans-Görg Roos. The ϵ-uniform convergence of a defect correction method on a Shishkin mesh. Applied Numerical Mathematics: Transactions of IMACS, 37(1–2): 79–94, April 2001. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/10/28/73/26/
                                        30/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/10/28/73/26/30/article.
                                        pdf.

Fortes:2014:NUM

[FR14] M. A. Fortes and M. L. Rodríguez. Non-uniform multiresolution analysis for surfaces and applications. Applied Numerical Mathematics: Transactions of IMACS, 75(??):123–135, January 2014. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413000834.

Filipuk:2018:DEF

[FR18] G. Filipuk and M. N. Rebocho. Differential equations for families of semiclassical orthogonal polynomials within class one. Applied Numerical Mathematics: Transactions of IMACS, 124(??):76–88, February 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417302192.

Franco:2004:RKM

[Fra04a] J. M. Franco. Runge–Kutta methods adapted to the numerical integration of oscillatory problems. Applied Numerical Mathematics: Transactions of IMACS, 50(3–4):427–443, September 2004. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Frank:2004:GST

[Fra04b]  Jason Frank.  Geomet-

ric space–time integration of ferromagnetic materials. Applied Numerical Mathematics: Transactions of IMACS, 48(3–4):307–322, March 2004. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Franco:2006:NMO

[Fra06] J. M. Franco. New methods for oscillatory systems based on ARKN methods. Applied Numerical Mathematics: Transactions of IMACS, 56(8):1040–1053, August 2006. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Franz:2014:SUP

[Fra14] Sebastian Franz. Superconvergence using pointwise interpolation in convectiondiffusion problems. Applied Numerical Mathematics: Transactions of IMACS, 76(??):132–144, February 2014. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0168927413001232.

Franz:2016:AFC

[Fra16] Sebastian Franz. Analysis of a family of continuousdiscontinuous Galerkin FEM for convection-diffusion problems. Applied Numerical Mathematics: Transactions of IMACS, 110(??):93– 109, December 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301453.

Freeman:1991:PUQ

[Fre91] T. L. Freeman. A parallel unconstrained quasi-Newton algorithm and its performance on a local memory parallel computer. Applied Numerical Mathematics: Transactions of IMACS, 7(5):369–378, June 1991. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/016892749190008N.

French:1998:DGF

[Fre98] Donald A. French. Discontinuous Galerkin finite element methods for a forward– backward heat equation. Applied Numerical Mathematics: Transactions of IMACS, 28(1):37–44, September 1, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.com/
cas/tree/store/apnum/sub/
1998/28/1/877.pdf.

French:2004:SEE

[Fre04] Ainslie D. French. Solution of the Euler equations on Cartesian grids. Applied Numerical Mathematics: Transactions of IMACS, 49(3–4):367–379, June 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Fauvet:2010:SPP

[FRRJT10] F. Fauvet, J.-P. Ramis, F. Richard-Jung, and J. Thomann. Stokes phenomenon for the prolate spheroidal wave equation. Applied Numerical Mathematics: Transactions of IMACS, 60(12): 1309–1319, December 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Fenu:2025:WCG

[FRRZ25] C. Fenu, L. Reichel, G. Rodriguez, and Y. Zhang. Weighted chained graphs and some applications. Applied Numerical Mathematics: Transactions of IMACS, 208 (part A)(??):232–245, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423003252.

Felix:2011:KPO

[FRV11] H. M. Felix, A. Sri Ranga, and D. O. Veronese. Kernel polynomials from Lorthogonal polynomials. Applied Numerical Mathematics: Transactions of IMACS, 61(5):651–665, May 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ford:1988:RAV

[FS88a] William F. Ford and Avram Sidi. Recursive algorithms for vector extrapolation methods. Applied Numerical Mathematics: Transactions of IMACS, 4(6):477– 489, November 1988. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Forsyth:1988:QCC

[FS88b] P. A. Forsyth, Jr. and P. H.

Sammon. Quadratic convergence for cell-centered grids. Applied Numerical Mathematics: Transactions of IMACS, 4(5):377– 394, July 1988. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ferracina:2005:SRT

[FS05] L. Ferracina and M. N. Spijker. Stepsize restrictions for total-variation-boundedness in general Runge–Kutta procedures. Applied Numerical Mathematics: Transactions of IMACS, 53(2–4): 265–279, May 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ferracina:2008:SSS

[FS08] L. Ferracina and M. N.

Spijker. Strong stability of singly-diagonally-implicit Runge–Kutta methods. Applied Numerical Mathematics: Transactions of IMACS, 58(11):1675–1686, November 2008. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Falletta:2015:NBE

[FS15] S. Falletta and L. Scuderi. A new boundary element integration strategy for retarded potential boundary integral equations. Applied Numerical Mathematics: Transactions of IMACS, 94(??):106– 126, August 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927415000586.

Fareed:2019:NIP

[FS19] Hiba Fareed and John R. Singler. A note on incremental POD algorithms for continuous time data. Applied Numerical Mathematics: Transactions of IMACS, 144(??):223–233, October 2019. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419301084.

Fan:2023:HON

[FS23a] Haitao Fan and Chi-Wang Shu. High order numerical methods for flows with hysteretic fluxes. Applied Numerical Mathematics: Transactions of IMACS, 184(??):234–252, February 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422002616.

Feng:2023:SPM

[FS23b] Mengya Feng and Tongjun Sun. Stochastic perturbation method for optimal control problem governed by parabolic PDEs with small uncertainties. Applied Numerical Mathematics: Transactions of IMACS, 185(??):483–502, March 2023. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422003300.

Frittelli:2024:MOF

[FS24] Massimo Frittelli and Ivonne Sgura. Matrix-oriented FEM formulation for reactiondiffusion PDEs on a large class of 2D domains. Applied Numerical Mathematics: Transactions of IMACS, 200(??):286–308, June 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001939.

Fish:1997:HCG

[FSB97] J. Fish, A. Suvorov, and V. Belsky. Hierarchical composite grid method for global-local analysis of laminated composite shells. Applied Numerical Mathematics: Transactions of IMACS, 23(2):241–258, March 21, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1997/23/
2/750.pdf.

Frank:1989:CSR

[FSU89] Reinhard Frank, Josef Schneid, and Christoph W. Ueberhuber. B-covariance: a survey. Applied Numerical Mathematics: Transactions of IMACS, 5(1–2):51– 61, February 1989. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Recent theoretical results in numerical ordinary differential equations.

Fu:2019:SCC

[FSWZ19] Hongfei Fu, Yanan Sun, Hong Wang, and Xiangcheng Zheng. Stability and convergence of a Crank–Nicolson finite volume method for space fractional diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 139(??): 38–51, May 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419300078.

Feng:2025:EAS

[FSX25] Jiaying Feng, Changtao Sheng, and Chenglong Xu. Exponentially accurate spectral Monte Carlo method for linear PDEs and their error estimates. Applied Numerical Mathematics: Transactions of IMACS, 217(??): 278–297, November 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425001278.

Feldstein:1996:OUM

[FT96] Alan Feldstein and Peter R.

Turner. Overflow and underflow in multiplication and division. Applied Numerical Mathematics: Transactions of IMACS, 21(3): 221–239, August 20, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=21&
issue=3&aid=692.

Feldstein:2006:GTO

[FT06] Alan Feldstein and Peter R.

Turner. Gradual and tapered overflow and underflow: a functional differential equation and its approximation. Applied Numerical Mathematics: Transactions of IMACS, 56(3–4): 517–532, March/April 2006.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Feehery:1997:ESA

[FTB97] William F. Feehery, John E. Tolsma, and Paul I. Barton. Efficient sensitivity analysis of large-scale differential–algebraic systems. Applied Numerical Mathematics: Transactions of IMACS, 25(1):41– 54, September 16, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=25&
                                        issue=1&aid=805.

Fuhrmann:2001:EUS

[Fuh01] Jürgen Fuhrmann. Existence and uniqueness of solutions of certain systems of algebraic equations with offdiagonal nonlinearity. Applied Numerical Mathematics: Transactions of IMACS, 37(3):359–370, May 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/10/28/73/31/
                                        33/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/10/28/73/31/33/article.
                                        pdf.

Fujino:1999:ECF

[Fuj99]  Seiji Fujino.  Estima-

tion of conflict-free matrix product preconditioner on vector supercomputers. Applied Numerical Mathematics: Transactions of IMACS, 30(2–3):257–266, June 10, 1999. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1999&volume=30&
issue=2-3&aid=979.

Fujino:2002:GHB

[Fuj02] Seiji Fujino. GPBiCG(m, ℓ):

a hybrid of BiCGSTAB and GPBiCG methods with efficiency and robustness. Applied Numerical Mathematics: Transactions of IMACS, 41(1):107–117, April 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/10/28/86/27/34/abstract.
html.

Funaro:1990:CAP

[Fun90] Daniele Funaro. Computational aspects of pseudospectral Laguerre approximations. Applied Numerical Mathematics: Transactions of IMACS, 6(6):447– 457, October 1990. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Funaro:1994:SEA

[Fun94] Daniele Funaro. Spectral elements in the approximation of boundary-value problems in complex geometries. Applied Numerical Mathematics: Transactions of IMACS, 15(2):201– 205, September 13, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1994&volume=15&
                                        issue=2&aid=494. Innova-

tive methods in numerical analysis (Bressanone, 1992).

Funaro:2004:SDI

[Fun04] Daniele Funaro. Superconsistent discretizations of integral type equations. Applied Numerical Mathematics: Transactions of IMACS, 48(1):1–11, January 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Faye:1985:SAP

[FV85] J.-P. Faye and J. Vignes. Stochastic approach of the permutationperturbation method for round-off error analysis. Applied Numerical Mathematics: Transactions of IMACS, 1(4):349–362, July 1985. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Frankel:1987:EMS

[FV87] J. I. Frankel and Brian Vick.

An exact methodology for solving nonlinear diffusion equations based on integral transforms. Applied Numerical Mathematics: Transactions of IMACS, 3(6):467– 477, November 1987. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Frank:1999:PIM

[FV99] J. Frank and C. Vuik. Parallel implementation of a multiblock method with approximate subdomain solution. Applied Numerical Mathematics: Transactions of IMACS, 30(4):403– 423, July 1, 1999. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1999&volume=30&
issue=4&aid=960.

Forsyth:2001:ISU

[FV01] P. A. Forsyth and K. R. Vetzal. Implicit solution of uncertain volatility/transaction cost option pricing models with discretely observed barriers. Applied Numerical Mathematics: Transactions of IMACS, 36(4): 427–445, March 2001. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/10/28/65/32/
                                        29/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/10/28/65/32/29/article.
                                        pdf.

Feng:2005:UAM

[FVB05] Huiyu Feng, Rob Van der Wijngaart, and Rupak Biswas. Unstructured adaptive meshes: bad for your memory? Applied Numerical Mathematics: Transactions of IMACS, 52(2– 3):153–173, February 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Fermo:2017:SDC

[FvdMS17] L. Fermo, C. van der Mee, and S. Seatzu. Scattering data computation for the Zakharov–Shabat system with non-smooth potentials. Applied Numerical Mathematics: Transactions of IMACS, 116(??): 195–203, June 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416301921.

Fermo:2020:NMC

[FvdMS20] Luisa Fermo, Cornelis van der Mee, and Sebastiano Seatzu. A numerical method to compute the scattering solution for the KdV equation. Applied Numerical Mathematics: Transactions of IMACS, 149(??):3–16, ???? 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419301734.

Falcone:2013:SLS

[FVGS13] M. Falcone, S. Finzi Vita, T. Giorgi, and R. G. Smits. A semi-Lagrangian scheme for the game p-Laplacian via p-averaging. Applied Numerical Mathematics: Transactions of IMACS, 73(??):63– 80, November 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412002012.

Fang:2007:NPE

[FW07] Yonglei Fang and Xinyuan Wu. A new pair of explicit ARKN methods for the numerical integration of general perturbed oscillators. Applied Numerical Mathematics: Transactions of IMACS, 57(2):166– 175, February 2007. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Fang:2008:TFE

[FW08] Yonglei Fang and Xinyuan Wu. A trigonometrically fitted explicit Numerov-type method for second-order initial value problems with oscillating solutions. Applied Numerical Mathematics: Transactions of IMACS, 58(3):341–351, March 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Feng:2022:RBC

[FW22] Bo Feng and Gang Wu. Refined bounds on the convergence of block Lanczos method for extended trustregion subproblem. Applied Numerical Mathematics: Transactions of IMACS, 181(??):388–402, November 2022. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742200174X.

Fei:2020:SOI

[FWHM20] Mingfa Fei, Nan Wang, Chengming Huang, and Xiaohua Ma. A second-order implicit difference scheme for the nonlinear time-space fractional Schrödinger equation. Applied Numerical Mathematics: Transactions of IMACS, 153(??): 399–411, July 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300714.

Fang:2018:ESC

[FWL18] Jinwei Fang, Boying Wu, and Wenjie Liu. An explicit spectral collocation method for the linearized Korteweg–de Vries equation on unbounded domain. Applied Numerical Mathematics: Transactions of IMACS, 126(??):34–52, April 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417302532.

Frontin:2021:FST

[FWW^{+}21] Cory V. Frontin, Gage S.

Walters, Freddie D. Witherden, Carl W. Lee, David M. Williams, and David L. Darmofal. Foundations of spacetime finite element methods: Polytopes, interpolation, and integration. Applied Numerical Mathematics: Transactions of IMACS, 166(??): 92–113, August 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000994.

Fu:2021:EEP

[FXCW21] Yayun Fu, Zhuangzhi Xu, Wenjun Cai, and Yushun Wang. An efficient energypreserving method for the two-dimensional fractional Schrödinger equation. Applied Numerical Mathematics: Transactions of IMACS, 165(??):232–247, July 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742100060X.

Feng:2022:UEB

[FXY22] Yue Feng, Zhiguo Xu, and Jia Yin. Uniform error bounds of exponential wave integrator methods for the long-time dynamics of the Dirac equation with small potentials. Applied Numerical Mathematics: Transactions of IMACS, 172(??):50– 66, February 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421002725.

Fardi:2025:FCD

[FZA25] Mojtaba Fardi, Mahmoud A. Zaky, and Babak Azarnavid. Fractional convection–diffusion systems in complex 2D and 3D geometries: a Bernoulli polynomial-based kernel method. Applied Numerical Mathematics: Transactions of IMACS, 218(??): 275–297, December 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425001680.

Fischer:2020:LES

[FZM20] C. Fischer, K. Zourmba, and A. Mohamadou. Lyapunov exponents spectrum estimation of fractional order nonlinear systems using Cloned Dynamics. Applied Numerical Mathematics: Transactions of IMACS, 154(??): 187–204, August 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300945.

Gabriel:2002:MFN

[Gab02]  Richard Gabriel.  Mini-

mization of the Frobenius norm of a complex matrix using planar similarities. Applied Numerical Mathematics: Transactions of IMACS, 40(3):391–414, February 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/85/33/28/abstract.
html.

Ghahremani:2025:ENM

[GAH25] Z. Ghahremani, A. Abdi, and G. Hojjati. Efficient numerical methods based on general linear methods for Volterra integral equations. Applied Numerical Mathematics: Transactions of IMACS, 216(??): 1–16, October 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425000984.

Galeao:2004:FEA

[GAML04] A. C. Galeão, R. C. Almeida, S. M. C. Malta, and A. F. D. Loula. Finite element analysis of convection dominated reaction–diffusion problems. Applied Numerical Mathematics: Transactions of IMACS, 48(2):205– 222, February 2004. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ganesh:1996:GCT

[Gan96] M. Ganesh. A general convergence theory for nonlinear equations with application to integro-differential equations. Applied Numerical Mathematics: Transactions of IMACS, 22(4):435– 449, December 16, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=22&
                                        issue=4&aid=734.

Gan:2009:DMN

[Gan09]  Siqing Gan.  Dissipativ-

ity of θ-methods for nonlinear delay differential equations of neutral type. Applied Numerical Mathematics: Transactions of IMACS, 59(6):1354–1365, June 2009.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Garcia-Archilla:2002:PFS

[GANT02] Bosco García-Archilla, Julia Novo, and Edriss S. Titi. Postprocessing Fourier spectral methods: The case of smooth solutions. Applied Numerical Mathematics: Transactions of IMACS, 43(3):191–209, November 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gumah:2020:SCT

[GAOB20] Ghaleb Gumah, Shrideh Al-Omari, and Dumitru Baleanu. Soft computing technique for a system of fuzzy Volterra integrodifferential equations in a Hilbert space. Applied Numerical Mathematics: Transactions of IMACS, 152(??):310–322, June 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419303289.

Garabedian:1987:ASW

[Gar87]  P. R. Garabedian.  As-

pects of supercritical wing theory. Applied Numerical Mathematics: Transactions of IMACS, 3(1–2): 49–57, May 1987. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0168927487900055.

Garabedian:1992:FUT

[Gar92] P. R. Garabedian. Failure of uniqueness in transonic flow. Applied Numerical Mathematics: Transactions of IMACS, 10(3–4): 231–234, 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Garcia:1996:HOI

[Gar96] Salvador Garcia. Higherorder incremental unknowns, hierarchical basis, and nonlinear dissipative evolutionary equations. Applied Numerical Mathematics: Transactions of IMACS, 19(4):467–494, February 12, 1996. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=19&
                                        issue=4&aid=645.

Garcia:2003:IUG

[Gar03]  Salvador Garcia.  Incre-

mental unknowns and graph techniques in three space dimensions. Applied Numerical Mathematics: Transactions of IMACS, 44(3):329– 365, February 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Garanzha:2005:QIS

[Gar05] V. A. Garanzha. Quasiisometric surface parameterization. Applied Numerical Mathematics: Transactions of IMACS, 55(3): 295–311, November 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Garrappa:2010:OCV

[Gar10]  Roberto Garrappa.  Or-

der conditions for Volterra Runge–Kutta methods. Applied Numerical Mathematics: Transactions of IMACS, 60(5):561–573, May 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gastaldi:1992:DDM

[Gas92] Lucia Gastaldi. A domain decomposition method associated with the streamline diffusion FEM for linear hyperbolic systems. Applied Numerical Mathematics: Transactions of IMACS, 10(5):357–380, October 1992. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gatski:1991:RIF

[Gat91] Thomas B. Gatski. Review of incompressible fluid flow computations using the vorticity-velocity formulation. Applied Numerical Mathematics: Transactions of IMACS, 7(3):227–239, March 1991. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Geranmayeh:2009:TDC

[GAW09] Amir Geranmayeh, Wolfgang Ackermann, and Thomas Weiland. Temporal discretization choices for stable boundary element methods in electromagnetic scattering problems. Applied Numerical Mathematics: Transactions of IMACS, 59(11): 2751–2773, November 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Guo:2024:EAR

[GAX24] Yayu Guo, Mejdi Azaïez, and Chuanju Xu. Error analysis of a reduced order method for the Allen–Cahn equation. Applied Numerical Mathematics: Transactions of IMACS, 203(??): 186–201, September 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000825.

Ghoreyshi:2025:FBM

[GAZD25] Amin Ghoreyshi, Mostafa Abbaszadeh, Mahmoud A. Zaky, and Mehdi Dehghan.

Finite block method for nonlinear time-fractional partial integro-differential equations: Stability, convergence, and numerical analysis. Applied Numerical Mathematics: Transactions of IMACS, 214(??): 82–103, August 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425000534.

Gomez-Bueno:2023:ISI

[GBBC^{+}23] I. Gómez-Bueno, S. Boscarino, M. J. Castro, C. Parés, and G. Russo. Implicit and semi-implicit wellbalanced finite-volume methods for systems of balance laws. Applied Numerical Mathematics: Transactions of IMACS, 184(??):18– 48, February 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422002562.

Gladwell:1997:FSH

[GBDB97] I. Gladwell, K. Bouas-Dockery, and R. W. Brankin. A Fortran 90 separable Hamiltonian system solver. Applied Numerical Mathematics: Transactions of IMACS, 25(2–3):207–217, October 17, 1997. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=25&
                                        issue=2-3&aid=815.

Gorgey:2015:CSE

[GC15] A. Gorgey and R. P. K. Chan. Choice of strategies for extrapolation with symmetrization in the constant stepsize setting. Applied Numerical Mathematics: Transactions of IMACS, 87(??):31–37, January 2015. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414001500.

Gao:2025:MPR

[GC25] Chen-Xiao Gao and Fang Chen. Modified partially randomized extended Kaczmarz method with residual for solving large sparse linear systems. Applied Numerical Mathematics: Transactions of IMACS, 212(??): 215–222, June 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425000285.

Garcia-Celayeta:2006:CMA

[GCHR06] Berta García-Celayeta, Inmaculada Higueras, and Teo Roldán. Contractivity/monotonicity for additive Runge–Kutta methods:

Inner product norms. Applied Numerical Mathematics: Transactions of IMACS, 56(6):862–878, June 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gatica:2012:P

[GCJ^{+}12] Gabriel N. Gatica, L. Pamela Cook, Kirk E. Jordan, Nilima Nigam, Olaf Steinbach, and Liwei Xu. Preface. Applied Numerical Mathematics: Transactions of IMACS, 62(6):665–666, June 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411002108.

Goovaerts:1991:ESD

[GCP91] D. Goovaerts, T. F. Chan, and Robert Piessens. The eigenvalue spectrum of domain decomposed preconditioners. Applied Numerical Mathematics: Transactions of IMACS, 8(4–5): 389–410, November 1991. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gu:2023:WGF

[GCZZ23] Shanshan Gu, Shimin Chai, Chenguang Zhou, and Jinhui Zhou. Weak Galerkin finite element method for linear poroelasticity problems. Applied Numerical Mathematics: Transactions of IMACS, 190(??): 200–219, August 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423001083.

Gheorghiu:2009:SML

[GD09] C. I. Gheorghiu and Florica-Ioana Dragomirescu. Spectral methods in linear stability. applications to thermal convection with variable gravity field. Applied Numerical Mathematics: Transactions of IMACS, 59(6):1290–1302, June 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gharibi:2021:CAW

[GD21] Zeinab Gharibi and Mehdi Dehghan. Convergence analysis of weak Galerkin flux-based mixed finite element method for solving singularly perturbed convection-diffusion-reaction problem. Applied Numerical Mathematics: Transactions of IMACS, 163(??): 303–316, May 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421000386.

Guo:2022:NAL

[GD22] Feng Guo and Weizhong Dai. A new absorbing layer for simulation of wave propagation based on a KdV model on unbounded domain. Applied Numerical Mathematics: Transactions of IMACS, 174(??):46–70, April 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422000071.

Gelu:2023:PUN

[GD23a] Fasika Wondimu Gelu and Gemechis File Duressa. A parameter-uniform numerical method for singularly perturbed Robin type parabolic convection– diffusion turning point problems. Applied Numerical Mathematics: Transactions of IMACS, 190(??): 50–64, August 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001009.

Guo:2023:NAL

[GD23b] Feng Guo and Weizhong Dai. A new absorbing layer approach for solving the nonlinear Schrödinger equation. Applied Numerical Mathematics: Transactions of IMACS, 189(??): 88–106, July 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742300096X.

Garres-Diaz:2023:GVD

[GDEdLD23] J. Garres-Díaz, C. Escalante, T. Morales de Luna, and M. J. Castro Díaz. A general vertical decomposition of Euler equations: Multilayer-moment models. Applied Numerical Mathematics: Transactions of IMACS, 183(??): 236–262, January 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422002288.

Gomes:2015:AMM

[GDS^{+}15] Anna Karina Fontes Gomes, Margarete Oliveira Domingues, Kai Schneider, Odim Mendes, and Ralf Deiterding. An adaptive multiresolution method for ideal magnetohydrodynamics using divergence cleaning with parabolichyperbolic correction. Applied Numerical Mathematics: Transactions of IMACS, 95(??):199–213, September 2015. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927415000173.

Gear:1993:MPA

[Gea93]  C. W. Gear.  Massive

parallelism across space in ODEs. Applied Numerical Mathematics: Transactions of IMACS, 11(1–3): 27–43, January 1993. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Parallel methods for ordinary differential equations (Grado, 1991).

Geba:2024:UWP

[Geb24]  Dan-Andrei Geba.  Un-

conditional well-posedness for the modified Kawahara equation. Applied Numerical Mathematics: Transactions of IMACS, 199(??): 2–25, May 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742300082X.

Garrappa:2020:SDS

[GEGG^{+}20] Roberto Garrappa, Cinzia Elia, Luca Gerardo-Giorda, Alessandro Pugliese, and Giuseppe Vacca. Structural dynamical systems: computational aspects. Applied Numerical Mathematics: Transactions of IMACS, 155(??):1–2, ???? 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742030091X.

Gemignani:2023:EIM

[Gem23]  Luca Gemignani.  Effi-

cient inversion of matrix φ-functions of low order. Applied Numerical Mathematics: Transactions of IMACS, 192(??):57–69, October 2023. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742300154X.

Genseberger:2010:IPP

[Gen10] Menno Genseberger. Improving the parallel performance of a domain decomposition preconditioning technique in the Jacobi– Davidson method for large scale eigenvalue problems. Applied Numerical Mathematics: Transactions of IMACS, 60(11):1083–1099, November 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Goligerdian:2024:NST

[GeO24] Arash Goligerdian and Mahmood Khaksar e Oshagh. The numerical solution of a time-delay model of population growth with immigration using Legendre wavelets. Applied Numerical Mathematics: Transactions of IMACS, 197(??): 243–257, March 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423002970.

Gerasoulis:1994:D

[Ger94] Apostolos Gerasoulis. Dedication. Applied Numerical Mathematics: Transactions of IMACS, 16(1– 2):1–2, December 1, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=16&
issue=1-2&aid=529.

Galan:1999:PSF

[GFB99] Santos Galán, William F.

Feehery, and Paul I. Barton. Parametric sensitivity functions for hybrid discrete/continuous systems. Applied Numerical Mathematics: Transactions of IMACS, 31(1): 17–47, September 1, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1999&volume=31&
issue=1&aid=990.

Grylonakis:2018:CUT

[GFPG18] E. N. G. Grylonakis, C. K.

Filelis-Papadopoulos, and G. A. Gravvanis. A class of unified transform techniques for solving linear elliptic PDEs in convex polygons. Applied Numerical Mathematics: Transactions of IMACS, 129(??): 159–180, July 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418300692.

Gerstberger:1995:COE

[GG95] R. Gerstberger and M. Günther. Charge-oriented extrapolation methods in digital circuit simulation. Applied Numerical Mathematics: Transactions of IMACS, 18(1–3):115–125, September 27, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=18&
issue=1-3&aid=597. Sev-

enth Conference on the Numerical Treatment of Differential Equations (Halle, 1994).

Ghaffari:2019:RSM

[GG19] Rezvan Ghaffari and Farideh Ghoreishi. Reduced spline method based on a proper orthogonal decomposition technique for fractional subdiffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 137(??):62–79, March 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418302630.

Garg:2022:OLP

[GG22] Deepika Garg and Sashikumaar Ganesan. An overlapping local projection stabilization for Galerkin approximations of Stokes and Darcy flow problems. Applied Numerical Mathematics: Transactions of IMACS, 171(??):106–127, January 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421002452.

Guillen-Gonzalez:2016:SAS

[GGG16] F. Guillén-González and J. R. Rodríguez Galván. On the stability of approximations for the Stokes problem using different finite element spaces for each component of the velocity. Applied Numerical Mathematics: Transactions of IMACS, 99(??):51–76, January 2016. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415001099.

Gaspar:2009:GMM

[GGLR09] F. J. Gaspar, J. L. Gracia, F. J. Lisbona, and C. Rodrigo. On geometric multigrid methods for triangular grids using threecoarsening strategy. Applied Numerical Mathematics: Transactions of IMACS, 59(7):1693–1708, July 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ghelardoni:1995:EEP

[GGM95] P. Ghelardoni, G. Gheri, and P. Marzulli. Error estimates for parallel shooting using initial or boundary value methods. Applied Numerical Mathematics: Transactions of IMACS, 18(1–3):127–139, September 27, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=18&
issue=1-3&aid=598. Sev-

enth Conference on the Numerical Treatment of Differential Equations (Halle, 1994).

Giraud:2007:ISP

[GGM07] L. Giraud, S. Gratton, and E. Martin. Incremental spectral preconditioners for sequences of linear systems. Applied Numerical Mathematics: Transactions of IMACS, 57(11– 12):1164–1180, November/ December 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Giannozzi:1988:OMC

[GGMP88] Paolo Giannozzi, Giuseppe Grosso, Saverio Moroni, and Giuseppe Pastori Parravicini. The ordinary and matrix continued fractions in the theoretical analysis of Hermitian and relaxation operators. Applied Numerical Mathematics: Transactions of IMACS, 4(2–4): 273–295, June 1988. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gemmrich:2012:CAM

[GGN12] S. Gemmrich, J. Gopalakrishnan, and N. Nigam. Convergence analysis of a multigrid algorithm for the acoustic single layer equation. Applied Numerical Mathematics: Transactions of IMACS, 62(6):767–786, June 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412000293.

Good:2002:IIE

[GGNP02] Marcel Good, Jean-Pierre Goux, Jorge Nocedal, and Victor Pereyra. iNEOS: an interactive environment for nonlinear optimization. Applied Numerical Mathematics: Transactions of IMACS, 40(1–2):49–57, January 2002. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/85/27/32/abstract.
html.

Giani:2012:BRT

[GGO12] Stefano Giani, Luka Grubisić, and Jeffrey S. Ovall. Benchmark results for testing adaptive finite element eigenvalue procedures. Applied Numerical Mathematics: Transactions of IMACS, 62(2):121–140, February 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0168927411001942.

Guo:2013:AAV

[GGO13] Shimin Guo, Lech A. Grzelak, and Cornelis W. Oosterlee. Analysis of an affine version of the Heston– Hull–White option pricing partial differential equation. Applied Numerical Mathematics: Transactions of IMACS, 72(??):143– 159, October 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741300086X.

Giani:2016:BRT

[GGO16] Stefano Giani, Luka Grubisić, and Jeffrey S. Ovall. Benchmark results for testing adaptive finite element eigenvalue procedures part 2 (conforming eigenvector and eigenvalue estimates). Applied Numerical Mathematics: Transactions of IMACS, 102(??):1–16, April 2016. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415001725.

Grauschopf:1997:AMP

[GGR97] T. Grauschopf, M. Griebel, and H. Regler. Additive multilevel preconditioners based on bilinear interpolation, matrix-dependent geometric coarsening and algebraic multigrid coarsening for second-order elliptic PDEs. Applied Numerical Mathematics: Transactions of IMACS, 23(1):63– 95, March 7, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1997/23/

1/744.pdf. Multilevel methods (Oberwolfach, 1995).

Guillen-Gonzalez:2022:CTF

[GGRBRG22] F. Guillén-González, M. A.

Rodríguez-Bellido, and D. A. Rueda-Gómez. Comparison of two finite element schemes for a chemo-repulsion system with quadratic production. Applied Numerical Mathematics: Transactions of IMACS, 173(??):

193–210, March 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421003366.

Guillen-Gonzalez:2017:CEE

[GGRN17] F. Guillén-González and M. V. Redondo-Neble. Convergence and error estimates of viscosity-splitting finite-element schemes for the primitive equations. Applied Numerical Mathematics: Transactions of IMACS, 111(??):219–245, January 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416301787.

Guilpin:2004:AAD

[GGS04] Christian Guilpin, Jacques Gacougnolle, and Yvan Simon. The ϵ-algorithm allows to detect Dirac delta functions. Applied Numerical Mathematics: Transactions of IMACS, 48(1):27– 40, January 2004. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gu:2016:SEM

[GGS16] Zhendong Gu, Xiaojing Guo, and Daochun Sun. Series expansion method for weakly singular Volterra integral equations. Applied Numerical Mathematics: Transactions of IMACS, 105(??): 112–123, July 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416300034.

Guillen-Gonzalez:2024:ESB

[GGT24] F. Guillén-González and G. Tierra. Energy-stable and boundedness preserving numerical schemes for the Cahn–Hilliard equation with degenerate mobility. Applied Numerical Mathematics: Transactions of IMACS, 196(??):62– 82, February 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742300260X.

Galanis:1991:SCR

[GH91] S. Galanis and A. Hadjidimos. On some convergence results of the kstep iterative methods. Applied Numerical Mathematics: Transactions of IMACS, 7(4):297–308, April 1991. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0168927491900669.

Grossmann:2002:CTS

[GH02] C. Grossmann and Z. Horváth.

Construction of two-sided bounds for initial-boundary value problems. Applied Numerical Mathematics: Transactions of IMACS, 42(1–3):177–187, August 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Guo:2007:MJS

[GH07] Ben-Yu Guo and Wei Huang. Mixed Jacobi-spherical harmonic spectral method for Navier–Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 57(8):939–961, August 2007. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Guo:2020:NPP

[GH20] Ke Guo and Deren Han. Nonsymmetric proximal point algorithm with moving proximal centers for variational inequalities: convergence analysis. Applied Numerical Mathematics: Transactions of IMACS, 147(??): 1–18, January 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419302107.

Gu:2021:INR

[GH21] Ruixue Gu and Bo Han. Inexact Newton regularization in Banach spaces based on two-point gradient method with uniformly convex penalty terms. Applied Numerical Mathematics: Transactions of IMACS, 160(??):122–145, February 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302944.

Ge:2025:NMF

[GH25] Zhihao Ge and Yanan He.

A new multiphysics finite element method for a quasi-static poroelasticity model. Applied Numerical Mathematics: Transactions of IMACS, 209(??): 1–19, March 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424002988.

Ghelardoni:1997:ASL

[Ghe97]  Paolo Ghelardoni.  Ap-

proximations of Sturm– Liouville eigenvalues using boundary value methods. Applied Numerical Mathematics: Transactions of IMACS, 23(3):311–325, April 25, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1997/23/
3/755.pdf.

Gustafsson:2000:IHO

[GHF00] Bertil Gustafsson and Lina Hemmingsson-Frändén. Implicit high-order difference methods and domain decomposition for hyperbolic problems. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4): 493–500, May 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/21/61/27/
                                        78/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/21/61/27/78/article.
                                        pdf.

Georgoulis:2009:DGM

[GHH09] Emmanuil H. Georgoulis, Edward Hall, and Paul Houston. Discontinuous Galerkin methods on hp-anisotropic meshes II: a posteriori error analysis and adaptivity. Applied Numerical Mathematics: Transactions of IMACS, 59(9):2179–2194, September 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ge:2020:LEO

[GHH20] Zhihao Ge, Yanan He, and Yinnian He. A lowest equalorder stabilized mixed finite element method based on multiphysics approach for a poroelasticity model. Applied Numerical Mathematics: Transactions of IMACS, 153(??):1–14, July 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300258.

Gao:2022:TMM

[GHHG22] Shuaibin Gao, Junhao Hu, Jie He, and Qian Guo. The truncated θ-Milstein method for nonautonomous and highly nonlinear stochastic differential delay equations. Applied Numerical Mathematics: Transactions of IMACS, 181(??): 234–254, November 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001635.

Garcke:2016:SLT

[GHK16] Harald Garcke, Michael Hinze, and Christian Kahle. A stable and linear time discretization for a thermodynamically consistent model for two-phase incompressible flow. Applied Numerical Mathematics: Transactions of IMACS, 99(??):151– 171, January 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415001324.

Gavrilyuk:2009:AAB

[GHKM09] I. P. Gavrilyuk, M. Hermann, M. V. Kutniv, and V. L. Makarov. Adaptive algorithms based on exact difference schemes for nonlinear BVPs on the half-axis. Applied Numerical Mathematics: Transactions of IMACS, 59(7):1529–1536, July 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gatica:2006:P

[GHR06] Gabriel N. Gatica, Norbert Heuer, and Rodolfo Rodríguez. Preface. Applied Numerical Mathematics: Transactions of IMACS, 56(10–11):1255, October/ November 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gao:2025:HOA

[GHS25] Yali Gao, Daozhi Han, and Sayantan Sarkar. A highorder accurate unconditionally stable bound-preserving numerical scheme for the Cahn–Hilliard–Navier–Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 217(??):96– 111, November 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425001217.

Grajewski:2005:DWP

[GHT05] Matthias Grajewski, Jaroslav Hron, and Stefan Turek. Dual weighted a posteriori error estimation for a new nonconforming linear finite element on quadrilaterals. Applied Numerical Mathematics: Transactions of IMACS, 54(3– 4):504–518, August 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gunther:2001:RMA

[GHW01] M. Günther, M. Hoschek, and R. Weiner. ROW methods adapted to a cheap Jacobian. Applied Numerical Mathematics: Transactions of IMACS, 37(1–2): 231–240, April 2001. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/73/26/
39/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/73/26/39/article.
pdf.

Gan:2020:TRK

[GHW20] Siqing Gan, Youzi He, and Xiaojie Wang. Tamed Runge–Kutta methods for SDEs with super-linearly growing drift and diffusion coefficients. Applied Numerical Mathematics: Transactions of IMACS, 152(??):

379–402, June 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419303228.

Giachetti:2012:SPD

[Gia12] Riccardo Giachetti. Spectral properties of the Dirac equation in unbounded vector potentials. Applied Numerical Mathematics: Transactions of IMACS, 62(9): 1119–1125, September 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927411000663.

Gill:1988:LPI

[Gil88] John Gill. Limit periodic iteration. Applied Numerical Mathematics: Transactions of IMACS, 4(2–4): 297–308, June 1988. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gill:1991:USC

[Gil91] John Gill. The use of the sequence F_{n}(z) = f_{n} ∘ ⋯ ∘ f_{1}(z) in computing fixed points of continued fractions, products, and series. Applied Numerical Mathematics: Transactions of IMACS, 8(6):469–476, December 1991. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gilewicz:2010:YIB

[Gil10] Jacek Gilewicz. 100 years of improvements of bounding properties of Padé approximants to the Stieltjes functions: One-point, twopoint and N-point Padé approximants. Applied Numerical Mathematics: Transactions of IMACS, 60(12): 1320–1331, December 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gatica:2023:NBS

[GIS23] Gabriel N. Gatica, Cristian Inzunza, and Filánder A. Sequeira. New Banach spaces-based fully-mixed finite element methods for pseudostress-assisted diffusion problems. Applied Numerical Mathematics: Transactions of IMACS, 193(??):148–178, November 2023. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742300199X.

Goeken:2000:RKH

[GJ00] David Goeken and Olin Johnson. Runge–Kutta with higher order derivative approximations. Applied Numerical Mathematics: Transactions of IMACS, 34(2–3):207–218, July 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/21/62/29/
                                        32/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/21/62/29/32/article.
                                        pdf.

Guo:2017:RWF

[GJ17] Jingyang Guo and Jae-Hun Jung. A RBF– WENO finite volume method for hyperbolic conservation laws with the monotone polynomial interpolation method. Applied Numerical Mathematics: Transactions of IMACS, 112(??):27– 50, February 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416301970.

Gjesdal:2007:IEM

[Gje07]  Thor Gjesdal.  Implicit–

explicit methods based on strong stability preserving multistep time discretizations. Applied Numerical Mathematics: Transactions of IMACS, 57(8):911– 919, August 2007. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Guo:2020:NAC

[GJLL20] Yayu Guo, Hongen Jia, Jichun Li, and Ming Li. Numerical analysis for the Cahn–Hilliard–Hele–Shaw system with variable mobility and logarithmic Flory– Huggins potential. Applied Numerical Mathematics: Transactions of IMACS, 150(??):206–221, April 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419302533.

Gong:2023:MBP

[GJlL23] Yuezheng Gong, Bingquan Ji, and Hong lin Liao. A maximum bound principle preserving iteration technique for a class of semilinear parabolic equations. Applied Numerical Mathematics: Transactions of IMACS, 184(??): 482–495, February 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422002896.

Guevara-Jordan:2003:MFS

[GJR03] J. M. Guevara-Jordan and S. Rojas. A method of fundamental solutions for modeling porous media advective fluid flow. Applied Numerical Mathematics: Transactions of IMACS, 47(3– 4):449–465, December 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Guillard:2008:AAP

[GJV08] Hervé Guillard, Aleš Janka, and Petr Vaněk. Analysis of an algebraic Petrov– Galerkin smoothed aggregation multigrid method. Applied Numerical Mathematics: Transactions of IMACS, 58(12):1861–1874, December 2008. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gohberg:1993:CNF

[GK93] I. Gohberg and I. Koltracht. Condition numbers for functions of matrices. Applied Numerical Mathematics: Transactions of IMACS, 12(1–3):107–117, May 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0168927493901147.

Special issue to honor Professor Saul Abarbanel on his sixtieth birthday (Neveh, 1992).

Geiser:2009:DDM

[GK09] Jürgen Geiser and Christos Kravvaritis. A domain decomposition method based on the iterative operator splitting method. Applied Numerical Mathematics: Transactions of IMACS, 59(3–4):608–623, March/ April 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Giuliani:2019:OCN

[GK19] Andrew Giuliani and Lilia Krivodonova. On the optimal CFL number of SSP methods for hyperbolic problems. Applied Numerical Mathematics: Transactions of IMACS, 135(??): 165–172, January 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418301910.

Gyulov:2022:PMI

[GK22] Tihomir B. Gyulov and Miglena N. Koleva. Penalty method for indifference pricing of American option in a liquidity switching market. Applied Numerical Mathematics: Transactions of IMACS, 172(??):525– 545, February 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421003093.

Grammont:2017:PIT

[GKA17] L. Grammont, H. Kaboul, and M. Ahues. A product integration type method for solving nonlinear integral equations in L^{1}. Applied Numerical Mathematics: Transactions of IMACS, 114(??):38–46, April 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416302318.

Gentili:2022:EMM

[GKB^{+}22] G. G. Gentili, M. Khosronejad, G. Bernasconi, S. Perotto, and S. Micheletti. Efficient modeling of multimode guided acoustic wave propagation in deformed pipelines by hierarchical model reduction. Applied Numerical Mathematics: Transactions of IMACS, 173(??): 329–344, March 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421003421.

Ghosh:2021:SAM

[GKKM21] Surath Ghosh, Snehasis Kundu, Sunil Kumar, and Emad E. Mahmoud. Spectral approximation methods for non equilibrium transport in turbulent channel flows using fADE. Applied Numerical Mathematics: Transactions of IMACS, 162(??):53–66, April 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303652.

Gremaud:2007:SNM

[GKL07] Pierre A. Gremaud, Christopher M. Kuster, and Zhilin Li. A study of numerical methods for the level set approach. Applied Numerical Mathematics: Transactions of IMACS, 57(5–7): 837–846, May 2007. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Georgiev:2009:LOM

[GKMS09] I. Georgiev, J. Kraus, S. Margenov, and J. Schicho. Locally optimized MIC(0) preconditioning of Rannacher–Turek FEM systems. Applied Numerical Mathematics: Transactions of IMACS, 59(10): 2402–2415, October 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gutknecht:2004:P

[GKS04] Martin H. Gutknecht, Erricos John Kontoghiorghes, and Valeria Simoncini. Preface. Applied Numerical Mathematics: Transactions of IMACS, 49(1):1, April 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gillis:2020:NAN

[GKS20] Nicolas Gillis, Michael Karow, and Punit Sharma. A note on approximating the nearest stable discrete-time descriptor systems with fixed rank. Applied Numerical Mathematics: Transactions of IMACS, 148(??): 131–139, February 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419302430.

Grajewski:2010:NAI

[GKT10] Matthias Grajewski, Michael Köster, and Stefan Turek. Numerical analysis and implementational aspects of a new multilevel grid deformation method. Applied Numerical Mathematics: Transactions of IMACS, 60(8):767–781, August 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gustafsson:1993:GMI

[GL93] Bertil Gustafsson and Per Lötstedt. The GMRES method improved by securing fast wave propagation. Applied Numerical Mathematics: Transactions of IMACS, 12(1–3): 135–152, May 1993. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0168927493901169.

Special issue to honor Professor Saul Abarbanel on his sixtieth birthday (Neveh, 1992).

Grune:2017:DSD

[GL17] Lars Grüne and Thuy T. T.

Le. A double-sided dynamic programming approach to the minimum time problem and its numerical approximation. Applied Numerical Mathematics: Transactions of IMACS, 121(??):68– 81, November 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417301447.

Glaister:1993:ENM

[Gla93] P. Glaister. An efficient numerical method for subcritical and supercritical open channel flows. Applied Numerical Mathematics: Transactions of IMACS, 11(6):497–508, April 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Glaister:1994:EFD

[Gla94] P. Glaister. An efficient flux difference splitting algorithm for unsteady duct flows of a real gas. Applied Numerical Mathematics: Transactions of IMACS, 15(1):27–52, August 1, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-

                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1994&volume=15&
                                        issue=1&aid=503.

Gao:2014:PCH

[GLLW14] Wei Gao, Hong Li, Yang Liu, and XiaoXi Wei. A Padé compact high-order finite volume scheme for nonlinear Schrödinger equations. Applied Numerical Mathematics: Transactions of IMACS, 85(??): 115–127, November 2014. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414001317.

Gong:2025:SCP

[GLLY25] Rongfang Gong, Xiaohui Liu, Catharine W. K. Lo, and Gaocheng Yue. A study of Cauchy problem of the Helmholtz equation based on a relaxation model: Regularization and analysis. Applied Numerical Mathematics: Transactions of IMACS, 216(??): 140–163, October 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425001059.

Gorbunov:2009:PMN

[GLM09] Vladimir K. Gorbunov, Igor V. Lutoshkin, and Yuliya V. Martynenko. A parametrization method for the numerical solution of singular differential equations. Applied Numerical Mathematics: Transactions of IMACS, 59(3–4): 639–655, March/April 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Guo:2018:NPT

[GLM18] Qian Guo, Wei Liu, and Xuerong Mao. A note on the partially truncated Euler–Maruyama method. Applied Numerical Mathematics: Transactions of IMACS, 130(??):157–170, August 2018. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741830093X.

Gao:2020:SOD

[GLML20] Yali Gao, Rui Li, Liquan Mei, and Yanping Lin. A second-order decoupled energy stable numerical scheme for Cahn– Hilliard-Hele–Shaw system. Applied Numerical Mathematics: Transactions of IMACS, 157(??):338–355, November 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420301926.

Guo:2017:PTE

[GLMY17] Qian Guo, Wei Liu, Xuerong Mao, and Rongxian Yue. The partially truncated Euler–Maruyama method and its stability and boundedness. Applied Numerical Mathematics: Transactions of IMACS, 115(??): 235–251, May 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417300259.

Gerisch:2009:HOL

[GLPW09] A. Gerisch, J. Lang, H. Podhaisky, and R. Weiner. Highorder linearly implicit twostep peer – finite element methods for time-dependent PDEs. Applied Numerical Mathematics: Transactions of IMACS, 59(3–4): 624–638, March/April 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Guo:2026:CDM

[GLQ26] Tao Guo, Yiqun Li, and Wenlin Qiu. Compact difference method for Euler– Bernoulli beams and plates with nonlinear nonlocal damping. Applied Numerical Mathematics: Transactions of IMACS, 219(??): 1–18, January 2026. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425001722.

Gambino:2009:VDM

[GLS09] G. Gambino, M. C. Lombardo, and M. Sammartino. A velocity–diffusion method for a Lotka–Volterra system with nonlinear cross and selfdiffusion. Applied Numerical Mathematics: Transactions of IMACS, 59(5): 1059–1074, May 2009. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gaspar:2003:FDA

[GLV03] F. J. Gaspar, F. J. Lisbona, and P. N. Vabishchevich. A finite difference analysis of Biot’s consolidation model. Applied Numerical Mathematics: Transactions of IMACS, 44(4):487– 506, March 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gaspar:2006:SGD

[GLV06] F. J. Gaspar, F. J. Lisbona, and P. N. Vabishchevich. Staggered grid discretizations for the quasi-static Biot’s consolidation problem. Applied Numerical Mathematics: Transactions of IMACS, 56(6):888–898, June 2006. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Guo:2025:EAB

[GLW^{+}25] Jie Guo, Ting Li, Zhong Wan, Jiaoyan Li, and Yamei Xiao. Efficient algorithms of box-constrained Nonnegative Matrix Factorization and its applications in image clustering. Applied Numerical Mathematics: Transactions of IMACS, 208 (part B) (??):176–188, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424002903.

Galeone:1985:SNS

[GM85] L. Galeone and C. Mastroserio. Stability of the numerical solution of the periodic reaction diffusion systems. Applied Numerical Mathematics: Transactions of IMACS, 1(4):315–324, July 1985. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gasca:1987:GSC

[GM87] M. Gasca and G. Mühlbach. Generalized Schur-complements and a test for total positivity. Applied Numerical Mathematics: Transactions of IMACS, 3(3): 215–232, June 1987. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0168927487900493.

Gasparo:1993:PQN

[GM93] M. Grazia Gasparo and M. Macconi. Parallel quasi-Newton techniques with applications to shooting methods. Applied Numerical Mathematics: Transactions of IMACS, 11(1–3):159–167, January 1993. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Parallel methods for ordinary differential equations (Grado, 1991).

Graves-Morris:1994:RPM

[GM94] P. R. Graves-Morris. A review of Padé methods for the acceleration of convergence of a sequence of vectors. Applied Numerical Mathematics: Transactions of IMACS, 15(2):153– 174, September 13, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=15&
issue=2&aid=495. Innova-

tive methods in numerical analysis (Bressanone, 1992).

Ghelardoni:1995:SSB

[GM95] Paolo Ghelardoni and Pietro Marzulli. Stability of some boundary value methods for IVPs. Applied Numerical Mathematics: Transactions of IMACS, 18(1–3):141–153, September 27, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=18&
issue=1-3&aid=596. Sev-

enth Conference on the Numerical Treatment of Differential Equations (Halle, 1994).

Gao:2008:SGD

[GM08] Zhiming Gao and Yichen Ma. Shape gradient of the dissipated energy functional in shape optimization for the viscous incompressible flow. Applied Numerical Mathematics: Transactions of IMACS, 58(11): 1720–1741, November 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gao:2010:NSF

[GM10] Zhiming Gao and Yichen Ma. A new stabilized finite element method for shape optimization in the steady Navier–Stokes flow. Applied Numerical Mathematics: Transactions of IMACS, 60(8):816–832, August 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gao:2016:IEM

[GM16] Yali Gao and Liquan Mei.

Implicit-explicit multistep methods for general twodimensional nonlinear Schrödinger equations. Applied Numerical Mathematics: Transactions of IMACS, 109(??):41– 60, November 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416300976.

Goyal:2017:AMS

[GM17] Kavita Goyal and Mani Mehra. An adaptive meshfree spectral graph wavelet method for partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 113(??):168–185, March 2017. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416302367.

Ge:2018:MIS

[GM18] Zhihao Ge and Mengxia Ma.

Multirate iterative scheme based on multiphysics discontinuous Galerkin method for a poroelasticity model. Applied Numerical Mathematics: Transactions of IMACS, 128(??):125–138, June 2018. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418300424.

Garcia:2002:NMO

[GMG02] Amelia García, Pablo Martín, and Ana B. González. New methods for oscillatory problems based on classical codes. Applied Numerical Mathematics: Transactions of IMACS, 42(1–3):141–157, August 2002. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Garcia:2004:HOR

[GMG04] Amelia García, Pablo Martín, and Ana B. González. High order Runge–Kutta– Nyström codes for the integration of oscillatory problems. Applied Numerical Mathematics: Transactions of IMACS, 48(1):13–25, January 2004. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ghanbari:2019:NSC

[GMG19] F. Ghanbari, P. Mokhtary, and K. Ghanbari. On the numerical solution of a class of linear fractional integrodifferential algebraic equations with weakly singular kernels. Applied Numerical Mathematics: Transactions of IMACS, 144(??): 1–20, October 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927419301175.

Gonzalez:2002:IAN

[GMGF02] Ana B. González, Pablo Martín, Amelia García, and José M. Farto. Increasing the accuracy in the numerical integration of perturbed oscillators with new methods. Applied Numerical Mathematics: Transactions of IMACS, 41(2): 295–304, May 2002. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gonzalez:2000:NIO

[GML00] Ana B. González, Pablo Martín, and David J. López. On the numerical integration of orbital problems with high order Runge–Kutta– Nyström methods. Applied Numerical Mathematics: Transactions of IMACS, 35(1):1–10, September 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/64/27/
27/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/64/27/27/article.
pdf.

Gatica:2009:ACB

[GMM09] Gabriel N. Gatica, Antonio Márquez, and Salim Meddahi. Analysis of the coupling of BEM, FEM and mixed-FEM for a twodimensional fluid–solid interaction problem. Applied Numerical Mathematics: Transactions of IMACS, 59(11):2735–2750, November 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Glafke:2012:CFB

[GMS12] Matthias Gläfke, Matthias Maischak, and Ernst P. Stephan. Coupling of FEM and BEM for a transmission problem with nonlinear interface conditions. Hierarchical and residual error indicators. Applied Numerical Mathematics: Transactions of IMACS, 62(6): 736–753, June 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927410002084.

Gao:2008:SOS

[GMZ08] Zhiming Gao, Yichen Ma, and Hongwei Zhuang. Shape optimization for Stokes flow. Applied Numerical Mathematics: Transactions of IMACS, 58(6):827–844, June 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Garau:2011:CAK

[GMZ11] Eduardo M. Garau, Pedro Morin, and Carlos Zuppa. Convergence of an adaptive Kačanov FEM for quasi-linear problems. Applied Numerical Mathematics: Transactions of IMACS, 61(4):512–529, April 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gunzburger:1986:SAS

[GN86] M. D. Gunzburger and R. A.

Nicolaides. On substructuring algorithms and solution techniques for the numerical approximation of partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 2(3–5):243–256, October 1986. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gumah:2020:NSH

[GNAS^{+}20] G. Gumah, M. F. M. Naser, M. Al-Smadi, S. K. Q. Al-Omari, and D. Baleanu. Numerical solutions of hybrid fuzzy differential equations in a Hilbert space. Applied Numerical Mathematics: Transactions of IMACS, 151(??):402–412, May 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742030009X.

Gunarathna:2019:EFH

[GND19] W. A. Gunarathna, H. M. Nasir, and W. B. Daundasekera. An explicit form for higher order approximations of fractional derivatives. Applied Numerical Mathematics: Transactions of IMACS, 143(??):51– 60, September 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419300704.

Gazzola:2019:ADG

[GNNR19] Silvia Gazzola, Silvia Noschese, Paolo Novati, and Lothar Reichel. Arnoldi decomposition, GMRES, and preconditioning for linear discrete ill-posed problems. Applied Numerical Mathematics: Transactions of IMACS, 142(??):102–121, August 2019. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419300479.

Garcia:2024:NSH

[GNUV24] A. García, M. Negreanu, F. Ureña, and A. M. Vargas. Numerical solution of a hydrodynamic model with cavitation using finite difference method at arbitrary meshes. Applied Numerical Mathematics: Transactions of IMACS, 205(??):

195–205, November 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001855.

Grigori:2019:SDF

[GNX19] Laura Grigori, Qiang Niu, and Yingxiang Xu. Stabilized dimensional factorization preconditioner for solving incompressible Navier– Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 146(??):309–327, December 2019. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927419301412.

Gomez:2021:DRA

[GNZ21] Alvaro Almeida Gomez, Antônio J. Silva Neto, and Jorge P. Zubelli. Diffusion representation for asymmetric kernels. Applied Numerical Mathematics: Transactions of IMACS, 166(??): 208–226, August 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742100101X.

Goncalves:2018:INL

[GO18] M. L. N. Gonçalves and F. R. Oliveira. An inexact Newton-like conditional gradient method for constrained nonlinear systems. Applied Numerical Mathematics: Transactions of IMACS, 132(??): 22–34, October 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418301168.

Gracia:2019:PUN

[GO19] J. L. Gracia and E. O’Riordan. Parameter-uniform numerical methods for singularly perturbed parabolic problems with incompatible boundary-initial data. Applied Numerical Mathematics: Transactions of IMACS, 146(??):436–451, December 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927419302077.

Glaubitz:2020:SDH

[GÖ20] Jan Glaubitz and Philipp Öffner. Stable discretisations of high-order discontinuous Galerkin methods on equidistant and scattered points. Applied Numerical Mathematics: Transactions of IMACS, 151(??): 98–118, May 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741930371X.

Gracia:2021:PUA

[GO21] Jose Luis Gracia and Eugene O’Riordan. Parameteruniform approximations for a singularly perturbed convection– diffusion problem with a discontinuous initial condition. Applied Numerical Mathematics: Transactions of IMACS, 162(??): 106–123, April 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742030386X.

Gracia:2023:SPC

[GO23] J. L. Gracia and E. O’Riordan.

A singularly perturbed convection– diffusion parabolic problem with incompatible boundary/initial data. Applied Numerical Mathematics: Transactions of IMACS, 190(??):168–186, August 2023. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001046.

Gomez:2003:RCT

[GOGF03] Susana Gomez, Michiyo Ono, Carlos Gamio, and Andres Fraguela. Reconstruction of capacitance tomography images of simulated two-phase flow regimes. Applied Numerical Mathematics: Transactions of IMACS, 46(2):197–208, August 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Goldstein:1986:WEM

[Gol86] Charles I. Goldstein. The weak element method applied to Helmholtz type equations. Applied Numerical Mathematics: Transactions of IMACS, 2(3–5):409– 426, October 1986. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Goldberg:2000:SCF

[Gol00] Moshe Goldberg. Stability criteria for finite difference approximations to parabolic systems. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4): 509–515, May 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/21/61/27/
                                        80/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/21/61/27/80/article.
                                        pdf.

Gu:2025:NSW

[GOLJ25] Jiaxi Gu, Daniel Olmos-Liceaga, and Jae-Hun Jung. A numerical study of WENO approximations to sharp propagating fronts for reaction– diffusion systems. Applied Numerical Mathematics: Transactions of IMACS, 211(??):1–16, May 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742400357X.

Gonzalez:2006:FDF

[Gon06] María González. Fully discrete FEM–BEM method for a class of exterior nonlinear parabolic–elliptic problems in 2D. Applied Numerical Mathematics: Transactions of IMACS, 56(10–11):1340– 1355, October/November 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Goodman:1990:SMR

[Goo90] Richard H. Goodman. Some models of relative error in products. Applied Numerical Mathematics: Transactions of IMACS, 6(3):209– 220, March 1990. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gracia:2006:PRS

[GOP06] J. L. Gracia, E. O’Riordan, and M. L. Pickett. A parameter robust second order numerical method for a singularly perturbed twoparameter problem. Applied Numerical Mathematics: Transactions of IMACS, 56(7):962–980, July 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gimperlein:2020:TDF

[GÖS20] Heiko Gimperlein, Ceyhun Özdemir, and Ernst P. Stephan. A time-dependent FEM-BEM coupling method for fluid-structure interaction in 3 d. Applied Numerical Mathematics: Transactions of IMACS, 152(??): 49–65, June 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420300246.

Gasca:1993:SPG

[GP93] M. Gasca and J. M. Peña. Scaled pivoting in Gauss and Neville elimination for totally positive systems. Applied Numerical Mathematics: Transactions of IMACS, 13(5):345–355, December 1993. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Givoli:1998:OLN

[GP98] Dan Givoli and Igor Patlashenko. Optimal local nonreflecting boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 27(4):367–384, August 15, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1998/27/
4/886.pdf.

Gerritsma:2000:CAS

[GP00] M. I. Gerritsma and T. N.

Phillips. Compatible approximation spaces for the velocity-pressure-stress formulation for creeping flows. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4):225–231, May 2000. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/29/
17/21/61/27/48/abstract.
html;   http://www.elsevier.
nl/gej-ng/29/17/21/61/
27/48/article.pdf.

Gobbert:2001:CCN

[GP01] Matthias K. Gobbert and Andreas Prohl. A comparison of classical and new finite element methods for the computation of laminate microstructure. Applied Numerical Mathematics: Transactions of IMACS, 36(2–3):155–178, February 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.nl../28/
abstract.html;   http:
//www.elsevier.nl/gej-

                                        ng/10/10/28/65/31/28/article.
                                        pdf.

Gori:2004:RFP

[GP04] Laura Gori and Francesca Pitolli. Refinable functions and positive operators. Applied Numerical Mathematics: Transactions of IMACS, 49(3–4):381–393, June 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gori:2017:TPR

[GP17] Laura Gori and Francesca Pitolli. Totally positive refinable functions with general dilation M. Applied Numerical Mathematics: Transactions of IMACS, 112(??):17– 26, February 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416301982.

Garg:2023:UDG

[GP23] Divay Garg and Kamana Porwal. Unified discontinuous Galerkin finite element methods for second order Dirichlet boundary control problem. Applied Numerical Mathematics: Transactions of IMACS, 185(??): 336–364, March 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422003233.

Gonzalez-Pinto:2006:SIM

[GPHA06] S. González-Pinto and D. Hernández-Abreu. Semi-implicit methods for differential systems with semi-stable equilibria. Applied Numerical Mathematics: Transactions of IMACS, 56(2):210– 221, February 2006. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gonzalez-Pinto:2016:SMB

[GPHA16] S. González-Pinto and D. Hernández-Abreu. Splitting-methods based on Approximate Matrix Factorization and Radau-IIA formulas for the time integration of advection diffusion reaction PDEs. Applied Numerical Mathematics: Transactions of IMACS, 104(??):166–181, June 2016. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415001464.

Gonzalez-Pinto:2022:CMN

[GPHA22] S. González-Pinto and D. Hernández-Abreu. Convergence in the maximum norm of ADItype methods for parabolic problems. Applied Numerical Mathematics: Transactions of IMACS, 171(??): 269–280, January 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421002610.

Gonzalez-Pinto:2025:BCS

[GPHA25] S. González-Pinto and D. Hernández-Abreu. Boundary corrections for splitting methods in the time integration of multidimensional parabolic problems. Applied Numerical Mathematics: Transactions of IMACS, 210(??): 95–112, April 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424003337.

Gonzalez-Pinto:2012:EFS

[GPHAM12] S. González-Pinto, D. Hernández-Abreu, and J. I. Montijano. An efficient family of strongly A-stable Runge– Kutta collocation methods for stiff systems and DAEs. Part II: Convergence results. Applied Numerical Mathematics: Transactions of IMACS, 62(10): 1349–1360, October 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927412000967.

Gonzalez-Pinto:2023:GTR

[GPHAPPR23] S. González-Pinto, D. Hernández-Abreu, G. Pagano, and S. Pérez-Rodríguez. Generalized TASE-RK methods for stiff problems. Applied Numerical Mathematics: Transactions of IMACS, 188(??):129–145, June 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423000685.

Guimond:2003:SPI

[GPíP03] Louis-Sébastien Guimond, Jan Patera, and Ji í Patera. Statistical properties and implementation of aperiodic pseudorandom number generators. Applied Numerical Mathematics: Transactions of IMACS, 46(3–4): 295–318, September 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gonzalez-Pinto:2003:VOS

[GPMPR03] S. González-Pinto, J. I.

Montijano, and S. Pérez-Rodríguez. Variable-order starting algorithms for implicit Runge–Kutta methods on stiff problems. Applied Numerical Mathematics: Transactions of IMACS, 44(1–2):77–94, January 2003. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gonzalez-Pinto:1995:IST

[GPMR95] S. González-Pinto, J. I. Montijano, and L. Rández. Iterative schemes for threestage implicit Runge–Kutta methods. Applied Numerical Mathematics: Transactions of IMACS, 17(4):363– 382, September 11, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=17&
issue=4&aid=618.

Gori:2004:RRW

[GPP04] Laura Gori, Laura Pezza, and Francesca Pitolli. Recent results on wavelet bases on the interval generated by GP refinable functions. Applied Numerical Mathematics: Transactions of IMACS, 51(4):549–563, December 2004. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gonzalez-Pinto:2012:VTS

[GPPR12] S. González-Pinto and S. Pérez-Rodríguez. A variable timestep-size code for advectiondiffusion-reaction PDEs. Applied Numerical Mathematics: Transactions of IMACS, 62(10):1447–1462, October 2012. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412001109.

Gastaldi:1989:CHP

[GQ89] Fabio Gastaldi and Alfio Quarteroni. On the coupling of hyperbolic and parabolic systems: analytical and numerical approach. Applied Numerical Mathematics: Transactions of IMACS, 6(1–2):3–31, December 1989. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Spectral multi-domain methods (Paris, 1988).

Grimm:2008:GIM

[GQ08] V. Grimm and G. R. W.

Quispel. Geometric integration methods that unconditionally contract volume. Applied Numerical Mathematics: Transactions of IMACS, 58(8):1103–1112, August 2008. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gunther:1993:MRM

[GR93] M. Günther and P. Rentrop. Multirate ROW methods and latency of electric circuits. Applied Numerical Mathematics: Transactions of IMACS, 13(1– 3):83–102, September 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Sixth Conference on the Numerical Treatment of Differential Equations (Halle, 1992).

Gutknecht:2002:FGC

[GR02] Martin H. Gutknecht and Miroslav Rozložník. A framework for generalized conjugate gradient methods — with special emphasis on contributions by Rüdiger Weiss. Applied Numerical Mathematics: Transactions of IMACS, 41(1):7– 22, April 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/86/27/28/abstract.
html.

Gonzalez-Requena:2002:NAD

[GRGJ02] Reinaldo J. González-Requena and Juan M. Guevara-Jordan. New approach for describing transient pressure response generated by horizontal wells of arbitrary geometry. Applied Numerical Mathematics: Transactions of IMACS, 40(4):433– 449, March 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/85/34/27/abstract.
html.

Gomes-Ruggiero:2001:CAM

[GRLL01] Márcia Ap. Gomes-Ruggiero, Orlando Francisco Lopes, and Véra Lucia Rocha Lopes. A comparative analysis of the monotone iteration method for elliptic problems. Applied Numerical Mathematics: Transactions of IMACS, 36(2– 3):231–248, February 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.nl../32/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/65/31/32/article.
pdf.

Grove:1994:AFT

[Gro94]  John W. Grove.  Ap-

plications of front tracking to the simulation of shock refractions and unstable mixing. Applied Numerical Mathematics: Transactions of IMACS, 14(1–3): 213–237, April 14, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=14&
issue=1-3&aid=455. Pro-

ceedings of the Third ARO Workshop on Adaptive Methods for Partial Differential Equations (Troy, NY, 1992).

Giovangigli:1989:ACA

[GS89] V. Giovangigli and M. D.

Smooke. Adaptive continuation algorithms with application to combustion problems. Applied Numerical Mathematics: Transactions of IMACS, 5(4):305– 331, July 1989. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Guo:1992:AHB

[GS92] Ren Li Guo and Robert D. Skeel. An algebraic hierarchical basis preconditioner. Applied Numerical Mathematics: Transactions of IMACS, 9(1):21–32, January 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Guo:1994:FEA

[GS94] Wen Guo and Martin Stynes. Finite element analysis of an exponentially fitted non-lumped scheme for advection-diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 15(4):375–393, November 1, 1994. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1994&volume=15&
                                        issue=4&aid=526.

Ganesh:1999:OOS

[GS99a] M. Ganesh and I. H. Sloan. Optimal order spline methods for nonlinear differential and integro–differential equations. Applied Numerical Mathematics: Transactions of IMACS, 29(4):445– 478, April 1, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1999/29/
4/933.pdf.

Goldfarb:1999:PSP

[GS99b] D. Goldfarb and K. Scheinberg. On parametric semidefinite programming. Applied Numerical Mathematics: Transactions of IMACS, 29(3):361–377, March 1, 1999. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1999/29/
3/967.pdf.

Gori:2005:CPC

[GS05] Laura Gori and Elisabetta Santi. Convergence properties of certain refinable quasi-interpolatory operators. Applied Numerical Mathematics: Transactions of IMACS, 55(3):312–321, November 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gutknecht:2008:UQD

[GS08] Martin H. Gutknecht and Thomas Schmelzer. Updating the QR decomposition of block tridiagonal and block Hessenberg matrices. Applied Numerical Mathematics: Transactions of IMACS, 58(6):871–883, June 2008.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gorbunov:2009:MNS

[GS09] Vladimir K. Gorbunov and Vyacheslav Yu. Sviridov. The method of normal splines for linear DAEs on the number semi-axis. Applied Numerical Mathematics: Transactions of IMACS, 59(3–4):656–670, March/ April 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Galiano:2015:DPM

[GS15a] Gonzalo Galiano and Virginia Selgas. Deterministic particle method approximation of a contact inhibition cross-diffusion problem. Applied Numerical Mathematics: Transactions of IMACS, 95(??): 229–237, September 2015. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414001871.

Goldman:2015:TEP

[GS15b] Ron Goldman and Plamen Simeonov. Two essential properties of (q, h)- Bernstein–Bézier curves. Applied Numerical Mathematics: Transactions of IMACS, 96(??):82–93, October 2015. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415000768.

Gaaf:2017:ALS

[GS17] Sarah W. Gaaf and Valeria Simoncini. Approximating the leading singular triplets of a large matrix function. Applied Numerical Mathematics: Transactions of IMACS, 113(??): 26–43, March 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416302161.

Gonzalez:2018:AVS

[GS18] María González and Virginia Selgas. Analysis of a velocity–stress–pressure formulation for a fluidstructure interaction problem. Applied Numerical Mathematics: Transactions of IMACS, 123(??): 275–299, January 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417302052.

Gao:2019:HON

[GS19] Wenwu Gao and Zhengjie Sun. High-order numerical solution of time-dependent differential equations with quasi-interpolation. Applied Numerical Mathematics: Transactions of IMACS, 146(??):276–290, December 2019. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927419301837.

Giani:2020:MHA

[GS20] Stefano Giani and Mohammed Seaid. Multihp adaptive discontinuous Galerkin methods for simplified P_{N} approximations of 3D radiative transfer in nongray media. Applied Numerical Mathematics: Transactions of IMACS, 150(??): 252–273, April 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419302570.

Guessab:2021:OGH

[GS21] Allal Guessab and Boris Semisalov. Optimal general Hermite–Hadamard-type inequalities in a ball and their applications in multidimensional numerical integration. Applied Numerical Mathematics: Transactions of IMACS, 170(??):83– 108, December 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421002075.

G:2024:IFW

[GS24] Manohara G. and Kumbinarasaiah S. An innovative Fibonacci wavelet collocation method for the numerical approximation of Emden– Fowler equations. Applied Numerical Mathematics: Transactions of IMACS, 201(??):347–369, July 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000771.

Ghosh:2025:GFG

[GS25a] Riya Ghosh and A. Antony Selvan. On Gabor frames generated by B-splines, totally positive functions, and Hermite functions. Applied Numerical Mathematics: Transactions of IMACS, 207(??):1–23, January 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424002253.

Gritsans:2025:DEE

[GS25b] Armands Gritsans and Felix Sadyrbaev. On differential equations with exponential nonlinearities. Applied Numerical Mathematics: Transactions of IMACS, 207(??):558–568, January 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424002241.

Godunov:2000:STC

[GSR00] S. K. Godunov, M. Sadkane, and M. Robbé. Smoothing techniques and computation of invariant subspaces of elliptic operators. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4):341–347, May 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/21/61/27/
                                        61/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/21/61/27/61/article.
                                        pdf.

Guo:2009:GJP

[GSW09] Ben-Yu Guo, Jie Shen, and Li-Lian Wang. Generalized Jacobi polynomials/ functions and their applications. Applied Numerical Mathematics: Transactions of IMACS, 59(5): 1011–1028, May 2009. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gottlieb:1993:ING

[GT93a] David Gottlieb and Roger Temam. Implementation of the nonlinear Galerkin method with pseudospectral (collocation) discretizations. Applied Numerical Mathematics: Transactions of IMACS, 12(1–3): 119–134, May 1993. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0168927493901158.

Special issue to honor Professor Saul Abarbanel on his sixtieth birthday (Neveh, 1992).

Gottlieb:1993:D

[GT93b] David Gottlieb and Eli Turkel. Dedication. Applied Numerical Mathematics: Transactions of IMACS, 12(1–3):1–2, May 1993. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0168927493901084.

Special issue to honor Professor Saul Abarbanel on his sixtieth birthday (Neveh, 1992).

Gelb:2000:ESV

[GT00] Anne Gelb and Eitan Tadmor. Enhanced spectral viscosity approximations for conservation laws. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4):3–21, May 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/61/27/
28/abstract.html;   http:

                                        //www.elsevier.nl/gej-
                                        ng/29/17/21/61/27/28/article.
                                        pdf.

Gould:2002:IWS

[GT02] Nicholas I. M. Gould and Philippe L. Toint. An iterative working-set method for large-scale nonconvex quadratic programming. Applied Numerical Mathematics: Transactions of IMACS, 43(1–2):109–128, October 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ganesh:2015:SEC

[GT15] M. Ganesh and T. Thompson. Schrödinger eigenbasis on a class of superconducting surfaces: Ansatz, analysis, FEM approximations and computations. Applied Numerical Mathematics: Transactions of IMACS, 89(??):45–75, March 2015. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414001883.

Guzel:2018:WSI

[GT18] Ahmet Guzel and Catalin Trenchea. The Williams step increases the stability and accuracy of the hoRA time filter. Applied Numerical Mathematics: Transactions of IMACS, 131(??): 158–173, September 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418301132.

Gerdjikov:2019:MMG

[GT19a] V. S. Gerdjikov and M. D.

Todorov. Manakov model with gain/loss terms and Nsoliton interactions: Effects of periodic potentials. Applied Numerical Mathematics: Transactions of IMACS, 141(??):62–80, ???? 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418301338.

Ghasemifard:2019:MPS

[GT19b] Azadeh Ghasemifard and Mahdieh Tahmasebi. Multilevel path simulation to jump-diffusion process with superlinear drift. Applied Numerical Mathematics: Transactions of IMACS, 144(??):176–189, October 2019. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927419301035.

Tang:2007:LSM

[gTpM07] Jian guo Tang and He ping Ma. A Legendre spectral method in time for first-order hyperbolic equations. Applied Numerical Mathematics: Transactions of IMACS, 57(1):1–11, January 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Glandon:2020:BRK

[GTS20] Ross Glandon, Paul Tranquilli, and Adrian Sandu. Biorthogonal Rosenbrock– Krylov time discretization methods. Applied Numerical Mathematics: Transactions of IMACS, 150(??): 233–251, April 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419302429.

Gu:2001:SCE

[Gu01] Gui-Ding Gu. Some conditions for existence and stability of relaxed incomplete LU factorizations. Applied Numerical Mathematics: Transactions of IMACS, 38(1–2):105–121, July 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/10/28/77/27/
                                        32/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/10/28/77/27/32/article.
                                        pdf.

Gu:2019:SCM

[Gu19] Zhendong Gu. Spectral collocation method for weakly singular Volterra integrodifferential equations. Applied Numerical Mathematics: Transactions of IMACS, 143(??):263–275, September 2019. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419300984.

Gu:2020:SCM

[Gu20] Zhendong Gu. Spectral collocation method for nonlinear Riemann–Liouville fractional differential equations. Applied Numerical Mathematics: Transactions of IMACS, 157(??): 654–669, November 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302038.

Guglielmi:2005:NIA

[Gug05] Nicola Guglielmi. On the Newton iteration in the application of collocation methods to implicit delay equations. Applied Numerical Mathematics: Transactions of IMACS, 53(2–4): 281–297, May 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gulin:2015:SCN

[Gul15] A. Gulin. Stability criteria for non-self-adjoint finite differences schemes in the subspace. Applied Numerical Mathematics: Transactions of IMACS, 93(??): 107–113, July 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414000968.

Gumgum:2020:TWS

[Güm20]  Sevin Gümgüm.  Taylor

wavelet solution of linear and nonlinear Lane–Emden equations. Applied Numerical Mathematics: Transactions of IMACS, 158(??):44– 53, December 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302191.

Guo:1996:IBF

[Guo96]  Chun-Hua Guo.  Incom-

plete block factorization preconditioning for linear systems arising in the numerical solution of the Helmholtz equation. Applied Numerical Mathematics: Transactions of IMACS, 19(4): 495–508, February 12, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=19&
issue=4&aid=642.

Guo:2000:SOS

[Guo00] Daniel X. Guo. A second order scheme for the Navier– Stokes equations: Application to the driven-cavity problem. Applied Numerical Mathematics: Transactions of IMACS, 35(4): 307–322, December 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.nl../28/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/64/32/28/article.
pdf.

Guofeng:2001:SIO

[Guo01] Zhang Guofeng. Stability of implicit one-block methods for delay differential equations. Applied Numerical Mathematics: Transactions of IMACS, 36(2– 3):275–279, February 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.nl../34/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/65/31/34/article.
pdf.

Guo:2015:MFD

[Guo15]  Daniel X. Guo.  Modi-

fied fully discretized projection method for the incompressible Navier–Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 96(??):187– 202, October 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927415000963.

Guo:2026:CTE

[Guo26] Chun-Hua Guo. Comments on: “Two efficient iteration methods for solving the absolute value equations”. Applied Numerical Mathematics: Transactions of IMACS, 219(??):96–98, January 2026. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425001801.

See [YW25].

Gustafson:1987:VSF

[Gus87] Karl Gustafson. Vortex separation and fine structure dynamics. Applied Numerical Mathematics: Transactions of IMACS, 3(1–2): 167–182, May 1987. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0168927487900109.

Gustafsson:1988:ICO

[Gus88] Bertil Gustafsson. Inhomogeneous conditions at open boundaries for wave propagation problems. Applied Numerical Mathematics: Transactions of IMACS, 4(1):3–19, March 1988. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gerisch:2002:OSA

[GV02] A. Gerisch and J. G. Verwer. Operator splitting and approximate factorization for taxis–diffusion–reaction models. Applied Numerical Mathematics: Transactions of IMACS, 42(1– 3):159–176, August 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gelli:2004:ELP

[GV04] Daniele Gelli and Domenico Vitulano. Exploiting light projection for Shape from Shading. Applied Numerical Mathematics: Transactions of IMACS, 51(4): 535–548, December 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Garau:2018:AIA

[GV18] Eduardo M. Garau and Rafael Vázquez. Algorithms for the implementation of adaptive isogeometric methods using hierarchical Bsplines. Applied Numerical Mathematics: Transactions of IMACS, 123(??): 58–87, January 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417301794.

Gonzalez-Vera:1993:TPP

[GVP93] P. González-Vera and M. Jiménez Paiz. Two-point partial Padé approximants. Applied Numerical Mathematics: Transactions of IMACS, 11(5):385–402, March 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gonzalez-Vera:1996:PCP

[GVSL96] P. González-Vera and J. C. Santos-León. Properties of certain piecewise polynomial product integration rules. Applied Numerical Mathematics: Transactions of IMACS, 21(3): 241–264, August 20, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=21&
                                        issue=3&aid=689.

Griffiths:2002:E

[GW02] D. F. Griffiths and G. A. Watson. Editorial. Applied Numerical Mathematics: Transactions of IMACS, 43(1–2):??, October 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Guo:2020:HVC

[GW20] Yuling Guo and Zhongqing Wang. An hp-version Chebyshev collocation method for nonlinear fractional differential equations. Applied Numerical Mathematics: Transactions of IMACS, 158(??):194–211, December 2020. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302282.

Gwinner:2009:VAB

[Gwi09] Joachim Gwinner. On the pversion approximation in the boundary element method for a variational inequality of the second kind modelling unilateral contact and given friction. Applied Numerical Mathematics: Transactions of IMACS, 59(11): 2774–2784, November 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Guan:2022:UOC

[GWLN22] Zhen Guan, Jungang Wang, Ying Liu, and Yufeng Nie. Unconditionally optimal convergence of a linearized Galerkin FEM for the nonlinear time-fractional mobile/immobile transport equation. Applied Numerical Mathematics: Transactions of IMACS, 172(??): 133–156, February 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421002798.

Guo:2025:RMS

[GWZ25] Yuling Guo, Zhongqing Wang, and Chao Zhang. A robust mapping spectral method for elastic equations in curved fan-shaped domains. Applied Numerical Mathematics: Transactions of IMACS, 211(??): 144–157, May 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425000078.

Gear:1993:PAT

[GX93] C. W. Gear and Xuhai Xu. Parallelism across time in ODEs. Applied Numerical Mathematics: Transactions of IMACS, 11(1–3): 45–68, January 1993. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Parallel methods for ordinary differential equations (Grado, 1991).

Gao:2011:FOA

[GX11] Zhen Gao and Shusen Xie. Fourth-order alternating direction implicit compact finite difference schemes for two-dimensional Schrödinger equations. Applied Numerical Mathematics: Transactions of IMACS, 61(4): 593–614, April 2011. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Gerasoulis:1994:PBC

[GY94] Apostolos Gerasoulis and Tao Yang. Performance bounds for column-block partitioning of parallel Gaussian elimination and Gauss– Jordan methods. Applied Numerical Mathematics: Transactions of IMACS, 16(1–2):283–297, December 1, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=16&
issue=1-2&aid=527.  A

Festschrift to honor Professor Robert Vichnevetsky on his 65th birthday.

Guo:2023:PEA

[GZHQ23] Tao Guo, Mahmoud A. Zaky, Ahmed S. Hendy, and Wenlin Qiu. Pointwise error analysis of the BDF3 compact finite difference scheme for viscous Burgers’ equations. Applied Numerical Mathematics: Transactions of IMACS, 185(??): 260–277, March 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422003105.

Gao:2023:HOU

[GZQS23] Zhongxiong Gao, Hong Zhang, Xu Qian, and Songhe Song. High-order unconditionally maximum-principle-preserving parametric integrating factor Runge– Kutta schemes for the nonlocal Allen–Cahn equation. Applied Numerical Mathematics: Transactions of IMACS, 194(??):97–114, December 2023. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423002349.

Ge:2022:FPI

[GZW22] Zhili Ge, Xin Zhang, and Zhongming Wu. A fast proximal iteratively reweighted nuclear norm algorithm for nonconvex low-rank matrix minimization problems. Applied Numerical Mathematics: Transactions of IMACS, 179(??):66– 86, September 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422001015.

Gao:2019:AMK

[GZZ19] Qinjiao Gao, Shenggang Zhang, and Jihong Zhang. Adaptive moving knots meshless method for Degasperis– Procesi equation with conservation laws. Applied Numerical Mathematics: Transactions of IMACS, 142(??): 90–101, August 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419300443.

Gao:2020:MWG

[GZZ20] Fuzheng Gao, Shangyou Zhang, and Peng Zhu. Modified weak Galerkin method with weakly imposed boundary condition for convectiondominated diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 157(??): 490–504, November 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302105.

Hosseini:2016:NSL

[HA16] S. A. Hosseini and A. Abdi.

On the numerical stability of the linear barycentric rational quadrature method for Volterra integral equations. Applied Numerical Mathematics: Transactions of IMACS, 100(??):1– 13, February 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415001476.

Hassani:2021:NOM

[HA21] H. Hassani and Z. Avazzadeh. Novel operational matrices for solving 2-dim nonlinear variable order fractional optimal control problems via a new set of basis functions. Applied Numerical Mathematics: Transactions of IMACS, 166(??): 26–39, August 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421000854.

Haase:1997:HEO

[Haa97]  G. Haase.  Hierarchical

extension operators plus smoothing in domain decomposition preconditioners. Applied Numerical Mathematics: Transactions of IMACS, 23(3):327–346, April 25, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1997/23/
                                        3/752.pdf.

Heydari:2021:OSD

[HAA21] M. H. Heydari, Z. Avazzadeh, and A. Atangana. Orthonormal shifted discrete Legendre polynomials for solving a coupled system of nonlinear variable-order time fractional reaction– advection–diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 161(??): 425–436, March 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303706.

Haber:2008:PML

[Hab08]  Eldad Haber.  A paral-

lel method for large scale time domain electromagnetic inverse problems. Applied Numerical Mathematics: Transactions of IMACS, 58(4):422–434, April 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Huang:2022:LNE

[HAC22] Chaobao Huang, Na An, and Hu Chen. Local H ^{1}-norm error analysis of a mixed finite element method for a time-fractional biharmonic equation. Applied Numerical Mathematics: Transactions of IMACS, 173(??): 211–221, March 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742100338X.

Hadjidimos:2013:I

[Had13] Apostolos Hadjidimos. Introduction. Applied Numerical Mathematics: Transactions of IMACS, 67(??): 1–3, May 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413000263.

Hagstrom:2015:HRD

[Hag15] Thomas Hagstrom. Highresolution difference methods with exact evolution for multidimensional waves. Applied Numerical Mathematics: Transactions of IMACS, 93(??):114–122, July 2015. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414001275.

Hairer:1997:VTS

[Hai97] E. Hairer. Variable time step integration with symplectic methods. Applied Numerical Mathematics: Transactions of IMACS, 25(2–3): 219–227, October 17, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=25&
issue=2-3&aid=816.

Hashemi:2021:FOA

[HAML21] M. S. Hashemi, E. Ashpazzadeh, M. Moharrami, and M. Lakestani. Fractional order Alpert multiwavelets for discretizing delay fractional differential equation of pantograph type. Applied Numerical Mathematics: Transactions of IMACS, 170(??):1– 13, December 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421002063.

Hanson:1987:CLL

[Han87] R. J. Hanson. On the constrained linear least-squares problem: a personal view. Applied Numerical Mathematics: Transactions of IMACS, 3(5):443–452, October 1987. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Han:1993:AEE

[Han93] Guo Qiang Han. Asymptotic error expansion of a collocation-type method for Volterra–Hammerstein integral equations. Applied Numerical Mathematics: Transactions of IMACS, 13(5):357–369, December 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Han:2006:QTP

[Han06] Xuli Han. Quadratic trigonometric polynomial curves concerning local control. Applied Numerical Mathematics: Transactions of IMACS, 56(1):105–115, January 2006. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Han:2019:TVB

[Han19] Huan Han. A tensor voting based fractional-order image denoising model and its numerical algorithm. Applied Numerical Mathematics: Transactions of IMACS, 145(??):133–144, November 2019. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892741930162X.

Haq:2023:ECT

[HAN23] Sirajul Haq, Shams Ul Arifeen, and Ayesha Noreen. An efficient computational technique for higher order KdV equation arising in shallow water waves. Applied Numerical Mathematics: Transactions of IMACS, 189(??):53–65, July 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423000831.

Harti:1993:DMR

[Har93] Ami Harti. Discrete multiresolution analysis and generalized wavelets. Applied Numerical Mathematics: Transactions of IMACS, 12(1–3):153–192, May 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/016892749390117A.

Special issue to honor Professor Saul Abarbanel on his sixtieth birthday (Neveh, 1992).

Harari:1998:UVA

[Har98]  Isaac Harari.  A uni-

fied variational approach to domain-based computation of exterior problems of time-harmonic acoustics. Applied Numerical Mathematics: Transactions of IMACS, 27(4):417–441, August 15, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1998/27/
4/889.pdf.

Harris:2000:SBE

[Har00] Frank E. Harris. Spherical Bessel expansions of sine, cosine, and exponential integrals. Applied Numerical Mathematics: Transactions of IMACS, 34(1):95–98, June 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/62/27/
32/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/62/27/32/article.
pdf;   http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0168927499000318.

Harbrecht:2009:NSP

[Har09]  Helmut Harbrecht.  On

the numerical solution of Plateau’s problem. Applied Numerical Mathematics: Transactions of IMACS, 59(11):2785–2800, November 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Harbrecht:2010:FEM

[Har10] Helmut Harbrecht. A finite element method for elliptic problems with stochastic input data. Applied Numerical Mathematics: Transactions of IMACS, 60(3): 227–244, March 2010. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Heydari:2021:VLP

[HAR21] M. H. Heydari, Z. Avazzadeh, and M. Razzaghi. Vieta–Lucas polynomials for the coupled nonlinear variable-order fractional Ginzburg–Landau equations. Applied Numerical Mathematics: Transactions of IMACS, 165(??):442–458, July 2021. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421000775.

Hascelik:2008:MAG

[Has08]  A. I. Hascelik.  Mod-

ified anti-Gauss and degree optimal average formulas for Gegenbauer measure. Applied Numerical Mathematics: Transactions of IMACS, 58(2):171–179, February 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hascelik:2009:SGF

[Has09] A. Ihsan Hascelik. Suitable Gauss and Filon-type methods for oscillatory integrals with an algebraic singularity. Applied Numerical Mathematics: Transactions of IMACS, 59(1):101– 118, January 2009. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hascelik:2013:AFT

[Has13]  A. Ihsan Hasçelik.  An

asymptotic Filon-type method for infinite range highly oscillatory integrals with exponential kernel. Applied Numerical Mathematics: Transactions of IMACS, 63(1):1–13, January 2013. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412001602.

Hascelik:2020:ECH

[Has20] A. Ihsan Hascelik. Efficient computation of highly oscillatory Fourier-type integrals with monomial phase functions and Jacobi-type singularities. Applied Numerical Mathematics: Transactions of IMACS, 150(??): 303–312, April 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741930282X.

Hautot:1988:CAC

[Hau88] A. Hautot. Convergence acceleration of continued fractions of Poincaré type. Applied Numerical Mathematics: Transactions of IMACS, 4(2–4):309–322, June 1988. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Huang:2020:LDG

[HAY20] Chaobao Huang, Na An, and Xijun Yu. A local discontinuous Galerkin method for time-fractional diffusion equation with discontinuous coefficient. Applied Numerical Mathematics: Transactions of IMACS, 151(??): 367–379, May 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419303113.

Hackbusch:2002:MAI

[HB02] Wolfgang Hackbusch and Steffen Börm. -matrix approximation of integral operators by interpolation. Applied Numerical Mathematics: Transactions of IMACS, 43(1–2):129–143, October 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Holtom:2020:EPA

[HB20] Theodore C. Holtom and Anthony C. Brooms. Error propagation analysis for a static convergent beam triple LIDAR. Applied Numerical Mathematics: Transactions of IMACS, 150(??): 1–17, April 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419302296.

Habibirad:2024:EMN

[HBA^{+}24] Ali Habibirad, Omid Baghani, Hadis Azin, Mehdi Zaferanieh, and Mustafa Inc. An efficient meshless numerical method with the error estimate for two-dimensional Schrödinger equation. Applied Numerical Mathematics: Transactions of IMACS, 202(??): 143–157, August 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001090.

Heydari:2025:PLC

[HBB25] M. H. Heydari, D. Baleanu, and M. Bayram. Piecewise logarithmic Chebyshev cardinal functions: Application for nonlinear integral equations with a logarithmic singular kernel. Applied Numerical Mathematics: Transactions of IMACS, 217(??): 355–371, November 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425001394.

Hubbard:2009:CDB

[HBJ09] M. E. Hubbard, M. J. Baines, and P. K. Jimack. Consistent Dirichlet boundary conditions for numerical solution of moving boundary problems. Applied Numerical Mathematics: Transactions of IMACS, 59(6): 1337–1353, June 2009. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Huang:2001:CIE

[HC01] Shi-Wei Huang and Tucker Carrington, Jr. Calculating interior eigenvalues and eigenvectors with an implicitly restarted and a filter diagonalization method. Applied Numerical Mathematics: Transactions of IMACS, 37(3):307–317, May 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/73/31/
29/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/73/31/29/article.
pdf.

He:2022:ECC

[HC22] Yuyu He and Hongtao Chen. Efficient and conservative compact difference scheme for the coupled Schrödinger–Boussinesq equations. Applied Numerical Mathematics: Transactions of IMACS, 182(??): 285–307, December 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422002197.

He:2024:PIP

[HCC24] Yuyu He, Hongtao Chen, and Bolin Chen. Physical invariants-preserving compact difference schemes for the coupled nonlinear Schrödinger–KdV equations. Applied Numerical Mathematics: Transactions of IMACS, 204(??): 162–175, October 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001478.

Hu:2022:EEP

[HCGW22] Dongdong Hu, Wenjun Cai, Xian-Ming Gu, and Yushun Wang. Efficient energy preserving Galerkin-Legendre spectral methods for fractional nonlinear Schrödinger equation with wave operator. Applied Numerical Mathematics: Transactions of IMACS, 172(??): 608–628, February 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421002981.

He:2024:SUR

[HCJW24] Xiaoxiao He, Yanping Chen, Haifeng Ji, and Haijin Wang. Superconvergence of unfitted Rannacher– Turek nonconforming element for elliptic interface problems. Applied Numerical Mathematics: Transactions of IMACS, 203(??):32– 51, September 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424001223.

He:2020:DCS

[HCS20] Zhu He, Jiaxiang Cai, and Bangyu Shen. Decoupled conservative schemes for computing dynamics of the strongly coupled nonlinear Schrödinger system. Applied Numerical Mathematics: Transactions of IMACS, 157(??): 276–290, November 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420301902.

Huang:2016:NTL

[HCW16] Peiqi Huang, Mingchao Cai, and Feng Wang. A Newton type linearization based two grid method for coupling fluid flow with porous media flow. Applied Numerical Mathematics: Transactions of IMACS, 106(??): 182–198, August 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416300411.

Huang:2003:CGL

[HCX03] Chengming Huang, Qianshun Chang, and Aiguo Xiao. B-convergence of general linear methods for stiff problems. Applied Numerical Mathematics: Transactions of IMACS, 47(1):31– 44, October 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hou:2018:TGM

[HCY18] Tianliang Hou, Luoping Chen, and Yin Yang. Twogrid methods for expanded mixed finite element approximations of semi-linear parabolic integro-differential equations. Applied Numerical Mathematics: Transactions of IMACS, 132(??): 163–181, October 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418301375.

Hariharan:1988:AGW

[HD88] S. I. Hariharan and P. K. Dutt. Acoustic gravity waves: a computational approach. Applied Numerical Mathematics: Transactions of IMACS, 4(6):491– 506, 1988. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Harari:2004:ASE

[HD04] Isaac Harari and Rabia Djellouli. Analytical study of the effect of wave number on the performance of local absorbing boundary conditions for acoustic scattering. Applied Numerical Mathematics: Transactions of IMACS, 50(1):15–47, July 2004. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

He:2022:IPP

[HD22] Xiaoxiao He and Weibing Deng. An interface penalty parameter free nonconforming cut finite element method for elliptic interface problems. Applied Numerical Mathematics: Transactions of IMACS, 173(??): 434–452, March 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421003457.

Hu:2023:TOS

[HD23] Ziyue Hu and Qiao-Li Dong.

A three-operator splitting algorithm with deviations for generalized DC programming. Applied Numerical Mathematics: Transactions of IMACS, 191(??):62– 74, September 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423000971.

Hosseinzadeh:2020:SSN

[HDS20] Hossein Hosseinzadeh, Mehdi Dehghan, and Zeynab Sedaghatjoo. The stability study of numerical solution of Fredholm integral equations of the first kind with emphasis on its application in boundary elements method. Applied Numerical Mathematics: Transactions of IMACS, 158(??):134–151, December 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420302117.

Henderson:2017:TGI

[HdSRI17] Nélio Henderson, Marroni de Sá Rêgo, and Janaína Imbiriba. Topographical global initialization for finding all solutions of nonlinear systems with constraints. Applied Numerical Mathematics: Transactions of IMACS, 112(??): 155–166, February 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892741630201X.

He:2021:MMB

[HDY21] Shangqin He, Congna Di, and Li Yang. The mollification method based on a modified operator to the ill-posed problem for 3D Helmholtz equation with mixed boundary. Applied Numerical Mathematics: Transactions of IMACS, 160(??):422– 435, February 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420303196.

He:2024:APD

[He24] Xin He. Accelerated primaldual methods with adaptive parameters for composite convex optimization with linear constraints. Applied Numerical Mathematics: Transactions of IMACS, 203(??):129–143, September 2024. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001272.

Hendricks:2016:HOA

[HEG16] Christian Hendricks, Matthias Ehrhardt, and Michael Günther. High-order ADI schemes for diffusion equations with mixed derivatives in the combination technique. Applied Numerical Mathematics: Transactions of IMACS, 101(??): 36–52, March 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415001609.

Hull:1996:RKR

[HEJ96] T. E. Hull, W. H. Enright, and K. R. Jackson. Runge–Kutta research at Toronto. Applied Numerical Mathematics: Transactions of IMACS, 22(1–3): 225–236, December 4, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=22&
issue=1-3&aid=716. Special

issue celebrating the centenary of Runge–Kutta methods.

Herbert:1991:ETC

[Her91]  Thorwald Herbert.  Ex-

ploring transition by computer. Applied Numerical Mathematics: Transactions of IMACS, 7(1):3– 25, January 1991. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0168927491901015.

Hesthaven:2000:SPM

[Hes00]  J. S. Hesthaven.  Spec-

tral penalty methods. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4):23–41, May 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/21/61/27/
                                        29/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/21/61/27/29/article.
                                        pdf.

Heuer:2000:SMS

[Heu00] Norbert Heuer. Schwarz method for systems of boundary integral equations. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4): 447–453, May 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/61/27/
72/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/61/27/72/article.
pdf.

Heyouni:2010:EAM

[Hey10]  M. Heyouni.  Extended

Arnoldi methods for large low-rank Sylvester matrix equations. Applied Numerical Mathematics: Transactions of IMACS, 60(11): 1171–1182, November 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Heydari:2019:CCW

[Hey19] Mohammad Hossein Heydari. Chebyshev cardinal wavelets for nonlinear variable-order fractional quadratic integral equations. Applied Numerical Mathematics: Transactions of IMACS, 144(??): 190–203, October 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927419301060.

Heydari:2020:CMC

[Hey20a] M. H. Heydari. A computational method for a class of systems of nonlinear variable-order fractional quadratic integral equations. Applied Numerical Mathematics: Transactions of IMACS, 153(??): 164–178, July 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420300532.

Heydari:2020:NSN

[Hey20b] M. H. Heydari. Numerical solution of nonlinear 2D optimal control problems generated by Atangana–Riemann– Liouville fractal–fractional derivative. Applied Numerical Mathematics: Transactions of IMACS, 150(??): 507–518, April 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419303071.

Huang:2012:TLS

[HFL12] Pengzhan Huang, Xinlong Feng, and Demin Liu. Twolevel stabilized method based on three corrections for the stationary Navier–Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 62(8):988– 1001, August 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927412000542.

Huang:2013:IAC

[HFL13] Y. Huang, P. A. Forsyth, and G. Labahn. Inexact arithmetic considerations for direct control and penalty methods: American options under jump diffusion. Applied Numerical Mathematics: Transactions of IMACS, 72(??):33–51, October 2013. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413000652.

Hagstrom:1998:EAR

[HG98] Thomas Hagstrom and John Goodrich. Experiments with approximate radiation boundary conditions for computational aeroacoustics. Applied Numerical Mathematics: Transactions of IMACS, 27(4):385–402, August 15, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1998/27/
4/887.pdf.

He:2024:ATD

[HGF24] Linshuang He, Jun Guo, and Minfu Feng. Analysis of two discontinuous Galerkin finite element methods for the total pressure formulation of linear poroelasticity model. Applied Numerical Mathematics: Transactions of IMACS, 204(??): 60–85, October 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424001454.

Hacioglu:2021:IAF

[HGM^{+}21] Emirhan Hacioglu, Faik Gürsoy, Samet Maldar, Yunus Atalan, and Gradimir V. Milovanović. Iterative approximation of fixed points and applications to twopoint second-order boundary value problems and to machine learning. Applied Numerical Mathematics: Transactions of IMACS, 167(??):143–172, September 2021. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421001197.

Hadizadeh:2011:JSS

[HGP11] M. Hadizadeh, F. Ghoreishi, and S. Pishbin. Jacobi spectral solution for integral algebraic equations of index-2. Applied Numerical Mathematics: Transactions of IMACS, 61(1):131– 148, January 2011. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hari:2001:COW

[HGR01] Lydia Peres Hari, Dan Givoli, and Jacob Rubinstein. Computation of open Willmore-type surfaces. Applied Numerical Mathematics: Transactions of IMACS, 37(1–2):257–269, April 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/73/26/
41/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/73/26/41/article.
pdf.

Hong:2021:AHO

[HGZW21] Qi Hong, Yuezheng Gong, Jia Zhao, and Qi Wang. Arbitrarily high order structurepreserving algorithms for the Allen–Cahn model with a nonlocal constraint. Applied Numerical Mathematics: Transactions of IMACS, 170(??):321–339, December 2021. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421002300.

Hagstrom:1998:FAE

[HH98] Thomas Hagstrom and S. I.

Hariharan. A formulation of asymptotic and exact boundary conditions using local operators. Applied Numerical Mathematics: Transactions of IMACS, 27(4):403–416, August 15, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/

                                        store/apnum/sub/1998/27/
                                        4/888.pdf.

Hewitt:2010:ASD

[HH10a] L. L. Hewitt and A. T. Hill. Algebraically stable diagonally implicit general linear methods. Applied Numerical Mathematics: Transactions of IMACS, 60(6):629–636, June 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Huang:2010:PPT

[HH10b] Chiau-Yu Huang and Feng-Nan Hwang. Parallel pseudotransient Newton–Krylov– Schwarz continuation algorithms for bifurcation analysis of incompressible sudden expansion flows. Applied Numerical Mathematics: Transactions of IMACS, 60(7):738–751, July 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Huang:2018:HVE

[HH18] Jianguo Huang and Xuehai Huang. An hp-version error analysis of the discontinuous Galerkin method for linear elasticity. Applied Numerical Mathematics: Transactions of IMACS, 133(??):83–99, ???? 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741730260X.

Hassan:2022:FEA

[HH22] Sattar M. Hassan and Akil J.

Harfash. Finite element analysis of a two-species chemotaxis system with two chemicals. Applied Numerical Mathematics: Transactions of IMACS, 182(??): 148–175, December 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001970.

Hamou:2022:MIT

[HHAA22] Abdelouahed Alla Hamou, Zakia Hammouch, Elhoussine Azroul, and Praveen Agarwal. Monotone iterative technique for solving finite difference systems of time fractional parabolic equations with initial/periodic conditions. Applied Numerical Mathematics: Transactions of IMACS, 181(??): 561–593, November 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001155.

He:2008:PFV

[HHC08] Guoliang He, Yinnian He, and Zhangxin Chen. A penalty finite volume method for the transient Navier– Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 58(11):1583–1613, November 2008. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hashemi:2023:ICF

[HHL23] A. S. Hashemi, M. Heydari, and G. B. Loghmani. Iterative compact finite difference method for the numerical study of fully wet porous fins with different profile shapes. Applied Numerical Mathematics: Transactions of IMACS, 186(??): 358–377, April 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423000302.

Hamkar:2012:SAR

[HHR12] Ahmad-Wahadj Hamkar, Stefan Hartmann, and Joachim Rang. A stiffly accurate Rosenbrock-type method of order 2 applied to FEanalyses in finite strain viscoelasticity. Applied Numerical Mathematics: Transactions of IMACS, 62(12): 1837–1848, December 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741200133X.

Hartung:1997:EUN

[HHT97] Ferenc Hartung, Terry L.

Herdman, and Janos Turi. On existence, uniqueness and numerical approximation for neutral equations with state-dependent delays. Applied Numerical Mathematics: Transactions of IMACS, 24(2–3): 393–409, June 19, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=24&
issue=2-3&aid=791. Volterra

centennial (Tempe, AZ, 1996).

Hadjidimos:2003:P

[HHV03] Apostolos Hadjidimos, Elias N.

Houstis, and Emmanuel Vavalis. Preface. Applied Numerical Mathematics: Transactions of IMACS, 45(1):1–2, April 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hong:2018:SRK

[HHW18] Jialin Hong, Chuying Huang, and Xu Wang. Symplectic Runge–Kutta methods for Hamiltonian systems driven by Gaussian rough paths. Applied Numerical Mathematics: Transactions of IMACS, 129(??):

120–136, July 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418300680.

Hammad:2020:EJG

[HHYD20] Magda Hammad, Ramy M. Hafez, Youssri H. Youssri, and Eid H. Doha. Exponential Jacobi–Galerkin method and its applications to multidimensional problems in unbounded domains. Applied Numerical Mathematics: Transactions of IMACS, 157(??):88– 109, November 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420301616.

Higham:1993:ECI

[Hig93a] Desmond J. Higham. Error control for initial value problems with discontinuities and delays. Applied Numerical Mathematics: Transactions of IMACS, 12(4): 315–330, June 1993. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Higueras:1993:CTE

[Hig93b] Immaculada Higueras. Coefficients of the Taylor expansion for the solution of differential-algebraic systems. Applied Numerical Mathematics: Transactions of IMACS, 12(6):497– 501, August 1993. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Higham:1996:RKT

[Hig96] Desmond J. Higham. Runge– Kutta type methods for orthogonal integration. Applied Numerical Mathematics: Transactions of IMACS, 22(1–3):217–223, December 4, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=22&
issue=1-3&aid=715. Special

issue celebrating the centenary of Runge–Kutta methods.

Higham:1997:RRK

[Hig97] Desmond J. Higham. Regular Runge–Kutta pairs. Applied Numerical Mathematics: Transactions of IMACS, 25(2–3):229–241, October 17, 1997. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=25&
issue=2-3&aid=817.

Hazanee:2013:ITD

[HILK13] A. Hazanee, M. I. Ismailov, D. Lesnic, and N. B. Kerimov. An inverse timedependent source problem for the heat equation. Applied Numerical Mathematics: Transactions of IMACS, 69(??):13–33, July 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927413000299.

Hindmarsh:1995:ABS

[Hin95] Alan C. Hindmarsh. Avoiding BDF stability barriers in the MOL solution of advection-dominated problems. Applied Numerical Mathematics: Transactions of IMACS, 17(3):311–318, September 4, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=17&
                                        issue=3&aid=584. Numeri-

cal methods for ordinary differential equations (Atlanta, GA, 1994).

Hinze:1997:NAC

[Hin97] Michael Hinze. On the numerical approximation and computation of minimal surface continua bounded by one-parameter families of polygonal contours. Applied Numerical Mathematics: Transactions of IMACS, 25(1):89–116, September 16, 1997. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=25&
issue=1&aid=800.

Hundsdorfer:2003:IET

[HJ03] Willem Hundsdorfer and Jérôme Jaffré. Implicit– explicit time stepping with spatial discontinuous finite elements. Applied Numerical Mathematics: Transactions of IMACS, 45(2–3): 231–254, May 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hayes:2005:SSM

[HJ05] Wayne Hayes and Kenneth R. Jackson. A survey of shadowing methods for numerical solutions of ordinary differential equations. Applied Numerical Mathematics: Transactions of IMACS, 53(2–4):299–321, May 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hoppensteadt:2006:NSP

[HJ06] F. C. Hoppensteadt and Z. Jackiewicz. Numerical solution of a problem in the theory of epidemics. Applied Numerical Mathematics: Transactions of IMACS, 56(3–4): 533–543, March/April 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hansbo:2009:WID

[HJ09] P. Hansbo and M. Juntunen. Weakly imposed Dirichlet boundary conditions for the Brinkman model of porous media flow. Applied Numerical Mathematics: Transactions of IMACS, 59(6): 1274–1289, June 2009. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Heris:2017:FBD

[HJ17] M. Saedshoar Heris and M. Javidi. On fractional backward differential formulas for fractional delay differential equations with periodic and anti-periodic conditions. Applied Numerical Mathematics: Transactions of IMACS, 118(??): 203–220, August 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417300673.

Hu:2021:ENM

[HJ21] Mingshang Hu and Lianzi Jiang. An efficient numerical method for forwardbackward stochastic differential equations driven by G-Brownian motion. Applied Numerical Mathematics: Transactions of IMACS, 165(??):578–597, July 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000829.

Hong:2017:OEE

[HJKW17] Jialin Hong, Lihai Ji, Linghua Kong, and Tingchun Wang. Optimal error estimate of a compact scheme for nonlinear Schrödinger equation. Applied Numerical Mathematics: Transactions of IMACS, 120(??): 68–81, October 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417301198.

Hong:2018:OEE

[HJL18] Jialin Hong, Lihai Ji, and Zhihui Liu. Optimal error estimate of conservative local discontinuous Galerkin method for nonlinear Schrödinger equation. Applied Numerical Mathematics: Transactions of IMACS, 127(??): 164–178, May 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418300138.

He:2010:ERT

[HJP10] Suyan He, Yuxi Jiang, and Shaohua Pan. An entropy regularization technique for minimizing a sum of Tchebycheff norms. Applied Numerical Mathematics: Transactions of IMACS, 60(3): 208–223, March 2010. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hamadi:2022:EEB

[HJR22] M. A. Hamadi, K. Jbilou, and A. Ratnani. An efficient extended block Arnoldi algorithm for feedback stabilization of incompressible Navier–Stokes flow problems. Applied Numerical Mathematics: Transactions of IMACS, 174(??): 142–162, April 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422000113.

Han:1997:KMS

[HJS97] Weimin Han, Søren Jensen, and Igor Shimansky. The Kačanov method for some nonlinear problems. Applied Numerical Mathematics: Transactions of IMACS, 24(1):57–79, June 9, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=24&
issue=1&aid=765.

Huang:2019:SWP

[HJX^{+}19] Qiumei Huang, Shanghui Jia, Fei Xu, Zhongwen Xu, and Changhui Yao. Solvability of wave propagation with Debye polarization in nonlinear dielectric materials and its finite element methods approximation. Applied Numerical Mathematics: Transactions of IMACS, 146(??): 145–164, December 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927419301746.

Hou:2019:TGM

[HJYL19] Tianliang Hou, Wenzhu Jiang, Yueting Yang, and Haitao Leng. Two-grid P_{0}^{2}–P_{1} mixed finite element methods combined with Crank–Nicolson scheme for a class of nonlinear parabolic equations. Applied Numerical Mathematics: Transactions of IMACS, 137(??): 136–150, March 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418302587.

Helmer:2023:AFV

[HJZ23] Christoph Helmer, Ansgar Jüngel, and Antoine Zurek. Analysis of a finite-volume scheme for a single-species biofilm model. Applied Numerical Mathematics: Transactions of IMACS, 185(??):386–405, March 2023. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422003245.

Hsiao:1985:IET

[HK85] George C. Hsiao and Rainer Kress. On an integral equation for the two-dimensional exterior Stokes problem. Applied Numerical Mathematics: Transactions of IMACS, 1(1):77–93, January 1985. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hermann:1993:RSP

[HK93] M. Hermann and D. Kaiser. RWPM: a software package of shooting methods for nonlinear two-point boundary value problems. Applied Numerical Mathematics: Transactions of IMACS, 13(1–3):103–108, September 1993. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Sixth Conference on the Numerical Treatment of Differential Equations (Halle, 1992).

Huttunen:2009:MRS

[HK09] Janne M. J. Huttunen and Jari P. Kaipio. Model reduction in state identification problems with an application to determination of thermal parameters. Applied Numerical Mathematics: Transactions of IMACS, 59(5):877–890, May 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hingyi:2022:DDC

[HK22] B. Hingyi and J. Karátson.

Detection of dead cores for reaction-diffusion equations with a non-smooth nonlinearity. Applied Numerical Mathematics: Transactions of IMACS, 177(??): 111–122, July 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422000654.

Hecht:2025:PEI

[HK25] F. Hecht and S.-M. Kaber.

A parallel exponential integrator scheme for linear differential equations. Applied Numerical Mathematics: Transactions of IMACS, 208 (part B)(??): 356–364, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424002940.

Hadjidimos:2016:PSI

[HKNV16] Apostolos Hadjidimos, Ilias S. Kotsireas, Dimitrios Noutsos, and Michael N. Vrahatis. Preface [Special issue: NumAn 2012]. Applied Numerical Mathematics: Transactions of IMACS, 104(??):1–2, June 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416000295.

Hansen:2012:SOP

[HKO12] Eskil Hansen, Felix Kramer, and Alexander Ostermann. A second-order positivity preserving scheme for semilinear parabolic problems. Applied Numerical Mathematics: Transactions of IMACS, 62(10):1428–1435, October 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892741200089X.

Hussaini:1989:SCM

[HKP89] M. Y. Hussaini, D. A. Kopriva, and A. T. Patera. Spectral collocation methods. Applied Numerical Mathematics: Transactions of IMACS, 5(3):177–208, May 1989. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hafez:1986:NSV

[HKS86] M. Hafez, G. Kuruvila, and M. D. Salas. Numerical study of vortex breakdown. Applied Numerical Mathematics: Transactions of IMACS, 2(3–5):291– 302, October 1986. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hartmann:2008:CED

[HKZ08] Judith Hartmann, Andreas Krahnke, and Christoph Zenger. Cache efficient data structures and algorithms for adaptive multidimensional multilevel finite element solvers. Applied Numerical Mathematics: Transactions of IMACS, 58(4):435–448, April 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hyman:1989:DRM

[HL89] James M. Hyman and Bernard Larrouturou. Dynamic rezone methods for partial differential equations in one space dimension. Applied Numerical Mathematics: Transactions of IMACS, 5(5):435– 450, September 1989. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hyman:1997:DTM

[HL97] James M. Hyman and Jia Li. Disease transmission models with biased partnership selection. Applied Numerical Mathematics: Transactions of IMACS, 24(2– 3):379–392, June 19, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=24&
issue=2-3&aid=790. Volterra

centennial (Tempe, AZ, 1996).

Hairer:1999:ITD

[HL99] Ernst Hairer and Christian Lubich. Invariant tori of dissipatively perturbed Hamiltonian systems under symplectic discretization. Applied Numerical Mathematics: Transactions of IMACS, 29(1):57–71, January 1, 1999. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1999/29/
1/895.pdf.

Hild:2002:QFE

[HL02a] Patrick Hild and Patrick Laborde. Quadratic finite element methods for unilateral contact problems. Applied Numerical Mathematics: Transactions of IMACS, 41(3):401–421, June 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hu:2002:ANS

[HL02b] Qiya Hu and Guoping Liang.

Acceleration of the nonsymmetrized two-level iteration. Applied Numerical Mathematics: Transactions of IMACS, 41(2): 305–323, May 2002. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Holstad:2003:TBC

[HL03] Astrid Holstad and Ivar Lie. Transparent boundary conditions for the shallow water equations with a mixed finite element formulation. Applied Numerical Mathematics: Transactions of IMACS, 44(1– 2):109–138, January 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

He:2008:SFE

[HL08] Yinnian He and Jian Li.

A stabilized finite element method based on local polynomial pressure projection for the stationary Navier– Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 58(10):1503–1514, October 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hou:2019:SAT

[HL19] Tianliang Hou and Haitao Leng. Superconvergence analysis and two-grid algorithms of pseudostressvelocity MFEM for optimal control problems governed by Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 138(??):78–93, April 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418302824.

Hou:2021:EPT

[HL21] Baohui Hou and Dong Liang. The energy-preserving time high-order AVF compact finite difference scheme for nonlinear wave equations in two dimensions. Applied Numerical Mathematics: Transactions of IMACS, 170(??):298–320, December 2021. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421002294.

Holst:2023:GTF

[HL23] Michael Holst and Martin Licht. Geometric transformation of finite element methods: Theory and applications. Applied Numerical Mathematics: Transactions of IMACS, 192(??): 389–413, October 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001861.

Hou:2024:TLE

[HL24] Baohui Hou and Huan Liu.

Two linear energy-preserving compact finite difference schemes for coupled nonlinear wave equations. Applied Numerical Mathematics: Transactions of IMACS, 201(??):531–549, July 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000874.

Huang:2001:MTM

[HLC01] Jianguo Huang, Likang Li, and Jinru Chen. On mortartype Morley element method for plate bending problem. Applied Numerical Mathematics: Transactions of IMACS, 37(4):519–533, June 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/73/33/
32/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/73/33/32/article.
pdf.

Hussein:2016:MTD

[HLIS16] M. S. Hussein, D. Lesnic, M. I. Ivanchov, and H. A. Snitko. Multiple timedependent coefficient identification thermal problems with a free boundary. Applied Numerical Mathematics: Transactions of IMACS, 99(??):24–50, January 2016. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927415001282.

Hu:2020:ESI

[HLJ20] Yunxia Hu, Hongwei Li, and Ziwen Jiang. Efficient semiimplicit compact finite difference scheme for nonlinear Schrödinger equations on unbounded domain. Applied Numerical Mathematics: Transactions of IMACS, 153(??):319–343, July 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742030060X.

Huang:2009:MQM

[HLL09] Jin Huang, Tao Lü, and Zi Cai Li. Mechanical quadrature methods and their splitting extrapolations for boundary integral equations of first kind on open arcs. Applied Numerical Mathematics: Transactions of IMACS, 59(12): 2908–2922, December 2009.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hong:2006:GCP

[HLMKZ06] Jialin Hong, Ying Liu, Hans Munthe-Kaas, and Antonella Zanna. Globally conservative properties and error estimation of a multi-symplectic scheme for Schrödinger equations with variable coefficients. Applied Numerical Mathematics: Transactions of IMACS, 56(6): 814–843, June 2006. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hartmann:2009:EPD

[HLMP09] R. Hartmann, M. Lukáčová-Medvid’ová, and F. Prill. Efficient preconditioning for the discontinuous Galerkin finite element method by low-order elements. Applied Numerical Mathematics: Transactions of IMACS, 59(8):1737–1753, August 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Holder:2001:EVS

[HLR01] Thomas Holder, Ben Leimkuhler, and Sebastian Reich. Explicit variable step-size and time-reversible integration. Applied Numerical Mathematics: Transactions of IMACS, 39(3–4):367–377, December 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        10/28/79/34/34/abstract.
                                        html.

Hazra:2018:NSB

[HLR18] Arijit Hazra, Gert Lube, and Hans-Georg Raumer. Numerical simulation of Bloch equations for dynamic magnetic resonance imaging. Applied Numerical Mathematics: Transactions of IMACS, 123(??): 241–255, January 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417302015.

He:2007:LTS

[HLT07] Yinnian He, Yunxian Liu, and Tao Tang. On large time-stepping methods for the Cahn–Hilliard equation. Applied Numerical Mathematics: Transactions of IMACS, 57(5–7):616–628, May 2007. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hou:2025:BCW

[HLW25] Yongli Hou, Yi Liu, and Yanqiu Wang. A boundarycorrected weak Galerkin mixed finite method for elliptic interface problems with curved interfaces. Applied Numerical Mathematics: Transactions of IMACS, 218(??):201–219, December 2025. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425001576.

Huang:2004:NEE

[HLY04] Hung-Tsai Huang, Zi-Cai Li, and Ningning Yan. New error estimates of Adini’s elements for Poisson’s equation. Applied Numerical Mathematics: Transactions of IMACS, 50(1):49–74, July 2004. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hu:2022:WCL

[HLY22] Ye Hu, Changpin Li, and Yubin Yan. Weak convergence of the L1 scheme for a stochastic subdiffusion problem driven by fractionally integrated additive noise. Applied Numerical Mathematics: Transactions of IMACS, 178(??): 192–215, August 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422000976.

Huang:2006:NEE

[HLZ06] Hung-Tsai Huang, Zi-Cai Li, and Aihui Zhou. New error estimates of biquadratic Lagrange elements for Poisson’s equation. Applied Numerical Mathematics: Transactions of IMACS, 56(5): 712–744, May 2006. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hansbo:2014:CFE

[HLZ14] Peter Hansbo, Mats G. Larson, and Sara Zahedi. A cut finite element method for a Stokes interface problem. Applied Numerical Mathematics: Transactions of IMACS, 85(??):90– 114, November 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414001184.

Hariharan:1986:LFA

[HM86] S. I. Hariharan and R. C. Maccamy. Low frequency acoustic and electromagnetic scattering. Applied Numerical Mathematics: Transactions of IMACS, 2(1):29– 35, February 1986. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hairer:1987:NMS

[HM87] E. Hairer and P. Maass. Numerical methods for singular nonlinear integro-differential equations. Applied Numerical Mathematics: Transactions of IMACS, 3(3):

243–256, June 1987. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0168927487900511.

Herrero:2000:NMC

[HM00] H. Herrero and A. M. Mancho. Numerical modeling in Chebyshev collocation methods applied to stability analysis of convection problems. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4): 161–166, May 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/61/27/
40/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/61/27/40/article.
pdf.

Hasanov:2001:NMB

[HM01] Alemdar Hasanov and Jennifer L. Mueller. A numerical method for backward parabolic problems with non-selfadjoint elliptic operators. Applied Numerical Mathematics: Transactions of IMACS, 37(1–2): 55–78, April 2001. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/73/26/
29/abstract.html;   http:

                                        //www.elsevier.nl/gej-
                                        ng/10/10/28/73/26/29/article.
                                        pdf.

Hermann:2009:CES

[HM09] M. Hermann and Th. Milde. Computing eigenfunctions of singular points in nonlinear parametrized two-point BVPs. Applied Numerical Mathematics: Transactions of IMACS, 59(3–4): 671–676, March/April 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hasanov:2015:FCA

[HM15] Alemdar Hasanov and Balgaisha Mukanova. Fourier collocation algorithm for identifying the spacewisedependent source in the advection–diffusion equation from boundary data measurements. Applied Numerical Mathematics: Transactions of IMACS, 97(??):1– 14, November 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927415000926.

Hasanov:2017:FCA

[HM17] Alemdar Hasanov and Balgaisha Mukanova. Fourier Collocation Algorithm for identification of a spacewise dependent source in wave equation from Neumanntype measured data. Applied Numerical Mathematics: Transactions of IMACS, 111(??):49–63, January 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301738.

Habib:2021:GLM

[HM21] Yousaf Habib and Lubna Mustafa. General linear methods with projection. Applied Numerical Mathematics: Transactions of IMACS, 161(??): 46–51, March 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303366.

Huang:2022:IAD

[HM22] Wei Huang and Michael Multerer. Isogeometric analysis of diffusion problems on random surfaces. Applied Numerical Mathematics: Transactions of IMACS, 179(??):50–65, September 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0168927422001076.

Hendy:2021:NEC

[HMD21] Ahmed S. Hendy and J. E.

Macías-Díaz. On a nonlinear energy-conserving scalar auxiliary variable (SAV) model for Riesz spacefractional hyperbolic equations. Applied Numerical Mathematics: Transactions of IMACS, 165(??): 339–347, July 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000635.

Hennart:2003:TON

[HMdV03] J.-P. Hennart, E. H. Mund, and E. del Valle. Third order nodal finite element methods with transverse and reduced integration for elliptic problems. Applied Numerical Mathematics: Transactions of IMACS, 46(2):209– 230, August 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hassani:2020:ENT

[HMN20] H. Hassani, J. A. Tenreiro Machado, and E. Naraghirad. An efficient numerical technique for variable order time fractional nonlinear Klein–Gordon equation. Applied Numerical Mathematics: Transactions of IMACS, 154(??): 260–272, August 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420301161.

Hiptmair:2014:TDG

[HMP14] Ralf Hiptmair, Andrea Moiola, and Ilaria Perugia. Trefftz discontinuous Galerkin methods for acoustic scattering on locally refined meshes. Applied Numerical Mathematics: Transactions of IMACS, 79(??):79–91, May 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413000044.

Higueras:2025:LSE

[HMR25] I. Higueras, J. I. Montijano, and L. Rández. Lowstorage exponentially fitted explicit Runge–Kutta methods. Applied Numerical Mathematics: Transactions of IMACS, 217(??): 372–389, November 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425001400.

Higueras:2003:SPIa

[HMT03a] I. Higueras, R. März, and C. Tischendorf. Stability preserving integration of index-1 DAEs. Applied Numerical Mathematics: Transactions of IMACS, 45(2–3):175–200, May 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Higueras:2003:SPIb

[HMT03b] I. Higueras, R. März, and C. Tischendorf. Stability preserving integration of index-2 DAEs. Applied Numerical Mathematics: Transactions of IMACS, 45(2–3):201–229, May 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hueber:2005:PEE

[HMW05] S. Hüeber, M. Mair, and B. I. Wohlmuth. A priori error estimates and an inexact primal-dual active set strategy for linear and quadratic finite elements applied to multibody contact problems. Applied Numerical Mathematics: Transactions of IMACS, 54(3– 4):555–576, August 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Himonas:2019:NSE

[HMY19] A. Alexandrou Himonas, Dionyssios Mantzavinos, and Fangchi Yan. The nonlinear Schrödinger equation on the half-line with Neumann boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 141(??):2–18, ???? 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418302241.

Hagstrom:2003:AEB

[HN03] Thomas Hagstrom and Jan Nordström. Analysis of extrapolation boundary conditions for the linearized Euler equations. Applied Numerical Mathematics: Transactions of IMACS, 44(1– 2):95–108, January 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hutzenthaler:2022:MPA

[HN22] Martin Hutzenthaler and Tuan Anh Nguyen. Multilevel Picard approximations of high-dimensional semilinear partial differential equations with locally monotone coefficient functions. Applied Numerical Mathematics: Transactions of IMACS, 181(??): 151–175, November 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001325.

Hofmann:2017:NBE

[HNP17] Michael Hofmann, Franziska Nestler, and Michael Pippig. NFFT based Ewald summation for electrostatic systems with charges and dipoles. Applied Numerical Mathematics: Transactions of IMACS, 122(??):39– 65, December 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417301666.

Hochbruck:2005:ERK

[HO05] Marlis Hochbruck and Alexander Ostermann. Exponential Runge–Kutta methods for parabolic problems. Applied Numerical Mathematics: Transactions of IMACS, 53(2–4):323–339, May 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hansen:2010:UCD

[HO10] Eskil Hansen and Alexander Ostermann. Unconditional convergence of DIRK schemes applied to dissipative evolution equations. Applied Numerical Mathematics: Transactions of IMACS, 60(1–2):55–63, January/February 2010. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hubenthal:2016:SAA

[HO16] M. Hubenthal and D. Onofrei. Sensitivity analysis for active control of the Helmholtz equation. Applied Numerical Mathematics: Transactions of IMACS, 106(??): 1–23, August 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416300332.

Hegarty:2024:FFE

[HO24a] A. F. Hegarty and E. O’Riordan.

Fitted finite element methods for singularly perturbed elliptic problems of convection-diffusion type. Applied Numerical Mathematics: Transactions of IMACS, 196(??):183–198, February 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423002751.

Hegarty:2024:HON

[HO24b] A. F. Hegarty and E. O’Riordan.

A higher order numerical method for singularly perturbed elliptic problems with characteristic boundary layers. Applied Numerical Mathematics: Transactions of IMACS, 201(??): 85–101, July 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000369.

Hoang:2015:FFE

[Hoa15] N. S. Hoang. Functionallyfitted explicit pseudo twostep Runge–Kutta–Nyström methods. Applied Numerical Mathematics: Transactions of IMACS, 92(??): 111–131, June 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927415000215.

Harten:1986:SRU

[HOEC86] Ami Harten, Stanley Osher, Björn Engquist, and Sukumar R. Chakravarthy. Some results on uniformly high-order accurate essentially nonoscillatory schemes. Applied Numerical Mathematics: Transactions of IMACS, 2(3–5):347–377, October 1986. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Holstad:2001:KUS

[Hol01] Astrid Holstad. The Koren upwind scheme for variable gridsize. Applied Numerical Mathematics: Transactions of IMACS, 37(4): 459–487, June 2001. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/10/28/73/33/
                                        29/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/10/28/73/33/29/article.
                                        pdf.

Hoppe:2023:ACI

[Hop23] R. H. W. Hoppe. An adaptive C^{0} Interior Penalty Discontinuous Galerkin method and an equilibrated a posteriori error estimator for the von Kármán equations. Applied Numerical Mathematics: Transactions of IMACS, 190(??):27–49, August 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423000041.

Horvath:1993:CSS

[Hor93] Zoltán Horváth. Consistency and stability for some nonnegativity-conserving methods. Applied Numerical Mathematics: Transactions of IMACS, 13(5):371–381, December 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Horvath:1998:PRK

[Hor98] Zoltán Horváth. Positivity of Runge–Kutta and diagonally split Runge–Kutta methods. Applied Numerical Mathematics: Transactions of IMACS, 28(2–4):309–326, October 1, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1998/28/
2-4/916.pdf.

Horvath:1999:MNC

[Hor99]  Róbert Horváth.  Maxi-

mum norm contractivity in the numerical solution of the one-dimensional heat equation. Applied Numerical Mathematics: Transactions of IMACS, 31(4): 451–462, December 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/10/28/46/20/
                                        20/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/10/28/46/20/20/article.
                                        pdf.

Horvath:2002:MCN

[Hor02] Róbert Horváth. On the monotonicity conservation in numerical solutions of the heat equation. Applied Numerical Mathematics: Transactions of IMACS, 42(1–3):189–199, August 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Horvath:2005:PSS

[Hor05] Zoltán Horváth. On the positivity step size threshold of Runge–Kutta methods. Applied Numerical Mathematics: Transactions of IMACS, 53(2–4):341–356, May 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hardy:1999:SVS

[HOS99] David J. Hardy, Daniel I. Okunbor, and Robert D. Skeel. Symplectic variable step size integration for N-body problems. Applied Numerical Mathematics: Transactions of IMACS, 29(1):19–30, January 1, 1999. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1999/29/
1/897.pdf.

Holst:2011:ERR

[HOS11] Michael Holst, Jeffrey S.

Ovall, and Ryan Szypowski. An efficient, reliable and robust error estimator for elliptic problems in R^{3}. Applied Numerical Mathematics: Transactions of IMACS, 61(5):675–695, May 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hou:2023:MSP

[Hou23]  Baohui Hou.  Meshless

structure-preserving GRBF collocation methods for stochastic Maxwell equations with multiplicative noise. Applied Numerical Mathematics: Transactions of IMACS, 192(??): 337–355, October 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001848.

Hafez:1985:MLS

[HP85] Mohamed M. Hafez and Timothy N. Phillips. A modified least squares formulation for a system of first-order equations. Applied Numerical Mathematics: Transactions of IMACS, 1(4):339–347, July 1985. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hoffmann:1991:ILA

[HP91] Walter Hoffmann and Kitty Potma. Implementing linear algebra algorithms on a Meiko Computing Surface. Applied Numerical Mathematics: Transactions of IMACS, 8(2):127– 148, September 1991. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0168927491900474.

Hemker:1997:APO

[HP97] P. W. Hemker and C. Pflaum. Approximation on partially ordered sets of regular grids. Applied Numerical Mathematics: Transactions of IMACS, 25(1):55– 87, September 16, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=25&
issue=1&aid=799.

Hasanov:2014:UAI

[HP14] Alemdar Hasanov and Burhan Pektas. A unified approach to identifying an unknown spacewise dependent source in a variable coefficient parabolic equation from final and integral overdeterminations. Applied Numerical Mathematics: Transactions of IMACS, 78(??):49–67, April 2014. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413001645.

Huang:2015:LSS

[HP15] Mingxia Huang and Dingguo Pu. Line search SQP method with a flexible step acceptance procedure. Applied Numerical Mathematics: Transactions of IMACS, 92(??):98–110, June 2015. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415000197.

Harbrecht:2018:SOP

[HP18] Helmut Harbrecht and Michael D.

Peters. The second order perturbation approach for elliptic partial differential equations on random domains. Applied Numerical Mathematics: Transactions of IMACS, 125(??): 159–171, March 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417302374.

He:2020:SFO

[HPH20] Dongdong He, Kejia Pan, and Hongling Hu. A spatial fourth-order maximum principle preserving operator splitting scheme for the multi-dimensional fractional Allen–Cahn equation. Applied Numerical Mathematics: Transactions of IMACS, 151(??):44–63, May 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419303691.

Harbrecht:2012:LRA

[HPS12] Helmut Harbrecht, Michael Peters, and Reinhold Schneider. On the low-rank approximation by the pivoted Cholesky decomposition. Applied Numerical Mathematics: Transactions of IMACS, 62(4): 428–440, April 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927411001814.

Hoppe:2021:NSF

[HPW21] Ronald H. W. Hoppe, Basanta R. Pahari, and James J. Winkle. Numerical simulation of the formation of spherulites in polycrystalline binary mixtures. Applied Numerical Mathematics: Transactions of IMACS, 166(??):61–75, August 2021.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000982.

Hadjidimos:1992:CMA

[HPY92] A. Hadjidimos, A. Psimarni, and A. K. Yeyios. On the convergence of the modified accelerated overrelaxation (MAOR) method. Applied Numerical Mathematics: Transactions of IMACS, 10(2):115–127, July 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hendy:2024:OSG

[HQAZ24] A. S. Hendy, L. Qiao, A. Aldraiweesh, and M. A. Zaky. Optimal spectral Galerkin approximation for time and space fractional reaction-diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 201(??): 118–128, July 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000357.

Huang:1996:MCM

[HR96] Weizhang Huang and Robert D.

Russell. A moving collocation method for solving time dependent partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 20(1–2): 101–116, February 29, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=1-2&aid=667. Work-

shop on the method of lines for time-dependent problems (Lexington, KY, 1995).

Huang:1997:HDM

[HR97] Weizhang Huang and Robert D. Russell. A high dimensional moving mesh strategy. Applied Numerical Mathematics: Transactions of IMACS, 26(1–2):63–76, December 22, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1997/26/
1-2/837.pdf.

Higueras:2006:SVP

[HR06] Inmaculada Higueras and Teo Roldán. Stage value predictors for additive and partitioned Runge–Kutta methods. Applied Numerical Mathematics: Transactions of IMACS, 56(1):1–18, January 2006. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hearn:2014:FCC

[HR14] Tristan A. Hearn and Lothar Reichel. Fast computation of convolution operations via low-rank approximation. Applied Numerical Mathematics: Transactions of IMACS, 75(??):136– 153, January 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413000822.

Hemker:1986:MGO

[HS86] P. W. Hemker and S. P.

Spekreijse. Multiple grid and Osher’s scheme for the efficient solution of the steady Euler equations. Applied Numerical Mathematics: Transactions of IMACS, 2(6):475–493, December 1986. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Higham:1995:ESF

[HS95] Desmond J. Higham and Tasneem Sardar. Existence and stability of fixed points for a discretised nonlinear reaction-diffusion equation with delay. Applied Numerical Mathematics: Transactions of IMACS, 18(1–3):155–173, September 27, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=18&
issue=1-3&aid=599. Sev-

enth Conference on the Numerical Treatment of Differential Equations (Halle, 1994).

Hosea:1996:AIT

[HS96] M. E. Hosea and L. F. Shampine. Analysis and implementation of TR-BDF2. Applied Numerical Mathematics: Transactions of IMACS, 20(1–2): 21–37, February 29, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=1-2&aid=663. Work-

shop on the method of lines for time-dependent problems (Lexington, KY, 1995).

Hyman:1997:AON

[HS97] James M. Hyman and Mikhail Shashkov. Adjoint operators for the natural discretizations of the divergence, gradient and curl on logically rectangular grids. Applied Numerical Mathematics: Transactions of IMACS, 25(4):413– 442, November 10, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=25&
issue=4&aid=853.

Hilden:1998:EDM

[HS98] Michael Hilden and Gerd Steinebach. ENO-discretizations in MOL-applications: some examples in river hydraulics. Applied Numerical Mathematics: Transactions of IMACS, 28(2–4):293–308, October 1, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1998/28/
2-4/915.pdf.

Hafner:2002:IDV

[HS02] Hartmut Häfner and Willi Schönauer. The integration of different variants of the (I)LU algorithm in the LINSOL program package. Applied Numerical Mathematics: Transactions of IMACS, 41(1):39– 59, April 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/86/27/30/abstract.
html.

Han:2007:DCP

[HS07] Weimin Han and M. Sofonea. On a dynamic contact problem for elasticvisco-plastic materials. Applied Numerical Mathematics: Transactions of IMACS, 57(5–7):498–509, May 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hoang:2009:FFE

[HS09a] Nguyen S. Hoang and Roger B. Sidje. Functionally fitted explicit pseudo twostep Runge–Kutta methods. Applied Numerical Mathematics: Transactions of IMACS, 59(1):39–55, January 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hundsdorfer:2009:AMT

[HS09b] Willem Hundsdorfer and Valeriu Savcenco. Analysis of a multirate theta-method for stiff ODEs. Applied Numerical Mathematics: Transactions of IMACS, 59(3–4): 693–706, March/April 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hansen:2011:ODC

[HS11] Anders C. Hansen and John Strain. On the order of deferred correction. Applied Numerical Mathematics: Transactions of IMACS, 61(8):961–973, August 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hessari:2017:FOS

[HS17] Peyman Hessari and Byeong-Chun Shin. First order system least squares pseudo-spectral method for Stokes–Darcy equations. Applied Numerical Mathematics: Transactions of IMACS, 120(??):35–52, October 2017. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741730106X.

Hesameddini:2019:SMP

[HS19a] Esmail Hesameddini and Mehdi Shahbazi. Solving multipoint problems with linear Volterra–Fredholm integro-differential equations of the neutral type using Bernstein polynomials method. Applied Numerical Mathematics: Transactions of IMACS, 136(??): 122–138, February 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418302289.

Huang:2019:DDG

[HS19b] Chaobao Huang and Martin Stynes. A direct discontinuous Galerkin method for a time-fractional diffusion equation with a Robin boundary condition. Applied Numerical Mathematics: Transactions of IMACS, 135(??):15–29, January 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418301740.

He:2020:EPF

[HS20] Mingyan He and Pengtao Sun. Energy-preserving finite element methods for a class of nonlinear wave equations. Applied Numerical Mathematics: Transactions of IMACS, 157(??): 446–469, November 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420301987.

Hadj:2021:IEM

[HS21a] Abdelhak Hadj and Hacene Saker. Integral equations method for solving a Biharmonic inverse problem in detection of Robin coefficients. Applied Numerical Mathematics: Transactions of IMACS, 160(??): 436–450, February 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303123.

See corrigendum [AS21].

Huang:2021:PBS

[HS21b] Xin Huang and Hai-Wei Sun. A preconditioner based on sine transform for two-dimensional semi-linear Riesz space fractional diffusion equations in convex domains. Applied Numerical Mathematics: Transactions of IMACS, 169(??): 289–302, November 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742100194X.

Heid:2022:AIL

[HS22] Pascal Heid and Endre Süli. An adaptive iterative linearised finite element method for implicitly constituted incompressible fluid flow problems and its application to Bingham fluids. Applied Numerical Mathematics: Transactions of IMACS, 181(??): 364–387, November 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001696.

Hosseinzadeh:2023:NML

[HS23] Hossein Hosseinzadeh and Ahmad Shirzadi. A new meshless local integral equation method. Applied Numerical Mathematics: Transactions of IMACS, 194(??):

44–58, December 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423002258.

Hofer:2024:HTG

[HS24] Jacob S. Hofer and Scott A. Strong. Hasimoto transformation of general flows expressed in the Frenet frame. Applied Numerical Mathematics: Transactions of IMACS, 199(??): 26–31, May 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423000120.

Huang:2025:CIF

[HSF25] Xinran Huang, Haiyan Su, and Xinlong Feng. H(div)- conforming IPDG FEM with pointwise divergence-free velocity field for the micropolar Navier–Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 218(??): 109–127, December 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425001485.

Hsiao:2006:BEM

[Hsi06] George C. Hsiao. Boundary element methods—an overview. Applied Numerical Mathematics: Transactions of IMACS, 56(10–11):1356– 1369, October/November 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hasler:2004:MFE

[HSS04] Urs Hasler, Anna Schneebeli, and Dominik Schötzau. Mixed finite element approximation of incompressible MHD problems based on weighted regularization. Applied Numerical Mathematics: Transactions of IMACS, 51(1):19–45, October 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Herty:2007:DDM

[HSS07] Michael Herty, Mohammed Seaïd, and Anita K. Singh. A domain decomposition method for conservation laws with discontinuous flux function. Applied Numerical Mathematics: Transactions of IMACS, 57(4):361–373, April 2007. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hussain:2014:ENM

[HST14] S. Hussain, F. Schieweck, and S. Turek. Efficient Newton-multigrid solution techniques for higher order space–time Galerkin discretizations of incompressible flow. Applied Numerical Mathematics: Transactions of IMACS, 83(??):51– 71, September 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414000713.

Hafner:1999:PPL

[HSW99] Hartmut Häfner, Willi Schönauer, and Rüdiger Weiss. The program package LINSOL: Basic concepts and realization. Applied Numerical Mathematics: Transactions of IMACS, 30(2– 3):213–224, June 10, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1999&volume=30&
issue=2-3&aid=976.

Hohenegger:2024:NTS

[HSWW24] Matthias Hohenegger, Giuseppina Settanni, Ewa B. Weinmüller, and Mered Wolde. Numerical treatment of singular ODEs using finite difference and collocation methods. Applied Numerical Mathematics: Transactions of IMACS, 205(??): 184–194, November 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001740.

He:2018:IWC

[HSX18] Xiaoming He, Jie Shen, and Chuanju Xu. 20th IMACS World Congress. Applied Numerical Mathematics: Transactions of IMACS, 133(??):1, ???? 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418301582.

Hwang:2018:IAM

[HSY18] Feng-Nan Hwang, Yi-Zhen Su, and Chien-Chou Yao. An iteratively adaptive multiscale finite element method for elliptic interface problems. Applied Numerical Mathematics: Transactions of IMACS, 127(??): 211–225, May 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418300187.

Hundsdorfer:1994:MLD

[HT94] W. Hundsdorfer and R. A.

Trompert. Method of lines and direct discretization: a comparison for linear advection. Applied Numerical Mathematics: Transactions of IMACS, 13(6):469– 490, February 1994. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Huang:2000:PSS

[HT00] Weizhang Huang and Tao Tang. Pseudospectral solutions for steady motion of a viscous fluid inside a circular boundary. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4):167–173, May 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/61/27/
41/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/61/27/41/article.
pdf.

Hosseini:2019:SSM

[HT19] Rasool Hosseini and Mehdi Tatari. Some splitting methods for hyperbolic PDEs. Applied Numerical Mathematics: Transactions of IMACS, 146(??):361–378, December 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927419301771.

Huang:2020:PDF

[HT20] Wenli Huang and Yuchao Tang. Primal-dual fixed point algorithm based on adapted metric method for solving convex minimization problem with application. Applied Numerical Mathematics: Transactions of IMACS, 157(??):

236–254, November 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420301872.

Halder:2024:HON

[HT24] Joydev Halder and Suman Kumar Tumuluri. A higher order numerical scheme to a nonlinear McKendrick– Von Foerster equation with singular mortality. Applied Numerical Mathematics: Transactions of IMACS, 202(??):21–41, August 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000941.

He:2023:ECF

[HTSZ23] Mingyan He, Jia Tian, Pengtao Sun, and Zhengfang Zhang. An energyconserving finite element method for nonlinear fourthorder wave equations. Applied Numerical Mathematics: Transactions of IMACS, 183(??):333–354, January 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422002513.

Hu:1999:MCD

[Hu99] Qiya Hu. Multilevel correction for discrete collocation solutions of Volterra integral equations with delay arguments. Applied Numerical Mathematics: Transactions of IMACS, 31(2): 159–171, October 23, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1999&volume=31&
issue=2&aid=992;   http:
//www.sciencedirect.com/
science/article/pii/S0168927498001275.

Huang:1998:EAS

[Hua98] Jie Huang. An efficient algorithm to solve a sequence of linear equations arising in nonlinear H^{∞} control. Applied Numerical Mathematics: Transactions of IMACS, 26(3):293–306, March 2, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1998/26/
3/857.pdf.

Huang:2000:DOL

[Hua00] Chengming Huang. Dissipativity of one-leg methods for dynamical systems with delays. Applied Numerical Mathematics: Transactions of IMACS, 35(1):11– 22, September 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/64/27/
28/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/64/27/28/article.
pdf.

Huang:2009:SSP

[Hua09] Chengming Huang. Strong stability preserving hybrid methods. Applied Numerical Mathematics: Transactions of IMACS, 59(5): 891–904, May 2009. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Huang:2017:ETL

[Hua17] Pengzhan Huang. An efficient two-level finite element algorithm for the natural convection equations. Applied Numerical Mathematics: Transactions of IMACS, 118(??):75–86, August 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417300570.

Huang:2019:ASW

[Hua19] Rong Huang. Accurate solutions of weighted least squares problems associated with rank-structured matrices. Applied Numerical Mathematics: Transactions of IMACS, 146(??): 416–435, December 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892741930203X.

Huang:2020:FTM

[Hua20] Aiqun Huang. A filter-type method for solving nonlinear semidefinite programming. Applied Numerical Mathematics: Transactions of IMACS, 158(??): 415–424, December 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420302476.

Huang:2021:PAM

[Hua21] Aiqun Huang. A proximal augmented method for semidefinite programming problems. Applied Numerical Mathematics: Transactions of IMACS, 169(??): 321–333, November 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421001987.

Huckle:1999:ASP

[Huc99]  Thomas Huckle.  Ap-

proximate sparsity patterns for the inverse of a matrix and preconditioning. Applied Numerical Mathematics: Transactions of IMACS, 30(2–3):291–303, June 10, 1999. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1999&volume=30&
issue=2-3&aid=982.

Hundsdorfer:2002:ASS

[Hun02] Willem Hundsdorfer. Accuracy and stability of splitting with stabilizing corrections. Applied Numerical Mathematics: Transactions of IMACS, 42(1– 3):213–233, August 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hussein:2020:WGF

[Hus20] Ahmed Jabbar Hussein. A weak Galerkin finite element method for solving timefractional coupled Burgers’ equations in two dimensions. Applied Numerical Mathematics: Transactions of IMACS, 156(??): 265–275, October 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420301318.

Hundsdorfer:1989:LSH

[HV89] W. H. Hundsdorfer and J. G. Verwer. Linear stability of the hopscotch scheme. Applied Numerical Mathematics: Transactions of IMACS, 5(5):423– 433, September 1989. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hundsdorfer:1995:NSE

[HV95] W. Hundsdorfer and J. G. Verwer. A note on splitting errors for advectionreaction equations. Applied Numerical Mathematics: Transactions of IMACS, 18(1–3):191–199, September 27, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=18&
issue=1-3&aid=617. Sev-

enth Conference on the Numerical Treatment of Differential Equations (Halle, 1994).

He:2022:FMB

[HV22] Jiewen He and Seakweng Vong. Fast modulusbased matrix splitting iteration methods for implicit complementarity problems. Applied Numerical Mathematics: Transactions of IMACS, 182(??):28– 41, December 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001866.

Hooghiemstra:2010:ASL

[HvdHV10] P. B. Hooghiemstra, D. R.

van der Heul, and C. Vuik. Application of the shifted-Laplace preconditioner for iterative solution of a higher order finite element discretisation of the vector wave equation: First experiences. Applied Numerical Mathematics: Transactions of IMACS, 60(11): 1157–1170, November 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Heroux:1991:PPC

[HVY91] Michael A. Heroux, Phuong Vu, and Chao Yang. A parallel preconditioned conjugate gradient package for solving sparse linear systems on a Cray Y-MP. Applied Numerical Mathematics: Transactions of IMACS, 8(2):93– 115, September 1991. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0168927491900452.

Hromadka:1993:ECA

[HW93] T. V. Hromadka II and R. J. Whitley. Expanding the CVBEM approximation in a series. Applied Numerical Mathematics: Transactions of IMACS, 11(6):509–516, 1993. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hoppe:1997:MIS

[HW97] Ronald H. W. Hoppe and Barbara Wohlmuth. Multilevel iterative solution and adaptive mesh refinement for mixed finite element discretizations. Applied Numerical Mathematics: Transactions of IMACS, 23(1):97– 117, March 7, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1997/23/

1/745. Multilevel methods (Oberwolfach, 1995).

Hill:2004:ANP

[HW04] A. T. Hill and W. L. Wan. Analysis and numerics for a parabolic equation with impulsive forcing. Applied Numerical Mathematics: Transactions of IMACS, 50(3–4):445–474, September 2004. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Heineken:2006:PMR

[HW06] Wolfram Heineken and Gerald Warnecke. Partitioning methods for reaction– diffusion problems. Applied Numerical Mathematics: Transactions of IMACS, 56(7):981–1000, July 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hong:2015:TSN

[HW15] Youngjoon Hong and Djoko Wirosoetisno. Timestepping schemes for the 3 d Navier–Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 96(??):153–164, October 2015. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741500094X.

Hu:2022:GCM

[HW22] Shenglong Hu and Qun Wang. A globally convergent method for solving a quartic generalized Markowitz portfolio problem. Applied Numerical Mathematics: Transactions of IMACS, 179(??): 255–272, September 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001234.

Hu:2015:AMM

[HWCF15] Fuxing Hu, Rong Wang, Xueyong Chen, and Hui Feng. An adaptive mesh method for 1D hyperbolic conservation laws. Applied Numerical Mathematics: Transactions of IMACS, 91(??):11–25, May 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414002098.

Hu:2020:SMH

[HWY20] Wujie Hu, Jinzhao Wu, and Gonglin Yuan. Some modified Hestenes–Stiefel conjugate gradient algorithms with application in image restoration. Applied Numerical Mathematics: Transactions of IMACS, 158(??): 360–376, December 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420302440.

Hou:2022:PPU

[HWZ22] Dianming Hou, Hui Wang, and Chao Zhang. Positivitypreserving and unconditionally energy stable numerical schemes for MEMS model. Applied Numerical Mathematics: Transactions of IMACS, 181(??): 503–517, November 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422001763.

Hsiao:2011:SBI

[HX11] George C. Hsiao and Liwei Xu. A system of boundary integral equations for the transmission problem in acoustics. Applied Numerical Mathematics: Transactions of IMACS, 61(9): 1017–1029, September 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hong:2015:PQI

[HXW15] Jialin Hong, Dongsheng Xu, and Peng Wang. Preservation of quadratic invariants of stochastic differential equations via Runge–Kutta methods. Applied Numerical Mathematics: Transactions of IMACS, 87(??): 38–52, January 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414001421.

Hu:2001:SSP

[HY01] Qiya Hu and Dehao Yu.

Solving singularity problems in unbounded domains by coupling of natural BEM and composite grid FEM. Applied Numerical Mathematics: Transactions of IMACS, 37(1–2):127–143, April 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/73/26/
33/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/73/26/33/article.
pdf.

Henson:2002:BPA

[HY02] Van Emden Henson and Ulrike Meier Yang. Boomer-AMG: a parallel algebraic multigrid solver and preconditioner. Applied Numerical Mathematics: Transactions of IMACS, 41(1):155– 177, April 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        10/28/86/27/37/abstract.
                                        html.

Himonas:2024:SKV

[HY24a] A. Alexandrou Himonas and Fangchi Yan. The Schrödinger–Korteweg-de Vries system on the half-line. Applied Numerical Mathematics: Transactions of IMACS, 199(??):32–58, May 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422003403.

Huang:2024:ILJ

[HY24b] Qiumei Huang and Huiting Yang. Implicitly linear Jacobi spectral-collocation methods for two-dimensional weakly singular Volterra– Hammerstein integral equations. Applied Numerical Mathematics: Transactions of IMACS, 201(??): 159–174, July 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000497.

Yang:1986:MLF

[hYK86] Zhong hua Yang and H. B.

Keller. Multiple laminar flows through curved pipes. Applied Numerical Mathematics: Transactions of IMACS, 2(3–5):257–271, October 1986. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Yu:2012:JSM

[hYqW12] Xu hong Yu and Zhong qing Wang. Jacobi spectral method with essential imposition of Neumann boundary condition. Applied Numerical Mathematics: Transactions of IMACS, 62(8):956–974, August 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412000529.

Hairer:1996:EGF

[HZ96] E. Hairer and M. Zennaro. On error growth functions of Runge–Kutta methods. Applied Numerical Mathematics: Transactions of IMACS, 22(1–3): 205–216, December 4, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-

                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=22&
                                        issue=1-3&aid=714. Special

issue celebrating the centenary of Runge–Kutta methods.

Huang:2002:CEE

[HZ02] Jianguo Huang and Jun Zou. Construction of explicit extension operators on general finite element grids. Applied Numerical Mathematics: Transactions of IMACS, 43(3): 211–227, November 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Han:2009:AFD

[HZ09] Houde Han and Zhiwen Zhang. An analysis of the finite-difference method for one-dimensional Klein– Gordon equation on unbounded domain. Applied Numerical Mathematics: Transactions of IMACS, 59(7):1568–1583, July 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Huang:2012:DFF

[HZ12] Jianguo Huang and Shangyou Zhang. A divergence-free finite element method for a type of 3D Maxwell equations. Applied Numerical Mathematics: Transactions of IMACS, 62(6):

802–813, June 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411001036.

Hendy:2020:GCA

[HZ20] Ahmed S. Hendy and Mahmoud A. Zaky. Global consistency analysis of L1– Galerkin spectral schemes for coupled nonlinear space– time fractional Schrödinger equations. Applied Numerical Mathematics: Transactions of IMACS, 156(??): 276–302, October 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420301458.

Hao:2021:FSP

[HZ21] Zhaopeng Hao and Zhongqiang Zhang. Fast spectral Petrov– Galerkin method for fractional elliptic equations. Applied Numerical Mathematics: Transactions of IMACS, 162(??):318–330, April 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000015.

Huang:2021:NMT

[HZAT21] Jianfei Huang, Jingna Zhang, Sadia Arshad, and Yifa Tang. A numerical method for two-dimensional multiterm time-space fractional nonlinear diffusion-wave equations. Applied Numerical Mathematics: Transactions of IMACS, 159(??): 159–173, January 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302798.

Han:2005:NSP

[HZBM05] Houde Han, Liang Zhu, Hermann Brunner, and Jingtang Ma. The numerical solution of parabolic Volterra integro-differential equations on unbounded spatial domains. Applied Numerical Mathematics: Transactions of IMACS, 55(1):83– 99, September 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Hu:2022:GCD

[HZC22] Qingjie Hu, Hongrun Zhang, and Yu Chen. Global convergence of a descent PRP type conjugate gradient method for nonconvex optimization. Applied Numerical Mathematics: Transactions of IMACS, 173(??): 38–50, March 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421003081.

Hu:2023:TTT

[HZCZ23] Qingjie Hu, Liping Zhu, Cuili Chang, and Wenqi Zhang. A truncated threeterm conjugate gradient method with complexity guarantees with applications to nonconvex regression problem. Applied Numerical Mathematics: Transactions of IMACS, 194(??):82– 96, December 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423002246.

Hendy:2021:GFN

[HZD21] Ahmed S. Hendy, Mahmoud A. Zaky, and Rob H. De Staelen. A general framework for the numerical analysis of high-order finite difference solvers for nonlinear multi-term timespace fractional partial differential equations with time delay. Applied Numerical Mathematics: Transactions of IMACS, 169(??): 108–121, November 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001781.

Hu:2024:NIR

[HZZ24] Xinlin Hu, Qisheng Zheng, and Kai Zhang. Nonnegative iterative reweighted method for sparse linear complementarity problem. Applied Numerical Mathematics: Transactions of IMACS, 203(??):16– 31, September 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001211.

Hou:2025:LSO

[HZZ25] Dianming Hou, Tianxiang Zhang, and Hongyi Zhu. A linear second order unconditionally maximum bound principle-preserving scheme for the Allen–Cahn equation with general mobility. Applied Numerical Mathematics: Transactions of IMACS, 207(??):222– 243, January 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424002423.

Ibrahim:2025:DFP

[IAH25] Abdulkarim Hassan Ibrahim and Suliman Al-Homidan. A derivative-free projection method with double inertial effects for solving nonlinear equations. Applied Numerical Mathematics: Transactions of IMACS, 209(??): 55–67, March 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424003015.

Izadi:2024:EQB

[IB24] Mohammad Izadi and Dumitru Baleanu. An effective QLM-based Legendre matrix algorithm to solve the coupled system of fractionalorder Lane–Emden equations. Applied Numerical Mathematics: Transactions of IMACS, 201(??): 608–627, July 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423003045.

Ibrahimoglu:2025:NFA

[Ibr25]  B. Ali Ibrahimoglu.  A

new fast algorithm for computing the mock-Chebyshev nodes. Applied Numerical Mathematics: Transactions of IMACS, 208 (part A) (??):246–255, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927424000527.

Ioriatti:2019:PSC

[ID19] Matteo Ioriatti and Michael Dumbser. A posteriori subcell finite volume limiting of staggered semi-implicit discontinuous Galerkin schemes for the shallow water equations. Applied Numerical Mathematics: Transactions of IMACS, 135(??): 443–480, January 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418301946.

Igarashi:1985:PPA

[Iga85] Masao Igarashi. Practical problems arising for finding roots of nonlinear equations. Applied Numerical Mathematics: Transactions of IMACS, 1(5):433– 455, September 1985. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Isic:2013:ASS

[IHS13] M. Isic, Z. Horvat, and M. Spasojevic. Advection step in the split-operator approach applied to river modeling. Applied Numerical Mathematics: Transactions of IMACS, 72(??): 1–18, October 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927413000664.

Imperiale:2014:MME

[IJ14] Sébastien Imperiale and Patrick Joly. Mathematical modeling of electromagnetic wave propagation in heterogeneous lossy coaxial cables with variable cross section. Applied Numerical Mathematics: Transactions of IMACS, 79(??): 42–61, May 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413001207.

Izzo:2017:GLM

[IJ17a] Giuseppe Izzo and Zdzislaw Jackiewicz. Generalized linear multistep methods for ordinary differential equations. Applied Numerical Mathematics: Transactions of IMACS, 114(??): 165–178, April 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416300575.

Izzo:2017:HSI

[IJ17b] Giuseppe Izzo and Zdzislaw Jackiewicz. Highly stable implicit-explicit Runge– Kutta methods. Applied Numerical Mathematics: Transactions of IMACS, 113(??):71–92, March 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416302203.

Izzo:2021:CSM

[IJ21] Giuseppe Izzo and Zdzislaw Jackiewicz. Construction of SDIRK methods with dispersive stability functions. Applied Numerical Mathematics: Transactions of IMACS, 160(??): 265–280, February 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420303172.

Ige:2024:PTW

[IK24] Olufemi Elijah Ige and Henrik Kalisch. Particle trajectories in a weakly nonlinear long-wave model on a shear flow. Applied Numerical Mathematics: Transactions of IMACS, 199(??): 59–72, May 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742300017X.

Issaoui:2023:ANS

[IKM23] M. Issaoui, M. Kzaz, and F. Maach. Asymptoticnumerical solvers for diffusion equation with timelike highly oscillatory forcing terms. Applied Numerical Mathematics: Transactions of IMACS, 186(??): 114–128, April 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423000053.

Izgin:2023:DFS

[IKMM23] Thomas Izgin, Stefan Kopecz, Angela Martiradonna, and Andreas Meister. On the dynamics of first and second order GeCo and gBBKS schemes. Applied Numerical Mathematics: Transactions of IMACS, 193(??):43– 66, November 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001976.

Ibrahim:2022:AMI

[IKR^{+}22] A. H. Ibrahim, P. Kumam, S. Rapajić, Z. Papp, and A. B. Abubakar. Approximation methods with inertial term for large-scale nonlinear monotone equations. Applied Numerical Mathematics: Transactions of IMACS, 181(??): 417–435, November 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001738.

Iordanis:2025:ECP

[IKS25] Ioannis Iordanis, Christos Koukouvinos, and Iliana Silou. On the efficacy of conditioned and progressive Latin hypercube sampling in supervised machine learning. Applied Numerical Mathematics: Transactions of IMACS, 208 (part A) (??):256–270, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423003240.

Ionescu:2019:BVM

[ILS19] Ioan R. Ionescu and Oana Lupascu-Stamate. Boundary variation method for the generalized Cheeger problem. Applied Numerical Mathematics: Transactions of IMACS, 140(??): 199–214, June 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892741930039X.

Iavernaro:1998:SOD

[IM98] F. Iavernaro and F. Mazzia. Solving ordinary differential equations by generalized Adams methods: properties and implementation techniques. Applied Numerical Mathematics: Transactions of IMACS, 28(2–4):107–126, October 1, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1998/28/
                                        2-4/905.pdf.

Ivanenko:2000:AGS

[IM00] Sergey A. Ivanenko and Galina V. Muratova. Adaptive grid shallow water modeling. Applied Numerical Mathematics: Transactions of IMACS, 32(4): 447–482, April 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/59/32/
34/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/59/32/34/article.
pdf.

Iavernaro:2002:PIP

[IM02] F. Iavernaro and F. Mazzia.

Parallel implicit predictor corrector methods. Applied Numerical Mathematics: Transactions of IMACS, 42(1–3):235–250, August 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

intHout:2013:SAS

[iM13] K. J. in ’t Hout and C. Mishra. Stability of ADI schemes for multidimensional diffusion equations with mixed derivative terms. Applied Numerical Mathematics: Transactions of IMACS, 74(??):83– 94, December 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413001141.

Ion:2022:NSS

[IMC22] Stelian Ion, Dorin Marinescu, and Stefan-Gicu Cruceanu. Numerical scheme for solving a porous Saint-Venant type model for water flow on vegetated hillslopes. Applied Numerical Mathematics: Transactions of IMACS, 172(??):67– 98, February 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421002737.

Ivanenko:2004:NNA

[IMM04] Sergey A. Ivanenko, Stanislav S. Makhanov, and Mud-Armeen Munlin. New numerical algorithms to optimize cutting operations of a fiveaxis milling machine. Applied Numerical Mathematics: Transactions of IMACS, 49(3–4):395–413, June 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Iavernaro:2020:CSP

[IMMS20] F. Iavernaro, F. Mazzia, M. S. Mukhametzhanov, and Ya.D. Sergeyev. Conjugatesymplecticity properties of Euler–Maclaurin methods and their implementation on the Infinity Computer. Applied Numerical Mathematics: Transactions of IMACS, 155(??):58–72, ???? 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419301631.

Imoro:2000:DOP

[Imo00] Baba Imoro. Discretized obstacle problems with penalties on nested grids. Applied Numerical Mathematics: Transactions of IMACS, 32(1):21–34, January 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/59/28/
28/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/59/28/28/article.
pdf.

intHout:1992:SCR

[in 92] K. J. in ’t Hout. The stability of a class of Runge– Kutta methods for delay differential equations. Applied Numerical Mathematics: Transactions of IMACS, 9(3–5):347–355, April 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/016892749290026A.

International Conference on the Numerical Solution of Volterra and Delay Equations (Tempe, AZ, 1990).

intHout:1995:CWR

[in 95] K. J. in ’t Hout. On the convergence of waveform relaxation methods for stiff nonlinear ordinary differential equations. Applied Numerical Mathematics: Transactions of IMACS, 18(1–3):175–190, September 27, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=18&
issue=1-3&aid=600. Sev-

enth Conference on the Numerical Treatment of Differential Equations (Halle, 1994).

intHout:1996:SAR

[in 96] K. J. in ’t Hout. On the stability of adaptations of Runge–Kutta methods to systems of delay differential equations. Applied Numerical Mathematics: Transactions of IMACS, 22(1–3): 237–250, December 4, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=22&
issue=1-3&aid=717. Special

issue celebrating the centenary of Runge–Kutta methods.

intHout:2002:CIE

[in 02] K. J. in ’t Hout. On the contractivity of implicit–explicit linear multistep methods. Applied Numerical Mathematics: Transactions of IMACS, 42(1–3):201–212, August 2002. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Iserles:1989:OSR

[IN89] A. Iserles and S. P. Nørsett.

Order stars and rational approximants to exp(z). Applied Numerical Mathematics: Transactions of IMACS, 5(1–2):63–70, February 1989. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Recent theoretical results in numerical ordinary differential equations.

Iserles:2001:F

[IN01] A. Iserles and S. P. Nørsett.

Foreword. Applied Numerical Mathematics: Transactions of IMACS, 39(3–4): 239–244, December 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/10/28/79/34/27/abstract.
html.

Iserles:2001:TSH

[INR01] A. Iserles, S. P. Nørsett, and A. F. Rasmussen. Time symmetry and high-order Magnus methods. Applied Numerical Mathematics: Transactions of IMACS, 39(3–4):379–401, December 2001. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/

10/28/79/34/35/abstract.
html.

Ideon:2018:QLR

[IO18] Erge Ideon and Peeter Oja. Quadratic/linear rational spline collocation for linear boundary value problems. Applied Numerical Mathematics: Transactions of IMACS, 125(??): 143–158, March 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417302404.

Ioakimidis:1989:KGQ

[Ioa89] Nikolaos I. Ioakimidis. On Kutt’s Gaussian quadrature rule for finite-part integrals. Applied Numerical Mathematics: Transactions of IMACS, 5(3):209– 213, May 1989. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Isenberg:2002:ODS

[IPL02] J. Isenberg, V. Pereyra, and D. Lawver. Optimal design of steel frame structures. Applied Numerical Mathematics: Transactions of IMACS, 40(1–2):59–71, January 2002. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/85/27/33/abstract.
html.

Izadi:2022:HAC

[IR22] Mohammad Izadi and Pradip Roul. A highly accurate and computationally efficient technique for solving the electrohydrodynamic flow in a circular cylindrical conduit. Applied Numerical Mathematics: Transactions of IMACS, 181(??): 110–124, November 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001477.

Izzo:2012:HSR

[IRC12] G. Izzo, E. Russo, and C. Chiapparelli. Highly stable Runge–Kutta methods for Volterra integral equations. Applied Numerical Mathematics: Transactions of IMACS, 62(8):1002– 1013, August 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412000633.

Iyiola:2022:CRT

[IS22] Olaniyi S. Iyiola and Yekini Shehu. Convergence results of two-step inertial proximal point algorithm. Applied Numerical Mathematics: Transactions of IMACS, 182(??):57–75, December 2022. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422001878.

Iyiola:2023:FAM

[IS23] Olaniyi S. Iyiola and Yekini Shehu. A fully adaptive method for variational inequalities with quasimonotonicity. Applied Numerical Mathematics: Transactions of IMACS, 194(??): 18–43, December 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423002234.

Iserles:1994:CAD

[Ise94] A. Iserles. Convergence acceleration as a dynamical system. Applied Numerical Mathematics: Transactions of IMACS, 15(2):101– 121, September 13, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1994&volume=15&
                                        issue=2&aid=496. Innova-

tive methods in numerical analysis (Bressanone, 1992).

Iserles:1997:EDS

[Ise97] Arieh Iserles. Exact and discretized stability of the pantograph equation. Applied Numerical Mathematics: Transactions of IMACS, 24(2–3):295–308, June 19, 1997. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=24&
issue=2-3&aid=783. Volterra

centennial (Tempe, AZ, 1996).

Iserles:2002:TGA

[Ise02] Arieh Iserles. Think globally, act locally: Solving highly-oscillatory ordinary differential equations. Applied Numerical Mathematics: Transactions of IMACS, 43(1–2):145–160, October 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Iskandar:1989:NNS

[Isk89] Labib Iskandar. New numerical solution of the Korteweg– de Vries equation. Applied Numerical Mathematics: Transactions of IMACS, 5(3):215–221, May 1989. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ito:2007:FIS

[IT07] K. Ito and J. Toivanen.

A fast iterative solver for scattering by elastic objects in layered media. Applied Numerical Mathematics: Transactions of IMACS, 57(5–7):811–820, May 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ismailov:2016:ICP

[IT16] Mansur I. Ismailov and Ibrahim Tekin. Inverse coefficient problems for a first order hyperbolic system. Applied Numerical Mathematics: Transactions of IMACS, 106(??): 98–115, August 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416300320.

Hout:2018:ASV

[itHT18] Karel J. in ’t Hout and Jari Toivanen. ADI schemes for valuing European options under the Bates model. Applied Numerical Mathematics: Transactions of IMACS, 130(??):143–156, August 2018. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418300928.

Ito:2017:EEE

[Ito17] Kazuho Ito. An error estimate for an energy conserving spectral scheme approximating the dynamic elastica with free ends. Applied Numerical Mathematics: Transactions of IMACS, 120(??):1–20, October 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417301095.

Ito:2022:CSG

[Ito22] Kazuho Ito. Convergence of a spectral Galerkin scheme for the dynamic elastica with periodical or hinged boundary. Applied Numerical Mathematics: Transactions of IMACS, 176(??): 56–82, June 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422000435.

Isik:2017:SOT

[ITZ17] Osman Rasit Isik, Aziz Takhirov, and Haibiao Zheng. Second order time relaxation model for accelerating convergence to steadystate equilibrium for Navier– Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 119(??):67–78, September 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0168927417300909.

Iusem:1997:IPM

[Ius97] Alfredo Noel Iusem. An interior point method for the nonlinear complementarity problem. Applied Numerical Mathematics: Transactions of IMACS, 24(4): 469–482, August 12, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=24&
issue=4&aid=794.

intHout:2009:SCF

[iV09] K. J. in ’t Hout and K. Volders. Stability of central finite difference schemes on non-uniform grids for the Black–Scholes equation. Applied Numerical Mathematics: Transactions of IMACS, 59(10):2593–2609, October 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

I:2016:ANS

[IV16] Arun I. and Murugesan Venkatapathi. Analysis of numerical solutions to Sommerfeld integral relation of the half-space radiator problem. Applied Numerical Mathematics: Transactions of IMACS, 106(??): 79–97, August 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741630037X.

Israeli:1993:DDM

[IVA93] M. Israeli, L. Vozovoi, and A. Averbuch. Domain decomposition methods for solving parabolic PDEs on multiprocessors. Applied Numerical Mathematics: Transactions of IMACS, 12(1–3):193–212, May 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/016892749390118B.

Special issue to honor Professor Saul Abarbanel on his sixtieth birthday (Neveh, 1992).

Ivan:2007:NHI

[Iva07] Mircea Ivan. A note on the Hermite interpolation polynomial for rational functions. Applied Numerical Mathematics: Transactions of IMACS, 57(2):230– 233, February 2007. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

intHout:2007:SAS

[iW07] K. J. in ’t Hout and B. D. Welfert. Stability of ADI schemes applied to convection–diffusion equations with mixed derivative terms. Applied Numerical Mathematics: Transactions of IMACS, 57(1):19– 35, January 2007. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

intHout:2009:USS

[iW09] K. J. in ’t Hout and B. D. Welfert. Unconditional stability of second-order ADI schemes applied to multidimensional diffusion equations with mixed derivative terms. Applied Numerical Mathematics: Transactions of IMACS, 59(3–4): 677–692, March/April 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Jackiewicz:1987:VSV

[Jac87] Zdzisław Jackiewicz. Variablestep variable-order algorithm for the numerical solution of neutral functional differential equations. Applied Numerical Mathematics: Transactions of IMACS, 3(4):317–329, August 1987. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Jacobsen:1988:MCF

[Jac88] Lisa Jacobsen. Meromorphic continuation of functions given by limit k-periodic continued fractions. Applied Numerical Mathematics: Transactions of IMACS, 4(2–4):323–336, June 1988. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Jackiewicz:1993:SAT

[Jac93] Zdzisław Jackiewicz. Stability analysis of time-point relaxation Runge–Kutta methods with respect to tridiagonal systems of differential equations. Applied Numerical Mathematics: Transactions of IMACS, 11(1– 3):189–209, January 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Parallel methods for ordinary differential equations (Grado, 1991).

Jackson:1996:NSL

[Jac96] Kenneth R. Jackson. The numerical solution of large systems of stiff IVPs for ODEs. Applied Numerical Mathematics: Transactions of IMACS, 20(1–2): 5–20, February 29, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=1-2&aid=662. Work-

shop on the method of lines for time-dependent problems (Lexington, KY, 1995).

Jackiewicz:2002:IDS

[Jac02] Z. Jackiewicz. Implementation of DIMSIMs for stiff differential systems. Applied Numerical Mathematics: Transactions of IMACS, 42(1–3):251–267, August 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Jackiewicz:2006:P

[Jac06] Zdzislaw Jackiewicz. Preface. Applied Numerical Mathematics: Transactions of IMACS, 56(3–4):269–270, March/April 2006. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Jadic:1994:NMA

[Jad94] Ioan Jadic. A numerical method for the approximation of singular integrals of functions with a pole of order two. Applied Numerical Mathematics: Transactions of IMACS, 15(4): 395–406, November 1, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1994&volume=15&
                                        issue=4&aid=516.

Jalilian:2021:VSS

[JAH21] A. Jalilian, A. Abdi, and G. Hojjati. Variable stepsize SDIMSIMs for ordinary differential equations. Applied Numerical Mathematics: Transactions of IMACS, 168(??):115–126, October 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001574.

Jameson:1993:CAA

[Jam93]  Antony Jameson.  Com-

putational algorithms for aerodynamic analysis and design. Applied Numerical Mathematics: Transactions of IMACS, 13(5): 383–422, December 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Jameson:1995:SBW

[Jam95] Leland Jameson. On the spline-based wavelet differentiation matrix. Applied Numerical Mathematics: Transactions of IMACS, 17(1):33–45, May 1, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=17&
issue=1&aid=553.

Jay:1995:CRK

[Jay95] Laurent Jay. Convergence of Runge–Kutta methods for differential-algebraic systems of index 3. Applied Numerical Mathematics: Transactions of IMACS, 17(2):97– 118, June 13, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=17&
                                        issue=2&aid=561.

Jbilou:2025:SVN

[Jbi25] Khalide Jbilou. Special volume on numerical analysis and scientific computation with applications. Applied Numerical Mathematics: Transactions of IMACS, 208 (part A)(??):1, February 2025. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424002952.

Jiang:2011:AEA

[JBLC11] Hao Jiang, Roberto Barrio, Xiangke Liao, and Lizhi Cheng. Accurate evaluation algorithm for bivariate polynomial in Bernstein– Bézier form. Applied Numerical Mathematics: Transactions of IMACS, 61(11): 1147–1160, November 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Jong:2021:NAS

[JCJP21] KumSong Jong, HuiChol Choi, KyongJun Jang, and SunAe Pak. A new approach for solving one-dimensional fractional boundary value problems via Haar wavelet collocation method. Applied Numerical Mathematics: Transactions of IMACS, 160(??):313–330, February 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303263.

Ji:2018:HOS

[JCL18] Haifeng Ji, Jinru Chen, and Zhilin Li. A high-order source removal finite element method for a class of elliptic interface problems. Applied Numerical Mathematics: Transactions of IMACS, 130(??): 112–130, August 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418300898.

Jodar:1994:LMP

[JCN94] L. Jódar, R. Company, and E. Navarro. Laguerre matrix polynomials and systems of second-order differential equations. Applied Numerical Mathematics: Transactions of IMACS, 15(1):53–63, August 1, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=15&
issue=1&aid=488.

Jodar:2003:ANS

[JCSR03] L. Jódar, J. I. Castaño, J. A. Sánchez, and G. Rubio. Accurate numerical solution of coupled time dependent parabolic initial value problems. Applied Numerical Mathematics: Transactions of IMACS, 47(3– 4):467–476, December 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Jiang:2025:NSS

[JCY25] Guo Jiang, Yuanqin Chen, and Jiayi Ying. Numerical solutions of stochastic delay integro-differential equations by block pulse functions. Applied Numerical Mathematics: Transactions of IMACS, 208 (part B) (??):214–230, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424002939.

Jung:2009:IAM

[JD09] Jae-Hun Jung and Vincent R. Durante. An iterative adaptive multiquadric radial basis function method for the detection of local jump discontinuities. Applied Numerical Mathematics: Transactions of IMACS, 59(7):1449–1466, July 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Jiang:2010:AGM

[JEG10] Lijian Jiang, Yalchin Efendiev, and Victor Ginting. Analysis of global multiscale finite element methods for wave equations with continuum spatial scales. Applied Numerical Mathematics: Transactions of IMACS, 60(8):862–876, August 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Jeon:2009:MCB

[Jeo09] Youngmok Jeon. A multiscale cell boundary element method for elliptic problems. Applied Numerical Mathematics: Transactions of IMACS, 59(11): 2801–2813, November 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Jespersen:1985:EDS

[Jes85] Dennis C. Jespersen. Enthalpy damping for the steady Euler equations. Applied Numerical Mathematics: Transactions of IMACS, 1(5):417–432, September 1985. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Jessup:1993:CAD

[Jes93] E. R. Jessup. A case against a divide and conquer approach to the nonsymmetric eigenvalue problem. Applied Numerical Mathematics: Transactions of IMACS, 12(5):403–420, July 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/016892749390101V.

Jezequel:1999:VPA

[Jéz99]  Fabienne Jézéquel.  A

validated parallel across time and space solution of the heat transfer equation. Applied Numerical Mathematics: Transactions of IMACS, 31(1): 65–79, September 1, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1999&volume=31&
issue=1&aid=987.

Jezequel:2004:DCC

[Jéz04] Fabienne Jézéquel. Dynamical control of converging sequences computation. Applied Numerical Mathematics: Transactions of IMACS, 50(2):147–164, August 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Jung:2011:IAR

[JGK11] Jae-Hun Jung, Sigal Gottlieb, and Saeja Oh Kim. Iterative adaptive RBF methods for detection of edges in two-dimensional functions. Applied Numerical Mathematics: Transactions of IMACS, 61(1):77–91, January 2011. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Jian:2020:FCI

[JHGZ20] Huan-Yan Jian, Ting-Zhu Huang, Xian-Ming Gu, and Yong-Liang Zhao. Fast compact implicit integration factor method with non-uniform meshes for the two-dimensional nonlinear Riesz space-fractional reaction-diffusion equation. Applied Numerical Mathematics: Transactions of IMACS, 156(??):346–363, October 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420301495.

Jia:2000:RSI

[Jia00] Zhongxiao Jia. A refined subspace iteration algorithm for large sparse eigenproblems. Applied Numerical Mathematics: Transactions of IMACS, 32(1):35– 52, January 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/59/28/
29/abstract.html;   http:

                                        //www.elsevier.nl/gej-
                                        ng/29/17/21/59/28/29/article.
                                        pdf.

Jia:2002:RHA

[Jia02] Zhongxiao Jia. The refined harmonic Arnoldi method and an implicitly restarted refined algorithm for computing interior eigenpairs of large matrices. Applied Numerical Mathematics: Transactions of IMACS, 42(4):489–512, September 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Jiang:2012:CBF

[Jia12] Qingtang Jiang. Correspondence between frame shrinkage and high-order nonlinear diffusion. Applied Numerical Mathematics: Transactions of IMACS, 62(1):51– 66, January 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927411001826.

Janik:1994:SPV

[JJ94] Tadeusz J. Janik and Søren Jensen. On the stability and performance of the p-version of the finite element method for first-order hyperbolic problems. Applied Numerical Mathematics: Transactions of IMACS, 15(4): 407–437, November 1, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=15&
issue=4&aid=504.

Johnston:2015:QOR

[JJ15] S. J. Johnston and K. Jordaan. Quasi-orthogonality and real zeros of some _{2}F_{2} and _{3}F_{2} polynomials. Applied Numerical Mathematics: Transactions of IMACS, 90(??):1–8, April 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414001913.

Jiang:2024:HER

[JJL^{+}24] Peng Jiang, Hongen Jia, Liang Liu, Chenhui Zhang, and Danxia Wang. Highly efficient, robust and unconditionally energy stable second order schemes for approximating the Cahn–Hilliard– Brinkman system. Applied Numerical Mathematics: Transactions of IMACS, 201(??):175–186, July 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000503.

Jeon:2014:FPI

[JK14] Youngmok Jeon and Do Young Kwak. A flux preserving immersed nonconforming finite element method for elliptic problems. Applied Numerical Mathematics: Transactions of IMACS, 81(??):94–104, July 2014. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414000476.

Jeong:2017:PFM

[JK17] Darae Jeong and Junseok Kim. Phase-field model and its splitting numerical scheme for tissue growth. Applied Numerical Mathematics: Transactions of IMACS, 117(??): 22–35, July 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417300351.

Junghanns:2020:NKM

[JK20] P. Junghanns and R. Kaiser. A note on Kalandiya’s method for a crack problem. Applied Numerical Mathematics: Transactions of IMACS, 149(??): 52–64, ???? 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419301096.

Jamshidi:2021:CNS

[JK21] Nahid Jamshidi and Minoo Kamrani. Convergence of a numerical scheme associated to stochastic differential equations with fractional Brownian motion. Applied Numerical Mathematics: Transactions of IMACS, 167(??):108–118, September 2021. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001227.

Jackson:1994:UBS

[JKN94] K. R. Jackson, A. Kværnø, and S. P. Nørsett. The use of Butcher series in the analysis of Newton-like iterations in Runge–Kutta formulas. Applied Numerical Mathematics: Transactions of IMACS, 15(3): 341–356, October 5, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=15&
issue=3&aid=507. Interna-

tional Conference on Scientific Computation and Differential Equations (Auckland, 1993).

Jebens:2012:LIP

[JKW12] Stefan Jebens, Oswald Knoth, and Rüdiger Weiner. Linearly implicit peer methods for the compressible Euler equations. Applied Numerical Mathematics: Transactions of IMACS, 62(10): 1380–1392, October 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927412000992.

Jameson:1986:CCM

[JL86] Antony Jameson and Peter D. Lax. Conditions for the construction of multipoint total variation diminishing difference schemes. Applied Numerical Mathematics: Transactions of IMACS, 2(3–5):335–345, October 1986. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). See corrigendum [JL87].

Jameson:1987:CCC

[JL87] A. Jameson and P. D. Lax. Corrigendum: “Conditions for the construction of multipoint total variation diminishing difference schemes”. Applied Numerical Mathematics: Transactions of IMACS, 3(3):289, June 1987. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0168927487900535.

See [JL86].

Jackiewicz:1991:NSN

[JL91] Z. Jackiewicz and E. Lo. The numerical solution of neutral functional-differential equations by Adams predictorcorrector methods. Applied Numerical Mathematics: Transactions of IMACS, 8(6):477–491, December 1991. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Jorge:1994:CRA

[JL94] J. C. Jorge and F. Lisbona.

Contractivity results for alternating direction schemes in Hilbert spaces. Applied Numerical Mathematics: Transactions of IMACS, 15(1):65–75, August 1, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=15&
issue=1&aid=485.

Jiang:2017:NMS

[JL17] Wei Jiang and Na Liu. A numerical method for solving the time variable fractional order mobile-immobile advection-dispersion model. Applied Numerical Mathematics: Transactions of IMACS, 119(??):18–32, September 2017. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417300880.

Jiang:2021:STA

[JL21] Zhuling Jiang and Jicheng Li. Solving tensor absolute value equation. Applied Numerical Mathematics: Transactions of IMACS, 170(??):255–268, December 2021. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421002117.

Jiang:2023:FPT

[JL23a] Yi Jiang and Jun Liu. Fast parallel-in-time quasiboundary value methods for backward heat conduction problems. Applied Numerical Mathematics: Transactions of IMACS, 184(??): 325–339, February 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422002768.

Jiang:2023:NPA

[JL23b] Zhuling Jiang and Jicheng Li. A new preconditioned AOR-type method for M-tensor equation. Applied Numerical Mathematics: Transactions of IMACS, 189(??):39–52, July 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423000843.

Jiao:2024:MCM

[JL24] Caiyu Jiao and Changpin Li.

Monte Carlo method for the Cauchy problem of fractional diffusion equation concerning fractional Laplacian. Applied Numerical Mathematics: Transactions of IMACS, 201(??):20–40, July 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000308.

Jeong:2025:OCD

[JL25] SeongHee Jeong and Sanghyun Lee. Optimal control for Darcy’s equation in a heterogeneous porous media. Applied Numerical Mathematics: Transactions of IMACS, 207(??):303–322, January 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424002319.

Jimenez:2013:VRM

[JLH13] Edwin Jimenez, Yaning Liu, and M. Yousuff Hussaini. Variance reduction method based on sensitivity derivatives, Part 2. Applied Numerical Mathematics: Transactions of IMACS, 74(??):151–159, December 2013. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927412001778.

Joseph:1990:RNR

[JLL90] G. Joseph, A. Levine, and J. Liukkonen. Randomized Newton–Raphson. Applied Numerical Mathematics: Transactions of IMACS, 6(6):459–469, October 1990. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Jin:2025:SGM

[JLWY25] Lixiang Jin, Zhaoxiang Li, Peipei Wang, and Lijun Yi. Spectral-Galerkin methods for the fully nonlinear Monge–Ampère equation. Applied Numerical Mathematics: Transactions of IMACS, 207(??): 621–641, January 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424002654.

Jia:2020:NAM

[JLZ20] Lueling Jia, Huiyuan Li, and Zhimin Zhang. Numerical analysis on the mortar spectral element methods for Schrödinger eigenvalue problem with an inverse square potential. Applied Numerical Mathematics: Transactions of IMACS, 158(??):54– 84, December 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420301975.

Jiang:2025:GAD

[JLZ25] Kai Jiang, Shifeng Li, and Juan Zhang. A general alternating-direction implicit Newton method for solving continuous-time algebraic Riccati equation. Applied Numerical Mathematics: Transactions of IMACS, 207(??):642–656, January 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424002666.

Johnson:1994:NSM

[JM94] Christopher R. Johnson and Robert S. MacLeod. Nonuniform spatial mesh adaptation using a posteriori error estimates: applications to forward and inverse problems. Applied Numerical Mathematics: Transactions of IMACS, 14(1–3): 311–326, April 14, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=14&
issue=1-3&aid=466. Pro-

ceedings of the Third ARO Workshop on Adaptive Methods for Partial Differential Equations (Troy, NY, 1992).

Jokar:2005:BAS

[JM05] Sadegh Jokar and Bahman Mehri. The best approximation of some rational functions in uniform norm. Applied Numerical Mathematics: Transactions of IMACS, 55(2):204–214, October 2005. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Jun:2006:AMD

[JM06] Younbae Jun and Tsun-Zee Mai. ADI method

— domain decomposition. Applied Numerical Mathematics: Transactions of IMACS, 56(8):1092–1107, August 2006. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Jbilou:2016:BEM

[JM16] K. Jbilou and A. Messaoudi. Block extrapolation methods with applications. Applied Numerical Mathematics: Transactions of IMACS, 106(??): 154–164, August 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416300393.

Jackiewicz:2017:CID

[JM17] Z. Jackiewicz and H. Mittelmann. Construction of IMEX DIMSIMs of high order and stage order. Applied Numerical Mathematics: Transactions of IMACS, 121(??):234–248, November 2017. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417301629.

Jay:2025:ASA

[JM25] Laurent O. Jay and Juan I.

Montijano. Analysis of starting approximations for implicit Runge–Kutta methods applied to ODEs based on the reverse method. Applied Numerical Mathematics: Transactions of IMACS, 215(??):1–14, September 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0168927425000868.

Jabeen:2022:DCA

[JMDN^{+}22] Saudia Jabeen, Jorge E.

Macías-Díaz, Muhammad Aslam Noor, Muhammad Bilal Khan, and Khalida Inayat Noor. Design and convergence analysis of some implicit inertial methods for quasi-variational inequalities via the Wiener–Hopf equations. Applied Numerical Mathematics: Transactions of IMACS, 182(??):76– 86, December 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422001957.

Juncu:2006:NEM

[JMP06] Gh. Juncu, E. Mosekilde, and C. Popa. Numerical experiments with MG continuation algorithms. Applied Numerical Mathematics: Transactions of IMACS, 56(6):844–861, June 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Jokar:2010:SAS

[JMPY10] Sadegh Jokar, Volker Mehrmann, Marc E. Pfetsch, and Harry Yserentant. Sparse approximate solution of partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 60(4):452–472, April 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Jbilou:1999:GFG

[JMS99] K. Jbilou, A. Messaoudi, and H. Sadok. Global FOM and GMRES algorithms for matrix equations. Applied Numerical Mathematics: Transactions of IMACS, 31(1): 49–63, September 1, 1999.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1999&volume=31&
issue=1&aid=959.

Jabeen:2025:NIP

[JMU^{+}25] Saudia Jabeen, Siegfried Macías, Saleem Ullah, Muhammad Aslam Noor, and Jorge E. Macías-Díaz. Note on an inertial projectionbased approach for solving extended general quasivariational inequalities and its convergence analysis. Applied Numerical Mathematics: Transactions of IMACS, 217(??):190–198, November 2025. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425001291.

Jackson:2002:SRA

[JN02] Kenneth R. Jackson and Nedialko S. Nedialkov. Some recent advances in validated methods for IVPs for ODEs. Applied Numerical Mathematics: Transactions of IMACS, 42(1– 3):269–284, August 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Jiang:2007:SOM

[JN07] Song Jiang and Guoxi Ni. A second-order γ-model BGK scheme for multimaterial compressible flows. Applied Numerical Mathematics: Transactions of IMACS, 57(5–7):597–608, May 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Jodar:2003:CSS

[JNPC03] L. Jódar, E. Navarro, A. E. Posso, and M. C. Casabán. Constructive solution of strongly coupled continuous hyperbolic mixed problems. Applied Numerical Mathematics: Transactions of IMACS, 47(3–4): 477–492, December 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

John:2001:RPE

[Joh01] Volker John. Residual a posteriori error estimates for two-level finite element methods for the Navier– Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 37(4):503–518, June 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/73/33/
31/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/73/33/31/article.
pdf.

Johnson:2005:HOB

[Joh05] Richard W. Johnson. Higher order B-spline collocation at the Greville abscissae. Applied Numerical Mathematics: Transactions of IMACS, 52(1):63–75, January 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Jung:2023:CMA

[JOL23] Jae-Hun Jung and Daniel Olmos-Liceaga. A Chebyshev multidomain adaptive mesh method for reaction– diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 190(??):283–302, August 2023. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001204.

Jordan:2011:ORA

[Jor11] Stephen A. Jordan. Optimization, resolution and application of composite compact finite difference templates. Applied Numerical Mathematics: Transactions of IMACS, 61(1):108– 130, January 2011. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Joulak:2005:CQO

[Jou05] H. Joulak. A contribution to quasi-orthogonal polynomials and associated polynomials. Applied Numerical Mathematics: Transactions of IMACS, 54(1): 65–78, June 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Jones:1993:LAG

[JP93] Mark T. Jones and Merrell L. Patrick. Lanczos algorithm for the generalized symmetric eigenproblem on sharedmemory architectures. Applied Numerical Mathematics: Transactions of IMACS, 12(5):377–389, July 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/016892749390099D.

Jakubik:2008:TIS

[JP08a] Peter Jakubik and Roland Potthast. Testing the integrity of some cavity — the Cauchy problem and the range test. Applied Numerical Mathematics: Transactions of IMACS, 58(6): 899–914, June 2008. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Jeon:2008:NCB

[JP08b] Youngmok Jeon and Eun-Jae Park. Nonconforming cell boundary element methods for elliptic problems on triangular mesh. Applied Numerical Mathematics: Transactions of IMACS, 58(6):800–814, June 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Jiang:2017:HSI

[JP17] Qingtang Jiang and Dale K.

Pounds. Highly symmetric \sqrt{3}-refinement Bi-frames for surface multiresolution processing. Applied Numerical Mathematics: Transactions of IMACS, 118(??): 1–18, August 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417300417.

Jarlebring:2019:IMD

[JP19] Elias Jarlebring and Federico Poloni. Iterative methods for the delay Lyapunov equation with T -Sylvester preconditioning. Applied Numerical Mathematics: Transactions of IMACS, 135(??): 173–185, January 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418301879.

Jackaman:2019:DCF

[JPP19] James Jackaman, Georgios Papamikos, and Tristan Pryer. The design of conservative finite element discretisations for the vectorial modified KdV equation. Applied Numerical Mathematics: Transactions of IMACS, 137(??): 230–251, March 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418302320.

Jiang:2022:AHO

[JQSC22] Chaolong Jiang, Xu Qian, Songhe Song, and Jin Cui. Arbitrary high-order linear structure-preserving schemes for the regularized long-wave equation. Applied Numerical Mathematics: Transactions of IMACS, 174(??): 89–111, April 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422000101.

Jian:2023:TCN

[JQYM23] Jinbao Jian, Yao Qin, Jianghua Yin, and Guodong Ma. Theoretical characteristics and numerical methods for a class of special piecewise quadratic optimization. Applied Numerical Mathematics: Transactions of IMACS, 184(??): 340–356, February 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422002811.

Jackiewicz:2000:DIM

[JR00] Z. Jackiewicz and R. A.

Renaut. Diagonally implicit multistage integration methods for pseudospectral solutions of the wave equation. Applied Numerical Mathematics: Transactions of IMACS, 34(2–3): 219–229, July 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/62/29/
33/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/62/29/33/article.
pdf.

Jodar:2002:ANS

[JR02] L. Jódar and M. D. Roselló.

Analytic-numerical solutions with a priori error bounds of initial value problems for the continuous coefficient wave equation. Applied Numerical Mathematics: Transactions of IMACS, 40(1–2):151–166, January 2002. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/85/27/40/abstract.
html.

Jarrad:2018:SSF

[JR18] G. A. Jarrad and A. J.

Roberts. Smooth subgrid fields underpin rigorous closure in spatial discretisation of reaction– advection–diffusion PDEs. Applied Numerical Mathematics: Transactions of IMACS, 132(??):91–110, October 2018. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418301211.

Jbilou:2014:P

[JRS14] Khalide Jbilou, Lothar Reichel, and Hassane Sadok. Preface. Applied Numerical Mathematics: Transactions of IMACS, 75(??): 1, January 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927413001499.

Jannelli:2020:NSS

[JRS20] Alessandra Jannelli, Marianna Ruggieri, and Maria Paola Speciale. Numerical solutions of space-fractional advection–diffusion equations with nonlinear source term. Applied Numerical Mathematics: Transactions of IMACS, 155(??): 93–102, ???? 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420300179.

Jauberteau:1990:NGM

[JRT90] F. Jauberteau, C. Rosier, and R. Temam. The nonlinear Galerkin method in computational fluid dynamics. Applied Numerical Mathematics: Transactions of IMACS, 6(5):361–370, July 1990. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Jackiewicz:2006:NSF

[JRW06] Z. Jackiewicz, M. Rahman, and B. D. Welfert. Numerical solution of a Fredholm integro-differential equation modelling neural networks. Applied Numerical Mathematics: Transactions of IMACS, 56(3–4): 423–432, March/April 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Jeon:2009:CBC

[JS09] Youngmok Jeon and Francisco-Javier Sayas. The CBEM– BEM coupling for elliptic problems. Applied Numerical Mathematics: Transactions of IMACS, 59(10): 2374–2387, October 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Jesus:2024:HON

[JS24] Carla Jesus and Ercília Sousa. High order numerical method for a subdiffusion problem. Applied Numerical Mathematics: Transactions of IMACS, 205(??):

169–183, November 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424001788.

Juhl:2025:ACS

[JS25] Austin Juhl and David Shirokoff. Algebraic conditions for stability in Runge–Kutta methods and their certification via semidefinite programming. Applied Numerical Mathematics: Transactions of IMACS, 207(??): 136–155, January 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424002113.

Jones:1988:CFN

[JT88] William B. Jones and W. J. Thron. Continued fractions in numerical analysis. Applied Numerical Mathematics: Transactions of IMACS, 4(2–4):143–230, June 1988. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Jungel:2002:RSH

[JT02] Ansgar Jüngel and Shaoqiang Tang. A relaxation scheme for the hydrodynamic equations for semiconductors. Applied Numerical Mathematics: Transactions of IMACS, 43(3): 229–252, November 2002.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Jung:2006:IMM

[JT06a] M. Jung and T. D. Todorov.

Isoparametric multigrid method for reaction–diffusion equations on two-dimensional domains. Applied Numerical Mathematics: Transactions of IMACS, 56(12): 1570–1583, December 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Jungel:2006:NAV

[JT06b] Ansgar Jüngel and Shaoqiang Tang. Numerical approximation of the viscous quantum hydrodynamic model for semiconductors. Applied Numerical Mathematics: Transactions of IMACS, 56(7):899–915, July 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Jordaan:2009:ITZ

[JT09] Kerstin Jordaan and Ferenc Toókos. Interlacing theorems for the zeros of some orthogonal polynomials from different sequences. Applied Numerical Mathematics: Transactions of IMACS, 59(8): 2015–2022, August 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Jeon:2018:UHD

[JT18] Youngmok Jeon and Mai Lan Tran. The upwind hybrid difference methods for a convection diffusion equation. Applied Numerical Mathematics: Transactions of IMACS, 133(??):69–82, ???? 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417302507.

Jannoun:2015:CTD

[JTB15] G. Jannoun, R. Touma, and F. Brock. Convergence of two-dimensional staggered central schemes on unstructured triangular grids. Applied Numerical Mathematics: Transactions of IMACS, 92(??):1–20, June 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927415000136.

Jan:2022:NSH

[JUAZ22] Hameed Ullah Jan, Marjan Uddin, Thabet Abdeljawad, and Muhammad Zamir. Numerical study of high order nonlinear dispersive PDEs using different RBF approaches. Applied Numerical Mathematics: Transactions of IMACS, 182(??): 356–369, December 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422002008.

Jung:1997:PMG

[Jun97] M. Jung. On the parallelization of multi-grid methods using a non-overlapping domain decomposition data structure. Applied Numerical Mathematics: Transactions of IMACS, 23(1):119– 137, March 7, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1997/23/

1/746.pdf. Multilevel methods (Oberwolfach, 1995).

Jung:2006:NAC

[Jun06]  Chang-Yeol Jung.  Nu-

merical approximation of convection–diffusion equations in a channel using boundary layer elements. Applied Numerical Mathematics: Transactions of IMACS, 56(6):756–777, June 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Jung:2007:NGP

[Jun07] Jae-Hun Jung. A note on the Gibbs phenomenon with multiquadric radial basis functions. Applied Numerical Mathematics: Transactions of IMACS, 57(2):213– 229, February 2007. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Jovanovic:2009:FDA

[JV09] Boško S. Jovanović and Lubin G. Vulkov. Finite difference approximations for some interface problems with variable coefficients. Applied Numerical Mathematics: Transactions of IMACS, 59(2):349– 372, February 2009. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Jackiewicz:1995:VSD

[JVZ95] Z. Jackiewicz, R. Vermiglio, and M. Zennaro. Variable stepsize diagonally implicit multistage integration methods for ordinary differential equations. Applied Numerical Mathematics: Transactions of IMACS, 16(3): 343–367, February 16, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=16&
issue=3&aid=533.

Jackiewicz:1996:RPR

[JVZ96] Z. Jackiewicz, R. Vermiglio, and M. Zennaro. Regularity properties of Runge– Kutta methods for ordinary differential equations. Applied Numerical Mathematics: Transactions of IMACS, 22(1–3):251–262, December 4, 1996. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=22&
issue=1-3&aid=718. Special

issue celebrating the centenary of Runge–Kutta methods.

Jackiewicz:1997:RPR

[JVZ97] Z. Jackiewicz, R. Vermiglio, and M. Zennaro. Regularity properties of Runge–Kutta methods for delay differential equations. Applied Numerical Mathematics: Transactions of IMACS, 24(2– 3):265–278, June 19, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=24&
issue=2-3&aid=781. Volterra

centennial (Tempe, AZ, 1996).

Jiang:2001:NCC

[JW01] Yao-Lin Jiang and Omar Wing. A note on convergence conditions of waveform relaxation algorithms for nonlinear differential-algebraic equations. Applied Numerical Mathematics: Transactions of IMACS, 36(2– 3):281–297, February 2001.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.nl../35/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/10/28/65/31/35/article.
                                        pdf.

Wang:2015:MSM

[jW15] Tian jun Wang. Mixed spectral method for heat transfer with inhomogeneous Neumann boundary condition in an infinite strip. Applied Numerical Mathematics: Transactions of IMACS, 92(??):82–97, June 2015. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927415000203.

Wang:2022:FDP

[jWC22] Tian jun Wang and Guo Chai. A fully discrete pseudospectral method for the nonlinear Fokker–Planck equations on the whole line. Applied Numerical Mathematics: Transactions of IMACS, 174(??): 17–33, April 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422000034.

Jiang:2020:AHO

[JWG20] Chaolong Jiang, Yushun Wang, and Yuezheng Gong.

Arbitrarily high-order energypreserving schemes for the Camassa–Holm equation. Applied Numerical Mathematics: Transactions of IMACS, 151(??):85–97, May 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0168927419303678.

Wang:2017:FDP

[jWjJ17] Tian jun Wang and Yu jian Jiao. A fully discrete pseudospectral method for Fisher’s equation on the whole line. Applied Numerical Mathematics: Transactions of IMACS, 120(??): 243–256, October 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417301381.

Wang:2009:EAL

[jWqW09] Tian jun Wang and Zhong qing Wang. Error analysis of Legendre spectral method with essential imposition of Neumann boundary condition. Applied Numerical Mathematics: Transactions of IMACS, 59(10): 2444–2451, October 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wang:2020:SMS

[jWS20] Tian jun Wang and Tao Sun. A spectral method for solving nonhomogeneous Neumann boundary value problems on quadrilaterals. Applied Numerical Mathematics: Transactions of IMACS, 157(??):1– 18, November 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420301690.

Wang:2008:CGL

[jWyG08] Tian jun Wang and Ben yu Guo. Composite generalized Laguerre–Legendre pseudospectral method for Fokker–Planck equation in an infinite channel. Applied Numerical Mathematics: Transactions of IMACS, 58(10):1448–1466, October 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Jia:2021:PFF

[JWZ21] Jinhong Jia, Hong Wang, and Xiangcheng Zheng. A preconditioned fast finite element approximation to variable-order timefractional diffusion equations in multiple space dimensions. Applied Numerical Mathematics: Transactions of IMACS, 163(??): 15–29, May 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000027.

Jia:2020:JBB

[JY20] Zhongxiao Jia and Yanfei Yang. A joint bidiagonalization based iterative algorithm for large scale general-form Tikhonov regularization. Applied Numerical Mathematics: Transactions of IMACS, 157(??): 159–177, November 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420301732.

Jiang:2023:USS

[JY23] Nan Jiang and Huanhuan Yang. Unconditionally stable, second order, decoupled ensemble schemes for computing evolutionary Boussinesq equations. Applied Numerical Mathematics: Transactions of IMACS, 192(??):241–260, October 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001769.

Jentschura:2010:CCF

[JZJ10] Ulrich D. Jentschura and Jean Zinn-Justin. Calculation of the characteristic functions of anharmonic oscillators. Applied Numerical Mathematics: Transactions of IMACS, 60(12): 1332–1341, December 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Jackiewicz:2006:SCW

[JZK06] Z. Jackiewicz and B. Zubik-Kowal. Spectral collocation and waveform relaxation methods for nonlinear delay partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 56(3–4): 433–443, March/April 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ji:2020:DFD

[JZS20] Bingquan Ji, Luming Zhang, and Qihang Sun. A dissipative finite difference Fourier pseudo-spectral method for the symmetric regularized long wave equation with damping mechanism. Applied Numerical Mathematics: Transactions of IMACS, 154(??):90–103, August 2020. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420300891.

Jia:2021:ESO

[JZXJ21] Junqing Jia, Hui Zhang, Huanying Xu, and Xiaoyun Jiang. An efficient second order stabilized scheme for the two dimensional time fractional Allen–Cahn equation. Applied Numerical Mathematics: Transactions of IMACS, 165(??): 216–231, July 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000660.

Jiang:2022:HOW

[JZZH22] Yan-Qun Jiang, Shu-Guang Zhou, Xu Zhang, and Ying-Gang Hu. High-order weighted compact nonlinear scheme for one- and two-dimensional Hamilton– Jacobi equations. Applied Numerical Mathematics: Transactions of IMACS, 171(??):353–368, January 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742100266X.

Khosravian-Arab:2024:SCD

[KAD24] Hassan Khosravian-Arab and Mehdi Dehghan. The sine and cosine diffusive representations for the Caputo fractional derivative. Applied Numerical Mathematics: Transactions of IMACS, 204(??):265–290, October 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742400165X.

Kallinderis:1996:DFV

[Kal96]  Y. Kallinderis.  A 3-D

finite-volume method for the Navier–Stokes equations with adaptive hybrid grids. Applied Numerical Mathematics: Transactions of IMACS, 20(4):387– 406, June 3, 1996. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=20&
                                        issue=4&aid=650. Adaptive

mesh refinement methods for CFD applications (Atlanta, GA, 1994).

Kaluza:2022:OGA

[Kał22]  Andrzej Kałuza.  Op-

timal global approximation of systems of jumpdiffusion SDEs on equidistant mesh. Applied Numerical Mathematics: Transactions of IMACS, 179(??):1– 26, September 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422001027.

Kamenetskii:2000:OFR

[Kam00] D. S. Kamenetskii. On one form of representing the difference potentials. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4):501–508, May 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/61/27/
79/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/61/27/79/article.
pdf.

Kamenetskiy:2016:RED

[Kam16] D. S. Kamenetskiy. On the relation of the Embedded Discontinuous Galerkin method to the stabilized residual-based finite element methods. Applied Numerical Mathematics: Transactions of IMACS, 108(??): 271–285, October 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416000283.

Kamandi:2025:NPB

[Kam25] Ahmad Kamandi. A novel projection-based method for monotone equations with Aitken Δ^{2} acceleration and its application to sparse signal restoration. Applied Numerical Mathematics: Transactions of IMACS, 213(??):1–11, July 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425000376.

Kaneko:1989:PMS

[Kan89] Hideaki Kaneko. A projection method for solving Fredholm integral equations of the second kind. Applied Numerical Mathematics: Transactions of IMACS, 5(4):333–344, July 1989. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kang:2004:CBI

[Kan04] Kab Seok Kang. Covolumebased intergrid transfer operator in P _{1} nonconforming multigrid method. Applied Numerical Mathematics: Transactions of IMACS, 51(1):47–67, October 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Karniadakis:1989:SES

[Kar89] George Em. Karniadakis. Spectral element simulations of laminar and turbulent flows in complex geometries. Applied Numerical Mathematics: Transactions of IMACS, 6(1–2):85– 105, December 1989. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Spectral multi-domain methods (Paris, 1988).

Kurkcu:2017:NMS

[KAS17] Ömür Kivanç Kürkçü, Ersin Aslan, and Mehmet Sezer. A numerical method for solving some model problems arising in science and convergence analysis based on residual function. Applied Numerical Mathematics: Transactions of IMACS, 121(??): 134–148, November 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417301514.

Kedia:2022:SNS

[KAS22] Nikki Kedia, Anatoly A.

Alikhanov, and Vineet Kumar Singh. Stable numerical schemes for timefractional diffusion equation with generalized memory kernel. Applied Numerical Mathematics: Transactions of IMACS, 172(??): 546–565, February 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421003147.

Katti:1989:NFO

[Kat89]  C. P. Katti.  A new

fourth-order finite difference method for a class of singular two-point boundary value problems. Applied Numerical Mathematics: Transactions of IMACS, 5(5):

451–457, September 1989. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kauthen:1993:IRK

[Kau93] Jean-Paul Kauthen. Implicit Runge–Kutta methods for some integrodifferentialalgebraic equations. Applied Numerical Mathematics: Transactions of IMACS, 13(1–3):125–134, September 1993. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Sixth Conference on the Numerical Treatment of Differential Equations (Halle, 1992).

Kauthen:1995:IRK

[Kau95]  J.-P. Kauthen.  Implicit

Runge–Kutta methods for singularly perturbed integrodifferential systems. Applied Numerical Mathematics: Transactions of IMACS, 18(1–3):201–210, September 27, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=18&
issue=1-3&aid=601. Sev-

enth Conference on the Numerical Treatment of Differential Equations (Halle, 1994).

Kauthen:1997:SSP

[Kau97]  J.-P. Kauthen.  A sur-

vey of singularly perturbed Volterra equations. Applied Numerical Mathematics: Transactions of IMACS, 24(2–3):95–114, June 19, 1997. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=24&
issue=2-3&aid=770. Volterra

centennial (Tempe, AZ, 1996).

Khalil:2021:CRE

[KB21] Omar A. Khalil and Gerd Baumann. Convergence rate estimation of poly-Sinc-based discontinuous Galerkin methods. Applied Numerical Mathematics: Transactions of IMACS, 165(??):527–552, July 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000805.

Kokiopoulou:2004:CSS

[KBG04] E. Kokiopoulou, C. Bekas, and E. Gallopoulos. Computing smallest singular triplets with implicitly restarted Lanczos bidiagonalization. Applied Numerical Mathematics: Transactions of IMACS, 49(1):39–61, April 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kakuba:2021:ACS

[KBK21] Godwin Kakuba, Fredrik Berntsson, and Vladimir Kozlov. An algorithm for computing a stationary flow in a binary bifurcation tree. Applied Numerical Mathematics: Transactions of IMACS, 159(??): 125–137, January 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742030252X.

Karatay:2011:IDA

[KBS11] Ibrahim Karatay, Serife R. Bayramoglu, and Ali Sahin. Implicit difference approximation for the time fractional heat equation with the nonlocal condition. Applied Numerical Mathematics: Transactions of IMACS, 61(12):1281–1288, December 2011. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927411001528.

Kennedy:1994:SNN

[KC94] Christopher A. Kennedy and Mark H. Carpenter. Several new numerical methods for compressible shear-layer simulations. Applied Numerical Mathematics: Transactions of IMACS, 14(4): 397–433, June 10, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=14&
issue=4&aid=480.

Kennedy:2003:ARK

[KC03] Christopher A. Kennedy and Mark H. Carpenter. Additive Runge–Kutta schemes for convection–diffusion– reaction equations. Applied Numerical Mathematics: Transactions of IMACS, 44(1–2):139–181, January 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kennedy:2019:DIR

[KC19a] Christopher A. Kennedy and Mark H. Carpenter. Diagonally implicit Runge– Kutta methods for stiff ODEs. Applied Numerical Mathematics: Transactions of IMACS, 146(??): 221–244, December 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927419301801.

Kennedy:2019:HOA

[KC19b] Christopher A. Kennedy and Mark H. Carpenter. Higher-order additive Runge–Kutta schemes for ordinary differential equations. Applied Numerical Mathematics: Transactions of IMACS, 136(??): 183–205, February 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418302332.

Krukier:2002:TSS

[KCB02] L. A. Krukier, L. G. Chikina, and T. V. Belokon. Triangular skewsymmetric iterative solvers for strongly nonsymmetric positive real linear system of equations. Applied Numerical Mathematics: Transactions of IMACS, 41(1):89– 105, April 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/86/27/33/abstract.
html.

Kojouharov:2004:NEL

[KCC04] Hristo V. Kojouharov and Benito M. Chen-Charpentier. Nonstandard Eulerian–Lagrangian methods for multi-dimensional reactive transport problems. Applied Numerical Mathematics: Transactions of IMACS, 49(2):225–243, May 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kim:2003:EQR

[KCI03] KyungJoong Kim, Ronald Cools, and L. Gr. Ixaru. Extended quadrature rules for oscillatory integrands. Applied Numerical Mathematics: Transactions of IMACS, 46(1):59–73, July 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kim:2001:AAC

[KCJP01] Seong-Lyong Kim, Jeong-Yeol Choi, In-Seuck Jeung, and Yang-Ho Park. Application of approximate chemical Jacobians for constant volume reaction and shock-induced combustion. Applied Numerical Mathematics: Transactions of IMACS, 39(1):87–104, October 2001. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/79/27/32/abstract.
html.

Kennedy:2000:LSE

[KCL00] Christopher A. Kennedy, Mark H. Carpenter, and R. Michael Lewis. Lowstorage, explicit Runge– Kutta schemes for the compressible Navier–Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 35(3): 177–219, November 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/10/28/64/31/
                                        26/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/10/28/64/31/26/article.
                                        pdf.

Karageorghis:2007:MDR

[KCS07] Andreas Karageorghis, C. S. Chenand, and Yiorgos-Sokratis Smyrlis. A matrix decomposition RBF algorithm: Approximation of functions and their derivatives. Applied Numerical Mathematics: Transactions of IMACS, 57(3):304–319, March 2007. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kordy:2016:AMO

[KCW16] M. Kordy, E. Cherkaev, and P. Wannamaker. Adaptive model order reduction for the Jacobian calculation in inverse multi-frequency problem for Maxwell’s equations. Applied Numerical Mathematics: Transactions of IMACS, 109(??):1– 18, November 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416300903.

Kong:2019:NKE

[KCY19] Linghua Kong, Meng Chen, and Xiuling Yin. A novel kind of efficient symplectic scheme for Klein– Gordon–Schrödinger equation. Applied Numerical Mathematics: Transactions of IMACS, 135(??): 481–496, January 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418302022.

Kazolea:2013:WBS

[KD13] M. Kazolea and A. I.

Delis. A well-balanced shock-capturing hybrid finite volume-finite difference numerical scheme for extended 1D Boussinesq models. Applied Numerical Mathematics: Transactions of IMACS, 67(??):167–186, May 2013. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411001127.

Kumar:2025:UCA

[KD25] Shridhar Kumar and Pratibhamoy Das. A uniformly convergent analysis for multiple scale parabolic singularly perturbed convectiondiffusion coupled systems: Optimal accuracy with less computational time. Applied Numerical Mathematics: Transactions of IMACS, 207(??):534–557, January 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424002575.

Kampanis:2013:FED

[KDAK13] N. A. Kampanis, A. I. Delis, D. C. Antonopoulou, and G. Kozyrakis. A finite element discretization of the standard parabolic equation in generalized boundary fitting coordinates. Applied Numerical Mathematics: Transactions of IMACS, 67(??):152–166, May 2013. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927411000882.

Kozyrakis:2016:NMS

[KDAK16] G. V. Kozyrakis, A. I. Delis, G. Alexandrakis, and N. A. Kampanis. Numerical modeling of sediment transport applied to coastal morphodynamics. Applied Numerical Mathematics: Transactions of IMACS, 104(??): 30–46, June 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414001603.

Kumar:2023:EWG

[KDD23] Naresh Kumar, Jogen Dutta, and Bhupen Deka. L ^{2} estimates for weak Galerkin finite element methods for second-order wave equations with polygonal meshes. Applied Numerical Mathematics: Transactions of IMACS, 192(??):84–103, October 2023. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001290.

Kheybari:2020:SAA

[KDH20] Samad Kheybari, Mohammad Taghi Darvishi, and Mir Sajjad Hashemi. A semi-analytical approach to Caputo type time-fractional modified anomalous subdiffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 158(??):103–122, December 2020. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302233.

Kozyrakis:2017:FDS

[KDK17] G. V. Kozyrakis, A. I. Delis, and N. A. Kampanis. A finite difference solver for incompressible Navier–Stokes flows in complex domains. Applied Numerical Mathematics: Transactions of IMACS, 115(??):275–298, May 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416301374.

Kutniv:2020:NAC

[KDKW20] M. V. Kutniv, B. Y. Datsko, A. V. Kunynets, and A. Włoch. A new approach to constructing of explicit one-step methods of high order for singular initial value problems for nonlinear ordinary differential equations. Applied Numerical Mathematics: Transactions of IMACS, 148(??): 140–151, February 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419302454.

Kumar:2022:HAA

[KDS22] Devendra Kumar, Komal Deswal, and Satpal Singh. A highly accurate algorithm for retrieving the predicted behavior of problems with piecewise-smooth initial data. Applied Numerical Mathematics: Transactions of IMACS, 173(??): 279–294, March 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421003391.

Kamranian:2017:AML

[KDT17] Maryam Kamranian, Mehdi Dehghan, and Mehdi Tatari. An adaptive meshless local Petrov–Galerkin method based on a posteriori error estimation for the boundary layer problems. Applied Numerical Mathematics: Transactions of IMACS, 111(??):181–196, January 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741630174X.

Kelmanson:1985:CAN

[Kel85] M. A. Kelmanson. A consistency analysis for the numerical solution of boundary integral equations. Applied Numerical Mathematics: Transactions of IMACS, 1(5):381–393, September 1985. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Khalil:2023:APP

[KESYB23] Omar A. Khalil, Hany A.

El-Sharkawy, Maha Youssef, and Gerd Baumann. Adaptive piecewise Poly-Sinc methods for function approximation. Applied Numerical Mathematics: Transactions of IMACS, 186(??): 1–18, April 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422003385.

Kim:1997:BCF

[KF97] Young Ik Kim and Alan Feldstein. Bifurcation and kcycles of a finite-dimensional iterative map, with applications to logistic delay equations. Applied Numerical Mathematics: Transactions of IMACS, 24(2–3): 411–424, June 19, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=24&
                                        issue=2-3&aid=792. Volterra

centennial (Tempe, AZ, 1996).

Korneev:2002:ASA

[KFOF02] V. Korneev, J. E. Flaherty, J. T. Oden, and J. Fish. Additive Schwarz algorithms for solving hp-version finite element systems on triangular meshes. Applied Numerical Mathematics: Transactions of IMACS, 43(4): 399–421, December 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kolezas:2025:CDI

[KFR25] Georgios D. Kolezas, George Fikioris, and John A. Roumeliotis. Convergence, divergence, and inherent oscillations in MAS solutions of 2D Laplace–Neumann problems. Applied Numerical Mathematics: Transactions of IMACS, 209(??): 171–186, March 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424003209.

Keyes:1990:DDT

[KG90] David E. Keyes and William D.

Gropp. Domain decomposition techniques for the parallel solution of nonsymmetric systems of elliptic boundary value problems. Applied Numerical Mathematics: Transactions of IMACS, 6(4):281–301, May 1990. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kahl:2008:SPS

[KGR08] C. Kahl, M. Günther, and T. Rossberg. Structure preserving stochastic integration schemes in interest rate derivative modeling. Applied Numerical Mathematics: Transactions of IMACS, 58(3):284–295, March 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kumar:2024:ECT

[KGZ24] Saurabh Kumar, Vikas Gupta, and Dia Zeidan. An efficient collocation technique based on operational matrix of fractionalorder Lagrange polynomials for solving the spacetime fractional-order partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 204(??):249–264, October 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001600.

Koren:1991:DDD

[KH91] B. Koren and P. W. Hemker. Damped, directiondependent multigrid for hypersonic flow computations. Applied Numerical Mathematics: Transactions of IMACS, 7(4): 309–328, April 1991. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/016892749190067A.

Khayat:2012:SAT

[KHA12] Roger E. Khayat, Michael A. N. Hanyk, and Moinuddin Ahmed. A spectral approach to thin jet flow. Applied Numerical Mathematics: Transactions of IMACS, 62(9): 1187–1196, September 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927410002011.

Khan:2021:HOP

[Kha21] Feroz Khan. Higher order pathwise approximation for the stochastic Burgers’ equation with additive noise. Applied Numerical Mathematics: Transactions of IMACS, 162(??):67–80, April 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303846.

Kheirkhah:2022:PNS

[KHB22] Farnaz Kheirkhah, Mojtaba Hajipour, and Dumitru Baleanu. The performance of a numerical scheme on the variable-order time-fractional advection– reaction–subdiffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 178(??): 25–40, August 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422000782.

Khelifi:1991:LMA

[Khe91] Mohamed Khelifi. Lanczos maximal algorithm for unsymmetric eigenvalue problems. Applied Numerical Mathematics: Transactions of IMACS, 7(2):179– 193, February 1991. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Karamollahi:2022:ERT

[KHLV22] Nasibeh Karamollahi, Mohammad Heydari, Ghasem Barid Loghmani, and Amit K. Verma. An explicit representation of the threepoint Hermite interpolant for the numerical solution of singular boundary value problems. Applied Numerical Mathematics: Transactions of IMACS, 182(??): 265–284, December 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422002021.

Kabanikhin:2014:VAR

[KHM^{+}14] Sergey Kabanikhin, Alemdar Hasanov, Igor Marinin, Olga Krivorotko, and David Khidasheli. A variational approach to reconstruction of an initial tsunami source perturbation. Applied Numerical Mathematics: Transactions of IMACS, 83(??):22– 37, September 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414000683.

Khan:2019:NAC

[KHM^{+}19] Mair Khan, Arif Hussain, M. Y. Malik, T. Salahuddin, and Shaban Aly. Numerical analysis of Carreau fluid flow for generalized Fourier’s and Fick’s laws. Applied Numerical Mathematics: Transactions of IMACS, 144(??): 100–117, October 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927419301333.

Kim:2021:TOW

[KHYY21] Chang Ho Kim, Youngsoo Ha, Hyoseon Yang, and Jungho Yoon. A thirdorder WENO scheme based on exponential polynomials for Hamilton–Jacobi equations. Applied Numerical Mathematics: Transactions of IMACS, 165(??): 167–183, July 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000428.

Kida:1990:PTP

[Kid90a]  Sôsaburô Kida.  Padé-

type and Padé approximants in several variables. Applied Numerical Mathematics: Transactions of IMACS, 6(5):371–391, July 1990. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kida:1990:RBP

[Kid90b] Sôsaburô Kida. Relation between Padé-type approximation and polynomial interpolation in several variables. Applied Numerical Mathematics: Transactions of IMACS, 6(5):393– 404, July 1990. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kiehl:1995:POS

[Kie95] Martin Kiehl. Parallel onestep methods with minimal parallel stages. Applied Numerical Mathematics: Transactions of IMACS, 17(4):397–409, September 11, 1995. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=17&
                                        issue=4&aid=619.

Kieri:2015:SCF

[Kie15] Emil Kieri. Stiff convergence of force-gradient operator splitting methods. Applied Numerical Mathematics: Transactions of IMACS, 94(??):33–45, August 2015. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927415000525.

Kieu:2017:MFE

[Kie17]  Thinh Kieu.  A mixed

finite element approximation for Darcy–Forchheimer flows of slightly compressible fluids. Applied Numerical Mathematics: Transactions of IMACS, 120(??): 141–164, October 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417301216.

Kim:1994:PIP

[Kim94]  Seongjai Kim.  A par-

allelizable iterative procedure for the Helmholtz problem. Applied Numerical Mathematics: Transactions of IMACS, 14(4): 435–449, June 10, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=14&
issue=4&aid=482.

Kim:1995:PMI

[Kim95]  Seongjai Kim.  Parallel

multidomain iterative algorithms for the Helmholtz wave equation. Applied Numerical Mathematics: Transactions of IMACS, 17(4): 411–429, September 11, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/
apnum/cas_sub/browse/browse.
cgi?year=1995&volume=17&
issue=4&aid=587.

Kim:2007:PEE

[Kim07] Kwang Y. Kim. A posteriori error estimators for locally conservative methods of nonlinear elliptic problems. Applied Numerical Mathematics: Transactions of IMACS, 57(9): 1065–1080, September 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kim:2012:FRP

[Kim12] Kwang-Yeon Kim. Flux reconstruction for the P2 nonconforming finite element method with application to a posteriori error estimation. Applied Numerical Mathematics: Transactions of IMACS, 62(12): 1701–1717, December 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927412001158.

Kim:2014:CPA

[Kim14] Seungil Kim. Cartesian PML approximation to resonances in open systems in R^{2}. Applied Numerical Mathematics: Transactions of IMACS, 81(??):50–75, July 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892741400049X.

Kim:2019:UGO

[Kim19] Yunho Kim. An unconstrained global optimization framework for real symmetric eigenvalue problems. Applied Numerical Mathematics: Transactions of IMACS, 144(??): 253–275, October 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927419301138.

Kim:2021:GAE

[Kim21] Kwang-Yeon Kim. Guaranteed and asymptotically exact a posteriori error estimator for lowest-order Raviart– Thomas mixed finite element method. Applied Numerical Mathematics: Transactions of IMACS, 165(??): 357–375, July 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000726.

Kincaid:1994:SSD

[Kin94] David R. Kincaid. Stationary second-degree iterative methods. Applied Numerical Mathematics: Transactions of IMACS, 16(1–2): 227–237, December 1, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/

                                        cas_sub/browse/browse.
                                        cgi?year=1994&volume=16&
                                        issue=1-2&aid=513.  A

Festschrift to honor Professor Robert Vichnevetsky on his 65th birthday.

Korneev:1999:DDP

[KJ99] Vadim G. Korneev and Søren Jensen. Domain decomposition preconditioning in the hierarchical pversion of the finite element method. Applied Numerical Mathematics: Transactions of IMACS, 29(4):479– 518, April 1, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1999/29/
                                        4/942.pdf.

Karageorghis:2012:MFS

[KJL12] A. Karageorghis, B. T. Johansson, and D. Lesnic. The method of fundamental solutions for the identification of a sound-soft obstacle in inverse acoustic scattering. Applied Numerical Mathematics: Transactions of IMACS, 62(12): 1767–1780, December 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927412001201.

Kreiss:1986:CSS

[KK86] Gunilla Kreiss and Heinz-Otto Kreiss. Convergence to steady state of solutions of Burgers’ equation. Applied Numerical Mathematics: Transactions of IMACS, 2(3–5):161–179, October 1986. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kress:2002:LSM

[KK02] R. Kress and L. Kühn.

Linear sampling methods for inverse boundary value problems in potential theory. Applied Numerical Mathematics: Transactions of IMACS, 43(1–2):161–173, October 2002. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kim:2006:PIN

[KK06] S. Kim and K. L. Kreider. Parameter identification for nonlinear elastic and viscoelastic plates. Applied Numerical Mathematics: Transactions of IMACS, 56(12):1538–1554, December 2006. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kim:2009:NMT

[KK09a] Junseok Kim and Kyungkeun Kang. A numerical method for the ternary Cahn– Hilliard system with a degenerate mobility. Applied Numerical Mathematics: Transactions of IMACS, 59(5):1029–1042, May 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kim:2009:PHO

[KK09b] Seonhee Kim and Sang Dong Kim. Preconditioning on high-order element methods using Chebyshev–Gauss– Lobatto nodes. Applied Numerical Mathematics: Transactions of IMACS, 59(2): 316–333, February 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kashima:2011:OAW

[KK11] Kenji Kashima and Reiichiro Kawai. An optimization approach to weak approximation of stochastic differential equations with jumps. Applied Numerical Mathematics: Transactions of IMACS, 61(5):641–650, May 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kulikov:2017:ACE

[KK17] G. Yu. Kulikov and M. V. Kulikova. Accurate cubature and extended Kalman filtering methods for estimating continuous-time nonlinear stochastic systems with discrete measurements. Applied Numerical Mathematics: Transactions of IMACS, 111(??): 260–275, January 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741630191X.

Kazmi:2020:ESS

[KK20a] Kamran Kazmi and Abdul Q. M. Khaliq. An efficient split-step method for distributed-order spacefractional reaction–diffusion equations with time-dependent boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 147(??):142–160, January 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419302284.

Kulikov:2020:NBC

[KK20b] G. Yu. Kulikov and M. V.

Kulikova. NIRK-based Cholesky-factorized squareroot accurate continuousdiscrete unscented Kalman filters for state estimation in nonlinear continuous-time stochastic models with discrete measurements. Applied Numerical Mathematics: Transactions of IMACS, 147(??):196–221, January 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419302302.

Kumar:2020:PUN

[KK20c] Devendra Kumar and Parvin Kumari. Parameter-uniform numerical treatment of singularly perturbed initialboundary value problems with large delay. Applied Numerical Mathematics: Transactions of IMACS, 153(??):412–429, July 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420300672.

Kroner:2022:PEE

[KK22a] Axel Kröner and Heiko Kröner. A priori error estimates for finite element approximations of regularized level set flows in higher norms. Applied Numerical Mathematics: Transactions of IMACS, 171(??): 307–328, January 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421002622.

Kulikova:2022:SRF

[KK22b] Maria V. Kulikova and Gennady Yu. Kulikov. Squareroot filtering via covariance SVD factors in the accurate continuous-discrete extended-cubature Kalman filter. Applied Numerical Mathematics: Transactions of IMACS, 171(??): 32–44, January 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421002415.

Kumari:2023:STO

[KK23] Archna Kumari and Vijay Kumar Kukreja. Study of 4-th order Kuramoto– Sivashinsky equation by septic Hermite collocation method. Applied Numerical Mathematics: Transactions of IMACS, 188(??): 88–105, June 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423000624.

Kreiss:2016:ASG

[KKE16] G. Kreiss, B. Krank, and G. Efraimsson. Analysis of stretched grids as buffer zones in simulations of wave propagation. Applied Numerical Mathematics: Transactions of IMACS, 107(??):1– 17, September 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416300563.

Kumar:2025:SSS

[KKK25] Rakesh Kumar, Kapil Kant, and B. V. Rathish Kumar. Superconvergent scheme for a system of green Fredholm integral equations. Applied Numerical Mathematics: Transactions of IMACS, 207(??):254–271, January 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424002459.

Koukoudakis:2025:NMM

[KKL^{+}25] N. Koukoudakis, C. Koukouvinos, A. Lappa, M. Mitrouli, and A. Psitou. Numerical methods in modeling with supersaturated designs. Applied Numerical Mathematics: Transactions of IMACS, 208 (part A)(??): 271–283, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927424000254.

Koorapetse:2021:DFR

[KKLD21] M. Koorapetse, P. Kaelo, S. Lekoko, and T. Diphofu. A derivative-free RMIL conjugate gradient projection method for convex constrained nonlinear monotone equations with applications in compressive sensing. Applied Numerical Mathematics: Transactions of IMACS, 165(??):431–441, July 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000751.

Karchevsky:2013:KMG

[KKN^{+}13] Andrey L. Karchevsky, Michael V.

Klibanov, Lam Nguyen, Natee Pantong, and Anders Sullivan. The Krein method and the globally convergent method for experimental data. Applied Numerical Mathematics: Transactions of IMACS, 74(??): 111–127, December 2013. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413001219.

Kolesov:2017:SME

[KKN^{+}17] Aleksandr E. Kolesov, Michael V.

Klibanov, Loc H. Nguyen, Dinh-Liem Nguyen, and Nguyen T. Thành. Single measurement experimental data for an inverse medium problem inverted by a multi-frequency globally convergent numerical method. Applied Numerical Mathematics: Transactions of IMACS, 120(??): 176–196, October 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417301228.

Kahou:2007:EFM

[KKP07] Guy Antoine Atenekeng Kahou, Emmanuel Kamgnia, and Bernard Philippe. An explicit formulation of the multiplicative Schwarz preconditioner. Applied Numerical Mathematics: Transactions of IMACS, 57(11– 12):1197–1213, November/ December 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Krol:2017:DSS

[KKP17] M. Król, M. V. Kutniv, and O. I. Pazdriy. Difference schemes for systems of second order nonlinear ODEs on a semi-infinite interval. Applied Numerical Mathematics: Transactions of IMACS, 119(??):33– 50, September 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417300867.

Kupper:2015:SAC

[KKR15] Dominique Küpper, Anne Kværnø, and Andreas Rößler. Stability analysis and classification of Runge–Kutta methods for index 1 stochastic differential-algebraic equations with scalar noise. Applied Numerical Mathematics: Transactions of IMACS, 96(??):24–44, October 2015. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741500063X.

Kaur:2024:OAN

[KKR24] Anurag Kaur, V. Kanwar, and Higinio Ramos. An optimized algorithm for numerical solution of coupled Burgers equations. Applied Numerical Mathematics: Transactions of IMACS, 204(??):352–361, October 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001673.

Krejic:2025:ALA

[KKR25] N. Krejić, E. H. M. Krulikovski, and M. Raydan. An augmented Lagrangian approach for cardinality constrained minimization applied to variable selection problems. Applied Numerical Mathematics: Transactions of IMACS, 208 (part A) (??):284–296, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927423003148.

Kadalbajoo:2016:RBF

[KKT16] Mohan K. Kadalbajoo, Alpesh Kumar, and Lok Pati Tripathi. A radial basis function based implicitexplicit method for option pricing under jump-diffusion models. Applied Numerical Mathematics: Transactions of IMACS, 110(??): 159–173, December 2016. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301568.

Koch:2000:IEM

[KKW00] Othmar Koch, Peter Kofler, and Ewa B. Weinmüller. The implicit Euler method for the numerical solution of singular initial value problems. Applied Numerical Mathematics: Transactions of IMACS, 34(2–3): 231–252, July 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/62/29/
34/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/62/29/34/article.
pdf.

Kumar:1987:SPO

[KL87] Ashwani Kumar and Angelo Lucia. Separation process optimization calculations. Applied Numerical Mathematics: Transactions of IMACS, 3(5):409– 425, October 1987. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kacur:1998:ADP

[KL98] J. Kačur and S. Luckhaus. Approximation of degenerate parabolic systems by nondegenerate elliptic and parabolic systems. Applied Numerical Mathematics: Transactions of IMACS, 26(3):307–326, March 2, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1998/26/
3/829.pdf.

Kim:2007:HOS

[KL07] Seongjai Kim and Hyeona Lim. High-order schemes for acoustic waveform simulation. Applied Numerical Mathematics: Transactions of IMACS, 57(4): 402–414, April 2007. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Knobloch:2009:LPS

[KL09] Petr Knobloch and Gert Lube. Local projection stabilization for advection– diffusion–reaction problems: One-level vs. two-level approach. Applied Numerical Mathematics: Transactions of IMACS, 59(12): 2891–2907, December 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kandel:2021:MPS

[KL21] Hom N. Kandel and Dong Liang. The mass-preserving solution-flux scheme for multi-layer interface parabolic equations. Applied Numerical Mathematics: Transactions of IMACS, 160(??):42– 64, February 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420302907.

Kaur:2023:STS

[KL23a] Avleen Kaur and S. H. Lui. Space-time spectral method for the Stokes problem. Applied Numerical Mathematics: Transactions of IMACS, 187(??):206–234, May 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423000442.

Kelly:2023:ASM

[KL23b] Cónall Kelly and Gabriel J. Lord. An adaptive splitting method for the Cox– Ingersoll–Ross process. Applied Numerical Mathematics: Transactions of IMACS, 186(??):252–273, April 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423000144.

Klibanov:2015:CER

[Kli15] Michael V. Klibanov. Carleman estimates for the regularization of ill-posed Cauchy problems. Applied Numerical Mathematics: Transactions of IMACS, 94(??): 46–74, August 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415000380.

Khan:2025:PCM

[KLR25] Ghulam Abbas Khan, Kaido Lätt, and Magda Rebelo. A polynomial collocation method for a class of singular fractional differential equations. Applied Numerical Mathematics: Transactions of IMACS, 207(??): 45–57, January 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424002216.

Krejic:2013:GMU

[KLS13] Natasa Krejić, Zorana Luzanin, and Irena Stojkovska. A gradient method for unconstrained optimization in noisy environment. Applied Numerical Mathematics: Transactions of IMACS, 70(??):1–21, August 2013. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927413000445.

Kucharik:2006:OSS

[KLSW06] Milan Kuchařík, Richard Liska, Stanly Steinberg, and Burton Wendroff. Optimallystable second-order accurate difference schemes for non-linear conservation laws in 3D. Applied Numerical Mathematics: Transactions of IMACS, 56(5): 589–607, May 2006. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kuo:2010:HCC

[KLSW10] Yueh-Cheng Kuo, Wen-Wei Lin, Shih-Feng Shieh, and Weichung Wang. A hyperplane-constrained continuation method for near singularity in coupled nonlinear Schrödinger equations. Applied Numerical Mathematics: Transactions of IMACS, 60(5):513–526, May 2010. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kim:2005:AFO

[KLY05] Sang Dong Kim, Yong Hun Lee, and Suh-Yuh Yang. Analysis of [H^{−1}, L^{2}, L^{2}] first-order system least squares for the incompressible Oseen type equations. Applied Numerical Mathematics: Transactions of IMACS, 52(1):77–88, January 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kacur:1995:SND

[KM95] Jozef Kačur and Karol Mikula. Solution of nonlinear diffusion appearing in image smoothing and edge detection. Applied Numerical Mathematics: Transactions of IMACS, 17(1):47– 59, May 1, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=17&
issue=1&aid=556.

Krause:2011:LSB

[KM11] Rolf Krause and Christina Mohr. Level set based multi-scale methods for large deformation contact problems. Applied Numerical Mathematics: Transactions of IMACS, 61(4):428–442, April 2011. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ke:2016:NDS

[KM16] Yifen Ke and Changfeng Ma. A note on “A dimensional split preconditioner for Stokes and linearized Navier–Stokes equations”. Applied Numerical Mathematics: Transactions of IMACS, 108(??): 223–225, October 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416300939.

Kang:2017:QRA

[KM17] Hongchao Kang and Junjie Ma. Quadrature rules and asymptotic expansions for two classes of oscillatory Bessel integrals with singularities of algebraic or logarithmic type. Applied Numerical Mathematics: Transactions of IMACS, 118(??):277–291, August 2017. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417300831.

Kopecz:2018:OCM

[KM18] S. Kopecz and A. Meister. On order conditions for modified Patankar–Runge–Kutta schemes. Applied Numerical Mathematics: Transactions of IMACS, 123(??): 159–179, January 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417301861.

Kaur:2019:TLI

[KM19] Deepti Kaur and R. K. Mohanty. Two-level implicit high order method based on half-step discretization for 1D unsteady biharmonic problems of first kind. Applied Numerical Mathematics: Transactions of IMACS, 139(??):1–14, May 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419300029.

Ke:2021:PCM

[KM21] Yifen Ke and Changfeng Ma. Prediction-correction matrix splitting iteration algorithm for a class of large and sparse linear systems. Applied Numerical Mathematics: Transactions of IMACS, 169(??): 256–272, November 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001951.

Kayal:2025:SMW

[KM25] Arnab Kayal and Moumita Mandal. Superconvergent method for weakly singular Fredholm–Hammerstein integral equations with nonsmooth solutions and its application. Applied Numerical Mathematics: Transactions of IMACS, 207(??): 24–44, January 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424002228.

Karimi:2020:NSS

[KME20] Akram Karimi, Khosrow Maleknejad, and Reza Ezzati. Numerical solutions of system of twodimensional Volterra integral equations via Legendre wavelets and convergence. Applied Numerical Mathematics: Transactions of IMACS, 156(??): 228–241, October 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420301471.

Ketcheson:2009:OIS

[KMG09] David I. Ketcheson, Colin B. Macdonald, and Sigal Gottlieb. Optimal implicit strong stability preserving Runge–Kutta methods. Applied Numerical Mathematics: Transactions of IMACS, 59(2):373–392, February 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Karimi:2021:IPA

[KMH21] Milad Karimi, Fridoun Moradlou, and Mojtaba Hajipour. On the ill-posed analytic continuation problem: an order optimal regularization scheme. Applied Numerical Mathematics: Transactions of IMACS, 161(??): 311–332, March 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303603.

Karlsen:2009:SGS

[KMR09] K. H. Karlsen, S. Mishra, and N. H. Risebro. Semi-Godunov schemes for multiphase flows in porous media. Applied Numerical Mathematics: Transactions of IMACS, 59(9):2322–2336, September 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Krebs:2010:AFB

[KMS10] Andreas Krebs, Matthias Maischak, and Ernst P. Stephan. Adaptive FEM– BEM coupling with a Schur complement error indicator. Applied Numerical Mathematics: Transactions of IMACS, 60(8):798–808, August 2010. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kalaiselvan:2019:PUN

[KMS19] Saravana Sankar Kalaiselvan, John J. H. Miller, and Valarmathi Sigamani. A parameter uniform numerical method for a singularly perturbed two-parameter delay differential equation. Applied Numerical Mathematics: Transactions of IMACS, 145(??):90–110, November 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        https://www.sciencedirect.
                                        com/science/article/pii/
                                        S0168927419301515.

Kumar:1993:SUS

[KN93] B. V. Ratish Kumar and K. B. Naidu. A streamline upwinding streamfunctionvorticity finite element analysis of Navier–Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 13(4): 335–344, November 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kenne:2008:SCP

[KN08] J. P. Kenne and L. J. Nkeungoue. Simultaneous control of production, preventive and corrective maintenance rates of a failureprone manufacturing system. Applied Numerical Mathematics: Transactions of IMACS, 58(2):180–194, February 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kaplan:2019:NLS

[KN19] Matthew Kaplan and David P. Nicholls. A nonlinear least squares framework for periodic grating identification with a high-order perturbation of surfaces implementation. Applied Numerical Mathematics: Transactions of IMACS, 143(??):20– 34, September 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419300650.

Knight:1994:FIN

[Kni94] Doyle D. Knight. A fully implicit Navier–Stokes algorithm using an unstructured grid and flux difference splitting. Applied Numerical Mathematics: Transactions of IMACS, 16(1–2): 101–128, December 1, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=16&
issue=1-2&aid=519.  A

Festschrift to honor Professor Robert Vichnevetsky on his 65th birthday. See erratum [Kni95].

Knight:1995:EFI

[Kni95] Doyle D. Knight. Erratum to: “A fully implicit Navier–Stokes algorithm using an unstructured grid and flux difference splitting” [Appl. Numer. Math.

16 (1994), no. 1–2, 101– 128, MR 95j:76073]. Applied Numerical Mathematics: Transactions of IMACS, 17(2):187, June 13, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=17&
                                        issue=2&aid=574.  See

[Kni94].

Kaneko:2003:WAP

[KNN03] Hideaki Kaneko, Richard D. Noren, and Boriboon Novaprateep. Wavelet applications to the Petrov– Galerkin method for Hammerstein equations. Applied Numerical Mathematics: Transactions of IMACS, 45(2–3):255–273, May 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kvaerno:1996:RKR

[KNO96] A. Kværnø, S. P. Nørsett, and B. Owren. Runge– Kutta research in Trondheim. Applied Numerical Mathematics: Transactions of IMACS, 22(1–3): 263–277, December 4, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.

cgi?year=1996&volume=22&
issue=1-3&aid=719. Special

issue celebrating the centenary of Runge–Kutta methods.

Klibanov:2016:NIN

[KNP16] Michael V. Klibanov, Loc H.

Nguyen, and Kejia Pan. Nanostructures imaging via numerical solution of a 3-D inverse scattering problem without the phase information. Applied Numerical Mathematics: Transactions of IMACS, 110(??): 190–203, December 2016. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301647.

Koch:2013:EEO

[KNT13] O. Koch, Ch. Neuhauser, and M. Thalhammer. Embedded exponential operator splitting methods for the time integration of nonlinear evolution equations. Applied Numerical Mathematics: Transactions of IMACS, 63(1):14–24, January 2013. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412001614.

Khajah:1992:NAS

[KO92] H. G. Khajah and E. L.

Ortiz. Numerical approximation of solutions of functional equations using the tau method. Applied Numerical Mathematics: Transactions of IMACS, 9(6):461– 474, May 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Khajah:1996:DDE

[KO96] H. G. Khajah and E. L. Ortiz. On a differential-delay equation arising in number theory. Applied Numerical Mathematics: Transactions of IMACS, 21(4):431– 437, November 22, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=21&
                                        issue=4&aid=702.

Kangro:2008:CSC

[KO08] Raul Kangro and Peeter Oja. Convergence of spline collocation for Volterra integral equations. Applied Numerical Mathematics: Transactions of IMACS, 58(10): 1434–1447, October 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kong:2025:CHO

[KOGL25] Linghua Kong, Songpei Ouyang, Rong Gao, and Haiyan Liang. Combined high order compact schemes for non-self-adjoint nonlinear Schrödinger equations. Applied Numerical Mathematics: Transactions of IMACS, 209(??): 242–257, March 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424002782.

Koko:2008:CAO

[Kok08] Jonas Koko. Convergence analysis of optimizationbased domain decomposition methods for a bonded structure. Applied Numerical Mathematics: Transactions of IMACS, 58(1):69– 87, January 2008. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Komori:2007:MCR

[Kom07]  Yoshio Komori.  Multi-

colored rooted tree analysis of the weak order conditions of a stochastic Runge–Kutta family. Applied Numerical Mathematics: Transactions of IMACS, 57(2):147– 165, February 2007. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kopriva:1986:SMM

[Kop86] David A. Kopriva. A spectral multidomain method for the solution of hyperbolic systems. Applied Numerical Mathematics: Transactions of IMACS, 2(3–5):221– 241, October 1986. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kopriva:1989:DDB

[Kop89] David A. Kopriva. Domain decomposition with both spectral and finite difference methods for the accurate computation of flows with shocks. Applied Numerical Mathematics: Transactions of IMACS, 6(1–2): 141–151, December 1989. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Spectral multi-domain methods (Paris, 1988).

Koren:1995:CIP

[Kor95] Barry Koren. Condition improvement for point relaxation in multigrid, subsonic Euler-flow computations. Applied Numerical Mathematics: Transactions of IMACS, 16(4):457– 469, April 6, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=16&
                                        issue=4&aid=551.

Keshavarz:2018:TWM

[KOR18] E. Keshavarz, Y. Ordokhani, and M. Razzaghi. The Taylor wavelets method for solving the initial and boundary value problems of Bratu-type equations. Applied Numerical Mathematics: Transactions of IMACS, 128(??): 205–216, June 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418300400.

Koschinski:2002:NMA

[Kos02]  Claus Koschinski.  New

methods for adapting and for approximating inverses as preconditioners. Applied Numerical Mathematics: Transactions of IMACS, 41(1):179–218, April 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/10/28/86/27/38/abstract.
html.

Koster:2012:NRN

[KOS^{+}12] M. Köster, A. Ouazzi, F. Schieweck, S. Turek, and P. Zajac. New robust nonconforming finite elements of higher order. Applied Numerical Mathematics: Transactions of IMACS, 62(3):166–184, March 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411002078.

Kirsiaed:2020:RSH

[KOS20] Evely Kirsiaed, Peeter Oja, and Gul Wali Shah. Rational spline histopolation of convex data. Applied Numerical Mathematics: Transactions of IMACS, 158(??): 377–391, December 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420302464.

Karakoc:2021:NIS

[KOS21] Seydi Battal Gazi Karakoc, Khaled Omrani, and Derya Sucu. Numerical investigations of shallow water waves via generalized equal width (GEW) equation. Applied Numerical Mathematics: Transactions of IMACS, 162(??):249–264, April 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420304062.

Knaub:2005:SMS

[KOW05] Karl R. Knaub, Robert E. O’Malley, Jr., and David B. Williams. The slow motion of shock layers for advection–diffusion–reaction equations. Applied Numerical Mathematics: Transactions of IMACS, 52(2– 3):299–310, February 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kozlovsky:1994:SPD

[Koz94] Gregory Kozlovsky. Solving partial differential equations using recursive grids. Applied Numerical Mathematics: Transactions of IMACS, 14(1–3):165–181, April 14, 1994. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=14&
issue=1-3&aid=468. Pro-

ceedings of the Third ARO Workshop on Adaptive Methods for Partial Differential Equations (Troy, NY, 1992).

Karageorghis:1992:CDE

[KP92] Andreas Karageorghis and Timothy N. Phillips. On the coefficients of differentiated expansions of ultraspherical polynomials. Applied Numerical Mathematics: Transactions of IMACS, 9(2):133–141, February 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kouibia:2001:ASP

[KP01] A. Kouibia and M. Pasadas.

Approximation by shape preserving interpolation splines. Applied Numerical Mathematics: Transactions of IMACS, 37(3):271–288, May 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/10/28/73/31/
                                        27/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/10/28/73/31/27/article.
                                        pdf.

Kouibia:2003:VBI

[KP03a] A. Kouibia and M. Pasadas. Variational bivariate interpolating splines with positivity constraints. Applied Numerical Mathematics: Transactions of IMACS, 44(4):507–526, March 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Koulisianis:2003:IPS

[KP03b] M. D. Koulisianis and T. S. Papatheodorou. Improving projected successive overrelaxation method for linear complementarity problems. Applied Numerical Mathematics: Transactions of IMACS, 45(1): 29–40, April 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kontoghiorghes:2007:SIN

[KP07] Erricos John Kontoghiorghes and Bernard Philippe. 2nd special issue on numerical algorithms, parallelism and applications. Applied Numerical Mathematics: Transactions of IMACS, 57(11–12): 1163, November/December 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kugelmann:2015:EUO

[KP15] Bernd Kugelmann and Roland Pulch. Existence and uniqueness of optimal solutions for multirate partial differential algebraic equations. Applied Numerical Mathematics: Transactions of IMACS, 97(??):69– 87, November 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415001117.

Kaluza:2018:OGA

[KP18] Andrzej Kaluza and Pawel Przybylowicz. Optimal global approximation of jump-diffusion SDEs via path-independent step-size control. Applied Numerical Mathematics: Transactions of IMACS, 128(??): 24–42, June 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418300333.

Kumar:2019:NSS

[KP19] Sushil Kumar and Cécile Piret. Numerical solution of space–time fractional PDEs using RBF-QR and Chebyshev polynomials. Applied Numerical Mathematics: Transactions of IMACS, 143(??): 300–315, September 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419300996.

Khenous:2006:HDS

[KPR06] Houari Boumediène Khenous, Julien Pommier, and Yves Renard. Hybrid discretization of the Signorini problem with Coulomb friction. Theoretical aspects and comparison of some numerical solvers. Applied Numerical Mathematics: Transactions of IMACS, 56(2):163– 192, February 2006. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kouibia:2012:NAD

[KPR12a] A. Kouibia, M. Pasadas, and M. L. Rodríguez. Numerical approximation by discrete interpolating variational splines. Applied Numerical Mathematics: Transactions of IMACS, 62(9):1109–1118, September 2012. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927411000717.

Kouibia:2012:VMS

[KPR12b] A. Kouibia, M. Pasadas, and M. L. Rodríguez. A variational method for solving Fredholm integral systems. Applied Numerical Mathematics: Transactions of IMACS, 62(9): 1041–1049, September 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741100208X.

Kapetina:2020:AUV

[KPRU20] M. N. Kapetina, A. Pisano, M. R. Rapaić, and E. Usai. Adaptive unit-vector law with time-varying gain for finite-time parameter estimation in LTI systems. Applied Numerical Mathematics: Transactions of IMACS, 155(??):16–28, ???? 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419302521.

Karaa:2015:PHE

[KPY15] Samir Karaa, Amiya K.

Pani, and Sangita Yadav. A priori hp-estimates for discontinuous Galerkin approximations to linear hyperbolic integro-differential equations. Applied Numerical Mathematics: Transactions of IMACS, 96(??): 1–23, October 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892741500077X.

Krivodonova:2013:ASDa

[KQ13a] Lilia Krivodonova and Ruibin Qin. An analysis of the spectrum of the discontinuous Galerkin method. Applied Numerical Mathematics: Transactions of IMACS, 64(1):1–18, February 2013. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927412001742.

Krivodonova:2013:ASDb

[KQ13b] Lilia Krivodonova and Ruibin Qin. An analysis of the spectrum of the discontinuous Galerkin method II: Nonuniform grids. Applied Numerical Mathematics: Transactions of IMACS, 71(??):41– 62, September 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927413000639.

Kumar:2010:LSH

[KR10] N. Kishore Kumar and G. Naga Raju. Leastsquares hp. Applied Numerical Mathematics: Transactions of IMACS, 60(1–2):38– 54, January/February 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kroker:2012:FVS

[KR12] I. Kröker and C. Rohde. Finite volume schemes for hyperbolic balance laws with multiplicative noise. Applied Numerical Mathematics: Transactions of IMACS, 62(4):441–456, April 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741100033X.

Kunisch:2015:GTS

[KR15] Karl Kunisch and Stefan H.

Reiterer. A Gautschi timestepping approach to optimal control of the wave equation. Applied Numerical Mathematics: Transactions of IMACS, 90(??): 55–76, April 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414001937.

Kulkarni:2018:DMP

[KR18] Rekha P. Kulkarni and Gobinda Rakshit. Discrete modified projection method for Urysohn integral equations with smooth kernels. Applied Numerical Mathematics: Transactions of IMACS, 126(??):180–198, April 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417302581.

Ku:2020:NIM

[KR20] JaEun Ku and Lothar Reichel. A novel iterative method for discrete Helmholtz decomposition. Applied Numerical Mathematics: Transactions of IMACS, 151(??):161–171, May 2020. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419303769.

Kovtunenko:2025:FAD

[KR25] Victor A. Kovtunenko and Yves Renard. FEM approximation of dynamic contact problem for fracture under fluid volume control using generalized HHT-α and semi-smooth Newton methods. Applied Numerical Mathematics: Transactions of IMACS, 218(??): 148–158, December 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425001503.

Kraaijevanger:1992:MNC

[Kra92] J. F. B. M. Kraaijevanger. Maximum norm contractivity of discretization schemes for the heat equation. Applied Numerical Mathematics: Transactions of IMACS, 9(6):475–492, May 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Koulouri:2016:CDM

[KRBK16] Alexandra Koulouri, Ville Rimpiläinen, Mike Brookes, and Jari P. Kaipio. Compensation of domain modelling errors in the inverse source problem of the Poisson equation: Application in electroencephalographic imaging. Applied Numerical Mathematics: Transactions of IMACS, 106(??): 24–36, August 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416300307.

Kress:2007:EED

[Kre07] Wendy Kress. Error estimates for deferred correction methods in time. Applied Numerical Mathematics: Transactions of IMACS, 57(3):335–353, March 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Krukier:1999:CAT

[Kru99] L. A. Krukier. Convergence acceleration of triangular iterative methods based on the skew-symmetric part of the matrix. Applied Numerical Mathematics: Transactions of IMACS, 30(2– 3):281–290, June 10, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1999&volume=30&
issue=2-3&aid=981.

Kraaijevanger:1989:ASE

[KS89] J. F. B. M. Kraaijevanger and M. N. Spijker. Algebraic stability and error propagation in Runge–Kutta methods. Applied Numerical Mathematics: Transactions of IMACS, 5(1–2):71– 87, February 1989. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Recent theoretical results in numerical ordinary differential equations.

Kaminski:1991:SIP

[KS91] Michael Kaminski and Avram Sidi. Solution of an integer programming problem related to convergence of rows of Padé approximants. Applied Numerical Mathematics: Transactions of IMACS, 8(3):217–223, October 1991. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Karamanos:2000:HOS

[KS00] G.-S. Karamanos and S. J. Sherwin. A high order splitting scheme for the Navier–Stokes equations with variable viscosity. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4):455–462, May 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/61/27/
73/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/61/27/73/article.
pdf.

Kopteva:2001:ADC

[KS01] Natalia Kopteva and Martin Stynes. Approximation of derivatives in a convection-diffusion twopoint boundary value problem. Applied Numerical Mathematics: Transactions of IMACS, 39(1):47–60, October 2001. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/79/27/30/abstract.
html.

Kim:2002:LSM

[KS02] Sang Dong Kim and Byeong Chun Shin. H^{1} least-squares method for the velocitypressure-stress formulation of Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 40(4):451–465, March 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/10/28/85/34/28/abstract.
html.

Kopteva:2004:NAS

[KS04] Natalia Kopteva and Martin Stynes. Numerical analysis of a singularly perturbed nonlinear reaction–diffusion problem with multiple solutions. Applied Numerical Mathematics: Transactions of IMACS, 51(2– 3):273–288, November 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Klapper:2007:LJA

[KS07] I. Klapper and T. Shaw. A large jump asymptotic framework for solving elliptic and parabolic equations with interfaces and strong coefficient discontinuities. Applied Numerical Mathematics: Transactions of IMACS, 57(5–7):657–671, May 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kim:2008:LSR

[KS08] Dongjin Kim and Dan Stanescu. Low-storage Runge–Kutta methods for stochastic differential equations. Applied Numerical Mathematics: Transactions of IMACS, 58(10): 1479–1502, October 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kim:2009:APL

[KS09a] Sang Dong Kim and Byeong-Chun Shin. Adjoint pseudospectral least-squares methods for an elliptic boundary value problem. Applied Numerical Mathematics: Transactions of IMACS, 59(2):334–348, February 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kozakevicius:2009:EAM

[KS09b] A. J. Kozakevicius and L. C. C. Santos. ENO adaptive method for solving one-dimensional conservation laws. Applied Numerical Mathematics: Transactions of IMACS, 59(9): 2337–2355, September 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kulikov:2009:ANI

[KS09c] G. Yu. Kulikov and S. K.

Shindin. Adaptive nested implicit Runge–Kutta formulas of Gauss type. Applied Numerical Mathematics: Transactions of IMACS, 59(3–4):707–722, March/ April 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Koukouvinos:2010:ILB

[KS10] Christos Koukouvinos and Dimitris E. Simos. Improving the lower bounds on inequivalent Hadamard matrices through orthogonal designs and meta-programming techniques. Applied Numerical Mathematics: Transactions of IMACS, 60(4): 370–377, April 2010. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Klein:2022:SAK

[KS22] Christian Klein and Nikola Stoilov. Spectral approach to Korteweg–de Vries equations on the compactified real line. Applied Numerical Mathematics: Transactions of IMACS, 177(??): 160–170, July 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422000514.

Kashif:2025:EUU

[KS25] Mohd Kashif and Manpal Singh. Existence, uniqueness and Ulam–Hyers stability result for variable order fractional predator–prey system and it’s numerical solution. Applied Numerical Mathematics: Transactions of IMACS, 207(??): 193–209, January 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742400223X.

Khader:2021:SCM

[KSHB21] M. M. Khader, Khaled M.

Saad, Zakia Hammouch, and Dumitru Baleanu. A spectral collocation method for solving fractional KdV and KdV–Burgers equations with non-singular kernel derivatives. Applied Numerical Mathematics: Transactions of IMACS, 161(??): 137–146, March 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303317.

Kumar:2025:EEP

[KSJY25] Naresh Kumar, Ajeet Singh, Ram Jiwari, and J. Y. Yuan. Error estimates with polynomial growth O(ϵ^{−∞}) for the HHO method on polygonal meshes of the Allen–Cahn model. Applied Numerical Mathematics: Transactions of IMACS, 211(??): 78–102, May 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425000017.

Kiyadeh:2025:NAM

[KSK^{+}25] Sanaz Hami Hassan Kiyadeh, Hosein Saadat, Ramin Goudarzi Karim, Ali Safaie, and Fayyaz Khodadosti. A novel amplifying methodology in Gauss–Legendre IRK integrations to cope with high-frequency stiff problems. Applied Numerical Mathematics: Transactions of IMACS, 218(??):43– 57, December 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425001473.

Khodayari-Samghabadi:2016:EEP

[KSMMM16] S. Khodayari-Samghabadi, S. H. Momeni-Masuleh, and A. Malek. An explicitextended penalty Galerkin method for solving an incompressible two-phase flow. Applied Numerical Mathematics: Transactions of IMACS, 107(??):48– 63, September 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416300551.

Kundakcioglu:2010:MIL

[KSP10] O. Erhun Kundakcioglu, Onur Seref, and Panos M. Pardalos. Multiple instance learning via margin maximization. Applied Numerical Mathematics: Transactions of IMACS, 60(4): 358–369, April 2010. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kogan:2016:FFI

[KSSS16] Tamara Kogan, Luba Sapir, Amir Sapir, and Ariel Sapir. The Fibonacci family of iterative processes for solving nonlinear equations. Applied Numerical Mathematics: Transactions of IMACS, 110(??): 148–158, December 2016. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301623.

Karasozen:2005:CPF

[KT05] B. Karasözen and V. G. Tsybulin. Cosymmetry preserving finite-difference methods for convection equations in a porous medium. Applied Numerical Mathematics: Transactions of IMACS, 55(1):69–82, September 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kamranian:2020:ASE

[KTD20] Maryam Kamranian, Mehdi Tatari, and Mehdi Dehghan. Analysis of the stabilized element free Galerkin approximations to the Stoke’s equations. Applied Numerical Mathematics: Transactions of IMACS, 150(??): 325–340, April 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419302697.

Kanbar:2020:WBC

[KTK20] F. Kanbar, R. Touma, and C. Klingenberg. Wellbalanced central schemes for the one and two-dimensional Euler systems with gravity. Applied Numerical Mathematics: Transactions of IMACS, 156(??):608– 626, October 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420301549.

Koomullil:2003:IAS

[KTS03] R. P. Koomullil, D. S. Thompson, and B. K. Soni. Iced airfoil simulation using generalized grids. Applied Numerical Mathematics: Transactions of IMACS, 46(3–4):319–330, September 2003. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kumar:2024:CNW

[KTYY24] N. Kumar, S. Toprakseven, N. Singh Yadav, and J. Y. Yuan. A Crank–Nicolson WG-FEM for unsteady 2D convection-diffusion equation with nonlinear reaction term on layer adapted mesh. Applied Numerical Mathematics: Transactions of IMACS, 201(??): 322–346, July 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000618.

Kurgan:1998:BES

[Kur98] E. Kurgan. A boundary element solution of the inhomogeneous magnetostatic problems. Applied Numerical Mathematics: Transactions of IMACS, 28(2–4):343–358, October 1, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1998/28/
2-4/918.pdf.

Kurkcu:2023:ESC

[Kür23] Ömür Kivanç Kürkçü. An exclusive spectral computational approach based on quadratic orthoexponential polynomials for solving integro-differential equations with delays on the real line. Applied Numerical Mathematics: Transactions of IMACS, 184(??):1– 17, February 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422002604.

Kuznetsov:1990:MLD

[Kuz90] Yu. A. Kuznetsov. Multilevel domain decomposition methods. Applied Numerical Mathematics: Transactions of IMACS, 6(4):303– 314, May 1990. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Karoui:1995:NMV

[KV95] Abderrazek Karoui and Rémi Vaillancourt. A numerical method for vanishinglag delay differential equations. Applied Numerical Mathematics: Transactions of IMACS, 17(4):383– 395, September 11, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=17&
                                        issue=4&aid=573.

Kandilarov:2007:IIM

[KV07] Juri D. Kandilarov and Lubin G. Vulkov. The immersed interface method for two-dimensional heatdiffusion equations with singular own sources. Applied Numerical Mathematics: Transactions of IMACS, 57(5–7):486–497, May 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Khademi:2020:EIE

[KV20] Ali Khademi and Jon Eivind Vatne. Estimation of the interpolation error for semiregular prismatic elements. Applied Numerical Mathematics: Transactions of IMACS, 156(??): 174–191, October 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420301422.

Kolomenskiy:2015:ADV

[KvyS15] Dmitry Kolomenskiy, Romain Nguyen van yen, and Kai Schneider. Analysis and discretization of the volume penalized Laplace operator with Neumann boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 95(??): 238–249, September 2015. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414000166.

Kreiss:1993:SDD

[KW93] Heinz-O. O. Kreiss and Lixin Wu. On the stability definition of difference approximations for the initial boundary value problem. Applied Numerical Mathematics: Transactions of IMACS, 12(1–3):213–227, May 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/016892749390119C.

Special issue to honor Professor Saul Abarbanel on his sixtieth birthday (Neveh, 1992).

Knoth:1995:NMS

[KW95] Oswald Knoth and Ralf Wolke. Numerical methods for the solution of large kinetic systems. Applied Numerical Mathematics: Transactions of IMACS, 18(1–3):211–221, September 27, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=18&
                                        issue=1-3&aid=602. Sev-

enth Conference on the Numerical Treatment of Differential Equations (Halle, 1994).

Knoth:1998:IER

[KW98] Oswald Knoth and Ralf Wolke. Implicit–explicit Runge–Kutta methods for computing atmospheric reactive flows. Applied Numerical Mathematics: Transactions of IMACS, 28(2–4):327–341, October 1, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1998/28/
                                        2-4/917.pdf.

Kucharik:2010:MFS

[KW10] Milan Kucharik and Burton Wendroff. Multidimensional first and second order symmetric Strang splitting for hyperbolic systems. Applied Numerical Mathematics: Transactions of IMACS, 60(1–2):74–82, January/February 2010. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Krause:2012:PCS

[KW12] Rolf Krause and Mirjam Walloth. Presentation and comparison of selected algorithms for dynamic contact based on the Newmark scheme. Applied Numerical Mathematics: Transactions of IMACS, 62(10): 1393–1410, October 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412001006.

Kulikov:2020:VSD

[KW20] G. Yu. Kulikov and R. Weiner.

Variable-stepsize doubly quasiconsistent singly diagonally implicit two-step peer pairs for solving stiff ordinary differential equations. Applied Numerical Mathematics: Transactions of IMACS, 154(??): 223–242, August 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420301185.

Kang:2021:NEE

[KW21] Hongchao Kang and Hong Wang. Numerical evaluation and error analysis of many different oscillatory Bessel transforms via confluent hypergeometric function. Applied Numerical Mathematics: Transactions of IMACS, 160(??):23– 41, February 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420302932.

Kwan:2009:ESG

[Kwa09]  Y.-Y. Kwan.  Efficient

spectral-Galerkin methods for polar and cylindrical geometries. Applied Numerical Mathematics: Transactions of IMACS, 59(1):170– 186, January 2009. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kweon:2000:HBP

[Kwe00] Jae Ryong Kweon. Hierarchical basis a posteriori error estimates for compressible Stokes flows. Applied Numerical Mathematics: Transactions of IMACS, 32(1):53–68, January 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/59/28/
30/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/59/28/30/article.
pdf.

Kweon:2001:MFE

[Kwe01] Jae Ryong Kweon. A mixed finite element method for a compressible Stokes problem with high Reynolds number. Applied Numerical Mathematics: Transactions of IMACS, 38(1–2): 87–103, July 2001. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/77/27/
31/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/77/27/31/article.
pdf.

Kweon:2003:OEE

[Kwe03] Jae Ryong Kweon. Optimal error estimate for a mixed finite element method for compressible Navier–Stokes system. Applied Numerical Mathematics: Transactions of IMACS, 45(2–3): 275–292, May 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kirby:2000:DGS

[KWLK00] R. M. Kirby, T. C. Warburton, I. Lomtev, and G. E. Karniadakis. A discontinuous Galerkin spectral/hp method on hybrid grids. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4):393–405, May 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/61/27/
67/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/61/27/67/article.
pdf.

Kim:2019:HOD

[KwS19] Mi-Young Kim and Dong wook Shin. A high order discontinuous Galerkin method with Lagrange multipliers for second-order elliptic problems. Applied Numerical Mathematics: Transactions of IMACS, 135(??): 47–68, January 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418301600.

Kang:2024:UES

[KWY24] Yuanyuan Kang, Jindi Wang, and Yin Yang. Unconditionally energy stable highorder BDF schemes for the molecular beam epitaxial model without slope selection. Applied Numerical Mathematics: Transactions of IMACS, 206(??): 190–209, December 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424002010.

Kaneko:1991:NSW

[KX91] Hideaki Kaneko and Yuesheng Xu. Numerical solutions for weakly singular Fredholm integral equations of the second kind. Applied Numerical Mathematics: Transactions of IMACS, 7(2):167–177, February 1991. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kurtz:2003:EUC

[KX03] Jason Kurtz and Christos Xenophontos. On the effects of using curved elements in the approximation of the Reissner–Mindlin plate by the p version of the finite element method. Applied Numerical Mathematics: Transactions of IMACS, 46(2):231–246, August 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kaneko:1992:DKM

[KXK92] Hideaki Kaneko, Yuesheng Xu, and Gilbert Kerr. Degenerate kernel method for multi-variable Hammerstein equations. Applied Numerical Mathematics: Transactions of IMACS, 10(6): 473–479, November 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Krivodonova:2004:SDL

[KXR^{+}04] L. Krivodonova, J. Xin, J.- F. Remacle, N. Chevaugeon, and J. E. Flaherty. Shock detection and limiting with discontinuous Galerkin methods for hyperbolic conservation laws. Applied Numerical Mathematics: Transactions of IMACS, 48(3–4): 323–338, March 2004. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kincaid:2003:VGI

[KYC03] David R. Kincaid, David M. Young, and Jen-Yuan Chen. Variations of the GMRES iterative method. Applied Numerical Mathematics: Transactions of IMACS, 45(1):3–10, April 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kyncheva:2017:CNH

[KYI17] Veselina K. Kyncheva, Viktor V. Yotov, and Stoil I. Ivanov. Convergence of Newton, Halley and Chebyshev iterative methods as methods for simultaneous determination of multiple polynomial zeros. Applied Numerical Mathematics: Transactions of IMACS, 112(??): 146–154, February 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416302148.

Kossioris:2013:FEA

[KZ13] Georgios T. Kossioris and Georgios E. Zouraris. Finite element approximations for a linear fourth-order parabolic SPDE in two and three space dimensions with additive space–time white noise. Applied Numerical Mathematics: Transactions of IMACS, 67(??):243–261, May 2013. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412000116.

Kumar:2021:EML

[KZ21] Sachin Kumar and Dia Zeidan. An efficient Mittag-Leffler kernel approach for time-fractional advectionreaction–diffusion equation. Applied Numerical Mathematics: Transactions of IMACS, 170(??):190–207, December 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421002221.

Kzaz:1992:CAS

[Kza92] M. Kzaz. Convergence acceleration of some Gaussian quadrature formulas for analytic functions. Applied Numerical Mathematics: Transactions of IMACS, 10(6):481–496, November 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Kzaz:1999:CAG

[Kza99] M. Kzaz. Convergence acceleration of the Gauss– Laguerre quadrature formula. Applied Numerical Mathematics: Transactions of IMACS, 29(2):201–220, February 1, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1999/29/
                                        2/940.pdf.

Li:2011:TLP

[LA11] Yuan Li and Rong An. Two-level pressure projection finite element methods for Navier–Stokes equations with nonlinear slip boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 61(3):285–297, March 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Lang:2012:NAD

[LA12] Holger Lang and Martin Arnold. Numerical aspects in the dynamic simulation of geometrically exact rods. Applied Numerical Mathematics: Transactions of IMACS, 62(10): 1411–1427, October 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412000979.

Li:2021:TEA

[LA21] Yuan Li and Rong An. Temporal error analysis of Euler semi-implicit scheme for the magnetohydrodynamics equations with variable density. Applied Numerical Mathematics: Transactions of IMACS, 166(??): 146–167, August 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001057.

Laburta:1998:SAI

[Lab98] M. P. Laburta. Starting algorithms for implicit Runge– Kutta–Nyström methods. Applied Numerical Mathematics: Transactions of IMACS, 27(3):233–251, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Laburta:1999:SAI

[Lab99]  M. P. Laburta.  Start-

ing algorithms for the iterations of the RKN–Gauss methods. Applied Numerical Mathematics: Transactions of IMACS, 31(1): 81–101, September 1, 1999.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1999&volume=31&
                                        issue=1&aid=989.

Loubenets:2009:HAF

[LAH09] A. Loubenets, T. Ali, and M. Hanke. Highly accurate finite element method for one-dimensional elliptic interface problems. Applied Numerical Mathematics: Transactions of IMACS, 59(1):119–134, January 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Lai:2009:IRL

[Lai09] Yongzeng Lai. Intermediate rank lattice rules and applications to finance. Applied Numerical Mathematics: Transactions of IMACS, 59(1):1–20, January 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Lamichhane:2013:TSF

[Lam13] Bishnu P. Lamichhane. Two simple finite element methods for Reissner– Mindlin plates with clamped boundary condition. Applied Numerical Mathematics: Transactions of IMACS, 72(??):91–98, October 2013.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413000755.

Lamura:2024:TFP

[Lam24] A. Lamura. Tethered flexible polymer under oscillatory linear flow. Applied Numerical Mathematics: Transactions of IMACS, 205(??): 206–214, November 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001879.

Lang:1995:TDF

[Lan95] Jens Lang. Two-dimensional fully adaptive solutions of reaction-diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 18(1–3):223–240, September 27, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=18&
issue=1-3&aid=605. Sev-

enth Conference on the Numerical Treatment of Differential Equations (Halle, 1994).

Lang:1997:AFR

[Lan97] Jens Lang. Adaptive FEM for reaction–diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 26(1–2):105–116, December 22, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1997/26/
1-2/840.pdf.

Lopez-Agredo:2023:TNA

[LARGVR23] Jorge L. López-Agredo, Diego A. Rueda-Gómez, and Élder J. Villamizar-Roa. Theoretical and numerical analysis of a parabolic system with chemoattraction modeling the growth of glioma cells. Applied Numerical Mathematics: Transactions of IMACS, 186(??): 143–163, April 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423000089.

Laurita:2017:NMSb

[Lau17a]  C. Laurita.  A numeri-

cal method for the solution of exterior Neumann problems for the Laplace equation in domains with corners. Applied Numerical Mathematics: Transactions of IMACS, 119(??): 248–270, September 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417300314.

Laurita:2017:NMSa

[Lau17b] Concetta Laurita. A numerical method for the solution of integral equations of Mellin type. Applied Numerical Mathematics: Transactions of IMACS, 116(??): 215–229, June 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417300296.

Lavastre:1994:SAS

[Lav94] Hélène Lavastre. On the stochastic acceleration of sequences of random variables. Applied Numerical Mathematics: Transactions of IMACS, 15(1):77– 98, August 1, 1994. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=15&
issue=1&aid=502.

Layton:2008:CCS

[Lay08]  Anita T. Layton.  On

the choice of correctors for semi-implicit Picard deferred correction methods. Applied Numerical Mathematics: Transactions of IMACS, 58(6):845–858, June 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Layton:2009:ESD

[Lay09] Anita T. Layton. On the efficiency of spectral deferred correction methods for time-dependent partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 59(7):1629–1643, July 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Liang:2020:EIS

[LAZ20] Zhao-Zheng Liang, Owe Axelsson, and Guo-Feng Zhang. Efficient iterative solvers for a complex valued two-by-two block linear system with application to parabolic optimal control problems. Applied Numerical Mathematics: Transactions of IMACS, 152(??):422–445, June 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419303198.

Lu:2021:MOM

[LB21] Min Lu and Zheng-Jian Bai. A modified optimization method for robust partial quadratic eigenvalue assignment using receptances and system matrices. Applied Numerical Mathematics: Transactions of IMACS, 159(??):73–92, January 2021.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302531.

Li:2023:LFD

[LB23] Yanjun Li and Hai Bi. A locking-free discontinuous Galerkin method for linear elastic Steklov eigenvalue problem. Applied Numerical Mathematics: Transactions of IMACS, 188(??): 21–41, June 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423000600.

Lynov:2000:AES

[LBCN00] J. P. Lynov, K. Bergeron, E. A. Coutsias, and A. H. Nielsen. An accurate and efficient spectral method for studies of the dynamical properties of forced, circular shear layers. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4):175–181, May 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/61/27/
42/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/61/27/42/article.
pdf.

Lee:2013:MMA

[LBLT13] T. E. Lee, M. J. Baines, S. Langdon, and M. J. Tindall. A moving mesh approach for modelling avascular tumour growth. Applied Numerical Mathematics: Transactions of IMACS, 72(??):99–114, October 2013. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927413000780.

Lima:1999:AEN

[LC99] P. M. Lima and M. P. Carpentier. Asymptotic expansions and numerical approximation of non-linear degenerate boundary-value problems. Applied Numerical Mathematics: Transactions of IMACS, 30(1):93– 111, May 10, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1999/30/
                                        1/952.pdf.

Li:2002:MFT

[LC02] Yaqing Li and Benito Chen. Modeling of flow and transport at the microscale. Applied Numerical Mathematics: Transactions of IMACS, 40(1–2):245–259, January 2002. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/85/27/47/abstract.
html.

Li:2019:DSS

[LC19] Shenghao Li and Min Chen.

Diamond shaped standing wave patterns of a twodimensional Boussinesq system. Applied Numerical Mathematics: Transactions of IMACS, 141(??): 91–101, ???? 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418301521.

Liu:2020:AFE

[LC20] D. J. Liu and Z. R. Chen.

The adaptive finite element method for the P -Laplace problem. Applied Numerical Mathematics: Transactions of IMACS, 152(??): 323–337, June 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419303277.

Liu:2021:NLD

[LC21] Zhengguang Liu and Shuangshuang Chen. Novel linear decoupled and unconditionally energy stable numerical methods for the modified phase field crystal model. Applied Numerical Mathematics: Transactions of IMACS, 163(??): 1–14, May 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000064.

Li:2024:IUE

[LC24] Jiyong Li and Qianyu Chen. Improved uniform error bounds on an exponential wave integrator method for the nonlinear Schrödinger equation with wave operator and weak nonlinearity. Applied Numerical Mathematics: Transactions of IMACS, 201(??):488–513, July 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000850.

Li:2025:NMA

[LC25] Hongwei Li and Xinyue Chen. Numerical method and analysis for fluidstructure model on unbounded domains. Applied Numerical Mathematics: Transactions of IMACS, 217(??):255–277, November 2025. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425001266.

Lin:2020:GSA

[LCH20] Xiuxiu Lin, Yanping Chen, and Yunqing Huang. Galerkin spectral approximation of optimal control problems with L^{2}-norm control constraint. Applied Numerical Mathematics: Transactions of IMACS, 150(??): 418–432, April 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419302892.

Lang:2003:TDM

[LCHR03] Jens Lang, Weiming Cao, Weizhang Huang, and Robert D. Russell. A two-dimensional moving finite element method with local refinement based on a posteriori error estimates. Applied Numerical Mathematics: Transactions of IMACS, 46(1): 75–94, July 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Li:2020:TGM

[LCHW20] Qingfeng Li, Yanping Chen, Yunqing Huang, and Yang Wang. Two-grid methods for semilinear time fractional reaction diffusion equations by expanded mixed finite element method. Applied Numerical Mathematics: Transactions of IMACS, 157(??):38–54, November 2020. CODEN ANMAEL.

ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0168927420301689.

Liu:2012:MSF

[LCJQ12] Yang G. Liu, Weng Cho Chew, Lijun Jiang, and Zhiguo Qian. A memory saving fast A-EFIE solver for modeling low-frequency large-scale problems. Applied Numerical Mathematics: Transactions of IMACS, 62(6):682–698, June 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927410002163.

Lee:2022:ESF

[LCK22] Chaeyoung Lee, Yongho Choi, and Junseok Kim. An explicit stable finite difference method for the Allen–Cahn equation. Applied Numerical Mathematics: Transactions of IMACS, 182(??):87–99, December 2022. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742200201X.

Liu:2018:NFD

[LCL18] Zhengguang Liu, Aijie Cheng, and Xiaoli Li. A novel finite difference discrete scheme for the time fractional diffusion-wave equation. Applied Numerical Mathematics: Transactions of IMACS, 134(??):17– 30, December 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418301491.

Liu:2017:FST

[LCLW17] Zhengguang Liu, Aijie Cheng, Xiaoli Li, and Hong Wang. A fast solution technique for finite element discretization of the space–time fractional diffusion equation. Applied Numerical Mathematics: Transactions of IMACS, 119(??):146–163, September 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417300995.

Lovisetto:2022:OIF

[LCM22] M. Lovisetto, D. Clamond, and B. Marcos. Optimized integrating factor technique for Schrödinger-like equations. Applied Numerical Mathematics: Transactions of IMACS, 178(??): 329–336, August 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001040.

Lovisetto:2024:IFT

[LCM24] Martino Lovisetto, Didier Clamond, and Bruno Marcos. Integrating factor techniques applied to the Schrödinger-like equations. Comparison with Split-Step methods. Applied Numerical Mathematics: Transactions of IMACS, 197(??): 258–271, March 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423002957.

Li:2019:PTL

[LCS19] Shishun Li, Rongliang Chen, and Xinping Shao. Parallel two-level space–time hybrid Schwarz method for solving linear parabolic equations. Applied Numerical Mathematics: Transactions of IMACS, 139(??): 120–135, May 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419300273.

Ladeinde:2001:TSC

[LCVG01] Foluso Ladeinde, Xiaodan Cai, Miguel R. Visbal, and Datta V. Gaitonde. Turbulence spectra characteristics of high order schemes for direct and large eddy simulation. Applied Numerical Mathematics: Transactions of IMACS, 36(4):

447–474, March 2001. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/65/32/
30/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/65/32/30/article.
pdf.

Li:2020:OCM

[LCW20] Xiang-Gui Li, Yongyong Cai, and Pengde Wang. Operator-compensation methods with mass and energy conservation for solving the Gross–Pitaevskii equation. Applied Numerical Mathematics: Transactions of IMACS, 151(??): 337–353, May 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300052.

Liu:2021:PAT

[LCZ21] Qiaohua Liu, Cuiping Chen, and Qian Zhang. Perturbation analysis for total least squares problems with linear equality constraint. Applied Numerical Mathematics: Transactions of IMACS, 161(??):69–81, March 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303329.

Liu:2023:SOS

[LCZ23] Yufen Liu, Wanrong Cao, and Zhongqiang Zhang. Strong 1.5 order scheme for second-order stochastic differential equations without Levy area. Applied Numerical Mathematics: Transactions of IMACS, 184(??): 273–284, February 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742200277X.

Liu:2025:EAA

[LCZ25] Weizhi Liu, Hu Chen, and Mahmoud Zaky. Error analysis of an ADI scheme for the two-dimensional fractal mobile/immobile transport equation with weakly singular solutions. Applied Numerical Mathematics: Transactions of IMACS, 210(??):113–122, April 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424003325.

Lima:1997:EMV

[LD97] Pedro Lima and Teresa Diogo. An extrapolation method for a Volterra integral equation with weakly singular kernel. Applied Numerical Mathematics: Transactions of IMACS, 24(2– 3):131–148, June 19, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=24&
issue=2-3&aid=772. Volterra

centennial (Tempe, AZ, 1996).

Lucht:2002:QLP

[LD02] W. Lucht and K. Debrabant. On quasi-linear PDAEs with convection: Applications, indices, numerical solution. Applied Numerical Mathematics: Transactions of IMACS, 42(1– 3):297–314, August 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Li:2010:ANI

[LD10] Chong-Jun Li and Catterina Dagnino. An adaptive numerical integration algorithm for polygons. Applied Numerical Mathematics: Transactions of IMACS, 60(3):165–175, March 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Liang:2021:EPT

[LD21] Zhao-Zheng Liang and Yan Dou. Efficient preconditioning techniques for velocity tracking of Stokes control problem. Applied Numerical Mathematics: Transactions of IMACS, 165(??):

322–338, July 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421000702.

Long:2022:CAC

[LD22] Xiaonian Long and Qianqian Ding. Convergence analysis of a conservative finite element scheme for the thermally coupled incompressible inductionless MHD problem. Applied Numerical Mathematics: Transactions of IMACS, 182(??): 176–195, December 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422001945.

Li:2025:ADF

[LD25] Xiaolin Li and Haiyun Dong. Analysis of a divergencefree element-free Galerkin method for the Navier– Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 217(??):73–95, November 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0168927425001199.

Li:2024:SCT

[LDC24] Yan Li, Nan Deng, and Wanrong Cao. Strong convergence of the tamed Euler–Maruyama method for stochastic singular initial value problems with nonglobally Lipschitz continuous coefficients. Applied Numerical Mathematics: Transactions of IMACS, 205(??):60– 86, November 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001739.

Lima:2020:P

[LDF^{+}20] Pedro Lima, Teresa Diogo, Luisa Fermo, Neville J. Ford, and Yuesheng Xu. Preface. Applied Numerical Mathematics: Transactions of IMACS, 149(??): 1–2, ???? 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419303265.

Lu:2024:PTM

[LDH^{+}24] Yue Lu, Zheng-Peng Dong, Zhi-Qiang Hu, Hong-Min Ma, and Dong-Yang Xue. A penalty-type method for solving inverse optimal value problem in second-order conic programming. Applied Numerical Mathematics: Transactions of IMACS, 198(??):419–427, April 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424000266.

Leibinger:2016:PCO

[LDIW16] Julia Leibinger, Michael Dumbser, Uwe Iben, and Isabell Wayand. A pathconservative Osher-type scheme for axially symmetric compressible flows in flexible visco-elastic tubes. Applied Numerical Mathematics: Transactions of IMACS, 105(??):47–63, July 2016. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416000234.

Limem:2014:NNM

[LDP^{+}14] A. Limem, G. Delmaire, M. Puigt, G. Roussel, and D. Courcot. Nonnegative matrix factorization under equality constraints

— a study of industrial source identification. Applied Numerical Mathematics: Transactions of IMACS, 85(??):1–15, November 2014. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414001007.

LeRoux:2012:NSP

[Le 12] Marie-Noelle Le Roux. Numerical solution of a parabolic problem arising in finance. Applied Numerical Mathematics: Transactions of IMACS, 62(7):815–832, July 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S016892741200030X.

Lottati:1994:SOG

[LE94] Itzhak Lottati and Shmuel Eidelman. A secondorder Godunov scheme on a spatial adapted triangular grid. Applied Numerical Mathematics: Transactions of IMACS, 14(1–3): 353–365, April 14, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=14&
issue=1-3&aid=465. Pro-

ceedings of the Third ARO Workshop on Adaptive Methods for Partial Differential Equations (Troy, NY, 1992).

Lee:1994:HON

[Lee94] Ding Lee. A high-order numerical method for solving ocean acoustic problems. Applied Numerical Mathematics: Transactions of IMACS, 16(1–2):271–281, December 1, 1994. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.

                                        cgi?year=1994&volume=16&
                                        issue=1-2&aid=517.  A

Festschrift to honor Professor Robert Vichnevetsky on his 65th birthday.

Lee:2010:CMO

[Lee10] D. W. Lee. Classical multiple orthogonal polynomials of Angelesco system. Applied Numerical Mathematics: Transactions of IMACS, 60(12):1342–1351, December 2010. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Lee:2023:FEA

[Lee23] Eunjung Lee. An L^{2}-finite element approximation of the solution to div/curlsystems with nonempty null space. Applied Numerical Mathematics: Transactions of IMACS, 192(??): 70–83, October 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423001678.

Leimkuhler:1999:RAR

[Lei99] Benedict Leimkuhler. Reversible adaptive regularization methods for atomic Nbody problems in applied fields. Applied Numerical Mathematics: Transactions of IMACS, 29(1):31–43, January 1, 1999. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1999/29/
1/935.pdf.

Leimkuhler:2002:EMT

[Lei02] Ben Leimkuhler. An efficient multiple time-scale reversible integrator for the gravitational N-body problem. Applied Numerical Mathematics: Transactions of IMACS, 43(1–2):175–190, October 2002. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Lembarki:1988:CAL

[Lem88] Alami Lembarki. Convergence acceleration of limit k-periodic continued fractions. Applied Numerical Mathematics: Transactions of IMACS, 4(2–4):337–349, June 1988. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Lembarki:2002:QLS

[Lem02] A. Lembarki. Quasi-linear sequence transformations and their construction. Applied Numerical Mathematics: Transactions of IMACS, 41(4):481–497, July 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Lenferink:2000:PCA

[Len00]  Wim Lenferink.  Point-

wise convergence of approximations to a convectiondiffusion equation on a Shishkin mesh. Applied Numerical Mathematics: Transactions of IMACS, 32(1): 69–86, January 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/59/28/
31/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/59/28/31/article.
pdf.

Leontyev:2010:DTI

[Leo10a] V. A. Leontyev. Direct time integration algorithm with controllable numerical dissipation for structural dynamics: Two-step Lambda method. Applied Numerical Mathematics: Transactions of IMACS, 60(3): 277–292, March 2010. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Leopold:2010:ABG

[Leo10b] Elie Leopold. Asymptotic behaviours and general recurrence relations. Applied Numerical Mathematics: Transactions of IMACS, 60(12):1352–1363, December 2010. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Levrie:1991:CFC

[Lev91a] Paul Levrie. G-continued fractions and convergence acceleration in the solution of third-order linear recurrence relations of Poincaré type. Applied Numerical Mathematics: Transactions of IMACS, 8(3):225–242, October 1991. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Levrie:1991:CAF

[Lev91b] Paul Levrie. Convergence acceleration for n-fractions. Applied Numerical Mathematics: Transactions of IMACS, 7(6):481–492, June 1991. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Li:2000:CNO

[LFB00] Bingkun Li, Graeme Fairweather, and Bernard Bialecki. A Crank–Nicolson orthogonal spline collocation method for vibration problems. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4):299–306, May 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/61/27/
57/abstract.html;   http:
//www.elsevier.nl/gej-

                                        ng/29/17/21/61/27/57/article.
                                        pdf.

Lima:2014:CMM

[LFL14] Pedro M. Lima, Neville J. Ford, and Patricia M. Lumb. Computational methods for a mathematical model of propagation of nerve impulses in myelinated axons. Applied Numerical Mathematics: Transactions of IMACS, 85(??):38– 53, November 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414001056.

Lopez-Fernandez:2004:NIL

[LFP04] M. López-Fernández and C. Palencia. On the numerical inversion of the Laplace transform of certain holomorphic mappings. Applied Numerical Mathematics: Transactions of IMACS, 51(2–3):289–303, November 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Li:2021:RPC

[LFQH21] Wei Li, Jilin Fang, Yi Qin, and Pengzhan Huang. Rotational pressure-correction method for the Stokes/ Darcy model based on the modular grad–div stabilization. Applied Numerical Mathematics: Transactions of IMACS, 160(??):

451–465, February 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303287.

Lopez-Fernandez:2015:GCQ

[LFS15] Maria Lopez-Fernandez and Stefan Sauter. Generalized convolution quadrature with variable time stepping. Part II: Algorithm and numerical results. Applied Numerical Mathematics: Transactions of IMACS, 94(??):88– 105, August 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415000513.

Liu:2021:DPS

[LFS21] Kai Liu, Ting Fu, and Wei Shi. A dissipation-preserving scheme for damped oscillatory Hamiltonian systems based on splitting. Applied Numerical Mathematics: Transactions of IMACS, 170(??):242–254, December 2021. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421002282.

Lu:2015:EMC

[LfX15] Xin Lu and Shu fang Xu. An eigenspace method for computing derivatives of semisimple eigenvalues and corresponding eigenvectors of quadratic eigenvalue problems. Applied Numerical Mathematics: Transactions of IMACS, 94(??):201– 208, August 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415000604.

Lepora:1987:ASS

[LG87] Paolo Lepora and Bruno Gabutti. An algorithm for the summation of series. Applied Numerical Mathematics: Transactions of IMACS, 3(6):523–528, November 1987. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ly:2002:OCM

[LG02] Le Ngoc Ly and Roland W. Garwood, Jr. An ocean circulation model with surface wave parameterization. Applied Numerical Mathematics: Transactions of IMACS, 40(1–2):351–366, January 2002. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/85/27/54/abstract.
html.

Liu:2019:VGR

[LG19] Yong Liu and Chuan-Qing Gu. Variant of greedy randomized Kaczmarz for ridge regression. Applied Numerical Mathematics: Transactions of IMACS, 143(??): 223–246, September 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419300959.

Lafuente:2021:LSB

[LG21] A. Castejón Lafuente and J. R. Illán González. A least squares-based approach to evaluating strongly singular integrals. Applied Numerical Mathematics: Transactions of IMACS, 162(??): 225–234, April 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303950.

Li:2025:CAW

[LGC25a] Chenxing Li, Fuzheng Gao, and Jintao Cui. Convergence analysis of weak Galerkin finite element variable-time-step BDF2 implicit scheme for parabolic equations. Applied Numerical Mathematics: Transactions of IMACS, 212(??):333–343, June 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425000455.

Li:2025:EAI

[LGC25b] Wenjuan Li, Fuzheng Gao, and Jintao Cui. Error analysis of implicit– explicit weak Galerkin finite element method for time-dependent nonlinear convection–diffusion problem. Applied Numerical Mathematics: Transactions of IMACS, 209(??): 232–241, March 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424003246.

Liu:2011:ASG

[LGH11] Meilin Liu, Zhen Gao, and Jan S. Hesthaven. Adaptive sparse grid algorithms with applications to electromagnetic scattering under uncertainty. Applied Numerical Mathematics: Transactions of IMACS, 61(1):24– 37, January 2011. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Li:2021:PNS

[LGS21] Jian Li, Jiawei Gao, and Yu Shu. A parallel, nonspatial iterative, and rotational pressure projection method for the nonlinear fluid-fluid interaction. Applied Numerical Mathematics: Transactions of IMACS, 165(??):119–136, July 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000441.

Li:2025:LDG

[LGSW25] Jia Li, Wei Guan, Shengzhu Shi, and Boying Wu. A local discontinuous Galerkin methods with local Lax– Friedrichs flux and modified central flux for one dimensional nonlinear convectiondiffusion equation. Applied Numerical Mathematics: Transactions of IMACS, 209(??):124–139, March 2025. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424003167.

Liu:2024:ASO

[LGZH24] Mengmeng Liu, Tao Guo, Mahmoud A. Zaky, and Ahmed S. Hendy. An accurate second-order ADI scheme for three-dimensional tempered evolution problems arising in heat conduction with memory. Applied Numerical Mathematics: Transactions of IMACS, 204(??):111–129, October 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001466.

Li:2002:GSS

[LH02] Z. C. Li and H. T. Huang. Global superconvergence of simplified hybrid combinations of the Ritz– Galerkin and FEMs for elliptic equations with singularities II. Lagrange elements and Adini’s elements. Applied Numerical Mathematics: Transactions of IMACS, 43(3): 253–273, November 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Li:2009:ECN

[LH09] Zi-Cai Li and Hung-Tsai Huang. Effective condition number for the finite element method using local mesh refinements. Applied Numerical Mathematics: Transactions of IMACS, 59(8): 1779–1795, August 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Lahivaara:2011:NUB

[LH11] T. Lähivaara and T. Huttunen. A non-uniform basis order for the discontinuous Galerkin method of the acoustic and elastic wave equations. Applied Numerical Mathematics: Transactions of IMACS, 61(4):473–486, April 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Li:2020:EES

[LH20] Yi Li and Yanren Hou. Error estimates of a second-order decoupled scheme for the evolutionary Stokes–Darcy system. Applied Numerical Mathematics: Transactions of IMACS, 154(??): 129–148, August 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300921.

Li:2021:STD

[LH21] Hu Li and Jin Huang. Solution of two-dimensional elasticity problems using a highaccuracy boundary element method. Applied Numerical Mathematics: Transactions of IMACS, 161(??): 52–68, March 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303494.

Li:2023:TOT

[LH23] Wei Li and Pengzhan Huang.

On a two-order temporal scheme for Navier– Stokes/Navier–Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 194(??):1– 17, December 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423002222.

Li:2024:ASH

[LH24] Xin Li and Xiuling Hu. Algebraically stable high-order multi-physical property-preserving methods for the regularized long-wave equation. Applied Numerical Mathematics: Transactions of IMACS, 203(??):144–159, September 2024. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424001363.

Lai:2009:VAP

[LHC09] Junjiang Lai, Jianguo Huang, and Chuanmiao Chen. Vibration analysis of plane elasticity problems by the C^{0} continuous time stepping finite element method. Applied Numerical Mathematics: Transactions of IMACS, 59(5):905–919, May 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Li:2023:NAH

[LHC23] Hu Li, Jin Huang, and Chong Chen. Numerical algorithm with high accuracy for the modified Helmholtz equation with Robin boundary value problem. Applied Numerical Mathematics: Transactions of IMACS, 187(??):107–119, May 2023.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423000387.

Lai:1996:GSS

[LHH96] Yu-Ling Lai, Apostolos Hadjidimos, and Elias N. Houstis. A generalized Schwarz splitting method based on Hermite collocation for elliptic boundary value problems. Applied Numerical Mathematics: Transactions of IMACS, 21(3): 265–290, August 20, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=21&
issue=3&aid=696.

Li:2008:ECN

[LHH08] Zi Cai Li, Hung-Tsai Huang, and Jin Huang. Effective condition number of the Hermite finite element methods for biharmonic equations. Applied Numerical Mathematics: Transactions of IMACS, 58(9): 1291–1308, September 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Lai:1994:GIH

[LHHR94] Yu-Ling Lai, Apostolos Hadjidimos, Elias N. Houstis, and John R. Rice. General interior Hermite collocation methods for second-order elliptic partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 16(1–2): 183–200, December 1, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=16&
issue=1-2&aid=531.  A

Festschrift to honor Professor Robert Vichnevetsky on his 65th birthday.

Le:2024:FDS

[LHN24] Anh Ha Le, Toan T. Huynh, and Quan M. Nguyen. A finite difference scheme for (2 + 1)D cubic-quintic nonlinear Schrödinger equations with nonlinear damping. Applied Numerical Mathematics: Transactions of IMACS, 205(??): 215–239, November 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001867.

Liu:2013:OMC

[LHÖ13] Yaning Liu, M. Yousuff Hussaini, and Giray Ökten. Optimization of a Monte Carlo variance reduction method based on sensitivity derivatives. Applied Numerical Mathematics: Transactions of IMACS, 72(??):160– 171, October 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413000871.

Lepsky:2000:ADG

[LHS00] Olga Lepsky, Changqing Hu, and Chi-Wang Shu. Analysis of the discontinuous Galerkin method for Hamilton–Jacobi equations. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4): 423–434, May 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/61/27/
70/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/61/27/70/article.
pdf.

Long:2020:PRG

[LHT20] Haie Long, Bo Han, and Shanshan Tong. A proximal regularized Gauss–Newton– Kaczmarz method and its acceleration for nonlinear ill-posed problems. Applied Numerical Mathematics: Transactions of IMACS, 151(??):301–321, May 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420300039.

Li:2017:GFE

[LHW17] Meng Li, Chengming Huang, and Nan Wang. Galerkin finite element method for the nonlinear fractional Ginzburg–Landau equation. Applied Numerical Mathematics: Transactions of IMACS, 118(??):131–149, August 2017. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417300636.

Li:2008:SSD

[LHWF08] Zi-Cai Li, Hsin-Yun Hu, Song Wang, and Qing Fang. Superconvergence of solution derivatives of the Shortley– Weller difference approximation to Poisson’s equation with singularities on polygonal domains. Applied Numerical Mathematics: Transactions of IMACS, 58(5):689–704, May 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Liu:2020:NAO

[LHX20] Wenjia Liu, Yanren Hou, and Dandan Xue. Numerical analysis of a 4th-order time parallel algorithm for the time-dependent Navier– Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 150(??):361–383, April 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419302855.

Li:2024:TGC

[LHY24] Ao Li, Jian Huang, and Nianyu Yi. A two-grid characteristic block-centered finite difference method for Darcy–Forchheimer slightly compressible miscible displacement problem. Applied Numerical Mathematics: Transactions of IMACS, 202(??):189–208, August 2024. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001156.

Li:1998:IIM

[Li98] Zhilin Li. The immersed interface method using a finite element formulation. Applied Numerical Mathematics: Transactions of IMACS, 27(3):253–267, July 1, 1998. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1998/27/
3/881.pdf.

Li:2000:MMF

[Li00a] Jichun Li. Multiblock mixed finite element methods for singularly perturbed problems. Applied Numerical Mathematics: Transactions of IMACS, 35(2):157– 175, October 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/64/30/
30/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/64/30/30/article.
pdf.

Li:2000:PRM

[Li00b] Zhiping Li. A periodic relaxation method for computing microstructures. Applied Numerical Mathematics: Transactions of IMACS, 32(3):291–303, March 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/59/30/
30/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/59/30/30/article.
pdf.

Li:2001:CSA

[Li01a]  Jichun Li.  Convergence

and superconvergence analysis of finite element methods on highly nonuniform anisotropic meshes for singularly perturbed reactiondiffusion problems. Applied Numerical Mathematics: Transactions of IMACS, 36(2–3):129–154, February 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.nl../27/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/65/31/27/article.
pdf.

Li:2001:CNL

[Li01b] Zhiping Li. Computations of needle-like microstructures. Applied Numerical Mathematics: Transactions of IMACS, 39(1):1–15, October 2001. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/79/27/27/abstract.
html.

Li:2005:BTG

[Li05] Shoufu Li. B-theory of general linear methods for Volterra functional differential equations. Applied Numerical Mathematics: Transactions of IMACS, 53(1):57–72, April 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Li:2008:TMH

[Li08] Z. C. Li. The Treffetz method for the Helmholtz equation with degeneracy. Applied Numerical Mathematics: Transactions of IMACS, 58(2):131–159, February 2008. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Li:2011:MGA

[Li11]  Xiaolin Li.  Meshless

Galerkin algorithms for boundary integral equations with moving least square approximations. Applied Numerical Mathematics: Transactions of IMACS, 61(12):1237–1256, December 2011. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927411001346.

Li:2012:NCC

[Li12] Hengguang Li. A note on the conditioning of a class of generalized finite element methods. Applied Numerical Mathematics: Transactions of IMACS, 62(6): 754–766, June 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927411000870.

Li:2016:EEM

[Li16]  Xiaolin Li.  Error esti-

mates for the moving leastsquare approximation and the element-free Galerkin method in n-dimensional spaces. Applied Numerical Mathematics: Transactions of IMACS, 99(??):77– 97, January 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415001294.

Li:2019:SST

[Li19] Jiyong Li. Symplectic and symmetric trigonometricallyfitted ARKN methods. Applied Numerical Mathematics: Transactions of IMACS, 135(??):381–395, January 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418302095.

Li:2022:EPE

[Li22] Jiyong Li. Energy-preserving exponential integrator Fourier pseudo-spectral schemes for the nonlinear Dirac equation. Applied Numerical Mathematics: Transactions of IMACS, 172(??):1– 26, February 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421002609.

Li:2023:UAN

[Li23] Jiyong Li. Uniformly accurate nested Picard iterative schemes for nonlinear Schrödinger equation with highly oscillatory potential. Applied Numerical Mathematics: Transactions of IMACS, 192(??): 132–151, October 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423001708.

Li:2025:NCS

[Li25] Zhen Li. Nonstaggered central scheme with steady-state discretization for solving the open channel flows via the flux globalization. Applied Numerical Mathematics: Transactions of IMACS, 207(??):58–85, January 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424002277.

Liang:2022:DGA

[Lia22] Hui Liang. Discontinuous Galerkin approximations to second-kind Volterra integral equations with weakly singular kernel. Applied Numerical Mathematics: Transactions of IMACS, 179(??): 170–182, September 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742200112X.

Lie:2001:WPT

[Lie01] Ivar Lie. Well-posed transparent boundary conditions for the shallow water equations. Applied Numerical Mathematics: Transactions of IMACS, 38(4):445–474, September 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/77/34/30/abstract.
html.

Linss:2001:SCU

[Lin01]  Torsten Linß.  Sufficient

conditions for uniform convergence on layer-adapted grids. Applied Numerical Mathematics: Transactions of IMACS, 37(1–2): 241–255, April 2001. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/73/26/
40/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/73/26/40/article.
pdf.

Lindner:2010:FSM

[Lin10] Marko Lindner. The finite section method and stable subsequences. Applied Numerical Mathematics: Transactions of IMACS, 60(4):501–512, April 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Loh:2018:NPC

[LIPT18] Jian Rong Loh, Abdulnasir Isah, Chang Phang, and Yoke Teng Toh. On the new properties of Caputo– Fabrizio operator and its application in deriving shifted Legendre operational matrix. Applied Numerical Mathematics: Transactions of IMACS, 132(??): 138–153, October 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741830134X.

Liu:1997:NIP

[Liu97] Yunkang Liu. Numerical investigation of the pantograph equation. Applied Numerical Mathematics: Transactions of IMACS, 24(2–3):309–317, June 19, 1997. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=24&
issue=2-3&aid=784. Volterra

centennial (Tempe, AZ, 1996).

Liu:2002:CGR

[Liu02]  Zhongyun Liu.  Conver-

gence of generalized relaxed multisplitting methods for symmetric positive definite matrices. Applied Numerical Mathematics: Transactions of IMACS, 42(4):

513–527, September 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Liu:2009:EAF

[Liu09] Biyue Liu. An error analysis of a finite element method for a system of nonlinear advection–diffusion–reaction equations. Applied Numerical Mathematics: Transactions of IMACS, 59(8): 1947–1959, August 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Liu:2021:FCP

[Liu21] Hai-Feng Liu. Frame completion with prescribed norms via alternating projection method. Applied Numerical Mathematics: Transactions of IMACS, 164(??): 161–174, ???? 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303330.

Liu:2024:AGR

[Liu24]  Yong Liu.  Accelerated

greedy randomized augmented Kaczmarz algorithm for inconsistent linear systems. Applied Numerical Mathematics: Transactions of IMACS, 195(??): 142–156, January 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423002568.

Luo:2020:ROEa

[LJ20a] Zhendong Luo and Wenrui Jiang. A reducedorder extrapolated Crank– Nicolson finite spectral element method for the 2D nonstationary Navier–Stokes equations about vorticitystream functions. Applied Numerical Mathematics: Transactions of IMACS, 147(??):161–173, January 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419302090.

Luo:2020:ROEc

[LJ20b] Zhendong Luo and Wenrui Jiang. A reducedorder extrapolated technique about the unknown coefficient vectors of solutions in the finite element method for hyperbolic type equation. Applied Numerical Mathematics: Transactions of IMACS, 158(??): 123–133, December 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420302257.

Liu:2020:RTL

[LJYS20] Qiaohua Liu, Shufang Jin, Lei Yao, and Dongmei Shen. The revisited total least squares problems with linear equality constraint. Applied Numerical Mathematics: Transactions of IMACS, 152(??):275–284, June 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419303307.

Liu:2007:AMG

[LK07] T. G. Liu and B. C. Khoo.

The accuracy of the modified ghost fluid method for gas– gas Riemann problem. Applied Numerical Mathematics: Transactions of IMACS, 57(5–7):721–733, May 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Li:2014:USH

[LK14] Yibao Li and Junseok Kim.

An unconditionally stable hybrid method for image segmentation. Applied Numerical Mathematics: Transactions of IMACS, 82(??): 32–43, August 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414000518.

Lim:2007:NMV

[LKJ07] Hyeona Lim, Seongjai Kim, and Jim Douglas Jr. Numerical methods for viscous and nonviscous wave equations. Applied Numerical Mathematics: Transactions of IMACS, 57(2):194– 212, February 2007. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Lin:2020:BIA

[LKJ20] Aihua Lin, Anastasiia Kuzmina, and Per Kristen Jakobsen. A boundary integral approach to linear and nonlinear transient wave scattering. Applied Numerical Mathematics: Transactions of IMACS, 147(??): 277–300, January 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419302351.

Lastdrager:2001:SGC

[LKV01] Boris Lastdrager, Barry Koren, and Jan Verwer. The sparse-grid combination technique applied to timedependent advection problems. Applied Numerical Mathematics: Transactions of IMACS, 38(4):377–401, September 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/77/34/27/abstract.
html.

Lee:1998:EES

[LL98] Hyun Young Lee and Jang Rae Lee. Error estimates for a single-phase quasilinear Stefan problem in one space dimension. Applied Numerical Mathematics: Transactions of IMACS, 26(3):327– 342, March 2, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1998/26/
3/856.pdf.

Lu:2002:SES

[LL02] Tao Lü and Jing Lu. Splitting extrapolation for solving second order elliptic systems with curved boundary in R^{d} by using d-quadratic isoparametric finite element. Applied Numerical Mathematics: Transactions of IMACS, 40(4):467–481, March 2002. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/85/34/29/abstract.
html.

Liu:2006:RPC

[LL06] M. B. Liu and G. R. Liu.

Restoring particle consistency in smoothed particle hydrodynamics. Applied Numerical Mathematics: Transactions of IMACS, 56(1):19–36, January 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Lee:2015:SIM

[LL15] Jaewook Lee and Younhee Lee. Stability of an implicit method to evaluate option prices under local volatility with jumps. Applied Numerical Mathematics: Transactions of IMACS, 87(??):20–30, January 2015. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414001251.

Li:2019:SOS

[LL19] Yuan Li and Xuelan Luo. Second-order semi-implicit Crank–Nicolson scheme for a coupled magnetohydrodynamics system. Applied Numerical Mathematics: Transactions of IMACS, 145(??):48–68, November 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        https://www.sciencedirect.
                                        com/science/article/pii/
                                        S0168927419301539.

Li:2020:CNF

[LL20a] Hengguang Li and Xun Lu. Condition numbers of finite element methods on a class of anisotropic meshes. Applied Numerical Mathematics: Transactions of IMACS, 158(??):22– 43, December 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742030218X.

Liu:2020:TFE

[LL20b] Zhengguang Liu and Xiaoli Li. Two fast and efficient linear semi-implicit approaches with unconditional energy stability for nonlocal phase field crystal equation. Applied Numerical Mathematics: Transactions of IMACS, 150(??):491–506, April 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741930306X.

Liu:2021:HON

[LL21] Haiyu Liu and Shujuan Lü. A high-order numerical scheme for solving nonlinear time fractional reaction–diffusion equations with initial singularity. Applied Numerical Mathematics: Transactions of IMACS, 169(??):32–43, November 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0168927421001811.

Li:2023:EEN

[LL23] Xiaolin Li and Shuling Li.

Effect of an efficient numerical integration technique on the element-free Galerkin method. Applied Numerical Mathematics: Transactions of IMACS, 193(??):

204–225, November 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742300209X.

Li:2024:EEF

[LL24] Minghao Li and Zhenzhen Li. Error estimates for the finite element method of the Navier–Stokes–Poisson– Nernst–Planck equations. Applied Numerical Mathematics: Transactions of IMACS, 197(??):186–209, March 2024. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423002908.

Liu:2018:SEB

[LLD18] Linna Liu, Mengling Li, and Feiqi Deng. Stability equivalence between the neutral delayed stochastic differential equations and the Euler–Maruyama numerical scheme. Applied Numerical Mathematics: Transactions of IMACS, 127(??): 370–386, May 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418300254.

Lin:2025:BCB

[lLH25] Xue lei Lin and Sean Hon. A block α-circulant based preconditioned MIN-RES method for wave equations. Applied Numerical Mathematics: Transactions of IMACS, 209(??): 258–274, March 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424002964.

Li:2017:NPL

[LLHC17] Zi-Cai Li, Ming-Gong Lee, Hung-Tsai Huang, and John Y. Chiang. Neumann problems of 2D Laplace’s equation by method of fundamental solutions. Applied Numerical Mathematics: Transactions of IMACS, 119(??):126–145, September 2017. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417301009.

Li:2018:SFV

[LLHC18] Rui Li, Jian Li, Xiaoming He, and Zhangxin Chen. A stabilized finite volume element method for a coupled Stokes–Darcy problem. Applied Numerical Mathematics: Transactions of IMACS, 133(??):2–24, ???? 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417302076.

Li:2020:DGR

[LLJY20] Shan Li, Zhenyan Lai, Lusha Jin, and Xuhong Yu. Diagonalized Gegenbauer rational spectral methods for secondand fourth-order problems on the whole line. Applied Numerical Mathematics: Transactions of IMACS, 151(??):494–516, May 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419303356.

Lee:2021:FHO

[LLKJ21] Seyeon Lee, Junseo Lee, Hyunju Kim, and Bongsoo Jang. A fast and high-order numerical method for nonlinear fractional-order differential equations with non-singular kernel. Applied Numerical Mathematics: Transactions of IMACS, 163(??):57–76, May 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421000210.

Leriche:2008:SEC

[LLL08] E. Leriche, P. Lallemand, and G. Labrosse. Stokes eigenmodes in cubic domain: primitive variable and Lattice Boltzmann formulations. Applied Numerical Mathematics: Transactions of IMACS, 58(7):

935–945, July 2008. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Liu:2012:MTP

[LLL12] Hongwei Liu, Xinze Liu, and Changhe Liu. Mehrotratype predictor-corrector algorithms for sufficient linear complementarity problem. Applied Numerical Mathematics: Transactions of IMACS, 62(12): 1685–1700, December 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412001146.

Liu:2019:SBS

[LLM19] Bin Liu, Lu Li, and Boris A.

Malomed. Stability band structure for periodic states in periodic potentials. Applied Numerical Mathematics: Transactions of IMACS, 141(??):44–53, ???? 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741830117X.

Lin:2025:ADF

[LLN25] Lei Lin, Junliang Lv, and Tian Niu. An adaptive DtN–FEM for the scattering problem from orthotropic media. Applied Numerical Mathematics: Transactions of IMACS, 209(??):

140–154, March 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424003210.

Lin:2021:PMT

[lLNW21] Xue lei Lin, Micheal K. Ng, and Andy Wathen. Preconditioners for multilevel Toeplitz linear systems from steady-state and evolutionary advection–diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 161(??): 469–488, March 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303792.

Liu:1996:OOP

[LLS^{+}96] Jinn-Liang Liu, Ing-Jer Lin, Miin-Zhih Shih, Ren-Chuen Chen, and Mao-Chung Hsieh. Object-oriented programming of adaptive finite element and finite volume methods. Applied Numerical Mathematics: Transactions of IMACS, 21(4):439– 467, November 22, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=21&
issue=4&aid=726.

Liang:2007:NSQ

[LLT07] K. W. Liang, P. Lin, and R. C. E. Tan. Numerical solution of quenching problems using mesh-dependent variable temporal steps. Applied Numerical Mathematics: Transactions of IMACS, 57(5–7):791–800, May 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Li:2020:IMP

[LLT20a] Yung-Ta Li, Po-Yu Lin, and Chun-Hao Teng. Inverse matrices for pseudospectral differentiation operators in polar coordinates by stepwise integrations and low-rank updates. Applied Numerical Mathematics: Transactions of IMACS, 150(??): 519–535, April 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419302910.

Lin:2020:MPI

[LLT20b] Po-Yu Lin, Yung-Ta Li, and Chun-Hao Teng. Multidomain pseudospectral integration preconditioning matrices for the advection and the diffusion operators. Applied Numerical Mathematics: Transactions of IMACS, 157(??): 505–521, November 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742030204X.

Li:2018:CRS

[LLV18] Wen Li, Dongdong Liu, and Seak-Weng Vong. Comparison results for splitting iterations for solving multi-linear systems. Applied Numerical Mathematics: Transactions of IMACS, 134(??): 105–121, December 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418301570.

Li:2020:MPU

[LLVX20] Wen Li, Dongdong Liu, Seak-Weng Vong, and Mingqing Xiao. Multilinear PageRank: Uniqueness, error bound and perturbation analysis. Applied Numerical Mathematics: Transactions of IMACS, 156(??): 584–607, October 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420301665.

Lyu:2020:FLF

[LLW20] Pin Lyu, Yuxiang Liang, and Zhibo Wang. A fast linearized finite difference method for the nonlinear multi-term timefractional wave equation. Applied Numerical Mathematics: Transactions of IMACS, 151(??):448–471, May 2020. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419303204.

Liu:2022:CNM

[LLW22] Qiaohua Liu, Chuge Li, and Yimin Wei. Condition numbers of multidimensional mixed least squares-total least squares problems. Applied Numerical Mathematics: Transactions of IMACS, 178(??):52–68, August 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422000769.

Luo:2022:RCM

[lLX22] Xin long Luo and Hang Xiao. The regularization continuation method with an adaptive time step control for linearly constrained optimization problems. Applied Numerical Mathematics: Transactions of IMACS, 181(??): 255–276, November 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001660.

Luo:2021:EPT

[lLXhLZ21] Xin long Luo, Hang Xiao, Jia hui Lv, and Sen Zhang. Explicit pseudo-transient continuation and the trustregion updating strategy for unconstrained optimization. Applied Numerical Mathematics: Transactions of IMACS, 165(??): 290–302, July 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000696.

Li:2011:SGT

[LLY11] M. Li, L.-Z. Liao, and X. M. Yuan. Some Goldstein’s type methods for cocoercive variant variational inequalities. Applied Numerical Mathematics: Transactions of IMACS, 61(2):216– 228, February 2011. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Li:2021:NAF

[LLY21] Jian Li, Qian Liu, and Jing Yue. Numerical analysis of fully discrete finite element methods for the stochastic Navier–Stokes equations with multiplicative noise. Applied Numerical Mathematics: Transactions of IMACS, 170(??): 398–417, December 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421002099.

Liao:2024:ATK

[LLY24] Yimou Liao, Wen Li, and Dan Yang. The accelerated tensor Kaczmarz algorithm with adaptive parameters for solving tensor systems. Applied Numerical Mathematics: Transactions of IMACS, 202(??): 100–119, August 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001053.

Liu:2019:ADI

[LLZ19] Zeting Liu, Fawang Liu, and Fanhai Zeng. An alternating direction implicit spectral method for solving two dimensional multiterm time fractional mixed diffusion and diffusion-wave equations. Applied Numerical Mathematics: Transactions of IMACS, 136(??): 139–151, February 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418302307.

Liu:2022:ASR

[LLZ^{+}22] Hongliang Liu, Haodong Li, Yameng Zhang, Shoufu Li, and Shuiping Yang. Asymptotic stability of Runge– Kutta method for solving nonlinear functional differential-algebraic equations. Applied Numerical Mathematics: Transactions of IMACS, 181(??): 277–292, November 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001659.

Li:2024:PAT

[LLZL24] Yuejie Li, Huanrong Li, Yihui Zeng, and Zhendong Luo. A preserving accuracy twogrid reduced-dimensional Crank–Nicolson mixed finite element method for nonlinear wave equation. Applied Numerical Mathematics: Transactions of IMACS, 202(??):1–20, August 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000953.

Laminie:2000:TDC

[LM00] J. Laminie and S. M. Mefire. Three-dimensional computation of a magnetic field by mixed finite elements and boundary elements. Applied Numerical Mathematics: Transactions of IMACS, 35(3):221–244, November 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/64/31/

27/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/64/31/27/article.
pdf.

Liao:2021:EEL

[LM21] Huiqing Liao and Heping Ma. Error estimate of a Legendre–Galerkin Chebyshev collocation method for a class of parabolic inverse problem. Applied Numerical Mathematics: Transactions of IMACS, 170(??): 179–189, December 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421002142.

Liu:2022:CBV

[LM22a] Ling Liu and Junjie Ma.

Collocation boundary value methods for auto-convolution Volterra integral equations. Applied Numerical Mathematics: Transactions of IMACS, 177(??):1–17, July 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0168927422000666.

Lopez:2022:NEL

[LM22b] L. Lopez and S. Maset.

Numerical event location techniques in discontinuous differential algebraic equations. Applied Numerical Mathematics: Transactions of IMACS, 178(??):

98–122, August 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422000733.

Lardas:2025:GFH

[LM25] Emmanouil Lardas and Marilena Mitrouli. On the growth factor of Hadamard matrices of order 20. Applied Numerical Mathematics: Transactions of IMACS, 208 (part A)(??):310–316, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424000199.

Li:2018:DSI

[LMA18] Yuan Li, Yanjie Ma, and Rong An. Decoupled, semiimplicit scheme for a coupled system arising in magnetohydrodynamics problem. Applied Numerical Mathematics: Transactions of IMACS, 127(??):142–163, May 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892741830014X.

Lopez:2002:VSV

[LMG02] David J. López, Pablo Martín, and Amelia García. A variable-stepsize variableorder multistep method for the integration of perturbed linear problems. Applied Numerical Mathematics: Transactions of IMACS, 42(1–3):285–295, August 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Laudadio:2024:CIE

[LMO24] Teresa Laudadio, Nicola Mastronardi, and Donatella Occorsio. Computing integrals with an exponential weight on the real axis in floating point arithmetic. Applied Numerical Mathematics: Transactions of IMACS, 200(??): 309–317, June 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001551.

Lopes:1999:LCQ

[LMP99] Vera Lúcia Rocha Lopes, José Mario Martínez, and Rosana Pérez. On the local convergence of quasi-Newton methods for nonlinear complementarity problems. Applied Numerical Mathematics: Transactions of IMACS, 30(1):3–22, May 10, 1999. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1999/30/
1/945.pdf.

Lattanzio:2019:KSA

[LMPS19] Corrado Lattanzio, Corrado Mascia, Ramón G. Plaza, and Chiara Simeoni. Kinetic schemes for assessing stability of traveling fronts for the Allen–Cahn equation with relaxation. Applied Numerical Mathematics: Transactions of IMACS, 141(??):234–247, ???? 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418302368.

Li:2018:SMT

[LMQZ18] Jingliang Li, Heping Ma, Yonghui Qin, and Shuaiyin Zhang. A spectral method for triangular prism. Applied Numerical Mathematics: Transactions of IMACS, 129(??):26–38, July 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418300540.

Lopez:2008:DAG

[LMS08] J. I. H. Lopez, J. R. Meneghini, and F. Saltara. Discrete approximation to the global spectrum of the tangent operator for flow past a circular cylinder. Applied Numerical Mathematics: Transactions of IMACS, 58(8): 1159–1167, August 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Lamnii:2009:QSS

[LMS09] A. Lamnii, H. Mraoui, and D. Sbibih. Quadratic spherical spline quasi-interpolants on Powell–Sabin partitions. Applied Numerical Mathematics: Transactions of IMACS, 59(5):1094–1109, May 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Lima:2017:NCA

[LMSW17] P. M. Lima, M. L. Morgado, M. Schöbinger, and E. B. Weinmüller. A novel computational approach to singular free boundary problems in ordinary differential equations. Applied Numerical Mathematics: Transactions of IMACS, 114(??): 97–107, April 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301933.

Liu:2020:TEM

[LMTW20] Wei Liu, Xuerong Mao, Jingwen Tang, and Yue Wu. Truncated Euler– Maruyama method for classical and time-changed nonautonomous stochastic differential equations. Applied Numerical Mathematics: Transactions of IMACS, 153(??):66–81, July 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420300428.

Laudadio:2017:NIC

[LMV17] Teresa Laudadio, Nicola Mastronardi, and Paul Van Dooren. Numerical issues in computing the antitriangular factorization of symmetric indefinite matrices. Applied Numerical Mathematics: Transactions of IMACS, 116(??):204–214, June 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892741630188X.

Laudadio:2025:FRA

[LMV25] Teresa Laudadio, Nicola Mastronardi, and Paul Van Dooren. Fast and reliable algorithms for computing the zeros of Althammer polynomials. Applied Numerical Mathematics: Transactions of IMACS, 207(??): 210–221, January 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424002356.

Liu:2023:BEM

[LMW23] Wei Liu, Xuerong Mao, and Yue Wu. The backward Euler–Maruyama method for invariant measures of stochastic differential equations with super-linear coefficients. Applied Numerical Mathematics: Transactions of IMACS, 184(??): 137–150, February 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422002574.

Lukacova-Medvidova:2007:FVE

[LMWZ07] M. Lukáčová-Medvid’ová, G. Warnecke, and Y. Zahaykah. Finite volume evolution Galerkin (FVEG) methods for three-dimensional wave equation system. Applied Numerical Mathematics: Transactions of IMACS, 57(9):1050–1064, September 2007. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Li:2010:OPE

[LMWZ10] Jingzhi Li, Jens Markus Melenk, Barbara Wohlmuth, and Jun Zou. Optimal a priori estimates for higher order finite elements for elliptic interface problems. Applied Numerical Mathematics: Transactions of IMACS, 60(1–2):19–37, January/February 2010. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Li:2018:SOU

[LMY18] Qi Li, Liquan Mei, and Bo You. A secondorder, uniquely solvable, energy stable BDF numerical scheme for the phase field crystal model. Applied Numerical Mathematics: Transactions of IMACS, 134(??):46–65, December 2018. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892741830151X.

Louter-Nool:1992:BCP

[LN92] Margreet Louter-Nool. Block-Cholesky for parallel processing. Applied Numerical Mathematics: Transactions of IMACS, 10(1): 37–57, June 1992. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Liu:2008:MFM

[LN08] Jijun Liu and Ming Ni. A model function method for determining the regularizing parameter in potential approach for the recovery of scattered wave. Applied Numerical Mathematics: Transactions of IMACS, 58(8):1113–1128, August 2008. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Liu:2021:MNV

[LN21] Xin Liu and Yufeng Nie.

A modified nonconforming virtual element with BDM-like reconstruction for the Navier–Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 167(??): 375–388, September 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001434.

Lin:2024:MSC

[LN24] Xuelei Lin and Michael K.

Ng. A matching Schur complement preconditioning technique for inverse source problems. Applied Numerical Mathematics: Transactions of IMACS, 201(??): 404–418, July 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000795.

Le:2026:CNF

[LN26] Anh Ha Le and Quan M.

Nguyen. A Crank–Nicolson finite difference scheme for coupled nonlinear Schrödinger equations with saturable nonlinearity and nonlinear damping. Applied Numerical Mathematics: Transactions of IMACS, 219(??): 219–238, January 2026. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425001825.

Long:2012:IFM

[LNZ12] Guangqing Long, Gnaneshwar Nelakanti, and Xiaohua Zhang. Iterated fast multiscale Galerkin methods for Fredholm integral equations of second kind with weakly singular kernels. Applied Numerical Mathematics: Transactions of IMACS, 62(3):201–211, March 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741100211X.

Lubich:1995:IET

[LO95] Christian Lubich and Alexander Ostermann. Interior estimates for time discretizations of parabolic equations. Applied Numerical Mathematics: Transactions of IMACS, 18(1–3):241–251, September 27, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=18&
issue=1-3&aid=604. Sev-

enth Conference on the Numerical Treatment of Differential Equations (Halle, 1994).

Lubich:1996:RKT

[LO96] Christian Lubich and Alexander Ostermann. Runge– Kutta time discretization of reaction-diffusion and Navier–Stokes equations: Nonsmooth-data error estimates and applications to long-time behaviour. Applied Numerical Mathematics: Transactions of IMACS, 22(1–3):279–292, December 4, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=22&
issue=1-3&aid=720. Special

issue celebrating the centenary of Runge–Kutta methods.

Lima:2003:NSS

[LO03] P. M. Lima and A. M.

Oliveira. Numerical solution of a singular boundary value problem for a generalized Emden–Fowler equation. Applied Numerical Mathematics: Transactions of IMACS, 45(4): 389–409, June 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Lo:2006:PTD

[Lo06]  Edisanter Lo.  Paramet-

ric target detection algorithm based on ISLIM for hyperspectral imaging. Applied Numerical Mathematics: Transactions of IMACS, 56(3–4):544–558, March/ April 2006. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Labidi:2022:NCF

[LO22] Samira Labidi and Khaled Omrani. A new conservative fourth-order accurate difference scheme for the nonlinear Schrödinger equation with wave operator. Applied Numerical Mathematics: Transactions of IMACS, 173(??):1–12, March 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421003135.

Labidi:2023:NAN

[LO23] Samira Labidi and Khaled Omrani. A new approach for numerical solution of Kuramoto–Tsuzuki equation. Applied Numerical Mathematics: Transactions of IMACS, 184(??): 527–541, February 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422002926.

Logg:2004:MAT

[Log04]  Anders Logg.  Multi-

adaptive time integration.

Applied Numerical Mathematics: Transactions of IMACS, 48(3–4):339–354, March 2004. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Lube:1998:NOD

[LOM98] G. Lube, F. C. Otto, and H. Müller. A nonoverlapping domain decomposition method for parabolic initial-boundary value problems. Applied Numerical Mathematics: Transactions of IMACS, 28(2–4):359–369, October 1, 1998. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1998/28/
2-4/919.pdf.

Longman:1986:SS

[Lon86] I. M. Longman. The summation of series. Applied Numerical Mathematics: Transactions of IMACS, 2(2):135–141, April 1986. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Longman:1988:NSP

[Lon88] I. M. Longman. A note on the summation of power series. Applied Numerical Mathematics: Transactions of IMACS, 4(5):431– 437, July 1988. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Lorentzen:2010:PAC

[Lor10] Lisa Lorentzen. Padé approximation and continued fractions. Applied Numerical Mathematics: Transactions of IMACS, 60(12): 1364–1370, December 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Lee:2003:ANC

[LOS03] Y. Lee, M. Oh, and V. I. Shin. Adaptive nonlinear continuous-discrete filtering. Applied Numerical Mathematics: Transactions of IMACS, 47(1):45– 56, October 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Lotfi:2019:EPM

[Lot19] Ali Lotfi. Epsilon penalty method combined with an extension of the Ritz method for solving a class of fractional optimal control problems with mixed inequality constraints. Applied Numerical Mathematics: Transactions of IMACS, 135(??): 497–509, January 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418302010.

Leca:1993:I

[LP93] Pierre Leca and Serge G. Petiton. Introduction. Applied Numerical Mathematics: Transactions of IMACS, 12(5):373–375, July 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/016892749390098C.

Lopez:1997:NPB

[LP97] L. Lopez and T. Politi. Numerical procedures based on Runge–Kutta methods for solving isospectral flows. Applied Numerical Mathematics: Transactions of IMACS, 25(4):443– 459, November 10, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=25&
issue=4&aid=806.

Lopez:2000:ACA

[LP00] L. Lopez and T. Politi. Applications of the Cayley approach in the numerical solution of matrix differential systems on quadratic groups. Applied Numerical Mathematics: Transactions of IMACS, 36(1):35– 55, November 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.nl../27/
                                        abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/10/28/65/26/27/article.
                                        pdf.

Luce:2001:FVS

[LP01] R. Luce and S. Perez. A finite volume scheme for an elliptic equation with heterogeneous coefficients. Application to a homogenization problem. Applied Numerical Mathematics: Transactions of IMACS, 38(4):427–444, September 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        10/28/77/34/29/abstract.
                                        html.

Langer:2005:DSA

[LP05] U. Langer and D. Pusch. Data-sparse algebraic multigrid methods for large scale boundary element equations. Applied Numerical Mathematics: Transactions of IMACS, 54(3–4):406–424, August 2005. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Latt:2024:NSC

[LP24] Kaido Lätt and Arvet Pedas. Numerical schemes for a class of singular fractional integro-differential equations. Applied Numerical Mathematics: Transactions of IMACS, 200(??):

331–343, June 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423002088.

Li:2025:PFE

[LP25] Jie Li and Lifang Pei. Parametric finite element method for a nonlocal curvature flow. Applied Numerical Mathematics: Transactions of IMACS, 212(??): 197–214, June 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425000273.

Levy:2000:BTV

[LPR00a] Doron Levy, Gabriella Puppo, and Giovanni Russo. On the behavior of the total variation in CWENO methods for conservation laws. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4):407–414, May 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/61/27/
68/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/61/27/68/article.
pdf.

Levy:2000:TOC

[LPR00b] Doron Levy, Gabriella Puppo, and Giovanni Russo. A third order central WENO scheme for 2D conservation laws. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4): 415–421, May 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/61/27/
69/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/61/27/69/article.
pdf.

Latt:2025:SFI

[LPS25] Kaido Lätt, Arvet Pedas, and Hanna Britt Soots. Singular fractional integrodifferential equations with non-constant coefficients. Applied Numerical Mathematics: Transactions of IMACS, 211(??):179–192, May 2025. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425000042.

Laminie:1994:INA

[LPT94] Jacques Laminie, Frédéric Pascal, and Roger Temam. Implementation and numerical analysis of the nonlinear Galerkin methods with finite elements discretization. Applied Numerical Mathematics: Transactions of IMACS, 15(2):219–246, September 13, 1994. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=15&
issue=2&aid=497. Innova-

tive methods in numerical analysis (Bressanone, 1992).

Landi:2016:SCI

[LPT16] G. Landi, E. Loli Piccolomini, and I. Tomba. A stopping criterion for iterative regularization methods. Applied Numerical Mathematics: Transactions of IMACS, 106(??):53–68, August 2016. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416300368.

Lillemae:2024:CPI

[LPV24] Margus Lillemäe, Arvet Pedas, and Mikk Vikerpuur. Central part interpolation schemes for fractional differential equations. Applied Numerical Mathematics: Transactions of IMACS, 200(??):318–330, June 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001770.

Li:2000:SAD

[LPZ00] Shengtai Li, Linda Petzold, and Wenjie Zhu. Sensitivity analysis of differentialalgebraic equations: a comparison of methods on a special problem. Applied Numerical Mathematics: Transactions of IMACS, 32(2):161–174, February 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/59/27/
29/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/59/27/29/article.
pdf.

Lin:2021:IWM

[LQS21] Fu-Rong Lin, Yi-Feng Qiu, and Zi-Hang She. IRK-WSGD methods for space fractional diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 164(??): 222–244, ???? 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303627.

Li:2023:SOU

[LQXK23] Yibao Li, Kang Qin, Qing Xia, and Junseok Kim. A second-order unconditionally stable method for the anisotropic dendritic crystal growth model with an orientation-field. Applied Numerical Mathematics: Transactions of IMACS, 184(??):512–526, February 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422002938.

Liu:2025:DSF

[LQY25] Liya Liu, Xiaolong Qin, and Jen-Chih Yao. A distributed stochastic forwardbackward-forward self-adaptive algorithm for Cartesian stochastic variational inequalities. Applied Numerical Mathematics: Transactions of IMACS, 211(??): 17–41, May 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425000030.

Loh:1987:NPV

[LR87] C. Y. Loh and H. Rasmussen. A numerical procedure for viscous free surface flows. Applied Numerical Mathematics: Transactions of IMACS, 3(6):479– 495, November 1987. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Leimkuhler:1993:RCW

[LR93] B. Leimkuhler and A. Ruehli.

Rapid convergence of waveform relaxation. Applied Numerical Mathematics: Transactions of IMACS, 11(1–3):211–224, January 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Parallel methods for ordinary differential equations (Grado, 1991).

Louchart:2000:SID

[LR00] Olivier Louchart and Anthony Randriamampianina. A spectral iterative domain decomposition technique for the incompressible Navier– Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4):233–240, May 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/61/27/
49/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/61/27/49/article.
pdf.

Lin:2001:SMA

[LR01] Tao Lin and D. L. Russell. A superconvergent method for approximating the bending moment of elastic beams with hysteresis damping. Applied Numerical Mathematics: Transactions of IMACS, 38(1–2):145–165, July 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/77/27/
35/abstract.html;   http:
//www.elsevier.nl/gej-

ng/10/10/28/77/27/35/article.
pdf.

Liu:2003:EMB

[LR03] I-Shih Liu and M. A. Rincon. Effect of moving boundaries on the vibrating elastic string. Applied Numerical Mathematics: Transactions of IMACS, 47(2): 159–172, November 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Li:2018:BCF

[LR18a] Xiaoli Li and Hongxing Rui. A block-centered finite difference method for the distributed-order time-fractional diffusionwave equation. Applied Numerical Mathematics: Transactions of IMACS, 131(??): 123–139, September 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418301107.

Li:2018:HOF

[LR18b] Xiaoli Li and Hongxing Rui.

A high-order fully conservative block-centered finite difference method for the time-fractional advection– dispersion equation. Applied Numerical Mathematics: Transactions of IMACS, 124(??):89–109, February 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0168927417302210.

Ludu:2019:RHP

[LR19] A. Ludu and A. Raghavendra. Rotating hollow patterns in fluids. Applied Numerical Mathematics: Transactions of IMACS, 141(??):167–184, ???? 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418302678.

Li:2020:SCB

[LR20a] Xiaoli Li and Hongxing Rui. Stability and convergence based on the finite difference method for the nonlinear fractional cable equation on non-uniform staggered grids. Applied Numerical Mathematics: Transactions of IMACS, 152(??): 403–421, June 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419303216.

Luo:2020:ROEb

[LR20b] Zhendong Luo and Hulin Ren. A reduced-order extrapolated finite difference iterative method for the Riemann–Liouville tempered fractional derivative equation. Applied Numerical Mathematics: Transactions of IMACS, 157(??): 307–314, November 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420301720.

Liang:2024:PRR

[LR24] Hao Liang and Hongxing Rui. A parameter robust reconstruction nonconforming virtual element method for the incompressible poroelasticity model. Applied Numerical Mathematics: Transactions of IMACS, 202(??): 127–142, August 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001077.

Li:2019:SSE

[LRC19] Xiaoli Li, Hongxing Rui, and Shuangshuang Chen. Stability and superconvergence of efficient MAC schemes for fractional Stokes equation on non-uniform grids. Applied Numerical Mathematics: Transactions of IMACS, 138(??):30–53, April 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418302848.

Luzyanina:2004:NSA

[LRE04] T. Luzyanina, D. Roose, and K. Engelborghs. Numerical stability analysis of steady state solutions of integral equations with distributed delays. Applied Numerical Mathematics: Transactions of IMACS, 50(1): 75–92, July 2004. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Levin:2025:DDM

[LRRÁY25] David Levin, José M. Ramón, Juan Ruiz-Álvarez, and Dionisio F. Yáñez. Datadependent moving least squares. Applied Numerical Mathematics: Transactions of IMACS, 216(??): 56–75, October 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742500100X.

Laliena:2009:SBI

[LRS09] A. R. Laliena, M.-L. Rapún, and F.-J. Sayas. Symmetric boundary integral formulations for Helmholtz transmission problems. Applied Numerical Mathematics: Transactions of IMACS, 59(11):2814–2823, November 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Lahmi:2023:NAN

[LRS23] B. Lahmi, M. Rhoudaf, and N. Staïli. Numerical analysis of a nonlinear discrete duality finite volume scheme for Leray–Lions type elliptic problems in Orlicz spaces. Applied Numerical Mathematics: Transactions of IMACS, 185(??): 406–433, March 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422003129.

Luo:1999:CAL

[LRT99] Zhi-Quan Luo, C. Roos, and T. Terlaky. Complexity analysis of logarithmic barrier decomposition methods for semi-infinite linear programming. Applied Numerical Mathematics: Transactions of IMACS, 29(3):379–394, March 1, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1999/29/
3/968.pdf.

Lepe:2025:FEA

[LRV25] Felipe Lepe, Gonzalo Rivera, and Jesus Vellojin. Finite element analysis for the Navier–Lamé eigenvalue problem. Applied Numerical Mathematics: Transactions of IMACS, 208 (part B)(??):1–20, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424002629.

Likeness:1986:EOC

[LS86] Barry Karl Likeness and Arthur David Snider. Efficient optimization of certain functionals in a hypercube. Applied Numerical Mathematics: Transactions of IMACS, 2(1):21– 27, February 1986. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Loze:1993:TSA

[LS93] Michael K. Loze and R. Saunders. Two simple algorithms for constructing a two-dimensional constrained Delaunay triangulation. Applied Numerical Mathematics: Transactions of IMACS, 11(5):403–418, 1993. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Lucht:1998:DBI

[LS98] W. Lucht and K. Strehmel. Discretization based indices for semilinear partial differential algebraic equations. Applied Numerical Mathematics: Transactions of IMACS, 28(2–4):371–386, October 1, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/

store/apnum/sub/1998/28/
2-4/920.pdf.

LeCalvez:1999:MKA

[LS99a] Caroline Le Calvez and Yousef Saad. Modified Krylov acceleration for parallel environments. Applied Numerical Mathematics: Transactions of IMACS, 30(2–3):191–212, June 10, 1999. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1999&volume=30&
issue=2-3&aid=975.

Linss:1999:HDS

[LS99b] Torsten Linß and Martin Stynes. A hybrid difference scheme on a Shishkin mesh for linear convectiondiffusion problems. Applied Numerical Mathematics: Transactions of IMACS, 31(3):255–270, November 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/46/19/
18/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/46/19/18/article.
pdf.

Lunk:2005:RKN

[LS05] Christoph Lunk and Bernd Simeon. Runge–Kutta– Nyström methods with maximized stability domain in structural dynamics. Applied Numerical Mathematics: Transactions of IMACS, 53(2–4):373–389, May 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Lin:2007:VSI

[LS07a] Youdong Lin and Mark A. Stadtherr. Validated solutions of initial value problems for parametric ODEs. Applied Numerical Mathematics: Transactions of IMACS, 57(10):1145–1162, October 2007. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Liu:2007:LDG

[LS07b] Yunxian Liu and Chi-Wang Shu. Local discontinuous Galerkin methods for moment models in device simulations: Performance assessment and two-dimensional results. Applied Numerical Mathematics: Transactions of IMACS, 57(5–7): 629–645, May 2007. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Lera:2010:LHG

[LS10] D. Lera and Ya. D. Sergeyev. Lipschitz and Hölder global optimization using spacefilling curves. Applied Numerical Mathematics: Transactions of IMACS, 60(1–2): 115–129, January/February 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Liao:2012:SYE

[LS12] Qifeng Liao and David Silvester. A simple yet effective a posteriori estimator for classical mixed approximation of Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 62(9): 1242–1256, September 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927410000814.

Lin:2013:MRM

[LS13] Yiding Lin and Valeria Simoncini. Minimal residual methods for large scale Lyapunov equations. Applied Numerical Mathematics: Transactions of IMACS, 72(??):52–71, October 2013. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413000743.

LeBorne:2016:AHW

[LS16] Sabine Le Borne and Lusine Shahmuradyan. Algorithms for the Haar wavelet based fast evaluation of aggregation integrals in population balance equations. Applied Numerical Mathematics: Transactions of IMACS, 108(??):1–20, October 2016. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741630068X.

Liu:2020:NEA

[LS20] Qian Liu and Dongyang Shi. New error analysis of a second order BDF scheme for unsteady natural convection problem. Applied Numerical Mathematics: Transactions of IMACS, 154(??): 243–259, August 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420301197.

Lin:2021:SCP

[LS21] Fu-Rong Lin and Zi-Hang She. Stability and convergence of 3-point WSGD schemes for two-sided space fractional advection-diffusion equations with variable coefficients. Applied Numerical Mathematics: Transactions of IMACS, 167(??): 281–307, September 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001367.

Li:2023:EEC

[LS23] Xiaoli Li and Jie Shen. Error estimate of a consistent splitting GSAV scheme for the Navier–Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 188(??): 62–74, June 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423000661.

Liu:2024:LTA

[LS24a] Baoping Liu and Avy Soffer. The large time asymptotic solutions of nonlinear Schrödinger type equations. Applied Numerical Mathematics: Transactions of IMACS, 199(??): 73–84, May 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423000582.

Liu:2024:NMF

[LS24b] Qian Liu and Dongyang Shi. Nonconforming mixed finite element method for the diffuse interface model of two-phase fluid flows. Applied Numerical Mathematics: Transactions of IMACS, 197(??):272–287, March 2024. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742300291X.

Liu:2024:MGI

[LS24c] Xinwu Liu and Ting Sun. Mixed Gaussian-impulse noise removal using non-convex high-order TV penalty. Applied Numerical Mathematics: Transactions of IMACS, 201(??):72–84, July 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424000345.

Lee:2025:NLS

[LS25] Eunjung Lee and Youngmin Shin. A novel least squares approach generating approximations orthogonal to the null space of the operator. Applied Numerical Mathematics: Transactions of IMACS, 207(??): 370–383, January 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424002526.

Liang:2024:SAT

[LSG24] Conggang Liang, Dongyang Shi, and Longfei Guo. Superconvergence analysis of a two-grid BDF2-FEM for nonlinear dispersive wave equation. Applied Numerical Mathematics: Transactions of IMACS, 201(??): 419–430, July 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000801.

Liu:2015:ASD

[LSGK15] Ping Liu, Giovanni Samaey, C. William Gear, and Ioannis G. Kevrekidis. On the acceleration of spatially distributed agent-based computations: a patch dynamics scheme. Applied Numerical Mathematics: Transactions of IMACS, 92(??): 54–69, June 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414002086.

Li:2012:ENE

[LSK12] Wenbin Li, Sven Simon, and Steffen Kieß. On the estimation of numerical error bounds in linear algebra based on discrete stochastic arithmetic. Applied Numerical Mathematics: Transactions of IMACS, 62(5): 536–555, May 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412000086.

Liang:2011:SAN

[LSL11] Hui Liang, M. H. Song, and M. Z. Liu. Stability of the analytic and numerical solutions for impulsive differential equations. Applied Numerical Mathematics: Transactions of IMACS, 61(11):1103–1113, November 2011. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Li:2020:CSA

[LSP20] Meng Li, Dongyang Shi, and Lifang Pei. Convergence and superconvergence analysis of finite element methods for the time fractional diffusion equation. Applied Numerical Mathematics: Transactions of IMACS, 151(??):141–160, May 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419303745.

Li:2007:PSI

[LST07] Zhilin Li, Yongzhong Song, and Tao Tang. Preface of the special issue of AP-NUM International Conference on Scientific Computing in Nanjing, China. Applied Numerical Mathematics: Transactions of IMACS, 57(5–7):473–474, May 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Le:2022:STD

[LSV22] Van Chien Le, Marián Slodicka, and Karel Van Bockstal. A space-time discretization for an electromagnetic problem with moving non-magnetic conductor. Applied Numerical Mathematics: Transactions of IMACS, 173(??):345– 364, March 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421003433.

Lemos-Silva:2025:ESD

[LSVT25] Márcia Lemos-Silva, Sandra Vaz, and Delfim F. M. Torres. Exact solution for a discrete-time SIR model. Applied Numerical Mathematics: Transactions of IMACS, 207(??): 339–347, January 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424002514.

Liu:2023:TEM

[LSW23] Huagui Liu, Banban Shi, and Fuke Wu. Tamed Euler– Maruyama approximation of McKean–Vlasov stochastic differential equations with super-linear drift and Hölder diffusion coefficients. Applied Numerical Mathematics: Transactions of IMACS, 183(??):56–85, January 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422002173.

Li:2019:USA

[LSWM19] Meng Li, Dongyang Shi, Junjun Wang, and Wanyuan Ming. Unconditional superconvergence analysis of the conservative linearized Galerkin FEMs for nonlinear Klein–Gordon–Schrödinger equation. Applied Numerical Mathematics: Transactions of IMACS, 142(??): 47–63, August 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419300418.

Liu:2022:TTC

[LSWW22] Pengjie Liu, Hu Shao, Yun Wang, and Xiaoyu Wu. A three-term CGPM-based algorithm without Lipschitz continuity for constrained nonlinear monotone equations with applications. Applied Numerical Mathematics: Transactions of IMACS, 175(??):98–107, May 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422000277.

Li:2017:PWM

[LSY17] Wen Li, Guohui Song, and Guangming Yao. Piece-wise moving least squares approximation. Applied Numerical Mathematics: Transactions of IMACS, 115(??): 68–81, May 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417300016.

Liu:2021:RWF

[LSY21] J. J. Liu, C. L. Sun, and M. Yamamoto. Recovering the weight function in distributed order fractional equation from interior measurement. Applied Numerical Mathematics: Transactions of IMACS, 168(??):84– 103, October 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001550.

Lopez:1993:BVM

[LT93] L. Lopez and D. Trigiante.

Boundary value methods and BV-stability in the solution of initial value problems. Applied Numerical Mathematics: Transactions of IMACS, 11(1– 3):225–239, January 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Parallel methods for ordinary differential equations (Grado, 1991).

Liu:2000:NSV

[LT00] Biyue Liu and Dalin Tang.

A numerical simulation of viscous flows in collapsible tubes with stenoses. Applied Numerical Mathematics: Transactions of IMACS, 32(1):87–101, January 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/59/28/
32/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/59/28/32/article.
pdf.

Liu:2001:EAG

[LT01] Wenbin Liu and Tao Tang. Error analysis for a Galerkinspectral method with coordinate transformation for solving singularly perturbed problems. Applied Numerical Mathematics: Transactions of IMACS, 38(3):315– 345, August 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/77/33/31/abstract.
html.

Loncaric:2005:QOP

[LT05] J. Lončarić and S. V. Tsynkov. Quadratic optimization in the problems of active control of sound. Applied Numerical Mathematics: Transactions of IMACS, 52(4):381–400, March 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Lewis:2007:RTI

[LT07] B. M. Lewis and H. T.

Tran. Real-time implementation of a frequency shaping controller on a cantilever beam. Applied Numerical Mathematics: Transactions of IMACS, 57(5–7): 778–790, May 2007. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Layton:2012:STI

[LT12] W. Layton and C. Trenchea.

Stability of two IMEX methods, CNLF and BDF2– AB2, for uncoupling systems of evolution equations. Applied Numerical Mathematics: Transactions of IMACS, 62(2):112– 120, February 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411001930.

Lord:2019:SEI

[LT19] Gabriel J. Lord and Antoine Tambue. Stochastic exponential integrators for a finite element discretisation of SPDEs with additive noise. Applied Numerical Mathematics: Transactions of IMACS, 136(??): 163–182, February 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418302344.

Liu:2023:ADF

[LT23] Anne Liu and Thomas Trogdon. An artificiallydamped Fourier method for dispersive evolution equations. Applied Numerical Mathematics: Transactions of IMACS, 192(??): 19–40, October 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423001526.

Larrazabal:2003:ERA

[LTC03] G. Larrazábal, C. R. Torres, and J. Castillo. An efficient and robust algorithm for 2D stratified fluid flow calculations. Applied Numerical Mathematics: Transactions of IMACS, 47(3– 4):493–502, December 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Li:2024:EIM

[LTDY24] Chunmei Li, Dan Tian, Xuefeng Duan, and Naya Yang. An efficient iterative method for solving the graph regularization Q-weighted nonnegative matrix factorization problem in multi-view clustering. Applied Numerical Mathematics: Transactions of IMACS, 205(??): 255–266, November 2024.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001909.

Lteif:2024:OSA

[Lte24] Ralph Lteif. An operatorsplitting approach with a hybrid finite volume/finite difference scheme for extended Boussinesq models. Applied Numerical Mathematics: Transactions of IMACS, 196(??):159– 182, February 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423002726.

Liu:2019:ANS

[LTT19] Zhongli Liu, Tianhai Tian, and Hongjiong Tian. Asymptoticnumerical solvers for highly oscillatory second-order differential equations. Applied Numerical Mathematics: Transactions of IMACS, 137(??):184–202, March 2019. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418302538.

Lu:1998:CCR

[Lu98a]  Ya Yan Lu.  A com-

plex coefficient rational approximation of ROOT(1+x). Applied Numerical Mathematics: Transactions of IMACS, 27(2):141–154, June 10, 1998. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1998/27/
                                        2/875.pdf.

Lu:1998:CLS

[Lu98b] Ya Yan Lu. Computing the logarithm of a symmetric positive definite matrix. Applied Numerical Mathematics: Transactions of IMACS, 26(4):483–496, April 1, 1998. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1998/26/
                                        4/859.pdf.

Luan:2017:FOT

[Lua17] Vu Thai Luan. Fourthorder two-stage explicit exponential integrators for time-dependent PDEs. Applied Numerical Mathematics: Transactions of IMACS, 112(??):91–103, February 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0168927416302021.

Lubich:1992:CQH

[Lub92] C. Lubich. On convolution quadrature and Hille– Phillips operational calculus. Applied Numerical Mathematics: Transactions of IMACS, 9(3–5): 187–199, April 1992. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0168927492900145.

International Conference on the Numerical Solution of Volterra and Delay Equations (Tempe, AZ, 1990).

Lubich:2004:VSI

[Lub04] Christian Lubich. A variational splitting integrator for quantum molecular dynamics. Applied Numerical Mathematics: Transactions of IMACS, 48(3–4): 355–368, March 2004. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Lucht:1995:FEM

[Luc95] W. Lucht. A finite element method for an ill-posed problem. Applied Numerical Mathematics: Transactions of IMACS, 18(1–3):253–266, September 27, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=18&
issue=1-3&aid=603. Sev-

enth Conference on the Numerical Treatment of Differential Equations (Halle, 1994).

Lucht:2005:PDA

[Luc05] W. Lucht. Partial differentialalgebraic systems of second order with symmetric convection. Applied Numerical Mathematics: Transactions of IMACS, 53(2–4): 357–371, May 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Luo:2018:SVM

[Luo18]  Xue Luo.  Spectral vis-

cosity method with generalized Hermite functions for nonlinear conservation laws. Applied Numerical Mathematics: Transactions of IMACS, 123(??): 256–274, January 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417302088.

Lampe:2012:EDH

[LV12] J. Lampe and H. Voss. Efficient determination of the hyperparameter in regularized total least squares problems. Applied Numerical Mathematics: Transactions of IMACS, 62(9): 1229–1241, September 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927410001170.

Li:2014:EPB

[LVfP14] Wen Li, Seak-Weng Vong, and Xiao fei Peng. On eigenvalue perturbation bounds for Hermitian block tridiagonal matrices. Applied Numerical Mathematics: Transactions of IMACS, 83(??):38–50, September 2014. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0168927414000701.

Liska:2021:LWM

[LVW21] Richard Liska, Pavel Váchal, and Burton Wendroff. Lax– Wendroff methods on highly non-uniform meshes. Dedicated to the memory of Blair Swartz (1932–2019). Applied Numerical Mathematics: Transactions of IMACS, 163(??):167–181, May 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000325.

Linz:1992:EBS

[LW92a] Peter Linz and Richard L. C. Wang. Error bounds for the solution of Volterra and delay equations. Applied Numerical Mathematics: Transactions of IMACS, 9(3–5):201–207, April 1992.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0168927492900156.

International Conference on the Numerical Solution of Volterra and Delay Equations (Tempe, AZ, 1990).

Lioen:1992:SLD

[LW92b] Walter M. Lioen and Dik T. Winter. Solving large dense systems of linear equations on systems with virtual memory and with cache. Applied Numerical Mathematics: Transactions of IMACS, 10(1):73–85, June 1992. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Lang:1993:ARM

[LW93] Jens Lang and Artur Walter. An adaptive Rothe method for nonlinear reaction-diffusion systems. Applied Numerical Mathematics: Transactions of IMACS, 13(1–3): 135–146, September 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Sixth Conference on the Numerical Treatment of Differential Equations (Halle, 1992).

Lube:1995:SFE

[LW95] G. Lube and D. Weiss. Stabilized finite element methods for singularly perturbed parabolic problems. Applied Numerical Mathematics: Transactions of IMACS, 17(4):431–459, September 11, 1995. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=17&
issue=4&aid=572.

Li:2004:MAY

[LW04] Zhiping Li and Xiaonan Wu.

Multi-atomic Young measure and artificial boundary in approximation of micromagnetics. Applied Numerical Mathematics: Transactions of IMACS, 51(1):69– 88, October 2004. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Liu:2007:CMV

[LW07] Hailiang Liu and Zhongming Wang. Computing multivalued velocity and electric fields for 1D Euler– Poisson equations. Applied Numerical Mathematics: Transactions of IMACS, 57(5–7):821–836, May 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Li:2017:MSR

[LW17] Jiyong Li and Xianfen Wang.

Multi-step Runge–Kutta– Nyström methods for special second-order initial value problems. Applied Numerical Mathematics: Transactions of IMACS, 113(??): 54–70, March 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416302215.

Li:2018:PCN

[LW18a] Hanyu Li and Shaoxin Wang. On the partial condition numbers for the indefinite least squares problem. Applied Numerical Mathematics: Transactions of IMACS, 123(??): 200–220, January 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417301885.

Liu:2018:SBH

[LW18b] Jijun Liu and Bingxian Wang. Solving the backward heat conduction problem by homotopy analysis method. Applied Numerical Mathematics: Transactions of IMACS, 128(??): 84–97, June 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418300412.

Li:2019:LDG

[LW19a] Changpin Li and Zhen Wang. The local discontinuous Galerkin finite element methods for Caputo-type partial differential equations: Numerical analysis. Applied Numerical Mathematics: Transactions of IMACS, 140(??):1–22, June 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419300182.

Li:2019:EPC

[LW19b] Jiyong Li and Xinyuan Wu.

Energy-preserving continuous stage extended Runge– Kutta–Nyström methods for oscillatory Hamiltonian systems. Applied Numerical Mathematics: Transactions of IMACS, 145(??): 469–487, November 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927419301163.

Li:2020:LDG

[LW20a] Changpin Li and Zhen Wang. The local discontinuous Galerkin finite element methods for Caputo-type partial differential equations: mathematical analysis. Applied Numerical Mathematics: Transactions of IMACS, 150(??):587–606, April 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419303150.

Liu:2020:NSC

[LW20b] Changying Liu and Xinyuan Wu. Nonlinear stability and convergence of ERKN integrators for solving nonlinear multi-frequency highly oscillatory second-order ODEs with applications to semilinear wave equations. Applied Numerical Mathematics: Transactions of IMACS, 153(??):352–380, July 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420300660.

Li:2021:OPW

[LW21a] Jiyong Li and Tingchun Wang. Optimal point-wise error estimate of two conservative fourth-order compact finite difference schemes for the nonlinear Dirac equation. Applied Numerical Mathematics: Transactions of IMACS, 162(??): 150–170, April 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420303834.

Li:2021:SKF

[LW21b] X. Y. Li and B. Y. Wu. Superconvergent kernel functions approaches for the second kind Fredholm integral equations. Applied Numerical Mathematics: Transactions of IMACS, 167(??): 202–210, September 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001318.

Li:2022:GCS

[LW22] Ting Li and Bin Wang. Geometric continuous-stage exponential energy-preserving integrators for chargedparticle dynamics in a magnetic field from normal to strong regimes. Applied Numerical Mathematics: Transactions of IMACS, 181(??):1–22, November 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001441.

Lakshmi:2025:NPS

[LW25] Mayur V. Lakshmi and Joab R. Winkler. Numerical properties of solutions of LASSO regression. Applied Numerical Mathematics: Transactions of IMACS, 208 (part A) (??):297–309, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927424000576.

Li:2024:NAF

[LWaZ24] Haozheng Li, Bo Wang, and Guang an Zou. Numerical analysis of finite element method for a stochastic active fluids model. Applied Numerical Mathematics: Transactions of IMACS, 201(??):217–246, July 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424000540.

Liu:2024:LDG

[LWC24] Yanhua Liu, Xuesong Wang, and Yao Cheng. Local discontinuous Galerkin method for a singularly perturbed fourth-order problem of convection–diffusion type. Applied Numerical Mathematics: Transactions of IMACS, 205(??):16– 37, November 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424001715.

Li:2019:MFS

[LWCH19] Zi-Cai Li, Yimin Wei, Yunkun Chen, and Hung-Tsai Huang. The method of fundamental solutions for the Helmholtz equation. Applied Numerical Mathematics: Transactions of IMACS, 135(??): 510–536, January 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418302058.

Lui:2007:LCG

[LWCT07] Lok Ming Lui, Yalin Wang, Tony F. Chan, and Paul Thompson. Landmark constrained genus zero surface conformal mapping and its application to brain mapping research. Applied Numerical Mathematics: Transactions of IMACS, 57(5–7): 847–858, May 2007. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Lew:2009:ART

[LWD^{+}09] S. Lew, C. H. Wolters, T. Dierkes, C. Röer, and R. S. MacLeod. Accuracy and run-time comparison for different potential approaches and iterative solvers in finite element method based EEG source analysis. Applied Numerical Mathematics: Transactions of IMACS, 59(8): 1970–1988, August 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Li:2018:CCM

[LWL18] Yiqun Li, Boying Wu, and Melvin Leok. Construction and comparison of multidimensional spectral variational integrators and spectral collocation methods. Applied Numerical Mathematics: Transactions of IMACS, 132(??):35–50, October 2018. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741830120X.

Long:2024:BPS

[LWLL24] Xian Jun Long, Xiao Ting Wang, Gao Xi Li, and Geng Hua Li. A Bregman proximal subgradient algorithm for nonconvex and nonsmooth fractional optimization problems. Applied Numerical Mathematics: Transactions of IMACS, 202(??): 209–221, August 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001120.

Li:2024:PBH

[LWLW24] Dandan Li, Songhua Wang, Yong Li, and Jiaqi Wu. A projection-based hybrid PRP-DY type conjugate gradient algorithm for constrained nonlinear equations with applications. Applied Numerical Mathematics: Transactions of IMACS, 195(??):105–125, January 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423002532.

Lai:2007:ESC

[LWT07] Ming-Chih Lai, Chin-Tien Wu, and Yu-Hou Tseng. An efficient semi-coarsening multigrid method for variable diffusion problems in cylindrical coordinates. Applied Numerical Mathematics: Transactions of IMACS, 57(5–7):801–810, May 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Li:2020:SGA

[LWW20] Dan Li, Chunmei Wang, and Junping Wang. Superconvergence of the gradient approximation for weak Galerkin finite element methods on nonuniform rectangular partitions. Applied Numerical Mathematics: Transactions of IMACS, 150(??): 396–417, April 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419302880.

Li:2022:SET

[LWW22] X. Y. Li, H. L. Wang, and B. Y. Wu. A stable and efficient technique for linear boundary value problems by applying kernel functions. Applied Numerical Mathematics: Transactions of IMACS, 172(??): 206–214, February 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421002932.

Li:2023:VTS

[LWW23] Meng Li, Lingli Wang, and Nan Wang. Variable-time-step BDF2 nonconforming VEM for coupled Ginzburg– Landau equations. Applied Numerical Mathematics: Transactions of IMACS, 186(??):378–410, April 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423000314.

Li:2010:ILQ

[LWWX10] Jianbing Li, Xuesong Wang, Tao Wang, and Shunping Xiao. An improved Levin quadrature method for highly oscillatory integrals. Applied Numerical Mathematics: Transactions of IMACS, 60(8):833– 842, August 2010. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Li:2025:VTS

[LWWZ25] Meng Li, Dan Wang, Junjun Wang, and Xiaolong Zhao. Variable-time-step weighted IMEX FEMs for nonlinear evolution equations. Applied Numerical Mathematics: Transactions of IMACS, 211(??):123–143, May 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425000054.

Li:2020:HVC

[LWY20] Shuangshuang Li, Lina Wang, and Lijun Yi. An hpversion of the C^{0}-continuous Petrov–Galerkin method for second-order Volterra integro-differential equations. Applied Numerical Mathematics: Transactions of IMACS, 152(??): 84–104, June 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300180.

Li:2022:FDS

[LWYG22] Can Li, Haihong Wang, Hongyun Yue, and Shimin Guo. Fast difference scheme for the reaction–diffusionadvection equation with exact artificial boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 173(??): 395–417, March 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421003470.

Lin:2022:SSP

[LWZ22] Yiding Lin, Xiang Wang, and Lei-Hong Zhang. Solving symmetric and positive definite second-order cone linear complementarity problem by a rational Krylov subspace method. Applied Numerical Mathematics: Transactions of IMACS, 176(??): 104–117, June 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422000484.

Li:2008:CCP

[LX08] Min Li and Aiguo Xiao. Characterizations and construction of Poisson/symplectic and symmetric multi-revolution implicit Runge–Kutta methods of high order. Applied Numerical Mathematics: Transactions of IMACS, 58(6):915–930, June 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Lin:2009:AEE

[LX09] Qun Lin and Hehu Xie. Asymptotic error expansion and Richardson extrapolation of eigenvalue approximations for second order elliptic problems by the mixed finite element method. Applied Numerical Mathematics: Transactions of IMACS, 59(8): 1884–1893, August 2009.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Li:2021:SDD

[LX21] Wuyang Li and Yingxiang Xu. Schwarz domain decomposition methods for the fluid-fluid system with friction-type interface conditions. Applied Numerical Mathematics: Transactions of IMACS, 166(??): 114–126, August 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001045.

Linss:2024:HWG

[LX24] Torsten Linß and Christos Xenophontos. An hp Weak Galerkin FEM for singularly perturbed problems. Applied Numerical Mathematics: Transactions of IMACS, 197(??):1–14, March 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423002763.

Lu:2021:DHS

[LXCM21] Kang-Ya Lu, Dong-Xiu Xie, Fang Chen, and Galina V. Muratova. Dominant Hermitian splitting iteration method for discrete spacefractional diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 164(??): 15–28, ???? 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420300726.

Luo:2025:HBV

[LXLY25] Qian Luo, Aiguo Xiao, Xuqiong Luo, and Xiaoqiang Yan. Hamiltonian boundary value methods applied to KdV–KdV systems. Applied Numerical Mathematics: Transactions of IMACS, 214(??):1–27, August 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425000303.

Liu:2025:SOS

[LXPY25] Xin Liu, Dandan Xue, Shuaichao Pei, and Hong Yang. A second-order semiimplicit spectral deferred correction scheme for Cahn– Hilliard–Navier–Stokes equation. Applied Numerical Mathematics: Transactions of IMACS, 216(??): 39–55, October 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425001011.

Li:2021:HOC

[LXZ21] Tingyue Li, Dinghua Xu, and Qifeng Zhang. Highorder compact schemes for semilinear parabolic moving boundary problems. Applied Numerical Mathematics: Transactions of IMACS, 161(??):452–468, March 2021. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303779.

Li:2022:NTL

[LXZS22] Shuguang Li, Da Xu, Jie Zhang, and Chengjiao Sun. A new three-level fourthorder compact finite difference scheme for the extended Fisher–Kolmogorov equation. Applied Numerical Mathematics: Transactions of IMACS, 178(??): 41–51, August 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422000721.

Li:2001:GSB

[LY01] Zi Cai Li and Ningning Yan. Global superconvergence for blending surfaces by boundary penalty plus hybrid FEMs. Applied Numerical Mathematics: Transactions of IMACS, 39(1):61– 85, October 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/

                                        10/28/79/27/31/abstract.
                                        html.

Liu:2003:PEE

[LY03] Wenbin Liu and Ningning Yan. A posteriori error estimates for control problems governed by nonlinear elliptic equations. Applied Numerical Mathematics: Transactions of IMACS, 47(2):173–187, November 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Li:2008:DDC

[LY08] Changfeng Li and Yirang Yuan. Domain decomposition with characteristic finite difference method for two-phase displacement problems. Applied Numerical Mathematics: Transactions of IMACS, 58(9): 1262–1273, September 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Li:2009:MUD

[LY09] Changfeng Li and Yirang Yuan. A modified upwind difference domain decomposition method for convection–diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 59(7):1584–1598, July 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Lee:2010:NSS

[LY10] Yeon Ju Lee and Jungho Yoon. Non-stationary subdivision schemes for surface interpolation based on exponential polynomials. Applied Numerical Mathematics: Transactions of IMACS, 60(1–2):130–141, January/ February 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Lin:2016:HPV

[LY16] Po-Hsien Lin and S.-T. John Yu. Hyperbolicity of pressure-velocity equations for computational hydro acoustics. Applied Numerical Mathematics: Transactions of IMACS, 109(??):73– 84, November 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416300861.

Liu:2024:ESE

[LY24a] Shiqin Liu and Haijun Yu.

Energetic spectral-element time marching methods for phase-field nonlinear gradient systems. Applied Numerical Mathematics: Transactions of IMACS, 205(??):38– 59, November 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001697.

Lorin:2024:CWD

[LY24b] Emmanuel Lorin and Xu Yang. Computational wavefunction dynamics in photonic graphene with symmetry breaking. Applied Numerical Mathematics: Transactions of IMACS, 199(??): 85–104, May 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423001514.

Luo:2019:GEB

[LYA^{+}19] Yongbing Luo, Yanbing Yang, Md Salik Ahmed, Tao Yu, Mingyou Zhang, Ligang Wang, and Huichao Xu. Global existence and blow up of the solution for nonlinear Klein–Gordon equation with general powertype nonlinearities at three initial energy levels. Applied Numerical Mathematics: Transactions of IMACS, 141(??):102–123, ???? 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418301363.

Long:2024:FML

[LYC24] Guangqing Long, Huanfeng Yang, and Dongsheng Cheng. Fast multi-level iteration schemes with compression technique for eigenproblems of compact integral operators. Applied Numerical Mathematics: Transactions of IMACS, 198(??): 448–460, April 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000278.

Li:2017:EEH

[LYF17] Zhiqiang Li, Yubin Yan, and Neville J. Ford. Error estimates of a high order numerical method for solving linear fractional differential equations. Applied Numerical Mathematics: Transactions of IMACS, 114(??):201–220, April 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416300587.

Liu:2017:DGT

[LYK17] Yanmei Liu, Yubin Yan, and Monzorul Khan. Discontinuous Galerkin time stepping method for solving linear space fractional partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 115(??):200–213, May 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417300247.

Liu:2023:CES

[LYLL23] Hongliang Liu, Yilin You, Haodong Li, and Shoufu Li. Canonical Euler splitting method for parabolic partial functional differential algebraic equations. Applied Numerical Mathematics: Transactions of IMACS, 190(??):65–83, August 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423001034.

Liu:2024:SPL

[LYMD24] Xuewei Liu, Zhanwen Yang, Qiang Ma, and Xiaohua Ding. A structure-preserving local discontinuous Galerkin method for the stochastic KdV equation. Applied Numerical Mathematics: Transactions of IMACS, 204(??):1–25, October 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424001417.

Lynch:1992:FSN

[Lyn92] Robert E. Lynch. Fundamental solutions of ninepoint discrete Laplacians. Applied Numerical Mathematics: Transactions of IMACS, 10(3–4):325–334, September 1992. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). A Festschrift to honor Professor Garrett Birkhoff on his eightieth birthday.

Lynch:1999:LAI

[Lyn99] Mark A. M. Lynch. Large amplitude instability in finite difference approximations to the Klein–Gordon equation. Applied Numerical Mathematics: Transactions of IMACS, 31(2): 173–182, October 23, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1999&volume=31&
issue=2&aid=993;   http:
//www.sciencedirect.com/
science/article/pii/S0168927498001287.

Lyon:2012:AER

[Lyo12] Mark Lyon. Approximation error in regularized SVD-based Fourier continuations. Applied Numerical Mathematics: Transactions of IMACS, 62(12): 1790–1803, December 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412001225.

Lohner:1999:UGB

[LYOI99] Rainald Löhner, Chi Yang, Eugenio Oñate, and Sergio Idelssohn. An unstructured grid-based, parallel free surface solver. Applied Numerical Mathematics: Transactions of IMACS, 31(3): 271–293, November 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/46/19/
19/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/46/19/19/article.
pdf.

Liu:2015:URI

[LYY15] J. J. Liu, M. Yamamoto, and L. Yan. On the uniqueness and reconstruction for an inverse problem of the fractional diffusion process. Applied Numerical Mathematics: Transactions of IMACS, 87(??):1–19, January 2015. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414001408.

Lin:2025:GCC

[LYZ25] Fu-Rong Lin, Xi Yang, and Gui-Rong Zhang. Generalized Clenshaw–Curtis quadrature method for systems of linear ODEs with constant coefficients. Applied Numerical Mathematics: Transactions of IMACS, 217(??):112–125, November 2025. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425001205.

Li:2025:HOS

[LYZF25] Zhi-Chen Li, Yang-Wen Yu, Xiao-Yu Zhang, and Qing Fang. A higher-order solver for the FitzHugh– Nagumo equation by combining nonstandard and compact finite difference scheme. Applied Numerical Mathematics: Transactions of IMACS, 217(??): 436–450, November 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742500145X.

Li:2023:OEE

[LYZJ23] Yingyuan Li, Wenjing Yan, Shengfeng Zhu, and Feifei Jing. Optimal error estimates of the discrete shape gradients for shape optimizations governed by the Stokes–Brinkman equations. Applied Numerical Mathematics: Transactions of IMACS, 190(??): 220–253, August 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001186.

Liu:2022:IWC

[LYZW22] Can Liu, Zhe Yu, Xinming Zhang, and Boying Wu. An implicit wavelet collocation method for variable coefficients space fractional advection-diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 177(??): 93–110, July 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422000691.

Leng:2013:IAS

[LZ13] Huinan Leng and Qi Zhang. An iterative algorithm for the stability analysis of dynamic interval systems. Applied Numerical Mathematics: Transactions of IMACS, 69(??):73–77, July 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413000421.

Li:2014:IEE

[LZ14] Dongfang Li and Chengjian Zhang. L^{∞} error estimates of discontinuous Galerkin methods for delay differential equations. Applied Numerical Mathematics: Transactions of IMACS, 82(??): 1–10, August 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414000488.

Li:2017:BBV

[LZ17] Cui Li and Chengjian Zhang.

Block boundary value methods applied to functional differential equations with piecewise continuous arguments. Applied Numerical Mathematics: Transactions of IMACS, 115(??): 214–224, May 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417300272.

Li:2018:EET

[LZ18] Xin Li and Luming Zhang.

Error estimates of a trigonometric integrator sine pseudospectral method for the extended Fisher–Kolmogorov equation. Applied Numerical Mathematics: Transactions of IMACS, 131(??):39– 53, September 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418300990.

Li:2020:MFV

[LZ20] Hengguang Li and Qinghui Zhang. A meshfree finite volume method with optimal numerical integration and direct imposition of essential boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 153(??):98–113, July 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420300404.

Li:2022:BGS

[LZ22] Cui Li and Yongtao Zhou. Block generalized Störmer– Cowell methods applied to second order nonlinear delay differential equations. Applied Numerical Mathematics: Transactions of IMACS, 178(??): 296–303, August 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422000988.

Li:2024:CSP

[LZ24] Xin Li and Luming Zhang. Conformal structure-preserving SVM methods for the nonlinear Schrödinger equation with weakly linear damping term. Applied Numerical Mathematics: Transactions of IMACS, 205(??): 120–136, November 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424001727.

Liu:2021:CNA

[LZCF21] Jun Liu, Chen Zhu, Yanping Chen, and Hongfei Fu. A Crank–Nicolson ADI quadratic spline collocation method for two-dimensional Riemann–Liouville spacefractional diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 160(??): 331–348, February 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303226.

Li:2019:NSR

[LZH19] Hongwei Li, Xin Zhao, and Yunxia Hu. Numerical solution of the regularized logarithmic Schrödinger equation on unbounded domains. Applied Numerical Mathematics: Transactions of IMACS, 140(??):91–103, June 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419300352.

Lin:2025:IPP

[LZH^{+}25] Si-Da Lin, Ya-Jing Zhang, Ming Huang, Jin-Long Yuan, and Hong-Han Bei. Inexact proximal penalty alternating linearization decomposition scheme of nonsmooth convex constrained optimization problems. Applied Numerical Mathematics: Transactions of IMACS, 211(??):42–60, May 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424003593.

Liu:2023:EEF

[LZIZ23] Wenju Liu, Tengjin Zhao, Kazufumi Ito, and Zhiyue Zhang. Error estimates of Fourier finite volume element method for parabolic Dirichlet boundary optimal control problems on complex connected domains. Applied Numerical Mathematics: Transactions of IMACS, 186(??):164–201, April 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423000065.

Li:2021:MPS

[LZJ21] Shuai Li, Zhixin Zhang, and Wei Jiang. Multiple positive solutions for fourpoint boundary value problem of fractional delay differential equations with p-Laplacian operator. Applied Numerical Mathematics: Transactions of IMACS, 165(??):348–356, July 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421000714.

Liu:2014:ANE

[LZL14] X. Liu, G. L. Zhang, and M. Z. Liu. Analytic and numerical exponential asymptotic stability of nonlinear impulsive differential equations. Applied Numerical Mathematics: Transactions of IMACS, 81(??): 40–49, July 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414000440.

Lan:2022:ESE

[LZQ22] Guangqiang Lan, Mei Zhao, and Siyuan Qi. Exponential stability of θ-EM method for nonlinear stochastic Volterra integro-differential equations. Applied Numerical Mathematics: Transactions of IMACS, 172(??): 279–291, February 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742100297X.

Liao:2017:NAC

[LZW17] Feng Liao, Luming Zhang, and Shanshan Wang. Numerical analysis of cubic orthogonal spline collocation methods for the coupled Schrödinger–Boussinesq equations. Applied Numerical Mathematics: Transactions of IMACS, 119(??): 194–212, September 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417301034.

Liao:2019:UCC

[LZW19] Feng Liao, Luming Zhang, and Tingchun Wang. Unconditional L_{∞} convergence of a conservative compact finite difference scheme for the N-coupled Schrödinger– Boussinesq equations. Applied Numerical Mathematics: Transactions of IMACS, 138(??):54–77, April 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418302836.

Liu:2020:STT

[LZW20] J. K. Liu, Y. X. Zhao, and X. L. Wu. Some three-term conjugate gradient methods with the new direction structure. Applied Numerical Mathematics: Transactions of IMACS, 150(??):433–443, April 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419302867.

Luo:2009:RFE

[LZY09] Zhendong Luo, Yanjie Zhou, and Xiaozhong Yang. A reduced finite element formulation based on proper orthogonal decomposition for Burgers’ equation. Applied Numerical Mathematics: Transactions of IMACS, 59(8):1933–1946, August 2009. CODEN ANMAEL.

ISSN 0168-9274 (print), 1873-5460 (electronic).

Liu:2024:NTS

[LZY24] X. Liu, M. Zhang, and Z. W.

Yang. Numerical threshold stability of a nonlinear age-structured reaction diffusion heroin transmission model. Applied Numerical Mathematics: Transactions of IMACS, 204(??): 291–311, October 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001648.

Li:2018:VLD

[LZZ18] Zhe Li, Kai Zheng, and Shugong Zhang. Verification of a low-degree polynomial vanishing at empirical points. Applied Numerical Mathematics: Transactions of IMACS, 134(??): 122–131, December 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418301612.

Li:2022:CLO

[LZZ22] Dongping Li, Yue Zhang, and Xiuying Zhang. Computing the Lyapunov operator ϕ-functions, with an application to matrixvalued exponential integrators. Applied Numerical Mathematics: Transactions of IMACS, 182(??): 330–343, December 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422002033.

Li:2025:BSB

[LZZ25] Wenbin Li, Tinggang Zhao, and Zhenyu Zhao. A biharmonic solver based on Fourier extension with oversampling technique for arbitrary domain. Applied Numerical Mathematics: Transactions of IMACS, 218(??):261–274, December 2025. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425001692.

Ma:2003:ERN

[Ma03] T. F. Ma. Existence results and numerical solutions for a beam equation with nonlinear boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 47(2): 189–196, November 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Markov:2004:SAA

[MA04] Svetoslav Markov and Rene Alt. Stochastic arithmetic: Addition and multiplication by scalars. Applied Numerical Mathematics: Transactions of IMACS, 50(3–4): 475–488, September 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Makai:2009:AFG

[MA09] Mihály Makai and Miklós Antal. Applications of finite groups to iterative problems in reactor physics. Applied Numerical Mathematics: Transactions of IMACS, 59(6):1237–1257, June 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ma:2024:TRN

[Ma24] Wen-Xiu Ma. Type (λ^{∗}, λ)

reduced nonlocal integrable AKNS equations and their soliton solutions. Applied Numerical Mathematics: Transactions of IMACS, 199(??):105–113, May 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422003294.

MacGiollaMhuiris:1986:CSS

[Mac86] Nessan Mac Giolla Mhuiris.

Calculations of the stability of some axisymmetric flows proposed as a model of vortex breakdown. Applied Numerical Mathematics: Transactions of IMACS, 2(3–5):273–290, October 1986. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

MacLeod:1992:CEA

[Mac92] Allan J. MacLeod. Chebyshev expansions for Abramowitz functions. Applied Numerical Mathematics: Transactions of IMACS, 10(2): 129–137, July 1992. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Mohebalizadeh:2023:FLS

[MAD23] Hamed Mohebalizadeh, Hojatollah Adibi, and Mehdi Dehghan. On the fractional Laplacian of some positive definite kernels with applications in numerically solving the surface quasi-geostrophic equation as a prominent fractional calculus model. Applied Numerical Mathematics: Transactions of IMACS, 188(??):75–87, June 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423000648.

Moradi:2020:SSP

[MAF20] A. Moradi, A. Abdi, and J. Farzi. Strong stability preserving second derivative diagonally implicit multistage integration methods. Applied Numerical Mathematics: Transactions of IMACS, 150(??):536–558, April 2020.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419303095.

Magolu:1991:CAS

[Mag91] Monga-Made Magolu. Conditioning analysis of sparse block approximate factorizations. Applied Numerical Mathematics: Transactions of IMACS, 8(1):25– 42, August 1991. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Michaeli:2013:NTM

[MAG13] Michael Michaeli, Franck Assous, and Anatoly Golubchik. A Nitsche type method for stress fields calculation in dissimilar material with interface crack. Applied Numerical Mathematics: Transactions of IMACS, 67(??):187–203, May 2013. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411001516.

Mahdi:2018:EGL

[MAH18] H. Mahdi, A. Abdi, and G. Hojjati. Efficient general linear methods for a class of Volterra integrodifferential equations. Applied Numerical Mathematics: Transactions of IMACS, 127(??):95–109, May 2018.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418300102.

Moradi:2022:IES

[MAH22] A. Moradi, A. Abdi, and G. Hojjati. Implicit–explicit second derivative general linear methods with strong stability preserving explicit part. Applied Numerical Mathematics: Transactions of IMACS, 181(??):23– 45, November 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742200143X.

Ma:2021:MDS

[MAHZ21] Zheng Ma, Anatoly A. Alikhanov, Chengming Huang, and Guoyu Zhang. A multidomain spectral collocation method for Volterra integral equations with a weakly singular kernel. Applied Numerical Mathematics: Transactions of IMACS, 167(??): 218–236, September 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421001355.

Maischak:2006:MSA

[Mai06] Matthias Maischak. Multiplicative Schwarz algorithms for the p-version Galerkin boundary element method in 3D. Applied Numerical Mathematics: Transactions of IMACS, 56(10–11):1370– 1382, October/November 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Maischak:2009:MAS

[Mai09]  Matthias Maischak.  A

multilevel additive Schwarz method for a hypersingular integral equation on an open curve with graded meshes. Applied Numerical Mathematics: Transactions of IMACS, 59(9): 2195–2202, September 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Majidian:2014:CQR

[Maj14] Hassan Majidian. Composite quadrature rules for a class of weakly singular Volterra integral equations with noncompact kernels. Applied Numerical Mathematics: Transactions of IMACS, 83(??):1– 11, September 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414000452.

Majidian:2017:FCC

[Maj17a] Hassan Majidian. Filon– Clenshaw–Curtis formulas for highly oscillatory integrals in the presence of stationary points. Applied Numerical Mathematics: Transactions of IMACS, 117(??):87–102, July 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417300399.

Majidian:2017:DRC

[Maj17b] Hassan Majidian. On the decay rate of Chebyshev coefficients. Applied Numerical Mathematics: Transactions of IMACS, 113(??): 44–53, March 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416302239.

Majic:2020:SIA

[Maj20]  Matt Majic.  A surface

integral approach to Poisson’s equation and analytic expressions for the gravitational field of toroidal mass distributions. Applied Numerical Mathematics: Transactions of IMACS, 148(??):98–108, February 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0168927419302156.

Mannix:1996:PFC

[Man96] C. E. Mannix, Jr. Potential function construction by use of an adaptive mesh algorithm for a class of singular integral equations. Applied Numerical Mathematics: Transactions of IMACS, 22(4):451– 475, December 16, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=22&
issue=4&aid=727.

Mansky:1997:OPN

[Man97] E. J. Mansky. An open problem in the numerical analysis of systems of Volterra integral equations of the second kind in two space-like variables. Applied Numerical Mathematics: Transactions of IMACS, 24(2– 3):149–151, June 19, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=24&
issue=2-3&aid=773. Volterra

centennial (Tempe, AZ, 1996).

Maponi:2005:NMD

[Map05]  P. Maponi.  A numeri-

cal method for a direct obstacle scattering problem. Applied Numerical Mathematics: Transactions of IMACS, 55(3):322–333, November 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Marcos:1993:NCB

[Mar93] M. Á. López Marcos. A note on the computation of bifurcation diagrams of the Kuramoto–Sivashinsky equation by pseudospectral methods. Applied Numerical Mathematics: Transactions of IMACS, 13(1–3): 147–154, September 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Sixth Conference on the Numerical Treatment of Differential Equations (Halle, 1992).

Maroni:1994:MCO

[Mar94] P. Maroni. Modified classical orthogonal polynomials associated with oscillating functions — open problems. Applied Numerical Mathematics: Transactions of IMACS, 15(2):259– 283, September 13, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1994&volume=15&
                                        issue=2&aid=498. Innova-

tive methods in numerical analysis (Bressanone, 1992).

Marz:1995:LDA

[Mär95] Roswitha März. On linear differential-algebraic equations and linearizations. Applied Numerical Mathematics: Transactions of IMACS, 18(1–3):267–292, September 27, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=18&
issue=1-3&aid=606. Sev-

enth Conference on the Numerical Treatment of Differential Equations (Halle, 1994).

Markov:1999:IMA

[Mar99a] Svetoslav Markov. An iterative method for algebraic solution to interval equations. Applied Numerical Mathematics: Transactions of IMACS, 30(2– 3):225–239, June 10, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1999&volume=30&
issue=2-3&aid=977.

Maroni:1999:SCC

[Mar99b] P. Maroni. Semi-classical character and finite-type relations between polynomial sequences. Applied Numerical Mathematics: Transactions of IMACS, 31(3): 295–330, November 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/10/28/46/19/
                                        20/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/10/28/46/19/20/article.
                                        pdf.

Marz:2002:DAS

[Mär02] Roswitha März. Differential algebraic systems anew. Applied Numerical Mathematics: Transactions of IMACS, 42(1–3):315–335, August 2002. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Marek:2003:IAD

[Mar03] Ivo Marek. Iterative aggregation/disaggregation methods for computing some characteristics of Markov chains. II. Fast convergence. Applied Numerical Mathematics: Transactions of IMACS, 45(1):11–28, April 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Martin:2005:OSW

[Mar05] Véronique Martin. An optimized Schwarz waveform relaxation method for the unsteady convection diffusion equation in two dimensions. Applied Numerical Mathematics: Transactions of IMACS, 52(4): 401–428, March 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Marcia:2008:SSS

[Mar08] Roummel F. Marcia. On solving sparse symmetric linear systems whose definiteness is unknown. Applied Numerical Mathematics: Transactions of IMACS, 58(4):449–458, April 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Marcozzi:2009:VIR

[Mar09] Michael D. Marcozzi. On the valuation of interest rate products under multi-factor HJM term-structures. Applied Numerical Mathematics: Transactions of IMACS, 59(12):2873–2890, December 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Mattheij:1986:CSS

[Mat86] R. M. M. Mattheij. On computing smooth solutions of problems with large Lipschitz constants. Applied Numerical Mathematics: Transactions of IMACS, 2(2):119–134, April 1986. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Matos:1991:CNT

[Mat91] Ana C. Matos. Construction of new transformations for lacunary power series based on the Cauchy-type approximants. Applied Numerical Mathematics: Transactions of IMACS, 7(6):493– 507, June 1991. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Matsekh:2005:GII

[Mat05] Anna M. Matsekh. The Godunov-inverse iteration: a fast and accurate solution to the symmetric tridiagonal eigenvalue problem. Applied Numerical Mathematics: Transactions of IMACS, 54(2):208–221, July 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Mattheij:2008:NVP

[Mat08] R. M. M. Mattheij. Numerical volume preservation of a divergence free fluid under symmetry. Applied Numerical Mathematics: Transactions of IMACS, 58(12): 1773–1788, December 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Matthies:2009:LPM

[Mat09]  Gunar Matthies.  Lo-

cal projection methods on layer-adapted meshes for higher order discretisations of convection–diffusion problems. Applied Numerical Mathematics: Transactions of IMACS, 59(10): 2515–2533, October 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Maubach:2008:DDF

[Mau08] J. Maubach. Dirichlet degrees of freedom need not be eliminated. Applied Numerical Mathematics: Transactions of IMACS, 58(12): 1852–1860, December 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Maes:2008:SAB

[MB08] Jan Maes and Adhemar Bultheel. Stability analysis of biorthogonal multiwavelets whose duals are not in L_{2} and its application to local semiorthogonal lifting. Applied Numerical Mathematics: Transactions of IMACS, 58(8): 1186–1211, August 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Maes:2010:PSS

[MB10] Jan Maes and Adhemar Bultheel. Powell–Sabin spline based multilevel preconditioners for the biharmonic equation. Applied Numerical Mathematics: Transactions of IMACS, 60(5):527–530, May 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ma:2020:SPO

[MB20] Ru-Ru Ma and Zheng-Jian Bai. A structurepreserving one-sided Jacobi method for computing the SVD of a quaternion matrix. Applied Numerical Mathematics: Transactions of IMACS, 147(??): 101–117, January 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419302260.

Ma:2023:IAP

[MBS23] Yuxue Ma, Jianchao Bai, and Hao Sun. An inexact ADMM with proximalindefinite term and larger stepsize. Applied Numerical Mathematics: Transactions of IMACS, 184(??): 542–566, February 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422002859.

Montagnac:2000:DCB

[MC00] Marc Montagnac and Jean-Marie Chesneaux. Dynamic control of a BICGSTAB algorithm. Applied Numerical Mathematics: Transactions of IMACS, 32(1):103– 117, January 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/59/28/
33/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/59/28/33/article.
pdf.

McKee:2017:NVP

[MC17] S. McKee and Jose A. Cuminato. A novel variant of a product integration method and its relation to discrete fractional calculus. Applied Numerical Mathematics: Transactions of IMACS, 114(??):179–187, April 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301908.

Mu:2021:NWW

[MC21] Lin Mu and Zheng Chen. A new WENO weak Galerkin finite element method for time dependent hyperbolic equations. Applied Numerical Mathematics: Transactions of IMACS, 159(??): 106–124, January 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420302634.

Messina:2020:NAD

[MCBV20] Eleonora Messina, Eugenio Chioccarelli, Georgios Baltzopoulos, and Antonia Vecchio. Numerical analysis of the dynamics of rigid blocks subjected to support excitation. Applied Numerical Mathematics: Transactions of IMACS, 155(??): 29–37, ???? 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419301345.

Martiradonna:2020:GGC

[MCD20] Angela Martiradonna, Gianpiero Colonna, and Fasma Diele. GeCo: Geometric Conservative nonstandard schemes for biochemical systems. Applied Numerical Mathematics: Transactions of IMACS, 155(??): 38–57, ???? 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419303368.

Montenegro:2009:ASA

[MCE^{+}09] R. Montenegro, J. M. Cascón, J. M. Escobar, E. Rodríguez, and G. Montero. An automatic strategy for adaptive tetrahedral mesh generation. Applied Numerical Mathematics: Transactions of IMACS, 59(9):2203–2217, September 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Muhieddine:2012:HTM

[MCM12] Mohamad Muhieddine, Édouard Canot, and Ramiro March. Heat transfer modeling in saturated porous media and identification of the thermophysical properties of the soil by inverse problem. Applied Numerical Mathematics: Transactions of IMACS, 62(9): 1026–1040, September 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412000736.

Mao:2006:QAS

[MCS06] Shipeng Mao, Shaochun Chen, and Huixia Sun. A quadrilateral, anisotropic, superconvergent, nonconforming double set parameter element. Applied Numerical Mathematics: Transactions of IMACS, 56(7): 937–961, July 2006. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Mao:2016:EAS

[MCS16] Zhiping Mao, Sheng Chen, and Jie Shen. Efficient and accurate spectral method using generalized Jacobi functions for solving Riesz fractional differential equations. Applied Numerical Mathematics: Transactions of IMACS, 106(??): 165–181, August 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741630040X.

Moore:1996:CPS

[MD96] Peter K. Moore and Robert H. Dillon. A comparison of preconditioners in the solution of parabolic systems in three space dimensions using DASPK and a high order finite element method. Applied Numerical Mathematics: Transactions of IMACS, 20(1–2): 117–128, February 29, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=1-2&aid=668. Work-

shop on the method of lines for time-dependent problems (Lexington, KY, 1995).

Magere:2000:STC

[MD00] E. Magère and M. O. Deville. Simulation of the Taylor– Couette flow in a finite geometry by spectral element method. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4): 241–249, May 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/61/27/
50/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/61/27/50/article.
pdf.

Maire:2006:QMC

[MD06] Sylvain Maire and Christophe De Luigi. Quasi-Monte Carlo quadratures for multivariate smooth functions. Applied Numerical Mathematics: Transactions of IMACS, 56(2):146–162, February 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Mirzaei:2010:MBM

[MD10] Davoud Mirzaei and Mehdi Dehghan. A meshless based method for solution of integral equations. Applied Numerical Mathematics: Transactions of IMACS, 60(3):245–262, March 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Macias-Diaz:2019:NSP

[MD19a]  J. E. Macías-Díaz.  On

the numerical and structural properties of a logarithmic scheme for diffusionreaction equations. Applied Numerical Mathematics: Transactions of IMACS, 140(??):104–114, June 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419300261.

Maleki:2019:FRD

[MD19b] Mohammad Maleki and Ali Davari. Fractional retarded differential equations and their numerical solution via a multistep collocation method. Applied Numerical Mathematics: Transactions of IMACS, 143(??): 203–222, September 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419300960.

Macias-Diaz:2020:ESE

[MD20a] J. E. Macías-Díaz. Existence of solutions of an explicit energy-conserving scheme for a fractional Klein–Gordon– Zakharov system. Applied Numerical Mathematics: Transactions of IMACS, 151(??):40–43, May 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419303721.

Macias-Diaz:2020:FEV

[MD20b] J. E. Macías-Díaz. A fully explicit variational integrator for multidimensional systems of coupled nonlinear fractional hyperbolic equations. Applied Numerical Mathematics: Transactions of IMACS, 154(??): 149–171, August 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300647.

Mohammadi:2020:MTB

[MD20c] Vahid Mohammadi and Mehdi Dehghan. A meshless technique based on generalized moving least squares combined with the secondorder semi-implicit backward differential formula for numerically solving timedependent phase field models on the spheres. Applied Numerical Mathematics: Transactions of IMACS, 153(??):248–275, July 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300581.

Mohammadi:2021:DFG

[MD21] Vahid Mohammadi and Mehdi Dehghan. A divergencefree generalized moving least squares approximation with its application. Applied Numerical Mathematics: Transactions of IMACS, 162(??):374–404, April 2021.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420303901.

Macias-Diaz:2022:DMD

[MD22] J. E. Macías-Díaz. On a discrete model that dissipates the free energy of a time-space fractional generalized nonlinear parabolic equation. Applied Numerical Mathematics: Transactions of IMACS, 172(??): 215–223, February 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421002944.

Mohammad:2023:SSA

[MD23a] Hassan Mohammad and Kamaluddeen Umar Danmalam. Structured spectral algorithm with a nonmonotone line search for nonlinear least squares. Applied Numerical Mathematics: Transactions of IMACS, 184(??):285–300, February 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422002781.

Mohammadi:2023:EED

[MD23b] Vahid Mohammadi and Mehdi Dehghan. Error estimates of divergence-free generalized moving least squares (Div-Free GMLS) derivatives approximations in Sobolev spaces. Applied Numerical Mathematics: Transactions of IMACS, 192(??): 373–388, October 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001824.

Mohapatra:2025:NEA

[MD25] Achyuta Ranjan Dutta Mohapatra and Bhupen Deka. A new error analysis of an explicit skeletal discontinuous Galerkin scheme for timedependent Maxwell equations. Applied Numerical Mathematics: Transactions of IMACS, 217(??):18– 42, November 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425001175.

MohammadiArani:2024:SDK

[MDA24] Reza MohammadiArani, Mehdi Dehghan, and Mostafa Abbaszadeh. Solving 2D damped Kuramoto–Sivashinsky with multiple relaxation time lattice Boltzmann method. Applied Numerical Mathematics: Transactions of IMACS, 196(??):83– 103, February 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423002593.

Momani:2021:NIC

[MDASAO21] Shaher Momani, Nadir Djeddi, Mohammed Al-Smadi, and Shrideh Al-Omari. Numerical investigation for Caputo– Fabrizio fractional Riccati and Bernoulli equations using iterative reproducing kernel method. Applied Numerical Mathematics: Transactions of IMACS, 170(??): 418–434, December 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421002336.

Mounim:2004:NMF

[MdD04] Abdellatif Serghini Mounim and Bernard M. de Dormale. A note on Mickens’ finite-difference scheme for the Lotka–Volterra system. Applied Numerical Mathematics: Transactions of IMACS, 51(2–3):341–344, November 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ma:2014:MSD

[MDD14] Qiang Ma, Deqiong Ding, and Xiaohua Ding. Meansquare dissipativity of several numerical methods for stochastic differential equations with jumps. Applied Numerical Mathematics: Transactions of IMACS, 82(??):44–50, August 2014. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414000464.

Morshed:2018:MMN

[MDD18] Adnan Morshed, Prashanta Dutta, and Robert H. Dillon. Mathematical modeling and numerical simulation of the TGF-β /Smad signaling pathway in tumor microenvironments. Applied Numerical Mathematics: Transactions of IMACS, 133(??):41–51, ???? 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417302398.

Malinen:2006:FFC

[MDHK06] Matti Malinen, Stephen R.

Duncan, Tomi Huttunen, and Jari P. Kaipio. Feedforward and feedback control of ultrasound surgery. Applied Numerical Mathematics: Transactions of IMACS, 56(1):55–79, January 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Macias-Diaz:2010:BPF

[MDP10] J. E. Macías-Díaz and A. Puri. A boundednesspreserving finite-difference scheme for a damped nonlinear wave equation. Applied Numerical Mathematics: Transactions of IMACS, 60(9):934–948, September 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Moradi:2023:VSM

[MDP23] A. Moradi, R. D’Ambrosio, and B. Paternoster. Variable stepsize multivalue collocation methods. Applied Numerical Mathematics: Transactions of IMACS, 190(??):1–14, August 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423000697.

Matos:2005:FPA

[MdR05] José M. A. Matos and Zélia da Rocha. Frobenius– Padé approximants for dorthogonal series: Theory and computational aspects. Applied Numerical Mathematics: Transactions of IMACS, 52(1):89–112, January 2005. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Macias-Diaz:2011:NSS

[MDRR11] J. E. Macías-Díaz and J. Ruiz-Ramírez. A nonstandard symmetry-preserving method to compute bounded solutions of a generalized Newell–Whitehead–Segel equation. Applied Numerical Mathematics: Transactions of IMACS, 61(4):630–640, April 2011. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Mandel:2005:ATP

[MDT05] Jan Mandel, Clark R.

Dohrmann, and Radek Tezaur. An algebraic theory for primal and dual substructuring methods by constraints. Applied Numerical Mathematics: Transactions of IMACS, 54(2): 167–193, July 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Medjo:1996:NSE

[Med96] Theodore Tachim Medjo.

The Navier–Stokes equations in the vorticity-velocity formulation: The twodimensional case. Applied Numerical Mathematics: Transactions of IMACS, 21(2):185–206, July 8, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=21&
issue=2&aid=681.

Muniz:2023:SSR

[MEGW23] Michelle Muniz, Matthias Ehrhardt, Michael Günther, and Renate Winkler. Strong stochastic Runge–Kutta– Munthe–Kaas methods for nonlinear Itô SDEs on manifolds. Applied Numerical Mathematics: Transactions of IMACS, 193(??): 196–203, November 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423002076.

Mehne:2008:SCS

[Meh08] H. H. Mehne. On solving constrained shape optimization problems for finding the optimum shape of a bar cross-section. Applied Numerical Mathematics: Transactions of IMACS, 58(8):1129–1141, August 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Mehrez:2022:NSF

[Meh22] Khaled Mehrez. New summation formulas of Fox– Wright-type series containing the polygamma functions. Applied Numerical Mathematics: Transactions of IMACS, 173(??): 158–179, March 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421003330.

Meng:2023:VEM

[Men23] Jian Meng. Virtual element method for the modified transmission eigenvalue problem in inverse scattering theory. Applied Numerical Mathematics: Transactions of IMACS, 192(??): 356–372, October 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001915.

Meurant:1991:DDM

[Meu91] Gérard A. Meurant. A domain decomposition method for parabolic problems. Applied Numerical Mathematics: Transactions of IMACS, 8(4–5):427–441, November 1991. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Meurant:2014:NSC

[Meu14]  Gérard Meurant.  Nec-

essary and sufficient conditions for GMRES complete and partial stagnation. Applied Numerical Mathematics: Transactions of IMACS, 75(??):100– 107, January 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413000846.

Martin:1999:NIP

[MF99] Pablo Martín and José M. Ferrándiz. Numerical integration of perturbed linear systems. Applied Numerical Mathematics: Transactions of IMACS, 31(2):183– 189, October 1999. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927498001263.

Min:2023:SMF

[MF23] Ya Min and Minfu Feng. Stabilized mixed finite element method for a quasistatic Maxwell viscoelastic model. Applied Numerical Mathematics: Transactions of IMACS, 193(??):22– 42, November 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001940.

Mohammadi-Firouzjaei:2023:SBD

[MFAD23] Hadi Mohammadi-Firouzjaei, Hojatollah Adibi, and Mehdi Dehghan. Study of the backward difference and local discontinuous Galerkin (LDG) methods for solving fourth-order partial integrodifferential equations (PI-DEs) with memory terms: Stability analysis. Applied Numerical Mathematics: Transactions of IMACS, 184(??):567–580, February 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422002914.

McLachlan:1997:OSP

[MG97] Robert I. McLachlan and Stephen K. Gray. Optimal stability polynomials for splitting methods, with application to the timedependent Schrödinger equation. Applied Numerical Mathematics: Transactions of IMACS, 25(2–3):275–286, October 17, 1997. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=25&
issue=2-3&aid=820.

Manservisi:2000:VIF

[MG00] S. Manservisi and M. Gunzburger. A variational inequality formulation of an inverse elasticity problem. Applied Numerical Mathematics: Transactions of IMACS, 34(1):99–126, June 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/62/27/
33/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/62/27/33/article.
pdf.

Mezzadri:2018:LDM

[MG18] Francesco Mezzadri and Emanuele Galligani. A lagged diffusivity method for reaction–convection–diffusion equations with Dirichlet boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 123(??):300–319, January 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417302039.

Mohanty:2022:HRO

[MG22] R. K. Mohanty and Bishnu Pada Ghosh. High resolution operator compact implicit half-step approximation for 3D quasi-linear hyperbolic equations and ADI method for 3D telegraphic equation on an irrational domain. Applied Numerical Mathematics: Transactions of IMACS, 172(??): 446–474, February 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421003019.

Macaraeg:1989:IAM

[MH89] M. G. Macaraeg and M. Y. Hussaini, editors. 12th International Association for Mathematics and Computation in Simulation (IMACS)

World Congress on Scientific Computation, volume 6(1–2) of Applied Numerical Mathematics: Transactions of IMACS. North-Holland Publishing Co., Amsterdam, The Netherlands, December 1989. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Melbo:2004:NSL

[MH04] Aslaug H. Strømmen Melbø and Desmond J. Higham. Numerical simulation of a linear stochastic oscillator with additive noise. Applied Numerical Mathematics: Transactions of IMACS, 51(1):89–99, October 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Maier:2014:DNR

[MH14] I. Maier and B. Haasdonk. A Dirichlet–Neumann reduced basis method for homogeneous domain decomposition problems. Applied Numerical Mathematics: Transactions of IMACS, 78(??): 31–48, April 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413001657.

Ma:2016:MMB

[MH16a] Changfeng Ma and Na Huang.

Modified modulus-based matrix splitting algorithms for a class of weakly nondifferentiable nonlinear complementarity problems. Applied Numerical Mathematics: Transactions of IMACS, 108(??):116–124, October 2016. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416300721.

Mohapatra:2016:LSS

[MH16b] Subhashree Mohapatra and Akhlaq Husain. Leastsquares spectral element method for three dimensional Stoke’s equations. Applied Numerical Mathematics: Transactions of IMACS, 102(??):31–54, April 2016. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415001713.

Mahdi:2019:NSG

[MHA19] H. Mahdi, G. Hojjati, and A. Abdi. On the numerical stability of the general linear methods for Volterra integro-differential equations. Applied Numerical Mathematics: Transactions of IMACS, 142(??): 139–150, August 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741930056X.

Mollapourasl:2018:LKB

[MHL18] Reza Mollapourasl, Majid Haghi, and Ruihua Liu. Localized kernel-based approximation for pricing financial options under regime switching jump diffusion model. Applied Numerical Mathematics: Transactions of IMACS, 134(??):81– 104, December 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418301569.

Makhanov:2003:GGA

[MI03] S. S. Makhanov and S. A.

Ivanenko. Grid generation as applied to optimize cutting operations of the five-axis milling machine. Applied Numerical Mathematics: Transactions of IMACS, 46(3–4): 331–351, September 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Michielse:1995:PMU

[Mic95]  Peter Michielse.  Paral-

lel multigrid using PVM. Applied Numerical Mathematics: Transactions of IMACS, 19(1–2):63–69, December 4, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=19&
                                        issue=1-2&aid=566.

Mickens:2003:NFD

[Mic03]  Ronald E. Mickens.  A

nonstandard finite-difference scheme for the Lotka– Volterra system. Applied Numerical Mathematics: Transactions of IMACS, 45(2–3):309–314, May 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Miel:1989:PSF

[Mie89] George Miel. Parallel solution of Fredholm integral equations of the second kind by orthogonal polynomial expansions. Applied Numerical Mathematics: Transactions of IMACS, 5(4):345– 361, July 1989. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Mehdiyeva:2013:OAF

[MII13] G. Yu. Mehdiyeva, M. N. Imanova, and V. R. Ibrahimov. On one application of forward jumping methods. Applied Numerical Mathematics: Transactions of IMACS, 72(??):234– 245, October 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741100184X.

Mikula:1997:SNC

[Mik97] Karol Mikula. Solution of nonlinear curvature driven evolution of plane convex curves. Applied Numerical Mathematics: Transactions of IMACS, 23(3):347– 360, April 25, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1997/23/
3/754.pdf.

Milovanovic:2017:GWB

[Mil17] Gradimir V. Milovanović.

Generalized weighted Birkhoff– Young quadratures with the maximal degree of exactness. Applied Numerical Mathematics: Transactions of IMACS, 116(??): 238–255, June 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301064.

Minkoff:1987:AOS

[Min87] Michael Minkoff. Approaches to optimization/simulation problems. Applied Numerical Mathematics: Transactions of IMACS, 3(5):453– 466, October 1987. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Minion:2004:SIP

[Min04] Michael L. Minion. Semiimplicit projection methods for incompressible flow based on spectral deferred corrections. Applied Numerical Mathematics: Transactions of IMACS, 48(3–4): 369–387, March 2004. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Mirzaei:2020:AKC

[Mir20] Davoud Mirzaei. On analysis of kernel collocation methods for spherical PDEs. Applied Numerical Mathematics: Transactions of IMACS, 150(??):222–232, April 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419302685.

Mitchell:1997:FDP

[Mit97] William F. Mitchell. The full domain partition approach to distributing adaptive grids. Applied Numerical Mathematics: Transactions of IMACS, 26(1–2):265–275, December 22, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1997/26/
1-2/851.pdf.

Mittal:2022:EAA

[Mit22] Avinash Kumar Mittal. Error analysis and approximation of Jacobi pseudospectral method for the integer and fractional order integro-differential equation. Applied Numerical Mathematics: Transactions of IMACS, 171(??): 249–268, January 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421002555.

Mitsotakis:2024:IMB

[Mit24a] Dimitrios Mitsotakis. On iterative methods based on Sherman–Morrison–Woodbury splitting. Applied Numerical Mathematics: Transactions of IMACS, 201(??): 282–289, July 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000564.

Mittal:2024:STP

[Mit24b] A. K. Mittal. A spacetime pseudospectral method for solving multi-dimensional quasi-linear parabolic partial differential (Burgers’) equations. Applied Numerical Mathematics: Transactions of IMACS, 195(??): 39–53, January 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423002441.

Mora:2023:WVS

[MJS23] Carlos M. Mora, Juan Carlos Jimenez, and Monica Selva. Weak variable stepsize schemes for stochastic differential equations based on controlling conditional moments. Applied Numerical Mathematics: Transactions of IMACS, 187(??): 235–261, May 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423000399.

Munthe-Kaas:1999:HOR

[MK99] Hans Munthe-Kaas. High order Runge–Kutta methods on manifolds. Applied Numerical Mathematics: Transactions of IMACS, 29(1):115–127, January 1, 1999. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1999/29/
                                        1/896.pdf.

Mahmood:2014:SDN

[MK14] Mohammed Shuker Mahmood and Karel Kovárik. Solution of double nonlinear problems in porous media by a combined finite volumefinite element algorithm. Applied Numerical Mathematics: Transactions of IMACS, 82(??):11–31, August 2014. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414000385.

Mtagulwa:2019:EMP

[MK19] Peter Mtagulwa and P. Kaelo.

An efficient modified PRP-FR hybrid conjugate gradient method for solving unconstrained optimization problems. Applied Numerical Mathematics: Transactions of IMACS, 145(??): 111–120, November 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927419301552.

Ma:2020:FEC

[MK20] Junjie Ma and Hongchao Kang. Frequency-explicit convergence analysis of collocation methods for highly oscillatory Volterra integral equations with weak singularities. Applied Numerical Mathematics: Transactions of IMACS, 151(??): 1–12, May 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419303459.

Maleknejad:2021:ASS

[MK21] Khosrow Maleknejad and Hamed Shahi Kalalagh. Approximate solution of some nonlinear classes of Abel integral equations using hybrid expansion. Applied Numerical Mathematics: Transactions of IMACS, 159(??): 61–72, January 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420302051.

Manohar:2026:ASM

[MKBS26] Ram Manohar, B. V. Rathish Kumar, Kedarnath Buda, and Rajen Kumar Sinha. Adaptive SIPG method for approximations of parabolic boundary control problems with bilateral box constraints on Neumann boundary. Applied Numerical Mathematics: Transactions of IMACS, 219(??):170– 201, January 2026. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425001588.

Malav:2025:EMP

[MKDC25] Krishna Murari Malav, Kapil Kant, Joydip Dhar, and Samiran Chakraborty. An efficient multi projection methods for systems of Fredholm integral equations with mixed weakly singular kernels: a superconvergence approach. Applied Numerical Mathematics: Transactions of IMACS, 217(??): 172–189, November 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425001230.

Mesri:2016:OSM

[MKH16] Y. Mesri, M. Khalloufi, and E. Hachem. On optimal simplicial 3D meshes for minimizing the Hessian-based errors. Applied Numerical Mathematics: Transactions of IMACS, 109(??): 235–249, November 2016. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301337.

Molabahrami:2023:DCM

[MKJ23] Ahmad Molabahrami, Boualem Khouider, and Mohsen Jalalian. A discrete collocation method based on the radial basis functions for solving system of integral equations of the second kind. Applied Numerical Mathematics: Transactions of IMACS, 185(??): 278–294, March 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422002999.

Mandal:2023:PMA

[MKN23] Moumita Mandal, Arnab Kayal, and Gnaneshwar Nelakanti. Projection methods for approximate solution of a class of nonlinear Fredholm integro-differential equations. Applied Numerical Mathematics: Transactions of IMACS, 184(??):49– 76, February 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422002598.

Munthe-Kaas:2012:MPS

[MKS12] Hans Munthe-Kaas and Tor Sørevik. Multidimensional pseudo-spectral methods on lattice grids. Applied Numerical Mathematics: Transactions of IMACS, 62(3): 155–165, March 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927411002042.

Mandel:1991:DDP

[ML91] Jan Mandel and G. Scott Lett. Domain decomposition preconditioning for p-version finite elements with high aspect ratios. Applied Numerical Mathematics: Transactions of IMACS, 8(4–5): 411–425, November 1991.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Manapova:2016:NSO

[ML16] A. R. Manapova and F. V.

Lubyshev. Numerical solution of optimization problems for semi-linear elliptic equations with discontinuous coefficients and solutions. Applied Numerical Mathematics: Transactions of IMACS, 104(??): 182–203, June 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741600026X.

Morlet:1997:SAS

[MLB97] Anne C. Morlet, Nancy J.

Lybeck, and Kenneth L. Bowers. The Schwarz alternating sinc domain decomposition method. Applied Numerical Mathematics: Transactions of IMACS, 25(4):461–483, November 10, 1997. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=25&
issue=4&aid=824.

Moon:2019:MHM

[MLJ19] Minam Moon, Raytcho Lazarov, and Hyung Kyu Jun. Multiscale HDG model reduction method for flows in heterogeneous porous media. Applied Numerical Mathematics: Transactions of IMACS, 140(??): 115–133, June 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419300224.

Makroglou:2006:MMS

[MLK06] Athena Makroglou, Jiaxu Li, and Yang Kuang. Mathematical models and software tools for the glucoseinsulin regulatory system and diabetes: an overview. Applied Numerical Mathematics: Transactions of IMACS, 56(3–4):559–573, March/April 2006. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Maroni:2002:QCO

[MM02a] P. Maroni and M. Mejri. The I ( q, ) classical orthogonal polynomials. Applied Numerical Mathematics: Transactions of IMACS, 43(4):423–458, December 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Meddahi:2002:CSF

[MM02b] Salim Meddahi and Antonio Márquez. A combination of spectral and finite elements for an exterior problem in the plane. Applied Numerical Mathematics: Transactions of IMACS, 43(3): 275–295, November 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Miranville:2002:BIU

[MM02c] A. Miranville and A. C.

Muresan. Block incremental unknowns for anisotropic elliptic equations. Applied Numerical Mathematics: Transactions of IMACS, 42(4):529–543, September 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Meseguer:2007:SFC

[MM07] A. Meseguer and F. Mellibovsky. On a solenoidal Fourier–Chebyshev spectral method for stability analysis of the Hagen–Poiseuille flow. Applied Numerical Mathematics: Transactions of IMACS, 57(8):920–938, August 2007. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Marti:2014:STI

[MM14] M. C. Martí and P. Mulet.

Some techniques for improving the resolution of finite difference componentwise WENO schemes for polydisperse sedimentation models. Applied Numerical Mathematics: Transactions of IMACS, 78(??): 1–13, April 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927413001633.

Michailidis:2016:SCM

[MM16] Panagiotis D. Michailidis and Konstantinos G. Margaritis. Scientific computations on multi-core systems using different programming frameworks. Applied Numerical Mathematics: Transactions of IMACS, 104(??):62–80, June 2016. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892741400213X.

Ma:2018:PMR

[MM18] Yunyun Ma and Fuming Ma. A projection method with regularization for Cauchy problem of the time-harmonic Maxwell equations. Applied Numerical Mathematics: Transactions of IMACS, 129(??): 71–82, July 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892741830059X.

Mehandiratta:2020:DST

[MM20a] Vaibhav Mehandiratta and Mani Mehra. A difference scheme for the timefractional diffusion equation on a metric star graph. Applied Numerical Mathematics: Transactions of IMACS, 158(??): 152–163, December 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302221.

Murphy:2020:MGF

[MM20b] Laura Murphy and Anotida Madzvamuse. A moving grid finite element method applied to a mechanobiochemical model for 3D cell migration. Applied Numerical Mathematics: Transactions of IMACS, 158(??): 336–359, December 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302294.

Manzini:2022:CVE

[MM22] Gianmarco Manzini and Annamaria Mazzia. Conforming virtual element approximations of the twodimensional Stokes problem. Applied Numerical Mathematics: Transactions of IMACS, 181(??): 176–203, November 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422001520.

Monch:2025:CBO

[MM25] Marius Mönch and Nicole Marheineke. Commutatorbased operator splitting for linear port-Hamiltonian systems. Applied Numerical Mathematics: Transactions of IMACS, 210(??): 25–38, April 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424003489.

Martin:2007:CSP

[MMBB07] Dian I. Martin, John C. Martin, Michael W. Berry, and Murray Browne. Out-of-core SVD performance for document indexing. Applied Numerical Mathematics: Transactions of IMACS, 57(11–12):1230– 1239, November/December 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Martinez:2020:EPE

[MMD20] Romeo Martínez and Jorge E. Macías-Díaz. An energypreserving and efficient scheme for a double-fractional conservative Klein–Gordon– Zakharov system. Applied Numerical Mathematics: Transactions of IMACS, 158(??):292–313, December 2020. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302373.

Martinez:2019:TAE

[MMDH19] Romeo Martínez, J. E.

Macías-Díaz, and Ahmed S. Hendy. Theoretical analysis of an explicit energyconserving scheme for a fractional Klein–Gordon– Zakharov system. Applied Numerical Mathematics: Transactions of IMACS, 146(??):245–259, December 2019. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927419301709.

Martinez:2021:ISL

[MMDS21] Romeo Martínez, Jorge E.

Macías-Díaz, and Qin Sheng. An implicit semi-linear discretization of a bi-fractional Klein–Gordon–Zakharov system which conserves the total energy. Applied Numerical Mathematics: Transactions of IMACS, 169(??): 179–200, November 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001823.

Mahara:2017:IAM

[MMKN17] Hitoshi Mahara, Koshiro Mizobe, Katsuyuki Kida, and Kazuaki Nakane. Image analyzing method to detect vague boundaries by using reaction–diffusion system. Applied Numerical Mathematics: Transactions of IMACS, 114(??): 124–131, April 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301660.

Moghaddam:2019:NAC

[MMM19] B. P. Moghaddam, J. A. Tenreiro Machado, and M. L. Morgado. Numerical approach for a class of distributed order time fractional partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 136(??): 152–162, February 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418302319.

Mendez:2002:TVD

[MMP02a] E. Méndez, A. Müller, and M. Paluszny. Three views of Dupin cyclides and blending of cones. Applied Numerical Mathematics: Transactions of IMACS, 40(1–2):39– 47, January 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.

elsevier.com/gej-ng/10/
10/28/85/27/31/abstract.
html.

Mendez:2002:TJC

[MMP02b] E. Mendez, A. Müller, and M. Paluszny. Tubelike joints: a classical geometry perspective. Applied Numerical Mathematics: Transactions of IMACS, 40(1–2):33– 38, January 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/85/27/30/abstract.
html.

Macconi:2009:TRQ

[MMP09] Maria Macconi, Benedetta Morini, and Margherita Porcelli. Trust-region quadratic methods for nonlinear systems of mixed equalities and inequalities. Applied Numerical Mathematics: Transactions of IMACS, 59(5): 859–876, May 2009. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Monteghetti:2020:TLD

[MMP20] Florian Monteghetti, Denis Matignon, and Estelle Piot. Time-local discretization of fractional and related diffusive operators using Gaussian quadrature with applications. Applied Numerical Mathematics: Transactions of IMACS, 155(??):

73–92, ???? 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418302666.

Matos:2020:ASI

[MMRV20] João C. Matos, José A. Matos, Maria João Rodrigues, and Paulo B. Vasconcelos. Approximating the solution of integrodifferential problems via the spectral Tau method with filtering. Applied Numerical Mathematics: Transactions of IMACS, 149(??): 164–175, ???? 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419301400.

McCormick:1990:CCF

[MMT90] Steve F. McCormick, Steven M. McKay, and J. W. Thomas. Computational complexity of the fast adaptive composite grid (FAC) method. Applied Numerical Mathematics: Transactions of IMACS, 6(4):315–327, May 1990. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Maroni:2003:IPP

[MN03] P. Maroni and I. Nicolau. On the inverse problem of the product of a form by a polynomial: The cubic case. Applied Numerical Mathematics: Transactions of IMACS, 45(4):419–451, June 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Moret:2008:RKM

[MN08] I. Moret and P. Novati.

A rational Krylov method for solving time-periodic differential equations. Applied Numerical Mathematics: Transactions of IMACS, 58(3):212–222, March 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Mastroianni:2020:PAP

[MN20] Giuseppe Mastroianni and Incoronata Notarangelo. Polynomial approximation with Pollaczek-type weights. A survey. Applied Numerical Mathematics: Transactions of IMACS, 149(??): 83–98, ???? 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419301618.

Maji:2023:ANS

[MN23] Sandip Maji and Srinivasan Natesan. Analytical and numerical solutions of time-fractional advection-diffusion-reaction equation. Applied Numerical Mathematics: Transactions of IMACS, 185(??):

549–570, March 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742200335X.

Meliani:2024:MAN

[MN24] Mostafa Meliani and Vanja Nikolić. Mixed approximation of nonlinear acoustic equations: Well-posedness and a priori error analysis. Applied Numerical Mathematics: Transactions of IMACS, 198(??): 94–111, April 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742300301X.

Marion:2014:NSS

[MNR14] P. Marion, K. Najib, and C. Rosier. Numerical simulations for a seawater intrusion problem in a free aquifer. Applied Numerical Mathematics: Transactions of IMACS, 75(??):48–60, January 2014. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927412001985.

Ma:2022:EHC

[MNSS22] Xiaoxia Ma, Hemant Kumar Nashine, Sourav Shil, and Fazlollah Soleymani. Exploiting higher computational efficiency index for computing outer generalized inverses. Applied Numerical Mathematics: Transactions of IMACS, 175(??): 18–28, May 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422000253.

Marthinsen:2001:QMB

[MO01] Arne Marthinsen and Brynjulf Owren. Quadrature methods based on the Cayley transform. Applied Numerical Mathematics: Transactions of IMACS, 39(3–4):403–413, December 2001. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/79/34/36/abstract.
html.

Maleknejad:2017:NSS

[MO17] K. Maleknejad and A. Ostadi. Numerical solution of system of Volterra integral equations with weakly singular kernels and its convergence analysis. Applied Numerical Mathematics: Transactions of IMACS, 115(??):82–98, May 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417300041.

Moebs:1998:MMR

[Moe98] Guy Moebs. A multilevel method for the resolution of a stochastic weakly damped nonlinear Schrödinger equation. Applied Numerical Mathematics: Transactions of IMACS, 26(3):353–375, March 2, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1998/26/
                                        3/831.pdf.

Mokhtary:2017:NAO

[Mok17]  P. Mokhtary.  Numeri-

cal analysis of an operational Jacobi Tau method for fractional weakly singular integro-differential equations. Applied Numerical Mathematics: Transactions of IMACS, 121(??):52– 67, November 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417301460.

Molenaar:1995:AMA

[Mol95] J. Molenaar. Adaptive multigrid applied to a bipolar transistor problem. Applied Numerical Mathematics: Transactions of IMACS, 17(1):61–83, May 1, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-

bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=17&
issue=1&aid=537.

Moniot:2009:DLS

[Mon09] Robert K. Moniot. Deming least-squares fit to multiple hyperplanes. Applied Numerical Mathematics: Transactions of IMACS, 59(1):135–150, January 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Montecinos:2021:ISF

[Mon21] Gino I. Montecinos. An iterative scaling function procedure for solving scalar non-linear hyperbolic balance laws. Applied Numerical Mathematics: Transactions of IMACS, 162(??): 35–52, April 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303822.

Moore:1995:GMC

[Moo95a] Gerald Moore. Geometric methods for computing invariant manifolds. Applied Numerical Mathematics: Transactions of IMACS, 17(3):319–331, September 4, 1995. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=17&
issue=3&aid=585. Numeri-

cal methods for ordinary differential equations (Atlanta, GA, 1994).

Moore:1995:CAM

[Moo95b] Peter K. Moore. Comparison of adaptive methods for one-dimensional parabolic systems. Applied Numerical Mathematics: Transactions of IMACS, 16(4):471– 488, April 6, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=16&
issue=4&aid=550.

Moorthy:1995:NIT

[Moo95c]  M. V. Moorthy.  Nu-

merical inversion of twodimensional Laplace transforms — Fourier series representation. Applied Numerical Mathematics: Transactions of IMACS, 17(2): 119–127, June 13, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=17&
issue=2&aid=563.

Moore:2004:IEB

[Moo04] Peter K. Moore. Interpolation error-based a posteriori error estimation for hprefinement using first and second derivative jumps. Applied Numerical Mathematics: Transactions of IMACS, 48(1):63–82, January 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Morgan:2005:RBG

[Mor05] Ronald B. Morgan. Restarted block-GMRES with deflation of eigenvalues. Applied Numerical Mathematics: Transactions of IMACS, 54(2):222–236, July 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

MacMullen:2002:CCS

[MOS02] H. MacMullen, E. O’Riordan, and G. I. Shishkin. The convergence of classical Schwarz methods applied to convection–diffusion problems with regular boundary layers. Applied Numerical Mathematics: Transactions of IMACS, 43(3): 297–313, November 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Maischak:2012:LSF

[MOS12] M. Maischak, S. Oestmann, and E. P. Stephan. A least-squares fem–bem coupling method for linear elasticity. Applied Numerical Mathematics: Transactions of IMACS, 62(4): 457–472, April 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411002133.

Miller:2000:PUS

[MOSW00] J. J. H. Miller, E. O’Riordan, G. I. Shishkin, and S. Wang. A parameter-uniform Schwarz method for a singularly perturbed reaction-diffusion problem with an interior layer. Applied Numerical Mathematics: Transactions of IMACS, 35(4): 323–337, December 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.nl../29/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/64/32/29/article.
pdf.

Motygin:2017:NAO

[Mot17] Oleg V. Motygin. Numerical approximation of oscillatory integrals of the linear ship wave theory. Applied Numerical Mathematics: Transactions of IMACS, 115(??):99–113, May 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741730003X.

Moussaoui:2003:UAI

[Mou03] Farid Moussaoui. A unified approach for inviscid compressible and nearly incompressible flow by leastsquares finite element method. Applied Numerical Mathematics: Transactions of IMACS, 44(1–2):183–199, January 2003. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Mikula:2014:IIO

[MOU14] Karol Mikula, Mario Ohlberger, and Jozef Urbán. Inflowimplicit/outflow-explicit finite volume methods for solving advection equations. Applied Numerical Mathematics: Transactions of IMACS, 85(??):16– 37, November 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414001032.

Mezzanotte:2024:ALM

[MOV24] Domenico Mezzanotte, Donatella Occorsio, and Ezio Venturino. Analysis of a line method for reactiondiffusion models of nonlocal type. Applied Numerical Mathematics: Transactions of IMACS, 203(??): 255–268, September 2024.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742400117X.

Melander:1987:CVD

[MOZ87] Mogens V. Melander, Edward A. Overman II, and Norman J. Zabusky. Computational vortex dynamics in two and three dimensions. Applied Numerical Mathematics: Transactions of IMACS, 3(1–2): 59–80, May 1987. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0168927487900067.

Monaghan:1985:AVP

[MP85] J. J. Monaghan and H. Pongracic. Artificial viscosity for particle methods. Applied Numerical Mathematics: Transactions of IMACS, 1(3):187–194, May 1985. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Moore:1994:SCS

[MP94] Peter K. Moore and Linda R. Petzold. A stepsize control strategy for stiff systems of ordinary differential equations. Applied Numerical Mathematics: Transactions of IMACS, 15(4): 449–463, November 1, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=15&
issue=4&aid=518.

Maly:1996:NMS

[MP96] Timothy Maly and Linda R.

Petzold. Numerical methods and software for sensitivity analysis of differentialalgebraic systems. Applied Numerical Mathematics: Transactions of IMACS, 20(1–2):57–79, February 29, 1996. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=1-2&aid=665. Work-

shop on the method of lines for time-dependent problems (Lexington, KY, 1995).

Melendo:1997:NAC

[MP97] Begoña Melendo and Manuel Palacios. A new approach to the construction of multirevolution methods and their implementation. Applied Numerical Mathematics: Transactions of IMACS, 23(2):259–274, March 21, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/

                                        store/apnum/sub/1997/23/
                                        2/737.pdf.

Martinez:1998:FAB

[MP98] J. J. Martínez and J. M. Peña. Fast algorithms of Björck–Pereyra type for solving Cauchy–Vandermonde linear systems. Applied Numerical Mathematics: Transactions of IMACS, 26(3): 343–352, March 2, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/apnum/sub/1998/
                                        26/3/858.pdf.

Mandel:2005:ISC

[MP05] Jan Mandel and Mirela O. Popa. Iterative solvers for coupled fluid–solid scattering. Applied Numerical Mathematics: Transactions of IMACS, 54(2):194–207, July 2005. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Malakuti:2011:SSF

[MP11] Kamyar Malakuti and Evgueni Parilov. A split-step finite difference method for nonparaxial nonlinear Schrödinger equation at critical dimension. Applied Numerical Mathematics: Transactions of IMACS, 61(7):891–899, July 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Mitrano:2015:NSD

[MP15] Arthur A. Mitrano and Rodrigo B. Platte. A numerical study of divergencefree kernel approximations. Applied Numerical Mathematics: Transactions of IMACS, 96(??):94–107, October 2015. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415000872.

Mascotto:2020:ENT

[MP20] Lorenzo Mascotto and Alexander Pichler. Extension of the nonconforming Trefftz virtual element method to the Helmholtz problem with piecewise constant wave number. Applied Numerical Mathematics: Transactions of IMACS, 155(??): 160–180, ???? 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419300868.

Mudheher:2024:HRC

[MPDB24] Aws Mushtaq Mudheher, S. Pishbin, P. Darania, and Shadi Malek Bagomghaleh. High-rate convergent multistep collocation techniques to a first-kind Volterra integral equation along with the proportional vanishing delay. Applied Numerical Mathematics: Transactions of IMACS, 204(??): 188–205, October 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001612.

Miyaji:2016:SRO

[MPG^{+}16] Tomoyuki Miyaji, Pawel Pilarczyk, Marcio Gameiro, Hiroshi Kokubu, and Konstantin Mischaikow. A study of rigorous ODE integrators for multi-scale setoriented computations. Applied Numerical Mathematics: Transactions of IMACS, 107(??):34–47, September 2016. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0168927416300435.

Martinez:2023:SAM

[MPHFP23] D. Martínez, F. Pla, H. Herrero, and A. Fernández-Pérez. A Schwarz alternating method for an evolution convection problem. Applied Numerical Mathematics: Transactions of IMACS, 192(??): 179–196, October 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001721.

Munoz-Perez:2021:ICM

[MPMD21] Luis F. Muñoz-Pérez and J. E. Macías-Díaz. An implicit and convergent method for radially symmetric solutions of Higgs’ boson equation in the de Sitter spacetime. Applied Numerical Mathematics: Transactions of IMACS, 165(??): 270–289, July 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000684.

Mauser:2022:UWP

[MPPR22] Norbert J. Mauser, Carl-Martin Pfeiler, Dirk Praetorius, and Michele Ruggeri. Unconditional wellposedness and IMEX improvement of a family of predictor-corrector methods in micromagnetics. Applied Numerical Mathematics: Transactions of IMACS, 180(??):33–54, October 2022. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001313.

Mantzavinos:2016:FTM

[MPSS16] D. Mantzavinos, M. G. Papadomanolaki, Y. G. Saridakis, and A. G. Sifalakis. Fokas transform method for a brain tumor invasion model with heterogeneous diffusion in 1 + 1 dimensions. Applied Numerical Mathematics: Transactions of IMACS, 104(??):47–61, June 2016. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414001597.

Mohaghegh:2016:MOR

[MPtM16] Kasra Mohaghegh, Roland Pulch, and Jan ter Maten. Model order reduction using singularly perturbed systems. Applied Numerical Mathematics: Transactions of IMACS, 103(??): 72–87, May 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892741600012X.

Magura:2017:HON

[MPTT17] S. Magura, S. Petropavlovsky, S. Tsynkov, and E. Turkel. High-order numerical solution of the Helmholtz equation for domains with reentrant corners. Applied Numerical Mathematics: Transactions of IMACS, 118(??):87–116, August 2017. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417300582.

Meletiou:2025:SRS

[MPTV25] Gerasimos C. Meletiou, Nikolaos K. Papadakis, Dimitrios S. Triantafyllou, and Michael N. Vrahatis. Structured ramp secret sharing schemata over rings of real polynomials. Applied Numerical Mathematics: Transactions of IMACS, 208 (part A)(??):317–339, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927424001430.

Messina:2024:LTB

[MPV24] Eleonora Messina, Mario Pezzella, and Antonia Vecchio. A long-time behavior preserving numerical scheme for age-of-infection epidemic models with heterogeneous mixing. Applied Numerical Mathematics: Transactions of IMACS, 200(??): 344–357, June 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001022.

McLachlan:2000:NIC

[MQ00] Robert I. McLachlan and G. R. W. Quispel. Numerical integrators that contract volume. Applied Numerical Mathematics: Transactions of IMACS, 34(2–3):

253–260, July 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/21/62/29/
                                        35/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/21/62/29/35/article.
                                        pdf.

McLachlan:2003:GIC

[MQ03] Robert I. McLachlan and G. R. W. Quispel. Geometric integration of conservative polynomial ODEs. Applied Numerical Mathematics: Transactions of IMACS, 45(4):411–418, June 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ma:2017:MLG

[MQO17] Heping Ma, Yonghui Qin, and Qiuli Ou. Multidomain Legendre–Galerkin Chebyshev-collocation method for one-dimensional evolution equations with discontinuity. Applied Numerical Mathematics: Transactions of IMACS, 111(??): 246–259, January 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416301775.

McFaddin:1992:CPS

[MR92] H. S. McFaddin and J. R. Rice. Collaborating PDE solvers. Applied Numerical Mathematics: Transactions of IMACS, 10(3–4): 279–295, 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

McCormick:1994:FVC

[MR94] Steve McCormick and Ulrich Rüde. A finite volume convergence theory for the fast adaptive composite grid methods. Applied Numerical Mathematics: Transactions of IMACS, 14(1– 3):91–103, April 14, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=14&
issue=1-3&aid=461. Pro-

ceedings of the Third ARO Workshop on Adaptive Methods for Partial Differential Equations (Troy, NY, 1992).

Mellado:2001:ESF

[MR01] Mario Mellado and Rodolfo Rodríguez. Efficient solution of fluid-structure vibration problems. Applied Numerical Mathematics: Transactions of IMACS, 36(4): 389–400, March 2001. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/65/32/

                                        27/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/10/28/65/32/27/article.
                                        pdf.

Merkle:2006:CDP

[MR06] C. Merkle and C. Rohde. Computation of dynamical phase transitions in solids. Applied Numerical Mathematics: Transactions of IMACS, 56(10–11):1450– 1463, October/November 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Maqbul:2020:TDS

[MR20] Md. Maqbul and A. Raheem. Time-discretization schema for a semilinear pseudo-parabolic equation with integral conditions. Applied Numerical Mathematics: Transactions of IMACS, 148(??):18–27, February 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419302417.

Mulet:2024:NSP

[MRÁSY24] Pep Mulet, Juan Ruiz-Álvarez, Chi-Wang Shu, and Dionisio F. Yáñez. A nonseparable progressive multivariate WENO-2r point value. Applied Numerical Mathematics: Transactions of IMACS, 204(??): 26–47, October 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001399.

Manzi:2000:FAT

[MRF00] Cristina Manzi, Francesca Rapetti, and Luca Formaggia. Function approximation on triangular grids: some numerical results using adaptive techniques. Applied Numerical Mathematics: Transactions of IMACS, 32(4):389–399, April 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/59/32/
30/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/59/32/30/article.
pdf.

Morgado:2017:NSD

[MRFF17] Maria Luísa Morgado, Magda Rebelo, Luis L. Ferrás, and Neville J. Ford. Numerical solution for diffusion equations with distributed order in time using a Chebyshev collocation method. Applied Numerical Mathematics: Transactions of IMACS, 114(??):108–123, April 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416302197.

Morikuni:2014:FLD

[MRH14] Keiichi Morikuni, Lothar Reichel, and Ken Hayami. FGMRES for linear discrete ill-posed problems. Applied Numerical Mathematics: Transactions of IMACS, 75(??):175–187, January 2014. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0168927413001050.

Manickam:2003:FFB

[MRS03] S. A. V. Manickam, Ch. Radha, and V. D. Sharma. Far field behaviour of waves in a vibrationally relaxing gas. Applied Numerical Mathematics: Transactions of IMACS, 45(2–3): 293–307, May 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Morigi:2010:CMM

[MRS10] S. Morigi, L. Reichel, and F. Sgallari. Cascadic multilevel methods for fast nonsymmetric blur- and noiseremoval. Applied Numerical Mathematics: Transactions of IMACS, 60(4): 378–396, April 2010. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Marinescu:1993:PIM

[MRV93] D. C. Marinescu, J. R. Rice, and E. A. Vavalis. Performance of iterative methods for distributed memory machines. Applied Numerical Mathematics: Transactions of IMACS, 12(5): 421–430, July 1993. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/016892749390102W.

Macaraeg:1986:ISC

[MS86] Michèle G. Macaraeg and Craig L. Streett. Improvements in spectral collocation discretization through a multiple domain technique. Applied Numerical Mathematics: Transactions of IMACS, 2(2):95–108, April 1986. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Mattheij:1990:SCL

[MS90] R. M. M. Mattheij and M. D. Smooke. Stability and convergence of linear parabolic mixed initialboundary value problems on nonuniform grids. Applied Numerical Mathematics: Transactions of IMACS, 6(6):471–485, October 1990. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Macaraeg:1991:LSH

[MS91] Michele G. Macaraeg and Craig L. Streett. Linear stability of high-speed mixing layers. Applied Numerical Mathematics: Transactions of IMACS, 7(1):93– 127, January 1991. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0168927491901059.

Meerbergen:1999:UKA

[MS99a] Karl Meerbergen and Miloud Sadkane. Using Krylov approximations to the matrix exponential operator in Davidson’s method. Applied Numerical Mathematics: Transactions of IMACS, 31(3):331–351, November 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/46/19/
21/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/46/19/21/article.
pdf.

Mikula:1999:SNC

[MŠ99b] Karol Mikula and Daniel Ševčovič. Solution of nonlinearly curvature driven evolution of plane curves. Applied Numerical Mathematics: Transactions of IMACS, 31(2):191–207, October 23, 1999. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-

bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1999&volume=31&
issue=2&aid=995;   http:
//www.sciencedirect.com/
science/article/pii/S0168927498001305.

Matache:2000:HFP

[MS00] A.-M. Matache and Ch.

Schwab. Homogenization via p-FEM for problems with microstructure. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4):43–59, May 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/61/27/
30/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/61/27/30/article.
pdf.

Mazzia:2002:NAN

[MS02] F. Mazzia and I. Sgura.

Numerical approximation of nonlinear BVPs by means of BVMs. Applied Numerical Mathematics: Transactions of IMACS, 42(1– 3):337–352, August 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Makhanov:2003:SNS

[MS03] Stanislav S. Makhanov and Andrei Yu. Semenov. Six numerical schemes for parabolic initial boundary value problems with a priori bounded solution. Applied Numerical Mathematics: Transactions of IMACS, 46(3–4): 353–377, September 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Maischak:2005:AH

[MS05] Matthias Maischak and Ernst P. Stephan. Adaptive hp. Applied Numerical Mathematics: Transactions of IMACS, 54(3– 4):425–449, August 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Mandal:2008:LBW

[MS08a] J. C. Mandal and J. Subramanian. On the link between weighted least-squares and limiters used in higherorder reconstructions for finite volume computations of hyperbolic equations. Applied Numerical Mathematics: Transactions of IMACS, 58(5):705–725, May 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Meddahi:2008:BFB

[MS08b] Salim Meddahi and Virginia Selgas. An H-based FEM–BEM formulation for a time dependent eddy current problem. Applied Numerical Mathematics: Transactions of IMACS, 58(8): 1061–1083, August 2008.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Mirzaei:2013:DML

[MS13] Davoud Mirzaei and Robert Schaback. Direct Meshless Local Petrov–Galerkin (DMLPG) method: a generalized MLS approximation. Applied Numerical Mathematics: Transactions of IMACS, 68(??): 73–82, June 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413000056.

Maleknejad:2019:AHF

[MS19] K. Maleknejad and M. Shahabi. Application of hybrid functions operational matrices in the numerical solution of two-dimensional nonlinear integral equations. Applied Numerical Mathematics: Transactions of IMACS, 136(??):46– 65, February 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418302113.

Manikantan:2024:MSH

[MS24] Arjun Thenery Manikantan and Jochen Schütz. Multi-step Hermite–Birkhoff predictor-corrector schemes. Applied Numerical Mathematics: Transactions of IMACS, 205(??):281–295, November 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001910.

Mashayekhi:2025:NSF

[MS25] Somayeh Mashayekhi and Salameh Sedaghat. Numerical solution of the forward Kolmogorov equations in population genetics using Eta functions. Applied Numerical Mathematics: Transactions of IMACS, 208 (part B)(??):160–175, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424002861.

Moradi:2020:TIE

[MSA20] A. Moradi, M. Sharifi, and A. Abdi. Transformed implicit–explicit second derivative diagonally implicit multistage integration methods with strong stability preserving explicit part. Applied Numerical Mathematics: Transactions of IMACS, 156(??): 14–31, October 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420301227.

MaraBeh:2023:ALM

[MSGM23] Raed Ali Mara’Beh, Raymond J. Spiteri, P. González, and José M. Mantas. 3additive linear multi-step methods for diffusion–reaction– advection models. Applied Numerical Mathematics: Transactions of IMACS, 183(??):15–38, January 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422002203.

Milovanovic:2010:BEG

[MSP10] Gradimir V. Milovanović, Miodrag M. Spalević, and Miroslav S. Pranić. Bounds of the error of Gauss– Turán-type quadratures, II. Applied Numerical Mathematics: Transactions of IMACS, 60(1–2):1–9, January/February 2010. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Mikula:2015:NPB

[MSS^{+}15] Karol Mikula, Róbert Spir, Michal Smísek, Emmanuel Faure, and Nadine Peyriéras. Nonlinear PDE based numerical methods for cell tracking in zebrafish embryogenesis. Applied Numerical Mathematics: Transactions of IMACS, 95(??): 250–266, September 2015. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414001548.

Mirzaee:2021:MLS

[MSS21] Farshid Mirzaee, Erfan Solhi, and Nasrin Samadyar. Moving least squares and spectral collocation method to approximate the solution of stochastic Volterra– Fredholm integral equations. Applied Numerical Mathematics: Transactions of IMACS, 161(??): 275–285, March 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420303639.

Mazroui:2007:RCB

[MST07] A. Mazroui, D. Sbibih, and A. Tijini. Recursive computation of bivariate Hermite spline interpolants. Applied Numerical Mathematics: Transactions of IMACS, 57(8):962–973, August 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Mazzia:2009:CEB

[MST09] Francesca Mazzia, Alessandra Sestini, and Donato Trigiante. The continuous extension of the Bspline linear multistep methods for BVPs on non-uniform meshes. Applied Numerical Mathematics: Transactions of IMACS, 59(3–4): 723–738, March/April 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Manimaran:2024:NDF

[MSZ^{+}24] J. Manimaran, L. Shangerganesh, M. A. Zaky, A. Akgül, and A. S. Hendy. Numerical discretization for Fisher–Kolmogorov problem with nonlocal diffusion based on mixed Galerkin BDF2 scheme. Applied Numerical Mathematics: Transactions of IMACS, 201(??): 145–158, July 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000485.

Martins:1994:CIM

[MT94] M. M. Martins and M. E.

Trigo. On the convergence of the interval MAOR method. Applied Numerical Mathematics: Transactions of IMACS, 15(4): 439–448, November 1, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=15&
issue=4&aid=487.

Mitchell:2005:AGR

[MT05] William F. Mitchell and Eite Tiesinga. Adaptive grid refinement for a model of two confined and interacting atoms. Applied Numerical Mathematics: Transactions of IMACS, 52(2– 3):235–250, February 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Meerschaert:2006:FDA

[MT06] Mark M. Meerschaert and Charles Tadjeran. Finite difference approximations for two-sided spacefractional partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 56(1):80– 90, January 2006. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Matthes:2011:CAP

[MT11] Michael Matthes and Caren Tischendorf. Convergence analysis of a partial differential algebraic system from coupling a semiconductor model to a circuit model. Applied Numerical Mathematics: Transactions of IMACS, 61(3):382–394, March 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Mukam:2020:SCL

[MT20] Jean Daniel Mukam and Antoine Tambue. Strong convergence of the linear implicit Euler method for the finite element discretization of semilinear nonautonomous SPDEs driven by multiplicative or additive noise. Applied Numerical Mathematics: Transactions of IMACS, 147(??): 222–253, January 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419302119.

Mulder:1999:SMF

[Mul99]  W. A. Mulder.  Spuri-

ous modes in finite-element discretizations of the wave equation may not be all that bad. Applied Numerical Mathematics: Transactions of IMACS, 30(4):425– 445, July 1, 1999. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1999&volume=30&
issue=4&aid=943.

Multerer:2019:NDM

[Mul19] M. D. Multerer. A note on the domain mapping method with rough diffusion coefficients. Applied Numerical Mathematics: Transactions of IMACS, 145(??): 283–296, November 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419301655.

Mund:2000:SSP

[Mun00] E. H. Mund. A short survey on preconditioning techniques in spectral calculations. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4): 61–70, May 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/21/61/27/
                                        31/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/21/61/27/31/article.
                                        pdf.

Muoi:2023:WRP

[Muo23]  Pham Quy Muoi.  l^{1}-

weighted regularization for the problem of recovering sparse initial conditions in parabolic equations from final measurements. Applied Numerical Mathematics: Transactions of IMACS, 183(??): 301–316, January 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422002537.

Murua:1998:RKN

[Mur98] A. Murua. Runge–Kutta– Nyström methods for general second order ODEs with application to multi-body systems. Applied Numerical Mathematics: Transactions of IMACS, 28(2–4):387–399, October 1, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1998/28/
2-4/921.pdf.

Murua:1999:FSNa

[Mur99a]  A. Murua.  Formal se-

ries and numerical integrators. part I: Systems of ODEs and symplectic integrators. Applied Numerical Mathematics: Transactions of IMACS, 29(2):221–251, February 1, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1999/29/
2/930.pdf.

Murua:1999:FSNb

[Mur99b] A. Murua. Formal series and numerical integrators. Part II: Application to index-2 differential-algebraic systems. Applied Numerical Mathematics: Transactions of IMACS, 29(1):99–113, January 1, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1999/29/
1/936.pdf.

Muravleva:2015:ULM

[Mur15] Larisa Muravleva. Uzawalike methods for numerical modeling of unsteady viscoplastic Bingham medium flows. Applied Numerical Mathematics: Transactions of IMACS, 93(??): 140–149, July 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414001020.

Murashige:2019:LWA

[Mur19]  Sunao Murashige.  Long

wave approximation using conformal mapping for largeamplitude internal waves in a two-fluid system. Applied Numerical Mathematics: Transactions of IMACS, 141(??):133–143, ???? 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418300916.

Mustapha:2011:VDL

[Mus11] Kassem Mustapha. The hpand h-versions of the discontinuous and local discontinuous Galerkin methods for one-dimensional singularly perturbed models. Applied Numerical Mathematics: Transactions of IMACS, 61(12):1223–1236, December 2011. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411001309.

Messina:2017:SBN

[MV17] Eleonora Messina and Antonia Vecchio. Stability and boundedness of numerical approximations to Volterra integral equations. Applied Numerical Mathematics: Transactions of IMACS, 116(??):230–237, June 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417300260.

Minev:2018:SSU

[MV18] Peter Minev and Petr N.

Vabishchevich. Splitting schemes for unsteady problems involving the grad–div operator. Applied Numerical Mathematics: Transactions of IMACS, 124(??): 130–139, February 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417302222.

Mishonov:2020:RFP

[MV20] Todor M. Mishonov and Albert M. Varonov. Robust formula for N-point Padé approximant calculation based on Wynn identity. Applied Numerical Mathematics: Transactions of IMACS, 157(??): 291–306, November 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420301896.

Moron-Vidal:2025:ESM

[MVBS25] Jorge Morón-Vidal, Francisco Bernal, and Atsushi Suzuki. An explicit substructuring method for overlapping domain decomposition based on stochastic calculus. Applied Numerical Mathematics: Transactions of IMACS, 208 (part A) (??):340–355, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424000333.

Melchior:2014:MRL

[MVG14] Samuel A. Melchior, Paul Van Dooren, and Kyle A. Gallivan. Model reduction of linear time-varying systems over finite horizons. Applied Numerical Mathematics: Transactions of IMACS, 77(??):72–81, March 2014. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927413001414.

Martin-Vaquero:2009:NET

[MVVA09a] J. Martín-Vaquero and J. Vigo-Aguiar. A note on efficient techniques for the second-order parabolic equation subject to non-local conditions. Applied Numerical Mathematics: Transactions of IMACS, 59(6): 1258–1264, June 2009. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Martin-Vaquero:2009:NSH

[MVVA09b] J. Martín-Vaquero and J. Vigo-Aguiar. On the numerical solution of the heat conduction equations subject to nonlocal conditions. Applied Numerical Mathematics: Transactions of IMACS, 59(10): 2507–2514, October 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Mattheij:1993:PSC

[MW93] R. M. M. Mattheij and S. J.

Wright. Parallel stable compactification for ODEs with parameters and multipoint conditions. Applied Numerical Mathematics: Transactions of IMACS, 13(4): 305–333, November 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Mao:2024:PEE

[MW24a] Mengli Mao and Wansheng Wang. A posteriori error estimates for fully discrete finite difference method for linear parabolic equations. Applied Numerical Mathematics: Transactions of IMACS, 206(??): 111–140, December 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424002022.

Mickens:2024:CAD

[MW24b] Ronald Mickens and Talitha Washington. Construction and analysis of a discrete heat equation using dynamic consistency: the mesoscale limit. Applied Numerical Mathematics: Transactions of IMACS, 199(??): 114–122, May 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742300137X.

Mao:2021:PEE

[MWC21] Wenting Mao, Huasheng Wang, and Chuanjun Chen. A-posteriori error estimations based on postprocessing technique for two-sided fractional differential equations. Applied Numerical Mathematics: Transactions of IMACS, 167(??):73– 91, September 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001185.

Wang:2000:PGM

[mWyG00] Yuan ming Wang and Ben yu Guo. Petrov–Galerkin methods for nonlinear systems without monotonicity. Applied Numerical Mathematics: Transactions of IMACS, 36(1):57– 78, November 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.nl../28/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/65/26/28/article.
pdf.

Mu:2018:DDF

[MWYZ18] Lin Mu, Junping Wang, Xiu Ye, and Shangyou Zhang. A discrete divergence free weak Galerkin finite element method for the Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 125(??): 172–182, March 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417302416.

Miao:2017:SNM

[MYSC17] Xin-He Miao, Jian-Tao Yang, B. Saheya, and Jein-Shan Chen. A smoothing Newton method for absolute value equation associated with second-order cone. Applied Numerical Mathematics: Transactions of IMACS, 120(??): 82–96, October 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417301083.

Mulvey:1987:RTO

[MZ87] John M. Mulvey and Stavros A. Zenios. Real-time operational planning for the U. S. Air Traffic System. Applied Numerical Mathematics: Transactions of IMACS, 3(5):427–441, October 1987. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Melnik:2004:MEW

[MZ04] R. V. N. Melnik and K. N. Zotsenko. Mixed electroelastic waves and CFL stability conditions in computational piezoelectricity. Applied Numerical Mathematics: Transactions of IMACS, 48(1):41–62, January 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ma:2025:IHD

[MZJ^{+}25] Guodong Ma, Wei Zhang, Jinbao Jian, Zefeng Huang, and Jingyi Mo. An inertial hybrid DFPM-based algorithm for constrained nonlinear equations with applications. Applied Numerical Mathematics: Transactions of IMACS, 209(??):

100–123, March 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424003179.

Mead:2005:IPM

[MZK05] J. Mead and B. Zubik-Kowal. An iterated pseudospectral method for delay partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 55(2):227– 250, October 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Meng:2020:VEM

[MZM20] Jian Meng, Yongchao Zhang, and Liquan Mei. A virtual element method for the Laplacian eigenvalue problem in mixed form. Applied Numerical Mathematics: Transactions of IMACS, 156(??):1–13, October 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300933.

Ma:2021:MRM

[MZN21] Wenhua Ma, Zhongshu Zhao, and Guoxi Ni. A multiresolution method for twophase fluids with complex equations of state by binomial solvers in three space dimensions. Applied Numerical Mathematics: Transactions of IMACS, 167(??): 92–107, September 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421001215.

Mao:2010:CSA

[MZS10] Shipeng Mao, Xuying Zhao, and Zhongci Shi. Convergence of a standard adaptive nonconforming finite element method with optimal complexity. Applied Numerical Mathematics: Transactions of IMACS, 60(7): 673–688, July 2010. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Mao:2024:DRO

[MZXX24] Wanying Mao, Qifeng Zhang, Dinghua Xu, and Yinghong Xu. Double reduction order method based conservative compact schemes for the Rosenau equation. Applied Numerical Mathematics: Transactions of IMACS, 197(??):15–45, March 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742300274X.

Ma:2017:ISL

[MZZ17] Zheng Ma, Yong Zhang, and Zhennan Zhou. An improved semi-Lagrangian time splitting spectral method for the semi-classical Schrödinger equation with vector potentials using NUFFT. Applied Numerical Mathematics: Transactions of IMACS, 111(??):144–159, January 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301659.

Nikan:2021:NSF

[NA21] O. Nikan and Z. Avazzadeh. Numerical simulation of fractional evolution model arising in viscoelastic mechanics. Applied Numerical Mathematics: Transactions of IMACS, 169(??): 303–320, November 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001999.

Nemati:2025:HON

[NAD25] Maedeh Nemati, Mostafa Abbaszadeh, and Mehdi Dehghan. High-order numerical solution for solving multidimensional Schrödinger– Poisson equation. Applied Numerical Mathematics: Transactions of IMACS, 217(??):43–72, November 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0168927425001023.

Nabelek:2024:PFG

[NAF24] Patrik Nabelek, Alexis Arlen, and Tanner Fromcke. On periodic and finite genus solutions to the integrable Kaup– Broer system for capillary waves. Applied Numerical Mathematics: Transactions of IMACS, 199(??): 123–135, May 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423001393.

Nagarajan:2022:PRF

[Nag22] Shivaranjani Nagarajan. A parameter robust fitted mesh finite difference method for a system of two reaction– convection–diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 179(??): 87–104, September 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742200109X.

Najafi:2020:STN

[Naj20] Esmaeil Najafi. Smoothing transformation for numerical solution of nonlinear weakly singular Volterra integral equations using quasilinearization and product integration methods. Applied Numerical Mathematics: Transactions of IMACS, 153(??):540–557, July 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300817.

Nakajima:2005:TLH

[Nak05] Kengo Nakajima. Threelevel hybrid vs. flat MPI on the Earth Simulator: Parallel iterative solvers for finite-element method. Applied Numerical Mathematics: Transactions of IMACS, 54(2):237–255, July 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Nakatsukasa:2012:EPB

[Nak12] Yuji Nakatsukasa. Eigenvalue perturbation bounds for Hermitian block tridiagonal matrices. Applied Numerical Mathematics: Transactions of IMACS, 62(1):67–78, January 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411001838.

Nakazono:2024:CAC

[Nak24] Nobutaka Nakazono. Consistency around a cube property of Hirota’s discrete KdV equation and the lattice sine-Gordon equation. Applied Numerical Mathematics: Transactions of IMACS, 199(??):136–152, May 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423001381.

Napoli:2016:SLS

[Nap16]  Anna Napoli.  Solutions

of linear second order initial value problems by using Bernoulli polynomials. Applied Numerical Mathematics: Transactions of IMACS, 99(??):109– 120, January 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927415001348.

Nordstrom:2001:FVA

[NB01] Jan Nordström and Martin Björck. Finite volume approximations and strict stability for hyperbolic problems. Applied Numerical Mathematics: Transactions of IMACS, 38(3):237–255, August 2001. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        10/28/77/33/27/abstract.
                                        html.

Nguyen-Ba:2011:SSP

[NBNTGV11] Truong Nguyen-Ba, Huong Nguyen-Thu, Thierry Giordano, and Rémi Vaillancourt. Strong-stability-preserving 3-stage Hermite– Birkhoff time-discretization methods. Applied Numerical Mathematics: Transactions of IMACS, 61(4): 487–500, April 2011. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Noor:1994:HAM

[NBP94] Ahmed K. Noor, W. Scott Burton, and Jeanne M. Peters. Hierarchical adaptive modeling of structural sandwiches and multilayered composite panels. Applied Numerical Mathematics: Transactions of IMACS, 14(1–3):69–90, April 14, 1994. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=14&
issue=1-3&aid=457. Pro-

ceedings of the Third ARO Workshop on Adaptive Methods for Partial Differential Equations (Troy, NY, 1992).

Nijimbere:2016:NTD

[NC16] V. Nijimbere and L. J.

Campbell. A nonlinear time-dependent radiation condition for simulations of internal gravity waves in geophysical fluid flows. Applied Numerical Mathematics: Transactions of IMACS, 110(??):75– 92, December 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416301428.

Noorizadegan:2022:ECN

[NCYC22] Amir Noorizadegan, Chuin-Shan Chen, D. L. Young, and C. S. Chen. Effective condition number for the selection of the RBF shape parameter with the fictitious point method. Applied Numerical Mathematics: Transactions of IMACS, 178(??): 280–295, August 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422000964.

Nowak:1985:NIS

[ND85] U. Nowak and P. Deuflhard. Numerical identification of selected rate constants in large chemical reaction systems. Applied Numerical Mathematics: Transactions of IMACS, 1(1):59–75, January 1985. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Nedaiasl:2020:HVC

[NDM20] Khadijeh Nedaiasl, Raziyeh Dehbozorgi, and Khosrow Maleknejad. hp-version collocation method for a class of nonlinear Volterra integral equations of the first kind. Applied Numerical Mathematics: Transactions of IMACS, 150(??): 452–477, April 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419302818.

Neumeyer:1995:NBC

[NER95] T. Neumeyer, G. Engl, and P. Rentrop. Numerical benchmark for the charge cycle in a combustion engine. Applied Numerical Mathematics: Transactions of IMACS, 18(1–3):293–305, September 27, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=18&
issue=1-3&aid=607.

Nestler:2016:NBA

[Nes16] Franziska Nestler. An NFFT based approach to the efficient computation of dipole– dipole interactions under various periodic boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 105(??): 25–46, July 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416000131.

Neubauer:1988:PPC

[Neu88] A. Neubauer. An a posteriori parameter choice for Tikhonov regularization in the presence of modeling error. Applied Numerical Mathematics: Transactions of IMACS, 4(6):507– 519, November 1988. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Neytcheva:1995:EIA

[Ney95] Maya G. Neytcheva. Experience in implementing the Algebraic Multilevel Iteration method on a SIMD-type computer. Applied Numerical Mathematics: Transactions of IMACS, 19(1–2): 71–90, December 4, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=19&
                                        issue=1-2&aid=567.

Nordstrom:2003:FVM

[NFAE03] Jan Nordström, Karl Forsberg, Carl Adamsson, and Peter Eliasson. Finite volume methods, unstructured meshes and strict stability for hyperbolic problems. Applied Numerical Mathematics: Transactions of IMACS, 45(4):453–473, June 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Nguyen:2015:VIE

[Ngu15]  Dinh-Liem Nguyen.  A

volume integral equation method for periodic scattering problems for anisotropic Maxwell’s equations. Applied Numerical Mathematics: Transactions of IMACS, 98(??):59–78, December 2015. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415001178.

Norton:2015:ISM

[NH15] Terence J. T. Norton and Adrian T. Hill. An iterative starting method to control parasitism for the Leapfrog method. Applied Numerical Mathematics: Transactions of IMACS, 87(??):145– 156, January 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414001585.

Nikooeinejad:2024:SIA

[NH24] Z. Nikooeinejad and M. Heydari. A spectral iterative algorithm for solving constrained optimal control problems with nonquadratic functional. Applied Numerical Mathematics: Transactions of IMACS, 201(??): 387–403, July 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424000606.

Nicolaides:1986:CHO

[Nic86] R. A. Nicolaides. Construction of higher order accurate vortex and particle methods. Applied Numerical Mathematics: Transactions of IMACS, 2(3–5):313– 320, October 1986. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Njaastad:1988:SSH

[Njå88] Olav Njåstad. Solution of the strong Hamburger moment problem by Laurent continued fractions. Applied Numerical Mathematics: Transactions of IMACS, 4(2–4):351–360, June 1988. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Nagaiah:2011:HOO

[NK11] Chamakuri Nagaiah and Karl Kunisch. Higher order optimization and adaptive numerical solution for optimal control of monodomain equations in cardiac electrophysiology. Applied Numerical Mathematics: Transactions of IMACS, 61(1):53–65, January 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ngoc:2024:EFK

[NK24a] Nguyen Thi Yen Ngoc and Vo Anh Khoa. An explicit Fourier–Klibanov method for an age-dependent tumor growth model of Gompertz type. Applied Numerical Mathematics: Transactions of IMACS, 198(??): 401–418, April 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000205.

Norevik:2024:FEC

[NK24b] Anders M. Norevik and Henrik Kalisch. On the formulation of energy conservation in the eeKdV equation. Applied Numerical Mathematics: Transactions of IMACS, 199(??):153–164, May 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742300140X.

Nkemzi:2007:OCR

[Nke07]  Boniface Nkemzi.  Opti-

mal convergence recovery for the Fourier-finite-element approximation of Maxwell’s equations in nonsmooth axisymmetric domains. Applied Numerical Mathematics: Transactions of IMACS, 57(9):989–1007, September 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Nguyen:2025:FST

[NLHT25] Quang Huy Nguyen, Van Chien Le, Phuong Cuc Hoang, and Thi Thanh Mai Ta. A fitted space-time finite element method for an advection-diffusion problem with moving interfaces. Applied Numerical Mathematics: Transactions of IMACS, 211(??):61–77, May 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425000029.

Ng:2020:SSC

[NLLG20] Sio Wan Ng, Siu-Long Lei, Juan Lu, and Zhiguo Gong. Speeding up Sim-Rank computations by polynomial preconditioners. Applied Numerical Mathematics: Transactions of IMACS, 153(??):147–163, July 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420300453.

Nemati:2018:ENM

[NLS18] S. Nemati, P. Lima, and S. Sedaghat. An effective numerical method for solving fractional pantograph differential equations using modification of hat functions. Applied Numerical Mathematics: Transactions of IMACS, 131(??): 174–189, September 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418301156.

Nemati:2020:LWC

[NLS20] S. Nemati, P. M. Lima, and S. Sedaghat. Legendre wavelet collocation method combined with the Gauss– Jacobi quadrature for solving fractional delay-type integrodifferential equations. Applied Numerical Mathematics: Transactions of IMACS, 149(??):99–112, ???? 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419301394.

Nour:2023:NUU

[NLZB23] M.-Y. Nour, A. Lamnii, A. Zidna, and D. Barrera. Non-uniform UEspline quasi-interpolants and their application to the numerical solution of integral equations. Applied Numerical Mathematics: Transactions of IMACS, 191(??):29– 44, September 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423001265.

Nourtier-Mazauric:2010:TEI

[NMB10] Elise Nourtier-Mazauric and Eric Blayo. Towards efficient interface conditions for a Schwarz domain decomposition algorithm for an advection equation with biharmonic diffusion. Applied Numerical Mathematics: Transactions of IMACS, 60(1–2):83–93, January/ February 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Nassif:2013:GAE

[NMKE13] Nabil R. Nassif, Noha Makhoul-Karam, and Jocelyne Erhel. A globally adaptive explicit numerical method for exploding systems of ordinary differential equations. Applied Numerical Mathematics: Transactions of IMACS, 67(??): 204–219, May 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927411001723.

Negron-Marrero:1994:NSP

[NMSF94] Pablo V. Negrón-Marrero and Bárbara L. Santiago-Figueroa. Novel stability patterns for the large necking of plates in tension: a numerical study. Applied Numerical Mathematics: Transactions of IMACS, 13(6):491– 512, February 1994. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Napov:2010:CBC

[NN10] Artem Napov and Yvan Notay. Comparison of bounds for V -cycle multigrid. Applied Numerical Mathematics: Transactions of IMACS, 60(3):176–192, March 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Notay:2013:FCA

[NN13] Yvan Notay and Artem Napov. Further comparison of additive and multiplicative coarse grid correction. Applied Numerical Mathematics: Transactions of IMACS, 65(??): 53–62, March 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412001973.

Nahid:2020:CAG

[NN20] Nilofar Nahid and Gnaneshwar Nelakanti. Convergence analysis of Galerkin and multi-Galerkin methods on unbounded interval using Hermite polynomials. Applied Numerical Mathematics: Transactions of IMACS, 152(??):66–83, June 2020.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420300234.

Naranjo-Noda:2023:JFH

[NNJ23] F. S. Naranjo-Noda and J. C. Jimenez. Jacobian-free High Order Local Linearization methods for large systems of initial value problems. Applied Numerical Mathematics: Transactions of IMACS, 187(??):158–175, May 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423000405.

Nool:1995:EPB

[Noo95] Margreet Nool. Explicit parallel block Cholesky algorithms on the Cray APP. Applied Numerical Mathematics: Transactions of IMACS, 19(1–2):91–114, December 4, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=19&
                                        issue=1-2&aid=626.

Nordstrom:1997:EPA

[Nor97] Jan Nordström. On extrapolation procedures at artificial outflow boundaries for the time-dependent Navier–Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 23(4):457–468, May 26, 1997. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1997/23/
4/760.pdf.

Nordstrom:1999:FES

[Nor99] Jan Nordström. On fluxextrapolation at supersonic outflow boundaries. Applied Numerical Mathematics: Transactions of IMACS, 30(4):447–457, July 1, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1999&volume=30&
issue=4&aid=937.

Notay:1992:CSM

[Not92] Yvan Notay. Conditioning of Stieltjes matrices by S/P consistently ordered approximate factorizations. Applied Numerical Mathematics: Transactions of IMACS, 10(5):381–396, October 1992. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Notay:1999:MBI

[Not99]  Y. Notay.  A multi-

level block incomplete factorization preconditioning. Applied Numerical Mathematics: Transactions of IMACS, 31(2):209–225, October 23, 1999. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1999&volume=31&
                                        issue=2&aid=998;   http:
                                        //www.sciencedirect.com/
                                        science/article/pii/S0168927498001330.

Novati:2003:PMB

[Nov03] Paolo Novati. A polynomial method based on Fejèr points for the computation of functions of unsymmetric matrices. Applied Numerical Mathematics: Transactions of IMACS, 44(1– 2):201–224, January 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Novati:2008:SSA

[Nov08] Paolo Novati. Some secant approximations for Rosenbrock W -methods. Applied Numerical Mathematics: Transactions of IMACS, 58(3):195–211, March 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Nowak:1996:FAM

[Now96] U. Nowak. A fully adaptive MOL-treatment of parabolic 1-D problems with extrapolation techniques. Applied Numerical Mathematics: Transactions of IMACS, 20(1–2):129–141, February 29, 1996. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=1-2&aid=669. Work-

shop on the method of lines for time-dependent problems (Lexington, KY, 1995).

Nandal:2021:NTF

[NP21] Sarita Nandal and Dwijendra Narain Pandey. Numerical technique for fractional variable-order differential equation of fourth-order with delay. Applied Numerical Mathematics: Transactions of IMACS, 161(??): 391–407, March 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303718.

Neumeyer:1997:DAC

[NR97] T. Neumeyer and P. Rentrop. The DAE-aspect of the charge cycle in a combustion engine. Applied Numerical Mathematics: Transactions of IMACS, 25(2–3): 287–295, October 17, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=25&
                                        issue=2-3&aid=821.

Noschese:2014:NSG

[NR14] Silvia Noschese and Lothar Reichel. A note on superoptimal generalized circulant preconditioners. Applied Numerical Mathematics: Transactions of IMACS, 75(??):188–195, January 2014. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        S0168927413001153.

Noelle:2006:ACS

[NRR06] S. Noelle, W. Rosenbaum, and M. Rumpf. 3D adaptive central schemes: Part I. algorithms for assembling the dual mesh. Applied Numerical Mathematics: Transactions of IMACS, 56(6):778–799, June 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Nagaiah:2008:ANS

[NRWF08] Ch. Nagaiah, S. Rüdiger, G. Warnecke, and M. Falcke. Adaptive numerical simulation of intracellular calcium dynamics using domain decomposition methods. Applied Numerical Mathematics: Transactions of IMACS, 58(11): 1658–1674, November 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Novati:2012:PLS

[NRZR12] P. Novati, M. Redivo-Zaglia, and M. R. Russo. Preconditioning linear systems via matrix function evaluation. Applied Numerical Mathematics: Transactions of IMACS, 62(12): 1804–1818, December 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412001237.

Natividad:2003:REC

[NS03] Maria Caridad Natividad and Martin Stynes. Richardson extrapolation for a convection–diffusion problem using a Shishkin mesh. Applied Numerical Mathematics: Transactions of IMACS, 45(2–3):315–329, May 2003. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Nesemann:2012:NSA

[NS12] Leo Nesemann and Ernst P.

Stephan. Numerical solution of an adhesion problem with FEM and BEM. Applied Numerical Mathematics: Transactions of IMACS, 62(5):606–619, May 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927412000244.

Norton:2013:PEM

[NS13] R. A. Norton and R. Scheichl. Planewave expansion methods for photonic crystal fibres. Applied Numerical Mathematics: Transactions of IMACS, 63(1):88– 104, January 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927412001687.

Nurnberg:2016:FFE

[NS16] Robert Nürnberg and Andrea Sacconi. A fitted finite element method for the numerical approximation of void electro-stress migration. Applied Numerical Mathematics: Transactions of IMACS, 104(??): 204–217, June 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927415001300.

Narayanamurthi:2020:EIP

[NS20]  Mahesh  Narayanamurthi

and Adrian Sandu. Efficient implementation of partitioned stiff exponential Runge–Kutta methods. Applied Numerical Mathematics: Transactions of IMACS, 152(??):141–158, June 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300118.

Nan:2021:EEL

[NS21a] Caixia Nan and Huailing Song. Error estimates of local discontinuous Galerkin method with implicit-explicit Runge Kutta for two-phase miscible flow in porous media. Applied Numerical Mathematics: Transactions of IMACS, 169(??): 334–350, November 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421002002.

Narayanamurthi:2021:PEM

[NS21b]  Mahesh  Narayanamurthi

and Adrian Sandu. Partitioned exponential methods for coupled multiphysics systems. Applied Numerical Mathematics: Transactions of IMACS, 161(??): 178–207, March 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303275.

Nguyen:2019:USF

[NSCC19] Chieu Thanh Nguyen, B. Saheya, Yu-Lin Chang, and Jein-Shan Chen. Unified smoothing functions for absolute value equation associated with second-order cone. Applied Numerical Mathematics: Transactions of IMACS, 135(??): 206–227, January 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418301958.

Noeiaghdam:2023:NNO

[NSD23] Samad Noeiaghdam, Denis Sidorov, and Aliona Dreglea. A novel numerical optimality technique to find the optimal results of Volterra integral equation of the second kind with discontinuous kernel. Applied Numerical Mathematics: Transactions of IMACS, 186(??): 202–212, April 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423000119.

Neves:1992:SNS

[NT92] K. W. Neves and S. Thompson. Software for the numerical solution of systems of functional differential equations with state-dependent delays. Applied Numerical Mathematics: Transactions of IMACS, 9(3–5): 385–401, April 1992. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/016892749290029D.

International Conference on the Numerical Solution of Volterra and Delay Equations (Tempe, AZ, 1990).

Nicholls:2016:HOP

[NT16] David P. Nicholls and Venu Tammali. A high-order perturbation of surfaces (HOPS) approach to Fok’s integral equations: Vector electromagnetic scattering by periodic crossed gratings. Applied Numerical Mathematics: Transactions of IMACS, 101(??): 1–17, March 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415001610.

Nicholls:2020:HOP

[NT20] David P. Nicholls and Xin Tong. High-order perturbation of surfaces algorithms for the simulation of localized surface plasmon resonances in graphene nanotubes. Applied Numerical Mathematics: Transactions of IMACS, 157(??): 544–562, November 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420302129.

Ngoc:2021:SCP

[NTHC21] Tran Bao Ngoc, Vo Viet Tri, Zakia Hammouch, and Nguyen Huu Can. Stability of a class of problems for time-space fractional pseudoparabolic equation with datum measured at terminal time. Applied Numerical Mathematics: Transactions of IMACS, 167(??): 308–329, September 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421001380.

North:2022:NID

[NTT22] Evan North, Semyon Tsynkov, and Eli Turkel. Noniterative domain decomposition for the Helmholtz equation with strong material discontinuities. Applied Numerical Mathematics: Transactions of IMACS, 173(??):51–78, March 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421003172.

Ntumy:2015:ASC

[NU15] Emmanuel A. Ntumy and Sergey V. Utyuzhnikov. Active sound control in composite regions. Applied Numerical Mathematics: Transactions of IMACS, 93(??): 242–253, July 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741400066X.

Nudo:2024:TOP

[Nud24] Federico Nudo. Two oneparameter families of nonconforming enrichments of the Crouzeix–Raviart finite element. Applied Numerical Mathematics: Transactions of IMACS, 203(??): 160–172, September 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001375.

Nurnberg:2009:NSI

[Nür09] Robert Nürnberg. Numerical simulations of immiscible fluid clusters. Applied Numerical Mathematics: Transactions of IMACS, 59(7):1612–1628, July 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Nhan:2023:ASO

[NV23] Thái Anh Nhan and Relja Vulanović. Analysis of a second-order hybrid scheme on Bakhvalov-type meshes: the truncation-error and barrier-function approach. Applied Numerical Mathematics: Transactions of IMACS, 186(??):84–99, April 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0168927423000016.

Nash:2009:HAR

[NW09] Patrick L. Nash and J. A. C. Weideman. High accuracy representation of the free propagator. Applied Numerical Mathematics: Transactions of IMACS, 59(12): 2937–2949, December 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Niu:2022:LDG

[NWL^{+}22] Yuxuan Niu, Jinfeng Wang, Yang Liu, Hong Li, and Zhichao Fang. Local discontinuous Galerkin method based on a family of secondorder time approximation schemes for fractional mobile/immobile convection– diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 179(??):149–169, September 2022. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001131.

Na:2022:ORR

[NX22] Xuyang Na and Xuejun Xu. An optimal Robin-Robin domain decomposition method for Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 171(??): 426–441, January 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421002695.

Nie:2013:SEE

[NY13] Cunyun Nie and Haiyuan Yu. Some error estimates on the finite element approximation for twodimensional elliptic problem with nonlocal boundary. Applied Numerical Mathematics: Transactions of IMACS, 68(??):31–38, June 2013. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413000172.

Nanta:2021:INT

[NYPW21] Supawan Nanta, Suriyon Yimnet, Kanyuta Poochinapan, and Ben Wongsaijai. On the identification of nonlinear terms in the generalized Camassa–Holm equation involving dual-power law nonlinearities. Applied Numerical Mathematics: Transactions of IMACS, 160(??): 386–421, February 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420303135.

Niu:2021:SSG

[NZY21] Ben-Xia Niu, Bing Zheng, and Zhan-Shan Yang. A self-scaling G-transformation for weighted least squares problems. Applied Numerical Mathematics: Transactions of IMACS, 166(??): 333–347, August 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421001136.

Oulmelk:2022:OCA

[OAHN22] A. Oulmelk, L. Afraites, A. Hadri, and M. Nachaoui. An optimal control approach for determining the source term in fractional diffusion equation by different cost functionals. Applied Numerical Mathematics: Transactions of IMACS, 181(??): 647–664, November 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422001830.

Odibat:2020:NSI

[OB20] Zaid Odibat and Dumitru Baleanu. Numerical simulation of initial value problems with generalized Caputotype fractional derivatives. Applied Numerical Mathematics: Transactions of IMACS, 156(??):94–105, October 2020. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420301306.

Odibat:2024:NSN

[OB24] Zaid Odibat and Dumitru Baleanu. Numerical simulation of nonlinear fractional delay differential equations with Mittag-Leffler kernels. Applied Numerical Mathematics: Transactions of IMACS, 201(??): 550–560, July 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000904.

Oberhuber:2015:NSA

[Obe15]  Tomás Oberhuber.  Nu-

merical solution for the anisotropic Willmore flow of graphs. Applied Numerical Mathematics: Transactions of IMACS, 88(??):1– 17, February 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414001627.

Ou:2022:MAN

[OCVW22] Caixia Ou, Dakang Cen, Seakweng Vong, and Zhibo Wang. Mathematical analysis and numerical methods for Caputo–Hadamard fractional diffusion-wave equations. Applied Numerical Mathematics: Transactions of IMACS, 177(??): 34–57, July 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422000538.

Odibat:2019:OSL

[Odi19] Zaid Odibat. On the optimal selection of the linear operator and the initial approximation in the application of the homotopy analysis method to nonlinear fractional differential equations. Applied Numerical Mathematics: Transactions of IMACS, 137(??): 203–212, March 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418302526.

Odibat:2024:NSI

[Odi24] Zaid Odibat. Numerical simulation for an initialboundary value problem of time-fractional Klein– Gordon equations. Applied Numerical Mathematics: Transactions of IMACS, 206(??):1–11, December 2024. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001958.

Ouardghi:2021:EGC

[OEAS21] Abdelouahed Ouardghi, Mofdi El-Amrani, and Mohammed Seaid. An enriched Galerkincharacteristics finite element method for convectiondominated and transport problems. Applied Numerical Mathematics: Transactions of IMACS, 167(??): 119–142, September 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001173.

Okawa:2023:WMN

[OFY^{+}23] Hirotada Okawa, Kotaro Fujisawa, Yu Yamamoto, Ryosuke Hirai, Nobutoshi Yasutake, Hiroki Nagakura, and Shoichi Yamada. The W4 method: a new multidimensional root-finding scheme for nonlinear systems of equations. Applied Numerical Mathematics: Transactions of IMACS, 183(??): 157–172, January 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422002240.

Oosterlee:2008:MRM

[OG08] C. W. Oosterlee and F. J.

Gaspar. Multigrid relaxation methods for systems of saddle point type. Applied Numerical Mathematics: Transactions of IMACS, 58(12):1933–1950, December 2008. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Oloniiju:2020:CPS

[OGS20] Shina D. Oloniiju, Sicelo P. Goqo, and Precious Sibanda. A Chebyshev pseudo-spectral method for the multidimensional fractional Rayleigh problem for a generalized Maxwell fluid with Robin boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 152(??):253–266, June 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419303320.

Orive:1992:FMT

[OGV92a] Ramón Orive and Pablo González-Vera. On a family of multivariate two-point rational approximants. Applied Numerical Mathematics: Transactions of IMACS, 9(6):493–510, May 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Orive:1992:CBT

[OGV92b] Ramón Orive and Pablo González-Vera. On the convergence of bivariate twopoint Padé-type approximants. Applied Numerical Mathematics: Transactions of IMACS, 10(6): 497–508, November 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ong:2020:OSF

[OH20] Thanh Hai Ong and Thi-Thao-Phuong Hoang. Optimized Schwarz and finite element cell-centered method for heterogeneous anisotropic diffusion problems. Applied Numerical Mathematics: Transactions of IMACS, 151(??):380–401, May 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300106.

Ordokhani:2025:NSD

[OHA25] Yadollah Ordokhani, Alireza Hosseinian, and Pouria Assari. Numerical simulation of a dynamic human capital model with demographic delays via the local discrete Galerkin method. Applied Numerical Mathematics: Transactions of IMACS, 217(??):234–254, November 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425001242.

Ojika:1988:NMB

[Oji88] Takeo Ojika. A numerical method for branch points of a system of nonlinear algebraic equations. Applied Numerical Mathematics: Transactions of IMACS, 4(5):419–430, July 1988. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Okongo:1998:IUN

[OK98] Nora Okong’o and Doyle D. Knight. Implicit unstructured Navier–Stokes simulation of leading edge separation over a pitching airfoil. Applied Numerical Mathematics: Transactions of IMACS, 27(3):269–308, July 1, 1998. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1998/27/
                                        3/878.pdf.

Osei-Kuffuor:2010:PHL

[OKS10] Daniel Osei-Kuffuor and Yousef Saad. Preconditioning Helmholtz linear systems. Applied Numerical Mathematics: Transactions of IMACS, 60(4):420–431, April 2010. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

OLeary:1987:NCM

[O’L87]  Dianne P. O’Leary.  A

note on the capacitance matrix algorithm, substructuring, and mixed or Neumann boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 3(4):339–345, August 1987. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Oliveira:2018:EAS

[OL18] Saulo Pomponet Oliveira and Stela Angelozi Leite. Error analysis of the spectral element method with Gauss– Lobatto–Legendre points for the acoustic wave equation in heterogeneous media. Applied Numerical Mathematics: Transactions of IMACS, 129(??):39–57, July 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418300527.

Olver:1992:SES

[Olv92] Peter J. Olver. Symmetry and explicit solutions of partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 10(3–4): 307–324, September 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). A Festschrift to honor Professor Garrett Birkhoff on his eightieth birthday.

Oberai:1998:IDN

[OMP98] Assad A. Oberai, Manish Malhotra, and Peter M. Pinsky. On the implementation of the Dirichlet-to-Neumann radiation condition for iterative solution of the Helmholtz equation. Applied Numerical Mathematics: Transactions of IMACS, 27(4):443–464, August 15, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1998/27/
                                        4/890.pdf.

Oyewole:2025:ATA

[OMSS25] Olawale K. Oyewole, Seithuti P. Moshokoa, Sani Salisu, and Yekini Shehu. Analysis of Tseng algorithm with inertial extrapolation step for stochastic variational inequality problem. Applied Numerical Mathematics: Transactions of IMACS, 217(??): 199–215, November 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425001254.

Odeh:1989:MCM

[ONL89] F. Odeh, O. Nevanlinna, and W. Liniger. Multiplier and contractivity methods for linear multistep methods. Applied Numerical Mathematics: Transactions of IMACS, 5(1–2):89–103, February 1989. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Recent theoretical results in numerical ordinary differential equations.

Oosterlee:1995:CPM

[Oos95] C. W. Oosterlee. The convergence of parallel multiblock multigrid methods. Applied Numerical Mathematics: Transactions of IMACS, 19(1–2):115–128, December 4, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=1-2&aid=568. Mas-

sively parallel computing and applications (Amsterdam, 1993–1994).

Oliveira:2004:RPA

[OP04] Carlos A. S. Oliveira and Panos M. Pardalos. Randomized parallel algorithms for the multidimensional assignment problem. Applied Numerical Mathematics: Transactions of IMACS, 49(1):117–133, April 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Offner:2025:AII

[ÖPT25] Philipp Öffner, Louis Petri, and Davide Torlo. Analysis for implicit and implicit– explicit ADER and DeC methods for ordinary differential equations, advection– diffusion and advection– dispersion equations. Applied Numerical Mathematics: Transactions of IMACS, 212(??):110–134, June 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424003568.

ORiordan:2015:LSP

[OQ15] E. O’Riordan and J. Quinn. A linearised singularly perturbed convection-diffusion problem with an interior layer. Applied Numerical Mathematics: Transactions of IMACS, 98(??):1– 17, December 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927415001142.

Occorsio:2018:NQS

[OR18] Donatella Occorsio and Maria Grazia Russo. A new quadrature scheme based on an Extended Lagrange Interpolation process. Applied Numerical Mathematics: Transactions of IMACS, 124(??):57– 75, February 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417302180.

Occorsio:2020:MSP

[OR20] Donatella Occorsio and Maria Grazia Russo. A mixed scheme of product integration rules in (−1, 1). Applied Numerical Mathematics: Transactions of IMACS, 149(??): 113–123, ???? 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419302661.

Orel:1993:PRK

[Ore93] Bojan Orel. Parallel Runge– Kutta methods with real eigenvalues. Applied Numerical Mathematics: Transactions of IMACS, 11(1– 3):241–250, January 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Parallel methods for ordinary differential equations (Grado, 1991).

Ortleb:2020:SAI

[Ort20] Sigrun Ortleb. L^{2}-stability analysis of IMEX-(σ, μ)DG schemes for linear advection– diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 147(??):43–65, January 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419302181.

Occorsio:2024:SSC

[ORT24] Donatella Occorsio, Maria Grazia Russo, and Woula Themistoclakis. On solving some Cauchy singular integral equations by de la Vallée Poussin filtered approximation. Applied Numerical Mathematics: Transactions of IMACS, 200(??): 358–378, June 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423002052.

ORiordan:2008:PUN

[OS08] E. O’Riordan and G. I. Shishkin. Parameter uniform numerical methods for singularly perturbed elliptic problems with parabolic boundary layers. Applied Numerical Mathematics: Transactions of IMACS, 58(12): 1761–1772, December 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ostermann:2012:EAS

[OS12] Alexander Ostermann and Katharina Schratz. Error analysis of splitting methods for inhomogeneous evolution equations. Applied Numerical Mathematics: Transactions of IMACS, 62(10): 1436–1446, October 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412000888.

Ostermann:1993:CHE

[Ost93] Alexander Ostermann. A class of half-explicit Runge– Kutta methods for differentialalgebraic systems of index 3. Applied Numerical Mathematics: Transactions of IMACS, 13(1–3): 165–179, September 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Sixth Conference on the Numerical Treatment of Differential Equations (Halle, 1992).

Ostermann:2002:SWM

[Ost02] Alexander Ostermann. Stability of W-methods with applications to operator splitting and to geometric theory. Applied Numerical Mathematics: Transactions of IMACS, 42(1–3):353–366, August 2002. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Oswald:1997:ITO

[Osw97]  Peter Oswald.  Intergrid

transfer operators and multilevel preconditioners for nonconforming discretizations. Applied Numerical Mathematics: Transactions of IMACS, 23(1):139–158, March 7, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1997/23/

1/747.pdf. Multilevel methods (Oberwolfach, 1995).

Ostermann:2002:CRK

[OT02] Alexander Ostermann and Mechthild Thalhammer. Convergence of Runge–Kutta methods for nonlinear parabolic equations. Applied Numerical Mathematics: Transactions of IMACS, 42(1– 3):367–380, August 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Offner:2020:AHO

[ÖT20] Philipp Öffner and Davide Torlo. Arbitrary highorder, conservative and positivity preserving Patankartype deferred correction schemes. Applied Numerical Mathematics: Transactions of IMACS, 153(??): 15–34, July 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742030026X.

Occorsio:2021:FPI

[OT21] Donatella Occorsio and Woula Themistoclakis. On the filtered polynomial interpolation at Chebyshev nodes. Applied Numerical Mathematics: Transactions of IMACS, 166(??): 272–287, August 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001124.

Oanh:2022:AAR

[OT22] Dang Thi Oanh and Ngo Manh Tuong. An approach to adaptive refinement for the RBF-FD method for 2D elliptic equations. Applied Numerical Mathematics: Transactions of IMACS, 178(??):123–154, August 2022. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422000770.

Occorsio:2024:AHT

[OT24] Donatella Occorsio and Woula Themistoclakis. Approximation of the Hilbert transform on the half-line. Applied Numerical Mathematics: Transactions of IMACS, 205(??):101–119, November 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424001764.

Occorsio:2025:VPF

[OT25] Donatella Occorsio and Woula Themistoclakis. De la Vallée Poussin filtered polynomial approximation on the half-line. Applied Numerical Mathematics: Transactions of IMACS, 207(??): 569–584, January 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424002332.

Ostermann:2004:SLM

[OTK04] A. Ostermann, M. Thalhammer, and G. Kirlinger. Stability of linear multistep methods and applications to nonlinear parabolic problems. Applied Numerical Mathematics: Transactions of IMACS, 48(3–4): 389–407, March 2004. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ou:2011:HTR

[Ou11] Yigui Ou. A hybrid trust region algorithm for unconstrained optimization. Applied Numerical Mathematics: Transactions of IMACS, 61(7):900–909, July 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ouyang:2025:GEA

[OXW25] Du Ouyang, Jichang Xiao, and Xiaoqun Wang. Generalization error analysis of deep backward dynamic programming for solving nonlinear PDEs. Applied Numerical Mathematics: Transactions of IMACS, 217(??): 319–339, November 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425001308.

Oliger:1996:SEE

[OZ96] Joseph Oliger and Xiaolei Zhu. Stability and error estimation for component adaptive grid methods. Applied Numerical Mathematics: Transactions of IMACS, 20(4):407–426, June 3, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=4&aid=648. Adaptive

mesh refinement methods for CFD applications (Atlanta, GA, 1994).

Omran:2023:NAG

[OZHP23] A. K. Omran, M. A.

Zaky, A. S. Hendy, and V. G. Pimenov. Numerical algorithm for a generalized form of Schnakenberg reaction-diffusion model with gene expression time delay. Applied Numerical Mathematics: Transactions of IMACS, 185(??): 295–310, March 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422003117.

Perez-Acosta:1991:BCD

[PA91] Francisco Pérez-Acosta. On the best choice of denominators in rational collocation methods for nonlinear problems. Applied Numerical Mathematics: Transactions of IMACS, 7(2):195– 200, February 1991. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Pfeiffer:2005:NAS

[PA05] Andreas Pfeiffer and Martin Arnold. Numerical analysis of structure preserving Nyström methods for Hamiltonian systems. Applied Numerical Mathematics: Transactions of IMACS, 53(2–4):391–408, May 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Perez-Arancibia:2018:PWS

[PA18] Carlos Pérez-Arancibia. A plane-wave singularity subtraction technique for the classical Dirichlet and Neumann combined field integral equations. Applied Numerical Mathematics: Transactions of IMACS, 123(??): 221–240, January 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417302027.

Pagano:2025:SEP

[Pag25] Giovanni Pagano. Stabilized explicit peer methods with parallelism across the stages for stiff problems. Applied Numerical Mathematics: Transactions of IMACS, 207(??): 156–173, January 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424002265.

Portero:2012:VSS

[PAJ12] L. Portero, A. Arrarás, and J. C. Jorge. Variable stepsize fractional step Runge– Kutta methods for timedependent partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 62(10): 1463–1476, October 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412001018.

Pantano:2007:ASI

[Pan07] C. Pantano. An additive semi-implicit Runge–Kutta family of schemes for nonstiff systems. Applied Numerical Mathematics: Transactions of IMACS, 57(3): 297–303, March 2007. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Panigrahi:2021:MFL

[Pan21] Bijaya Laxmi Panigrahi. Mixed Fourier Legendre spectral Galerkin methods for two-dimensional Fredholm integral equations of the second kind. Applied Numerical Mathematics: Transactions of IMACS, 168(??):235–250, October 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001719.

Papini:1995:ACD

[Pap95] Alessandra Papini. About the central difference method for singularly perturbed boundary value problems. Applied Numerical Mathematics: Transactions of IMACS, 17(3):333–346, September 4, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=17&

issue=3&aid=586. Numeri-

cal methods for ordinary differential equations (Atlanta, GA, 1994).

Patel:2017:SLM

[PAP17] Ajit Patel, Sanjib Kumar Acharya, and Amiya Kumar Pani. Stabilized Lagrange multiplier method for elliptic and parabolic interface problems. Applied Numerical Mathematics: Transactions of IMACS, 120(??): 287–304, October 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417301356.

Park:2004:PMD

[Par04] Jungho Park. A primal mixed domain decomposition procedure based on the nonconforming streamline diffusion method. Applied Numerical Mathematics: Transactions of IMACS, 50(2):165–181, August 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Park:2014:AMF

[Par14] Won-Kwang Park. Analysis of a multi-frequency electromagnetic imaging functional for thin, crack-like electromagnetic inclusions. Applied Numerical Mathematics: Transactions of IMACS, 77(??):31–42, March 2014.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927413001505.

Park:2021:CSN

[Par21] R. Park. A comparison of six numerical methods for integrating a compartmental Hodgkin–Huxley type model. Applied Numerical Mathematics: Transactions of IMACS, 168(??): 201–220, October 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421001690.

Pasquarelli:1991:DDS

[Pas91]  Franco Pasquarelli.  Do-

main decomposition for spectral approximation to Stokes equations via divergence-free functions. Applied Numerical Mathematics: Transactions of IMACS, 8(6):493– 514, December 1991. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Paternoster:1998:RKN

[Pat98]  B. Paternoster.  Runge–

Kutta(–Nyström) methods for ODEs with periodic solutions based on trigonometric polynomials. Applied Numerical Mathematics: Transactions of IMACS, 28(2–4):401–412, October 1, 1998. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1998/28/
2-4/922.pdf.

Paternoster:2000:PFC

[Pat00] B. Paternoster. A phasefitted collocation-based Runge– Kutta–Nyström method. Applied Numerical Mathematics: Transactions of IMACS, 35(4):339–355, December 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.nl../30/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/64/32/30/article.
pdf.

Paul:1992:DDD

[Pau92] Christopher A. H. Paul. Developing a delay differential equation solver. Applied Numerical Mathematics: Transactions of IMACS, 9(3–5):403–414, April 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/016892749290030H.

Pavarino:2000:DDM

[Pav00] Luca F. Pavarino. Domain decomposition methods with small overlap for Q_{n}–Q_{n−2} spectral elements. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4):463–470, May 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/61/27/
74/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/61/27/74/article.
pdf.

Poullet:2010:EBI

[PB10] Pascal Poullet and Amir Boag. Equation-based interpolation and incremental unknowns for solving the Helmholtz equation. Applied Numerical Mathematics: Transactions of IMACS, 60(11):1148–1156, November 2010. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Patel:2021:EMA

[PB21] Vijay Kumar Patel and Dhirendra Bahuguna. An efficient matrix approach for the numerical solutions of electromagnetic wave model based on fractional partial derivative. Applied Numerical Mathematics: Transactions of IMACS, 169(??):1– 20, November 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001756.

Pasquetti:2008:PPM

[PBC08] R. Pasquetti, R. Bwemba, and L. Cousin. A pseudopenalization method for high Reynolds number unsteady flows. Applied Numerical Mathematics: Transactions of IMACS, 58(7):946–954, July 2008. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Plaza:2000:LRS

[PC00] A. Plaza and G. F. Carey.

Local refinement of simplicial grids based on the skeleton. Applied Numerical Mathematics: Transactions of IMACS, 32(2):195– 218, February 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/59/27/
31/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/59/27/31/article.
pdf.

Porras:2010:SGL

[PCA10] Ignacio Porras and Francisco Cordobés-Aguilar. Study of a generalized Levin–Weniger convergence accelerator of numerical series: Particular cases and applications to multi-electron integrals in atomic physics. Applied Numerical Mathematics: Transactions of IMACS, 60(12):1371–1381, December 2010. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Park:2025:TOF

[PCL^{+}25] Kwanghyuk Park, Xinjuan Chen, Dongjin Lee, Jiaxi Gu, and Jae-Hun Jung. A third-order finite difference weighted essentially non-oscillatory scheme with shallow neural network. Applied Numerical Mathematics: Transactions of IMACS, 218(??):1– 21, December 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425001461.

Porta:2017:RKL

[PCR17] Federica Porta, Anastasia Cornelio, and Valeria Ruggiero. Runge–Kutta-like scaling techniques for firstorder methods in convex optimization. Applied Numerical Mathematics: Transactions of IMACS, 116(??): 256–272, June 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416301611.

Puelz:2017:CRM

[PCRR17] Charles Puelz, Suncica Canić, Béatrice Rivière, and Craig G. Rusin. Comparison of reduced models for blood flow using Runge–Kutta discontinuous Galerkin methods. Applied Numerical Mathematics: Transactions of IMACS, 115(??):114–141, May 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417300077.

Palmerio:1996:MMA

[PD96] Bernadette Palmerio and Alain Dervieux. Multimesh and multiresolution analysis for mesh adaptive interpolation. Applied Numerical Mathematics: Transactions of IMACS, 22(4):477– 493, December 16, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=22&
issue=4&aid=732.

Pelloni:2001:EEF

[PD01] Beatrice Pelloni and Vassilios A. Dougalis. Error estimates for a fully discrete spectral scheme for a class of nonlinear, nonlocal dispersive wave equations. Applied Numerical Mathematics: Transactions of IMACS, 37(1–2):95–107, April 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/10/28/73/26/
                                        31/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/10/28/73/26/31/article.
                                        pdf.

PantaPazos:1999:CTT

[PdV99] Rubén Panta Pazos and Marco Tullio de Vilhena. Convergence in transport theory. Applied Numerical Mathematics: Transactions of IMACS, 30(1):79– 92, May 10, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1999/30/
                                        1/951.pdf.

Pearson:2016:FIS

[Pea16] John W. Pearson. Fast iterative solvers for large matrix systems arising from timedependent Stokes control problems. Applied Numerical Mathematics: Transactions of IMACS, 108(??):87– 101, October 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416300708.

Pechstein:2009:BET

[Pec09] Clemens Pechstein. Boundary element tearing and interconnecting methods in unbounded domains. Applied Numerical Mathematics: Transactions of IMACS, 59(11):2824–2842, November 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Pellegrino:2015:NEN

[Pel15] E. Pellegrino. Numerical evaluation of new quadrature rules using refinable operators. Applied Numerical Mathematics: Transactions of IMACS, 90(??): 132–145, April 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414001986.

Pellegrino:2020:IFV

[Pel20] Sabrina Francesca Pellegrino. On the implementation of a finite volumes scheme with monotone transmission conditions for scalar conservation laws on a star-shaped network. Applied Numerical Mathematics: Transactions of IMACS, 155(??):181–191, ???? 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419302508.

Pereyra:1988:NMI

[Per88]  V. Pereyra.  Numerical

methods for inverse problems in three-dimensional geophysical modeling. Applied Numerical Mathematics: Transactions of IMACS, 4(1):97–139, March 1988. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Pereyra:1999:ADS

[Per99] V. Pereyra. Asynchronous distributed solution of large scale nonlinear inversion problems. Applied Numerical Mathematics: Transactions of IMACS, 30(1):31– 40, May 10, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1999/30/
                                        1/947.pdf.

Pereira:2003:SMP

[Per03] Edgar Pereira. On solvents of matrix polynomials. Applied Numerical Mathematics: Transactions of IMACS, 47(2): 197–208, November 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Petzold:1987:OAM

[Pet87] Linda R. Petzold. Observations on an adaptive moving grid method for onedimensional systems of partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 3(4):347–360, August 1987. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Petiton:1992:PSM

[Pet92] Serge G. Petiton. Parallel subspace method for non-Hermitian eigenproblems on the Connection Machine (CM2). Applied Numerical Mathematics: Transactions of IMACS, 10(1): 19–35, June 1992. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Petropoulos:2000:RSL

[Pet00] P. G. Petropoulos. Reflectionless sponge layers for the numerical solution of Maxwell’s equations in cylindrical and spherical coordinates. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4): 517–524, May 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/61/27/
81/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/61/27/81/article.
pdf.

Peyraut:2009:RNA

[PFHL09] F. Peyraut, Z.-Q. Feng, Q.-C. He, and N. Labed. Robust numerical analysis of homogeneous and nonhomogeneous deformations. Applied Numerical Mathematics: Transactions of IMACS, 59(7):1499–1514, July 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Pflaum:2008:MCG

[Pfl08]  Christoph Pflaum.  A

multigrid conjugate gradient method. Applied Numerical Mathematics: Transactions of IMACS, 58(12): 1803–1817, December 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Pulch:2002:MCS

[PG02] R. Pulch and M. Günther. A method of characteristics for solving multirate partial differential equations in radio frequency application. Applied Numerical Mathematics: Transactions of IMACS, 42(1– 3):397–409, August 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Papadimitropoulos:2021:DAB

[PG21] Symeon Papadimitropoulos and Dan Givoli. The double absorbing boundary method for the Helmholtz equation. Applied Numerical Mathematics: Transactions of IMACS, 168(??): 182–200, October 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001665.

Palencia:1993:SLM

[PGA93] C. Palencia and B. García-Archilla. Stability of linear multistep methods for sectorial operators in Banach spaces. Applied Numerical Mathematics: Transactions of IMACS, 12(6):503– 520, August 1993. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Polit:2001:FLO

[PGC01] Monique Polit, Antoine Genovesi, and Bernard Claudet. Fuzzy logic observers for a biological wastewater treatment process. Applied Numerical Mathematics: Transactions of IMACS, 39(2):173–180, November 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/79/33/29/abstract.
html.

Peer:2008:NFO

[PGDB08] A. A. I. Peer, A. Gopaul, M. Z. Dauhoo, and M. Bhuruth. A new fourthorder non-oscillatory central scheme for hyperbolic conservation laws. Applied Numerical Mathematics: Transactions of IMACS, 58(5):674–688, May 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Power:1986:IES

[PGM86] Henry Power, Reinaldo García, and Guillermo Miranda. Integral equation solution for the flow due to the motion of a body of arbitrary shape near a plane interface at small Reynolds number. Applied Numerical Mathematics: Transactions of IMACS, 2(2):79– 94, April 1986. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Pieraccini:2003:GNT

[PGP03] S. Pieraccini, M. G. Gasparo, and A. Pasquali. Global Newton-type methods and semismooth reformulations for NCP. Applied Numerical Mathematics: Transactions of IMACS, 44(3):367– 384, February 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Parodi:2010:PLA

[PGS10] Mauro Parodi, Mauro Gaggero, and Marco Storace. Piecewise linear approximations of multivariate functions: a multiresolutionbased compression algorithm suitable for circuit implementation. Applied Numerical Mathematics: Transactions of IMACS, 60(9): 924–933, September 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Peng:2020:PPF

[PGYF20] Gang Peng, Zhiming Gao, Wenjing Yan, and Xinlong Feng. A positivity-preserving finite volume scheme for three-temperature radiation diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 152(??):125–140, June 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300143.

Petkovic:1991:HIA

[PH91] Miodrag S. Petković and Jürgen Herzberger. Hybrid inclusion algorithms for polynomial multiple complex zeros in rectangular arithmetic. Applied Numerical Mathematics: Transactions of IMACS, 7(3):241– 262, March 1991. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Perline:2015:HLS

[PH15] K. R. Perline and B. T. Helenbrook. A hybrid level-set/ moving-mesh interface tracking method. Applied Numerical Mathematics: Transactions of IMACS, 92(??): 21–39, June 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927415000148.

Pimenov:2017:BTS

[PH17] V. G. Pimenov and A. S. Hendy. BDF-type shifted Chebyshev approximation scheme for fractional functional differential equations with delay and its error analysis. Applied Numerical Mathematics: Transactions of IMACS, 118(??): 266–276, August 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417300879.

Phillips:1987:SPA

[Phi87] Timothy N. Phillips. The smoothing properties of the alternating direction implicit method in multigrid iterations. Applied Numerical Mathematics: Transactions of IMACS, 3(6):513– 522, November 1987. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Phillips:1991:MID

[Phi91] Timothy N. Phillips. On methods of incomplete LU decompositions for solving Poisson’s equation in annular regions. Applied Numerical Mathematics: Transactions of IMACS, 8(6):515– 531, December 1991. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Pei:2019:LSO

[PHY19] Shuaichao Pei, Yanren Hou, and Bo You. A linearly second-order energy stable scheme for the phase field crystal model. Applied Numerical Mathematics: Transactions of IMACS, 140(??):134–164, June 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419300285.

Picasso:2005:AAS

[Pic05] M. Picasso. An adaptive algorithm for the Stokes problem using continuous, piecewise linear stabilized finite elements and meshes with high aspect ratio. Applied Numerical Mathematics: Transactions of IMACS, 54(3–4):470–490, August 2005. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Pirozzi:2009:HRC

[Pir09] Maria Antonietta Pirozzi.

High-resolution compact upwind finite difference methods for linear wave phenomena. Applied Numerical Mathematics: Transactions of IMACS, 59(8): 1905–1921, August 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Pishbin:2022:SIA

[Pis22] S. Pishbin. Solving integralalgebraic equations with non-vanishing delays by Legendre polynomials. Applied Numerical Mathematics: Transactions of IMACS, 179(??):221–237, September 2022. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422001167.

Portero:2004:AOR

[PJB04] L. Portero, J. C. Jorge, and B. Bujanda. Avoiding order reduction of fractional step Runge–Kutta discretizations for linear time dependent coefficient parabolic problems. Applied Numerical Mathematics: Transactions of IMACS, 48(3–4):409–424, March 2004. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Phillips:1991:CSC

[PK91] Timothy N. Phillips and Andreas Karageorghis. A conforming spectral collocation strategy for Stokes flow through a channel contraction. Applied Numerical Mathematics: Transactions of IMACS, 7(4): 329–345, April 1991. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/016892749190068B.

Patra:2021:LMO

[PK21] Suchismita Patra and V. V.

K. Srinivas Kumar. A local min-orthogonal based numerical method for computing multiple coexisting solutions to cooperative p-Laplacian systems. Applied Numerical Mathematics: Transactions of IMACS, 169(??):221–242, November 2021. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001963.

Pargaei:2023:HWM

[PK23] Meena Pargaei and B. V. Rathish Kumar. A 3D Haar wavelet method for a coupled degenerate system of parabolic equations with nonlinear source coupled with non-linear ODEs. Applied Numerical Mathematics: Transactions of IMACS, 185(??):141–164, March 2023. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422002987.

Pedretscher:2019:PIU

[PKP19] B. Pedretscher, B. Kaltenbacher, and O. Pfeiler. Parameter identification and uncertainty quantification in stochastic state space models and its application to texture analysis. Applied Numerical Mathematics: Transactions of IMACS, 146(??):38– 54, December 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419301722.

Pashos:2010:AIN

[PKSB10] George Pashos, Eleni D. Koronaki, Antony N. Spyropoulos, and Andreas G. Boudouvis. Accelerating an inexact Newton/GMRES scheme by subspace decomposition. Applied Numerical Mathematics: Transactions of IMACS, 60(4):397–410, April 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Pan:2020:NCM

[PL20] Jiajia Pan and Huiyuan Li. A new collocation method using near-minimal Chebyshev quadrature nodes on a square. Applied Numerical Mathematics: Transactions of IMACS, 154(??): 104–128, August 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300842.

Pan:2025:WGS

[PL25] Jiajia Pan and Huiyuan Li. Weak Galerkin spectral element methods for elliptic eigenvalue problems: Lower bound approximation and superconvergence. Applied Numerical Mathematics: Transactions of IMACS, 218(??): 182–200, December 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425001515.

Plata:2008:RPP

[Pla08] Sergio Amat Plata. A review on the piecewise polynomial harmonic interpolation. Applied Numerical Mathematics: Transactions of IMACS, 58(8):1168– 1185, August 2008. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Pepin:2022:HDS

[PLB22] Alexandre Pepin, Sophie Léger, and Normand Beaudoin. High-degree splines from discrete Fourier transforms: Robust methods to obtain the boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 181(??): 594–617, November 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422001805.

Pereyra:2003:AOD

[PLI03] V. Pereyra, D. Lawver, and J. Isenberg. An algorithm for optimal design of steel frame structures. Applied Numerical Mathematics: Transactions of IMACS, 47(3– 4):503–514, December 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Plociniczak:2022:LGL

[Pło22] Lukasz Płociniczak. Linear Galerkin–Legendre spectral scheme for a degenerate nonlinear and nonlocal parabolic equation arising in climatology. Applied Numerical Mathematics: Transactions of IMACS, 179(??): 105–124, September 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422001106.

Plociniczak:2023:LGN

[Pło23] Lukasz Płociniczak. A linear Galerkin numerical method for a quasilinear subdiffusion equation. Applied Numerical Mathematics: Transactions of IMACS, 185(??): 203–220, March 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422003075.

Pathria:1991:VDS

[PM91] D. Pathria and J. L. Morris. A variable degree spectral collocation algorithm for the solution of nonlinear evolutionary equations. Applied Numerical Mathematics: Transactions of IMACS, 8(3):243–256, October 1991. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Plessix:2003:SVP

[PM03] R. E. Plessix and W. A.

Mulder. Separation-of-variables as a preconditioner for an iterative Helmholtz solver. Applied Numerical Mathematics: Transactions of IMACS, 44(3):385– 400, February 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Parumasur:2005:ASM

[PM05] Nabendra Parumasur and Janusz R. Mika. Amplitudeshape method for solving partial differential equations of chemical kinetics.

Applied Numerical Mathematics: Transactions of IMACS, 55(4):473–479, December 2005. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Przybylowicz:2014:SAS

[PM14] Pawel Przybylowicz and Pawel Morkisz. Strong approximation of solutions of stochastic differential equations with time-irregular coefficients via randomized Euler algorithm. Applied Numerical Mathematics: Transactions of IMACS, 78(??):80–94, April 2014. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413001736.

Priyadarshana:2023:SOF

[PMP23] S. Priyadarshana, J. Mohapatra, and S. R. Pattanaik. A second order fractional step hybrid numerical algorithm for time delayed singularly perturbed 2D convection– diffusion problems. Applied Numerical Mathematics: Transactions of IMACS, 189(??):107–129, July 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423000958.

Pasha:2021:TOA

[PNA21] Syed Ahmed Pasha, Yasir Nawaz, and Muhammad Shoaib Arif. A third-order accurate in time method for boundary layer flow problems. Applied Numerical Mathematics: Transactions of IMACS, 161(??):13–26, March 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303305.

Pohl:1993:CDM

[Poh93] Bert Pohl. On the convergence of the discretized multi-splitting waveform relaxation algorithm. Applied Numerical Mathematics: Transactions of IMACS, 11(1–3):251–258, January 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Parallel methods for ordinary differential equations (Grado, 1991).

Porwal:2017:DGM

[Por17] Kamana Porwal. Discontinuous Galerkin methods for a contact problem with Tresca friction arising in linear elasticity. Applied Numerical Mathematics: Transactions of IMACS, 112(??): 182–202, February 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416302136.

Potempa:1985:IIM

[Pot85] Thom Potempa. Improved implementation of the Mc-Cracken and Yanosik nine point finite difference procedure. Applied Numerical Mathematics: Transactions of IMACS, 1(3):261– 272, May 1985. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Potapov:1997:NMI

[Pot97] V. D. Potapov. Numerical method for investigation of stability of stochastic integro-differential equations. Applied Numerical Mathematics: Transactions of IMACS, 24(2–3): 191–201, June 19, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=24&
                                        issue=2-3&aid=776. Volterra

centennial (Tempe, AZ,

1996).

Poullet:2000:SIU

[Pou00]  Pascal Poullet.  Stag-

gered incremental unknowns for solving Stokes and generalized Stokes problems. Applied Numerical Mathematics: Transactions of IMACS, 35(1):23– 41, September 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/64/27/
29/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/64/27/29/article.
pdf.

Powell:1994:TBA

[Pow94] Kenneth G. Powell. A treebased adaptive scheme for solution of the equations of gas dynamics and magnetohydrodynamics. Applied Numerical Mathematics: Transactions of IMACS, 14(1–3):327–352, April 14, 1994. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=14&
issue=1-3&aid=464. Pro-

ceedings of the Third ARO Workshop on Adaptive Methods for Partial Differential Equations (Troy, NY, 1992).

Petzold:1992:OMN

[PP92] Linda R. Petzold and Florian A. Potra. ODAE methods for the numerical solution of Euler–Lagrange equations. Applied Numerical Mathematics: Transactions of IMACS, 10(5):397– 413, October 1992. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Phillips:2000:FPC

[PP00] R. M. Phillips and T. N. Phillips. Flow past a cylinder using a semi-Lagrangian spectral element method. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4): 251–257, May 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/21/61/27/
                                        51/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/21/61/27/51/article.
                                        pdf.

Patel:2024:DLS

[PP24] Subhashree Patel and Bijaya Laxmi Panigrahi. Discrete Legendre spectral projection-based methods for Tikhonov regularization of first kind Fredholm integral equations. Applied Numerical Mathematics: Transactions of IMACS, 198(??): 75–93, April 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423003161.

Plaza:2000:RDA

[PPC00] Angel Plaza, Miguel A.

Padrón, and Graham F. Carey. A 3D refinement/derefinement algorithm for solving evolution problems. Applied Numerical Mathematics: Transactions of IMACS, 32(4): 401–418, April 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/59/32/
31/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/59/32/31/article.
pdf.

Plaza:2005:NDS

[PPS05] Angel Plaza, Miguel A.

Padrón, and José P. Suárez. Non-degeneracy study of the 8-tetrahedra longest-edge partition. Applied Numerical Mathematics: Transactions of IMACS, 55(4): 458–472, December 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Papadomanolaki:2010:CSB

[PPS10] M. G. Papadomanolaki, E. P.

Papadopoulou, and Y. G. Saridakis. p-cyclic SOR for BVPs with periodic boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 60(4):411–419, April 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Paluszny:2002:SPA

[PPT02] M. Paluszny, R. Patterson, and F. Tovar. The singular point of an algebraic cubic. Applied Numerical Mathematics: Transactions of IMACS, 40(1–2):23–31, January 2002. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        10/28/85/27/29/abstract.
                                        html.

Peng:2025:WGF

[PQ25] Hui Peng and Wenya Qi. Weak Galerkin finite element method with the total pressure variable for Biot’s consolidation model. Applied Numerical Mathematics: Transactions of IMACS, 207(??):450–469, January 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742400254X.

Parihar:1989:NSP

[PR89] K. S. Parihar and M. P. Ramachandran. On the numerical solution of Prandtl’s integral equation. Applied Numerical Mathematics: Transactions of IMACS, 5(3):223–235, May 1989. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Perkins:1990:DDM

[PR90] A. Louise Perkins and Garry Rodrigue. A domain decomposition method for solving a two-dimensional viscous Burgers’ equation. Applied Numerical Mathematics: Transactions of IMACS, 6(4):329–340, May 1990. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Pasadas:2009:CAS

[PR09] M. Pasadas and M. L.

Rodríguez. Construction and approximation of surfaces by discrete PDE splines on a polygonal domain. Applied Numerical Mathematics: Transactions of IMACS, 59(1):205–218, January 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Philippe:2012:GKS

[PR12] Bernard Philippe and Lothar Reichel. On the generation of Krylov subspace bases. Applied Numerical Mathematics: Transactions of IMACS, 62(9): 1171–1186, September 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411000080.

Pes:2022:DRM

[PR22] Federica Pes and Giuseppe Rodriguez. A doubly relaxed minimal-norm Gauss– Newton method for underdetermined nonlinear leastsquares problems. Applied Numerical Mathematics: Transactions of IMACS, 171(??):233–248, January 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421002464.

Prevost:1990:DEP

[Pre90] Marc Prevost. Determinantal expression for partial Padé approximants. Applied Numerical Mathematics: Transactions of IMACS, 6(3):221–224, March 1990. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Prevost:1995:RCP

[Pré95] M. Prévost. Rate of convergence of Padé approximants for a particular Wynn series. Applied Numerical Mathematics: Transactions of IMACS, 17(4):461– 469, September 11, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.

cgi?year=1995&volume=17&
issue=4&aid=622.

Prevost:2010:RVZ

[Pré10] Marc Prévost. Recurrence for values of the zeta function. Applied Numerical Mathematics: Transactions of IMACS, 60(12): 1382–1394, December 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Pinto:2016:RIS

[PRGO16] M. A. V. Pinto, C. Rodrigo, F. J. Gaspar, and C. W. Oosterlee. On the robustness of ILU smoothers on triangular grids. Applied Numerical Mathematics: Transactions of IMACS, 106(??):37–52, August 2016. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416300319.

Provenzi:2024:AQI

[Pro24] Edoardo Provenzi. Advances in a quantum informationbased color perception theory. Applied Numerical Mathematics: Transactions of IMACS, 203(??): 269–275, September 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001181.

Praetorius:2020:SAY

[PRS20] Dirk Praetorius, Michele Ruggeri, and Ernst P. Stephan. The saturation assumption yields optimal convergence of two-level adaptive BEM. Applied Numerical Mathematics: Transactions of IMACS, 152(??): 105–124, June 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300155.

Pourfattah:2023:EAB

[PRS23] Elnaz Pourfattah, Mohammad Jahangiri Rad, and Behzad Nemati Saray. An efficient algorithm based on the Pseudospectral method for solving Abel’s integral equation using Hermite cubic spline scaling bases. Applied Numerical Mathematics: Transactions of IMACS, 185(??): 434–445, March 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422003270.

Pejcev:2024:NCQ

[PRSS24] Aleksandar V. Pejcev, Lothar Reichel, Miodrag M. Spalević, and Stefan M. Spalević. A new class of quadrature rules for estimating the error in Gauss quadrature. Applied Numerical Mathematics: Transactions of IMACS, 204(??): 206–221, October 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742400151X.

Pauca:2002:MTA

[PRST02] V. P. Pauca, A. F. Rodriguez, X. Sun, and K. S. Trivedi. A methodology towards automatic implementation of N-body algorithms. Applied Numerical Mathematics: Transactions of IMACS, 40(1–2):3– 21, January 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/85/27/28/abstract.
html.

Pruess:2000:QCS

[Pru00]  Steven Pruess.  Qual-

itatively correct Sturm– Liouville eigenfunctions. Applied Numerical Mathematics: Transactions of IMACS, 34(1):127–141, June 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/62/27/
34/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/62/27/34/article.
pdf.

Pfabe:2000:NMI

[PS00] Kristin Pfabe and Thomas S. Shores. Numerical methods for an ion transport problem. Applied Numerical Mathematics: Transactions of IMACS, 32(2):175– 193, February 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/10/28/59/27/
                                        30/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/10/28/59/27/30/article.
                                        pdf.

Pereyra:2002:LSS

[PS02] V. Pereyra and G. Scherer. Least squares scattered data fitting by truncated SVDs. Applied Numerical Mathematics: Transactions of IMACS, 40(1–2):73–86, January 2002. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        10/28/85/27/34/abstract.
                                        html.

Pereyra:2003:LSL

[PS03] V. Pereyra and G. Scherer. Large scale least squares scattered data fitting. Applied Numerical Mathematics: Transactions of IMACS, 44(1–2):225–239, January 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Pistella:2004:P

[PS04] Francesca Pistella and Rosa Maria Spitaleri. Preface. Applied Numerical Mathematics: Transactions of IMACS, 51(4):407–408, December 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Pistella:2005:P

[PS05] Francesca Pistella and Rosa Maria Spitaleri. Preface. Applied Numerical Mathematics: Transactions of IMACS, 55(3):251–252, November 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Pennacchio:2009:AMP

[PS09] Micol Pennacchio and Valeria Simoncini. Algebraic multigrid preconditioners for the bidomain reaction–diffusion system. Applied Numerical Mathematics: Transactions of IMACS, 59(12):3033–3050, December 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Pramanick:2019:EET

[PS19] Tamal Pramanick and Rajen Kumar Sinha. Error estimates for two-scale composite finite element approximations of parabolic equations with measure data in time for convex and nonconvex polygonal domains. Applied Numerical Mathematics: Transactions of IMACS, 143(??): 112–132, September 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419300625.

Porwal:2021:FEM

[PS21] Kamana Porwal and Pratibha Shakya. A finite element method for an elliptic optimal control problem with integral state constraints. Applied Numerical Mathematics: Transactions of IMACS, 169(??): 273–288, November 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001938.

Paalvast:1991:BHL

[PSB91] Edwin M. Paalvast, Henk J. Sips, and Leo C. Breebaart. Booster: a high-level language for portable parallel algorithms. Applied Numerical Mathematics: Transactions of IMACS, 8(2): 177–192, September 1991. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/016892749190050A.

Pulch:2018:IAM

[PSL18] Roland Pulch, Diana Estévez Schwarz, and René Lamour. Index-analysis for a method of lines discretising multirate partial differential algebraic equations. Applied Numerical Mathematics: Transactions of IMACS, 130(??):51–69, August 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418300758.

Shen:2016:NTM

[pSLqJcY16] Wei ping Shen, Chong Li, Xiao qing Jin, and Jen chih Yao. Newton-type methods for inverse singular value problems with multiple singular values. Applied Numerical Mathematics: Transactions of IMACS, 109(??): 138–156, November 2016. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301027.

Padron:2004:ATU

[PSP04a] Miguel A. Padrón, José P.

Suárez, and Ángel Plaza. Adaptive techniques for unstructured nested meshes. Applied Numerical Mathematics: Transactions of IMACS, 51(4):565–579, December 2004. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Plaza:2004:NEP

[PSP04b] Ángel Plaza, José P. Suárez, and Miguel A. Padrón. Nonequivalent partitions of dtriangles with Steiner points. Applied Numerical Mathematics: Transactions of IMACS, 49(3–4):415–430, June 2004. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Padron:2005:CSB

[PSP05] Miguel A. Padrón, José P. Suárez, and Ángel Plaza. A comparative study between some bisection based partitions in 3D. Applied Numerical Mathematics: Transactions of IMACS, 55(3): 357–367, November 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Petkovic:2004:FSM

[PSR04] M. S. Petković, T. Sakurai, and L. Rančić. Family of simultaneous methods of Hansen–Patrick’s type. Applied Numerical Mathematics: Transactions of IMACS, 50(3–4):489–510, September 2004. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Podhaisky:2002:DAT

[PSW02] H. Podhaisky, B. A. Schmitt, and R. Weiner. Design, analysis and testing of some parallel two-step W-methods for stiff systems. Applied Numerical Mathematics: Transactions of IMACS, 42(1– 3):381–395, August 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Peng:2021:ABP

[PSWZ21] Jie Peng, Shi Shu, Junxian Wang, and Liuqiang Zhong. An adaptive BDDC preconditioner for advectiondiffusion problems with a stabilized finite element discretization. Applied Numerical Mathematics: Transactions of IMACS, 165(??): 184–197, July 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000623.

Pomeranz:1995:CSI

[PT95] Shirley B. Pomeranz and Travis A. Tull. Construction of some iterative methods for solving boundary element linear systems. Applied Numerical Mathematics: Transactions of IMACS, 16(3):369–381, February 16, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=16&
                                        issue=3&aid=534.

Patel:2009:NTP

[PT09] Manisha Patel and M. G. Timol. Numerical treatment of Powell–Eyring fluid flow using Method of Satisfaction of Asymptotic Boundary Conditions (MSABC). Applied Numerical Mathematics: Transactions of IMACS, 59(10):2584–2592, October 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Pedas:2011:DCM

[PT11] Arvet Pedas and Enn Tamme. A discrete collocation method for Fredholm integro-differential equations with weakly singular kernels. Applied Numerical Mathematics: Transactions of IMACS, 61(6):738–751, June 2011. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Potts:2015:FEA

[PT15] Daniel Potts and Manfred Tasche. Fast ESPRIT algorithms based on partial singular value decompositions. Applied Numerical Mathematics: Transactions of IMACS, 88(??):31– 45, February 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414001640.

Petrov:2019:RSR

[PT19] Miroslav S. Petrov and Todor D. Todorov. Refinement strategies related to cubic tetrahedral meshes. Applied Numerical Mathematics: Transactions of IMACS, 137(??):169–183, March 2019. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418302551.

Petrov:2023:INI

[PT23] Miroslav S. Petrov and Todor D. Todorov. Isoparametric numerical integration on enriched 4D simplicial elements. Applied Numerical Mathematics: Transactions of IMACS, 183(??): 221–235, January 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422002318.

Pedas:2016:SCF

[PTV16] Arvet Pedas, Enn Tamme, and Mikk Vikerpuur. Spline collocation for fractional weakly singular integrodifferential equations. Applied Numerical Mathematics: Transactions of IMACS, 110(??):204–214, December 2016. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301386.

Pedas:2020:NSL

[PTV20] Arvet Pedas, Enn Tamme, and Mikk Vikerpuur. Numerical solution of linear fractional weakly singular integro-differential equations with integral boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 149(??): 124–140, ???? 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419301862.

Petrosyan:2019:RJS

[PTW19] Armenak Petrosyan, Hoang Tran, and Clayton Webster. Reconstruction of jointly sparse vectors via manifold optimization. Applied Numerical Mathematics: Transactions of IMACS, 144(??):140–150, October 2019. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927419301370.

Pulliam:1986:ISM

[Pul86] Thomas H. Pulliam. Implicit solution methods in computational fluid dynamics. Applied Numerical Mathematics: Transactions of IMACS, 2(6):441–474, December 1986. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Pulch:2005:MTS

[Pul05] R. Pulch. Multi time scale differential equations for simulating frequency modulated signals. Applied Numerical Mathematics: Transactions of IMACS, 53(2–4): 421–436, May 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Pulch:2009:PCM

[Pul09] Roland Pulch. Polynomial chaos for multirate partial differential algebraic equations with random parameters. Applied Numerical Mathematics: Transactions of IMACS, 59(10): 2610–2624, October 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Pulch:2012:PCB

[Pul12]  Roland Pulch.  Polyno-

mial chaos for boundary value problems of dynamical systems. Applied Numerical Mathematics: Transactions of IMACS, 62(10): 1477–1490, October 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412001031.

Philippe:1993:PIS

[PV93] Bernard Philippe and Brigitte Vital. Parallel implementations for solving generalized eigenvalue problems with symmetric sparse matrices. Applied Numerical Mathematics: Transactions of IMACS, 12(5): 391–402, July 1993. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0168927493901006.

Pedas:2024:RSC

[PV24] Arvet Pedas and Mikk Vikerpuur. On the regularity of solutions to a class of nonlinear Volterra integral equations with singularities. Applied Numerical Mathematics: Transactions of IMACS, 204(??): 176–187, October 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742400148X.

Protopapas:2025:MMS

[PVH25] Eleftherios Protopapas, Panayiotis Vafeas, and Maria Hadjinicolaou. A mathematical model for studying the Red Blood Cell magnetic susceptibility. Applied Numerical Mathematics: Transactions of IMACS, 208 (part A) (??):356–365, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S016892742400120X.

Perez:2022:NSB

[PVM22] Aroldo Pérez and José Villa-Morales. A numerical scheme for the blow-up time of solutions of a system of nonlinear ordinary differential equations. Applied Numerical Mathematics: Transactions of IMACS, 171(??): 442–452, January 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421002713.

Podhaisky:1998:NEK

[PWS98] H. Podhaisky, R. Weiner, and B. A. Schmitt. Numerical experiments with Krylov integrators. Applied Numerical Mathematics: Transactions of IMACS, 28(2–4):413–425, October 1, 1998. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/

                                        store/apnum/sub/1998/28/
                                        2-4/923.pdf.

Podhaisky:2005:RTP

[PWS05] H. Podhaisky, R. Weiner, and B. A. Schmitt. Rosenbrocktype ‘peer’ two-step methods. Applied Numerical Mathematics: Transactions of IMACS, 53(2–4):409–420, May 2005. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Podhaisky:2006:LIT

[PWS06] H. Podhaisky, R. Weiner, and B. A. Schmitt. Linearlyimplicit two-step methods and their implementation in Nordsieck form. Applied Numerical Mathematics: Transactions of IMACS, 56(3–4):374–387, March/ April 2006. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Przybylowicz:2024:ASS

[PWX24] Paweł Przybyłowicz, Yue Wu, and Xinheng Xie. On approximation of solutions of stochastic delay differential equations via randomized Euler scheme. Applied Numerical Mathematics: Transactions of IMACS, 197(??):143–163, March 2024. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423002817.

Pan:2021:DMI

[PWY21] Yueyue Pan, Lifei Wu, and Xiaozhong Yang. A difference method with intrinsic parallelism for the variablecoefficient compound KdV– Burgers equation. Applied Numerical Mathematics: Transactions of IMACS, 169(??):201–220, November 2021. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742100177X.

Pan:2020:TLL

[PXHZ20] Kejia Pan, Junyi Xia, Dongdong He, and Qifeng Zhang. A three-level linearized difference scheme for nonlinear Schrödinger equation with absorbing boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 156(??): 32–49, October 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420301239.

Pugh:2021:SNS

[PYD21] M. C. Pugh, David Yan, and F. P. Dawson. A study of the numerical stability of an ImEx scheme with application to the Poisson–Nernst– Planck equations. Applied Numerical Mathematics: Transactions of IMACS, 163(??):239–253, May 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421000416.

Pollock:2020:DCP

[PZ20] Sara Pollock and Yunrong Zhu. Discrete comparison principles for quasilinear elliptic PDE. Applied Numerical Mathematics: Transactions of IMACS, 156(??): 106–124, October 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420301288.

Panaseti:2016:CCH

[PZMX16] Pandelitsa Panaseti, Antri Zouvani, Niall Madden, and Christos Xenophontos. A C^{1}- conforming hp finite element method for fourth order singularly perturbed boundary value problems. Applied Numerical Mathematics: Transactions of IMACS, 104(??):81–97, June 2016. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416000246.

Qamar:2009:NAP

[QAE^{+}09] S. Qamar, I. Angelov, M. P.

Elsner, A. Ashfaq, A. Seidel-Morgenstern, and G. Warnecke. Numerical approximations of a population balance model for coupled batch preferential crystallizers. Applied Numerical Mathematics: Transactions of IMACS, 59(3–4): 739–753, March/April 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Qiu:2017:TSA

[QAMX17] Hailong Qiu, Rong An, Liquan Mei, and Changfeng Xue. Two-step algorithms for the stationary incompressible Navier– Stokes equations with friction boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 120(??):97–114, October 2017. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417301186.

Qin:2012:ISP

[QC12] Hai-Hua Qin and David Colton. The inverse scattering problem for cavities. Applied Numerical Mathematics: Transactions of IMACS, 62(6):699–708, June 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927410001959.

Qin:2023:ATS

[QCW^{+}23] Yi Qin, Lele Chen, Yang Wang, Yi Li, and Jian Li. An adaptive time-stepping DLN decoupled algorithm for the coupled Stokes–Darcy model. Applied Numerical Mathematics: Transactions of IMACS, 188(??): 106–128, June 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742300065X.

Qin:2026:CRP

[QDA^{+}26] Zikang Qin, Xiaoyu Duan, Hengbin An, Haoxuan Zhang, and Shaoliang Hu. CSCMO: Relative permittivity-based complex shifted operator preconditioning method for solving time-harmonic Maxwell equations. Applied Numerical Mathematics: Transactions of IMACS, 219(??): 145–169, January 2026. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425001734.

Qin:2019:TFN

[QH19] Yi Qin and Yanren Hou. The time filter for the nonstationary coupled Stokes/ Darcy model. Applied Numerical Mathematics: Transactions of IMACS, 146(??): 260–275, December 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927419301874.

Qi:2022:UES

[QH22] Longzhao Qi and Yanren Hou. An unconditionally energy-stable linear Crank–Nicolson scheme for the Swift–Hohenberg equation. Applied Numerical Mathematics: Transactions of IMACS, 181(??):46– 58, November 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001490.

Qi:2024:NAS

[Qi24]  Longzhao Qi.  Numeri-

cal analysis of a secondorder energy-stable finite element method for the Swift–Hohenberg equation. Applied Numerical Mathematics: Transactions of IMACS, 197(??):119–142, March 2024. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423002933.

Qiu:2023:OAS

[Qiu23]  Hailong Qiu.  An op-

timally accurate secondorder time-stepping algorithm for the nonstationary magneto-hydrodynamics equations. Applied Numerical Mathematics: Transactions of IMACS, 184(??): 151–170, February 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422002586.

Qin:2015:IIS

[QL15] Hai-Hua Qin and Xiaodong Liu. The interior inverse scattering problem for cavities with an artificial obstacle. Applied Numerical Mathematics: Transactions of IMACS, 88(??):18– 30, February 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414001639.

Qin:2016:LSM

[QL16] Haihua Qin and Xiaodong Liu. The linear sampling method for inhomogeneous medium and buried objects from far field measurements. Applied Numerical Mathematics: Transactions of IMACS, 105(??): 82–95, July 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416300010.

Qaddouri:2008:OSM

[QLL^{+}08] Abdessamad Qaddouri, Lahcen Laayouni, Sébastien Loisel, Jean Côté, and Martin J. Gander. Optimized Schwarz methods with an overset grid for the shallowwater equations: preliminary results. Applied Numerical Mathematics: Transactions of IMACS, 58(4): 459–471, April 2008. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Quatember:2003:GCM

[QM03] Bernhard Quatember and Hannes Mühlthaler. Generation of CFD meshes from biplane angiograms: an example of image-based mesh generation and simulation. Applied Numerical Mathematics: Transactions of IMACS, 46(3–4): 379–397, September 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Qamar:2010:AVC

[QM10] Shamsul Qamar and Sidrah Mudasser. On the application of a variant CE/ SE method for solving two-dimensional ideal MHD equations. Applied Numerical Mathematics: Transactions of IMACS, 60(6): 587–606, June 2010. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Qiu:2019:MLS

[QM19] Hailong Qiu and Liquan Mei. Multi-level stabilized algorithms for the stationary incompressible Navier– Stokes equations with damping. Applied Numerical Mathematics: Transactions of IMACS, 143(??):188–202, September 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419300856.

Qin:2020:LTG

[QM20] Yonghui Qin and Heping Ma. Legendre-tau-Galerkin and spectral collocation method for nonlinear evolution equations. Applied Numerical Mathematics: Transactions of IMACS, 153(??): 52–65, July 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420300362.

Qiu:2015:DCS

[QMLC15] Hailong Qiu, Liquan Mei, Hui Liu, and Stephen Cartwright. A defectcorrection stabilized finite element method for Navier– Stokes equations with friction boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 90(??):9–21, April 2015. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0168927414001925.

Qiao:2023:SEN

[QNA23] L. Qiao, O. Nikan, and Z. Avazzadeh. Some efficient numerical schemes for approximating the nonlinear two-space dimensional extended Fisher–Kolmogorov equation. Applied Numerical Mathematics: Transactions of IMACS, 185(??): 466–482, March 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422003282.

Quenjel:2023:FID

[QPT23] El-Houssaine Quenjel, Patrick Perré, and Ian Turner. A 3D face interpolated discretisation method for simulating anisotropic diffusive processes on meshes coming from wood morphology. Applied Numerical Mathematics: Transactions of IMACS, 192(??): 280–296, October 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001757.

Qin:2003:ESU

[QR03] Yuming Qin and Jaime E. Muñoz Rivera. Exponential stability and universal attractors for the Navier–Stokes equations of compressible fluids between two horizontal parallel plates in. Applied Numerical Mathematics: Transactions of IMACS, 47(2):209–235, November 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Qiao:2024:FOE

[QR24] Zhijun Qiao and Enrique G. Reyes. Fifth-order equations of Camassa–Holm type and pseudo-peakons. Applied Numerical Mathematics: Transactions of IMACS, 199(??):165–176, May 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422003373.

Quenjel:2021:NFV

[Que21] El Houssaine Quenjel. Nonlinear finite volume discretization for transient diffusion problems on general meshes. Applied Numerical Mathematics: Transactions of IMACS, 161(??): 148–168, March 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303512.

Quirk:1996:PAG

[Qui96] James J. Quirk. A parallel adaptive grid algorithm for computational shock hydrodynamics. Applied Numerical Mathematics: Transactions of IMACS, 20(4):427– 453, June 3, 1996. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=4&aid=653. Adaptive

mesh refinement methods for CFD applications (Atlanta, GA, 1994).

Quyen:2019:FEA

[Quy19] Tran Nhan Tam Quyen. Finite element analysis for identifying the reaction coefficient in PDE from boundary observations. Applied Numerical Mathematics: Transactions of IMACS, 145(??):297–314, November 2019. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927419301679.

Qamar:2004:SMF

[QW04] Shamsul Qamar and Gerald Warnecke. Simulation of multicomponent flows using high order central schemes. Applied Numerical Mathematics: Transactions of IMACS, 50(2):183–201, August 2004. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Qi:2025:SCR

[QW25] Ruisheng Qi and Xiaojie Wang. Strong convergence rates of Galerkin finite element methods for SWEs with cubic polynomial nonlinearity. Applied Numerical Mathematics: Transactions of IMACS, 215(??): 112–137, September 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425000728.

Quan:2024:REP

[QWW24] Chaoyu Quan, Shijie Wang, and Xu Wu. Roundoff error problems in interpolation methods for time-fractional problems. Applied Numerical Mathematics: Transactions of IMACS, 203(??): 202–224, September 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424000928.

Qiao:2020:ADI

[QWX20] Leijie Qiao, Zhibo Wang, and Da Xu. An alternating direction implicit orthogonal spline collocation method for the two dimensional multiterm time fractional integrodifferential equation. Applied Numerical Mathematics: Transactions of IMACS, 151(??):199–212, May 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300040.

Qiu:2021:CNT

[QXG21] Wenlin Qiu, Da Xu, and Jing Guo. The Crank–Nicolsontype Sinc–Galerkin method for the fourth-order partial integro-differential equation with a weakly singular kernel. Applied Numerical Mathematics: Transactions of IMACS, 159(??): 239–258, January 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302877.

Qiao:2022:FSO

[QXQ22] Leijie Qiao, Da Xu, and Wenlin Qiu. The formally second-order BDF ADI difference/compact difference scheme for the nonlocal evolution problem in three-dimensional space. Applied Numerical Mathematics: Transactions of IMACS, 172(??): 359–381, February 2022.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421003068.

Qian:2025:EAP

[QZH25] Yanxia Qian, Yongchao Zhang, and Yunqing Huang. Error analysis of positivitypreserving energy stable schemes for the modified phase field crystal model. Applied Numerical Mathematics: Transactions of IMACS, 207(??):470–498, January 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424002472.

Qiu:2025:NAH

[QZM25] Wenlin Qiu, Xiangcheng Zheng, and Kassem Mustapha. Numerical approximations for a hyperbolic integrodifferential equation with a non-positive variablesign kernel and nonlinearnonlocal damping. Applied Numerical Mathematics: Transactions of IMACS, 213(??):61–76, July 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425000509.

Reksoprodjo:2003:DHO

[RA03] H. S. R. Reksoprodjo and R. K. Agarwal. Development of a higher-order accurate kinetic wave/particle flux-splitting algorithm for the Euler equations. Applied Numerical Mathematics: Transactions of IMACS, 46(3–4):399–410, September 2003. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Rang:2005:PIL

[RA05] J. Rang and L. Angermann.

Perturbation index of linear partial differential-algebraic equations. Applied Numerical Mathematics: Transactions of IMACS, 53(2–4): 437–456, May 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Russo:2009:FEA

[RA09] Anahí Dello Russo and Ana Alonso. Finite element approximation of Maxwell eigenproblems on curved Lipschitz polyhedral domains. Applied Numerical Mathematics: Transactions of IMACS, 59(8):1796– 1822, August 2009. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Roberts:2017:IDD

[RA17] J. A. Roberts and A. Al Themairi. Introducing delay dynamics to Bertalanffy’s spherical tumour growth model. Applied Numerical Mathematics: Transactions of IMACS, 114(??): 154–164, April 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416302033.

Rabinowitz:1994:AAS

[Rab94]  Philip Rabinowitz.  Ap-

plication of approximating splines for the solution of Cauchy singular integral equations. Applied Numerical Mathematics: Transactions of IMACS, 15(2):285– 297, September 13, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=15&
issue=2&aid=500. Innova-

tive methods in numerical analysis (Bressanone, 1992).

Rabago:2025:LTT

[Rab25] Julius Fergy Tiongson Rabago. Localization of tumor through a non-conventional numerical shape optimization technique. Applied Numerical Mathematics: Transactions of IMACS, 217(??): 135–171, November 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425001229.

Rahimkhani:2025:NMF

[Rah25]  Parisa Rahimkhani.  A

numerical method for Ψfractional integro-differential equations by Bell polynomials. Applied Numerical Mathematics: Transactions of IMACS, 207(??): 244–253, January 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424002484.

Ramakrishnan:1994:SUG

[Ram94] R. Ramakrishnan. Structured and unstructured grid adaptation schemes for numerical modeling of field problems. Applied Numerical Mathematics: Transactions of IMACS, 14(1–3): 285–310, April 14, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=14&
issue=1-3&aid=463. Pro-

ceedings of the Third ARO Workshop on Adaptive Methods for Partial Differential Equations (Troy, NY, 1992).

Ramage:1996:SMP

[Ram96] Alison Ramage. A structured matrix problem in dynamical systems. Applied Numerical Mathematics: Transactions of IMACS, 21(3): 291–301, August 20, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=21&
issue=3&aid=695.

Ramachandran:2012:FDC

[Ram12] M. P. Ramachandran. Fast derivative computation using smooth X-splines. Applied Numerical Mathematics: Transactions of IMACS, 62(11):1654–1662, November 2012. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412000761.

Rebollo:2023:CVS

[RÁM23] Tomás Chacón Rebollo, Enrique Delgado Ávila, and Macarena Gómez Mármol. On a certified VMS–Smagorinsky reduced basis model with LPS pressure stabilisation. Applied Numerical Mathematics: Transactions of IMACS, 185(??):365–385, March 2023. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422003257.

Rang:2015:APR

[Ran15] Joachim Rang. An analysis of the Prothero–Robinson example for constructing new adaptive ESDIRK methods of order 3 and 4. Applied Numerical Mathematics: Transactions of IMACS, 94(??):75–87, August 2015. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415000501.

Rang:2016:PRE

[Ran16] Joachim Rang. The Prothero and Robinson example: Convergence studies for Runge– Kutta and Rosenbrock– Wanner methods. Applied Numerical Mathematics: Transactions of IMACS, 108(??):37–56, October 2016. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416300605.

Ranner:2020:SFE

[Ran20] Thomas Ranner. A stable finite element method for low inertia undulatory locomotion in three dimensions. Applied Numerical Mathematics: Transactions of IMACS, 156(??): 422–445, October 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420301537.

Rocha:2018:NAC

[RAOC18] Adson M. Rocha, Juarez S. Azevedo, Saulo P. Oliveira, and Maicon R. Correa. Numerical analysis of a collocation method for functional integral equations. Applied Numerical Mathematics: Transactions of IMACS, 134(??):31–45, December 2018. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418301508.

Rahibe:1999:BPP

[RAS99] M. Rahibe, N. Aubry, and G. I. Sivashinsky. Bifurcations in a planar propagating flame as the size of the domain increases. Applied Numerical Mathematics: Transactions of IMACS, 31(1):103–115, September 1, 1999. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1999&volume=31&
                                        issue=1&aid=971.

Rathsfeld:2013:SDS

[Rat13] Andreas Rathsfeld. Shape derivatives for the scattering by biperiodic gratings. Applied Numerical Mathematics: Transactions of IMACS, 72(??):19–32, October 2013. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413000676.

Rodriguez:2012:RRE

[RB12] Marcos Rodríguez and Roberto Barrio. Reducing rounding errors and achieving Brouwer’s law with Taylor Series Method. Applied Numerical Mathematics: Transactions of IMACS, 62(8):1014–1024, August 2012. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412000645.

Rehm:1985:FDC

[RBBC85] Ronald G. Rehm, P. Darcy Barnett, Howard R. Baum, and Daniel M. Corley. Finite difference calculations of buoyant convection in an enclosure: Verification of the nonlinear algorithm. Applied Numerical Mathematics: Transactions of IMACS, 1(6):515– 529, November 1985. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Rubio:2002:TDC

[RBC02] Obidio Rubio, Elba Bravo, and Julio R. Claeyssen. Thermally driven cavity flow with Neumann condition for the pressure. Applied Numerical Mathematics: Transactions of IMACS, 40(1–2):327–336, January 2002. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        10/28/85/27/52/abstract.
                                        html.

Rodrigues:2025:RBI

[RBdM25] Savio B. Rodrigues, Giovanni Belloni Fernandes Braga, and Marcello Augusto Faraco de Medeiros. The residual balanced IMEX decomposition for singlydiagonally-implicit schemes. Applied Numerical Mathematics: Transactions of IMACS, 208 (part B)(??): 58–78, February 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424002678.

Ruiz-Baier:2015:NSM

[RBT15] Ricardo Ruiz-Baier and Héctor Torres. Numerical solution of a multidimensional sedimentation problem using finite volume-element methods. Applied Numerical Mathematics: Transactions of IMACS, 95(??): 280–291, September 2015. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414000294.

Rao:2018:NSS

[RC18] S. Chandra Sekhara Rao and Sheetal Chawla. Numerical solution of singularly perturbed linear parabolic system with discontinuous source term. Applied Numerical Mathematics: Transactions of IMACS, 127(??): 249–265, May 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418300151.

Rebollo:1998:MCF

[RCGM98] Tomás Chacón Rebollo, Daniel Franco Coronil, Francisco Ortegón Gallego, and Isabel Sánchez Muñoz. Modelling of compressible flows with highly oscillating initial data by homogenization. Applied Numerical Mathematics: Transactions of IMACS, 26(4):435–464, April 1, 1998. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1998/26/
4/825.pdf.

Rufai:2024:NPB

[RCR24] Mufutau Ajani Rufai, Bruno Carpentieri, and Higinio Ramos. A new pair of block techniques for direct integration of third-order singular IVPs. Applied Numerical Mathematics: Transactions of IMACS, 204(??): 222–231, October 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424001594.

Ranga:1996:ASQ

[RdAP96] A. Sri Ranga, E. X. L. de Andrade, and G. M. Phillips. Associated symmetric quadrature rules. Applied Numerical Mathematics: Transactions of IMACS, 21(2):175–183, July 8, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=21&
                                        issue=2&aid=690.

Razzaq:2012:FMT

[RDH^{+}12] M. Razzaq, H. Damanik, J. Hron, A. Ouazzi, and S. Turek. FEM multigrid techniques for fluidstructure interaction with application to hemodynamics. Applied Numerical Mathematics: Transactions of IMACS, 62(9):1156–1170, September 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411000092.

Rakhshan:2019:GLG

[RE19] Seyed Ali Rakhshan and Sohrab Effati. A generalized Legendre–Gauss collocation method for solving nonlinear fractional differential equations with time varying delays. Applied Numerical Mathematics: Transactions of IMACS, 146(??): 342–360, December 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927419301886.

Reed:2003:HFV

[Ree03] M. B. Reed. An H-form variant of the partitioned QN method. Applied Numerical Mathematics: Transactions of IMACS, 45(1): 79–85, April 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Reinhardt:1985:PEE

[Rei85] Hans-Jürgen Reinhardt. A posteriori error estimation for finite element modifications of line methods applied to singularly perturbed partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 1(2):145– 176, March 1985. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Reich:1999:PAI

[Rei99] Sebastian Reich. Preservation of adiabatic invariants under symplectic discretization. Applied Numerical Mathematics: Transactions of IMACS, 29(1):45– 55, January 1, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1999/29/
                                        1/898.pdf.

Rendl:1999:SPC

[Ren99] Franz Rendl. Semidefinite programming and combinatorial optimization. Applied Numerical Mathematics: Transactions of IMACS, 29(3):255–281, March 1, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.com/
                                        cas/tree/store/apnum/sub/
                                        1999/29/3/962.pdf.

Renka:2013:NLS

[Ren13] Robert J. Renka. Nonlinear least squares and Sobolev gradients. Applied Numerical Mathematics: Transactions of IMACS, 65(??): 91–104, March 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412002024.

Renka:2014:TRM

[Ren14] Robert J. Renka. A trust region method for constructing triangle-mesh approximations of parametric minimal surfaces. Applied Numerical Mathematics: Transactions of IMACS, 76(??):93– 100, February 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413001396.

Reisinger:2016:PCP

[RF16] C. Reisinger and P. A.

Forsyth. Piecewise constant policy approximations to Hamilton–Jacobi– Bellman equations. Applied Numerical Mathematics: Transactions of IMACS, 103(??):27–47, May 2016. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416000027.

Rodriguez-Fernandez:2024:SMD

[RFMV24] Alejandro Rodríguez-Fernández and Jesús Martín-Vaquero. On the stability of θmethods for DDEs and PDDEs. Applied Numerical Mathematics: Transactions of IMACS, 204(??): 312–328, October 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424001661.

Rollin:2002:VZL

[RG02] Stefan Röllin and Martin H. Gutknecht. Variations of Zhang’s Lanczos-type product method. Applied Numerical Mathematics: Transactions of IMACS, 41(1):119– 133, April 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        10/28/86/27/35/abstract.
                                        html.

Rebollo:2005:NSP

[RG05] T. Chacón Rebollo and D. Rodríguez Gómez. A numerical solver for the primitive equations of the ocean using term-by-term stabilization. Applied Numerical Mathematics: Transactions of IMACS, 55(1):1– 31, September 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Rahmoune:2020:SNM

[RG20] Azedine Rahmoune and Ahmed Guechi. Sinc– Nyström methods for Fredholm integral equations of the second kind over infinite intervals. Applied Numerical Mathematics: Transactions of IMACS, 157(??): 579–589, November 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302130.

Roul:2021:CFD

[RG21] Pradip Roul and V. M.

K. Prasad Goura. A compact finite difference scheme for fractional Black–Scholes option pricing model. Applied Numerical Mathematics: Transactions of IMACS, 166(??):40–60, August 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000878.

Ranjan:2022:UCN

[RG22] Kumar Rajeev Ranjan and S. Gowrisankar. Uniformly convergent NIPG method for singularly perturbed convection diffusion problem on Shishkin type meshes. Applied Numerical Mathematics: Transactions of IMACS, 179(??): 125–148, September 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001088.

Roul:2019:NHO

[RGA19] Pradip Roul, V. M. K. Prasad Goura, and Ravi Agarwal. A new high order numerical approach for a class of nonlinear derivative dependent singular boundary value problems. Applied Numerical Mathematics: Transactions of IMACS, 145(??): 315–341, November 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419301680.

Rathan:2020:SSI

[RGB20] Samala Rathan, Naga Raju Gande, and Ashlesha A. Bhise. Simple smoothness indicator WENO-Z scheme for hyperbolic conservation laws. Applied Numerical Mathematics: Transactions of IMACS, 157(??): 255–275, November 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420301884.

Roul:2021:HAN

[RGK21] Pradip Roul, V. M. K. Prasad Goura, and Klaus Kassner. A high accuracy numerical approach for electrohydrodynamic flow of a fluid in an ion-drag configuration in a circular cylindrical conduit. Applied Numerical Mathematics: Transactions of IMACS, 165(??): 303–321, July 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000611.

Rodrigo:2016:LFA

[RGL16] C. Rodrigo, F. J. Gaspar, and F. J. Lisbona. On a local Fourier analysis for overlapping block smoothers on triangular grids. Applied Numerical Mathematics: Transactions of IMACS, 105(??):96–111, July 2016. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416300022.

Roul:2019:DSA

[RGMO19] Pradip Roul, V. M. K. Prasad Goura, Harshita Madduri, and K. Obaidurrahman. Design and stability analysis of an implicit nonstandard finite difference scheme for fractional neutron point kinetic equation. Applied Numerical Mathematics: Transactions of IMACS, 145(??): 201–226, November 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927419301527.

Ranocha:2018:SAD

[RGÖS18] Hendrik Ranocha, Jan Glaubitz, Philipp Öffner, and Thomas Sonar. Stability of artificial dissipation and modal filtering for flux reconstruction schemes using summation-by-parts operators. Applied Numerical Mathematics: Transactions of IMACS, 128(??): 1–23, June 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892741830028X.

Rahhali:2025:PCS

[RGS25] M. A. Rahhali, T. Garcia, and P. Spiteri. Parallel cloud solution of large algebraic multivalued systems. Applied Numerical Mathematics: Transactions of IMACS, 208 (part A) (??):366–389, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742400062X.

Rich:1992:ADM

[RH92] Lawrence C. Rich and David R. Hill. Automatic differentiation in MATLAB. Applied Numerical Mathematics: Transactions of IMACS, 9(1):33–43, January 1992. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Rhanizar:1997:EMO

[Rha97] B. Rhanizar. On extrapolation methods in optimization. Applied Numerical Mathematics: Transactions of IMACS, 25(4):485– 498, November 10, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=25&
issue=4&aid=804.

Rhanizar:1999:HPS

[Rha99] B. Rhanizar. Hybrid procedures for solving some unconstrained nonlinear optimization problems. Applied Numerical Mathematics: Transactions of IMACS, 30(4):459–474, July 1, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1999&volume=30&
issue=4&aid=934.

Ramirez:2002:NSH

[RI02] Isabel Ramirez and Jorg Imberger. The numerical simulation of the hydrodynamics of Barbamarco Lagoon, Italy. Applied Numerical Mathematics: Transactions of IMACS, 40(1–2):273–289, January 2002. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        10/28/85/27/49/abstract.
                                        html.

Riahi:2022:PPT

[Ria22] Mohamed Kamel Riahi. PiTSBiCG: Parallel in time Stable Bi-Conjugate gradient algorithm. Applied Numerical Mathematics: Transactions of IMACS, 181(??): 225–233, November 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422001623.

Richter:1991:EFE

[Ric91] Gerard R. Richter. Explicit finite element methods for scalar hyperbolic equations. Applied Numerical Mathematics: Transactions of IMACS, 7(6):509– 521, June 1991. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Richter:1994:EFE

[Ric94] Gerard R. Richter. An explicit finite element method for the wave equation. Applied Numerical Mathematics: Transactions of IMACS, 16(1–2):65–80, December 1, 1994. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=16&
issue=1-2&aid=524.  A

Festschrift to honor Professor Robert Vichnevetsky on his 65th birthday.

Rickard:2008:EWB

[Ric08]  Yotka Rickard.  An ef-

ficient wavelet-based solution of electromagnetic field problems. Applied Numerical Mathematics: Transactions of IMACS, 58(4): 472–485, April 2008. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Riley:1992:NSV

[Ril92] Bruce V. Riley. The numerical solution of Volterra integral equations with nonsmooth solutions based on sinc approximation. Applied Numerical Mathematics: Transactions of IMACS, 9(3–5):249–257, April 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/016892749290019A.

International Conference on the Numerical Solution of Volterra and Delay Equations (Tempe, AZ, 1990).

Risch:2005:SNL

[Ris05]  Uwe Risch.  Supercon-

vergence of a nonconforming low order finite element. Applied Numerical Mathematics: Transactions of IMACS, 54(3– 4):324–338, August 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Rivara:2009:LBA

[Riv09] Maria-Cecilia Rivara. Leppbisection algorithms, applications and mathematical properties. Applied Numerical Mathematics: Transactions of IMACS, 59(9): 2218–2235, September 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Rogers:1991:UDS

[RK91] Stuart E. Rogers and Dochan Kwak. An upwind differencing scheme for the incompressible Navier–Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 8(1):43– 64, August 1991. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Rao:2008:EBS

[RK08] S. Chandra Sekhara Rao and Mukesh Kumar. Exponential B-spline collocation method for self-adjoint singularly perturbed boundary value problems. Applied Numerical Mathematics: Transactions of IMACS, 58(10): 1572–1581, October 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Roberts:2020:DMM

[RKR20] J. A. Roberts, N. I. Kavallaris, and A. P. Rowntree. A discrete mutualism model: analysis and exploration of a financial application. Applied Numerical Mathematics: Transactions of IMACS, 149(??):141–152, ???? 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419302478.

Reshniak:2015:SSM

[RKVZ15] V. Reshniak, A. Q. M.

Khaliq, D. A. Voss, and G. Zhang. Split-step Milstein methods for multi-channel stiff stochastic differential systems. Applied Numerical Mathematics: Transactions of IMACS, 89(??): 1–23, March 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414001767.

Ramirez:1986:RQS

[RL86] V. Ramírez and J. Lorente.

C^{1} rational quadratic spline interpolation to convex data. Applied Numerical Mathematics: Transactions of IMACS, 2(1):37–42, February 1986. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Rincon:2006:EEV

[RL06] M. A. Rincon and I.-Shih Liu. Error estimates of vibrations of elastic string with moving ends. Applied Numerical Mathematics: Transactions of IMACS, 56(6):745–755, June 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ran:2021:FDS

[RL21] Maohua Ran and Xiaojuan Lei. A fast difference scheme for the variable coefficient timefractional diffusion wave equations. Applied Numerical Mathematics: Transactions of IMACS, 167(??):31– 44, September 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421001203.

Rincon:2019:NAN

[RLHC19] M. A. Rincon, I.-S. Liu, W. R. Huarcaya, and B. A. Carmo. Numerical analysis for a nonlinear model of elastic strings with moving ends. Applied Numerical Mathematics: Transactions of IMACS, 135(??): 146–164, January 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418301909.

Roccia:2024:HTP

[RLMG24] Bruno A. Roccia, Carmina Alturria Lanzardo, Fernando D. Mazzone, and Cristian G. Gebhardt. On the homogeneous torsion problem for heterogeneous and orthotropic cross-sections: Theoretical and numerical aspects. Applied Numerical Mathematics: Transactions of IMACS, 201(??): 579–607, July 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000783.

Rocca:2006:STC

[RLSS06] Giulio Della Rocca, Maria Carmela Lombardo, Marco Sammartino, and Vincenzo Sciacca. Singularity tracking for Camassa–Holm and Prandtl’s equations. Applied Numerical Mathematics: Transactions of IMACS, 56(8):1108–1122, August 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Roy:2025:RVI

[RM25] Tapas Roy and Dilip Kumar Maiti. Refined variational iteration method through optimal linear operator and Lagrange multiplier: Local/ global convergence. Applied Numerical Mathematics: Transactions of IMACS, 212(??):1–28, June 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425000133.

Radid:2004:ART

[RMCG04] A. Radid, J. C. Miellou, A. Chambaudet, and M. Grivet. About a rock thermochronologycal model: Laslett’s law study and time equivalent method. Applied Numerical Mathematics: Transactions of IMACS, 50(2):203–226, August 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ramezani:2020:SHO

[RMH20] M. Ramezani, R. Mokhtari, and G. Haase. Some high order formulae for approximating Caputo fractional derivatives. Applied Numerical Mathematics: Transactions of IMACS, 153(??): 300–318, July 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300611.

Roy:2009:NMH

[RMK09] Suman Roy, A. S. Vasudeva Murthy, and Ramesh B. Kudenatti. A numerical method for the hyperbolic-heat conduction equation based on multiple scale technique. Applied Numerical Mathematics: Transactions of IMACS, 59(6):1419–1430, June 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Rebollo:2012:NSP

[RMM12] Tomás Chacón Rebollo, Macarena Gómez Mármol, and Isabel Sánchez Muñoz. Numerical solution of the Primitive Equations of the ocean by the Orthogonal Sub-Scales VMS method. Applied Numerical Mathematics: Transactions of IMACS, 62(4):342–359, April 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0168927411001061.

Roman:2017:NAG

[RMS17] Fabio Roman, Carla Manni, and Hendrik Speleers. Numerical approximation of GB-splines by a convolutional approach. Applied Numerical Mathematics: Transactions of IMACS, 116(??):273–285, June 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416302185.

Rotaru:2004:FAF

[RN04] Tiberiu Rotaru and Hans-Heinrich Nägeli. Fast algorithms for fair dynamic load redistribution in heterogeneous environments. Applied Numerical Mathematics: Transactions of IMACS, 49(1):81–95, April 2004. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Rezazadeh:2022:JCM

[RN22] Tohid Rezazadeh and Esmaeil Najafi. Jacobi collocation method and smoothing transformation for numerical solution of neutral nonlinear weakly singular Fredholm integro-differential equations. Applied Numerical Mathematics: Transactions of IMACS, 181(??): 135–150, November 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422001507.

Raina:2025:WGF

[RN25] Aayushman Raina and Srinivasan Natesan. A weak Galerkin finite element method for fourth-order parabolic singularly perturbed problems on layer adapted Shishkin mesh. Applied Numerical Mathematics: Transactions of IMACS, 207(??): 520–533, January 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424002563.

Ramos:2022:RES

[RNG22] Eduardo Ramos, Victor Nolasco, and Marcio Gameiro. Rigorous enclosures of solutions of Neumann boundary value problems. Applied Numerical Mathematics: Transactions of IMACS, 180(??):104–119, October 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001349.

Ruijter:2016:NFM

[RO16] M. J. Ruijter and C. W.

Oosterlee. Numerical Fourier method and second-order Taylor scheme for backward SDEs in finance. Applied Numerical Mathematics: Transactions of IMACS, 103(??):1–26, May 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416000040.

Roberts:2001:HDE

[Rob01] A. J. Roberts. Holistic discretization ensures fidelity to Burgers’ equation. Applied Numerical Mathematics: Transactions of IMACS, 37(3):371–396, May 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/10/28/73/31/
                                        34/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/10/28/73/31/34/article.
                                        pdf.

Roberts:2010:CIE

[Rob10] A. J. Roberts. Choose interelement coupling to preserve self-adjoint dynamics in multiscale modelling and computation. Applied Numerical Mathematics: Transactions of IMACS, 60(10):949– 973, October 2010. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Rahimkhani:2017:FOB

[ROB17] Parisa Rahimkhani, Yadollah Ordokhani, and Esmail Babolian. Fractionalorder Bernoulli functions and their applications in solving fractional Fredholm– Volterra integro-differential equations. Applied Numerical Mathematics: Transactions of IMACS, 122(??):66– 81, December 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741730168X.

Rogov:2019:DDP

[Rog19] B. V. Rogov. Dispersive and dissipative properties of the fully discrete bicompact schemes of the fourth order of spatial approximation for hyperbolic equations. Applied Numerical Mathematics: Transactions of IMACS, 139(??):136–155, May 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419300194.

Rahimkhani:2019:ICC

[ROL19] P. Rahimkhani, Y. Ordokhani, and P. M. Lima. An improved composite collocation method for distributedorder fractional differential equations based on fractional Chelyshkov wavelets. Applied Numerical Mathematics: Transactions of IMACS, 145(??):1–27, November 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927419301382.

Roop:2020:DFB

[Roo20] John P. Roop. Discretization of fractional boundary value problems using split operator local extension problems. Applied Numerical Mathematics: Transactions of IMACS, 152(??): 267–274, June 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419303319.

Rorro:2006:ASE

[Ror06] M. Rorro. An approximation scheme for the effective Hamiltonian and applications. Applied Numerical Mathematics: Transactions of IMACS, 56(9): 1238–1254, September 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Rose:1993:ESS

[Ros93] Milton E. Rose. An enthalpy scheme for Stefan problems in several dimensions. Applied Numerical Mathematics: Transactions of IMACS, 12(1–3):229–238, May 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/016892749390120G.

Special issue to honor Professor Saul Abarbanel on his sixtieth birthday (Neveh, 1992).

Roul:2020:FON

[Rou20a] Pradip Roul. A fourthorder non-uniform mesh optimal B-spline collocation method for solving a strongly nonlinear singular boundary value problem describing electrohydrodynamic flow of a fluid. Applied Numerical Mathematics: Transactions of IMACS, 153(??): 558–574, July 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300854.

Roul:2020:HAN

[Rou20b] Pradip Roul. A high accuracy numerical method and its convergence for timefractional Black–Schole’s equation governing European options. Applied Numerical Mathematics: Transactions of IMACS, 151(??):472–493, May 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419303125.

Rozza:2005:RBM

[Roz05] Gianluigi Rozza. Reducedbasis methods for elliptic equations in sub-domains with a posteriori error bounds and adaptivity. Applied Numerical Mathematics: Transactions of IMACS, 55(4):403–424, December 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Raha:2001:CPS

[RP01] Soumyendu Raha and Linda R. Petzold. Constraint partitioning for structure in pathconstrained dynamic optimization problems. Applied Numerical Mathematics: Transactions of IMACS, 39(1):105–126, October 2001. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        10/28/79/27/33/abstract.
                                        html.

Rad:2017:NPA

[RP17] Jamal Amani Rad and Kourosh Parand. Numerical pricing of American options under two stochastic factor models with jumps using a meshless local Petrov– Galerkin method. Applied Numerical Mathematics: Transactions of IMACS, 115(??):252–274, May 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417300302.

Reuter:2000:HOV

[RR00] Jörg Reuter and Dietmar Rempfer. High order vorticity-velocity method for the simulation of pipe flow transition. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4): 105–111, May 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/61/27/
34/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/61/27/34/article.
pdf.

Rapetti:2014:QSM

[RR14] Francesca Rapetti and Germain Rousseaux. On quasistatic models hidden in Maxwell’s equations. Applied Numerical Mathematics: Transactions of IMACS, 79(??):92–106, May 2014. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412002036.

Rabiei:2021:FOB

[RR21] Kobra Rabiei and Mohsen Razzaghi. Fractional-order Boubaker wavelets method for solving fractional Riccati differential equations. Applied Numerical Mathematics: Transactions of IMACS, 168(??):221–234, October 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742100146X.

Roul:2020:DAN

[RREP^{+}20] Pradip Roul, Vikas Rohil, Gilberto Espinosa-Paredes, V. M. K. Prasad Goura, R. S. Gedam, and K. Obaidurrahman. Design and analysis of a numerical method for fractional neutron diffusion equation with delayed neutrons. Applied Numerical Mathematics: Transactions of IMACS, 157(??): 634–653, November 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302075.

Rivara:2012:MPL

[RRMJ12] Maria-Cecilia Rivara, Pedro Rodriguez, Rafael Montenegro, and Gaston Jorquera. Multithread parallelization of Lepp-bisection algorithms. Applied Numerical Mathematics: Transactions of IMACS, 62(4): 473–488, April 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411001292.

Ryabenkii:2000:DPH

[RS00] Victor S. Ryaben’kii and Ivan L. Sofronov. Difference potentials for the Helmholtz equation in exterior domains. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4): 533–540, May 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/61/27/
83/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/61/27/83/article.
pdf.

Ragulskis:2008:GCT

[RS08a] M. Ragulskis and L. Saunoriene.

Generalisations of the compound trapezoidal rule. Applied Numerical Mathematics: Transactions of IMACS, 58(1):40–58, January 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Robbe:2008:UNB

[RS08b] M. Robbé and M. Sadkane.

Use of near-breakdowns in the block Arnoldi method for solving large Sylvester equations. Applied Numerical Mathematics: Transactions of IMACS, 58(4): 486–498, April 2008. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Radha:2009:LRF

[RS09] R. Radha and S. Sivananthan. Local reconstruction of a function from a nonuniform sampled data. Applied Numerical Mathematics: Transactions of IMACS, 59(2):393–403, February 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ryckelynck:2014:QC

[RS14] P. Ryckelynck and L. Smoch. Quadratic choreographies. Applied Numerical Mathematics: Transactions of IMACS, 75(??):108–122, January 2014. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413000901.

Ray:2020:AFE

[RS20] Tanushree Ray and Rajen Kumar Sinha. An adaptive finite element method for semilinear parabolic interface problems with nonzero flux jump. Applied Numerical Mathematics: Transactions of IMACS, 153(??):381–398, July 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300684.

Reichel:2021:NRG

[RS21] Lothar Reichel and Miodrag M. Spalević. A new representation of generalized averaged Gauss quadrature rules. Applied Numerical Mathematics: Transactions of IMACS, 165(??): 614–619, July 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303664.

Risthaus:2022:SPF

[RS22] Lennart Risthaus and Matti Schneider. Solving phasefield models in the tensor train format to generate microstructures of bicontinuous composites. Applied Numerical Mathematics: Transactions of IMACS, 178(??): 262–279, August 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422000952.

Ruuskanen:2006:UPT

[RSD^{+}06] A. R. Ruuskanen, A. Seppänen, S. Duncan, E. Somersalo, and J. P. Kaipio. Using process tomography as a sensor for optimal control. Applied Numerical Mathematics: Transactions of IMACS, 56(1):37–54, January 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Rozloznik:2014:NIR

[RSK14] Miroslav Rozlozník, Alicja Smoktunowicz, and Jirí Kopal. A note on iterative refinement for seminormal equations. Applied Numerical Mathematics: Transactions of IMACS, 75(??):167– 174, January 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413001190.

Rathi:2024:VMS

[RSK24] Anil Rathi, Dipak Kumar Sahoo, and B. V. Rathish Kumar. Variational multiscale stabilized finite element analysis of transient MHD Stokes equations with application to multiply driven cavity flow. Applied Numerical Mathematics: Transactions of IMACS, 198(??): 43–74, April 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742300315X.

Rameau:1989:PBP

[RSL89] J. F. Rameau and C. Schmidt-Lainé. Perturbed bifurcation phenomenon in a diffusion flame problem. Applied Numerical Mathematics: Transactions of IMACS, 5(3):237–255, May 1989. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Rathore:2023:FNM

[RSR23] Ajay Singh Rathore, Vembu Shanthi, and Higinio Ramos. A fitted numerical method for a singularly perturbed Fredholm integro-differential equation with discontinuous source term. Applied Numerical Mathematics: Transactions of IMACS, 185(??): 88–100, March 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422003051.

Reichel:2012:TRB

[RSY12] Lothar Reichel, Fiorella Sgallari, and Qiang Ye. Tikhonov regularization based on generalized Krylov subspace methods. Applied Numerical Mathematics: Transactions of IMACS, 62(9): 1215–1228, September 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927410001765.

Ryabenkii:1995:ENT

[RT95] V. S. Ryaben′kiĭ and S. V.

Tsynkov. An effective numerical technique for solving a special class of ordinary difference equations. Applied Numerical Mathematics: Transactions of IMACS, 18(4):489–501, November 9, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=18&
issue=4&aid=629. See er-

rata [Tsy96].

Rogava:2014:CSD

[RT14] J. Rogava and M. Tsiklauri. Convergence of a semidiscrete scheme for an abstract nonlinear second order evolution equation. Applied Numerical Mathematics: Transactions of IMACS, 75(??):22–36, January 2014. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741200164X.

Ren:2020:CMM

[RT20] Quanwei Ren and Hongjiong Tian. Compensated θ-Milstein methods for stochastic differential equations with Poisson jumps. Applied Numerical Mathematics: Transactions of IMACS, 150(??):27–37, April 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741930248X.

Roul:2019:NOF

[RTA19] Pradip Roul, Kiran Thula, and Ravi Agarwal. Nonoptimal fourth-order and optimal sixth-order B-spline collocation methods for Lane–Emden boundary value problems. Applied Numerical Mathematics: Transactions of IMACS, 145(??): 342–360, November 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927419301114.

Reich:2023:VIS

[RTH23] Simeon Reich, Truong Minh Tuyen, and Nguyen Song Ha. Variational inequalities over the solution sets of split variational inclusion problems. Applied Numerical Mathematics: Transactions of IMACS, 192(??): 319–336, October 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001812.

Ryabenkii:2001:LTN

[RTT01] V. S. Ryaben’kii, S. V.

Tsynkov, and V. I. Turchaninov. Long-time numerical computation of wave-type solutions driven by moving sources. Applied Numerical Mathematics: Transactions of IMACS, 38(1–2): 187–222, July 2001. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/77/27/
37/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/77/27/37/article.
pdf.

Roos:2015:UCD

[RTU15] Hans-Goerg Roos, Ljiljana Teofanov, and Zorica Uzelac. Uniformly convergent difference schemes for a singularly perturbed third order boundary value problem. Applied Numerical Mathematics: Transactions of IMACS, 96(??):108– 117, October 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415000938.

Rice:2000:IRM

[RTV00] J. R. Rice, P. Tsompanopoulou, and E. Vavalis. Interface relaxation methods for elliptic differential equations. Applied Numerical Mathematics: Transactions of IMACS, 32(2):219– 245, February 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/59/27/
32/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/59/27/32/article.
pdf.

Rice:2002:FTI

[RTV02] J. R. Rice, P. Tsompanopoulou, and E. Vavalis. Fine tuning interface relaxation methods for elliptic differential equations. Applied Numerical Mathematics: Transactions of IMACS, 43(4):459–481, December 2002. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Rauch:2021:SCM

[RTW21] Reuben Rauch, Manfred R.

Trummer, and J. F. Williams. A spectral collocation method for mixed functional differential equations. Applied Numerical Mathematics: Transactions of IMACS, 161(??):101–110, March 2021. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303184.

Ryabenkii:2007:DFD

[RU07] V. S. Ryaben’kii and S. V.

Utyuzhnikov. Differential and finite-difference problems of active shielding. Applied Numerical Mathematics: Transactions of IMACS, 57(4):374–382, April 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ryabenkii:2015:AMD

[RU15] V. S. Ryaben’kii and S. V.

Utyuzhnikov. An algorithm of the method of difference potentials for domains with cuts. Applied Numerical Mathematics: Transactions of IMACS, 93(??): 254–261, July 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414000233.

Reichel:2021:TKS

[RU21] Lothar Reichel and Ugochukwu O. Ugwu. Tensor Krylov subspace methods with an invertible linear transform product applied to image processing. Applied Numerical Mathematics: Transactions of IMACS, 166(??): 186–207, August 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421001069.

Rihan:2025:SSM

[RU25] Fathalla A. Rihan and K. Udhayakumar. Splitstep θ-method for stochastic pantograph differential equations: Convergence and mean-square stability analysis. Applied Numerical Mathematics: Transactions of IMACS, 217(??):1– 17, November 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425001163.

Rump:1987:SLS

[Rum87] Siegfried M. Rump. Solution of linear systems with verified accuracy. Applied Numerical Mathematics: Transactions of IMACS, 3(3):233–241, June 1987. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/016892748790050X.

Russell:1995:PIS

[Rus95] William S. Russell. Polynomial interpolation schemes for internal derivative distributions on structured grids. Applied Numerical Mathematics: Transactions of IMACS, 17(2):129– 171, June 13, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=17&
issue=2&aid=562.

Rebollo:2004:SNO

[RV04] Tomás Chacón Rebollo and Eliseo Chacón Vera. Study of a non-overlapping domain decomposition method: Poisson and Stokes problems. Applied Numerical Mathematics: Transactions of IMACS, 48(2):169– 194, February 2004. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Rebollo:2005:SNO

[RV05a] Tomás Chacón Rebollo and Eliseo Chacón Vera. Study of a non-overlapping domain decomposition method: Steady Navier–Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 55(1): 100–124, September 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Roland:2005:NIS

[RV05b] Ch. Roland and R. Varadhan. New iterative schemes for nonlinear fixed point problems, with applications to problems with bifurcations and incomplete-data problems. Applied Numerical Mathematics: Transactions of IMACS, 55(2):215– 226, October 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Rodriguez:2009:FBA

[RV09] Ana Alonso Rodríguez and Alberto Valli. A FEM– BEM approach for electromagnetostatics and timeharmonic eddy-current problems. Applied Numerical Mathematics: Transactions of IMACS, 59(9): 2036–2049, September 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Rodriguez:2015:FEP

[RV15] Ana Alonso Rodríguez and Alberto Valli. Finite element potentials. Applied Numerical Mathematics: Transactions of IMACS, 95(??):2– 14, September 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741400124X.

Rayal:2022:TDG

[RV22] Ashish Rayal and Sag Ram Verma. Two-dimensional Gegenbauer wavelets for the numerical solution of tempered fractional model of the nonlinear Klein–Gordon equation. Applied Numerical Mathematics: Transactions of IMACS, 174(??): 191–220, April 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422000216.

Riemslagh:2000:ALE

[RVD00] Kris Riemslagh, Jan Vierendeels, and Erik Dick. An arbitrary Lagrangian–Eulerian finite-volume method for the simulation of rotary displacement pump flow. Applied Numerical Mathematics: Transactions of IMACS, 32(4):419–433, April 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/59/32/
32/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/59/32/32/article.
pdf.

Retamoso:2002:EBC

[RVdCVR02] Mário R. Retamoso, Marco Túlio Vilhena, Haroldo F. de Campos Velho, and Fernando M. Ramos. Estimation of boundary condition in hydrologic optics. Applied Numerical Mathematics: Transactions of IMACS, 40(1– 2):87–100, January 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/gej-
                                        ng/10/10/28/85/27/35/abstract.
                                        html.

Rozema:2023:QSC

[RVM23] Gerk Rozema, Arthur E. P. Veldman, and Natasha M. Maurits. Quasi-simultaneous coupling methods for partitioned problems in computational hemodynamics. Applied Numerical Mathematics: Transactions of IMACS, 184(??):461–481, February 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422002884.

Robinson:1987:HOO

[RW87] Allan R. Robinson and Leonard J. Walstad. The Harvard open ocean model: calibration and application to dynamical process, forecasting, and data assimilation studies. Applied Numerical Mathematics: Transactions of IMACS, 3(1–2): 89–131, May 1987. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0168927487900080.

Rong:2008:NAA

[RX08] Zhijian Rong and Chuanju Xu. Numerical approximation of acoustic waves by spectral element methods. Applied Numerical Mathematics: Transactions of IMACS, 58(7):999–1016, July 2008. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Rathinasamy:2013:SSC

[RY13] A. Rathinasamy and Baojian Yin. Strong solutions of a class of hybrid diffusion processes with state-dependent regime-switching. Applied Numerical Mathematics: Transactions of IMACS, 72(??):72–90, October 2013. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413000640.

Ryabenkii:2000:NTD

[Rya00] Victor Ryaben’kii. Nonreflecting time-dependent boundary conditions on artificial boundaries of varying location and shape. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4):481–492, May 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/61/27/
77/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/61/27/77/article.
pdf.

Radvogin:2000:LIS

[RZ00] Yulian B. Radvogin and Nikolai A. Zaitsev. A locally implicit second order accurate difference scheme for solving 2D time-dependent hyperbolic systems and Euler equations. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4):525–532, May 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/61/27/
82/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/61/27/82/article.
pdf.

Redivo-Zaglia:2004:PSC

[RZ04] M. Redivo-Zaglia. Pseudo-Schur complements and their properties. Applied Numerical Mathematics: Transactions of IMACS, 50(3–4): 511–519, September 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Rohde:2015:RRS

[RZ15] Christian Rohde and Christoph Zeiler. A relaxation Riemann solver for compressible two-phase flow with phase transition and surface tension. Applied Numerical Mathematics: Transactions of IMACS, 95(??): 267–279, September 2015. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414000725.

Ran:2018:NCD

[RZ18] Maohua Ran and Chengjian Zhang. New compact difference scheme for solving the fourth-order time fractional sub-diffusion equation of the distributed order. Applied Numerical Mathematics: Transactions of IMACS, 129(??):58–70, July 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418300679.

Ran:2021:PFE

[RZS21] Hongtao Ran, Bo Zheng, and Yueqiang Shang. A parallel finite element variational multiscale method for the Navier–Stokes equations with nonlinear slip boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 168(??):

274–292, October 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421001720.

Sallam:1990:NSG

[SA90] S. Sallam and W. Ameen. Numerical solution of general n th-order differential equations via splines. Applied Numerical Mathematics: Transactions of IMACS, 6(3):225–238, March 1990. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Sherwin:2000:UNS

[SA00] Spencer J. Sherwin and Mark Ainsworth. Unsteady Navier–Stokes solvers using hybrid spectral/hp element methods. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4): 357–363, May 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/21/61/27/
                                        63/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/21/61/27/63/article.
                                        pdf.

Szabo:2005:IUF

[SA05] Barna Szabó and Ricardo Actis. On the importance and uses of feedback information in FEA. Applied Numerical Mathematics: Transactions of IMACS, 52(2– 3):219–234, February 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Siomina:2008:LOU

[SA08] Iana Siomina and Sven Ahlinder. Lean optimization using supersaturated experimental design. Applied Numerical Mathematics: Transactions of IMACS, 58(1):1–15, January 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Schierz:2012:SOM

[SA12a] Tom Schierz and Martin Arnold. Stabilized overlapping modular time integration of coupled differentialalgebraic equations. Applied Numerical Mathematics: Transactions of IMACS, 62(10):1491–1502, October 2012. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412001067.

Schiller:2012:CCA

[SA12b] H. Schiller and M. Arnold.

Convergence of continuous approximations for discontinuous ODEs. Applied Numerical Mathematics: Transactions of IMACS, 62(10):1503–1514, October 2012. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412001079.

Soori:2018:SON

[SA18] Z. Soori and A. Aminataei. Sixth-order nonuniform combined compact difference scheme for multi-term time fractional diffusion-wave equation. Applied Numerical Mathematics: Transactions of IMACS, 131(??):72–94, September 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0168927418300953.

Soori:2019:NAC

[SA19] Z. Soori and A. Aminataei. A new approximation to Caputo-type fractional diffusion and advection equations on non-uniform meshes. Applied Numerical Mathematics: Transactions of IMACS, 144(??):21–41, October 2019. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927419301217.

Singh:2020:NAC

[SA20] Brajesh Kumar Singh and Saloni Agrawal. A new approximation of conformable time fractional partial differential equations with proportional delay. Applied Numerical Mathematics: Transactions of IMACS, 157(??): 419–433, November 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302014.

Santos:2021:CAE

[SA21] Ricardo Santos and Leonardo Alves. A comparative analysis of explicit, IMEX and implicit strong stability preserving Runge–Kutta schemes. Applied Numerical Mathematics: Transactions of IMACS, 159(??): 204–220, January 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742030283X.

Sahihi:2020:SSS

[SAA20] Hussein Sahihi, Tofigh Allahviranloo, and Saeid Abbasbandy. Solving system of second-order BVPs using a new algorithm based on reproducing kernel Hilbert space. Applied Numerical Mathematics: Transactions of IMACS, 151(??): 27–39, May 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892741930340X.

Sacco:1993:CSO

[Sac93] Riccardo Sacco. Convergence of a second-order accurate Petrov–Galerkin scheme for convection-diffusion problems in semiconductors. Applied Numerical Mathematics: Transactions of IMACS, 11(6):517–528, April 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Sadaka:1996:PAV

[Sad96] R. Sadaka. Padé approximation of vector functions. Applied Numerical Mathematics: Transactions of IMACS, 21(1):57–70, June 6, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=21&
                                        issue=1&aid=684.

Sadaka:1997:ACC

[Sad97] R. Sadaka. An algorithm for constructing a class of Padé approximants of vector functions. Applied Numerical Mathematics: Transactions of IMACS, 24(4): 483–499, August 12, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-

bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=24&
issue=4&aid=764.

Saeed:2014:CTB

[Sae14] Khalid Saeed. Carathéodory– Toeplitz based mathematical methods and their algorithmic applications in biometric image processing. Applied Numerical Mathematics: Transactions of IMACS, 75(??):2–21, January 2014. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741200075X.

Salas:1986:MSS

[SAG86] M. D. Salas, S. Abarbanel, and David Gottlieb. Multiple steady states for characteristic initial value problems. Applied Numerical Mathematics: Transactions of IMACS, 2(3–5):193–210, October 1986. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Sharifi:2024:CDI

[SAH24] M. Sharifi, A. Abdi, and G. Hojjati. On the construction of diagonally implicit two-step peer methods with RK stability. Applied Numerical Mathematics: Transactions of IMACS, 198(??):138–147, April 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423003227.

Salane:1989:IPC

[Sal89] Douglas E. Salane. Improving the performance of a code for solving stiff systems of ODEs. Applied Numerical Mathematics: Transactions of IMACS, 5(4):363– 373, July 1989. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Salas:1991:I

[Sal91]  ManuelD. Salas.  Intro-

duction. Applied Numerical Mathematics: Transactions of IMACS, 7(1): 1, January 1991. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/016892749190100E.

Salas:1993:SWI

[Sal93] M. D. Salas. Shock wave interaction with an abrupt area change. Applied Numerical Mathematics: Transactions of IMACS, 12(1–3): 239–256, May 1993. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0168927493901217.

Salah:2003:TDC

[Sal03] I. Ben Salah. Third degree classical forms. Applied Numerical Mathematics: Transactions of IMACS, 44(4):433–447, March 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Samarskii:1994:TSR

[Sam94] Aleksander A. Samarskii.

Theory of stability and regularization of difference schemes and its application to ill-posed problems of mathematical physics. Applied Numerical Mathematics: Transactions of IMACS, 16(1–2):51–64, December 1, 1994. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=16&
issue=1-2&aid=520.  A

Festschrift to honor Professor Robert Vichnevetsky on his 65th birthday.

Sweilam:2020:NSN

[SAMSB20a] Nasser Sweilam, Seham Al-Mekhlafi, Salma Shatta, and Dumitru Baleanu. Numerical study for a novel variable-order multiple time delay awareness programs mathematical model. Applied Numerical Mathematics: Transactions of IMACS, 158(??):212–235, December 2020. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420302166.

Sweilam:2020:NST

[SAMSB20b] Nasser Sweilam, Seham Al-Mekhlafi, Salma Shatta, and Dumitru Baleanu. Numerical study for two types variable-order Burgers’ equations with proportional delay. Applied Numerical Mathematics: Transactions of IMACS, 156(??):364– 376, October 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420301501.

Sand:1989:CCT

[San89] Jørgen Sand. Choices in contractivity theory. recent theoretical results in numerical ordinary differential equations. Applied Numerical Mathematics: Transactions of IMACS, 5(1–2):105– 115, February 1989. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Sand:2002:IEH

[San02] J. Sand. On implicit Euler for high-order high-index DAEs. Applied Numerical Mathematics: Transactions of IMACS, 42(1– 3):411–424, August 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Santos:2003:PCG

[San03] Reginaldo J. Santos. Preconditioning conjugate gradient with symmetric algebraic reconstruction technique (ART) in computerized tomography. Applied Numerical Mathematics: Transactions of IMACS, 47(2):255–263, November 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Sanders:2018:PSH

[San18] Toby Sanders. Parameter selection for HOTV regularization. Applied Numerical Mathematics: Transactions of IMACS, 125(??): 1–9, March 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417302271.

Sandu:2020:CRI

[San20]  Adrian Sandu.  Conver-

gence results for implicitexplicit general linear methods. Applied Numerical Mathematics: Transactions of IMACS, 156(??): 242–264, October 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420301203.

Sarra:2005:ARB

[Sar05] Scott A. Sarra. Adaptive radial basis function methods for time dependent partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 54(1):79–94, June 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Sauter:2000:CIO

[Sau00] Thilo Sauter. Computation of irregularly oscillating integrals. Applied Numerical Mathematics: Transactions of IMACS, 35(3): 245–264, November 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/10/28/64/31/
                                        28/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/10/28/64/31/28/article.
                                        pdf.

Sazaklioglu:2022:NSS

[Saz22] Ali Ugur Sazaklioglu. On the numerical solutions of some identification problems for one- and multidimensional parabolic equations backward in time. Applied Numerical Mathematics: Transactions of IMACS, 181(??):76–93, November 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0168927422001465.

Sazaklioglu:2024:INM

[Saz24] Ali Ugur Sazaklioglu. An iterative numerical method for an inverse source problem for a multidimensional nonlinear parabolic equation. Applied Numerical Mathematics: Transactions of IMACS, 198(??):428–447, April 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000242.

Sandu:2003:FNT

[SB03] Adrian Sandu and Christian Borden. A framework for the numerical treatment of aerosol dynamics. Applied Numerical Mathematics: Transactions of IMACS, 45(4):475–497, June 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Seydaoglu:2014:HOS

[SB14] M. Seydaoglu and S. Blanes.

High-order splitting methods for separable nonautonomous parabolic equations. Applied Numerical Mathematics: Transactions of IMACS, 84(??):22– 32, October 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414000750.

Sarra:2018:RRB

[SB18] Scott A. Sarra and Yikun Bai. A rational radial basis function method for accurately resolving discontinuities and steep gradients. Applied Numerical Mathematics: Transactions of IMACS, 130(??): 131–142, August 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418300904.

Soleymani:2019:POU

[SB19] F. Soleymani and M. Barfeie. Pricing options under stochastic volatility jump model: a stable adaptive scheme. Applied Numerical Mathematics: Transactions of IMACS, 145(??):69– 89, November 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419301424.

Smotzer:2025:WHE

[SB25] Thomas Smotzer and John Buoni. A Wachspress– Habetler extension to the HSS iteration method in R^{n×n}. Applied Numerical Mathematics: Transactions of IMACS, 215(??):49– 58, September 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425000819.

Seydaoglu:2021:CMS

[SBBC21] Muaz Seydaoglu, Philipp Bader, Sergio Blanes, and Fernando Casas. Computing the matrix sine and cosine simultaneously with a reduced number of products. Applied Numerical Mathematics: Transactions of IMACS, 163(??):96–107, May 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000118.

Striebel:2009:MRS

[SBG09] Michael Striebel, Andreas Bartel, and Michael Günther. A multirate ROW-scheme for index-1 network equations. Applied Numerical Mathematics: Transactions of IMACS, 59(3–4): 800–814, March/April 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Singh:2020:SMD

[SBS^{+}20] Harendra Singh, D. Baleanu, H. M. Srivastava, Hemen Dutta, and Navin Kumar Jha. Solution of multidimensional Fredholm equations using Legendre scaling functions. Applied Numerical Mathematics: Transactions of IMACS, 150(??): 313–324, April 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892741930279X.

Song:2024:CCF

[SBS24] Yanlai Song, Mahdiar Barfeie, and Fazlollah Soleymani. Computing compact finite difference formulas under radial basis functions with enhanced applicability. Applied Numerical Mathematics: Transactions of IMACS, 201(??): 370–386, July 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424000643.

Slodicka:2003:NSN

[SC03] Marián Slodićka and Ivan Cimrák. Numerical study of nonlinear ferromagnetic materials. Applied Numerical Mathematics: Transactions of IMACS, 46(1): 95–111, July 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Shi:2008:AMI

[SC08] Yuying Shi and Qianshun Chang. Acceleration methods for image restoration problem with different boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 58(5):602–614, May 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Sahlodin:2011:DTR

[SC11] Ali M. Sahlodin and Benoît Chachuat. Discretize-then-relax approach for convex/concave relaxations of the solutions of parametric ODEs. Applied Numerical Mathematics: Transactions of IMACS, 61(7): 803–820, July 2011. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Sarkar:2019:SDG

[SC19] Tanmay Sarkar and Praveen Chandrashekar. Stabilized discontinuous Galerkin scheme for the magnetic induction equation. Applied Numerical Mathematics: Transactions of IMACS, 137(??):116–135, March 2019. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418302599.

Shi:2020:SOA

[SC20] Jiankang Shi and Minghua Chen. A second-order accurate scheme for twodimensional space fractional diffusion equations with time Caputo–Fabrizio fractional derivative. Applied Numerical Mathematics: Transactions of IMACS, 151(??): 246–262, May 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300088.

Suparatulatorn:2022:SCA

[SC22] Raweerote Suparatulatorn and Khuanchanok Chaichana. A strongly convergent algorithm for solving common variational inclusion with application to image recovery problems. Applied Numerical Mathematics: Transactions of IMACS, 173(??): 239–248, March 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421003408.

Shao:2025:FOR

[SC25a] Wenting Shao and Cheng Chen. A fourth order Runge–Kutta type of exponential time differencing and triangular spectral element method for two dimensional nonlinear Maxwell’s equations. Applied Numerical Mathematics: Transactions of IMACS, 207(??): 348–369, January 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424002447.

Shen:2025:EET

[SC25b] Xiaojuan Shen and Yongyong Cai. Error estimates of time discretizations for a Cahn–Hilliard phase-field model for the two-phase magnetohydrodynamic flows. Applied Numerical Mathematics: Transactions of IMACS, 207(??): 585–607, January 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424002642.

Scalone:2022:PPS

[Sca22]  Carmela Scalone.  Posi-

tivity preserving stochastic θ-methods for selected SDEs. Applied Numerical Mathematics: Transactions of IMACS, 172(??): 351–358, February 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421003020.

Schwandt:1987:IAM

[Sch87] Hartmut Schwandt. Interval arithmetic methods for systems of nonlinear equations arising from discretizations of quasilinear elliptic and parabolic partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 3(3):257–287, June 1987. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0168927487900523.

Schwandt:1989:TIA

[Sch89] Hartmut Schwandt. Truncated interval arithmetic block cyclic reduction. Applied Numerical Mathematics: Transactions of IMACS, 5(6):495–527, October 1989. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Schafer:1991:PAN

[Sch91]  Michael Schäfer.  Paral-

lel algorithms for the numerical simulation of threedimensional natural convection. Applied Numerical Mathematics: Transactions of IMACS, 7(4): 347–365, April 1991. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/016892749190069C.

Schild:1993:SPN

[Sch93] K.-H. Schild. Stability properties of non-rationally stabilized collocation schemes for stiff boundary value problems. Applied Numerical Mathematics: Transactions of IMACS, 13(1–3): 191–198, September 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Sixth Conference on the Numerical Treatment of Differential Equations (Halle, 1992).

Schropp:1995:UDS

[Sch95a] Johannes Schropp. Using dynamical systems methods to solve minimization problems. Applied Numerical Mathematics: Transactions of IMACS, 18(1–3):321–335, September 27, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=18&
issue=1-3&aid=613. Sev-

enth Conference on the Numerical Treatment of Differential Equations (Halle, 1994).

Schwitzer:1995:MSP

[Sch95b] Felix Schwitzer. W -methods for semilinear parabolic equations. Applied Numerical Mathematics: Transactions of IMACS, 18(1– 3):351–366, September 27, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.com/
cgi-bin/cas/tree/store/

                                        apnum/cas_sub/browse/browse.
                                        cgi?year=1995&volume=18&
                                        issue=1-3&aid=610. Sev-

enth Conference on the Numerical Treatment of Differential Equations (Halle, 1994).

Schiesser:1996:PBC

[Sch96]  W. E. Schiesser.  PDE

boundary conditions from minimum reduction of the PDE. Applied Numerical Mathematics: Transactions of IMACS, 20(1–2): 171–179, February 29, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=20&
                                        issue=1-2&aid=672. Work-

shop on the method of lines for time-dependent problems (Lexington, KY, 1995).

Scholz:1998:ENS

[Sch98] Hans-Eberhard Scholz. The examination of nonlinear stability and solvability of the algebraic equations for the implicit Taylor series method. Applied Numerical Mathematics: Transactions of IMACS, 28(2–4):439–458, October 1, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/

store/apnum/sub/1998/28/
2-4/925.pdf.

Schonauer:1999:ESD

[Sch99] Willi Schönauer. Experiments with search directions for a generalized CG method. Applied Numerical Mathematics: Transactions of IMACS, 30(2– 3):241–256, June 10, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1999&volume=30&
issue=2-3&aid=978.

Schropp:2002:BRK

[Sch02] Johannes Schropp. Behavior of Runge–Kutta discretizations near equilibria of index 2 differential algebraic systems. Applied Numerical Mathematics: Transactions of IMACS, 42(1– 3):425–435, August 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Schlier:2004:DBN

[Sch04]  Ch. Schlier.  Discrep-

ancy behaviour in the nonasymptotic regime. Applied Numerical Mathematics: Transactions of IMACS, 50(2):227–238, August 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Schaback:2008:RFW

[Sch08a] Robert Schaback. Recovery of functions from weak data using unsymmetric meshless kernel-based methods. Applied Numerical Mathematics: Transactions of IMACS, 58(5):726–741, May 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Schlier:2008:SHS

[Sch08b] Christoph Schlier. On scrambled Halton sequences. Applied Numerical Mathematics: Transactions of IMACS, 58(10):1467–1478, October 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Schittkowski:2009:ASS

[Sch09] Klaus Schittkowski. An active set strategy for solving optimization problems with up to 200,000,000 nonlinear constraints. Applied Numerical Mathematics: Transactions of IMACS, 59(12): 2999–3007, December 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Schmitt:2012:ASG

[Sch12] B. A. Schmitt. On algebraic stability of general linear methods and peer methods. Applied Numerical Mathematics: Transactions of IMACS, 62(10):

1544–1553, October 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412000918.

Schleper:2016:HTR

[Sch16a] Veronika Schleper. A HLLtype Riemann solver for two-phase flow with surface forces and phase transitions. Applied Numerical Mathematics: Transactions of IMACS, 108(??): 256–270, October 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416300897.

Schlottbom:2016:EAD

[Sch16b] Matthias Schlottbom. Error analysis of a diffuse interface method for elliptic problems with Dirichlet boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 109(??): 109–122, November 2016. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741630099X.

Schmid:2023:CEA

[Sch23]  Harald Schmid.  Com-

putation of the eigenvalues for the angular and Coulomb spheroidal wave equation. Applied Numerical Mathematics: Transactions of IMACS, 185(??): 101–119, March 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422003063.

Smyl:2021:NCF

[SCLL21] Danny Smyl, Liang Chen, Li Lai, and Dong Liu. Noncooperative finite element games. Applied Numerical Mathematics: Transactions of IMACS, 167(??): 273–280, September 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421001409.

St-Cyr:2005:NOI

[SCT05] Amik St-Cyr and Stephen J. Thomas. Nonlinear operator integration factor splitting for the shallow water equations. Applied Numerical Mathematics: Transactions of IMACS, 52(4): 429–448, March 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Sommeijer:1992:SPB

[SCvdH92] B. P. Sommeijer, W. Couzy, and P. J. van der Houwen. A-stable parallel block methods for ordinary and integrodifferential equations. Applied Numerical Mathematics: Transactions of IMACS, 9(3–5):267–281, April 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0168927492900215.

International Conference on the Numerical Solution of Volterra and Delay Equations (Tempe, AZ, 1990).

Simon:1993:EUS

[SD93] Horst D. Simon and Leonardo Dagum. Experience in using SIMD and MIMD parallelism for computational fluid dynamics. Applied Numerical Mathematics: Transactions of IMACS, 12(5):431–442, July 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/016892749390103X.

Sinha:2009:FEM

[SD09] Rajen K. Sinha and Bhupen Deka. Finite element methods for semilinear elliptic and parabolic interface problems. Applied Numerical Mathematics: Transactions of IMACS, 59(8): 1870–1883, August 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Shakeri:2011:FVS

[SD11] Fatemeh Shakeri and Mehdi Dehghan. A finite volume spectral element method for solving magnetohydrodynamic (MHD) equations. Applied Numerical Mathematics: Transactions of IMACS, 61(1):1–23, January 2011. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Salehi:2013:MLS

[SD13a] Rezvan Salehi and Mehdi Dehghan. A moving least square reproducing polynomial meshless method. Applied Numerical Mathematics: Transactions of IMACS, 69(??):34–58, July 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892741300038X.

Shakeri:2013:HOF

[SD13b] Fatemeh Shakeri and Mehdi Dehghan. A high order finite volume element method for solving elliptic partial integro-differential equations. Applied Numerical Mathematics: Transactions of IMACS, 65(??): 105–118, March 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927412001791.

Shirilord:2022:CRI

[SD22a] Akbar Shirilord and Mehdi Dehghan. Combined real and imaginary parts method for solving generalized Lyapunov matrix equation. Applied Numerical Mathematics: Transactions of IMACS, 181(??):94–109, November 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0168927422001337.

Shirilord:2022:DPS

[SD22b] Akbar Shirilord and Mehdi Dehghan. Double parameter splitting (DPS) iteration method for solving complex symmetric linear systems. Applied Numerical Mathematics: Transactions of IMACS, 171(??): 176–192, January 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421002385.

Shirilord:2024:IMC

[SD24a] Akbar Shirilord and Mehdi Dehghan. Iterative method for constrained systems of conjugate transpose matrix equations. Applied Numerical Mathematics: Transactions of IMACS, 198(??): 474–507, April 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000163.

Song:2024:IAR

[SD24b] Zefang Song and Huafei Di. Instability analysis and regularization approximation to the forward/backward problems for fractional damped wave equations with random noise. Applied Numerical Mathematics: Transactions of IMACS, 199(??): 177–212, May 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422003397.

Shirilord:2025:GDB

[SD25] Akbar Shirilord and Mehdi Dehghan. Gradient descentbased parameter-free methods for solving coupled matrix equations and studying an application in dynamical systems. Applied Numerical Mathematics: Transactions of IMACS, 212(??): 29–59, June 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742500011X.

Shayanfard:2020:CMA

[SDG20] F. Shayanfard, H. Laeli Dastjerdi, and F. M. Maalek Ghaini. Collocation method for approximate solution of Volterra integro-differential equations of the thirdkind. Applied Numerical Mathematics: Transactions of IMACS, 150(??): 139–148, April 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419302673.

Sofronov:2015:ATB

[SDK15] I. L. Sofronov, L. Dovgilovich, and N. Krasnov. Application of transparent boundary conditions to high-order finite-difference schemes for the wave equation in waveguides. Applied Numerical Mathematics: Transactions of IMACS, 93(??): 195–205, July 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741400107X.

Saini:2024:PUH

[SDK24] Sumit Saini, Pratibhamoy Das, and Sunil Kumar. Parameter uniform higher order numerical treatment for singularly perturbed Robin type parabolic reaction diffusion multiple scale problems with large delay in time. Applied Numerical Mathematics: Transactions of IMACS, 196(??):1–21, February 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742300257X.

SouzadeCursi:1999:AGD

[SdSC99] J. E. Souza de Cursi and M. B. de Souza Cortes. Approximate Gaussian distributions in optimization by random perturbation methods. Applied Numerical Mathematics: Transactions of IMACS, 30(1):23–30, May 10, 1999. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1999/30/
                                        1/946.pdf.

Sallam:1993:QSI

[SE93] S. Sallam and M. N. El Tarazi. Quadratic spline interpolation on uniform meshes. Applied Numerical Mathematics: Transactions of IMACS, 11(5): 419–427, March 1993. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Seaid:2009:ELM

[Sea09]  Mohammed Seaïd.  An

Eulerian–Lagrangian method for coupled parabolic-hyperbolic equations. Applied Numerical Mathematics: Transactions of IMACS, 59(3–4): 754–768, March/April 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Slayi:2021:SGU

[SED21] Sarah Slayi, Toufic El Arwadi, and Séréna Dib. Stabilized Gauge Uzawa scheme for an incompressible micropolar fluid flow. Applied Numerical Mathematics: Transactions of IMACS, 167(??):45–72, September 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0168927421001021.

Scolnik:2002:COR

[SEGV02] H. Scolnik, N. Echebest, M. T. Guardarucci, and M. C. Vacchino. A class of optimized row projection methods for solving large nonsymmetric linear systems. Applied Numerical Mathematics: Transactions of IMACS, 41(4): 499–513, July 2002. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Seiler:2002:CIS

[Sei02] Werner M. Seiler. Completion to involution and semidiscretisations. Applied Numerical Mathematics: Transactions of IMACS, 42(1–3):437–451, August 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Selgas:2014:SBF

[Sel14] V. Selgas. A symmetric BEM–FEM method for an axisymmetric eddy current problem. Applied Numerical Mathematics: Transactions of IMACS, 79(??): 107–123, May 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414000142.

Settanni:2024:PHC

[Set24] Giuseppina Settanni. Potentiality of the HOFiD bvp code in solving different kind of second order boundary value problems. Applied Numerical Mathematics: Transactions of IMACS, 200(??):379–388, June 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742300226X.

Shephard:2005:AMG

[SFJ^{+}05] Mark S. Shephard, Joseph E. Flaherty, Kenneth E. Jansen, Xiangrong Li, Xiaojuan Luo, Nicolas Chevaugeon, Jean-François Remacle, Mark W. Beall, and Robert M. O’Bara. Adaptive mesh generation for curved domains. Applied Numerical Mathematics: Transactions of IMACS, 52(2– 3):251–271, February 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Su:2021:PDI

[SFZ21] Haiyan Su, Xinlong Feng, and Jianping Zhao. Penalty decoupled iterative methods for the stationary natural convection equations with different Rayleigh numbers. Applied Numerical Mathematics: Transactions of IMACS, 163(??): 270–291, May 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742100012X.

Stewart:1992:NEN

[SG92] Kris Stewart and Tunc Geveci. Numerical experiments with a nonlinear evolution equation which exhibits blow-up. Applied Numerical Mathematics: Transactions of IMACS, 10(2):139–147, July 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Shampine:1996:SBE

[SG96] L. F. Shampine and I. Gladwell. Software based on explicit RK formulas. Applied Numerical Mathematics: Transactions of IMACS, 22(1–3):293–308, December 4, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=22&
issue=1-3&aid=721. Special

issue celebrating the centenary of Runge–Kutta methods.

Sidi:2000:SAO

[SG00] Avram Sidi and Dan Givoli. Stability and accuracy of optimal local non-reflecting boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4):327–340, May 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/61/27/
60/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/61/27/60/article.
pdf.

Schonauer:2002:P

[SG02] Willy Schönauer and Martin H. Gutknecht. Preface. Applied Numerical Mathematics: Transactions of IMACS, 41(1):1–6, April 2002. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/86/27/27/abstract.
html.

Sen:2004:GCB

[SG04] Rabindranath Sen and Pulak Guhathakurta. A globally convergent ball Stirling method. Applied Numerical Mathematics: Transactions of IMACS, 51(2– 3):329–340, November 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Striebel:2005:COM

[SG05] Michael Striebel and Michael Günther. A charge oriented mixed multirate method for a special class of index-1 network equations in chip design. Applied Numerical Mathematics: Transactions of IMACS, 53(2–4): 489–507, May 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Shampine:2006:DDA

[SG06] L. F. Shampine and P. Gahinet.

Delay-differential-algebraic equations in control theory. Applied Numerical Mathematics: Transactions of IMACS, 56(3–4):574–588, March/April 2006. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Sinha:2007:EEF

[SG07] Rajen K. Sinha and Jürgen Geiser. Error estimates for finite volume element methods for convection–diffusion– reaction equations. Applied Numerical Mathematics: Transactions of IMACS, 57(1):59–72, January 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Singer:2009:UAP

[SG09] Michael A. Singer and William H. Green. Using adaptive proper orthogonal decomposition to solve the reaction–diffusion equation. Applied Numerical Mathematics: Transactions of IMACS, 59(2):272–279, February 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Shi:2016:MPS

[SG16] YuFeng Shi and Yan Guo. A maximum-principle-satisfying finite volume compact-WENO scheme for traffic flow model on networks. Applied Numerical Mathematics: Transactions of IMACS, 108(??): 21–36, October 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416300691.

Sun:2017:MSH

[SG17] Zhengjie Sun and Wenwu Gao. A meshless scheme for Hamiltonian partial differential equations with conservation properties. Applied Numerical Mathematics: Transactions of IMACS, 119(??):115–125, September 2017. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417301010.

Svard:2006:SAD

[SGN06] Magnus Svärd, Jing Gong, and Jan Nordström. Stable artificial dissipation operators for finite volume schemes on unstructured grids. Applied Numerical Mathematics: Transactions of IMACS, 56(12): 1481–1490, December 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Svard:2008:AEU

[SGN08] Magnus Svärd, Jing Gong, and Jan Nordström. An accuracy evaluation of unstructured node-centred finite volume methods. Applied Numerical Mathematics: Transactions of IMACS, 58(8):1142–1158, August 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Sandu:2021:LIG

[SGR21] Adrian Sandu, Michael Günther, and Steven Roberts. Linearly implicit GARK schemes. Applied Numerical Mathematics: Transactions of IMACS, 161(??): 286–310, March 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303640.

Schotzau:2000:SSF

[SGS00] Dominik Schötzau, Klaus Gerdes, and Christoph Schwab. Stable and stabilized hp-finite element methods for the Stokes problem. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4): 349–356, May 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/61/27/
62/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/61/27/62/article.
pdf.

Sharma:2020:CMN

[SGS20] Deepika Sharma, Kavita Goyal, and Rohit Kumar Singla. A curvelet method for numerical solution of partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 148(??):28– 44, February 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419302387.

Shen:2022:MPP

[SGY22] Yuan Shen, Qianming Gao, and Xue Yin. A multiparameter parallel ADMM for multi-block linearly constrained separable convex optimization. Applied Numerical Mathematics: Transactions of IMACS, 171(??):369–388, January 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421002658.

Streett:1991:NSA

[SH91] C. L. Streett and M. Y.

Hussaini. Numerical simulation of the appearance of chaos in finite-length Taylor–Couette flow. Applied Numerical Mathematics: Transactions of IMACS, 7(1):41–71, January 1991. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0168927491901037.

Sardar:1997:DCV

[SH97] Tasneem K. Sardar and Desmond J. Higham. Dynamics of constant and variable stepsize methods for a nonlinear population model with delay. Applied Numerical Mathematics: Transactions of IMACS, 24(2– 3):425–438, June 19, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=24&
                                        issue=2-3&aid=793. Volterra

centennial (Tempe, AZ, 1996).

Schonauer:2002:NEO

[SH02] Willi Schönauer and Hartmut Häfner. Numerical experiments to optimize the use of (I)LU preconditioning in the iterative linear solver package LIN-SOL. Applied Numerical Mathematics: Transactions of IMACS, 41(1):23– 37, April 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        10/28/86/27/29/abstract.
                                        html.

Scheffel:2009:SSN

[SH09] Jan Scheffel and Cristian Håkansson. Solution of systems of nonlinear equations – a semi-implicit approach. Applied Numerical Mathematics: Transactions of IMACS, 59(10): 2430–2443, October 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Shang:2010:PIF

[SH10] Yueqiang Shang and Yinnian He. Parallel iterative finite element algorithms based on full domain partition for the stationary Navier–Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 60(7): 719–737, July 2010. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Song:2021:MSL

[SH21a] Sheng-Zhong Song and Zheng-Da Huang. A modified SSOR-like preconditioner for non-Hermitian positive definite matrices. Applied Numerical Mathematics: Transactions of IMACS, 164(??):175–189, ???? 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303585.

Sun:2021:MBM

[SH21b] Li Sun and Yu-Mei Huang.

A modulus-based multigrid method for image retinex. Applied Numerical Mathematics: Transactions of IMACS, 164(??): 199–210, ???? 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303615.

Shampine:1985:MS

[Sha85a] L. F. Shampine. Measuring stiffness. Applied Numerical Mathematics: Transactions of IMACS, 1(2):107– 119, March 1985. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Shampine:1985:SSU

[Sha85b] Lawrence F. Shampine. The step sizes used by onestep codes for ODEs. Applied Numerical Mathematics: Transactions of IMACS, 1(1):95–106, January 1985. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Shampine:1987:SVO

[Sha87] L. F. Shampine. Starting variable-order Adams and BDF codes. Applied Numerical Mathematics: Transactions of IMACS, 3(4):331– 337, August 1987. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Shapira:1998:MMD

[Sha98]  Yair Shapira.  Multi-

grid methods for 3-D definite and indefinite problems. Applied Numerical Mathematics: Transactions of IMACS, 26(3):377–398, March 2, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1998/26/
                                        3/826.pdf.

Shampine:2005:SOD

[Sha05] L. F. Shampine. Solving ODEs and DDEs with residual control. Applied Numerical Mathematics: Transactions of IMACS, 52(1):113– 127, January 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Shadimetov:2012:OQF

[SHA12] Kh. M. Shadimetov, A. R.

Hayotov, and S. S. Azamov. Optimal quadrature formula in K_{2}(P_{2}) space. Applied Numerical Mathematics: Transactions of IMACS, 62(12):1893–1909, December 2012. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412001456.

Sharma:2021:RPE

[Sha21] Natasha Sharma. Robust a-posteriori error estimates for weak Galerkin method for the convection-diffusion problem. Applied Numerical Mathematics: Transactions of IMACS, 170(??): 384–397, December 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742100235X.

Sidi:2025:IPR

[SHB^{+}25] H. Ould Sidi, A. S. Hendy, M. M. Babatin, L. Qiao, and M. A. Zaky. An inverse problem of Robin coefficient identification in parabolic equations with interior degeneracy from terminal observation data. Applied Numerical Mathematics: Transactions of IMACS, 212(??): 242–253, June 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425000315.

Shen:1996:NPM

[She96] Jie Shen. On a new pseudocompressibility method for the incompressible Navier– Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 21(1):71–90, June 6, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=21&
issue=1&aid=680.

Shen:2000:NFC

[She00]  Jie Shen.  A new fast

Chebyshev–Fourier algorithm for Poisson-type equations in polar geometries. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4):183–190, May 2000. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.nl/gej-ng/29/
17/21/61/27/43/abstract.
html;   http://www.elsevier.
nl/gej-ng/29/17/21/61/
27/43/article.pdf.

South:1986:SAI

[SHG86] Jerry C. South, Jr., Mohamed M. Hafez, and David Gottlieb. Stability analysis of intermediate boundary conditions in approximate factorization schemes. Applied Numerical Mathematics: Transactions of IMACS, 2(3–5):181–192, October 1986. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Shivanian:2020:CRH

[Shi20] Elyas Shivanian. Completion of right-hand side in the frame of inverse Cauchy problem of elliptic type equation through homogenization meshless collocation method. Applied Numerical Mathematics: Transactions of IMACS, 156(??): 493–513, October 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420301628.

Shiue:2025:ULT

[Shi25] Ming-Cheng Shiue. On the unconditional long-time L^{2}- stability of the BDF2 time stepping scheme for the twodimensional Navier–Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 214(??): 104–109, August 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425000704.

Saffarzadeh:2019:CAI

[SHL19] M. Saffarzadeh, M. Heydari, and G. B. Loghmani. Convergence analysis of an iterative numerical algorithm for solving nonlinear stochastic Itô-Volterra integral equations with m-dimensional Brownian motion. Applied Numerical Mathematics: Transactions of IMACS, 146(??): 182–198, December 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419301825.

Shen:2019:CNT

[SHLY19] Weiping Shen, Yaohua Hu, Chong Li, and Jen-Chih Yao. Convergence of the Newton-type methods for the square inverse singular value problems with multiple and zero singular values. Applied Numerical Mathematics: Transactions of IMACS, 143(??): 172–187, September 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419300649.

Shyy:1986:AGM

[Shy86] Wei Shyy. An adaptive grid method for Navier-Stokes flow computation. II. grid addition. Applied Numerical Mathematics: Transactions of IMACS, 2(1):9– 19, February 1986. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Shyy:1991:SAGa

[Shy91a] W. Shyy. Structure of an adaptive grid computational method from the viewpoint of dynamic chaos. Applied Numerical Mathematics: Transactions of IMACS, 7(3):263–285, March 1991. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Shyy:1991:SAGb

[Shy91b] Wei Shyy. Structure of an adaptive grid computational method from the viewpoint of dynamic chaos. II. grid addition and probability distribution. Applied Numerical Mathematics: Transactions of IMACS, 7(6):523–545, June 1991. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Shehu:2020:PMA

[SI20] Yekini Shehu and Olaniyi S. Iyiola. Projection methods with alternating inertial steps for variational inequalities: Weak and linear convergence. Applied Numerical Mathematics: Transactions of IMACS, 157(??): 315–337, November 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420301914.

Sidi:1990:AVE

[Sid90] Avram Sidi. Application of vector extrapolation methods to consistent singular linear systems. Applied Numerical Mathematics: Transactions of IMACS, 6(6):487– 500, October 1990. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Sidilkover:2002:FSE

[Sid02] David Sidilkover. Factorizable schemes for the equations of fluid flow. Applied Numerical Mathematics: Transactions of IMACS, 41(3):423–436, June 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Sidi:2010:SNS

[Sid10]  Avram Sidi.  Survey of

numerical stability issues in convergence acceleration. Applied Numerical Mathematics: Transactions of IMACS, 60(12):1395–1410, December 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Sidi:2014:AES

[Sid14] Avram Sidi. Analysis of errors in some recent numerical quadrature formulas for periodic singular and hypersingular integrals via regularization. Applied Numerical Mathematics: Transactions of IMACS, 81(??): 30–39, July 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414000403.

Sidi:2023:SAN

[Sid23] Avram Sidi. Spectrally accurate numerical quadrature formulas for a class of periodic Hadamard Finite Part integrals by regularization. Applied Numerical Mathematics: Transactions of IMACS, 184(??): 171–196, February 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422002306.

Simos:1991:NTM

[Sim91] T. E. Simos. A Numerovtype method for the numerical solution of the radial Schrödinger equation. Applied Numerical Mathematics: Transactions of IMACS, 7(2):201–206, February 1991. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Simeon:1993:EDF

[Sim93] B. Simeon. An extended descriptor form for the numerical integration of multibody systems. Applied Numerical Mathematics: Transactions of IMACS, 13(1–3): 209–220, September 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Sixth Conference on the Numerical Treatment of Differential Equations (Halle, 1992).

Simkani:1994:LLB

[Sim94a] M. Simkani. Local limiting behavior of the zeros of approximating polynomials. Applied Numerical Mathematics: Transactions of IMACS, 14(4): 451–456, June 10, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1994&volume=14&
                                        issue=4&aid=481.

Simpson:1994:AMT

[Sim94b] R. B. Simpson. Anisotropic mesh transformations and optimal error control. Applied Numerical Mathematics: Transactions of IMACS, 14(1–3):183–198, April 14, 1994. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=14&
issue=1-3&aid=467. Pro-

ceedings of the Third ARO Workshop on Adaptive Methods for Partial Differential Equations (Troy, NY, 1992).

Simeon:1998:ORS

[Sim98] B. Simeon. Order reduction of stiff solvers at elastic multibody systems. Applied Numerical Mathematics: Transactions of IMACS, 28(2–4):459–475, October 1, 1998. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1998/28/
2-4/926.pdf.

Simoncini:2004:BTP

[Sim04] V. Simoncini. Block triangular preconditioners for symmetric saddle-point problems. Applied Numerical Mathematics: Transactions of IMACS, 49(1):

63–80, April 2004. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Simoncini:2010:EKS

[Sim10] V. Simoncini. Extended Krylov subspace for parameter dependent systems. Applied Numerical Mathematics: Transactions of IMACS, 60(5):550–560, May 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Singh:2023:RCT

[Sin23] Mehakpreet Singh. Rate of convergence of two moments consistent finite volume scheme for non-classical divergence coagulation equation. Applied Numerical Mathematics: Transactions of IMACS, 187(??): 120–137, May 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742300034X.

Singh:2024:DOW

[Sin24] Maneesh Kumar Singh. A discrete-ordinate weak Galerkin method for radiative transfer equation. Applied Numerical Mathematics: Transactions of IMACS, 201(??):628–641, July 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742400031X.

Shin:2011:SCR

[SJ11] Byeong-Chun Shin and Jae-Hun Jung. Spectral collocation and radial basis function methods for one-dimensional interface problems. Applied Numerical Mathematics: Transactions of IMACS, 61(8):911–928, August 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Shivanian:2018:SMR

[SJ18] Elyas Shivanian and Ahmad Jafarabadi. The spectral meshless radial point interpolation method for solving an inverse source problem of the time-fractional diffusion equation. Applied Numerical Mathematics: Transactions of IMACS, 129(??):1–25, July 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418300539.

Shi:2020:SAT

[SJ20] Dongyang Shi and Xu Jia.

Superconvergence analysis of two-grid finite element method for nonlinear Benjamin–Bona–Mahony equation. Applied Numerical Mathematics: Transactions of IMACS, 148(??):45– 60, February 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419302363.

Singh:1991:ISS

[SK91] G. S. Singh and S. S. Kulkarni. An integrated solution strategy on parallel computers. Applied Numerical Mathematics: Transactions of IMACS, 8(2):163– 175, September 1991. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0168927491900496.

Sommeijer:1996:SMT

[SK96] B. P. Sommeijer and J. Kok. Splitting methods for threedimensional bio-chemical transport. Applied Numerical Mathematics: Transactions of IMACS, 21(3): 303–320, August 20, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=21&
issue=3&aid=697.

Saltiel:1997:PRT

[SK97] Craig Saltiel and Joseph Kolibal. Parallel radiative transport using a unified matrix approach. Applied Numerical Mathematics: Transactions of IMACS, 25(4):499–509, November 10, 1997. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=25&
issue=4&aid=802.

Stoffer:2001:VCB

[SK01] Daniel Stoffer and Urs Kirchgraber. Verification of chaotic behaviour in the planar restricted three body problem. Applied Numerical Mathematics: Transactions of IMACS, 39(3–4):415–433, December 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/79/34/37/abstract.
html.

Smyrlis:2010:UDV

[SK10] Yiorgos-Sokratis Smyrlis and Andreas Karageorghis. The under-determined version of the MFS: Taking more sources than collocation points. Applied Numerical Mathematics: Transactions of IMACS, 60(4): 337–357, April 2010. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Schutz:2016:NSS

[SK16] Jochen Schütz and Klaus Kaiser. A new stable splitting for singularly perturbed ODEs. Applied Numerical Mathematics: Transactions of IMACS, 107(??):18– 33, September 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416300423.

Shao:2022:PBS

[SK22] Xin-Hui Shao and Chong-Bo Kang. A preconditioner based on sine transform for space fractional diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 178(??): 248–261, August 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422000940.

Schlegel:2012:NSM

[SKAW12] Martin Schlegel, Oswald Knoth, Martin Arnold, and Ralf Wolke. Numerical solution of multiscale problems in atmospheric modeling. Applied Numerical Mathematics: Transactions of IMACS, 62(10): 1531–1543, October 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927412001092.

Scherer:2008:NTF

[SKBAS08] R. Scherer, S. L. Kalla, L. Boyadjiev, and B. Al-Saqabi. Numerical treatment of fractional heat equations. Applied Numerical Mathematics: Transactions of IMACS, 58(8):1212– 1223, August 2008. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Skeel:1989:SOB

[Ske89a] Robert D. Skeel. The secondorder backward differentiation formula is unconditionally zero-stable. Applied Numerical Mathematics: Transactions of IMACS, 5(1–2):145–149, February 1989. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Recent theoretical results in numerical ordinary differential equations.

Skelboe:1989:SPB

[Ske89b]  Stig Skelboe.  Stability

properties of backward differentiation multirate formulas. Applied Numerical Mathematics: Transactions of IMACS, 5(1–2):151– 160, February 1989. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Recent theoretical results in numerical ordinary differential equations.

Skeel:1999:SIF

[Ske99]  Robert D. Skeel.  Sym-

plectic integration with floating-point arithmetic and other approximations. Applied Numerical Mathematics: Transactions of IMACS, 29(1):3–18, January 1, 1999. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1999/29/
1/899.pdf.

Smoktunowicz:2019:ENA

[SKO19] Alicja Smoktunowicz, Ryszard Kozera, and Gianluca Oderda. Efficient numerical algorithms for constructing orthogonal generalized doubly stochastic matrices. Applied Numerical Mathematics: Transactions of IMACS, 142(??):16–27, August 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419300455.

Sigalotti:2016:KPC

[SKR^{+}16] Leonardo Di G. Sigalotti, Jaime Klapp, Otto Rendón, Carlos A. Vargas, and Franklin Peña-Polo. On the kernel and particle consistency in smoothed particle hydrodynamics. Applied Numerical Mathematics: Transactions of IMACS, 108(??):242–255, October 2016. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416300836.

Singh:2023:UCS

[SKS23] Satpal Singh, Devendra Kumar, and Vembu Shanthi. Uniformly convergent scheme for fourthorder singularly perturbed convection-diffusion ODE. Applied Numerical Mathematics: Transactions of IMACS, 186(??):334–357, April 2023. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742300020X.

Saadat:2025:NAF

[SKS^{+}25a] Hosein Saadat, Sanaz Hami Hassan Kiyadeh, Ali Safaie, Ramin Goudarzi Karim, and Fayyaz Khodadosti. A new amplification-fitting approach in Newton–Cotes rules to tackling the highfrequency IVPs. Applied Numerical Mathematics: Transactions of IMACS, 207(??):86–96, January 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424002289.

Salihu:2025:RMR

[SKS^{+}25b] Nasiru Salihu, Poom Kumam, Sani Salisu, Lin Wang, and Kanokwan Sitthithakerngkiet. A revised MR-MIL Riemannian conjugate gradient method with simplified global convergence properties. Applied Numerical Mathematics: Transactions of IMACS, 214(??): 110–126, August 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425000698.

Soleimani:2017:IPM

[SKW17] Behnam Soleimani, Oswald Knoth, and Rüdiger Weiner. IMEX peer methods for fast-wave-slow-wave problems. Applied Numerical Mathematics: Transactions of IMACS, 118(??): 221–237, August 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892741730065X.

Santos-Leon:2001:CIU

[SL01a] J. C. Santos-León. Computation of integrals over the unit circle with nearby poles. Applied Numerical Mathematics: Transactions of IMACS, 36(2– 3):179–195, February 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.nl../29/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/65/31/29/article.
pdf.

Sun:2001:IAN

[SL01b] W. Sun and K. M. Liu. Iterative algorithms for nonlinear ordinary differential eigenvalue problems. Applied Numerical Mathematics: Transactions of IMACS, 38(3):361–376, August 2001. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/77/33/33/abstract.
html.

Su:2008:MHM

[SL08] Menglong Su and Zhenxin Liu. Modified homotopy methods to solve fixed points of self-mapping in a broader class of nonconvex sets. Applied Numerical Mathematics: Transactions of IMACS, 58(3):236–248, March 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Shen:2009:KTC

[SL09] Weiping Shen and Chong Li. Kantorovich-type convergence criterion for inexact Newton methods. Applied Numerical Mathematics: Transactions of IMACS, 59(7):1599–1611, July 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Sousa:2015:WFD

[SL15] Ercília Sousa and Can Li. A weighted finite difference method for the fractional diffusion equation based on the Riemann–Liouville derivative. Applied Numerical Mathematics: Transactions of IMACS, 90(??): 22–37, April 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414001901.

Sun:2017:FEI

[SL17] Xuenan Sun and Xuezhang Liang. Fast and exact 2d image reconstruction based on Hakopian interpolation. Applied Numerical Mathematics: Transactions of IMACS, 121(??): 185–197, November 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417301605.

Shi:2020:USAa

[SL20] Dongyang Shi and Qian Liu. Unconditional superconvergent analysis of a linearized finite element method for Ginzburg–Landau equation. Applied Numerical Mathematics: Transactions of IMACS, 147(??):118–128, January 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419302326.

Shin:2021:LHO

[SL21] Jaemin Shin and Hyun Geun Lee. A linear, high-order, and unconditionally energy stable scheme for the epitaxial thin film growth model without slope selection. Applied Numerical Mathematics: Transactions of IMACS, 163(??):30–42, May 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420304050.

Shi:2022:NCS

[SL22] Dongyang Shi and Chaoqun Li. A new combined scheme of H^{1}-Galerkin FEM and TGM for bacterial equations. Applied Numerical Mathematics: Transactions of IMACS, 171(??): 23–31, January 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421002397.

Schmidt-Laine:1986:LSF

[SLJ86] C. Schmidt-Lainé and D. Jeandel. A least squares formulation of one-dimensional turbulence models. Applied Numerical Mathematics: Transactions of IMACS, 2(1):43–56, February 1986. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Shen:2011:ULM

[SLJ11] W. P. Shen, C. Li, and X. Q. Jin. A Ulm-like method for inverse eigenvalue problems. Applied Numerical Mathematics: Transactions of IMACS, 61(3):356–367, March 2011. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Shehu:2021:AVR

[SLMD21] Yekini Shehu, Lulu Liu, Xuewen Mu, and Qiao-Li Dong. Analysis of versions of relaxed inertial projection and contraction method. Applied Numerical Mathematics: Transactions of IMACS, 165(??):1–21, July 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421000465.

Shan:2017:STL

[SLW17] Yingying Shan, Wenjie Liu, and Boying Wu. Space-time Legendre–Gauss–Lobatto collocation method for twodimensional generalized sine-Gordon equation. Applied Numerical Mathematics: Transactions of IMACS, 122(??):92–107, December 2017. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417301769.

Si:2024:RVC

[SLW24] Zhiyong Si, Ziyi Li, and Leilei Wei. A rotational velocity-correction projection method for the Micropolar Navier–Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 205(??): 267–280, November 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001934.

Sun:2010:SRF

[SLZ10] Ping Sun, Zhendong Luo, and Yanjie Zhou. Some reduced finite difference schemes based on a proper orthogonal decomposition technique for parabolic equations. Applied Numerical Mathematics: Transactions of IMACS, 60(1–2):154–164, January/February 2010. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Smooke:1985:SNT

[SM85] M. D. Smooke and R. M. M.

Mattheij. On the solution of nonlinear twopoint boundary value problems on successively refined grids. Applied Numerical Mathematics: Transactions of IMACS, 1(6):463– 487, November 1985. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Streett:1989:SMD

[SM89] C. L. Streett and M. G. Macaraeg. Spectral multidomain for large-scale fluid dynamic simulations. Applied Numerical Mathematics: Transactions of IMACS, 6(1–2):123–139, December 1989. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Schmitt:1993:CHO

[SM93] B. A. Schmitt and Z. Mei. Construction of higher-order algebraic one-step schemes in stiff BVPs. Applied Numerical Mathematics: Transactions of IMACS, 13(1–3): 199–208, September 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Sixth Conference on the Numerical Treatment of Differential Equations (Halle, 1992).

Salcedo:2008:LIT

[SM08] J. V. Salcedo and M. Martínez. LPV identification of a turbocharged diesel engine. Applied Numerical Mathematics: Transactions of IMACS, 58(10):1553–1571, October 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Serna:2013:FPS

[SM13] Susana Serna and Antonio Marquina. Fronts propagating with signal dependent speed in limited diffusion and related Hamilton–Jacobi formulations. Applied Numerical Mathematics: Transactions of IMACS, 73(??):48– 62, November 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741200147X.

Shukla:2020:CFR

[SM20] Ankita Shukla and Mani Mehra. Compact filtering as a regularization technique for a backward heat conduction problem. Applied Numerical Mathematics: Transactions of IMACS, 153(??):82–97, July 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300386.

Spitaleri:2001:MFD

[SMA01] Rosa Maria Spitaleri, Riccardo March, and Daniele Arena. A multigrid finitedifference method for the solution of Euler equations of the variational image segmentation. Applied Numerical Mathematics: Transactions of IMACS, 39(2): 181–189, November 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/gej-
ng/10/10/28/79/33/30/abstract.
html.

Sato:2023:HOL

[SMB23] Shun Sato, Yuto Miyatake, and John C. Butcher. High-order linearly implicit schemes conserving quadratic invariants. Applied Numerical Mathematics: Transactions of IMACS, 187(??): 71–88, May 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423000363.

Schilders:2008:BMU

[SMC08] W. H. A. Schilders, P. B. L. Meijer, and E. Ciggaar. Behavioural modelling using the MOESP algorithm, dynamic neural networks and the Bartels–Stewart algorithm. Applied Numerical Mathematics: Transactions of IMACS, 58(12): 1972–1993, December 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Schlegel:2004:SAL

[SMEN04] Martin Schlegel, Wolfgang Marquardt, Rainald Ehrig, and Ulrich Nowak. Sensitivity analysis of linearlyimplicit differential–algebraic systems by one-step extrapolation. Applied Numerical Mathematics: Transactions of IMACS, 48(1):83– 102, January 2004. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Saadvandi:2012:DPM

[SMJ12] Maryam Saadvandi, Karl Meerbergen, and Elias Jarlebring. On dominant poles and model reduction of second order time-delay systems. Applied Numerical Mathematics: Transactions of IMACS, 62(1):21– 34, January 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411001681.

Stanic:2024:QRG

[SMJ24] Marija P. Stanić, Tatjana V. Tomović Mladenović, and Aleksandar Ne. Jovanović. Quadrature rules of Gaussian type for trigonometric polynomials with preassigned nodes. Applied Numerical Mathematics: Transactions of IMACS, 200(??):399–408, June 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001356.

Saberi-Movahed:2022:SIAb

[SMTHE22a] Farid Saberi-Movahed, Azita Tajaddini, Mohammed Heyouni, and Lakhdar Elbouyahyaoui. Some iterative approaches for Sylvester tensor equations, Part I: a tensor format of truncated Loose Simpler GM-RES. Applied Numerical Mathematics: Transactions of IMACS, 172(??): 428–445, February 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421003056.

Saberi-Movahed:2022:SIAa

[SMTHE22b] Farid Saberi-Movahed, Azita Tajaddini, Mohammed Heyouni, and Lakhdar Elbouyahyaoui. Some iterative approaches for Sylvester tensor equations, Part II: a tensor format of Simpler variant of GCRO-based methods. Applied Numerical Mathematics: Transactions of IMACS, 172(??): 413–427, February 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742100307X.

Singh:2021:TMC

[SMW21] Mehakpreet Singh, Themis Matsoukas, and Gavin Walker. Two moments consistent discrete formulation for binary breakage population balance equation and its convergence. Applied Numerical Mathematics: Transactions of IMACS, 166(??): 76–91, August 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001008.

Svard:2004:SFV

[SN04] Magnus Svärd and Jan Nordström. Stability of finite volume approximations for the Laplacian operator on quadrilateral and triangular grids. Applied Numerical Mathematics: Transactions of IMACS, 51(1):101– 125, October 2004. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Sapagovas:2022:ADM

[SN22] Mifodijus Sapagovas and Jurij Novickij. Alternating direction method for the wave equation with integral boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 182(??):1– 13, December 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742200191X.

Sun:2019:FAC

[SND19] Jing Sun, Daxin Nie, and Weihua Deng. Fast algorithms for convolution quadrature of Riemann– Liouville fractional derivative. Applied Numerical Mathematics: Transactions of IMACS, 145(??): 384–410, November 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419301072.

Sun:2021:HOB

[SND21] Jing Sun, Daxin Nie, and Weihua Deng. High-order BDF fully discrete scheme for backward fractional Feynman–Kac equation with nonsmooth data. Applied Numerical Mathematics: Transactions of IMACS, 161(??):82–100, March 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420303342.

Sekine:2021:NPA

[SNOK21] Kouta Sekine, Mitsuhiro T. Nakao, Shin’ichi Oishi, and Masahide Kashiwagi. A numerical proof algorithm for the non-existence of solutions to elliptic boundary value problems. Applied Numerical Mathematics: Transactions of IMACS, 169(??):87–107, November 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0168927421001793.

Singh:2022:CAV

[SNW22] Mehakpreet Singh, R. K.

Nayak, and Gavin Walker. Convergence analysis of volume preserving scheme for mass based coalescence equation. Applied Numerical Mathematics: Transactions of IMACS, 173(??): 365–379, March 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421003469.

Srati:2025:IPD

[SOA^{+}25] M. Srati, A. Oulmelk, L. Afraites, A. Hadri, M. A. Zaky, A. Aldraiweesh, and A. S. Hendy. An inverse problem of determining the parameters in diffusion equations by using fractional physics-informed neural networks. Applied Numerical Mathematics: Transactions of IMACS, 208 (part B) (??):189–213, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424002927.

Shahmorad:2020:TLN

[SOB20] Sedaghat Shahmorad, M. H. Ostadzad, and D. Baleanu. A Tau-like numerical method for solving fractional delay integro-differential equations. Applied Numerical Mathematics: Transactions of IMACS, 151(??): 322–336, May 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420300076.

Sobieraj:2024:MMC

[Sob24] Michał Sobieraj. A multilevel Monte Carlo algorithm for stochastic differential equations driven by countably dimensional Wiener process and Poisson random measure. Applied Numerical Mathematics: Transactions of IMACS, 206(??): 141–160, December 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424002034.

Sod:1991:CVM

[Sod91] Gary A. Sod. A compressible vortex method with application to the interaction of an oblique shock wave with a boundary layer. Applied Numerical Mathematics: Transactions of IMACS, 8(3):257–273, October 1991. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Soderlind:2006:TSS

[Söd06] Gustaf Söderlind. Time-step selection algorithms: Adaptivity, control, and signal processing. Applied Numerical Mathematics: Transactions of IMACS, 56(3–4): 488–502, March/April 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Sofronov:2017:TTB

[Sof17]  Ivan Sofronov.  Trun-

cated transparent boundary conditions for second order hyperbolic systems. Applied Numerical Mathematics: Transactions of IMACS, 120(??): 115–124, October 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417301101.

Solovev:2015:FEA

[Sol15]  Sergey I. Solov’ëv.  Fi-

nite element approximation with numerical integration for differential eigenvalue problems. Applied Numerical Mathematics: Transactions of IMACS, 93(??): 206–214, July 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414000257.

Sommeijer:1986:EER

[Som86] B. P. Sommeijer. On the economization of explicit Runge–Kutta methods. Applied Numerical Mathematics: Transactions of IMACS, 2(1):57–68, February 1986. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Sommeijer:1993:EHO

[Som93]  B. P. Sommeijer.  Ex-

plicit, high-order Runge– Kutta–Nyström methods for parallel computers. Applied Numerical Mathematics: Transactions of IMACS, 13(1–3):221–240, September 1993. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Sixth Conference on the Numerical Treatment of Differential Equations (Halle, 1992).

Song:1991:NSF

[Son91] Yu Song. Numerical simulation of flame propagations in circular cylinders. Applied Numerical Mathematics: Transactions of IMACS, 8(3):275–288, October 1991. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Soni:2000:GGP

[Son00] Bharat K. Soni. Grid generation: Past, present, and future. Applied Numerical Mathematics: Transactions of IMACS, 32(4): 361–369, April 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/59/32/
28/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/59/32/28/article.
pdf.

Sousa:2009:ESA

[Sou09] Ercília Sousa. On the edge of stability analysis. Applied Numerical Mathematics: Transactions of IMACS, 59(6):1322–1336, June 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Sadjadi:1999:ATR

[SP99] Seyed Jafar Sadjadi and K. Ponnambalam. Advances in trust region algorithms for constrained optimization. Applied Numerical Mathematics: Transactions of IMACS, 29(3):423– 443, March 1, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1999/29/
3/970.pdf.

Shokri:2022:CAT

[SP22] Javad Shokri and Saeed Pishbin. On the convergence analysis of the Tau method applied to fourth-order partial differential equation based on Volterra–Fredholm integral equations. Applied Numerical Mathematics: Transactions of IMACS, 173(??): 144–157, March 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421003329.

Speleers:2012:IQP

[Spe12]  Hendrik Speleers.  In-

terpolation with quintic Powell–Sabin splines. Applied Numerical Mathematics: Transactions of IMACS, 62(5):620–635, May 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927412000256.

Spigler:1990:NSD

[Spi90] Renato Spigler. Numerical simulation for degenerate diffusions. Applied Numerical Mathematics: Transactions of IMACS, 6(5):405– 413, July 1990. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Spijker:1993:NRS

[Spi93] M. N. Spijker. Numerical ranges and stability estimates. Applied Numerical Mathematics: Transactions of IMACS, 13(1–3): 241–249, September 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Sixth Conference on the Numerical Treatment of Differential Equations (Halle, 1992).

Spijker:1995:ESN

[Spi95] M. N. Spijker. The effect of the stopping of the Newton iteration in implicit linear multistep methods. Applied Numerical Mathematics: Transactions of IMACS, 18(1–3):367–386, September 27, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=18&
issue=1-3&aid=612. Sev-

enth Conference on the Numerical Treatment of Differential Equations (Halle, 1994).

Spijker:1996:EPR

[Spi96] M. N. Spijker. Error propagation in Runge–Kutta methods. Applied Numerical Mathematics: Transactions of IMACS, 22(1–3): 309–325, December 4, 1996.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=22&
issue=1-3&aid=722. Special

issue celebrating the centenary of Runge–Kutta methods.

Spijker:1997:NSR

[Spi97] M. N. Spijker. Numerical stability, resolvent conditions and delay differential equations. Applied Numerical Mathematics: Transactions of IMACS, 24(2– 3):233–246, June 19, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=24&
issue=2-3&aid=779. Volterra

centennial (Tempe, AZ, 1996).

Spitaleri:1999:CVA

[Spi99] R. M. Spitaleri. Computational and visual advances in multigrid grid generation. Applied Numerical Mathematics: Transactions of IMACS, 30(1):141–152, May 10, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/

store/apnum/sub/1999/30/
1/956.pdf.

Spitaleri:2000:FFM

[Spi00] R. M. Spitaleri. Full-FAS multigrid grid generation algorithms. Applied Numerical Mathematics: Transactions of IMACS, 32(4): 483–494, April 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/59/32/
35/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/59/32/35/article.
pdf.

Spitaleri:2003:P

[Spi03] Rosa Maria Spitaleri. Preface. Applied Numerical Mathematics: Transactions of IMACS, 46(3–4):247–248, September 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Spitaleri:2004:P

[Spi04] Rosa Maria Spitaleri. Preface. Applied Numerical Mathematics: Transactions of IMACS, 49(3–4):261–262, June 2004. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Spijker:2013:ESC

[Spi13] M. N. Spijker. The existence of stepsize-coefficients for boundedness of linear multistep methods. Applied Numerical Mathematics: Transactions of IMACS, 63(1):45–57, January 2013. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412001651.

Sanders:2020:ENM

[SPS20] Toby Sanders, Rodrigo B. Platte, and Robert D. Skeel. Effective new methods for automated parameter selection in regularized inverse problems. Applied Numerical Mathematics: Transactions of IMACS, 152(??): 29–48, June 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300167.

Singh:2024:WSV

[SPYS24] Aman Singh, Eugene B. Postnikov, Poonam Yadav, and Vineet Kumar Singh. Weakly singular Volterra integral equation with combined logarithmicpower-law kernel: Analytical and computational consideration. Applied Numerical Mathematics: Transactions of IMACS, 197(??): 164–185, March 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423002805.

Shang:2017:PFE

[SQ17] Yueqiang Shang and Jin Qin.

Parallel finite element variational multiscale algorithms for incompressible flow at high Reynolds numbers. Applied Numerical Mathematics: Transactions of IMACS, 117(??):1–21, July 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417300338.

Schuck:1988:NSS

[SR88a] P. L. Schuck and H. Rasmussen. Numerical solution of a subsidence mound problem in a porous medium. Applied Numerical Mathematics: Transactions of IMACS, 4(6):521– 546, November 1988. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Sorauf:1988:NSC

[SR88b] C. M. Sorauf and J. A.

Rial. Numerical study of caustics and polarizations of seismic waves interacting with the three-dimensional geological structures. Applied Numerical Mathematics: Transactions of IMACS, 4(1):71–81, March 1988. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Stynes:1997:MUS

[SR97] Martin Stynes and Hans-Görg Roos. The midpoint upwind scheme. Applied Numerical Mathematics: Transactions of IMACS, 23(3):361–374, April 25, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.com/
cas/tree/store/apnum/sub/
1997/23/3/753.pdf.

Shanazari:2009:TDA

[SR09] Kamal Shanazari and Nima Rabie. A three dimensional adaptive nodes technique applied to meshless-type methods. Applied Numerical Mathematics: Transactions of IMACS, 59(6):1187–1197, June 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Song:2024:ROT

[SR24] Junpeng Song and Hongxing Rui. A reduced-order two-grid method based on POD technique for the semilinear parabolic equation. Applied Numerical Mathematics: Transactions of IMACS, 205(??): 240–254, November 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001922.

Safdari:2021:SNL

[SRK21] Hamid Safdari, Majid Rajabzadeh, and Moein Khalighi. Solving a nonlinear fractional convectiondiffusion equation using local discontinuous Galerkin method. Applied Numerical Mathematics: Transactions of IMACS, 165(??): 22–34, July 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000453.

Safdari:2022:LAN

[SRK22] Hamid Safdari, Majid Rajabzadeh, and Moein Khalighi. LDG approximation of a nonlinear fractional convection-diffusion equation using B-spline basis functions. Applied Numerical Mathematics: Transactions of IMACS, 171(??): 45–57, January 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421002427.

Serna-Reyes:2023:ASW

[SRMDRL23] Adán J. Serna-Reyes, J. E.

Macías-Díaz, and Nuria Reguera-López. Analysis of a scheme which preserves the dissipation and positivity of Gibbs’ energy for a nonlinear parabolic equation with variable diffusion. Applied Numerical Mathematics: Transactions of IMACS, 183(??): 355–368, January 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422002550.

Sanz-Serna:1994:USI

[SS94a]  J. M. Sanz-Serna.  An

unconventional symplectic integrator of W. Kahan. Applied Numerical Mathematics: Transactions of IMACS, 16(1–2):245–250, December 1, 1994. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=16&
issue=1-2&aid=506.  A

Festschrift to honor Professor Robert Vichnevetsky on his 65th birthday.

Sei:1994:EAN

[SS94b] Alain Sei and W. W. Symes. Error analysis of numerical schemes for the wave equation in heterogeneous media. Applied Numerical Mathematics: Transactions of IMACS, 15(4): 465–480, November 1, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=15&
issue=4&aid=512.

Schmidt:1999:CSO

[SS99] Jens G. Schmidt and Gerhard Starke. Coarse space orthogonalization for indefinite linear systems of equations arising in geometrically nonlinear elasticity. Applied Numerical Mathematics: Transactions of IMACS, 30(2–3): 267–280, June 10, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1999&volume=30&
issue=2-3&aid=980.

Schmidt:2000:PEV

[SS00] Alfred Schmidt and Kunibert G. Siebert. A posteriori estimators for the h-p version of the finite element method in 1D. Applied Numerical Mathematics: Transactions of IMACS, 35(1):43– 66, September 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/64/27/
30/abstract.html;   http:
//www.elsevier.nl/gej-

                                        ng/10/10/28/64/27/30/article.
                                        pdf.

Sousa:2002:INB

[SS02] Ercília Sousa and Ian Sobey. On the influence of numerical boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 41(2):325–344, May 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Sachs:2008:ESP

[SS08a] E. W. Sachs and A. K. Strauss. Efficient solution of a partial integrodifferential equation in finance. Applied Numerical Mathematics: Transactions of IMACS, 58(11): 1687–1703, November 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Soheili:2008:MMM

[SS08b] Ali Reza Soheili and John M. Stockie. A moving mesh method with variable mesh relaxation time. Applied Numerical Mathematics: Transactions of IMACS, 58(3):249–263, March 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Stavroyiannis:2009:OFP

[SS09] S. Stavroyiannis and T. E. Simos. Optimization as a function of the phase-lag order of nonlinear explicit twostep P -stable method for linear periodic IVPs. Applied Numerical Mathematics: Transactions of IMACS, 59(10):2467–2474, October 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Slevinsky:2010:RAT

[SS10] Richard M. Slevinsky and Hassan Safouhi. A recursive algorithm for the G transformation and accurate computation of incomplete Bessel functions. Applied Numerical Mathematics: Transactions of IMACS, 60(12): 1411–1417, December 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Stykel:2012:KSM

[SS12] T. Stykel and V. Simoncini.

Krylov subspace methods for projected Lyapunov equations. Applied Numerical Mathematics: Transactions of IMACS, 62(1):35– 50, January 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741100170X.

Smoktunowicz:2013:IRT

[SS13a] Alicja Smoktunowicz and Agata Smoktunowicz. Iterative refinement techniques for solving block linear systems of equations. Applied Numerical Mathematics: Transactions of IMACS, 67(??):220–229, May 2013. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927411002066.

Steihaug:2013:RCH

[SS13b] Trond Steihaug and Sara Suleiman. Rate of convergence of higher order methods. Applied Numerical Mathematics: Transactions of IMACS, 67(??): 230–242, May 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892741100122X.

Sizikov:2016:GQS

[SS16] Valery Sizikov and Denis Sidorov. Generalized quadrature for solving singular integral equations of Abel type in application to infrared tomography. Applied Numerical Mathematics: Transactions of IMACS, 106(??):69–78, August 2016. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416300344.

Siskova:2017:RTD

[SS17] K. Sisková and M. Slodicka. Recognition of a time-dependent source in a time-fractional wave equation. Applied Numerical Mathematics: Transactions of IMACS, 121(??):1– 17, November 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417301411.

Shakya:2019:PEA

[SS19] Pratibha Shakya and Rajen Kumar Sinha. A posteriori error analysis for finite element approximations of parabolic optimal control problems with measure data. Applied Numerical Mathematics: Transactions of IMACS, 136(??):23– 45, February 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418302125.

Schutz:2021:APS

[SS21] Jochen Schütz and David C.

Seal. An asymptotic preserving semi-implicit multiderivative solver. Applied Numerical Mathematics: Transactions of IMACS, 160(??):84–101, February 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/pii/
                                        S0168927420302804.

Sadkane:2025:EGS

[SS25] Miloud Sadkane and Roger B. Sidje. Estimating the growth of solutions of linear delayed difference and differential equations by alternating maximization. Applied Numerical Mathematics: Transactions of IMACS, 212(??):254–267, June 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425000339.

Shloof:2022:SFF

[SSA^{+}22] A. M. Shloof, N. Senu, A. Ahmadian, N. M. A. Nik Long, and S. Salahshour. Solving fractal-fractional differential equations using operational matrix of derivatives via Hilfer fractalfractional derivative sense. Applied Numerical Mathematics: Transactions of IMACS, 178(??):386–403, August 2022. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422000393.

Sreelakshmi:2024:ICT

[SSA24] A. Sreelakshmi, V. P. Shyaman, and Ashish Awasthi. An interwoven composite tailored finite point method for two dimensional unsteady Burgers’ equation. Applied Numerical Mathematics: Transactions of IMACS, 197(??):71–96, March 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423002799.

Shahni:2023:BCM

[SSC23] Julee Shahni, Randhir Singh, and Carlo Cattani. Bernoulli collocation method for the third-order Lane–Emden– Fowler boundary value problem. Applied Numerical Mathematics: Transactions of IMACS, 186(??): 100–113, April 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423000077.

Srivastava:2021:ENA

[SSKS21] Nikhil Srivastava, Aman Singh, Yashveer Kumar, and Vineet Kumar Singh. Efficient numerical algorithms for Riesz-space fractional partial differential equations based on finite difference/operational matrix. Applied Numerical Mathematics: Transactions of IMACS, 161(??): 244–274, March 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420303500.

Sanz-Serna:1993:SCS

[SSL93] J. M. Sanz-Serna and S. Larsson. Shadows, chaos, and saddles. Applied Numerical Mathematics: Transactions of IMACS, 13(1–3): 181–190, September 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Sixth Conference on the Numerical Treatment of Differential Equations (Halle, 1992).

Sultana:2020:NSC

[SSPZ20] Farheen Sultana, Deeksha Singh, Rajesh K. Pandey, and Dia Zeidan. Numerical schemes for a class of tempered fractional integrodifferential equations. Applied Numerical Mathematics: Transactions of IMACS, 157(??):110–134, November 2020. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420301707.

Shiromani:2023:NTS

[SSR23] Ram Shiromani, Vembu Shanthi, and Higinio Ramos. Numerical treatment of a singularly perturbed 2-D convection-diffusion elliptic problem with Robin-type boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 187(??):176–191, May 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423000417.

Swati:2021:AAH

[SSS21] Swati, Mandeep Singh, and Karanjeet Singh. An advancement approach of Haar wavelet method and Bratu-type equations. Applied Numerical Mathematics: Transactions of IMACS, 170(??):74–82, December 2021. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421002051.

Sabiu:2023:MOP

[SSS^{+}23] Jamilu Sabi’u, Abdullah Shah, Predrag S. Stanimirović, Branislav Ivanov, and Mohammed Yusuf Waziri. Modified optimal Perry conjugate gradient method for solving system of monotone equations with applications. Applied Numerical Mathematics: Transactions of IMACS, 184(??): 431–445, February 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422002860.

Saad:2004:VAR

[SST04] Yousef Saad, Azzeddine Soulaimani, and Ridha Touihri. Variations on algebraic recursive multilevel solvers (ARMS) for the solution of CFD problems. Applied Numerical Mathematics: Transactions of IMACS, 51(2–3):305–327, November 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Sbibih:2009:PFQ

[SST09] D. Sbibih, A. Serghini, and A. Tijini. Polar forms and quadratic spline quasiinterpolants on Powell–Sabin partitions. Applied Numerical Mathematics: Transactions of IMACS, 59(5): 938–958, May 2009. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Sablonniere:2012:HOQ

[SST12] P. Sablonnière, D. Sbibih, and M. Tahrichi. Highorder quadrature rules based on spline quasi-interpolants and application to integral equations. Applied Numerical Mathematics: Transactions of IMACS, 62(5): 507–520, May 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927411002121.

Sbibih:2015:SQS

[SST15] D. Sbibih, A. Serghini, and A. Tijini. Superconvergent quadratic spline quasiinterpolants on Powell–Sabin partitions. Applied Numerical Mathematics: Transactions of IMACS, 87(??): 74–86, January 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414001524.

Sanz-Serna:1989:SCP

[SSV89] J. M. Sanz-Serna and J. G.

Verwer. Stability and convergence at the PDE/stiff ODE interface. Applied Numerical Mathematics: Transactions of IMACS, 5(1–2):117– 132, February 1989. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Recent theoretical results in numerical ordinary differential equations.

Sanz-Serna:1997:SIT

[SSV97] J. M. Sanz-Serna and J. G.

Verwer. Special issue on time integration. Applied Numerical Mathematics: Transactions of IMACS, 25(2–3): 135–320, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Sleijpen:2010:BCI

[SSvG10] Gerard L. G. Sleijpen, Peter Sonneveld, and Martin B. van Gijzen. Bi-CGSTAB as an induced dimension reduction method. Applied Numerical Mathematics: Transactions of IMACS, 60(11):1100–1114, November 2010. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Sommeijer:2002:P

[SSW02] B. P. Sommeijer, J. G. VerwerK. Strehmel, and R. Weiner. Preface. Applied Numerical Mathematics: Transactions of IMACS, 42(1– 3):??, August 2002. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Strauss:2004:PDT

[SSW04] Daniel J. Strauß, Gabriele Steidl, and Udo Welzel. Parameter detection of thin films from their X-ray reflectivity by support vector machines. Applied Numerical Mathematics: Transactions of IMACS, 48(2):223– 236, February 2004. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Sabiu:2020:TOH

[SSW20] Jamilu Sabi’u, Abdullah Shah, and Mohammed Yusuf Waziri. Two optimal Hager–Zhang conjugate gradient methods for solving monotone nonlinear equations. Applied Numerical Mathematics: Transactions of IMACS, 153(??): 217–233, July 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300635.

Soodhalter:2014:KSR

[SSX14] Kirk M. Soodhalter, Daniel B.

Szyld, and Fei Xue. Krylov subspace recycling for sequences of shifted linear systems. Applied Numerical Mathematics: Transactions of IMACS, 81(??): 105–118, July 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414000208.

Saito:2016:UPS

[SSZ16] Norikazu Saito, Yoshiki Sugitani, and Guanyu Zhou. Unilateral problem for the Stokes equations: the wellposedness and finite element approximation. Applied Numerical Mathematics: Transactions of IMACS, 105(??):124–147, July 2016. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416300290.

Swanson:1986:PTA

[ST86] R. C. Swanson and E. Turkel.

Pseudo-time algorithms for the Navier–Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 2(3–5):321–333, October 1986. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Scherer:1989:ACS

[ST89] R. Scherer and H. Türke. Algebraic characterization of A-stable Runge–Kutta methods. Applied Numerical Mathematics: Transactions of IMACS, 5(1–2):133– 144, February 1989. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Recent theoretical results in numerical ordinary differential equations.

Shampine:2001:SDM

[ST01] L. F. Shampine and S. Thompson. Solving DDEs in MATLAB. Applied Numerical Mathematics: Transactions of IMACS, 37(4): 441–458, June 2001. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/10/28/73/33/
                                        28/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/10/28/73/33/28/article.
                                        pdf.

Soto:2005:NAD

[ST05] Monica Selva Soto and Caren Tischendorf. Numerical analysis of DAEs from coupled circuit and semiconductor simulation. Applied Numerical Mathematics: Transactions of IMACS, 53(2–4):471–488, May 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Stynes:2008:URE

[ST08] Martin Stynes and Lutz Tobiska. Using rectangular Q_{p} elements in the SD-FEM for a convection diffusion problem with a boundary layer. Applied Numerical Mathematics: Transactions of IMACS, 58(12): 1789–1802, December 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Salkuyeh:2009:OIP

[ST09a] Davod Khojasteh Salkuyeh and Faezeh Toutounian. Optimal iterate of the power and inverse iteration methods. Applied Numerical Mathematics: Transactions of IMACS, 59(7):1537–1548, July 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Salkuyeh:2009:SSI

[ST09b] Davod Khojasteh Salkuyeh and Faezeh Toutounian. A sparse-sparse iteration for computing a sparse incomplete factorization of the inverse of an SPD matrix. Applied Numerical Mathematics: Transactions of IMACS, 59(6):1265–1273, June 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Salmi:2011:IMP

[ST11] Santtu Salmi and Jari Toivanen. An iterative method for pricing American options under jump-diffusion models. Applied Numerical Mathematics: Transactions of IMACS, 61(7): 821–831, July 2011. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Salmi:2014:ISP

[ST14a] Santtu Salmi and Jari Toivanen. IMEX schemes for pricing options under jumpdiffusion models. Applied Numerical Mathematics: Transactions of IMACS, 84(??):33–45, October 2014. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414000981.

Steinbach:2014:FFT

[ST14b] O. Steinbach and L. Tchoualag. Fast Fourier Transform for efficient evaluation of Newton potential in BEM. Applied Numerical Mathematics: Transactions of IMACS, 81(??):1–14, July 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414000397.

Senosiain:2019:NIU

[ST19] M. J. Senosiain and A. Tocino. On the numerical integration of the undamped harmonic oscillator driven by independent additive Gaussian white noises. Applied Numerical Mathematics: Transactions of IMACS, 137(??):49–61, March 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418302642.

Simos:2020:ENO

[ST20] T. E. Simos and Ch. Tsitouras. Explicit, ninth order, two step methods for solving inhomogeneous linear problems x′′(t) = Λx(t) + f(t). Applied Numerical Mathematics: Transactions of IMACS, 153(??): 344–351, July 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300702.

Song:2025:SAE

[ST25] Huailing Song and Qingkui Tan. Stability analysis and error estimates of implicitexplicit Runge–Kutta least squares RBF-FD method for time-dependent parabolic equation. Applied Numerical Mathematics: Transactions of IMACS, 207(??): 499–519, January 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424002551.

Stevenson:1997:ETP

[Ste97] Rob Stevenson. Experiments in 3D with a three-point hierarchical basis preconditioner. Applied Numerical Mathematics: Transactions of IMACS, 23(1):159–175, March 7, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1997/23/

1/748.pdf. Multilevel methods (Oberwolfach, 1995).

Stefanica:2005:PFA

[Ste05a] Dan Stefanica. Parallel FETI algorithms for mortars. Applied Numerical Mathematics: Transactions of IMACS, 54(2):266–279, July 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Steinbach:2005:NDD

[Ste05b]  O. Steinbach.  A natu-

ral domain decomposition method with non-matching grids. Applied Numerical Mathematics: Transactions of IMACS, 54(3– 4):362–377, August 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Stevenson:2008:MOP

[Ste08] Rob Stevenson. More on the order of prolongations and restrictions. Applied Numerical Mathematics: Transactions of IMACS, 58(12): 1875–1880, December 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Stiriba:2003:LPP

[Sti03] Youssef Stiriba. A local piecewise parabolic method for Hamilton–Jacobi equations. Applied Numerical Mathematics: Transactions of IMACS, 45(4):499–511, June 2003. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Stojanovic:1996:SDM

[Sto96] Mirjana Stojanović. Splines difference methods for a singular perturbation problem. Applied Numerical Mathematics: Transactions of IMACS, 21(3):321–333, August 20, 1996. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-

                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=21&
                                        issue=3&aid=693.

Straetemans:1998:RCD

[Str98a] F. A. J. Straetemans. Resolvent conditions for discretizations of diffusion– convection–reaction equations in several space dimensions. Applied Numerical Mathematics: Transactions of IMACS, 28(1):45–67, September 1, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1998/28/
                                        1/883.pdf.

Strand:1998:NSH

[Str98b]  Bo Strand.  Numerical

studies of hyperbolic IBVP with high-order finite difference operators satisfying a summation by parts rule. Applied Numerical Mathematics: Transactions of IMACS, 26(4):497–521, April 1, 1998. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1998/26/
                                        4/852.pdf.

Spotz:2000:FHO

[STS00] William F. Spotz, Mark A. Taylor, and Paul N. Swarztrauber. Fast and high-order solutions to the spherical shallow water equations. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4):191–197, May 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/61/27/
44/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/61/27/44/article.
pdf.

Su:1994:DDM

[Su94]  Chao Wei Su.  A do-

main decomposition method for determining the diffusion coefficient of a twodimensional linear diffusion equation in the time domain. Applied Numerical Mathematics: Transactions of IMACS, 15(4):481–493, November 1, 1994. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=15&
issue=4&aid=511.

Subasi:2004:ASB

[Sub04] Murat Subaşi. Approximations to the solution of boundary value problems of high order. Applied Numerical Mathematics: Transactions of IMACS, 48(1):103– 111, January 2004. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Schindler:2012:SSP

[SUP^{+}12] Thorsten Schindler, Heinz Ulbrich, Friedrich Pfeiffer, Arie van der Velde, and Arjen Brandsma. Spatial simulation of pushbelt CVTs with time-stepping schemes. Applied Numerical Mathematics: Transactions of IMACS, 62(10): 1515–1530, October 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927412001080.

Suslov:2010:SCA

[Sus10] I. M. Suslov. Strong coupling asymptotics of the βfunction in ϕ^{4} theory and QED. Applied Numerical Mathematics: Transactions of IMACS, 60(12): 1418–1428, December 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Sleijpen:1995:OAS

[Sv95] Gerard L. G. Sleijpen and Henk A. van der Vorst. An overview of approaches for the stable computation of hybrid BiCG methods. Applied Numerical Mathematics: Transactions of IMACS, 19(3):235–254, December 15, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=19&
issue=3&aid=633;   http:
//www.sciencedirect.com/
science/article/pii/0168927495000852.

Special issue on iterative methods for linear equations (Atlanta, GA, 1994).

Sharp:2000:EEB

[SV00] P. W. Sharp and J. H.

Verner. Extended explicit Bel’tyukov pairs of orders 4 and 5 for Volterra integral equations of the second kind. Applied Numerical Mathematics: Transactions of IMACS, 34(2–3): 261–274, July 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/62/29/
36/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/62/29/36/article.
pdf.

Sommariva:2024:TTF

[SV24] Alvise Sommariva and Marco Vianello. TetraFreeQ: Tetrahedra-free quadrature on polyhedral elements. Applied Numerical Mathematics: Transactions of IMACS, 200(??):389–398, June 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742300185X.

Sommeijer:2009:P

[SVAW09] Ben Sommeijer, Jan Verwer, Martin Arnold, and Rüdiger Weiner. Preface. Applied Numerical Mathematics: Transactions of IMACS, 59(3–4):405–409, March/April 2009. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Shoja:2017:SIM

[SVB17] A. Shoja, A. R. Vahidi, and E. Babolian. A spectral iterative method for solving nonlinear singular Volterra integral equations of Abel type. Applied Numerical Mathematics: Transactions of IMACS, 112(??):79– 90, February 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416301751.

Sommeijer:1994:TIT

[SvdHK94] B. P. Sommeijer, P. J. van der Houwen, and J. Kok. Time integration of threedimensional numerical transport models. Applied Numerical Mathematics: Transactions of IMACS, 16(1–2): 201–225, December 1, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=16&
issue=1-2&aid=525.  A

Festschrift to honor Professor Robert Vichnevetsky on his 65th birthday.

Sommeijer:1986:SLM

[SvdHN86] B. P. Sommeijer, P. J.

van der Houwen, and B. Neta. Symmetric linear multistep methods for second-order differential equations with periodic solutions. Applied Numerical Mathematics: Transactions of IMACS, 2(1):69– 77, February 1986. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Sudirham:2006:STD

[SvdVvD06] J. J. Sudirham, J. J. W.

van der Vegt, and R. M. J. van Damme. Space–time discontinuous Galerkin method for advection–diffusion problems on time-dependent domains. Applied Numerical Mathematics: Transactions of IMACS, 56(12): 1491–1518, December 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Sommeijer:2005:P

[SVSW05] Ben P. Sommeijer, Jan G.

Verwer, Karl Strehmel, and Rüdiger Weiner. Preface. Applied Numerical Mathematics: Transactions of IMACS, 53(2–4):89, May 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Stevenson:2022:OPS

[SvV22] Rob Stevenson and Raymond van Venetië. Operator preconditioning: the simplest case. Applied Numerical Mathematics: Transactions of IMACS, 172(??): 292–299, February 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421002701.

Stephan:1985:AGP

[SW85] E. Stephan and W. L. Wendland. An augmented Galerkin procedure for the boundary integral method applied to mixed boundary value problems. Applied Numerical Mathematics: Transactions of IMACS, 1(2):121–143, March 1985. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Swartz:1986:AFT

[SW86] Blair K. Swartz and Burton Wendroff. AZTEC: a front tracking code based on Godunov’s method. Applied Numerical Mathematics: Transactions of IMACS, 2(3–5):385–397, October 1986. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Shephard:1994:TIA

[SW94] Mark S. Shephard and Rolf Wentorf. Toward the implementation of automated analysis idealization control. Applied Numerical Mathematics: Transactions of IMACS, 14(1–3): 105–124, April 14, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=14&
issue=1-3&aid=451.

Saad:1995:DIS

[SW95a] Yousef Saad and Kesheng Wu. Design of an iterative solution module for a parallel sparse matrix library (P SPARSLIB). Applied Numerical Mathematics: Transactions of IMACS, 19(3):343–357, December 15, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=19&
issue=3&aid=638;   http:
//www.sciencedirect.com/
science/article/pii/0168927495000909.

Special issue on iterative methods for linear equations (Atlanta, GA, 1994).

Schmitt:1995:MFM

[SW95b] Bernhard A. Schmitt and Rüdiger Weiner. Matrixfree W -methods using a multiple Arnoldi iteration. Applied Numerical Mathematics: Transactions of IMACS, 18(1–3):307–320, September 27, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=18&
                                        issue=1-3&aid=608. Sev-

enth Conference on the Numerical Treatment of Differential Equations (Halle, 1994).

Schonauer:1995:P

[SW95c] W. Schönauer and R. Weiss. Preface. Applied Numerical Mathematics: Transactions of IMACS, 19(3): 173–174, December 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0168927495900055.

Schonauer:1995:EAG

[SW95d] Willi Schönauer and Rüdiger Weiss. An engineering approach to generalized conjugate gradient methods and beyond. Applied Numerical Mathematics: Transactions of IMACS, 19(3):175– 206, December 15, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=19&
issue=3&aid=631;   http:
//www.sciencedirect.com/
science/article/pii/0168927495000836.

Special issue on iterative methods for linear equations (Atlanta, GA, 1994).

Schmitt:1998:PPK

[SW98] Bernhard A. Schmitt and Rüdiger Weiner. Polynomial preconditioning in Krylov–ROW-methods. Applied Numerical Mathematics: Transactions of IMACS, 28(2–4):427–437, October 1, 1998. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1998/28/
2-4/924.pdf.

Song:2003:SEI

[SW03] Yongzhong Song and Li Wang.

On the semiconvergence of extrapolated iterative methods for singular linear systems. Applied Numerical Mathematics: Transactions of IMACS, 44(3):401– 413, February 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Sun:2005:DGM

[SW05] Shuyu Sun and Mary F. Wheeler. Discontinuous Galerkin methods for coupled flow and reactive transport problems. Applied Numerical Mathematics: Transactions of IMACS, 52(2–3):273–298, February 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Schneider:2006:WDM

[SW06] Reinhold Schneider and Toralf Weber. Wavelets for density matrix computation in electronic structure calculation. Applied Numerical Mathematics: Transactions of IMACS, 56(10–11):1383– 1396, October/November 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Shen:2007:FLG

[SW07] Jie Shen and Li-Lian Wang. Fourierization of the Legendre– Galerkin method and a new space–time spectral method. Applied Numerical Mathematics: Transactions of IMACS, 57(5–7):710–720, May 2007. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Schlenkrich:2009:GCQ

[SW09a] Sebastian Schlenkrich and Andrea Walther. Global convergence of quasi-Newton methods based on adjoint Broyden updates. Applied Numerical Mathematics: Transactions of IMACS, 59(5):1120–1136, May 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Skaflestad:2009:SMS

[SW09b] B. Skaflestad and W. M.

Wright. The scaling and modified squaring method for matrix functions related to the exponential. Applied Numerical Mathematics: Transactions of IMACS, 59(3–4):783–799, March/ April 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Schroder:2011:EEE

[SW11] Andreas Schröder and Sebastian Wiedemann. Error estimates in elastoplasticity using a mixed method. Applied Numerical Mathematics: Transactions of IMACS, 61(10):1031–1045, October 2011. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Steinebach:2012:PMO

[SW12] Gerd Steinebach and Rüdiger Weiner. Peer methods for the one-dimensional shallow-water equations with CWENO space discretization. Applied Numerical Mathematics: Transactions of IMACS, 62(10): 1567–1578, October 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412000955.

Song:2013:NSR

[SW13] Lunji Song and Yujiang Wu. Nonlinear stability of reaction-diffusion equations using wavelet-like incremental unknowns. Applied Numerical Mathematics: Transactions of IMACS, 68(??):83–107, June 2013. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413000020.

Sun:2017:IZO

[SW17] Liangliang Sun and Ting Wei. Identification of the zeroth-order coefficient in a time fractional diffusion equation. Applied Numerical Mathematics: Transactions of IMACS, 111(??): 160–180, January 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301726.

Soleimani:2018:SIP

[SW18] Behnam Soleimani and Rüdiger Weiner. Superconvergent IMEX peer methods. Applied Numerical Mathematics: Transactions of IMACS, 130(??): 70–85, August 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741830076X.

Shi:2020:USAb

[SW20a] Dongyang Shi and Ran Wang. Unconditional superconvergence analysis of a two-grid finite element method for nonlinear wave equations. Applied Numerical Mathematics: Transactions of IMACS, 150(??): 38–50, April 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741930251X.

Sun:2020:DLM

[SW20b] Pengtao Sun and Cheng Wang. Distributed Lagrange multiplier/fictitious domain finite element method for Stokes/parabolic interface problems with jump coefficients. Applied Numerical Mathematics: Transactions of IMACS, 152(??): 199–220, June 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419303411.

Shi:2021:QUC

[SW21] Dongyang Shi and Yanmi Wu. Quasi-uniform convergence analysis of rectangular Morley element for the singularly perturbed Bi-wave equation. Applied Numerical Mathematics: Transactions of IMACS, 161(??): 169–177, March 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420303524.

Su:2024:SCF

[SW24] Shuai Su and Jiming Wu. A symmetric and coercive finite volume scheme preserving the discrete maximum principle for anisotropic diffusion equations on star-shaped polygonal meshes. Applied Numerical Mathematics: Transactions of IMACS, 198(??):217–235, April 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424000084.

Shiri:2020:CMT

[SWB20] Babak Shiri, Guo-Cheng Wu, and Dumitru Baleanu. Collocation methods for terminal value problems of tempered fractional differential equations. Applied Numerical Mathematics: Transactions of IMACS, 156(??): 385–395, October 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420301513.

Shiri:2021:TVP

[SWB21] Babak Shiri, Guo-Cheng Wu, and Dumitru Baleanu. Terminal value problems for the nonlinear systems of fractional differential equations. Applied Numerical Mathematics: Transactions of IMACS, 170(??): 162–178, December 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001835.

Sun:2015:EEI

[SWCH15] F. X. Sun, J. F. Wang, Y. M.

Cheng, and A. X. Huang. Error estimates for the interpolating moving least-squares method in n-dimensional space. Applied Numerical Mathematics: Transactions of IMACS, 98(??):79– 105, December 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415001129.

Schmitt:2005:IPP

[SWE05] B. A. Schmitt, R. Weiner, and K. Erdmann. Implicit parallel peer methods for stiff initial value problems. Applied Numerical Mathematics: Transactions of IMACS, 53(2–4):457–470, May 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Shankar:2013:SDM

[SWFK13] Varun Shankar, Grady B. Wright, Aaron L. Fogelson, and Robert M. Kirby. A study of different modeling choices for simulating platelets within the immersed boundary method. Applied Numerical Mathematics: Transactions of IMACS, 63(1):58–77, January 2013. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927412001663.

Schmitt:2009:POE

[SWJ09] B. A. Schmitt, R. Weiner, and S. Jebens. Parameter optimization for explicit parallel peer two-step methods. Applied Numerical Mathematics: Transactions of IMACS, 59(3–4): 769–782, March/April 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Sun:2024:TCP

[SWJ^{+}24] Yanxiao Sun, Jiang Wu, Maosheng Jiang, Steven M. Wise, and Zhenlin Guo. A thermodynamically consistent phase-field model and an entropy stable numerical method for simulating two-phase flows with thermocapillary effects. Applied Numerical Mathematics: Transactions of IMACS, 206(??):161–189, December 2024. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742400206X.

Shi:2020:EIA

[SWL20] Wei Shi, Xinyuan Wu, and Kai Liu. Efficient implementation of the ARKN and ERKN integrators for multi-frequency oscillatory systems with multiple time scales. Applied Numerical Mathematics: Transactions of IMACS, 151(??): 13–26, May 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419303460.

Sulman:2011:EAN

[SWR11] Mohamed M. Sulman, J. F.

Williams, and Robert D. Russell. An efficient approach for the numerical solution of the Monge–Ampère equation. Applied Numerical Mathematics: Transactions of IMACS, 61(3): 298–307, March 2011. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Su:2011:CDM

[SWW11] Lijuan Su, Wenqia Wang, and Hong Wang. A characteristic difference method for the transient fractional convection–diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 61(8):946– 960, August 2011. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Shi:2016:CRM

[SWW16] Cong Shi, Chen Wang, and Ting Wei. Convolution regularization method for backward problems of linear parabolic equations. Applied Numerical Mathematics: Transactions of IMACS, 108(??):143–156, October 2016. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416300885.

Shao:2017:NSA

[SWW17] Wenting Shao, Xionghua Wu, and Cheng Wang. Numerical study of an adaptive domain decomposition algorithm based on Chebyshev tau method for solving singular perturbed problems. Applied Numerical Mathematics: Transactions of IMACS, 118(??):19–32, August 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417300429.

Shen:2000:FEM

[SWX00] Jie Shen, Feng Wang, and Jinchao Xu. A finite element multigrid preconditioner for Chebyshev-collocation methods. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4): 471–477, May 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/61/27/
75/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/61/27/75/article.
pdf.

Shen:2023:VEM

[SWY^{+}23] Xiaoqin Shen, Chen Wang, Qian Yang, Jikun Zhao, and Zhiming Gao. Virtual element method for linear elastic clamped plate model. Applied Numerical Mathematics: Transactions of IMACS, 191(??):1– 16, September 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423001198.

Shi:2022:FOB

[SXL22] Yilei Shi, Shusen Xie, and Dong Liang. A fourthorder block-centered compact difference scheme for nonlinear contaminant transport equations with adsorption. Applied Numerical Mathematics: Transactions of IMACS, 171(??): 212–232, January 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421002476.

Shen:2009:GCT

[SXP09] Chungen Shen, Wenjuan Xue, and Dingguo Pu. Global convergence of a tri-dimensional filter SQP algorithm based on the line search method. Applied Numerical Mathematics: Transactions of IMACS, 59(2):235–250, February 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Santiago:2003:MSA

[SY03] Cosmo D. Santiago and Jin-Yun Yuan. Modified ST algorithms and numerical experiments. Applied Numerical Mathematics: Transactions of IMACS, 47(2): 237–253, November 2003.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Shen:2005:NMS

[SY05] Yun-Qiu Shen and Tjalling J.

Ypma. Newton’s method for singular nonlinear equations using approximate left and right nullspaces of the Jacobian. Applied Numerical Mathematics: Transactions of IMACS, 54(2): 256–265, July 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Shen:2007:SRD

[SY07] Yun-Qiu Shen and Tjalling J.

Ypma. Solving rank-deficient separable nonlinear equations. Applied Numerical Mathematics: Transactions of IMACS, 57(5–7): 609–615, May 2007. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Song:2008:UMM

[SY08] Huailing Song and Yirang Yuan. An upwind-mixed method on changing meshes for two-phase miscible flow in porous media. Applied Numerical Mathematics: Transactions of IMACS, 58(6):815–826, June 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Shi:2018:SAN

[SY18] Dongyang Shi and Huaijun Yang. Superconvergence analysis of a new low order nonconforming MFEM for time-fractional diffusion equation. Applied Numerical Mathematics: Transactions of IMACS, 131(??): 109–122, September 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418301120.

Shih:2005:GMG

[SYG^{+}05] Alan M. Shih, Tzu-Yi Yu, Sankarappan Gopalsamy, Yasushi Ito, and Bharat Soni. Geometry and mesh generation for high fidelity computational simulations using non-uniform rational B-splines. Applied Numerical Mathematics: Transactions of IMACS, 55(3): 368–381, November 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Shen:2020:NAD

[SYL^{+}20] Xiaoqin Shen, Qian Yang, Linjin Li, Zhiming Gao, and Tiantian Wang. Numerical approximation of the dynamic Koiter’s model for the hyperbolic parabolic shell. Applied Numerical Mathematics: Transactions of IMACS, 150(??):

194–205, April 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419302788.

Steele:2018:QFC

[SYW18] A. Steele, J. Yalim, and B. Welfert. QX factorization of centrosymmetric matrices. Applied Numerical Mathematics: Transactions of IMACS, 134(??):11– 16, December 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741830148X.

Sheng:2022:RGA

[SYW22] Zhou Sheng, Weiwei Yang, and Jie Wen. A Riemannian gradient ascent algorithm with applications to orthogonal approximation problems of symmetric tensors. Applied Numerical Mathematics: Transactions of IMACS, 182(??): 235–247, December 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001994.

Sheng:2020:NCC

[SYY20] Zhiqiang Sheng, Guangwei Yuan, and Jingyan Yue. A nonlinear convex combination in the construction of finite volume scheme satisfying maximum principle. Applied Numerical Mathematics: Transactions of IMACS, 156(??): 125–139, October 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742030129X.

Skeel:1997:CIS

[SZ97] Robert D. Skeel and Meiqing Zhang. Cheap implicit symplectic integrators. Applied Numerical Mathematics: Transactions of IMACS, 25(2–3):297–302, October 17, 1997. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=25&
issue=2-3&aid=822.

Sturm:1999:SPD

[SZ99] Jos F. Sturm and Shuzhong Zhang. Symmetric primal– dual path-following algorithms for semidefinite programming. Applied Numerical Mathematics: Transactions of IMACS, 29(3):301– 315, March 1, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1999/29/
3/964.pdf.

Schotzau:2009:RPE

[SZ09] Dominik Schötzau and Liang Zhu. A robust a-posteriori error estimator for discontinuous Galerkin methods for convection–diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 59(9): 2236–2255, September 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Shi:2012:EAG

[SZ12] Wenjie Shi and Chengjian Zhang. Error analysis of generalized polynomial chaos for nonlinear random ordinary differential equations. Applied Numerical Mathematics: Transactions of IMACS, 62(12): 1954–1964, December 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412001511.

Shi:2017:GPC

[SZ17] Wenjie Shi and Chengjian Zhang. Generalized polynomial chaos for nonlinear random delay differential equations. Applied Numerical Mathematics: Transactions of IMACS, 115(??): 16–31, May 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892741630246X.

Shan:2022:PTS

[SZ22a] Li Shan and Haicheng Zhang. Partitioned time stepping method with different time scales for a dual-porosity-Stokes model. Applied Numerical Mathematics: Transactions of IMACS, 171(??):281–306, January 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421002634.

Shi:2022:USF

[SZ22b] Dongyang Shi and Sihui Zhang. Unconditional superconvergence of the fullydiscrete schemes for nonlinear prey-predator model. Applied Numerical Mathematics: Transactions of IMACS, 172(??):118–132, February 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421002774.

Szafraniec:1994:LIF

[Sza94] F. H. Szafraniec. Localization in the implicit function theorem: the method and an application. Applied Numerical Mathematics: Transactions of IMACS, 15(2):299– 303, September 13, 1994.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=15&
issue=2&aid=501. Innova-

tive methods in numerical analysis (Bressanone, 1992).

Shu:1992:HOE

[SZE^{+}92] Chi-Wang Shu, Thomas A.

Zang, Gordon Erlebacher, David Whitaker, and Stanley Osher. High-order ENO schemes applied to two-and three-dimensional compressible flow. Applied Numerical Mathematics: Transactions of IMACS, 9(1):45– 71, January 1992. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Sayari:2020:AHM

[SZE20] Sayed Sayari, Abdelhamid Zaghdani, and Miled El Hajji. Analysis of HDG method for the reaction–diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 156(??): 396–409, October 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420301562.

Song:2018:RWG

[SZL18] Lunji Song, Shan Zhao, and Kaifang Liu. A relaxed weak Galerkin method for elliptic interface problems with low regularity. Applied Numerical Mathematics: Transactions of IMACS, 128(??):65–80, June 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418300308.

Sidi:2023:IUS

[SZQH23] Hamed Ould Sidi, Mahmoud A. Zaky, Wenlin Qiu, and Ahmed S. Hendy. Identification of an unknown spatial source function in a multidimensional hyperbolic partial differential equation with interior degeneracy. Applied Numerical Mathematics: Transactions of IMACS, 192(??):1–18, October 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001411.

Sun:2019:RTD

[SZW19] Liangliang Sun, Yun Zhang, and Ting Wei. Recovering the time-dependent potential function in a multi-term time-fractional diffusion equation. Applied Numerical Mathematics: Transactions of IMACS, 135(??):228–245, January 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418301971.

Shen:2021:PPA

[SZY21] Yuan Shen, Yannian Zuo, and Aolin Yu. A partially proximal S-ADMM for separable convex optimization with linear constraints. Applied Numerical Mathematics: Transactions of IMACS, 160(??):65– 83, February 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302920.

TaAsan:1986:MMV

[Ta’86] Shlomo Ta’Asan. Multigrid method for a vortex breakdown simulation. Applied Numerical Mathematics: Transactions of IMACS, 2(3–5):303–311, October 1986. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Tadmor:1986:MEP

[Tad86] Eitan Tadmor. A minimum entropy principle in the gas dynamics equations. Applied Numerical Mathematics: Transactions of IMACS, 2(3–5):211–219, October 1986. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Taha:1996:IST

[Tah96] Thiab R. Taha. Inverse scattering transform numerical schemes for nonlinear evolution equations and the method of lines. Applied Numerical Mathematics: Transactions of IMACS, 20(1–2):181–187, February 29, 1996. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=1-2&aid=673. Work-

shop on the method of lines for time-dependent problems (Lexington, KY, 1995).

Taha:2019:F

[Tah19]  Thiab R. Taha.  Fore-

word. Applied Numerical Mathematics: Transactions of IMACS, 141(??): 1, ???? 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419300558.

Taha:2024:F

[Tah24]  Thiab R. Taha.  Fore-

word. Applied Numerical Mathematics: Transactions of IMACS, 199 (??):1, May 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423002118.

Tan:1987:CDE

[Tan87] Roger C. E. Tan. Computing derivatives of eigensystems by the topological ϵalgorithm. Applied Numerical Mathematics: Transactions of IMACS, 3(6):539– 550, November 1987. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Tang:1993:FDS

[Tan93] Tao Tang. A finite difference scheme for partial integro-differential equations with a weakly singular kernel. Applied Numerical Mathematics: Transactions of IMACS, 11(4):309–319, February 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Tang:2001:NSR

[Tan01] Hua-Zhong Tang. Nonlinear stability of the relaxing schemes for scalar conservation laws. Applied Numerical Mathematics: Transactions of IMACS, 38(3):347– 359, August 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/

                                        10/28/77/33/32/abstract.
                                        html.

Tang:2023:IFO

[Tan23] Shujiang Tang. Improvements of the fifth-order WENO-JS-type scheme with normalized smoothing factor for gas dynamic Euler equations. Applied Numerical Mathematics: Transactions of IMACS, 184(??): 301–324, February 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742200280X.

Tang:2024:NSL

[Tan24]  Shujiang Tang.  New

simple local smoothness indicators for fifth-order WENO schemes simulating compressible flows. Applied Numerical Mathematics: Transactions of IMACS, 197(??):46–70, March 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423002775.

Tarrazo:1998:CUI

[Tar98] Manuel Tarrazo. Calculating uncertainty intervals in approximate equation systems. Applied Numerical Mathematics: Transactions of IMACS, 27(1):59–67, May 1, 1998. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1998/27/
1/867.pdf.

Tausch:2009:NDP

[Tau09] Johannes Tausch. Nyström discretization of parabolic boundary integral equations. Applied Numerical Mathematics: Transactions of IMACS, 59(11):2843–2856, November 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Tian:2001:ITM

[TB01] Tianhai Tian and Kevin Burrage. Implicit Taylor methods for stiff stochastic differential equations. Applied Numerical Mathematics: Transactions of IMACS, 38(1–2):167–185, July 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/77/27/
36/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/77/27/36/article.
pdf.

Temimi:2020:TBN

[TBRBM20] H. Temimi, M. Ben-Romdhane, M. Baccouch, and M. O. Musa. A two-branched numerical solution of the twodimensional Bratu’s problem. Applied Numerical Mathematics: Transactions of IMACS, 153(??): 202–216, July 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420300520.

Torres:2003:SRF

[TC03] C. R. Torres and J. E. Castillo. Stratified rotating flow over complex terrain. Applied Numerical Mathematics: Transactions of IMACS, 47(3–4):531–541, December 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Turk:2019:CSC

[TC19] Önder Türk and Ramon Codina. Chebyshev spectral collocation method approximations of the Stokes eigenvalue problem based on penalty techniques. Applied Numerical Mathematics: Transactions of IMACS, 145(??):188–200, November 2019. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419301576.

Taraschi:2022:CPH

[TC22] Giovanni Taraschi and Maicon R. Correa. On the convergence of the primal hybrid finite element method on quadrilateral meshes. Applied Numerical Mathematics: Transactions of IMACS, 181(??): 552–560, November 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001799.

Tang:1989:GPS

[TCCW89] Y. N. Tang, Y. M. Chen, W. H. Chen, and M. L. Wasserman. Generalized pulse-spectrum technique for 2-D and 2-phase history matching. Applied Numerical Mathematics: Transactions of IMACS, 5(6):529– 539, October 1989. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Tunga:2009:BAU

[TD09] M. Alper Tunga and Metin Demiralp. Bound analysis in univariately truncated Generalized High Dimensional Model Representation for random-data partitioning: Interval GHDMR. Applied Numerical Mathematics: Transactions of IMACS, 59(6):1431–1448, June 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Tavelli:2013:HRM

[TDC13] M. Tavelli, M. Dumbser, and V. Casulli. High resolution methods for scalar transport problems in compliant systems of arteries. Applied Numerical Mathematics: Transactions of IMACS, 74(??):62–82, December 2013. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413001049.

Thang:2021:IUS

[TDMT21] Nguyen Van Thang, Nguyen Van Duc, Luong Duy Nhat Minh, and Nguyen Trung Thành. Identifying an unknown source term in a timespace fractional parabolic equation. Applied Numerical Mathematics: Transactions of IMACS, 166(??): 313–332, August 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742100115X.

Toumi:2017:SOA

[TDPU17] Asma Toumi, Guillaume Dufour, Ronan Perrussel, and Thomas Unfer. Second order accurate asynchronous scheme for modeling linear partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 121(??): 115–133, November 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417301502.

Tian:2023:HOW

[TDW23] Kang-Bo Tian, Wai Sun Don, and Bao-Shan Wang. High order WENO finite difference scheme with adaptive dual order ideal weights for hyperbolic conservation laws. Applied Numerical Mathematics: Transactions of IMACS, 187(??): 50–70, May 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423000351.

Temimi:2015:SDG

[Tem15]  H. Temimi.  Supercon-

vergence of discontinuous Galerkin solutions for higherorder ordinary differential equations. Applied Numerical Mathematics: Transactions of IMACS, 88(??):46– 65, February 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414001652.

Temimi:2023:EAN

[Tem23] Helmi Temimi. Error analysis of a novel discontinuous Galerkin method for the two-dimensional Poisson’s equation. Applied Numerical Mathematics: Transactions of IMACS, 189(??):

130–150, July 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423000983.

Tamagno:2003:NST

[TER03] José P. Tamagno, Sergio A. Elaskar, and Gustavo A. Ríos. Numerical simulation of time-dependent reacting flows in pulse facilities. Applied Numerical Mathematics: Transactions of IMACS, 47(3– 4):515–530, December 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Terekhov:2022:TDL

[Ter22] Andrew V. Terekhov. A three-dimensional Laguerre one-way wave equation solver. Applied Numerical Mathematics: Transactions of IMACS, 173(??):380–394, March 2022. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422000010.

Tangman:2008:ETI

[TGB08] D. Y. Tangman, A. Gopaul, and M. Bhuruth. Exponential time integration and Chebychev discretisation schemes for fast pricing of options. Applied Numerical Mathematics: Transactions of IMACS, 58(9): 1309–1319, September 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Thong:2022:EAS

[TGV22] Duong Viet Thong, Aviv Gibali, and Phan Tu Vuong. An explicit algorithm for solving monotone variational inequalities. Applied Numerical Mathematics: Transactions of IMACS, 171(??): 408–425, January 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421002671.

Toro:2009:AFV

[TH09] Eleuterio F. Toro and Arturo Hidalgo. ADER finite volume schemes for nonlinear reaction–diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 59(1):73– 100, January 2009. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Tang:2018:SCA

[TH18] Qili Tang and Yunqing Huang. Stability and convergence analysis of a Crank– Nicolson leap-frog scheme for the unsteady incompressible Navier–Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 124(??): 110–129, February 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417302064.

Torkaman:2023:INB

[TH23] Soraya Torkaman and Mohammad Heydari. An iterative Nyström-based method to solve nonlinear Fredholm integral equations of the second kind. Applied Numerical Mathematics: Transactions of IMACS, 194(??):59– 81, December 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423002271.

Themistoclakis:2017:SEB

[The17] Woula Themistoclakis. Some error bounds for Gauss– Jacobi quadrature rules. Applied Numerical Mathematics: Transactions of IMACS, 116(??):286–293, June 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417300545.

Thompson:1985:SDA

[Tho85] Joe F. Thompson. A survey of dynamically-adaptive grids in the numerical solution of partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 1(1):3– 27, January 1985. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Tang:2019:PCI

[THW19] Jingyong Tang, Chengdai Huang, and Yongli Wang. Predictor–corrector inexact smoothing algorithm for symmetric cone complementarity problems with Cartesian P_{0}-property. Applied Numerical Mathematics: Transactions of IMACS, 143(??):146–158, September 2019. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741930087X.

Tiago:2015:NSS

[Tia15] Jorge Tiago. Numerical simulations for the stabilization and estimation problem of a semilinear partial differential equation. Applied Numerical Mathematics: Transactions of IMACS, 98(??):18– 37, December 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415001154.

Touma:2012:NOC

[TJ12] R. Touma and G. Jannoun. Non-oscillatory central schemes on unstructured grids for two-dimensional hyperbolic conservation laws. Applied Numerical Mathematics: Transactions of IMACS, 62(8):941–955, August 2012. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412000517.

Tcheutia:2018:MRE

[TJK18] D. D. Tcheutia, A. S. Jooste, and W. Koepf. Mixed recurrence equations and interlacing properties for zeros of sequences of classical qorthogonal polynomials. Applied Numerical Mathematics: Transactions of IMACS, 125(??):86–102, March 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417302386.

Timonov:2005:NIP

[TK05] Alexandre Timonov and Michael V. Klibanov. A new iterative procedure for the numerical solution of coefficient inverse problems. Applied Numerical Mathematics: Transactions of IMACS, 54(2):280–291, July 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Touma:2015:WBC

[TK15] R. Touma and C. Klingenberg. Well-balanced central finite volume methods for the Ripa system. Applied Numerical Mathematics: Transactions of IMACS, 97(??):42– 68, November 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415001002.

Trandal:2019:LWD

[TK19] Krister J. Trandal and Henrik Kalisch. Long wave dynamics for a liquid CO_{2} lake in the deep ocean. Applied Numerical Mathematics: Transactions of IMACS, 141(??):144–157, ???? 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419300236.

Thanh:2011:NAB

[TKN11] Mai Duc Thanh, Dietmar Kröner, and Nguyen Thanh Nam. Numerical approximation for a Baer–Nunziato model of two-phase flows. Applied Numerical Mathematics: Transactions of IMACS, 61(5):702–721, May 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Tai:2007:PCL

[TL07] Xue-Cheng Tai and Hongwei Li. A piecewise constant level set method for elliptic inverse problems. Applied Numerical Mathematics: Transactions of IMACS, 57(5–7):686–696, May 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Teng:2020:MSS

[TLG20] Long Teng, Aleksandr Lapitckii, and Michael Günther. A multi-step scheme based on cubic spline for solving backward stochastic differential equations. Applied Numerical Mathematics: Transactions of IMACS, 150(??): 117–138, April 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419302557.

Tang:2022:CAA

[TLGC22] Yan Tang, Honghua Lin, Aviv Gibali, and Yeol Je Cho. Convergence analysis and applications of the inertial algorithm solving inclusion problems. Applied Numerical Mathematics: Transactions of IMACS, 175(??):1–17, May 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422000241.

Ta:2018:CT

[TLP18a] Thi Thanh Mai Ta, Van Chien Le, and Ha Thanh Pham. Corrigendum to “Shape optimization for Stokes flows using sensitivity analysis and finite element method” [Appl. Numer. Math. 126 (April 2018) 160–179]. Applied Numerical Mathematics: Transactions of IMACS, 129(??):192, July 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418300564.

See [TLP18b].

Ta:2018:SOS

[TLP18b] Thi Thanh Mai Ta, Van Chien Le, and Ha Thanh Pham. Shape optimization for Stokes flows using sensitivity analysis and finite element method. Applied Numerical Mathematics: Transactions of IMACS, 126(??): 160–179, April 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417302593.

See corrigendum [TLP18a].

Tan:2021:SAI

[TLQ21] Bing Tan, Songxiao Li, and Xiaolong Qin. Selfadaptive inertial single projection methods for variational inequalities involving non-Lipschitz and Lipschitz operators with their applications to optimal control problems. Applied Numerical Mathematics: Transactions of IMACS, 170(??): 219–241, December 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421002130.

Tran:2009:BIE

[TLSS09] T. Tran, Q. T. Le Gia, I. H. Sloan, and E. P. Stephan. Boundary integral equations on the sphere with radial basis functions: error analysis. Applied Numerical Mathematics: Transactions of IMACS, 59(11): 2857–2871, November 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Thompson:1992:DAM

[TLV92] C. P. Thompson, G. K. Leaf, and J. Van Rosendale. A dynamically adaptive multigrid algorithm for the incompressible Navier–Stokes equations validation and model problems. Applied Numerical Mathematics: Transactions of IMACS, 9(6):511–532, May 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Tsamasphyros:2004:IPP

[TM04] G. Tsamasphyros and S. Markolefas. Integration pointwise pollution error estimates in the finite element method in one dimension. Applied Numerical Mathematics: Transactions of IMACS, 51(2–3): 345–360, November 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Tsamasphyros:2005:SPE

[TM05] G. Tsamasphyros and S. Markolefas. Some a priori error estimates with respect to H^{θ} norms, 0 < θ < 1, for the h-extension of the finite element method in two dimensions. Applied Numerical Mathematics: Transactions of IMACS, 52(4):449–458, March 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Truhar:2015:RPT

[TM15] Ninoslav Truhar and Suzana Miodragović. Relative perturbation theory for definite matrix pairs and hyperbolic eigenvalue problem. Applied Numerical Mathematics: Transactions of IMACS, 98(??): 106–121, December 2015. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415001269.

Tolle:2021:EGF

[TM21] Kevin Tolle and Nicole Marheineke. Extended group finite element method. Applied Numerical Mathematics: Transactions of IMACS, 162(??):1–19, April 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420303810.

Tang:2024:LTE

[TM24] Yiyi Tang and Xuerong Mao. The logarithmic truncated EM method with weaker conditions. Applied Numerical Mathematics: Transactions of IMACS, 198(??): 258–275, April 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424000102.

Tang:1992:PIM

[TMD92] Tao Tang, Sean McKee, and Teresa Diogo. Product integration methods for an integral equation with logarithmic singular kernel. Applied Numerical Mathematics: Transactions of IMACS, 9(3–5):259–266, April 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/016892749290020E.

International Conference on the Numerical Solution of Volterra and Delay Equations (Tempe, AZ, 1990).

Tuncer:2015:PFE

[TMM15] N. Tuncer, A. Madzvamuse, and A. J. Meir. Projected finite elements for reactiondiffusion systems on stationary closed surfaces. Applied Numerical Mathematics: Transactions of IMACS, 96(??):45–71, October 2015. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415000550.

Takagi:1987:NAE

[TMS87] Toshiyuki Takagi, Kazuyoshi Miki, and Hiroki Sano. New approach to electron beam analysis using a domain decomposition and overlapping method in a threedimensional boundary-fitted coordinate system. Applied Numerical Mathematics: Transactions of IMACS, 3(4):305–316, August 1987. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Tambue:2016:WCS

[TN16] Antoine Tambue and Jean Medard T. Ngnotchouye. Weak convergence for a stochastic exponential integrator and finite element discretization of stochastic partial differential equation with multiplicative & additive noise. Applied Numerical Mathematics: Transactions of IMACS, 108(??): 57–86, October 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416300617.

Tobiska:2014:TLH

[Tob14] Lutz Tobiska. A two-level higher order local projection stabilization on hexahedral meshes. Applied Numerical Mathematics: Transactions of IMACS, 86(??):74– 80, December 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414001391.

Tocci:2001:SAL

[Toc01] Michael D. Tocci. Sensitivity analysis of large-scale time dependent PDEs. Applied Numerical Mathematics: Transactions of IMACS, 37(1–2):109–125, April 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/10/28/73/26/
                                        32/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/10/28/73/26/32/article.
                                        pdf.

Torres:2002:IFF

[TOCV02] Carlos R. Torres, J. Ochoa, J. Castillo, and M. Van Woert. Initial flow field of stratified flow past an impulsively started sphere. Applied Numerical Mathematics: Transactions of IMACS, 40(1–2):235–244, January 2002. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/85/27/46/abstract.
html.

Ternat:2011:TSM

[TOD11] Fabien Ternat, Oscar Orellana, and Prabir Daripa. Two stable methods with numerical experiments for solving the backward heat equation. Applied Numerical Mathematics: Transactions of IMACS, 61(2):266– 284, February 2011. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Tolstykh:2003:MPC

[Tol03] Andrei I. Tolstykh. On multioperators principle for constructing arbitrary-order difference schemes. Applied Numerical Mathematics: Transactions of IMACS, 46(3–4):411–423, September 2003. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Tolstykh:2004:CPO

[Tol04] Andrei I. Tolstykh. Centered prescribed-order approximations with structured grids and resulting finite-volume schemes. Applied Numerical Mathematics: Transactions of IMACS, 49(3–4): 431–440, June 2004. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Tomanovic:2024:IEZ

[Tom24] Jelena Tomanović. Incorporating the external zeros and poles of the integrand into Gauss-type quadrature rules. Applied Numerical Mathematics: Transactions of IMACS, 200(??): 409–428, June 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423001228.

Tone:2004:EAS

[Ton04]  Florentina Tone.  Error

analysis for a second order scheme for the Navier–Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 50(1): 93–119, July 2004. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Toprakseven:2021:WGF

[Top21]  Suayip Toprakseven.  A

weak Galerkin finite element method for time fractional reaction–diffusion– convection problems with variable coefficients. Applied Numerical Mathematics: Transactions of IMACS, 168(??):1–12, October 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001501.

Toro:2006:MMS

[Tor06] E. F. Toro. MUSTA: a multistage numerical flux. Applied Numerical Mathematics: Transactions of IMACS, 56(10–11):1464–1479, October/November 2006. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Torlo:2022:IPP

[TÖR22] Davide Torlo, Philipp Öffner, and Hendrik Ranocha. Issues with positivity-preserving Patankar-type schemes. Applied Numerical Mathematics: Transactions of IMACS, 182(??):117–147, December 2022. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742200188X.

Toutounian:1997:UCL

[Tou97] Faezeh Toutounian. The use of the CADNA library for validating the numerical results of the hybrid GMRES algorithm. Applied Numerical Mathematics: Transactions of IMACS, 23(2): 275–289, March 21, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cas/
                                        tree/store/apnum/sub/1997/
                                        23/2/734.pdf.

Touma:2010:UCS

[Tou10] R. Touma. Unstaggered central schemes with constrained transport treatment for ideal and shallow water magnetohydrodynamics. Applied Numerical Mathematics: Transactions of IMACS, 60(7):752–766, July 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Toulopoulos:2025:UST

[Tou25]  I. Toulopoulos.  A uni-

fied space-time finite element scheme for a quasilinear parabolic problem. Applied Numerical Mathematics: Transactions of IMACS, 214(??):127–142, August 2025. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425000686.

Towers:2016:FGS

[Tow16] John D. Towers. A fixed grid, shifted stencil scheme for inviscid fluid-particle interaction. Applied Numerical Mathematics: Transactions of IMACS, 110(??):26– 40, December 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741630143X.

Tassaddiq:2025:CSF

[TQA^{+}25] Asifa Tassaddiq, Sania Qureshi, Ioannis K. Argyros, Francisco I. Chicharro, Amanullah Soomro, Paras Nizamani, and Evren Hincal. A convergent and stable fourth-order iterative procedure based on Kung– Traub conjecture for nonlinear systems. Applied Numerical Mathematics: Transactions of IMACS, 214(??): 54–79, August 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425000546.

Tan:2024:EAS

[TQY24] Bing Tan, Xiaolong Qin, and Jen-Chih Yao. Extragradient algorithms for solving equilibrium problems on Hadamard manifolds. Applied Numerical Mathematics: Transactions of IMACS, 201(??):187–216, July 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000588.

teRiele:1991:Pa

[tRDvdV91a] Herman J. J. te Riele, T. J. Dekker, and H. A. van der Vorst. Preface. Applied Numerical Mathematics: Transactions of IMACS, 7(5):367–368, June 1991. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/016892749190007M.

teRiele:1991:Pb

[tRDvdV91b] Herman J. J. te Riele, T. J. Dekker, and H. A. van der Vorst. Preface. Applied Numerical Mathematics: Transactions of IMACS, 8(2):91–92, September 1991. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/016892749190044Z.

Trompert:1993:LUG

[Tro93] Ron Trompert. Local uniform grid refinement and systems of coupled partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 12(4):331–355, June 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Trotta:1996:MFE

[Tro96] R. L. Trotta. Multidomain finite elements for advection– diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 21(1):91–118, June 6, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=21&
issue=1&aid=679.

Trummer:2000:SMC

[Tru00] Manfred R. Trummer. Spectral methods in computing invariant tori. Applied Numerical Mathematics: Transactions of IMACS, 34(2–3):275–292, July 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/62/29/
37/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/62/29/37/article.
pdf.

Thompson:2006:FFD

[TS06] S. Thompson and L. F.

Shampine. A friendly Fortran DDE solver. Applied Numerical Mathematics: Transactions of IMACS, 56(3–4):503–516, March/ April 2006. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Tougas:2008:TUU

[TS08] Jane E. Tougas and Raymond J. Spiteri. Two uses for updating the partial singular value decomposition in latent semantic indexing. Applied Numerical Mathematics: Transactions of IMACS, 58(4):499–510, April 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Tahir:2023:NAC

[TS23] Shko Ali Tahir and Murat Sari. A new approach for the coupled advection–diffusion processes including source effects. Applied Numerical Mathematics: Transactions of IMACS, 184(??): 391–405, February 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422002847.

Tameh:2024:DPM

[TS24a] Mahboubeh Tavakoli Tameh and Fatemeh Shakeri. Difference potentials method for the nonlinear convection– diffusion equation with interfaces. Applied Numerical Mathematics: Transactions of IMACS, 202(??): 246–263, August 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424000230.

Toprakseven:2024:EWG

[TS24b] Suayip Toprakseven and Natesan Srinivasan. An efficient weak Galerkin FEM for third-order singularly perturbed convection-diffusion differential equations on layer-adapted meshes. Applied Numerical Mathematics: Transactions of IMACS, 204(??):130–146, October 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001491.

Tsao:1991:AST

[Tsa91] Nai Kuan Tsao. On the accuracy of solving triangular systems in parallel. Applied Numerical Mathematics: Transactions of IMACS, 7(2):207–215, February 1991. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Tsao:1992:AST

[Tsa92] Nai Kuan Tsao. On the accuracy of solving triangular systems in parallel. II. Applied Numerical Mathematics: Transactions of IMACS, 9(1):73–89, January 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Tauber:2010:IMA

[TSB10] Clovis Tauber, Pierre Spiteri, and Hadj Batatia. Iterative methods for anisotropic diffusion of speckled medical images. Applied Numerical Mathematics: Transactions of IMACS, 60(11): 1115–1130, November 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Tse:2000:SPC

[Tse00] K. L. Tse. Simulation of penetrative convection with adaptive spectral domain decomposition method. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4):267–274, May 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/61/27/
53/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/61/27/53/article.
pdf.

Takacs:2024:HOR

[TSF24] Bálint Máté Takács, Gabriella Svantnerné Sebestyén, and István Faragó. High-order reliable numerical methods for epidemic models with nonconstant recruitment rate. Applied Numerical Mathematics: Transactions of IMACS, 206(??):75–93, December 2024. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424002046.

Tsiantos:2001:CEW

[TSFB01] Vassilios D. Tsiantos, Thomas Schrefl, Josef Fidler, and Athanasios Bratsos. Costeffective way to speed up micromagnetic simulations in granular media. Applied Numerical Mathematics: Transactions of IMACS, 39(2):191–204, November 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/79/33/31/abstract.
html.

Tsitouras:2001:OER

[Tsi01] Ch. Tsitouras. Optimized explicit Runge–Kutta pair of orders 9(8). Applied Numerical Mathematics: Transactions of IMACS, 38(1–2): 123–134, July 2001. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/77/27/
33/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/77/27/33/article.
pdf.

Sheng:2016:HSC

[tSqWyG16] Chang tao Sheng, Zhong qing Wang, and Ben yu Guo. An hp-spectral collocation method for nonlinear Volterra functional integrodifferential equations with delays. Applied Numerical Mathematics: Transactions of IMACS, 105(??): 1–24, July 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416000143.

Tsynkov:1996:EEN

[Tsy96] Semyon V. Tsynkov. Errata to: “An effective numerical technique for solving a special class of ordinary difference equations” [Appl. Numer. Math. 18 (1995), no. 4, 489–501, MR 96g:65124] by V. S. Ryaben′kiĭ and Semyon V. Tsynkov. Applied Numerical Mathematics: Transactions of IMACS, 19(4):509, February 12, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=19&
issue=4&aid=660.  See

[RT95].

Tsynkov:1998:NSP

[Tsy98] Semyon V. Tsynkov. Numerical solution of problems on unbounded domains. A review. Applied Numerical Mathematics: Transactions of IMACS, 27(4):465–532, August 15, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.

elsevier.com/cas/tree/
store/apnum/sub/1998/27/
4/891.pdf.

Torres:2003:MSL

[TT03] Germán Torres and Cristina Turner. Method of straight lines for a Bingham problem in cylindrical pipes. Applied Numerical Mathematics: Transactions of IMACS, 47(3–4):543–558, December 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Tamasan:2020:MRS

[TT20] A. Tamasan and A. Timonov. The method of regularized successive iterations for coupled physics conductivity imaging from a single internal data. Applied Numerical Mathematics: Transactions of IMACS, 147(??):19–30, January 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419302089.

Tomar:2024:OEE

[TTP24] Aditi Tomar, Lok Pati Tripathi, and Amiya K. Pani. Optimal error estimates of a non-uniform IMEX-L1 finite element method for time fractional PDEs and PIDEs. Applied Numerical Mathematics: Transactions of IMACS, 205(??): 137–168, November 2024.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424001752.

Turkel:1986:ASN

[Tur86] Eli Turkel. Accuracy of schemes with nonuniform meshes for compressible fluid flows. Applied Numerical Mathematics: Transactions of IMACS, 2(6):529– 550, December 1986. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Turkel:1993:RPM

[Tur93] E. Turkel. Review of preconditioning methods for fluid dynamics. Applied Numerical Mathematics: Transactions of IMACS, 12(1–3): 257–284, May 1993. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0168927493901228.

Special issue to honor Professor Saul Abarbanel on his sixtieth birthday (Neveh, 1992).

tenThijeBoonkkamp:1987:OEH

[tV87] J. H. M. ten Thije Boonkkamp and J. G. Verwer. On the odd-even hopscotch scheme for the numerical integration of time-dependent partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 3(1–2):183–193, May 1987. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0168927487900110.

See corrigendum [BtTBV87].

Trompert:1991:SRM

[TV91] R. A. Trompert and J. G.

Verwer. A static-regridding method for two-dimensional parabolic partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 8(1):65– 90, August 1991. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Taylor:2000:GDM

[TW00] Mark A. Taylor and B. A.

Wingate. A generalized diagonal mass matrix spectral element method for nonquadrilateral elements. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4):259–265, May 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/61/27/
52/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/61/27/52/article.
pdf.

Tong:2023:DDK

[TWD23] Shanshan Tong, Wei Wang, and Chaofeng Dong. A data-driven Kaczmarz iterative regularization method with non-smooth constraints for ill-posed problems. Applied Numerical Mathematics: Transactions of IMACS, 192(??):152–178, October 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742300171X.

Tong:2021:AHP

[TWH21] Shanshan Tong, Wei Wang, and Bo Han. Accelerated homotopy perturbation iteration method for a nonsmooth nonlinear ill-posed problem. Applied Numerical Mathematics: Transactions of IMACS, 169(??): 122–145, November 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001768.

Tan:2023:SCR

[TWL23] Li Tan, Shengrong Wang, and Liangqing Luo. Strong convergence rate of implicit Euler scheme to a CIR model with delay. Applied Numerical Mathematics: Transactions of IMACS, 190(??): 15–26, August 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423000995.

Tamang:2020:NAB

[TWMP20] N. Tamang, B. Wongsaijai, T. Mouktonglang, and K. Poochinapan. Novel algorithm based on modification of Galerkin finite element method to general Rosenau– RLW equation in (2 + 1)- dimensions. Applied Numerical Mathematics: Transactions of IMACS, 148(??): 109–130, February 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741930193X.

Tang:2018:PSC

[TX18] Xiaojun Tang and Heyong Xu. A pseudospectral scheme and its convergence analysis for high-order integro-differential equations. Applied Numerical Mathematics: Transactions of IMACS, 125(??): 51–67, March 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417302209.

Turkel:1998:APB

[TY98] E. Turkel and A. Yefet.

Absorbing PML boundary layers for wave-like equations. Applied Numerical Mathematics: Transactions of IMACS, 27(4):533–557, August 15, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1998/27/
                                        4/892.pdf.

Turkel:2000:CHO

[TY00] E. Turkel and A. Yefet. On the construction of a high order difference scheme for complex domains in a Cartesian grid. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4): 113–124, May 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/21/61/27/
                                        35/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/21/61/27/35/article.
                                        pdf.

Trovon:2003:NHK

[TY03] Alexandre Trovon and Yanlin Yu. A note on heat kernel and approximation. Applied Numerical Mathematics: Transactions of IMACS, 47(2):265–271, November 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Tian:2011:CBI

[TYJ11] Hongjiong Tian, Quanhong Yu, and Cilai Jin. Continuous block implicit hybrid one-step methods for ordinary and delay differential equations. Applied Numerical Mathematics: Transactions of IMACS, 61(12): 1289–1300, December 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411001632.

Tang:2001:EBN

[TYKK01a] Dalin Tang, Chun Yang, Shunichi Kobayashi, and David N. Ku. Experimentbased numerical simulation of unsteady viscous flow in stenotic collapsible tubes. Applied Numerical Mathematics: Transactions of IMACS, 36(2– 3):299–320, February 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.nl../36/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/65/31/36/article.
pdf.

Tang:2001:GFD

[TYKK01b] Dalin Tang, Chun Yang, Shunichi Kobayashi, and David N. Ku. Generalized finite difference method for 3-D viscous flow in stenotic tubes with large wall deformation and collapse. Applied Numerical Mathematics: Transactions of IMACS, 38(1–2):49–68, July 2001.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/10/28/77/27/
                                        29/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/10/28/77/27/29/article.
                                        pdf.

Tangborn:2000:WTA

[TZ00] Andrew Tangborn and Sara Q. Zhang. Wavelet transform adapted to an approximate Kalman filter system. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4):307–316, May 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/21/61/27/
                                        58/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/21/61/27/58/article.
                                        pdf.

Tan:2021:DMP

[TZ21] Zengqiang Tan and Chengjian Zhang. The discrete maximum principle and energy stability of a new secondorder difference scheme for Allen–Cahn equations. Applied Numerical Mathematics: Transactions of IMACS, 166(??):227–237, August 2021. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001094.

Tocino:2013:LMS

[TZA13] A. Tocino, R. Zeghdane, and L. Abbaoui. Linear mean-square stability analysis of weak order 2.0 semiimplicit Taylor schemes for scalar stochastic differential equations. Applied Numerical Mathematics: Transactions of IMACS, 68(??): 19–30, June 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413000159.

Uddin:2020:SPM

[Udd20] M. Monir Uddin. Structure preserving model order reduction of a class of secondorder descriptor systems via balanced truncation. Applied Numerical Mathematics: Transactions of IMACS, 152(??):185–198, June 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419303423.

Usman:2024:ICS

[UHLZ24] Muhammad Usman, Muhammad Hamid, Dianchen Lu, and Zhengdi Zhang. Innovative coupling of s-stage one-step and spectral methods for non-smooth solutions of nonlinear problems. Applied Numerical Mathematics: Transactions of IMACS, 204(??): 329–351, October 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001405.

Usman:2021:LNO

[UHUL21] Muhammad Usman, Muhammad Hamid, Rizwan Ul Haq, and Moubin Liu. Linearized novel operational matricesbased scheme for classes of nonlinear time-space fractional unsteady problems in 2D. Applied Numerical Mathematics: Transactions of IMACS, 162(??): 351–373, April 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303949.

Siraj-ul-Islam:2013:LRB

[uIVS13] Siraj ul Islam, R. Vertnik, and B. Sarler. Local radial basis function collocation method along with explicit time stepping for hyperbolic partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 67(??): 136–151, May 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411001620.

Urquiza:2008:CSD

[UNGD08] J. M. Urquiza, D. N’Dri, A. Garon, and M. C. Delfour. Coupling Stokes and Darcy equations. Applied Numerical Mathematics: Transactions of IMACS, 58(5): 525–538, May 2008. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Ushakova:2018:CHC

[Ush18] Olga V. Ushakova. Criteria for hexahedral cell classification. Applied Numerical Mathematics: Transactions of IMACS, 127(??): 18–39, May 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417302623.

Ungor:2003:LBS

[ÜSHT03] Alper Üngör, Alla Sheffer, Robert B. Haber, and Shang-Hua Teng. Layer based solutions for constrained space–time meshing. Applied Numerical Mathematics: Transactions of IMACS, 46(3–4):425–443, September 2003. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Usmani:1997:ASF

[Usm97] Asif S. Usmani. An hadaptive SUPG-FEM solution of the pure advection equation. Applied Numerical Mathematics: Transactions of IMACS, 26(1–2):193–202, December 22, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1997/26/
1-2/846.pdf.

Utyuzhnikov:2008:RTW

[Uty08] S. V. Utyuzhnikov. Robintype wall functions and their numerical implementation. Applied Numerical Mathematics: Transactions of IMACS, 58(10): 1521–1533, October 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Umeda:2022:ECF

[UWY22] Akihiro Umeda, Yuta Wakasugi, and Shuji Yoshikawa. Energy-conserving finite difference schemes for nonlinear wave equations with dynamic boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 171(??): 1–22, January 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421002373.

Valko:2005:NID

[VA05] Peter P. Valkó and Joseph Abate. Numerical inversion of 2-D Laplace transforms applied to fractional diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 53(1):73–88, April 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Vaish:2021:HVI

[VA21] Rajat Vaish and Md. Kalimuddin Ahmad. Hybrid viscosity implicit scheme for variational inequalities over the fixed point set of an asymptotically nonexpansive mapping in the intermediate sense in Banach spaces. Applied Numerical Mathematics: Transactions of IMACS, 160(??): 296–312, February 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742030324X.

Vabishchevich:2021:SSN

[Vab21] Petr N. Vabishchevich. Splitting schemes for non-stationary problems with a rational approximation for fractional powers of the operator. Applied Numerical Mathematics: Transactions of IMACS, 165(??):414–430, July 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421000763.

Vabishchevich:2022:NSC

[Vab22] P. N. Vabishchevich. Numerical solution of the Cauchy problem for Volterra integrodifferential equations with difference kernels. Applied Numerical Mathematics: Transactions of IMACS, 174(??):177–190, April 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422000137.

vanIseghem:1986:ECR

[van86a] Jeannette van Iseghem. An extended cross rule for vector Padé approximants. Applied Numerical Mathematics: Transactions of IMACS, 2(2):143–155, April 1986. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

vanLeer:1986:NDU

[van86b] Bram van Leer. On numerical dispersion by upwind differencing. Applied Numerical Mathematics: Transactions of IMACS, 2(3–5):379– 384, October 1986. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Vanaja:1992:NSS

[Van92] V. Vanaja. Numerical solution of a simple Fokker– Planck equation. Applied Numerical Mathematics: Transactions of IMACS, 9(6):533–540, May 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

vanderHouwen:1993:PSS

[van93] P. J. van der Houwen. Parallel step-by-step methods. Applied Numerical Mathematics: Transactions of IMACS, 11(1–3):69–81, January 1993. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Parallel methods for ordinary differential equations (Grado, 1991).

vanGijzen:1995:LSF

[van95] M. B. van Gijzen. Large scale finite element computations with GMRESlike methods on a Cray Y-MP. Applied Numerical Mathematics: Transactions of IMACS, 19(1–2): 51–62, December 4, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=1-2&aid=569.

vanderHouwen:1996:DRK

[van96] P. J. van der Houwen. The development of Runge– Kutta methods for partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 20(3):261–272, May 22, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=3&aid=657. Selected

keynote papers presented at 14th IMACS World Congress (Atlanta, GA, 1994).

vanLeersum:1998:NMH

[van98] J. van Leersum. A numerical model of a high performance two-stroke engine. Applied Numerical Mathematics: Transactions of IMACS, 27(1):83–108, May 1, 1998. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1998/27/
1/864.pdf.

VandenHeuvel:2000:NSR

[Van00] Guido Van den Heuvel. New stability results for Runge– Kutta methods adapted to delay differential equations. Applied Numerical Mathematics: Transactions of IMACS, 34(2–3):

293–308, July 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/62/29/
38/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/62/29/38/article.
pdf.

Varga:1992:HHP

[Var92] Richard S. Varga. How highprecision calculations can stimulate mathematical research. Applied Numerical Mathematics: Transactions of IMACS, 10(3–4): 177–193, September 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). A Festschrift to honor Professor Garrett Birkhoff on his eightieth birthday.

Vasconcelos:2017:DSA

[Vas17] Paulo B. Vasconcelos. Datasparse approximation on the computation of a weakly singular Fredholm equation: a stellar radiative transfer application. Applied Numerical Mathematics: Transactions of IMACS, 114(??): 55–62, April 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741630160X.

vonSchwerin:1995:RKS

[vB95] Reinhold von Schwerin and Hans Georg Bock. A Runge– Kutta starter for a multistep method for differentialalgebraic systems with discontinuous effects. Applied Numerical Mathematics: Transactions of IMACS, 18(1–3):337–350, September 27, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=18&
                                        issue=1-3&aid=614. Sev-

enth Conference on the Numerical Treatment of Differential Equations (Halle, 1994).

Vandenberghe:1999:ASP

[VB99] Lieven Vandenberghe and Stephen Boyd. Applications of semidefinite programming. Applied Numerical Mathematics: Transactions of IMACS, 29(3):283– 299, March 1, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1999/29/
                                        3/963.pdf.

VanCriekingen:2007:MHD

[VB07] S. Van Criekingen and R. Beauwens. Mixed-hybrid discretization methods for the P_{1} equations. Applied Numerical Mathematics: Transactions of IMACS, 57(2):117–130, February 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

VandenBerghe:1993:NSF

[VBD93] G. Vanden Berghe, P. Bocher, and H. De Meyer. Numerical solution of Fredholm equations based on mixed interpolation. Applied Numerical Mathematics: Transactions of IMACS, 13(1– 3):15–22, September 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Sixth Conference on the Numerical Treatment of Differential Equations (Halle, 1992).

Verwer:1996:IEA

[VBH96] J. G. Verwer, J. G. Blom, and W. Hundsdorfer. An implicit-explicit approach for atmospheric transportchemistry problems. Applied Numerical Mathematics: Transactions of IMACS, 20(1–2):191–209, February 29, 1996. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=20&
issue=1-2&aid=674. Work-

shop on the method of lines for time-dependent problems (Lexington, KY, 1995).

Viguerie:2022:TNA

[VBVA22] Alex Viguerie, Silvia Bertoluzza, Alessandro Veneziani, and Ferdinando Auricchio. A theoretical and numerical analysis of a Dirichlet– Neumann domain decomposition method for diffusion problems in heterogeneous media. Applied Numerical Mathematics: Transactions of IMACS, 173(??): 94–111, March 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421003317.

vanBrummelen:2008:STM

[vBvdZdB08] E. H. van Brummelen, K. G. van der Zee, and R. de Borst. Space/time multigrid for a fluid–structure-interaction problem. Applied Numerical Mathematics: Transactions of IMACS, 58(12): 1951–1971, December 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

vanderHouwen:1993:PBP

[vC93] P. J. van der Houwen and Nguyen Huu Cong. Parallel block predictor-corrector methods of Runge–Kutta type. Applied Numerical Mathematics: Transactions of IMACS, 13(1–3):

109–123, September 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Sixth Conference on the Numerical Treatment of Differential Equations (Halle, 1992).

Varga:2010:ZPS

[VC10] Richard S. Varga and Amos J. Carpenter. Zeros of the partial sums of cos(z) and sin(z), III. Applied Numerical Mathematics: Transactions of IMACS, 60(4):298–313, April 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Vazquez-Cendon:2012:ANK

[VCC12] M. Elena Vázquez-Cendón and Luis Cea. Analysis of a new Kolgan-type scheme motivated by the shallow water equations. Applied Numerical Mathematics: Transactions of IMACS, 62(4): 489–506, April 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411000961.

Varma:2020:DGS

[VCN20] V. Dhanya Varma, Nagaiah Chamakuri, and Suresh Kumar Nadupuri. Discontinuous Galerkin solution of the convection–diffusion– reaction equations in fluidized beds. Applied Numerical Mathematics: Transactions of IMACS, 153(??): 188–201, July 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742030043X.

vandenEshof:2004:ACG

[vdES04] Jasper van den Eshof and Gerard L. G. Sleijpen. Accurate conjugate gradient methods for families of shifted systems. Applied Numerical Mathematics: Transactions of IMACS, 49(1):17–37, April 2004. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

vanderHouwen:1999:PSC

[vdHMdS99] P. J. van der Houwen, E. Messina, and J. J. B. de Swart. Parallel Störmer– Cowell methods for highprecision orbit computations. Applied Numerical Mathematics: Transactions of IMACS, 31(3): 353–374, November 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/46/19/
22/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/46/19/22/article.
pdf.

vanderHouwen:2001:FBT

[vdHS01] P. J. van der Houwen and B. P. Sommeijer. Factorization in block-triangularly implicit methods for shallow water applications. Applied Numerical Mathematics: Transactions of IMACS, 36(2–3):113–128, February 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.nl../26/
abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/65/31/26/article.
pdf.

vanderHouwen:1998:SPC

[vdHSW98] P. J. van der Houwen, K. Strehmel, and R. Weiner. Selected papers from the 8th conference on the numerical treatment of differential equations, September 1–5, 1997, Alexisbad, Germany. Applied Numerical Mathematics: Transactions of IMACS, 28(2–4): 91–94, 1998. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

vanderHeul:2001:SAS

[vdHVW01] D. R. van der Heul, C. Vuik, and P. Wesseling. Stability analysis of segregated solution methods for compressible flow. Applied Numerical Mathematics: Transactions of IMACS, 38(3):257– 274, August 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/77/33/28/abstract.
html.

vanderVeen:1995:CAS

[vdSvdH95] W. A. van der Veen, J. J. B. de Swart, and P. J. van der Houwen. Convergence aspects of step-parallel iteration of Runge–Kutta methods. Applied Numerical Mathematics: Transactions of IMACS, 18(1–3):397–411, September 27, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=18&
issue=1-3&aid=611. Sev-

enth Conference on the Numerical Treatment of Differential Equations (Halle, 1994).

vanderVegt:2008:STD

[vdVS08] J. J. W. van der Vegt and J. J. Sudirham. A space–time discontinuous Galerkin method for the time-dependent Oseen equations. Applied Numerical Mathematics: Transactions of IMACS, 58(12): 1892–1917, December 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

VanDaele:1998:CSM

[VDVV98] M. Van Daele, H. De Meyer, T. Van Hecke, and G. Vanden Berghe. On a class of P - stable mono-implicit Runge– Kutta–Nyström methods. Applied Numerical Mathematics: Transactions of IMACS, 27(1):69–82, May 1, 1998. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1998/27/
1/866.pdf.

Vejchodsky:2010:HOD

[Vej10] Tomáš Vejchodský. Higherorder discrete maximum principle for 1D diffusion– reaction problems. Applied Numerical Mathematics: Transactions of IMACS, 60(4):486–500, April 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Venutelli:2015:NOF

[Ven15] Maurizio Venutelli. New optimized fourth-order compact finite difference schemes for wave propagation phenomena. Applied Numerical Mathematics: Transactions of IMACS, 87(??): 53–73, January 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414001512.

Verlan:1993:AFS

[Ver93] Igor I. Verlan. About a family of C^{4} splines with one free generating function. Applied Numerical Mathematics: Transactions of IMACS, 13(5):423–435, December 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Verner:1996:HOE

[Ver96a]  J. H. Verner.  High-

order explicit Runge–Kutta pairs with low stage order. Applied Numerical Mathematics: Transactions of IMACS, 22(1–3):345–357, December 4, 1996. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=22&
                                        issue=1-3&aid=723. Special

issue celebrating the centenary of Runge–Kutta methods.

Verwer:1996:ERK

[Ver96b]  J. G. Verwer.  Explicit

Runge–Kutta methods for parabolic partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 22(1–3): 359–379, December 4, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=22&
issue=1-3&aid=704. Special

issue celebrating the centenary of Runge–Kutta methods.

Verner:2006:ISM

[Ver06] J. H. Verner. Improved starting methods for twostep Runge–Kutta methods of stage-order p − 3. Applied Numerical Mathematics: Transactions of IMACS, 56(3–4):388–396, March/April 2006. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Vainikko:2004:PSP

[VG04] Eero Vainikko and Ivan G.

Graham. A parallel solver for PDE systems and application to the incompressible Navier–Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 49(1): 97–116, April 2004. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Vasconcelos:2024:LRA

[VGL24] Paulo B. Vasconcelos, Laurence Grammont, and Nilson J. Lima. Low rank approximation in the computation of first kind integral equations with Tau-Toolbox. Applied Numerical Mathematics: Transactions of IMACS, 205(??):1– 15, November 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424001703.

vanVeldhuizen:1998:NCH

[vHA98] M. van Veldhuizen, J. A. Hendriks, and C. A. J. Appelo. Numerical computation in heterovalent chromatography. Applied Numerical Mathematics: Transactions of IMACS, 28(1):69–89, September 1, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1998/28/
                                        1/882.pdf.

Verwer:2012:P

[VHAW12] Jan Verwer, Willem Hundsdorfer, Martin Arnold, and Rüdiger Weiner. Preface. Applied Numerical Mathematics: Transactions of IMACS, 62(10):1257– 1258, October 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927412000876.

vanIseghem:1987:LTI

[vI87] Jeannette van Iseghem. Laplace transform inversion and Padé-type approximants. Applied Numerical Mathematics: Transactions of IMACS, 3(6):529– 538, November 1987. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Vichnevetsky:1987:F

[Vic87a] Robert Vichnevetsky. Foreword. Applied Numerical Mathematics: Transactions of IMACS, 3(1– 2):1–2, May 1987. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/016892748790002X.

Vichnevetsky:1987:WPR

[Vic87b] Robert Vichnevetsky. Wave propagation and reflection in irregular grids for hyperbolic equations. Applied Numerical Mathematics: Transactions of IMACS, 3(1–2):133–166, May 1987. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0168927487900092.

Vichnevetsky:1992:WPA

[Vic92] Robert Vichnevetsky. On wave propagation in almost periodic structures. Applied Numerical Mathematics: Transactions of IMACS, 10(3–4):195–229, September 1992. CODEN ANMAEL.

ISSN 0168-9274 (print), 1873-5460 (electronic). A Festschrift to honor Professor Garrett Birkhoff on his eightieth birthday.

Vichnevetsky:2005:E

[Vic05] Robert Vichnevetsky. Editorial. Applied Numerical Mathematics: Transactions of IMACS, 55(4): 383, December 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Vass:2017:FSD

[VK17] József Vass and Sergey N. Krylov. A fast stable discretization of the Constant-Convection–Diffusion-Reaction equations of Kinetic Capillary Electrophoresis (KCE). Applied Numerical Mathematics: Transactions of IMACS, 122(??):82–91, December 2017. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417301678.

Veldman:2008:SPU

[VL08] Arthur E. P. Veldman and Ka-Wing Lam. Symmetrypreserving upwind discretization of convection on nonuniform grids. Applied Numerical Mathematics: Transactions of IMACS, 58(12): 1881–1891, December 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Vong:2019:SOS

[VL19] Seakweng Vong and Pin Lyu.

On a second order scheme for space fractional diffusion equations with variable coefficients. Applied Numerical Mathematics: Transactions of IMACS, 137(??): 34–48, March 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418302654.

Vrahatis:2003:LNI

[VMP03] M. N. Vrahatis, G. D.

Magoulas, and V. P. Plagianakos. From linear to nonlinear iterative methods. Applied Numerical Mathematics: Transactions of IMACS, 45(1):59–77, April 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Vohralik:2007:MNF

[VMS07] M. Vohralík, J. Maryška, and O. Severýn. Mixed and nonconforming finite element methods on a system of polygons. Applied Numerical Mathematics: Transactions of IMACS, 57(2):176– 193, February 2007. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Vulanovic:2021:IKT

[VN21] Relja Vulanović and Thái Anh Nhan. An improved Kellogg– Tsan solution decomposition in numerical methods for singularly perturbed convection-diffusion problems. Applied Numerical Mathematics: Transactions of IMACS, 170(??): 128–145, December 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421002105.

Varma:2021:PEE

[VNC21] V. Dhanya Varma, Suresh Kumar Nadupuri, and Nagaiah Chamakuri. A posteriori error estimates and an adaptive finite element solution for the system of unsteady convection-diffusion-reaction equations in fluidized beds. Applied Numerical Mathematics: Transactions of IMACS, 163(??): 108–125, May 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000209.

Valenciano:2000:ASE

[VO00a] José Valenciano and Robert G. Owens. An h-p adaptive spectral element method for Stokes flow. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4):

365–371, May 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/61/27/
64/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/61/27/64/article.
pdf.

Valenciano:2000:NAM

[VO00b] José Valenciano and Robert G.

Owens. A new adaptive modification strategy for numerical solutions to elliptic boundary value problems. Applied Numerical Mathematics: Transactions of IMACS, 32(3):305– 329, March 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/59/30/
31/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/59/30/31/article.
pdf.

Vandewalle:1991:NEN

[VP91] Stefan Vandewalle and Robert Piessens. Numerical experiments with nonlinear multigrid waveform relaxation on a parallel processor. Applied Numerical Mathematics: Transactions of IMACS, 8(2):149– 161, September 1991. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0168927491900485.

Vera:2001:CSA

[VR01] E. Chacón Vera and T. Chacón Rebollo. On cubic spline approximations for the vortex patch problem. Applied Numerical Mathematics: Transactions of IMACS, 36(4):359–387, March 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/10/28/65/32/
                                        26/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/10/28/65/32/26/article.
                                        pdf.

vanRaalte:2004:FSD

[vR04] M. H. van Raalte. A feasibility study for discontinuous Galerkin discretization with embedded Dirichlet boundary condition. Applied Numerical Mathematics: Transactions of IMACS, 51(2–3):361–383, November 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Verma:2021:NSC

[VRC21] Amit Kumar Verma, Mukesh Kumar Rawani, and Carlo Cattani. A numerical scheme for a class of generalized Burgers’ equation based on Haar wavelet nonstandard finite difference method. Applied Numerical Mathematics: Transactions of IMACS, 168(??): 41–54, October 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001483.

Vedral:2025:SCL

[VRK25] Joshua Vedral, Andreas Rupp, and Dmitri Kuzmin. Strongly consistent lowdissipation WENO schemes for finite elements. Applied Numerical Mathematics: Transactions of IMACS, 210(??):64–81, April 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424003490.

Vichnevetsky:1991:NNI

[VS91] Robert Vichnevetsky and Thomas Scheidegger. The nonlocal nature of internal reflection in computational fluid dynamics with spectral methods. Applied Numerical Mathematics: Transactions of IMACS, 8(6):533– 539, December 1991. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

vanderHouwen:1993:APD

[vS93] P. J. van der Houwen and B. P. Sommeijer. Analysis of parallel diagonally implicit iteration of Runge–Kutta methods. Applied Numerical Mathematics: Transactions of IMACS, 11(1– 3):169–188, January 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Parallel methods for ordinary differential equations (Grado, 1991).

VanderHouwen:1994:BKM

[VS94] P. J. Van der Houwen and B. P. Sommeijer. Butcher– Kuntzmann methods for nonstiff problems on parallel computers. Applied Numerical Mathematics: Transactions of IMACS, 15(3): 357–374, October 5, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1994&volume=15&
issue=3&aid=508. Interna-

tional Conference on Scientific Computation and Differential Equations (Auckland, 1993).

Verwer:1995:EMS

[VS95] J. G. Verwer and D. Simpson. Explicit methods for stiff ODEs from atmospheric chemistry. Applied Numerical Mathematics: Transactions of IMACS, 18(1–3):413–430, September 27, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=18&
issue=1-3&aid=616.

vanderHouwen:1996:CCD

[vS96] P. J. van der Houwen and B. P. Sommeijer. CWI contributions to the development of parallel Runge– Kutta methods. Applied Numerical Mathematics: Transactions of IMACS, 22(1–3):327–344, December 4, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=22&
issue=1-3&aid=729. Special

issue celebrating the centenary of Runge–Kutta methods.

vanderHouwen:1997:ECM

[vS97] P. J. van der Houwen and B. P. Sommeijer. Euler– Chebyshev methods for integro-differential equations. Applied Numerical Mathematics: Transactions of IMACS, 24(2–3): 203–218, June 19, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/

                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=24&
                                        issue=2-3&aid=777. Volterra

centennial (Tempe, AZ, 1996).

Verwer:2004:P

[VS04] Jan G. Verwer and Ben P. Sommeijer. Preface. Applied Numerical Mathematics: Transactions of IMACS, 48(3–4):253, March 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

vanderHouwen:1992:PDI

[vSC92] P. J. van der Houwen, B. P. Sommeijer, and Nguyen Huu Cong. Parallel diagonally implicit Runge–Kutta– Nyström methods. Applied Numerical Mathematics: Transactions of IMACS, 9(2):111–131, February 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Vuik:2002:CVD

[VSeYD02] C. Vuik, A. Segal, L. el Yaakoubi, and E. Dufour. A comparison of various deflation vectors applied to elliptic problems with discontinuous coefficients. Applied Numerical Mathematics: Transactions of IMACS, 41(1):219– 233, April 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/

10/28/86/27/39/abstract.
html.

VanBockstal:2017:IMK

[VSG17] K. Van Bockstal, M. Slodicka, and F. Gistelinck. Identification of a memory kernel in a nonlinear integrodifferential parabolic problem. Applied Numerical Mathematics: Transactions of IMACS, 120(??): 305–323, October 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741730140X.

vanderHouwen:1997:ISF

[vSK97] P. J. van der Houwen, B. P.

Sommeijer, and J. Kok. The iterative solution of fully implicit discretizations of three-dimensional transport models. Applied Numerical Mathematics: Transactions of IMACS, 25(2–3): 243–256, October 17, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=25&
issue=2-3&aid=818.

vanderHouwen:1990:ASO

[vSW90] P. J. van der Houwen, B. P.

Sommeijer, and F. W. Wubs. Analysis of smoothing operators in the solution of partial differential equations by explicit difference schemes. Applied Numerical Mathematics: Transactions of IMACS, 6(6):501–521, October 1990. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Vichnevetsky:1991:SSD

[VT91] R. Vichnevetsky and L. H. Turner. Spurious scattering from discontinuously stretching grids in computational fluid dynamics. Applied Numerical Mathematics: Transactions of IMACS, 8(3):289–299, October 1991. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Verwer:1993:ALU

[VT93] J. G. Verwer and R. A. Trompert. Analysis of local uniform grid refinement. Applied Numerical Mathematics: Transactions of IMACS, 13(1–3): 251–270, September 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). Sixth Conference on the Numerical Treatment of Differential Equations (Halle, 1992).

Vulanovic:1992:SSS

[Vul92] R. Vulanović. Some switching schemes for quasilinear singular perturbation problems. Applied Numerical Mathematics: Transactions of IMACS, 9(2):143– 155, February 1992. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Vulanovic:1995:UCR

[Vul95] Relja Vulanović. A uniform convergence result for semilinear singular perturbation problems violating the standard stability condition. Applied Numerical Mathematics: Transactions of IMACS, 16(3): 383–399, February 16, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=16&
issue=3&aid=536.

Verstappen:1995:DNS

[VV95] R. W. C. P. Verstappen and A. E. P. Veldman. Direct numerical simulation of turbulence on a Connection Machine CM-5. Applied Numerical Mathematics: Transactions of IMACS, 19(1–2): 147–158, December 4, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=1-2&aid=570.

VanDaele:2002:ACH

[VV02] M. Van Daele and T. Van Hecke. Algebraic conditions for high-order convergent deferred correction schemes based on Runge– Kutta–Nyström methods for second order boundary value problems. Applied Numerical Mathematics: Transactions of IMACS, 42(1– 3):453–464, August 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

VanDaele:2005:EFQ

[VV05] M. Van Daele and G. Vanden Berghe and H. Vande Vyver. Exponentially fitted quadrature rules of Gauss type for oscillatory integrands. Applied Numerical Mathematics: Transactions of IMACS, 53(2–4):509–526, May 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

VanDaele:2007:GNI

[VV07] M. Van Daele and G. Vanden Berghe. Geometric numerical integration by means of exponentially-fitted methods. Applied Numerical Mathematics: Transactions of IMACS, 57(4):415–435, April 2007. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

VandenBerghe:2009:EFO

[VV09] G. Vanden Berghe and M. Van Daele. Exponentiallyfitted Obrechkoff methods for second-order differential equations. Applied Numerical Mathematics: Transactions of IMACS, 59(3–4): 815–829, March/April 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

VandenBerghe:1995:MDS

[VVD95] G. Vanden Berghe, M. Van Daele, and H. De Meyer. A modified difference scheme for periodic and semiperiodic Sturm–Liouville problems. Applied Numerical Mathematics: Transactions of IMACS, 18(1–3):69–78, September 27, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=18&
issue=1-3&aid=615. Sev-

enth Conference on the Numerical Treatment of Differential Equations (Halle, 1994).

vanderHouwen:1997:SID

[vvdV97] P. J. van der Houwen and W. A. van der Veen. The solution of implicit differential equations on parallel computers. Applied Numerical Mathematics: Transactions of IMACS, 25(2–3): 257–274, October 17, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=25&
issue=2-3&aid=827.

Vandekerckhove:2008:ALB

[VVR08] Christophe Vandekerckhove, Pieter Van Leemput, and Dirk Roose. Acceleration of lattice Boltzmann models through state extrapolation: a reaction–diffusion example. Applied Numerical Mathematics: Transactions of IMACS, 58(11): 1742–1757, November 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

VanBarel:2024:AMR

[VVV24] Marc Van Barel, Niel Van Buggenhout, and Raf Vandebril. Algorithms for modifying recurrence relations of orthogonal polynomial and rational functions when changing the discrete inner product. Applied Numerical Mathematics: Transactions of IMACS, 200(??): 429–452, June 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001927.

Vermiglio:1993:MNC

[VZ93] Rossana Vermiglio and Marino Zennaro. Multistep natural continuous extensions of Runge–Kutta methods: the potential for stable interpolation. Applied Numerical Mathematics: Transactions of IMACS, 12(6):521– 546, August 1993. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Waadeland:1988:CCF

[Waa88] Haakon Waadeland. Computation of continued fractions by square-root modification: reflections and examples. Applied Numerical Mathematics: Transactions of IMACS, 4(2–4): 361–375, June 1988. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wagatha:1985:BCN

[Wag85] Ludwig Wagatha. On boundary conditions for the numerical simulation of wave propagation. Applied Numerical Mathematics: Transactions of IMACS, 1(4):309– 314, July 1985. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wagner:1998:MGA

[Wag98] Yvonne Wagner. A moving grid algorithm for a heat exchanger with phase changes. Applied Numerical Mathematics: Transactions of IMACS, 28(2–4):477–491, October 1, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1998/28/
2-4/927.pdf.

Wain:1998:SFP

[Wai98] Richard J. Wain. Spurious fixed points of a variable step size, variable order, predictor corrector algorithm. Applied Numerical Mathematics: Transactions of IMACS, 27(2):155– 169, June 10, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1998/27/
2/872.pdf.

Walz:1990:ICM

[Wal90] Guido Walz. Increasing the convergence modulus of an asymptotic expansion. Applied Numerical Mathematics: Transactions of IMACS, 6(5):415–423, July 1990. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Walker:1995:RSP

[Wal95] Homer F. Walker. Residual smoothing and peak/plateau behavior in Krylov subspace methods. Applied Numerical Mathematics: Transactions of IMACS, 19(3):279– 286, December 15, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=19&
issue=3&aid=635;   http:
//www.sciencedirect.com/
science/article/pii/0168927495000879.

Special issue on iterative methods for linear equations (Atlanta, GA, 1994).

Walden:2000:FBS

[Wal00a] Johan Waldén. Filter bank subdivisions of bounded domains. Applied Numerical Mathematics: Transactions of IMACS, 32(3):331– 357, March 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/59/30/
32/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/59/30/32/article.
pdf.

Walden:2000:GAS

[Wal00b] Johan Waldén. A general adaptive solver for hyperbolic PDEs based on filter bank subdivisions. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4):317–325, May 2000.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/21/61/27/
                                        59/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/21/61/27/59/article.
                                        pdf.

Walloth:2019:REL

[Wal19] Mirjam Walloth. A reliable, efficient and localized error estimator for a discontinuous Galerkin method for the Signorini problem. Applied Numerical Mathematics: Transactions of IMACS, 135(??): 276–296, January 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418301983.

Wang:1996:CMI

[Wan96]  Xinmin Wang.  Conver-

gence for the MSOR iterative method applied to Hmatrices. Applied Numerical Mathematics: Transactions of IMACS, 21(4):469– 479, November 22, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=21&
                                        issue=4&aid=698.

Wang:2001:ABN

[Wan01] Yuan-Ming Wang. Asymptotic behavior of the numerical solutions for a system of nonlinear integrodifferential reaction-diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 39(2):205–223, November 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/79/33/32/abstract.
html.

Wang:2005:OSP

[Wan05]  Yanqiu Wang.  Overlap-

ping Schwarz preconditioner for the mixed formulation of plane elasticity. Applied Numerical Mathematics: Transactions of IMACS, 54(2):292–309, July 2005. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wang:2007:ENS

[Wan07a] Yuan-Ming Wang. The extrapolation of Numerov’s scheme for nonlinear twopoint boundary value problems. Applied Numerical Mathematics: Transactions of IMACS, 57(3):253–269, March 2007. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wang:2007:MIT

[Wan07b] Yuan-Ming Wang. Monotone iterative technique for numerical solutions of fourth-order nonlinear elliptic boundary value problems. Applied Numerical Mathematics: Transactions of IMACS, 57(10): 1081–1096, October 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wang:2009:NSM

[Wan09] Yuan-Ming Wang. Numerical solutions of a Michaelis– Menten-type ratio-dependent predator–prey system with diffusion. Applied Numerical Mathematics: Transactions of IMACS, 59(5): 1075–1093, May 2009. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wang:2011:NMC

[Wan11] Yuan-Ming Wang. On Numerov’s method for a class of strongly nonlinear twopoint boundary value problems. Applied Numerical Mathematics: Transactions of IMACS, 61(1):38–52, January 2011. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wang:2017:FGM

[Wan17a] Wansheng Wang. Fullygeometric mesh one-leg methods for the generalized pantograph equation: Approximating Lyapunov functional and asymptotic contractivity. Applied Numerical Mathematics: Transactions of IMACS, 117(??): 50–68, July 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741730034X.

Wang:2017:EAC

[Wan17b] Yuan-Ming Wang. Error analysis of a compact finite difference method for fourth-order nonlinear elliptic boundary value problems. Applied Numerical Mathematics: Transactions of IMACS, 120(??): 53–67, October 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417301071.

Wang:2020:SAE

[Wan20] Junjun Wang. Superconvergence analysis of an energy stable scheme for nonlinear reaction–diffusion equation with BDF mixed FEM. Applied Numerical Mathematics: Transactions of IMACS, 153(??):457–472, July 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742030074X.

Wang:2021:HOC

[Wan21] Junjie Wang. High-order conservative schemes for the space fractional nonlinear Schrödinger equation. Applied Numerical Mathematics: Transactions of IMACS, 165(??):248–269, July 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421000672.

Wang:2023:CST

[Wan23] Xiaoliang Wang. A class of spectral three-term descent Hestenes–Stiefel conjugate gradient algorithms for large-scale unconstrained optimization and image restoration problems. Applied Numerical Mathematics: Transactions of IMACS, 192(??):41–56, October 2023. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423001538.

Wan:2025:NSI

[Wan25] Yifei Wan. A new scaleinvariant hybrid WENO scheme for steady Euler and Navier–Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 210(??):

177–198, April 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424003556.

Ward:1992:NGB

[War92] William A. Ward, Jr. Numerically generated basis functions for elliptic boundary value problems. Applied Numerical Mathematics: Transactions of IMACS, 10(5):415–433, October 1992. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Weber:2012:QAS

[WAV12] Steffen Weber, Martin Arnold, and Michael Valásek. Quasistatic approximations for stiff second order differential equations. Applied Numerical Mathematics: Transactions of IMACS, 62(10): 1579–1590, October 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412001195.

Wei:2024:SDG

[WaZ24] Yuanhong Wei and Guang an Zou. A splitting discontinuous Galerkin projection method for the magnetohydrodynamic equations. Applied Numerical Mathematics: Transactions of IMACS, 197(??):363–388, March 2024. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423003033.

Wang:2021:SPP

[WaZW21] Xue Wang, Guang an Zou, and Bo Wang. The stabilized penalty-projection finite element method for the Navier– Stokes–Cahn–Hilliard–Oono system. Applied Numerical Mathematics: Transactions of IMACS, 165(??): 376–413, July 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742100074X.

Wang:2023:SDF

[WaZW23] Xuyang Wang, Guang an Zou, and Bo Wang. A stabilized divergence-free virtual element scheme for the nematic liquid crystal flows. Applied Numerical Mathematics: Transactions of IMACS, 192(??): 104–131, October 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423001691.

Wimp:1990:FSA

[WB90] Jet Wimp and Bob Boyer. Formal series and an algorithm for computing some special determinants with elements in a ring. Applied Numerical Mathematics: Transactions of IMACS, 6(3):239–248, March 1990. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wille:1992:DNC

[WB92a] David R. Willé and Christopher T. H. Baker. DELSOL

— a numerical code for the solution of systems of delaydifferential equations. Applied Numerical Mathematics: Transactions of IMACS, 9(3–5):223–234, April 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0168927492900178.

Wille:1992:TDD

[WB92b] David R. Willé and Christopher T. H. Baker. The tracking of derivative discontinuities in systems of delaydifferential equations. Applied Numerical Mathematics: Transactions of IMACS, 9(3–5):209–222, April 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0168927492900167.

International Conference on the Numerical Solution of Volterra and Delay Equations (Tempe, AZ, 1990).

Witelski:2003:ASH

[WB03] T. P. Witelski and M. Bowen. ADI schemes for higherorder nonlinear diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 45(2–3): 331–351, May 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Williams:2002:FSI

[WBCK02] Russell Williams, Kevin Burrage, Ian Cameron, and Minnie Kerr. A four-stage index 2 Diagonally Implicit Runge–Kutta method. Applied Numerical Mathematics: Transactions of IMACS, 40(3):415– 432, February 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/gej-ng/10/
                                        10/28/85/33/29/abstract.
                                        html.

Wensch:2002:ELG

[WC02] Jörg Wensch and Fernando Casas. Extrapolation in Lie groups with approximated BCH-formula. Applied Numerical Mathematics: Transactions of IMACS, 42(1–3):465–472, August 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wang:2011:MSS

[WC11] Wenqiang Wang and Yanping Chen. Mean-square stability of semi-implicit Euler method for nonlinear neutral stochastic delay differential equations. Applied Numerical Mathematics: Transactions of IMACS, 61(5):696–701, May 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wei:2014:LSC

[WC14] Yunxia Wei and Yanping Chen. Legendre spectral collocation method for neutral and high-order Volterra integro-differential equation. Applied Numerical Mathematics: Transactions of IMACS, 81(??): 15–29, July 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414000415.

Wang:2024:PPW

[WC24a] Xue Wang and Guoxian Chen. A positivitypreserving well-balanced wetdry front reconstruction for shallow water equations on rectangular grids. Applied Numerical Mathematics: Transactions of IMACS, 198(??):295–317, April 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000126.

Wang:2024:IPE

[WC24b] Xuesong Wang and Yao Cheng. An improved pointwise error estimate of the LDG method for 1-d singularly perturbed reaction– diffusion problem. Applied Numerical Mathematics: Transactions of IMACS, 196(??):199–217, February 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423002556.

Wang:2025:WBP

[WC25] Xue Wang and Guoxian Chen. Well-balanced and positivity-preserving wet-dry front reconstruction scheme for Ripa models. Applied Numerical Mathematics: Transactions of IMACS, 213(??):38–60, July 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425000388.

Wolters:1995:DPN

[WCGW95] Lex Wolters, Gerard Cats, Nils Gustafsson, and Tomas Wilhelmsson. Data-parallel numerical methods in a weather forecast model. Applied Numerical Mathematics: Transactions of IMACS, 19(1–2):159–171, December 4, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1995&volume=19&
issue=1-2&aid=571.

Wang:2023:UOE

[WCJ23] Tingchun Wang, Yue Cheng, and Lihai Ji. Unconditional optimal error estimates of conservative methods for Klein–Gordon–Dirac system in two dimensions. Applied Numerical Mathematics: Transactions of IMACS, 183(??):263–278, January 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422002501.

Wang:2022:SPG

[WCL22] Yibo Wang, Wanrong Cao, and Shengyue Li. A spectral Petrov–Galerkin method for optimal control problem governed by a fractional ordinary differential equation. Applied Numerical Mathematics: Transactions of IMACS, 177(??): 18–33, July 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422000642.

Wang:2021:SEE

[WCM21] Zhibo Wang, Dakang Cen, and Yan Mo. Sharp error estimate of a compact L1-ADI scheme for the two-dimensional timefractional integro-differential equation with singular kernels. Applied Numerical Mathematics: Transactions of IMACS, 159(??): 190–203, January 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302828.

Wang:2023:EIM

[WCM23] Wansheng Wang, Qiong Chen, and Mengli Mao. An efficient IMEX method for nonlinear functional differential equations with statedependent delay. Applied Numerical Mathematics: Transactions of IMACS, 185(??):56–71, March 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742200294X.

Wang:2021:RNC

[WCS21] Yue Wang, Hu Chen, and Tao Sun. α-robust H^{1}- norm convergence analysis of ADI scheme for twodimensional time-fractional diffusion equation. Applied Numerical Mathematics: Transactions of IMACS, 168(??):75–83, October 2021. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001549.

Wang:2018:CEF

[WCSQ18] Hua Wang, Jinru Chen, Pengtao Sun, and Fangfang Qin. A conforming enriched finite element method for elliptic interface problems. Applied Numerical Mathematics: Transactions of IMACS, 127(??):1–17, May 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417302611.

Wang:2014:MAS

[WCW14] Xin Wang, Liqun Cao, and Yaushu Wong. Multiscale approach for stochastic elliptic equations in heterogeneous media. Applied Numerical Mathematics: Transactions of IMACS, 85(??):54–76, November 2014. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414001093.

Wang:2009:ASP

[WCXL09] Feng Wang, Jinru Chen, Wei Xu, and Zhilin Li. An additive Schwarz preconditioner for the mortar-type rotated Q_{1} FEM for elliptic problems with discontinuous coefficients. Applied Numerical Mathematics: Transactions of IMACS, 59(7):1657–1667, July 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wang:2022:AIW

[WD22] Bao-Shan Wang and Wai Sun Don. Affine-invariant WENO weights and operator. Applied Numerical Mathematics: Transactions of IMACS, 181(??):630–646, November 2022. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422001817.

Wubs:1992:EIM

[WdG92] F. W. Wubs and E. D. de Goede. An explicitimplicit method for a class of time-dependent partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 9(2):157–181, February 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wang:2020:CFS

[WDH20] Junjie Wang, Hongbin Dai, and Yuanxian Hui. Conservative Fourier spectral scheme for higher order Klein–Gordon–Schrödinger equations. Applied Numerical Mathematics: Transactions of IMACS, 156(??): 446–466, October 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420301598.

Wang:2023:FOW

[WDL23] Bao-Shan Wang, Wai Sun Don, and Peng Li. Fifthorder well-balanced positivitypreserving finite difference AWENO scheme with hydrostatic reconstruction for hyperbolic chemotaxis models. Applied Numerical Mathematics: Transactions of IMACS, 186(??): 41–56, April 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422003427.

Wang:2021:HOA

[WDU21] Xiaofeng Wang, Weizhong Dai, and Muhammad Usman. A high-order accurate finite difference scheme for the KdV equation with time-periodic boundary forcing. Applied Numerical Mathematics: Transactions of IMACS, 160(??):102– 121, February 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302968.

Wan:2021:SOT

[WDZS21] Jianjun Wan, Wen Dong, Yongcheng Zhao, and Shicang Song. Second-order two-scale analysis for heating effect of periodical composite structure. Applied Numerical Mathematics: Transactions of IMACS, 163(??): 204–220, May 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421000374.

Warnock:1999:SIP

[WE99] Robert L. Warnock and James A. Ellison. From symplectic integrator to Poincaré map: spline expansion of a map generator in Cartesian coordinates. Applied Numerical Mathematics: Transactions of IMACS, 29(1):89– 98, January 1, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

                                        tronic). URL http://www.
                                        elsevier.com/cas/tree/
                                        store/apnum/sub/1999/29/
                                        1/900.pdf.

Weiner:2012:EPM

[WEA12] Rüdiger Weiner and Tamer El-Azab. Exponential peer methods. Applied Numerical Mathematics: Transactions of IMACS, 62(10): 1335–1348, October 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412001055.

Weekes:2001:NCW

[Wee01] Suzanne L. Weekes. Numerical computation of wave propagation in dynamic materials. Applied Numerical Mathematics: Transactions of IMACS, 37(4): 417–440, June 2001. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/73/33/
27/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/73/33/27/article.
pdf.

Weiss:1995:TOK

[Wei95] Rüdiger Weiss. A theoretical overview of Krylov subspace methods. Applied Numerical Mathematics: Transactions of IMACS, 19(3):207– 233, December 15, 1995. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=19&
issue=3&aid=632;   http:
//www.sciencedirect.com/
science/article/pii/0168927495000844.

Special issue on iterative methods for linear equations (Atlanta, GA, 1994).

Weiser:2009:PNS

[Wei09]  Martin Weiser.  Point-

wise nonlinear scaling for reaction–diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 59(8):1858–1869, August 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Weise:2018:REE

[Wei18] Michael Weise. Residual error estimation for anisotropic Kirchhoff plates. Applied Numerical Mathematics: Transactions of IMACS, 125(??):10–22, March 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417302246.

Welfert:2010:NCG

[Wel10a] B. D. Welfert. A note on classical Gauss–Radau and Gauss–Lobatto quadratures. Applied Numerical Mathematics: Transactions of IMACS, 60(6):637–644, June 2010. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Welfert:2010:WCG

[Wel10b] B. D. Welfert. The weights of classical Gauss–Radau quadratures with double end-point. Applied Numerical Mathematics: Transactions of IMACS, 60(5): 574–586, May 2010. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wensch:1998:ESF

[Wen98]  J. Wensch.  An eight

stage fourth order partitioned Rosenbrock method for multibody systems in index-3 formulation. Applied Numerical Mathematics: Transactions of IMACS, 27(2):171–183, June 10, 1998. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1998/27/
2/873.pdf.

Wensch:2005:KRM

[Wen05] Jörg Wensch. Krylov-ROW methods for DAEs of index 1 with applications to viscoelasticity. Applied Numerical Mathematics: Transactions of IMACS, 53(2–4): 527–541, May 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Weniger:2010:ITP

[Wen10a] Ernst Joachim Weniger. An introduction to the topics presented at the conference “Approximation and extrapolation of convergent and divergent sequences and series”

CIRM Luminy: September 28, 2009–October 2, 2009. Applied Numerical Mathematics: Transactions of IMACS, 60(12): 1184–1187, December 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Weniger:2010:SDP

[Wen10b] Ernst Joachim Weniger. Summation of divergent power series by means of factorial series. Applied Numerical Mathematics: Transactions of IMACS, 60(12): 1429–1441, December 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wen:2025:MDG

[Wen25] Jing Wen. A mixed discontinuous Galerkin method for the Biot equations. Applied Numerical Mathematics: Transactions of IMACS, 212(??):283–299, June 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425000352.

Wang:2010:CST

[WG10] Xiaojie Wang and Siqing Gan. Compensated stochastic theta methods for stochastic differential equations with jumps. Applied Numerical Mathematics: Transactions of IMACS, 60(9):

877–887, September 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wu:2011:NPS

[WG11] Jiming Wu and Zhiming Gao. A nine-point scheme with explicit weights for diffusion equations on distorted meshes. Applied Numerical Mathematics: Transactions of IMACS, 61(7): 844–867, July 2011. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wu:2018:MBM

[WG18] Shi-Liang Wu and Peng Guo. Modulus-based matrix splitting algorithms for the quasi-complementarity problems. Applied Numerical Mathematics: Transactions of IMACS, 132(??): 127–137, October 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418301351.

Wang:2019:PEA

[WG19] Zhoufeng Wang and Rui Guo. A posteriori error analysis for the scattering by obstacles in a homogeneous chiral environment. Applied Numerical Mathematics: Transactions of IMACS, 135(??): 246–263, January 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418301934.

Wu:2022:SST

[WG22] Xiaojuan Wu and Siqing Gan. Split-step theta Milstein methods for SDEs with non-globally Lipschitz diffusion coefficients. Applied Numerical Mathematics: Transactions of IMACS, 180(??):16–32, October 2022. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422001283.

Wang:2023:ITO

[WG23a] Yahui Wang and Cheng Guo. Improved third-order WENO scheme with a new reference smoothness indicator. Applied Numerical Mathematics: Transactions of IMACS, 192(??): 454–472, October 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423001897.

Wu:2023:STS

[WG23b] Hua Wu and Qiyi Gao. A space-time spectral method for solving the nonlinear Klein–Gordon equation. Applied Numerical Mathematics: Transactions of IMACS, 190(??): 110–137, August 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001113.

Walther:1999:MVJ

[WGB99] Andrea Walther, Andreas Griewank, and André Best. Multiple vector–Jacobian products are cheap. Applied Numerical Mathematics: Transactions of IMACS, 30(2–3):367–377, June 10, 1999. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1999&volume=30&
issue=2-3&aid=985.

Wandelt:2012:SPR

[WGKS12] M. Wandelt, M. Günther, F. Knechtli, and M. Striebel. Symmetric partitioned Runge– Kutta methods for differential equations on Lie groups. Applied Numerical Mathematics: Transactions of IMACS, 62(12): 1740–1748, December 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412001171.

Wang:2015:MMN

[WGW15] Xiaojie Wang, Siqing Gan, and Desheng Wang. θ-Maruyama methods for nonlinear stochastic differential delay equations. Applied Numerical Mathematics: Transactions of IMACS, 98(??):38–58, December 2015. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927415001166.

Wang:2013:AVP

[WH13] Fei Wang and Weimin Han. Another view for a posteriori error estimates for variational inequalities of the second kind. Applied Numerical Mathematics: Transactions of IMACS, 72(??):225– 233, October 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927413000937.

Wang:2018:SPN

[WH18] Pengde Wang and Chengming Huang. Structurepreserving numerical methods for the fractional Schrödinger equation. Applied Numerical Mathematics: Transactions of IMACS, 129(??): 137–158, July 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418300709.

Wang:2019:CCN

[WH19a] Pengfei Wang and Pengzhan Huang. Convergence of the Crank–Nicolson extrapolation scheme for the Korteweg–de Vries equation. Applied Numerical Mathematics: Transactions of IMACS, 143(??):88– 96, September 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419300820.

Wang:2019:TGE

[WH19b] Wansheng Wang and Qingguo Hong. Two-grid economical algorithms for parabolic integro-differential equations with nonlinear memory. Applied Numerical Mathematics: Transactions of IMACS, 142(??):28–46, August 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419300388.

Wang:2023:GIP

[WH23] Qingsong Wang and Deren Han. A generalized inertial proximal alternating linearized minimization method for nonconvex nonsmooth problems. Applied Numerical Mathematics: Transactions of IMACS, 189(??): 66–87, July 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423000855.

Wen:2019:SAR

[WHL19] Jiao Wen, Chengming Huang, and Min Li. Stability analysis of Runge–Kutta methods for Volterra integrodifferential equations. Applied Numerical Mathematics: Transactions of IMACS, 146(??):73–88, December 2019. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892741930176X.

Wang:2021:TDE

[WHW21] Yifei Wang, Jin Huang, and Xiaoxia Wen. Twodimensional Euler polynomials solutions of twodimensional Volterra integral equations of fractional order. Applied Numerical Mathematics: Transactions of IMACS, 163(??): 77–95, May 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742100009X.

Winter:1992:IMS

[Win92] Dik T. Winter. Influence of memory systems on vector processor performance. Applied Numerical Mathematics: Transactions of IMACS, 10(1):59–72, June 1992. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Winkler:2001:CNN

[Win01] Joab R. Winkler. Condition numbers of a nearly singular simple root of a polynomial. Applied Numerical Mathematics: Transactions of IMACS, 38(3):275– 285, August 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/77/33/29/abstract.
html.

Winkler:2004:TCM

[Win04] Joab R. Winkler. The transformation of the companion matrix resultant between the power and Bernstein polynomial bases. Applied Numerical Mathematics: Transactions of IMACS, 48(1):113– 126, January 2004. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wang:2025:TBA

[WJF25] Yi-Fan Wang, Ji-Teng Jia, and Xin Fan. A tridiagonalizationbased arbitrary-stride reduction approach for (p, q)- pentadiagonal linear systems. Applied Numerical Mathematics: Transactions of IMACS, 213(??): 77–87, July 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425000479.

Wang:2022:TDD

[WJM22] Chen-Ye Wang, Yao-Lin Jiang, and Zhen Miao. Time domain decomposition of parabolic control problems based on discontinuous Galerkin semi-discretization. Applied Numerical Mathematics: Transactions of IMACS, 176(??):118–133, June 2022. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422000526.

Wang:2019:EAO

[WJW19] Yongxing Wang, Peter K. Jimack, and Mark A. Walkley. Energy analysis for the one-field fictitious domain method for fluid-structure interactions. Applied Numerical Mathematics: Transactions of IMACS, 140(??): 165–182, June 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419300406.

Wilhelm:2000:SUF

[WK00] Dirk Wilhelm and Leonhard Kleiser. Stable and unstable formulations of the convection operator in spectral element simulations. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4):275–280, May 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/61/27/
54/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/61/27/54/article.
pdf.

Wolke:2002:TIM

[WK02] Ralf Wolke and Oswald Knoth. Time-integration of multiphase chemistry in size-resolved cloud models. Applied Numerical Mathematics: Transactions of IMACS, 42(1–3):473–487, August 2002. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wang:2004:HOG

[WKM04] R. Wang, P. Keast, and P. Muir. A high-order global spatially adaptive collocation method for 1-D parabolic PDEs. Applied Numerical Mathematics: Transactions of IMACS, 50(2):239–260, August 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Watanabe:2020:SII

[WKN20] Yoshitaka Watanabe, Takehiko Kinoshita, and Mitsuhiro T. Nakao. Some improvements of invertibility verifications for second-order linear elliptic operators. Applied Numerical Mathematics: Transactions of IMACS, 154(??):36–46, August 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420300830.

Weiner:2012:VSD

[WKP12] R. Weiner, G. Yu. Kulikov, and H. Podhaisky. Variablestepsize doubly quasi-consistent parallel explicit peer methods with global error control. Applied Numerical Mathematics: Transactions of IMACS, 62(10): 1591–1603, October 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927412001043.

Wang:2026:PWE

[WKWC26] Jialing Wang, Anxin Kong, Tingchun Wang, and Wenjun Cai. Point-wise error estimates of two mass- and energy-preserving schemes for two-dimensional Schrödinger– Poisson equations. Applied Numerical Mathematics: Transactions of IMACS, 219(??):202–218, January 2026. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425001813.

Wang:2009:SSB

[WL09a] Peng Wang and Zhenxin Liu. Split-step backward balanced Milstein methods for stiff stochastic systems. Applied Numerical Mathematics: Transactions of IMACS, 59(6):1198–1213, June 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wang:2009:HOM

[WL09b] Yuan-Ming Wang and Xiao-Lin Lan. Higher-order monotone iterative methods for finite difference systems of nonlinear reaction– diffusion–convection equations. Applied Numerical Mathematics: Transactions of IMACS, 59(10): 2677–2693, October 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wang:2009:NMF

[WL09c] Zewen Wang and Jijun Liu.

New model function methods for determining regularization parameters in linear inverse problems. Applied Numerical Mathematics: Transactions of IMACS, 59(10):2489–2506, October 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wang:2010:NNM

[WL10] Jin Wang and Anita Layton. New numerical methods for Burgers’ equation based on semi-Lagrangian and modified equation approaches. Applied Numerical Mathematics: Transactions of IMACS, 60(6): 645–657, June 2010. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wang:2016:ISP

[WL16] Haibing Wang and Jijun Liu. An inverse scattering problem with generalized oblique derivative boundary condition. Applied Numerical Mathematics: Transactions of IMACS, 108(??): 226–241, October 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416300915.

Wang:2018:FDS

[WL18] Bo Wang and Dong Liang. The finite difference scheme for nonlinear Schrödinger equations on unbounded domain by artificial boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 128(??): 183–204, June 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418300503.

Wu:2021:REC

[WL21] Gang Wu and Fei Li.

A randomized exponential canonical correlation analysis method for data analysis and dimensionality reduction. Applied Numerical Mathematics: Transactions of IMACS, 164(??): 101–124, ???? 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302890.

Wang:2022:GFE

[WL22] Lingli Wang and Meng Li. Galerkin finite element method for damped nonlinear Schrödinger equation. Applied Numerical Mathematics: Transactions of IMACS, 178(??): 216–247, August 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422000861.

Wright:2024:NSE

[WL24] Sarah D. Wright and James V.

Lambers. Numerical solution of an extra-wide angle parabolic equation through diagonalization of a 1-D indefinite Schrödinger operator with a piecewise constant potential. Applied Numerical Mathematics: Transactions of IMACS, 199(??): 227–247, May 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423000594.

Wu:2025:NPD

[WL25] Renkai Wu and Zexian Liu. A new primal-dual hybrid gradient scheme for solving minimax problems with nonlinear term. Applied Numerical Mathematics: Transactions of IMACS, 210(??): 147–164, April 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424003519.

Wang:2026:HAS

[WLCZ26] Lan Wang, Yiyang Luo, Meng Chen, and Pengfei Zhu. A highly accurate symplectic-preserving scheme for Gross–Pitaevskii equation. Applied Numerical Mathematics: Transactions of IMACS, 219(??): 41–52, January 2026. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425001709.

Wang:2022:HOW

[WLG22] Bao-Shan Wang, Peng Li, and Zhen Gao. Highorder well-balanced and positivity-preserving finitedifference AWENO scheme with hydrostatic reconstruction for shallow water equations. Applied Numerical Mathematics: Transactions of IMACS, 181(??): 483–502, November 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001647.

Wang:2021:NHO

[WLM21] Zhibo Wang, Yuxiang Liang, and Yan Mo. A novel high order compact ADI scheme for two dimensional fractional integro-differential equations. Applied Numerical Mathematics: Transactions of IMACS, 167(??): 257–272, September 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001379.

Wang:2024:NLS

[WLY24] Lele Wang, Xin Liao, and Huaijun Yang. A new linearized second-order energy-stable finite element scheme for the nonlinear Benjamin–Bona–Mahony– Burgers equation. Applied Numerical Mathematics: Transactions of IMACS, 201(??):431–445, July 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000813.

Wang:2025:GES

[WLZ25] Dongmei Wang, Hengguang Li, and Qinghui Zhang. General enrichments of stable GFEM for interface problems: Theory and extreme learning machine construction. Applied Numerical Mathematics: Transactions of IMACS, 214(??): 143–159, August 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425000716.

Wang:2007:MRA

[WM07] L. X. Wang and Roderick V. N. Melnik. Model reduction applied to square to rectangular martensitic transformations using proper orthogonal decomposition. Applied Numerical Mathematics: Transactions of IMACS, 57(5–7):510–520, May 2007. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wang:2008:SPC

[WM08] Linxiang Wang and Roderick V. N. Melnik. Simulation of phase combinations in shape memory alloys patches by hybrid optimization methods. Applied Numerical Mathematics: Transactions of IMACS, 58(4):511–524, April 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wang:2022:CCT

[WM22] Shaoxin Wang and Lingsheng Meng. A contribution to the conditioning theory of the indefinite least squares problems. Applied Numerical Mathematics: Transactions of IMACS, 177(??): 137–159, July 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422000745.

Wu:2009:CIR

[WMC09] Fuke Wu, Xuerong Mao, and Kan Chen. The Cox– Ingersoll–Ross model with delay and strong convergence of its Euler–Maruyama approximate solutions. Applied Numerical Mathematics: Transactions of IMACS, 59(10):2641–2658, October 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wang:2017:EIR

[WMF17] Bin Wang, Fanwei Meng, and Yonglei Fang. Efficient implementation of RKN-type Fourier collocation methods for secondorder differential equations. Applied Numerical Mathematics: Transactions of IMACS, 119(??):164–178, September 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417301046.

Wang:2019:SSS

[WMLB19] Ying Wang, Liquan Mei, Qi Li, and Linlin Bu. Split-step spectral Galerkin method for the two-dimensional nonlinear space-fractional Schrödinger equation. Applied Numerical Mathematics: Transactions of IMACS, 136(??):257–278, February 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418302393.

Wang:2012:IEM

[WN12] Haibing Wang and Gen Nakamura. The integral equation method for electromagnetic scattering problem at oblique incidence. Applied Numerical Mathematics: Transactions of IMACS, 62(7):860–873, July 2012.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412000499.

Wang:2025:ANB

[WOLY25] Yijia Wang, Chen Ouyang, Liangfu Lv, and Gonglin Yuan. Analysis of a new BFGS algorithm and conjugate gradient algorithms and their applications in image restoration and machine learning. Applied Numerical Mathematics: Transactions of IMACS, 210(??): 199–221, April 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424003581.

Wong:2008:GFE

[Won08] J. C.-F. Wong. The Galerkin finite element method for the solution of some spatiotemporally dependent reaction diffusion systems. Applied Numerical Mathematics: Transactions of IMACS, 58(3):352–375, March 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wozny:2010:EAS

[Woź10] Paweł Woźny. Efficient algorithm for summation of some slowly convergent series. Applied Numerical Mathematics: Transactions of IMACS, 60(12):1442–1453, December 2010. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Weiss:1999:ONS

[WP99] Rüdiger Weiss and Isabella Podgajezki. Overview on new solvers for nonlinear systems. Applied Numerical Mathematics: Transactions of IMACS, 30(2– 3):379–391, June 10, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1999&volume=30&
                                        issue=2-3&aid=986.

Wang:2024:PSB

[WPAZ24] Yongshuai Wang, Zhenjiang Peng, Md. Abdullah Al Mahbub, and Haibiao Zheng. Partitioned schemes for the blood solute dynamics model by the variational multiscale method. Applied Numerical Mathematics: Transactions of IMACS, 198(??): 318–345, April 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424000060.

Wolf:2016:AIL

[WPL16] Thomas Wolf, Heiko K. F. Panzer, and Boris Lohmann.

ADI iteration for Lyapunov equations: a tangential approach and adaptive shift selection. Applied Numerical Mathematics: Transactions of IMACS, 109(??):85– 95, November 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301003.

Wang:2024:HAW

[WPN^{+}24] Jiaqun Wang, Guanxu Pan, Mengdie Niu, Youhe Zhou, and Xiaojing Liu. Highly accurate wavelet solution for the two-dimensional Bratu’s problem. Applied Numerical Mathematics: Transactions of IMACS, 203(??):52– 68, September 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001193.

Wu:2018:SBF

[WPS18] Gang Wu, Hong-Kui Pang, and Jiang-Li Sun. A shifted block FOM algorithm with deflated restarting for matrix exponential computations. Applied Numerical Mathematics: Transactions of IMACS, 127(??): 306–323, May 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418300242.

Wu:2019:ECA

[WPT19] S. Wu, B. Peng, and Z. F. Tian. Exponential compact ADI method for a coupled system of convectiondiffusion equations arising from the 2D unsteady magnetohydrodynamic (MHD) flows. Applied Numerical Mathematics: Transactions of IMACS, 146(??):89– 122, December 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419301758.

Wang:2017:NAT

[WQ17] Zewen Wang and Shufang Qiu. A numerical approximation of the two-dimensional elastic wave scattering problem via integral equation method. Applied Numerical Mathematics: Transactions of IMACS, 113(??): 156–167, March 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416302343.

Wang:2024:TRC

[WQS^{+}24] Wenli Wang, Gangrong Qu, Caiqin Song, Youran Ge, and Yuhan Liu. Tikhonov regularization with conjugate gradient least squares method for large-scale discrete ill-posed problem in image restoration. Applied Numerical Mathematics: Transactions of IMACS, 204(??):147–161, October 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001508.

Wang:2020:AHO

[WR20] Yuan-Ming Wang and Lei Ren. Analysis of a highorder compact finite difference method for Robin problems of time-fractional sub-diffusion equations with variable coefficients. Applied Numerical Mathematics: Transactions of IMACS, 156(??):467–492, October 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420301677.

Wu:2004:CDS

[WS04] Xiaonan Wu and Zhi-Zhong Sun. Convergence of difference scheme for heat equation in unbounded domains using artificial boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 50(2):261– 277, August 2004. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wang:2021:IAD

[WS21] Wenli Wang and Caiqin Song. Iterative algorithms for discrete-time periodic Sylvester matrix equations and its application in antilinear periodic system. Applied Numerical Mathematics: Transactions of IMACS, 168(??): 251–273, October 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421001744.

Wei:2022:QPP

[WS22] Siqi Wei and Raymond J. Spiteri. Qualitative property preservation of high-order operator splitting for the SIR model. Applied Numerical Mathematics: Transactions of IMACS, 172(??): 332–350, February 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421002786.

Wimmer:2009:RRR

[WSC09] Georg Wimmer, Thorsten Steinmetz, and Markus Clemens. Reuse, recycle, reduce (3R) – strategies for the calculation of transient magnetic fields. Applied Numerical Mathematics: Transactions of IMACS, 59(3–4): 830–844, March/April 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wu:2021:NFD

[WSC21] Tingting Wu, Yuran Sun, and Dongsheng Cheng. A new finite difference scheme for the 3D Helmholtz equation with a preconditioned iterative solver. Applied Numerical Mathematics: Transactions of IMACS, 161(??):348–371, March 2021. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303731.

Wen:2020:DGM

[WSHC20] Jing Wen, Jian Su, Yinnian He, and Hongbin Chen. Discontinuous Galerkin method for the nonlinear Biot’s model. Applied Numerical Mathematics: Transactions of IMACS, 151(??): 213–228, May 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419303708.

Shin:2015:HDG

[wSJP15] Dong wook Shin, Youngmok Jeon, and Eun-Jae Park. A hybrid discontinuous Galerkin method for advectiondiffusion-reaction problems. Applied Numerical Mathematics: Transactions of IMACS, 95(??):292–303, September 2015. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741400186X.

Weiner:1997:RRC

[WSP97] R. Weiner, B. A. Schmitt, and H. Podhaisky. ROWMAP— a ROW-code with Krylov techniques for large stiff ODEs. Applied Numerical Mathematics: Transactions of IMACS, 25(2–3): 303–319, October 17, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=25&
issue=2-3&aid=823.

Weiner:2004:PPT

[WSP04] R. Weiner, B. A. Schmitt, and H. Podhaisky. Parallel ‘peer’ two-step Wmethods and their application to MOL-systems. Applied Numerical Mathematics: Transactions of IMACS, 48(3–4):425–439, March 2004. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wouwer:1997:SUO

[WSS97] A. Vande Wouwer, P. Saucez, and W. E. Schiesser. Some user-oriented comparisons of adaptive grid methods for partial differential equations in one space dimension. Applied Numerical Mathematics: Transactions of IMACS, 26(1–2):49–62, December 22, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1997/26/
1-2/836.pdf.

Wensch:1996:CLI

[WSW96] J. Wensch, K. Strehmel, and R. Weiner. A class of linearly-implicit Runge– Kutta methods for multibody systems. Applied Numerical Mathematics: Transactions of IMACS, 22(1–3): 381–398, December 4, 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=22&
issue=1-3&aid=724. Special

issue celebrating the centenary of Runge–Kutta methods.

Wang:2018:SAN

[WSY18] Peizhen Wang, Ming Sun, and Changhui Yao. ℓ^{2} superconvergence analysis of nonconforming element approximation for 3D time-harmonic Maxwell’s equations. Applied Numerical Mathematics: Transactions of IMACS, 127(??):40–55, May 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417302659.

Wang:1993:IMR

[WT93] Ren Hong Wang and Jie Qing Tan. On interpolating multivariate rational splines. Applied Numerical Mathematics: Transactions of IMACS, 12(4):357–372, June 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wingate:2008:PNC

[WT08] Beth A. Wingate and Mark A. Taylor. Performance of numerically computed quadrature points. Applied Numerical Mathematics: Transactions of IMACS, 58(7):1030–1041, July 2008. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Westerkamp:2017:CII

[WT17] Armin Westerkamp and Manuel Torrilhon. Curvatureinduced instability of a Stokes-like problem with non-standard boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 121(??):96– 114, November 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417301484.

Wang:2020:EES

[WT20] Wansheng Wang and Jiao Tang. Efficient exponential splitting spectral methods for linear Schrödinger equation in the semiclassical regime. Applied Numerical Mathematics: Transactions of IMACS, 153(??): 132–146, July 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300416.

Walton:2024:DLA

[WTB24] Steven Walton, Minh-Binh Tran, and Alain Bensoussan. A deep learning approximation of non-stationary solutions to wave kinetic equations. Applied Numerical Mathematics: Transactions of IMACS, 199(??): 213–226, May 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422003324.

Wang:2025:TGH

[WTF25] Lu Wang, Youjun Tan, and Minfu Feng. The two-grid hybrid high-order method for the nonlinear strongly damped wave equation on polygonal mesh and its reduced-order model. Applied Numerical Mathematics: Transactions of IMACS, 210(??):1–24, April 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424003477.

Wang:2021:HVD

[WTY21] Lina Wang, Hongjiong Tian, and Lijun Yi. An hpversion of the discontinuous Galerkin time-stepping method for Volterra integral equations with weakly singular kernels. Applied Numerical Mathematics: Transactions of IMACS, 161(??): 218–232, March 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303561.

Wu:2003:ZPV

[Wu03] Xinyuan Wu. On zeros of polynomial and vector solutions of associated polynomial system from Vieta theorem. Applied Numerical Mathematics: Transactions of IMACS, 44(3):415– 423, February 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wu:2009:NSS

[Wu09] Chunhong Wu. Numerical solution for Stokes’ first problem for a heated generalized second grade fluid with fractional derivative. Applied Numerical Mathematics: Transactions of IMACS, 59(10):2571–2583, October 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wojcik:1988:LSE

[WVBM88] G. L. Wojcik, D. K.

Vaughan, M. Barenberg, and J. Mould. Large-scale, explicit wave simulations on the Cray-2. Applied Numerical Mathematics: Transactions of IMACS, 4(1): 47–70, March 1988. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wagner:1999:MPS

[WW99] Boris Wagner and Rüdiger Weiss. Minimization properties and short recurrences. Applied Numerical Mathematics: Transactions of IMACS, 30(2–3): 175–190, June 10, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/

cas_sub/browse/browse.
cgi?year=1999&volume=30&
issue=2-3&aid=974.

Wang:2005:PIP

[WW05] Ren-Hong Wang and Jing-Xin Wang. The poisedness of interpolation problem for splines. Applied Numerical Mathematics: Transactions of IMACS, 54(1): 95–103, June 2005. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wei:2014:MQB

[WW14] Ting Wei and Jungang Wang. A modified quasiboundary value method for an inverse source problem of the time-fractional diffusion equation. Applied Numerical Mathematics: Transactions of IMACS, 78(??): 95–111, April 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413001724.

Wang:2019:SAN

[WW19] Bin Wang and Xinyuan Wu. A symplectic approximation with nonlinear stability and convergence analysis for efficiently solving semilinear Klein–Gordon equations. Applied Numerical Mathematics: Transactions of IMACS, 142(??): 64–89, August 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419300467.

Wang:2024:SAE

[WW24] Guangyan Wang and Tongke Wang. Singular asymptotic expansion and Legendre collocation method for two-term weakly singular Volterra integral equation of the second kind. Applied Numerical Mathematics: Transactions of IMACS, 197(??):344–362, March 2024. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423003021.

Wang:2025:NAM

[WWC^{+}25] Zhaowei Wang, Danxia Wang, Yanping Chen, Chenhui Zhang, and Hongen Jia. Numerical approximation for the MHD equations with variable density based on the Gauge–Uzawa method. Applied Numerical Mathematics: Transactions of IMACS, 207(??): 272–302, January 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424002435.

Wu:2020:NFT

[WWF20] Jilian Wu, Leilei Wei, and Xinlong Feng. Novel fractional time-stepping algorithms for natural convection problems with variable density. Applied Numerical Mathematics: Transactions of IMACS, 151(??): 64–84, May 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419303447.

Wu:2024:VEM

[WWH24] Bangmin Wu, Fei Wang, and Weimin Han. The virtual element method for a contact problem with wear and unilateral constraint. Applied Numerical Mathematics: Transactions of IMACS, 206(??):29– 47, December 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424002009.

Wang:2021:LTS

[WWL21] Wansheng Wang, Zheng Wang, and Zhaoxiang Li. Long time H_{α}^{s} stability of a classical scheme for Cahn– Hilliard equation with polynomial nonlinearity. Applied Numerical Mathematics: Transactions of IMACS, 165(??):35–55, July 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000477.

Wang:2023:LDG

[WWLL23] Nian Wang, Jinfeng Wang, Yang Liu, and Hong Li. Local discontinuous Galerkin method for a nonlocal viscous water wave model. Applied Numerical Mathematics: Transactions of IMACS, 192(??):431–453, October 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001903.

Wu:2008:TTA

[WWLS08] Jiming Wu, Yingxi Wang, Wen Li, and Weiwei Sun. Toeplitz-type approximations to the Hadamard integral operator and their applications to electromagnetic cavity problems. Applied Numerical Mathematics: Transactions of IMACS, 58(2):101–121, February 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wang:2022:LIV

[WWM22] Wansheng Wang, Zheng Wang, and Mengli Mao. Linearly implicit variable step-size BDF schemes with Fourier pseudospectral approximation for incompressible Navier–Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 172(??): 393–412, February 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421003044.

Wendroff:1993:BPC

[WWS^{+}93] Burton Wendroff, Tony Warnock, Lewis Stiller, Dean Mayer, and Ralph Brickner. Bits and pieces: constructing chess endgame databases on parallel and vector architectures. Applied Numerical Mathematics: Transactions of IMACS, 12(1–3): 285–295, May 1993. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0168927493901239.

Wu:2007:PAI

[WWS07] Meijun Wu, Li Wang, and Yongzhong Song. Preconditioned AOR iterative method for linear systems. Applied Numerical Mathematics: Transactions of IMACS, 57(5–7):672–685, May 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wu:2025:EET

[WWWW25] Peixuan Wu, Xiaohui Wu, Yang Wang, and Ruqing Wang. Error estimates of a two-grid BDF2 virtual element scheme for semilinear parabolic equation. Applied Numerical Mathematics: Transactions of IMACS, 217(??): 451–475, November 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425001424.

Wang:2013:EBE

[WWX13] Bin Wang, Xinyuan Wu, and Jianlin Xia. Error bounds for explicit ERKN integrators for systems of multifrequency oscillatory secondorder differential equations. Applied Numerical Mathematics: Transactions of IMACS, 74(??):17–34, December 2013. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413000962.

Wang:2025:LDU

[WWZ25] Qi Wang, Kun Wang, and Guanyu Zhou. A lineardecoupled and unconditionally energy stable fully discrete scheme for Peterlin viscoelastic model. Applied Numerical Mathematics: Transactions of IMACS, 217(??):412–435, November 2025. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425001436.

Wang:2022:USS

[WWZJ22] Danxia Wang, Xingxing Wang, Ran Zhang, and Hongen Jia. An unconditionally stable second-order linear scheme for the Cahn– Hilliard–Hele–Shaw system. Applied Numerical Mathematics: Transactions of IMACS, 171(??):58–75, January 2022. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421002403.

Wang:2025:ESE

[WWZS25] Haifeng Wang, Yan Wang, Hong Zhang, and Songhe Song. Energy stability and error estimate of the RKMK2e scheme for the extended Fisher–Kolmogorov equation. Applied Numerical Mathematics: Transactions of IMACS, 212(??): 60–76, June 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425000145.

Wu:2006:ERK

[WX06] Xinyuan Wu and Jianlin Xia. Extended Runge–Kutta-like formulae. Applied Numerical Mathematics: Transactions of IMACS, 56(12):

1584–1605, December 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wu:2022:GRE

[WX22] Nianci Wu and Hua Xiang.

On the generally randomized extended Gauss–Seidel method. Applied Numerical Mathematics: Transactions of IMACS, 172(??): 382–392, February 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421003032.

Wang:2024:LTB

[WXY24] Xingchang Wang, Runzhang Xu, and Yanbing Yang. Long-time behavior for fourth order nonlinear wave equations with dissipative and dispersive terms. Applied Numerical Mathematics: Transactions of IMACS, 199(??): 248–265, May 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423000107.

Wang:2002:SAN

[WY02] Xiaoshen Wang and Xiu Ye. Superconvergence analysis for the Navier–Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 41(4):

515–527, July 2002. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wang:2022:DSE

[WY22] Shuai Wang and Guangwei Yuan. Discrete strong extremum principles for finite element solutions of diffusion problems with nonlinear corrections. Applied Numerical Mathematics: Transactions of IMACS, 174(??): 1–16, April 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422000058.

Wen:2011:DRK

[WYL11] Liping Wen, Yuexin Yu, and Shoufu Li. Dissipativity of Runge–Kutta methods for Volterra functional differential equations. Applied Numerical Mathematics: Transactions of IMACS, 61(3):368–381, March 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wu:2012:MFP

[WYP12] Ting-Ting Wu, Yu-Fei Yang, and Zhi-Feng Pang. A modified fixed-point iterative algorithm for image restoration using fourth-order PDE model. Applied Numerical Mathematics: Transactions of IMACS, 62(2):79– 90, February 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411001851.

Wu:2020:ALS

[WYY20] Xiaolei Wu, Yuyuan Yan, and Yubin Yan. An analysis of the L1 scheme for stochastic subdiffusion problem driven by integrated space–time white noise. Applied Numerical Mathematics: Transactions of IMACS, 157(??):69– 87, November 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420301586.

Wu:2019:FVE

[WYYL19] Dan Wu, Jingyan Yue, Guangwei Yuan, and Junliang Lv. Finite volume element approximation for nonlinear diffusion problems with degenerate diffusion coefficients. Applied Numerical Mathematics: Transactions of IMACS, 140(??): 23–47, June 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419300248.

Wang:2002:MTR

[WZ02] Kai Wang and Jun Zhang.

Multigrid treatment and robustness enhancement for factored sparse approximate inverse preconditioning. Applied Numerical Mathematics: Transactions of IMACS, 43(4):483–500, December 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wang:2014:ACN

[WZ14] Yong Wang and Jian-Xun Zhao. An algorithm for a class of nonlinear complementarity problems with non-Lipschitzian functions. Applied Numerical Mathematics: Transactions of IMACS, 82(??):68–79, August 2014. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414000439.

Wang:2016:ING

[WZ16] Jueyu Wang and Detong Zhu. The inexact-Newton via GMRES subspace method without line search technique for solving symmetric nonlinear equations. Applied Numerical Mathematics: Transactions of IMACS, 110(??):174–189, December 2016. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301635.

Wu:2017:UCM

[WZ17] Jinbiao Wu and Hui Zheng.

Uniform convergence of multigrid methods for adaptive meshes. Applied Numerical Mathematics: Transactions of IMACS, 113(??): 109–123, March 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416302306.

Wang:2019:SIF

[WZ19] Quanxiang Wang and Zhiyue Zhang. A stabilized immersed finite volume element method for elliptic interface problems. Applied Numerical Mathematics: Transactions of IMACS, 143(??):75– 87, September 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419300637.

Wang:2022:CNT

[WZ22] Yuan-Ming Wang and Yu-Jia Zhang. A Crank– Nicolson-type compact difference method with the uniform time step for a class of weakly singular parabolic integro-differential equations. Applied Numerical Mathematics: Transactions of IMACS, 172(??):

566–590, February 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421003111.

Wang:2008:NSO

[WZL08] Wan-Sheng Wang, Yuan Zhang, and Shou-Fu Li. Nonlinear stability of oneleg methods for delay differential equations of neutral type. Applied Numerical Mathematics: Transactions of IMACS, 58(2):122– 130, February 2008. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Wang:2013:FFV

[WZL13] Quanxiang Wang, Zhiyue Zhang, and Zhilin Li. A Fourier finite volume element method for solving two-dimensional quasigeostrophic equations on a sphere. Applied Numerical Mathematics: Transactions of IMACS, 71(??):1– 13, September 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927413000536.

Wang:2013:TRM

[WZW13] Jun-Gang Wang, Yu-Bin Zhou, and Ting Wei. Two regularization methods to identify a space-dependent source for the time-fractional diffusion equation. Applied Numerical Mathematics: Transactions of IMACS, 68(??):39–57, June 2013. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413000032.

Wang:2021:EEI

[WZZ21] Fangyuan Wang, Xiangcheng Zheng, and Zhaojie Zhou. Error estimate for indirect spectral approximation of optimal control problem governed by fractional diffusion equation with variable diffusivity coefficient. Applied Numerical Mathematics: Transactions of IMACS, 170(??): 146–161, December 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421002129.

Xiao:2014:PRB

[XB14] Jianping Xiao and John P.

Boyd. Periodized radial basis functions, part I: Theory. Applied Numerical Mathematics: Transactions of IMACS, 86(??):43– 73, December 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414001305.

Xie:1985:MPS

[XC85] G. Q. Xie and Y. M. Chen. A modified pulse-spectrum technique for solving inverse problems of two-dimensional elastic wave equation. Applied Numerical Mathematics: Transactions of IMACS, 1(3):217–237, May 1985. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Xie:2020:IMM

[XC20] Huiqing Xie and Meng Chen. An incomplete modal method for eigenvector derivatives of polynomial eigenvalue problems. Applied Numerical Mathematics: Transactions of IMACS, 156(??): 322–345, October 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420301483.

Xu:2022:EIP

[XCHW22] Zhuangzhi Xu, Wenjun Cai, Dongdong Hu, and Yushun Wang. Exponential integrator preserving mass boundedness and energy conservation for nonlinear Schrödinger equation. Applied Numerical Mathematics: Transactions of IMACS, 173(??): 308–328, March 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742100341X.

Xin:2006:VSD

[XF06] Jianguo Xin and Joseph E.

Flaherty. Viscous stabilization of discontinuous Galerkin solutions of hyperbolic conservation laws. Applied Numerical Mathematics: Transactions of IMACS, 56(3–4):444–458, March/April 2006. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Xie:2022:ENM

[XF22] Changping Xie and Shaomei Fang. Efficient numerical methods for Riesz spacefractional diffusion equations with fractional Neumann boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 176(??):1–18, June 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422000083.

Xu:2019:FCB

[XFG19] Zhenhua Xu, Chunhua Fang, and Hongrui Geng. Fast computation of Bessel transform with highly oscillatory integrands. Applied Numerical Mathematics: Transactions of IMACS, 145(??): 121–132, November 2019.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892741930159X.

Xie:2022:HHO

[XFL22] Chun-Mei Xie, Min-Fu Feng, and Yan Luo. A hybrid highorder method for the Sobolev equation. Applied Numerical Mathematics: Transactions of IMACS, 178(??): 84–97, August 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742200068X.

Xiao:2000:RPG

[XFLC00] Aiguo Xiao, Hongyuan Fu, Shoufu Li, and Guangnan Chen. Regularity properties of general linear methods for initial value problems of ordinary differential equations. Applied Numerical Mathematics: Transactions of IMACS, 34(4):405– 420, August 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/10/28/62/32/
                                        32/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/10/28/62/32/32/article.
                                        pdf.

Xie:2022:DLI

[XG22] Zuodi Xie and Tieqiang Gang. Discrete LQR and ILQR methods based on high order Runge–Kutta discretizations. Applied Numerical Mathematics: Transactions of IMACS, 182(??): 196–213, December 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001891.

Xu:2022:EMM

[XGHM22] Fei Xu, Yasai Guo, Qiumei Huang, and Hongkun Ma. An efficient multigrid method for semilinear interface problems. Applied Numerical Mathematics: Transactions of IMACS, 179(??):238–254, September 2022. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001246.

Xiang:2008:NQB

[XGM08] Shuhuang Xiang, Weihua Gui, and Pinghua Mo. Numerical quadrature for Bessel transformations. Applied Numerical Mathematics: Transactions of IMACS, 58(9):1247–1261, September 2008. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Xu:2020:TTG

[XGQ20] Da Xu, Jing Guo, and Wenlin Qiu. Time twogrid algorithm based on finite difference method for two-dimensional nonlinear fractional evolution equations. Applied Numerical Mathematics: Transactions of IMACS, 152(??): 169–184, June 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419303435.

Xu:2021:LPM

[XHJM21] Fei Xu, Qiumei Huang, Kun Jiang, and Hongkun Ma. Local and parallel multigrid method for semilinear elliptic equations. Applied Numerical Mathematics: Transactions of IMACS, 162(??):20–34, April 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420303755.

Xu:2022:MCG

[XHYM22] Fei Xu, Qiumei Huang, Huiting Yang, and Hongkun Ma. Multilevel correction goal-oriented adaptive finite element method for semilinear elliptic equations. Applied Numerical Mathematics: Transactions of IMACS, 172(??): 224–241, February 2022.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421002762.

Xiu:2008:NIF

[Xiu08] Dongbin Xiu. Numerical integration formulas of degree two. Applied Numerical Mathematics: Transactions of IMACS, 58(10): 1515–1520, October 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Xiao:2025:USS

[XK25] Jingjing Xiao and Desong Kong. An unconditionally stable second-order scheme for Maxwell’s equations in the Cole–Cole dispersive medium. Applied Numerical Mathematics: Transactions of IMACS, 211(??): 211–227, May 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425000091.

Xie:2009:FNS

[XL09a] Wen-Jing Xie and Fu-Rong Lin. A fast numerical solution method for two dimensional Fredholm integral equations of the second kind. Applied Numerical Mathematics: Transactions of IMACS, 59(7): 1709–1719, July 2009. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Xu:2009:NCF

[XL09b] Xianmin Xu and Zhiping Li. Non-conforming finite element and artificial boundary in multi-atomic Young measure approximation for micromagnetics. Applied Numerical Mathematics: Transactions of IMACS, 59(5):920–937, May 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Xiong:2011:PEE

[XL11] Chunguang Xiong and Yuan Li. A posteriori error estimates for optimal distributed control governed by the evolution equations. Applied Numerical Mathematics: Transactions of IMACS, 61(2):181– 200, February 2011. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Xie:2023:FBG

[XL23] Jiangming Xie and Maojun Li. A fast BDF2 Galerkin finite element method for the one-dimensional timedependent Schrödinger equation with artificial boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 187(??): 89–106, May 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423000375.

Xu:2025:SNM

[XLHL25] Lei Xu, Li-Bin Liu, Zaitang Huang, and Guangqing Long. Supercloseness of the NIPG method on a Bakhvalov-type mesh for a singularly perturbed problem with two small parameters. Applied Numerical Mathematics: Transactions of IMACS, 207(??): 431–449, January 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424002538.

Xie:2007:SCF

[XLK07] W. F. Xie, T. G. Liu, and B. C. Khoo. The simulation of cavitating flows induced by underwater shock and free surface interaction. Applied Numerical Mathematics: Transactions of IMACS, 57(5–7):734–745, May 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Xu:2019:FOW

[XLKY19] Runzhang Xu, Wei Lian, Xiangkun Kong, and Yanbing Yang. Fourth order wave equation with nonlinear strain and logarithmic nonlinearity. Applied Numerical Mathematics: Transactions of IMACS, 141(??): 185–205, ???? 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418301466.

Xie:2020:TNE

[XLZ20] Jianqiang Xie, Dong Liang, and Zhiyue Zhang. Two novel energy dissipative difference schemes for the strongly coupled nonlinear space fractional wave equations with damping. Applied Numerical Mathematics: Transactions of IMACS, 157(??):178–209, November 2020. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742030180X.

Xu:2023:AMS

[XLZ23] Minqiang Xu, Kai Liu, and Lei Zhang. Analysis of the MAC scheme for the three dimensional Stokes problem. Applied Numerical Mathematics: Transactions of IMACS, 193(??): 131–147, November 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423002040.

Xu:2023:USA

[XP23] Chao Xu and Lifang Pei.

Unconditional superconvergence analysis of two modified finite element fully discrete schemes for nonlinear Burgers’ equation. Applied Numerical Mathematics: Transactions of IMACS, 185(??):1–17, March 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422002963.

Xu:2013:CAS

[Xu13] Guoliang Xu. Consistent approximations of several geometric differential operators and their convergence. Applied Numerical Mathematics: Transactions of IMACS, 69(??):1–12, July 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413000184.

Xu:2016:CPF

[Xu16] Kuan Xu. The Chebyshev points of the first kind. Applied Numerical Mathematics: Transactions of IMACS, 102(??):17–30, April 2016. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416000039.

Xu:2021:PEE

[Xu21] Shipeng Xu. A posteriori error estimates for weak Galerkin methods for second order elliptic problems on polygonal meshes. Applied Numerical Mathematics: Transactions of IMACS, 161(??):510–524, March 2021. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303780.

Xu:2023:SML

[Xu23] Yuesheng Xu. Sparse machine learning in Banach spaces. Applied Numerical Mathematics: Transactions of IMACS, 187(??): 138–157, May 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423000429.

Xiao:2019:SSS

[XW19] Aiguo Xiao and Junjie Wang. Symplectic scheme for the Schrödinger equation with fractional Laplacian. Applied Numerical Mathematics: Transactions of IMACS, 146(??): 469–487, December 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927419302041.

Xiong:2025:FAC

[XWC25] Wenzhuo Xiong, Xiao Wang, and Xiujun Cheng. Fast algorithms of compact scheme for solving parabolic equations and their application. Applied Numerical Mathematics: Transactions of IMACS, 215(??): 90–111, September 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425000832.

Xu:2026:SLD

[XWM26] Yan Xu, Boying Wu, and Xiong Meng. Superconvergence of the local discontinuous Galerkin method with generalized numerical fluxes for onedimensional linearized KdV equations. Applied Numerical Mathematics: Transactions of IMACS, 219(??):99– 121, January 2026. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425001758.

Xiao:2019:CLI

[XWW19] Aiguo Xiao, Chenxi Wang, and Junjie Wang. Conservative linearly-implicit difference scheme for a class of modified Zakharov systems with high-order space fractional quantum correction. Applied Numerical Mathematics: Transactions of IMACS, 146(??): 379–399, December 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927419301916.

Xing:2021:FOC

[XWX21] Zhiyong Xing, Liping Wen, and Hanyu Xiao. A fourth-order conservative difference scheme for the Riesz space-fractional Sine– Gordon equations and its fast implementation. Applied Numerical Mathematics: Transactions of IMACS, 159(??):221–238, January 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302853.

Xie:2021:LIF

[XWZ21] Jianqiang Xie, Quanxiang Wang, and Zhiyue Zhang. Linear implicit finite difference methods with energy conservation property for space fractional Klein– Gordon–Zakharov system. Applied Numerical Mathematics: Transactions of IMACS, 167(??):389–419, September 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001458.

Xu:2022:TGB

[XXF22] Jie Xu, Shusen Xie, and Hongfei Fu. A two-grid block-centered finite difference method for the nonlinear regularized long wave equation. Applied Numerical Mathematics: Transactions of IMACS, 171(??): 128–148, January 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421002361.

Xiong:2017:MIR

[XXQ17] Xiangtuan Xiong, Xuemin Xue, and Zhi Qian. A modified iterative regularization method for ill-posed problems. Applied Numerical Mathematics: Transactions of IMACS, 122(??): 108–128, December 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417301770.

Xia:2024:UES

[XXYZ24] Binhu Xia, Xiaojian Xi, Rongrong Yu, and Peijun Zhang. Unconditional energy-stable method for the Swift–Hohenberg equation over arbitrarily curved surfaces with second-order accuracy. Applied Numerical Mathematics: Transactions of IMACS, 198(??): 192–201, April 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000059.

Xie:2019:VCP

[XY19] Jinghua Xie and Lijun Yi. An h–p version of the continuous Petrov–Galerkin time stepping method for nonlinear second-order delay differential equations. Applied Numerical Mathematics: Transactions of IMACS, 143(??):1–19, September 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0168927419300686.

Xiao:2024:RHR

[XY24] Chuanfu Xiao and Chao Yang. RA-HOOI: Rankadaptive higher-order orthogonal iteration for the fixed-accuracy low multilinearrank approximation of tensors. Applied Numerical Mathematics: Transactions of IMACS, 201(??): 290–300, July 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000539.

Xia:2025:ETG

[XY25] Yande Xia and Yun-Bo Yang. An efficient two-grid algorithm based on Newton iteration for the stationary inductionless magnetohydrodynamic system. Applied Numerical Mathematics: Transactions of IMACS, 212(??):312–332, June 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425000327.

Xu:2020:AEP

[XYHM20] Fei Xu, Meiling Yue, Qiumei Huang, and Hongkun Ma. An asymptotically exact a posteriori error estimator for non-selfadjoint Steklov eigenvalue problem. Applied Numerical Mathematics: Transactions of IMACS, 156(??): 210–227, October 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742030146X.

Xu:2024:MBH

[XYZ24] Xuejun Xu, Xiu Ye, and Shangyou Zhang. A macro BDM H-div mixed finite element on polygonal and polyhedral meshes. Applied Numerical Mathematics: Transactions of IMACS, 206(??):283–297, December 2024. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424002095.

Xie:2019:CSO

[XZ19] Ying Xie and Chengjian Zhang. A class of stochastic one-parameter methods for nonlinear SFDEs with piecewise continuous arguments. Applied Numerical Mathematics: Transactions of IMACS, 135(??): 1–14, January 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418301752.

Xie:2022:SWM

[XZ22] Shenglan Xie and Peng Zhu. Superconvergence of a WG method for the Stokes equations with continuous pressure. Applied Numerical Mathematics: Transactions of IMACS, 179(??):27– 38, September 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422001052.

Xu:2019:NKB

[XZH19] Liwei Xu, Shangyou Zhang, and George C. Hsiao. Nonsingular kernel boundary integral and finite element coupling method. Applied Numerical Mathematics: Transactions of IMACS, 137(??):80–90, March 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418302617.

Xu:2007:MAC

[XZL07] Zhenli Xu, Peng Zhang, and Ruxun Liu. δ-mapping algorithm coupled with WENO reconstruction for nonlinear elasticity in heterogeneous media. Applied Numerical Mathematics: Transactions of IMACS, 57(1):103– 116, January 2007. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Xie:2019:CSD

[XZL19] Jianqiang Xie, Zhiyue Zhang, and Dong Liang. A conservative splitting difference scheme for the fractionalin-space Boussinesq equation. Applied Numerical Mathematics: Transactions of IMACS, 143(??):61– 74, September 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419300662.

Xu:2021:FOL

[XZT21] Minqiang Xu, Lufang Zhang, and Emran Tohidi. A fourthorder least-squares based reproducing kernel method for one-dimensional elliptic interface problems. Applied Numerical Mathematics: Transactions of IMACS, 162(??):124–136, April 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303883.

Xie:2019:EAW

[XZW19] Shenglan Xie, Peng Zhu, and Xiaoshen Wang. Error analysis of weak Galerkin finite element methods for time-dependent convection– diffusion equations. Applied Numerical Mathematics: Transactions of IMACS, 137(??):19–33, March 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418302691.

Xu:2019:EAL

[XZZ19] Yang Xu, Yanming Zhang, and Jingjun Zhao. Error analysis of the Legendre– Gauss collocation methods for the nonlinear distributedorder fractional differential equation. Applied Numerical Mathematics: Transactions of IMACS, 142(??): 122–138, August 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419300583.

Xu:2015:AUF

[XZZL15] Yinghong Xu, Lipu Zhang, Jing Zhang, and Hua Luo. An adaptive updating full-Newton step interior-point algorithm with modified Newton direction. Applied Numerical Mathematics: Transactions of IMACS, 91(??):98–106, May 2015. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415000124.

Yamada:2018:WMS

[Yam18] Toshihiro Yamada. Weak Milstein scheme without commutativity condition and its error bound. Applied Numerical Mathematics: Transactions of IMACS, 131(??):95–108, September 2018. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418300965.

Yamada:2023:NAC

[Yam23] Toshihiro Yamada. A new algorithm for computing path integrals and weak approximation of SDEs inspired by large deviations and Malliavin calculus. Applied Numerical Mathematics: Transactions of IMACS, 187(??):192–205, May 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423000430.

Yang:2018:MCC

[Yan18] Changqing Yang. Modified Chebyshev collocation method for pantograph-type differential equations. Applied Numerical Mathematics: Transactions of IMACS, 134(??):132–144, December 2018. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418301594.

Yang:2021:SEE

[Yan21a] Huaijun Yang. Superconvergence error estimate of Galerkin method for Sobolev equation with Burgers’ type nonlinearity. Applied Numerical Mathematics: Transactions of IMACS, 168(??): 13–22, October 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421001471.

Yang:2021:GPR

[Yan21b] Xi Yang. A geometric probability randomized Kaczmarz method for large scale linear systems. Applied Numerical Mathematics: Transactions of IMACS, 164(??): 139–160, ???? 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303238.

Yang:2022:NAS

[Yan22] Huaijun Yang. A novel approach of superconvergence analysis of the bilinearconstant scheme for timedependent Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 173(??): 180–192, March 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421003378.

Yang:2023:UOE

[Yan23] Huaijun Yang. Unconditionally optimal error estimate of mass- and energystable Galerkin method for Schrödinger equation with cubic nonlinearity. Applied Numerical Mathematics: Transactions of IMACS, 183(??):39–55, January 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422002215.

Yao:2024:PFM

[Yao24] Hui Yao. A phase field method for convective phase change problem preserving maximum bound principle. Applied Numerical Mathematics: Transactions of IMACS, 204(??): 232–248, October 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424001582.

Yang:2010:LSB

[YB10] Yidu Yang and Hai Bi. Lower spectral bounds by Wilson’s brick discretization. Applied Numerical Mathematics: Transactions of IMACS, 60(8):782–787, August 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Yang:2013:ENE

[YBL13] Yidu Yang, Hai Bi, and Sirui Li. The extrapolation of numerical eigenvalues by finite elements for differential operators. Applied Numerical Mathematics: Transactions of IMACS, 69(??):59–72, July 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927413000408.

Yilmaz:2020:SOC

[YBW20] Fikriye Yilmaz, Hacer Öz Bakan, and Gerhard-Wilhelm Weber. Strong-order conditions of Runge–Kutta method for stochastic optimal control problems. Applied Numerical Mathematics: Transactions of IMACS, 157(??):470–489, November 2020. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302026.

Yang:2000:SGG

[YC00] J. C. Yang and T. C. Chen.

The structured grid generation using Beta-spline curve/surface modeling. Applied Numerical Mathematics: Transactions of IMACS, 32(4):435–445, April 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/59/32/
33/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/59/32/33/article.
pdf.

You:2013:DIR

[YC13] Xiong You and Zhaoxia Chen. Direct integrators of Runge–Kutta type for special third-order ordinary differential equations. Applied Numerical Mathematics: Transactions of IMACS, 74(??): 128–150, December 2013. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927413001177.

Yilmaz:2016:PBV

[YÇ16] Fikriye Yilmaz and Aytekin Çibik. A projectionbased variational multiscale method for the optimal control problems governed by the stationary Navier–Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 106(??): 116–128, August 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416300356.

Yi:2023:TGI

[YCWH23] Huaming Yi, Yanping Chen, Yang Wang, and Yunqing Huang. A twogrid immersed finite element method with the Crank–Nicolson time scheme for semilinear parabolic interface problems. Applied Numerical Mathematics: Transactions of IMACS, 189(??):1–22, July 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423000818.

Yang:2012:HMM

[YCY12] Fenlin Yang, Ke Chen, and Bo Yu. Homotopy method for a mean curvature-based denoising model. Applied Numerical Mathematics: Transactions of IMACS, 62(3):185–200, March 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411002091.

Yuan:2007:IPS

[YD07] Yongxin Yuan and Hua Dai.

Inverse problems for symmetric matrices with a submatrix constraint. Applied Numerical Mathematics: Transactions of IMACS, 57(5–7):646–656, May 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Yang:2022:ISP

[YD22] Dandan Yang and Zhiwen Duan. The inverse source problem of Cherenkov radiation model. Applied Numerical Mathematics: Transactions of IMACS, 182(??):42– 56, December 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001969.

Duan:2009:ALS

[yDqGnJT09] Huo yuan Duan, Shao qin Gao, Bo nan Jiang, and Roger C. E. Tan. Analysis of a least-squares finite element method for the thin plate problem. Applied Numerical Mathematics: Transactions of IMACS, 59(5):976–987, May 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Yu:2017:TOD

[YDWW17] Yanyan Yu, Weihua Deng, Yujiang Wu, and Jing Wu. Third order difference schemes (without using points outside of the domain) for one sided space tempered fractional partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 112(??): 126–145, February 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416302124.

Ye:2004:NBL

[Ye04] Mao Ye. Numerical boundary layers of conservation laws with relaxation extension. Applied Numerical Mathematics: Transactions of IMACS, 51(2–3): 385–405, November 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Younes:2025:TWS

[YEZ25] G. Abi Younes, N. El Khatib, and M. Zaydan. Traveling wave solutions for the accelerated Frenkel–Kontorova model: the monostable cases. Applied Numerical Mathematics: Transactions of IMACS, 215(??):25– 48, September 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425000820.

Yin:2024:EHA

[YF24] Fengli Yin and Yayun Fu. Explicit high accuracy energy-preserving Lie-group sine pseudo-spectral methods for the coupled nonlinear Schrödinger equation. Applied Numerical Mathematics: Transactions of IMACS, 195(??):1–16, January 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423002404.

Yao:2020:CGC

[YFLX20] Shengwei Yao, Qinliang Feng, Lue Li, and Jieqiong Xu. A class of globally convergent three-term Dai-Liao conjugate gradient methods. Applied Numerical Mathematics: Transactions of IMACS, 151(??): 354–366, May 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300015.

Yang:1995:NFP

[YG95] Ulrike Meier Yang and Kyle A. Gallivan. A new family of preconditioned iterative solvers for nonsymmetric linear systems. Applied Numerical Mathematics: Transactions of IMACS, 19(3):287–317, December 15, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=19&
issue=3&aid=636;   http:
//www.sciencedirect.com/
science/article/pii/0168927495000887.

Special issue on iterative methods for linear equations (Atlanta, GA, 1994).

Yang:1999:NFB

[YG99] U. Meier Yang and K. A. Gallivan. A new family of block methods. Applied Numerical Mathematics: Transactions of IMACS, 30(2–3):155–173, June 10, 1999. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1999&volume=30&
issue=2-3&aid=973.

Guo:2009:LGC

[yGpY09] Ben yu Guo and Jian ping Yan. Legendre–Gauss collocation method for initial value problems of second order ordinary differential equations. Applied Numerical Mathematics: Transactions of IMACS, 59(6): 1386–1408, June 2009. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Guo:2005:SOJ

[yGqWsWC05] Ben yu Guo, Zhong qing Wang, Zheng su Wan, and Delin Chu. Second order Jacobi approximation with applications to fourthorder differential equations. Applied Numerical Mathematics: Transactions of IMACS, 55(4):480–502, December 2005. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Yan:2015:FEM

[YGY15] Ming Yan, Wei Gong, and Ningning Yan. Finite element methods for elliptic optimal control problems with boundary observations. Applied Numerical Mathematics: Transactions of IMACS, 90(??):190–207, April 2015. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414002116.

Guo:2007:SMD

[yGyZ07] Ben yu Guo and Xiao yong Zhang. Spectral method for differential equations of degenerate type on unbounded domains by using generalized Laguerre functions. Applied Numerical Mathematics: Transactions of IMACS, 57(4):455–471, April 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Yang:2000:MPC

[YH00] B. Yang and J. S. Hesthaven. Multidomain pseudospectral computation of Maxwell’s equations in 3-D general curvilinear coordinates. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4):281–289, May 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/61/27/
55/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/61/27/55/article.
pdf.

Yuan:2007:PID

[YH07] Guangwei Yuan and Xudeng Hang. Parallel iterative difference schemes based on prediction techniques for S_{n} transport method. Applied Numerical Mathematics: Transactions of IMACS, 57(5–7):746–752, May 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Yang:2018:ANE

[YH18] Haijian Yang and Feng-Nan Hwang. An adaptive nonlinear elimination preconditioned inexact Newton algorithm for highly local nonlinear multicomponent PDE systems. Applied Numerical Mathematics: Transactions of IMACS, 133(??): 100–115, ???? 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418300175.

Yin:2023:GPI

[YHT23] Lina Yin, Yunqing Huang, and Qili Tang. A generalized parametric iterative finite element method for the 2D/3D stationary incompressible magnetohydrodynamics. Applied Numerical Mathematics: Transactions of IMACS, 193(??): 242–261, November 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423002210.

Yi:2012:AEE

[Yi12] Lijun Yi. On the asymptotic exactness of error estimators based on the equilibrated residual method for quadrilateral finite elements. Applied Numerical Mathematics: Transactions of IMACS, 62(12):

1925–1937, December 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927412001481.

Yang:2021:UOE

[YJ21] Yun-Bo Yang and Yao-Lin Jiang. Unconditional optimal error estimates of linearized second-order BDF Galerkin FEMs for the Landau–Lifshitz equation. Applied Numerical Mathematics: Transactions of IMACS, 159(??):21–45, January 2021. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742030249X.

Yang:2023:OEE

[YJ23] Yun-Bo Yang and Yao-Lin Jiang. Optimal error estimates of a lowestorder Galerkin-mixed FEM for the thermoviscoelastic Joule heating equations. Applied Numerical Mathematics: Transactions of IMACS, 183(??):86– 107, January 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422002227.

Yin:2024:PFP

[YJJ^{+}24] Jianghua Yin, Jinbao Jian, Xianzhen Jiang, Jiansheng Wu, and Guodong Ma. A proximal fully parallel splitting method with a relaxation factor for separable convex programming. Applied Numerical Mathematics: Transactions of IMACS, 195(??):17–38, January 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423002416.

Yang:2018:TSS

[YJZ18] Xiu Yang, Xiaoyun Jiang, and Hui Zhang. A timespace spectral tau method for the time fractional cable equation and its inverse problem. Applied Numerical Mathematics: Transactions of IMACS, 130(??): 95–111, August 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418300886.

Yanev:2004:EAB

[YK04a] Petko Yanev and Erricos J.

Kontoghiorghes. Efficient algorithms for block downdating of least squares solutions. Applied Numerical Mathematics: Transactions of IMACS, 49(1): 3–15, April 2004. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Yaremchuk:2004:POP

[YK04b] Svetlana Yaremchuk and Vyacheslav Kryzhanivskyy. Problem of optimal placement of discrete physical field sources. Applied Numerical Mathematics: Transactions of IMACS, 50(1):121–131, July 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Yanev:2007:CEM

[YK07] Petko I. Yanev and Erricos J. Kontoghiorghes. Computationally efficient methods for estimating the updated-observations SUR models. Applied Numerical Mathematics: Transactions of IMACS, 57(11– 12):1245–1258, November/ December 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Yang:2013:SFV

[YL13] Ming Yang and Jijun Liu. Solving a final value fractional diffusion problem by boundary condition regularization. Applied Numerical Mathematics: Transactions of IMACS, 66(??): 45–58, April 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892741200222X.

Yi:2026:RBI

[YL26] Yingdi Yi and Jijun Liu.

On the recovery of boundary impedance for 3-dimensional obstacle by acoustic wave scattering using modified Kaczmarz iteration algorithm. Applied Numerical Mathematics: Transactions of IMACS, 219(??): 53–72, January 2026. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425001746.

Yang:2020:ENM

[YLFT20] Shuiping Yang, Fawang Liu, Libo Feng, and Ian W. Turner. Efficient numerical methods for the nonlinear two-sided spacefractional diffusion equation with variable coefficients. Applied Numerical Mathematics: Transactions of IMACS, 157(??):55– 68, November 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420301604.

Yuan:2020:CGA

[YLH20] Gonglin Yuan, Tingting Li, and Wujie Hu. A conjugate gradient algorithm for large-scale nonlinear equations and image restoration problems. Applied Numerical Mathematics: Transactions of IMACS, 147(??): 129–141, January 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419302314.

Yang:2009:NFE

[YLL09] Yidu Yang, Qin Li, and Sirui Li. Nonconforming finite element approximations of the Steklov eigenvalue problem. Applied Numerical Mathematics: Transactions of IMACS, 59(10): 2388–2401, October 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Yao:2021:RNS

[YLL21] Teng-Teng Yao, Fang Lu, and Wei Li. A Riemannian nonmonotone spectral method for self-adjoint tangent vector field. Applied Numerical Mathematics: Transactions of IMACS, 161(??):208–217, March 2021. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742030355X.

Yin:2021:CET

[YLLZ21] Baoli Yin, Yang Liu, Hong Li, and Fanhai Zeng. A class of efficient time-stepping methods for multi-term timefractional reaction–diffusionwave equations. Applied Numerical Mathematics: Transactions of IMACS, 165(??):56–82, July 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000490.

Yu:2009:SGP

[YLS^{+}09] Zhensheng Yu, Ji Lin, Jing Sun, Yunhai Xiao, Liying Liu, and Zhanhui Li. Spectral gradient projection method for monotone nonlinear equations with convex constraints. Applied Numerical Mathematics: Transactions of IMACS, 59(10): 2416–2423, October 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Yao:2020:NSF

[YLW20a] Qingyun Yao, Yi Li, and Haibing Wang. Numerical solutions of the forward and inverse problems arising in diffuse optical tomography. Applied Numerical Mathematics: Transactions of IMACS, 154(??): 70–89, August 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300878.

Yuan:2020:PCG

[YLW20b] Gonglin Yuan, Junyu Lu, and Zhan Wang. The PRP conjugate gradient algorithm with a modified WWP line search and its application in the image restoration problems. Applied Numerical Mathematics: Transactions of IMACS, 152(??): 1–11, June 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300209.

Yu:2021:EST

[YLW21] Xuhong Yu, Long Li, and Zhongqing Wang. Efficient space-time Legendre rational spectral method for parabolic problems in unbounded domains. Applied Numerical Mathematics: Transactions of IMACS, 170(??):39–54, December 2021. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421002087.

Yang:2021:FPM

[YLX21] Yin Yang, Xueyang Li, and Aiguo Xiao. Fourier pseudospectral method for fractional stationary Schrödinger equation. Applied Numerical Mathematics: Transactions of IMACS, 165(??): 137–151, July 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000507.

Yang:2019:QAS

[YLY19] Yanbing Yang, Jiaheng Li, and Tao Yu. Qualitative analysis of solutions for a class of Kirchhoff equation with linear strong damping term, nonlinear weak damping term and power-type logarithmic source term. Applied Numerical Mathematics: Transactions of IMACS, 141(??):263–285, ???? 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419300042.

Yadav:2024:CAH

[YM24] Narendra Singh Yadav and Kaushik Mukherjee. Convergence analysis of higherorder approximation of singularly perturbed 2D semilinear parabolic PDEs with non-homogeneous boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 206(??): 210–246, December 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001971.

Yan:2021:CAH

[YMD21] Rian Yan, Qiang Ma, and Xiaohua Ding. Convergence analysis of the hpversion spectral collocation method for a class of nonlinear variable-order fractional differential equations. Applied Numerical Mathematics: Transactions of IMACS, 170(??): 269–297, December 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001422.

Yosibash:2000:CSS

[Yos00] Zohar Yosibash. Computing singular solutions of elliptic boundary value problems in polyhedral domains using the p-FEM. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4):71–93, May 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/61/27/
32/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/61/27/32/article.
pdf.

Yu:2018:CT

[YP18a] Yongchao Yu and Jigen Peng. Corrigendum to “The matrix splitting based proximal fixed-point algorithms for quadratically constrained l 1 minimization and Dantzig selector” [Appl. Numer. Math. 125 (2018)

23–50]. Applied Numerical Mathematics: Transactions of IMACS, 126(??): 199, April 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417302635.

See [YP18b].

Yu:2018:MSB

[YP18b] Yongchao Yu and Jigen Peng. The matrix splitting based proximal fixed-point algorithms for quadratically constrained l_{1} minimization and Dantzig selector. Applied Numerical Mathematics: Transactions of IMACS, 125(??):23–50, March 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417302283.

See corrigendum [YP18a].

Yan:2021:ATS

[YPD21] David Yan, M. C. Pugh, and F. P. Dawson. Adaptive time-stepping schemes for the solution of the Poisson– Nernst–Planck equations. Applied Numerical Mathematics: Transactions of IMACS, 163(??):254–269, May 2021. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000404.

Yang:2022:SOB

[YQCZ22] Xuehua Yang, Wenlin Qiu, Haifan Chen, and Haixiang Zhang. Second-order BDF ADI Galerkin finite element method for the evolutionary equation with a nonlocal term in three-dimensional space. Applied Numerical Mathematics: Transactions of IMACS, 172(??): 497–513, February 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421003123.

Young:1992:URI

[YR92] David M. Young and Bi Roubolo Vona. On the use of rational iterative methods for solving large sparse linear systems. Applied Numerical Mathematics: Transactions of IMACS, 10(3–4): 261–278, September 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). A Festschrift to honor Professor Garrett Birkhoff on his eightieth birthday.

Yun:2009:CLT

[YR09] Beong In Yun and Kyung Soo Rim. Construction of Lanczos type filters for the Fourier series approximation. Applied Numerical Mathematics: Transactions of IMACS, 59(2):280–300, February 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Yousefi:2022:WPG

[YR22] Hassan Yousefi and Timon Rabczuk. Wave propagation in generalized thermo-poro-elastic media via wavelet-based celladaptive central high resolution schemes using UNO limiters. Applied Numerical Mathematics: Transactions of IMACS, 173(??): 112–143, March 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421003196.

Yuttanan:2021:LWM

[YRV21a] Boonrod Yuttanan, Mohsen Razzaghi, and Thieu N. Vo. Legendre wavelet method for fractional delay differential equations. Applied Numerical Mathematics: Transactions of IMACS, 168(??): 127–142, October 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001537.

Yuttanan:2021:NMB

[YRV21b] Boonrod Yuttanan, Mohsen Razzaghi, and Thieu N. Vo. A numerical method based on fractional-order generalized Taylor wavelets for solving distributed-order fractional partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 160(??): 349–367, February 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420303251.

Yang:2009:PFV

[YS09] Min Yang and Huailing Song. A postprocessing finite volume element method for time-dependent Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 59(8): 1922–1932, August 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Yang:2022:OEE

[YS22] Huaijun Yang and Dongyang Shi. Optimal error estimates of Galerkin method for a nonlinear parabolic integro-differential equation. Applied Numerical Mathematics: Transactions of IMACS, 181(??):403–416, November 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422001726.

Yang:2024:EMJ

[YS24a] Xiu Yang and Changtao Sheng. Efficient mapped Jacobi spectral method for integral equations with twosided singularities. Applied Numerical Mathematics: Transactions of IMACS, 206(??):94–110, December 2024. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001995.

Yousif:2024:AMV

[YS24b] Osman Omer Osman Yousif and Mohammed A. Saleh. Another modified version of RMIL conjugate gradient method. Applied Numerical Mathematics: Transactions of IMACS, 202(??): 120–126, August 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000989.

Yang:2014:NRT

[YSBL14] Yidu Yang, Lingling Sun, Hai Bi, and Hao Li. A note on the residual type a posteriori error estimates for finite element eigenpairs of nonsymmetric elliptic eigenvalue problems. Applied Numerical Mathematics: Transactions of IMACS, 82(??): 51–67, August 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414000506.

Yang:2025:SMD

[YSZG25] Qian Yang, Xiaoqin Shen, Jikun Zhao, and Zhiming Gao. A surface mesh DG-VEM for elliptic membrane shell model. Applied Numerical Mathematics: Transactions of IMACS, 217(??): 298–318, November 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742500131X.

Yefet:2000:FOC

[YT00] A. Yefet and E. Turkel. Fourth order compact implicit method for the Maxwell equations with discontinuous coefficients. Applied Numerical Mathematics: Transactions of IMACS, 33(1–4): 125–134, May 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/29/17/21/61/27/
                                        36/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/29/17/21/61/27/36/article.
                                        pdf.

Yasir:2003:CSI

[YT03] Kamal H. Yasir and Yun Tang. Computation of the singularity induced bifurcation points in DAEs via extended system reduction. Applied Numerical Mathematics: Transactions of IMACS, 44(3):425– 431, February 2003. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Yang:2021:MSC

[YT21] Yin Yang and Zhuyan Tang.

Mapped spectral collocation methods for Volterra integral equations with noncompact kernels. Applied Numerical Mathematics: Transactions of IMACS, 160(??): 166–177, February 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302981.

Yang:2025:PEE

[YT25] Shuo Yang and Hongjiong Tian. A posteriori error estimates and time adaptivity for fully discrete finite element method for the incompressible Navier–Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 216(??): 17–38, October 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425000996.

Yan:2024:WBP

[YTC24] Ruifang Yan, Wei Tong, and Guoxian Chen. A wellbalanced positivity-preserving multidimensional central scheme for shallow water equations. Applied Numerical Mathematics: Transactions of IMACS, 197(??): 97–118, March 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423002830.

Yuan:2018:CMS

[YTZZ18] Xianju Yuan, Tianyu Tian, Hongni Zhou, and Jiwei Zhou. Comparisons of methods for solving static deflections of a thin annular plate. Applied Numerical Mathematics: Transactions of IMACS, 127(??): 266–279, May 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418300217.

Yu:1999:SPP

[Yu99] Hongyi Yu. Solving parabolic problems with different time steps in different regions in space based on domain decomposition methods. Applied Numerical Mathematics: Transactions of IMACS, 30(4):475–491, July 1, 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/

cas_sub/browse/browse.
cgi?year=1999&volume=30&
issue=4&aid=957.

Yu:2008:SBC

[Yu08]  Zhensheng Yu.  Solving

bound constrained optimization via a new nonmonotone spectral projected gradient method. Applied Numerical Mathematics: Transactions of IMACS, 58(9): 1340–1348, September 2008. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Yuan:1993:CTB

[Yua93]  J. Y. Yuan.  The con-

vergence of the two-block SAOR method for leastsquares problems. Applied Numerical Mathematics: Transactions of IMACS, 11(5):429–441, March 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Yuan:2020:CSL

[Yua20] Long Yuan. A combined scheme of the local spectral element method and the generalized plane wave discontinuous Galerkin method for the anisotropic Helmholtz equation. Applied Numerical Mathematics: Transactions of IMACS, 150(??): 341–360, April 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419302879.

Yuzbasi:2022:NBF

[Yüz22] Suayip Yüzbasi. A new Bell function approach to solve linear fractional differential equations. Applied Numerical Mathematics: Transactions of IMACS, 174(??):221–235, April 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422000149.

Yang:2017:EEC

[YV17] Huanhuan Yang and Alessandro Veneziani. Efficient estimation of cardiac conductivities via POD–DEIM model order reduction. Applied Numerical Mathematics: Transactions of IMACS, 115(??):180–199, May 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417300089.

Yang:2008:TFA

[YW08] Hongli Yang and Xinyuan Wu. Trigonometricallyfitted ARKN methods for perturbed oscillators. Applied Numerical Mathematics: Transactions of IMACS, 58(9):1375–1395, September 2008. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Yang:2019:TUT

[YW19] Tinggan Yang and Yihong Wang. Two uniform tailored finite point method for strongly anisotropic and discontinuous diffusivity. Applied Numerical Mathematics: Transactions of IMACS, 139(??):156–171, May 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419300170.

Yu:2024:ESS

[YW24] Xuhong Yu and Mengyao Wang. Efficient spectral and spectral element methods for Sobolev equation with diagonalization technique. Applied Numerical Mathematics: Transactions of IMACS, 201(??):265–281, July 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742400059X.

Yu:2025:TEI

[YW25] Xiaohui Yu and Qingbiao Wu. Two efficient iteration methods for solving the absolute value equations. Applied Numerical Mathematics: Transactions of IMACS, 208 (part B) (??):148–159, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424002769.

See comments [Guo26].

Yao:2020:PPT

[YWH20] Changhui Yao, Yifan Wei, and Qiumei Huang. Postprocessing technique of twogrid algorithm for wave propagation with Debye polarization in nonlinear dielectric materials. Applied Numerical Mathematics: Transactions of IMACS, 157(??): 405–418, November 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420301999.

Yu:2020:GJF

[YWSL20] Zhe Yu, Boying Wu, Jiebao Sun, and Wenjie Liu. A generalized-Jacobi-function spectral method for space– time fractional reaction– diffusion equations with viscosity terms. Applied Numerical Mathematics: Transactions of IMACS, 152(??):355–378, June 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892741930323X.

Yao:2023:EFL

[YWW23] G-Q. Yao, X. Wen, and Z-Q. Wang. An efficient Fourier– Laguerre spectral-Galerkin method for exterior problems of two-dimensional complex obstacles. Applied Numerical Mathematics: Transactions of IMACS, 193(??):93– 108, November 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001988.

Yan:2025:AAB

[YX25] Fengna Yan and Yinhua Xia.

Analysis of average bound preserving time-implicit discretizations for convection– diffusion–reaction equation. Applied Numerical Mathematics: Transactions of IMACS, 211(??):103–122, May 2025. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425000066.

Young:1995:PGI

[YXB95] David M. Young, Shengyou Xiao, and Karen R. Baker. Periodically generated iterative methods for solving elliptic equations. Applied Numerical Mathematics: Transactions of IMACS, 19(3):375–387, December 15, 1995. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-

                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1996&volume=19&
                                        issue=3&aid=640;   http:
                                        //www.sciencedirect.com/
                                        science/article/pii/0168927495000925.

Special issue on iterative methods for linear equations (Atlanta, GA, 1994).

Yan:2025:SAF

[YXF25] Qiuhui Yan, Xufeng Xiao, and Xinlong Feng. Stabilization and adaptive FEM for optimal control problems of stationary convectiondominated diffusion equations on surfaces. Applied Numerical Mathematics: Transactions of IMACS, 215(??):157–176, September 2025. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425000960.

Yang:2021:IFT

[YXN21] Shuping Yang, Xiangtuan Xiong, and Yan Nie. Iterated fractional Tikhonov regularization method for solving the spherically symmetric backward time-fractional diffusion equation. Applied Numerical Mathematics: Transactions of IMACS, 160(??):217–241, February 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303159.

Yan:2017:SCS

[YXT17] Zhiping Yan, Aiguo Xiao, and Xiao Tang. Strong convergence of the split-step theta method for neutral stochastic delay differential equations. Applied Numerical Mathematics: Transactions of IMACS, 120(??): 215–232, October 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741730123X.

Yue:2019:CMM

[YXX19] Meiling Yue, Hehu Xie, and Manting Xie. A cascadic multigrid method for nonsymmetric eigenvalue problem. Applied Numerical Mathematics: Transactions of IMACS, 146(??):55– 72, December 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927419301795.

Yao:2024:SSD

[YXX24] Lin Yao, Yinhua Xia, and Yan Xu. L-stable spectral deferred correction methods and applications to phase field models. Applied Numerical Mathematics: Transactions of IMACS, 197(??): 288–306, March 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423002994.

Yuan:2018:EAH

[YXZ18] Xinpeng Yuan, Chunguang Xiong, and Guoqing Zhu. Error analysis of the high order scheme for homogenization of Hamilton–Jacobi equation. Applied Numerical Mathematics: Transactions of IMACS, 126(??): 138–159, April 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417302544.

Yang:2024:VSS

[YXZL24] Zhen-Ping Yang, Shuilian Xie, Yong Zhao, and Gui-Hua Lin. Variable samplesize operator extrapolation algorithm for stochastic mixed variational inequalities. Applied Numerical Mathematics: Transactions of IMACS, 196(??):22– 44, February 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423002581.

Yang:2013:ASD

[YY13] Qing Yang and Yirang Yuan. An approximation of semiconductor device of heat conduction by mixed finite element method and characteristics-mixed finite element method. Applied Numerical Mathematics: Transactions of IMACS, 70(??):42–57, August 2013. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413000391.

Yu:2024:RTD

[YY24] Shuang Yu and Hongqi Yang.

Regularization with two differential operators and its application to inverse problems. Applied Numerical Mathematics: Transactions of IMACS, 201(??): 129–144, July 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000473.

Yang:2023:HOU

[YYZ23] Jun Yang, Nianyu Yi, and Hong Zhang. High-order, unconditionally maximumprinciple preserving finite element method for the Allen–Cahn equation. Applied Numerical Mathematics: Transactions of IMACS, 188(??):42–61, June 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423000636.

Yang:2017:DGM

[YZ17] Yubo Yang and Peng Zhu. Discontinuous Galerkin methods with interior penalties on graded meshes for 2D singularly perturbed convectiondiffusion problems. Applied Numerical Mathematics: Transactions of IMACS, 111(??):36–48, January 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416301714.

Yin:2019:SDF

[YZ19] Yunhui Yin and Peng Zhu. The streamline-diffusion finite element method on graded meshes for a convectiondiffusion problem. Applied Numerical Mathematics: Transactions of IMACS, 138(??):19–29, April 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418302861.

Ye:2021:NSD

[YZ21] Xiu Ye and Shangyou Zhang. A numerical scheme with divergence free H-div triangular finite element for the Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 167(??):211–217, September 2021. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742100132X.

Ye:2022:WDC

[YZ22] Xiu Ye and Shangyou Zhang. A weak divergence CDG method for the biharmonic equation on triangular and tetrahedral meshes. Applied Numerical Mathematics: Transactions of IMACS, 178(??): 155–165, August 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742200085X.

Yang:2024:SOS

[YZ24a] Hui Yang and Shengfeng Zhu. Shape optimization of Stokes flows by a penalty method. Applied Numerical Mathematics: Transactions of IMACS, 196(??): 133–158, February 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423002738.

Yang:2024:MVE

[YZ24b] Minghui Yang and Zhaojie Zhou. Mixed virtual element methods for elliptic optimal control problems with boundary observations in L^{2}(Γ). Applied Numerical Mathematics: Transactions of IMACS, 203(??): 97–112, September 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424001259.

Yang:2021:NAL

[YZC21] Zhanwen Yang, Tianqing Zuo, and Zhijie Chen. Numerical analysis of linearly implicit Euler-Riemann method for nonlinear Gurtin– MacCamy model. Applied Numerical Mathematics: Transactions of IMACS, 163(??):147–166, May 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420303913.

Yang:2023:HOC

[YZG23] Xiaojia Yang, Lin Zhang, and Yongbin Ge. Highorder compact finite difference schemes for solving the regularized long-wave equation. Applied Numerical Mathematics: Transactions of IMACS, 185(??): 165–187, March 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422003026.

Yang:2019:CAU

[YZH19a] Xiaofeng Yang, Guo-Dong Zhang, and Xiaoming He.

Convergence analysis of an unconditionally energy stable projection scheme for magneto-hydrodynamic equations. Applied Numerical Mathematics: Transactions of IMACS, 136(??): 235–256, February 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741830240X.

Yuan:2019:WNI

[YZH19b] Beibei Yuan, Mingwang Zhang, and Zhengwei Huang. A wide neighborhood interiorpoint algorithm with arcsearch for P_{∗}(κ) linear complementarity problem. Applied Numerical Mathematics: Transactions of IMACS, 136(??):293–304, February 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418302423.

Youssri:2024:RJS

[YZH24] Y. H. Youssri, M. A.

Zaky, and R. M. Hafez. Romanovski–Jacobi spectral schemes for high-order differential equations. Applied Numerical Mathematics: Transactions of IMACS, 198(??):148–159, April 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423003239.

Yan:2022:NRM

[YZQ^{+}22] Jingye Yan, Hong Zhang, Xu Qian, Xiaowei Chen, and Songhe Song. A novel regularized model for the logarithmic Klein–Gordon equation. Applied Numerical Mathematics: Transactions of IMACS, 176(??): 19–37, June 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742200040X.

Yan:2025:HOM

[YZWQ25] Jingye Yan, Hong Zhang, Yabing Wei, and Xu Qian. High-order and mass-conservative regularized implicit-explicit relaxation Runge–Kutta methods for the low regularity Schrödinger equations. Applied Numerical Mathematics: Transactions of IMACS, 216(??):210–221, October 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425001151.

Yang:2025:WGF

[YZZ25] Lin Yang, Qilong Zhai, and Ran Zhang. The weak Galerkin finite element method for Stokes interface problems with curved interface. Applied Numerical Mathematics: Transactions of IMACS, 208 (part B) (??):98–122, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742400271X.

Zmievskaya:2015:NSS

[ZAB15] G. I. Zmievskaya, T. A.

Averina, and A. L. Bondareva. Numerical solution of stochastic differential equations in the sense of Stratonovich in an amorphization crystal lattice model. Applied Numerical Mathematics: Transactions of IMACS, 93(??): 15–29, July 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741400097X.

Zaky:2021:USC

[ZAED21] Mahmoud A. Zaky, Ibrahem G. Ameen, Nermeen A. Elkot, and Eid H. Doha. A unified spectral collocation method for nonlinear systems of multi-dimensional integral equations with convergence analysis. Applied Numerical Mathematics: Transactions of IMACS, 161(??):27–45, March 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420303354.

Zaky:2019:EUN

[Zak19] Mahmoud A. Zaky. Existence, uniqueness and numerical analysis of solutions of tempered fractional boundary value problems. Applied Numerical Mathematics: Transactions of IMACS, 145(??): 429–457, November 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419301151.

Zaky:2020:ASC

[Zak20] Mahmoud A. Zaky. An accurate spectral collocation method for nonlinear systems of fractional differential equations and related integral equations with nonsmooth solutions. Applied Numerical Mathematics: Transactions of IMACS, 154(??):205–222, August 2020. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420301173.

Zang:1991:RSS

[Zan91] Thomas A. Zang. On the rotation and skew-symmetric forms for incompressible flow simulations. Applied Numerical Mathematics: Transactions of IMACS, 7(1):27– 40, January 1991. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0168927491901026.

Zanna:2001:FET

[Zan01]  Antonella Zanna.  The

Fer expansion and timesymmetry: A Strang-type approach. Applied Numerical Mathematics: Transactions of IMACS, 39(3–4):435–459, December 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/79/34/38/abstract.
html.

Zarantonello:1999:SPT

[Zar99] Sergio E. Zarantonello. SGI parallel technology for graphics and visualization. Applied Numerical Mathematics: Transactions of IMACS, 30(1):113–123, May 10, 1999. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1999/30/
1/953.pdf.

Zarin:2017:EGM

[Zar17] Helena Zarin. Exponentially graded mesh for a singularly perturbed problem with two small parameters. Applied Numerical Mathematics: Transactions of IMACS, 120(??): 233–242, October 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417301393.

Zhang:2007:DCM

[ZB07] Yanzhi Zhang and Weizhu Bao. Dynamics of the center of mass in rotating Bose– Einstein condensates. Applied Numerical Mathematics: Transactions of IMACS, 57(5–7):697–709, May 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Zhang:2019:RTF

[ZB19a] Xiaolong Zhang and John P. Boyd. Revisiting the Thomas–Fermi equation: Accelerating rational Chebyshev series through coordinate transformations. Applied Numerical Mathematics: Transactions of IMACS, 135(??):186–205, January 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418301922.

Zmour:2019:NAW

[ZB19b] Lhassane Zmour and Abderrahim Bouidi. A numerical approximation with WLS/ SUPG algorithm for solving White–Metzner viscoelastic flows. Applied Numerical Mathematics: Transactions of IMACS, 141(??): 206–219, ???? 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741830268X.

Zampa:2024:DFH

[ZBD24] E. Zampa, S. Busto, and M. Dumbser. A divergencefree hybrid finite volume/ finite element scheme for the incompressible MHD equations based on compatible finite element spaces with a posteriori limiting. Applied Numerical Mathematics: Transactions of IMACS, 198(??):346–374, April 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742400014X.

Zhang:2025:NAE

[ZBNC25] Zhizhuo Zhang, Mikaël Barboteu, Xiaobing Nie, and Jinde Cao. Numerical analysis of evolutionary mixed variational problems: Applications in modeling asphalt pavements with interlayer frictional contact conditions. Applied Numerical Mathematics: Transactions of IMACS, 209(??): 208–231, March 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424003234.

Zhang:2019:TGD

[ZBY19] Yu Zhang, Hai Bi, and Yidu Yang. The two-grid discretization of Ciarlet– Raviart mixed method for biharmonic eigenvalue problems. Applied Numerical Mathematics: Transactions of IMACS, 138(??): 94–113, April 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892741830271X.

Zenios:1991:PCB

[ZC91] Stavros A. Zenios and Yair Censor. Parallel computing with block-iterative image reconstruction algorithms. Applied Numerical Mathematics: Transactions of IMACS, 7(5): 399–415, June 1991. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/016892749190010W.

Zhu:1992:MGM

[ZC92] Jian-Ping Zhu and Yung Ming Chen. Multilevel grid method for history matching multidimensional multiphase reservoir models. Applied Numerical Mathematics: Transactions of IMACS, 10(2):159–174, July 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Zingg:1999:RKM

[ZC99] D. W. Zingg and T. T.

Chisholm. Runge–Kutta methods for linear ordinary differential equations. Applied Numerical Mathematics: Transactions of IMACS, 31(2):227–238, October 23, 1999. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1999&volume=31&
issue=2&aid=994;   http:
//www.sciencedirect.com/
science/article/pii/S0168927498001299.

Zhang:2010:BBV

[ZC10] Chengjian Zhang and Hao Chen. Block boundary value methods for delay differential equations. Applied Numerical Mathematics: Transactions of IMACS, 60(9): 915–923, September 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Zunino:2011:UIP

[ZCC11] Paolo Zunino, Laura Cattaneo, and Claudia Maria Colciago. An unfitted interface penalty method for the numerical approximation of contrast problems. Applied Numerical Mathematics: Transactions of IMACS, 61(10):1059–1076, October 2011. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Zweck:2021:SCL

[ZCGS21] John Zweck, Yanping Chen, Matthew J. Goeckner, and Yannan Shen. Spectral computation of low probability tails for the homogeneous Boltzmann equation. Applied Numerical Mathematics: Transactions of IMACS, 162(??):301–317, April 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000040.

Zhang:2025:EEI

[ZCQ25] Qifeng Zhang, Haiyan Cao, and Hongyu Qin. Error estimates of implicit-explicit compact BDF2 schemes for the pseudo parabolic equations with logarithmic nonlinearity. Applied Numerical Mathematics: Transactions of IMACS, 212(??): 135–154, June 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425000108.

Zhu:2011:ESM

[ZCSH11a] Huajun Zhu, Yaming Chen, Songhe Song, and Huayu Hu. Erratum to “Symplectic and multi-symplectic wavelet collocation methods for two-dimensional Schrödinger equations” [Applied Numerical Mathematics 61 (3) (2011) 308–321]. Applied Numerical Mathematics: Transactions of IMACS, 61(8):974–976, August 2011. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). See [ZCSH11b].

Zhu:2011:SMS

[ZCSH11b] Huajun Zhu, Yaming Chen, Songhe Song, and Huayu Hu. Symplectic and multisymplectic wavelet collocation methods for twodimensional Schrödinger equations. Applied Numerical Mathematics: Transactions of IMACS, 61(3):308– 321, March 2011. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). See erratum [ZCSH11a].

Zhang:2020:EES

[ZCY20] Jun Zhang, Chuanjun Chen, and Xiaofeng Yang. Efficient and energy stable method for the Cahn– Hilliard phase-field model for diblock copolymers. Applied Numerical Mathematics: Transactions of IMACS, 151(??):263–281, May 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419303381.

Zhang:2015:GPE

[ZCZ15] Bei Zhang, Shaochun Chen, and Jikun Zhao. Guaranteed a posteriori error estimates for nonconforming finite element approximations to a singularly perturbed reaction-diffusion problem. Applied Numerical Mathematics: Transactions of IMACS, 94(??): 1–15, August 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892741500029X.

Zhang:2020:GCM

[ZD20] Xiaoguang Zhang and Hong Du. A generalized collocation method in reproducing kernel space for solving a weakly singular Fredholm integro-differential equations. Applied Numerical Mathematics: Transactions of IMACS, 156(??): 158–173, October 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420301434.

Zhang:2021:RPG

[ZD21] Hui Zhang and Hua Dai.

The regularizing properties of global GMRES for solving large-scale linear discrete ill-posed problems with several right-hand sides. Applied Numerical Mathematics: Transactions of IMACS, 164(??):57–71, ???? 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302567.

Zhang:2025:IIF

[ZD25a] Yuxin Zhang and Hengfei Ding. Implicit integration factor method coupled with Padé approximation strategy for nonlocal Allen–Cahn equation. Applied Numerical Mathematics: Transactions of IMACS, 213(??): 88–107, July 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425000510.

Zhou:2025:NAA

[ZD25b] Hao Zhou and Duy-Minh Dang. Numerical analysis of American option pricing in a two-asset jump-diffusion model. Applied Numerical Mathematics: Transactions of IMACS, 216(??):98– 126, October 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425000674.

ZavaletaC:2003:ASS

[ZdBT03] A. U. Zavaleta C., A. L. de Bortoli, and M. Thompson. Analysis and simulation for a system of chemical reaction equations with a vortex formulation. Applied Numerical Mathematics: Transactions of IMACS, 47(3– 4):559–573, December 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Zaky:2018:SFF

[ZDM18] M. A. Zaky, E. H. Doha, and J. A. Tenreiro Machado. A spectral framework for fractional variational problems based on fractional Jacobi functions. Applied Numerical Mathematics: Transactions of IMACS, 132(??): 51–72, October 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418301193.

Zegeling:1997:REP

[Zeg97]  Paul A. Zegeling.  r-

refinement for evolutionary PDEs with finite elements or finite differences. Applied Numerical Mathematics: Transactions of IMACS, 26(1–2):97–104, December 22, 1997. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1997/26/
1-2/839.pdf.

Zennaro:1993:CRK

[Zen93] M. Zennaro. Contractivity of Runge–Kutta methods with respect to forcing terms. Applied Numerical Mathematics: Transactions of IMACS, 11(4):321– 345, February 1993. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Zeng:2021:CMB

[Zen21] Min-Li Zeng. A circulantmatrix-based new accelerated GSOR preconditioned method for block two-by-two linear systems from image restoration problems. Applied Numerical Mathematics: Transactions of IMACS, 164(??):245–257, ???? 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000076.

Zheng:2020:WTS

[ZEW20] Xiangcheng Zheng, V. J.

Ervin, and Hong Wang. Wellposedness of the twosided variable coefficient Caputo flux fractional diffusion equation and error estimate of its spectral approximation. Applied Numerical Mathematics: Transactions of IMACS, 153(??): 234–247, July 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300659.

Zhang:2020:KSM

[ZFC20] Liping Zhang, Hung-Yuan Fan, and Eric King-Wah Chu. Krylov subspace methods for discrete-time algebraic Riccati equations. Applied Numerical Mathematics: Transactions of IMACS, 152(??):499–510, June 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419303149.

Zhang:2024:FDL

[ZFS24] Yue Zhang, Xinlong Feng, and Haiyan Su. Fully decoupled, linear and unconditionally energy stable time discretization scheme for solving the unsteady thermally coupled magnetohydrodynamic equations with variable density. Applied Numerical Mathematics: Transactions of IMACS, 197(??): 210–229, March 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423002921.

Zhang:2020:LTE

[ZFW20] Huili Zhang, Xinlong Feng, and Kun Wang. Long time error estimates of IFE methods for the unsteady multi-layer porous wall model. Applied Numerical Mathematics: Transactions of IMACS, 156(??): 303–321, October 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420301446.

Zhao:2017:DDS

[ZFX17] Jingjun Zhao, Yan Fan, and Yang Xu. Delaydependent stability of symmetric Runge–Kutta methods for second order delay differential equations with three parameters. Applied Numerical Mathematics: Transactions of IMACS, 117(??):103–114, July 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417300661.

Zhou:2025:DLM

[ZFX25] Wenrong Zhou, Hongfei Fu, and Shusen Xie. A decoupled linear, mass- and energyconserving relaxation-type high-order compact finite difference scheme for the nonlinear Schrödinger equation. Applied Numerical Mathematics: Transactions of IMACS, 215(??):59– 89, September 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425000856.

Zhang:2019:GDF

[ZFZ19] Li Zhang, Minfu Feng, and Jian Zhang. A globally divergence-free weak Galerkin method for Brinkman equations. Applied Numerical Mathematics: Transactions of IMACS, 137(??): 213–229, March 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418302514.

Zhang:1992:BPD

[ZG92a] Wen Zhang and Ian Gladwell. Bifurcation phenomena in a detonation problem. Applied Numerical Mathematics: Transactions of IMACS, 9(3–5):427–445, April 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/0168927492900329.

Zhang:1992:BPS

[ZG92b] Wen Zhang and Ian Gladwell. Bifurcation phenomena in singular free boundary value problems of ordinary differential equations. Applied Numerical Mathematics: Transactions of IMACS, 10(2):149–158, July 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Zhang:2020:RGR

[ZG20] Jianhua Zhang and Jinghui Guo. On relaxed greedy randomized coordinate descent methods for solving large linear least-squares problems. Applied Numerical Mathematics: Transactions of IMACS, 157(??): 372–384, November 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420301963.

Zhang:2021:NSN

[ZG21] Lin Zhang and Yongbin Ge.

Numerical solution of nonlinear advection diffusion reaction equation using highorder compact difference method. Applied Numerical Mathematics: Transactions of IMACS, 166(??): 127–145, August 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421001033.

Zhu:2017:WBH

[ZGDL17] Qiangqiang Zhu, Zhen Gao, Wai Sun Don, and Xianqing Lv. Well-balanced hybrid compact-WENO scheme for shallow water equations. Applied Numerical Mathematics: Transactions of IMACS, 112(??):65–78, February 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301957.

Zhang:2025:FLM

[ZGF^{+}25] Xiaoyan Zhang, Guangyu Gao, Zhenwu Fu, Yang Li, and Bo Han. A frozen Levenberg–Marquardt–Kaczmarz method with convex penalty terms and two-point gradient strategy for ill-posed problems. Applied Numerical Mathematics: Transactions of IMACS, 209(??): 187–207, March 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424003222.

Zhang:1998:EPG

[ZGL98] T. Zhang, G. H. Golub, and K. H. Law. Eigenvalue perturbation and generalized Krylov subspace method. Applied Numerical Mathematics: Transactions of IMACS, 27(2):185– 202, June 10, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1998/27/
2/869.pdf.

Zhang:2012:EPC

[ZGO12] B. Zhang, L. A. Grzelak, and C. W. Oosterlee. Efficient pricing of commodity options with early-exercise under the Ornstein–Uhlenbeck process. Applied Numerical Mathematics: Transactions of IMACS, 62(2):91– 111, February 2012. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411001863.

Zhao:2023:WGM

[ZGR23] Jijing Zhao, Fuzheng Gao, and Hongxing Rui. The weak Galerkin method for the miscible displacement of incompressible fluids in porous media on polygonal mesh. Applied Numerical Mathematics: Transactions of IMACS, 185(??): 530–548, March 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422003348.

Zhang:2009:EOL

[ZH09] Chengjian Zhang and Yaoyao He. The extended one-leg methods for nonlinear neutral delay-integro-differential equations. Applied Numerical Mathematics: Transactions of IMACS, 59(6): 1409–1418, June 2009. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Zhou:2015:NAT

[ZH15] Qunyan Zhou and Dan Hang. Nonmonotone adaptive trust region method with line search based on new diagonal updating. Applied Numerical Mathematics: Transactions of IMACS, 91(??): 75–88, May 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927414002141.

Zare:2020:DRP

[ZH20] Hossein Zare and Masoud Hajarian. Determination of regularization parameter via solving a multiobjective optimization problem. Applied Numerical Mathematics: Transactions of IMACS, 156(??): 542–554, October 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420301653.

Zaky:2021:EDP

[ZH21] Mahmoud A. Zaky and Ahmed S. Hendy. An efficient dissipation-preserving Legendre–Galerkin spectral method for the Higgs boson equation in the de Sitter spacetime universe. Applied Numerical Mathematics: Transactions of IMACS, 160(??):281–295, February 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303202.

Zhao:2024:TEA

[ZH24] Guomei Zhao and Shuaifei Hu. Temporal error analysis of an unconditionally energy stable secondorder BDF scheme for the square phase-field crystal model. Applied Numerical Mathematics: Transactions of IMACS, 202(??): 222–245, August 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424001144.

Zhang:1996:UMS

[Zha96] Wen Zhang. Using MOL to solve a high order nonlinear PDE with a moving boundary in the simulation of a sintering process. Applied Numerical Mathematics: Transactions of IMACS, 20(1–2):235–244, February 29, 1996. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1996&volume=20&

                                        issue=1-2&aid=677. Work-

shop on the method of lines for time-dependent problems (Lexington, KY, 1995).

Zhang:1997:MIM

[Zha97]  Jun Zhang.  Multigrid

with inexact minimal residual smoothing acceleration. Applied Numerical Mathematics: Transactions of IMACS, 24(4):501–512, August 12, 1997. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

                                        //www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=24&
                                        issue=4&aid=766.

Zhang:2000:SAI

[Zha00] Jun Zhang. Sparse approximate inverse and multilevel block ILU preconditioning techniques for general sparse matrices. Applied Numerical Mathematics: Transactions of IMACS, 35(1):67– 86, September 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

                                        http://www.elsevier.nl/
                                        gej-ng/10/10/28/64/27/
                                        31/abstract.html;   http:
                                        //www.elsevier.nl/gej-
                                        ng/10/10/28/64/27/31/article.
                                        pdf.

Zhao:2001:EIS

[Zha01] Shuangsuo Zhao. An efficient implementation scheme of the simplified Newton iteration for block systems of nonlinear equations. Applied Numerical Mathematics: Transactions of IMACS, 39(2):225–237, November 2001. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/79/33/33/abstract.
html.

Zhao:2007:ENI

[Zha07]  Y. B. Zhao.  Enlarging

neighborhoods of interiorpoint algorithms for linear programming via least values of proximity measure functions. Applied Numerical Mathematics: Transactions of IMACS, 57(9): 1033–1049, September 2007. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Zhang:2009:FCD

[Zha09] Shangyou Zhang. A family of 3D continuously differentiable finite elements on tetrahedral grids. Applied Numerical Mathematics: Transactions of IMACS, 59(1):219–233, January 2009. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Zhang:2014:FOA

[Zha14] Qinghai Zhang. A fourthorder approximate projection method for the incompressible Navier–Stokes equations on locally-refined periodic domains. Applied Numerical Mathematics: Transactions of IMACS, 77(??):16–30, March 2014. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413001517.

Zhan:2019:SOC

[Zha19a]  Qingyi Zhan.  Shadow-

ing orbits of a class of random differential equations. Applied Numerical Mathematics: Transactions of IMACS, 136(??): 206–214, February 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741830237X.

Zhang:2019:SCF

[Zha19b] Yunzhang Zhang. Stability and convergence of first order time discrete linearized pressure correction projection method for the diffusive Peterlin viscoelastic model. Applied Numerical Mathematics: Transactions of IMACS, 139(??):93–114, May 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741830285X.

Zhang:2020:CBE

[Zha20a] Wei Zhang. Convergence of the balanced Euler method for a class of stochastic Volterra integro-differential equations with non-globally Lipschitz continuous coefficients. Applied Numerical Mathematics: Transactions of IMACS, 154(??): 17–35, August 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300775.

Zhang:2020:TNA

[Zha20b]  Wei Zhang.  Theoreti-

cal and numerical analysis of a class of stochastic Volterra integro-differential equations with non-globally Lipschitz continuous coefficients. Applied Numerical Mathematics: Transactions of IMACS, 147(??): 254–276, January 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419302272.

Zhang:2021:IBS

[Zha21a] Ju-Li Zhang. Inexact block SSOR-like preconditioners for non-Hermitian positive definite linear systems of strong Hermitian parts. Applied Numerical Mathematics: Transactions of IMACS, 165(??):598–613, July 2021.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421000647.

Zhao:2021:HEM

[Zha21b] Zhenyu Zhao. A Hermite extension method for numerical differentiation. Applied Numerical Mathematics: Transactions of IMACS, 159(??):46–60, January 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420302518.

Zheng:2007:EAC

[Zhe07] Chunxiong Zheng. Error analysis of a coupled finite– infinite element method for an exterior Poisson problem. Applied Numerical Mathematics: Transactions of IMACS, 57(3):270–280, March 2007. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Zheng:2019:SBP

[Zhe19] Guang-Hui Zheng. Solving the backward problem in Riesz–Feller fractional diffusion by a new nonlocal regularization method. Applied Numerical Mathematics: Transactions of IMACS, 135(??):99–128, January 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0168927418301764.

Zhang:2014:MAP

[ZHJ14] Wenxing Zhang, Deren Han, and Suoliang Jiang. A modified alternating projection based predictioncorrection method for structured variational inequalities. Applied Numerical Mathematics: Transactions of IMACS, 83(??):12– 21, September 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414000671.

Zhou:2003:ODD

[ZHL03] X. Zhou, Y. C. Hon, and Jichun Li. Overlapping domain decomposition method by radial basis functions. Applied Numerical Mathematics: Transactions of IMACS, 44(1–2):241–255, January 2003. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Zhou:2008:EMM

[ZHL08] Tianhe Zhou, Danfu Han, and Ming-Jun Lai. Energy minimization method for scattered data Hermite interpolation. Applied Numerical Mathematics: Transactions of IMACS, 58(5):

646–659, May 2008. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Zhang:2022:SEA

[ZHL22] Li Zhang, Jin Huang, and Hu Li. Splitting extrapolation algorithms for solving linear delay Volterra integral equations with a spatial variable. Applied Numerical Mathematics: Transactions of IMACS, 178(??): 372–385, August 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927422001039.

Zhou:2017:FDM

[Zho17] Guanyu Zhou. The fictitious domain method with penalty for the parabolic problem in moving-boundary domain: the error estimate of penalty and the finite element approximation. Applied Numerical Mathematics: Transactions of IMACS, 115(??):42–67, May 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417300028.

Zhou:2018:FDM

[Zho18] Guanyu Zhou. The fictitious domain method with H^{1}- penalty for the Stokes problem with Dirichlet boundary condition. Applied Numerical Mathematics: Transactions of IMACS, 123(??): 1–21, January 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417301782.

Zaky:2022:LJC

[ZHS22] Mahmoud A. Zaky, Ahmed S.

Hendy, and D. Suragan. Logarithmic Jacobi collocation method for Caputo– Hadamard fractional differential equations. Applied Numerical Mathematics: Transactions of IMACS, 181(??):326–346, November 2022. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001714.

Zioko:1999:SMC

[Zió99]  Mariusz Zióko.  Stabil-

ity of method of characteristics. Applied Numerical Mathematics: Transactions of IMACS, 31(4): 463–486, December 1999. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/46/20/
21/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/46/20/21/article.
pdf.

Zinn-Justin:2010:SDS

[ZJ10]  Jean Zinn-Justin.  Sum-

mation of divergent series: Order-dependent mapping. Applied Numerical Mathematics: Transactions of IMACS, 60(12):1454–1464, December 2010. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Zhang:2019:UCN

[ZJ19a] Hui Zhang and Xiaoyun Jiang. Unconditionally convergent numerical method for the two-dimensional nonlinear time fractional diffusion-wave equation. Applied Numerical Mathematics: Transactions of IMACS, 146(??):1–12, December 2019. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419301710.

Zhang:2019:SMS

[ZJ19b] Liying Zhang and Lihai Ji. Stochastic multi-symplectic Runge–Kutta methods for stochastic Hamiltonian PDEs. Applied Numerical Mathematics: Transactions of IMACS, 135(??):396–406, January 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418302083.

Zhou:2019:NSH

[ZJ19c] Shaobo Zhou and Hai Jin. Numerical solution to highly nonlinear neutraltype stochastic differential equation. Applied Numerical Mathematics: Transactions of IMACS, 140(??): 48–75, June 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741930025X.

Zhang:2025:TFO

[ZJ25] Xin Zhang and Yuanfeng Jin. Two fourth-order conservative compact difference schemes for the generalized Korteweg-de Vries– Benjamin Bona Mahony equation. Applied Numerical Mathematics: Transactions of IMACS, 212(??): 223–241, June 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425000297.

Zhang:2018:CNA

[ZJH18] Tong Zhang, JiaoJiao Jin, and YuGao HuangFu. The Crank–Nicolson/Adams–Bashforth scheme for the Burgers equation with H_{2} and H_{1} initial data. Applied Numerical Mathematics: Transactions of IMACS, 125(??): 103–142, March 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892741730226X.

Zhang:2023:FNS

[ZJH^{+}23] Meihui Zhang, Jinhong Jia, Ahmed S. Hendy, Mahmoud A. Zaky, and Xiangcheng Zheng. Fast numerical scheme for the time-fractional option pricing model with asset-price-dependent variable order. Applied Numerical Mathematics: Transactions of IMACS, 192(??):414–430, October 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423001794.

Zheng:2022:MTT

[ZJLA22] Minling Zheng, Zhengmeng Jin, Fawang Liu, and Vo Anh. Matrix transfer technique for anomalous diffusion equation involving fractional Laplacian. Applied Numerical Mathematics: Transactions of IMACS, 172(??): 242–258, February 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421002919.

rightsarereserved:2025:MAM

[ZJR25] Alaeddine Zahir, Khalide Jbilou, and Ahmed Ratnani. Multilinear algebra methods for higher-dimensional graphs. Applied Numerical Mathematics: Transactions of IMACS, 208 (part A) (??):390–407, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927423002829.

Zubik-Kowal:2000:CPM

[ZK00] Barbara Zubik-Kowal. Chebyshev pseudospectral method and waveform relaxation for differential and differentialfunctional parabolic equations. Applied Numerical Mathematics: Transactions of IMACS, 34(2–3): 309–328, July 2000. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/29/17/21/62/29/
39/abstract.html;   http:
//www.elsevier.nl/gej-
ng/29/17/21/62/29/39/article.
pdf.

Zhou:2021:APN

[ZKO^{+}21] Guanyu Zhou, Takahito Kashiwabara, Issei Oikawa, Eric Chung, and Ming-Cheng Shiue. An analysis on the penalty and Nitsche’s methods for the Stokes–Darcy system with a curved interface. Applied Numerical Mathematics: Transactions of IMACS, 165(??):83–118, July 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421000489.

Zhang:2011:HIS

[ZL11a] Ming Zhang and Xue-Zhang Liang. On a Hermite interpolation on the sphere. Applied Numerical Mathematics: Transactions of IMACS, 61(5):666–674, May 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Zhao:2011:HSP

[ZL11b] Zhenyu Zhao and Junfeng Liu. Hermite spectral and pseudospectral methods for numerical differentiation. Applied Numerical Mathematics: Transactions of IMACS, 61(12): 1322–1330, December 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927411001693.

Zeng:2017:NCN

[ZL17] Fanhai Zeng and Changpin Li. A new Crank–Nicolson finite element method for the time-fractional subdiffusion equation. Applied Numerical Mathematics: Transactions of IMACS, 121(??):82– 95, November 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417301472.

Zhang:2018:SCI

[ZL18a] Jin Zhang and Xiaowei Liu.

Supercloseness of the continuous interior penalty method for singularly perturbed problems in 1D: Vertexcell interpolation. Applied Numerical Mathematics: Transactions of IMACS, 123(??):88–98, January 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417301848.

Zhang:2018:MCA

[ZL18b] Tingting Zhang and Hui Liang. Multistep collocation approximations to solutions of first-kind Volterra integral equations. Applied Numerical Mathematics: Transactions of IMACS, 130(??): 171–183, August 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418300941.

Zhang:2021:UCF

[ZL21] Jin Zhang and Xiaowei Liu.

Uniform convergence of finite element methods on Bakhvalov-type meshes in the case of N^{−1} ≤ ϵ. Applied Numerical Mathematics: Transactions of IMACS, 165(??):519–526, July 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000830.

Zhang:2022:SFE

[ZL22] Jin Zhang and Yanhui Lv. Supercloseness of finite element method on a Bakhvalov-type mesh for a singularly perturbed problem with two parameters. Applied Numerical Mathematics: Transactions of IMACS, 171(??):329–352, January 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421002646.

Zhang:2023:USE

[ZL23] Xiaodi Zhang and Xiaonian Long. Unconditional stability and error analysis of an Euler IMEX-SAV scheme for the micropolar Navier–Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 192(??): 214–240, October 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423001563.

Zlotnik:2024:SEB

[ZL24] Alexander Zlotnik and Timofey Lomonosov. On stability and error bounds of an explicit in time higher-order vector compact scheme for the multidimensional wave and acoustic wave equations. Applied Numerical Mathematics: Transactions of IMACS, 195(??): 54–74, January 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742300243X.

Zlatev:1985:GSS

[Zla85a]  Zahari Zlatev.  General

scheme for solving linear algebraic problems by direct methods. Applied Numerical Mathematics: Transactions of IMACS, 1(2):177– 186, March 1985. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Zlatev:1985:VSV

[Zla85b] Zahari Zlatev. Variable stepsize variable formula methods based on predictorcorrector schemes. Applied Numerical Mathematics: Transactions of IMACS, 1(5):395–416, September 1985. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Zhang:2020:LET

[ZLCH20] Li-Ping Zhang, Zi-Cai Li, Zhen Chen, and Hung-Tsai Huang. The Laplace equation in three dimensions by the method of fundamental solutions and the method of particular solutions. Applied Numerical Mathematics: Transactions of IMACS, 154(??):47–69, August 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300751.

Zheng:2015:UCH

[ZLG15] Quan Zheng, Xuezheng Li, and Yue Gao. Uniformly convergent hybrid schemes for solutions and derivatives in quasilinear singularly perturbed BVPs. Applied Numerical Mathematics: Transactions of IMACS, 91(??):46–59, May 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927415000021.

Zhao:2024:STP

[ZLG24] Zhihui Zhao, Hong Li, and Wei Gao. A space-time Petrov–Galerkin method for the two-dimensional regularized long-wave equation. Applied Numerical Mathematics: Transactions of IMACS, 198(??):276–294, April 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000072.

Zhang:2019:MMF

[ZLHW19] Li-Ping Zhang, Zi-Cai Li, Hung-Tsai Huang, and Yimin Wei. The modified method of fundamental solutions for exterior problems of the Helmholtz equation; spurious eigenvalues and their removals. Applied Numerical Mathematics: Transactions of IMACS, 145(??): 236–260, November 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927419301606.

Zhang:2020:EBB

[ZLJ20] Chengjian Zhang, Cui Li, and Jingyao Jiang. Extended block boundary value methods for neutral equations with piecewise constant argument. Applied Numerical Mathematics: Transactions of IMACS, 150(??): 182–193, April 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419302806.

Zhao:2022:FEF

[ZLL22] Xuan Zhao, Xiaoli Li, and Ziyan Li. Fast and efficient finite difference method for the distributed-order diffusion equation based on the staggered grids. Applied Numerical Mathematics: Transactions of IMACS, 174(??):34–45, April 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742200006X.

Zheng:2020:NMA

[ZLS20] Yaqin Zheng, Yingzhen Lin, and Yang Shen. A new multiscale algorithm for solving second order boundary value problems. Applied Numerical Mathematics: Transactions of IMACS, 156(??): 528–541, October 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420301641.

Zhang:2022:SFD

[ZLSZ22] Fan Zhang, Dongfang Li, Hai-Wei Sun, and Jia-Li Zhang. A stabilized fully-discrete scheme for phase field crystal equation. Applied Numerical Mathematics: Transactions of IMACS, 178(??): 337–355, August 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001003.

Zhang:2020:SCA

[ZLW20a] Jun Zhang, Shimin Lin, and JinRong Wang. Stability and convergence analysis of Fourier pseudo-spectral method for FitzHugh–Nagumo model. Applied Numerical Mathematics: Transactions of IMACS, 157(??): 563–578, November 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302099.

Zhang:2020:IFE

[ZLW20b] Zhiyue Zhang, Dong Liang, and Quanxiang Wang. Immersed finite element method and its analysis for parabolic optimal control problems with interfaces. Applied Numerical Mathematics: Transactions of IMACS, 147(??): 174–195, January 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419302338.

Zhang:2022:EST

[ZLW22] Tiangong Zhang, Huiyuan Li, and Zhongqing Wang. Efficient space–time Jacobi rational spectral methods for second order time-dependent problems on unbounded domains. Applied Numerical Mathematics: Transactions of IMACS, 176(??):

159–181, June 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422000496.

Zheng:2021:OOF

[ZLWF21] Xiangcheng Zheng, Huan Liu, Hong Wang, and Hongfei Fu. Optimal-order finite element approximations to variable-coefficient two-sided space-fractional advection–reaction–diffusion equations in three space dimensions. Applied Numerical Mathematics: Transactions of IMACS, 161(??): 1–12, March 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303299.

Zhao:2019:KPI

[ZLX19] Jingjun Zhao, Yu Li, and Yang Xu. A kind of product integration scheme for solving fractional ordinary differential equations. Applied Numerical Mathematics: Transactions of IMACS, 136(??): 279–292, February 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418302411.

Zhang:2022:EAI

[ZLX22] Tao Zhang, Xiaolin Li, and Liwei Xu. Error analysis of an implicit Galerkin meshfree scheme for general second-order parabolic problems. Applied Numerical Mathematics: Transactions of IMACS, 177(??): 58–78, July 2022. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422000678.

Zha:2023:SPC

[ZLY23] Yuanyuan Zha, Zhe Li, and Lijun Yi. Superconvergent postprocessing of the C^{1}-conforming finite element method for fourthorder boundary value problems. Applied Numerical Mathematics: Transactions of IMACS, 193(??):67– 82, November 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423002027.

Zhang:2017:OPS

[ZM17] Zhongqiang Zhang and Heping Ma. Order-preserving strong schemes for SDEs with locally Lipschitz coefficients. Applied Numerical Mathematics: Transactions of IMACS, 112(??):1– 16, February 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927416301891.

Zafarghandi:2019:NAR

[ZM19] Fahimeh Saberi Zafarghandi and Maryam Mohammadi. Numerical approximations for the Riesz space fractional advection-dispersion equations via radial basis functions. Applied Numerical Mathematics: Transactions of IMACS, 144(??): 59–82, October 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419301187.

Zhang:2025:OEB

[ZM25] Teng Zhang and Ying Ma. Optimal error bounds of the time-splitting sinepseudospectral method for the biharmonic nonlinear Schrödinger equation. Applied Numerical Mathematics: Transactions of IMACS, 207(??):414–430, January 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424002460.

Zhang:2013:PES

[ZMC13] Jin Zhang, Liquan Mei, and Yaping Chen. Pointwise estimates of the SD-FEM for convection-diffusion problems with characteristic layers. Applied Numerical Mathematics: Transactions of IMACS, 64(1):19– 34, February 2013. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927412001754.

Zhang:2025:RGD

[ZMF25] Xian Zhang, Ya Min, and Minfu Feng. Robust globally divergence-free weak Galerkin variational data assimilation method for convection-dominated Oseen equations. Applied Numerical Mathematics: Transactions of IMACS, 218(??):22– 42, December 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425001527.

Zharovsky:2012:NSC

[ZML^{+}12] Evgeniy Zharovsky, Amin Moosaie, Anne Le Duc, Michael Manhart, and Bernd Simeon. On the numerical solution of a convectiondiffusion equation for particle orientation dynamics on geodesic grids. Applied Numerical Mathematics: Transactions of IMACS, 62(10):1554–1566, October 2012. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927412000906.

Zhang:2021:ECH

[ZMY21] Mingzhu Zhang, Xinyu Mao, and Lijun Yi. Exponential convergence of the hp-version of the composite Gauss– Legendre quadrature for integrals with endpoint singularities. Applied Numerical Mathematics: Transactions of IMACS, 170(??): 340–352, December 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421002312.

Zhang:2021:EAD

[ZN21] Jianjun Zhang and James G. Nagy. An effective alternating direction method of multipliers for color image restoration. Applied Numerical Mathematics: Transactions of IMACS, 164(??): 43–56, ???? 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420302087.

Zhang:2002:IOP

[ZNK02] Shao-Liang Zhang, Kazuhide Nakata, and Masakazu Kojima. Incomplete orthogonalization preconditioners for solving large and dense linear systems which arise from Semidefinite Programming. Applied Numerical Mathematics: Transactions of IMACS, 41(1):235–245, April 2002. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/gej-ng/10/
10/28/86/27/40/abstract.
html.

Zhang:2014:PEE

[ZO14] B. Zhang and C. W. Oosterlee. Pricing of early-exercise Asian options under Lévy processes based on Fourier cosine expansions. Applied Numerical Mathematics: Transactions of IMACS, 78(??):14–30, April 2014. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927413001621.

Zou:2010:HEE

[Zou10] Qingsong Zou. Hierarchical error estimates for finite volume approximation solution of elliptic equations. Applied Numerical Mathematics: Transactions of IMACS, 60(1–2):142–153, January/February 2010. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Zou:2011:ERH

[Zou11]  Qingsong Zou.  Efficient

and reliable hierarchical error estimates for an elliptic obstacle problem. Applied Numerical Mathematics: Transactions of IMACS, 61(3):344–355, March 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Zhang:2009:TLE

[ZOZ09] Lin Zhang, Jie Ouyang, and Xiao-Hua Zhang. On a twolevel element-free Galerkin method for incompressible fluid flow. Applied Numerical Mathematics: Transactions of IMACS, 59(8): 1894–1904, August 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Zhu:1997:ASL

[ZP97] Wenjie Zhu and Linda R. Petzold. Asymptotic stability of linear delay differentialalgebraic equations and numerical methods. Applied Numerical Mathematics: Transactions of IMACS, 24(2–3):247–264, June 19, 1997. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:

//www.elsevier.com/cgi-
bin/cas/tree/store/apnum/
cas_sub/browse/browse.
cgi?year=1997&volume=24&
issue=2-3&aid=780. Volterra

centennial (Tempe, AZ, 1996).

Zhu:1998:ASH

[ZP98] Wenjie Zhu and Linda R.

Petzold. Asymptotic stability of Hessenberg delay differential–algebraic equations of retarded or neutral type. Applied Numerical Mathematics: Transactions of IMACS, 27(3):309– 325, July 1, 1998. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (elec-

tronic). URL http://www.
elsevier.com/cas/tree/
store/apnum/sub/1998/27/
3/874.pdf.

Zhu:2012:LSF

[ZP12] Xiaojing Zhu and Dingguo Pu. A line search filter algorithm with inexact step computations for equality constrained optimization. Applied Numerical Mathematics: Transactions of IMACS, 62(3):212–223, March 2012. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927411002145.

Zampieri:2024:NCG

[ZP24] Elena Zampieri and Luca F.

Pavarino. A numerical comparison of Galerkin and Collocation Isogeometric approximations of acoustic wave problems. Applied Numerical Mathematics: Transactions of IMACS, 200(??):453–465, June 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423001575.

Zheng:1992:TDI

[ZPT92] R. Zheng and N. Phan-Thien. Transforming the domain integrals to the boundary using approximate particular solutions: a boundary element approach for nonlinear problems. Applied Numerical Mathematics: Transactions of IMACS, 10(5):435– 445, October 1992. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Zhu:2011:RMW

[ZQLK11] Jun Zhu, Jianxian Qiu, Tiegang Liu, and Boo Cheong Khoo. RKDG methods with WENO type limiters and conservative interfacial procedure for one-dimensional compressible multi-medium flow simulations. Applied Numerical Mathematics: Transactions of IMACS, 61(4):554–580, April 2011. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Zhang:2018:FDF

[ZQY18] Tong Zhang, Yanxia Qian, and JinYun Yuan. The fully discrete fractional-step method for the Oldroyd model. Applied Numerical Mathematics: Transactions of IMACS, 129(??): 83–103, July 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418300588.

Zhang:2023:MEE

[ZQZ23] Jiyuan Zhang, Yifan Qin, and Qifeng Zhang. Maximum error estimates of two linearized compact difference schemes for two-dimensional nonlinear Sobolev equations. Applied Numerical Mathematics: Transactions of IMACS, 184(??): 253–272, February 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422002744.

Zlotnik:2015:NCN

[ZR15] A. Zlotnik and A. Romanova.

On a Numerov–Crank– Nicolson–Strang scheme with discrete transparent boundary conditions for the Schrödinger equation on a semi-infinite strip. Applied Numerical Mathematics: Transactions of IMACS, 93(??):279–294, July 2015. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414000749.

Zhang:2021:MMB

[ZR21] Li-Li Zhang and Zhi-Ru Ren. A modified modulus-based multigrid method for linear complementarity problems arising from free boundary problems. Applied Numerical Mathematics: Transactions of IMACS, 164(??): 89–100, ???? 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927420302841.

Zara:2023:KSM

[ZRA23] Aiman Zara, Shafiq Ur Rehman, and Fayyaz Ahmad. Kernel smoothing method for the numerical approximation of Benjamin–Bona–Mahony– Burgers’ equation. Applied Numerical Mathematics: Transactions of IMACS, 186(??):320–333, April 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927423000168.

Zieniuk:2018:NCD

[ZS18] Eugeniusz Zieniuk and Krzysztof Szerszeń. NURBS curves in direct definition of the shapes of the boundary for 2D Stokes flow problems in modified classical BIE. Applied Numerical Mathematics: Transactions of IMACS, 132(??):

111–126, October 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418301235.

Zheng:2021:PSF

[ZS21a] Bo Zheng and Yueqiang Shang. A parallel stabilized finite element variational multiscale method based on fully overlapping domain decomposition for the incompressible Navier–Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 159(??): 138–158, January 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302622.

Zheng:2021:TLD

[ZS21b] Bo Zheng and Yueqiang Shang. Two-level defectcorrection stabilized algorithms for the simulation of 2D/3D steady Navier–Stokes equations with damping. Applied Numerical Mathematics: Transactions of IMACS, 163(??):182–203, May 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000106.

Zhang:2020:CSM

[ZSG^{+}20] Jiansong Zhang, Xiaomang Shen, Hui Guo, Hongfei Fu, and Huiran Han. Characteristic splitting mixed finite element analysis of compressible wormhole propagation. Applied Numerical Mathematics: Transactions of IMACS, 147(??): 66–87, January 2020. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419302144.

Zheng:2004:LSM

[ZSJ04] Jianmin Zheng, Thomas W. Sederberg, and Richard W. Johnson. Least squares methods for solving differential equations using Bézier control points. Applied Numerical Mathematics: Transactions of IMACS, 48(2):237–252, February 2004. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Zhu:2020:HOR

[ZSQ20] Jun Zhu, Chi-Wang Shu, and Jianxian Qiu. Highorder Runge–Kutta discontinuous Galerkin methods with a new type of multiresolution WENO limiters on triangular meshes. Applied Numerical Mathematics: Transactions of IMACS, 153(??):519–539, July 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420300805.

Zhu:2021:HOR

[ZSQ21] Jun Zhu, Chi-Wang Shu, and Jianxian Qiu. High-order Runge–Kutta discontinuous Galerkin methods with multi-resolution WENO limiters for solving steady-state problems. Applied Numerical Mathematics: Transactions of IMACS, 165(??): 482–499, July 2021. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927421000817.

Zhang:2023:USA

[ZSS23] Sihui Zhang, Xiangyu Shi, and Dongyang Shi. Unconditional superconvergence analysis of nonconforming EQ_{1}^{rot} finite element method for the nonlinear coupled predator-prey equations. Applied Numerical Mathematics: Transactions of IMACS, 185(??): 571–591, March 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422003361.

Sun:2006:FDD

[zSW06] Zhi zhong Sun and Xiaonan Wu. A fully discrete difference scheme for a diffusion-wave system. Applied Numerical Mathematics: Transactions of IMACS, 56(2):193–209, February 2006. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Zhou:2020:FVM

[ZSY20] Huifang Zhou, Zhiqiang Sheng, and Guangwei Yuan. A finite volume method preserving maximum principle for the diffusion equations with imperfect interface. Applied Numerical Mathematics: Transactions of IMACS, 158(??): 314–335, December 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302439.

Zhou:2020:IET

[ZSZZ20] Yongtao Zhou, Jorge L. Suzuki, Chengjian Zhang, and Mohsen Zayernouri. Implicit-explicit time integration of nonlinear fractional differential equations. Applied Numerical Mathematics: Transactions of IMACS, 156(??):555–583, October 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420301215.

Zhang:2006:CGM

[ZT06] Yunxin Zhang and Yongji Tan. Convergence of general meshless Schwarz method using radial basis functions. Applied Numerical Mathematics: Transactions of IMACS, 56(7):916–936, July 2006. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Zadorin:2015:TGM

[ZTZ15] A. I. Zadorin, S. V. Tikhovskaya, and N. A. Zadorin. A two-grid method for elliptic problem with boundary layers. Applied Numerical Mathematics: Transactions of IMACS, 93(??): 270–278, July 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414001044.

Zullo:2025:LFP

[Zul25] Federico Zullo. Lommel functions, Padé approximants and hypergeometric functions. Applied Numerical Mathematics: Transactions of IMACS, 209(??): 275–284, March 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742400299X.

Zuppa:2003:GQP

[Zup03] Carlos Zuppa. Good quality point sets and error estimates for moving least square approximations. Applied Numerical Mathematics: Transactions of IMACS, 47(3–4):575–585, December 2003. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Zuppa:2004:EEM

[Zup04] Carlos Zuppa. Error estimates for modified local Shepard’s interpolation formula. Applied Numerical Mathematics: Transactions of IMACS, 49(2):245–259, May 2004. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Zuur:1995:TSS

[Zuu95] E. A. H. Zuur. Time-step sequences for parabolic differential equations. Applied Numerical Mathematics: Transactions of IMACS, 17(2):173–186, June 13, 1995. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1995&volume=17&
                                        issue=2&aid=560.

Zamani:1987:AIC

[ZW87] N. G. Zamani and S. H.

Woodside. Algorithm for imposing the circulation constraint in a finite element code. Applied Numerical Mathematics: Transactions of IMACS, 3(1–2): 209–213, May 1987. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/0168927487900134.

Zhang:2009:CSU

[ZW09] Kai Zhang and Song Wang.

A computational scheme for uncertain volatility model in option pricing. Applied Numerical Mathematics: Transactions of IMACS, 59(8):1754–1767, August 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Zhao:2019:FFD

[ZW19a] Meng Zhao and Hong Wang.

Fast finite difference methods for space–time fractional partial differential equations in three space dimensions with nonlocal boundary conditions. Applied Numerical Mathematics: Transactions of IMACS, 145(??): 411–428, November 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S016892741930114X.

Zhong:2019:RMA

[ZW19b] Min Zhong and Wei Wang. A regularizing multilevel approach for nonlinear inverse problems. Applied Numerical Mathematics: Transactions of IMACS, 135(??): 297–315, January 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418302034.

Zhao:2024:WAS

[ZW24] Yuying Zhao and Xiaojie Wang. Weak approximation schemes for SDEs with super-linearly growing coefficients. Applied Numerical Mathematics: Transactions of IMACS, 198(??): 176–191, April 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424000035.

Zhang:2022:MNP

[ZWFX22] Yuanyuan Zhang, Qingbiao Wu, Yuye Feng, and Yao Xiao. Modified Newton-PSBTS method for solving complex nonlinear systems with symmetric Jacobian matrices. Applied Numerical Mathematics: Transactions of IMACS, 182(??): 308–329, December 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001982.

Zhang:2025:OEE

[ZWG25] Houchao Zhang, Junjun Wang, and Xueran Gong. Optimal L^{2} error estimates of two structurepreserving finite element methods for Schrödinger– Boussinesq equations. Applied Numerical Mathematics: Transactions of IMACS, 211(??):193–210, May 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425000121.

Zhang:2017:SOO

[ZWH^{+}17] Cheng Zhang, Hui Wang, Jingfang Huang, Cheng Wang, and Xingye Yue. A second order operator splitting numerical scheme for the “good” Boussinesq equation. Applied Numerical Mathematics: Transactions of IMACS, 119(??): 179–193, September 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417301022.

Zhang:2018:AAC

[ZWJ18] Xiaoping Zhang, Jiming Wu, and Lili Ju. An accurate and asymptotically compatible collocation scheme for nonlocal diffusion problems. Applied Numerical Mathematics: Transactions of IMACS, 133(??):52–68, ???? 2018. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417302428.

Zheng:2015:AME

[ZWK15] Mengdi Zheng, Xiaoliang Wan, and George Em Karniadakis. Adaptive multielement polynomial chaos with discrete measure: Algorithms and application to SPDEs. Applied Numerical Mathematics: Transactions of IMACS, 90(??): 91–110, April 2015. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414001895.

Zhu:2011:STO

[ZWL11] Shengfeng Zhu, Qingbiao Wu, and Chunxiao Liu. Shape and topology optimization for elliptic boundary value problems using a piecewise constant level set method. Applied Numerical Mathematics: Transactions of IMACS, 61(6): 752–767, June 2011. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Zhang:2023:PPR

[ZWN23] Jiexing Zhang, Yi Wang, and Guoxi Ni. A positivitypreserving RD-FV scheme for diffusion problems on triangular meshes. Applied Numerical Mathematics: Transactions of IMACS, 185(??):221–235, March 2023. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422003087.

Zhu:2009:PTS

[ZX09] Qiao Zhu and Aiguo Xiao.

Parallel two-step ROWmethods for stiff delay differential equations. Applied Numerical Mathematics: Transactions of IMACS, 59(8):1768–1778, August 2009. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Zhu:2014:HOU

[ZX14] Peng Zhu and Shenglan Xie. Higher order uniformly convergent continuous/discontinuous Galerkin methods for singularly perturbed problems of convectiondiffusion type. Applied Numerical Mathematics: Transactions of IMACS, 76(??): 48–59, February 2014. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927413001244.

Zhu:2022:SWG

[ZX22] Peng Zhu and Shenglan Xie. Superconvergent weak Galerkin methods for nonself adjoint and indefinite elliptic problems. Applied Numerical Mathematics: Transactions of IMACS, 172(??):300–314, February 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927421002993.

Zhu:2017:NDE

[ZXW17] Peng Zhu, Shenglan Xie, and Xiaoshen Wang. Nonsmooth data error estimates for FEM approximations of the time fractional cable equation. Applied Numerical Mathematics: Transactions of IMACS, 121(??): 170–184, November 2017. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927417301630.

Zeng:2022:ELF

[ZXYW22] Shilin Zeng, Ziqing Xie, Xiaofeng Yang, and Jiangxing Wang. Efficient, linear and fast numerical algorithm for the volume conserved nonlocal Allen–Cahn equation. Applied Numerical Mathematics: Transactions of IMACS, 181(??): 204–224, November 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422001519.

Zhang:2014:DPI

[ZY14] Tie Zhang and Shun Yu.

The derivative patch interpolation recovery technique and superconvergence for the discontinuous Galerkin method. Applied Numerical Mathematics: Transactions of IMACS, 85(??): 128–141, November 2014. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414001299.

Zhang:2019:ENS

[ZY19] Jun Zhang and Xiaofeng Yang. On efficient numerical schemes for a twomode phase field crystal model with face-centered-cubic (FCC) ordering structure. Applied Numerical Mathematics: Transactions of IMACS, 146(??):13– 37, December 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927419301692.

Zhang:2023:SFP

[ZY23] Jie Zhang and Xinmin Yang. Smoothing fast proximal gradient algorithm for the relaxation of matrix rank regularization problem. Applied Numerical Mathematics: Transactions of IMACS, 190(??): 303–320, August 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742300123X.

Zhang:2025:MQQ

[ZY25] JiHong Zhang and JiaLi Yu. A Multi-Quadrics quasiinterpolation scheme for numerical solution of Burgers’ equation. Applied Numerical Mathematics: Transactions of IMACS, 208 (part B) (??):38–44, February 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424002617.

Zhang:2022:NAB

[ZYC22] Man Zhang, Xiaozhong Yang, and Yanhua Cao. Numerical analysis of blockby-block method for a class of fractional relaxationoscillation equations. Applied Numerical Mathematics: Transactions of IMACS, 176(??):38–55, June 2022. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422000411.

Zhao:2023:PGL

[ZYH23] Hong-Jie Zhao, Haijian Yang, and Jizu Huang. Parallel generalized Lagrange– Newton method for fully coupled solution of PDEconstrained optimization problems with bound-constraints. Applied Numerical Mathematics: Transactions of IMACS, 184(??):219–233, February 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422002756.

Zhao:2025:NDP

[ZYJD25] Zhenyu Zhao, Kai Yu, Xianzheng Jia, and Zhihong Dou. Numerical differentiation of the piecewise smooth function by using Fourier extension method. Applied Numerical Mathematics: Transactions of IMACS, 208 (part B)(??): 45–57, February 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424002630.

Zhang:2023:DGM

[ZYJZ23] Min Zhang, Wenjing Yan, Feifei Jing, and Haixia Zhao. Discontinuous Galerkin method for the diffusive-viscous wave equation. Applied Numerical Mathematics: Transactions of IMACS, 183(??): 118–139, January 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927422002185.

Zhang:2020:ECW

[ZYLL20] Haixiang Zhang, Xuehua Yang, Yanling Liu, and Yuan Liu. An extrapolated CN–WSGD OSC method for a nonlinear time fractional reaction–diffusion equation. Applied Numerical Mathematics: Transactions of IMACS, 157(??): 619–633, November 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420302178.

Zhang:2021:NAA

[ZYQS21] Hong Zhang, Jingye Yan, Xu Qian, and Songhe Song. Numerical analysis and applications of explicit high order maximum principle preserving integrating factor Runge–Kutta schemes for Allen–Cahn equation. Applied Numerical Mathematics: Transactions of IMACS, 161(??):372–390, March 2021. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892742030372X.

Zhang:2023:THO

[ZYQS23] Hong Zhang, Jingye Yan, Xu Qian, and Songhe Song. Temporal high-order, unconditionally maximumprinciple-preserving integrating factor multi-step methods for Allen–Cahntype parabolic equations. Applied Numerical Mathematics: Transactions of IMACS, 186(??):19–40, April 2023. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
S0168927422003415.

Zheng:2017:UEE

[ZYS17] Haibiao Zheng, Jiaping Yu, and Li Shan. Unconditional error estimates for time dependent viscoelastic fluid flow. Applied Numerical Mathematics: Transactions of IMACS, 119(??):1– 17, September 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741730082X.

Zhang:2014:NMM

[ZYSZ14] Jiansong Zhang, Danping Yang, Shuqian Shen, and Jiang Zhu. A new MMOCAA– MFE method for compressible miscible displacement in porous media. Applied Numerical Mathematics: Transactions of IMACS, 80(??):65–80, June 2014. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927414000270.

Zhao:2020:SCS

[ZYX20] Jingjun Zhao, Yulian Yi, and Yang Xu. Strong convergence and stability of the split-step theta method for highly nonlinear neutral stochastic delay integro differential equation. Applied Numerical Mathematics: Transactions of IMACS, 157(??):385–404, November 2020. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420301951.

Zhang:2024:HMD

[ZYZJ24] Jiansong Zhang, Yun Yu, Jiang Zhu, and Maosheng Jiang. Hybrid mixed discontinuous Galerkin finite element method for incompressible miscible displacement problem. Applied Numerical Mathematics: Transactions of IMACS, 198(??): 122–137, April 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423003203.

Zheng:2017:DGF

[ZZ17] Yunying Zheng and Zhengang Zhao. The discontinuous Galerkin finite element method for fractional cable equation. Applied Numerical Mathematics: Transactions of IMACS, 115(??): 32–41, May 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927417300053.

Zhan:2018:AOS

[ZZ18] Rui Zhan and Jingjun Zhao. The analysis of operator splitting methods for the Camassa–Holm equation. Applied Numerical Mathematics: Transactions of IMACS, 130(??): 1–22, August 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927418300746.

Zhang:2019:SGA

[ZZ19a] Lu Zhang and Zhaojie Zhou.

Spectral Galerkin approximation of optimal control problem governed by Riesz fractional differential equation. Applied Numerical Mathematics: Transactions of IMACS, 143(??): 247–262, September 2019.

CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419300844.

Zhou:2019:CSB

[ZZ19b] Yongtao Zhou and Chengjian Zhang. Convergence and stability of block boundary value methods applied to nonlinear fractional differential equations with Caputo derivatives. Applied Numerical Mathematics: Transactions of IMACS, 135(??): 367–380, January 2019. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927418302071.

Zheng:2020:TDS

[ZZ20] Yunying Zheng and Zhengang Zhao. The time discontinuous space–time finite element method for fractional diffusion-wave equation. Applied Numerical Mathematics: Transactions of IMACS, 150(??):105–116, April 2020. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419302466.

Zhang:2024:BDW

[ZZ24] Shangyou Zhang and Peng Zhu. BDM H (div) weak Galerkin finite element methods for Stokes equations. Applied Numerical Mathematics: Transactions of IMACS, 197(??): 307–321, March 2024. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423003008.

Zhu:2025:RAE

[ZZ25] Kai Zhu and Min Zhong.

2D and 3D reconstructions in acousto-electric tomography via two-point gradient Kaczmarz-type algorithm. Applied Numerical Mathematics: Transactions of IMACS, 218(??): 220–237, December 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927425001564.

Zhou:2018:WGM

[ZZC^{+}18] Chenguang Zhou, Yongkui Zou, Shimin Chai, Qian Zhang, and Hongze Zhu. Weak Galerkin mixed finite element method for heat equation. Applied Numerical Mathematics: Transactions of IMACS, 123(??): 180–199, January 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741730185X.

Zhang:2025:NEG

[ZZF25] Bingyin Zhang, Chengxi Zhou, and Hongfei Fu. A novel efficient generalized energy-optimized exponential SAV scheme with variable-step BDF k method for gradient flows. Applied Numerical Mathematics: Transactions of IMACS, 210(??):39–63, April 2025. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927424003465.

Zhang:2025:FEE

[ZZH25] Xindan Zhang, Jianping Zhao, and Yanren Hou. Finite element error estimation for parabolic optimal control problems with time delay. Applied Numerical Mathematics: Transactions of IMACS, 212(??): 176–196, June 2025. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425000261.

Zhang:2018:OEE

[ZZHS18] Yuhong Zhang, Haibiao Zheng, Yanren Hou, and Li Shan. Optimal error estimates of both coupled and two-grid decoupled methods for a mixed Stokes–Stokes model. Applied Numerical Mathematics: Transactions of IMACS, 133(??):

116–129, ???? 2018. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S016892741830031X.

Zheng:2021:LSM

[ZZJ21] Rumeng Zheng, Hui Zhang, and Xiaoyun Jiang. Legendre spectral methods based on two families of novel secondorder numerical formulas for the fractional activatorinhibitor system. Applied Numerical Mathematics: Transactions of IMACS, 162(??):235–248, April 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303871.

Zhao:2024:SCB

[ZZjQ24] Xuan Zhao, Haifeng Zhang, and Ren jun Qi. Stability and convergence of BDF2-ADI schemes with variable step sizes for parabolic equation. Applied Numerical Mathematics: Transactions of IMACS, 203(??): 173–185, September 2024. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927424000436.

Zhang:2001:DCG

[ZZL01] C. J. Zhang, S. Z. Zhou, and X. X. Liao. Dconvergence and GDNstability of Runge–Kutta methods for a class of delay systems. Applied Numerical Mathematics: Transactions of IMACS, 37(1–2): 161–170, April 2001. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL

http://www.elsevier.nl/
gej-ng/10/10/28/73/26/
35/abstract.html;   http:
//www.elsevier.nl/gej-
ng/10/10/28/73/26/35/article.
pdf.

Zheng:2017:NFD

[ZZL17] Quan Zheng, Xin Zhao, and Yufeng Liu. A novel finite difference scheme for Burgers’ equation on unbounded domains. Applied Numerical Mathematics: Transactions of IMACS, 111(??): 1–16, January 2017. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301696.

Zhou:2021:ROE

[ZZLL21] Yanjie Zhou, Yanan Zhang, Ye Liang, and Zhendong Luo. A reduced-order extrapolated model based on splitting implicit finite difference scheme and proper orthogonal decomposition for the fourth-order nonlinear Rosenau equation. Applied Numerical Mathematics: Transactions of IMACS, 162(??):192–200, April 2021. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420303937.

Zhao:2016:SAE

[ZZO16] Jingjun Zhao, Rui Zhan, and Alexander Ostermann. Stability analysis of explicit exponential integrators for delay differential equations. Applied Numerical Mathematics: Transactions of IMACS, 109(??):96– 108, November 2016. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927416301076.

Zhang:2023:HMF

[ZZPJ23] Jiansong Zhang, Jiang Zhu, Héctor Andrés Vargas Poblete, and Maosheng Jiang. A hybrid mixed finite element method for convection– diffusion–reaction equation with local exponential fitting technique. Applied Numerical Mathematics: Transactions of IMACS, 189(??): 23–38, July 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927423000703.

Zhao:1997:CTN

[ZZW97] Shuangsuo Zhao, Guofeng Zhang, and Changyin Wang. The convergence of two Newton-like methods for solving block nonlinear equations and a class of rpoint (r +1)^{st}-order A-stable one-block methods. Applied Numerical Mathematics: Transactions of IMACS, 25(1):117–133, September 16, 1997. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.elsevier.com/cgi-
                                        bin/cas/tree/store/apnum/
                                        cas_sub/browse/browse.
                                        cgi?year=1997&volume=25&
                                        issue=1&aid=801.

Zhang:2019:PDM

[ZZX19a] Benxin Zhang, Zhibin Zhu, and Chuanpei Xu. A primaldual multiplier method for total variation image restoration. Applied Numerical Mathematics: Transactions of IMACS, 145(??): 145–158, November 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927419301588.

Zhao:2019:AOS

[ZZX19b] Jingjun Zhao, Rui Zhan, and Yang Xu. The analysis of operator splitting for the Gardner equation. Applied Numerical Mathematics: Transactions of IMACS, 144(??):151–175, October 2019. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL https:/

/www.sciencedirect.com/
science/article/pii/S0168927419301047.

Zhao:2020:IRK

[ZZX20] Jingjun Zhao, Yanming Zhang, and Yang Xu. Implicit Runge–Kutta and spectral Galerkin methods for the two-dimensional nonlinear Riesz space distributedorder diffusion equation. Applied Numerical Mathematics: Transactions of IMACS, 157(??):223–235, November 2020. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927420301859.

Zhao:2019:NVE

[ZZZ19] Jikun Zhao, Bei Zhang, and Xiaopeng Zhu. The nonconforming virtual element method for parabolic problems. Applied Numerical Mathematics: Transactions of IMACS, 143(??): 97–111, September 2019. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.sciencedirect.com/
science/article/pii/S0168927419300832.

Zhang:2023:CAH

[ZZZ23] Zeze Zhang, Hongchan Zheng, and Jie Zhou. Convergence analysis of Hermite subdivision schemes of any arity. Applied Numerical Mathematics: Transactions of IMACS, 183(??): 279–300, January 2023. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S016892742200232X.

Zhang:2026:RGD

[ZZZX26] Min Zhang, Zimo Zhu, Qijia Zhai, and Xiaoping Xie. Robust globally divergencefree HDG finite element method for steady thermally coupled incompressible MHD flow. Applied Numerical Mathematics: Transactions of IMACS, 219(??): 19–40, January 2026. CO-DEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

                                        /www.sciencedirect.com/
                                        science/article/pii/S0168927425001710.